A Bibliography of Publications about Supercomputing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

15 August 2025 Version 2.182

Title word cross-reference

#46 [Ano98e].

(b, k) [AC84a]. (r_{∞}, n_{1/2}, s_{1/2}) [Hoc85]. (r_{∞}, n_{1/2}) [Tem89a]. 1/f [PN96]. $100 [And10]. $100M [Mac96b]. $125.00 [Ano00a]. 16 × 16 [GJW91, Has84, LH86a, LH86b, LH86c, LH86d]. $20 [Ano88r]. 20 [NPS^{+}20]. 3 [ACK^{+}95, CGLY96, CS90, CMAS11, DGO90, EFR^{+}05, FDM07, IHE^{+}00, JB90, KSM^{+}08, KTN^{+}14, MKDY90, Mir88, OPR01, PLS20, Pau08, PPM90, WLH00]. 4 [WLH00]. $55.00 [Ano96c]. $6 [Ano95v]. 6 [FMD07, RWL^{+}98]. TM [BE92, Blu92, Cyb91b, SSRL91]. γ [Her94]. K [YZL^{+}20, BY21, OGR95]. μ [AT93a, AT93b]. N [Ano94-59, Ano94-141, BAD01, SHMH97, Ano94-116]. Π [Rau91]. R [SB81, SB82a, Rav92, Rav95]. s [SC92]. S_{n} [ARW93a]. SU(3) [MHP84, KM85]. θ [Che91]. × [FT93a]. X + Y [AG94]. Z [IMA93].

-Adjacent [AC84b, AC84a]. -Body

[Ano94-116, Ano94-59, Ano94-141, BAD01, SHMH97]. -CLF [Her94]. -D [KTN^{+}14, CS90, FMD07, IHE^{+}00, Mir88, Pau08]. -Dimensional [JB90, RWL^{+}98, Rav92, Rav95]. -Dimensions [OGR95]. -Lattice [GAW96b, GAW96a]. -matrix [SB81, SB82a]. -means [BY21, YZL^{+}20]. -Qubit [NPS^{+}20]. -Step [SC92]. -ultimate [BDRR94].

/NII [Voi94].

0*T [ACA94]. 0-8493-4417-4 [Ano94p].

1 [Ano94h, Ano94-135, Ano94-130, Asa93a, AG90, Bak10, BK77, BCK13, Cal81, Dic81, Dic82, DFS93, DR81, DR82, Duf82, ER94, EM78, FR81, Fin82, Gin82, Hus86a, KJ85, Kol81, MSAD91, McB92a, McB92b, Mes93a, MW81, Mon93, OLLG96, PK80, PCM84, Pet83, Rus78, SG81, SG82, SMFG85, SB81, SB82a, SB82b, SBHW80, hTD88, Tem83, Tem88, Tem89a, Tem89b, WG82, WSL88, WS84d, WL83, Yuv77]. 1-D [Ano94-62]. 1-Gbit [Ano94-135]. 1-Gbit/ [Ano94-135]. 1-S [WL83]. 1-Teraflop [Ano94-130]. 1.3-ms [Ano94-135]. 10 [Ano88h]. 10-11 [Vag88]. 100 [HT72, IMP93, Str94]. 100/200 [TKI85]. 1003.10 [IEE95a]. 1003.10-1995 [IEE95a]. 1012 [CK92a]. 10th [IEE96c, Str03, VAS82]. 11 [Ara14]. 11-15 [Sig90a]. 1100/90 [Hod87]. 11th [Ano92a, FJSP95]. 12 [OHIB93, OHHIB94]. 13-16 [C^{+}97]. 13-18 [L^{+}95]. 14-18 [Soc94]. 14th [Ano93c, Ano94-108]. 15 [Ano94h, DHT89]. 15-17 [LCV90b]. 15-19 [IEE93d]. 15th [Ano93d]. 16-20 [CBCH93]. 16-bit [Has84]. 164 [Nor03, WG82]. 164/MAX [CG86]. 17-22 [IEE96d]. 17th [Ano95u, FL92]. 187th [Ano95q]. 19 [BCH^{+}22, Kra20, VSB^{+}21]. 1980’s [Sci86]. 1983 [Fer83]. 1984 [Por86]. 1985 [Dup87, IEE85, KK85]. 1985-1995 [Mac96c]. 1986 [Joh86c, TR86]. 1986-1990 [Meu92b]. 1987 [Ano97b, Sch88a]. 1987-1989 [Man89a]. 1988 [DHT89]. 1989 [Ano88d, BP89b, LCV90a, MB94b, Tec89]. 1989/90 [Eid91]. 1990 [ACM90, Ano90g, GL90]. 1990s [Nat89a, SC91b, TAM^{+}91]. 1991 [Ano91q, Ano94-126, BH92b, LLR93b, MN91, NBC92]. 1992 [Ano92a, Ano93b, HS94b, HK93b, KH93, KK93, KWW92, Meu92c, PEH93, Tho93c, Uni92a, Zyg93]. 1993 [Ano93n, BP93, Bro93, HK93a, JPTE94, Kah93a, Meu93, NAS93, Pel93a, SR93b, ZAS94]. 1994 [Ano94-107, Ano94-108, GT94, Ham94, HWP95, IEE94e, Kho94, M^{+}94, VV95].

1995 [ACM95a, B^{+}95, DDC96, FJSP95, IEE95a, IEE95d, KMG96]. 1996 [Cra96]. 1997 [Uni98]. 1998 [Lid99]. 1999 [Ano98b, Ano98c]. 19th [Ano96-43]. 1M [HL88a]. 1S [SA82, SA83, THH81, THH82]. 1st [Ano96k, GLH94, Guo94, HPP88, KFF93b, LM92, Pra95].

2 [AABK95, Ano97i, Asl91a, Bai88, BCM94, BHM94a, BHS^{+}02, BS04, BB13, BÇG14, But92, Cal86a, Cal86b, Cal88, Car91, Car94b, Cha84, CYXL18, CDH84, CGS91, DCG90, DD87, DD90, DL90, Dub87, DS94c, EJL90, Elm93, Ess90, FG87, FK04, FSY88, Gis86, GD94a, HB93, Hel92, HS94d, Ho88, Hoc94, jJ88, KN88, Lar84, LMP^{+}90, LXW^{+}16, LMM85b, LMM85a, LM90b, LSK04, MLR90b, MLR90a, MNV93, Mon88, MDW93, NSH95, PCY^{+}19, PZS^{+}20, Pol88d, PO88, PTS93, PK89, PK94, Ric90b, Ric91a, Ric91b, SN95a, SN95b, SI90, SI91a, SI91b, Tze88, VDK91, WCZ^{+}18, WQS92, WFT93, Wil90a, WVBM88a, WVBM88b, Yau88, ZH88, ALPP00]. 2- [Sus93]. 2-CPU [Hoc85]. 2-D [AABK95, DS94c, Elm93, GD94a, HS94d, MDW93, WQS92]. 2-Stroke [HB93]. 2.44 [IHE^{+}00]. 2/400 [MM91b]. 20 [DH86b, LMM85b, LMM85a, LMM86]. 200 [DH86b, HL88a, LMM85b, LMM85a, LMM86, McB93, MU83, TKI85]. 2000 [LSK04, PIH04]. 2001 [Coc02c, Coc02d, Pin01]. 2003 [BCCP05, Stu03]. 2010 [War10]. 2021 [Ano22]. 205 [Dic81, Dic82, Mil88b, Tem83, Uni87b, WL83]. 2051-02 [Bur93]. 210th [Cul95a]. 21164-Based [Was96b]. 216-Processor [MDH00]. 21st [Bel92a, Bor01, Joh97]. 2230-12 [YW94]. 22nd [ACM95b, Ano97-30]. 23-27 [IEE94b]. 23rd [Gra93c]. 24 [GKL^{+}87, LMM86]. 24-28 [SEA84]. 24.-26 [Meu93]. 2435-71 [HCPS95]. 25-27 [Bup87]. 25.-27 [Meu92c]. 255.00 [Ano00a]. 25nm [Ano03a]. 25th [Ras91]. 2656-26 [Che96]. 2692-04 [BBBC96]. 26th [Ano93i, Ano93-31, Isk96]. 27th [Ano94a, Ano94-75, EP 97]. 2969-56 [SSSSE96]. 29th [Ano96a, Rol96]. 2D [BTV96, RDHC94, SB94b, SJPS94, SJPS96, TM94a, TM94b]. 2D-Position [RDHC94]. 2D/3D [SB94b]. 2nd [Ahm92, AGP96, AB94, HS^{+}91, IEE93b, LCV90b, RMO96, LCV90a].

3 [AGZ94a, Bac88, CGL92, DMPR93, Elm93, Iwa90, KLY94, KG95, KBLD08, MAA93b, RRSS93, Sch97c, SHZK94, TW92, VNB93, VTTS98, VH93b, Wat93]. 3-7 [Sig95]. 3-94 [BBDS94]. 3-D [AGZ94a, DMPR93, Elm93, KLY94, MAA93b, Sch97c, SHZK94, VTTS98, VH93b]. 3-D-spectral [DP90]. 3-D-spectral/finite [DP90]. 3-Dimensional [Sus93]. 3.0 [CSFS00]. 3.06 [Ano03a]. 3.8GB [RWNJ94]. 3.8GB/ [RWNJ94]. 300 [IHE^{+}00]. 300-MHz [IHE^{+}00]. 3090 [Kha91]. 31st [Ano94a]. 32 [FG87, Gri86, Kah92]. 32-bit [Has84, Wei91]. 32-GFLOP [Kah92]. 32nd [AIA94]. 3300-Itanium-Prozessoren [Ano01b]. 369 [GKS09]. 36th [Ano98a]. 370 [MPSB87]. 3800 [IAKH92, KISY94, SKIY94]. 3D [CGLY96, CGLxx, GE12, Ano94-89, Ano97-29, BV96, CS89, CGLxx, Chexx, GD94b, HB93, jJ88, KTKK93, LM93, LXW^{+}16, PS88, SB94b, SRBL94, WH93, WLKI95, Whe83]. 3D-CAD [KTKK93]. 3rd [Ano96a, Hel93, PH95, SC93, LLR93b].

4 [Ano03a, BJ95, BAM93, DH86b, DH86a, HLPP97, HMKI97, Hor97b, Hor97a, KE93, STSK95, TOY96, Wat93, YSK^{+}96, YMY92]. 4.0 [Mon88]. 4.5 [TMT^{+}20]. 40 [DAC^{+}18, Hab86, WSL88]. 400 [MM91b]. 416 [VY88]. 48 [Ber90a, CK90, HFH86, HFH87, Meu87, Nag88, VM87]. 4D [Kau93a]. 4M [DTV00]. 4th [Ano94a, Goo97, IEE97a, Pow97, SJD96, USE00a, USE00b]. 4WA [FT93a]. 4WB [FT93a]. 4WD [FT93a]. 4WS [FT93a, YOY97].

5 [CS93a, HT93, HP95, KC95, KR94d, Lee96, LW94, Mar95, McB93, PTC^{+}93, PW94, RYYT89, SNS^{+}97]. 5-9 [Ano97q]. 500 [CP13, FB91b, Heg96]. 500-MHz [FB91b]. 512 [CKT21]. 5d [GE12]. 5th [Ano01a, IEE96b, ML95b, NBC92, USE01].

600 [LSK04]. 6000 [Yuv77, MSAD91]. 622 [Lev98, RRW84]. 64 [SCV01]. 64-Bit [AT93a, AT93b, Ano01c, Hir92c, KM89, Mac90]. 6D [FDM07]. 6SF [Ano94p]. 6th [Ano96a, DLM99, GT94, Hen97, IEE96a].

7-11 [ACM97]. 737-LGA [Cha94b]. 758 [BTV96]. 759 [BV96]. 76 [HL88a]. 7600 [Bro17a]. ’77 [KK89c, HWS^{+}88]. 7F [Par90b]. 7th [Ano93a, LM92, ME96].

8 [ASM86, Ano87e, WSL88]. 80

[DD90, GB90]. 80860 [IAIK92]. 810/20 [DH86b]. ’84 [Ano85b]. 8400 [Was96b]. 860 [Ano94-117, HL96, PL94, Rot94, VSB94, YSS94, YR93]. 864 [Cho90a, SO91]. ’87 [KK87]. ’88 [Ano88t, Ano88u, KK88, ML89]. 881 [FGG09]. ’89 [ACM89a, KK89a, Meu89a, UL89, Voe89a, Voe89b, Zor89b, Zor89a, Ano88e]. 8th [Ano91m, BBM96, De 96].

’90 [Ano90f, IEE90, Meu90, PL91c, SIG90b, Zor90, CK92a, Car94a, CWLT97, DP96, Fox90a, Lee96, Mar92]. 900 [HE98]. 90D [Ano94-69]. 90er [Kro92]. ’90’s [Con91, Con90]. ’91 [BG91, Meu91, Ano94-126, Mye92a]. ’92 [Hel93, Kah92, KK93, Meu92c, PEH93, Nat91a]. ’93 [Ano93-39, Ano93n, IEE93a, IEE93d, KSW93, Meu93, Pel93a, SC93, VO93, Dra94b, Lew93]. 93SC001 [OK93]. 93SC003 [HTI93]. 93SC005 [KC93c]. 93SC006 [LCVR93].

93SC007 [GP93b]. 93SC010 [KD93]. 93SC011 [KC93a]. 93SC012 [GBF93]. 93SC014 [PCK93]. 93SC016 [Vui93]. 93SC018 [HB93]. 93SC019 [GMBW93]. 93SC021 [Maa93a]. 93SC022 [GWG93]. 93SC023 [BCW93]. 93SC024 [CZRB93]. 93SC025 [BD93b]. 93SC026 [Moi93]. 93SC028 [HMS93]. 93SC029 [BIRB93]. 93SC031 [PBDM93]. 93SC035 [BE93a]. 93SC038 [FS93b]. 93SC040 [VF93]. 93SC041 [Gle93]. 93SC043 [Jab93]. 93SC044 [MI93]. 93SC045 [Him93]. 93SF017 [AVS93]. 93SF055 [Van93]. ’94 [Ano94-134, DJM94, GT94, IEE94a, IEE94e, Kho94, Soc94, Dra94a, Hol95, Qui95]. ’95 [ACM95c, Ece96, HBCN95, IEE95c, ML95b, Ano95-34, Dra96a, IS95]. ’96 [ACM96, Ano96b, De 96, IEE96d, Ano96t, Ano97i, Dra96b]. ’97 [IEE97b, Ano97i, Ano97l, JLC98]. 9th [Ano93g].

= [Ano93i, Lin83].

A&M [Nas91]. A/E/C [Kho94]. A2Cloud [SKB^{+}20]. A2Cloud-RF [SKB^{+}20]. Aachen [Ano93i, Ano93-31, Ano94a, Ano94-75]. Abarbanel [Por86]. AbBT [IK91]. ABCL [YMY92]. ABCL/onEM [YMY92]. ABCL/onEM-4 [YMY92]. absolute [Gre90b]. Absorbers [FSGS93]. Absorption [WRW93]. Absorptive [Bla97]. Abstract [Ano94-41, Mas91, Smi96b, AH90, CKM88, SKP91, SG92d]. Abstraction [Mas92, ML97]. Abstracts [Cor89b, Nor93a, PP91]. AC [CS93a, CD95a, CD95b]. Academia [Coh91, Fed96]. Academic [Ano87a, Ano00a, McD88]. Academic/Plenum [Ano00a]. Accelerated [BHEG94, MDW93, SCSL12, BWV^{+}17, KSP13, LSS^{+}20]. Accelerating [PMS^{+}08, SGIS93]. Acceleration [AM93a, Now93, War93a, Bra89a, SWS^{+}12]. Accelerator [Ano90o, Ano09, HBKR96, PSS^{+}19]. Accelerators [BKK11, DF12, KWB^{+}10, cFM07, Ipe19, KBD10]. Accelerometer [LKH94]. Accelerometers [UU94]. acceptance [HLA^{+}18, Lie20]. Access [ACKW01, Ano94-37, Ano97p, Bar00c, Bar00d, BGH^{+}02, CMHK92, Dal84, Hay84, Ho88, Lew94b, LLY92, MOWW96, MH96, MP92, MK07, Nat87b, Nat87c, OM91, PVA94, Pro01, SKIY94, SKIY97, VLA92, AS99, Ano97-27, CD95b, Cyr86, EE93, EHHS89, Gra92, Jay87, Joh86c, Lee86, LC91, Lim91b, SFL^{+}94, SS07, Yau88]. accesses [GV92, LB96]. acceSX [Ho88]. Accident [DM93, Gon93, IK93, JR94]. Accidents [PA93b]. accommodate [Dra89]. Accuracy [Bal93, KT94, GB22, SKB89]. Accurate [VTSM12, ZEC^{+}17, ZBN^{+}19]. Ace [RR99]. ACh [UR95]. ACh-Driven [UR95]. Achieve [EKZ90, Bre87]. Achievement [Ano94-32, Coc01, War03]. Achievements [Pin01, Tho93a, Ano95l, Gil94a]. Achieving [Eck93, GGG^{+}98, GMSB93, KCZJ14, MBSK92, MBK^{+}92, WWJ09]. Acid [TYKE93]. acids [MW88]. ACM [Ano98b, Ano98c, FJSP95, Ano96b, Coc02a, Coc02b, IEE96d]. ACM/IEEE [ACM95c, ACMxx, Ano96b, IEE96d]. Acoustic [AHH94, BS97, JBWB97, NNSY94, LCV90b]. Acoustics [IEE95b, Ano96q, App96, HL88b, LLR93b, LCV90b, LCV90a]. Acousto [Ano90a]. Acousto/optical [Ano90a]. Acquasparta [Pow97]. Acquisition [Bel93, BK91a]. Acquisitions [DCWH07]. across [Ano95-31, MD04, vL99]. ACS [Con11]. Action [Lat16]. Active [CMAS11, HU93, Lie93, SQM94]. Activities [Ano97-28, Van91a, JG99]. Activity [HV95, JBI91, KVP95, RSRG95, VT95, CHWW13, JR91]. Acts [AHFK93, Kon96]. Actuators [KP94]. Acyclic [Ano94-56, BP92]. Ad [YFY^{+}13, Ano94h]. Ad-hoc [YFY^{+}13]. ADA [Ano86a, LPS90, SB94c]. Ada95 [Sti98a, Sti98b]. Adam [Ano95v]. Adaptable [VKK80]. Adaptation [HTV88, Sch88b, GFBR10, Kar13]. Adapting [Ano93b, MRSB94]. Adaptive [ABCH97, Ano94c, Ano94d, Ano94-53, Ano94-112, BA95, CC94b, CB00, EHS94, Gal96, GCS94, HS93a, JP94, KV96, KCZJ14, LN94, LL08, MS94c, Ost94, RE94, Sch97c, TMAS97, VTTS98, WMR96, Zas93, ZM94, Aba09, BTV96, BV96, Bru91, Fuj11, MV16, Use93, MD04]. adaptive-grid [BTV96, BV96]. Adaptivity [PDR94]. Add [CKS99, PSS^{+}19]. addendum [Ano91n, Ano91o]. Adding [The90b, The91, BJ95]. Addison [Sch88a]. Addison-Wesley [Sch88a]. Additive [Alu96, BHW98, Mas94b]. Address [KNS95, SLB93, TAAL95, YQTV12]. Addressing [HG02, OM91, Pau08, RG92, YTL87]. Adds [Smi95, Ano94-120]. ADEAS [AHH94]. Adelaide [NBC92]. Adhara [AZ94]. Adjacent [AC84b, AC84a]. Adjoint [GT91, MF93, TYK93]. adjusted [TDBL13]. Adjustment [IHK93, IHSK93, OSKO95]. adl [SK93a]. Adleman [Bas95b]. Administration [Ano93b, Per83, Uni92b, Uni92a, Mob12]. Admission [PH11]. Adopts [Bar01]. advance [Ano92l]. Advanced [AM93b, And90b, Ano94-136, Ano97o, Bha94, Cra96, D^{+}95, DMKW93, FWWD95, GY93b, HS94b, HNST93, KT94, KHC14, KWW92, Kow89b, Lag89, MS94b, OS93, OMM93, PW86b, PW86c, PW86a, Pol87a, SKVZ93, SLRP95, Vuj93, Ano90j, AG90, Asl91a, BMW91, FMT91, GBB^{+}05, LEY86, Pol87c]. Advances [ALPP00, Ano90m, Bor92, GK18, MLY10, DDJ98b, BBM19, COS89, DLM99, OMM93, San91, HBCN95]. Advancing [Ano00b]. advantage [PL91c]. Advantages [DT96, VM94]. Advection [CT94, LC97b, LS93b]. Adventures [ORS94, HS96, HSxx, OHIB93, ORSS94, OHHIB94, SS95, SGH97]. Advisor [Per83]. advisory [Joh88]. aeroacoustics [L^{+}95]. AERODAYS [Pel93a]. Aerodynamic [GW93c, Him93, OK93, YF95, BBC^{+}99]. Aerodynamical [PPM90]. aerodynamics [BPM^{+}89, HP88b, PB88, SD88]. Aeronautical [Pel93b]. Aeronautics [Pel93a]. aerosol [Ano97c, Pan97]. Aerospace [AIA93, AIA94, Ano98a, IEE94b, LPC^{+}95, PC94b, Pet89b, RG94, SHMR93, AU87, Uni87a, VVH95]. AES [LFJ^{+}20]. affect [Kra97]. affected [WH94]. Affiliates [Ano87a, Fer83]. affine [CK90, Kor93]. Affinity [Ano94e, LS94, Ste94a]. affordable [AGEL13, Ano88a]. African [New93]. After [Ano92b, Ano95w]. Afterword [DM88a, DM88b]. again [Ano00b, Sch12]. Against [Ano95-46, VSB^{+}21]. Agarose [HPLC93]. Age [Fox89, Gha96, Ren97, PMC22, Rya90, CCKSS90]. Agency [Ano93-29]. Agenda [Ano94w, Inf86]. agents [SNEP14]. Ages [Opp95b]. aggregate [FGC06, YFY^{+}13]. Aggregation [MS96]. agree [Ano93b]. agreement [Uni92c]. agreements [KW11]. Ahead [Bel99, Jon96, Zim96, Ano98f, CSFS00, New14]. AHS [DCG93, DCGxx]. AI [Bar00c, Bar00d, HHT^{+}94, Hug93, LQFC18, Sri94, Ull84, WZ87]. AI-based [Sri94]. Aid [FNK93, SPK94, Ano94-36, KK89c]. Aided [KC93a, KD93, MM90, RC94, RLC91]. Aiding [TSSK94, VRSG93]. aids [Ano95i, Ano96-33, HSW^{+}90]. Aims [Ano93o]. Air [AABB93, ABCE97, Ano93-46, Ano94-48, Car94a, Cha94b, EDJ^{+}10, Fie93, FA93, Hai97, HCV97, KY90, LYKM97, SSKR97, SLS96, TMAS97, Zla01, Ano89f, KfGERJxx, ODAZ15]. Air-Cleaner [LYKM97]. Air-Cooled [Car94a].

Air-Cooling [Cha94b]. air-sea [KfGERJxx]. Airborne [Rhe90, SKSD94]. Aircraft [Law90, RG94, Riz94]. Aircrafts [NSF90]. Aizu [M^{+}95, Ike95]. Aizu-Wakamatsu [M^{+}95]. Alabama [Alaxx]. Alamos [AB94, BBB^{+}91, CKS99, Met86a, Ano99, Lew17, Mac91b]. Alan [Ano00a]. ALAS [Mil97a]. Alaskan [OLLG96]. Albatross [KBM^{+}02]. Albuquerque [Ano94-126, NAS93, Bor92]. Alfred [War03]. Algebra [ALPP00, Cal81, CDH84, CDW94, Dem91, Don91, Dub87, Ede94a, NJL94, NGDH96, DHD89, DH86a, DS86a, Don93a, GJM86, Gal87, GPS90, Hak89, Ipe19, Sim00]. algebraic [Ano87e, ARW92, EGK87b, FRW92, Ked94]. Algebrain [Moi93]. AlGoAs [MS84]. Algorithm [AABB93, AGZ94a, AG94, Ano92e, Ano94-42, Ano94-62, Ano94-92, Ano94-95, Ano94-115, AZ99, Bak93, BE93a, Ber90a, BS94d, CJ93, Cha84, CJHH94, CS93b, DA94, DE84, ES92, EH97c, FC93, Hun94, KMNT95, KMNT96, KK95a, KESH94, LS93b, Mik94, Moh94, MBW01, MM94c, MM94d, NK94, OIY91, OS94, PC93, PK94, RCS21, Sch96, Str97, SA94, TL96, TH94, TW92, VR94, Vog93, WMKS96, WLKI95, WK95, WM91, XCLW93, KC93b, AGZ94b, ARW93a, Amm90, Amm92, ARE95, AT89, BY21, Bis94c, Bra92, Bru91, Cha92b, CK90, CFP20, Dav86b, Dav89, DY90, DS86b, Dra90a, EB18, FMD07, FB91a, Gal87, GS87a, GS88b, GJ87, GS89d, Hog02, Hol90b, HLTZ93, JP90, KNHN16, KFML20, KTN^{+}14, Kor93, KM85, KESH95, LS93a, LPS86, LYC93, Luc91, MHP84, Nee90a, Par90c]. algorithm [PS98, Pin91, PW23, PK89, Sar91, Sch88b, SCK^{+}00, TCM95, Tem88, WLLZ20, YW94, BTV96, BV96, FGG09, Lev98, RRW84, SZG95]. Algorithmic [Ber89a, SHMR94, SHMR96, HKS93]. Algorithms [ALPP00, Ano94-44, Ano94-51, Ano94-53, Ano94-57, Ano94-84, Ano94-91, Ano94-141, BOS93, BP86, BHEG94, BS01, Bro96, Cal86a, Cal86b, Car93, CHL93, CDMW94, CDH84, Che92a, CMF94, CCSM97, DHD89, DLMW95, DT96, Elm95b, FR96b, FR96c, FT94, FWWD95, GS89c, Gen94, HL93a, HS94c, HCL94, Hem84, HHK94, JM90, JP94, KN88, LH94, LPV94, Maj94, MP94, MT86, MTH88, M^{+}95, MRAR95, NJL94, NGLPJ96, OLLG96, Pap92, PP93, PS94b, RS94a, Rie93, SH90, Sha94b, Sin94b, SJPS96, TS94, TZY88, WG91, tDv87, Ano97-29, Bad08, BR95, Ber86b, Ber86a, BS88a, Ber90c, Bis93, BS90a, Bru91, Cal96, Car89a, CBCJ92, CLR09, Cho90a, Con87a, CB89, DH86a, DGH20, Dra91b, GPS90, Hea91, JM89b, Kar89, Kha95, Kon91b, Kra93, Lag89, LD90, Li91, LHPG21, Mag10, MS88, ME91]. algorithms [MO88, MP87a, MP87b, Mik89, MRSB94, MKfDA96, Nat88a, Noo95, ODAZ15, OW94, Qui87, Ram86, Sam85, SL88, TR86, TT93, TB89, Woo92, Woo94, YFY^{+}13, YHA93, Gig94, Sam91]. Algorithms/Architecture [M^{+}95]. Alias [WGW04, Woo05]. Alignment [SSM93, BS04]. Alignments [GSB95]. Alive [Ano96-38]. All-Port [TM94a, TM94b, WK95]. ALLIANT [Ano87e, DD90, ASM86, Ano88b, Ano90b, Bie88, Gan86, GB90, WSL88]. Allocation [Ano94-84, BBD^{+}08, KKF96, KCZJ14, MD94, TF92, WMKS96, ZYL^{+}16, ADG^{+}05, PLC^{+}19, PB87, WWJ09]. Alloy [TSCG94]. Almost [Ano95z, GP93a]. Alone [DDJ98b]. Alpha [Was96b, SW94]. Alphabet [Voi94]. AlphaServer [Was96b]. already [Ano94-121, Tri95a, Tri95b]. also [Wat92b]. alteration [Ano90o]. Alternating [YA93]. Alternative [HCV97, Sti98a, Sti98b, EKTB99, TF15]. Alternatives [Bel96, EHG01]. Although [Rya90]. Altix [Fat10]. always [Par94]. Amazon [MDH^{+}16]. AMBER [MM91b]. ambiguity [Car94b]. Ambiguous [JC94b, YOY97]. Amdahl [dRSGS16, Man92]. America [ANS92, Gra93c, AS93, GGN20, Smi95]. American [Kho94, Ano97j, Coh91, Dal84, GGN20, JT87, Men84]. Ames [AU87, BPM^{+}89, Uni87a, Sim92a]. Ametek [Ano88h]. Amfisbaena [VVKB96]. Amino [TYKE93]. AMLCD [KFJB94]. Amon [KMNT95]. Amon2 [KMNT96]. among [LZF16, NMS93]. AMR [NĆ02]. Amsterdam [ACM90, Emm85, Sig90a, Tru88, tDv87]. Analog [Lan94, MHW94, PB98]. Analogic [Ros93b, Ros95]. Analyser [PNK93]. Analyses [San93]. Analysis [AKDM93, ADLL01, AH93, Ano94-50, Ano94-96, Ano94-114, ASNT91, BGS^{+}12, BGIM90, BSB93, BBC92, BW94, BJH97, BCCG97, BK95b, Cal85a, CP25, CH94, Che94b, CS84, DZM^{+}13, DH93, DT96, ER94, Eij90a, ES92, ES96, EO91, EHG01, FBH93, FV94, GG96, GCB92, Glo84, GA84, GW93b, Gua87a, Har94a, Hei89, HL91, HLxx, Hel92, HB08, HMBS93, HHGS93, HK97, Isa93, JB90, JM89a, KKDO97, KKKP93, KBG^{+}13, KSTF94, Koc93, KA91, KO93b, Kra01a, KK96b, LS92a, LD90, LR92, LPLP97, Lie93, L^{+}93, Lim93, LEMS95, MS97, Mah94c, MM93b, Mas92, Mas95, MOOK94, MT91, MC10, MTHP93, MDH00, Mil97a, NU91, NSF90, Now93, PC94a, PYTL97, Par86, PS88, Psa92, RCK97, Ros93a, SES94, Sch19, SKVZ93, Sei94, SS94, SKSD94, SAGS93, SVD96, Sob93a]. Analysis [SA82, SA83, SB94d, SHB^{+}13, TK93, TD90, TP95, WR95, WGW04, Woo05, XOZ^{+}20, Yan94, ZX95, ZS93, Ade92, Ano93s, Ban88, Bli89, BE92, Bra91a, Bru88, Cha92a, Che88, CH89b, CV90, CD09, CGLY96, CGLxx, Chexx, CV88c, Che99, CH92a, Com98, DI88, Din92, Dra89, EGP88, FL92, Gal88b, Gal91, GCP90, Gor89, GTV91, GE96, HRC09, Hag90, HP91, HP92, JP90, JM89c, KCG08, KSB^{+}19, KSM^{+}19, KAMB19, Kwo87, LP94, LY88a, LY88b, LY88c, Li89, Lin89, LHPG21, MP88, Mar95, MK92a, MK92b, Mit88, Ng95, ODAZ15, Par90b, PP92a, PO88, QB92, Ram88, RGH17, Row86, SNS^{+}97, Sch90a, Scr88, SL88, Sob92, Tri85, Xu91, YH92, HS93b]. analysts [Ano94-121]. Analytic [Vog93]. Analytical [GGW93a, JWG93, JCJY94, SS96a, AP91, Ons88]. Analytics [PN13, Ano96n, BWHS18, CP13]. Analyze [Cla96, Ano95-39]. Analyzing [Che90e, CBLS13, ER96, FLP^{+}07, HAG^{+}13, SLS96, Che90d, CH92b, Gua88d]. and/or [Ano94-138]. Anemometry [L^{+}95]. Angle [BSJW96, CHL93, PB94a]. Anglo [Ano86a]. Angular [SB81, SB82a, FSY88, KMB09, SB82b]. Angulary [VMS93]. Angulary-Dependent [VMS93]. Anicol [Uth94]. Animal [Mis90]. Animated [Hug93, KC93c]. Animation [BS91, CMPR93, FD97, NT89]. Animations [FCGG90]. Anisotropic [AFML93]. Annealing [Abr92, Rig93, SB96, Bra88, Jab88, RCS21]. Annecy [Ham94]. Anniversary [SEA84, Str03, Ano93d, Ras91]. Announced [Coc02c, Coc02d, Ano87d, Ano94-86, Ano97b]. Announcement [Mon88]. Annual [Ano93a, BG91, Ann92, EP 97, FJSP95, Gra93c, IEE96a, Isk96, LM92, Min88, MP92, Supxxb, SR93b, USE00a, USE01, ACM95b, JC87a, Cha94a, Gra94, Hel93, JC87b, KWW92, NN87, San86, TfGERJxx]. Annular [KS93b]. ANSI [KK89c]. ansto [Aus93]. answer [Wal81]. ANSYS [PHVJ95]. ante [Ano94-83, Ano03a]. Antenna [MF92]. Antennas [FJSP95]. Anti [Dum97, MS94b]. Anti-dumping [Dum97]. Anti-Skid [MS94b]. Antibacterial [JBI91]. Anticipated [Glo84]. Antifuse [CP94a]. Antilles [HS94b]. Anto [CZRB93]. Anto-ignition [CZRB93]. Anupam [DRRM94]. ANURAG [She93, Nee94]. Anwendungen [Meu89a, Meu90, Meu91, Meu92b, Meu92c, Meu93, Meu95, Sch92a]. Any [FCD97]. Anymore [TF95]. AP1000 [IHIS91, SIDH95, SWJ95, SB96]. APE [IMP93, Sch95c]. APE-100 [IMP93]. APE100 [SPGD98]. APE100/ [SPGD98]. APEmille [Ano98g]. Aperture [YWD94, YWDxx, Gig94, YW94]. APEX [SS07]. APEX-Map [SS07]. APL [Ano94-49]. APOLLO2 [CTRR93]. App [PSO12, Ros09, Car93, Mil93, Noo95]. Appendages [HGC94]. appendix [Kue87]. APPI [Men87]. Apple [Cop93, Gar99]. Application [Ano93n, Ano94f, Ano94-74, AVS93, BMD^{+}20, BJ93, BSJ^{+}13, BP96, Bur94b, CDMW94, CT93b, DGO90, DMPR93, DDB^{+}10, ES96, EH97a, FMD07, Gon93, HTI93, IEE95a, IK93, IK91, JBWB97, Kad94, KTKK93, KDBG95, Kos89, LB94b, LMP^{+}90, LS93b, Lie93, MMK97, Nai94, Now93, PP93, Ram94, Rig93, SkLC^{+}03, Sho91, SLML93, SSBS99, Sug94, TGL96, WP94, YK94, Bur91, BMS92, Bli91, BM85, Bru88, Bru91, CS86b, Che90b, Fat10, FKL^{+}08, FGM^{+}03, Fra90, GP90, Gri86, Hab92, HS93a, KSM^{+}08, Kin96, LC91, Lim91b, LSS^{+}20, MLWC20, MLR90b, MLR90a, MKHY95, Mir88, Ng95, PSG03, Ver95, BBC^{+}99, GM93b]. Application-based [IK91]. Application-focused [BMD^{+}20]. Application-level [BSJ^{+}13]. Application-Specific [Ano94f]. Applications [Abr94, ASS94, ATL90, Ano88e, Ano88d, Ano88c, Ano90f, Ano93c, Ano93t, Ano94g, Ano94h, Ano94-38, Ano94-54, Ano94-103, Ano96e, Ano96f, Ano96i, Ano96-39, Ano97o, Ano97q, Ano97w, Ano97-28, Ano22, AJ93, Ara97, AZ94, Bar93a, Ber90b, Bha94, BP93, BBC^{+}05, CGFT05, Che94b, CFS95, Chr93, Cig97, Cla96, CM95, Edw97, FR98, FLP^{+}07, GS01, Gen97, GHWZ94, GGC^{+}11, Gun88, HL95, Him93, HK97, Hwa84, Hwa85, KGKa93, KHC14, KTG08, KMT94, KC93c, KLD95, KS90, KSW93, LK93, LB94a, LL08, Law00, LCP^{+}11, MKG90, Mil97a, MVS94, MBSW01, MK07, Natxxa, Nag94, NB93, Ope96, PN13, PT93, Res01, RL90a, RCR93, Sam91, SG81, SG82, SBZ^{+}08, SKC02, SZ96, SJR05, SK93b, TSCG94, Tho93a, TA94, TY96, Uni86b, Uni86a, Uni86c, VH93a]. Applications [VD94, VWC96, WAM^{+}01, Wei90, Wes89, Wil91, YSS94, Abe90, AB01, ABB^{+}13, Ano85b, Ano92y, Ano93i, Ano93-31, Ano94a, Ano94-75, Ano96a, Ano03a, Ara96, Bad08, BLW11, Ber82, Ber90c, BBC^{+}89, BPD06, CS82, Cal11a, Car89a, CBCJ92, CC88a, Che83, CKL^{+}13, CP92a, DJM94, De 96, Deg90, DSZ96, DT08, DM96c, D^{+}95, EKTB99, Ede92, Eig01, EMS11, Elm93, Emm84, Emm85, EWS^{+}13, GFBR10, Gan88, Gin93, Gra91, Gua88b, HG88, HKN89, IEE91, IEE96a, Joh92, JPTE94, Kon91b, Kow85, LAdS^{+}15, Lan92, LW11, IJS94, M.I87, MD04, Mar86, Mar88b, ML89, ME96, McC88, McN87, MDH^{+}16, MO88, Mil90, Mil91, ML95b, Uni91b, NBC92, Num85, OGO^{+}20, Por89, R^{+}00, Rol96, Rol97, Sim92a, Smi81, UL89, Ull84, WJC09, W^{+}12, Wie94, Wil10, WLN^{+}96a, WLN^{+}96b, Woo92]. applications [Woo94, WT13, YFY^{+}13, Zec93, tDv87, Ano94-79, BP89b, BP93, Hab89, Nat86f, WZ97, Ano00a]. Applied [Ano91c, Ano94v, ALMS92, Fie93, Ham94, HHGS93, OGOR97, RG94, WJ94, GL90, Kav92, LM92, Mil88a, PGK^{+}10]. Applying [Ano94-70, Fox90a, OMR93]. Appreciation [Pin99]. Approach [ABBB94, Ana94, Ano94-74, Ano96s, AM93c, App95, ACL93, AFT97, BS94c, BHLST94, BCCG97, Bos94b, CJ93, CCZ93, CH94, CP94b, Che92b, CSG99, Dic94, GM94a, HP03, HP93, JC94b, Jia94, JKNK93, KHSJ94, KV96, KDLS86, LR92, LC95, OH92, OD01, Opp95a, PCK93, Pas95, Pau09, Rei85, Rui91, Sch95b, Sob93b, TGV08, TY96, TM94b, AP91, Ano93u, Bis94b, BHS92, CCG^{+}17, CV89b, Che92c, EGK87b, Kuc87, LFJ^{+}20, McD90, PB98, RMM87, RM88, Rob87, SEV^{+}09, SA10a, SA10b, SB18, WF08, WD94]. Approaches [Bar93b, DDLV93, NS93, Rot94, SSM93, WABD97]. approaching [DH86a]. Approximate [Gur88, IJY^{+}14, PPP94, RT93]. Approximating [Phi85]. Approximation [Glo89, LM93, GS90, GS92a, ST90]. Approximations [BWGG94, CHL93, Cyb89a, Joh92]. April [CL91, Chi90, DP91, Elm95a, GH94a, GH94b, GH94c, Gig94, Hen97, IEE94a, IEE96c, IEE97b, KK89a, KSW93, Lid96, LCHS96, McC88, Sie94, VO93, Joh86c]. APS [GT94]. arbitrarily [LP94]. Arbitrary [DLPQ94, Lan94, Ara14]. arbitrary-rank [Ara14]. Arbitrating [SKIY94, SKIY97]. Arc [Nor97b]. ARC2D [BB91a]. Arch [Kel91]. archaeologist [Ano91t]. Architect [War03, War10]. Architectural [Bad99, Mir92, Nor84, SE92, KC95, Kwo87]. Architecture [ACM95b, Abr92, AU91, Ahm92, Ano94i, Ano94-127, AK94, BBH95, BA95, Bha94, CSG99, DHM^{+}88, DVWW05, EH97b, FB91b, GBG89, GS94d, HP03, HF94, HMNN91, HHOM91, HHOM92, Iwa90, JTX^{+}22, Joh97, KFB91, KRJ93, Kum94, MGA94, Meh94, MB12, M^{+}95, NB94, OGR95, Pel94, RL77, RL78, Rat87, RS93, Sah95, Smi81, Ste96, SC91b, TMHH95, VPDA93, Wat87, Wei89, YMY92, Yew88, KC93b, ZS94a, Abr90, AU90, Ano91f, Ano98g, Asa93a, BDM94, Bec90, Bhu95, Bur93, Bur94b, CS93a, Chu87, CRA10, Con00, CP93b, Cyr86, DRAB08, GBC^{+}05, GHS86, Haw86, Hog02, JS86, KHS88, Kha93, Kog91, KS88, KGLA85, KAMB19, LR89, MP88, MO88, MPSB87, PT92, PS88, Pol88a, Pop92, RR89, RGL^{+}15, Ros95, RCS21, Sca92, SK93a, Sch87b, Sho91, SMM88, SA83]. architecture [TS88, Van86, Vei85, VSM^{+}07a, VSM^{+}07b, Wat72, Sch88a]. Architecture-Independent [OGR95]. Architectures [And90b, Ano94-139, BIR94, Bar93a, BE93b, CPS96a, CO94, DXJM93, DT96, Ede94b, Gan94b, GVBC95, GG95, HS94a, HHK94, KTG08, LA94, MT86, MTH88, MS94c, Nar95, OH92, Rie93, Sah94b, SH90, SH93, SH94a, SG94b, VSM96, VKK80, VPGG01, WB85, YAG93, YAGxx, ZWP03, Abr88, AJFH86, BP86, BH95, CPS96b, Clo96, CP92c, DM96c, Don87, Erc88, FR95, GS89b, GMSS^{+}11, Gil94b, Hor90, Hor93, KK89b, KK90, Kon91b, Kra93, McA92, Nat88a, SKS04, SH94b, Sel95, Tan89a, VFK^{+}04, Vet12, McD88, Ano94p]. Architekturen [Meu89a, Meu90, Meu91, Meu92b, Meu92c, Meu93, Meu95]. Archivable [VV94]. archive [JR91]. Archiving [HM93c]. Ardent [LM90b]. Area [Ano92-47, BGH^{+}02, HNST93, PPP94, VW95, WCG94, YJD93, Ano95-27, Ano95w, GB91, Kah91, Kon87]. areas [Sha95b]. Arena [Ano94-72]. Aren’t [Coc02a, Coc02b, Ano95v]. Argonne [Ano85b, CKL^{+}13]. Argument [MS94a]. Argus [FGC06]. arising [Lou92]. Arithmetic [AFF93, Dun92, Gol91a, Gol91b, Wic92, BW88, LD90, Sch87d, Wai05]. Arithmetics [FGG09]. Arlington [HS^{+}91]. ARM [Els21, MSPPD20]. Arm-based [MSPPD20]. Armies [Bar00a, Bar00b]. Army [Ano90r, Ano95w]. Arnoldi [JV93]. ARPA [Gla93]. ARPS [DXJM93]. Arrangement [SAGS93]. Array [AW94, Ano94-41, Cha94b, GHK^{+}91, GMSB93, IGH95, Kun84, Li92, LLY92, Mas91, Mas92, Mas95, Meh94, MBSK92, MBK^{+}92, OM91, RP94, TCF94, Gok90b, Gok91, Gok92, GS94c, GHI95, GV92, Rob87, SLY89]. Arrays [Ano94-58, Chi95, LCH87, MM94b, VPDA93, WH93, WRW93, Mal88b, Row86, WB88]. arrested [Ano96-34]. Arsenide [Bac88, FB91b, Ano94-55, Dey95, Zho88]. Art [BBC^{+}05, KB19, OT07, Pay97, DDC96, Jet91, Jet92, LS87, LB94c]. ARTful [SFC^{+}21]. Artificial [COC93, MPH93, Opp95a, PD94, Pet97, Ram94, SC97, Tho93a, WWY93, CC88b, Has17, HD89]. Artist [Cox88, Pic92]. ARTS [BNSP99]. artwork [Lie90]. ary [DT96]. ASC [Wat72]. ASCE [Kho94]. Ascended [Ano92-43]. ASCI [BBH^{+}00, MSW96]. ASE [BP93, Ano88e]. ASE/93 [BP93]. Asia [IEE97b]. ASIC [PBK96]. Asilomar [Sin94a]. asking [DMW87]. ASME [L^{+}95, Suh97]. ASME/JSME [L^{+}95]. ASMO [NAAW97]. ASMO-II [NAAW97]. ASP [Pot88]. Aspartic [HGB90]. Aspects [HU93, Her89, HA93, LC93, Pli97, SHMR96, TP97, WS84d, tDv87, Rei88, SHMR94]. Aspiration [JC94d]. assembler [Kue87]. Assemblies [BMP93, CLPV93, GA97, OMR93]. Assembling [FC93, BPD06]. Assembly [HM93b, TSSK94, UEGM93, Vuj93]. Assembly/Disassembly [TSSK94]. Assertions [TG94]. Assess [MF97]. Assessment [AH93, Fat10, HS96, HSxx, OHHIB94, Sol93, TSSK94, DGG18]. assets [ZCPT00]. Assignment [CC94a, YKK96, Abr88]. Assimilation [DLLG98, HBDS93, Kau93a, WLH00]. assistance [CV91b]. Assisted [GMG94, SR94, SYMT92, Gal89a, LY90b]. Associate [Sul97]. Associated [GMW94, HM97]. Associative [BF92, HHT^{+}94, Kok94, Sch90b, WSP95, Pot87]. Asteroid [Ano96t]. Asteroid-Impact [Ano96t]. Astrid [GMS97a, GMS97b]. Astro [Sti98a, Sti98b]. Astrocomp [CBM^{+}05]. Astronautics [MSAD91]. astronomical [Nat85]. Astronomy [MN91]. Astrophysical [Ano94-59]. Astrophysics [Mal88a, MN91, Fra94, NN90, Vag88]. Asymmetric [Mil88b]. Asymptotic [LMM93, PH97, Scr88]. Asynchronous [MM93b, RW94a, SNS95, DuB90, DR91, JG99, VO93]. ATExpert [KW93]. Athens [HPP88]. Atlanta [Ano94-108, USE00a]. ATLAS [Uch86]. ATM [Ano93-40, Ano94-31, Ano94-35, Ano94-135, KG95]. ATM-Based [Ano94-31]. Atmosphere [CSRB90]. Atmospheric [ABCH97, DS94b, DLLG98, FA93, KB93, LS93b, TM88, Ano94k, HLDS95, HK93b, KH93, KK93, Kin96, Pan97]. atom [Ano95l]. Atomic [BB90, M^{+}94, IMF91]. Atomic-Scale [IMF91]. Atomique [Pre93b]. Atomistic [GIF^{+}12]. Atomistic-Continuum [GIF^{+}12]. Atos [Els21]. atria [KSM^{+}08]. Attack [Pau08]. attacking [Sha95b]. attain [And20]. Attaining [CMAS11]. Attempt [Ano92h]. Attenuating [BS97]. Attraction [GB96]. attractor [Ree88]. audit [Fin94, Aus93, MSCxx, Uni96]. Auditor [Aus93]. Auditor-General [Aus93]. Aufprall [Ano97d]. Aug [DW97]. augmented [LL88]. August [AB94, Ano92g, Ara96, Bro93, Bup87, Cha94a, CBCH93, C^{+}97, Cul95a, Dup86, Dup87, FL92, GT94, GP93c, HKR94, IEE96b, KK93, Lag89, L^{+}95, Met86a, ML95b, NN90, OMM93, Sha89, IEE94d, TC94, Uni87c, USE90, USE00b, VV95, VAS82]. Australia [Ano92g, KMG96, ME96]. Australian [IS95, Ill96]. Austria [ACM97]. authentication [Coc01]. Author [Ano90c]. Authority [Alaxx]. Auto [PRS94, Pol90, Pol88e]. Auto-Radiograms [PRS94]. auto-scheduling [Pol88e]. Automata [SS96a, Wag96]. Automated [And90a, Ano94l, Gal96, JTX^{+}22, Mil97a, WD93b, VM07]. Automatic [AK87, AH90, Ano89b, BEH^{+}94, BCR96, Cig97, CKM88, DMCK92, DCG93, DCGxx, FBZ92, HKP88, KRVJ94, Kar89, McK94, PE88, PRS94, Pol87d, Seh92, SSKa93, VNB93, Who92, XZC^{+}20, YZL^{+}20, Bab90, BMS92, Bli89, Blu92, Cre91, Eig91, Fea94, HP88a, HA90b, Pet89a]. Automating [KK96b]. Automation [Ano93i, Ano93-31, Ano94-75, Ano96a, Bos94a, Nag96b, Sha94a, Elm95a, Rol96]. automative [Mar86]. Automaton [ER96, LUT96, ZH88]. Automobile [HFT94, HCV97, OK93, TD90, YS94]. automobiles [Moi93]. Automotive [Ano88q, Ano94-75, Ano95j, Ano96a, AJ93, Bal93, BMCA93, BJH97, CT93a, Cig97, DHLS97, Del97, Div97, Fry97, Gib93, HMBS93, Him93, HNST93, KC93c, Mar86, Raa97, Rit97, Rol96, Tak94, VD94, VF93, Vro94, Ano93i, Ano93-31, ES88, Gal93, Gin93, Mar88b, Rol97]. Autonoma [C^{+}97]. autonomic [AZC13]. Autonomous [ALPP00, Die95, ESMH93, GIBGA93, HRG93, MGA94]. Autorenverzeichnis [Ano92c]. Autotasking [Dic90, EO91, LS92b, Nag90]. autotuning [BWV^{+}17]. Auvil [Ano22]. auxiliary [NNS^{+}90]. Available [CE18, KNYT95, Ano92-44, Ano94-86]. avalanche [PIH04]. Avenue [Nat86g]. Average [BBD^{+}08, YH90]. Avionics [Bar93a]. avoid [DMW87]. AVX [CKT21]. AVX-512 [CKT21]. awaiting [Gro92c]. Award [Ano22, Pin99, SC99, Ano93b, Str94, Ano97j, Ano98d, Ano11a, Coc02a, Coc02b, Coc02c, Coc02d, Pin01, War09, War10]. awarded [Ano95w]. Awarding [Ano88v]. Awards [Pin99, Ano87d, Ano94-86, Ano16, War03, Ano93b, Str94]. Aware [BBD^{+}08, CGFT05, CTD^{+}16, CLY^{+}19, HNS^{+}10, LLGS09, NPS^{+}20, TGV08, A^{+}12]. away [Uni92b, Uni92a]. awesome [Day12]. Axial [VH93b]. AXP [Was96b]. Aylesford [Ano94p].

B [McD88, AC84b]. B.C [CCKSS90]. B2 [FL92]. Babcock [Bab94]. Back [Sin08a, Ano89n, Ano95r, Cou13, Sch18]. Backed [CCZ93]. Background [CDA94, Hos95]. Backplane [RWNJ94]. Backpropagation [XCLW93, LW94, TCM95]. backs [Ano95x]. bacteria [Wal17]. bad [Moo08]. Baden [Sch94b]. Baden-Wurttemberg [Sch94b]. bag [WLLZ20]. bailout [Ano89i]. Baking [Jan96]. Balance [BOS97, GD97, War93a, BAAD^{+}97, CD09]. balanced [TRLD13]. Balancing [CRY94, CMF94, FT94, HL95, Kar94, KK95b, LH94, RS94c, Tri95b, Tri95c, VR94, YKB^{+}00, Cat92a, CG87, EB18, MD04, Tri95a, Tri95a]. Ball [DSB96]. Ballistics [Deh90, Chi90]. Balloting [Ano97b]. Band [BSL94, GZA86, WNKS96]. Banded [Li95, WN92, Bis94c, Con86, Con87a, Gan86, Sam85, TT93]. Bandwagon [Bel96]. Bandwidth [ET96, Hic18, PIH04]. bang [Gur94]. Bangalore [Bal94, IEE97a, Pra95]. banks [Use93]. Barcelona [ACM95a, DLM99, RMO96, Sig95]. Bari [De 96]. Barnes [Ano94-116]. Barrier [Gra93a, Ano97h, BP89a, BP90]. Barrier-breaking [Gra93a]. Base [JML95, RS94a, Sch94b, SK93b, Pit89]. Based [Ano94-74, AZ94, BCHH94, Bar00c, Bar00d, BCC^{+}09, BEK02, BHLST94, BS94d, BKM93, CRV94, CGW05, CP94a, Chr90, CM95, DA94, DS94a, DCWH07, DGJG93, FT96b, GHWZ94, Geu97, GCS94, HRG93, HMS^{+}86a, HMS^{+}86b, HCL94, HHOM91, HHOM92, IEE95a, Kar93, Koc93, LM93, LB93, MSGW94, MNB94, MTLL94, NB93, OIY91, OP96, PCK93, RRSS93, SEH98, SBJ90, SBW^{+}19, Sei94, STN93, SB94b, SSM93, Ste96, TAKB06, TSCG94, VVKB96, Was96b, AP90, AISS97, Ano90l, Ano91f, Ano94-100, Ano96h, Ano03a, AJFH86, BS00, Bar88, BWV^{+}17, BY21, Bis94b, BF92, Cal86a, Cal86b, Che90d, Che90e, CV92a, Chexx, Che92c, Che92b, CDS98, DM96c, DAC^{+}18, EB18, GP91b, GR91, HM93a, Her94, Hsi91, HGS91, Isk96, IK91, KP94, KS94a, Kue93, KNYT95, KAMB19, LG87, LY90b, LAL02]. based [LLSR02, Mal91, Mar90, McG87, MSPPD20, Mit88, MRSB94, PB98, PGK^{+}10, RWCA94, Ram88, RFS87, Rob87, Roj19, SKB^{+}20, SEH99a, SEH99b, Sch87a, SSLR90, SSS90, TS90, Tur79, TV89, WLLZ20, WG91, YFY^{+}13, Ano92h, Ano94-31, KGB^{+}96, TOY96, TP95, VWC96, Ano94-120, BBWR90, Sri94]. Basel [Ano96-43, LM92]. Baseline [SBW^{+}19]. Basic [CDW94, GW91, Joh91, LM90b, Wie94, Gal91, GJW91, SZG95, Dub87]. Basis [Ano94t, Ano01b, CCKSS90, FBA93, Nal94, SF93b]. Bastion [Pou94a, Pou94b]. Batch [GBF93, WMKS96, ZYL^{+}16]. Bath [BP93]. bathymetry [MMG^{+}18]. Battery [CSPJ97]. Battles [Gar99]. Battling [SHS15]. baut [Ano96-35]. Bay [DW97, PD94, App95, Uni96]. Bayer [ZBLZ95]. Baysal [Whe83]. BBM [PW23]. BBN [WGR93]. Be [Cze93, Sul97, Ano90n, Ano92-44, Ano97u, Bau96, LQFC18, Nag96a, Ruh95, Sch22, SDFP93, Vro94, Win02, EMS11]. Beach [L^{+}93, Lim93]. beamforming [GMF00]. beat [Gep00]. beats [Ada95b]. becomes [Hel93]. Beds [NCVG96]. beef [Ano96-34]. beer [Gep99]. before [Ano94k, Bro91b, Tec89, Uni92b, Uni86b, Uni89a, Uni86a, Uni86c, Uni92a]. begin [Ano97b]. beginning [Sha95b]. Begins [Pin01]. Behavior [DA97, FZM91, Gib01, LMY88, Lim91a, MS88, Mil90, Mil91]. Behavioral [Edw97, Gal88a, LY91b]. Behaviour [DDF93, Sch90c, VA94, Wie96, WQS92]. Behind [Mur97, Nor97a]. Beijing [Guo94]. Being [RDZ93, Ano88q, Ano92l, Ren97]. Belgium [DDC96, LCHS96]. Bell [Ano97b, Ano00a, KHHS95]. belong [Tho93b]. Benard [GW93b]. Bench [Gru97]. Benchmark [AGZ94a, AYL^{+}18, BBDS94, CLPV93, EGJ^{+}02, FY92, GGW93b, HLPP97, IK91, LMM85b, LMM85a, LMxx, MK07, SW88, SPS91, Van91a, Was96b, WGOY91, WOK^{+}00, WF94, Ano85a, Ano03a, BCL91, DMW87, GS06, Gib95, GREC91, KB18, LM90b, Oed92a, Oed92b, Eig91, Nai94]. Benchmarker [BBC^{+}00]. Benchmarking [BGQ19, DM96c, HHOM92, Jar12, LW94, MSPPD20, Mur91b, RK22, UT91, WL83, Wri19, WHMA97, XZC^{+}20, And11, DMW87, Eig01, GCP90, HL88a, Hoc91, Hoc96]. Benchmarks [Ano94m, Ano94-118, AHOK02, BE92, Blu92, CP94b, Cyb91b, DAF^{+}90, EK96, FBGM93, GGW93a, MNV93, Men84, SCG^{+}08, SSRL91, Ste94c, WOG94, Wor84, Ber89b, But92, Fat10, Hae91, VSH91, WT11, Yi11, CKPK90a, Cyb90, CKPK90b, Cyb91a, CBHS91, Rau91]. benefit [Ber82]. Benefits [Ano94-110, FG92, Ano00b, BFS11]. Benz [Kad94]. Beowulf [AV02, Ano98e, Bec01, Bro00, Bro01, DDJ98a, DWM^{+}01, FDD02, MCB^{+}01, MBR05, MDH00, NĆ02, OCVA01, Ote02, Spe00, SSBS99, Ste00, Ste01b, Ste01c, Ste01a, Ste02, UP01, VPGG01, WAM^{+}01, YKB^{+}00]. Beowulf-Class [NĆ02, Ste00]. berechnen [Ano97d]. Berechnung [Wat95]. Berkeley [Ano94n]. Berlin [Stu95]. Berrington [Ano00a]. Best [Ano94-118, Bas95b, SA10a, SA10b]. Better [And11, Ano93-43, MHE97, SKSD94, Str94, Ano08b]. Between [Bel93, Lu93, SH93, Tre97, WD93b, GL89, GE12, HS94a, MT91, RE94, RSRG95, SH94b]. Beware [Eij90b, Eij91]. Beyond [ABCE97, Ano94-110, CCKSS90, Fos96, Get15, Lee89, LCP^{+}11, Mil93, PN13, SF82, Sin18, Ano18, Bel92b, TG95]. beyonds [ARF12]. BFC [HP88b]. Bhabha [M^{+}94]. Bi [JML95, Cha92b, FZM91, Van91b]. Bi-Base [JML95]. Bi-CG [FZM91]. Bi-CGSTAB [Cha92b, Van91b]. Biasing [VNB93]. Bibliography [Ros93a, Lay91a, Mac92, Mac96c]. Bicyclic [JBI91]. bid [Dal96]. bidding [Ano92o]. Bidiagonal [LPNRJ94]. bidimensional [Mil87]. bids [Ano96j]. Biennial [ME96]. Bifurcation [BK95b]. Big [Ano87a, BvRS^{+}11, Dav92, GGN20, LQFC18, PN13, Ano97y, Gur94, SSP93, Str94, HAG^{+}13, MBM^{+}20]. Big-Time [Dav92]. biggest [Ano97-27, Sha95b]. Bijker [CCKSS90]. Bilinear [MDW93]. Bilinear-Discontinuous [MDW93]. Bill [Ano91s]. billing [CK92c]. Billion [Ano93-34, ARF12]. Bills [SW10a]. Binaries [Mil88b]. Binary [AFAGR96, OIY91, PG93, GE12, HM93a, SAB^{+}05]. Binding [Ano94-137]. bio [HR04, RD07]. bio-computing [HR04]. bio-molecules [RD07]. Bioattenuation [WWKR97]. Biochemical [ATL90, Kaz93]. Biocomputer [Wal17]. Bioelectric [FWWD95]. Biographies [Wei88]. BioinfoPortal [OGO^{+}20]. Bioinformatics [L^{+}93, Lim93, SJR05, OGO^{+}20]. Biological [CC88b, CV93, Cra96, FCGG90, MC10, NB94, STN93, Gre89a, OMM93]. Biologically [Lie93]. Biologists [Cra96]. Biology [AAB06, DLMW95, Fox90a, SGIS93, SR93b, AB03, Ano92l, Bad04, BA08, Gib01, MW88, SSS92]. Biomagnetic [FWWD95]. Biomechanical [FCGG90]. Biomechanics [HTV88, RM92]. Biomedical [Ano94-136, MKHY95, KG95, Ros89]. biomedicine [PH95, Pow97]. Biomembranes [SABK94]. biophysical [FMD07]. biosciences [CCC^{+}89]. Bipartitioning [Pel94]. Biped [KT93b]. bird [Ano97m]. Birdstrike [Sch90c]. Birth [ABHS89a, Coc02a, Coc02b, Rya90, ABHS89b]. birthdays [Rya90]. BIS [MMR96]. Bisection [VR94, CP92c]. Bisectional [GBG89]. Bit [AT93a, AT93b, SI90, SI91a, SI91b, Ano01c, GS93, Has84, Hir92c, KM89, Kra93, Mac90, TS91, Wei91, YFY^{+}13]. bit-oriented [TS91]. bit-parallel [Kra93]. Bit-serial [SI90, SI91a, SI91b, GS93]. bit-store [YFY^{+}13]. Bits [Ano94o, Ano02a, Ano02b, SHS15]. BKL [BBS94]. Black [WS93, BK89]. Black-Box [WS93]. Blade [Sch90c, SR93a]. BLAS [CDW94, DD90, DD99, GP93a, Wij89a]. BLAS3 [GJM86]. Blast [VTTS98]. blatant [Ano88y]. blend [Ano96-28]. blends [Ano96-28]. Blind [SR94]. Blitz [HN90]. Block [ASS94, Ano94-51, Bra89a, Cal86a, Cal86b, CDW94, Coc02c, Coc02d, DD93, HL96, HKMCS94, KNS95, NJL94, NGLPJ96, OB94, Rot94, Tsu91, VAGRMVA90, VB90, BS88b, CH98, Fra90, GS87a, GS88b, GP93a, GPS86, LY90c, Mac97a, Noo95, SZG95, Sch87d, SE98, WB88, JP90]. Block-Cyclic [HKMCS94, KNS95]. Block-oriented [Cal86a, Cal86b]. Block-Recursive [Ano94-51]. block-sparse [SZG95]. Block-Structured [Tsu91]. block-tridiagonal [SE98]. Blocked [DD99]. Blocking [AD88, NGDH96]. blocks [HOSZ97]. Blood [GIF^{+}12, CRA10]. Blooming [Sch92b]. Blowing [MFK94]. Blue [ABB^{+}13, BSJ^{+}13, BBK^{+}08, BCK13, CCD^{+}13, CP13, CEH^{+}12, CRA10, CKL^{+}13, CNC^{+}08, CHT^{+}13, DT08, DLJ^{+}08, EO13, EWS^{+}13, FKL^{+}08, HOF^{+}12, KHZ^{+}08, OWG^{+}13, PMS^{+}08, RIB^{+}13, SCG^{+}13b, War03, Bro17a, ABC^{+}05, AAC^{+}05, Ano96-30, Ano96-44, AUW08, ADG^{+}05, BGH^{+}05, BHD^{+}05, BJV^{+}16, CBB^{+}05, CSFS00, CBC^{+}05, EMS11, EFR^{+}05, Eva97, FGM^{+}03, GBC^{+}05, GS06, GZE^{+}05, GBB^{+}05, HBB^{+}05, HCH95, IBM01a, IBC^{+}11, IBP^{+}05, KBG^{+}13, KBVH14, KHV11, KB18, LKFU05, LM13, MSW^{+}05, Mor01, MAA^{+}05, MSA^{+}07, OBB^{+}05, RGL^{+}15, SWG06, SAB^{+}05, SPP^{+}05, Tan95, War10, War00, WAB^{+}05, ZYL^{+}16, IBM01b]. blue-glass [Bro17a]. BlueGene [ARF12, ABB^{+}03, CD09, KCM02a, KCM02b]. BlueGene/L [ABB^{+}03, KCM02a, KCM02b]. BlueGene/P [ARF12]. Blueprint [FK04, FK99]. Bn [RMH93]. Board [Sul97, Ano94-135, Ano97b, Lie90, TMHH95]. Bodies [HGC94]. Body [Ano92-38, Ano94-116, BJH97, BPUS94, EFIM91, HTI93, HF93, NAAW97, PYTL97, RCK97, Swe94, Ano93u, Ano94-59, Ano94-141, BAAD^{+}97, BAD01, SHMH97]. Boiling [SKAT93]. BOINC [GHdF10]. Boltzmann [MKND97, MF93, PMS94]. Boltzmann/finite [CRA10]. Boltzmann/Spencer [MF93]. Bomb [Tsy94]. Bombay [M^{+}94]. Bombs [San95]. Bombsight [CCKSS90]. Bone [HTV88, HHTD90]. Book [Ach99, ALPP00, Ano94p, Ano96c, Ano00a, Bra94, Bue86a, Dun99, Haw88, Kaz92, Kow86, MM94a, McD88, Nor97a, Pap97, Por86, Sch88a, Sim00, Tru88, Wen94, CCKSS90]. Books [CCKSS90]. Bookshelf [Wil96]. Boolean [CP94a, OIY91]. boom [Ano94s]. Boondocks [Dal95]. Boost [Ano03b, Coo95, Ano91g, Ano92-46, Mac97b]. boosts [Ano89h]. Bootstrapping [Law89]. borked [SHS15]. boss [Ano94s]. Boston [Ano88w, Ano92y, GL92, Lun94]. Bother [Mul96]. Bottleneck [Ano90e, Par90a, RG92]. Bottlenecks [SH94a, XOZ^{+}20]. Boulder [Bro93, Nat84]. Bound [Ano94y, LL94, MD04]. Boundary [Ano94g, GL88, Ji91, JKNK93, MF92, Rul93, TK93, Che94a, DL92, LG87, Mac92, Vez95]. Boundary-Element [MF92]. Bounds [Fos03, LZ95, MRL^{+}17, PB87]. Bowdoin [TC94]. Box [OK93, WS93, Way96, Bau96]. Boyer [HA90b]. BPPS [PRSS94, RKDM94, RRMD94]. BPPS-Linpack [RRMD94]. Brain [AKDM93, Ano91k, Bar00a, Bar00b, Bar01, BBL95, DLJ^{+}08, DP91, GIF^{+}12, HWP95, KVP95, Pas95, SVML95, SYMT92, Tay95b, Ade10, BR95, EFH^{+}00, Fur12, Has17]. Brain-scale [DLJ^{+}08]. Braking [FCD97]. Branch [Ano94y, JKL19, Lil88, MD04]. branch-and-bound [MD04]. branched [LP94]. Branches [EBS88]. Brazil [SN96]. Brazilian [OGO^{+}20]. Breaking [QD91, Gra93a]. Breakthrough [Ano97-29]. breast [Str94]. Brew [Gar99]. Bridge [WD93b]. Brief [FG93]. Briefs [Gar99, Gar01, Pau08, Pau09]. Bright [Lew94a, Lew94c]. bring [HBB^{+}05]. bring-up [HBB^{+}05]. Brings [Bor92, Ano91t, Bor93]. brink [Com92b]. Broad [Lew94b]. Broadband [Bür88, Clo96, NGPH99]. Broadcast [BP91a, Bue91a, EFH^{+}00, Kah97, Asi98, TS94, TM94a, Yan94]. Broadcasting [ABP92, WK95]. Broadcasting/forwarding [ABP92]. Broker [Sol94]. Brokerage [BJ93]. brother [SSP93]. Brought [Ano95-40, Bro17a]. Brownian [RRSG96]. Browser [Jab90, Ham90]. Brunswick [PEH93, TC94]. Brussels [Lid96, LCHS96]. brutal [Ano96u]. BSD [BE88]. BSP [GRRM99]. BT [Ano86a, WT11]. bubble [Ano95w, WQS92]. Bucharest [Ano93g]. budget [Ano94s, Ano95l, Gep99, LC12]. budgets [Ano94-27]. Buffer [TY96, CV91a]. Buffers [Ano94-30, Lee87b]. Build [Ano03b, Pal15, SSBS99, Fur12]. Building [Ano94q, Ano00b, Dip96, Fos93, Gok91, GHK^{+}91, IS20, KS94a, Per04, Per06, AV02, Ano94n, AUW08, CGLY96, CGLxx, Chexx, Faz87, Gan88, MP92, SNEP14, Kos95]. builds [Per87]. Built [OBR94, SDFP93, DK01, Wal17]. Built-in-Self-Test [OBR94]. Bulk [SNS95]. bulletin [Nat86c, CSRxx]. Bunching [Nag96b]. burden [Ano89i]. Buried [BS97]. Burkhardt [Zor93a]. Burst [TV89, GTV91]. Bursty [YYW^{+}20]. Bus [Sah94b, SDK10]. Business [Ano95-42, Cor89a, Don92b, Joh94, Ano97y, Hel93, Nat92a, Ano85b]. Bussed [Fid90, Fid91]. butterflies [Bue91b]. Butterfly [WGR93]. buys [Ano96-29]. BWR [RMPW93]. bye [Str12]. Bytes [Ano94o].

C [Ano96c, Kho94, KSW93, Nor03, YGSB94, Ano94p, AC91, Ara14, Car88, CS93a, CC94b, Cra91, Ede94b, Gis86, GS93, GM93a, GS94c, GR91, Gua87a, Gua87b, Gua88a, HC99, Lee96, Mac90, Mac91a, TFVK94, WSL88, WL96]. C-1 [WSL88]. C-90 [Lee96]. C. [Dro95]. c/o [Ano94p]. C0* [TMP94]. C2 [Jon89]. C2G2 [KS94b]. C3I [Coo95]. C4 [Cha94b]. C90 [MSTK93, Oed92a, Oed92b, RM96, SWSR97]. C90-T3D [SWSR97]. C916 [Gra93a]. C98 [GMS97a, GMS97b]. CA [IEE95c, AIA93, Ara96, CG96, Ece96, KK89a, USE01]. Cable [RCK97, And10, LP94]. Cache [Ano94u, Ano94r, Ano94-43, Ano94-139, Che93b, Che92b, DMCK92, GAV95, GM94b, KG96, KB96, LS94, Li95, MTLL94, NB93, OA94, SM94, TNIA92, TA94, Yan93, Bre87, CV91a, CV88a, Che89c, CV89a, CV89b, Che92c, CDS98, LYL87a, LMY88, LY90b, LY90c, LY91a, Mar88a, Por89, SKS04, WFJ^{+}17, WHL93, BJ95]. cache-based [CDS98]. Cache-Coherent [KB96, TA94]. Cached [GS94e, HS93c]. Caches [GS94b, TY96, CV92a, CV88b, CKM88, Kar89, Lee87b]. Caching [GAV95, Hic18, MM93b]. CAD [Ano96d, Fie86, GY93b, GM93b, KTKK93, Lie90, RSB94, YKY90, YSL97]. CAD/CAE [GY93b]. CAD/CAM [Ano96d, GM93b, Lie90, YSL97]. CAD/CAM-Application [GM93b]. CADSOL [GRSS93]. CAE [GY93b, HTI93]. CAL [Kel85]. Calandria [DB94]. calcolo [LP90]. Calculate [SHG95]. calculated [Ano95w]. Calculating [Mar90, WRW93]. Calculation [CLP93, EBS88, SB81, SB82a, SAGS93, WFT93, FSY88, Gri86, KNHN16, MB97, SB82b]. Calculational [BMP93, TKI93]. Calculations [Ada93, CLPV93, CPR93, CNGR90, DD87, DMKW93, FSGS93, HS94d, HFH86, HFH87, INKN01, IMA93, KGKa93, LB82, MKND97, Mon93, NM93, Now93, PNK93, SG81, SG82, SR93a, VM87, Vuj93, WD93a, Ano94-122, Ano94-123, For93, MR87, NPS93, PIH04, RGL^{+}15, SKB89]. Calibrating [SNEP14]. Calibration [KT94]. Calif [B^{+}95]. California [ACM95c, ACM03, AU87, Ano93n, Bel86, Gra94, IEE95d, Uni87a, SR93b, Clo96, GE96, JD95, Sin94a, IEE94d]. Call [Ano95b, Ano96e, Ano96f, Ano98b, Ano98c, Ano09, KL99, Lat16]. Caltech [Jet91, Ano90n, Fox89, Jet92]. CAM [Ano96d, Lie90, YSL97]. CAM-Application [GM93b]. Cambridge [Ano94-107, Ano96c, LLR93b]. Camera [SVML95, Ano08b]. campus [BBBC96, Sca92, Jet92]. Can [Has17, PD94, Pic91b, Pop91, SDFP93, Spe97, Ano97u, Bab94, DMW87, HHS01b, MAFW08, Sha95b, Sug80a]. Canada [Ano88t, Ano88u, BG91, Goo97, VAS82]. Canadian [BG91]. cancer [Ano97-29, Law89, OMA^{+}96, Str94]. Cancun [Sie94]. CANS [WWTE92]. Cap [MAT85]. Capabilities [Ano90k, Bur93, WAD^{+}89a, Nat88b, Wil88b, WAD^{+}89b]. Capability [EDJ^{+}10, Chi20, Mac97b, Sha95b, WZB86]. Capable [Ano97s]. capacity [Bla97, CSFS00, Hsi91]. capillarity [Gre88a]. Capri [Sch97a]. captain [Per87]. Car [Ano92h, LB94b, Nag96b, NW97, Rit97, BBC^{+}99, HG88, SD88, BBK^{+}08]. Carbon [CSRB90]. Carburettor [SS94]. Career [CCKSS90]. CARL [Bec90]. Carlo [ALM93, Ano87e, AHAM93, Ask93, Bak93, BL93, BPJ94, BJLW95, BBS94, BLFT84, Bro96, Cha84, DKS93, Dec90, Din93, FBA93, Gri88, Gup88, HEJM95, HAAS93, IK91, KY93, MZ95, MNR86, MMRL93, MNV93, MS94c, MBN93, NM93, PB88, Rie93, SF93a, Sol84, TW92, Uen93, VNB93, YFOT93]. Carlo/Sn [FBA93]. Carnegie [Ano88n]. Carnegie-Mellon [Ano88n]. Carolina [LC90, Nor89, Nor93b, Nor93a, L^{+}95, VO93]. Carolyn [Ano96c]. Carried [RDZ93]. carries [Ano89i]. Carrying [BPUS94, Min86]. Cars [AC93, Ano96u, Str94]. Cartridge [SCV01]. CASA [Mes93b]. Cascade [FI91]. Case [Ano94-104, DBK09, cF03, Joh97, JML95, KNS97, SSG93, Bis94a, Bla97, DI88, Gal88a, Gan86, HS93a, JY92, Mal91, NW03, PGK^{+}10, RR89, RGL^{+}15, Sie90, WvTB^{+}07, Wij89a, ZH88]. cash [Ano92-27]. CASMO [KE93]. CASMO-4 [KE93]. Castine [Wuo94]. Casting [Ano92v]. Cat [Ade10]. Cat-brain [Ade10]. catalog [Natxxf, Nor93a]. Cathedral [Bel86]. Catone [Vag88]. Causes [YYW^{+}20]. Cavanaugh [IEE93b]. Cavitation [KKDO97]. CBI [Nor03]. CBN [TF97]. CBN-comb [TF97]. CC [GB96]. CC-COMA [GB96]. CCA [AKM^{+}06]. CCM2 [SL93]. CCN [SQS^{+}19]. CD [BGIM90, BIRB93]. CDC [Bro17a]. cDNA [AGD93]. cDNAs [AKDM93]. CEA [PPR95]. CEA/CEL [PPR95]. CEA/CEL-V [PPR95]. Cedar [DGG93, Ano91a, Ber86b, Ber86a, Ber89a, Dav86a, DGG92b, Gal88b, GB90, ASK85, Ano94-105, Ano94-113, BB91a, DKLS86, De 91a, De 91b, DGG92a, Eig90a, Eig90b, Fra90, FGM90, Gal91, GJW91, Gan86, Guz86, Guz87, Guz88, KDLS86, Kuc87, LM90a, Lav89, Mal86a, Mal86b, ME87, ME91, PJ90, War89, Yew88]. CEL-V [PPR95]. Celebrates [Str03]. Celerity [Ano88f]. Cell [ASSW93, Ano92h, FG87, Man90, RCR93, WMMC10, Fuj11, Gib01, LLDF95, Ano02a, Ano02b, EMS11]. Cellar [Cia88d, Cia88e, Cia88f, Cia88a]. Cells [Gib01, Hug93]. Cellular [DGJG93, ER96, Nag96b, SS96a, Wag96, CD08, CRA10, ZH88]. Census [BF91, Cre91]. Center [Fin94, AU87, Ano87a, Ano88w, Ano93-46, Ano94-79, Ano94-107, Ano94-129, Ano94-135, Ano22, Cra96, DRB^{+}20, MSCxx, Gro90, Hab89, IEE90, IEE93a, IEE94b, Kah97, KK89a, Lay91b, Mic90, Uni87a, Nat86f, Nor89, Nor93b, Nor93a, Norxx, Pro94, Pit88, PBK91, SN89, Sim97, Uni92e, WZ97, WL94, Ano94k, Ano95v, Ano96-37, And89, AW93, Ano97f, Ano97r, Bac88, BPM^{+}89, BBW90, Cul95b, Deu86, Joh86a, KNS97, Lay91a, LC90, Mir90, Nas91, San91, Sim92a, Tay94, Wes96, Joh86b]. centered [HS96, HSxx]. Centers [Ano88r, Ano95-35, Ano95-45, Ano95-49, Ano96t, Ano97n, Ano97x, Ano97z, BvRS^{+}11, Bra91b, Dau96, Egg94, FG92, Hir94, Lew94a, Lew94c, LZF16, TGV08, Web91, Ano94-86, Ano95t, Ano97y, Han03, IEE92, Nat92a, Pou88, Red91, San90, Mar85a, Mar85b]. Central [BK95b, SA10a]. Centre [Ano92-46, Ano94-33, Bha94, Coo95, Hos95, IEE94a, KSW93, Man89a, M^{+}94, Ano95-43, Cou90, Att96, Man92]. centres [Mac97c]. Century [Bor01, Bel92a, Joh97]. Ceramic [SKK^{+}90]. ceramics [Ano95w]. CERBERUS [MM93a]. Ceremony [Pin99]. CERN [VV95, Ano94s, Ano95v]. Cernobbio [DJM94]. certain [Ked92, Rob89]. Ceruzzi [CCKSS90]. Cetraro [D^{+}95]. CFD [PEH93, Ano94-32, Ano94-54, Cig97, Ece96, EK96, FR95, FK93, Gen94, HB93, Him93, Lan93, LM92, Sch97a, Sim92a, Uth94, WvTB^{+}07]. CFD-Applications [Cig97]. CFTP [KK89c]. CG [Bra89b, FZM91, SZG95]. CG-algorithm [SZG95]. CGCM [TfGERJxx]. CGSTAB [Cha92b, Van91b]. Chain [BM93b]. Chaining [Che89b, TYZ88]. chairman [Ano93b]. Challenge [BEH^{+}94, Ste94b, Tho93a, Ano92l, Hoc94, New95, Rya13, New91].

Challenges [Ano97-28, Bos94a, BBC^{+}05, Con91, Eck92b, Fry97, Gen92, Get15, Gin93, HM97, Rob93, Sha95a, Sha95b, UEGM93, Con90, DSZ96, Eck92a, Gro92c, KCG08, SR10, SMDL90, Woo92, Woo94]. Chamber [BD93b]. Champagne [Gep99]. Champaign [Ano22]. champion [Ano96z, KW92]. Champions [Lew94b]. Chandy [RM88]. Chandy-Misra [RM88]. Change [SGH97, Woo96b, YOY97, EP 97]. Changes [Dav92, RSRG95, DDJ98b, Sul97, Ano88y]. Changing [Fry97, Gar01, Ano93b, Pan93]. Channel [Ano94-104, GFM96, MS96, RE94, Sco96, XB96, Gre88b, Wic96]. Channels [KB96, Pan96]. Chant [Ano94-87]. Chaos [DP91, Kha91, Sto95]. Chaotic [WM91]. Chapter [Ano95c, DLPQ94, Dic94, Gen94, GD94b, HVZ94, HGC94, HFT94, KMT94, LC94, Riz94, SHZK94, Sha94b, Soe94, Tak94]. Characteristics [Ano94-37, LTD^{+}93, PCK93, Sus93, DAC^{+}18, EWS^{+}13, LC91, Lim91b, LF03, PIH04, WT11]. Characterization [Ano94-38, Cal88, Lim91a, BSJ^{+}13, Gal88a, JY92, LLDF95, PBK91, RGL^{+}15, Vaj91, WMK90]. characterizations [BCL91]. Characterizing [Che93a, Gle91, ST92, UU94, XOZ^{+}20, Sch94a]. Charge [BSB93, Ano96o, Ulr12]. Charging [BK97, Div97]. Charity [Ano96-34]. Charles [CBCH93, Nor97a, Bab94, Lew96a]. Cheap [Bar00a, Bar00b]. cheaper [HHS01b]. check [Bin88]. checking [McT96]. Checkpoint [BIB^{+}18]. Checkpointer [PL94]. Checkpointing [CCR11, MVS94]. Chem [GAB^{+}96]. Chemical [Ano89r, DS94b, DAF^{+}90, Maa93a, Sta94, WBP87, CHWW13, CKT21, C^{+}97, Heh86, JT87, MCH91, WCHK91, ZMDS96]. Chemically [Ano94b, WABD97]. Chemicals [GD97]. Chemistry [ATL90, Ano96i, FJSP95, Fox90a, SG81, SG82, War93b, Zey91, ARE95, BS00, Bup87, C^{+}97, CCC^{+}89, Dup86, Dup87, GAB^{+}96, Har90, Har91, JT87, Kin96, NRN00, R^{+}00, TF15]. chemists [Ano90l, Ano02a, Ano02b]. Chen [Ano95p]. Chesapeake [App95]. Chess [HN90, Tan95, Ano96z]. Chicago [Cul95a, Ano96u]. Chief [Ano94-39, Ano94-40, YVC89, Ano96-34, Bru90a, Pin01, Sul97]. China [And20, Cal11b, Cal11a, Cal12, Coc02a, Coc02b, LQFC18, Moo11, ZLC21]. Chinese [QD91]. Chip [Bar01, CMAS11, HOF^{+}12, HHOM91, Pau05, Pau08, Ros93b, Ano89h, Ano91h, Ano91l, Ano94-122, Ano94-123, Ano95w, BHD^{+}05, HBB^{+}05, IBP^{+}05, Joe87, KFN02, Moo06, OBB^{+}05, Pou94b, Ros95, IBM13a, WAB^{+}05, Hay89, She90, Tra89]. Chips [Ano93o, DM88a, DM88b, For02, MD88]. choices [Ano92-28]. choke [SF82]. Cholesky [Ano94-65, Ano94-115, Con86, EHS94, GHNL87, KESH94, KESH95, Noo95, Rot94]. Chooses [Ano97k]. Chromodynamics [AGLL98]. Chromosome [FC93, Rig93]. Chronology [CCKSS90]. Chuck [Ano11a]. Church [Mul08]. Ciarcia [Cia88a]. CIM [KWW92]. Circle [Sul97, Mer86]. Circuit [BAT99, Cia88d, Cia88e, Cia88f, GI93, GS94e, LWY^{+}20, RLC91, SO95, Cha92a, Cia88a, Hun90, HGS91, Hun91, Lie90, Sal89, Xia88, YKY90, Yan90c, Yan90b, YH90]. Circuit-switched [GS94e]. Circuits [BS94c, PDR91, RL90b, SJA94, Wuo94]. Circulation [AD97, CSRB90, DGG92b, DGG93, De 91a, De 91b, DGG92a]. Circulative [Che91]. CISC [SE92]. CISS [HKR94]. City [ANS92, Ano95-38, DW97, Fra94, Isk96, Uni96]. civil [Kho94]. Claim [Ewa96]. Claims [Bar00c, Bar00d]. Clamping [KTK94]. Clara [Ano93n, KK89a]. Class [DT96, EBS88, Fer86, KN88, NJL94, NĆ02, Ste00, Ano96r, Arn88, AI92, ARW93b, JOK^{+}18, Lou90a, MKfDA96, Pop92, MRS88]. Classes [JML95, Che90a]. classical [Gup88, KA92]. Classification [Kon93, MSGW94, MH96, RPY94, SM94, WWY93, Cre91, Wes89]. Classifier [GCS94]. Classifying [ML95a]. classroom [DS89]. Cleaner [LYKM97]. cleans [Ano97t]. cleanups [Ren97]. Clear [Ano90e, Ano94-48, Par90a, Ano96u]. CLF [Her94]. Client [Ana94, Hic18]. Client-side [Hic18]. Client/Server [Ana94]. Climate [ABBB94, Che96, Dic90, DFWW93, FT94, Gue90, Hun94, Per93b, SBW^{+}19, WMBC97, Ano95w, DT08, Mec95, Per93a, SD88, Str94]. climbing [SL88]. clinical [JD95, KSM^{+}08]. Clinton [Ano93b]. Clock [CWLT97, KS95]. Clone [Rig93]. Clones [LD93a]. Closed [BD93b, GI93]. closely [Ano91h]. closet [Gei16]. Cloth [FDD02]. Cloud [BY21, BCW20, CHWW13, dRSGS16, DXJM93, FSC18, IJY^{+}14, KB19, KHC14, KCZJ14, MPT12, MT13, PN13, Ros09, Rya13, W^{+}12, WOS09, ZLRC20, AZC13, BWHS18, CBKA09, CBLS13, DLS93, EB18, LC12, MBM^{+}20, MMG^{+}18, SKB^{+}20, Som13, TF15, WLLZ20, XZC^{+}20, YFY^{+}13, Kar13, MDH^{+}16]. Cloud/Mesoscale [DXJM93]. clouds [Ros08]. Cloudscheduler [BCW20]. Cloudy [SS09]. clout [Tri95a, Tri95b]. Club [Ber89b, VSH91]. Clubs [Ano92z]. Cluster [Ano94-78, Ano94-138, Ano99, Ano01a, Bae01, BB99, BIRB93, BSD^{+}20, CKS99, DD05, DD02, DXJM93, FBJ94a, FDD02, HMS93, Kop00, KR94c, MCB^{+}01, MDH00, PBDM93, Sch19, Ste01b, Ste01c, Ste01a, US01, WAM^{+}01, YKB^{+}00, AV02, Ano02a, Ano02b, ABMN02, DWM^{+}01, Fat10, Gan86, GB90, IS20, KG95, LM90a, Liu95, MBR05, SNS^{+}97, STH^{+}98, Ste02, SJR05, VSM^{+}07b, TMP94]. Cluster-C0* [TMP94]. cluster-supercomputing [Ano02a, Ano02b]. clustered [HRC09]. Clustering [DDJ98a, Lum01, YZL^{+}20, KESH95, Sch88b, ZEC^{+}17]. Clusters [ABGL96, Chi00, DSSS05, GBF93, KMKD97, Kra01a, LC97a, NĆ02, OCVA01, PBM95, Spe00, SSBS99, Ste00, UP01, Ano93v, Ano93-41, BGKR99, EKTB99, GKSR14, Hol93, SA10b, SD92]. Clutch [RCK97]. CM [But92, CS93a, Hel92, HT93, HP95, KC95, KR94d, Lee96, LW94, Mar95, McB93, PTC^{+}93, PW94, Ric90b, Ric91a, Ric91b, SNS^{+}97, SI90, SI91a, SI91b, Ste92, VDK91]. CM-2 [But92, Hel92, Ric90b, Ric91a, Ric91b, SI90, SI91a, SI91b, VDK91]. CM-200 [McB93]. CM-5 [CS93a, HT93, HP95, KC95, KR94d, Lee96, LW94, Mar95, McB93, PTC^{+}93, PW94, SNS^{+}97, Ste92]. CMOS [BS94c, Lan94, TOY96]. CMOS-Based [TOY96]. CMS [FSC18]. CMU [Ras91]. CNC [KTK94]. CNN [Ros93b, RCZ93, Ros95]. CNS [Asa93a]. CNS-1 [Asa93a]. CNSF [Cor87]. Co [BCK13, GL93b, SES94, YMZ90, CCG^{+}17, IBM13c, vL99]. Co-design [BCK13, IBM13c]. co-designed [CCG^{+}17]. co-operating [vL99]. Co-operative [GL93b]. Co-rotating [YMZ90]. Co-Scheduling [SES94]. CO2 [KfGERJxx, TfGERJxx]. coalescing [Pol87b]. Coarse [CWW94, IMA93, Man91, EAMS95a, EAMS95b]. Coarse-grain [CWW94]. Coarse-Mesh [IMA93, EAMS95a, EAMS95b]. Coastal [DA97]. Coatings [Rit97]. COBOL [SPK94]. COBSQL [SPK94]. Cockpit [Kun95]. Code [ASSW93, AJ93, AGLL98, AK93, CS90, CTRR93, ER94, Gon93, GMS97a, GMS97b, GD94b, KLD95, KS94b, MJRS94, MNV93, MAA93b, MBN93, MWO95, QD91, RCR93, RDZ93, SBHW80, Swe94, VNB93, Vuj93, WW92, WFT93, WF94, YFOT93, YG92, Ano90o, BWV^{+}17, BAAD^{+}97, BAD01, BV96, BHS92, CLmWH91, DS96b, EAMS95a, EAMS95b, EY91, FR95, GAW96a, GAW96b, HJZ94, KK89c, Mir88, PLS20, PO88, SNK^{+}93, SHMH97, NB92]. Code-Breaking [QD91]. coder [Win02]. Codes [BCR96, Cal81, DMKW93, DR81, DR82, GL93a, KO93a, Kon93, MKB87, PG93, SKVZ93, Smi93, VWC96, GSZ91, LS92b, LLDF95, Mer86, Par90c, RCB03]. Codevelop [Ano87a, Ano87c]. Coding [MKRI93, SLML93, Use93]. Coefficient [Che92a, Gie96]. coefficients [HA91]. Coevolution [Die95]. Cognition [MH95]. Coherence [Ano94u, Ano94-43, Che92b, DMCK92, KG96, LS94, MTLL94, OA94, SM94, CV91b, CV92b, CV88a, CV88c, Che89c, CV89b, Che92c, LY90b, LY91a]. Coherent [KB96, TA94]. Coincident [ADGA95, Mit88]. Coincidizing [CS91]. Cold [Ano97t]. collaborate [Ano97j]. Collaborated [PCY^{+}19, PZS^{+}20]. Collaborating [Cra96]. Collaboration [WCZ^{+}18, JG99]. Collaborations [Cop93]. Collaborative [ABSS94, PC94b, LPC^{+}95]. Collapse [Gie96, CK90]. Collection [ALPP00, Ano94-46, AP87a, Ano85a, IEE89a]. Collective [Ano94q, Ano94-124, BGPS94, TM94b, CC96]. College [TC94, AV02]. Colleges [Mur06]. Collegiate [Coc02c, Coc02d]. colliding [GH90, Gre90a, GH91]. Collision [MBN93, OMR93, Ste94d, VMS93, Ano00a, CKT21, Gre89b]. collisionless [SHMH97]. Colloquium [Ano97g, tDv87]. Cologne [ACM91]. Colombella [Lag89]. Color [Ano89c, Bar00c, Bar00d, FG87, KFF93a, KFF93b]. Colorado [Bro93, McC88, Nat84]. Colossal [Lin83]. Colour [Ano90d]. Colt [WN10]. Columbia [ACC^{+}96]. column [WQS92]. COMA [GB96, XB96]. comb [TF97]. Combat [Har94a]. Combination [FCD97, JCJY94, Gri92, UPK87]. Combinational [BS94c]. Combined [RPY94, Ano97-29]. Combining [LY91a, MH95, PR94b, Sch96, Lee86]. Combustion [AGH^{+}90, BSB93, GWG93, HK97, Maa93a, MA97, MJRS94, War93b, BCW93, CYXL18]. Combustion-Flow [MJRS94]. come [Ano96u, Pol90]. Comes [Fox89, Ano89l, Ano91v]. Comet [Ste94d]. comfort [WvTB^{+}07]. Comfortable [Bra93]. coming [KW92]. Command [Mor92c]. commemorative [Ras91]. Commentary [Ano96g]. Comments [Gup88, Str94]. commerce [AS99]. Commercial [Ano92-44, DM93, Don92a, Joh97, KP95, Kun95, NCDS97, Lop89, TF15]. commercialization [AG90, Asl91a]. commercializes [Ano95v]. Commercially [KNYT95]. commerical [Ano92-44]. Commissariat [Pre93b]. Committee [Ano88p, Bro91b, CKS99, Kho94, Uni92b, Uni86b, Uni89a, Uni86a, Uni86c, Uni92a, Nat88b, Uni98]. COMMIX [Asl91b, WFT93]. COMMIX-2 [WFT93]. Commodity [Ber95b, Dav00, BB99, DK01, WTC^{+}02]. Common [Ano93-38, KHMD94, Mah94b, Ano85b]. Communicating [LS92a]. Communication [Ano94q, Ano94t, Ano94-44, Ano94-117, BGPS94, BBD^{+}08, BEK02, BC95, BCR96, BD94, CP96, DDHK94, EVM^{+}98, EHG95, EH97a, FGKT97, GFB^{+}03, GSB95, GBG89, GBK^{+}96, GB92, HPPF94, HNS^{+}10, IK82, Jay88a, KHSJ94, KP96, LB94a, Mac91a, MKSF96, NSP94, RWNJ94, RW94a, SNS95, SC20, TZ94, TM94b, VSB94, Abr88, AGZ94b, ABMN02, Cal96, CD09, DDT95, FTT97, Hae91, Hoc94, HEB96, KC89, KG03, NRM^{+}09, Oed92a, PGK^{+}10, Pol86, Pol89, Suz89, TTD^{+}11, WTC^{+}02, YQTV12]. Communication-Aware [BBD^{+}08, HNS^{+}10]. communication-based [PGK^{+}10]. Communication-Efficient [KHSJ94]. Communications [Ano88p, Ano94-60, Ano94-124, BS92, FBJ^{+}94b, Her89, JM93, Mic90, MLY10, SNS95, SDK98, Ano96p, Ano96-31, BBW19, CC96, GB91, LEY86, NGPH99, Ste94f]. communicators [Ano97g]. Community [BCC^{+}08, DFWW93, Mur06, Pel93a, PMP21, Ano99, CKS99, Win02]. Como [DJM94]. Compact [Ano95j, Ano95-28, PW23]. Compacting [Ano94-30]. Companies [Vro94]. Company [Pau08, Ano95x, Ano95-37]. Comparable [Ano92h]. Comparative [Bie88, DG95, HC93, MTHP93, SL88, TIOK94, Feo92]. Compared [Tho93d, HL88a]. Comparing [LY90a, LKJ03, VY88]. Comparison [ADLL01, AABK95, BIR94, BCHH94, BHW98, Bot96, CHL93, CV90, DH86b, EJ97, HS93b, HLJT93, JWG93, JW98, KN88, KJ85, KV96, LMM85b, LMM85a, MSAD91, Smi93, SSM93, VW95, Wag96, BB13, But92, CP25, CKD^{+}19, HA90b, Jor87, KC95, KB88, KBVH14, Lee96, Lop89, Mar95, MP91d, SWS^{+}12, SWL^{+}91, SWL^{+}92, TT93, Tur90]. Comparisons [Ma99, Pap92, SM94, Tem83]. Compatibility [Ohr86]. Compatible [CWLT97]. Compcon [Bel86, IEE95c]. Compensated [Sch92b]. Compete [Ano88g, Ano97e]. Competition [GE12, Kah94, Pin01, Ano96z, Dum97, Gra94]. competitiveness [Ano93b, Asp93, Els89]. competitors [Ano94-121]. Competiveness [Bro91b]. Compilation [Fah94, FXAC94, FY96, HKMCS94, HLB94, LS93c, NPS^{+}20, SLRP95, TAAL95, WS90, GC92, WS87a, WS87b, WS87c]. Compile [Ano94-112, CH92a, Gor89, LS92a, Pol89, MP90, YH92]. Compile-Time [LS92a, Ano94-112, CH92a, Pol89, MP90, YH92]. Compiler [Ano94u, Ano94-49, Ano94-66, BGS94, BWV^{+}17, CCSR92, CWW94, Che89c, CV89a, Che92b, CV95, EJL90, Gis86, GMG94, GGV90, HKT92, Hus86b, LB96, Li91, Li95, MP87a, MP87b, MTLL94, OB95, PW86b, PW86a, Pol88a, PE95, Rue92, SNS95, SIDH95, SM94, SLRP95, TYZ90, TBC94, Tsu91, Vei85, ACK^{+}95, BMS92, BGS82, Car88, CS93a, Che92c, Eig90a, Eig90b, EB91, KN86, LY90b, PW86c, Pol87b, Rob87, RCZ93, TMT^{+}20]. compiler-assisted [LY90b]. Compiler-based [BWV^{+}17, Che92b, Che92c]. Compiler-Directed [Ano94u, CV95, SM94, Che89c, CV89a, GGV90]. compiler-driven [KN86]. Compilers [TP95, UZ95, Ano87d, BE92, Chi20, Gua88d, Hag90, HP91, HP92, Leu90, PHK88, PP92b, Pol88e, Sch90a, SLY90]. Compiling [BCR96, Dra88, Har86, MPC89, OH92, OSKO95, Van94]. complaint [Mac96a]. Complementarity [FJSP95]. Complementary [DDLV93, AT89]. Complete [KS94b, RG94, SJPS94, SJPS96, TS94, Car89b]. completes [Ano02a, Ano02b]. Completion [LL94]. Complex [BGQ19, CH94, DL96, DDLV93, EKZ90, FSGS93, Fie93, FGG09, KO90, KO93b, L^{+}93, Lim93, ZS93, AB01, Ano99, Bar88, BMD^{+}20, CKS99, DSZ96, GJG88, Gib01, Heg96, SSLR90, Sug80a, Wai05]. complex-arithmetic [Wai05]. Complexes [CS94b]. complexities [GS89b]. Complexity [BWGG94, GIBGA93, Kon93, PPP94, RBK95, Amm90, Amm92]. compliant [MNY09]. Complicated [KMT94]. Component [BEK02, Ber07, CJHH94, Dav00, TK93, Wil94, AKM^{+}06, Mil88a]. Component-Based [BEK02]. Components [BMCA93, Gui96, Mis90, Raa97, SL99, DK01, HPS88]. composite [YB86]. Composition [BLO94]. Compositional [Bri90, Kon93]. Compositions [KT93a]. compound [Hus86b, HKP88]. Compounds [BCW93, Kaz93]. Comprehensive [LR92, AV02, ALPP00, CDO90]. Compressed [PB94b]. Compressible [HS94d, GB22, Ram86, TR86]. Compression [BB93, SC93, NRM^{+}09]. computable [FRW92]. Computation [ALPP00, Ano93-34, AT93a, AT93b, CRY94, CPR93, GS94a, HGC94, HLB94, HP88b, JW98, LB94a, MT86, MTH88, MFK94, NdMM09, PPM90, Saa93b, SJDV09, SRBL94, WKHS97, YKK96, AJFH86, Chi81, DRAB08, EGK89a, FKL^{+}08, Fuj11, GS88a, Jab88, Jay88a, Joh91, Kow85, LP94, RJ13, RD07, TS91, Tho93b, Kow86, Wen94, Bue86a]. Computational [AAB06, ALPP00, Ano94v, Ano94-34, Ano94-71, Bad04, BS01, CDO90, C^{+}97, CP93b, DDF93, Duk91, FD93, FJSP95, FK98, GI93, Got91a, HW97, HMBS93, HK93a, Hos95, HM97, KO93b, Kra01b, LC93, LLR93b, Lay91a, LCV90b, LCV90a, ME96, Mil17, MBW01, NSH95, NS93, NRN00, NBC92, OHIB93, OHHIB94, OGR95, PPP94, Pli97, Por86, Pro01, R^{+}00, Ric90b, Ric91a, San91, SHMR96, Sim92b, SGIS93, TP97, VS99, Web93, WvTB^{+}07, WABD97, Wil90a, WS84d, A^{+}02, AB03, BA08, BPM^{+}89, BB91b, Buc83, BPD06, COS89, DGL89, DGH20, Ece96, Fox97, FW90, GCS20, Hab92, Han03, HHS01b, HP95, HKS93, LM92, Lou92, MSK^{+}02, MA85, PSG03, PEH93, RCB03, SHMR94, SMDL90, Str94, TF15, Vez95, WHBH93, Wil91, Woo92, Woo94, Dra94a, L^{+}95, NBC92]. computationally [MKfDA96]. computationally-efficient [MKfDA96]. Computations [Ano94-32, Ano94-140, Ano95-41, ACL93, CC94b, CDC^{+}87, Duk91, DP91, Fox89, FY96, GML90, JKNK93, KB94, Lin83, MR95, MRL^{+}17, NRS95, NGLPJ96, Riz94, SkLC^{+}03, Ste90, TFO94, WR97, WG91, XL94, ZW02, B^{+}89, CC88a, GPS90, GPS86, Ji91, KK89a, Kra90, SW88, SN95a, SN95b, SC04, Van95a, Woo92, Woo94, YF98, van95b]. Compute [HOF^{+}12, MDH^{+}16, Sch19, IBM13a, BHD^{+}05, Gui10, HBB^{+}05, OBB^{+}05]. Computed [Kuw94]. Computer [ACM89b, ACM95b, Ahm92, AW93, Ano85b, Ano91c, Ano91b, Ano93l, Ano94w, Ano94-59, Ano94-89, Ano94-111, Ano95c, Ano95-34, Ara97, Bal94, Bel86, BP84, Bru88, CCZ93, Car94a, CCKSS90, CG86, CS90, Cra96, CSG99, DGO90, Dec90, DRRM94, DMW87, Dun92, DS94c, EFIM91, Edw97, EGJ^{+}02, FCGG90, Fie93, FR81, FNK93, Geu97, Gil93, Gis86, Gol91a, Gol91b, Gre91a, Hab86, Hal87, HHTD90, Hay86, Hel92, HGB90, HP03, HH93, ITOK93, IHIS91, Jab93, KC93a, KD93, KKKP93, KMNT96, Kho94, KHC14, KGS93, Kol81, Kra97, KB97, KZ94, LB82, MK93, Maj94, MA97, MM90, MB12, Moh94, Nag94, Nal94, NW97, Ohr86, PT93, Pin99, PH95, Pow97, PC93, Rie93, Rus78, SEH98, SW10a, Sch88a, Sch95a, SKVZ93, SBHW80, SB94a, SC99, SBN82, SS96c, SR94]. Computer [SC91b, Tan95, TVT^{+}16, Van93, WOS09, Wic92, WF93, WCG94, XMR92, YFOT93, Zho88, AGEL13, AGZ94b, Ano85a, Ano89d, Ano90n, Ano91l, Ano93a, Ano93g, Ano94-82, AUW08, Asp93, Bas95b, Bhu95, BS91, Cha94a, CH89b, CDS98, CK92c, De 96, EAMS95a, EAMS95b, EY91, GS89d, GREC91, Hen91, Hoc96, Hog02, Hsi91, IEE89b, IU87, Jon03, KHS88, KMB09, KTN^{+}14, LC12, McA92, MSW91, OYK^{+}14, Par90b, Pic92, PMS94, Ras91, Rei88, RD94, Ros08, SEH99a, SEH99b, SW91, SKB89, Str94, Sug80a, Van86, WHL93, Yew88, Pin99, Mah94a, Mut94, Pin01, Pou94b, PH95]. Computer-Aided [KC93a, KD93, MM90]. Computerised [RS94a]. Computerized [OCVG93]. Computers [ADLL01, AFF93, Ano90u, Ano94-46, Ano94-58, Ano94-126, Ano97a, ACL93, AFT97, Att96, BGMR96, Ber82, CCKSS90, Com98, DDHK94, DJSP93, Dem91, DKH86, EAMEG11, EBS88, EK96, FGG09, FHM99, Gen94, GL93a, GS94d, Hag01, HQ91, HK93a, HML^{+}21, JC94c, Kau93b, KH98, Kop00, LPV94, Man91, Mes97a, Mes97b, MS93, MBSW01, Rei85, Res01, RK22, Sah94a, SPM^{+}10, SF82, Sin94a, SSOH95, Sug80a, TF92, TL96, Tho96a, Tho96b, VB90, WMBC97, Who92, Yan93, ZM94, Afu90, AS88, AP91, Ano97-27, BCC^{+}05, BB92, BGKR99, DL92, Dic81, Dic82, Don85, DS86a, Don86, Don93a, Ece96, Eig92, GMF00, Hae91, Hak89, Kan15, KK82, Khe94, Khe95, KA92, Kog91, Kra92, KC92, Lan92, Lee90, Mik89, NP90, PEH93, Pol87a, Pol89, Qui87, Sch89b, Sch87c]. computers [Sch87d, Sha89, Sha95b, SPS91, VdSK^{+}05, Voe89a, VVH95, WLCG02, Way96, Zor93b, Zor93c, dRC94, dC94, tDv87, ALPP00, Sim00, Tru88]. Computers/Software [Ano97a]. Computervision [WG94]. Computerworld [Ano97j]. ComputeServer [BBWR90]. Computing [AIA93, APK^{+}12, AS98, Ana94, Ano88g, Ano90g, Ano93r, Ano94-31, Ano94-35, Ano94-33, Ano94-60, Ano94-70, Ano94-71, Ano94-103, Ano94-110, Ano95-34, Ano96t, Ano97i, Ano97k, Ano08a, Ano09, Ara97, BGS94, Bad99, BKK11, Bae01, B^{+}95, Bai97, Bak10, BCC^{+}08, BGS^{+}12, BCC^{+}09, BCW20, BS98a, BS94a, Ber82, BEK02, Ber07, BGM^{+}11, Bha94, BS92, BBHL01, BJS02, BEH^{+}94, BIB^{+}18, BCH^{+}22, BH17, BEGGK07, BGH^{+}02, BNSP99, CLB19, CGFT05, CC94a, CCYT05, CH94, CH10, CDPW94, CFS95, CK90, Chr93, Coc02a, Coc02b, CMAS11, CBHS91, Cze16, DD05, DCWH07, dRSGS16, Din91, DT97, DFF^{+}02, DSSS05, DMT^{+}21, Ede94a, EBS02, EAGEG09, EGEAH^{+}08, EDJ^{+}10, Els02, EHG01, FSC18, For02, FJSD96, Fos96, FGKT97, Fos03, Fox89, FBJ^{+}94b, FLP^{+}07, FS93b, Gar01, GSG^{+}94, GS01, GBF93, Gen97, GL90]. Computing [GCY^{+}08, Gre94, GT94, Han90, HC99, HHS01a, HB08, Hof94, HG02, HP04, Hol94, Hol95, HNS^{+}10, HHK19, HML^{+}21, IEE93b, IEE94c, IJY^{+}14, IHE^{+}00, IH94, IBBA20, IJM14, Jar12, JM93, Joh97, JPMG08, Jon96, Kah94, KMKD97, Kau93a, KB19, KHC14, KTG08, KWB^{+}10, KT11, KLQ19, KRS13, KCZJ14, KBLD08, LL08, LM08, Lat16, Lew93, Lew17, LLGS09, LMM^{+}21, Lid99, LCP^{+}11, LB94c, Lum01, Man92, MPT12, MM93a, MB20, ML97, MB12, MC10, Mes17, MKRI93, Mil97b, MLY10, Moh94, MBD99, MS94d, MRGR12, Mur06, Mur07, Nat91a, OLWW94, OD01, OPR01, OHIB93, OT07, PN13, Pap16, PH11, Pel93b, PW05a, Pin01, PMP21, PSO12, Pro01, RS93, Ros09, Sah94b, Sak02, SBZ^{+}08, Sch94b, SKC02, SBW^{+}19, SN96, SCSL12, SZ11, SL99, S^{+}93, Smi96b, SP12, SW10b]. Computing [SDK98, SMS95, Ste00, Ste01b, Ste01c, Ste01a, SS09, IEE94d, Str10, Str03, SMDS15, SLS96, SHB^{+}13, TGV08, TAKB06, TPJ^{+}19, Van13, VV95, Voi94, WNKS96, WKL^{+}16, WN10, WP94, Wil96, WG93b, WGW04, Woo05, Wri19, XZC^{+}20, YJD93, YMY92, ZLRC20, ZWP03, Zim96, ACM89b, Aba09, Abe91, AGEL13, AB01, A^{+}12, AISS97, AZC13, Ano89d, The90a, Ano91x, Hig92, Ano93c, Ano93d, Ano93-39, Ano95u, Ano95-38, Ano96a, Ano96-43, Ano97-30, Ano99, Ano03a, AUW08, AB96, Ara14, AKM^{+}06, ABMN02, Bad04, Bad08, BB99, BBC^{+}99, BG02, BY21, BS94b, BBB^{+}20a, BBM19, BBM96, Bor93, Bro17b, BPD06, CHWW13, CP25, CBKA09, CS91, Car92, CWD^{+}08, Cat92b, CCG^{+}17, CF12, COT21, CB99, CKS99, CDG^{+}06, CBM^{+}05, CBLS13, Cyb91a, DDC96, Dam11, DHA^{+}13, DF12, Don87]. computing [D^{+}95, DvdS12, DRSS99, Duf00, EB18, EKTB99, EHF^{+}97, FEK20, cFM07, Fly66, FK98, FK99, FK04, Fox98, FP00, GFBR10, GH94a, GH94b, GH94c, GCS20, GMSS^{+}11, GHdF10, GKSR14, GM93a, GYL00, GLH94, GAB^{+}96, HW11, HS94b, Han94, HPPF94, HBKR96, Haw86, Hen97, HS95b, HS95c, HR04, IEE89b, IEE96b, IEE97b, IEE97a, Ipe19, JPTE94, Kar13, KG98, KDP^{+}14, KFML20, KSM^{+}08, KHBB01, KBM^{+}02, KCG08, KG03, KFN02, Kos95, KW11, KBD10, LAdS^{+}15, LW11, LY88c, Li89, Lid96, LCHS96, LC12, Liu12, LAL02, LG03, LM13, LLSR02, IJS94, MAFW08, MNY09, M^{+}09, Mag10, MD04, MBM^{+}20, MM91a, Mar96, MI01, May01, Mec95, MDH^{+}16, MUKX06, MT13, MMG^{+}18, MMG^{+}00, Mye92b, NSW08, NSR90, NRM^{+}09, OGO^{+}20, Pit87, PHK88, PT92, PLC^{+}19, Pei17, PSS^{+}19, PGK^{+}10, PW05b, PW23]. computing [PMC22, Poz13, PMS94, RAG11, Rol96, Rya13, SCV01, SCG13a, SKB^{+}20, SSS92, SKS04, SFC^{+}21, STH^{+}98, Sch87c, SEV^{+}09, SD92, Sha95b, SQS^{+}19, Som13, SW99, Ste02, Ste94f, SDMS99, SS07, SJR05, TTD^{+}11, TF15, TS90, TAM^{+}91, Tru88, Uni93, Vag88, Vet12, WJC09, W^{+}12, WFJ^{+}17, War03, WLLZ20, WF08, Wil10, YFY^{+}13, Zen99, Zor92a, Bra94, CCKSS90, GMSS^{+}11, GL90, Kah97, Kar13, Kho94, KRS13, Rep92, Pap97, Ano96c]. COMS [CWLT97]. concentrate [Mac97c]. Concept [FT93a, Pet97, Tur86, Ano96u]. Concepts [EJ97, Hel96, MPSB87, Ruh95]. Conceptual [ESMH93, GY93b]. concerto [Ano93y]. Conchology [Ill96]. Concord [Lee94]. Concurrency [NMS93, EBB^{+}20, McC94, McG87, SMR10, UPK87]. Concurrent [AK95, Ano87a, Ano87c, Ara91, BBC^{+}89, EGH^{+}06, Lie20, Ste94a, Swe94, YMY92, Bra88, Dav86b, HP88a, LA93, Mes93a, Rat87, Sha90, Wij89a]. concurrent/vector [Wij89a]. Condition [FCD97, KSTB94, NG92]. Conditional [EBS88]. Conditions [BBL95, JKNK93, Rul93, GP91c, NG92, Per87]. Conductance [KW95]. conductivity [CNC^{+}08]. Conference [ACM88, ACM90, ACM91, ACM92b, ACM92a, ACM93, ACM94, ACM95a, ACM95c, ACM96, ACM97, ACMxx, ACM03, AIA93, AB94, Ano88w, Ano90f, Ano90g, Ano93n, Ano94-107, Ano97k, Ano97q, Ano98b, Ano98c, Ano01a, B^{+}95, Bel86, BP93, BH92b, CL91, CBCH93, CG96, C^{+}97, DDJ98a, DP91, EP 97, Ece96, FJSP95, GG^{+}97a, GH94a, GH94b, GH94c, GL90, GT94, GP93c, HKR94, Ham94, HS95b, HS95c, Hol94, HPP88, HBCN95, IEE85, IEE93a, IEE93c, IEE94b, IEE94c, IEE95b, IEE96d, IEE97b, Kah92, KK87, KK88, KK89a, KWW92, KSW93, L^{+}93, Lim93, L^{+}95, Mar86, Mar88b, MW88, McC88, NBC92, Pel93a, PEH93, PL91c, Pit90, PH95, Sch97a, Sha89, Sig89, Sig90a, Sig95, S^{+}93, Sin94a, SC93, SR93b, Tho93c, Uni87c, USE00a, USE01, Wuo94, Zyg93, AU87, AGP96, Ano92y, Ano93a, Ano93g, Ano93i]. conference [Ano93-39, Ano94-108, Ano94-134, Ano96k, Ano96l, Ano97c, Ara96, Asp93, Bal94, BBM96, BP89b, Bro93, Cha94a, DDC96, De 96, DSZ96, EM94b, Goo97, Gra93c, Gra94, HS^{+}91, Hel93, IEE96a, IEE97a, Isk96, JPTE94, KK85, KFF93b, Lid96, LCHS96, LM92, Lun94, ME96, Met86a, Pow97, RD94, RMO96, SN96, SJD96, Suh97, ACM94, ACM95a, IEE95b, Ill96, Sin94a]. Conferences [Ano93f, Ano94x, Dra94a, Dra94b, Dra96b, Dra96a, Edw97, FJSP95, Qui95, Ano94a]. conferencing [DR98]. Confidence [Moh94]. Configurable [TSCG94]. Configuration [BMSD94, Els89, SKB89, Abr88, GJW91]. Configuration-space [SKB89]. Conflict [GL93b, VLA92, Cal88, Fuj99]. Conflict-Free [VLA92, Fuj99]. Conflicts [DL96]. Conformal [CT93b]. Conformations [Che93a]. confuse [DMW87]. Congestion [Sug96]. Congress [FL92, IEE93a, Kho94, KSW93, Tec89, Uni92b, Uni86b, Uni89a, Uni86a, Uni86c, Uni92a, Uni98, VAS82, DJM94, Uni92c]. Congressi [GT94]. Congruences [Mas92]. Congruential [BB92]. conjecture [Wun89]. Conjugate [Ano94-45, Che90b, FR96b, FR96c, Gre90c, HFH86, HFH87, JM89a, KJ85, LPV94, Man91, Meu87, MT97, Nat86h, Sea86, VAGRMVA90, And88, Bau88, Gib95, HVY91, JM89c, MS88, ME91, Meu89b, SZ89, Yan92]. Conjugate-Gradients [FR96b, FR96c]. conjunction [Ano94-126, Kho94]. Connect [NBKP95a, HHS01b, NBKP95b]. connected [Bue91b]. Connecticut [Ano90l]. Connection [HT93, Pin99, Pin01, DFSZ88, Ste92, VDK91, Ano94-99, Hil91, Hil92, LTD^{+}93, LB94c, Mar95, SL93, Sha94b, WZ87, Wun89]. Connectionist [ABMW93, Asa93a, Asa93b]. Connectivity [Fri95, GS87c]. Conquer [DT96, AT89, Don93c, LR88b]. CONSENSUS [BSKJ93]. Conserved [HC93]. conserving [SG92d]. consider [SCV01]. Consideration [SNS95, ES88, Lil91]. Considerations [Chr90, Lan92, LYL87a]. Consistency [GM94b, Str97]. Consistent [PL94, RMH93]. Consolidation [Smi96b]. consortium [Ano90n, Mes93a, BCH^{+}22]. Constant [SS94, Bue86b]. Constellations [DSSS05]. constrained [Ram88]. Constraint [AP90, SB94b, YSL97]. Constraint-Based [SB94b]. Constraints [Cla96, DDF93, FD97, WN92, ZYL^{+}16, HY91, yHY92]. Construct [SNS95]. Constructing [CDPW94, LB93, SA94]. Construction [AFAGR96, AM93b, CCKSS90, SSM93, CKS99]. Constructive [SB94b]. Constructs [Zey91, Gal89b]. consuming [GBS18]. Consumption [NW97, SB18]. Contact [FD93]. containers [MGS^{+}20]. Containing [FSGS93, WNKS96]. Contaminant [AD97, Ewi97, Ver97, YCC97, Chi86]. Contaminants [ZL97]. Conte [RD94]. contemporary [Asp93, Vet12]. Content [Rig93]. Contention [CP96, PDR94]. Contest [Coc02a, Coc02b, DDJ98a, BBD92]. Context [GT97, JC94b, KR94a]. Context-Free [JC94b]. Contigs [SSKa93]. contiguous [Ano94-84, WMKS96]. continues [CKS99]. Continuing [Ede94a]. Continuity [Poe95]. Continuous [DNV93, GGW93b, IHK93, Uen93]. Continuum [GIF^{+}12, KY93, Woo96a, KG98]. contract [CKS99]. contracts [Dam11]. Contribution [Cox88, HB93, Pin99, Gal93, RLKW93]. Contributions [WG93a]. Control [AJ97, Ano93g, Ano94y, BWO96, BK97, DNV93, DHLS97, EHG01, GHWZ94, HRG93, HBB^{+}05, HHGS93, HED93, Hug93, KISY94, KWW92, MGA94, MPH93, MS94b, Ost94, PH11, RE94, SBJ90, SR93a, Sta94, WSP95, YOY97, YK94, ZS94a, Amm89, Amm90, Amm92, CV89b, DCG90, DGL89, GP91b, GP91c, HC91, Ji91, Joh88, McC94, MP91b, Pol90, SFL^{+}94, Uni98, HT72]. control-flow [Amm90, Amm92]. Controlled [FT93a, IHK93, VT95, KFN02]. Controller [HU93, DuB90, DR91, Kon87]. Controlling [Ano94-52, GCS94, KB97, LMD98, Sto95]. Controls [Ram94, Woo96a, Har95]. Controversial [Gar01]. Convection [GWH93, NU91, WG93a, Ha88, Ha90a, NSB96]. Convections [GW93b]. Convective [CS90]. Convention [Ano88w, IEE94a, IEE94b, KK89a]. Convergence [FR96a, FZM91, WR97]. converging [MKfDA96]. Conversion [Sha94a, SMH91]. Convertible [Raw97]. CONVEX [BMS92, Ano88g, Cha94b, Jon89, NBGS96, WSL88]. Convexity [EH97c]. Convey [Bak10]. Convolution [Meh94, MB97]. Convolutional [ZFF^{+}18]. Conway [Pev93]. Coolants [VM94]. Cooled [Car94a, CSB89]. cooler [Ano96u]. Cooling [Cha94b, CMAS11, Lam14, TTM^{+}20]. cooperation [Str94, YQTV12]. Cooperative [RWCA94, RLW93, Ano93b]. Coordinate [TK93]. coordinates [Hun92].

Coordination [Kho94, BMD^{+}20]. coordinator [Sch90b]. coping [Hil97]. Copper [McC88]. Coprocessor [AT93a, AT93b]. Coprocessors [Mar90]. Copy [OA94, Has17]. Copyright [Waz89]. CORAL [Han20, IS20, MLWC20, PLS20, RMM20]. CORBA [LLSR02, CSFS00]. CORBA-based [LLSR02]. Core [BCR96, FBGM93, HM93c, JKNK93, JR94, MTK93, MAA93b, NM93, PP93, PNK93, RMPW93, RRSS93, Smi93, TBC94, WD93b, YZL^{+}20, BC95, CLF^{+}19, KNHN16, Ano98d]. Core-Reflector [JKNK93]. Cores [IHSK93, TKI93, Ano00b, HCD^{+}18]. Cori [DAC^{+}18, HCD^{+}18]. Cornell [Cor89b, Ano98f, Ano02a, Ano02b, Bro91d, BK91b, Cor89b, CR89, Lee89]. Corner [War93a]. cornerstones [Cat92b]. cornerturn [Hol90a]. Corporate [Rei93, AS93]. Corporation [Hab86, Was96b, CCKSS90]. Corrected [JBWB97]. correcting/detecting [AC84a, AC84b]. Correction [DS94a, OLLG96, Tze88]. Corrections [Ano95-34]. correlated [Shi95]. Correlates [BBL95]. Correlation [Bel93, FBB97, WAM^{+}01, KMB09]. Corrigenda [RL78]. Corrigendum [Gol91a]. Corrosion [VM94]. Corrupted [MRL^{+}17]. Corruption [LMM^{+}21]. Cortex [BPL95, KDBG95, RSRG95]. Cortical [ADGA95, HHTD90, HV95, Sal95, VT95]. Cosmic [Nor96, Rud90, BP84]. COSMIC/NASTRAN [BP84]. Cosmological [DS96a]. Cost [GSG^{+}94, GGBR95, HW96, KV96, Mah94c, SW10a, AP91, Ano94-82, Ano95-28, Ano97-29, Poi90, Rob87, Sam85, Sma95, TF15]. Cost-Comparison [KV96]. cost-effective [Poi90, Rob87, Sam85]. Costs [FG92, SNS95]. Could [Sch12, Ano90e, Gib01, Kra97, Par90a, Poo96a]. Council [Kho94]. Counting [Cla96, AL92b, SM89].

Country [GV96a, KSB^{+}19]. country-scale [KSB^{+}19]. counts [Sit78]. Coupled [ADG^{+}08, CCSR92, CSRB90, GIF^{+}12, Ano91h, BDM94, Har86, KDK89, PW23, SA90, vL99]. Coupling [ADGA95, DDF93, GD97, RAES96, RSRG95, War93b]. Course [Jia94, Nar95, WF08]. Cover [Poo96b, Van97, Ano87d, Ano17, SL90, Tru88]. covering [Joh86c]. covers [Way96]. COVID [BCH^{+}22, Kra20, VSB^{+}21]. COVID-19 [BCH^{+}22, Kra20, VSB^{+}21]. Cox [Law89]. cozy [Ano97j]. CP [Asi98]. CP-PACS [Asi98]. CPAR [CC94b]. CPROF [SSS90]. CPROP [LB93]. CPU [Ano01c, EBB^{+}20, GB22, Hoc85, PCY^{+}19, WCZ^{+}18]. CPU/GPU [EBB^{+}20]. CPU/MIC [WCZ^{+}18]. CPUs [Ano91h, Nag88]. Cracks [LDMC96, Ano97h]. Cracow [BBM96]. CRAFT [SZG95]. Crandall [Wen94]. Cranking [Div97]. Crankshafts [GA97, TF97]. Crash [AKT90, Ano97m, Jab93, Ano88q, Bin88, Bru88, HG88, HPS88]. Crashes [AVS93]. crashing [SSP93]. Crashworthiness [LCVR93]. CRAY [Hoc85, Ano89e, Ano91d, Ano93h, Ano94-27, Ano95c, Ano95d, Ano98d, CB99, DBK09, KH87, KG95, LH87, Nor97a, PBK91, RR89, War10, AFT96, ARW93a, ACK^{+}95, AGLL98, ALN^{+}01, BAAD^{+}97, Ber90a, BH92a, Car91, Cha84, C^{+}97, DGO90, DD90, Dec90, DH91a, Dic90, DE84, Duf82, Ent99, EM78, FR96b, FR96c, Fra94, GG^{+}97a, GRRM99, HL88a, HKN89, KLN90a, LG97, Law90, LMP^{+}90, LMM86, Ma99, PSG03, RL90b, Rus78, Sar91, Sch96, SZG95, SB82b, SPS90, SO91, TM88, Tem88, Tem89a, VSH91, VDK91, Vez95, VY88, WLH00, WW92, Wil90a, WKHS97, Ale90, AHP97, AGK^{+}87, AKG87, And90a, AFF93, AF97, Ano88g, Ano88l, Ano89b, Ano89l, Ano89n, Ano92-29, Ano93s, Ano94z, Ano94-83, Ano94-119, Ano95j, Ano95e, Ano95g, Ano95i, Ano95h, Ano95p, Ano95s].

Cray [Ano95f, Ano96-28, Ano96-29, Ano96-42, Ano11a, Ano16, ARE95, ABGL96, AZ95, AZ99, ABHS89b, ABHS89a, Bac88, Bai88, BK77, BOS93, BAT99, BMT96, Bea90, BJ95, Bel98, Ber96, Bie88, BS04, BB13, BÇG14, Bow88, BL91, BBC^{+}00, Bro17a, Bro97, BP84, BWHS18, BBW19, BBB^{+}20b, But92, Cra92, Cal81, Cal85a, Cal85b, Cal86a, Cal86b, Cal88, CC96, Car93, Car94a, CM84, CM86, CDH84, CWLT97, Che89b, CS84, CS86a, Chi20, Cho90a, CK90, CRM94, CS93b, CS95, Cla18, CCSM97, CCSS98, CKD^{+}19, CP96, Cra96, Dan91, DCG90, Dao88, DO89, DD87, DP90, DH91b, Dic81, Dic82, Din92, DAC^{+}18, DL90, DH86b, DH86a, DR93, Dow98, DR81, DR82, DVWW05, DAKM98, EE93, EJL90, EAMS95a, EAMS95b, EBS88, Erw84, EO91, ET96, Ess90, EK96, Ewa89, Ewa96, EY91]. Cray [FG87, Fat10, FR81, Fin82, Fon85, FSY88, Gib95, Gin82, Gis86, GP93a, GYL00, Gra93a, GMS97a, GMS97b, GKL^{+}87, GS89d, GZA86, Gur88, HPH^{+}20, Hak89, Hal87, HS94c, Has84, Hay89, HCD^{+}18, HVY91, Hic18, Ho91, Hol90b, HE98, HES93, HKSY94, HSKY95, HP95, HEB96, HPLT01, HFH86, HFH87, Hug94, Hus86a, HN90, JKL19, Jen87, KN88, KC95, KBVH14, KS93c, Kha93, KJ85, KK89c, KLN90b, KFW94, Kol81, KMRR20, Kra88, KM85, KSM^{+}19, KZ94, KB18, KAMB19, LH86a, LH86b, LH86c, LH86d, LR90, Lar84, Law89, Lea09, Lee96, LS92b, LS93a, LW94, LB82, LMM85b, LMM85a, LKJ03, LSK04, Mac90, Mac91a, Mac96a, MW82, MLR90b, MLR90a, Mar95, Mar92, MWRK18, MGS^{+}20, McT96, MPG96, Meu87, MH18, Mil93, MW81, MRS88, Mir88, MKB87, Mon88, MSW91, MSTK93]. Cray [MT97, MWO95, Mur97, Nat89b, NSH95, Nag88, Nag90, NR86, Noo95, NSP94, OL86, Oed92a, Oed92b, Ohr86, OLLG96, OD88, Pap92, PS94a, Par94, Par90c, PPR95, PK80, PCM84, Pet83, Pin99, Pin01, Pin91, PBK96, PO88, PHVJ95, QB92, RM96, Rei85, Rei88, RS85, Rit88b, Rit88a, RT20, RT97, SSSR20, SG81, SG82, SW91, SMFG85, Sch89b, SB81, SB82a, ST94, Sea86, SWS^{+}12, SBHW80, SPS91, SCK^{+}00, SC99, SHMH97, SSLR90, SWL^{+}91, Smi95, Smi96a, Smi96d, Smi01, SI90, SI91a, SI91b, SA82, SA83, SS90c, Str97, Stu95, Sul97, Svo93, hTD88, THH81, THH82, Tem83, Tem89b, TH19, VSH90, Vaj91, Van95a, VM87, VTSM12, WZ87, WG82, WHBH93, War03, War09, Wat93, Way96, Wes89, WWTE92, WB85, Wil88a, WMK90, WS84d, WLN^{+}96a, WLN^{+}96b, WVBM88a, WVBM88b]. Cray [WOK^{+}00, WL83, YQTV12, Yuv77, ZH88, ZCPT00, ZM86, Zey91, van95b, Chi88]. CRAY-1 [Duf82, EM78, Rus78, SB82b, Tem88, Tem89a, BK77, Cal81, Dic81, Dic82, DR81, DR82, FR81, Fin82, Gin82, Hus86a, KJ85, Kol81, MW81, PK80, PCM84, Pet83, SG81, SG82, SMFG85, SB81, SB82a, SBHW80, hTD88, Tem89b, WG82, WS84d, Yuv77]. CRAY-1M [HL88a]. Cray-1S [SA82, SA83, THH81, THH82]. CRAY-2 [Car91, DD90, LMP^{+}90, VDK91, Wil90a, Bai88, But92, Cal86a, Cal86b, Cal88, DCG90, DD87, DL90, Ess90, FSY88, KN88, PO88, SI91a, SI91b, WVBM88a, WVBM88b, ZH88]. Cray-3 [KH87, Wat93]. Cray-4 [BJ95, Wat93]. Cray-class [MRS88]. Cray-on-a-chip [Hay89]. CRAY-System [Ent99]. CRAY-T [ACK^{+}95]. CRAY-T3D [FR96b, FR96c, Sch96]. CRAY-T3E [Che99, Ma99]. CRAY-X [HL88a]. CRAY-X/MP [HL88a]. CRAY-YMP [Car91, HP95]. Cray/2 [FG87]. Cray/SGI [Smi95]. Crayons [Sin08a]. CrayPat [KAMB19]. Crays [Ano89k, BRL^{+}20]. Craystack [Bur00, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Bur01a]. CRC [WD94]. Create [Law00, Van97, Ano93v]. Created [Kra20]. Creating [KC93c, MB94a, Ano95-31]. Creation [Lin82]. Creativity [Pin01]. Creator [Coc01]. Creveld [CCKSS90]. critic [Ano96o]. critical [Ana91]. Criticality [CLP93, CPR93, LUT96, MNV93, Mon93, PN96, SR93a]. Cross [Ano94-50, App95, Car94b, FH95, GD97, IHSK93]. Cross-Loop [Ano94-50]. Cross-Media [App95]. Cross-Platform [FH95]. Cross-Section [IHSK93]. crossbar [DuB90, DR91, HM93a]. crossbar-based [HM93a]. Crothers [Ano00a]. Crown [Ano88g, Per02]. Cruncher [Ano95-31, Ano97d]. Crunchers [WGOY91]. Crunching [Fri91]. Cryogenic [Ano96u]. Cryptochip [Ano96-31]. Cryptography [DDJ98a]. Crystal [Ano90e, Par90a]. Crystallography [CDMW94, HGB90]. CS [Ano97b, Ano98d, BCM94, BHM94a, Hoc94, SN95a, SN95b, Win02]. CS-2 [BCM94, BHM94a, Hoc94, SN95a, SN95b]. CSE [Kah94, Voi94]. CSIDC [Pin01]. CSRD [CSR89, CSRxx]. CTADEL [VWC96]. CTM [WLH00]. CTSS [Mir88]. Cuban [CE18]. cube [Bue91b, HCL88]. cube-connected-cycles [Bue91b]. cubes [DT96]. cubic [BE93c]. CUG [MH18, TH19]. Cul [Bel96]. Cul-de-Sac [Bel96]. Culler [Pin01]. Cup [Smi95]. Cure [LHLM95]. Current [AD97, Bos94a, KF95, LEMS95, MG95, TYK93, GKSR14, KK93, KBVH14]. Currently [RDZ93]. Currents [FA93]. Curriculum [HS96, HSxx, Sub94, AGEL13, Win02]. Curtain [Ano95k]. Curves [Mil88b, BB91b]. Curvilinear [PA93a]. Custom [EFG^{+}05, KC93c, Mil97b]. Customer [Bel93]. customers [CS82, CS86b]. customizable [RR99]. Customized [Ano00b]. Customizing [GZW^{+}22]. cuts [Ano95l, Ano96u]. cutting [Wat92b]. cutting-edge [Wat92b]. CV [Sch97b]. CV-Joints [Sch97b]. Cyber [TGV08, HL88a, Whe83, Dic81, Dic82, Mil88b, Tem83, Uni87b, WL83]. Cyber-Physical [TGV08]. cyberinfrastructure [NSW08]. Cybernetic [Gib01]. Cycle [Bel93, RCK97]. Cycles [Ano94-96, HB93, Mye86, Ano96u, Bue91b]. Cyclic [ACL93, HKMCS94, KNS95, SSG93, Fra90, GS87a, GS88b, Sch87d]. Cyclically [GT91]. Cydra [RYYT89]. Cylinder [AGH^{+}90, YF95, YYK93, Xu91]. cylinder-to-cylinder [Xu91]. Cystic [MHKY97].

D [Ach99, Ano96c, KTN^{+}14, Kow86, RD94, AGZ94a, AABK95, Ano94-28, Ano94-62, ACK^{+}95, BJ95, Bel99, CGLY96, CS90, CMAS11, DGO90, DP90, DMPR93, DS94c, EFR^{+}05, Elm93, FMD07, FDM07, GD94a, HKT92, HS94d, IHE^{+}00, KLY94, KSM^{+}08, MKDY90, MAA93b, Mir88, MDW93, OPR01, PLS20, Pau08, PPM90, Sch97c, SHZK94, VTTS98, VH93b, WLH00, WQS92]. D-cache [BJ95]. D2 [SVML95]. DaDianNao [LLL^{+}17]. DAGS’94 [Dra94a]. Daimler [Kad94]. Daimler-Benz [Kad94]. Damage [HMS93, Ano95w]. Damped [Man91]. Damping [IHK93]. Damping-force [IHK93]. DAMQ [Ano94-30]. dannykh [BKM88]. Daresbury [THH81, THH82]. Dark [Ano95l]. DARPA [Ano89f, Coc01]. Darwin [Tre97]. DASH [GM94b]. Data [AW94, A^{+}02, Ano92k, Ano92q, Ano94c, Ano94t, Ano94-29, Ano94-50, Ano94-98, Bel93, BMSP94, Cal11a, Che90e, CMHK92, Che94b, CCSS98, Con94, Den80, DGT84, DLLG98, Dra95, ES92, EBS02, FHM95, Fei05, FK93, Fru93, GPKK82, GG96, GS94c, GCY^{+}08, Gol99, GAV95, HMM94, HL95, HB96, HQ91, HBDS93, HM93c, HHGS93, HT72, IK82, IM96, IBC^{+}11, Jac85, JS86, JHGLG93, Jia94, JB90, JM89b, JM89a, Joh90, JM90, KN88, KHSJ94, KG96, KM92, Ker94, KCPT95, Kun95, LYL87b, LR92, LZF16, LMM^{+}21, MPT12, Mas91, MTHP93, Natxxe, NGDH96, PN13, Pit89, Pot87, PW94, Psa92, RAES96, RS94a, RWNJ94, Rei85, RW94a, SKSD94, SkLC^{+}03, SIKD94, Smi93, SC93, SSOH95, SHB^{+}13, TZ94, TG95, TGV08, TA94, TY96, TC93, UZ95, VW95, VV94, Who92, XL94]. Data [Y^{+}92, YKK96, Ano90a, AJFH86, Bab90, BAAD^{+}97, BCH^{+}93, BB91a, BWHS18, BJV^{+}16, BF92, CD95b, CP13, Che90d, Che96, CLF^{+}19, CV88c, CFP20, DGT82, ESTA94, FL92, GJG88, GB91, GZE^{+}05, GP91b, GP91c, GS93, GM93a, Gor89, GGV90, GV92, GE96, Jéz00, JM89c, KFML20, KB18, Law89, Lee87b, LL88, LEY86, Li89, LQFC18, LLDF95, MP88, MWRK18, Moo11, NS88, ODAZ15, Pan96, PP91, PP92a, Pol90, Sci86, SSSR20, SF82, SLY89, SC04, SR10, SB18, SGB91, Su92, TZY88, Tan89b, TYZ90, The90b, The91, TJC91a, TJC91b, Tuc91, VM07, WLH00, Wil90b, Woo92, Woo94, YH92, GS92b, HAG^{+}13, MBM^{+}20]. Data-Acquisition [Bel93]. Data-Communication [RWNJ94]. data-dependent [Woo92, Woo94]. data-driven [AJFH86, SB18]. Data-flow [Rei85, GS92b]. Data-Intensive [GCY^{+}08, MPT12, SC04]. data-level [TZY88]. Data-Localization [YKK96]. Data-Parallel [HMM94, HQ91, RAES96, UZ95, Con94, BCH^{+}93, GS93]. data/control [GP91c]. Database [ALPP00, Bar01, CK92a, CGHL94, CV93, FH95, LS93c, SSS94, Tak93, WWY93, Hsi91, McC94, Wie87, Win02]. databases [Ano99, Mou89, Mou90]. Datacenters [IBBA20]. Dataflow [KNYT95, Mas95, MNB94, Rui91, SA82, YMY92, YKK96, Ana91, Bro86, tHd90, Joe87, PSS^{+}19, Sch90b, Sch95c, SMM17, TS90, TS91, Van86]. datasets [CKS99]. daunting [Per87]. David [Win02]. day [Kar10]. Days [Ano95l, LM92]. Dayton [IEE94b]. DB [Win02]. DBC [CK92a, SG92b, GS93]. DBC/1012 [CK92a]. DC [IEE94e, Kho94, ACM92b, ACM92a, Ano96q, App96, FL92, Gra93c, Soc94]. DCC [SC93]. Deadlock [Tan87]. Death [Ano92d, Smi96a]. Debate [Can92, Con91, Con90]. debis [Har91]. Debugger [Ano94-106, SABJ94, EM91]. Debugging [AP87b, KC93c, Kra01b, LAdS^{+}15, LR88a, MB94a, BHS92, EGP88, FP00]. Debuts [Ano93z]. Dec [LP90, Uni91b]. Decades [Woo96b]. December [ACM95c, Fox97, HHK94, IEE85, IEE97a, KK85, Nat84, PL91c, Pra95, Sam91]. Decentralized [XL94, FEK20]. Decision [ABCE97, BCCG97, BK91a, DGJG93, EFPSS93, OIY91, VRSG93, Ano94k]. Decision-Machine [DGJG93]. decisions [RYYT89]. decodable [Mer86]. Decomposition [ABBB94, BHLST94, CHMS94, Kar94, KDBG95, MM94b, TD96, AT89, Bab90, BS87a, Bis94c, Bra91a, BS90a, CS88, Che88, CS89, Che90b, Chi81, Fra90, GL88, HLTZ93, LG87, Lou90b, Nee90a, Scr88, Bis94b]. Decoupled [HT94]. Decoupling [SSOH95]. Dedicated [Ano93i, Rit97, Ano94a]. dedication [Deu86]. deduction [LMD98]. Deductive [Tak93]. Deep [DA97, Han89, CFP20, KSB^{+}19, MBM^{+}20, Ano96-30, Ano96-44, Eva97, HCH95, Tan95]. deep-learner [CFP20]. defeated [Eva97]. Defense [CKS99, HG02]. defined [SFC^{+}21]. Defining [Lun94]. Definition [Sak02, Hus86b, HKP88]. Deflation [Man91]. deflection [Smi92]. Deformation [FDD02, Ver95]. Degradation [WWKR97]. degree [Che91]. delays [Gra92]. Delft [DSZ96, DFSZ88]. deliver [Ano96u, Mar90]. Delivering [TC93]. Delivery [Ano97p, MW81, MKHY95, MHKY97]. Delta [MFK94, HL93b, KRJ93, Pad89, TFB94a, TFB94b]. Demand [Ano92j, Mas95, TP95, AJFH86, Bab94, FK98, SCH94d, VM07, Wat92b]. Demand-Driven [TP95, AJFH86]. demonstrates [Ano92-42]. Demonstration [JR94, PT93]. Denelcor [Hay86]. denies [Ano96o]. Denneau [War03]. denominators [ARW92]. Dense [BCCG97, NJL94, Pop97, BS88a, DH86a, GPS90, GKL^{+}87, HOSZ97, HLTZ93, Ipe19]. denser [Bas95b]. Density [DAF^{+}90, HFNP96, RHH96, VD96, Ano89h, FGC06, WM92]. Deoxyribonucleotide [HL91, HLxx]. Department [Ano94-107, CKS99, Ano91v]. departmental [Ano88f, RYYT89]. Departments [Vro94]. departures [Bel92a]. Dependable [Bar93a]. Dependence [Ano94-50, Ano94-109, Ban88, ES92, Mas95, MT91, Psa92, Dra89, Hag90, HP91, Li89, PP91, PP92a]. Dependences [CWW94, TG95, GP91b, SLY89]. Dependencies [Ano94-50, Mas91, Mil87]. Dependent [GH93, GWH93, KG94, Now93, RDZ93, VMS93, Woo92, Woo94]. Depletion [DLG93]. Deploy [PSS^{+}19]. Deployment [JTX^{+}22, KSB^{+}19, SR10]. Depression [SS94]. Depth [RS94c]. Depth-First [RS94c]. Derivation [BS01, KRS94, LMM93, Rul93]. Derivatives [SF93b, SA10a]. Derived [WWKR97]. Deriving [KKB92]. Derrick [Ano00a]. described [Kah92]. Description [MGA94, BV96, Gok89, Gok90a, SK93a, Sch90b]. Design [AAC^{+}05, AH93, Ano94-87, Ano94-30, Ano94-31, Ano94-107, AJ93, ATSA93, BMCA93, BGM^{+}11, BJH97, Bos94a, CBB^{+}05, CV92a, Cig97, CGHL94, CW89, CF94, DPS97, DMPR93, DuB90, DR91, ES96, EAMEG11, ESMH93, EH97a, FNP^{+}84, FB91b, Gib93, GY93b, GTV91, GD94b, HB93, HHSW93, HU93, HW97, HSW^{+}90, HS96, HSxx, HGB90, HCV97, HK97, Hwa84, IK82, IBC^{+}11, IHSK93, Jia94, Joe87, KC93a, KWH94, KTK94, KP94, KTNM93, Kul94, Kum91, KT93b, LN94, Leg94, LMP^{+}90, Lin82, LJ97, LYKM97, MWB95, MCB^{+}01, MOWW96, MHE97, MM94b, MBB^{+}91, Nag94, NW97, PPP94, PCK93, PS96, PSS^{+}19, PPG94, Pin01, Raa97, RRSS93, Ric90a, RL90b, RLC91, RGL^{+}15, Row86, RW94b, Sei94, Smi93, Smi89, SBW94, Str97, SCG^{+}13b, IBM13a, Tsu91, VHJB94, WKL^{+}16, WD93b, WJ94, YS94, YSL97, ZS93, Ano96u]. design [Ano97m, Ano02a, Ano02b, BBC^{+}99, BHM94a, Bel89, Ber89a, BGT90, BCK13, BHD^{+}05, Car93, CV91a, Che93b, DS96b, EM94b, Gal87, GL91, GREC91, HCL88, HT72, Hin93, Jon89, JD95, KK90, LD90, Lav89, LYL87a, Loo84, NAS93, Pad89, PGK^{+}10, PBK96, Pol88a, Ram88, RYYT89, Ren97, RMM20, Saa87, Sam85, SCK^{+}00, IBM13c, TB89, Vei85, VFK^{+}04, Wie87, YKY90]. design-space [SCK^{+}00]. designed [Ano94-27, CCG^{+}17]. designer [Ano95p]. designers [Bel92a, DWV92]. Designing [Ano94-51, ABMW93, Cyb89b, Geu97, GV96a, GIBGA93, Jen87, KK82, MMR96, Qui87]. Designs [SW10a, Wea97, Ano96u, Leu90]. desk [Ano92-47, Kah91]. Desktop [Ano90p, BB98, CF12, PSB01, SB01, Ano90q, Ano94-82, DR98, Gui06, Mar90, TMHH95, Vol89]. Desktops [BH93]. despite [Ano96z]. Destination [RFS87]. Detailed [War93b, ZX95, CD09]. Details [HN90]. Detect [Mas91, Str94]. Detecting [EGP92, GV92, Her95, AC84a, AC84b]. Detection [BEH^{+}94, DKF94, FBB97, GMG94, GL93b, MRL^{+}17, SKN96, Ver97, CV88c, NG92, PE88, WLLZ20]. Determinant [Mur91a]. Determination [Ano94-89, Ano94-98, EWS^{+}13, HCV97, NBGS96]. Determine [NW97]. Determining [GS94a]. deterministic [DY90]. Detonics [CNGR90]. Detroit [IEE95b]. Deukmejian’s [Deu86]. Developed [AHSS93, Ano92h, RS94a, Ano90l, Kel85]. developers [Str94]. Developing [Ano92e, CCSM97, DP90, GV96a, PL91a, PL91b, PRS94, DWV92, SF82, Woo92, Woo94]. Development [ASSW93, AKT90, Ano94-32, Ano94-33, ATSA93, Asa93b, AA93, Bha94, Cho90a, Cig97, Dav86a, DGO90, Dav00, EM94a, FNP^{+}84, FBJ^{+}94b, FNT93, HM93b, Him93, HNST93, ITOK93, KGKa93, KTNM93, LC90, LKH94, Mac91b, Maj94, MM90, MAA93b, Mit96, MMK97, MT96, NNSY94, New93, Nor97b, Oya99, PZA87, PRSS94, Raa97, Roh94, RDZ93, SN89, Sma93, Swe94, TK93, Uni86b, Uni86a, VD94, VF93, VB90, Y^{+}92, AG90, Asl91a, CSFS00, Ele93, Gil88, GM87, Gua88c, IKM85, Kha91, Kin96, LSS^{+}20, Mar86, Mar88b, MT13, Nat91a, PATT12, R^{+}00]. Developments [Ano89o, Ano93j, Ano95n, Ano95o, BM96, Fer83, Mar85a]. Develops [Ano97o]. Device [Ano91c, Ano94-94, KTK94, TTD^{+}11, Ano90e, MS84, Par90a, YKY90]. Devices [AHSS93, Hes90, Man90, BMW91, Bur93, SF91]. DeWitt [Win02]. DeWiz [Kra01b]. Dfl.180.00 [Tru88]. DFS [SSH96]. Diacid [VM94]. diagnoses [Tze86]. Diagnosis [Sei94, Rol97, TYZ85]. Diagnostic [KB97, Chi20]. Diagnostics [Hei90, OBR94, GBB^{+}05]. diagonal [GP93a]. diagonalizable [LTT92]. Diagonalization [Ber90a]. diagram [SCK^{+}00]. Diagrams [OIY91]. dialects [Guz88, KB88]. Dialing [Ano93-43, DDJ98b]. DIALOG [Bau96]. Dialogue [Kar93]. Dialogue-Based [Kar93]. Diameter [CJ94, DF90a]. Diamond [MF93]. Diaphragm [SP94]. Diaphragm-Disc [SP94]. DIBU [AK93]. Did [Bel96]. Diego [ACM95c, AIA93, Ano97f, Ano97r, Deu86, JD95, Lay91a, PBK91, San91, SR93b, Tay94, Ano97z, Dau97]. dielectric [GKR91, Lee84]. Dies [Coc01]. Diesel [BK97, BPW97, CZRB93, GP93b, KKDO97, KR94b, OGOR97]. Difference [FBH93, KLY94, LC97b, NU91, Bue86b, Che94a, DP90, PW23, SA10a, Str94, Vaf88]. Differenced [MF93]. differences [EGK89a]. Differencing [MDW93, War93a]. Different [GAV95, MMRL93, WSP95, Sne94a, Sne94b]. Differential [CSSY92, Gal96, GRSS93, Wat91, WS93, Cha90, DGL89, Pet89a, Scr88, TFB94a, TFB94b]. difficult [HHS01b]. Diffraction [BHMH98, For93, PB94a]. Diffusion [AK93, CFV^{+}90, GWG93, IMA93, JV93, Koh96, LM93, MDW93, NU91, Now93, PA93a, RDZ93, SMFG85, War93a, WR95, KC93b, Zas93, CGM91, EAMS95a, EAMS95b, Hun92, KfGERJxx, SM92]. Diffusion-Accelerated [MDW93]. Diffusion-Synthetic [War93a]. Diffusive [SBY93]. Digest [Bel86, IEE95c, Wuo94]. digit [AW91]. Digital [Ano91e, Ano91h, CCKSS90, DGJG93, DM88a, DM88b, HU93, LKH94, MI93, MD88, MBK^{+}92, NAAW97, PRS94, Sin08a, Wad86, Ada95b, Bar88, Wes89, Was96b]. Dimension [DT96, XL94, KS93c, Ree88]. Dimensional [AD97, ACG^{+}90, AK93, BCM90, BSJW96, BY96, DLPQ94, GMBW93, GWH93, HFT94, Hun94, JB90, KO93b, Man90, OK93, Soe94, Sus93, TKM96, VD96, WD93a, Wea97, WFT93, Ach99, BR95, BS91, HP88b, KJ85, LB82, LLDF95, MB97, Nee90b, Nix92, RWL^{+}98, Rav92, Rav95, SBHW80, SM92]. Dimensions [Ano92-38, Cox88, KT93a, OGR95, OMR93]. dimeric [PB98]. dinosaur [Tay95a]. Dioxide [CSRB90]. dip [jJ88]. Direct [BPJ94, BJLW95, Ger90, HVSB93, KY93, KSP13, LN94, MS94c, RT97, Taf96, TYK93, WG93a, Abr90, DD88, GSZ91, GW95, HPH^{+}20, HJZ94, JHZ95, LA93, Sam85, ST90, Wil90b, Yan90a]. Directed [Ano94u, Ano94-56, CV95, SM94, WTC^{+}02, Che89c, CV89a, GGV90, PL91a, PL91b].

Direction [YA93, BMS92]. Directional [FCD97]. Directions [Bos94a, CCKSS90, Chi00, DSSS05, KB19, MMW86, MSK^{+}02]. Director [Bro91b, Lew94b, Ano95v]. directories [CV91b]. Directory [CS82, CS86b, MTLL94, MH94, Tra89, Ano88s, CV90, CV92a, LY90b, She90]. directory-based [CV92a, LY90b]. Dirt [Lew96a, Lew96b]. Dirtology [Jan96]. disappear [WZB86]. disappearance [Tay95a]. Disassembly [TSSK94]. Disaster [RWCA94]. Disc [SP94, Ano02a, Ano02b]. Disciplines [Pet89b]. Discontinuity [GH93]. Discontinuous [MDW93, YA93]. discoveries [PMS^{+}08]. discovery [Mit88, Nat91b, Nat92b]. Discrete [KGKa93, Meh94, MRL^{+}17, Was96a, AZC13, Joh91, Kon91a, KY91a, KY91b, RMM87, RM88]. Discretisation [GW93a]. Discretizations [PA93a]. Discretized [Vui93]. Discs [YMZ90]. Discussion [Kau93b, Mur91b]. diseased [MKHY95]. disjoint [NS88]. Disk [KRJ93, Ano95w, BJ84]. dispersion [WQS92]. displacement [CK90]. Display [ABM88, JB90, KFJB94, Kue93, SDB94, Ano90a, Ano97m, Mal89]. Displays [Bar00c, Bar00d, Ano97m]. Disposed [DA97]. dispute [Ano96-34]. Disruptive [JR94]. Dissent [Lew17]. Dissipation [GML90]. Distance [AM94, BBD^{+}08, KHN89]. Distillation [ZBLZ95]. Distinct [ER94]. Distinguished [Pin01]. Distress [COC93]. Distributed [AW94, Abr94, ADLL01, Ana94, Ano94-34, Ano94-35, Ano94-49, Ano94-45, Ano94-58, Ano94-84, Ano94-85, Ano94-106, Ano94-103, Ano94-90, Ano94-143, Ano95w, Ano01a, ABSS94, ASNT91, ALMS92, AHH94, AZ94, BAAD92, BIR94, BCHH94, Bec89a, BD93a, Ber95a, BSKJ93, BC95, BNSP99, CGFT05, CD95b, CC94a, CGSG94, CV95, DLLG98, DHHW93, DVWW05, EBS02, EKZ90, FBJ94a, Fos96, FS93b, GY92, GY93a, GM94a, GMSS^{+}11, GHdF10, Gol99, GMG94, GL93a, Gra93b, GL89, GS94d, GR94, HL95, HKT92, Hun94, IEE93b, IK82, Jay88b, KK95a, KV96, KISY94, Kon96, KRS13, Kue93, LK93, Law00, Lee94, LPV94, LCVR93, LL94, MWB95, Mah94b, Mes93b, Mes00, MS94c, MRAR95, MS94d, OH92, PR94b, Rag94, RW94a, SEA84, SNS95, Sch94c, SSKR97, Sho91, SG94a, IEE94d, SLRP95, SO95, SKN96, TH94]. Distributed [TG94, TAAL95, WP94, Who92, YFOT93, AGZ94b, Ara96, AM96, BBH^{+}00, BY21, BGKR99, Cal96, Car92, DL92, Dra90a, DuB90, DR91, GMF00, Hab92, HPPF94, IEE96b, Ji91, KHS88, Kha95, Kim96, KG03, KA96, KG95, Kre95, KSM^{+}19, Lil89, Liu95, LA93, PLC^{+}19, RLKW93, SFL^{+}94, SD92, SC04, SY91, War03, WvTB^{+}07, ZEC^{+}17, ZGL14, HB89]. distributed-concurrent [LA93]. Distributed-Memory [Ano94-85, DLLG98, GMG94, GS94d, HKT92, PR94b, SLRP95, Who92, AGZ94b]. Distributes [Kun95]. Distributing [YTL87]. Distribution [Ano94-65, Ano94-141, CWLT97, FSGS93, IK82, KKPR93, LMH90, PG93, Ano87d, BAAD^{+}97, BB91a, Fea94, KS95, Rob89, Whe89]. Distribution-Independent [Ano94-141]. Distributions [CLPV93, GG96, HKMCS94, KKKP93, KNS95, LD93b, SHG95, USZS96, VW95]. diverged [MT13]. diverse [Kim96]. diversity [Zor92a]. Divide [DT96, AT89, Don93c, LR88b]. Divide-and-Conquer [DT96, Don93c, LR88b]. divided [EGK89a]. diving [Wie94]. Division [Bro91b, Has84, Lee94, Age05, Mas93, Nat91a]. dizzying [WZ87]. dlia [BKM88]. DMBC [Sah95]. DME [GR94]. DNA [Bar93b, Bas95b, Bau96, BM93b, CGW05, Coc02a, Coc02b, Hei89, HL91, HLxx, HH93, JHGLG93, KGS93, KKPR93, KT93a, LD93a, Lu93, Pev93, PRS94, Poo96b, Poo96a, PG93, SD93, SSKa93, Tri93]. DNA-Based [CGW05]. DNN [XZC^{+}20]. DNNTune [XZC^{+}20]. DO [Day12, Ano94-110, Ano94-118, GP06, HHS01a, HHS01b, JA92a, PB90, Rag06, Ada95b, MAFW08, Win02, JA92b]. Do-It-Yourself [HHS01a, HHS01b, JA92a, JA92b]. Doacross [TZ94, SY91, Ano94-42, CY91, OSKO95]. doall [Jac85, RP94]. Documentation [BKM93, Ano90t, Nat89b]. DOD [Fed96, Ano95v, Ano99]. DOE [Ano94-36, Ano97-27]. Does [Gui10, JA92b, Tho93b, JA92a]. doesn’t [Win02]. Domain [ABBB94, BS90a, CBT91, Div97, KDBG95, KRS13, Pau08, Bab90, Che88, CS89, Chi81, Fra90, GL88, LG87, Scr88]. Domain-Specific [KRS13]. domains [Ano90l, AGD93]. Domik [Ano96c]. Dominating [TM94a, TM94b]. Dongarra [SB94d]. Dongarra-Sorensen [SB94d]. Don’t [Bar00c, Bar00d]. Door [OT07]. Doorframe [JC94d]. Doors [Ano92-39, Bro17a, Ano92-40]. Dopamine [SVML95, WR95]. Doppler [RCR93]. Dopplergrams [KRJ93]. dose [MB97]. dosimetry [Ano96w]. Double [Ano93k, Ano95w, NNSY94, Ban90, LKFU05, RR95]. double-edged [RR95]. Doubling [CSRB90]. Dowd [Bra94]. Down [Ano95-40, JWG93, Kra20, Ano88y, Str94, Way96, WF08]. Downturn [Gar01]. DPS [Tra89]. DPS-chip [Tra89]. Draft [DHHW93]. DRAM [HS93c, IBP^{+}05]. Drama [Smi95]. Drawing [SHA^{+}92]. dream [Ano96u, Bro16]. Dredged [DA97]. Drive [BPU94, MS94b, Ste94e]. Driveline [AM93b]. Driven [BISB96, HLB94, TD96, TP95, AJFH86, Che89a, HS93a, KN86, SN95a, SN95b, SB18, WGS91, UR95]. driven/dataflow [TS91]. Drivers [FT93a]. Drives [Ano95y, DDJ98a, Ano95w].

Driving

[HRG93, Kad94, MF97, NCDS97, Wie96]. Drop [Gre91b, Gre89b]. Drops [Ano95k, Gre88c]. Drosophila [HKG90, Hun92, Hun93]. Drug [HSW^{+}90, Ric90a, MHKY97, PB98, PMS^{+}08]. drugs [Str94]. Drum [BSJW96, Koh96]. Dry [HFNP96, RHH96, Was96a]. DSM [Ano94-66]. DSMC [BJLW95]. DSNP [AA93]. DSP [KG03, MBSK92, She90, Wei91]. DSP-chip [She90]. DSPACK [Yan90a]. Dual [EHG95, IEE96a, LM93, Pel94]. Dual-Processor [EHG95]. Dual-use [IEE96a]. due [DT96]. Dummies [TD90]. dummy [Ano97m]. dumping [Ano96o, Dum97, Mac96a]. Duplication [DA94]. Durability [MMK97]. During [Bel93, FCD97, Sch90c, DGG18, Got91b, MB93, MB94b, Pit90]. Dust [Ano89g, Ano91d]. Dutch [HS94b]. DVS [Hic18]. Dynamic [AC93, AJ97, Ano94-38, Ano94-37, AZ94, DD93, DDF93, EJ97, Fan87, FI93, GZA86, HL95, HC91, HK96, HMC94, KK95b, KB94, Mah94c, MOWW96, OP96, PZS^{+}20, PP92a, PR94b, RCK97, VR94, VV94, WQS92, ZBLZ95, BAAD^{+}97, BP92, CGLY96, CGLxx, Chexx, Chu91, DCG90, Fin82, JG88, KFML20, Mil87, Ng95, PBK91, Ram88, Roj19, SAB^{+}05, WWJ09, ZCPT00]. Dynamical [BY96, BPU94, KLN90a, Nag96c, Pas95, Sug96, KLN90b]. Dynamically [TSCG94]. Dynamics [ADGA95, ATL90, Ano88h, Ano94-137, Ano97k, ABGL96, BHEG94, CFV^{+}90, CH10, CHMS94, DAKM98, ES96, FR81, Fra94, GI93, HP93, Kue93, KK92, LD93b, Law90, LB94c, MAA93b, NCDS97, NS93, Por86, SKVZ93, Sim92b, SFF94, VVKB96, Web93, WKFFK97, Wil90a, ARF12, BBK^{+}08, COS89, COT21, DGL89, DB95, Ece96, Elm93, Elm95a, EFG^{+}05, GKS09, Hea91, Hua92, HKS93, Kha91, KHZ^{+}08, Lag89, LM92, Lou92, MCH91, MA85, OYK^{+}14, Ons88, PEH93, PZGL91, PS98, RCB03, Sch89b, SCK^{+}00, Ske89, SPP^{+}05, WCHK91].

E-mail [PA92]. E. [Ach99]. Earll [Por86]. Early [ABB^{+}13, GZE^{+}05, HLPP97, Kaz92, RS85, EFR^{+}05, Oed92a, Oed92b, SPP^{+}05]. Earth [Ano93t, Bla93, Bro91d, Fos93, GS87b, NS86, Sci86, TP97, Zyg93, GG^{+}97a, BS98a, BBB^{+}20b, Kra20, Zyg93]. ease [TKI85]. Eastern [Chi90]. Easy [SBF94]. Easyflow [AJFH86]. Eating [Ano91d]. EBE [HFH86, HFH87]. EC [Ano94s, Pel93b]. ECCO [RCR93]. ECL [LH86a, LH86b, LH86c, LH86d, LH87]. Eclipsing [Mil88b]. ECMWF [HK93b, KH93, DTV00, Isa93]. Econometric [BBC92]. Economic [LC90, NDLV88]. Ecosystem [Ano94c, GGN20]. ECU [VD94]. Ed [Wei90, Ano94p, Kow86]. Eddy [KS93b, PSB01]. eden [Bro01]. Edge [MFK94, TL96, TPJ^{+}19, ZLRC20, PMC22, Wat92b]. Edge-Cloud [ZLRC20]. edged [RR95]. Edinburgh [Ano94-33, BBC^{+}89, Swe94]. Edited [Ano00a]. Editing [Pau08, Sky94]. Editor [Ano94-28, Ano94-39, Ano94-40, Ano98e, Ken92, AB03, BKK11, Pan93, Pin01, Sul97]. Editor-in [Ano94-39]. Editor-in-Chief [Ano94-40, Sul97]. Editorial [AAB06, Bad99, Fox98, Jar12, KL99, MH18, Sul97, AB01, DF12]. Editors [AP93, BEGGK07, CFS95, HG02, HP04]. Edmonton [Ano88t, Ano88u]. EDN [She90, Tra89]. edn. [Ahm92]. eds [CCKSS90, Por86]. educating [Gra94]. Education [Ano90g, Ano91j, Ano94w, Ano94-71, Bae01, EP 97, Gar01, Joh94, JPMG08, LM08, Mah94a, Mur06, SB94a, Sub94, Sun94, Bal94, Gra93c, Gra94, Isk96, NCKMM88, SC91a]. educational [CBKA09]. Educator [Ano92f]. EEC [Rep92, Ste85]. EEG [KVP95]. EEGs [Her95]. EFDC [HW97, WHMA97]. EFDC/HEM3D [HW97, WHMA97]. Effect [BJ95, BMP93, BPU94, LMH90, Sei94, VSM96, WMR96, Woo96a, BP89a, BP91b, HEB96, NW03, Wal81]. Effective [FSGS93, GSG^{+}94, HAAS93, VH93a, Ber89b, CV92b, GV91, Poi90, Rob87, Sam85]. effectively [FRW92]. Effectiveness [DP96, GS94b, NRS95, UU94, EB91, Lee87b]. Effects [HVZ94, KBD97, KW95, SS94, TYKE93, KPS88, PLC^{+}19, Seh88, Whe89]. efficiencies [DZM^{+}13]. Efficiency [FBCB18, GW93a, INKN01, MP94, Aus93, Mir88, New14]. Efficient [AGZ94a, ASS94, AZC13, AG94, Ano94-41, Ano94-42, Ano94-44, Ano94-45, Ano94-46, Ano94-43, CP94a, CH94, CS93b, CS95, EBS02, EH97b, cF03, GS90, GS92a, GGZ^{+}20, GW91, Gre90c, HS95a, HE98, IJY^{+}14, JTX^{+}22, KHSJ94, KNS95, Kra93, KM85, LMT95, Lei85, LS93a, LY88a, LFJ^{+}20, Lil89, Mah94c, MCW98, McK94, MHP84, MRGR12, NB93, NG92, NB94, NR86, OA94, Pau08, Sob93b, SY91, TGV08, TL96, TF94, UZ95, USZS96, VV94, ZLRC20, Zla01, Aba09, AM15, ABMN02, BR95, Bis94a, LYC93, MKfDA96, Par90c, Qui87, Ren97, SQS^{+}19, SHMH97, TCM95, Woo92, Woo94, YF98]. Efficiently [AGLL98, MSTK93]. Effort [Fah94, Ano92-42, Ano99, CKS99]. Efforts [MB20, Nee94, Pin99, Ano00b, Com92a, Hib01, SS10]. EFR [DMPR93]. Ehrhart [Cla96]. EIC [Sak02]. eigenproblem [AT89]. Eigenproblems [LO96, ALN^{+}01]. eigensolution [Pin91]. Eigensolver [BHLST94, CKD^{+}19, HV94, SB94d, Bis94b, LTT92]. eigensolvers [AT91, Bis93]. Eigenvalue [DKH86, HE98, BS88a, DS86b, GKR91, GKL^{+}87, LPS86, LP86]. Eigenvalues [WKHS97, CP92c, LTT92]. Eight [Ano95p, MP92]. Eight-Way [Ano95p]. eighteenth [KWW92]. Eighth [Sie94, Sin94a]. Einstein [AGL^{+}99].

EISPACK [McD85]. Elan [BHM94a]. Elan-Elite [BHM94a]. Elastic [BS98b, DDF93, KB19, MDH^{+}16]. elasticity [CS88, CS89, CBLS13]. Elastomers [MHE97]. elect [Ano97b]. Election [Pin01]. Electric [Ano92h, FS93a, LPLP97, Ano03a]. Electrical [MS94b, NBGS96, Asp93, CNC^{+}08]. electrical-conductivity [CNC^{+}08]. Electro [EH97b, LEMS95, BBBC96]. Electro- [LEMS95]. electro-optic [BBBC96]. Electro-Optical [EH97b]. Electrochemical [Ano95q]. Electrode [GML90, RSRG95]. Electrodes [LMH90]. electroluminescent [Ano02a, Ano02b]. Electromagnetic [AFT96, ASSW93, AHSS93, Ano92g, DNV93, JBWB97, Man90, SE90, Fuj11]. Electromagnetics [SHMR96, Taf96, Ano87d, Ano96q, App96, CP93b, EY91, PSG03, SHMR94, Vez95]. Electron [BHMH98, Hes90, Sil91, Ano02a, Ano02b, EBS88, For93, PB94a, Zho88]. Electronic [Bos94a, PA92, Ris94, Suh97, AS99, Gla93, HBCN95, NPS93, RGL^{+}15]. electronic-structure [NPS93]. Electronica [MMR96]. Electronics [Ano92h, Ano94-108, IEE94b, MS94a, Ada95a, Ano97-29, Art93, Asp93]. electrons [BDM94]. Electrophoresis [CL91, SBY93]. Electrophysiology [MH95, KSM^{+}08]. Electrostatic [RD07]. electrostatics [TFVK94]. electrotechnical [De 96]. Elegans [Dro95]. Element [Ano94g, Ano94b, BK95b, Bro97, BV93, CSSY92, ER94, FB91b, Glo84, Glo89, IS95, JM90, LM93, MF92, Nag94, NBGS96, OD01, RDZ93, SMFG85, Sch94b, Sha94b, TD90, TOWC15, AS99, Ano95v, Arb92, Che94a, CRA10, DL92, FR95, Hea91, Ji91, JM89b, Mac96c, Ram86, SSLR90, TR86, Van95a, Vez95, Was96a, YYYS93, YHA93, van95b]. Elementary [WG91]. Elemente [Wat95]. elements [Mac92]. eleventh [Ano96l]. eliminate [Ano90e, Par90a]. Eliminates [Ano95d]. Elimination [MM94c, MM94d, Pap92, Rag94, WL83, CG87, Gan86]. Elite [BHM94a]. Elizabeth [Ano96c]. Elliptic [RT97, BB91b, BJ84, FGM90, GS87a, GS88b, GS89b, Lee96]. Ellis [Coc01]. EM-4 [BAM93, Ano94-85]. Embeddable [AK94]. Embedded [Ano95r, Ano00c, Ano00d, CSFS00, DDJ98a, EVM^{+}98, GSG^{+}94, IBP^{+}05, MRGR12, OGR95, SDK98]. Embedding [AM93c]. Embeddings [HS95a]. Emergence [AU91, AU90]. Emergencies [EFPSS93]. Emergency [IK93, ITOK93, RWCA94, TIOK94, VRSG93]. Emergent [BWO96]. Emerging [Mar85a, Sah94a, Sch93a, Cat92b]. Emerson [Kaz92]. Emission [HEJM95]. Emissions [BK97]. Emitting [Bar00c, Bar00d, Ano92h]. Emphasis [Sch93a, Kah92]. Emphasizes [Hol94]. Empirical [ACK^{+}95, LH94, Wie96, CY91, SLY89, SLY90, VSH91]. Employment [Gar01, Ano94s]. Empowering [Mor92b, KWW92]. emulate [WZ87]. emulation [ZS94b, ZS94c]. Enable [Ano94-143, HC99]. Enabled [GK18, Stu97]. enablement [MLWC20]. enables [AB01]. Enabling [APK^{+}12, Ano94-47, Ano96h, BCH12, FT96b, KHBB01, SMS95, ZEC^{+}17]. Enchancing [EE93]. enclosure [Ha88, Ha90a]. Encoder [TCJS93]. encoding [BR95, WD94]. encompassing [Ano95-32]. Encouraging [cFC07]. Encryption [WM91]. End [DM88a, DM88b, GF90, MD88, Mou89, Mou90, Ano89k]. End-user [Mou89, Mou90]. ended [Fin94, MSCxx, TR86]. Energie [Pre93b]. energies [Ano94s]. Energy [CTD^{+}16, FSGS93, FBCB18, FLP^{+}07, GGW93b, JBWB97, JR94, Mac97b, Mir90, MRGR12, Nat90, Pau08, SW10a, SHG95, TGV08, Uni86b, Uni86a, Uen93, A^{+}12, BMR85, For93, KNHN16, KFML20, MWRK18, RLKW93, Roj19, SNEP14, SN96, SQS^{+}19, Uni93, Uni95]. energy-aware [A^{+}12]. Energy-Efficient [MRGR12, Pau08, TGV08]. Energy-Time [FLP^{+}07]. enforcement [CV88c, Dam11]. Engine [BCW93, BK97, BPW97, GWG93, GP93b, HK97, KLSC97, KB97, Law90, OGOR97, PB94b, Tak93, TCJS93, VM94, VF93, BCK13, PSO12]. Engineer [BCC^{+}08, Bor01, MM94a, Per87, Wen94, Hil97]. Engineer-manager [Per87]. Engineering [AS98, Ano90g, Ano94-107, BGS^{+}12, GT97, Got91a, GK18, Gro90, HF93, Hwa85, Jal94, Jon89, KS94a, LJ97, L^{+}95, LCD97, MW81, MBW01, Nas91, Pin01, Pit90, SC99, Str94, SR93b, Vro94, Ade92, Ano88e, Ano88d, Ano88c, Ano89r, BP89b, BP93, C^{+}97, CCC^{+}89, Cre91, FK98, Fox97, GL90, HS^{+}91, Hen91, IAIK92, Isk96, JT87, JD95, Kho94, LP90, Mar86, Mar88b, MB89, Som13, SPK94, Gra93c, Pin01]. engineering-90 [HS^{+}91]. Engineers [Gar99, Kho94, Gra94, HW11]. Engines [AABB93, AGH^{+}90, BSB93, HB93, Pay97, Ano96u, AJFH86, Mar90]. England [ACM94, Ano88d, Ano94p, OMM93, Pit90]. English [NSW08]. Enhance [MNZ^{+}15, TJ94, SC91a]. Enhanced [LYKM97, MM93a, McC94, EFH^{+}00, SC20]. Enhancement [Ano88k, yHYZ87]. Enhancements [Iwa92]. Enhancing [Ano96i, BKM93, Mas95, Pol88a]. Enormous [Lin83]. enough [Ano95v, Bab94, SCG13a, Win02]. Ensemble [Kau93a, WLLZ20]. ensuring [Ano97u]. Enterprise [Ano92i]. Enters [Ano90p, CCKSS90, Ren97, Opp95b]. entities [Ano96z]. Entrepreneur [CCKSS90]. Entropy [Lu93, BB87]. entry [Ano94-83, Ano94-120]. entry-level [Ano94-120]. Enumeration [SVD96]. Environment [AW93, Ano94-51, Ano94-114, Ano94-138, ASNT91, AHH94, Bae01, Bha94, CWD^{+}08, DPS97, DHHW93, EGH^{+}06, FCGG90, FG93, Gil93, Gin82, Gol99, Gue90, GBK^{+}96, IEE95a, IJM14, JAB92, JTX^{+}22, KK95a, Kue93, KNWB93, MM90, NU91, OCVG93, OPR01, PS96, Par86, PL91a, PL91b, Per93a, PNK93, RL90a, RLW93, Sat93, SES94, SBF94, SLB93, SkLC^{+}03, Ste94a, WKFFK97, YJD93, Ano85b, Ano90t, ABGL96, BBB^{+}20b, Bru90a, Cho90b, Def87, Don85, EB18, FEK20, GGJ89, GBB^{+}05, Gok92, Gua88b, Gua88c, GG88, Hab92, IEE91, JG88, JR91, Joh86c, Kha93, Kha95, Kos95, KW11, Lev89, LC12, MP88, McC94, MK92a, MK92b, Mir88, MSW91, MAA^{+}05, NRN00, PZGL91, Pol88d, Rob85, SCV01, Sch95c, Ser98, SW99, Tur90, War89, YVC89, YB90, GR94, HB89]. environment/application [Mir88]. Environmental [ABCE97, Ano93-29, Ano94-48, Ano94-110, Ash93, BAAD92, Cul95b, JBWB97, Koo97, SWSR97, Uni92e, Uni96, GG^{+}97a, MWRK18, NS86, Sci86, Sch94b]. Environments [Ano94-102, Ano94-103, CDPW94, CM93, DT97, GFB^{+}03, MS94d, Ste94e, A^{+}02, Bla97, Bri90, JS86, KBM^{+}02, LL88, LLSR02, MGS^{+}20, McN87, PGL87, PSM93, Pei17, Saa87]. Envisioning [Llo94]. EPA [Cul95b]. epacts [Whe89]. Epilogue [Bur01a]. EPL [ZW02]. EPL-Julia [ZW02]. EPS [GT94]. EPS-APS [GT94]. Equalization [Pan96]. Equalizing [MD04]. Equation [AFML93, Ano94-61, Cal86a, Cal86b, DMPR93, Gui08, JR94, Pev93, SMFG85, Sus93, KC93b, DGL89, Gao86, Gri86, Scr88, Sta95a, Sta95b]. Equations [AM93a, AGL^{+}99, Ano94-140, CSSY92, Dic94, Duf82, Duf91, Gal96, Glo89, GW93a, GRSS93, HO92b, LMM93, McB93, MF93, MDW93, MM94c, MM94d, Rul93, Sha94b, SC92, Taf96, TYK93, VAGRMVA90, Vui93, War93a, Wat91, WS93, Ach99, And88, Ano87e, BS87b, Cha90, Che99, Dav86b, DD88, Don85, GS87a, GS88b, GS89a, GS90, GS92a, Kam86, Ked92, LM90a, Pet89a, PO88, Sch87a, SM92, TFB94a, TFB94b, vdV91, McB92a, McB92b]. Equilibrium [HVZ94, NSH95]. Equipment [Bar88, CCKSS90, Was96b]. Equivalence [McT96, SZ11]. equivalents [Bru91]. Era [Bro91c, dRSGS16, DM88a, DM88b, MD88, SGH97, Ano89k, EP 97, KS87a, SA10a, SA10b, Zor92a, Sch22, Haw88]. ERANOS [DMKW93]. Erba [DJM94]. Ergonoic [Geu97]. Erie [MKDY90]. Erratum [Ano18]. Error [NPS^{+}20, AC84a, Bli89, AC84b]. Error-Aware [NPS^{+}20]. error-correcting [AC84a, AC84b]. error-correcting/detecting [AC84a, AC84b]. errors [Mit88]. ERS [OLLG96]. ERS-1 [OLLG96]. Esprit [Ano85b, Hey94]. ESS [IS20]. Essay [Smi96d, Smi96a]. Essential [GP85]. EST [FAKD93]. established [Bla97]. Establishes [Ano93-46]. Establishing [MS94a]. establishment [Uni96]. estate [Ano89i]. Estet [GMS97a, GMS97b]. Estet-Astrid [GMS97a, GMS97b]. Estimates [KB93, SH91, TDBL13]. Estimating [Gre94, RDHC94, SK92]. Estimation [BB87, Fuj99, Mis90, PC97, RMPW93, Wil94, YOY97, YAG93, YAGxx]. Estonia [KK93]. ETA [Ano88h, CSB89]. ETA10 [Car89b]. ETH [HKR94]. Ethernet [Kon87, OBB^{+}05, WTC^{+}02]. ethical [Chr92]. EU [Ano03a]. Euclidean [Gur88, RW89]. Euclidian [FRW92]. EUDOC [PMS^{+}08]. Euler [Dic94, Gri86, GD94b, Sha94b]. Eulerian [BK93]. EuroBen [Van91a]. europaeische [Wac92]. Europe [Ano93-39, Lid96, Lid99, LCHS96, Ano14, Duf84, Duf85, Kir89]. European [Ano92j, Ano97c, DLM99, RMO96, Ano85b, AGL11, DMKW93, Els21, LPC^{+}95, LMP^{+}90, Pel93b, PC94b, RCR93, SS90a, SS90b, SSxx, Sta88]. Europort [SS96b]. Eutrophication [HW97, WHMA97]. EUVL [Bar01].

Evacuation [ITOK93, TIOK94]. Evaluate [Bur94a]. Evaluating [DAC^{+}18, EAMEG11, GB96, Ked94, KB96, McK94, PSO12, Sri94, VdSK^{+}05]. Evaluation [All93, Ano94-30, Ano94-54, Ano94-73, Ano94-102, Ano94-103, AC84b, AHOK02, AK93, BK77, BBH95, BS94a, Bie88, BD94, CGSG94, CP94b, COC93, CMF94, DVWW05, FR95, GA95, Gin82, HS96, HSxx, HKT92, IBC^{+}11, IK91, KC93c, LMH90, LBT94, LMM86, MOWW96, MTLL94, MH94, NH91, PTC^{+}93, RRMD94, RLKW93, Rue92, TNIA92, TGL96, USZS96, WOG94, AP90, Ano94-100, AC84a, ACK^{+}95, BS94b, Ber89b, Bli89, CC96, CKPK90a, Cyb90, CKPK90b, Cyr86, DR91, Eig01, Har86, HP88a, HY92, Inf86, JOK^{+}18, JD95, KS87b, LS92b, LC12, LYC93, Mal86a, MMW86, Mar88c, MI01, MDH^{+}16, OW94, Par90b, PPR95, PP91, PP92a, PP92b, Poi90, RGH17, SCG^{+}08, SNS^{+}97, SWS^{+}12, SHB91, hTD88, Tan89a, TC94, WHL93, YB90, YHA93, Yi11]. Evangelizing [Coc01]. Eve [Ano95v]. Even [AFML93, Ano89n, Gib01]. Event [EGP88, Mal91, AZC13, Kon91a, KY91a, KY91b, Li91, Mal89, RMM87, RM88, TS91, WGS91, Zor89b]. Event-based [Mal91]. event-driven [TS91, WGS91]. event-driven/dataflow [TS91]. events [SM89]. ever [Ano96u, DWV92]. ever-higher [DWV92]. Every [Dun92, Gol91a, Gol91b, Wic92, Ano92-47, Ano94k, Kah91]. everything [Way96]. Evidence [RSRG95]. Evolution [CCKSS90, San93, Ano89d, CK92b, IEE89b]. Evolutionary [ZW02, CFP20]. evolve [Ano94-131]. Ewald [FMD07, Fin82]. Exact [MF93, Psa92]. exaflop [Gel11]. Examination [MH96, SMH91]. example [DFSZ88]. Examples [FD93, SS96b, SBN82, WBP87, IBM13c]. Exascale [Ano08a, DMT^{+}21, Get15, KLQ19, MRL^{+}17, SBW^{+}19, YWXZ12, And20, AGL11, BBB^{+}20a, CCG^{+}17, Hsu15, Hsu16, LSS^{+}20, PLC^{+}19, PMC22, Vet12, MB20, Mes17, Sch22]. Excel [Ano94p]. excellence [BBD92, DRB^{+}20]. Exchange [PPG94, SJPS94, SJPS96, TS94, XL94, Bue91a]. Exchanger [MS97]. excitation [CGLY96, CGLxx]. excitations [Chexx]. Excitatory [KW95]. excited [RLKW93]. exclusive [Per83]. Execute [GS94a]. Executing [LZ95, Chu87, HC91]. Execution [Col94, GHWZ94, JAB92, KR94a, Li92, Mah94c, MM93b, McK94, MNB94, VSM96, YSKS95, YAG93, YAGxx, ZX95, Che89a, Cho90b, GP88, KPS88, MLR90b, MLR90a, OWG^{+}13, Pol87d, Sch88b, Seh88, SY91, YF98]. executive [Bli91]. Exegesis [CK92a]. Exercise [HWS^{+}88]. Exercises [PT93]. Exhaust [OGOR97, WJ94]. Exhaustive [Lu93, QD91]. Exhibit [AIA94]. Exhibition [Ano88w, GH94a, GH94b, GH94c, HS95b, HS95c, IEE94a, IS95, KK89a, KSW93, L^{+}95, Qui95, Ano93-39, Hel93, LCHS96, Ill96]. Existing [STN93]. Exons [HH93]. Expanding [Coo95, HS94a, Pap16]. Expands [Ano87a, Ano00b]. Expansion [Mur91a, Ano02a, Ano02b]. expansive [PZ89]. Expectations [Ano94j, Lin83]. Experience [BS94a, Chi00, DR81, DR82, Eig91, GK92, JM89c, JM89a, Kau93b, MJH90, Mes93b, RS85, Rit88a, Sim92a, BS94b, HKN89, Kar89, LSS^{+}20, NSH95]. Experiences [ASSW93, Ano94-78, BMT96, CDH84, DFS93, DXJM93, EHG95, GB96, GL93a, HL88a, Hay86, Kah93a, MMRL93, Nag88, Rit88b, SMFG85, SSH96, ABB^{+}13, Sch94c]. Experiment [Fos96, UU94, BCCP05, TfGERJxx]. Experimental [ASM86, Ano90a, Ano94-49, Ano94-68, BMP93, CAB93, Don87, DDF93, Gis86, Ha88, Ha90a, HGS88, HHK19, JCJY94, LGG^{+}87, OMM93, PP91, PTC^{+}93, TGL96, USZS96, MS84]. Experimentalist [Gha96]. experimentation [GGJ89]. Experimenting [EO13]. Experiments [ASNT91, Asl91b, AK93, BD94, CRM94, DCG90, DGG92a, DGG92b, DAKM98, Fra90, FGM90, Gri88, KK92, RT97, BP86, Kor93, SZG95, VDK91]. Expert [IK93, Dan91, Joh88]. Expertise [Pin99]. Experts [PD94, Ano94-119, Ano95l, Ano97j, PZ89]. Explain [Bis94d]. Explaining [SH93, SH94b]. Explanatory [FNK93]. Explicit [Gri86, Noo95, Sch93b, WVBM88a, WVBM88b]. Exploit [Rie93]. Exploitation [TJ94, VSH90, CBB^{+}05, Lee86]. Exploiting [AACK92, EAGEG09, GW95, LS94, Nag90, NMS93, SWG06, WBP87, FMT91]. Exploration [BGM^{+}11, Che90f, DMT^{+}21, SGH97, BRL^{+}20, GE96, SGB91]. Exploratorium [Ano94-34]. explore [Ano02a, Ano02b]. Exploring [Bro91d, WKL^{+}16]. Exponential [ALM93, BL93]. exponentials [Ked94]. exponentiation [Joh92]. Exponentiations [NdMM09]. Export [Ano90s, Ano95k, EHG01, Har95, MWRK18, Uni98]. Express [Ano01b]. Expressed [AKDM93]. Expressibility [Bis94d]. Expressing [Ano94-50]. Expression [BGS^{+}12, Ede94b]. Expressions [Ano94-109]. Extended [SYMT92, TM94b, YMY92, CKM88, Dub87, GV92, PP91, SKP91, SG92d]. Extending [dRSGS16, Dra89, Gua88a, MR95]. Extensible [MCLK07]. Extension [KG96, KNYT95]. Extensions [AHOK02, CC94b, Ho92a, Wai05]. EXTENT [Ano94-51]. Exterior [PK94, PK89]. Extra [Bai88, VHJB94]. Extracellular [WR95]. Extracting [NPS^{+}20]. Extraction [CDA94, LY90a, UPK87, YKY90]. extremal [BGT90]. Extreme [BCH12, CCR11, WKL^{+}16, ZLC21]. Extreme-Scale [BCH12, WKL^{+}16, ZLC21]. Extremely [LHLM95, Luc91]. Extruder [YMZ90]. Eye [BBL95, HA92]. eyes [Ano95-32, Str94].

F [Ano00a, Wei90, BE93a]. F.E.M. [AJ93]. F.S.U. [LGG^{+}87]. Fabric [CEH^{+}12, CCKSS90]. Face [Ano95-35, Ano91t, Pan93]. face-to-face [Ano91t]. faced [Ano96z]. Faces [Ano97u, Ano97-28]. Facial [RM92]. Facilities [BB98, KA93a, Mon93, Pap16, Fed96, Jor87, Sci86, Sha87]. Facility [Ano88l, Cor89b, Rit88c, SW10a, And90c, MRM87, Yau88, BK91b, CR89, Lee89, Uni96]. FACOM [MU83, TKI85]. Factor [Els21, Moh94, RSB94, Tem88]. Factorial [AH93]. Factoring [BtR95, CB89, Luc91]. Factorization [Ano94-65, Ano94-115, EHS94, GMW94, KESH94, KSH94, MP94, Rag94, Rot94, AZ95, Con86, Con94, Dav89, DY90, DD90, DDT95, GHNL87, Kra92, KC92, KESH95]. factorizations [Eij90a, Eij90b, Eij91]. Factors [Ano94-52, DCWH07, GH93]. factory [KWW92]. faculty [Pan96]. Faddeev [SKB89]. Fail [Bar00c, Bar00d]. Failure [Sei94, WH94]. Failures [Ano94-139, HRC09]. fair [Ano96-41]. Fall [Mic90, CKS99, Dal96, Gre89b]. Falling [LDMC96]. falter [Ano95a]. Family [LS93c, NU91, AJFH86, BE93c, WZ87]. farming [Str94]. fashion [Ano99, CKS99]. Fast [Ano92k, Ano92q, BP90, BHS^{+}02, EGK87a, EGK89a, Elm95b, EH97c, Joh92, LH86a, LH86b, LH86c, LH86d, LH87, LM93, LG87, MOOK94, Mik94, Mik89, NS93, RMPW93, RT97, SKVZ93, SAGS93, TKI93, UP01, VTSM12, WG91, ABP92, BP91a, CHWW13, CC88a, Cat92b, CV88a, Dra90a, EGK89b, GS87a, GS88b, GS89b, Gut95, Heg96, Mas94a, MB97, OYK^{+}14, Sta95a, Sta95b, NR86]. Faster [Ano94-110, BGQ19, BBS94, Nag94, Ano91l, Bas95b, Ber82, BE88]. Fastest [Ano93l, Bar00c, Bar00d, Tho96a, Tho96b, Ano90n, Ano97u, Ano97s, Ano00b]. Fat [Lei85]. Fat-Trees [Lei85]. Fate [ZL97]. father [JNM^{+}98]. Fatigue [Bel93, ES88, JCJY94]. Fault [Ano94-53, BOS93, CRV94, CB94, CJ94, DO89, EVM^{+}98, GFB^{+}03, GBG89, GMG94, LL08, LLGS09, LBT94, MNZ^{+}15, TYZ85, Tze86, Con00, Dao88, HCL88, Mit88, OD88, SHL^{+}20, SO91, The90b, The91, TYZ88]. Fault-Aware [LLGS09]. fault-diagnoses [Tze86]. Fault-diagnosis [TYZ85]. fault-tolerance [The90b, The91]. Fault-Tolerant [Ano94-53, CRV94, CB94, EVM^{+}98, GFB^{+}03, GBG89, LBT94, Tze86, SHL^{+}20, TYZ88]. Faults [LMH90]. Faust [Gua88b, Gua88c, Ham90]. FCI [ARE95]. FCRC [ACM96]. fears [Ano96-34]. Feature [SCV01, Ano97j]. featured [Bro17a]. Features [Ara91, AGD93, KZ94, MTHP93, NW97, Oed92a, Oed92b]. Feb [B^{+}95, Zyg93]. February [Ano96l, Clo96, Don92a, GL90, GE96, JD95, M^{+}94, Wuo94]. Federal [Ano95l, MP92, Uni86b, Uni86a, Uni86c, Waz89, Ano95h]. Federated [CGHL94]. Feedback [PH11]. Feeder [SS94]. Fees [Ano94-129]. Feet [Ano95s, Bar00c, Bar00d]. FEL [SNK^{+}93]. Fellows [Pin01]. Fellowship [Kah93a]. FEM [HS93b]. FEM-Analysis [HS93b]. FEM5 [KA91]. Fermat [YB86]. Fermilab [Fer83]. Fermions [KLN90a, KLN90b]. Fernbach [Ano16]. fever [Ade10]. FEVS [SZ11]. Fewer [Ano95-45]. FFMachine [Wun89]. FFT [AGZ94a, AABK95, Ano94-62, Bue91b, Cal96, DWM^{+}01, FDM07, OLWW94, Tem88]. FFTs [Car91, EFR^{+}05, GJ87, Swa86]. Fibonacci [Alu96, AM15, Mas94a, Mas94b]. Fibre [Ano94-104, Gre91a]. Fibre-Optic [Gre91a]. fibrosis [MHKY97]. FIDAP [Web93]. FIDISOL [GRSS93]. FIDISOL/CADSOL [GRSS93]. Field [AU87, Bai92, Ewa89, KMG96, Uni87a, Pop97, Sat93, SBY93, Tho93a, VH93b, WWKR97, Ano92l, Ano95w, CR94, MR86, MR87, Ano97j, ML95b]. Field-Derived [WWKR97]. Field-programmable [ML95b]. Fields [KMT94, MF92, Ano87d, Arn88, ARW93b, BE93c, CK90, Nee90b]. FIFO [HHOM91, HHOM92]. Fifth [ANS92, BG91, Pit90, Uni98, Ano93-39, HK93b, KH93, NAS93]. Fight [Ano92-30]. fights [Ano92-29]. File [Ano94-37, CF94, HERC95, JR91, KR94d, NK96, Wie87, Bin88, MK92a, MK92b, SC04]. File-Access [Ano94-37]. files [Ano95c, Hib01, Mac96a]. Filesystem [SSH96]. film [Ano95w, Hua92]. Filter [WWY93, Use93]. filtering [Use93]. Final [DHT89, Uni96, Ede92, Joh86c, Uni91b]. finalist [New91, New95]. Finalists [Lew93]. finance [Zen99]. Financial [Bro91b, Fin94, Ano96n, MSCxx]. find [Ano98f]. Finding [BBD^{+}08, GM93b, SW10a, CP92c, GS87c]. Findings [KF95]. Fine [BBK^{+}08, BL91, CWW94, INKN01, MKSF96, SFL^{+}94, ZL97, FMD07, TS91]. Fine-grain [CWW94, SFL^{+}94]. Fine-Grained [INKN01, ZL97, BBK^{+}08, BL91]. Fingering [BCM90]. Finite [Ano94b, BV93, CSSY92, FBH93, Glo84, Glo89, JM90, KLY94, LM93, LC97b, Mac92, Nag94, NBGS96, OD01, RDZ93, SMFG85, Sha94b, SSLR90, TD90, TOWC15, Ano95v, Arb92, CH89b, Che94a, CRA10, DP90, Gri86, Hea91, JM89b, Mac96c, PSG03, Ram86, TR86, Vaf88, Van95a, YYYS93, YHA93, van95b]. Finite-Difference [KLY94]. Finite-Element [BV93, RDZ93, Ano95v]. finite-volume [PSG03]. Finiten [Wat95]. Fink [Hil97, MM94a, Wen94]. Finned [MS97]. Finnish [Ale90]. Fire [LC95, Ano89e, Ano95x, Ano96j]. firm [Ano93-40]. firms [Ano95x, Ano96n, Bla97]. First [Ano90f, CL91, CCKSS90, DH93, GJS93, HKR94, Kho94, MBR05, M^{+}95, Pau08, RS94c, Tec89, SSJL94, Uni89a, Ano94-27, Bel86, BP89b, Fra94, Hor82a, Hor82b, KK85, MSPPD20, Men87, Sch22, Uni86b, Uni86a, Uni86c, Uni91a, Uni98, Ano90g, FJSP95, IEE85, Ill96]. FISHNET [KHS88]. Fitted [TK93]. Fitting [WS99]. Five [Ano93-34, Ano95v, Ano97-27]. five-year [Ano95v]. Fixational [BBL95]. Fixed [Ano94-54, GREC91]. fixed-time [GREC91]. Fixpoints [CH94]. FL [Ano94-100, DP91]. Flame [BD93b, CGM91]. Flamelet [GWG93]. Flames [GWG93, HVSB93]. flash [Per04, Per06]. FLASH3 [FKL^{+}08]. Flat [Ano97m]. Flat-panel [Ano97m]. Flattening [GF95]. FLEX [FG87]. FLEX/32 [FG87]. Flexible [FGG09, LA94, PYTL97, Pau09, Wat72]. Flexibly [SA90]. FLICC [MP92]. Flight [CCKSS90, Ano91x]. FLITE3D} [BMT96]. Floating [Bal93, Dun92, Gol91a, Gol91b, IHE^{+}00, MD88, Wic92, Ano94-122, Ano94-123, Ano97s, CBB^{+}05, Wei91, DM88a, DM88b]. Floating-Point [Dun92, Gol91a, Gol91b, IHE^{+}00, MD88, Wic92, CBB^{+}05, Wei91, DM88a, DM88b]. Floor [Qui95]. flops [Kog11]. Florence [Ano96a, Rol96]. Florida [CL91, DP91, IEE85, KK85, L^{+}93, Lim93, Tho93c, Gig94]. Flosolver [Sin94c]. Flow [AGH^{+}90, Ano94-114, AAS88, AFT97, BM96, BP96, Den80, DGT84, DB94, ER96, Fru93, GPKK82, GW93a, Gra93b, GIF^{+}12, GW93b, GWH93, HVZ94, Hai97, Hal96, Har94b, HFNP96, HK96, HS94d, Jon19, KO93a, KKDO97, KS93b, KLSC97, KMT94, KY90, KR94c, Kuw92, Kuw94, LR92, Leu96, LDMC96, MS97, MS96, MKND97, MJRS94, MFK94, MMHM93, Nag96c, NAAW97, PB94b, PC97, PC93, RE94, RRSG96, RG94, Riz94, Saa93b, SS96a, SW96, Soe94, TK93, TOWC15, VM87, VD96, VF93, VH93b, VB90, War93b, WR97, Woo96a, YMZ90, YCC97, YYK93, Amm89, Amm90, Amm92, CV88c, CRA10, DGT82, DRAB08, Gri86, LXW^{+}16, MB93, MB94b, Nee90b, Pol90, Rei85, The90b, The91, Woo93, YH92, GS92b]. Flowfield [MKG90]. Flows [Ano94b, Ano94-140, BPJ94, DLPQ94, GFM96, Ger90, Gol96, HGC94, HFT94, KY93, KO90, PPM90, PSB01, RHH96, SHZK94, Tak94, TFO94, BS91, GB22, KfGERJxx, Ram86, TR86, ZBN^{+}19]. flowsheet [Har89]. fluent [HP88b]. fluent/BFC [HP88b]. Fluid [Ano88h, Ano94-114, Ano97k, DD93, GI93, GW93a, Har94b, HGC94, Jon19, KLSC97, KK92, MKB87, MI93, MMHM93, Nag96c, NS93, Por86, PC93, RT93, Sch93b, Sim92b, Soe94, VM87, VF93, Web93, WKFFK97, Wil90a, COS89, DGL89, Ece96, HKS93, LM92, Lou92, MA85, MB93, MB94b, PEH93, PZGL91, RCB03, Woo93, Ano96u]. Fluid-Dynamical [Nag96c]. Fluidized [NCVG96]. Fluids [Glo89, L^{+}95, Gup88]. Flux [FBH93, Ull84]. fly [YH90, Yi90, BAD01]. flyer [Norxx]. FM [LC97a]. Focus [Ano97e, Cla98, Dav87, HTV88, Tay95a, Voi94]. focused [BMD^{+}20]. Focuses [Pin01]. Folding [Ess90, IMP93, XCLW93, Mil87]. Food [Hae91]. Food-web [Hae91]. Foods [KS90]. Force [Gro90, Bel92a, IHK93, RD07, Ano93-46]. forced [BJZfDA96]. Ford [Ano96u]. Forecast [BCHH94, Dic90, GJS94, SS09, VW95, Com98, DTV00]. Forecasting [Dic81, Dic82, Kau93a, WCG94, Sel95]. Forecasts [Koo97]. Forefront [DR93, GLS11, IEE95d]. Foresees [Lew94c]. forest [SKB^{+}20]. Forester. [CCKSS90]. Foreword [MH18, RT20]. forges [Fed96]. Forget [Poo96b, Poo96a]. Forging [BMCA93]. Form [AK87, FDD02]. Formal [PGK^{+}10, Roh94, Mac96a]. Formalism [CTRR93, JC94a]. Formalizing [GP91a]. Format [EBS02]. Formation [Gre91b, Pan97, Ste94b, SBSR96, Gre89b, Hun92, OMM93, Yos09]. Former [WG93b]. Forming [KD93, Ano93-37]. Forms [NJL94, Ano90n]. Formula [BGQ19]. formulated [PW23]. Formulation [Ano94-91, TYK93]. Formulations [Ano94-116]. Forsees [Lew94a]. Fortran [Ano85b, KK89c, DE84, Don85, KK89c, KBC^{+}74, LK93, Sch89b, WW92, AK87, AP87b, Ano94-69, Bli89, Can92, Chi20, DP96, Eig90a, Eig90b, FBZ92, Fah94, Fos93, FXAC94, Guz87, Guz88, HWS^{+}88, HKT92, KB88, KZ94, Mac91a, Mar92, MWO95, MR95, Pet83, Pol87d, SKP91, SLY90, hTD88, YGSB94, YKK96]. Fortran-style [SKP91]. Fortran/ANSI [KK89c]. Fortran/PVM [MWO95]. Forum [Ano97-28, MP92, Dun92, Str94, Wic92]. Forward [Bar00c, Bar00d, Bro17b]. Forwarding [KCPT95, ABP92]. Fosdick [Ano96c]. Foster [Stu03]. found [HHS01b]. Foundation [Bor94, Jan96, Web91, Ano96-37, NN87, Nat92a, Red91, San86, San90]. foundations [Gib01, Gir91]. Four [FG87, Eig91, EY91, MP91d, SWL^{+}92]. four-processor [EY91]. Fourier [NR86, CC88a, Heg96, HA91, MB97]. Fourteenth [IEE95d]. Fourth [KK89a, Ano93n, Gra94, RLKW93]. fourth-order [RLKW93]. FPGA [BCC^{+}09, BB92, CP94a, GM93a, PSS^{+}19]. FPGA-Based [BCC^{+}09]. FPGAs [BS94d, Van13]. FPL [ML95b]. FPS [Ano91f, CG86, WG82]. FPS-164 [CG86, WG82]. FPS-164/MAX [CG86]. FPU [LKFU05, Wai05]. Fractal [KT93a, SLML93]. fraction [Ano94-82]. Fractional [AH93]. Fragment [INKN01]. Framework [Ber07, Bis94d, EAMEG11, GJP96a, MCLK07, PCY^{+}19, PZS^{+}20, SE92, WCZ^{+}18, Abe91, BPD06, CH92b, CH98, EFR^{+}05, FKL^{+}08, FGM^{+}03, GZE^{+}05, GJP96b, Jéz00, KCG08, MV16, RGH17, SKB^{+}20]. Frameworks [Ano94-102, FV94, KRS13]. France [ACM88, GL90, Ham94, JPTE94, Ano96-34]. Francisco [ACM03, B^{+}95, Bel86, IEE95c, IEE94d]. Frank [Mur10, MP91a, MP91b, Mur10]. Frank-Wolfe [MP91a, MP91b]. Frank-Wolfe/gradient [MP91b]. Fredericton [BG91]. Free [BBL95, Coc01, FDD02, JC94b, DDJ98b, TK93, VLA92, Fuj99]. Free-Form [FDD02]. FreeBSD [Coc01]. Freeway [HK96, Sug96, WMR96, Gou90]. Freeways [Hal96]. French [Ano86a, Ano96j]. Frequency [AM93b, CBT91, Ano96w]. Fresh [Ano95t]. Friction [Ano92r, KB97]. Fridge [Ano95d]. Friendly [ATSA93, OS93]. Fringe [DR93]. Front [Ano17, GF90, Van93]. Frontal [GF90, CDS98, DS96b, ZMDS96]. Frontier [Ano92r, GIBGA93, Sch22]. Frontiers [AB94, Cha94a, EP 97, Isk96, Met86b, Met86a, Gra93c, Gra94]. Fronts [Gar99]. frozen [Bro16]. fsck [BE88, Bin88]. Fuel [KRVJ94, MTK93, MA97, Mon93, NW97, PP93, SS94, SAGS93, Sol93]. Fuels [VA94]. FUJITSU [HL88a, AHOK02, BHS^{+}02, BCHJ94, DTV00, DH86b, Heg96, IU87, LMM85b, LMM85a, LMM86, MHP84, NRN00, R^{+}00, SWL^{+}91, SWJ95, SB96, SE98, Uch86, WLKI95]. Fukuoka [Ano91q]. Fukushima [M^{+}95]. Full [DLPQ94, FT93a, KRJ93, Gep01]. Full-Disk [KRJ93]. Full-Time [FT93a]. Fully [HR94, DS86b]. fun [Faz87, Mur10]. Function [GGBR95, Lie93, MBN93, RMPW93, Sal95, WG91, XMR92, Cyb89a, Hus86b, HKP88, KMB09, NNS^{+}90, ZAS94]. Functional [DAF^{+}90, Fri95, Gir91, HTV88, KKKP93, KF95, KKPR93, Mah94c, PT92, RSRG95, Sei94, SLML93, SB94d, BH92a, EFH^{+}00, GP91a, Gok89, SKP91, SZ11]. Functional-link [PT92]. functionalities [PT92]. Functionality [Wea97]. Functionally [Ano92h]. Functions [CP94a, HM93c, Lan94, NH91, OIY91, Car94b, EFG^{+}05, SA10a]. Fundamental [MR90b, Sah95]. Fundamentals [Mag10, PATT12, WP94]. Funding [Ano88r, Ano92l, Ano93l, Ano97z, Ano86a, Win02]. Funds [Coc01]. Further [Tem89a, Tem89b, Ano81, Ano92-45]. Fusion [DH93, YMT93, Ano94s]. Future [Ano92m, Ano93p, Ano95-35, Ano97k, Bos94a, Chi00, DSSS05, Els02, Gal93, GA84, Gre91a, Hin93, HF93, IBBA20, KB19, Lew94a, Lew94c, LEMS95, MSK^{+}02, Oya02, Pay97, RG94, RSB94, VSM^{+}07a, VSM^{+}07b, Wat93, Ano93d, Ano94-119, Ano96w, Hey90, Hol84, Kon91b, KAMB19, Mye92b]. Fuzzy [BJ93, CJ93, CCSS98, DGJG93, HRG93, HHGS93, Her95, Rol96, VRSG93, Ano96a]. Fuzzy-Logic [DGJG93]. FX [ASM86, Ano87e, DD90, GB90, WSL88]. FX/8 [ASM86, Ano87e, WSL88]. FX/80 [DD90, GB90]. FY [Ano93t]. FY1989 [Ano88r]. FY90 [Nat91a]. FY91 [Nat91a].

G [GLS11]. GaAs

[Cha94b, KH87, MS84, Wat93]. GAIA [Yi11]. Gain [Sch95b, Ano97-27]. Gaining [Buz84]. Gains [Ano93m, Ano92w, Hsi91]. Gaithersburg [Uni91a]. Galaxies [Ste94b]. galaxy [CKT21]. Galley [NK96]. Gallium [Ano94-55, Bac88, Dey95, FB91b, Zho88]. Gallium-Arsenide [FB91b]. gamble [Cal11b]. game [Sne94a, Sne94b]. Games [Coc02a, Coc02b]. Gaming [Ros09]. Gamma [BKT94, SVML95, FA93]. Gamma-Camera [SVML95]. Gamma-Ray [BKT94]. gamut [Gro92c]. Gap [HS94a, SH93, SS94, SH94b]. Gara [War10]. Garbage [Ano94-46, GK92, AP87a]. Garmisch [SEA84]. Garmisch-Partenkirchen [SEA84]. Garry [Eva97]. Gas [KY93, KK96a, KR94c, MKG90, Nor97b, Ris94, Tho93c, Ano02a, Ano02b, BRL^{+}20, COT21, WQS92]. gas-liquid [WQS92]. Gate [Cha94b, Sch90b]. Gated [TP95]. gateway [OGO^{+}20]. Gatlingburg [SJD96]. Gauge [Dec90, GAW96a, GAW96b, ALN^{+}01, KM85, MHP84, MSTK93]. Gauss [Ano94-92]. GAUSSIAN [GYL00, CG87, Fox90a, Gan86, Pap92, Rag94]. Gaze [BBL95]. Gaze-Free [BBL95]. Gbit/ [Ano94-135]. Gbit/sec [CWLT97]. GBP [Ano00a]. gcs [Dra91b]. Geert [MM94a]. Gel [HPLC93, SBY93]. Gell [Mur10]. Gell-Man [Mur10]. Gemini [SWS^{+}12]. GenBank [Kar93]. Gene [Bas95a, Bas95b, BGS^{+}12, BM93b, BCK13, LS93c, UEGM93, ABC^{+}05, ABB^{+}13, AAC^{+}05, AUW08, ADG^{+}05, BSJ^{+}13, BGH^{+}05, BBK^{+}08, BHD^{+}05, BJV^{+}16, CCD^{+}13, CBB^{+}05, CP13, CEH^{+}12, CSFS00, CRA10, CKL^{+}13, CNC^{+}08, CBC^{+}05, CHT^{+}13, DT08, DLJ^{+}08, EO13, EMS11, EFR^{+}05, EWS^{+}13, FKL^{+}08, FGM^{+}03, GBC^{+}05, GS06, GBB^{+}05, HBB^{+}05, HOF^{+}12, IBM01a, IBC^{+}11, IBP^{+}05, KBG^{+}13, KBVH14, KHV11, KHZ^{+}08, LKFU05, LM13, MSW^{+}05, Mor01, MAA^{+}05, MSA^{+}07, OBB^{+}05, OWG^{+}13, PMS^{+}08, RGL^{+}15, RIB^{+}13, SWG06, SAB^{+}05, SCG^{+}13b, SPP^{+}05, IBM13a, IBM13b, IBM13c, War03, War10, War00, WAB^{+}05, ZYL^{+}16, IBM01b]. Gene/L [ABC^{+}05, AAC^{+}05, ADG^{+}05, BGH^{+}05, BBK^{+}08, BHD^{+}05, CBB^{+}05, CNC^{+}08, CBC^{+}05, DT08, DLJ^{+}08, EMS11, EFR^{+}05, FKL^{+}08, GBC^{+}05, GS06, GBB^{+}05, HBB^{+}05, IBP^{+}05, KHZ^{+}08, LKFU05, MSW^{+}05, MAA^{+}05, MSA^{+}07, OBB^{+}05, PMS^{+}08, SAB^{+}05, WAB^{+}05]. Gene/P [IBM08, AUW08, CRA10, KHV11, RGL^{+}15]. Gene/Q [BCK13, ABB^{+}13, BSJ^{+}13, BJV^{+}16, CCD^{+}13, CP13, CEH^{+}12, CKL^{+}13, CHT^{+}13, EO13, EWS^{+}13, HOF^{+}12, KBVH14, LM13, OWG^{+}13, RIB^{+}13, SCG^{+}13b, IBM13a, IBM13c, ZYL^{+}16]. Genecrunch [SS96b]. GeneID [KGS93]. GenEng [Kar93]. General [AHP97, ADLL01, ES92, JML95, MKSF96, OSKO95, YFOT93, Abe91, CGLY96, CGLxx, Chexx, CH92b, CCC^{+}89, DY90, Fan87, GSZ91, Gup88, Lee86, SZ89, Aus93]. general-purpose [Lee86]. generalization [LMD98]. Generalized [Ano94-56, Ede94b, LPNRJ94, MTK93, XL94, YOY97, Bra92, GKL^{+}87, HCL88, Ked92, MKfDA96]. Generate [Bar01]. Generated [Fru93, Rue92, Vro94]. generates [MS84]. generating [TZY88]. Generation [Ano88b, Ano94-52, Ano94-76, ACA94, BKT94, BMSP94, BS94c, Cla98, FBCB18, HM93b, IHIS91, KNS95, Kok94, KT80, Mes97a, Mes97b, Moi93, OYWK91, Spe97, SJA94, WMMC10, YG92, YYK93, ANS92, Ano95g, Ano95h, Ano02a, Ano02b, BWV^{+}17, BF92, Com92a, Gep00, Gha84, Lan92, Lil89, Mas94b, MSPPD20, VSM^{+}07a, VSM^{+}07b, Wei92, YK87]. Generations [KBG^{+}13]. Generator [Ano94-64, AAS88, Ent99, IK91, KS94b, VWC96, WW92, CDG^{+}06, Gok89, Gok90a, Gut95, Mas94a, Pry94]. Generators [Alu96, And90b, Bro96, AM15, AI92, CMP94, KA92]. Generic [SL99, VVKB96, AUW08, CDG^{+}06]. Genes [FAKD93, KGS93, Ada95b]. GENESIS [TP97, Hey94]. Genetic [Ano94-57, BS94d, CJ93, CDMW94, Dip96, FC93, GCS94, LB93, PP93, PC93, Ros93a, SLML93]. Geneva [Hen97]. Genie [Bas95a, Bas95b]. Genius [Ano90h]. Genome [BGS^{+}12, CL91, Eck93, L^{+}93, Lim93, Rob93, SGIS93, SLML93, SSS92, LC93]. Genome-Wide [BGS^{+}12]. Genomes [Cha93, DDLV93]. Genomic [FAKD93, Gil93, HH93, MTHP93, San93]. Gent [DDC96]. Genviewer [MKRI93]. GeoComputations [Ano96k]. Geographical [RMO96]. geographically [BGKR99]. geography [Sha95b]. Geologic [Ano96l]. Geomagnetic [Ano95w, BBB^{+}20b]. Geomechanics [BJS02, OD01]. Geometric [PW94, BWV^{+}17]. Geometries [DLPQ94, GMBW93, KO93b, ML93, ZEC^{+}17]. Geometry [ALM93, Ano89c, FSGS93, GGW93a, IMA93, Moi93, MF93, PA93a, Rul93, Sus93, Wil91, FJSP95]. Geophysical [CU90, Chr93, SKSD94, WR97]. Geophysics [Gui08]. George [Per83]. Georgia [Ano94-108, USE00a]. Geoscience [LCP^{+}11]. Geosciences [PW05a]. German [Ano97j, Got91b]. Germany [ACM91, Ano93i, Ano93-31, Ano94a, Ano94-75, Ano97c, GH94a, GH94b, GH94c, HK93a, HWP95, SEA84, WSB96, KSW93]. Get [Bac88, BBWR90, Bur00, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Bur01a, Coc01, Coc02c, Coc02d, DDJ98a, Pau08, SDK10, Ano97j, AL92b, Mac96b, Str94]. Gets [Ano91s, Ano91g]. Getting [Ano96x, Cor87, Mye86]. GF11 [BDW85]. GFLOP [Kah92, Ano94b]. Gflop/s [Ano94b]. GFLOPS [IHE^{+}00, Ano92p, TC93, KG95]. gfortran [Chi20]. GHz [Ano03a]. Giant [Ano92-29, ALN^{+}01, Cat92b]. Giants [Ano92-30, Dey95]. Gigabit [Mes93b, Ano90i, Ber95a, NĆ02, WTC^{+}02]. Gigaflop [ACSH90, Ano88a, DH86a]. GigaRing [Sco96, Wic96]. GIS [CCSM97, Ope96, SCH94d]. git [JKL19]. Gitta [Ano96c]. give [Tri95a, Tri95b]. Given [SNS95, Chi16]. gives [Ano90a, Ano90o]. Giving [Boy15, Uni92b, Uni92a]. Glacier [CCG^{+}17]. glass [ARF12, Bro17a, SKK^{+}90]. Glass-Ceramic [SKK^{+}90]. Glen [Pin01]. Glenda [SBF94]. Glimpse [Egg94]. Glimpses [Sin18, Ano18]. Global [Ano94-46, Ano94-58, BL93, BK93, Ber95a, Con91, CSRB90, DJSP93, DS94b, EFH^{+}00, HV95, Hun94, Kah97, KGKa93, Mas95, MTHP93, Mil97a, Asi98, SLB93, SKN96, TAKB06, Tay95b, TG94, Uni92c, AISS97, AUW08, Bur91, Con90, Con88, DGG18, GBS18, HS^{+}91, Kin96, Lun94, STH^{+}98, Str94, YQTV12]. Global-Local [KGKa93]. global-scale [AUW08]. Globalizer [GBS18]. Globally [BHS^{+}02]. glossary [Ins87a, Ins87b, Ins90]. Glow [Coc02a, Coc02b, KG96]. Glueballs [Ano96m]. GMB [Jab90]. GMRES [FGG09, van95b]. GMRES-like [van95b]. GNU [Coc01]. Go [Han89, Bab94]. Goal [SBW^{+}19]. Goals [Eck93, Gil94a]. Godson [Cal11b]. Goede [Ach99]. Goes [Ano87a, Ano05, Bar01, DDJ98b, Ano97v, Cou11, Jam95, Wal09]. Going [Bar01, Chr90, HWG98, Com92a, Sch19]. Golden [Ano98d]. Goldmine [Ano97f]. GONG [Bro93]. Good [KKB92, SCG13a, Str12, Cla97, Win02]. Good-bye [Str12]. Good-enough [SCG13a]. Google [PSO12]. Gorden [Ano97b]. Gordon [KHHS95]. Govermnent [Per83]. Government [Ano89i, Bar01, Bro91b, Coc03a, Coc03b, Joh94, Spe97]. Governor [Deu86]. GPAW [RGL^{+}15]. GPGPU [MMG^{+}18, WFJ^{+}17]. GPU [AM15, BWV^{+}17, EBB^{+}20, GB22, KSP13, Roj19, SC20, SCSL12, TMT^{+}20]. GPU-Accelerated [SCSL12, BWV^{+}17]. GPU-based [Roj19]. GPUs [ZBN^{+}19]. Gradient [Ano94-45, Gre90c, JM89a, KJ85, LPV94, Man91, Meu87, MT97, Sea86, VAGRMVA90, And88, Bau88, Gib95, HVY91, JM89c, MS88, ME91, Meu89b, MP91a, MP91b, MP91c, Nat86h, SZ89, SM92, Yan92]. gradient-like [Yan92]. gradient-type [SZ89]. Gradients [FR96b, FR96c, HFH86, HFH87]. Graduate [Pan96]. GRAIL [UEGM93]. Grain [MKSF96, CWW94, GW95, Mar91, SFL^{+}94, TS91]. Grained [INKN01, ZL97, BBK^{+}08, BL91, FMD07]. Gram [JP90, Poo96a]. Grammar [JC94a].

Grammars [JC94b]. Grand [BEH^{+}94, Con91, Get15, Sha95a, Sha95b, SMDL90, Ano92l, Con90, Eck92a, Gro92c, Eck92b, Gen92]. Grant [Ano90r, Ano92l, Ano97a]. Grantees [Ano94-107]. grants [Ano95v]. Granular [Gol96, HFNP96, KK96a, KK95b, LDMC96, RHH96, RRSG96, SW96, SBSR96, VD96, Was96a]. Granularity [Cal85b, CDH84, CSY89, HC91, MRSB94]. GRAPE [Ano94-59, EFIM91, Ste94b]. GRAPE-4 [Ano94-59]. Graph [Ano94-57, Ano94-93, CJ94, GZW^{+}22, HB08, Jab90, Wal92, FP91, JG88, Sar91, Sch90b, CP13]. Graph500 [GZW^{+}22]. Graphic [Gon93, PZGL91]. Graphical [Hug93, NB92, OCVG93, RRSS93]. Graphics [AW93, Ano94-120, ABM88, Bor89, CB99, DHM^{+}88, Don93b, GS87b, IHE^{+}00, IS95, Ill96, KCY93, KP95, LMP^{+}90, LM90b, LMxx, MP91e, RL96, SDB94, SJDV09, ACM89b, And89, Ano93-27, Ano93-28, EM78, tHd90, Kha91, Lev89, Pic88, Pic89, Pic91a, Pic92, PF90, WQS92, Ano94-121, Ano96-28, SS96b, SMM88]. Graphics/Cray [CB99]. Graphoelements [Her95]. Graphs [Ano94-56, OGR95, Pel94, SA82, BP92, GS87c, GP91c, Mil87, ZLC21]. gravitating [KMN^{+}05]. Gravity [BISB96]. Great [Con91, BB91b, Con90]. Greater [Dal84]. Greatest [Ano96u]. Greece [HPP88]. greedy [SQS^{+}19]. Green [FBCB18, TKM96, A^{+}12, Ano94s, SQS^{+}19]. Green500 [cFC07]. Grenoble [JPTE94]. Grey [Now93]. Grid [Ano94-94, Ano99, BCHH94, BHW98, Fos03, GHWZ94, Kac02, MAFW08, Man91, Mes00, Moi93, Sch97c, WJC09, Wil10, A^{+}02, BTV96, BV96, CF12, Gri92, KDP^{+}14, KTN^{+}14, MNY09, M^{+}09, Mag10, WF08, CKS99, Coc02a, Coc02b, FK99, FK04, GFB^{+}03, IJM14, KBM^{+}02, KHC14, KCZJ14, MNY09, PW05b]. Grid-Based [BCHH94, GHWZ94]. grid-computing [KDP^{+}14]. GRIDHPC [FEK20]. Grids [Ano94-114, BS94e, Taf96, TMAS97, FK98, GKSR14, Nee90a, Kra01b]. Grinding [TF97]. GROMACS [AMS^{+}15, PCY^{+}19]. Grossrechner [Ano96-35]. Ground [GML90, Ver97, Chi86]. Grounding [AFT96]. Groundwater [AFT97, Ewi97, HM97, Sch97c]. Group [DLM99, JV93, MH18, Pau08, Rep92, TH19, Van91a, RT20, Sam91]. Grove [Sin94a]. Grow [Ano96-39]. Growing [Ano92j, Ano92-30]. grown [Ano95-32]. grows [WZ87]. growth [Gur94, Hsi91, Hua92, PZ89]. GS1000 [ABM88, LM90b]. GSQL [MWB95]. GTRAN2 [Vuj93]. Guaranteed [CDR96]. Guarded [KKF96]. Guest [AB01, Ken92, Rit88c, AP93, AB03, BKK11, BEGGK07, CFS95, HG02, HP04, Pan93]. Guide [Ano85b, BBC^{+}00, Fah94, MM94a, Sny99, SSBS99, Wen94, Car89b, EM91, GB91, Hil97, Nor89, Nor93b, PW05b, SKB^{+}20, Wom90]. Guided [CWLT97, HSW^{+}90, PK87]. Guided-wave [CWLT97]. Guides [BS98b]. guitars [Ano95w]. Guys [BvRS^{+}11].

H [Ahm92, Hil97, Kaz92, MM94a, Wen94, San95]. H-Bombs [San95]. H0* [GMW94]. Hackers [Yuv77]. Hacking [Coc02a, Coc02b]. Hafnian [BGQ19]. Half [GGW93a, Ano96u]. Half-Space [GGW93a]. halo [BBW19]. halo-swapping [BBW19]. halting [Ano95c]. Hamburg [Ano97c]. Hamming [AW91]. Hampton [HKS93]. hand [Ano94-63, Che90b, KK89c, SG92c, SG92d]. hand-parallelizing [KK89c]. Handbook [A^{+}12, Guz87]. handle [Ano97u, Ano01c]. Handling [TSSK94, HD88]. handmade [Ano89k]. Handover [DGJG93]. Hands [BBHL01]. Hands-On [BBHL01]. hang [Bab94]. Happened [Ano94-133]. Happens [Mur10]. Hard [PCK93, Tru88]. hardly [Sug80a]. Hardware [Ano94-110, Ano94-127, CGSG94, CSG99, Dra95, Fer86, GCY^{+}08, Gru97, HW96, Lei85, MOWW96, MSA^{+}07, NdMM09, N^{+}95, OGY91, Sch95a, STSK95, Smi96b, Uch96, Uch97, VH93a, WG91, YSK^{+}96, YMY92, And90c, Ano88s, Ano89p, Ano97y, BP90, CCG^{+}17, Gok89, GV91, Lav89, LY91a, OGY90, SWS^{+}12, UPK87, Ano97d, Ano95u, NRN00, Oya99]. Hardware-based [WG91]. Hardware-Efficient [Lei85]. Hardware/Software [CSG99, GV91]. Harmonic [DC93, DNV93]. harness [Ano97m]. Harnessing [Sun94]. Harold [Sch88a]. Harrar [CCKSS90]. hash [Kha95]. Hashing [PW94]. HASP [AHFK93]. Haus [Ano95w]. Hawaii [HBCN95]. Hawkhill [CCKSS90]. Hazards [DM93, RWCA94, Law89]. HC [Bak10]. HC-1 [Bak10]. HDTV [Ano90j]. Head [L^{+}95]. hearing [Tec89, Uni86b, Uni89a, Uni86a, Uni86c, Uni98]. Hearings [Ano88i, Uni92a, Uni92b]. Heart [Coc03a, Coc03b, Ada95b]. Heat [Cha94b, GML90, MS97, MS84, Sha89, WH94]. heating [Ha88, Ha90a]. Heats [Ano95-35]. Hedy [Bar00a, Bar00b]. height [BB91b]. Heights [Ano92-43]. held [AU87, AB94, Asp93, App96, BP93, Bro93, Bup87, Dup86, Dup87, Fra94, HS94b, IEE94b, Kho94, Kow89b, LP90, Lun94, MB93, MB94b, Uni87a, NBC92, OMM93, Pit90, RD94, TC94, Uni98, Vag88, VO93, ZAS94]. HELIOS [VMS93]. Helioseismology [KRJ93]. helium [CKT21]. Helmholtz [Sta95a, Sta95b]. Help [Ano94l, AL92b]. Helps [Ano94-48]. helvetische [Ano95m]. HEM3D [HW97, WHMA97]. HEMT [AM91, MS84]. Henry [Zor93a]. HEP [BBC^{+}89, Hay86, Kow85, LGG^{+}87, Smi81]. Here [Ano88q, Sha95b, DCG93, DCGxx, Ano89l]. Herman [Ano95w]. HERMES [LMP^{+}90]. Hermite [EGK87b, EGK89a]. Heterogene [Meu92a]. Heterogeneity [Erc88, FBCB18].

Heterogeneous

[AACK92, Bak10, BD93a, BMP93, DCG93, DCGxx, GY93a, GB22, Hen97, KK95a, MWB95, MC10, NRS95, SWSR97, SDFP93, WRW93, YAG93, YAGxx, YZL^{+}20, ALPP00, CYXL18, CLF^{+}19, GY92, KFML20, Kha95, Kim96, Kos95, Sch94a]. Heteropolymer [IMP93]. HeteroSort [YKB^{+}00]. Heuristic [CDR96]. Heuristics [ET96, WD93b]. Hewlett [Pin01]. Hexagonal [CT93b, IMA93, OMR93]. Hexagonal- [IMA93]. HI [IEE96c]. Hibbert [Ano00a]. HiCOO [YQTV12]. Hidden [DS94a, ZS94a]. HIDM [Wat91]. Hierarchical [CD92, KKB92, SS96b, Wal92, YQTV12, BMD^{+}20, BB91a, FP91, GJM86, Gal87, Gal89b, GL88, GKSR14, HY91, yHY92, HY92, Hun90, MS88, ME91, RG92, ZS94b, ZS94c]. hierarchical-memory [Gal89b]. hierarchies [GGV90]. Hierarchy [HKG90, Koc93, GJ87]. High [APK^{+}12, AMS^{+}15, Abr92, AS98, AB01, Ahm92, ABCH97, AAB06, ABBB94, ALPP00, Ano88v, Ano89h, Hig92, Ano94q, Ano94-34, Ano94-31, Ano94-51, Ano94j, Ano94-60, Ano94-61, Ano94-62, Ano94-70, Ano94-71, Ano94-104, Ano94-110, Ano94-114, Ano94-143, Ano09, BCH12, Ara97, AT93a, AT93b, BGMR96, BGS94, Bad99, BA08, BKK11, Bae01, Bak10, Bar01, BCM90, BCC^{+}08, BBC^{+}99, BGS^{+}12, BCC^{+}09, BY21, BCW20, BS98a, BEK02, Ber07, BGM^{+}11, Bhu95, BS92, BBHL01, BJS02, BEH^{+}94, BS01, BIB^{+}18, BCH^{+}22, BH17, BEGGK07, BGH^{+}02, BNSP99, CLB19, CGFT05, CCKSS90, CCYT05, CH10, CDPW94, CFS95, CB99, CMAS11, DDHK94, Dao88, DD05, DCWH07, Dem91, DKH86, Don91, DSSS05, DvdS12, DT96, Ede94b, ORS94, EGJ^{+}02, EBS02, EAGEG09, EAMEG11, EGEAH^{+}08, EHF^{+}97, EDJ^{+}10, Els02, EHG01, cFM07, FSC18]. High [For02, FJSD96, Fos96, FGKT97, FBJ^{+}94b, FGG09, FLP^{+}07, FJP94, FHM99, Gar01, GSG^{+}94, GS01, GH94b, GH94c, Gen97, Ger90, GCS20, GCY^{+}08, GYL00, GS94d, GW93c, GAB^{+}96, GB92, GMMT91, Hag01, HC99, HBKR96, HNS94, HB08, HS95c, Hof94, Hog02, HG02, HP04, Hol95, HR04, HNS^{+}10, HERC95, HHK19, HML^{+}21, IEE93b, IEE94c, IEE96b, IEE97b, IEE97a, IHE^{+}00, IH94, IBBA20, IM96, Jar12, JPMG08, Jon96, Kah94, Kah97, KT94, KMKD97, KB19, KH98, KBD97, KTG08, KWB^{+}10, KT11, KLM94, KRS13, KBLD08, LL08, Lan93, LM08, Lat16, LC97a, LW11, LLGS09, LMM^{+}21, Lid96, LCHS96, LCP^{+}11, Liu12, Lum01, MD04, Mar96, MCW98, ML97, Mec95, MB12, MC10, Mes97a, Mes97b, MLY10, MBW01, MRGR12, MS93, MBSW01, Mur06, Mur07, Mye92b, NGLPJ96, NdMM09, NGPH99, OD01]. High [OPR01, ORSS94, OT07, Pap16, Pap97, PH11, Pel93b, PZS^{+}20, PW05a, Pin01, PMP21, PW23, Pre93b, PSO12, Pro01, PMS94, Rag06, Rep92, Res01, RCB03, RG94, RS93, Sak02, SPM^{+}10, SEH98, SEH99a, SEH99b, SVML95, SW10a, SBZ^{+}08, Sch88a, Sch94b, Sch19, SKC02, SkLC^{+}03, SCSL12, Shi95, SZ11, SL99, Smi96b, Smi96c, SP12, SW10b, SJDV09, SDK98, SS09, Ste94f, IEE94d, Ste94e, Str10, Str03, SMDS15, SSGH94, SLS96, SHB^{+}13, TKI93, TFO94, TGV08, TF97, TPJ^{+}19, VWC96, Van13, VS99, Voi94, WKL^{+}16, WN10, WP94, Wil96, WG93b, WGW04, Woo05, Wri19, ZS94a, Zec93, Zen99, ZS94b, ZS94c, ZWP03, Zim96, ZW02, ZBN^{+}19, dRC94, dC94, Aba09, AGEL13, AGZ94b, A^{+}02, AB03, The90a, Ano91x, Ano93b, Ano93-39, Ano94s, Ano94-122, Ano94-123, Ano95u]. high [Ano96c, Ano03a, AB96, Ara14, AKM^{+}06, ABMN02, Bad04, Bai88, BBBC96, BMR85, BG02, BP86, BBM19, Bor93, BGKR99, BPD06, CP25, Car91, CWD^{+}08, CBB^{+}05, Che83, COT21, CDS98, CDG^{+}06, CBM^{+}05, Dam11, DHA^{+}13, DRAB08, DF12, DS86a, Don86, Don93a, DGH20, DRSS99, Duf00, EGK89b, Eig92, EM78, FEK20, FGC06, Fly66, Fox98, FP00, Fuj11, GFBR10, GMF00, GB91, GMSS^{+}11, GKSR14, GB22, GV91, GV96b, GL96a, GL96b, GL97, HW11, Hag90, HP91, HPPF94, Ipe19, IS20, KHS88, KG98, KFML20, KSM^{+}08, KHBB01, KBM^{+}02, KMB09, KCG08, KG03, KFN02, KW11, Kum91, LAdS^{+}15, Lee84, LHPG21, LAL02, LG03, LLSR02, IJS94, MBM^{+}20, Mas94a, MI01, May01, MDH^{+}16, MUKX06, MMG^{+}18, MMG^{+}00, New95, NRM^{+}09, NP90, OGO^{+}20, ODAZ15, Pan96, Pei17, PSS^{+}19, PGK^{+}10, RAG11]. high [Ram86, Ros95, SSSR20, SCV01, SKB^{+}20, SSS92, SKS04, SFC^{+}21, STH^{+}98, Sch90a, SEV^{+}09, SD92, SN96, SC04, She93, Sim00, Smi89, SHB91, SHL^{+}20, SW99, SDMS99, SS07, SGS^{+}20, Sug80a, SO91, TTD^{+}11, TR86, VdSK^{+}05, Vet12, WH94, WWJ09, WFJ^{+}17, War03, WHL93, Zor92a, Bra94, D^{+}95, Edw97, FJSP95, FXAC94, GH94a, GBK^{+}96, HS95b, KA91, Lid99, MHW94, MR95, YGSB94]. high-accuracy [GB22]. high-density [FGC06]. High-Dimension [DT96]. high-energy [BMR85]. High-Level [EAMEG11, IM96, KRS13, Rag06, GB92]. High-order [ZBN^{+}19, EGK89b]. High-Performance [APK^{+}12, Ahm92, AAB06, ABBB94, ALPP00, Ano94q, Ano94-51, Ano94-114, Ano09, Ara97, AT93a, AT93b, BGMR96, BGS94, BKK11, Bae01, Bak10, BGS^{+}12, BCC^{+}09, BS98a, BEK02, Ber07, BGM^{+}11, BBHL01, BJS02, BEH^{+}94, BS01, BIB^{+}18, BCH^{+}22, BH17, BEGGK07, BGH^{+}02, BNSP99, CGFT05, CCYT05, CH10, CDPW94, CFS95, CMAS11, DDHK94, DD05, DCWH07, DKH86, DSSS05, Ede94b, EGJ^{+}02, EBS02, EAGEG09, EAMEG11, EGEAH^{+}08, EDJ^{+}10, Els02, EHG01, For02, FJSD96, Fos96, FGKT97, FBJ^{+}94b, FGG09, FLP^{+}07, FHM99, Gar01, GSG^{+}94, Gen97, GCY^{+}08, Hag01, HC99, HNS94, HB08, Hof94, HG02, HP04, Hol95, HNS^{+}10, HHK19, HML^{+}21, IEE94c, IHE^{+}00, IH94, IBBA20, Jar12, JPMG08, Jon96, Kah94, Kah97, KMKD97, KH98, KTG08, KWB^{+}10, KT11, KRS13, KBLD08, LL08, LM08, Lat16, LLGS09, LMM^{+}21]. High-Performance [LCP^{+}11, Lum01, MCW98, ML97, MB12, MC10, Mes97a, Mes97b, MLY10, MBD99, MBW01, MRGR12, MBSW01, Mur06, Mur07, NdMM09, OD01, OPR01, OT07, Pap16, PH11, PW05a, Pin01, PMP21, PSO12, Pro01, Rep92, Res01, RS93, Sak02, SPM^{+}10, SEH98, SW10a, SBZ^{+}08, Sch88a, Sch94b, Sch19, SKC02, SkLC^{+}03, SCSL12, SZ11, Smi96b, Smi96c, SP12, SW10b, SJDV09, SDK98, SS09, Ste94e, Str10, SMDS15, SSGH94, SLS96, SHB^{+}13, TGV08, TPJ^{+}19, VS99, Voi94, WKL^{+}16, WN10, WP94, WG93b, WGW04, Woo05, ZWP03, Zim96, ZW02, AB01, Hig92, Bad99, BA08, BBC^{+}99, BY21, Bhu95, CLB19, CB99, Dem91, DvdS12, EHF^{+}97, cFM07, GH94b, GH94c, GCS20, GAB^{+}96, HBKR96, HS95c, Hog02, HR04, Lid96, LCHS96, Liu12, MD04, Mar96, Mec95, Mye92b, Pap97, PW23]. High-performance [RCB03, SEH99a, SEH99b, SL99, Ste94f, Van13, Wil96, Zen99, ZS94b, ZS94c, dRC94, dC94, Aba09, AGEL13, AGZ94b, A^{+}02, AB03, Ano91x, Ano93-39, Ano95u, Ano96c, Ano03a, AB96, Ara14, AKM^{+}06, ABMN02, Bad04, BG02, BBM19, Bor93, BGKR99, BPD06, CP25, CWD^{+}08, CBB^{+}05, COT21, CDS98, CDG^{+}06, Dam11, DHA^{+}13, DRAB08, DF12, DS86a, Don93a, DGH20, DRSS99, Duf00, Eig92, Fox98, FP00, Fuj11, GFBR10, GMF00, GMSS^{+}11, GL96a, GL96b, GL97, HP91, HPPF94, Ipe19, IS20, KHS88, KG98, KFML20, KSM^{+}08, KHBB01, KBM^{+}02, KMB09, KCG08, KG03, KFN02, KW11, LAdS^{+}15, LAL02, LG03, LLSR02, MBM^{+}20, MI01, MDH^{+}16, MUKX06, MMG^{+}00, NRM^{+}09, NP90, OGO^{+}20, ODAZ15, Pei17, PSS^{+}19, PGK^{+}10, RAG11, SCV01, SKB^{+}20, SKS04, SFC^{+}21, STH^{+}98, SEV^{+}09, SC04, Sim00]. high-performance [SHB91, SHL^{+}20, SW99, SDMS99, SS07, TTD^{+}11, VdSK^{+}05, WWJ09, WFJ^{+}17, War03, WHL93, Zor92a, Bra94, Edw97, FJSP95, GBK^{+}96, Lid99, GH94a, HS95b]. High-Performance-Computing [BCC^{+}08]. High-Precision [TKI93]. High-resolution [PMS94, LHPG21]. High-Scalable [PZS^{+}20]. High-Speed [Ano94-104, GS94d, GMMT91, TFO94, Dao88, NGPH99, Shi95, BBBC96, Che83, Fly66, GB91, SGS^{+}20, Sug80a, SO91, KA91, MHW94]. High-Speed-Grinding [TF97]. High-Tech [Bar01, CCKSS90, Ano93b, She93]. High-Temperature [FJP94, COT21, WH94]. High-Throughput [BCW20, FSC18, HHK19]. Highconverter [AK93]. Higher [Cox88, ML95a, SE92, DWV92, NCKMM88]. Highlights [Ano93t, Waz89, Nat91b, Nat92b, Natxxc]. Highly [DO89, GW91, GHK^{+}91, NK94, RS94c, SO95, VH93a, WKFFK97, YFOT93, BWHS18, Gok91, MCH91, Wat72, WCHK91]. Highway [Jan96]. Highways [Cze93]. Hijacks [Pau08]. Hill [Bel86, SL88]. Hilton [IEE90, L^{+}95]. HiPER [MCW98]. HiPER-P [MCW98]. HIPPI [Kum94, KNWB93, TF95, VDK91]. HIRLAM [WC93]. Hist [Gar01]. Historians [BF91]. historical [Asp93]. History [Bra91b, CCKSS90, Leu96, SR93a]. Hit [Ano92z, Ano93o, Ano95-30]. HITACHI [INKN01, Ano03a, DH86b, IAKH92, Kah92, KISY94, LMM85b, LMM85a, LMM86]. hits [Gep01]. hitting [Ano91h]. HLRZ [HK93a, HWP95, Att96]. Hnet [SHB91]. hoard [Ano89i]. hoc [YFY^{+}13]. holding [Ano94-63]. Hole [Ano92n, KKDO97]. Holes [FSGS93]. holistic [SEV^{+}09]. Holland [Tru88]. holographic [Ano90a]. Holography [NNSY94, Ano91e]. holostore [Ano90e, Par90a]. Holzmann [Coc02a, Coc02b]. Home [Ano92-41, Bec01, Bro01, Ano95-32, FSC18, TAKB06]. home-grown [Ano95-32]. homegrown [Cal12]. Homogeneity [Poe95]. Homogeneous [ALMS92, SW96, TGV08, GHS86, Haw86]. Homogenization [HM93b]. Homologous [LS93c, Lu93]. homology [SNS^{+}97]. Honeycomb [RS93]. Honeywell [Ho88]. Hong [IEE94a]. Honolulu [IEE96c]. honor [RD94, Str94]. Honoring [GSB95]. Hook [Ede94b]. Hop [DG95]. hopes [Sin08b]. Hopping [BWO96, Nag96c]. Hord [Ano94p]. Horizon [Car88, Kue87, TS88, Pit89, CC88a, Dra88, Gle88, Kop88, KS88]. Horizons [Cer95]. Horizontal [Koh96, Saa93b, SBSR96, PB87]. horns [Ano95l]. Horovod [KSM^{+}19]. Horse [Kop88]. Hospital [Ano91g]. Host [Kum94, Vol89]. Hot [ZLC21, Gle91, YTL87]. hot-spot [YTL87]. Hotel [ANS92, Bel86, IEE95b]. hots [Ano96-34]. House [Bro91b, Uni92b, Uni86b, Uni89a, Uni86a, Uni86c, Uni92a, Uni98, Ano88i, Ano95l]. Householder [Ber90a, CB89, MM94c, MM94d]. Howard [Per87]. HP [Ano93m, Ano97j]. HP/UX [Ano93m]. HPC [Ano97j, IEE97b, KBVH14, MJRS94, Stu97, Wac92]. HPC-Enabled [Stu97]. HPC-initiative [Wac92]. HPC2002 [Ano03a, EGJ^{+}02]. HPCC [Ano94-60, Ano95v, GS06, SCG^{+}08, Voi94]. HPCC/IITA [Voi94]. HPCG [AYL^{+}18, KB18]. HPCN [DSZ96, Lid96, LCHS96]. HPF [Ano94-69, AHOK02, PSG03, SZG95, SIDH95, SVD96, Str94, TBC94, TCF94]. HPF/Fortran [Ano94-69]. HPF/JA [AHOK02]. HPGMG [KB18]. HTG [GP91b]. HTMT [Gar01]. Huge [Com92a, TOWC15, Ano89i, Poo96a]. Hughes [CCKSS90]. Human [AKDM93, AHFK93, Ano91k, Ano92h, Ano92-38, ATSA93, Bar00a, Bar00b, CL91, CCKSS90, DP91, KF95, MH95, OS93, Sal95, SSS92, EFH^{+}00, KSM^{+}08, KWW92, Eck93, LC93]. Human-Friendly [ATSA93]. Humans [Ada95b, Ano96z]. Hundred [Tec89, Uni92b, Uni89a, Uni92a, Uni98, Bas95b]. Hungarian [Kac02]. Hungary [VV95]. Hurricane [GP06, Str11]. Hut [Ano94-116]. HVDC [GML90, LMH90]. Hybrid [AJ97, BS87b, CGSG94, EBB^{+}20, FBA93, KHV11, LB96, Ris94, Sch95b, SJPS96, GKS09, GKSR14, GB22, KB18, Luc91, RCS21, SG92d, SB18, TS91, WT11, WT13, ZBN^{+}19]. hybrid-node [ZBN^{+}19]. Hybridized [LD93a]. Hydraulic [KO93b]. Hydraulics [ANS92, Asl91b, Woo92, Woo94]. Hydrocarbon [CFV^{+}90, War93b]. hydrocarbons [CNC^{+}08]. Hydrodynamic [BOS97, CKT21, HW97, MS93, WH93, WHMA97, DP90]. Hydrogen [CKT21]. Hydrogen-helium [CKT21]. Hydrolysis [JBI91]. Hydrophobic [ZL97]. Hydropneumatic [SQM94]. Hynes [Ano88w]. Hyperbolic [CB89, GW93a]. Hypercube [Ano94-43, AM93c, CJHH94, HL96, HS95a, HMS^{+}86a, HMS^{+}86b, NB94, yHYZ87, ZS94b, ZS94c, JHZ95]. Hypercubes [PDR91, WK95, Fid90, HM93a]. Hyperplane [Ano94-50]. Hypersonic [HVZ94]. Hypersphere [AM93c]. Hypertasking [Bab90]. Hypertexi [Ros93a]. Hypertext [SSS94]. Hypertree [LTD^{+}93]. Hypothetical [Pet97].

I-Caching [MM93b]. I-WAY [Fos96]. i.e [Nee90a]. I/O [Ano90e, Ano94-38, Ano94-41, Ano94-104, BBH95, BIB^{+}18, CP94b, Fei94, FCBH95b, FCBH95a, GS94d, HNS94, Hic18, May01, MS94d, NNS^{+}90, TGL96, YYW^{+}20, Zec93, dRC94, dC94, Par90a]. i860 [Ano91p, BBWR90, Fri91, Int91, KM89, VVKB96]. i860-based [BBWR90]. IA [SCV01]. IA-64 [SCV01]. IABG [Har90]. IADM [RFS87]. Ian [Stu03]. IASTED [Ham94]. IBM [Age05, Ano96-35, War03, ABB^{+}13, Ano94-111, Ano95-32, Ano02a, Ano02b, Ano03a, Bar00c, Bar00d, BIRB93, Ber95b, BSJ^{+}13, BBD92, BBK^{+}08, BCK13, CCD^{+}13, CP13, CEH^{+}12, Che89b, Chr93, CRA10, Coc03a, Coc03b, CKL^{+}13, CNC^{+}08, Con11, CHT^{+}13, DD90, DT08, DLJ^{+}08, EO13, Erw84, ET96, EWS^{+}13, Eva97, FKL^{+}08, Gar99, GYL00, Gui05, Hak89, HJZ94, HOF^{+}12, HMS93, HAG^{+}13, IS20, JHZ95, Kha91, KHZ^{+}08, MP84, MSAD91, Nai94, NPS^{+}20, OWG^{+}13, PMS^{+}08, PBDM93, PMS94, RAG11, RMM20, RIB^{+}13, SS10, SCG^{+}13b, TGL96, Wai05, WOK^{+}00, Mob12, Kaz92]. IBM-ACS [Con11]. IBM-PC [MP84]. IC [FNP^{+}84, PBK96]. ICAP [HGS88]. ICCG [RG92]. ickp [PL94]. ICON [WCZ^{+}18]. ICON-MIC [WCZ^{+}18]. Icosahedral [Ano94-89, KTN^{+}14]. ICS [KK88]. ICSPAT [Ano93n]. Idaho [Str94]. Ideal [ACG^{+}90]. IDEF [BJH97]. Identification [AGD93, FAKD93, FBJ94a, Ram94, RP94, SC97]. Idiom [PE95]. Idling [Kra20]. IEEE [Cha94a, FJSP95, IEE94b, IEE95b, IEE95a, IEE95d, IEE94d, SR93b, Wuo94, ACM95c, ACMxx, Ano85b, Ano87d, Ano88p, Ano93b, Ano96b, Ano97b, Bel86, CCKSS90, CKS99, Pin01, Ano17]. if [Ano97u, Bab94, CK92c]. IFS [Den93, DTV00, GJS93]. Ignition [BCW93, BSB93, GMBW93, CZRB93]. Igtlib [Dec90]. IH [RAG11]. II [AB94, ABD92, ABMN02, AK93, BV96, Fin82, Mes97b, NAAW97, Per83, USE90, VSM^{+}07b, WGR93, YR93]. III [GE96, AVS93, Mir92, Zor93a]. IIs [Ano97m]. IITA [Voi94]. IJCNN [IEE93a]. IKBS [Ano86a]. Ill [LHLM95, SA10a].

Ill-Posed [LHLM95, SA10a]. Illiac

[Hor82a, Hor82b]. Illinois

[Ano22, Ano97z, Dau97, Goo88]. Illustrated [CCKSS90]. I’m [HWG98]. IMACS [LLR93b, LCV90a, LCV90b, VAS82]. Image [Ano92-38, Ano94-143, Ara97, BMSD94, CJHH94, EMS11, EH97c, Gan94b, Gol99, GS89d, HCL94, IEE94a, IJM14, JB90, OYWK91, PRSS94, SVML95, Sch93a, Sch19, SKSD94, WN10, WGR93, Y^{+}92, ZM94, Ano90a, BÇG14, HCPS95, LHPG21, OMA^{+}96, SA90, Wes89, RRSS93]. imagery [Gig94]. Images [CDA94, Hes90, OLLG96, BR95]. Imaginary [ARW93b, Cra96, Arn88]. Imagine [Ano94-63]. Imaging [Ano94-136, Egg94, JBWB97, Pet97, Ano91t, Ano95w, CNC^{+}08, JD95, KFF93b, LM13, Nat85, YW94]. IMB [SCG^{+}08]. Imitations [Pet97]. Immersion [JLC98]. Immersive [Coc01]. Impact [Ano96t, FNP^{+}84, FBCB18, Glo84, GF90, Her89, HPLT01, LC90, Nat88b, Pic91b, Pou88, RK22, Sma93, TD90, CSY89, CV92a, CBLS13, Duf00, Gal87, WAD^{+}89a, yHY92, MT13, Nat90, NDLV88, Pol88a, Pol88b, Sta88, Uni92c, Wil88b, WAD^{+}89b]. Impacts [YYW^{+}20, Str94]. imperfect [Gib01]. imperil [Ano95l]. Implement [HCL94, CDG^{+}06]. Implementation [ASSW93, AHP97, AKG87, Ano94-31, Ano94-45, Ano94-64, Ano94-74, AC84a, AC84b, AT93a, AT93b, Bak93, BAT99, BE93b, BS92, BCH^{+}93, BCHJ94, Cal96, CLF^{+}19, CGHL94, CF94, CBT91, DL92, DJSP93, DD99, DLMW95, Dra95, EFR^{+}05, FG87, FHKT97, GG96, HLDS95, HGS88, HF94, Hof93, HE98, Jab93, KS95, KMB09, KNYT95, LO96, LLDF95, MWB95, Meh94, MS94b, MT97, MS93, NR86, PBDM93, PR94a, Pry94, Sar91, SL93, SF93a, SWJ95, SSBS99, SO95, TCJS93, Tem88, TW92, Tsu91, Tze88, VPGG01, Wic96, WC93, YW94, AAC^{+}05, And88, Ara14, Bar88, Cyr86, GV96b, GL96a, GL96b, GL97, Ho91, Hol90b, HLTZ93, Kan15, KNHN16, LFJ^{+}20, Mik89, MHP84, PSG03, Ram86, RCS21, Sam85, SO91, TCM95, TR86, TMHH95, ZCPT00]. Implementations [AABK95, BHEG94, DL90, MRAR95, Saa93a, Sim92b, SB94d, CGL92, Ece96, Gir91, PEH93, Sul91, WT11]. Implemented [KH87, PK94, PK89, Whe83]. Implementing [AGEL13, AGK^{+}87, Ano94-51, DH86a, GL93a, GL89, Mac96c, PCM84, SHMH97, WLKI95, WCZ^{+}18, Wij89a]. Implications [GS94d, MG95, MK07, XOZ^{+}20, Kel85]. Implicit [AJ93, BAM93, HFH86, HFH87]. Importance [MBN93, CLmWH91]. Important [Pet89b]. imprecise [Pal15]. Improve [CT93a, CB02, MF97, Ano94-132, Ano95w, dCCF01]. Improved [Ano90k, Ano94u, Ano94-65, Ano97m, BB98, GW93a, IMA93, VR94, Ano97-29, CV91a, EB18, Inf86, Nag90]. Improvement [BKT94, IHSK93, WLH00, JP90, NNS^{+}90, Por89]. Improvements [AF97, CTRR93]. Improves [Jon19]. Improving [Ano94-66, BJV^{+}16, FT96a, Hic18, LCD97, MTLL94, OGR95, ZYL^{+}16, SL92]. IMS [HMSS87]. In-Core [MTK93, PP93]. In-Cycle [RCK97]. In-Cylinder [AGH^{+}90, YYK93]. Inaugural [Pin99]. Inclined [BISB96, DSB96]. Included [TD90]. including [Rol97]. incomplete [Eij90a, Eij90b, Eij91]. Incomprehensible [Vui93]. Incompressible [DLPQ94, Glo89, HGC94, ZBN^{+}19]. Incorporating [WFJ^{+}17]. Increase [Ano88r, Jan96]. Increased [INKN01, Dak90]. Incredible [Mur10]. Incremental [Ano94-57, Ano94-93]. indefinite [Lou90a, Nan86, Saa88]. Independent [Ano94-141, CGW05, NMS93, OGR95, PP92b]. Independent-set [CGW05]. Index [Ano90c, EBS88, VV94, CCKSS90].

Indexing [VV94]. India [Bal94, IEE97a, M^{+}94, Pil93, Ano86a, Mah94a, Pra95, SR94]. Indian [Ano96o, DB94, SB94a]. Indiana [KWW92]. Indices [SYG94]. Indigenous [Nee94]. Indirect [OM91, UT91]. Indirections [UZ95]. indispensable [Mil88a]. Individual [Ano93o]. Indoor [WvTB^{+}07]. Inductances [NBGS96]. induction [AH90]. Industrial [BGIM90, CK92a, GSG^{+}94, Gin82, KFB91, Ker94, LMP^{+}90, Nat86d, Ano95g, De 96, EKTB99, Gra93a]. Industries [CT93a, FD93, Ano93i, Ano93-31, Ano94a, Ano94-75, Ano96a, Asp93, Rol96]. Industry [Ano88q, Ano94-67, Ano95-33, Ano96p, Cla97, DHLS97, Del97, DAF^{+}90, Fei05, Fry97, Gal93, Gib93, IH94, Jan96, Mar86, Mar88b, New93, PC94b, SHMR93, Sne94a, Spe97, Tak94, VN04, Afu90, Bla97, Bro91a, Bro91b, CDO90, EHF^{+}97, Els89, EM94b, Fed87a, Fed87b, HG88, Hol84, LPC^{+}95, Rol97, Tec89, Ste85, Uni92b, Uni89a, Uni92a, VVH95]. Industry-Government [Spe97]. Inelastic [Gie96, CGLY96, CGLxx, Chexx]. Inexact [FFM95]. Infantile [Bar00a, Bar00b]. Inference [Ano94-95, ML95a, Sch95b]. Inferior [BPL95]. InfiniBand [KBVH14]. Infinite [GGW93b, GNJW93, JWG93]. Influence [BK97, BSB93, Ede94a, HPS88, JR94, KZ94, Mac91b, RMPW93, GJ87]. Influenced [HV95, JC94d]. InfoMall [FBJ^{+}94b]. inform [DMW87]. Informal [Pic91b, Sun94, Ins87a, Ins87b, Ins90]. informatici [LP90]. Informatics [Ham94, Rob93]. Information [Ano88p, Ano92-47, Ano94-60, Ber90b, BKM93, Bro91b, Fos03, GL93b, GP93c, IEE95a, KG94, MP92, Mye96, Pug94, RWCA94, Sch95b, Sch94b, SG94a, Sub94, Sun94, TF94, Vro94, Zag82, AGP96, Ano96n, Bur93, Bur94b, IEE95c, IEE95d, IEE97b, Jam95, Kah91, Lun94, RMO96, Sha87, Zyg93]. Informative [PG93]. InfoWorld’s [Ano93d]. infrared [And90a]. Infrastructure [Ano94-60, BEK02, IH94, Bis93, BFS11, FK99, FK04, HBB^{+}05, HCPS95, Lun94, WJC09]. infrastructures [IEE95d]. inhaled [MKHY95]. inherent [SK92]. Inherently [AA93]. Inhibitor [HGB90]. Inhibitors [VM94]. Inhibitory [KW95]. Initial [GCP90, HCV97]. Initiative [Ano94-73, Coc01, Ano93b, Cat92b, Wac92, Pel93b]. Initio [Las92, BS00, BBK^{+}08]. Injected [MA97]. Injection [JJYL94, MHE97, Chi86]. injection-pumping [Chi86]. Inn [IEE93b]. Inner [BJS02]. Inner/Outer [BJS02]. Innovation [Mil17]. Innovative [ORS94, ORSS94, SS95, SCV01, Ano96u]. Innovator [Stu03]. Input [Mil90, Mil91, ODAZ15]. Input/output [Mil90, Mil91]. INRESB [CGLY96, CGLxx, Chexx]. INRESB-3D-SUPII [CGLY96, CGLxx, Chexx]. insatiable [Ano96n]. Insidious [Ano88y]. Insight [BCH12, Buz84, TC93]. Insights [Ano95r, Ano96z, HGB90]. Inspection [DS94c, KRVJ94, KT94, Leg94]. Instability [Bot96, NCVG96]. Install [Ano88l]. installation [HLA^{+}18, Lie20]. installations [Ano96-36]. installed [Ano90n]. Installs [Ano87a]. Institute [Ano94-107, B^{+}89, DP91, HS94b, Min88, Ano95-41, Duk91, Ste90]. Instruction [Ano96i, Bro86, HMNN91, KA93a, RF93, Vaj91, WS84b, WS84a, WS84c, Dra88, FMT91, LLDF95, PJ90]. Instruction-level [RF93, Vaj91]. Instructions [UT91, TZY88]. Instructor [LJ97]. Instrumentation [Bli89, GP90, GM87, HMC94, SAB^{+}05]. instrumented [Rau91, SSRL91]. Intake [Hai97, LYKM97, OGOR97, WJ94]. INTBIS [HKSY94, HSKY95]. Integer [Ano94y, Bue86b, Ano97j, ARW92, BW88, FB91a, Ked94, RW89]. integer-sorting [Ano97j]. integers [BtR95]. Integral [DD93, HS^{+}91, MRAR95, GL88, LG87]. integral-based [LG87]. Integrals [EBS88, SB81, SB82a, FSY88, RWL^{+}98, SB82b]. Integrate [LC95]. Integrate-and-Fire [LC95]. Integrated [EFPSS93, KWH94, KSTF94, NW97, OMA^{+}96, PDR91, PL91a, PL91b, RL90b, SB94b, BHM94b, GV91, Gua88c, HCD^{+}18, Hod87, JG88, PZGL91, YKY90]. Integrates [FXAC94]. Integrating [Ano94-102, Bae01, DPS97, FSC18, HB96, LCH87, PS96, YSL97, OGO^{+}20]. Integration [Ano94-68, ATSA93, CV93, DGBE96, DDLV93, Fri95, Gil93, JG99, Lan93, Leg94, RSB94, Taf96, Par90b, And89]. INTEL [Eck93, AABK95, Ano90n, Ano91p, Ano94g, Ano94h, Ano94b, Ano94-117, Ano97j, Ano01c, Ano03a, Bab90, Chi20, CM95, Cou13, DAC^{+}18, Fri91, GFM96, GGG^{+}98, Gro93, Gut95, GAW96a, GAW96b, HL96, Hay89, HCD^{+}18, HFCM98, Hoc94, HL93b, IAIK92, JHZ95, KM89, KRJ93, LHPG21, Mas93, MH98, Mit98, Per89, RLKW93, SNS^{+}97, SZG95, TGL96, TFB94a, TFB94b, WMKS96, Wat95]. Intel/Paragon [Wat95]. Intelligence [Ano93b, SEA84, Sri94, Tho93a, CC88b, HD89]. Intelligent [ATSA93, CSPJ97, Cze93, Dil93, FT93a, FNK93, GY93b, GL93b, GBK^{+}96, KA93a, KG94, KTNM93, Sri94, Chu87]. Intensive [Ano94-38, GCY^{+}08, MPT12, MH96, SMM17, SC04, UL89, YFY^{+}13]. Inter [DL96, Hae91]. Inter-processor [Hae91]. Inter-Vector-Conflicts [DL96]. Interaction [Dip96, RE94, BB87, LY90c]. Interactions [ML95a]. Interactive [Ano94-28, Bli91, FCGG90, FK93, HED93, LMP^{+}90, RRSS93, SGB91, TC21, WKFFK97, Wil92a, YF95, Ano89h, Kha91, MB97, SCH94d, WvTB^{+}07, WWTE92, Wil90b]. InterCom [BGPS94]. Interconnect [JS95, KGB^{+}96, BHM94a, SWS^{+}12]. Interconnection [CJ94, CEH^{+}12, CTD^{+}16, GVBC95, GS94e, MCW98, OA94, Sie90, TVT^{+}16, Abr90, ABC^{+}05, Fid91, yHYZ87, Tur89, TYZ85, Tze86, TYZ88]. Interconnections [DG95, Goo97, Wat92a]. Interconnects [ADG^{+}08, CG96, Clo96, PLC^{+}19, Wic96]. Intercooler [BPW97]. Interdependence [Mit96]. Interest [DAF^{+}90]. Interface [Ano94-41, Ano94-74, Ano94-124, Ano97o, BB98, FBA93, Gle93, Gon93, Kar93, Kum94, PR94a, RRSS93, SP94, TYK93, Y^{+}92, YGSB94, Ano94-135, Ano95v, Bru90b, BHM94b, DLM99, FTT97, GG88, KDP^{+}14, KfGERJxx, Nag90, SWJ95, Suz89]. Interfaces [Smi95, Smi96b, Smi96c]. interference [OL86]. Interferometric [SPGD98, YWD94, YWDxx]. interim [Sar90]. Interior [Ano94-91, Bro91d, LB94b, SD88]. Interlanguage [Mac91a]. Interleaved [SL92]. interlinked [Kra88]. intermediate [GP91b]. intermediates [RLKW93]. Internal [Ano90l, GWG93, HK97, Tak94]. International [ACM88, ACM90, ACM91, ACM92b, ACM92a, ACM93, ACM94, ACM95a, ACM95b, ACM96, ACM97, ACM03, ANS92, AGP96, Ano88w, Ano90f, Ano91m, Ano91q, Ano92g, Ano92y, Ano93g, Ano93i, Ano93-31, Ano93n, Ano94p, Ano94-75, Ano94-108, Ano94-134, Ano96k, Ano96a, Ano97q, Ano98b, Ano98c, Ara96, Bal94, Bel86, BP93, BH92b, CL91, CBCH93, DSZ96, Elm95a, Emm84, Emm85, FJSP95, GH94a, GH94b, GH94c, GL90, Goo97, GT94, Guo94, HS^{+}91, HKR94, Ham94, HS95b, HS95c, HPP88, HBCN95, IEE85, IEE93a, IEE93b, IEE94a, IEE95b, IEE96b, IEE96c, IEE97a, Ill96, KK87, KK88, KK89a, KK93, KMG96, KSW93, KK92, LP90, Lid96, L^{+}93, Lim93, M.I87, Mar86, Mar88b, MM91a, M^{+}95, MB93, MB94b, NBC92, Pit90, PH95, Pow97, Pra95, Rol96, SN96, Sie94, Sig90a, Sig95, SJD96, IEE94d, SR93b]. International [Tho93c, Uni87c, WG93b, Wuo94, ZAS94, Zyg93, Ano93-39, BBM96, BP89b, Cal11a, Cha94a, DDC96, IEE95d, JPTE94, KK85, LCHS96, IJS94, ML95b, Pou88, Suh97, GP93c, HHK94]. Internet [Gro90, Ano95v, Ano97f, Ano97i, Bar00a, Bar00b, Bar00c, Bar00d, BB98, JNM^{+}98, Nor17, Opp95b, Pau08]. Internet-Addressing [Pau08]. INTERpack [HBCN95]. interpersonal [Ano97g]. Interpolation [AM94, Isa93, JB90, KH98, AM96, EGK87a, EGK87b, EGK89a, EGK89b, Par90c]. interpret [SM89]. Interpretation [Mas91, TC93, AH90, Arb92]. Interpreting [Ano94-69]. Interprocedural [LY88b, LY88c, Tri85, YH92, LY88a, Li89, Sch90a]. Interprocessor [BGPS94, NSP94, SKIY94, SKIY97, Abr88, Pol89]. Interrogation [HHSW93]. Interrupts [VSM96]. intersections [CBA90, Xu91]. Intersociety [HBCN95, Suh97]. Intertask [Ano94u]. Interval [Sch87d]. Interview [Hay84, Lew94a, Lew96a]. intractable [RJ13]. Intraprocedural [Li89]. Intrinsic [BS01]. introduced [WZB86]. Introducing [Ano89j, Cra92, CG86, KM89, Nat86e, Ker94]. Introduction [AP93, AB03, BKK11, BEGGK07, CFS95, Cze16, FJSD96, HW11, HG02, HP04, Ken92, M^{+}09, Nor84, Pap97, SS94, SMR10, Tor87, Ano96c, GC92, KA92, Pan93, RJ13, Rot92, SMM88, Uni87b]. intrusion [WLLZ20]. invalidation [CV88a]. Invariant [BHLST94, AT89, Bis94c, HLTZ93, Bis94b]. invented [Bas95b]. Inventing [Taf96, Faz87]. invention [Ano92-42, Rei93]. inventory [Ano88y]. Inverse [AM94, JBWB97, Phi85, EM94b]. Inverse-Distance-Weighted [AM94]. Inversion [BS97, Mis90]. Investigation [WGR93, Bro93, Wun89]. Investigations [BMP93, WM91]. Invited [AHAM93, Ask93, Bur93, DKS93, Dil93, ESMH93, Fie93, FBGM93, FNT93, FNK93, IHSK93, KJ94, NS93, OS93, PSB01, Pre93b, SH93, Smi93, TKI93, Tho93a, TP93, Ano87b]. Involvement [Ano94-66, Ewi97, Joh94, Gla93]. IofNEWT [Ano94-68]. Ionized [BPJ94]. ions [Fin82]. IOS [JS95]. Iowa [Ano87a]. IP [Ano00b]. iPSC [Ano94-117, HL96, PL94, Rot94, VSB94, YSS94, YR93]. iPSC-860 [YR93]. iPSC/860 [Ano94-117, HL96, PL94, Rot94, VSB94, YSS94]. IPSN [CLP93, CPR93]. Ipswitch [Ano95-32]. IQ [Ano03b]. IR [Ano96u]. IRDS [KTNM93]. Iron [BvRS^{+}11]. Irradiated [Mon93]. Irregular [Ano94-94, Ano94-112, Ano96e, Ano96f, CC94b, FR98, Gal96, LB94a, LB96, Sus93, TA94, GF95, LG87, Nee90a, YF98]. Irregularly [CCSR92]. IRS [BMSP94, KK93]. ISATA [Ano93-31, Ano94-75, Ano93i, Ano94a]. ISBN [Ano94p]. ISE [SDB94]. ISI [AK94]. Ising [IK91]. Islands [Max81]. isn’t [Win02]. ISODATA [Sch88b]. isolating [Ho92a]. Isovalue [SA94]. ISP [Hod87]. ISPAN [HHK94]. Israel [Por86, MA85]. ISSCC [Ano91h]. ISSIPNN [IEE94a]. Issue [AP93, Ano96e, Ano96f, Ano09, BKK11, Ber07, FR98, GMSS^{+}11, KRS13, MB12, MD94, MLY10, PW05a, RT20, TH19, Tor87, WS84a, AB03, Ano94-126, Ano95l, DF12, yFH89, Kar13, MH18, RF93, WS84b, WS84c]. Issues [Abr90, Ano94-70, Ano94-71, Ano97g, Bro91b, Dra95, GP85, GL91, Leu90, MP90, PC97, Waz89, WR97, Wil90b, AISS97, CV92a, CDS98, DRSS99, GC92, HLDS95, JD95, LR90, Lee86, Mil87, Vei85, WLN^{+}96a, WLN^{+}96b]. Issuing [HMNN91]. ISVAS [FK93]. Italy [Ano96a, De 96, D^{+}95, HS95b, HS95c, LP90, Lag89, MM91a, PH95, Pow97, Rol96, Sch97a, ACM95b, DJM94]. Itanium [Ano01b, Ano01c, SCV01]. Iterated [AH93, CF92]. Iteration [Man91, Vog93, Ske87]. Iterative [AFT96, Ano94-94, BJS02, CMF94, CCSM97, DD93, DL90, GKR91, GT91, JC94c, KLY94, LB94a, PHVJ95, Sob92, Sob93a, Vui93, WSP95, vdV91, B^{+}89, Bru91, Ho91, HEB96, KA96, Lou90a, Rob85, Sch87a, SG92c, Van95a, ZGL14]. ITT [MAT85]. IV [Hor82a, Hor82b]. IXM2 [HHT^{+}94].

J [Ano96c, Hil97, Kow86, Nor97a, Pre93a, Wen94, Car94a]. J-90 [Car94a]. J90 [PBK96, WLH00]. JA [AHOK02]. Jack [Bro91b]. Jacobi [Man91, SO95]. JAERI [AHAM93, HAAS93]. JAERIPULSE [SNK^{+}93]. Jahre [Kro92]. Jams [Nag96b]. Jan [Uni91b]. Janeiro [SN96]. January [AIA94, Ano94-107, Ano98a, CG96, Clo96, GL90, GE96, ZAS94]. Janus [BCC^{+}09]. Japan [ACM93, Ano90f, Ano90m, Ano91q, Ano92o, Her90b, KK92, MN91, NN90, Ano91i, Ano93b, Ano94-72, Ano95x, Ano96o, HHK94, IEE93a, KW92, Kah92, Kah93b, Kah93c, Kah94, Men87, M^{+}95, Oya99, Ste85, Uni92c]. Japanese [Ano92h, BEW82, Jen88, Men84, MO88, Mor92c, Nat86h, Asi98, YK87]. Java [Fox97, AISS97, Ano00c, Ano00d, Bam97, BBHL01, CSFS00, DDJ98a, FTT97, Fox98, Gar99, KHBB01, LAL02, MI01, MMG^{+}00, RCB03, SEH98, SEH99a, SEH99b, TTD^{+}11, Van97, WN10, WGW04, Woo05]. Java-based [AISS97]. JED [Mal89]. Jerk [YK94]. Jessup [Ano96c]. Jets [CAB93]. Jim [Coc01]. Job [Dow98, FR96a, KCM02a, KCM02b, TDBL13, Ano97u, Joh88, MWRK18, MSW91, SHL^{+}20, TRLD13, vL99]. Jobs [Bar01, CB02, dCCF01, LF03]. John [Joh86a, Kaz92, Joh86b]. Johnson [Kaz92]. join [Hug94, Kha95]. Joining [SSKa93]. joins [Ano92o, Dey95]. Joint [Ano94-73, Ano97q, Ano97-30, BBM96, GT94, HKR94, IEE93a, KSW93, LPLP97, RMO96, Ano03a, COS89]. Joints [Sch97b]. Jones [XMR92]. Jose [CG96, Clo96, Gra94, GE96, Hel93]. Josephson [Ano91c, GL91, Wal81]. Journal [RF93, Sup87b, IJS94, M.I87, GMSS^{+}11, KRS13]. journey [Pic92]. JPL [Jet91, Jet92]. Jr [Bro91b, MM94a]. Jr. [Hil97, Wen94]. JSME [L^{+}95]. Juelich [Hos95, MG95]. Julia [ZW02]. Julich [HK93a, HWP95, WSB96]. July [ACM88, ACM92b, ACM92a, ACM93, ACM94, ACM95a, ACM97, Ano94-134, EM94b, Fra94, IEE93b, KMG96, ME96, NBC92, Sam91, Sig95, Uni92b, Uni92a]. Jun [Suh97]. Junction [Ano91c, UR95]. June [ACM90, ACM91, ACM95b, ACM03, Fin94, AGP96, Ano88t, Ano88u, Ano94-100, Ano96a, Bal94, BG91, DSZ96, D^{+}95, Ece96, MSCxx, Goo97, HK93a, HPP88, IEE96a, Jet90, Kho94, Kow89b, LLR93b, L^{+}93, Lim93, PH95, Pow97, Rol96, Tec89, Sig90a, SJD96, Uni86b, Uni89a, Uni86a, Uni86c, TR86]. Juni [Meu89a, Meu90, Meu91, Meu92c, Meu93]. Jupiter [Str94]. Jupyter [TC21]. Just [Coc02a, Coc02b, TF95, DMW87, Mal89, MS84, Rya90, Way96].

k-ary [DT96]. Kanai [War03]. Kanazawa [HHK94]. Karachi [ZAS94]. Karin [Bro91c, Haw88]. Karlsruhe [KSW93]. Kasetsart [US01]. Kasparov [Eva97]. Kaufmann [Pre93a]. KBS [BSKJ93]. Keep [Bar01, SGIS93, Ano92-29]. keeping [Ano92-42]. keeps [Ano95a, Ano96u]. Keith [Ano00a]. KEK [NN90]. Ken [CCKSS90, Ano16]. Kendall [Rot92]. Kennedy [Ano16]. Kenneth [Ano00a, Ano97b]. Kent [Ano94p, Ano93b]. Kerberos [Coc01]. Kernel [HWS^{+}88]. Kevin [Bra94]. key [Ano94-122, Ano94-123, MP92]. Keynote [Mes93b]. Keyworth [Per83]. KFA [HK93a, HWP95]. KfK [AHSS93]. Kick [Ano91i]. Killer [Ros09]. Kilometrage [Koo97]. Kind [Sil91]. Kinematic [DDF93, VD96]. Kinetics [Gol96, HV95, Koc93, Maa93a, Lag89]. Kingdom [ML95b, OMM93, ACM94]. KISMET [Kue93]. kit [Ano95-31]. Kittyhawk [AUW08]. KIVA [YR93]. KIVA-II [YR93]. Kiwis [Smi95]. Kluwer [Ano00a, McD88]. Knapsack [MRAR95]. Kneading [YMZ90]. Knights [DAC^{+}18]. KNL [HLA^{+}18]. Knock [Ano92-39, Ano92-40, BCW93]. Know [Dun92, GP06, Gol91a, Gol91b, Tri95c, Wic92, WZB86]. Knowledge [KS94a, RC94, SBJ90, STN93, SK93b, YSL97, Bar88, Cre91, Das94]. Knowledge-Based [STN93, Bar88]. Knoxville [IEE94c]. Kock [HMS93]. Kohonen [AGD93]. Kometen [Ano97d]. kommerzielle [Sch92a]. Kong [IEE94a, IEE94a]. Konrad [Stu95]. Konrad-Zuse-Zentrum [Stu95]. Konvektionsstroemungen [Wat95]. Korea [Ano92-30]. Kosloff [Whe83]. Kosloff/Baysal [Whe83]. kph [Ano92h]. kph/Light [Ano92h]. kph/Light-Emitting [Ano92h]. Kriging [KH98]. Kruh [CCKSS90]. Krylov [BS87b, GS92a, Saa89, Saa93a]. KS [Fox97]. KSR [Ano92p, BIR94, BM93a, DFS93, ER94]. KSR-1 [DFS93, ER94]. KSR1 [BD94, Her94, NK94, Rot92]. Kuba [MSTK93]. Kuba-Moriarty [MSTK93]. Kudos [DDJ98b]. Kurzweil [Bar00a, Bar00b].

L

[Ano00a, Bro91b, Pre93a, Ano95w, ABC^{+}05, ABB^{+}03, AAC^{+}05, ADG^{+}05, BGH^{+}05, BBK^{+}08, BHD^{+}05, CBB^{+}05, CNC^{+}08, CBC^{+}05, DT08, DLJ^{+}08, EMS11, EFR^{+}05, FKL^{+}08, GBC^{+}05, GS06, GBB^{+}05, HBB^{+}05, IBP^{+}05, KCM02a, KCM02b, KHZ^{+}08, LKFU05, MSW^{+}05, MAA^{+}05, MSA^{+}07, OBB^{+}05, PMS^{+}08, SAB^{+}05, WAB^{+}05]. L. [Pre93a]. La-Grammar [JC94a]. Lab [Gil93, Str94]. Labeling [CJHH94]. Labor [Cop93, Lee94]. Laboratories [Mac91b]. Laboratory [AB94, Ano94-107, UU94, Pan97, PMS^{+}08, BBB^{+}91, CH89a, CKS99, THH81, THH82, WMBC97]. laborious [Emr89]. Labs [Hug94, Mac97b]. Lachesis [Dow98]. Lafayette [KWW92, RD94]. Lagged [Alu96, AM15, Mas94a, Mas94b]. lagged-Fibonacci [AM15, Mas94a, Mas94b]. Lagrangian [CT94]. Lahaina [HBCN95]. Lake [ANS92, Ano95-38, Isk96, BOS97, MKDY90]. Laker [Ano97b]. Lamarr [Bar00a, Bar00b]. lambda [Lee87a]. Laminar [CAB93]. Lanczos [AHP97, GG96, GZA86, HE98, LO96, Sch96]. Land [Mil97a, OLLG96]. Landau [Hil97, MM94a, Wen94]. Landing [DAC^{+}18]. Landing-based [DAC^{+}18]. Landmark [Cha93]. Langley [Gri86]. Language [Ano93-38, Ara91, AA93, CC94b, Ele93, Fos93, JAB92, Kar93, KNYT95, NB92, PCM84, Roh94, RCZ93, Tri93, Tsu91, Bec90, BCH^{+}93, CS93a, DHA^{+}13, Gok89, Gua87b, Gua88a, Han94, Joh88, Kel85, LG03, Nor17, RR99, Rob87, SK93a, Sch94a, Tur79, HAG^{+}13]. Languages [GPKK82, IKM85, JC94a, KRS13, PB90, PZA86, Zim96, Feo92, PHK88, SWS^{+}12, SMR10]. LANL [Ano95y]. LANs [MKSF96]. Lanthanide [CS94b]. LAPACK [AF97, Dem91, Don91, GB90]. Laplacian [Sat93]. laptops [AMS^{+}15]. Large [Ano96q, Ask93, BPJ94, BBC92, BBC^{+}89, Cap96, Che83, CDC^{+}87, DAKM98, Ede94a, GGG^{+}98, Gol99, GZA86, GK93, HWS^{+}88, HK93a, HFH86, HFH87, IHSK93, Iwa92, KS93b, Khe94, Khe95, Lu93, Ma99, MS97, Mar91, MR87, OS94, PZS^{+}20, PSB01, Rui91, SBJ90, SkLC^{+}03, SKK^{+}90, Sob93a, VAGRMVA90, WB85, WVBM88a, WVBM88b, YZL^{+}20, Zla01, Zor93b, van95b, BAAD^{+}97, BtR95, B^{+}89, BS90b, BJ84, BY88, CH87, Che90c, Che93b, Che89c, Che99, DSZ96, Duf90, GW95, GHdF10, GPS86, Gra92, GKL^{+}87, HRC09, HOSZ97, HY89, IU87, Jor87, Kos95, LPD^{+}11, Lee87b, LXW^{+}16, LW94, MP91a, MP91c, NNS^{+}90, NP90, Rob85, Sie90, Sob92, Sug80a, WT11, WT13, YTL87, Zor93c, vdV91, App96]. Large-Eddy [PSB01]. Large-Scale [CDC^{+}87, DAKM98, HWS^{+}88, OS94, PZS^{+}20, Rui91, SkLC^{+}03, WVBM88b, YZL^{+}20, Zla01, Ano96q, Che83, GZA86, HFH86, HFH87, MR87, WVBM88a, Che90c, Che93b, Che89c, Che99, DSZ96, Duf90, GHdF10, Gra92, Jor87, Lee87b, LXW^{+}16, NP90, Sie90, WT11, WT13, YTL87, App96]. Largest [Kra20]. Larkfield [Ano94p]. Laser [L^{+}95, Sch92b]. lasers [And10]. Last [Pou94a, Pou94b, Ano97t, Zen99]. late [DT96]. Latency [Ano94-124, CMHK92, Smi01, Ano94-135, Lil91, SSSR20]. Lateral [MFK94]. Latest [WJ94, Ano95h]. Latin [GGN20]. Lattice [AGLL98, Dec90, GAW96b, KK96a, KMG96, KR94c, MKND97, ALN^{+}01, CRA10, DM96c, KM85, MHP84, PMS94, GAW96a, KMG96]. lattice-based [DM96c]. lattice-Boltzmann [CRA10]. lattice-Boltzmann/finite [CRA10]. Lattice-Gas [KR94c]. Lattices [RMH93]. launches [Ano03a]. launching [Ano01c]. Lauritzen [Ano94-95]. Lauritzen-Spiegelhalter [Ano94-95]. Lausanne [Ano97-30]. Law [Gar01, Bar01, dRSGS16]. Lawrence [CH89a, WMBC97]. Laws [VMS93, Dum97]. Layer [BNSP99, Ano95w]. Layout [Kul94, SIKD94, BGH^{+}05]. Lead [Bel96, Dau97, Gui05]. leadership [JOK^{+}18]. Leading [Hei89]. Leads [MMRL93, MHE97]. Leakage [BMP93]. Learn [Bur94a]. Learned [Con11, MWO95, SB94c, DRB^{+}20, Gil94a]. learner [CFP20]. Learning [CCKSS90, Che93a, Die95, Eis95, GCS94, GGBR95, HS96, HSxx, KSTB94, KDBG95, MPH93, Opp95a, SR94, BRL^{+}20, EP 97, Ipe19, KMRR20, KSB^{+}19, MBM^{+}20, McA92, Roj19, SNEP14, TCM95]. Learning-centered [HS96, HSxx]. Least [OB94, Ano92-44, Duf90, GPS86, HOSZ97, Poo96a]. least-squares [Duf90, HOSZ97]. Lee [CCKSS90]. leech [Ano95w]. Leeds [Ano96k]. legacy [Ano96-42]. Legate [BLP^{+}21]. Legislation [Uni92b, Uni92a]. Length [FBB97, PC97]. Lennard [XMR92]. Lennard-Jones [XMR92]. Leonard [Bas95b]. Lessons [Con11, DRB^{+}20, MWO95, SB94c, Gil94a]. Let [DDJ98b]. Lethality [SKC02]. lets [Ano93v]. Letters [Ewa96]. Level [Ano94f, BCK13, EAMEG11, Fox89, IM96, IBC^{+}11, IMA93, Koc93, KCOP94, KRS13, Rag06, Wal92, AMS^{+}15, Ana91, Ano91h, Ano91v, Ano94-120, BSJ^{+}13, DD90, GP93a, GB92, KW11, Loo84, LM13, MAFW08, RF93, Sch90b, TTD^{+}11, TZY88, VSH90, Vaj91, YSKS95, DD99]. levels [FMT91]. Leveraging [BBW19]. Lewis [MF93]. lexically [BGS82]. lexicographic [RS94b]. LGA [Cha94b]. liabilities [ZCPT00]. Libfabric [SSSR20]. Libraries [CDPW94, IEE93c, JM93, PPG94, Bis94a, Don93a, HLJT93, MP92, STH^{+}98, TTD^{+}11]. Library [Ano87a, Ano94q, Ano94-74, BGPS94, Dec90, Dem91, Don91, EHG95, GFB^{+}03, Lay91b, RW94a, SL99, WN10, ZW02, AC91, ABMN02, Ham90, Mic90, AF97, BCHJ94]. Library-Based [Ano94-74]. Life [Che92c, Che92b, ES88, Poo96a, Str94, CCKSS90]. Lifetime [Coc01, Rit97]. Lifts [Bar00c, Bar00d]. ligand [ZEC^{+}17]. Light [Bar00c, Bar00d, Del97, Fei05, Mil88b, Ano94s, Ano02a, Ano02b]. Light-Emitting [Bar00c, Bar00d, Ano92h]. Ligure [ACM95b]. Like [Bar00a, Bar00b, Ano90o, Yan92, van95b, WB85]. Likelihood [Mis90, YOY97]. Limit [PA93a, Ber82, SA10a]. limitations [Blu92]. Limited [PS94b, VW95, WCG94, YJD93, HY92]. Limits [CCKSS90, EM94a, GB96, Moh94, RJ13, TMP94, ARF12, Bel92a]. Line [Bel93, EFPSS93, GSG^{+}94, HRG93, RW94b, TW92, Ano94-27, CKS99]. Linear [Ada93, ALPP00, Ano94-61, Ano94-93, Ano94-94, AJ93, Bea90, BCZ95, Cal81, Cal86a, Cal86b, CDH84, Che92a, CDW94, Cla96, Dem91, DS86a, Don91, Don93a, Dub87, Duf82, Duf91, Ede94a, GMW94, GT91, Hak89, HL96, HO92b, JML95, JC94c, Lan94, Ma99, MTK93, Meh94, MM94c, MM94d, NJL94, NGDH96, OS94, PS94a, PS94b, PK94, RCK97, SC92, Van91b, VF93, VAGRMVA90, WN92, WNKS96, YA93, AD88, And88, AS88, Ano87e, Ano94-109, BR95, Bra89c, BS90b, Che90b, Dav86b, DD88, DHD89, Don85, DH86a, Duf00, FJ91, GJM86, Gal87, GPS90, GMW91, Gao86, GW95, Gok90b, HOSZ97, HVY91, Ipe19, KS86a, Kor93, Lou90b, Lou92, McD85, PO88, PK89, Rob85, Saa88, Sam85, SZ89, Sch87a, SG92d, Sim00, SE98, TFB94a, TFB94b, Van95a, WHBH93, Wil90b, Yan92]. Lines [Ano03b, Ano05, TJ94]. Linguistic [Bar93b, CV93, MGA94]. Link [PDR94, Ano94-135, DFSZ88, Natxxe, PT92]. linked [KG95, TYZ89]. Linking [DLG93]. Links [Cra91, KNWB93]. Linpack [DZM^{+}13, RRMD94]. Linux [Ano01b, USE00a, USE01, Bae01, Bar01, DDJ98a, Luc01, Lum01, Ste01b, Ste02]. Linux-Basis [Ano01b]. Liquid [Gre91b, JJYL94, NCVG96, Pop97, WG93a, Gre89b, WQS92, CSB89]. Liquid-Fluidized [NCVG96]. Liquid-Nitrogen-Cooled [CSB89]. Liquids [BCCG97, Fin82]. LISP [BGS82, Cho90b, FP91, Har86, AGK^{+}87, AKG87, HP88a]. List [Ano91r, cFC07, RM96, SMDS15, Ano93b, TYZ89]. listing [CGLY96, CGLxx]. literature [Ana91].

Little [ANS92, BvRS^{+}11, MS84]. live [CK92c]. Livermore [CH89a, WMBC97, Mac91b, Mac96b, hTD88]. living [Gib01, Wal17, Wit89]. LLNL [CCD^{+}13]. Lloyd [Ano96c]. LMFBR [IHSK93, Mon93]. Load [Ano94d, Ano94-65, BBL95, CRY94, CMF94, FT94, HL95, HS93b, Kar94, KK95b, LH94, Raa97, RS94c, Tri95a, Tri95b, Tri95c, UT91, VR94, YKB^{+}00, BAAD^{+}97, CD09, CG87, EB18, FGC06, MD04, Tri95a]. Load-Balancing [CRY94, FT94, Tri95b, Tri95c, Tri95a, EB18, Tri95a]. Load/Store [UT91]. Loading [Bel93, DM93, Cal88]. Loadings [ACL93, BPU94, BPUS94]. loads [CGLY96, CGLxx, Chexx]. loan [Ano89i]. Local [ALM93, Ano94-35, Ano95-27, AM96, BL93, BWO96, BGM^{+}11, Cal86a, Cal86b, CCSM97, DAF^{+}90, Han03, HV95, HNST93, KGKa93, Kon93, KG95, MD94, SVD96, GHNL87, GB91, GV91, Han90, KHS88, Kon87]. local-area [GB91]. local-memory [GHNL87]. Local-memory-based [Cal86a, Cal86b]. Locality [Ano94u, GAV95, KM92, OGR95, TJ94, Yi90]. Locality-Improving [OGR95]. Localization [YKK96]. localized [Ha88, Ha90a]. Locally [BHS^{+}02]. Locating [BCCG97]. location [BGT90]. Locations [BM93b]. Lock [TNIA92, Ano95l]. Lockheed [Str94]. locking [PGK^{+}10]. Locomotion [KT93b]. logarithm [Joh91, Joh92]. Logging [RDZ93]. Logic [BOS93, BS98a, DGJG93, GHK^{+}91, HHGS93, WS84a, Cou11, Gok89, Gok90b, Gok91, GS94c, LMD98, ML95b, Shu88, WS84b, WS84c, Gok90a]. Logical [Ruh95]. Logistic [GGBR95]. Logistics [Ano94p]. LogP [LZ95]. LogP-Machine [LZ95]. London [Ano97g, Don92a, Pit90]. Long [HM93c, UHU09, Bab94]. Long-Term [HM93c]. Look [Ano89l, MTHP93, ALPP00, Ano88q, Ano93d, Asi98, Sha95b]. Looking [Bel99, Jon96, Zim96, Ano95w]. Looks [Ano97k]. Lookup [BS94d]. Loop [Ano94-50, BCZ95, Gru97, OSKO95, Pol87b, RP94, TZ94, TF92, BP91b, BL91, Gal89b, GF95, LY90c, Pol87a, Pol87c]. Loops [Ano94-42, Ano94-127, CWW94, Col94, KKB92, Li92, OP96, RAP95, Ban90, BP89a, Bec89b, CY91, CH90, Fan87, Gan86, HC91, Jac85, Tan87, hTD88, TYZ90]. loosely [vL99]. Loretta [Ano22]. Losses [Hai97]. Lost [Ano94-96]. Lot [HWG98]. Lotto [QD91]. loud [Din92]. love [Day12]. Lovelace [Coc02c, Coc02d]. Lovell [MB93, MB94b]. Low [AM93b, Ano94-124, Ano95-28, BJLW95, BS98a, FS93a, For93, Lee84, NW97, SKS04, Str94, Wal92, Ano97-29, Ano02a, Ano02b, Cou11, DDT95, EO13, KFN02, SSSR20, Sma95, SCG^{+}13b, TF15]. Low-cost [Ano95-28, Ano97-29, Sma95, TF15]. Low-energy [For93]. Low-Level [Wal92]. Low-life [Str94]. low-overhead [EO13]. Low-power [SKS04, Cou11, KFN02]. low-voltage [Ano02a, Ano02b]. lowers [Ano94-83]. LP [CK90]. LPAR [BK95a]. LSI [AM91, TOY96]. Ltd [Ano94p]. LU [AZ95, AZ99, Con94, Dav89, DY90, DD90, DDT95, Kra92]. lubricants [Ano95w]. Lubrication [VF93]. Lucid [AJFH86]. Lugano [Ano95u, GT94]. Lung [Ano97j]. lungs [MKHY95]. LWR [FBGM93]. Lyle [Kaz92].

M [KSW93, Mur10, Por86, Fos93, FXAC94]. MA [Ano96c, LLR93b, Ano88w, Ano92y, GL92]. Maarten [HS94b]. Maass [HA91]. Mac [Val94]. Machine [Ano92n, Ano94-49, AHAM93, Bak93, BRL^{+}20, DGJG93, GCS94, GMMT91, HAAS93, KSTB94, KTK94, KNYT95, LJ97, Mar95, MOOK94, MD94, NBGS96, OS93, PP92b, QD91, RC94, Roj19, Ros95, SO95, AP87b, Ano94-82, Che99, DLM99, Gle91, Ipe19, SNEP14, Sch94c, Sch90b, Sin08b, Ste92, Stu95, Tan89a, Uch86, WLH00, WLLZ20, Ano94-99, Elm95a, Hil91, Hil92, HT93, LTD^{+}93, LZ95, LB94c, SL93, Sha94b, WZ87]. Machine-independent [PP92b]. Machines [Ano94-45, BC95, Bur00, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Bur01a, CK92a, CM95, DKS93, DGT94, GPKK82, GB96, HKT92, IJM14, KKB92, LCVR93, NPS^{+}20, PHK88, SSG93, SNS95, SCSL12, SKN96, TAAL95, Ano93b, BS94e, BWHS18, BBW19, BBB^{+}20b, Cal96, CK92c, Dra89, Fat10, GC92, Gri92, Gua88a, KA96, LW94, LQFC18, PLS20, Per87, Pol87b, SF82, Shu88, Zor92a, Ano92-44]. Machining [KWH94]. Macintosh [DD02]. Macro [YKK96, MSW91]. Macro-Dataflow [YKK96]. macro-tasking [MSW91]. Macromolecular [Ske89, PB94a]. Macromolecules [Lie93]. Macroprocessor [RRW84, Lev98]. Macrotask [YSKS95]. Macrotask-level [YSKS95]. Made [SSxx, Ano94k, Bau96, Mur10, SS90a, SS90b]. Madeleine [ABMN02]. Magazine [Ano85b]. Magnetic [JB90, LB82, GE12]. magnetism [Ano95w]. Magnetoencephalographic [LEMS95]. Magnetohydrodynamic [ACG^{+}90, Pop92]. Magnetohydrodynamics [YMT93, MPG96, Mir88]. mail [PA92]. main [Tay95a, Zor89b]. Maine [TC94]. Mainframe [Ano92q, Ano93-38, Wal85, WZB86]. mainframes [Jor87, Wal85, WZB86, WZ87, Zor89a, Zor90, Zor91]. Mainstream [DD05, Ano95-44]. Maintaining [TG94]. Maintenance [KSTB94, LC12]. Major [Gar99, RSB94, Bur91, BMS92, Cou90]. make [Ano94-86, Bel92a]. maker [Ano96-29]. makers [Ano89a, Ano95a, Bab94, WZ87]. makes [Ano94-55, Ano02a, Ano02b, Str94]. Making [CE18, Erw84, cF03, Rag06, KWW92, Nat90]. Malo [ACM88]. Mammalian [Ano94-52, PG93]. Man [Bro17a, Mut94, OS93, Cap96, CFH^{+}01, Mur10]. Man-Machine [OS93]. manage [GHdF10]. managed [CV88b]. Management [AW94, ALPP00, Bro91b, Cou90, GGG^{+}98, GY93a, JHGLG93, KG94, Ker94, LL08, Leg94, MTK93, NW03, PP93, RWCA94, RS94a, Ris94, SEH98, SSS94, SK93b, Zen94, A^{+}02, BMD^{+}20, BSD^{+}20, CV89a, CKM88, FP91, GY92, GV91, Hus86b, Kar89, LR90, MNY09, SEH99a, SEH99b, Sha87, SR10, SHL^{+}20, SCG^{+}13b, YH92, ZCPT00]. Manager [CSPJ97, Jab90, Per87, PW05b]. Manager/Browser [Jab90]. managers [Ano93v]. Managing [BGKR99, FGKT97, JKL19, Kop00, Spe00, Tys91]. Manchester [ACM94, Man89a, Ano94-134, GS92b, Man92]. Manifolds [OGOR97]. Manipulating [OIY91]. Manipulation [EKZ90, Ano94-100, Pad89]. manipulator [LL88, NS88]. Manipulators [CMPR93]. Mannheim [Meu89a, Meu90, Meu91, Meu92c, Meu93]. Manoeuvring [MGA94]. Manual [TSSK94, WD93b, Bru90b, Chexx, Ham90, HA90b, JT91, Kue87, SG92b]. Manufacts [Ano95-29]. Manufacture [BJH97]. manufacturer [SH91]. Manufacturing [Ano93b, Ano94-107, Hug93, KWH94, KSTF94, LJ97, Raw97, Bra89d, HG88]. Manuscript [BF91]. Many [Cal85b, EFIM91, YZL^{+}20, Ano89a, Ano93u, Che90b, CLF^{+}19, HCD^{+}18, KNHN16, ZBN^{+}19]. Many-Body [EFIM91, Ano93u]. Many-Core [YZL^{+}20, CLF^{+}19, KNHN16]. Many-processor [Cal85b]. Map [OCVG93, RPY94, Rig93, SHA^{+}92, Gla93, Has17, SS07]. Maple [Mon88]. Mapper [AM93c]. Mapping [Ano94t, BS94d, CT93b, Cha93, CP94a, CM93, DDLV93, HPLT01, IM96, KKB92, KESH94, LC93, NB93, Pel94, SH90, SGIS93, WAM^{+}01, Who92, Ana91, CD08, KTN^{+}14, KSB^{+}19, SL88, Whe89]. MapReduce [GGZ^{+}20]. Maps [AGD93, FC93, LB93, Din92]. March [AU87, Ano90f, Ano93a, Ano95u, Ano95-38, Ano96-43, Bel86, HBCN95, IEE95c, JT87, Joh86c, JD95, LCV90b, LCV90a, M^{+}95, MP92, Uni87a, RMO96, S^{+}93, SC93, Vag88]. Margherita [ACM95b]. Marine [MMG^{+}18]. Mark [Ano94-130]. Market [Ano90m, Ano90p, Ano92-30, Ano00b, Gre94, Her90b, NDLV88, Ano92w, Ano92-29, Ano95h, Bra89d, Her90a, Int81, PF90]. marketplace [SDMS99]. markets [Uni92c]. Markov [BM93b, DS94a]. Mars [Pic92]. Maryland [Uni91a]. Mask [Kok94]. MasPar [Car94b, PS98]. Mass [Ano93t, BOS97, GD97, Hal87, IEE95d, Ano90e, Nat87d, Par90a, SSSSE96]. mass-parallel [SSSSE96]. Massachusetts [Ano94-107, MB93, MB94b, Lun94]. Masses [DD02]. massiv [Wat95]. Massive [ARF12, BS98b, CP13, LHLM95, PT93, KTN^{+}14, LAdS^{+}15, Sim92a]. Massively [AK95, ASSW93, AABK95, Ano94-46, AFT97, Bak93, BJLW95, BÇG14, Bur94a, CNC^{+}08, CS94a, DKS93, DXJM93, Eck93, FBZ92, Fei94, FCBH95b, FCBH95a, FM93, Gok92, Goo97, GD97, GVBC95, HL93a, Hel92, HK93a, Hil91, Hil92, HYL^{+}20, Ike95, IGH95, Jab93, JA92a, JA92b, JM93, KNHN16, KC93c, KRJ93, Kra90, MM93b, MOOK94, Mor01, NB94, SABJ94, SB96, Tan95, WMBC97, WCG94, XMR92, YWD94, YWDxx, Afu90, Ano92l, Ano93c, Ano95p, Cre91, CK92c, Din93, EAMS95a, EAMS95b, GP90, GHI95, HLDS95, Kan15, Kra92, Loo84, LLDF95, PS98, RGL^{+}15, Sca92, Smi91, Uni93, Was96a, YW94, Zor92a, dRC94, dC94]. Massively-Parallel [ASSW93, SB96, Smi91]. masterPlan [Kul94]. Match [Ano93r]. Matching [Bel93, CP94a, Nat88a]. Matchup [Smi95]. Material [Ano92h, DA97, DH93, SBSR96, CK90, Was96a]. Materials [Ano94-107, Gal93, KK95b, SSJL94, Sil91, WAD^{+}89a, Nat88b, SCK^{+}00, WH94, Wil88b, WAD^{+}89b]. Math [DDJ98b, EFG^{+}05]. Mathematical [Ano97q, DC93, DLMW95, HM97, KSW93, NH91, SKVZ93, Soe94, TYK93, Poo96a]. Mathematics [ALPP00, JM93]. MATLAB [DP96]. Matrices [Ano94-92, Ano94-94, BGQ19, Che92a, ET96, GG96, ALPP00, Bis94c, Che90a, Che91, Con86, CB89, GSZ91, Gan86, HLTZ93, LTT92, Luc91, Pin91, Wij89b]. Matrix [ALPP00, CHL93, CLY^{+}19, CP93a, DD87, DDB^{+}10, DR81, DR82, FY96, GS88a, GS89c, GR94, HS95a, HL93a, HL93b, KSH94, LPV94, Li95, Mis90, NGLPJ96, SL99, TB89, USZS96, WAM^{+}01, AGZ94b, Bai88, CC88a, CS88, Fuj99, HLJT93, Kra90, KC92, LD90, Phi85, SW88, SB81, SB82a, SB82b, TT93, Yan90a]. matrix-multiplication [AGZ94b]. Matrix-Vector [DDB^{+}10, LPV94]. Matter [DCWH07, FGM^{+}03, GZE^{+}05]. Maui [HBCN95, Ano94-77]. MAX [CG86]. maxflow [BÇG14]. maximally [Gao86]. maximin [LR88b]. Maximizing [Bro00, CWW94]. Maximum [Mis90, BB87]. MaxPar [Che89a, Ho92a]. Maxwell [Taf96]. May [ACM96, Ano88r, Ano88w, Ano93g, Ano95q, Ano97g, Bar01, COS89, Cra96, De 96, DHT89, Fer83, Gro90, HS^{+}91, Ham94, HS95b, HS95c, IEE94b, IEE94c, IEE95b, KK89a, LM92, PEH93, Sch97a, Uni91a, Ano95l, Ano95w, Bau96, Gib01, Sch22, WZB86]. Mazda [AKT90]. Maze [Mik94]. MC2 [DTV00]. Mccormick [Wei90]. MCNP4 [SF93a, YFOT93]. MCSPARSE [GMW91]. MCU [Bal93]. mdb [DKF94, EM91]. MDIONS [Fin82]. ME [Wuo94, Bar00c, Bar00d]. ME20 [Ano94p]. mean [DF90a, Mer01, TfGERJxx]. Meaning [Bar93b]. Means [Hel96, Pay97, SZ98, YZL^{+}20, Bel93, BY21]. Measure [GA95, Lu93, YH90, Yi90]. Measured [Moh94, Smi93]. Measurement [KT94, KNWB93, Mit88, NSP94, Wil88a, EHHS89, McG87, OL86]. Measurement-based [Mit88, McG87]. Measurements [CU90, DCW93, KBC^{+}74, RCR93, Ano87d, BL91, EFR^{+}05, GJW91, Hoc85, Mal86b, Riv90, SZG95, Tem89a, Tem89b]. Measuring [Ano88j, DP91, SBZ^{+}08, DMW87]. MECA [Sol84]. Mechanical [BPU94, KC93a, KWH94, KA93b, Shi95]. Mechanics [Bra93, CNGR90, HFH86, HFH87, MKB87, Nat84, Opp95a, Sch97c, BP86, Hab92, WHBH93, Dra94a]. Mechanism [BCW93, MTLL94, TNIA92]. Mechanisms [Gre88a, GW93b, GWH93, Con00]. Mechanization [Hal87]. Mechatronic [HHGS93, KP94, RW94b, SP94, YS94]. Mechatronically [FT93a]. mechatronics [Ano94a, Ano94-75]. MED [Sal95]. Medal [Ano95w]. Media [Ano93q, App95, GD97, KK96a, OS93, PH97, PC97, WABD97, Bas95b, KMN^{+}05]. Medical [Ada95a, HCPS95, LLSR02, OMA^{+}96, JD95]. medical-image [OMA^{+}96]. Medicine [Ano94v, SR93b]. Meditation [Sul97]. Mediterranean [De 96]. Medium [Bur91, GJS94, GGW93b, GNJW93, JWG93, MH94]. Medium-Range [GJS94, Bur91]. Medium-Scale [MH94]. Meeting [AIA94, ANS92, Ano95q, Ano97l, Bor92, DLM99, Fry97, SEA84, Ano95-38, Ano98a, Cul95a, FJSP95]. meetings [DR98]. Megaflops [Lee89]. MEIKO [SN95a, SN95b, BCM94, BHM94a, Hoc94]. Melbourne [KMG96, ME96]. Melecon [De 96]. Mellon [Ano88n]. Members [Ano97b]. Membership [Pin01]. Memberships [Ano98d]. Membrane [KW95]. Memories [WSP95, Yan93, Bre87, Cat92b]. Memorium [Ber96]. Memory [Abr94, ADLL01, AM15, Ano88k, Ano94f, Ano94t, Ano94-49, Ano94-45, Ano94-43, Ano94-58, Ano94-84, Ano94-85, Ano94-90, Ano94-139, AZ94, BIR94, BCHH94, BGM^{+}11, BC95, Cal85a, CGSG94, CS84, CS86a, CV95, CW89, DL96, DS96a, DLLG98, DHHW93, DVWW05, EJL90, EHS94, FBJ94a, Fri94, GGZ^{+}20, GM94a, GB96, GMG94, GL93a, GS94d, HKT92, IGH95, JML96, KABG95, KV96, KCPT95, KB96, Lee94, LPV94, LMY88, LCVR93, Mah94b, McK94, MH96, MS94c, MH94, MK07, OH92, OB95, OBB^{+}05, PBM95, PR94b, PWVH95, SKIY94, SKIY97, SNS95, ST92, SLRP95, SO95, TSCG94, TH94, VFK^{+}04, WAM^{+}01, Who92, XB96, YFOT93, AP87a, AGZ94b, AP87b, Ano91h, Ano97j, BMD^{+}20, BHM94b, BF92, Cal86a, Cal86b, Cal88, Cal96, Car93, CGL92, Che93b, Che89c, CH92a, CH92b, Con88, Cre91, Cyr86, DL92]. memory [DH91a, DH91b, DI88, EE93, EHHS89, GJM86, Gal87, Gal88a, GS88a, GJG88, Gal89b, Gal91, GL88, GJ87, GHNL87, Gle91, Gok92, GHI95, GGV90, GTV91, Gra92, GL96a, GL96b, GL97, Hir92c, Hus86b, KFW94, Kon91a, KY91a, KY91b, KFN02, KA96, Lee86, LYL87b, Lee87b, LR88a, Lil91, Lim91a, MRM87, ME87, MS88, ME91, Mit88, NG92, OL86, OWG^{+}13, Par90c, PS88, RMM87, RLKW93, RG92, Saa87, SFL^{+}94, SL92, SG92d, SSS90, Ske89, SS07, SY91, TYZ89, Tho90, TV88, TV89, Tur89, Wal17, Yan90a, Yan90b, Yan91, YH92]. Memory-Adaptive [EHS94]. memory-conserving [SG92d]. Memory-Efficient [GGZ^{+}20]. memory-saving [Par90c]. Memristive [Ipe19]. mere [Poo96a, SF82]. merge [WZ87]. Merged [Coc03a, Coc03b]. merging [Ano95p]. Meritorious [Pin99]. Mesh [Ano94-53, BE93b, IMA93, TM94a, YYK93, EAMS95a, EAMS95b, FMD07, Fuj11]. Mesh-Generation [YYK93]. meshed [Wil90b, Wil92b, Wil92c]. Meshes [Ano94-76, CCSR92, Gal96, JP94, PPM90, SJPS94, SJPS96, TS94, TM94b]. Mesoscale [DXJM93, FA93, FM93, Gro92a]. Message [Age05, ABBB94, Ano94-39, Ano94-40, BCM94, DS96a, DHHW93, DFWW93, GB96, Gle93, HLB94, HPLT01, IHIS91, PDR94, PR94a, Sak02, SN95a, SN95b, SYG94, SABJ94, SSOH95, VSM96, YG92, AAC^{+}05, DLM99, MRM87, Saa87, SWJ95, CO94]. Message-driven [SN95a, SN95b]. Message-Ordering [PDR94]. Message-Passing [ABBB94, DS96a, HPLT01, SABJ94, VSM96, AAC^{+}05, CO94]. messaging [KC95]. Met. [Wil93]. Meta [Ano92s]. Metabolic [OCVG93]. Metabolism [Hei89, HL91, HLxx]. MetaCenter [SSH96, Bor94]. metacomputer [vL99]. Metacomputing [KNS97]. Metal [KD93, Nor97b, Ano93-37]. Metallic [BS97]. Metals [WG93a]. metamorphosis [Sha96]. Meteorological [BM93a, Gro92a]. Meteorology [Che94b, HK93b, Kau93b, KH93]. Method [AFT96, AHP97, ABCH97, Ano94-45, Ano94-116, BL93, BJLW95, BV93, DD93, DMPR93, EJ97, FSGS93, FBA93, FI93, FZM91, FHKT97, GG96, GW93a, Gre90c, GZA86, HL96, HM93b, HGS88, HC93, JV93, JM89a, JM90, JC94d, KY93, KGKa93, KO93b, LO96, MKND97, Meu87, ML93, MF92, MMK97, Nag94, NNSY94, NdMM09, NBGS96, Now93, OMR93, RMPW93, SMFG85, SSKa93, SAGS93, SO95, Sus93, TK93, Uen93, UU94, VAGRMVA90, Vog93, WD93a, WRW93, XL94, YA93, Zas93, Ano87e, Ano90l, BB87, Bau88, BGT90, BBK^{+}08, BB91b, CH87, CS88, Che88, CS89, CH89b, Che90c, Chi86, Chi81, DL92, EGK89b, GSZ91, Hea91, HP95, JM89b, JM89c, KS86a, Kan15, Meu89b, MP91a, MP91b, MP91c, Nat86h, OYK^{+}14, PP92a, Roj19, RCS21, Sch87a, SM92, SG92c, SG92d, Sta95a, Sta95b, Vez95]. method [Whe83, YYYS93]. Methode [Wat95]. Methodological [GY93b]. Methodologies [EAMEG11]. Methodology [ATSA93, GB92, HCV97, IK82, KWH94, NMS93, Eig92, JY92, TS90, W^{+}12]. Methods [ALM93, Ada93, AKT90, Ano94g, Ano94c, Ano97q, Ber90b, CT93b, CLP93, CPR93, DAF^{+}90, DL90, FS93a, FGKT97, FI93, GT91, IHSK93, JBWB97, KSW93, Las92, Lil88, MKDY90, ML95a, MS94c, MR90b, PHVJ95, RAP95, RCR93, Saa93a, Sch93b, Sob93a, SC92, TKI93, Vui93, War93a, Wei90, WD93b, Ach99, And88, B^{+}89, Bra89c, BS90b, BS87b, Bur94b, Car89a, D^{+}95, FFM95, Fra90, GS90, GS92a, GL90, HS^{+}91, Ho91, HGS91, Ji91, Joh91, LG87, Lou90a, Lou90b, Mac96c, McC88, PGK^{+}10, Por89, Rob85, Saa87, Saa89, SZ89, Sob92, Svo93, Tze88, UPK87, Vaf88, Van95a, WB88, Yan92, Yi90, van95b, vdV91]. metric [Mar88a]. Metrology [UU94]. metropolitan [BBBC96]. Mexican [Bar01]. Mexico [New91, Ano94-126, C^{+}97, Fra94, NAS93, Met86a, New95, Sie94]. MFE [Chi86]. MHD [AAS88]. MHz [FB91b, IHE^{+}00]. MI [DW97]. MIC [PCY^{+}19, WCZ^{+}18]. Michael [Ano94p]. Michigan [BOS97, IEE95b, Uni96]. Micro [Ano00b, Ano02a, Ano02b, Ano03a, EO91, BY21]. Micro- [EO91]. micro-services [BY21]. Microarchitecture [BP92]. microcode [SK94]. Microcomputers [WZB86, Che96]. microdrops [GH90, Gre90a, GH91]. Microelectronic [Guo94]. Microelectronics [CCKSS90, Mic90]. micrographs [PB94a]. Microlithics [CS94a]. Micromotors [VHJB94]. Microprocessor [Has84, HMS^{+}86a, HMS^{+}86b, MS94b, Mit96, Hsi91, Int91, KM89, Sug80a]. Microprocessor-Based [HMS^{+}86a, HMS^{+}86b, Hsi91]. Microprocessors [BH93, Gep00, LCP^{+}11, Asa98, Per87, WZB86]. Microscope [Sil91, Ano02a, Ano02b]. Microscopic [BM96, MJH90, BGMR96]. Microsoft [Ano01c, Mul08]. Microstructure [Sil91, TFVK94]. Microstructures [Gol96]. Microsupercomputer [MBB^{+}91]. Microsystems [OBR94]. microtasked [MSTK93]. Microtasking [MKB87, CH90]. Microwave [RL90b, RLC91]. mid [Ano88l, Ano93h, Per83]. mid-1989 [Ano88l]. mid-range [Ano93h]. mid-term [Per83]. Middle [Opp95b]. Middleware [Ano00c, Ano00d, BNSP99]. Midrange [Ano92q]. Midrange/Mainframe [Ano92q]. Midwest [Ano93a]. Migratable [MNZ^{+}15]. migrating [Ano96n]. Migration [Ano94d, CCR11, KLY94, LZF16, LCVR93, SE92, WGR93, LM13, MK92a, MK92b, MDP^{+}00, Tze88, Whe83]. Mikroelektronik [Gil92]. Milan [HS95b, HS95c, PH95]. Miles [Han89]. Milestone [Coc02a, Coc02b]. Military [Jon96]. millennium [Nat95]. Milliarden [Ano97d]. Milling [KTK94]. Million [Ano90r, Ano88r, Ano94-86, Ano95v, Wal85]. MIMD [Ano94p, AZ94, BPJ94, DFSZ88, GP90, Gle91, HQ91, HKT92, HS94d, Hor93, Kow85, LCVR93, RS85, TS91, TFVK94, VAGRMVA90, KC93b]. MIMD-supercomputers [DFSZ88]. MIN [CRV94, TV89]. MIN-Based [CRV94, TV89]. Mind [Ote02, Tay95b]. mine [Gal89a]. Mineral [Las92]. Mini [SS96c, WQS92]. Mini-Computer [SS96c]. mini-supercomputer [WQS92]. miniaturization [Ano97-29]. Miniaturized [SVML95]. minicomputer [Nix92, WZ87]. Minimal [LPD^{+}11, LL94, AW91, Cha92b, FRW92]. Minimal-overhead [LPD^{+}11]. minimisation [CP25]. Minimization [PPP94, CH87, Che90c, Gre90b]. Minimizing [KP96, MT96, XMR92, ZGL14]. Minimum [EDA95, Cal96]. Mining [Ano99, CKS99].

Minis [Wal85, WZB86, WZ87, Zor89a, Zor90, Zor91]. minisupercomputer [Rav92, Rav95]. Minisupercomputers [Ano88b, HB89, WSL88]. minisupers [WZ87]. Minneapolis [B^{+}89, JT87, MW88, SF91]. Minnesota [Fin94, B^{+}89, MSCxx, Min88, MW88, Pro94, WL94, Min92]. Miprac [HA92]. MIPS [Cre91, KFB91]. Mira [CKL^{+}13]. Miracle [Ano94-77]. Mirror [Ano94p]. Misleading [Bai92]. Misra [RM88]. missions [Ano97m]. Mississippi [IEE93c]. MIST [Ano93b]. misunderstanding [DMW87]. Mitaka [MN91]. MITI [NW03]. Mito [Ano90f]. Mitsubishi [Ano03a]. Miura [War09]. Mix [Ano93r]. Mix-and-Match [Ano93r]. Mixed [Ano94y, Div97, LM93, LG03, Ano93v, Roj19]. Mixed-language [LG03]. ML [KSM^{+}19]. MLSL [KSM^{+}19]. MN [JT87]. MOB [NJL94]. Mobile [ABM^{+}04, GIBGA93, MGA94, WMMC10, XZC^{+}20, ZLRC20, Liu12, MT13]. Mobile-cloud [XZC^{+}20]. MOC [Chi86]. MOD2.5 [MM93a]. Mode [Sei94, KB18, VO93]. Model [AM93b, AH93, Ano94z, Ano94v, Ano94-58, ABM88, BM93a, BSJW96, BMP93, Ber95b, BPW97, CU90, CGW05, Che90e, CLY^{+}19, CSRB90, DC93, DJSP93, DGG92b, DS94b, Den93, DFS93, DFWW93, DXJM93, DS94c, FM93, FI93, HTI93, HPLC93, HBDS93, HLxx, Hop93, Joh94, KFJB94, KW95, KB94, Mah94b, MKDY90, MNB94, Mis90, NW97, OK93, RWCA94, RT93, RR89, Ros93c, SPM^{+}10, SSKR97, Sei94, Sha94a, SR93a, SS96c, Sug96, TKM96, TM94a, VF93, WMR96, WFT93, WS84d, WC93, Woo96a, WF94, YJD93, Yan94, AGY^{+}11, AKM^{+}06, BGT90, CS91, CGM91, Che90d, CLF^{+}19, Chu87, DP90, DGG92a, DTV00, Fuj11, HLDS95, Kin96, KA96, Law89, MHKY97, NSH95, Nix92, ODAZ15, PSM93, RFS87, SFC^{+}21, Shu88, Smi91, Ste92, Str94, TS90, WM92, YKY90, ZCPT00, HL91].

Model-based [RWCA94]. Model-System [HLxx, HL91]. modeled [Ano95w]. Modeling [AD97, ABCH97, ABBB94, ABCE97, Ano94w, Ano94-48, Ano95w, AFT97, BOS97, CS94b, DBK09, Deh90, DGT84, DA97, GVBC95, Gun88, Heh86, Hun94, JBI91, KLY94, KFJB94, KR94b, LPS90, Mil88b, MRSB94, Mun04, Per93b, Pli97, Pop97, Sch97c, SWSR97, Ste94d, Str10, TMAS97, IBM13c, TP97, TF94, VA94, WWKR97, WMBC97, Wri19, YCC97, ZL97, ZBLZ95, AP91, CC96, De 91a, De 91b, Fox97, Gal89a, Gre88b, Gro92b, HPS88, Kin96, LF03, PLS20, Per93a, SB18, SCH94d, Sug80a, Was96a, WT13]. Modelled [RRSG96]. Modelling [AM93b, Ash93, BPUS94, BM96, BBB^{+}20b, CCSS98, Div97, EHHS89, Fra94, Geu97, GWG93, Hel96, Hey96, JJYL94, Jar12, KD93, KSTF94, KDBG95, LC94, LPLP97, LC95, Moi93, OL86, Pal15, PB94b, Pas95, RSB94, Ruh95, Sei94, Soe94, SB94b, Tay95b, WH93, Wie96, WG93a, BWHS18, TM88, WH94]. Models [Ano94-52, BCHH94, BK93, BBC92, BY96, BM93b, Bot96, BB93, BP96, DS94a, DGO90, Dic81, Dic82, Dic90, DH93, Dip96, Fie93, Fos93, FT94, GH93, GP93b, GD97, HW97, JW98, KB93, LS93b, Max81, MCB^{+}01, Nag96b, Nag96c, PPG94, San93, SKVZ93, Tay95b, Van94, WSP95, WHMA97, XZC^{+}20, Zla01, Ano94-120, DLS93, FRS^{+}88, Gib01, Gil94b, KSB^{+}19, LP94, LCV90b, Ons88, Par90b, Pop92, SNEP14, YQTV12]. Modem [Bar00c, Bar00d]. Moderator [DB94]. Modern [Lin82, RLC91, Smi93, Gil88, KK82]. Modernizing [Jon96]. Modes [GA97, KO93a, SSG93, GH90, GH91]. MODFLOW [MT97]. Modifications [Bin88]. Modified [BE93b, Chi86, Eij90b, Eij91]. MODTRAN [WLCG02]. Modular [BK97, GI93, Hus86a, Kra01b, NdMM09, OCVA01, VD94, Wat72]. Module [BS98a, CMPR93, CC94a, Hei90].

Modules [BLO94, Ano97-29, FGC06]. Modulo [EDA95, Rob89]. Moffett [AU87, Uni87a]. mold [Ano93s]. Moldability [CB02, dCCF01]. Moldflow [Ano93s]. moldmaking [Ano95i]. Molecular [ARF12, Ano87a, Ano92r, Ano94-137, ABGL96, BB90, BHEG94, CFV^{+}90, CH10, CHMS94, Cra96, DAF^{+}90, DAKM98, ES96, FR81, Gun88, INKN01, LB94c, PZS^{+}20, SFF94, VVKB96, AMS^{+}15, BBK^{+}08, DB95, EFG^{+}05, FGM^{+}03, GZE^{+}05, GKS09, Hua92, KHZ^{+}08, OYK^{+}14, PS98, PB88, SSS92, Sch89b, SCK^{+}00, SPP^{+}05]. molecular-dynamics [SCK^{+}00]. Molecules [Bos94b, DAF^{+}90, WKHS97, Ano02a, Ano02b, Lag89, RD07]. Molina [CCKSS90]. Moment [AFT96]. MOMI [DFSZ88]. MOMI-connection [DFSZ88]. Monaco [LMT95]. MONC [BBW19]. Money [SW10a, CKS99, Sin08b]. Monitor [Val94, Lav89, War89, Wil88a, WMK90]. Monitoring [Ano94c, GSG^{+}94, KSTB94, KB97, SKAT93, UP01, YSS94, Dan91]. Monoacid [VM94]. Monoacid/Diacid [VM94]. Monograph [SG94b]. monomial [CR94]. monster [Gei16, Moo06]. Monte [Ano87e, AHAM93, BBS94, HAAS93, IK91, VNB93, ALM93, Ask93, Bak93, BL93, BPJ94, BJLW95, BLFT84, Bro96, Cha84, DKS93, Dec90, Din93, FBA93, Gri88, Gup88, HEJM95, KY93, MZ95, MNR86, MMRL93, MNV93, MS94c, MBN93, NM93, PB88, Rie93, SF93a, Sol84, TW92, Uen93, YFOT93]. Monte-Carlo [MBN93]. Monteporizo [Vag88]. Monterey [IEE95d]. Montgomery [Alaxx]. month [Ano92-44]. months [Ano96u]. Montreal [Bup87, Dup86, Dup87, Goo97, VAS82]. Monty [War03]. moons [BK91b]. Moore [Bar01, HA90b]. Moriarty [MSTK93]. morphogenesis [Hun93, Hun92]. Mortem [KK96b]. Mortgage [CCZ93]. Mortgage-Backed [CCZ93]. Mosaic [MWB95, VSW94]. Mosaics [OLLG96]. Mosher [Lew96a]. Most [DE84, DKH86, US01, AL92b, DMW87, Gep01, Sha95b]. Motifs [HC93]. Motion [AABB93, BSB93, BISB96, DSB96, GGW93b, LJ97, MK93, Sat93, YK94, Ano90l, Ano96u, HAG^{+}13]. Motor [DNV93, KDBG95, Koo97, New93, RSRG95]. Motorola [Ano00b]. Motors [DC93, FS93a]. Motorways [Wie96]. Mould [MHE97]. Moulding [MHE97]. Mountain [McC88]. Move [Ano96-44, Ano94s, Ano98f, jJ88]. move-out [jJ88]. Movement [BBL95, BJV^{+}16, Jac85, Zho88]. Moves [Bar00c, Bar00d, Ano95-44, CSFS00]. Moving [Ano92s, Ano94-99]. MP [Cra92, CDH84, DH86b, DH86a, Lar84, MSTK93, Oed92a, Oed92b, WB85, ARW93a, And90a, Ano88l, ABHS89b, ABHS89a, BOS93, Bow88, BH92a, BL91, Cal85a, Cal85b, Car94b, CM84, CM86, Che89b, CS84, CS86a, CRM94, CS93b, CS95, Dan91, Dao88, DO89, DP90, DH91a, DH91b, Dic90, Din92, EE93, EY91, FSY88, GP93a, GS89d, GZA86, Gur88, HL88a, HVY91, Ho91, Hoc85, Hol90b, HKN89, HES93, HKSY94, HSKY95, KN88, Kha93, Kra88, KM85, LS92b, LS93a, MLR90b, MLR90a, MKB87, MSW91, Nag90, NR86, OL86, OD88, Pap92, PS94a, Par90c, PBK91, Pin91, QB92, Rei85, Rei88, RS85, Rit88b, Rit88a, Sar91, SW91, Sea86, SPS90, SSLR90, SWL^{+}91, SS90c, Svo93, hTD88, Tem89a, Tem89b, VSH90, VSH91, Vaj91, WHBH93, Wes89, Wil88a, WMK90, ZM86, van95b]. MP-2 [Car94b]. MP/2 [Cha84, LMM85b, LMM85a, NSH95]. MP/24 [GKL^{+}87, LMM86]. MP/416 [VY88]. MP/48 [Ber90a, CK90, HFH86, HFH87, Meu87, Nag88, VM87]. MP/Model [RR89]. MP1 [RBL94]. MP2 [KNHN16]. MP8/864 [Cho90a, SO91]. MPCU [Hei90]. MPEG [TCJS93]. MPF [MRM87]. MPI [Ano03a, BBB^{+}20a, Bis94a, BHS^{+}02, BBW19, CCSM97, DLM99, GGZ^{+}20, Gle93, GRRM99, GL96a, GL96b, GL97, LC97a, LKJ03, LSK04, PGK^{+}10, SPM^{+}10, SWS^{+}12, SWJ95, WT11, WT13]. MPI-2 [BHS^{+}02, LSK04]. MPI-FM [LC97a]. MPI/OpenMP [WT11, WT13]. MPICH [LKJ03]. MPMD [KB18]. MPMM [FM93]. MPP [Ano92-29, Ano94-118, Bel96, BD94, FG87, Hoc94, KLY94, KFW94, KG95, WMKS96]. MPPs [DSSS05]. MRI [EFH^{+}00, TfGERJxx]. MRI. [KF95]. ms [Ano94-135]. MSFV [HHOM91, HHOM92]. MTA [BS04, Smi01]. MTA-2 [BS04]. MTPPS [GJP94]. much [Faz87]. Multi [AACK92, BCM90, BHW98, CWLT97, IMA93, KB96, LM13, RCK97, RSRG95, VWC96, VB90, XB96, AMS^{+}15, BAD01, DHA^{+}13, LM90a, LXW^{+}16, MSW91, SY91, Yi11]. Multi-Block [VB90]. Multi-Body [RCK97]. Multi-Channel [XB96]. multi-cluster [LM90a]. Multi-Dimensional [BCM90]. Multi-Electrode [RSRG95]. Multi-Gbit [CWLT97]. Multi-Gbit/sec [CWLT97]. Multi-Grid [BHW98]. multi-job [MSW91]. Multi-Level [IMA93, LM13, AMS^{+}15]. multi-phase [LXW^{+}16]. Multi-Platform [VWC96, BAD01]. Multi-processors [KB96, SY91]. multi-stage [DHA^{+}13]. Multi-threaded [AACK92]. multi-zone [Yi11]. Multicast [Ano94-31, Ano94-88]. multichip [Ano97-29]. Multicluster [Che92a, CWD^{+}08, Fra90, FGM90]. Multicolored [FHKT97]. Multicomputer [AK94, MCW98, Rui91, AP90, SWJ95]. Multicomputers [Ano94-44, Ano94-84, CSSY92, GB92, LB96, Rue92, Ste96, SLRP95]. Multiconference [Chi90]. Multicore [Moo08, MRGR12, KBD10, PATT12, WT11, WT13]. Multicriterial [Sob93b]. multidatabases [ALPP00]. Multidimensional [AFAGR96, Ano94-41, GW93a, ML93, NR86, YYK93].

Multidisciplinary [BWGG94, Ewi97, Kue93, YSS94]. Multidomain [GD94b, LS93b]. multifractal [DLS93]. Multifrontal [PS94b, ZMDS96, Luc91]. Multigrid [Ano94-45, Dic94, Hem84, HGS88, McC88, VM87, WLKI95, Wei90, Zas93, BWV^{+}17, GKSR14, Kan15]. Multigroup [ALM93, AM93a, Rul93, Zas93]. Multijoin [KK95a]. Multilayer [RPY94, SKK^{+}90]. Multilevel [NJL94, NGDH96, AGY^{+}11]. multilevel-PGAS [AGY^{+}11]. Multimedia [CFS95, Int92, MWB95, MBM^{+}20, Moo06, Ste94e, TF94]. Multinode [Hor97b, Hor97a]. Multiobjective [CJ93]. multipipelines [GZR89]. Multiple [Ano96r, FGKT97, GSB95, GAV95, IBC^{+}11, MD94, Mor92a, PC97, SSM93, BJV^{+}16, KB18, LLDF95, MI01, Mit88, Nag88, SFC^{+}21, SG92c, SG92d, TYZ85, vL99]. multiple-instruction-multiple-data [LLDF95]. Multiple-Issue [MD94]. Multiple-Level [IBC^{+}11]. multiple-path [TYZ85]. multiples [RW89]. Multiplex [Gil93]. Multiplexing [HNST93]. Multiplication [CLY^{+}19, DDB^{+}10, Has84, LPV94, AGZ94b, Bai88, CP93a, HLJT93, HL93b, TT93]. Multiplicative [BHW98]. multiplied [LH87]. Multiplier [Has84, LH86a, LH86b, LH86c, LH86d, LH87]. multiply [Ano94-131]. Multipoint [BWGG94]. Multipole [BHEG94, OYK^{+}14, Sta95a, Sta95b]. Multipole-Accelerated [BHEG94]. Multiport [JML96, PDR94]. multiprecise [BW88]. Multiprocessing [CDH84, KABG95, KHMD94, Pol88c, And90c, Asl91b, Def87, Hol90b, JS86, KW92, Lee86, Mir88, SCV01, Sar91]. Multiprocessor [AACK92, Ano94-30, Ano94-56, Ano94-85, Ano95-30, Ber90c, Ber90b, EHG95, FBJ94a, GP85, Hwa85, KLN90a, Lar84, LYL87a, PC93, RWNJ94, SLB93, Sma95, Sob93a, SB96, Swa86, SO91, TF92, WF93, ZX95, ASK85, Bau88, BS87a, Ber89a, BB91a, Che83, Che93b, CV88b, Che89c, CG87, Con88, Dav86a, DI88, EO91, Gal88a, Gal89b, Gal91, GJ87, GHNL87, Gha84, Gri92, Guz86, Har86, HY89, Kam86, KLN90b, LMY88, Lim91a, LY91b, LPS86, LP86, Mar88a, McG87, ME91, Mil87, Mit88, RG92, SSS90, Ske89, Smi81, Sob92, Su92, Tan89b, Tze86]. Multiprocessors [AW94, Abr94, Ano94-43, Ano94-90, AZ94, DG95, DS96a, GM94a, HT94, Joh97, Jor86, Kir89, KCPT95, McK94, MH94, NB93, OA94, PVA94, PR94b, Rot94, ST92, SM94, TA94, YSKS95, YG92, Abr90, CGL92, CV89a, DD90, DDT95, GS88a, GL88, GGJ89, GGV90, Gra92, Hus86b, HKP88, KS86a, Kon91a, KY91a, KY91b, LYL87b, Lee87b, LY90a, Lil91, MRM87, Mir88, Pol88d, SMH91, SA90, TYZ89, TV88, Tur89, Vei85, Yan90a, Yan90b, Yan91, YTL87]. multiprogrammed [Ang91, Mil87]. multiprogramming [Pol88c]. Multirate [Yan94]. Multiresolution [ZM94]. Multiscale [SSKR97, TMAS97]. Multisplittings [HO92b]. Multistage [Ano94-88, Ano94-105, OA94, FJ91]. multistaged [Kra88]. multistep [Svo93]. Multisupercomputer [LZF16]. Multisupport [Ost94]. multitasked [Mil87]. Multitasking [Ber90a, CM84, Cha84, CM86, FSY88, Guz86, HKN89, Lar84, Meu87, Rei85, Ros93c, ZH88, DCG90, DH86a, KM85, Nag88, Nag90, WLH00]. Multithreaded [Ano94-126, BH95, FT96a, HMNN91, HHOM91, HHOM92, HLB94, VTSM12, BÇG14]. Multithreading [Smi01]. Multitoroidal [ADG^{+}08]. Multitransputer [GJP94]. Multivector [KISY94, MS88]. Multiway [JLC98]. mummy [Ano91t]. mundane [Gro92c]. Munich [GH94a, GH94b, GH94c]. MuPAD [SW99]. Murman [Por86]. Murray [Nor97a]. MuSE [DGJK93]. Music [Pau08, KLM94]. Music-Editing [Pau08]. Muster [Ano93o, Sha95b]. My [Jon03]. Myriad [CGHL94]. Mysteries [Nor96]. mythical [Cap96].

n [DT96, BAAD^{+}97, Swe94]. n$-$HLF$/ [Tem89b]. N-Body [Swe94, BAAD^{+}97]. n-cubes [DT96]. N3S [JY92]. nach [Wat95]. Nacional [C^{+}97]. NAECON [IEE94b]. Nag [Chi20]. Nagoya [IEE93a]. NAL [Ano94-32, Kah93a, Sin94c]. NAMD [KHZ^{+}08]. name [Sne94a, Sne94b]. Named [Pin01, Stu03]. Names [Coc01, Pau08]. nanocomputing [WWJ09]. Nanoelectromechanical [DDJ98b]. nanoparticles [GE12]. Nanotechnology [Mer01]. NAO [MN91]. Naples [Pel93a]. narrow [Con87a]. NAS [AGZ94a, Ano94-78, AHOK02, BBDS94, Gib95, Joh86c, Nai94, PO88, WT11]. NASA [NAS93, AU87, Ano89m, Ano92o, Ano93t, Ano94-73, Ano95v, BPM^{+}89, Gri86, MDH^{+}16, Uni87a, SF82, Sim92a]. Nashville [Chi90]. NASTRAN [BP84, GZA86]. National [Ano91j, Ano94-60, Ano22, BBB^{+}91, Bor94, BK91b, Cor89b, CH89a, CKS99, CR89, Cul95a, Cul95b, IEE94b, Joh86a, Lee89, Mac91b, Sha89, Str94, Uni92e, WMBC97, Joh86b, AB94, Han03, OGO^{+}20, Pou88, Ste90, YK87, Ano94-79, Ano95w, Ano96-37, Hab89, Joh94, Kah97, Mar85a, Mar85b, Mir90, Nat86f, Red91, San86, San90, Uni96, UU94, WZ97, Web91]. nations [Ste85]. Nationwide [Ano95-31, Ano93-40]. Native [EBS02]. NATO [HS94b, Coo95, Kow89b, Lag89, OMM93]. Natural [Ano91k, JC94a, Kar93, Max81, WWKR97, WG93a, Ha88, Ha90a, Kel85]. Nature [Ano94-80, PWVH95, Ano93e]. Navier [Ano87a, Ano87c, Ano92e, Ano94-140, Che99, DLPQ94, Dic94, FY92, Glo89, KR94c, LM90a, MFK94, Riz94, SBHW80, Vui93]. Navier/Stokes [FY92]. Navigation [MGA94]. NB [BG91]. NBS [Ano85a]. NC [KWH94, Way96]. NCAR [Nat87a, Nat91a, Nat86a, Nat84, SL93]. NCCS [Ano93t]. NCSA [Ano87a, Lew94b, Nat86e, Nat87d, Nat91b, Nat92b]. NCSC [Norxx]. NCUBE [PC93, Ano94-81]. Near [KY93, WK95, Mit88]. near-coincident [Mit88]. Near-Optimal [WK95]. Nearing [Coc02a, Coc02b]. NEC [Ho88, Ano92o, DTV00, Dub87, HLPP97, Hib01, Iwa90, jJ88, MM91b, SWL^{+}91, TW92, Tze88, Wat87]. Necessary [Poe95]. Need [Ano94-110, Coc01, Dal84, Ewi97, PC94a, Sub94, MS84, SSS94, VVH95]. needed [Ano92-42, CK92c]. Needleman [AFF93]. Needleman-Wunsch [AFF93]. Needles [Bur00, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Bur01a]. Needs [FT93a, Fry97, HG02, PB90]. Neighborly [Ano96s]. Nelson [Ano93b]. NEM [MTK93]. Nematode [Dro95]. neocortex [DLJ^{+}08]. NERSC [DAC^{+}18, HCD^{+}18, Sim97]. Nervous [Dro95]. NESC [Uni92d, Uni92e]. Nested [KKB92, Lou90a, TMAS97, BCH^{+}93, CH90, Fan87, Gan86, HC91, Tan87, TYZ90]. Nests [OSKO95, TZ94, GF95]. Net [GVBC95, KLM94, Law00, MBK^{+}92, SDK98, Her94, PT92]. Netherlands [DSZ96, Emm85, Sig90a, tDv87, ACM90]. Nets [BKT94, Nor97b, Cas01, Jab88, Str94]. Netsim [TIOK94]. Network [Ano90i, Ano94-30, Ano94-35, Ano94-52, Ano94-104, ABMW93, CCZ93, CPS96a, CS93b, CS95, CP96, GD94a, GS94e, HL95, Her95, HV95, HNST93, Ho88, KMNT96, KW95, LLR93a, LAPR94, LTD^{+}93, LLL^{+}17, LA93, MSGW94, NSP94, OA94, Per86, SF93a, Ste96, VDK91, VDK92, VT95, Waz89, XCLW93, ZS94a, ZM86, ZYL^{+}16, ABC^{+}05, Ano94-135, Ano95-37, Asa93a, Asa93b, BBBC96, BHM94b, CPS96b, Cho90a, Coc01, CBM^{+}05, DuB90, DR91, FDM07, FRS^{+}88, Fox98, HCL88, Her94, HY89, yHY92, KTN^{+}14, Kon87, KGLA85, Lee87a, LL88, LS93a, LEY86, LC12, LW94, LAL02, McD90, OGO^{+}20, PS88, RFS87, SBC91, Smi89, SHB91, Tur89, TYZ85, Way96, WWTE92, YYYS93, Yau88, Ano94-105]. Network-Based [Ste96, LAL02]. network-interface [Ano94-135]. Networked [FGKT97, Ano97g, DB95]. Networking [Ano95-32, Hof94, KEMB99, KNS97, Lid99, NĆ02, OPR01, Pel93b, PC94b, Pre93b, Sch94b, WP94, AB01, DAC^{+}18, Ede92, GH94a, GH94b, GH94c, HS95b, HS95c, LPC^{+}95, Lid96, LCHS96, Mac97c, Mec95, Uni91b]. Networks [ADGA95, Ano92i, Ano94-35, Ano94-53, Ano94-88, Ano94-143, ALMS92, BGMR96, BGS^{+}12, BA95, CJ94, CTD^{+}16, COC93, For02, Gre91a, GVBC95, HK96, Hol94, HHK94, HW96, IEE93a, IEE94a, IK82, LN94, Lan93, Lei85, LBT94, Lie93, MJH90, MPH93, MCW98, MMK97, MVS94, NRS95, NB94, NBKP95a, Opp95a, OCVA01, RE94, Ram94, RG94, STN93, SC97, TVT^{+}16, TM94a, TPJ^{+}19, XB96, Yan94, ZLRC20, ZFF^{+}18, AP90, ABP92, Ano93-41, Ano95-27, BP91a, Bue91a, Cap96, Cat92b, CF92, CD08, Chi86, Cyb89b, DF90a, Dra90a, Dra90b, Dra91a, Fid90, Fid91, FJ91, GB91, HWP95, Hol93, yHYZ87, HY92, KHS88, Mor92a, MP92, NS88, NGPH99, NBKP95b, Pou88, RD07, Sci86, Sie90, Smi92, SGS^{+}20, Tze86, TYZ88, VO93]. Netze [Meu92a]. neuen [Ano01b]. Neumann [Ano87e, Joh86a, Joh86b]. Neumann-Ulam [Ano87e]. Neural [ADGA95, Ano91l, BA95, BKT94, BK95b, COC93, CS93b, CS95, FRS^{+}88, GD94a, Her95, HV95, IEE93a, IEE94a, KLM94, Lie93, LLL^{+}17, ML95a, MPH93, MHW94, MSGW94, MMK97, MBK^{+}92, NB94, Nor97b, Opp95a, OCVA01, Ram94, STN93, SC97, Str94, VT95, WSP95, WWY93, XCLW93, ZS94a, ZFF^{+}18, Cho90a, Cyb89b, Her94, HWP95, Jab88, LS93a, LW94, McD90, SBC91, WWTE92]. Neurocognitive [Ruh95]. Neuroimaging [Fri95, MH95]. Neuromodules [Die95, Pas95, Sto95]. Neuromuscular [UR95]. Neuron [Ano92h, KDBG95, LP94]. Neuronal [AB95, RBK95, RSRG95]. Neurone [LC95]. Neurons [Eis95, RSRG95]. Neuroscience [KF95]. Neurosciences [MG95]. NeuSim [OCVA01]. Neutral [GGW93a, GGW93b]. Neutron [Bak93, DCW93, HL93a, JV93, JWG93, SMFG85, Uen93, WD93a, WRW93, KC93b, Zas93, EAMS95a, EAMS95b]. Neutrons [FSGS93, GNJW93]. Nevada [ACM89a, Ano96l]. never [Gib01]. News [Ano95v, Ano95-36, Ano95w, Ano96u, Ano97j, Ano97l, Ano97m, Ano99, Ano00b, Ano00c, Ano00d, Ano02a, Ano02b, Ano03a, Ano14, Bar00a, Bar00b, Bar00c, Bar00d, Bar01, Bor92, Boy15, Bra94, CKS99, CSFS00, Coc01, Coc02a, Coc02b, Coc02c, Coc02d, Coc03a, Coc03b, Cou13, DDJ98a, Gar99, Gar01, Hsu15, IS95, Lam14, Law00, Nat86b, Natxxb, Nor17, Pau08, Pau09, DDJ98b, Str94, Wal17, Sup88a, Ano93b, Ano94-55, Cla97, Moo08, Pit86, Ano95-34, Ano95-35, Ano96t, Ano97i, Ano97k, Ano97l]. Newsletter [Ano93-42, Ano95-47, Ano85b, Norxx]. Newton [Xia88, CH87, Che90c, EGK89b, FFM95]. Next [Ano94-76, Ano97n, Ano02a, Ano02b, ACA94, Cla98, EGJ^{+}02, FBCB18, Jen88, Mes97a, Mes97b, Spe97, WMMC10, ANS92, Ano95g, Ano95v, BG02, Gha84, Per02, Zen99]. Next-Generation [FBCB18, WMMC10, Ano02a, Ano02b]. Nexus [FTT97]. NII [Voi94]. Ninety [Uni86b, Uni86a, Uni86c]. Ninety-ninth [Uni86b, Uni86a, Uni86c]. Ninf [STH^{+}98]. Ninth [GL90, Cha94a, Uni86b, Uni86a, Uni86c].

Nissan [Him93]. NIST [Ano97o, Ano02a, Ano02b]. NIST’s [Ano94s]. Nitrogen [CSB89]. NJ [LCV90a, PEH93, COS89, LCV90b]. NLG [Ano00a]. NLMS [MKfDA96]. NMFECC [Fon85]. No [Cla97, Way96]. Nobeyama [KK92]. Nodal [CT93b, DD93, DMPR93, FBH93, GH93, JKNK93, Koc93, MTK93, TYK93, Vog93]. Node [Ano94-52, Ano94-139, TM94b, WMKS96, CS93a, Kum91, Smi89, VSM^{+}07a, ZBN^{+}19]. Nodes [CT93b, DJSP93, EHG95, KO93a]. NOGAPS [Ros93c]. Noise [FBB97, JC94d, LB94b, RK22, Wil94, YF95, PN96]. Nominations [Ano16]. Non [Ano94-53, Ano94-84, Ano94-109, AJ93, CCSM97, GT91, HVZ94, KB94, Lan94, MTK93, MZ95, McD85, OP96, RCK97, SJA94, SB94d, VF93, WMKS96, WRW93, DY90, HS93a, WHBH93]. Non-Adaptive [Ano94-53, HS93a]. Non-contiguous [Ano94-84, WMKS96]. non-deterministic [DY90]. Non-Equilibrium [HVZ94]. Non-Linear [AJ93, Lan94, MTK93, RCK97, VF93, Ano94-109, McD85, WHBH93]. Non-Local [CCSM97]. Non-scan [SJA94]. Non-Self-Adjoint [GT91]. Non-Spinule [MZ95]. Non-strict [SB94d]. Non-Uniform [KB94, OP96, WRW93]. Nonaqueous [BCCG97, Pop97]. Nonblocking [Yan94]. Nondestructive [Ano91m, TC94]. nondeterminacy [EGP92]. nondeterminancy [PE88]. Nondeterminism [Kra01a]. Nonlinear [Ada93, AM93c, Cla96, FD93, FV94, KFF93a, ZM86, BS87b, Gou90, Hea91, HGS91, Hun91]. Nonnumeric [Ano94-85]. Nonpreemptive [Mil87]. nonrect [Wil90b]. nonrectilinear [Wil92a]. Nonsymmetric [Ma99, SC92, BS90b, Cha92b, KS86a, Kam86, Saa88, SG92c, Yan92]. Nonzero [BW94]. Norden [CCKSS90]. Norfolk [S^{+}93]. norm [FRW92]. normal [LMD98].

Normalization [Amm89, Amm92]. normalizations [Amm90]. Normalized [SYG94]. norms [RW89]. Norris [Bro91c, Haw88, Nor03]. North [Nor89, Nor93b, Nor93a, Norxx, Tru88, VO93, LC90]. North-Holland [Tru88]. Norway [Kow89b]. Norwegian [SSH96]. Note [KCM02a, Arn89, Con86]. notebooks [Ano95-32]. notes [CSR89]. Novel [DLMW95, GMSS^{+}11, KTG08, Sha94a, UU94, VPDA93, GBS18, KFML20, PP92a, SQS^{+}19, VFK^{+}04]. November [ACM89a, Ano91q, Ano92y, Ano94a, Ano94-126, DHT89, EP 97, Emm85, Gra93c, Gra94, Har91, HWP95, HK93b, IEE90, IEE93d, IEE94e, IEE96d, Isk96, KH93, KFF93b, Lun94, NAS93, RD94, Soc94, SF91, Sin94a, Tho93c, Uni98, USE01, Pin99]. novices [NSW08]. Novo [GLS11]. Novo-G [GLS11]. Nozzle [KKDO97]. NPB [Yi11]. NRLM [UU94]. NSA [Bro16]. NSF [Ano94-107, Ano88i, Ano94-73, Ano94-86, Ano95t, Ano95x, Ano95v, Ano95-37, Ano95-45, Ano97k, Ano97z, Bor94, Bra91b, Dau96, Dau97, Dav87, Fat10, FG92, Hay84, Hir94, IEE92, Lew94a, Lew94c, Nat84]. NSF-NASA [Ano94-73]. NSFLEX [PBDM93]. NT [Ano95-32]. NTTC [Ano93b]. Nubira [Hai97]. Nuclear [Ano90f, Ano97q, Ano05, ATSA93, CU90, EFPSS93, ESMH93, FNT93, GY93b, GL93a, IHSK93, KA93a, KTKK93, KSW93, Mal88a, MTK93, PA93b, PP93, RDZ93, Tho93a, Tsy94, VRSG93, VA94, ANS92, Ano96o, Ano96-33, Ano97e, Ano97u, CKT21, Cla97, Kav92]. nucleic [MW88]. Nucleotide [Kon93, MKRI93, TYKE93]. Nucleus [RWCA94]. nukes [Cla97]. NUMA [AW94, WF93, XB96]. Number [Alu96, And90b, Ano94-64, Ano95-31, Ano97d, Bro96, Ent99, GS94a, IK91, LD93a, WGOY91, AM15, Arn88, AI92, ARW93b, CMP94, FRW92, Gut95, KA92, Mas94a, Mas94b, Pry94, YB86]. Number-Cruncher [Ano97d]. Numbers [GW93b, OGY91, Ano91h, ARW92, Fri91, Ked94, OGY90]. Numeric [MH96]. Numerical [Ach99, Ale90, AGKT02, ALPP00, Ano94-32, Ano94-68, Ano94-127, AT93a, AT93b, BK95a, BSB93, BD93b, BS98b, CAB93, CZRB93, Che90e, Che90f, COT21, Chi86, CDPW94, Dic81, Dic82, Don91, DGH20, Ede94a, FA93, Fra94, FI91, GG^{+}97a, GW91, Ger90, GHW93, GW93a, GF90, GW93c, HS94c, HMS93, HVSB93, Hof94, KY93, KBD97, KLSC97, KR94b, KT93b, MS97, MMHM93, PB94b, Pay97, RT97, Sch93b, SHZK94, Sha89, Sim00, Soe94, SZ98, Str97, SD88, Tak94, UR95, VW95, WH93, Wat91, WS84d, WG93a, WCG94, YCC97, ZL97, tDv87, Ber86a, Bru88, Bru91, CGM91, Che90d, CV88b, DHD89, EY91, Ha88, Ha90a, HPH^{+}20, HJZ94, HPS88, JHZ95, KSP13, MMG^{+}00, Nat88a, Pet83, PMC22, Saa87, SBC91, Scr88, Sta88, Svo93, TB89, Fra94]. Numerically [Fru93, FB91a, UL89]. Numerische [Gil92]. NURETH [ANS92]. NURETH-5 [ANS92]. NV [AIA94, Ano95q, Ano98a]. NVH [Pay97]. NW [GAB^{+}96]. NWT [Ano94-32]. NX [PR94a]. NY [IEE96a, IEE96b].

o [Ano94p, Ano90e, Ano94-38, Ano94-41, Ano94-104, BBH95, BIB^{+}18, CP94b, Fei94, FCBH95b, FCBH95a, GS94d, HNS94, Hic18, Hop93, May01, MS94d, NNS^{+}90, Par90a, TGL96, YYW^{+}20, Zec93, dRC94, dC94]. Oakland [USE01]. Object [Ano90s, Ano94-74, BLO94, BPL95, CSSY92, DS94c, GD94a, GJP96a, Gui96, HP93, JAB92, KWH94, KS94a, KNYT95, MBD99, PW94, SSS94, Sol94, Sti98a, Sti98b, SK93b, YMY92, CH98, GJP96b, Jéz00]. Object-based [KNYT95]. Object-Oriented [Ano94-74, CSSY92, GJP96a, Gui96, HP93, MBD99, SK93b, YMY92, CH98, GJP96b, Jéz00]. Objects [BS97, HB96, MNZ^{+}15]. obrabotki [BKM88]. observability [Mal90]. observation [AC91]. Observations [Bel96, Gin82]. Observatory [Ano97k, BK91b]. Obstacles [MMHM93]. Obtaining [ACSH90]. OC [KG95]. OC-3 [KG95]. occurred [Ano88y]. OCE [Cop93]. Ocean [Ano94-107, Ano94-140, BB93, Che90f, CSRB90, DGG92b, DGG93, De 91a, De 91b, DGG92a, LCV90b]. ocean-acoustic [LCV90b]. Oct [Asp93, WSB96]. October [AIA93, Ano90g, Ano93n, Ano94a, Ano97q, B^{+}89, GL92, HS94b, IEE93a, IEE93c, Mar86, Mar88b, MB93, MB94b, Pel93a, Pit90, Sin94a, SR93b, USE00a]. octree [BR95]. odd [ARW93b]. ODS [Tak93]. Off [Bar00c, Bar00d, FLP^{+}07, Ano95-31, DM93, Gep00]. off-beat [Gep00]. off-the-shelf [Ano95-31]. offer [Ano96r]. offerings [Ano95-32]. offers [ALPP00, Ano95v]. Office [Wil93]. officers [Ano97b]. official [Way96]. Offline [Dam11]. offloading [TMT^{+}20, VM07]. offs [RYYT89]. often [Per87]. often-daunting [Per87]. OH [IEE94b, Wei88]. Ohio [Ano88l, Ano92v, Ano97p, AA93, BBW90]. OhioLINK [Ano97p]. Oil [Gui08, RDZ93, Ano95w, BK89, BRL^{+}20]. OK [Ano91s]. OKs [Ano03a]. Old [Ano97m, Pou94b]. Oligomeric [LD93a]. Oligonucleotide [KKKP93, KKPR93, KT93a, Tak93]. Olsen [CCKSS90]. Omega [Mor92a]. omnipresent [Sug80a]. on-chip [Ano91h, KFN02]. On-Demand [Mas95, FK98, VM07]. On-Line [Bel93, EFPSS93, GSG^{+}94, HRG93, RW94b]. On-the [YH90]. On-the-fly [Yi90]. on/Roll [DM93]. Oncology [HSW^{+}90]. One [Ano94-59, Eck93, GMBW93, LB82, Mut94, Tec89, Uni92b, Uni89a, Uni92a, Uni98, Ano92-42, Ano94-121, Ano97t, Ano97s, BMR88, Faz87, LSK04, PGK^{+}10, Rob89]. One-dimensional [LB82]. one-sided [LSK04, PGK^{+}10]. One-Tflops [Ano94-59]. one-two [Ano94-121]. onEM-4 [YMY92]. Online [Nat89b, AZC13, Ano90t, Nat87c]. Only [Ano94-139]. onto [Pau05, WAM^{+}01]. Onward [Bai97]. Open [Ano92r, Coc03a, Coc03b, Her90b, IH94, OGOR97, Ano96j, BBE^{+}20, Her90a, Ano85b, CWD^{+}08]. Open-Source [Coc03a, Coc03b, BBE^{+}20]. opened [Ano95-43]. OpenMP [Ano03a, EO13, EBB^{+}20, TMT^{+}20, WT11, WT13]. OpenRTE [CWD^{+}08]. Opens [Ano94-72, Ano95v]. OpenStack [JTX^{+}22]. operate [Wal85]. Operated [RCK97]. Operating [Chr90, FG93, GGG^{+}98, Hus86a, Koe96, Koe97, Chu87, Kon87, MAA^{+}05, RCZ93, vL99]. Operation [Ano94-33, ESMH93, FCD97, VSH90]. operation-level [VSH90]. Operations [Ano94-31, KS90, Mas92, NJL94, NGDH96, Sah95, Sta94, SKN96, Uni92b, Ano97s, Sch22, Wal85]. operative [GL93b]. Operator [GW93a, Mor92b, KWW92]. operators [PW23]. Opinion [Bai97, GPKK82]. Oppenheimer [Mur10, Mur10]. Opportunities [New93]. Opteron [Fat10]. Optic [Gre91a, BBBC96]. Optical [AGP96, And10, CWLT97, DFSZ88, EH97a, EH97b, GP93c, KT94, McA92, SDK10, XB96, Ano86b, Ano89f, Ano90a, BF92, Bur93, Bur94b, DRAB08, Goo97, KS95, Wic96, WM92, Ano95-38, Clo96]. optical-flow [DRAB08]. Optimal [Ano94-88, DA94, DF90b, Gib93, KS86b, LN94, LZF16, LL94, MPH93, MD94, OSKO95, SAGS93, SBW94, Str97, WK95, EM94b, Fid90, MP91b, MRSB94]. Optimisation [BMSD94, EY91, GI93, LB82, RS94a, Van93, VHJB94]. optimised [BBC^{+}99]. Optimization [AK95, AKG87, AYL^{+}18, Ber95b, BCR96, Bro97, BWGG94, Chi95, Deg90, Ede94b, Fah94, GP91c, GM93b, HW97, HM97, IMA93, KR94a, LPLP97, MTK93, MTLL94, PW86a, Pay97, PP93, RL90b, RW94b, SWG06, SP12, Sob93b, SKN96, WD93b, BSJ^{+}13, BB91a, CLF^{+}19, Chu91, ES88, GBS18, HP92, HES93, KFML20, KSB^{+}19, McC92, MP91a, MP91c, MP90, MP91d, MM91b, Nix92, PB98, Ren97, RGH17, SC20, SSLR90, TMT^{+}20, Win02]. Optimizations [HKT92, KK96b, Li95, PW86b, PW86c, Pol87a, Pol87c, Pol88a, Vei85]. Optimize [CC94a, BBW19, BRL^{+}20, WH94]. Optimized [ST94, BHS92, EBB^{+}20, Sch89b]. Optimizing [AGK^{+}87, BGH^{+}05, Dic81, Dic82, EJL90, GS01, GS06, HKSY94, HSKY95, JCJY94, KM92, SNS95, TY96, ZFF^{+}18, ARE95, BGS82, DP90, Eig92, GJG88, HN90, LXW^{+}16]. Optimum [CS90, EDA95, GS94a, Isa93]. Option [Pin01]. OPTOCOM [SSSSE96]. Optoelectronic [CG96, Rui91, SSSSE96]. optoelectronics [Ano93b]. Or-Parallel [VPGG01, Seh88]. Oracle [Chi20]. Orbit [BS98a]. Orbitals [INKN01]. Order [GW93a, ML95a, EGK89b, Ram88, RLKW93, ZBN^{+}19]. Ordering [LD93a, MOWW96, PDR94, Rig93, GE12, Wil92b, Wil92c]. orderings [Wij89b]. Ordinary [KBC^{+}74, Ban79, HHS01b]. Ordinates [KGKa93]. Oregon [IEE93d, USE90]. Organic [Ver97]. Organisation [FBJ^{+}94b]. Organising [GD94a]. Organization [ABB^{+}03, CD92, Jia94, Pau08, XB96, GJ87, Ull83, Ull84, Wie87]. organizations [HS93c, KWW92]. Organized [LUT96, PN96, UU94]. Organizing [GY93b, RPY94]. organs [Ano97m]. Orientation [Ano94-89]. Oriented [Ano94-74, CSSY92, GJP96a, Gui96, HP93, JAB92, KWH94, KP95, KS94a, MBD99, SSS94, Sti98a, Sti98b, SK93b, YMY92, AGEL13, Cal86a, Cal86b, CH98, GJP96b, Jéz00, Kar13, TS91, Yau88, Pop92]. Origin [LSK04, PIH04]. original [Sch95c]. Orlando [Ano94-100, Gig94, Tho93c]. ORM [EH97a]. ORNL [DBK09, LSS^{+}20]. Orthogonal [FBA93, Rag94, SC92, Bra92]. Orthopaedic [HTV88]. orthotropic [CS88]. Oscillators [BK95b]. OSF [Ano94h]. OSF/1 [Ano94h]. Osservatorio [Vag88]. Other [And90b, Ano94-110, JBWB97, Ano93b, Ano94-119, Ano95a, Fid90, Guz88, Jon03, Sha95b, Ste85]. Ottobrunn [Har91]. our [Ano88y]. Out-of-Core [BCR96, TBC94, BC95]. Outcomes [Tys91]. Outer [BJS02]. outline [Ano94-119]. outlook [DvdS12]. Output [Che90e, XOZ^{+}20, Che90d, DGG18, Mil90, Mil91]. Outreach [JPMG08, WZ97]. Outstanding [Pin99]. overhauls [Ano96-37]. Overhead [MT96, TZ94, BP89a, Bec89b, BP91b, Cal96, DDT95, EO13, LPD^{+}11, Pol88b]. Overheads [KABG95]. Overlap [GF90, LB94a]. overlapped [AGZ94b]. Overlapping [Pev93, Tak93, WB88, Lou90b]. Overview [ABB^{+}03, BCC^{+}05, FG93, GBC^{+}05, Hir94, HCH95, IBM08, Int91, IAKH92, JML95, MH98, Mir92, Ste92, SPP^{+}05, Wil96, vdSD96a, vdSD96b, ALPP00, Dra91a, Hey94, Mar96, Sch90a]. own [Sug80b, Van97]. Oxford [ML95b, OMM93]. OZ [Kel85].

P [IBM08, ARF12, Ano94-36, AUW08, CK92a, CRA10, DM96a, DM96b, KHV11, MCW98, RGL^{+}15]. P-90 [CK92a]. P-Vision [DM96a, DM96b]. P03T [Fah94]. P2 [Rul93]. PA [EM94b, EP 97, Ras91]. Pace [Ano95-33, WZ87]. Pacemaker [Ano94-52]. Pacific [Sin94a, Suh97]. Package [Ano94-87, Cha94b, DGBE96, JC94c, KCOP94, Kul94, MRS88, BW88, Guo94, HVY91, SW88, Yan90a]. packages [Ano92-44, WLN^{+}96a, WLN^{+}96b]. Packaging [Ano94-108, Car94a, CBC^{+}05, CHT^{+}13, HMKI97, KP95, Tho93d, Wat92b, CG96, Gla93, HBCN95, SCV01, Suh97, Wat92a]. packer [Car88]. packet [DuB90, DR91, Joe87]. packs [Ano88f, Ano94-82]. PACS [JD95, HCPS95, Asi98]. Paddon [Kow86]. pages [Ano96c, Ano00a]. Paging [Ano94-90]. Painful [Smi96b]. Pairwise [DD88]. Pakistan [ZAS94]. Palazzo [GT94]. palm [Ano94-63, BPD06]. Palmer [Kaz92]. PALMNET [HNST93]. Palmtop [Ano96p]. PALS [VPGG01]. Panel [Ano91s, Ano95-45, DMT^{+}21, HP95, Mur91b, Rot94, Ano94-126, Ano97m]. Panels [BJH97]. Paper [Bur93, KJ94, Ano90t]. paperboard [Lin89]. Papers [AB94, Ano87b, Ano96e, Ano96f, Ano98b, Ano98c, Ano09, LP90, MB93, MB94b, Nor03, Uni87c, Ano93-39, Ano95-38, Bel86, BBD92, IEE95c, MW88, Nat84, Wuo94]. parabolic [Che94a, GS89a, GS90, GS92a, Sch87a, Scr88]. Paradigm [Bad99, FD97, CK92c, Kel85, VFK^{+}04, VSM^{+}07a, VSM^{+}07b, SLRP95]. Paradigms [MMRL93, VW95]. Paradox [Pev93]. parafrase [Tri85, Pol88d]. Parafrase-2 [Pol88d]. Paragon [Ano94h, Gro93, WMKS96, AABK95, Ano94g, Ano94b, Bem92, CM95, Eck93, GFM96, Gut95, GAW96a, GAW96b, Hoc94, PR94a, Rot94, SNS^{+}97, SZG95, TGL96, Wat95]. Paragon/XP [SNS^{+}97]. Paralex [BAAD92]. Parallel [AP87a, AK95, AGZ94a, AFAGR96, ASSW93, AHP97, Alu96, AABK95, AG94, And88, Ano89b, Ano90g, Ano90m, Ano92w, Ano92x, Ano93z, Ano94c, Ano94i, Ano94t, Ano94y, Ano94-28, Ano94-37, Ano94-33, Ano94-49, Ano94-46, Ano94j, Ano94-54, Ano94-61, Ano94-62, Ano94-65, Ano94-64, Ano94-74, Ano94-89, Ano94-106, Ano94-91, Ano94-92, Ano94-93, Ano94-94, Ano94-95, Ano94-96, Ano94-97, Ano94-98, Ano94-111, Ano94-115, Ano94-116, Ano94-143, Ano95-39, Ano95-40, Ano99, Ara96, AM94, ACL93, AHOK02, AZ99, AFT97, Ash93, Att96, BAAD92, BK95a, BBDS94, B^{+}95, BM93a, Bak93, BPJ94, BJLW95, BOS93, BAT99, BBH95, Ber95b, Ber86b, Ber86a, BS88a, Bha94, Bie88, Bis94b, BHLST94, BSL94, BJS02, BIB^{+}18, Bos94b, BCHJ94, Bro96, BV93, BS98b, BWGG94, Bur94a, BNSP99, Car89a, CLR09, CC94a, CTM94, CDMW94, Che92a, CBCH93]. Parallel [CC94b, CD92, Chi81, Chi95, CDW94, Cho90b, CMF94, Chr93, CKS99, CRA10, CCSM97, CDC^{+}87, CP92b, CP92a, CP94c, Con87a, CP92c, CF94, CM95, CT94, CO94, CS94a, CSG99, CS94b, Cyb91a, CBHS91, Cze16, DDHK94, DKS93, DJSP93, DGBE96, De 91a, De 91b, DGG93, DD02, Def87, DFSZ88, DRRM94, DLMW95, DLLG98, DL90, DGT94, DT97, DFF^{+}02, DFWW93, DXJM93, Dun99, DT96, EKTB99, Ece96, Eck93, Ede94a, EGK87b, ES96, EH97b, EH97c, ET96, EHS94, EK96, FBZ92, FHM95, Far90, Fei94, FCBH95b, FCBH95a, FR96a, FM93, FB91b, Fox89, FJP94, FS93b, FY96, GJS94, GPS90, GSZ91, GMW94, GFM96, GG96, GBF93, Gen94, GHWZ94, GJP94, GMSS^{+}11, Gil94b, GP93a, GKSR14, GHK^{+}91, Gol99, GPS86, GL93a, Gra91, GD97, GVBC95, GL94, Gui96, GMSB93, GBK^{+}96]. Parallel [GK93, GMMT91, HMM94, HL95, HL93a, HQ91, Hay86, HCL94, Hea91, HR94, Hel92, HV94, HGS88, HK93a, HHT^{+}94, HVSB93, HS94d, Ho91, HK93b, Hol90b, Hor90, Hor93, HMKI97, HHK94, HES93, HO92b, HERC95, HGS91, Hun91, HD89, IEE93c, Ike95, INKN01, IM96, Jab93, JA92a, JA92b, JC94a, JC94b, Jay87, JAB92, JM89a, JM90, JM93, JP94, JC94c, JW98, Kar94, Kau93b, KH93, KB93, Kau93a, KB19, KMNT95, KMNT96, KDBG95, KLM94, Koe96, Koe97, KC93c, KB94, Kon91a, KY91a, KRJ93, Kow85, Kow86, KRS13, KK96b, KHV11, KDLS86, Kuc87, KHMD94, KKB92, KESH94, KSH94, KNYT95, LL08, LPNRJ94, LMT95, LM93, LA94, LR92, Lei91, Lew93, Li92, LY91b, LLY92, LB94c, Lou90b, Ma99, Mah94c, Maj94, MM93a, MP94, MM93b, Man91, MM91a, MJRS94, MMRL93, Mas94b]. Parallel [MOOK94, MB94a, May01, McB93, Meh94, MPG96, M^{+}95, MBD99, MSW91, MRAR95, Mor01, MS94d, MKSF96, MM94c, MM94d, Nai94, NKTT95, Nar95, NMS93, NdMM09, NB92, NB94, NK96, N^{+}95, NĆ02, NK94, OS94, OD01, OLLG96, OB94, Ope96, OP96, OYWK91, PIH04, PB90, PC94a, PEH93, PCY^{+}19, PZS^{+}20, PCM84, PBDM93, PT93, Pin91, PK87, PHVJ95, PW94, RL96, Rag06, RKDM94, RAES96, RAP95, Rav92, Rav95, RS94a, RMM87, RM88, RS94c, RBL94, RT97, RG92, Rue92, Rui91, SSG93, Sal89, SH90, SG92a, SYG94, Sch97a, SF93a, Sch96, SBF94, SD92, Sha94a, SLB93, Sie94, Sim92b, S^{+}93, Sin94b, SABJ94, SFF94, SB96, SG94b, SHG95, Ste94c, SPGD98, SSOH95, Str94, SO95, SKN96, Sug94, SA94, SRBL94, Swe94, SLS96, TFO94, Tan95, TGL96]. Parallel [TP93, TY96, Uni87c, Uch96, Uch97, UZ95, VVKB96, VW95, Van94, Van95a, Vez95, VAGRMVA90, VPGG01, VB90, WLKI95, WCZ^{+}18, WMBC97, WGS91, Who92, Wil93, WB85, Wil95, WL96, WC93, WCG94, WF94, Xia88, XL94, XMR92, YFOT93, YJD93, YKB^{+}00, YMT93, YWD94, YWDxx, KC93b, Zen94, ZM94, Zim96, AD88, Abr88, Afu90, AGZ94b, AP87b, AS88, AP91, Ang91, Ano85a, Ano88s, Ano89h, Ano92l, Ano93c, Ano95p, Ano95u, Ano96-43, Ano97-30, AM96, Bab90, BS00, BBC^{+}99, BAAD^{+}97, BAD01, BP89a, Bec89b, BP91b, Bis93, BCH^{+}93, BS88b, Bra89b, BS90a, Bri90, Bru90a, Bue86a, Con87b, Cal96, CBCJ92, Cha90, CH87, CSY89, Che89a, CH89b, Che90c, CGL92, CV88b, Che99, CH90, CH92a, CH92b, Chu87, CNC^{+}08, Con86, Con94, Cre91, CMP94, CK92c, DD88, Dav89, DY90, DD90]. parallel [DSZ96, DM96c, Din91, Din93, DWM^{+}01, DS86b, Don87, DLM99, Don92a, EGK87a, EGK89a, EGK89b, Eij90a, ESTA94, EHF^{+}97, EGP88, EM91, EGP92, EAMS95a, EAMS95b, Fan87, FMD07, FDM07, Feo92, FR95, For93, Fra90, Fuj11, FMT91, GJM86, GMW91, GS87a, GS88b, GS89a, GS89b, Gib95, GP88, GP90, Gok91, GC92, Gok92, GS93, GM93a, GS94c, GHI95, Goo97, GYL00, GV96b, GM87, Gua87b, Gua88a, Gua88b, Gua88c, Gua88d, Guz88, Hae91, HLDS95, HJZ94, Han94, HC91, HVY91, Hil91, Hil92, Hor98, Hsi91, HR04, Hun90, HLJT93, HLTZ93, IEE96c, Jay88a, JM89b, JM89c, Joh90, JHZ95, KPS88, Kan15, KB88, KNHN16, Kha93, KTN^{+}14, KG03, KY91b, Kos95, Kra93, Kra90, Kra92, KC92, Kre95, KESH95, LD90, Lan92, LP94, Lee90, Lei89, LR88a, LY88b]. parallel [LY88c, Li89, LY90c, Lim91a, LHPG21, Loo84, LYC93, LM13, LF03, LLDF95, MD04, MCH91, Mar91, Mas94a, McB92a, McC92, MB97, Meu89b, MP91b, MP90, Mik89, Mil93, Mor92a, NNS^{+}90, NPS93, NRN00, Noo95, OW94, PE88, PHK88, PSG03, PTT89, PS98, Pol86, Pol87a, Pol87d, Pol87b, Pol89, Pry94, PMS94, Qui87, RR99, Rei88, R^{+}00, RGL^{+}15, Saa87, SNS^{+}97, SN95a, SN95b, Sar90, Sca92, Sch94c, Sel95, SL92, SC04, Shu88, Sie90, Smi91, Sta95a, Sta95b, SJR05, Sul91, SE98, Suz89, SSSSE96, Tan87, TYZ89, TYZ90, TCM95, TFB94a, TFB94b, TFVK94, UL89, Uni93, Vol89, WHBH93, WLCG02, Was96a, WD94, Woo92, Woo94, WCHK91, WT11, Supxxa, YYYS93, Yan90a, Yan90c, YF98, YW94, Yew88, YVC89, YB90, ZCPT00, ZBN^{+}19, Zor92a, dRC94, dC94]. parallel [tDv87, DDC96, HK93b, JPTE94, PEH93, Pra95, WN10, YGSB94, Seh88, Ano95z, Ano94p]. Parallel-Processing [Hay86]. Parallel-Vector [BCHJ94]. Parallel/Distributed [CC94a, SD92]. Parallel/High [MBD99]. Parallel/High-Performance [MBD99]. Parallel/Vector [Far90, PHVJ95]. parallelen [Wat95]. Paralleles [Kro92]. Parallelisation [ER94, Geo94, PRSS94].

Parallelising [CCSS98, BMT96].

Parallelism [AACK92, Ano94j, BAM93, BEH^{+}94, CWW94, GGG^{+}98, HB96, KP96, KM92, KBC^{+}74, Lee86, LPS90, SSG93, SWG06, Uen93, WBP87, AMS^{+}15, Ano91h, Dak90, FMT91, FP91, GW95, GP91a, Gir91, HC91, Jéz00, Jor86, Kim96, Kos95, LY90a, Lil91, MPC89, PB87, Pol88a, Pol88b, RF93, SK92, Sim92a, VSH90, Whe89, Zor92b]. Parallelizable [Dic94, AT91, LTT92]. Parallelization [Ano94-42, BCHH94, BK93, Ber95b, Den93, FBZ92, Fah94, GJS93, GMS97a, GMS97b, HBDS93, INKN01, McK94, ME91, MT96, OPR01, YFOT93, YR93, ARW93a, BMS92, Blu92, BBK^{+}08, Eig91, GB22, Gua87a, HA90b, Her94, LY88a, Seh92]. Parallelized [KR94c]. Parallelizing [ASS94, CHMS94, DS94b, Isa93, KLN90a, KLN90b, KLD95, LXW^{+}16, PE95, RAP95, Sea86, TP95, Yan91, BE92, EB91, GF95, Hag90, HP91, HP92, KK89c, Leu90, PP92b, Pol88d, Sch90a, SLY90]. parallelo [LP90]. Parallelrechner [Sch92a]. PARAM [Bha94]. Parameter [PC97, Ji91, YKY90]. parameterized [BE93c, SS07]. Parameters [AH93, PA93b, VT95, Hoc91]. Parametric [PPG94]. PARAMICS [Ano94-99]. Paramid [Ste94c]. PARASPICE [Yan90c, Yan90b]. PARC [Coc02a, Coc02b]. PARCEL [HP88a]. ParCo93 [JPTE94]. ParCo95 [DDC96]. Parellel [IGH95]. ParInt [DGBE96]. Paris [Ano90g, GL90, GLH94]. Parity [AFML93]. Park [IEE93b]. Parker [Bro91c, Haw88]. parllel [Yan90b]. PARMACS [Hof93]. parole [All93]. Parrinello [BBK^{+}08]. PARSIM [Bru90b]. Parsing [JC94b]. Part [BV96, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Cia88d, Cia88e, Cia88f, Jon96, Zim96, Sci86, AM93b, Mes97a, Mes97b]. Partenkirchen [SEA84]. Partial [Ano94-100, BS94b, BS94a, CSSY92, EAGEG09, Gal96, GRSS93, GF90, MT96, Wat91, WS93, YKK96, Cha90, CG87, DGL89, LMD98, Pet89a, Pin91, Scr88, TFB94a, TFB94b]. Partially [RAP95, CH87, Che90c]. Participation [Ano97b]. ParticLE [KDP^{+}14, ASSW93, ACSH90, BWO96, BD93a, GGW93a, GGW93b, Gre89a, Kel91, Koh96, Man90, MMRL93, MR90b, Nag96c, FMD07, Fuj11, LLDF95]. Particle-in-Cell [ASSW93, Man90, Fuj11, LLDF95]. Particles [RRSG96, Soe94, ARF12]. particular [CCC^{+}89, Kah92]. partioning [Sar91]. Partition [CB00, HL96]. Partitionable [NMS93]. Partitioning [Ano94-57, Ano94-93, CTD^{+}16, Gal96, GP88, JP94, WF93, KFML20, Pol88d]. partitions [BBWR90]. Partly [SS09]. Partners [Dal96, Str94]. Partnerships [Ano96t, Ano97k, Dau96, Dau97, Spe97, Ano98f, Fed96]. Parts [PPG94]. party [SSP93, WLN^{+}96a, WLN^{+}96b]. Pasadena [Ece96]. PASCAL [Tsu91, MW82, MT91, PK80]. Pascal-Plus [PK80]. Passby [Wil94]. Passenger [AC93, LB94b]. Passing [ABBB94, DS96a, DHHW93, DFWW93, GB96, Gle93, HPLT01, IHIS91, PR94a, SYG94, SABJ94, SSOH95, Sul97, VSM96, YG92, AAC^{+}05, BCM94, DLM99, MRM87, Saa87, SWJ95, CO94]. Past [Bro17b, DLPQ94, Els02, HF93, Fer84, Hey90]. patent [Cal11a]. Path [LZF16, AGL11, BJZfDA96, TYZ85]. Paths [DMW87, BJV^{+}16, Hsu16, NS88]. patient [PMS^{+}08]. Pattern [DB95, Hun92, KKKP93, Kok94, SD93, SBSR96, OMM93]. Patterns [DB94, KKPR93, MF92, MK07, VT95, XCLW93, PB94a]. Paul [Hil97, MM94a, Wen94, McD88]. Pavement [COC93]. Pay [HWG98, Lew96a, Lew96b]. PB [CDW94]. PB-BLAS [CDW94]. PBS [Cla18]. PC [Ano88x, Ano97l, Ano97v, BS00, Chexx, EKTB99, Fri94, MP84, SSBS99, Ste00].

PC-based [BS00]. PCB [Guo94]. PCBFC [TK93]. PCG [JC94c, Nat86h]. PCs [HHS01b, Mac96c, Rya90]. Pda [Ano94-101]. PDE [CMF94, MCB^{+}01, MRL^{+}17, Pet89a, VWC96]. PDE-Based [VWC96]. PDEs [BTV96, BV96, Gri92]. PDG [KKF96]. PE [KSM^{+}19, HPLT01]. PEACE [BNSP99]. Peak [HS94a, SH93, SH94b]. peer [Man89a, Man92]. Peering [Wit89]. Pell [Ked92]. Pellets [KRVJ94]. Pen [BDRR94]. penalty [Lil88]. pendent [Gre88c]. Penetrating [Ver97]. Penetration [CNGR90, Nor97b]. Pennsylvania [ACM96, Sha89]. Pentadiagonal [HL96]. Pentium [Ano97m, Ano03a]. People [CCKSS90, IS95]. Perception [Poe95]. Perceptron [RPY94]. PERCS [RAG11]. Perfect [FR91, Poi89, Use93, Rau91, Ber89b, Ano91n, Ano91o, BE92, Blu92, CKPK90a, Cyb90, CKPK90b, Cyb91a, CBHS91, Cyb91b, Eig91, Poi90, Rau91, SSRL91, VSH91]. Perfect-Benchmark [Eig91]. perfectly [Gib01]. Perform [Has84]. Performance [APK^{+}12, Abr94, ASK85, AS98, AP93, Ahm92, AAB06, ABBB94, ALPP00, ACSH90, AF97, Ano88j, Ano94h, Ano94q, Ano94r, Ano94-34, Ano94-31, Ano94-51, Ano94-54, Ano94-60, Ano94-61, Ano94-62, Ano94-70, Ano94-66, Ano94-69, Ano94-71, Ano94-105, Ano94-102, Ano94-103, Ano94-104, Ano94-96, Ano94-110, Ano94-114, Ano09, BCH12, AYL^{+}18, Ara97, Ata91, AT93a, AT93b, BGMR96, BGS94, BKK11, Bae01, Bai92, BLW11, Bak10, BCC^{+}08, BBH95, BGS^{+}12, BCC^{+}09, BK97, BS98a, BEK02, Ber07, BGM^{+}11, BS92, BHLST94, BBHL01, BJS02, BE92, BEH^{+}94, BS01, BIB^{+}18, BD94, BCH^{+}22, BCHJ94, BH17, Bro00, BEGGK07, BGH^{+}02, BNSP99, Cal81, CC96, CGFT05, CC94a, CGSG94, CCYT05, CH89b, CP94b, CLY^{+}19, Che99, CH10, Chi95, CDPW94, CFS95, CMF94, CS90, CB02, CDS98, CMAS11, DDHK94, DD05, DCWH07, DBK09, DTV00]. Performance [DDT95, DS96a, DI88, Don85, DKH86, Don91, DSSS05, DMT^{+}21, DVWW05, Ede94b, Eig01, EGJ^{+}02, EBS02, EAGEG09, EAMEG11, EGEAH^{+}08, EDJ^{+}10, Els02, EHG01, FBZ92, FDM07, FT96a, FCD97, For02, FJSD96, FXAC94, Fos96, FGKT97, FBJ^{+}94b, FGG09, FLP^{+}07, FHM99, Gal88b, Gal89b, GB90, Gar01, GSG^{+}94, GS01, Gen97, GA95, Gle88, GCY^{+}08, GCS94, GRRM99, GA84, GW93c, GMSB93, GK93, GS94e, HMM94, HPH^{+}20, HS94a, Hag01, HL93a, HC99, Har89, Har94a, HR94, Hel92, HNS94, HB08, HFCM98, HAAS93, Hoc91, Hof94, HG02, HP04, HMC94, Hol95, HY92, HS93c, HNS^{+}10, HERC95, HPLT01, HHK19, HML^{+}21, HW96, IEE93b, IEE94c, IHE^{+}00, IH94, IBBA20, IM96, IK91, Jar12, JPMG08, Jon96, JML96, Jor87, JY92, JCJY94, Kah94, Kah97, KN88, KMKD97, KB19, KBG^{+}13, KH98]. Performance [Kha95, KK95a, KTG08, KWB^{+}10, KT11, KLM94, KCG08, KC93c, KRS13, KSB^{+}19, KSM^{+}19, KBLD08, KZ94, KR94d, LL08, Lan92, LM08, Lat16, LC97a, LLGS09, LMM^{+}21, LBT94, LTD^{+}93, LCP^{+}11, LHPG21, LYC93, LMM86, Lum01, LCD97, MM93a, Mal90, MSAD91, Mar95, Mar88c, MCW98, McB92a, McB93, ML97, MDH^{+}16, MB12, MC10, Mes97a, Mes97b, MLY10, MBD99, MBW01, MTLL94, MBSK92, MR95, MBK^{+}92, MRGR12, MBSW01, Mur06, Mur07, NH91, Nai94, NGLPJ96, NdMM09, N^{+}95, NBKP95a, NK94, OD01, OPR01, OT07, OW94, Pap16, PH11, Par90b, Par86, Pel93b, PW05a, Pin01, PMP21, PL94, PTC^{+}93, PHVJ95, PSO12, Pro01, RMPW93, RRMD94, Rep92, Res01, RS93, Rot94, SSG93, SCG^{+}08, Sak02, SNS^{+}97, SWG06, SPM^{+}10, SEH98, SW10a, Sar90, SYG94, SBZ^{+}08, SES94, Sch88a]. Performance [Sch94b, SH93, Sch19, SH91, SKC02, SBHW80, SkLC^{+}03, SCSL12, SZ11, SE92, ST92, Smi96b, Smi96c, SP12, Smi92, SW10b, SA82, SJDV09, SDK98, SS09, IEE94d, Ste94e, Str10, Str97, Str03, SMDS15, SSGH94, SE98, SLS96, SHB^{+}13, Tan89a, TGV08, TPJ^{+}19, TV88, TMP94, VWC96, Van94, VS99, Voi94, WKL^{+}16, WN10, WP94, WG93b, WOG94, WGW04, Woo05, Wri19, WT11, WT13, YSS94, Yan93, YGSB94, YHA93, Yi11, ZWP03, Zim96, ZW02, Aba09, AMS^{+}15, AGEL13, AB01, AGZ94b, A^{+}02, AB03, AP91, Ano89h, Ano90e, The90a, Ano91h, Ano91x, Hig92, Ano93-39, Ano94-122, Ano94-123, Ano95u, Ano96c, Ano96r, Ano00c, Ano00d, Ano03a, AB96, Ara14, AKM^{+}06, ABMN02, Bad99, Bad04, BA08, BBC^{+}99, BP91b, BG02, BY21, BBE^{+}20, BP86, Ber89b, BSJ^{+}13, Bhu95, Bis94b]. performance [Bli89, BBM19, Bor93, BL91, BGKR99, BJV^{+}16, BPD06, CLB19, CP25, Car91, CWD^{+}08, CBB^{+}05, CV92a, Che93b, CV88b, COT21, CB99, CRA10, CDG^{+}06, CKD^{+}19, CBM^{+}05, CKPK90a, Cyb90, CKPK90b, Cyb91b, Cyr86, Dak90, Dam11, Dan91, DHA^{+}13, Dem91, DK01, DH91a, DH91b, DRAB08, DF12, DS86a, Don86, DMW87, Don93a, D^{+}95, DvdS12, DGH20, Don93b, DRSS99, Duf00, EE93, Eig92, EMS11, EHF^{+}97, EFR^{+}05, EWS^{+}13, EM78, FEK20, cFM07, dCCF01, Fox98, FP00, Fuj11, GFBR10, Gal91, GJW91, GCP90, GGJ89, GH94b, GH94c, GMF00, GZE^{+}05, GCS20, GMSS^{+}11, GKSR14, GYL00, Gra93a, GV91, Gra91, GV96b, GL96a, GL96b, GL97, GAB^{+}96, HM93a, HW11, Hag90, HP91, HLDS95, HPPF94, HBKR96, HS95c, Ho92a, Hog02, Hsi91, HY91, yHY92, HP95, HEB96, HR04, IEE96b, IEE97b]. performance [IEE97a, Inf86, Ipe19, IS20, KHS88, KG98, KK90, KBVH14, KFML20, KSM^{+}08, KHBB01, KBM^{+}02, KMB09, KG03, KFN02, Kos89, KW11, KS87b, Kum91, KAMB19, Kwo87, LAdS^{+}15, Lav89, LS92b, LW11, Lid96, LCHS96, LMY88, Liu12, LAL02, LG03, LLSR02, LM90b, LKJ03, LSK04, IJS94, MD04, MBM^{+}20, Mal86a, Mal86b, MP88, Mal91, Mar88a, Mar96, Mar90, MMW86, MSW^{+}05, MI01, May01, McB92b, Mec95, ME91, MUKX06, MMG^{+}18, MMG^{+}00, Mye92b, NRM^{+}09, NP90, NBKP95b, OGO^{+}20, ODAZ15, Pap97, Par90a, Pei17, PSS^{+}19, PGK^{+}10, Poi90, PW23, Por89, RAG11, R^{+}00, RCB03, Riv90, RGL^{+}15, Row86, SCV01, SKB^{+}20, SSS92, SEH99a, SEH99b, SKS04, SFC^{+}21, STH^{+}98, SH94b, Sch90a, SZG95, SEV^{+}09, SD92, SC20, SC04, SL99, SWL^{+}91, SWL^{+}92, Sim00, SWJ95, SHB91, SHL^{+}20, SW99]. performance [Ste94f, SDMS99, SS07, TTD^{+}11, hTD88, TF15, Tho90, Tri95a, Tri95b, Tur89, Van13, VdSK^{+}05, Vet12, Wal85, WWJ09, WFJ^{+}17, War03, WSL88, Wat87, Wil88a, WMK90, Wil96, WHL93, Zec93, Zen99, ZS94b, ZS94c, Zor92a, dRC94, dC94, Bra94, Edw97, FJSP95, GBK^{+}96, Lid99, GH94a, HS95b]. Performance-Aware [CLY^{+}19]. performance-evaluation [Cyr86]. performance/cost [AP91]. Performed [HS93b]. period [Joh86c, TR86]. Periodic [SE90]. Peripheral [Has84]. Perl [DDJ98a]. permutation [Lee87a]. permutations [FJ91]. Perrin [Arn89]. persevere [Ano92-42]. Personal [Ano91p, Ano95z, BBWR90, Fri91, Hir92a, Hir92b, Hir92c, IAIK92, MAT85, Pou86, Smi96d, DDJ98b, Ano93d, Don93a, SKB89, Sha89, Smi96a, Voe89a, Wal09]. Perspective [BCC^{+}08, Bel98, Gha96, Gup94, Hay84, Woo96b, Ano89d, ACK^{+}95, CCKSS90, DDC96, Don93a, GE12, IEE89b, Som13]. Perspectives [Ewa89, LPC^{+}95, PC94b, RLC91, WG93b, AL92a, Asp93]. Perspektiven [Kro92]. persuasive [PA92]. Perturbation [MTK93, Rie93, Mal91, RLKW93]. Perugia [Lag89]. pervasive [PA92]. peta [KNHN16]. peta-scale [KNHN16]. Petaflop [GIF^{+}12, GKS09, IBM01a, IBM01b]. Petaflops [Bai97, SMS95, CSFS00]. Petascale [Bad08, CYXL18, OYK^{+}14, TVT^{+}16, HPH^{+}20]. PETASYS [AI92, CP92a, CP93a]. PETASYS/TERASYS [CP93a]. Petersburg [AGP96, GP93c, IEE85, KK85, L^{+}93, Lim93]. Petri [GVBC95, Her94]. Petroleum [SPS90, WG82]. petrophysical [BRL^{+}20]. PETSc [BBB^{+}20b]. PGAS [AGY^{+}11, SWS^{+}12]. Phantom [HEJM95]. pharmaceuticals [MKHY95]. PHASE [AK93, BY96, BCCG97, CMF94, LUT96, OLWW94, Pop97, RWNJ94, Saa93b, SPGD98, TKM96, Ano02a, Ano02b, LXW^{+}16, Tze88]. Phase-Rotation [OLWW94]. Phase-Tolerant [RWNJ94]. Phenomena [AM93b, HMS93, MS96, Nat84, Gro92a, Sug80a]. Philadelphia [ACM96, EM94b, Sha89]. philosophies [RYYT89]. philosophy [Wor81]. Phone [WMMC10]. phoneme [McD90]. Photo [CTM94]. Photo-Realistic [CTM94]. photodiodes [PIH04]. Photon [MNR86, BLFT84]. photonic [Suh97]. photovoltaics [Lam14]. PHWRs [DB94, JJYL94]. Phylogenetic [MOOK94]. Phylogenetics [MBW01]. Physical [Bel92a, BS01, Cha93, FC93, FI93, GP93b, Hel96, LC93, Man90, Pet97, TK93, TGV08, WBP87, BHD^{+}05, Sti98a, Sti98b]. Physical-Component-Boundary-Fitted-Coordinate [TK93]. Physical-Space [FI93]. Physicists [Mor92c]. Physics [Ano88n, Ano95-34, BB90, BBM96, DMKW93, GT94, HK93a, Hes90, JA92a, JA92b, MI93, MR90b, NAAW97, WBP87, BMR85, MSK^{+}02, Ric90b, Ric91a, SN96, Wie94, Ano95-34, Ano96t, Ano97i]. physiology [Wit89]. PIC [Par90c]. Picard [Ske87]. picks [Ano95-37]. picoseconds [MS84]. Picture [Ano96x, Ano96u].

Picturing [Pic89, Pic91a]. Piece [Ano92-30]. Piezoelectric [KP94]. Piezoelectric-based [KP94]. pill [Ano08b]. PIM [GHI95, IGH95]. pin [Fid90]. pin-optimal [Fid90]. Pinch [CCKSS90]. Pinion [DDF93]. Pioneer [Bar00a, Bar00b, Lea09, Ano98d]. Pioneers [DDJ98b, Wei88]. PIOUS [MS94d]. Pipe [Cha94b, RHH96, SW96, Ano96u, WH94]. Pipeline [Ano94-44, RL77, VSM96, Ano93u, JS86, RL78]. Pipelined [KK95a, MSAD92, OLWW94, WS84a, WS90, Gao86, Lil88, LY90a, RR89, TK89, WS84b, WS84c, WS87a, WS87b, WS87c]. Pipelines [VSB^{+}21]. Pipelining [Che89b, CDR96, DF90b, JS86, KS86b]. piping [Ng95]. Piranha [GK92]. Pirelli [DM96a, DM96b]. PIRUN [US01]. Pisa [LP90]. pitfalls [DMW87]. Pittsburg [Ras91]. Pittsburgh [And89, EP 97, Gro90, IEE96d, Pit88, AW93, Ano98f, Bac88, Wes96]. pivoting [CG87]. Pivots [AZ99]. Placement [Geo94, Rav92, Rav95]. plain [NSW08]. Plan [Ano89m, Ano90s, Ano95v, Ano96-28, KS87b, Nat91a]. Planar [MSGW94, Rul93]. Plane [DSB96, NPS93, Ano96u]. Plane-wave [NPS93]. planet [Ano96z]. Planetary [Ano97m, BK91b, GS87b]. Plankton [Pli97]. Planning [Ano88i, Ano94w, DHLS97, Sat93, Ano94-121]. Plans [Coc03a, Coc03b, Ano94-119, Ano96j]. Plant [Kul94, PNK93, SKAT93]. Plants [ESMH93, FNT93, KTKK93]. Plantwide [ZBLZ95]. PLAPACK [vdG97]. Plasma [BJLW95, Sch93b, SJDV09, IAIK92, LG03, Ano95-34]. Plasmas [ACSH90, YMT93]. Plastic [HTI93]. plastics [Ano02a, Ano02b]. Plateau [Ste01a]. Platform [FH95, KHMD94, MC10, Mit98, VVKB96, VWC96, AUW08, BAD01, CBKA09, IBP^{+}05, Way96]. Platforms [GD97, SE92, KFML20, MI01]. Plausible [Kon93]. Play [DD05, Per87, Vol89]. player [Per02].

Playing [Coc02a, Coc02b]. PlayStation [KBLD08]. Plenum [Ano00a]. PLIM [Saa93b]. PlotTool [Wom90]. Plug [DD05]. Plug-and-Play [DD05]. plugin [MWRK18]. Plugs [Bed93]. plus [Tze88, PK80]. PMCommunication [STH^{+}98]. PMD [Che99]. PMDO [KGKa93]. PMS [CFH^{+}01]. PN [LMM93]. Pocket [AFF93, Chu89, Ano93d]. pocket-size [Ano93d]. Poincaré [CR94]. Point [Ano94-91, Ano94-111, Bal93, BBD^{+}08, Dun92, Gol91a, Gol91b, IHE^{+}00, MD88, PK94, Wic92, Ano94-122, Ano94-123, Ano94-135, Ano97s, CBB^{+}05, Eij90a, Eij90b, KMB09, PK89, RS94b, Wei91, DM88a, DM88b, WTC^{+}02]. point-to-point [Ano94-135]. Pointer [MT91, MH96, Gua88d]. Pointer-Intensive [MH96]. pointers [Gua87a, Mar92]. points [BB91b]. Poison [JJYL94]. Poisson [LG87, Lil89]. Poland [BBM96, Elm95a]. Polaris [PE95]. polarized [BDM94]. Policies [MP92, Uni86b, Uni86a, Uni86c]. Policy [Ano88p, Cra91, GV96a, Waz89, Aba09, CV92b, Per83, Uni98]. Political [LG97]. Pollutants [Fie93]. Pollution [FA93, SLS96, Zla01, ODAZ15]. Polygon [CO94]. polygonal [ST90]. polyhedra [Wil92b, Wil92c]. Polymere [KJ94]. Polymeric [Gal93]. Polymers [Bar00c, Bar00d]. polymorphism [GR91]. Polynomial [KKB92, Nan86, But92, EGK87a, FB91a, GS90, RW89, Sch87a]. Polynomials [Cla96, ARW92, Ked94, SA10b]. polyvariance [LMD98]. Ponders [Coc03a, Coc03b]. Pool [SES94, TY96]. Poor [CFH^{+}01, Cap96]. poorly [Win02]. POPLAS [KA91]. POPLAS/FEM5 [KA91]. popular [DMW87]. Population [Fry97, Ons88, Pli97, Kha91, RCS21]. Population-dynamics [Ons88]. Pore [MKND97]. Pore-Scale [MKND97]. Porous [PH97, PC97, WABD97]. Port [TM94a, TM94b, WK95, YYK93]. Portability [Ano85b, DMT^{+}21, HWS^{+}88, Hir92b, BBE^{+}20, GFBR10]. Portable [Ano94-51, Ano94-58, Ano94-64, Ano94-74, Ano94-106, BK95a, Dem91, DLMW95, Don91, GW04, HERC95, LMT95, MMRL93, SSKR97, WW92, ABMN02, Bis94a, BCH^{+}93, GG88, KA96, MRM87, Pry94, Yan90c, AKG87]. portably [Rau91]. Portal [Pro01, RW94a]. Portals [CLB19]. Porting [ARE95, BM93a, CM95, DFS93, EAMS95a, EAMS95b, MWO95, PLS20, WLN^{+}96a, WLN^{+}96b]. Portland [IEE93d, USE90, Bor93]. Posed [BM93a, LHLM95, SA10a]. Position [DNV93, RDHC94]. Positionspapiere [Kue92, Reu92]. Positron [HEJM95]. Posits [PMC22]. POSIX [IEE95a, Coc03a, Coc03b]. Possible [Gie96]. Post [KK96b, NG92, Car88]. Post-Mortem [KK96b]. post-packer [Car88]. Post/Wait [NG92]. Potassium [KW95]. Potential [Ano94j, BBL95, BM96, Sat93, XMR92, Din91, Par90b]. Potentials [Ano94-52]. Pounds [Ano94p]. Power [Ano94-92, AAS88, BPU94, CGFT05, CLPV93, CCKSS90, CBT91, ESMH93, FS93a, FNT93, GCB92, KC93a, Lew17, NM93, SWG06, SB01, VLA92, Yei92, ANS92, Ano87d, Ano88f, Ano93v, Ano94-63, Ano94-82, Ano94-121, Ano96n, Ano96u, BSJ^{+}13, BBWR90, Cou11, De 96, Deg90, DZM^{+}13, HPS88, KFN02, MWRK18, OW94, RMM20, SKS04, SB18, SCG^{+}13b, Uni91a, Zor92b]. Power-Aware [CGFT05]. Power-of-Two-Strides [VLA92]. Power2 [HF94]. POWER7 [RAG11]. POWER7-IH [RAG11]. Powered [Ris94]. Powerful [MW81, Mor92c, US01, VNB93, Zey91, Art93, Gep01, Sha95b, WZB86, GRSS93]. PowerPC [Wai05, HF94, SW94]. Powers [Ano93m]. PowerTools [Ano93v]. Powertrain [EJ97, GA97]. Poznan [KNS97]. PPFS [HERC95]. PR [AG90, Asl91a, CGS91, CGL92]. PR-1 [AG90]. PR-2 [Asl91a, CGS91]. PR-3 [CGL92]. Practical [B^{+}89, HA93, LC97b, Meu89b, PK87, SLB93, Ste96, Wil95, Ano96l, BH92a, EGK87a, EGK89b, KA92, KA96, McC92]. Practice [Ker94, PMP21, Zen94]. Practices [Kho94]. PRAM [LZ95]. PRAM-Programs [LZ95]. Prandtl [GW93b]. Pre [LSS^{+}20, PH97]. Pre-Asymptotic [PH97]. Pre-exascale [LSS^{+}20]. Precision [TKI93, Ano87d, Roj19]. Preconditioned [Ano94-45, DL90, HFH86, HFH87, KLY94, Nat86h, Saa93a, SO95, VAGRMVA90, Yan92, And88, Bau88, Bra89b, HVY91]. preconditioner [Fuj99]. Preconditioners [Ma99]. Preconditioning [HO92b, Man91, Saa88, CH98, Nan86]. Preconditionings [KJ85]. predator [Smi91]. predict [PZ89, Str11]. predictability [GRRM99]. Predictable [PH11]. Predicted [CP94b]. Predicting [Abr94, Mir88, Raw97, SSG93]. Prediction [Ale90, AGKT02, Ano94-96, BM93b, Che90e, DBK09, DXJM93, FBZ92, Hai97, Hof94, IK93, KGS93, MKRI93, TAKB06, WS84d, XCLW93, ASK85, Ata91, Bur91, CHWW13, Che90d, Gal89b, Gle88, MV16, Sar90]. Predictions [Jon19, WB85, ZX95]. Predictive [KSTB94]. Predictor [TAKB06]. Predicts [Bar00a, Bar00b, HH93]. Preface [Ano20, CGR05, HF94]. Prefetch [Ano94-127, BCK13, Lee87b]. Prefetching [Dra95, NGDH96, OP96, Bre87, Gor89, GGV90, LYL87b]. prefine [KK89c]. Prefrontal [KDBG95]. Preliminary [EHG95, Gal91, GK92, Men84, Par90b, YJD93, Pad89, SWS^{+}12]. Preloading [CMHK92]. Premixed [HVSB93]. Prepare [Lat16]. Preparing [HCD^{+}18, ODAZ15]. prepass [CLmWH91]. Prepattern [HKG90]. Prerequisites [Har94a, Poe95]. Presence [WN92, Wea97, Gua87a]. Present [Bur93, Els02, Ano96w, Fer84, Hey90, CCKSS90]. Presented [Pin99, Ano95-38, HBCN95, L^{+}95, MB93, MB94b, Pan96, Sha89, Uni87c]. Preserving [IJY^{+}14, KP96, PW23]. President [Age05, Ano97b]. president-elect [Ano97b]. Presidential [Per83]. Press [Ano96c]. Pressure [BJLW95, Hai97, RHH96]. prestack [Tze88]. prestigious [Ano87d]. Prevention [DM93, Tan87]. Preventive [CCR11]. Previews [Ano95-34, Ano97l]. prey [Smi91]. Price [Ano94p, Jan96]. priced [Ano89l]. Pricing [CCZ93]. primal [CK90, Kor93]. Primary [TY96]. prime [BtR95, Tem88]. primes [BY88]. primitives [Wij89a]. Princeton [COS89, LCV90b, LCV90a, Ano87a, Ano87c, Ano94s, TCJS93]. principle [Jor86]. Principles [Eis95, SSJL94, SBN82, BF92]. printed [Lie90]. Priori [MRL^{+}17]. PRISM [Bis93]. Privacy [IJY^{+}14]. Privacy-Preserving [IJY^{+}14]. private [Str94]. Privatization [Li92, RP94]. PRIVATIZING [RP94]. prize [BBD92, Ano97b, KHHS95]. Pro [Cla18]. Probabilistic [AH93, Ano94-95, KGKa93, LD93b, Sol93]. Probabilities [Ano96t, VMS93]. Probability [MBN93, OMR93]. probation [All93]. probation/parole [All93]. probe [Ano92-45, SS07]. probed [Ano90l]. Probes [LD93a]. Probing [Nor96]. Problem [Ano94-59, Ano94-141, AM93c, CJ93, CGW05, CDR96, DGJG93, FBH93, GNJW93, Gha96, Iwa92, KS94a, SF93b, BS88a, BGT90, Bra88, Bru91, BY88, DS86b, FGM90, GKR91, GJG88, LPS86, LP86, Nag88, PB88, SL90, SW99, Gur88, Pop92]. Problem-Oriented [Pop92]. Problems [ALM93, ALPP00, Ano94v, Ano94-110, Ano94-112, Ano96e, Ano96f, BK95a, BL93, DD93, FD93, FR98, FI93, Geo94, HHGS93, Ike95, KJ85, LHLM95, Li95, LS93b, ML93, MRAR95, MR90b, PRS94, PC93, RDZ93, VTTS98, WGOY91, Ano93u, Ano96l, Bra92, BJ84, CP25, CS88, CS89, Che94a, COT21, CFP20, DGL89, DHD89, Duf90, Duf00, EM94b, Feo92, GBS18, Gra93a, GKL^{+}87, HOSZ97, HHS01b, Ho92a, Ji91, KK93, Kin96, LD90, Lou92, Luc09, NP90, Pet83, PZGL91, Poo96a, SBC91, Sha95b, dRC94, dC94, FJSP95]. Procedural [Max81, TS90]. Procedure [JCJY94, AC91, Ked92, Sit78, Sta88]. Procedures [BJS02]. proceeding [VAS82]. Proceedings [ACM89a, ACMxx, Ano93n, Ano97q, Asp93, DT97, DP91, GL90, Gro90, HS94b, IEE90, IEE93b, IEE93c, IEE93d, Kow89b, KK92, Lag89, ML89, McC88, Meu90, Meu91, OMM93, Pit90, Por86, RD94, SEA84, Soc94, Uni91a, USE00a, USE00b, USE01, VO93, Zyg93, ACM90, ACM91, ACM92b, ACM93, ACM94, ACM95a, ACM95c, ACM96, ACM97, ACM03, AU87, Ano88t, Ano88u, Ano90f, Ano94a, Ano96l, App96, BG91, B^{+}89, BP89b, Bro93, Bup87, CL91, Chi90, C^{+}97, DDC96, DJM94, DLM99, Dup86, Dup87, EP 97, Emm85, EM94b, Fer83, Fra94, GG^{+}97a, GH94b, GH94c, Gra94, HKR94, Hel93, HS95c, HK93b, HHK94, HBCN95, IEE85, IEE93a, IEE94a, IEE95b, IEE96d, JPTE94, KK85, KK87, KK88, KK89a, KH93, KK93, Kho94, KMG96, KWW92, LLR93b, LCV90b, LCV90a, LCHS96, L^{+}93, Lim93, Lun94]. proceedings [M^{+}94, Mar86, Mar88b, Men87, Met86a, M^{+}95, MP92, MA85, Uni87a, NBC92, Pel93a, PEH93, SF91, Sig89, Sig90a, Sig95, TC94, USE90, Vag88, ZAS94, ACM95b, ANS92, Ano91q, Ano94-107, Ano94-108, B^{+}95, Cha94a, CBCH93, GT94, Ham94, IEE94b, IEE94c, IEE94e, IEE95d, KSW93, Meu89a, Sie94, S^{+}93, IEE94d, SR93b, Tho93c, GH94a, HS95b, HPP88, ML95b, VV95]. proceeedings [ACM88, ACM92a]. Process [CWLT97, FT96a, FCD97, GL89, GM93b, JML96, KWH94, KCOP94, MDP^{+}00, Nag94, Pel94, Pin01, Sch97b, Sta94, Wea97, ZBLZ95, Har89, Nat91b, Nat92b, Pol86, SF82, Shu88, ZMDS96]. Processes [GSG^{+}94, GMBW93, GM93b, Kaz93, KB97, LS92a, Maa93a, TSSK94, War93b, Ano97y, Ano00a, Sha90]. Processing [Ano89b, Ano90m, Ano93n, Ano94-107, Ano94-97, Ano94-130, Ano95p, Ara97, Ash93, B^{+}95, BMSD94, Bor89, BH17, Bro97, CTM94, CBCH93, DM88a, DM88b, Eck93, Fet95, FR81, FB91b, Gan94b, GJP94, GHI95, Gol99, GP93c, Hay86, HCL94, HAG^{+}13, HYL^{+}20, IEE94a, IEE95b, IHE^{+}00, IGH95, JC94a, KN88, Kue93, KS90, MD88, Mur91a, NMS93, PRSS94, Rui91, SH90, Sch93a, SKSD94, Sie94, S^{+}93, Ste95, Uni87c, WMBC97, WN10, Y^{+}92, YWD94, YWDxx, ZLC21, ZM94, ASM86, AGP96, Ano88s, Ano92w, Ano92y, Ano93v, Ano94-122, Ano94-123, Ano99, Ara96, Bar88, BF92, Bur93, Bur94b, Con87b, CP25, Cho90a, Don92a, DWV92, EMS11, FL92, FMT91, Goo97, HD89, IEE96c, Kha93, MBM^{+}20, McC94, Mil93, MU83, MSW91, MMG^{+}18, Mor92a, Pit89, Pra95, Rei88, RR89, Ros95, Sca92]. processing [Sch89b, SMM17, Sie90, SA90, WHBH93, Supxxa, YFY^{+}13, Zag82, CKS99]. Processor [Ano94-84, BK77, BBD^{+}08, EHG95, GMSB93, HHT^{+}94, HMNN91, HHOM91, HHOM92, Kue93, Lil91, MHW94, MDH00, MBSK92, NKTT95, Par86, PB87, SCV01, TF92, VPDA93, Web93, Ang91, BJ95, BH95, Cal85b, EY91, Fan87, GJM86, Gok92, Hae91, HT72, Loo84, Mal88b, MS88, MHP84, PJ90, Rob87, Sam85, SS10, Tan87, TS88, VSH90, VSH91, Vaj91, VFK^{+}04, Wat72, DFSZ88]. processor-based [Rob87]. processor-in-memory [Gok92]. Processors [BB94, Bra93, DD99, FT96a, GS94a, GS94b, GL94, HK93b, KH93, Kun84, MSAD92, MBK^{+}92, SES94, SJDV09, Wil93, tDv87, Bau96, CLmWH91, CLF^{+}19, CKS99, Int92, Jor87, KB96, Lil88, LY90a, NRN00, R^{+}00, Sch12, SCK^{+}00, SY91, Wei91, Wei92, Wij89a]. procurement [Uni92c]. Produce [OLLG96, Spe97]. Product [BMSP94, BKM93, GM93b, HM93c, Ker94, Mil17, RSB94, Sei94, Wea97, Ano96u, Fuj99]. Production [Ano94-37, Bae01, CT93a, DHLS97, EGH^{+}06, Sch97b, TGL96, UHU09, Wil93, XOZ^{+}20, And90a, Elm95a, MSPPD20, Stu95, TDBL13, TRLD13]. Productions [Wad86]. Productivity [Ano94-107, Bla93, Das94, KSM^{+}19]. Products [WWKR97, Bab94]. Professionals [Cra96]. professor [Ano92-42]. Profile [Kop91, TD96, IEE91, PF90, IEE95a]. profiler [SSS90]. Profiles [BGS^{+}12, Lea09]. Profiling [VSB94, Ano88s]. profit [CBLS13]. Program [AHFK93, ACG^{+}90, Ano87a, Ano91j, Ano94-60, Ano96x, AVS93, BE93a, Bra93, Bur94a, CH94, ORS94, Emr89, HMC94, HH93, IM96, ITOK93, Kah93a, KA91, LB93, MG95, OHIB93, ORSS94, OHHIB94, PDR91, PZA87, Roh94, SG92a, SS95, TA94, WF93, WWY93, ZX95, Amm89, Ano94-100, Ano96-37, Bin88, BM85, BH92a, Bre87, CGLY96, CGLxx, Chexx, CRM94, CDG^{+}06, Deg90, FP91, HP92, HA92, Kel85, KB18, LY88a, LY88b, Lim91a, MLR90b, MLR90a, MSTK93, Nat86d, Pad89, Pol86, SNS^{+}97, Sch89b, Tan89b, Tri85, TJC91a, Whe83, BS92, Joh94, Web91]. Programmable [Ano94-127, GHK^{+}91, MS94a, CKM88, Gok91, Kar89, ML95b]. Programme [Ano94-33]. Programmer [DDJ98a, Sny99, Guz87]. Programmers [Coc01, PB90]. Programming [Ano94z, Ano94-28, Ano94-51, Ano94-58, Ano94-74, Ano94-93, AM93c, Ara91, AK94, BAAD92, BK95a, Ber95b, Bha94, Coc02a, Coc02b, Dip96, FHM95, FH95, FI93, GJ87, GBK^{+}96, HC99, HQ91, Hop93, KBM^{+}02, KB94, KHV11, Mah94b, MMRL93, NB92, PL91a, PL91b, PBM95, PT93, PK94, Rag06, SPM^{+}10, SBF94, SL99, Sit78, Ste94a, Sug94, TP93, VW95, Wil95, WL96, WLN^{+}96a, WLN^{+}96b, YSL97, ZM86, AGY^{+}11, Ali86, Bea90, BM85, Chu91, DCG90, Ele93, Feo92, Gil94b, Gok92, Gua87b, Gua88b, GG88, JG88, KK82, Kor93, KA96, LW11, MGS^{+}20, Mil93, MMG^{+}00, MAA^{+}05, NRN00, PGL87, PK89, RR99, SW91, SMR10, TS90, YQTV12, Hil97]. Programs [AK87, Ano94t, Ano94-106, Bie88, BC95, Cla96, FBZ92, Fin82, HMM94, HLB94, KRS94, KBC^{+}74, Mah94c, MB94a, MH96, OS94, RAES96, Rue92, SYG94, SABJ94, SPK94, TBC94, TCF94, Van94, WZ97, WNKS96, WB85, Ana91, Ang91, Ata91, Ban79, Bli89, BE92, Blu92, CV88b, Cho90b, CH92a, CH92b, Eig91, Eig92, EGP88, EGP92, EO91, GP88, GP91b, GRRM99, Gua87a, GM87, Gua88c, KPS88, Kim96, LMD98, LR88a, LY88b, LY91b, McC92, MPC89, MP90, NG92, PE88, Par90b, Pol87d, Pol88d, Sar90, Seh88, SKP91, Seh92, SK92, SLY90, Shu88, Uni86b, Uni86a, Uni86c, Yi90, LZ95]. Progress [GS92b, MA85, Smi96b, SMDS15, UEGM93, AG90, Asl91a, CGS91, CGL92, TC94, TR86, Por86]. Progression [IK93]. PROGS [TF94]. Project [Ano97p, Con11, DMT^{+}21, GG97b, Rui91, TP97, Ano91a, Ano93b, Ano98g, AG90, Asl91a, BBH^{+}00, BBB^{+}20a, Bis93, CGS91, CGL92, CFH^{+}01, HP88a, Hey94, Hug94, Jet90, Jet91, Jet92, Mil90, NW03, SPP^{+}05, IBM13b, YK87, Ano98e, AUW08, BOS97, BBC^{+}89, Coc01, Eck93, Hab89, IBM08, KBM^{+}02, MB20, Mes17]. Projection [ABCH97, Bra89c, BS90a, BS90b, Bra92, KS86a, Lun90, MP91a, MP91b, MP91c, Wil90b]. projections [KAMB19]. Projects [JLC98, Pit87, Pop97, Ano98f, Eid91, Gha84, Man89a, Man92, Wen94]. Prolog [KPS88, Seh88, Seh92, SK92, VPGG01]. PrologX [Goo88]. Promise [EGEAH^{+}08, Ano97-29]. Promises [Gar01, Tri95a, Tri95b]. Propagating [BD93b]. Propagation [Ano94-52, DGT94, FJSP95, LMM^{+}21, Mas95, Sch92b, Vaf88]. Properties [BWO96, BS01, GK93, NS88, SSJL94, Pan97, Rob89]. property [WH94]. proportional [Law89]. proposal [Ano92l, Gua87b]. Proposals [OGY91, OGY90]. Proposed [Moh94, Waz89, Ano00b, Sta88]. Propulsion [AJ97]. Prospects [Ano86b, PRS94, Tho93a, Fer84, dRC94, dC94]. Prospectus [Bra91b]. prostate [Law89]. ProSTEP [Ker94]. ProSTEP-Introducing [Ker94]. protecting [CLF^{+}19]. Protection [Ano93-29]. Protein [Ano94-98, Ano95w, Che93a, MKRI93, TAKB06, WWY93, XCLW93, ZAS94, Ano90l, IBM01a, IBM01b, ZAS94]. Protein-Coding [MKRI93]. Proteinases [HGB90]. Proteins [HC93, MW88]. PROTEUS [AK93]. PROTEUS-PHASE-II [AK93]. Protocol [HNST93, KG96, OA94, SC20, SQS^{+}19]. Protocols [Ano94f, CGSG94, GM94b, MH94, RR99]. Prototype [Ano94-78, CGHL94, HHOM91, HHOM92, FGC06]. Prototypes [Mil17]. Prototyping [CPS96b, CPS96a, DGJK93, Gil93, Wea97, Hin93]. protsessory [BKM88]. proud [Win02]. prover [HA90b]. provide [Ano95-37]. Provided [THH81, CS82, CS86b, THH82]. providers [CBLS13]. provides [PF90]. Provisioning [BCW20]. Prowess [Moo11]. Prozessoren [Ano01b]. PSC [Pit86]. PScheD [LG97]. PSCs [DDJ98b]. Pseudo [OGY91, AI92, OGY90]. pseudo-random [AI92]. pseudoprimes [Arn89]. Pseudorandom [Ano94-64, KA92, CMP94, Mas94a, Mas94b, Pry94]. PSIM [RKDM94]. PSL [HA91]. PSM [FI93]. PSO [EB18]. PSO-based [EB18]. Psolve [Dav86b].

Psoul [Sha94a]. pt [FL92]. PTOPP [McC92]. Public [Bar01, Str94]. Public-private [Str94]. publication [Han94]. Publications [Ano85b, DGG18]. Publishers [Ano94p, Ano00a, McD88]. Pugh [Kaz92]. Pulsation [HS93b]. Pulsed [DCW93, SBY93]. Pump [Ren97]. Pump-and-Treat [Ren97]. pumping [Chi86]. punch [Ano94-121]. purchase [Aus93]. Purchasing [Lew17]. Purine [Hei89]. Purpose [Ano94-59, EFIM91, FHM99, MKSF96, YFOT93, Abe91, CGLY96, CGLxx, Chexx, Gup88, Lee86]. Purposeful [SAGS93]. purposes [Sha87]. push [Wal85]. pushes [Ano93b, Ano97m]. put [Ano97u]. Puts [Pau08, Ano90o]. Putting [CSFS00]. PVM [Ano94d, ABGL96, CT94, DKF94, DLM99, FS93b, GB96, KCOP94, MWO95, Sch94c, Sch96, Str94]. PVM/MPI [DLM99]. PWR [HM93b, MTK93, MAA93b, Ng95]. Pyramidal [Eis95]. Pyrimidine [Hei89]. PYRROS [YG92]. Python [BLP^{+}21].

Q [BCK13, ABB^{+}13, BSJ^{+}13, BJV^{+}16, CCD^{+}13, CP13, CEH^{+}12, CKL^{+}13, CHT^{+}13, EO13, EWS^{+}13, HOF^{+}12, KBVH14, LM13, OWG^{+}13, RIB^{+}13, SCG^{+}13b, IBM13a, IBM13c, ZYL^{+}16]. Q2R [ZH88]. QCD [Att96, BCK13, KLN90a, KLN90b, Tou87]. QCDOC [BCC^{+}05, FMD07, FDM07]. QCDSP [BCC^{+}05]. QCE [FNK93]. QED [KDK89]. Qespera [MV16]. QMRCGSTAB [Cha92b]. QoS [KCZJ14]. QR [MP94, MM94b]. quadratic [Arn88, ARW93b, BtR95, BE93c]. Quadrics [FWS96, SPGD98, TCM95]. Qualitative [KSTF94]. Quality [ABCE97, Koo97, KCZJ14, Leg94, LCD97, MD04, Rei93, SSS94, SSKR97, TMAS97, CMP94, HPS88, Mas94a, MSPPD20]. Quanta [Ano96v]. Quantifying [FBCB18]. Quantitative [BHMH98, HP03, JW98, KS94a, TC94, WHL93]. Quantization [BB93]. Quantum [AGLL98, BH17, CS94b, Fox90a, HML^{+}21, LWY^{+}20, Llo94, MCH91, MR86, RK22, SG81, SG82, TW92, WCHK91, ARE95, Art93, BS00, BDM94, Din93, Gui10, MR87, Ric91b, Shi95]. quantum-chemistry [ARE95]. quantum-mechanical [Shi95]. quark [BK89, Din91]. quasars [Din92]. Quasi [SF93b, Cha92b, CH87, Che90c]. quasi-minimal [Cha92b]. quasi-Newton [CH87, Che90c]. Quasi-Stationary [SF93b]. Quasidiffusion [ML93]. Quasimolecular [Gre88b]. Quasirandom [Bro96]. Qubit [NPS^{+}20]. quells [Ano96-34]. Queries [CCSS98]. Query [RS94a, Win02, YFY^{+}13]. Querying [TF94]. quest [Ano96n, DWV92]. Questions [Ano92x]. queue [MV16]. Queueing [SES94]. QUICK [OK93]. quicker [Bin88]. quickly [Ano95w]. quietly [Ano01c].

R [Ano94p, Ano96c, Kaz92, Pin01, BE93a, SB82b]. r$-$INF$/ [Tem89b]. R&D [Ano89n, Ano95v, NW03, Str94]. R&D-100 [Str94]. R-matrix [SB82b]. Race [DKF94, CSFS00, NG92, Sch89a]. races [Per87]. Rack [DDF93]. racks [ARF12]. Radar [Ver97, YWD94, YWDxx, Ano89f, Gig94, YW94]. Radiated [YF95]. Radiation [GHW93, MF92, Rud90, YMT93, Ano96w, Ano97-29, KK93]. Radiation-Magnetohydrodynamics [YMT93]. Radiative [Now93, DLS93]. radical [Bel92a]. Radio [Ano96w, DGJG93, Din92, Nat85]. radio-loud [Din92]. Radiograms [PRS94]. Radiographic [Ano95y]. Radiological [IK93]. Radiosity [RBL94]. RAE [Wil90a]. Raking [BCZ95]. Raleigh [VO93]. RAMONA [WFT93]. Random [Alu96, And90b, BA95, Ent99, Hey96, IK91, LD93a, OGY91, AC91, AM15, AI92, Gut95, Lil89, OGY90, SKB^{+}20, SM92, Whe89, Ano96x]. randomaccess [GS06]. Randomized [CRY94]. randomness [Pic89, Pic91a]. Range [Ano94-109, GJS94, Mil97a, Ano93h, Bur91]. ranging [Ano95l]. rank [Ara14, KTN^{+}14]. rank-mapping [KTN^{+}14]. Ranking [RM96]. Rapid [CP94a, DGJK93, Gol96, HLA^{+}18, Hin93]. rapidly [MKfDA96]. rapidly-converging [MKfDA96]. RAPS [Hop93]. Rarefied [BPJ94]. Rat [SVML95]. Rate [SGIS93]. Rates [JBI91, NSP94, WWKR97, DWV92]. rating [DMW87]. Ratio [YOY97]. Rational [HSW^{+}90]. Rationale [Joh97]. Ray [BKT94, GMMT91, SYMT92, Def87, CDMW94, PB94a]. Rayleigh [GW93b, PB88]. Rayleigh-Benard [GW93b]. RC [CGLY96, CGLxx, Chexx]. Re [JHGLG93, Lee94, Rei93, SPK94, Taf96, Isk96]. Re-engineering [SPK94, Isk96]. Re-Inventing [Taf96]. Re-invention [Rei93]. Re-Sequencing [JHGLG93]. Re-Thinking [Lee94]. reach [Sch22, Wal85, WZB86]. Reaching [BHS^{+}02, Ano95-31]. Reactabot [TSSK94]. Reacting [Ano94b]. Reaction [BCW93, Hun92, MCH91, SM92, WCHK91]. reaction-diffusion [Hun92, SM92]. Reactions [Las92]. Reactive [WABD97]. Reactivity [DCW93, SR93a, Heh86, Lag89]. Reactor [ANS92, ATSA93, AA93, DH93, DMKW93, FBH93, GY93b, IK93, JKNK93, Koc93, KTNM93, LD93b, NS93, PNK93, SKVZ93, SAGS93, Smi93, SKAT93, TKI93, WFT93]. Reactors [Ask93, BJLW95, NM93, ANS92, Kav92]. Readies [Ano91f]. Reading [HK93b, KH93, PRS94, Ano81]. ready [BAD01, Pou94b]. Reagan [Per83]. Real [AC93, Ano87a, Ano94-70, ACA94, Coc02c, Coc02d, Don92b, EFPSS93, EKZ90, FGG09, GSG^{+}94, GMF00, HS94b, Heg96, IK93, KFF93a, KLD95, LMP^{+}90, PCK93, SBZ^{+}08, SH93, Sha90, SB94c, UP01, WOS09, Ano89i, Ano96l, Ano96n, Dan91, LTT92, RGH17, SH94b, Ver95, HS94b]. Real-Time [Ano94-70, Don92b, EFPSS93, EKZ90, GSG^{+}94, LMP^{+}90, PCK93, SB94c, UP01, GMF00, Sha90, Ano96n, RGH17, Ver95]. Realistic [CTM94, CNGR90, FD97, Eig01]. Reality [CMPR93, Coc01, Dil93, DPS97, Gan94a, Hel93, Ike95, KTKK93, Kul94, PS96, TSCG94, Wag96, Gro92b, Rol97, Ver95]. Realizing [Ano94-110, TYZ88]. Really [Ano94-118, DDJ98a, Faz87, And10]. Realtime [Kue93]. Reasoning [Kaz93]. Reasons [RBK95]. reassessment [AB94]. Rebecca [Ren97]. Receives [Coc02a, Coc02b, War10, Ano93b]. recently [Eva97]. Receptors [SVML95]. recipe [Pic88]. Recipients [Coc01]. Reciprocating [AGH^{+}90, PB94b]. Reciprocity [Smi96c]. reclaims [Gui05]. Recognition [BPL95, GD94a, KKKP93, MSGW94, PE95, PW94, SD93, SHA^{+}92, SLML93, UEGM93, ZS94a, AH90, DB95, McD90]. recognize [War03]. Recognizes [Pin99, Pin01]. recommendations [Ano96w, AB96]. Reconfigurable [BS98a, BEGGK07, CH10, DDB^{+}10, EAGEG09, EAMEG11, EGEAH^{+}08, GLS11, GG95, GZR89, HNS^{+}10, IBBA20, Mil97b, MCLK07, Sah94b, SPM^{+}10, SP12, UHU09, YKB^{+}00, AGY^{+}11, Chu87, Gok90b, GS94c, HBKR96, KK89b, KMB09, KCG08, LR89, SKS04]. Reconfiguration [EAGEG09, HNS^{+}10]. Reconnaissance [Gri90]. Reconstruction [Ano94-89, FBH93, FWWD95, KVP95, LD93a, GS89d, Use93, VM07]. record [Ano94s, Ano97j, Bro91a, Bro91b, Sin94a]. Recording [RSRG95]. records [Ano91g]. Recovering [VM07]. recovery [Bin88]. Recurrence [CP94c, PS94a, AH90, CP92b, CP92a, LR88b]. Recurrences [BCZ95, WN92, WNKS96, McD85]. recurrent [GF95]. Recursion [OB94]. recursions [Mas94b]. Recursive [Ano94-51, Pel94]. Recycled [GK92]. Redefining [Ano93w, Ano96y, RMM20]. Redesigning [Far90]. Redistribution [KHSJ94, TCF94, RCS21]. Reduce [VRSG93, Shu88]. reduce-or [Shu88]. Reduced [Bau88, GT91]. Reducing [Abr88, Ano94-111, Bec89b, CTD^{+}16, DL96, GSB95, Gra92, KABG95, TZ94, Lil88]. Reduction [BL93, BSL94, Din92, ET96, HTI93, JC94d, KRJ93, LB94b, OB95, SSG93, SKN96, Bru88, DD88, Fra90, GS87a, GS88b, Lil91, Pol89, Roj19, Sch87d]. Reductions [MM94c, MM94d, RF94]. redundancy [WWJ09]. redundant [GV92]. REFAL [Tur79]. Reference [WGW04, Woo05, Bru90b, Kue87, PJ90, SG92b]. Reference-Set [WGW04, Woo05]. References [MT91, Gua88d]. referencing [LMY88]. Refine [STN93]. Refinement [JP94, Fuj11]. Reflecting [SBW^{+}19]. Reflections [Luc09]. Reflector [JKNK93]. Reflexive [Che92a]. Reform [WG93b]. Refraction [JBWB97]. regarding [Gup88]. Regime [SBY93]. Region [Cha94a, LQFC18, YMZ90, ZLC21]. Regional [ABCH97, GGN20]. Regions [FSGS93, Lu93, MKRI93, SLML93, LG87]. Register [CMHK92, EDA95, EJL90, KKF96, MSAD92, MD94, Loo84]. register-level [Loo84]. Regular [CCSR92, Mal88b]. Regulations [Ano92t]. Regulatory [BGS^{+}12, MTHP93]. Rehosting [SS96c]. reinforcement [KMRR20]. Reinventing [Sim97]. Rekindled [Can92]. RELAP5 [MM93a]. RELAP5/MOD2.5 [MM93a]. Related [GM93b, CP92a, PB87, SM89]. Relation [HL91, HLxx, FB91a]. relations [AH90]. relationship [ZAS94]. Relationships [Hal96]. Relaxation [FG87, Hun90, HGS91, Hun91, WGS91, Xia88]. Relaxing [ZYL^{+}16]. Release [JR94]. Released [DDJ98a, Bro17a]. Releases [Pau08, Ano94-81]. Relevant [BCW93]. Reliability [KSTF94, LC12, YWXZ12]. Reliable [DDHK94, WWJ09]. Relocation [Uni95]. remain [Ano96z]. Remapping [XL94]. Remark [Lev98]. remarkable [DWV92, Zor92a]. Remarks [Cra96, Leu96, LB82, Deu86]. Remediation [HM97]. Reminiscences [Con11]. Remodeling [HHTD90]. Remote [Joh86c, Luc01, Pro01, Ano96l, Han03, KG95, TJC91b, Tuc91]. Remotely [Y^{+}92]. render [Ano02a, Ano02b]. Rendering [Bos94b, CO94, GMMT91, Kar94, TH94, ST90, Wil90b]. Renewed [Ano97z]. renewing [Gra93c]. Renner [Ren97]. Reno [ACM89a, AIA94, Ano95q, Ano98a]. Renowned [Lin83]. Reordering [GMW94, KESH94, KSH94]. Repeatedly [Moh94]. Replace [Ano96t, Dau96]. Replacement [Mil97b]. replay [NRM^{+}09]. replica [RCS21]. Replicated [CG86]. Report [Ano85b, Ano88p, Ano01a, DGG93, FG93, GJS93, GS92b, Hof93, KHHS95, Mye92b, Sim97, Tay94, Uni96, Wes96, Ano89p, The90a, Ano91n, Ano91o, Ano98g, AG90, Asl91a, Ann92, Con87b, JC87a, CGS91, CGL92, DHT89, Ede92, Has17, Int81, Inf86, JC87b, Joh86c, Kah93b, Min88, Mal86a, Uni91b, NN87, Nat88b, Per83, Poi89, Poi90, San86, Supxxb, Sam91, Sar90, TR86, Uni92c, Wil88b]. Reports [Ano93e, Ano94-80, Ano85b, New91, IEE89a, NN90, New95, Ren97]. Repose [BSJW96]. Repository [Ano95y]. Representation [BE93a, EBS02, EKZ90, FD97, Pug94, RC94, SLML93, WGW04, Woo05, FP91, GP91b, WLLZ20]. Representations [PWVH95, AW91, PMC22]. Representatives [Tec89, Uni92b, Uni86b, Uni89a, Uni86a, Uni86c, Uni92a, Uni98, Ano95l, Bro91b]. Reprocessing [Sol93]. Reproducibility [PMP21]. Reproducible [Ano94-64, Mas94a, Pry94]. Repulsion [EBS88]. Request [Sol94]. Requests [CB00]. require [Sha95b]. Requirements [Ano94t, EDA95, LEMS95, MSAD92, Mar85a, Pet89b, Rob93, Uni93, Ver97, Bro86]. RES [Car92]. Research [Age05, Ano92x, Ano93t, Ano93-43, Ano94w, Ano94-135, Ano95-35, Ano95-41, BPM^{+}89, Ber95a, BIB^{+}18, BBW90, Cor89b, Car94a, Chr93, Duk91, DP91, Gin82, Gue90, HTV88, HWP95, KP95, Kir89, Kow89b, KJ94, KCM02a, Lag89, LZF16, M^{+}94, Mar85a, Mar85b, Mir90, Uni87a, OMM93, PC94b, SW10a, SN89, Sim92a, Sma93, Ste90, SK93b, Tho93c, Resxx, Uni86c, UU94, Wie96, AISS97, ALPP00, Ano85b, Ano90n, Ano94k, Ano96w, JC87a, DHT89, JC87b, JT87, Kra97, LPC^{+}95, Man89a, Man92, Min88, Mar86, Mar88b, Mec95, Mil88a, Mil90, Nor93a, Nat90, OMA^{+}96, SJD96, UL89, Uni91a, Uni93, VVH95, AFF93, Ano94z, Ano95e, Ano95g, Ano95f, Ano96-28, CB99, Ewa89, Ewa96, Hos95, KS93c, KFW94, LR90, McT96, Rot92]. Researchers [Ano97-27, AL92b, Bar00a, Bar00b, CE18, Sha95b, Ano95-31]. Reservoir [SPS90, WG82, Ano94-132, Bri90, Hen91, SPS91]. residence [Ano93b]. Residential [KY90]. residual [Cha92b]. residue [CR94]. resiliency [SEV^{+}09]. resilient [IS20]. Resistance [Gar01]. Resoltion [NAAW97]. Resolution [ABCH97, BCM90, Ger90, GL93b, MS93, SVML95, Str94, UZ95, LHPG21, PMS94]. resolutions [CR94]. Resolve [HC93]. resolving [GB22]. Resonance [JB90, WRW93]. Resource [ADG^{+}05, GGG^{+}98, KCZJ14, PS94b, WN92, ZLRC20, CS82, CS86b, MNY09, SKB^{+}20, Ste90]. Resources [Mor92c, SP12, Cat92a, Jet91, Jet92, Kos95, MBM^{+}20, Natxxf]. Respond [PB90]. Response [BS97, CHL93, HL93a, IK93, Smi01, AB95, Gou90, TfGERJxx]. Responses [Pli97]. responsibilities [SH91]. responsibility [Nat89a]. Restitution [Gie96]. restoration [LHPG21]. Restricted [Mis90]. Restrictions [Ano95k]. Restructuring [Gan86, Gen94, TA94, Eig90b, Gua88d, HP88a, LY88a, LY88b, Pol86, Pol87d, Tri85]. Results [Ano94-85, BBDS94, FBGM93, GJS93, Gro92a, HLPP97, KA93b, Men84, PC94a, PMP21, PL94, SS96a, Sim92b, Str97, YJD93, ASM86, Don93c, Ece96, Lou92, Oed92a, Oed92b, PEH93, SPP^{+}05, WLN^{+}96a, WLN^{+}96b]. RESY [EFPSS93]. rethinking [Win02]. Retina [WR95]. Retire [Can92]. Retrieval [Ber90b, IJM14, Pug94, SG94a, WSP95]. Returns [HED93, Cre91]. reunification [Got91b]. Reveal [Pop91]. Revealing [LMM^{+}21]. reveals [Ano88y]. reversals [Ano95w]. Reverse [BGS^{+}12, LM13]. Reversed [Ano95-35]. reversible [Jab88]. Review [Ahm92, Ano91h, Ano93x, Ano94p, Ano95-34, Ano96c, Ano00a, Bue86a, Dun99, Hal96, Har94b, Haw88, Hil97, Kaz92, Kow86, MM94a, Nor97a, Pap97, Por86, Ros93b, Tru88, Wen94, ACM89b, Sup88b, Hig92, McD88, Mur10, Nat91a, PSOM90, SIG90b, Sch88a, Sim00, Smi88, Resxx, Wal95, TC94]. reviewed [Man89a, Man92]. Reviewer [Pin99]. Reviews [Ach99, ALPP00, Ano93q, Bra94, BBC^{+}05, CCKSS90]. Revision [Chi20]. Revisited [XB96]. Revisiting [IBBA20]. Revive [VN04]. Revolution [Ano91u, Ano91v, CK92b, CCKSS90]. Reweaving [CCKSS90]. RF [SKB^{+}20]. RI [KNHN16]. RI-MP2 [KNHN16]. Richard [Wen94]. RICIS [UL89]. Ride [PYTL97, Ada95b]. Riemann [RT93]. Rifkin [CCKSS90]. right [Che90b, SG92c, SG92d]. right-hand [Che90b, SG92c]. Rigid [Fin82, LF03]. Rim [Suh97]. Rim/ASME [Suh97]. Rings [BK91b, CR94]. ringside [Ano94-121]. Rio [SN96]. RISC [Hil97, Ano94-120, Ano96-35, BB94, DD99, Mar90, SD92, Wei92]. RISC-based [Ano94-120].

RISC-Grossrechner [Ano96-35]. RISCs [SW94]. rise [Afu90, GA12]. rising [Cal11a]. Risk [AH93, Lin83, Sol93]. Rivalry [Lew17]. rivals [Don93b]. Rivers [Hab89]. RMA [BBW19]. RNA [Ess90, KKPR93]. Road [BGMR96, Koo97, Mye96, UHU09, BBC^{+}99, Bem92, Gla93, Has17, Vet12]. road-vehicle [BBC^{+}99]. Roadmap [SBW^{+}19]. ‘Roadrunner’ [GKS09]. roads [MT13]. Robin [UU94]. Robinson [The90a, Con87b]. Robocopter [Bar00c, Bar00d]. Robot [Ano93b, GCS94, GA12]. Robot-in-residence [Ano93b]. Robotic [CMPR93, Coc03a, Coc03b]. Robotics [SB94c]. Robots [Bar00a, Bar00b, Coc01, KA93a, KT93b, MGA94]. Robust [KB94, SHA^{+}92, BS88b]. ROC [COS89]. Rock [Jon19]. Rockefeller [IEE90]. Rockwell [GZA86]. RODOS [EFPSS93]. Role [BKT94, BPL95, ML97, Cal91, Lev89, NCKMM88]. Roles [Ano96-39, Ons88]. Roll [DM93]. rolls [Ano95e, Ano95f]. Roma [Vag88]. Roof [Raw97]. Roofline [KAMB19]. roofline-based [KAMB19]. Room [KLSC97, KY90]. Roon [Sei94]. root [Phi85]. rose [Par94]. Rotating [BSJW96, Koh96, YMZ90]. Rotation [OLWW94]. Rotor [VH93b]. Rotorcraft [Str10]. Rough [BISB96, DSB96]. Round [UU94]. roundoff [Bli89]. Roundtable [Fer83]. Routed [Ano94-88, PDR94, TM94a, TM94b, WK95]. Router [DDHK94, MCW98, RWNJ94]. Routers [LN94]. Routine [VW95, WL83]. Routines [FGG09, PRSS94, Sch96, LSK04]. Routing [Ano94-53, CRV94, CB94, Geo94, KMNT95, KMNT96, LBT94, Mik94, PDR94, RE94, ST94, Dra90a, HS93a, Hol90b, Joe87, RFS87, SQS^{+}19, Smi92]. rover [Ano97m]. Rovibrational [WKHS97]. Row [Bra89c, BS90b, Ano92-31, BS90a]. RS [BIRB93, HMS93, MSAD91, PBDM93]. RS/ [BIRB93, HMS93, PBDM93]. RS/6000 [MSAD91]. RTD [Pel93a]. Rubbia [Ano93y]. Rubin [Wen94, Hil97, MM94a]. Rule [LB93]. Rule-Based [LB93]. Rules [Tys91, Bel89]. Run [Ano94-42, Ano94-112, CWD^{+}08, DCG93, DCGxx, FY96, FP91, Hem84, RP94, RAP95, RW94a, TJC91a, Cho90b, GHdF10, Gro92c, Hab92, Lee90, Mil87, Pol88b, Str94]. Run-Time [CWD^{+}08, RP94, RAP95, RW94a, Ano94-42, Ano94-112, FY96, TJC91a, Hab92, Lee90, Mil87, Pol88b]. Running [Ano88h, AGLL98, BIRB93, Ano97j, Ano97s]. runs [Ano93s]. Runtime [ASS94, AZ94, CCSR92, EAGEG09, KK96b, LMT95, LLGS09, MS94c, RAES96, SLB93, TBC94, TCF94, WF94, EO13, GFBR10, Guz86, SFC^{+}21, TDBL13]. Runtime-Tunable [WF94]. Rural [SR94]. RUS [LLR93a]. Russia [AGP96, Ano93z, Ano94s, GP93c]. Rutgers [Asp93].

S [Ahm92, Ano00a, Por86, Sch88a, War03, Ano94-135, Lil91, RWNJ94, Bem92, Bra93, DH86b, Gro93, IAKH92, Kha93, KISY94, NBKP95b, NBKP95a, SKIY94, WL83, Ano94b, GKS09]. S-15 [Ano94h]. S-3800 [IAKH92, KISY94, SKIY94]. S-810 [DH86b]. S-810/20 [DH86b]. S-Connect [NBKP95a, NBKP95b]. S-MP [Kha93]. S. [Wei90]. S600 [Web93]. S600/ [Web93]. S810 [LMM85b, LMM85a, LMM86]. S810/20 [LMM85b, LMM85a, LMM86]. SA [Ano94-52]. Sac [Bel96]. Sachs [Per87]. Safe [AA93, BLO94, Sti98a, Sti98b]. Safety [CLP93, CPR93, FNT93, Koo97, Law90, MS94a, NS93, PA93b, SDK98, Ano97u]. sailboats [Per87]. Saint [CBCH93]. Salesman [Gur88, Bra88]. Salishan [Feo92]. Salt [ANS92, Ano95-38, Isk96]. same [Rya90, Sne94a, Sne94b]. sample [AS93]. SAMPLES [Ano96x]. Samuel [RD94]. San [ACM95c, ACM03, AIA93, B^{+}95, Bel86, CG96, Clo96, Gra94, GE96, Hel93, IEE95c, JD95, IEE94d, SR93b, Ano97f, Ano97r, Ano97z, Dau97, Deu86, Lay91a, PBK91, San91, Tay94]. sanctions [Kra97]. Sandia [Ano88m, Ano96-40, Ano97s]. Sandpile [LDMC96, LUT96]. Santa [ACM95b, Ano93n, KK89a]. SAR [CDA94, Gol99, OLLG96, SPGD98]. Saratoga [Ano97q]. SAS [Che88, Che92a]. Satan [RR95]. Satellite [BS98a, SF82]. Satellites [Bar01, Ber95a]. Satisfiability [SB96]. Saturated [TOWC15]. Saul [Por86]. save [CKS99]. Saving [App95, Hen91, Par90c]. Savings [SW10a, Ano89i]. Saviors [Ano96-32]. savvy [PW05b]. says [Bab94]. SBH [DDLV93]. SC’93 [Bor93]. Scalability [Ano94-113, HJZ94, HT94, JHZ95, KB93, Kre95, MP94, SSKR97, TOWC15, ZX95, Kwo87, LSK04, WLN^{+}96a, WLN^{+}96b]. Scalable [Ano94-43, Ano94-114, Ano94-115, Ano94-116, Ano01a, AFT97, BIR94, BHEG94, CSSY92, DWM^{+}01, DXJM93, EFR^{+}05, For02, FBJ^{+}94b, GS94a, GLS11, GHWZ94, GL94, HMM94, HNS94, HT93, HMC94, Hol94, IEE93c, IEE94c, JS95, Kan15, KTN^{+}14, KMRR20, KCPT95, KHZ^{+}08, KR94d, Mit98, MS94d, NRM^{+}09, PN13, PZS^{+}20, Pfe93, PW94, SAB^{+}05, SABJ94, Sob93a, SHL^{+}20, UP01, BY21, BWHS18, CKS99, GTV91, GREC91, Hsi91, HLJT93, IS20, KSB^{+}19, PSG03, PMS94, SCK^{+}00, Sob92, SS07, YQTV12, ZEC^{+}17, BCHJ94]. Scalable-Parallel [DXJM93]. Scalar [CDC^{+}87, Jor87, JC94c, NKTT95, VH93a, Wei89, WS90, BJ95, ST90, WS87a, WS87b, WS87c]. ScalaTrace [NRM^{+}09]. Scale [BCH12, ACL93, BBC92, CCR11, CDC^{+}87, DAKM98, GGG^{+}98, GK93, HWS^{+}88, HK93a, Iwa92, KSM^{+}19, MKND97, MH94, OS94, PZS^{+}20, Rui91, SkLC^{+}03, SKK^{+}90, Sob93a, WKL^{+}16, WVBM88b, YZL^{+}20, Zla01, Ano96q, AUW08, B^{+}89, BBC^{+}89, Che83, CH87, Che90c, Che93b, Che89c, Che99, DSZ96, DLJ^{+}08, Duf90, GHdF10, GPS86, GB22, Gra92, GZA86, HFH86, HFH87, IU87, Jor87, KNHN16, Kos95, KSB^{+}19, LPD^{+}11, Lee87b, LXW^{+}16, MP91a, MP91c, MR87, NNS^{+}90, NP90, Sie90, Sob92, WVBM88a, WT11, WT13, YTL87, ZLC21, van95b, DLG93, IMF91, App96]. scale-resolving [GB22]. scaled [MP91c]. Scales [Mil97a, PC97, Ano89n, Ano92p, LAdS^{+}15]. Scaling [CP94b, DT08, HYL^{+}20, LEY86, McB92b, ARF12, CK90, Kor93]. Scan [Bar01, RM96, SJA94]. Scanjet [PRS94]. Scanner [PRS94]. scanning [And90a]. Scattered [Kar94]. Scattering [AFML93, JBWB97, MF92, SE90, PB94a, Ric91b]. ‘Scavenger’ [SW91]. Scenario [Coo95, SB94a]. Scenarios [RG94]. scene [Wad86]. Scenes [EKZ90]. SCF [LA93]. Schauble [Ano96c]. Scheduler [Dow98, WMKS96]. Schedulers [BMSP94, FR96a, SFC^{+}21]. Schedules [EDA95]. Scheduling [Ang91, Ano94e, Ano94-56, Ano94-117, CD92, DA94, DRSS99, GM94a, HED93, HNS^{+}10, Kim96, KCM02a, KCM02b, KCZJ14, LG97, MPT12, OP96, PR94b, PK87, PS94b, RF94, SSSR20, SES94, TGV08, TF92, Wal92, YG92, ZYL^{+}16, Aba09, All93, BP89a, Bec89b, BP91b, BP92, Bro86, CLmWH91, Dan91, EB18, Fan87, LY90c, Mil87, Pol86, Pol88d, Pol88e, Pol90, Tan87, Tan89b, TDBL13, TRLD13, YF98]. Scheme [AM93a, Ano94u, AC84b, FG87, HBDS93, LC97b, MSGW94, MH96, OK93, PK87, RC94, SKIY94, SKIY97, SB94b, TF94, VV94, Wat91, AC84a, BS87a, CV91b, Che94a, CV88a, CH90, yHYZ87, LY90b, LA93]. Schemes [Ano94-43, JML95, LLY92, MNZ^{+}15, MS93, NU91, RBL94, SVD96, YS94, Abr90, Bra89b, CV92b, CDS98, GPS86, SL92]. Schmidt [JP90]. Schneck [McD88]. Scholarship [Ano94-33]. Schönauer [Tru88]. School [Ano88v, ORS94, ORSS94, Pan96, VV95, New95]. Schools [Mur07]. Schumpeterian [Bla97]. Schwarz [KC93b]. schweren [Ano97d]. sci [Mou89, Mou90]. sci-tech [Mou89, Mou90]. SCIDDLE [ABGL96]. SCIDDLE-PVM [ABGL96]. Science [Ano90u, Ano93t, Ano94w, Ano94-71, Ano95l, Ano96-27, Ano96z, Ano96-37, Bar01, Bor94, Bro91b, CCD^{+}13, JC87a, CR89, Cor89a, Cra91, Don94, Duk91, EW90, FHM99, Got91a, GK18, HHK19, JC87b, KS94a, Leu96, Nat92a, Natxxc, Nat95, Nas91, OHIB93, OHHIB94, Per83, Pit90, Pre93a, Red91, San86, San90, SHMR93, SC99, UHU09, Web91, Zen94, AGEL13, Ano93g, Ano95w, BFS11, CCC^{+}89, De 96, DGH20, Eck92a, Ede92, FK98, Fox97, FW90, GCS20, GGN20, WAD^{+}89a, HS^{+}91, Hoc96, HK93b, IBM01a, KS93a, KH93, LP90, Lay91a, ML89, McN87, Mur10, NS86, Uni91b, Nat91b, Nat92b, Ras91, RD94, Ros89, Sci86, San91, SMDL90, Wil88b, WAD^{+}89b, Zyg93, IBM01b, Ano95w, Coc02a, Coc02b, Eck92b, Mah94a, Mut94]. Science/Technology [Ano96-27]. Sciences [AIA94, Ano91k, Ano93t, Bla93, GS87b, Ano98a, GG^{+}97a, GL90, Han03, Nat90, Edw97]. Scientific [Abr94, AF97, Ano88p, Ano94-38, Ano94-37, Ano94-103, BCH12, B^{+}95, BPM^{+}89, BCC^{+}09, BS94a, BEK02, Ber07, BCL91, BBD92, Bla93, BCHJ94, CGFT05, CCKSS90, CG86, Che94b, Cla96, Con90, Con91, Cox88, CM93, DT97, FJSD96, Get15, GHWZ94, Gup94, Hab89, Hab86, HC99, HLB94, Joh97, KB94, KBLD08, KBD10, LLR93a, LAPR94, MT86, MTH88, MCB^{+}01, MBD99, Nat91a, OYWK91, PB90, PN13, Pap97, Pic91b, RL90a, Sch87c, SkLC^{+}03, SR10, SZ11, SL99, S^{+}93, SK93b, SGH97, TY96, Uth94, VWC96, WN10, Wen94, WB85, Woo93, Hwa85, ACM89b, ABB^{+}13, Ano96c, Ano99, AKM^{+}06, Bur91, BLW11, BS94b, Ber82, Che83, CKS99, EWS^{+}13, Gin93, GLH94, Gua88b, GG88, Han94, Haw86, Ipe19, Jor87, M^{+}94, MSK^{+}02, Nat90, NSR90, OGO^{+}20, Pit87, Poz13, Rya13, SKB^{+}20]. scientific [SPP^{+}05, Wil88a, WMK90, WT13, Tru88]. Scientist [Dun92, Gol91a, Gol91b, Hil97, MM94a, Wen94, Wic92, Ano11b]. Scientists [Ano88v, Coc01, HHS01b, HW11, Sha95b, Str94]. scientometric [DGG18]. Scope [Geo94, Pap16]. scoped [BGS82]. Scores [Ohr86]. Scottsdale [KFF93b]. scramjet [CYXL18]. Screen [AH93, Coc02a, Coc02b, Rud90]. Screw [YMZ90]. SCRI [Ano95-41, Car89b]. SCS [Chi90, Ano88k, Hab86, WSL88]. SCS-40 [Hab86, WSL88]. SDBS [DA94]. SDE [TS90]. SDEM [ER94]. SDSC [Ano87a]. Sea [DGO90, Sch19, KfGERJxx, LB94b]. Sea-going [Sch19]. Seamless [Hir92b, Kha93]. Search [AFAGR96, Ano94-85, BK91b, GZW^{+}22, IJY^{+}14, Lu93, RS94c, Tak93, VSB^{+}21, WWY93, CDG^{+}06, Svo93]. Searching [Coc02a, Coc02b, PTS93]. Searchlight [GNJW93]. Searle [CCKSS90]. Seas [Ano91w]. seashell [Pic88]. Season [GP06, Str11]. Seattle [USE00b, CBKA09]. sec [CWLT97]. Second [Ano88b, Ano88w, Ano93-34, GPKK82, GG^{+}97a, GW93a, Kah93a, KK87, KT80, L^{+}93, Lim93, Mar88b, Pel93a, Uni92b, Uni92a, Ano97s, Boy15, Bue86b, IEE95d, Lan92, Ram88, Sch22, Wal85]. Second-Generation [Ano88b]. Secondary [Ano96i, Pan97]. secret [Per89]. Section [AAB06, Ano90d, GS89c, GGN20, IHSK93, Ken92, LQFC18, ZLC21]. Sector [Sun94]. Secure [SDK98]. Securities [CCZ93]. Security [Bar00a, Bar00b, Chr90, Coc01, Pei17, Sub94, Uni98, USE90, Opp95b, Rya13]. Sediments [BS97, BCCG97, ZL97]. see [Ano97u, YB90]. Seek [Coc01]. Seeking [RDHC94]. Seeks [Dau96, Pin01, Ano95v]. SEFOR [RCR93]. Segmentation [HKG90, BÇG14]. segments [Bra89d]. segregation [GE12]. Seidel [Ano94-92]. Seismic [HYL^{+}20, KN88, Arb92, Bro93, CGLY96, CGLxx, Chexx, Gou90, LS92b, LM13, PLS20, Vaf88, HWS^{+}88]. Seitz [Ano11a]. Selected [Ano91r, Pet89b, Pin01, WZ97, Kin96, MW88, Ste85]. Selecting [Ano94-118, SSG93]. Selection [CB00, Lew17, MK07, SKB^{+}20, WH94]. selective [CV88a, WFJ^{+}17]. Self [Ano94-30, BA95, CP94b, CTRR93, FNK93, GT91, GD94a, LUT96, OBR94, PN96, PK87, RPY94, Sto95, Tan89b, Zas93, Fan87, KMN^{+}05, Tan87]. Self-Adaptive [BA95]. Self-Compacting [Ano94-30]. Self-Controlling [Sto95]. Self-Explanatory [FNK93]. self-gravitating [KMN^{+}05]. Self-Organized [LUT96, PN96]. Self-organizing [RPY94]. Self-Scaling [CP94b]. Self-Scheduling [PK87, Tan89b, Fan87, Tan87]. Self-Shielding [CTRR93]. Self-Stabilized [Zas93]. Semantic [DKF94, IJM14, Mas92, Pug94, Sha94a, Coc01]. semantics [Ano94-100]. semantics-based [Ano94-100]. Semi [CT94, GGW93b, GNJW93, JML95, LMP^{+}90, Rui91, SQM94, Car89b]. Semi-Active [SQM94]. semi-complete [Car89b]. Semi-Dataflow [Rui91]. Semi-Infinite [GGW93b, GNJW93]. Semi-Interactive [LMP^{+}90]. Semi-Lagrangian [CT94]. Semi-Linear [JML95]. semiannual [Sam91]. Semiconductor [Ano94-94, Ano03a, BMW91, DWV92, SF91]. Semiconductors [IMF91, Wat92b]. Seminar [Don92a, M^{+}94, Meu89a, Meu91, Meu92c, Meu93, MB93, MB94b]. sends [Ano93b]. Sensed [Y^{+}92]. Sensing [Luc01, Ano96l]. Sensitivity [AH93, BKT94, CSRB90, ODAZ15, Ram88]. Sensor [YK94, Ano02a, Ano02b]. Sensori [KDBG95]. Sensori-Motor [KDBG95]. sensors [Ano96u]. Seoul [IEE97b]. separable [CH87, Che90c]. Separated [Rya90]. Separation [SHZK94, ZS93, Cla18]. separators [Kra88]. September [ANS92, Ano88d, Ano91m, Ano93c, Ano93n, Ano94-108, Ano96q, Ano97c, Ano97-30, App96, BBM96, BP89b, BP93, DDC96, DJM94, DLM99, Guo94, Hel93, HKS93, IEE95d, JPTE94, Kah93a, KWW92, KK92, Lag89, MM91a, MW88, ML95b, NN90, SEA84, SN96, VV95, Ano96k]. Sequence [AKDM93, BS04, DLG93, EBS88, HC93, LD93a, Lop89, MC10, MDH00, PRS94, Ros93a, SSM93, Tak93, Tur90, DT96, Han90]. Sequences [FAKD93, KDBG95, KGS93, Kon93, KT93a, Lu93, Meh94, MKRI93, SD93, AGD93]. Sequencing [DDLV93, JHGLG93, SSKa93, SGIS93]. Sequential [BB91a, JAB92, Sha94a, SJA94, KK89c, KHN89]. Sequential-Type [KHN89]. SERC [THH81, THH82]. Serial [JML96, Bas95b, GS93, SI90, SI91a, SI91b, Suz89]. Series [Ano95-33, DMT^{+}21, Iwa90, Koe96, Koe97, STSK95, Uch96, Uch97, CR94, IAKH92, OW94, Tru88]. Server [Ana94, Ano94-143, KGS93, Ano95v]. servers [Ano96r, Kah91, Sma95, Ano92-47]. serves [BBWR90]. Service [APK^{+}12, Ano22, BCW20, KCZJ14, PD94, Pin99, Raa97, THH81, Kar13, KG03, KW11, THH82, WJC09]. service-oriented [Kar13]. Services [Bür88, Coc02c, Coc02d, Lay91b, Mit98, AAC^{+}05, Ano96n, BY21, Liu12, Lun94, Min92, MNY09]. sessile [Gre88c]. session [Tec89, Uni92b, Uni86b, Uni89a, Uni86a, Uni86c, Uni92a, Uni98]. sessions [Ano94-126]. Set [BJ93, CDW94, FGG09, TM94a, VRSG93, WGW04, Woo05, CGW05, Dra88, Kue87, PJ90, Poo96a, SL90, Way96, YH90]. set-top [Way96]. Sets [Ano90b, Ano94-130, Ano95-33, BBD^{+}08, Gol99, Ano95h, Ano97j, DD88, Wil90b]. settlement [KSB^{+}19]. Setup [Kad94]. Seventeenth [Gro90]. Seventh [B^{+}95, MP92]. Several [Gar99, But92]. Severe [Gon93]. Seville [Mar88b]. Seymour [Nor97a, Ano94-119, Ano96-42, Ano98d, Ano16, Bel98, Ber96, Bro17a, Cra96, Lea09, Mur97, SC99, Smi96a, Smi96d, Sul97]. SGI [Ano93-27, Ano93-28, Ano96-28, Ano96-29, Che99, Fat10, Gar99, LKJ03, LSK04, PIH04, Smi95]. SGI/CRAY [Che99]. SGI/CRAY-T3E [Che99]. shake [Gib01]. Shallow [BK93, DGO90, DA97, McB93, WF94, Ach99, QB92, McB92a, McB92b]. Shallow-Water [BK93]. Shannon [Cra96]. Shape [SBW94, TSCG94]. Shape-Memory [TSCG94]. Shapes [MSGW94]. share [CH92a]. Shared [Ano94f, Ano94-43, Ano94-58, Ano94-90, BIR94, BCHH94, CGSG94, CV95, DS96a, DVWW05, Fri94, GB96, GA84, GL93a, HB96, KABG95, KG96, KV96, KFW94, KCPT95, KB96, Lee94, McK94, MH94, OB95, OIY91, PBM95, SP12, TAAL95, Tur89, AP87b, Ano97j, Car93, CGL92, Che93b, Che89c, CH92b, DI88, GS88a, Gle91, GTV91, Gra92, GL96a, GL96b, GL97, Kon91a, KY91a, KY91b, Lee86, LYL87b, Lee87b, LR88a, LMY88, Lil91, Lim91a, LY91b, MRM87, Mit88, Nat89a, NG92, PS88, RMM87, Saa87, SFL^{+}94, SSS90, Ske89, SY91, TYZ89, TV88, TV89, Yan90a, Yan90b, Yan91]. Shared-Memory [Ano94-58, Ano94-90, DS96a, DVWW05, KABG95, KCPT95, KB96, McK94, MH94, Ano97j, Car93, CGL92, Che93b, Che89c, CH92b, GS88a, GTV91, Gra92, GL96a, GL96b, GL97, KY91a, KY91b, Lim91a, NG92, Ske89, SY91, TYZ89, TV89, Yan90a, Yan90b, Yan91]. Shares [Ano94z]. Sharing [Ano94-138, LHLM95, ZLRC20, Fon85, Mir88, Rya90]. Shear [Ano95-35]. Sheared [Ger90]. sheet [KD93]. sheet-Metal [KD93]. Sheldon [Ano94p]. shelf [Ano95-31]. shell [Xu91]. shells [QB92]. Shielding [CTRR93]. shift [Tze88]. shifted [Ske87]. shifters [FR91]. Ship [Ano88b]. SHiPCC [Sch19]. Ships [Ano91f, DM93]. SHMEM [AGY^{+}11, LSK04, Sch96]. Shopping [BM90]. Short [DA97, ESTA94, Oed92a, Pic88]. Short-Term [DA97]. Shortage [DDJ98a]. Shortchanging [Bar01]. Should [Bar00c, Bar00d, Cra96, Dun92, Gol91a, Gol91b, Wic92, Ruh95, Vro94, Win02]. Show [Pin99, Ano96u, Ano95-34]. Showcase [USE00a, USE01]. Showcases [Lew93]. Showdown [del89]. shows [Ano96u, New14]. shrink [Art93]. Shutdown [JJYL94]. Shuttle [LMP^{+}90]. Si [CWLT97]. SIAM [B^{+}95, FJSP95, S^{+}93, Dra94a]. sic [Yan90b, YH90]. Sick [Bur00, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Bur01a]. Side [KBD97, TD90, Hic18, KPS88, Seh88, Sta88]. sided [LSK04, MM94c, MM94d, PGK^{+}10]. Sidelights [CCKSS90]. sides [Che90b, SG92c, SG92d]. Sidney [Bro91c, Ano16, Haw88]. SIEMENS [EHHS89]. Sierpinski [BY88]. Sierra [Ano20, BBE^{+}20, DRB^{+}20, Lie20, SGS^{+}20, TTM^{+}20]. sieve [BtR95]. SIEVE.1 [SG92a]. sieves [BB92]. SIGGRAPH [ACM89b]. Sight [Ano96x]. sights [Ano95h]. sigmoidal [Cyb89a]. Sign [Ano89n]. Signal [Ano92y, Ano93n, CWLT97, DM88a, DM88b, IEE95b, MD88, MBK^{+}92, SH90, Sch93a, Tak93, Bar88, CP25, Cho90a, Sch12]. Signal-Processing [MD88]. Signals [KB97, Sin94a]. Signature [GR91]. Signature-based [GR91]. Signatures [CP94a]. Signed [AW91]. Sigrid [Roh94]. Silent [LMM^{+}21]. Silicon [Ano94-120, Ano02a, Ano02b, Poo96b, Poo96a, UHU09, Ano03a, Ano92h, Ano94-121, Ano96-28, CB99, SS96b]. Silicon-Based [Ano92h]. Silver [Ano02a, Ano02b]. SIMD [Bak93, BS94e, CBB^{+}05, CP92b, CP94c, CP92c, Hel92, Hor90, KR94a, Kra92, KC92, LL88, LPS90, PVA94, PL91a, PL91b, SI90, SI91a, SI91b, TL96, Tur90]. SIMD-Type [LPS90]. similarities [Han90]. Simple [RBK95, RRW84, Ste94b, WMR96, WD93b, Car92, HY89, Lev98, Sch89b]. simplest [Gib01]. Simplex [TL96]. Simplification [Maa93a, Ano90a]. Simplified [LMM93, Sol93]. simply [Faz87]. SIMTRAN [MAA93b]. simulate [Ano93-37, Ano97e, Gib01, RRSS93]. SIMULATE-3 [RRSS93]. Simulated [Abr92, HV95, Jab88, Rig93, SB96, VT95, Ano95w, Bra88]. Simulating [Ano96t, Dro95, FDD02, GGC^{+}11, ITOK93, PDR91]. Simulation [ALM93, AABB93, AC93, AGH^{+}90, AHFK93, ACSH90, AJ97, AKT90, Ano87a, Ano94b, AAS88, AHH94, Ask93, BCM90, BPJ94, BJLW95, BOS93, BAT99, BD93a, BS97, BCCG97, BD93b, BMW91, BS98b, CMPR93, Cal85a, CT93a, CZRB93, CS86a, CS90, CBT91, DHLS97, DO89, DKS93, Del97, DCW93, Elm95b, EJ97, FA93, FR81, FD97, FJP94, FNK93, GFM96, Ger90, Gil92, GHW93, GMBW93, GL93a, Gre91b, GF90, GW93c, HKR94, HHTD90, HMS93, HL91, HLxx, HGB90, HVSB93, HP93, HFT94, HS93b, Hug93, HKG90, Hun93, IK91, Jab93, Jar12, KY93, KS93b, KSM^{+}08, KBD97, KLSC97, KK95b, KLM94, KY90, Kue93, KT93b, Kuw92, Kuw94, LWY^{+}20, LCVR93, LB82, LCD97, Maa93a, MK93, MKG90, MJH90, MA97, MZ95, MJRS94, MKDY90, MHE97, MF92, MI93, MS94c, MBK^{+}92, MMHM93, NT89]. Simulation [NB94, PB94b, Pay97, PA93b, PZS^{+}20, PK80, Raa97, RG94, Rol97, RLW93, Sch97b, SHZK94, SSJL94, SPS90, Sta94, SZ98, Str10, Stu97, Tak94, TD90, TIOK94, Van93, VTSM12, VH93b, WG82, WG93a, WJ94, Wri19, YS94, YF95, YYK93, ZBLZ95, All93, AZC13, Ano88q, Ano92u, Ano94-132, Ano95w, Asl91b, BAAD^{+}97, Bec90, Bri90, Bru90a, CGM91, CYXL18, Chi86, Chi90, Cla97, Con88, Dao88, DLS93, DSZ96, DT08, DLJ^{+}08, DuB90, FMD07, Fin82, FGM^{+}03, Fox97, GZE^{+}05, Gre88a, Gre88c, Gre89a, Gre89b, GH90, Gre90a, GH91, Gup88, GAW96a, GAW96b, HPH^{+}20, Hae91, HJZ94, HG88, Hen91, HPS88, Hua92, Hun92, Hun90, HGS91, Hun91, KSP13, Kha91, Kon91a, KY91a, KY91b, Kop88, KMN^{+}05, MSTK93, Mun04, OL86, OYK^{+}14, Ons88, OD88, PB94a, PZGL91, PB88, RMM87]. simulation [RM88, Sal89, SNK^{+}93, SF91, Sta88, SD88, Suz89, Wad86, WQS92, Xia88, YKY90, YVC89, Zho88, ZMDS96, ZBN^{+}19]. Simulations [Ano88h, Ano92r, Ano94-116, ABGL96, BGMR96, BWO96, CH10, Dec90, DS96a, EFIM91, Ewi97, Gra93b, GIF^{+}12, HL88b, IMF91, KO90, KK96a, LD93a, Lan93, LDMC96, Man90, OK93, PSB01, SBJ90, Sch93b, SBW^{+}19, SZ96, SBY93, TOWC15, UR95, VVKB96, Wag96, WVBM88a, WVBM88b, YMZ90, AMS^{+}15, Ano97j, BBS94, BS91, Bup87, Che96, CKT21, Din93, DWM^{+}01, Dup86, Dup87, Fra94, GG^{+}97a, GKS09, GHdF10, GB22, Gro92a, JHZ95, LXW^{+}16, LG03, PH95, Pow97, SCK^{+}00, Tho90, WvTB^{+}07, Woo93, Yos09]. Simulator [AA93, Div97, HEJM95, Kad94, MF97, OCVA01, RKDM94, SO95, Che89a, Hog02, HD88, SPS91, SO91, Yan90c, Yan90b, KDP^{+}14]. Simultaneous [GT97, GM93b, HMNN91, Kra88, Ked92, NSH95]. Singapore [PL91c, Cha94a]. Single [Ano94-122, Ano94-123, Ano95w, BISB96, Bro97, CMF94, CMAS11, DG95, FT96a, Gil93, HHOM91, MRS88, OA94, Ros93b, BDM94, Moo06, SS10, TMHH95, Coc03a, Coc03b]. single-board [TMHH95]. Single-Chip [CMAS11, HHOM91, Ros93b].

Single-layer [Ano95w]. Single-Phase [CMF94]. Single-Process [FT96a]. single-processor [SS10]. Single-user [MRS88]. singular [AGD93, BS87a, BS88a]. Singularities [Ano18, Sin18]. Sinking [Ano96-30]. SINRAT [AVS93]. Sint [HS94b]. SIRIAT [CMPR93]. Sisal [OH92]. Sistemi [LP90]. Site [Sim97, Tay94, Wes96]. Sites [Dau97, KKKP93, KKPR93, vL99]. Siting [Ano97k]. situ [BWHS18]. Six [WGOY91, WOG94]. Sixth [S^{+}93]. Size [Ano95-40, CB00, OSKO95, Ano93d, LY90c, YH90]. Sized [Pic91b]. Skeleton [RM92]. Skew [GGZ^{+}20]. Skew-Tolerant [GGZ^{+}20]. Skid [MS94b]. Skinny [BvRS^{+}11]. Skyline [Far90]. Skystation [Ano90o]. Slab [RMH93]. Slant [ESTA94]. Sleeper [Ano93-33]. SLEPc [BBB^{+}20b]. SLICC [Ano94-124]. Slice [KMNT95]. Sliding [NdMM09]. Sliding-Window [NdMM09]. Slip [Gru97]. Slow [KW95]. Slowing [JWG93]. SM2 [JML96]. SMA [TSCG94]. Small [AV02, Art93, BBD^{+}08, CDH84, GW93b, KP94, LD93a, New14, VHJB94, Ano94-27, ARW93b, Lag89, PB94a]. Small-college [AV02]. Smaller [DDJ98a, DDJ98a]. smallest [Ano95w]. Smarr [Pre93a]. Smart [Ano95w, DDJ98b, Ano93b, Win02]. smartcard [Bam97]. smarter [Ano99, CKS99]. smartest [Ano96z]. smartphone [Ano14]. smashers [Ano95l]. Smith [Haw88, Bro91c]. Smithsonian [Ano97j, Str94]. Smog [MR90a]. smooth [FMD07]. Smoothness [RS94b]. SMP [Ano95p]. SN [MF93, War93a, YA93, AM93a, FBA93, MDW93]. SN/ [MF93]. SNA [KSW93, Ano90f]. Snapshot [Ano93-29]. SNI [Web93]. snooper [MP84]. Snooping [TNIA92]. Snow [DDJ98b]. Snowbird [SC93]. SOC [Ano00b]. Social [CCKSS90, EM94a, Pic91b]. Societies [HKR94]. Society [Ano95q, Bel86, FJSP95, JT87, Kho94, MR90a, Pin99, SR93b, Ano85b, CCKSS90, Pin01]. Sociology [CCKSS90]. Soda [Pic91b]. soft [Ano96a, Rol96]. Software [ABB^{+}03, Ano88p, Ano88o, Ano88m, Ano88n, Ano94-88, Ano97a, AB96, ATSA93, BCC^{+}08, BMSD94, CGSG94, CDPW94, Coc01, Coc02a, Coc02b, CDR96, CSG99, DMCK92, DGBE96, Fer86, Fer89, FBJ^{+}94b, Gar01, GW91, Gui96, Iwa90, Jal94, JC94c, KFN02, LS94, Mar85a, MM90, MW81, MP91e, MSA^{+}07, Natxxd, OGY91, Pau08, PL91a, PL91b, PRS94, Pin01, Por89, DDJ98b, Tho90, Tri95a, Tri95b, Tri95c, VSM96, WH93, WKL^{+}16, WGR93, Y^{+}92, YMY92, Ano88s, Ano92-44, Ano94-97, Ano95u, Ano97m, Ano97v, Ano98g, AG90, Asl91a, CS82, CS86b, CCG^{+}17, CV90, CV91b, CV92b, CV88b, Com92a, Com92b, Don85, Gan88, GBS18, GV91, Hak89, Hug94, IKM85, Kon87, Kue87, KS88, LY91a, MT13, Natxxa, Nag90, OGY90, Oya99, PATT12, PGK^{+}10, R^{+}00, RIB^{+}13, Sch94c, Som13, Str94, SMH91, UPK87]. software [Voe89b, Ano85b, Ano01a]. Software-controlled [KFN02]. software-managed [CV88b]. Software/Hardware [YMY92]. Soil [HM97]. solar [NSB96]. solicited [Ano16]. Solid [CJ93, Gre91b, HFH86, HFH87, Soe94, Wuo94, BJ84, WHBH93]. Solid-State [Wuo94]. Solidification [CS90]. solids [Fin82, Ano96u]. Solitons [BS98b]. Solo [Kop91]. Solution [Cal86a, Cal86b, Duf82, Duf91, FBH93, For02, GMW94, GSG^{+}94, GW93a, GRSS93, JC94c, Kam86, KE93, LR88b, MDW93, MR90b, MM94c, MM94d, PO88, Soe94, Sus93, Taf96, TYK93, Vui93, BBE^{+}20, BGT90, COT21, Con87a, DD88, Duf90, Duf00, GKR91, GS87a, GS88b, GS89a, GS90, GS92a, GP93a, GV91, Gri92, GKL^{+}87, Gur88, Kra88, Pet89a, SBC91, SCV01, Sam85, Scr88, SE98, TFB94a, TFB94b, Van95a, vdV91]. Solutions [Cla96, HMBS93, JWG93, KP95, KTK94, MRL^{+}17, YYW^{+}20, Ano96l, DGL89, Rya13, TTM^{+}20]. Solve [Dic94, DMPR93, DGJG93, PC93, Wat91, Che99, HHS01b, Poo96a, RG92]. Solver [Ano94-61, Ano94-94, BE93b, CSSY92, Far90, HR94, HE98, KR94c, LM93, MT97, RT93, RT97, VB90, BTV96, BV96, BS88b, CRA10, GMW91, Gao86, GW95, Gri86, LG87, Mar91]. Solvers [ADLL01, CMF94, CP94c, DLPQ94, DKH86, LPNRJ94, VY88, WS93, AD88, CP25, CP92b, CP92a, FGM90, GS89b, HEB96, KA96, Lee96, ZGL14]. Solving [AGL^{+}99, AS88, Ano94-94, Ano94-110, BCZ95, Bra88, BJ84, BV93, Cha90, CGW05, Che92a, Gal96, Gui08, GZA86, HL96, HOSZ97, HO92b, Iwa92, KS94a, LM90a, LO96, ML93, PS94a, SZ89, SMFG85, SL90, SF93b, TFVK94, VAGRMVA90, And88, Ano87e, Che90b, Che94a, Dav86b, DHD89, DS96b, GBS18, HVY91, KS86a, Ked92, Lou90b, Lou92, Nag88, Rob85, SG92d, SW99, WB88]. Some [BCW93, BY88, CCKSS90, DL90, FD93, GJW91, GS89b, Gin82, HU93, HG88, Her89, LPNRJ94, Lee86, Leu96, Lou92, LB82, Ull83, VM94, Woo92, Woo94, Ano88y, Bra89b, Hal96, HOSZ97, Ji91, Lee96, PP91, Tem83]. Something [Mur10]. sometimes [Win02]. sonar [GMF00]. soon [Mar90]. Sophisticated [Ano97m, WZB86]. Sopron [VV95]. SOR [FHKT97]. Sorensen [SB94d]. SORFIND [HH93]. Sorrowful [Ano94-125]. Sorting [AG94, Bue92, YKB^{+}00, Ano97j, BM87, Gre89a, LYC93, Ull84]. Sought [Coo95, Ano92l]. Soup [Voi94]. Source [Coc03a, Coc03b, DCW93, LEMS95, Ano02a, Ano02b, BBE^{+}20]. Sourcebook [DFF^{+}02]. Sources [Ano90q]. South [L^{+}95, New93]. Southampton [BP89b, Ano88d]. Southern [Ano91w]. Soviet [Ano89o, WG93b]. SP [TGL96, WOK^{+}00, WT11]. SP1 [Ber95b, HJZ94, JHZ95, Nai94, PMS94, SSGH94]. SP1/ [Ber95b]. SP2 [ET96, GYL00, JHZ95].

Space [Ano93t, Ano94-138, Ano18, AZ94, BGM^{+}11, Bro91b, CHL93, Col94, FSGS93, FI93, GGW93a, LMP^{+}90, Poe95, SLB93, Sin18, SF93b, Taf96, TAAL95, WR95, YF98, Ano96j, Ano11b, Ede92, Hun92, MBR05, SKB89, SCK^{+}00, Zyg93, YQTV12]. Space-Angle [CHL93]. Space-Based [AZ94]. Space-Energy [FSGS93]. Space-Time [Ano18, Col94, Sin18, SF93b]. Space/time [YF98]. Space/time-efficient [YF98]. Spaces [Pet97]. Spaceship [Rud90]. Spacetime [FD97]. Spain [ACM95a, DLM99, Mar88b, Sig95, Ano94s, RMO96]. Span [Che92b, Che92c]. Spang [The90a, Con87b]. spanning [BJZfDA96]. SPARC [Ano91f, Kha93]. SPARCstations [Ano90o]. Spark [Bed93, BSB93]. Sparse [ADLL01, Ano94-65, Ano94-92, Ano94-115, AZ95, Ber90b, CLY^{+}19, DD87, DDB^{+}10, DR81, DR82, Duf82, Duf91, ET96, EHS94, FY96, GMW94, GG96, GS89c, HR94, Kra92, KC92, KESH94, KSH94, Mis90, NGLPJ96, PS94b, Rag94, Rot94, UZ95, USZS96, WL83, AD88, And88, AS88, Ber90c, BJV^{+}16, CC88a, Con94, Dav86b, DD88, Dav89, DY90, DS96b, GMW91, GSZ91, GW95, Gri92, HOSZ97, HVY91, Ipe19, Kra90, KESH95, Luc91, Mar91, Pin91, Rob85, SW88, SZ89, SZG95, Wij89a, Wij89b, Yan90a, ZGL14]. sparse-matrix [Kra90]. Sparsity [NN88]. sparta [SO95]. Spatial [AM94, CHMS94, CCSS98, Dip96, GW93a, Gri90, HP93, Mil97a, War93a, HJZ94, JHZ95, Kha91, Smi91]. Spatially [YYW^{+}20]. Spatio [HV95, RBK95, VT95]. Spatio-Temporal [HV95, RBK95, VT95]. Speakout [Ano86c]. speaks [Win02]. SPEC [Ano03a, DDJ98a, EGJ^{+}02, GA95]. Special [AP93, AB03, AAB06, Spe87, Ano88p, Ano94-59, Ano94-126, Ano96e, Ano96f, Ano09, BKK11, Ber07, DF12, EFIM91, yFH89, FR98, FT93a, FHM99, GS89c, GMSS^{+}11, GGN20, Has17, Kah93b, KHHS95, Ken92, KRS13, LQFC18, MB12, MLY10, Mye92b, OS94, PW05a, RT20, SCV01, TH19, Tor87, Abe91, Che90a, Kar13, MH18, RF93, Tru88, ZLC21]. Special-Purpose [Ano94-59, FHM99, Abe91]. Specialised [Sub94]. specialist [Ano87d]. specialists [Ano93-41, Hol93]. Specialized [Ano97-29, Mik89]. Specific [Ano94f, EH97a, KRS13, MGA94]. Specification [BSKJ93, Coc03a, Coc03b, Asa93a]. Spector [War03]. spectra [BB87]. Spectral [Ano03b, Ano05, BK93, DJSP93, KO90, KB93, VR94, WF94, DWM^{+}01, FR95]. spectral/finite [DP90]. Spectrum [Bar00a, Bar00b, CCKSS90, Kad94, Ano89d, IEE89b, Ano17]. Speculation [Hal96]. Speculative [Col94, YSKS95, OWG^{+}13]. Speech [IEE94a, IEE95b, Mes93b, Ste95, ZS94a]. Speed [Abr92, Ano94-104, Ano94-143, Bal93, Bar00c, Bar00d, Che89b, EM94a, GS94d, GW93c, GMMT91, Hal96, KBD97, Lan93, Pre93b, RG94, TFO94, TF97, Woo96a, ZS94a, Ano94-132, Ano97m, Ano03a, Bai88, BBBC96, Ber82, Buc83, Che83, Dao88, Fly66, GB91, KW92, NGPH99, Pan96, Ram86, Ros95, Shi95, Smi89, SGS^{+}20, Sug80a, SO91, TR86, AM91, KA91, MHW94]. Speed-Flow [Hal96]. Speeds [Ano88m, Ano93-34, Ano94-29, Ano95-30, Wal85]. Speedup [Ban79, WN92, WB85, PB87]. Spelling [DS94a]. Spencer [MF93]. spent [Win02]. SPEOS [Del97]. Sperry [CCKSS90]. Spetsializirovannye [BKM88]. SPH [HM93b]. Sphere [BISB96, CT94, LC97b]. SPICE03 [PDR91]. Spice2 [Yan91]. Spiegelhalter [Ano94-95]. Spike [RBK95]. Spiking [ADGA95]. Spin [Pau08, Ano90o, BDM94, Poo96a]. spin-polarized [BDM94]. Spinule [MZ95]. Spinule- [MZ95]. SPLASH [ABD92, Bue92, Gok90b, Hol90a, PTS93]. splatting [Wil92a]. Split [TFVK94]. Split-C [TFVK94]. splitting [Ske87]. SPMD [MVS94]. Spokane [IEE93b]. sponsored [Kho94]. Sponsors [Coc02a, Coc02b]. Spot [Ano92z, YTL87]. Spotlight [Cla97]. Spotlights [Hol95]. spots [Gle91]. SpotSDC [LMM^{+}21]. spotting [Ano89f]. Spray [CZRB93, KR94b, MA97]. Spread [Bar00a, Bar00b]. Spring [Ano95q, Bel86]. Springer [Kar13]. Springs [Ano97q]. SPS [PA93b]. Spy [Bar01]. Square [Rot92, YF95, Phi85]. Squares [OB94, Bue86b, Duf90, GPS86, HOSZ97]. squeezed [Ano96o]. Squeezing [DE84, DKH86, MRS88, Pau05]. SR8000 [INKN01]. SRC [Wun89]. SS [MMR96]. SSA [TP95]. SSA-Based [TP95]. SSD [GKL^{+}87]. SSI [Ano93-30]. SSN [BBBC96]. St [ACM88, AGP96, GP93c, IEE85, L^{+}93, Lim93]. St. [KK85]. STA [Kah93a]. STAB [FZM91]. Stability [ACG^{+}90, CBT91, FCD97, JP90, NU91, Sug96, Deg90]. Stabilized [Zas93]. Stable [AABB93, DY90, FB91a]. stack [RIB^{+}13]. stacked [Ano97-29]. stacks [CH90]. Staff [Dal95]. Stage [FI93, HCV97, DHA^{+}13]. Staging [IBC^{+}11]. Stale [CV88c]. Stand [HS93b]. Standard [AKG87, Ano93k, Bal93, DHHW93, IEE95a, Pau08, Don85, FP00, Lee96, DDJ98a]. Standardization [Bar01]. Standards [Ano90b, Ano96-31, Bar00c, Bar00d]. Standoff [Ano93-35]. Stands [Ewa96]. Stanford [Coc02c, Coc02d]. Star [ACA94, CJ94, LBT94, BJZfDA96, BGIM90, HT72, BIRB93]. STAR-100 [HT72]. STAR-CD [BGIM90, BIRB93]. Stardent [Ano90p]. stars [Ano94-128, Bro93]. Start [Cra96, Pau09, Ano95t]. Start-up [Pau09]. started [Cor87]. State [Alaxx, AA93, Bar00a, Bar00b, BBC^{+}05, Cal91, CL91, DP91, IEE93c, IMA93, JR94, KB19, OT07, Ote02, Pay97, Ulr12, Wuo94, BJ84, DDC96, Gur94, Jet91, Jet92, JOK^{+}18, LS87, PMC22, RFS87, Ano97j, LB94c]. State-Of-The-Art [Pay97, KB19, OT07, DDC96, Jet91, Jet92, LS87, LB94c]. statement [Bro91a, Bro91b, Sit78]. States [Bro91b, Ste95, Tho93c, Fis83, KLQ19]. Static [CTD^{+}16, LH94, Pel94, PR94b, Sob93a, Woo05, YG92, YKK96, Ana91, CGLY96, CGLxx, Chexx, PBK91, Sob92, The90b, The91]. Stationary [SF93b]. Statistical [ML95a, Opp95a, SKB^{+}20]. Statistics [Gri90, Pev93]. Stator [VH93b]. Stator/Rotor [VH93b]. Status [Ano85b, ACC^{+}96, DGG93, DvdS12, FG93, GJS93, LEMS95, NN90, SKAT93, Ano96w, Ano98g, Mal86a, MMW86]. Staudenmaier [CCKSS90]. Steady [Dic94, IMA93, Soe94, Gur94]. Steady-State [IMA93]. Stealth [Coo95]. steel [CGLY96, CGLxx, Chexx]. Steepest [TL96]. Steepest-Edge [TL96]. Steering [Ano94v, MKDY90, VS99, YF95, Bli91, Hab92, WvTB^{+}07]. Stellar [ABM88, LM90b, SMM88]. Stencil [BCR96]. Step [BSL94, Koc93, SC92, BMS92, Kos95, SHL^{+}20, BKM93, Kue93]. Stereo [FI93]. Stereo-View [FI93]. Sterling [Ano94p]. Steve [Ano95p]. Steven [War03]. Stiefel [Bra89a]. Stiff [Kah94]. Stiffness [JC94d]. Stochastic [Gie96, JWG93, LC95, LUT96, Chu91, Smi91, ZCPT00]. Stock [Jon96, Zim96]. stockpile [Ano97u]. Stokes [Ano87a, Ano87c, Ano92e, Ano94-140, Bra92, Bru91, Che99, DLPQ94, Dic94, FY92, Glo89, KR94c, LM90a, MFK94, Riz94, SBHW80, Vui93]. Stokes-Flow [KR94c]. Stone [Ahm92, Sch88a]. Storage [Ano90e, Ano93t, GA84, Hal87, IEE95d, JML95, KISY94, LR90, LLY92, Mon93, Par90a, SVD96, Ano02a, Ano02b, Bro86, GV91, IS20, Nat87d, NNS^{+}90, PLC^{+}19]. Store [UT91, YFY^{+}13]. storge [Bas95b]. Storm [GP06]. Stormy [Tay95a, Str11]. Story [Ano94-118, CCKSS90, MSA^{+}07, Mur97, Nor97a, Poo96b, Van97, Ano92-42, Mal91]. Storytellers [CCKSS90]. Strain [BS98b]. strange [Ree88]. Strassen [GV96b]. strategic [PL91c]. Strategies [AJ97, Ano89e, Ano94y, BCHH94, BSKJ93, BC95, Gol99, GAV95, KF95, LLGS09, NP90, Sta94, Ste85, WMR96, ZS93, Che89c, LY91a, Nat86a, PLC^{+}19]. Strategy [Cha93, Che92b, Coc03a, Coc03b, Die95, Joh97, Lew17, Che92c, Hol84, LY90c, WLH00, WAB^{+}05]. Stratified [GFM96, Ger90, KfGERJxx]. Stream [Bot96]. Streaming [HHOM91, HHOM92, OPR01]. Streaming/FIFO [HHOM91, HHOM92]. Streamlined [OM91]. Streams [HAG^{+}13, MM93b, PVA94, SSSR20]. Streem [ITOK93, TIOK94]. Street [Ano94-128, Gre88b, SW10b]. Strength [JCJY94]. Stress [ER94, JM89a, WG93b, Xu91, JM89c]. Striatal [KW95]. strict [SB94d]. stride [Gep01]. Strides [VLA92]. strikes [Cou13, Sch18]. Striking [Lew96a, Lew96b]. Strip [JC94d, CH89b]. stripped [Way96]. stripped-down [Way96]. Stroke [HB93]. Strong [Sch92b, KW92]. stronger [Rob89]. Stronghold [Ano93-38]. strongly [KDK89, Shi95]. Structural [AK95, AS98, Bel93, Bra93, DH93, FV94, KA91, Law90, MTHP93, NSF90, QB92, SC97, Sug96, Ano92l, BP86, Che88, CH89b, ES88, Gou90, Hea91, HES93, Ng95, NP90, PO88]. Structure [ATL90, Ano94-98, BW94, HTI93, KA93b, Kuw94, Lie93, OS94, Sch89b, TAKB06, Bre87, Gua88d, KfGERJxx, NPS93, PW23, RGL^{+}15, Yos09, ZAS94]. structure-function [ZAS94]. structure-preserving [PW23]. Structured [ASS94, Ano85b, CRY94, Cyr86, Pli97, Tsu91, ALPP00, CB89, SMM17]. Structures [FCGG90, Ger90, Hun93, JM90, Raa97, SBN82, Sob93a, ZM94, Ano95w, Ano96q, App96, Gou90, JM89b, KK89a, Lin89, MB89, PB94a, Pot87, SSLR90, Sob92, ZS94b, ZS94c]. STS [Rig93]. STS-Content [Rig93]. Student [Coc02c, Coc02d, Kah93a]. Students [Ano88n]. Studies [BK97, Cal85b, CFV^{+}90, CLP93, Cra96, HB93, KA93b, LC94, Sil91, SABK94, WR97, Ano88q, Ano89f, FMD07, GB90, Gre90b, Ha88, Ha90a, Shi95, Sie90, WHBH93, RD94]. studio [Ano96u]. Study [AJ97, Ano94-104, CAB93, CS86a, DBK09, DG95, DS96a, DDF93, GY93b, HS94b, HCL94, HL91, HLxx, JML95, Kel91, KNS97, LC90, LH94, LYKM97, MH95, NAAW97, RCK97, SSG93, Sal95, TIOK94, TOWC15, WS90, YFOT93, ZM86, Bis94a, Bis94c, CC88a, CY91, Das94, DDT95, DI88, Dub87, EB91, EMS11, Feo92, Gal88a, Gan86, GKSR14, HPH^{+}20, HS93a, HJZ94, IEE92, JY92, KDK89, LY91b, McG87, NDLV88, NW03, PBK91, PGK^{+}10, RR89, RGL^{+}15, SLY89, SLY90, VSH91, WS87a, WS87b, WS87c, WvTB^{+}07, Wij89a, ZH88]. Studying [AM93b, Ano95w, Che89a, YB90]. stuff [Poo96a]. Style [Bro17a, Bro17b, SKP91]. Sub [GP93b]. Sub-Models [GP93b]. Subcomittee [Uni86a]. Subcommittee [Ano88p, Bro91b, CCKSS90, Uni92b, Uni92a, Ano88i]. Subcontracting [GT97]. subcycling [Bru88]. subdivision [CBA90]. Subject [DDF93, LC12]. submarines [Ano89f]. submission [DT96]. submissions [vL99]. Subprograms [CDW94, Dub87]. Subroutines [BCHJ94]. subscripts [SLY89]. Subspace [Bis94b, BHLST94, Saa93a, AT89, Bis94c, HLTZ93, Saa89]. subspaces [Che90b]. Substitution [TYKE93]. Substrate [DDHK94, KMKD97, SKK^{+}90, Lee84]. Subsurface [BCCG97, YCC97]. subsystem [Con88, OBB^{+}05, OWG^{+}13, WTC^{+}02]. Subsystems [Ano91b, Ano94-104, FCBH95b, FCBH95a]. succeed [KWW92]. Success [Blu92, NPS^{+}20]. sufficiently [ALPP00]. Suggests [Ano95-45]. Suitable [MM93a]. suite [Ano95-32, Ano03a, SCK^{+}00, SZ11]. Suited [ACL93]. sum [Fin82]. Summaries [MP92, Ano95-38]. Summary [Kau93b, Ano89p, Man89a, Man92, Ros95]. Summer [Ano94-33, Kah93a, Wun89]. Summit [Ano20, BBE^{+}20, CFP20, Lie20, LSS^{+}20, SGS^{+}20, TTM^{+}20]. sun [Ano94k, Bro93, Ano90o, Ano91f]. SUNDAS [Sho91]. Sunder [Ano94-127]. SUNMOS [Ano94h]. Sunnyvale [Ara96]. Sunset [Max81]. Sunway [AYL^{+}18, CLF^{+}19, CLY^{+}19, HYL^{+}20, LFJ^{+}20, LWY^{+}20, ZFF^{+}18]. Supeordenador [Pac85]. Super [Ano89j, Ano90q, Ano92z, Ano93m, Ano94-128, Ano96-32, Bar00a, Bar00b, Cas01, Hos95, Koc93, Pic91b, Vag88, VKK80, Ano89l, Bro17b, Don93c, Lee90, Lev89, LQFC18, VVH95, Wat72, LMT95, LQFC18]. super-AI [LQFC18]. Super-Computational [Hos95]. super-computers [Lee90]. Super-computing [Vag88, Bro17b]. super-data [LQFC18]. super-priced [Ano89l]. super-workstations [Lev89]. Supercheap [Cas03]. superchip [Ano96p]. superchips [Doz92]. Supercompiler [Tur86, Tur79]. supercomputadora [Pac86]. Supercomputational [BB90, EW90]. Supercomputations [Tou87]. ’supercomputer [Sug80b, Fin94, AM91, AK95, AU91, Alaxx, ATL90, AGKT02, AABK95, Ano87a, Ano87c, Ano88q, Ano88v, Ano88r, Ano88p, Ano88m, Ano88s, Ano88w, Ano88x, Ano89j, Ano89m, Ano89o, Ano89p, Ano90j, Ano90k, Ano90m, Ano90r, Ano90p, Ano90s, Ano91c, Ano91i, Ano91j, Ano91s, Ano91t, Ano91r, Ano91u, Ano91v, Ano91w, Ano92h, Ano92k, Ano92q, Ano92t, Ano92v, Ano92-27, Ano92-28, Ano92-29, Ano92-30, Ano92-31, Ano93z, Ano93w, Ano93-34, Ano93-33, Ano93-35, Ano93-36, Ano93-32, Ano93p, Ano94l, Ano94i, Ano94-61, Ano94-72, Ano94-97, Ano94-129, Ano94-130, Ano95j, Ano95k, Ano95x, Ano95y, Ano95-33, Ano95w, Ano95-42, Ano95-49, Ano96t, Ano96y, Ano96-33, Ano96-34, Ano96-36, Ano97a, Ano97f, Ano97m, Ano97t, Ano97r, Ano97u, Ano97v, Ano01b, Ano03b, ABM88, AC84b, AAS88, ABMW93]. Supercomputer [ABHS89a, BM93a, Bar00a, Bar00b, Bar00c, Bar00d, Bar01, BJLW95, BOS93, BR95, BS94a, Bha94, BBC92, BGQ19, B^{+}89, BBC^{+}89, Bor92, Bor89, Bra91a, BB93, Bra91b, BCHJ94, Bro91c, BBW90, Bro91d, Bup87, BK91a, BK91b, BEW82, Cor89b, Com89, CK92a, CSB89, Car94a, Cas03, CCKSS90, CPS96a, CGW05, Cha94b, Che90d, Che90e, CWLT97, CTD^{+}16, CLY^{+}19, Chr90, Chu91, CS93b, Chu89, Cia88d, Cia88e, Cia88f, CB00, CB02, CDC^{+}87, Coc01, Coc02c, Coc02d, CR89, CW89, Con11, Coo95, CM95, Cox88, CS94a, CBT91, CKPK90a, Cyb90, CKPK90b, Cyb91b, DO89, Dau96, DLS93, Dav00, DGL89, Deh90, DGT84, Deu86, Dip96, Don94, Dub87, DDB^{+}10, DP91, Dup86, Dup87, DM88a, DM88b, Ede92, Egg94, EFH^{+}00, Elm93, Elm95b, Emm85, MSCxx, Fer83, Fei05, FG92, FB91b, Fox89, Fox90b, FBCB18, FY92]. Supercomputer [FHKT97, Gal89a, GZW^{+}22, GFM96, GGG^{+}98, GBG89, GGC^{+}11, Gre88c, Gre89b, GH90, Gre90b, Gre90a, Gre91b, GH91, GML90, GL89, Gro93, Gro90, GG97b, Gun88, GMSB93, GBK^{+}96, HS94a, HLPP97, HHS01a, HHS01b, Har90, Har91, Har94a, Har95, HTV88, HSW^{+}90, Haw88, HMS^{+}86a, HMS^{+}86b, HCL94, Hei89, HL91, HLxx, HFCM98, HGS88, Hes90, HWG98, HT93, Hir94, HS94d, Ho88, Ho91, Hol90b, HMKI97, Hor97b, Hor97a, HYL^{+}20, HKG90, Hun92, Hun93, Int81, IMF91, Ike95, IS95, Ill96, INKN01, IMP93, IK82, Iwa90, Joh86a, JA92a, JA92b, JT87, Joh94, JS95, JM89a, JM90, JBI91, KK89a, KC89, KP95, KFF93a, KFJB94, Kel91, KFB91, KH87, KHN89, Kir89, KISY94, KGB^{+}96, KLM94, Koe96, Koe97, KA91, Kop91, KRJ93, KY90, Kra20, KMN^{+}05, Kun95, LMH90, Lag89, Lam14, Lan93]. Supercomputer [LLR93a, LAPR94, Lay91a, Lay91b, Lee89, Leg90, Lew96a, Lew96b, Lew17, LO96, Lin89, Lin82, Llo94, LMxx, LLL^{+}17, LCD97, MKG90, MJH90, Mar86, Mar88b, Mar85a, Mar85b, MNR86, MMW86, Mas93, MSW96, MH98, McD88, McN87, MM90, MMR96, Meu90, Meu91, Mik94, Mil97b, Mil88b, MRS88, Mir92, Mir90, Mis90, Mit98, Mit96, MF92, MSA^{+}07, MR90b, Mor92b, MD88, Mor92c, MAT85, MBSK92, MBK^{+}92, Mur91a, MK07, Mur97, Mye86, Nat85, Nat84, NKTT95, Nas91, Ng95, N^{+}95, NBGS96, NSB96, Nor96, NBKP95a, Num85, OLLG96, OD88, Pro94, PB84, PGL87, PB94a, PBK91, PCY^{+}19, PZS^{+}20, PZA86, PZA87, PL91a, PL91b, PZ89, Pet89b, Pic91b, PSB01, Pop91, PK94, Rag06, Ram86, Ram88, RWNJ94, Ren97, Rhe90, RL90a, Riv90, RL90b, Rol96, Ros93b, Rud90, RLW93, Sti84, San91]. Supercomputer [Saa93a, SKIY94, SKIY97, Sch89a, SKB89, Sch87b, SH93, SH94a, SB01, SSxx, STSK95, SKK^{+}90, SF91, SDFP93, Sil91, Sim97, SS96c, Sob93a, Ste96, Ste94c, SPGD98, SYMT92, Str94, Sum82, TAKB06, Tay94, Tho93d, TR86, TOY96, TW92, UU93, Uch96, Uch97, US01, Vaf88, VVKB96, VAGRMVA90, VTSM12, WL94, WLKI95, WCZ^{+}18, War89, Wat95, Waz89, Web91, Wei89, WM91, WS99, WGR93, WOG94, Wor84, Joh86b, XOZ^{+}20, YSK^{+}96, Y^{+}92, YAG93, YAGxx, YR93, YMY92, YWD94, YWDxx, Yos09, YZL^{+}20, del89, ASK85, Ada95b, AU90, Afu90, ABB^{+}13, Ali86, AAC^{+}05, And90c, Ano85b, Ano86a, Ano86b, Ano87d, Ano88f, Ano89e, Ano89h, Ano89l, Ano89a, Ano90a, Ano90e, Ano90l, Ano90n, Ano90o, Ano91f, Ano91g, Ano91h, Ano91l, Ano92b, Ano92d, Ano92l, Ano92o, Ano92p]. supercomputer [Ano92w, Ano92-42, Ano92-44, Ano92-47, Ano92-46, Ano93i, Ano93v, Ano93-27, Ano93-28, Ano94k, Ano94n, Ano94-63, Ano94-86, Ano94-83, Ano94-82, Ano94-81, Ano94-101, Ano94-122, Ano94-123, Ano94-121, Ano95e, Ano95i, Ano95r, Ano95p, Ano95t, Ano95f, Ano95-30, Ano95-31, Ano95-37, Ano95a, Ano96m, Ano96o, Ano96a, Ano96r, Ano96-29, Ano96z, Ano96-37, Ano97d, Ano97h, Ano97s, Ano97-29, Ano98g, Ano00b, Ano14, ADG^{+}05, AM96, AI92, ACC^{+}96, Art93, Asa93a, Asa93b, AJFH86, Ata91, ABHS89b, Aus93, Bur91, Bab94, Bab90, BB99, Bam97, Bar88, Bas95b, Bau96, BDW85, Bel89, Bel92a, BS94b, BGH^{+}05, BGT90, Bla97, BBK^{+}08, BM85, Bow88, Bra88, Bra89d, Bro86, BJ84, BJV^{+}16, Bur93, Bur94b, CS82, CS86b, Cra92, Cap96, CC88a, Car89b, Cha92a, Cha90, CPS96b, CYXL18, Che96, CD09, CLF^{+}19]. supercomputer [CGLY96, CGLxx, Chexx, Chi86, Cho90a, Cia88a, Cia88b, Cia88c, CNC^{+}08, Con00, CBC^{+}05, CHT^{+}13, Cre91, CFH^{+}01, CP93b, Dak90, Dal96, Dan91, Dao88, Dav86a, Deg90, DK01, DGT82, Din91, Din92, DLJ^{+}08, DAC^{+}18, Don93b, DHT89, DWV92, DuB90, DR91, Dum97, EFR^{+}05, Ele93, Els89, Emm84, Erc88, Eva97, Fed87a, Fed87b, FMD07, FDM07, Faz87, yFH89, FGC06, Fie86, dCCF01, FFM95, For93, Gan88, GS06, Gep99, Gep01, GBS18, GKS09, Gha84, Gib01, Gil88, Gil94b, Gil94a, GHdF10, Gle88, Gre88a, Gre88b, Gre89a, Gri86, GL96a, GL96b, GL97, Gro92a, Gui05, Gui06, GHS86, Gut95, GAW96a, GAW96b, Ha88, Ha90a, Han90, HJZ94, Han20, Har89, Hei90, Her94, Hib01, Hol84, Hor82a, Hor82b, Hsu15, Hua92, IBM01a, IEE92, Inf86, IAKH92, JKL19, JR91, Ji91, jJ88, Joe87]. supercomputer [Joh88, JM89b, JM89c, Jon03, Kah92, Kah91, KS87a, KN86, KS95, Kha91, KDK89, Kop88, Kor93, Kow85, Kra93, Kra88, KGLA85, KSB^{+}19, LPD^{+}11, Lav89, Law89, Lee84, LS93a, LEY86, LXW^{+}16, LFJ^{+}20, Lie90, LR89, LHPG21, Loo84, LYC93, LM90b, Man92, IJS94, M.I87, Mac97a, Mac96b, Mac97c, MP84, MCH91, Mar90, MKHY95, MHKY97, McB92a, McB92b, McC94, MB97, MM91b, Moo06, MV16, NW03, Nat88a, Nat86h, NN88, NPS93, New14, NGPH99, Nix92, NBKP95b, Ons88, PIH04, Pan97, Pan96, Par90a, Par90b, PTT89, Per04, Per06, PS98, Pic88, Pic89, Pic91a, PBK96, Poo96a, Pop92, PF90, Por89, Pou94b, Pou88, Pri12, PK89, QB92, RYYT89, Ric90b, Ric91a, Ric91b, Rob85, Rol97, RCZ93, Ros95, RCS21, Sci86, Tec89, SN95a, SN95b, Sam85, SNEP14, Sar91, Sar90]. supercomputer [SNK^{+}93, SF82, Sch22, SH94b, Sch88b, SS90a, SS90b, Sha87, Sha96, SMM17, Shi95, Sho91, SL90, SSLR90, Sin08b, SB18, SGB91, SS10, Sob92, Sol84, SCH94d, SMM88, Ste85, SSSSE96, Svo93, SO91, TK89, TCM95, TM88, The90b, The91, Tho93b, Tho90, TTM^{+}20, TMHH95, TFB94a, TFB94b, Tre97, Tri95a, Tri95b, Tze88, Uni87b, Uni92b, Uni86b, Uni89a, Uni86a, Uni86c, Uni92a, Uni92c, Uni91a, Uni95, Uni98, Ull83, Ull84, Van97, Vol89, Wal81, Wal85, WZ87, Wal09, WH94, Was96a, Wat87, WQS92, Wes89, Whe83, Wic96, Wit89, Woo92, Woo94, Woo93, WCHK91, Xu91, YYYS93, Yau88, YW94, YHA93, Yi11, Zag82, ZBN^{+}19, Mob12, IBM01b, Abe90, Abe91, Ano93-31, Ano95m, Ano95-41, Ano96-35, Ano97p, AYL^{+}18, BBH^{+}00, BBBC96, DHM^{+}88, Duk91]. Supercomputer [Gen92, Met86a, Meu89a, Meu92a, Meu92b, Meu92c, Meu93, Meu95, Uni91b, Nat86g, RLC91, Ste90, Swe94, TCJS93, ZFF^{+}18, Nor97a]. Supercomputer-Aided [RLC91]. Supercomputer-based [Che90d, Che90e, Ano90l]. supercomputer-class [Ano96r]. Supercomputer-enhanced [EFH^{+}00]. supercomputer-level [Ano91h]. supercomputer-like [Ano90o]. Supercomputer/Transistor [Ano92h]. Supercomputers [ACM89b, AP93, AGL^{+}99, And90b, Ano88k, Ano88j, Ano88p, Ano88n, Ano92e, Ano92j, Ano92x, Ano92-38, Ano92-39, Ano92-32, Ano92-33, Ano92-34, Ano92-35, Ano92-36, Ano92-37, Ano93-37, Ano93-38, Ano94r, Ano94-131, Ano94-132, Ano94-128, Ano94-133, Ano94-144, Ano96-39, Ano96-37, Ano18, Ano20, ADG^{+}08, AM94, AL92a, ABM^{+}04, BMR85, BPJ94, BH93, BAT99, BD93a, Bel92b, BBD^{+}08, Ben90b, Ben90a, BHMH98, BBWR90, Bro17a, BS98b, CCR11, CRV94, CS90, Coc03a, Coc03b, Com92b, CP94c, CS94b, CK92b, Dal84, DS89, DLPQ94, Den80, DLLG98, Doz92, Due89, Duf84, Duf91, EM94a, Far90, FCBH95b, FCBH95a, FR96a, Fer84, Fer86, FNP^{+}84, Gar99, GS89c, GW91, Gha96, Glo84, GHW93, Gre94, Gri88, GIF^{+}12, Gro92b, Hay84, Hem84, Her90b, Her89, Hwa84, Hwa85, Jen88, Joh97, KC93c, KT80, Kra01b]. Supercomputers [Kra01a, Kuw92, Kuw94, Law00, Lev82, Lew94b, Lin83, LMM85b, LMM85a, LMM86, Mal88a, MT86, MTH88, McB93, MR90a, Men84, MNZ^{+}15, MK97, Mul96, Nag96a, Nar95, NB94, Nor84, PW86b, PW86a, Pau05, Pau08, Pil93, Pit90, PK87, PHVJ95, Pou86, RL96, Ric90a, RG94, SSP93, Sin18, Sol94, SBW94, DDJ98b, SZ98, SHG95, Ste94b, Ste94e, Str11, TP93, TF95, TOWC15, Tsy94, WBP87, WNKS96, Wat92a, WS84a, WS90, WMMC10, YMT93, YYW^{+}20, ZM86, ZBLZ95, Zla01, Zor92b, vdSD96a, vdSD96b, AMS^{+}15, And11, Ano88e, Ano88d, Ano88c, Ano88o, Ano93h, Ano93s, Ano93d, Ano94-27, Ano94-36, Ano94-29, Ano94-120, Ano94-125, Ano95c, Ano95g, Ano95h, Ano95l, Ano95-28, Ano95-32, Ano95v, Ano95-39, Ano96n, Ano96u, Ano96-28, Ano96-41, Ano97e, Ano97j, AL92b, BS00, BWV^{+}17]. supercomputers [Bea90, BAD01, BM90, BBE^{+}20, Ber82, Ber89b, Boy15, BP89b, BP93, BMW91, Buc83, BF92, Cal12, Car91, Cas01, CD08, CKT21, Cla97, CCC^{+}89, Com92a, CP92b, CBM^{+}05, DFSZ88, DHD89, DZM^{+}13, Din93, Duf85, Duf90, DB95, EKTB99, EHHS89, Fer89, Fuj99, FT93b, Gin93, GB22, Gra91, Gro92c, Gup88, GZR89, HOSZ97, HL88a, Her90a, Hil97, Hoc91, Hos88, HS93c, HD89, HN90, IKM85, JOK^{+}18, Jor87, KNHN16, KSP13, KA92, Kos89, Kre95, KS86b, Lie20, Lop89, LQFC18, LF03, LM90b, LLDF95, Man89a, Mac96c, Mar88c, MSPPD20, Meu89b, MRSB94, Moo08, MR86, MR87, MS84, NCKMM88, ODAZ15, Oya99, PW86c, Pal15, Par90b, Per86, Per87, PZGL91, Pot87, RWL^{+}98, RLKW93, RD07, Roj19, RGL^{+}15, Rya90, Saa89, SCG^{+}08, Sal89, SH91, SL92, Sha95b, SWL^{+}91, SWL^{+}92, Sma95, Smi91]. supercomputers [SGS^{+}20, TKI85, TDBL13, TRLD13, TFVK94, Uni92c, VM07, VSM^{+}07a, VSM^{+}07b, Wal85, WZB86, WLCG02, Wat93, WS84b, WS84c, WS87a, WS87b, WS87c, WvTB^{+}07, WWTE92, WT11, WT13, ZEC^{+}17, ZLC21, ZGL14, ZS94b, ZS94c, ZMDS96, vdV91, Ano96-38, Ano90b, Ano92-40, Day12, MM94a, Wen94, Ach99, Bue86a, Kow86]. Supercomputing [ACM88, ACM89a, ACM90, ACM91, ACM92b, ACM92a, ACM93, ACM94, ACM95a, ACM95c, ACM96, ACM97, ACMxx, ACM03, AU87, Ade92, ACKW01, AS93, And89, AW93, Sup87a, Ano88i, Sup88a, Sup88b, Ano88t, Ano88u, Ano88w, Ano89q, Ano89r, Ano90f, New91, Ano91q, Ano91p, Ano91v, Ano92-41, Ano93b, Ano93o, Ano93t, Ano93-39, Ano93-40, Ano93-41, Ano93-46, Ano94-79, Ano94-126, Ano94-135, Ano94-136, Ano95-35, Ano95-40, Ano95-46, Ano95-43, Ano95-44, Ano95-45, Ano96e, Ano96f, Ano96i, Ano96j, Ano96s, Ano96x, Ano96-40, Ano97b, Ano97o, Ano97k, Ano97l, Ano97n, Ano97q, Ano97w, Ano97x, Ano97y, Ano97z, Ano98f, Ano99, Ano00c, Ano00d, Ano01a, Ano01c, Ano22, ASNT91, ALMS92, Att96, Bac88, Bai92, BDM94, BBB^{+}91, BM87, BLP^{+}21, BMR88, Bel96, Bel99, Ben90a, Ber95a, BvRS^{+}11, BG91, Bla93, BB94]. Supercomputing [Bro91a, Bro91b, BBC^{+}05, BB98, BF91, Buz84, CL91, CU90, Cat92a, CBA90, CG86, Che94b, Chi00, C^{+}97, Cla98, CKS99, CCC^{+}89, Coc02a, Coc02b, Coh91, Con91, Cor89a, Cra91, CE18, Cul95b, DM96a, DM96b, Dau97, DKLS86, Dav87, DD02, DGG18, DCG93, DCGxx, DPS97, Don92b, Due89, Eid91, Eck92a, Eck92b, ORS94, EVM^{+}98, EGH^{+}06, EKZ90, FCGG90, cF03, cFC07, FR98, FV94, FT96b, Fra94, FWS96, Fri91, Fri94, FI91, GGZ^{+}20, GG^{+}97a, Gei16, GS87b, GK92, Gel11, GLS11, Get15, GY93a, GCB92, Glo89, GV96a, GK18, Got91b, Gou90, Gri90, Gri93, GG95, Gue90, GJP96a, Gui96, Gup94, Gur94, HB89, Hab89, HVZ94, Har94b, HS96, HSxx, Hey90, Hir92b, Hol93, Hol94, HK97, Ins87a, IEE91, IEE85, Ins87b, IEE89a, IEE90, Ins90, IEE93d]. Supercomputing [IEE94e, IEE95a, IEE96d, Iwa92, Jet90, Jet91, Jet92, Jon19, Kac02, Kah92, Kah93b, Kah93c, KK85, KK87, KK88, KK89b, KK89a, KK90, KEMB99, KS93a, Kav92, Ken92, KK95a, Kok94, Kon96, Kon91b, Kow89a, Kow89b, KL99, KNS97, KR94c, KJ94, KDLS86, KS90, Kun84, KNWB93, KSW93, KK92, LP90, LW92, LC90, Lei85, Lei91, Lew94a, Lew94c, LS87, LC91, Lim91b, L^{+}93, Lim93, LL94, Lun90, Man89a, Man92, Min86, Mac91b, M^{+}94, Man88, Man89b, Mar88b, MW88, MB89, Men87, Mes93b, Mes00, Mil88a, Mil17, MP91e, MCLK07, Moo11, MB93, MB94b, Uni87a, Nat89a, Nat87a, Nat91a, Nat86f, Nor89, Nor93b, Nor93a, NU91, Nee94, NSW08, NSF90, OHIB93, ORSS94, OHHIB94, Ope96, OGY90, OGY91, Pit88, PB98, PS96, Pau09, Pel93b, PC94b, Pfe93, PL91c, Por86]. Supercomputing [Pre93a, RS84, Ros89, Rya92, Soc94, SIG90b, Sch95a, Sch93a, SS96b, SHMR93, SHMR96, SN89, Sig90a, Sig95, SZ96, Sma93, Smi88, SFF94, Sne94a, Sne94b, Sta94, Ste94e, Sti98a, Sti98b, SABK94, SS95, SGH97, Tho89, Taf96, TC21, Tor87, UL89, Uni92e, Uni96, Uni89b, UHU09, VN04, VPDA93, VSB^{+}21, VDK92, VTTS98, WZ97, Wal90, Wes96, Woo96b, Supxxa, YWXZ12, YF95, Zen94, ZS93, AV02, AGY^{+}11, AS99, AB94, AM15, And20, Spe87, Ano90q, Ano90t, Ano94p, Ano94s, Ano94a, Ano94-75, Ano94-119, Ano95v, Ano96h, Ano02a, Ano02b, AGL11, Ban88, Bec89a, Ben90b, BBD92, BFS11, Bor94, Bri90, Cal91, Con87b, Car89a, CBCJ92, Che94a, Chi16, CR89, Con90, Cou90, Das94, Doz92, Fed96, FTT97, Gar92, GW04, GY92, GGN20, WAD^{+}89a, GE12, GJP96b]. supercomputing [HRC09, Han03, Hil91, Hil92, Hir92a, Hir92c, HCPS95, Hod87, HK93b, Hor90, Hor93, Hor98, Hsu16, IAIK92, Int91, Int92, Joh86c, Joh90, KYS88, KH93, Kel85, Kha95, Kim96, Kin96, Kro92, Kuc87, KS87b, KS88, Kum91, LPC^{+}95, LCV90b, Lei89, Lev89, LGG^{+}87, Liu95, LA93, Mic90, Min92, Mac97b, Mac92, ML89, McC88, MSK^{+}02, MO88, Men90, Mes93a, Met86b, Met86a, Mil90, Mil91, MK92a, MK92b, MP92, Mun04, MA85, Nat86a, Nat86d, Nat90, Nat88b, NNS^{+}90, NN90, New95, Oya02, Pan93, PSM93, Asi98, Rat87, Red91, RGH17, Rob87, Sup87b, San90, Sch94c, Sch12, Sch18, SA10a, SA10b, Ser98, Sha89, SHMR94, Ste90, SC91a, SJR05, Tru88, Uni93, VFK^{+}04, Ver95, VDK91, Wal95, WTC^{+}02, Wil88b, WAD^{+}89b, Win02, Wor81, YK87, Zec93, Zhe97, vL99, Ano97-27]. SuperComputing [JTX^{+}22, Ano94-134, Ano95-34, Ano96b, Ano97i, Ano97-28, Ano98b, Ano98c, BCCP05, BH92b, Dra94a, Dra94b, Dra96b, Dra96a, HWP95, HPP88, JLC98, Lew93, M^{+}94, Men87, MN91, MB94b, Mye92a, Mye92b, Qui95, RF93, SSH96, Wei90, Dun99, Ano00a]. Supercomputing-based [PB98]. Supercomputing-Enabled [GK18]. supercomuputer [GL91]. Superconcurrent [NRS95]. Superconductor [Cou11, AC84a]. Superconductors [FJP94, MK93]. supercooled [ARF12]. SuperCPU [Ano91f]. superfast [MS84]. Superfluids [MK93]. Superhighway [Mye96, IEE95c, IEE97b]. superhighways [MP92]. Superhuman [Ano92-42]. Superimposed [SHA^{+}92]. Supermen [Mur97, Nor97a]. Superminicomputers [Wal85]. Superminis [Gre94]. Supernet [Ano95b, KGB^{+}96, Ano85b, Ano88m, BBBC96]. SUperNetwork [Sho91]. Superordenadorea [PBM87]. SuperPascal [Han94]. SUPERPHENIX [RCR93]. Superpipelined [DRAB08, CLmWH91]. superpositions [Cyb89a]. Superpower [Sug80a]. Superproblems [Nor84]. SuperQuest [Ano88v]. supersafe [Ano96-31]. superscalar [CLmWH91, LY90a]. Superseded [Ano96x]. Supersystem [LCH87]. Supertoroidal [Dra90b, Bue91a, DF90a, Dra90a, Dra91a]. Supertoroids [Bue91b]. supervised [McA92]. SuperWeb [AISS97]. Superworkstations [HB89]. SUPII [CGLY96, CGLxx, Chexx]. supplant [Sug80a]. Supplement [ACM89b]. Supplier [Vro94]. Suppliers [Ano91r]. supply [Ano92o]. Supplying [EDJ^{+}10]. Support [ASS94, ABCE97, Ano94-112, Ano94-136, ATSA93, AZ94, BK91a, CCSR92, Dil93, EFPSS93, FBZ92, FNT93, IJM14, Iwa92, KEMB99, Kue87, KCZJ14, LB96, MS94c, MR95, PC94a, TBC94, Ano86a, BP92, Das94, GTV91, KC95, Per86, WLLZ20].

Supported [Ano94-129]. Supporting [Ano93t, HHK19, KWH94, SK93b, Kon87]. Supremacy [LWY^{+}20]. Suprenum [PTT89, Gil88, Gil94a, SS90b, SSxx, McB92a, McB92b]. Suprenum-1 [McB92a, McB92b]. Surface [AD97, BISB96, Gre91b, HHSW93, HW97, KT94, Las92, Sky94, SHG95, TK93, TSCG94, WHMA97, ZL97, CBA90, Ver95]. Surfaces [SE90, SA94]. Surgeons [Coc03a, Coc03b]. Surplus [Gre94]. surprises [Ano88y]. surprising [PF90]. Survey [Ano94j, CLB19, Fet95, PC94a, Pic91b, PMP21, TVT^{+}16, AS93, Ana91, Ano93b, AB96, BBB^{+}20a, Don86, Gha84, LAL02, PF90, Rob85, Rya13]. Surveyor [Dun92, Wic92]. Survivability [SKC02]. Survivability-Lethality [SKC02]. survival [Bla97, Law89]. survive [WZ87]. suspect [Tay95a]. Suspension [JCJY94, RW94b, SQM94, SP94]. Suspensions [HU93, Ano95-39]. Sustainable [cFC07, TRLD13]. sustained [HPH^{+}20]. SV1 [BBC^{+}00]. SVD [Ber90c, Ber90b]. Svoboda [Ano98d]. SW4 [PLS20]. swapping [BBW19]. Sweeps [YA93]. Sweet [Ano92z]. Swimming [Ano95-46]. Swirling [Soe94]. Swiss [GG97b, New14]. Swiss-Tx [GG97b]. Switch [CH90, SSGH94, DuB90, DR91]. Switch-stacks [CH90]. switched [DuB90, DR91, GS94e, Joe87]. switches [MS84]. Switching [Ano94-105, Clo96]. Switzerland [Ano93c, Ano96-43, Ano97-30, GT94, Hen97, LM92, Mar86, Ano95u, HKR94]. sword [RR95]. SX [DTV00, Dub87, HLPP97, Ho88, HMKI97, Hor97b, Hor97a, Iwa90, jJ88, MM91b, PK89, PK94, STSK95, SWL^{+}91, TOY96, TW92, Tze88, Wat87, YSK^{+}96, Yau88]. SX-2 [Dub87, Ho88, jJ88, MM91b, PK89, PK94, Tze88, Yau88]. SX-2/400 [MM91b]. SX-3 [Iwa90, TW92, SWL^{+}91]. SX-4 [HLPP97, HMKI97, Hor97b, Hor97a, STSK95, TOY96, YSK^{+}96]. SX-4M [DTV00]. SX2 [Sha87]. Sydney [Ano92g]. Symbolic [Ano94-109, GHNL87, Hag90, HP91, HP92, Mur91a, Ste00, TP95, Kog91, PHK88, SW99, KDP^{+}14]. Symmetric [BHLST94, Joh97, Wij89b, BS88a, Bis94b, Bis94c, DS86b, HLTZ93, LPS86, LP86, Nan86, Pin91]. Symmetrized [ML93]. SYMPLER [KDP^{+}14]. Symposium [ACM95b, Ano88t, Ano88u, Ano91q, Ano91x, Ano93i, Ano93-31, Ano94-75, Ano96a, Ano96q, BG91, Bup87, Dup86, Emm85, FJSP95, HHK94, IEE93b, IEE94a, IEE95d, JT87, KK93, KMG96, LLR93b, LCV90b, LCV90a, M^{+}95, MN91, NAS93, SF91, Sie94, IEE94d, TC94, USE00b, ZAS94, Ano91m, Ano92g, App96, Emm84, Guo94, IEE96b, IEE96c, Uni87a, Ras91, Rol96, UL89, Uni91a, Dup87, FJSP95]. Synapse [UR95]. Synapses [MZ95]. Synaptic [KW95, MZ95]. synchonization [HY89]. Synchronisation [OB95]. Synchronization [BK95b, HW96, SSOH95, ABP92, AB95, BP89a, Bec89b, BP90, BP91a, BP91b, CSY89, EGP88, Jay87, Jay88a, JOK^{+}18, Li91, MP87a, MP87b, MP91d, PJ90, Su92, TZY88, Tan89b, TYZ90, TS91]. synchronizations [ZGL14]. Synchronized [PVA94, JS86]. synchronizers [Jay88b]. synchronizing [Pol88d]. Synchronous [Dra91b, GS94e, KY91b, SB96, Dra89, JG99]. Synergy [Coo95]. Syntactic [Pug94, SD93]. Syntax [PL91a, PL91b]. Syntax-directed [PL91a, PL91b]. Synthesis [BHD^{+}05, GB92, Hei89, HL91, HLxx, JBI91, LD93b, Lan94, M^{+}95, WGR93, Pic88, Scr88]. Synthetic [BD94, War93a, YWD94, YWDxx, Gig94, YW94]. Syracuse [Fox97, IEE96a, IEE96b]. System [ABB^{+}03, ABBB94, ABCE97, Ano94l, Ano94-92, Ano94-113, ABM88, ATSA93, BK95a, BCC^{+}09, Bel93, BK97, BMSD94, BMSP94, BS01, BK95b, BPU94, BPUS94, Bro97, Cal85a, CSB89, Car94a, CWLT97, CS84, CS86a, CV95, Chr90, Coo95, CF94, DHLS97, Div97, DMKW93, DGJG93, Dro95, DAKM98, Eck93, ESMH93, Ent99, Fah94, Fon85, Fos93, FXAC94, FNT93, GGG^{+}98, GY93a, GJP94, GCB92, GY93b, GCS94, Hai97, HL96, Har94a, HLxx, HNS94, HM93c, Ho88, Hor97b, Hor97a, HCH95, HERC95, Hus86a, Ike95, IHK93, IK93, JJYL94, JS95, JCJY94, KWH94, KRVJ94, KTKK93, KSTB94, KH87, KLN90a, Koe96, Koe97, KCM02a, KCM02b, KTNM93, KR94d, LMT95, Lee94, LCD97, LYKM97, MWB95, MSW96, Mil97a, MW81, Mor01, MS94d, NT89, NK96, Oed92a, Oed92b, OS93, OGOR97, PA93b, PZA87, RWCA94, Ram94]. System [RS94a, Ris94, Rus78, Sch94b, SD93, SWSR97, SHA^{+}92, SKAT93, SDB94, SG94a, Ste95, TF94, Uni89b, UEGM93, UP01, VD94, VVKB96, VF93, VT95, VY88, WMKS96, WKL^{+}16, WG94, WOK^{+}00, WJ94, WWY93, YOY97, YS94, Mob12, AP87a, AD88, Abr88, Ali86, ARF12, Ano90a, Ano91g, Ano93v, Ano93-36, Ano94-29, Ano94-63, BAAD^{+}97, Ber86b, Ber86a, BSD^{+}20, Bin88, BS87b, Cra92, Car93, Car92, Che83, Che93b, Chu87, Coc01, Dan91, DKLS86, DT08, Don93b, EE93, Ele93, EM78, FKL^{+}08, FTT97, Gal88b, GMW91, Gal91, GBC^{+}05, Gar92, GBS18, GW95, GY92, GTV91, HPPF94, HLA^{+}18, Har89, HCD^{+}18, Hod87, Joh88, KC93a, Kha91, KLN90b, KTN^{+}14, Kon87, KS88, KHZ^{+}08, KB18, LM90a, Mar91, MWRK18, ME87, MU83, Nat87d, Nat89b, Ng95, OMA^{+}96, Pad89, PSG03, PMS^{+}08, PT92, Pet89a]. system [PW23, RAG11, Rei88, RMM20, RCZ93, RIB^{+}13, SCV01, Sam85, Sha87, SC04, Smi81, SE98, Suz89, TS91, TJC91b, Tuc91, Tur79, TV88, Uch86, Wat87, WWTE92, YKY90, YB90, HWS^{+}88, HL91, MPSB87, Sho91]. System/370 [MPSB87]. Systematic [Jan96, Jia94, LFJ^{+}20]. Systems [Age05, AFT96, ASSW93, AJ97, Ano94d, Ano94-70, Ano94-66, Ano94-143, AC84b, AGLL98, BAAD92, Bec01, BKM93, BBC^{+}00, BP84, Bur94a, CCKSS90, CC94a, Che90e, Che92a, Che94b, CD92, CH10, CMAS11, CBT91, Cze93, DDJ98a, Dec90, DL96, Del97, Dil93, DAF^{+}90, EFIM91, EFPSS93, FBZ92, Fei94, Fer86, Fet95, FGKT97, FBJ^{+}94b, FL92, GJS94, GP85, GMW94, GGZ^{+}20, GHWZ94, GBG89, GT91, GMG94, GA84, GL93b, Gru97, GZA86, GK93, Hab86, Hal87, HHGS93, HO92b, HM97, IEE85, IEE95d, IK82, IBC^{+}11, Jar12, JC94c, KG94, KLY94, KB19, Kho94, Koe97, KS94a, KZ94, LPNRJ94, Leg94, LMM^{+}21, LJ97, LB82, Ma99, Mah94b, MM93b, MS94a, Mas93, MRAR95, NMS93, OB95, PDR94, PCK93, PS94a, Pas95, PS94b, Pug94, RDHC94, RCK97, Raw97, RSB94, Sin94a, SC97, Sri94, Ste94e]. Systems [SSOH95, SC92, Tan95, TGL96, TP97, TG94, TVT^{+}16, USE00b, Uch96, Uch97, VRSG93, Van91b, VKK80, VAGRMVA90, WMBC97, WP94, WF93, Wri19, YK94, YSL97, ZS93, And88, AS88, Ang91, Ano87e, Ano93c, Ano93g, Ano95-30, Ano96a, Ano96-28, Ano97j, AC84a, ALN^{+}01, Bal94, Bar88, Bau88, BSJ^{+}13, BBM19, BS88b, Bra89c, BS90b, Bru90a, CS91, Cha92b, CSY89, Che89a, Che90b, Che90d, CV92a, CGLY96, CGLxx, Chexx, Che89c, COT21, CC88b, Con87a, CKD^{+}19, Dav86b, DJM94, De 96, DSZ96, Din93, DvdS12, DS96b, Duf00, EO91, Fly66, Gal87, Gal88a, GJG88, Gal89b, GP93a, GYL00, GV91, Guz86, HRC09, HPH^{+}20, Han20, HVY91, HY89, HY91, Hun92, KS86a, Kam86, KK82, KK85, KK89b, KK90, Kel85, KBVH14, Kim96, KG03, Kra88, Kre95, Lee90, Lou90a, Lou90b, Lou92]. systems [Lun90, Lun94, MLWC20, MD04, MGS^{+}20, McA92, MV16, Nan86, PLC^{+}19, Pol86, PO88, Ram88, Rob85, Rol96, Sci86, Saa88, Sam85, SZ89, SFC^{+}21, Sch94a, SEV^{+}09, Shi95, SG92c, SG92d, SSS90, Smi91, SHL^{+}20, SS07, Sug80a, Sum82, Tan89b, Tho90, TV89, WB88, W^{+}12, War03, WvTB^{+}07, Wit89, WHL93, Yan92, YVC89, YB90, YQTV12, YK87, Zyg93, vdV91, ALPP00, Kaz92]. systems/soft [Ano96a, Rol96]. Systolic [BE93b, BE93a, KRS94, LCH87, Meh94, MM94b, MRSB94, Kun84]. Systolic/Wavefront [Kun84]. Szczecin [Elm95a].

T [Ano00a, ACA94, ACK^{+}95, CWLT97]. T-90 [CWLT97]. T21 [DJSP93]. T3D [KG95, AFT96, AHP97, ARE95, AZ95, BMT96, BAAD^{+}97, Bro97, CC96, CD95a, CCSM97, CCSS98, CP96, DR93, EAMS95a, EAMS95b, ET96, FR96b, FR96c, Gib95, HS94c, HEB96, KC95, KS93c, KLD95, Mar95, MPG96, MT97, MWO95, NSP94, PPR95, Sch96, SZG95, ST94, SWSR97, SHMH97, Str97, Stu95, Van95a, Vez95, WLN^{+}96a, WLN^{+}96b, WKHS97]. T3E [AZ99, AGLL98, ALN^{+}01, Che99, Dow98, DAKM98, GRRM99, GYL00, HE98, HPLT01, LG97, LSK04, Ma99, PSG03, RT97, SCK^{+}00, WOK^{+}00, ZCPT00]. T3E-600 [LSK04]. T3E-900 [HE98]. T800 [HMSS87]. Table [BS94d]. tables [Ano89p]. Tackle [Ano92x, Sha95b]. Tackling [MK97]. tactics [Don93c, Ste85]. tag [RFS87]. Tags [AKDM93]. TaihuLight [AYL^{+}18, ZFF^{+}18, CLF^{+}19, CLY^{+}19, HYL^{+}20, LFJ^{+}20, LWY^{+}20]. Taiwan [Chi16, Kah97]. take [Ano96-28]. Takes [Coc02a, Coc02b, Cop93]. Taking [Jon96, Zim96]. Tale [BvRS^{+}11, SW94]. Tales [CCKSS90]. Talk [Ano93-38, Erw84, PSB01, She93]. Talking [Ano94-87, Ano94-121]. talks [Ano97j]. Tallahassee [CL91, DP91]. Tallinn [KK93]. Tape [Chi95, Ano94-29]. Target [Ano93-38, Gel11]. targeted [MKHY95]. targeting [SFC^{+}21]. Targets [Ano92-44, Ano95g]. Task [ABBB94, Ano94t, Cal85b, CD08, DA94, GM94a, Gro90, HB96, HNS^{+}10, Ste94b, TGV08, Wal92, YG92, YKK96, Abr88, BP92, BGH^{+}05, EB18, FMT91, FP91, GP91c, Par90b]. tasking [MSW91]. Tasks [CD92, MGA94, NMS93, Chu87, EBB^{+}20, Mil87]. Tata [Sin08b]. Taxman [DDJ98b]. Taylor [Ano00a]. TB [Ano95y]. TC [Ano97w, Ano97-28]. TC2000 [WGR93]. TCGMSG [GB96]. Tcl [DDJ98b]. TCMP [SKIY97]. TDM [FI93]. Teachers [ORS94, ORSS94]. Teaching [EP 97, Roh94, Som13, AGEL13, NSW08, WF08]. teaching-oriented [AGEL13]. Team [Don93c, Hug94]. Teams [Pin01]. teaspoonful [Bas95b]. Tech [Bar01, CCKSS90, Gar01, Ano93b, Mou89, Mou90, RM12, She93]. Technical [AHSS93, Ano85b, Ano97-28, Bai97, CKS99, Kad94, Kho94, Mur97, Natxxf, Nor97a, Ras91, SBJ90, Ano94-108, Das94, Wuo94]. Technician [HED93]. Technique [KKF96, RP94, SKSD94, EB18, Gri92, Gua88d]. Techniques [AM93b, Ano94-34, Ano94-102, Ask93, BBC92, Bro97, BV93, CWW94, CMP94, EJL90, FV94, FJP94, Fru93, Gal96, GS01, HS96, HSxx, HKMCS94, Hug93, KABG95, KT94, LB94a, NM93, OPR01, OP96, Pay97, SM94, SLRP95, TAAL95, VS99, WS90, WJ94, YSL97, Ano89f, Ara96, Arb92, Bra88, EB91, Hun91, LY90a, LKFU05, ME96, MP91d, Nee90b, NBC92, RFS87, Ren97, RGH17, Saa88, Sch90a, SPP^{+}05, TYZ90, WS87a, WS87b, WS87c, Wil10, ZMDS96]. Technological [Asp93, CCKSS90, Pan96]. Technologies [Ano93v, Ano94-60, Bla93, BBC^{+}05, FT96b, FTT97, GCY^{+}08, KHC14, SKC02, SMS95, Tho93d, TOY96, YSK^{+}96, Ano92m, Ano96h, CBM^{+}05, DWV92, Hin93, IEE95c, IEE96a, Mag10, MBM^{+}20, Rat87, Wat92b]. technologist [Sug80b]. Technology [AM91, Ano91h, Ano92-43, Ano93i, Ano93-31, Ano93n, Ano93-42, Ano93-44, Ano94-75, Ano94-107, Ano95-47, Ano96a, Ano96-27, Ano97-29, Ber07, BS94d, BKM93, Bro91b, BEW82, Cer95, CCKSS90, CP94a, Com98, DM96a, DM96b, Egg94, FNP^{+}84, Fos03, HMKI97, IEE95a, Isk96, IBP^{+}05, KTKK93, Ker94, Kol81, Lan94, Law00, Lin82, MI93, Mor92b, OS93, Sch95a, STSK95, Sma93, DDJ98b, SK93b, Sun94, Voe89a, Voe89b, Zor89b, Zor89a, Zor90, Zor91, Zor93c, Ano85b, Ano92u, Ano92y, Ano93b, Ano94-119, Cha94a, D^{+}95, yFH89, WAD^{+}89a, Gra93c, Guo94, IU87, Leg90, Nat88b, Per83, PBK96, Rol96, SMM17, SR10, Sum82, Wil88b, WAD^{+}89b, CCKSS90, KWW92]. Technology-based [Isk96]. Technophile [Due89]. Tektronix [CCKSS90]. Tele [JLC98]. Tele-Immersion [JLC98]. telecom [DSZ96]. Telecommunications [Waz89, De 96, Mun04]. telecomp [DSZ96]. Telegraphos [KMKD97]. Telemicrocopy [Ano93-43]. Telerobot [Dil93]. Telescopes [BKT94]. Tell [Ano94-118]. Temperature [DB94, FJP94, COT21, Lee84, WH94]. Tempering [SPGD98]. Template [OP96, SL99, Sch90b]. Templates [HC99]. Temporal [BPL95, HV95, PCK93, RBK95, Ruh95, Ste95, VV94, VT95]. Ten [Tri95c]. Tennessee [Pan96, Chi90, IEE94c]. TensorFlow [KSM^{+}19]. tensors [Ara14]. Tentative [Eis95]. Tera [Ano95a, Ano97j]. Teraflop [Ano00c, Ano00d, Fox90b, Ano97h, Ano98g, Gar92, MSW96, Ano94-130]. Teraflops [Con00, ACC^{+}96, CSFS00, Ano97j, Bem92]. TeraGrid [BFS11]. TeraGyroid [BCCP05]. Terascale [FKL^{+}08]. Terasys [SK94, CP93a, GHI95, IGH95]. Term [DA97, HM93c, Per83]. Terminal [Deh90, Fei94]. terminals [Way96]. terms [Ano97b, Ins87a, Ins87b, Ins90].

TERPSICHORE [ACG^{+}90]. Terra [DHA^{+}13]. Terrain [Fie93, Max81, OLLG96]. Test [Ano94-109, Ano97u, Bel93, BS94c, Gru97, HTI93, HED93, OBR94, RP94, SB96, SJA94, CSFS00, HBB^{+}05, Sta88, IK91]. Testbed [KGB^{+}96, SHB91, Mes93b]. Testbeds [Ano90i]. Testing [Ano05, DAF^{+}90, FS93a, HS93b, MMK97, Raa97, Tay95b, Ano91m, Ano96-33, CMP94, Par90b]. Tests [Was96b, Wil94, KB18, PP91]. Tetrahydropyridazinones [JBI91]. TEX [And89]. Texas [HS^{+}91, Nas91, Uni89b]. Text [Deu86, PTS93]. textbook [Ano85b]. Texts [Pug94, SLML93]. Texture [CM93, RPY94]. TFLOP [MSW96, GKS09]. Tflop/s [GKS09]. Tflops [Ano94-59, GGG^{+}98, HFCM98, MH98, Mit98]. Thailand [US01]. Thalamo [VT95]. Thalamo-Cortical [VT95]. theater [SIG90b]. Their [Che94b, Her89, Pet97, Ano94-128, Bab94, CV92a, CK92c, DWV92, GS89b, Gil94b, LK93, Mik89, MRAR95, MS93, Pol88a, SH94a, Tze86]. them [AL92b, Bur94a, Sha95b, WZB86]. thematic [Ano96l]. theme [Cha94a]. theorem [HA90b]. Theoretical [HS94a, NU91, SH93, BMR85, Gir91, OMM93, SH94b, Sho91]. Theories [STN93]. Theory [ALPP00, BJ93, Dec90, GAW96a, KMG96, Lei91, MTK93, RMH93, Sch92b, VRSG93, Vuj93, Wei90, Ano92g, ALN^{+}01, Deg90, KM85, Lei89, Mag10, McC88, MHP84, MR86, MR87, MSTK93, Ons88, PT92, Poo96a, RLKW93, Sie90, Win02, GAW96b]. Therapeutics [VSB^{+}21]. therapies [Ano97-29]. There [Ano93p, DCG93, DCGxx, DT96, Ano89a]. Thermal [ANS92, KLSC97, KO93b, KS90, Sch92b, TGV08, Asl91b, Woo92, Woo94]. Thermal-Aware [TGV08]. Thermal-Hydraulic [KO93b]. thermal-hydraulics [Woo92, Woo94]. thermally [KfGERJxx]. thermally-stratified [KfGERJxx]. Thermo [NS93]. Thermo-Fluid [NS93]. Thermochemical [VA94]. Thermodynamic [OGOR97]. Thermofluids [BGIM90]. these [Sha95b]. thesis [Pan96]. Theta [HLA^{+}18]. they’re [Rya90]. Thick [PA93a]. Thin [Ano95w, Hua92, PIH04]. Thin-film [Ano95w]. Things [Tri95c, Nor17]. Thinking [Lee94, Ano92-44]. Third [Ano88w, BP93, C^{+}97, DT97, HM93b, IHIS91, KK88, KK89a, McC88, IEE94d, Wei92, WLN^{+}96a, WLN^{+}96b, Ste01a]. Third-generation [Wei92]. third-party [WLN^{+}96a, WLN^{+}96b]. Thirst [Dak90]. Thirteenth [TC94]. Thirty [Bel86]. Thirty-first [Bel86]. THOR [Dav00]. Thorp [Ano94s]. Thought [Han89]. thoughts [Ull83]. Threaded [EHG95, AACK92]. Threads [Ano94-87, FHM95]. threaten [Ano95l]. Three [ACG^{+}90, Ano92-38, Ano94-103, BvRS^{+}11, BB91b, DLPQ94, GWH93, HFT94, Hsu16, KBG^{+}13, KJ85, Koc93, KO93b, LMM85b, LMM85a, LMM86, Man90, OK93, OMR93, Rue92, SSG93, SBHW80, Soe94, WD93a, WFT93, Woo96b, Ach99, BR95, BS91, DD90, HL88a, HP88b, LLDF95, MB97, Nee90b, Nix92, SWL^{+}91]. Three-Dimensional [ACG^{+}90, DLPQ94, GWH93, HFT94, KO93b, Man90, OK93, Soe94, WD93a, KJ85, SBHW80, BR95, BS91, HP88b, LLDF95, MB97, Nee90b, Nix92]. Threshold [WSP95]. Throughput [BCW20, CCYT05, FSC18, GS94e, HHK19, SSSR20, SL92]. throw [Ano94-128, Ano97y]. TI [Ano96p, Wat72]. TianHe [JTX^{+}22, CYXL18, GZW^{+}22, LXW^{+}16, PCY^{+}19, PZS^{+}20, WCZ^{+}18]. Tianhe-2 [CYXL18, LXW^{+}16, PCY^{+}19, PZS^{+}20, WCZ^{+}18]. TianheGraph [GZW^{+}22]. tidal [Nix92]. Tide [Ano95-46]. tie [Ano93-40]. tier [Sca92]. Tight [Ano94-137]. Tightly [ADG^{+}08, GIF^{+}12, Har86]. Tile [OSKO95].

Tiling [BDRR94]. Time [AC93, Ano87a, Ano94-54, Ano94-70, Ano94-138, Ano96-41, Ano18, ACL93, CWD^{+}08, Col94, Dav92, Don92b, EFPSS93, EKZ90, FT93a, Fon85, FLP^{+}07, GSG^{+}94, GH93, GWH93, IK93, KG94, KFF93a, Koc93, KKB92, LS92a, LMP^{+}90, LZ95, LL94, NT89, Now93, PCK93, Poe95, RP94, RAP95, RW94a, RDZ93, Sin18, SB94c, SF93b, Taf96, TFB94a, TFB94b, UP01, VV94, Ano94-42, Ano94-112, Ano96n, Bru88, Cho90b, CH92a, Dan91, FB91a, FY96, FP91, GMF00, GBS18, Gor89, GREC91, Hab92, HS94b, Hen91, JOK^{+}18, Lee90, LM13, MP90, Mil87, Mir88, Pol88b, Pol89, RGH17, Sha90, TJC91a, Ver95, YH92, CCKSS90]. time-consuming [GBS18]. Time-Dependent [GH93, GWH93, Now93, RDZ93]. time-efficient [YF98]. Time-Life [CCKSS90]. Time-parallel [TFB94a, TFB94b]. Time-sharing [Fon85, Mir88]. Times [Ano94p, Ano89n, Bas95b, Ber82, MV16, YAG93, YAGxx]. timing [BHD^{+}05]. timings [Sit78]. Tire [Ano93-44]. Titan [DHM^{+}88, KSB^{+}19, LM90b, Mir92]. Titan-2 [LM90b]. TM [Gro93, PR94a]. TMAPOL [Rob87]. TMC [KC95, SNS^{+}97]. TN [SJD96]. Toaster [Sol94]. Today [Fru93, Gup94, Hag01, HG02, KDLS86, Bel92b, Ber82, Kuc87, Sha95b, Supxxa]. together [Ano96u, HHS01b, Pol90]. token [Bro86]. Tokyo [ACM93, Ano97l]. Tolerance [MNZ^{+}15, Con00, The90b, The91]. Tolerant [Ano94-53, CRV94, CB94, EVM^{+}98, GFB^{+}03, GGZ^{+}20, GBG89, LBT94, RWNJ94, HCL88, SHL^{+}20, Tze86, TYZ88]. Tolerating [Ano94-139, CMHK92]. Tomography [HEJM95, GS89d, HWP95]. Tomorrow [Bau96, Egg94, Fru93, Hag01, HG02, Gro92c, Min86]. Tonnen [Ano97d]. Too [Ano94j, Ano89a]. Tool [Ano94-28, CT93a, DKF94, Eck93, Geu97, KS94a, LLSR02, MKRI93, PC94a, PCK93, Rei93, Sei94, TSSK94, Uth94, Ver97, Gan88, JG88, KK89c, MHJ91, PSS^{+}19, SM89]. Tool-Based [PCK93]. Toolbox [BK97]. Toolkit [AGKT02, BS94a, MMRL93, BS94b, Hei90, Abe90, Abe91]. Tools [Ano94-102, Ber95b, DT97, GHdF10, GL94, GRSS93, HMC94, KEMB99, MM90, MP91e, Rag06, IBM13c, TF97, ASK85, AB96, CH89a, EO91, IKM85, Kin96, KAMB19, LP90, MLWC20, Mal86a, MSW^{+}05, Sit78]. toolset [McC92]. Top [Ano97d, RM12, Way96, WF08]. Top-500 [Ano97d]. top-down [WF08]. TOP500 [DZM^{+}13, Str03, SMDS15, Fei05]. Topic [MBSW01, Res01]. Topical [ANS92, Ano95-38]. topics [Spe87, Tru88, ZLC21]. Topographical [BM93a]. Topological [Ver95]. Topology [CTD^{+}16, EH97a, LA94]. Topology-Aware [CTD^{+}16]. tops [Kog11]. Tori [CB94]. toroidal [PS88]. Torque [DNV93]. Torus [BS94e, KMNT96, ABC^{+}05, FMD07, FDM07, KTN^{+}14, RD07]. Total [GF90, Rei93, Bue91a]. Touchstone [HL93b, KRJ93, TFB94a, TFB94b]. Tough [Ano92x, Ano96z]. Tour [Cra96]. tourist [Ano11b]. Tournament [Coc02c, Coc02d]. Tower [Yei92]. Toxic [GD97]. Toxicant [Pli97]. trace [MHJ91, SSS90]. Tracer [DS94b, SYMT92]. traces [NRM^{+}09]. TraceView [MHJ91]. Tracing [GMMT91, MLR90b, MLR90a, And90c, Def87]. Track [Ano92k, Ano92q]. Tracked [BPU94]. Tracks [Bel93]. Traction [MS94b]. Tractor [KB97]. Trade [Ber95b, FLP^{+}07, Smi96c, Tys91, RYYT89]. Trade-Off [FLP^{+}07]. trade-offs [RYYT89]. Tradeoffs [Ano94-90, Lin82, WKL^{+}16, DKLS86, Smi89]. Trading [Cre91, Ano96-41]. tradition [Min86]. Traditional [LCP^{+}11]. Traffic [BGMR96, BWO96, BY96, Bot96, BM96, BP96, ER96, Hel96, Hey96, HK96, Jan96, KNWB93, Leu96, MJH90, Nag96b, Nag96c, PN96, SS96a, Sug96, TKM96, Wag96, WMR96, Woo96a, GTV91, Par90b, TV89]. Trailing [HS93b, MFK94]. Trailing-Edge [MFK94]. train [HPS88]. trained [SNEP14]. Training [BBHL01, CS93b, CS95, JPMG08, MB20, CFP20, Han03, KSM^{+}19, LS93a, Mou89, Mou90, SBC91]. Trains [GW93c, KBD97, RBK95, TFO94]. Traleika [CCG^{+}17]. Trani [MM91a]. transactional [OWG^{+}13]. Transactions [Pin01]. transducer [Ano90a]. Transfer [MS97, MHE97, Now93, SSOH95, WG94, A^{+}02, DLS93, KfGERJxx, PMS^{+}08, Sha89, Su92, VO93]. Transfers [Ano94-111, GJG88]. Transform [ALM93, BL93, BB93, Cla96, WF94, CC88a, ESTA94, MB97, Sul91]. Transform-Vector [BB93]. Transformation [Col94, KHN89, MLWC20, Pre93a, VMS93, WWKR97, BM85, KS93a, Pol87b, Tan89b]. Transformational [OH92]. Transformations [Ano94-47, BGS94, HS95a, KFF93a, KKB92, OGR95, Ban90, CDG^{+}06, Poi90, SKP91]. transforming [Bru91]. Transforms [KFF93b, Heg96, NR86]. TRANSFUSION [RDZ93]. Transient [Cha92a, CBT91, FBGM93, GH93, KR94c, TfGERJxx, VTTS98, WD93a, WFT93, Deg90, Elm93, VM07]. Transients [AFT96, SW96]. Transistor [Ano92h, Ano03a, MS84]. transistors [Ano03a]. Transition [LC94, LUT96, Uni93]. Transitions [BY96, TKM96]. Translation [AK87, RC94, Uch86]. Translator [DP96, WG94]. Transmission [KC93a, MZ95, Ano87d]. Transmissions [Cig97]. transmittance [TM88]. Transonic [Riz94, SHZK94, Gri86]. transparent [CWD^{+}08]. Transport [AD97, ALM93, Ada93, AFML93, Bak93, BL93, DMPR93, Fie93, FA93, GHW93, HL93a, JWG93, KE93, MKND97, MS94a, MNR86, MMRL93, ML93, Nat84, PA93a, PH97, Sus93, TKI93, Uen93, Vuj93, WD93a, WABD97, YA93, YCC97, ZL97, ARW93a, BLFT84, Chi86]. transportability [KFF93b]. Transportation [CJ93, FD93, ITOK93, TIOK94, Ano94a, Ano94-75, Ano96a, NSH95, Rol96]. Transputer [CM95, MVS94, WH93, BK89, DJM94, HMSS87, IAIK92, MRSB94, WB88, DJM94, Row86]. Transputers [LW92, Vol89, BBC^{+}89, Hey90, Wal90]. Transtech [Ste94c]. TRANSYT [Par90b]. TRANSYT-7F [Par90b]. Trapezoid [Mas92]. trash [Dum97]. travel [SCH94d]. Traveling [Gur88, Bra88]. Treat [Ren97]. Treatement [BMP93]. Treatment [AFML93, Hey96, Zla01, MHKY97, OMA^{+}96]. Tree [CRY94, SSM93, BAAD^{+}97, BAD01]. Tree-Structured [CRY94]. Trees [AFAGR96, Lei85, BJZfDA96, Fid91]. tremendously [HHS01b]. Trend [Jal94, TK89]. Trends [Ano94-67, Ano95-36, Ano96p, AGL11, Gar01, GGN20, IU87, KYS88, KT11, Koo97, LQFC18, Meu89a, Meu90, Meu91, Meu92c, Meu93, MG95, Pay97, Sch95a, Sch93a, SHMR96, SJD96, Sne94a, VN04, Ano95u, Ano96u, Ano97m, Cyb91b, Duf00, HKS93, JPTE94, SHMR94, Meu92b, Meu95]. trendsetter [Gil88]. triangular [AS88, RG92]. Trichloroethene [WWKR97]. TRICOMM [VO93]. Tridiagonal [VY88, BS88b, Gao86, LPS86, SE98, WB88]. Tridiagonalization [BSL94]. tried [Ano94k]. trillion [Ano97s, Bas95b]. trim [Bro17a]. TRIMARAN [Mon93]. TRIMARAN-1 [Mon93]. triply [RLKW93]. TRIPOLI [MNV93, MBN93, VNB93]. TRIPOLI-2 [MNV93]. TRIPOLI-3 [VNB93]. Trondheim [Kow89b]. trouble [Ano95x].

Troubleshooting [CFP20]. Truck [BPUS94, HRG93]. Trucks [MS97]. Truly [Ano94-141]. Trust [Cla18]. Try [Poo96b, Poo96a, Ano93-41, Hol93]. trying [Ano95a]. TSMC [Ano03a]. Tsukuba [NN90]. Tube [KS93b]. Tunable [WF94, SEV^{+}09]. Tuned [GAV95]. tuning [BLW11, Con88]. Tunnel [Ano94-32, FI91, GI93]. Tunnels [Ano94-68]. Turbine [MKG90, Ris94, VH93b]. Turbocharged [BPW97]. Turbocharger [OGOR97]. Turbulence [BSB93, Gha96, KfGERJxx, LC94, MKDY90, MK97, OK93, PB94b, Pop91, Soe94, SJDV09, WG93a, BS91, FKL^{+}08, KSP13]. Turbulent [BD93b, CAB93, FA93, GFM96, GWG93, GWH93, HVSB93, KS93b, PSB01, Riz94, GB22, KfGERJxx, ZBN^{+}19]. Turing [Hun93]. Turn [Ano88x, Sch12]. Turnaround [Gar99]. Tutor [Ano88n, Pet89a]. Tutorial [FJSP95, Hwa84, Ano88w, Pot88]. TV [Ano90j]. twelfth [Ham94]. Twelve [Ano87d]. twentieth [YB86]. Twenty [Sin94a, Gra94]. Twenty-eighth [Sin94a]. twenty-fourth [Gra94]. Twin [YMZ90]. twinning [GE12]. TWIST [SK94]. Twisted [SIKD94]. Two [AD97, ADLL01, Ano95l, ACL93, AK93, BSJW96, BY96, BSL94, Che90a, DLMW95, EFPSS93, EBS88, FI93, GMBW93, Hun94, JV93, Lu93, MP94, MT13, MM94c, MM94d, OK93, OMR93, PC93, RT93, Rei88, RBL94, Saa93b, Sch93b, SW94, TKM96, VLA92, VD96, Fin94, Ano94-120, Ano94-121, BS91, CC88a, CGL92, MSCxx, KMB09, LM90b, Sca92, SM92, Wil88a, WMK90, WBP87]. Two-Box [OK93]. Two-Dimensional [AD97, AK93, BSJW96, Hun94, TKM96, VD96, BS91, SM92]. Two-electron [EBS88]. Two-Fluid [PC93, RT93, Sch93b]. Two-Group [JV93]. Two-Phase [Saa93b]. two-point [KMB09]. Two-sided [MM94c, MM94d]. Two-Stage [FI93].

Two-Step [BSL94]. two-tier [Sca92].

Two-Time [ACL93]. TX [HS^{+}91, GG97b]. Type [BLO94, SKIY97, SZ89, KHN89, LPS90]. Types [GAV95, Kim96, WHL93]. Tyre [DM96a, DM96b].

U [CP92c]. U.S [Por86, Tec89]. U.S. [Ano93-46, Ano97u, BS92, BF91, Fed87a, WAD^{+}89a, Hol84, MA85, Nat92a, Nat88b, Tec89, Sch18, Uni92a, Uni92c, Uni98, Wil88b, WAD^{+}89b]. U.S.-Israel [MA85]. U.S.-Japan [Uni92c]. U.S.A [KK85]. Ubiquitous [BGH^{+}02, FT96b, Ano96h, FTT97]. UbiWorld [DPS97, PS96]. UCAR [Ano97j, Nat86a]. UCITA [Gar01]. UIS [Sch94b]. UK [BP89b, BP93, HK93b, KH93, Ano92-42, Ano95v, Coo95]. Ulam [Ano87e]. Ultimate [CCKSS90, NW97, BDRR94]. Ultra [Car91, FBCB18]. Ultra-Green [FBCB18]. Ultrafast [Gar99, SF82]. Ultrahigh [AM91]. Ultrahigh-Speed [AM91]. Ultrasonic [DGT94]. Umpire [BMD^{+}20]. UMSI [UMS84]. UNAM [C^{+}97, Fra94, GG^{+}97a]. UNAM-CRAY [Fra94, C^{+}97, GG^{+}97a]. Unbalanced [Ano94e]. Unbroken [Sul97]. Uncertain [Ano94-98, Ano95-35, Sat93]. Uncertainties [VRSG93]. Uncertainty [GIBGA93, IK93]. Unconditionally [AABB93]. undecidable [RJ13]. undercarriage [HPS88]. Undergraduate [Ano94-71, MAFW08, SC91a, WF08]. Undergraduates [BBHL01]. Understanding [Bar93b, BCC^{+}08, CCKSS90, DDLV93, Dun99, Hor98, KA93a, TMP94, Wor84]. Underwater [Nor17]. Unesco [Ano90g]. UNI [VD94]. UNI-ECU [VD94]. UNICORE [AS99]. Unicos [Rit88a, Rit88c]. Unification [Sch95b, CRM94]. Unified [Dic94, DFS93, HBDS93, JC94b, Pet97, TAAL95, YS94, PLC^{+}19]. Uniform [CWW94, KB94, OP96, WRW93, AS99, Rob89]. unifying [CS91]. Unimodular [Ano94-47, Ban90]. Uninitiated [Wil96]. Union [Ano89o, WG93b]. Uniprocessor [Cal86a, Cal86b]. uniprocessors [KW92]. Uniquely [Mer86]. Unisys [Hod87]. Unit [GIBGA93, IHE^{+}00, KISY94, CBB^{+}05]. Unite [Ano97p]. United [ACM94, Bro91b, ML95b, OMM93, Tho93c, Fis83, KLQ19]. Units [BH17, KB97]. Universal [FJ91, GM94a, Ros95, WG94, Jor86, Lei85]. Universality [CF92]. Universe [Fra94, WS99, BAD01, Pri12, Yos09]. Universidad [C^{+}97]. Universities [Dal84, Fer83, Ano97-27]. University [Ano88d, Ano97e, Ano22, Asp93, CL91, Cra96, DP91, HS^{+}91, NAS93, Nas91, NBC92, Joh86c, AA93, Cor89b, Goo88, Min92, Pan96, Uni89b, US01]. Unix [Hil97, BE88, Bin88, Car89b, Chr90, Coc03a, Coc03b, USE90, GL94, MK92a, MK92b]. unleashes [Ano95-32]. Unlimited [YSKS95]. Unloading [DM93]. unperturbed [Eij91]. Unsteady [CAB93, HFT94, HS94d]. Unstructured [Ano94b, Ano94-114, Ano94-117, BS94e, JP94, MR95, NĆ02, NK94, PPM90, VTTS98]. Unsymmetric [GMW94, HV94, Van91b, DD88, Dav89, DY90, DS96b, GMW91, Wij89b]. Unveil [Gar99]. unveiling [Ano95h]. Unveils [Pau09, Ano93h, Ano93-27, Ano93-28, Ano94-27, Ano03a]. Unwrapping [SPGD98]. Update [Ano90j, Ano93-45, Ano94-142, Ano95-48, Ano95-49, Ano96-42, Ano97b, Ano97l, Ano98d, Ano08b, FY92, Ros08, Sin08b, Ano89p, Natxxa, Natxxd, Ano96d]. updated [ALPP00]. Upgrade [DBK09]. upgrades [Fed96]. Uploads [Bar00a, Bar00b]. Upon [Sch22, AJFH86]. Upper [LZ95, LL94]. ups [Ano03a, Ano93b]. Upscatter [AM93a]. Urban [SLS96]. Urbana [Ano22]. urged [Mac97b]. US$6 [Ano94-86]. US$67 [Ano90r]. US/ROC [COS89]. USA [Ano93n, IEE85, IEE95d, LLR93b, LCV90a, MB94b, PEH93, S^{+}93, SR93b, ACM96, ACM03, ANS92, Ano95-38, Ano96l, CBCH93, EM94b, HS^{+}91, IEE94b, IEE95b, IEE95c, KK89a, LCV90b, L^{+}93, Lim93, MW88, MB93, RD94, USE00a, USE00b, Wuo94]. usable [Pol88b]. Usage [EJL90, Nat84, VMS93, BBB^{+}20a, Pry94]. Use [Bel93, BHMH98, BS01, CU90, DD90, Dil93, EM94a, HTV88, HC93, HK93b, Hug93, KH93, LK93, MS94a, PC93, SCH94d, Ste94e, VRSG93, Van91b, Ver97, XCLW93, YS94, Ano90n, Ano92-46, Ano00b, Aus93, Bur93, Duf85, FGC06, GJM86, Gin93, IEE96a, JG88, MP88, Meu89b, Pou88, TKI85, WH94, Xu91]. Used [Ano97k, Che90e, CLP93, CPR93, MS97, Ano88q, Che90d, Ren97, Ruh95]. Useful [JA92b, JA92a]. Usenet [Coc01]. USENIX [Coc01]. User [Ano94f, Ano97o, DLM99, KCOP94, MH18, Nat86c, RRSS93, RT20, TH19, Y^{+}92, Bru90b, Chexx, EM91, GG88, Ham90, LM92, MPC89, MRS88, Mou89, Mou90, Nor89, Nor93b, SFC^{+}21, SH91, SS90c, Wom90]. user-defined [SFC^{+}21]. User-Interface [Y^{+}92]. User-Level [Ano94f, KCOP94]. USERNET [KGLA85]. Users [Ano90k, Ano94j, Erw84, PC94a, Ano92-44, Ano93-40, Ano94-27, Ano94-86, Ano94-131, Ano95g, HCD^{+}18, JT91, PF90, Uni91a]. Uses [Sah94a, Ano90a]. ushering [Zor92a]. Using [ABBB94, AJ97, AH93, Ano94-89, Ano94-93, Ano94-96, Ano94-140, Ano94-143, Ano95-31, Ano18, AAS88, AM94, AHOK02, AA93, BGMR96, BJLW95, BGS^{+}12, BS98a, Ber95b, BBHL01, BS94d, BSKJ93, BM93b, BD94, Bre87, BB98, BS98b, BK91b, Cha93, CTD^{+}16, Che99, CB02, CHMS94, CCSM97, Col94, CT94, Cox88, Dip96, DS94c, EJ97, Fah94, dCCF01, FR81, FG92, FV94, FS93b, GMW94, GSG^{+}94, GHK^{+}91, GI93, GD94a, GML90, GZA86, HS95a, HB96, Has84, INKN01, IJM14, IHSK93, IK91, JC94a, Joh97, JML96, Kau93b, KFF93a, KSTB94, Koc93, Kon96, KKF96, KB97, KK96b, KW11, KCZJ14, LMH90, LPLP97, Lu93, LCD97, MKND97, Mas91, Mas92, ME87, MNZ^{+}15, Mis90, Mon93, MM94c, MM94d, Nag88, NPS^{+}20, OK93, OPR01, OLLG96, PH11, PK80, PRS94, PW94, RS94a, RG94, SSG93]. Using [SG92a, SNS95, Sha87, STN93, SAGS93, Sil91, Sin18, Smi91, SA82, SHG95, SJDV09, SC91a, TJ94, TGL96, Van13, Wag96, WH93, Wal92, WBP87, Web93, WFT93, Wil93, WL96, YOY97, YKK96, YSL97, vdG97, AGZ94b, AZC13, AM93b, ABGL96, AM96, BBC92, Bli91, BMW91, BBB^{+}20b, BJV^{+}16, But92, CBA90, Che96, CV88c, Chi86, COC93, CS90, Cla18, CFP20, CNC^{+}08, DLPQ94, Din92, Don85, DH86a, EB18, Ece96, cFM07, Fin82, FKL^{+}08, For93, GP93a, Gok91, Goo97, HOSZ97, Han03, HBKR96, Hea91, Hun90, Hun91, HP88b, IBM01a, IAIK92, IS20, KWH94, KSP13, KDBG95, Kra88, Kue93, Lan93, Man89a, Man92, MWB95, MBM^{+}20, Mas94b, MB97, MSW91, NSH95, NU91, OYK^{+}14, ODAZ15, PPM90, PEH93, RPY94, RW94b, SCG^{+}08, SNS^{+}97, SSSR20, SZ89, SNEP14, Sat93, Smi92, SW99, Svo93]. using [TM88, TF97, TOWC15, VSH91, Vez95, WHBH93, Was96a, WQS92, Wil88a, WMK90, WOG94, Yi11, IBM01b]. USSR [Rya92]. UT [ANS92, Ano95-38, Isk96]. Utah [SC93]. UTCHEM [SPS91]. Utility [FHM95]. Utilization [WOK^{+}00, ADG^{+}05]. Utilizing [HFH86, HFH87, MTK93, Nor97b, SB01, Roj19]. UX [Ano93m].

V [WFT93, Tem83, Wom90, PPR95]. VA [S^{+}93, HKS93]. vacation [Pic92]. Validate [Wea97]. Validation [CPR93, FD97, GP93b, KE93, MNV93, MAA93b, Con00, IBM13c]. Validity [MF97]. Value [Mas95, BS87a, BS88a, Che94a]. Valve [SP94]. Vancouver [Ano91m]. Vaporization [KR94b]. Variability [FBCB18, Hey96]. Variable [BWGG94, Li91]. Variable-Complexity [BWGG94]. variables [AH90, Jay87, Lil89]. Variably [TOWC15]. Variance [Ano94-111, BL93, Mis90]. Variant [AK94, SAGS93, Cha92b]. Variation [Raw97, Wea97]. Variational [DMPR93, HHSW93, Kau93a, Rul93, WLH00]. Variations [RHH96, BtR95]. various [Don85, SPS91]. Varying [PCK93]. VAX [BMSD94, WZ87]. VAX/VMS [BMSD94]. VAXes [WZ87]. vBNS [JNM^{+}98]. vCUDA [SCSL12]. VEC [JML96]. VEC-SM2 [JML96]. VECFEM [Bra93, GRSS93]. VECFEM-Powerful [GRSS93]. VECFEM/S [Bra93]. VECTIS [MJRS94]. Vector [AK87, And90b, Ano94r, Ano94-122, Ano94-123, Ano94-130, Ano96-43, Ano97-30, Asa98, AT93a, AT93b, BOS93, BAT99, BB93, Bra93, BCHJ94, BS98b, CBCJ92, Che89b, CDC^{+}87, CP94c, DL96, DL90, DDB^{+}10, Far90, FR81, FHKT97, GS94b, Gen94, HHOM91, HHOM92, Hos88, IHE^{+}00, IJM14, JC94c, Koe96, Koe97, Kor93, KT80, KZ94, LPV94, MNR86, MNB94, MM91b, MS93, Mur91a, OIY91, PVA94, Pet83, PHVJ95, SKIY94, SKIY97, ST92, Sul91, Uch96, Uch97, UT91, WLKI95, WNKS96, Web93, Yan93, ZM86, ASM86, Ano93c, Ano95h, Ano95u, Ano95x, BJ95, Cha92a, CP92b, CP92c, DD90, Def87, Deg90, DB95, EE93, ESTA94, FFM95, Fuj99, GJ87, GL96a, GL96b, GL97, Gua87b, GS89d, GZR89, GHS86, Guz88, Haw86, Hea91, HS93c, Jor87, Kha93, McC94, Mil93, MU83, MPSB87, MHP84]. vector [NRN00, Par90c, Rav92, Rav95, R^{+}00, RR89, Sam85, Sch89b, Sch87c, Sch88b, Sch87d, Tan89a, Tho93b, Tru88, TV88, Tur89, WLH00, WLLZ20, Wij89a, tDv87]. vector-efficient [Par90c]. vector-multiprocessing [Def87].

Vector-Parallel [Koe96, Koe97, Uch96, Uch97, CBCJ92, NRN00, R^{+}00]. Vector-Processing [IHE^{+}00, McC94]. Vector-Processor [HHOM91, HHOM92]. Vector-supercomputer [Kor93]. Vector-supercomputers [Hos88]. Vectorial [JML96]. vectorised [Fin82]. Vectorizable [VH93a, BTV96, BV96, DO89, hTD88]. Vectorization [DCG90, DMCK92, Gre90c, LKFU05, TG95, WHBH93, Gri86]. Vectorized [AK93, BLFT84, Cha84, CS90, HFH86, HFH87, Max81, Mik94, OD88, PB88, SBHW80, VM87]. Vectorizing [EBS88, Tsu91]. Vectors [FBA93, VLA92, CKS99, Hil97]. Vegas [Ano96l]. Vehicle [AM93b, AJ97, AKT90, AVS93, BJ93, BPU94, Cze93, FT93a, HU93, HTI93, HP93, HD88, HF93, Koo97, LPLP97, MMK97, PYTL97, Van93, Wil94, YK94, BBC^{+}99, HP88b]. Vehicles [Ano94-99, Bel93, DM93, FCD97, Geu97, NCDS97, Ost94, Ris94, Spe97, Ano91x]. Velocity [BSB93]. vendor [Ano88s]. Vendors [Ano90s, Ano93-38, Ano88s]. Vents [HCV97]. venture [Ano03a]. VER.02.6 [MM93a]. Verification [GM94b, SZ11, WAB^{+}05, YYK93, Hin93]. Verified [UU94]. Versatile [Dav00, RDHC94, Vog93, SNK^{+}93]. Version [DFWW93, MM93a, ALPP00, CV89b, Fin82, GYL00, GV92, GS89d, MSTK93, SK93a, SSRL91, Str94]. versus [But92, HA90b, Pol88c]. Vertebrate [HH93]. vertex [GS87c]. Vertical [Fet95, LMH90, RHH96, PB87]. Vertically [Ano97-29]. Vertically-stacked [Ano97-29]. Very [Abr92, DA97, Fly66, NW97, BJ84, Ger90, LW94, Pan96, Ros95, Win02]. Vesta [CF94]. vestiges [Ano97t]. VF [DD90]. Vforce [MCLK07]. VGTC [Ano22]. VI [Fer86]. Via [DBK09, LMP^{+}90, All93, Ano93-40, Ano94-66, AT89, DB95, FJ91, GB96, Gre88a, KKKP93, Lou90b, Scr88, Taf96, TYZ88, WWJ09]. Viable [ZWP03, Ano03a]. Vibration [GA97, SBSR96]. Vibrations [BPU94, KB97, Ost94, RCK97, HL88b]. Vice [Age05]. victim [WFJ^{+}17]. Video [Ano90j, KFF93a, KFJB94, San95, TCJS93, ACM89b, GW04, Per02, SIG90b]. videotape [Mic90]. Vienna [ACM97, Fah94]. View [FI93, Kah93c, Mut94, Qui95, Wil93, HA92, Kah93b, Nat87a, Nat91a]. Viewpoint [Bel96, SS90c]. Viewpoints [TAM^{+}91]. Views [Bar00a, Bar00b, Bar00c, Bar00d, Bar01, Coc01, Coc02a, Coc02b, Coc02c, Coc02d, Coc03a, Coc03b, DDJ98a, IM96, KC93c, MB94a, DDJ98b]. Villa [DJM94]. Village [Con91, Con90]. Viper [Hof93]. Virginia [Cra96]. Virtual [Ano94-144, Ano05, CMPR93, Dil93, DPS97, DR98, Gan94a, Hel93, Ike95, KTKK93, Kul94, KB96, Law00, OB95, PS96, PBM95, RE94, SCSL12, SDFP93, Ste94e, TSCG94, TJ94, Wea97, WKFFK97, Zhe97, AP87a, Ano95-31, DLM99, GHdF10, Hir92c, MNY09, Rol97, Sch94c, Ver95]. Virtualization [MUKX06, LPD^{+}11, MMG^{+}18]. virtually [Ano94k]. Viruses [Ano94-89, Str94]. Vis [Hib01]. viscoelastic [YHA93]. Viscoplastic [ACL93, DH93]. Viscous [BCM90, Glo89, SHZK94]. Visibility [Wil92b, Wil92c]. Vision [Ara97, DM96a, DM96b, DS94c, KRVJ94, LPNRJ94, Maj94, Nal94, Gar92, IBM01a, IBM01b]. Visionary [Pin01]. visions [Ano94-119]. VisionSmart [KFB91]. Vista [JT91, Tuc91]. Visual [BPL95, BBC^{+}05, HC93, JR94, Ros93b, VA94, Ros95, GE96]. Visualisation [Bos94b, CTM94, DB94, GWH93, JJYL94, SKSD94, Uth94, M^{+}94]. Visualization [ACM89b, AHSS93, Ano90k, BCH12, Ano22, ABSS94, Arb92, BvRS^{+}11, CH89a, Che90e, Che94b, Cox88, CM93, EGH^{+}06, FK93, Fru93, FW90, FZM91, Ger90, Gra93b, HMM94, HFT94, KA93b, Kuw92, Kuw94, Lan93, LLR93a, LAPR94, MKDY90, Men90, Nee90b, NSR90, OPR01, OYWK91, RW94a, RLW93, SG92a, Soe94, SJDV09, SYMT92, SHB^{+}13, VH93b, WKFFK97, AL92b, AL92a, BPM^{+}89, Che90d, Che96, GL92, Gin93, Gri93, Hab89, Hab92, Hib01, Hin93, Kau91, KSM^{+}08, KG95, MHJ91, Nee90a, PSM93, SIG90b, Sar90, Sch95c, Sha90, SN89, TB89, TJC91a, TJC91b, Tuc91, Wil91, Woo93, Zho88, GLH94]. Visualizations [CGM91, TFO94]. Visualize [Cox88]. Visualizing [Ric91b, Nee90a]. vital [Ber82]. Vivo [SVML95]. VLA [Din92]. VLIW [Ano00b]. VLSI [NAS93, Ano95w, BM87, Bow88, CG86, DuB90, DR91, Fie86, Geo94, Lei89, Lei91, Man88, Man89b, Meh94, PPP94, VPDA93]. VLUGR2 [BTV96]. VLUGR3 [BV96]. VM [BCW20]. VMS [BMSD94]. VNM [DMPR93]. Void [BMP93]. Vol [CCKSS90, FL92]. Volcanic [VTTS98]. voltage [Ano02a, Ano02b, LMH90]. Volume [ALPP00, GL92, Kar94, Kau91, KCY93, SN89, TH94, Gri86, PSG03, ST90, Tru88, Wil90b, Wil91, WM92]. volumes [MBM^{+}20, Wil92a]. Volumic [Pet97]. Volunteer [FSC18, Kra20, Pin99]. Vortex [Ger90, MK93, Gre88b]. VOXAR [KG95]. VOXAR-All [KG95]. Voyages [Ano91w]. VP [AT93a, AT93b, DH86b, EHHS89, HL88a, LMM85b, LMM85a, SWL^{+}91, TKI85]. VP-100 [TKI85]. VP-100/200 [TKI85]. VP-200 [DH86b, HL88a, LMM85b, LMM85a]. VP-2600 [SWL^{+}91]. VP100 [MU83]. VP100/200 [MU83]. VP1200 [Man92]. VP200 [MHP84]. VPC [Gua87b]. VPP [Heg96, AHOK02]. VPP300 [Koe96, Koe97, Uch96, Uch97]. VPP500 [Ano94i, Ano94-61, BCHJ94, LO96, NKTT95, N^{+}95, SE98, UU93]. VPP700 [DTV00, Koe96, Koe97, Uch96, Uch97].

VPS [Gri86]. VPS-32 [Gri86]. VR

[Hel93, IS95, JLC98]. vs

[Ano94s, CCR11, CCKSS90, HS93a, Lee96, Mar92, Nag96c, ZMDS96]. vs. [HW96]. Vulcan [FG93]. VX [Koe96, Koe97, Uch96, Uch97]. VX/VPP300/VPP700 [Koe96, Koe97, Uch96, Uch97]. VXM [Ano93v]. Vying [Per02]. vysokoproizvoditelnoi [BKM88].

W [Kaz92, Tru88, YW94, Cha94b]. w-k [YW94]. W. [Pre93a]. W3C [Bar00c, Bar00d, Bar01]. WA [Ano91m]. Waals [WKHS97]. Wait [NG92]. Waiting [Ano93-32, MV16]. Wakamatsu [M^{+}95]. Wakeup [KL99]. walk [SM92]. Walking [Gil93]. Walkthrough [Kul94]. Wall [Ano94-128, DH93, YWXZ12, SW10b]. walnut [Bro17a]. Want [Rag06]. wants [Ano14]. War [CCKSS90, Ano97t]. warehouse [Jam95]. Warfare [EDJ^{+}10]. Waring’s [Wun89]. Warren [Ano92-46]. wars [Ano89e]. was [Eva97]. Washington [ACM92b, ACM92a, Ano96q, App96, FL92, Gra93c, IEE93b, IEE94e, Kho94, Soc94, USE00b, Cra91]. Watch [Ano93j, Ano95n, Ano95o]. Water [BK93, GI93, HW97, McB92a, McB92b, McB93, SKAT93, WF94, WHMA97, Ach99, Chi86, GH90, Gre90a, GH91, MW88]. Waters [AD97, DA97, ZL97, KB18]. Watershed [Mil97a]. Watson [Mob12]. Wave [Ano97n, BS98b, DGT94, Jen88, NCVG96, SE90, TKM96, WVBM88a, WVBM88b, BB87, CWLT97, NPS93, Vaf88]. Waveform [Ske87, WGS91, Xia88]. waveform-relaxation-Newton [Xia88]. waveforms [HA91]. Wavefront [Kun84]. waveguide [GKR91]. Wavelet [BB93, Sul91]. Wavelets [SRBL94]. Waves [HFNP96, Max81, RHH96, VD96, VTTS98]. Way [Ano94p, Ano95p, Gie96, Ano93d, CSFS00, DMW87, HHS01b, New14, Way96, Fos96]. ways [Sha95b]. WDM [DG95, KGB^{+}96]. WDM-Based [KGB^{+}96]. Weaknesses [BM96]. weapons [Ano97e, Cla97, Kra97]. Wear [KB97]. Weather [Ale90, AGKT02, BCHH94, Bar00a, Bar00b, Che90e, DGT84, Dic81, Dic82, GJS94, Hof94, JC94d, SBW^{+}19, Sel95, VW95, WS84d, WCG94, Ano97j, Bur91, Che90d, Che96, Sha95b, Tay95a]. web [Hae91, ACKW01, Ano95-32, Ano96h, CLB19, Coc01, Coc02c, Coc02d, CBM^{+}05, FT96b, Ser98, DDJ98b, VSW94, Zhe97]. Web-Based [FT96b, Ano96h]. Weight [HTI93, LHLM95, Ano94-128, AW91]. Weighted [AM94, TD96, PW23]. welcomes [Str94]. weld [WH94]. Welding [Nor97b, SZ96, SJD96]. weldments [WH94]. Well [Ano96-38, BM93a, RDZ93, Ano88y, DMW87]. Well-Posed [BM93a]. Weltrekordrechner [Ano01b]. Wenes [MM94a]. Wesley [Sch88a]. West [KWW92, RD94, SEA84]. Westin [IEE95b]. Wet [Gil93]. WHAMS3D [CGS91, CGL92, GCP90]. Wheel [Gru97]. wheelchair [Str12]. Where [GP06, BFS11]. Which [Ruh95, Sha95b]. While [Col94, KP96, Ano92-42, Ano96-28]. While-Loops [Col94]. Whirl [GA97]. Whistleblower [Ano96x]. Whitewater [Ano93a]. Who [Ano90q, Bro17a, Sch89a, Str03, Mur10]. Whole [Ano94-118, GGC^{+}11, NM93]. Whoop [RCK97]. Who’s [Str03]. WI [Ano93a]. Wicked [Luc09]. Wide [Ano95-32, BGS^{+}12, BGH^{+}02, Kah91, VSW94, Zhe97, Con86, Ano92-47]. Wide-Area [BGH^{+}02]. Widely [KG96]. Will [And20, Ano95l, Art93, LQFC18, SGIS93, Sul97, Ano89i, Ano92-44, Ano96-28, Ano96z, Ano97u, Ano97-27, Bab94, Bel92a, Hsu15, Mar90, Mer01, Wat92b]. William [Nor03, Pin01]. WIMS [DLG93]. Win [Ano97z, Smi95, Ano97j]. Wind [Ano94-32, Ano94-68, GI93, JC94d, KBD97, BB87, Boy15]. wind-wave [BB87]. Wind/Water [GI93]. Window [NdMM09, Mye92b, Y^{+}92]. Windowed [OB94]. Windows [USE00b, NSB96, Ano95-32, Ano01c, Chi00, Ste01c]. Winds [Coc02c, Coc02d]. Windshield [Jan96]. Wing [MFK94]. Wings [Riz94]. Winner [Gui08, Pin01]. Winners [Ano87d, KHHS95, Ano97b]. Winning [CCKSS90, Sch89a]. Wins [Ano11a, War09, Ano97b]. Winter [JT87]. Wire [EBS02, KMNT95, KMNT96, HY92, Pan96]. wire-limited [HY92]. wireles [Ano95e, Ano95f]. wireless [Ano96p, Ano00b]. Wires [TSCG94]. wiring [HY91, yHY92]. within [CLPV93, KKPR93, MGS^{+}20]. without [Ano90o, Fos03]. Wizards [Mur97, Nor97a]. Woes [Ano96x]. Wolfe [MP91a]. Wolfe/gradient [MP91b]. Women [Pin01]. Wonderful [Mur10]. Won’t [Ano93-30]. wood [Ano91m]. Word [Cra91, Pev93]. Work [PD94, RS94c, RDZ93, Ano95c, Das94, HHS01b, Per87, Win02]. Work-Load [RS94c]. worker [GA12]. workflows [BRL^{+}20]. Workforce [Lat16]. Working [Rep92, RG94, Ano90t, YH90]. Workload [Ano94-37, BD94, SWG06, VSB94, Bra91a, LF03, PBK91]. Workloads [Ano94e, Wil88a, WMK90]. Works [Bar01]. Workshop [Ano90t, Ano92a, Ano93c, Ano94-100, Ano94-126, Ano95u, Ano96-43, Ano01a, B^{+}89, DT97, Fox97, Fra94, GL92, GLH94, HWP95, HK93b, HKS93, KH93, Kow89b, KK92, LP90, Lag89, Men87, Nat84, OMM93, Por86, WSB96, Ano96w, Ano97-30, COS89, D^{+}95, Elm95a, Har90, Har91, Hen97, MM91a, ML95b, Pra95, USE90, Vag88, Wun89, FJSP95, HK93a, MA85]. Workshops [Dra94a, Dra94b, Dra96b, Dra96a, Edw97, FJSP95]. Workstation [Ano94-138, ASNT91, DXJM93, GBF93, GL89, HMS93, KMKD97, KR94c, LC97a, PBM95, SF93a, Web93, Ano94-82, Fie86, tHd90, KC89, KG95, Mac97a, SNS^{+}97, Tho93b]. Workstations [Ano94-78, Ano97m, BIRB93, Bra93, HW96, LMP^{+}90, Liu95, MM94a, NGLPJ96, NBKP95a, Wen94, Cap96, Che96, DB95, FT93b, Hil97, Lev89, Mac96c, NBKP95b, PZGL91, SD92, Tri95a, Tri95b]. World [Ano88w, Ano90g, ACA94, Bar00c, Bar00d, Che90f, Gep01, IS95, KK89a, Mac90, Mur10, Tho96a, Tho96b, TW92, VAS82, Ano90n, Ano93b, Ano96l, Ano97u, Ano97s, Ano00b, DJM94, DZM^{+}13, Gra94, LM92, Per93a, Pic92, Ano95-32, VSW94, Zhe97]. Wormhole [Ano94-53, Ano94-88, CB94, PDR94, RE94, TM94a, TM94b, WK95]. Wormhole-Routed [Ano94-88, PDR94, TM94a, TM94b, WK95]. Worth [Mul96]. Would [DCG93, DCGxx, Ber82, Poo96a, SF82]. write [CV91a, CV92b, MWRK18]. write-through [CV91a]. WSDL [Bar01]. Wunsch [AFF93]. Wurttemberg [Sch94b].

X [ALM93, Ano85b, ABHS89b, ABHS89a, Ber90a, BH92a, Cal85a, Cal85b, CDMW94, CM84, Cha84, CM86, CDH84, Che89b, CS84, CS86a, CK90, Dao88, DO89, DP90, DH91a, DH91b, DH86b, DH86a, EE93, EY91, FSY88, GKL^{+}87, GS89d, GZA86, Gur88, Ho91, Hoc85, HKN89, HES93, HFH86, HFH87, KN88, Kra88, KM85, Lar84, LMM85b, LMM85a, LMM86, MLR90b, MLR90a, Meu87, MKB87, MF93, Nag88, NR86, OL86, OD88, PB94a, Par90c, PBK91, Rei85, RS85, RRSS93, Rit88b, Rit88a, RR89, SW91, Sea86, SSLR90, Svo93, hTD88, Tem89a, Tem89b, VSH90, VM87, Vol89, VY88, WHBH93, Wes89, WB85, Wil88a, WMK90, Y^{+}92, ZM86]. X-IMAGE [RRSS93]. X-MP [ABHS89b, ABHS89a, Ber90a, BH92a, Cal85a, Cal85b, CM84, Cha84, CM86, Che89b, CS84, CS86a, CK90, DO89, DP90, DH91a, DH91b, EE93, EY91, FSY88, GKL^{+}87, GS89d, GZA86, Gur88, Ho91, Hoc85, HKN89, HES93, HFH86, HFH87, KN88, Kra88, KM85, LMM85b, LMM85a, LMM86, MLR90b, MLR90a, Meu87, MKB87, Nag88, NR86, OL86, OD88, Par90c, PBK91, Rei85, RS85, Rit88a, RR89, SW91, Sea86, SSLR90, Svo93, hTD88, Tem89a, Tem89b, VSH90, VM87, VY88, WHBH93, Wes89, Wil88a, WMK90, ZM86]. X-MP-2 [CDH84, Lar84]. X-MP-4 [DH86b, DH86a]. X-MP-like [WB85]. X-MP/2 [Cha84, LMM85b, LMM85a]. X-MP/24 [GKL^{+}87, LMM86]. X-MP/416 [VY88]. X-MP/48 [Ber90a, CK90, HFH86, HFH87, Meu87, Nag88, VM87]. X-MP/Model [RR89]. X-Ray [CDMW94, PB94a]. X-Window [Y^{+}92]. X-Y-Z [MF93]. X/MP [Dao88, HL88a, Rit88b]. X/OPEN [Ano85b]. X1 [DVWW05]. XC [CKD^{+}19, DAC^{+}18, KMRR20, MWRK18, MGS^{+}20]. XC-40 [DAC^{+}18]. XC40 [Cla18, HPH^{+}20, HLA^{+}18, HCD^{+}18, SSSR20]. Xcount [SM89]. XE6 [KBVH14, KB18]. XE6/XK7 [KB18]. XK7 [KB18]. XLink [Bar00c, Bar00d]. XML [Poz13]. XMT [BB13, BÇG14, VTSM12]. XMT-2 [BB13, BÇG14]. XP [Ano94h, Bem92, Gro93, Int91, LMM86, SNS^{+}97]. XP-200 [LMM86]. XP/S [Ano94h, Bem92, Gro93]. XP/S-15 [Ano94h]. xQard [SSS94]. XT [YQTV12]. XT4 [DBK09]. Xylem [EM91].

Y-Geometry [ALM93]. Y-MP [Cra92, MSTK93, Oed92a, Oed92b, And90a, Ano88l, BOS93, Bow88, BL91, CRM94, CS93b, CS95, Dan91, DH91a, DH91b, Dic90, Din92, GP93a, HVY91, Hol90b, HKSY94, HSKY95, LS92b, LS93a, MSW91, NSH95, Nag90, Pap92, PS94a, Pin91, QB92, Rei88, SPS90, SWL^{+}91, SS90c, VSH91, Vaj91, van95b]. Y-MP/2 [NSH95]. Y-MP8 [Cho90a, SO91]. Y-MP8/864 [Cho90a, SO91]. Y/MP [Sar91]. Y2K [Gar99]. Yao [War03]. Yardley [CCKSS90]. Year [Ano90u, Ano92f, Zyg93, Ano95v, Mes93a, Stu03]. Years [CCKSS90, Fin94, Ano88y, MSCxx, Zen99]. Yesterday [Hag01]. YH [HCL94, YJD93]. Yield [RL90b, Don93c]. Yields [Ano88m, Ano95w]. YMP [Car91, HP95]. York [Ano97q, Ano00a, IEE90]. Yourself [HHS01a, HHS01b, JA92a, JA92b].

Z [HA91, War03, MF93]. Zentrum [Stu95]. Zeolites [CFV^{+}90]. Zero [SA10b]. Zeroclusters [SA10b]. ZeroOne [Ano85b]. Zeuthen [FWS96]. zip [Ano90o]. zone [NSB96, WH94, Yi11]. ZPL [DLMW95, Sny99]. ZS [MSAD91]. ZS-1 [MSAD91]. Zurich [Ano93c, HKR94, Mar86]. Zuse [Stu95]. Zycad [TS90].

References

Allcock:2002:DMT

[A^{+}02] Bill Allcock et al. Data management and transfer in high-performance computational grid environments. Parallel Computing, 28(5): 749–771, May 2002. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
35/21/60/57/31/abstract.
html.

Ahmad:2012:HEA

[A^{+}12] Ishfaq Ahmad et al. Handbook of energy-aware and green computing. Chapman and Hall/CRC computer and information science series. CRC Press, 2000 N.W.

Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4665-0116-2 (set), 1-4665-0112-X (vol. 2), 1-4398-5040-2 (vol. 1). xvii + 1196 + 27 (two volumes) pp. LCCN ????

Aybar:1993:SDO

[AA93] H. S. Aybar and T. Aldemir. Simulator development for the Ohio State University inherently safe reactor using the DSNP language. In Kusters et al. [KSW93], pages 831–841. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Aluru:2006:ESS

[AAB06] Srinivas Aluru, Nancy M. Amato, and David A. Bader. Editorial: Special section on high-performance computational biology. IEEE Transactions on Parallel and Distributed Systems, 17(8):737– 739, August 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
td/2006/08/l0737.pdf.

Abrugia:1993:USA

[AABB93] G. Abrugia, U. Amato, P. Belardini, and C. Bertoli. An unconditionally stable algorithm for the simulation of air motion inside engines. In Anonymous [Ano93-31], pages 795–806. ISBN 0-947719-62-8. LCCN ????

An:1995:CFI

[AABK95] M. An, N. Anupindi, M. Bletsas, and G. Kechriotis. Comparison of 2-D FFT implementations on the Intel Paragon massively parallel supercomputer. In IEEE [IEE95b], pages 2755–2758. ISBN 0-7803-2432-3, 0-7803-2431-5, 0-7803-2433-1, 0-7803-2562-1. ISSN 0749-8411. LCCN TK 7882 S65 I16 1995. Five volumes.

Almasi:2005:DIM

[AAC^{+}05] G. Almási, C. Archer, J. G.

Castaños, J. A. Gunnels, C. C. Erway, P. Heidelberger, X. Martorell, J. E. Moreira, K. Pinnow, J. Ratterman, B. D. Steinmacher-Burow, W. Gropp, and B. Toonen. Design and implementation of message-passing services for the Blue Gene/L supercomputer. IBM Journal of Research and Development, 49(2/3):393–406, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/almasi.pdf.

Alverson:1992:EHP

[AACK92] G. A. Alverson, R. Alverson, D. Callahan, and B. Koblenz. Exploiting heterogeneous parallelism on a multi-threaded multiprocessor. In ACM [ACM92b], pages 188–197. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Ariel:1988:SMP

[AAS88] P. D. Ariel, B. D. Aggarwala, and M. Sezgin. Simulation of MHD power generator flow using a supercomputer. Zeitschrift für Angewandte Mathematik und Mechanik, 68(10):503–??, 1988. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Ames:1994:FSI

[AB94] Karyn R. Ames and Alan Brenner, editors. Frontiers of supercomputing II: a national reassessment: Papers from the 2nd Frontiers of Supercomputing Conference held at Los Alamos National Laboratory, August 20–24, 1990, volume 12 of Los Alamos series in basic and applied sciences 12. University of California Press, Berkeley, CA, USA, 1994. ISBN 0-520-08401-2. LCCN QA76.88 .F76 1994. Papers from the 2nd Frontiers of Supercomputing Conference held at Los Alamos National Laboratory, 8/20-24/90.

Arnoldi:1995:NRS

[AB95] H.-M. R. Arnoldi and W. Brauer. Neuronal response synchronization. In Herrmann et al. [HWP95], pages 397–406. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Appelbe:1996:STH

[AB96] Bill Appelbe and Donna Bergmark. Software tools for high-performance computing: survey and recommendations. Scientific Programming, 5(3):239–249, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Afsarmanesh:2001:GEH

[AB01] Hamideh Afsarmanesh and Marian Bubak. Guest editorial: High-performance computing and networking enables complex applications. Future Generation Computer Systems, 17(8):v– vi, June 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/25/abstract.
html.

Aluru:2003:GEI

[AB03] Srinivas Aluru and David A.

Bader. Guest Editor’s introduction: Special issue on high-performance computational biology. Journal of Parallel and Distributed Computing, 63(7–8):671–673, July/August 2003. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Abawajy:2009:EAS

[Aba09] J. H. Abawajy. An efficient adaptive scheduling policy for high-performance computing. Future Generation Computer Systems, 25(3):364–370, March 2009. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Almasi:2003:OBS

[ABB^{+}03] George Almási, Ralph Bellofatto, José Brunheroto, Călin Caşcaval, José G. Castaños, Paul Crumley, C. Christopher Erway, Derek Lieber, Xavier Martorell, José E. Moreira, Ramendra Sahoo, Alda Sanomiya, Luis Ceze, and Karin Strauss. An overview of the BlueGene/L system software organization. Parallel Processing Letters, 13(4):561–??, December 2003. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Alam:2013:EES

[ABB^{+}13] S. Alam, C. Bekas, H. Boettiger, A. Curioni, G. Fourestey, W. Homberg, M. Knobloch, T. Laino, T. Maurer, B. Mohr, D. Pleiter, A. Schiller, T. Schulthess, and V. Weber. Early experiences with scientific applications on the IBM Blue Gene/Q supercomputer. IBM Journal of Research and Development, 57(1/2):14:1– 14:9, January–March 2013. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Ambrosiano:1994:HCS

[ABBB94] J. J. Ambrosiano, J. Bolstad, A. J. Bourgeois, and J. C. Brown. High-performance climate system modeling using a domain and task decomposition message-passing approach. In IEEE [IEE94c], pages 397–405. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Adiga:2005:BGT

[ABC^{+}05] N. R. Adiga, M. A. Blumrich, D. Chen, P. Coteus, A. Gara, M. E. Giampapa, P. Heidelberger, S. Singh, B. D. Steinmacher-Burow, T. Takken, M. Tsao, and P. Vranas. Blue Gene/L torus interconnection network. IBM Journal of Research and Development, 49 (2/3):265–276, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/adiga.pdf.

Ambrosiano:1997:EDS

[ABCE97] J. Ambrosiano, R. Balay, C. Coats, and A. Eyth. The environmental decision support system: Air quality modeling and beyond. In Delic and Wheeler [DW97], pages 47–58. ISBN 0-89871-378-1. LCCN ????

Almgren:1997:HRA

[ABCH97] A. S. Almgren, J. B. Bell, P. Colella, and L. H. Howell. A high resolution adaptive projection method for regional atmospheric modeling. In Delic and Wheeler [DW97], pages 69–80. ISBN 0-89871-378-1. LCCN ????

Arnold:1992:SI

[ABD92] Jeffrey M. Arnold, Duncan A. Buell, and Elaine G. Davis. SPLASH II. Technical report SRC-TR-92-061, Supercomputing Research Center: IDA, Lanham, MD, USA, March 18, 1992. 11 pp.

Abelson:1990:STA

[Abe90] Harold Abelson. The Supercomputer Toolkit and its applications. A.I. memo 1249, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, 1990. 15 pp.

Abelson:1991:STG

[Abe91] Harold Abelson. The Supercomputer Toolkit: a general framework for specialpurpose computing. A.I. memo 1329, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, November 1991. 41 pp.

Arbenz:1996:MDS

[ABGL96] P. Arbenz, M. Billeter, P. Guentert, and P. Luginbuehl. Molecular dynamics simulations on Cray clusters using the SCIDDLE-PVM environment. Lecture Notes in Computer Science, 1156:142–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

August:1989:CXM

[ABHS89a] M. C. August, G. M. Brost, C. C. Hsiung, and A. J. Schiffleger. Cray X-MP: The birth of a supercomputer. Computer, 22(1):45– 52, January 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

August:1989:CXB

[ABHS89b] Melvin C. August, Gerald M.

Brost, Christopher C. Hsiung, and Alan J. Schiffleger. Cray X-MP: The birth of a supercomputer. Computer, 22(1):45–52, January 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Apgar:1988:DSS

[ABM88] Brian Apgar, Bret Bersack, and Abraham Mammen. A display system for the Stellar graphics supercomputer model GS1000. Computer Graphics, 22(4):255–262, August 1988. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        proceedings/graph/54852/
                                        p255-apgar/.

Austin:2004:MS

[ABM^{+}04] Todd Austin, David Blaauw, Scott Mahlke, Trevor Mudge, Chaitali Chakrabarti, and Wayne Wolf. Mobile supercomputers. Computer, 37 (5):81–??, May 2004. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        co/2004/05/r5081abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/co/2004/05/r5081.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/co/2004/05/
                                        r5081.pdf.

Aumage:2002:MIP

[ABMN02] Olivier Aumage, Luc Bougé, Jean-François Méhaut, and Raymond Namyst. Madeleine II: a portable and efficient communication library for high-performance cluster computing. Parallel Computing, 28(4):607–626, April 2002. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        35/21/60/39/29/abstract.
                                        html.

Asanovic:1993:DCNb

[ABMW93] K. Asanovic, J. Beck, N. Morgan, and J. Wawrzynek. Designing a connectionist network supercomputer. International Journal of Neural Systems, 4(4):317–326, December 1993. CODEN IJSZEG. ISSN 0129-0657.

Andrews:1992:BFN

[ABP92] John B. Andrews, Carl J.

Beckmann, and David K. Poulsen. Broadcasting/ forwarding networks for fast synchronization. Technical Report CSRD 1151, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. 32 pp.

Abraham:1988:RIC

[Abr88] Santosh G. Abraham. Reducing interprocessor communication in parallel architectures: system configuration and task assignment. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1988. x + 215 pp.

Abraham:1990:IAD

[Abr90] Seth Abraham. Issues in the architecture of direct interconnection schemes for multiprocessors. Technical report, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1990. ix + 202 pp.

Abramson:1992:VHS

[Abr92] David Abramson. A very high speed architecture for simulated annealing. Computer, 25(5):27–36, May 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abramson:1994:PPS

[Abr94] D. Abramson. Predicting the performance of scientific applications on distributed memory multiprocessors. In IEEE [IEE94c], pages 285– 292. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Anupam:1994:DCV

[ABSS94] Vinod Anupam, Chandrajit Bajaj, Daniel Schikore, and Matthew Schikore. Distributed and collaborative visualization. Computer, 27 (7):37–43, July 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ariat:1984:IEA

[AC84a] J. Ariat and W. C. Carter. Implementation and evaluation of a (b, k)-adjacent errorcorrecting/detecting scheme for superconductor systems. IBM Journal of Research and Development, 28(2):159–169, March 1984. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Ariat:1984:IEB

[AC84b] J. Ariat and W. C. Carter.

Implementation and evaluation of a (B,K)-adjacent Error-Correcting/Detecting scheme for supercomputer systems. IBM Journal of Research and Development, 28(2):159–169, March 1984. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Anastasio:1991:OCL

[AC91] Thomas A. Anastasio and William W. Carlson. An observation on the C library procedure random(). Technical report SRC-TR-91-044, Supercomputing Research Center: IDA, Lanham, MD, USA, September 25, 1991. 6 pp.

Acevedo:1993:RTD

[AC93] M. Acevedo and J. T.

Celigueta. Real time dynamic simulation of passenger cars. In Anonymous [Ano93-31], pages 559–566. ISBN 0-947719-62-8. LCCN ????

Arvind:1994:SNG

[ACA94] Arvind, D. Chiou, and Boon Seong Ang. 0*T (Star T) the next generation: In the real world. In Balakrishnan [Bal94], pages 400–406. ISBN 0-07-462044-4. LCCN ????

Arsenin:1996:STS

[ACC^{+}96] I. Arsenin, D. Chen, N. Christ, R. Edwards, A. Gara, S. Hanson, C. Jung, A. Kahler, A. D. Kennedy, G. Kilcup, Y. Luo, C. Malureanu, R. Mawhinney, J. Parsons, J. Sexton, C. Sui, and P. Vranas. Status of the 0.8 teraflops supercomputer at Columbia. In Kieu et al. [KMG96], pages 804– 807. ISBN ???? ISSN 0920-5632 (print), 1873-3832 (electronic). LCCN QC793.3.F5 I569 1995.

Anderson:1990:TTD

[ACG^{+}90] D. V. Anderson, W. A. Cooper, R. Gruber, S. Merazzi, and U. Schwenn. TERPSICHORE: a threedimensional ideal magnetohydrodynamic stability program. In Pitcher [Pit90], pages 47–56. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Achdou:1999:BRN

[Ach99] Yves Achdou. Book reviews: Numerical methods for the three dimensional shallow water equations on supercomputers, by E. D. de Goede. Mathematics of Computation, 68 (225):??, January 1999. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.
ams.org/jourcgi/jour-pbprocess?
fn=110&arg1=S0025-5718-
99-00992-8&u=/mcom/1999-
68-225/.

Arpaci:1995:EEC

[ACK^{+}95] Remzi H. Arpaci, David E.

Culler, Arvind Krishnamurthy, Steve G. Steinberg, and Katherine Yelick. Empirical evaluation of the CRAY-T3D: a compiler perspective. ACM SIGARCH Computer Architecture News, 23(2): 320–331, May 1995. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Aloisio:2001:WAS

[ACKW01] Giovanni Aloisio, Massimo Cafaro, Carl Kesselman, and Roy Williams. Web access to supercomputing. Computing in Science and Engineering, 3(6):66–72, November/December 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c6066abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c6066.pdf.

Arzt:1993:TTS

[ACL93] M. Arzt, J.-Y. Cognard, and P. Ladeveze. A two-time scale approach for viscoplastic computations under cyclic loadings suited to parallel computers. In Anonymous [Ano93-31], pages 693–700.

ISBN 0-947719-62-8. LCCN ????

ACM:1988:ICS

[ACM88] ACM, editor. 1988 International Conference on Supercomputing: July 4–8, 1988, St. Malo, France: conference proceeedings. ACM Press, New York, NY 10036, USA, 1988. ISBN 0-89791-272-1. LCCN QA76.5 .I547 1988. ACM order number: 415881.

ACM:1989:PSN

[ACM89a] ACM, editor. Proceedings, Supercomputing ’89 November 13–17, 1989, Reno, Nevada. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989.

ACM:1989:SVR

[ACM89b] ACM. SIGGRAPH video review: Computer graphics. supplement and visualization in scientific computing — supercomputers, 1989. 1 videocassette.

ACM:1990:CPI

[ACM90] ACM, editor. Conference proceedings, 1990 International Conference on Supercomputing: June 11– 15, 1990, Amsterdam, the Netherlands, volume 18(3) of Computer architecture news. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-369-8. LCCN QA1 .A85457. ACM order number: 415902.

ACM:1991:CPI

[ACM91] ACM, editor. Conference proceedings, 1991 International Conference on Supercomputing: June 17–21, 1991, Cologne, Germany. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-434-1. LCCN ????

ACM:1992:ICS

[ACM92a] ACM, editor. 1992 International Conference on Supercomputing: July 19–23, 1992, Washington, DC: conference proceeedings. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. ACM order number: 415921.

ACM:1992:CPI

[ACM92b] ACM, editor. Conference proceedings / 1992 International Conference on Supercomputing, July 19–23, 1992, Washington, DC. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

ACM:1993:ICS

[ACM93] ACM, editor. 1993 International Conference on Supercomputing: July 20–22, 1993, Tokyo, Japan: conference proceedings. ACM Press, New York, NY 10036, USA, 1993. ACM order number: 415931.

ACM:1994:CPI

[ACM94] ACM, editor. Conference proceedings: 1994 International Conference on Supercomputing: July 11–15, 1994, Manchester, England, United Kingdom. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-665-4. LCCN ????

ACM:1995:CPI

[ACM95a] ACM, editor. Conference proceedings of the 1995 International Conference on Supercomputing, Barcelona, Spain, July 3–7, 1995, CON-FERENCE PROCEEDINGS OF THE INTERNATIONAL CONFERENCE ON SU-PERCOMPUTING 1995; 9th. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

ACM:1995:PAI

[ACM95b] ACM, editor. Proceedings / the 22nd annual International Symposium on Computer Architecture, June 22– 24, 1995, Santa Margherita Ligure, Italy, volume 23(2) of Computer Architecture News. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-698-0. ISSN 0163-5964 (ACM), 0884-7495 (IEEE). LCCN QA 76.9 A73 I56 1995.

ACM:1995:SPA

[ACM95c] ACM, editor. Supercomputing ’95 proceedings of the 1995 ACM/IEEE Supercomputing Conference: December 3–8, 1995, San Diego, California. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-816-9. LCCN ???? CD-ROM for Windows, Macintosh, or UNIX: ACM order no. 415952. IEEE Computer Society Press order #FW07435.

ACM:1996:FCP

[ACM96] ACM, editor. FCRC ’96:

Conference proceedings of the 1996 International Conference on Supercomputing: Philadelphia, Pennsylvania, USA, May 25–28, 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

ACM:1997:CPI

[ACM97] ACM, editor. Conference proceedings of the 1997 International Conference on Supercomputing: Vienna, Austria, July 7-11, 1997. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-902-5. LCCN ???? ACM order number: 415971.

ACM:19xx:PLA

[ACMxx] ACM. Proceedings of the … ACM/IEEE supercomputing conference, 19xx. On CD-ROM.

ACM:2003:CPI

[ACM03] ACM, editor. Conference proceedings of the 2003 International Conference on Supercomputing: June 23–26, 2003, San Francisco, California, USA. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-733-8. LCCN QA76.5 .I547 2003. ACM order number 415031.

Anderson:1990:OGP

[ACSH90] D. V. Anderson, B. C. Curtis, D. E. Shumaker, and E. J. Horowitz. Obtaining gigaflop performance from particle simulation of plasmas. In Pitcher [Pit90], pages 297– 308. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Abraham:1988:BPS

[AD88] Santosh G. Abraham and Timothy Alden Davis. Blocking for parallel sparse linear system solvers. Technical Report CSRD 742, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 25 + [1] pp.

Abdelrhman:1997:TMC

[AD97] M. A. Abdelrhman and E. H.

Dettmann. Two-dimensional modeling of current circulation and contaminant transport in surface waters. In Delic and Wheeler [DW97], pages 117–124. ISBN 0-89871-378-1. LCCN ????

Adams:1993:NNM

[Ada93] M. L. Adams. New nonlinear methods for linear transport calculations. In Kusters et al. [KSW93], pages 683– 694. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Adam:1995:ME

[Ada95a] J. A. Adam. Medical electronics. IEEE Spectrum, 32 (1):80–83, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adam:1995:HDD

[Ada95b] John A. Adam. Humans do digital; genes ride up; heart beats on a supercomputer. IEEE Spectrum, 32(1): 80–83, January 1, 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adeli:1992:SEA

[Ade92] Hojjat Adeli. Supercomputing in engineering analysis, volume 1 of New generation computing; 1. M. Dekker, New York, NY, USA, 1992.

ISBN 0-8247-8559-2. xi + 362 pp. LCCN TA345 .S87 1992.

Adee:2010:CBF

[Ade10] S. Adee. Cat-brain fever. IEEE Spectrum, 47(1):16– 17, January 2010. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aridor:2005:RAU

[ADG^{+}05] Y. Aridor, T. Domany, O. Goldshmidt, J. E. Moreira, and E. Shmueli. Resource allocation and utilization in the Blue Gene/L supercomputer. IBM Journal of Research and Development, 49(2/3):425–436, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/aridor.pdf.

Aridor:2008:MIT

[ADG^{+}08] Yariv Aridor, Tamar Domany, Oleg Goldshmidt, Yevgeny Kliteynik, Edi Shmueli, and Jose E. Moreira. Multitoroidal interconnects for tightly coupled supercomputers. IEEE Transactions on Parallel and Distributed Systems, 19(1):52– 65, January 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
td/2008/01/ttd2008010052s.
zip.

Aertsen:1995:CDC

[ADGA95] A. Aertsen, M. Diesmann, S. Gruen, and M. Arndt. Coupling dynamics and coincident spiking in cortical neural networks. In Herrmann et al. [HWP95], pages 213– 224. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Amestoy:2001:ACT

[ADLL01] Patrick R. Amestoy, Iain S.

Duff, Jean-Yves L’Excellent, and Xiaoye S. Li. Analysis and comparison of two general sparse solvers for distributed memory computers. ACM Transactions on Mathematical Software, 27(4):388– 421, December 2001. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Anderson:1997:PIL

[AF97] E. Anderson and M. Fahey. Performance improvements to LAPACK for the Cray Scientific Library. LA-PACK Working Note 126, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April

1997. URL http://www.
netlib.org/lapack/lawns/
lawn126.ps;   http://www.
netlib.org/lapack/lawnspdf/
lawn126.pdf. UT-CS-97-

359, April 1997.

Al-Furaih:1996:PCM

[AFAGR96] I. Al-Furaih, S. Aluru, S. Goil, and S. Ranka. Parallel construction of multidimensional binary search trees. In ACM [ACM96], pages 205–212. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Anderson:1993:PAN

[AFF93] D. J. Anderson, J. H. Fry, and E. D. Fleischmann. Pocket arithmetic and Needleman-Wunsch on Cray Research computers. In Lim et al. [L^{+}93], pages 435– 444. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Akherraz:1993:AST

[AFML93] A. Akherraz, C. Fedon-Magnaud, and J. J. Lautard. An anisotropic scattering treatment for the even parity transport equation. In Kusters et al. [KSW93], pages 467–475. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Ala:1996:IMM

[AFT96] G. Ala, E. Francomano, and A. Tortorici. Iterative moment method for electromagnetic transients in grounding systems on CRAY T3D. Lecture Notes in Computer Science, 1041:9–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ashby:1997:SAM

[AFT97] S. F. Ashby, R. D. Falgout, and A. F. B. Tompson. A scalable approach to modeling groundwater flow on massively parallel computers. In Delic and Wheeler [DW97], pages 201–216. ISBN 0-89871-378-1. LCCN ????

Afuah:1990:RMP

[Afu90] Allan Nembo Afuah. The rise of massively parallel computers in the supercomputer industry. Thesis (M.S.), Massachusetts Institute of Technology, Sloan School of Management, Cambridge, MA, USA, 1990. 82 pp. Supervised by James M. Utterback.

Aslam:1990:ASD

[AG90] Sohail Aslam and E. J. (Efstratios J.) Gallopoulos. The advanced software development and commercialization project: progress report PR-1. Technical Report CSRD 1047, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1990. 49 pp.

Anand:1994:EPS

[AG94] T. Anand and P. Gupta. An efficient parallel sorting algorithm for X+Y. In Balakrishnan [Bal94], pages 284–289. ISBN 0-07-462044-4. LCCN ????

Arrigo:1993:ISD

[AGD93] P. Arrigo, F. Giuliano, and G. Damiani. Identification of singular domains on cDNA sequences by Kohonen’s features maps. In Lim et al. [L^{+}93], pages 479–486. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Agerwala:2005:MVP

[Age05] Tilak Agerwala. Message from the Vice President, Systems, IBM Research Division. IBM Journal of Research and Development, 49(2/3):??, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/message.
pdf.

Abuzaghleh:2013:IAH

[AGEL13] Omar Abuzaghleh, Kathleen Goldschmidt, Yasser Elleithy, and Jeongkyu Lee. Implementing an affordable highperformance computing for teaching-oriented computer science curriculum. ACM Transactions on Computing Education, 13(1):3:1–3:??, January 2013. CODEN ???? ISSN 1946-6226.

Adamson:1990:SCF

[AGH^{+}90] B. N. Adamson, A. D. Gosman, C. Hill, C. J. Marooney, B. Nasseri, M. Sarantinos, T. Theodoropoulos, and H. G. Weller. Simulation of in-cylinder flow and combustion in reciprocating engines. In Pitcher [Pit90], pages 253– 263. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Anderson:1987:IOL

[AGK^{+}87] J. Wayne Anderson, William F.

Galway, Robert R. Kessler, Herbert Melenk, and Winfried Neun. Implementing and optimizing Lisp for the Cray. IEEE Software, 4 (4):74–83, July 1987. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Alpert:2002:NWP

[AGKT02] Pinhas Alpert, Alexander Goikhman, Jacob Katzenelson, and Marina Tsidulko. Numerical weather prediction on the supercomputer toolkit. Lecture Notes in Computer Science, 2327:335– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2327/23270335.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270335.
pdf.

Allen:1999:SEE

[AGL^{+}99] Gabrielle Allen, Tom Goodale, Gerd Lanfermann, Thomas Radke, Edward Seidel, Werner Benger, Hans-Christian Hege, Andre Merzky, Joan Massó, and John Shalf. Solving Einstein’s equations on supercomputers. Computer, 32 (12):52–58, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //dlib.computer.org/co/
                                        books/co1999/pdf/rz052.
                                        pdf;   http://www.computer.
                                        org/computer/co1999/rz052abs.
                                        htm.

Attig:2011:TSE

[AGL11] N. Attig, P. Gibbon, and Th. Lippert. Trends in supercomputing: The European path to exascale. Computer Physics Communications, 182(9):2041–2046, September 2011. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465510004571.

Attig:1998:RCL

[AGLL98] N. Attig, S. Guesken, P. Lacock, and T. Lippert. Running a code for lattice quantum chromodynamics efficiently on CRAY T3E systems. Lecture Notes in Computer Science, 1401:183–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferov:1996:OIP

[AGP96] Z. I. Alferov, Y. V. Gulyaev, and D. R. Pape, editors. Optical information processing: International conference; 2nd — June 1996, St. Petersburg, Russia, number 2969 in PROCEEDINGS- SPIE THE INTERNATIONAL SOCI-ETY FOR OPTICAL EN-GINEERING 1996. SPIE Optical Engineering Press, Bellingham, WA, USA, 1996. ISBN 0-8194-2375-0. ISSN 0361-0748. LCCN ????

Aggarwal:2011:SMP

[AGY^{+}11] Vikas Aggarwal, Alan D.

George, Changil Yoon, Kishore Yalamanchili, and Herman Lam. SHMEM+: a multilevel-PGAS programming model for reconfigurable supercomputing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 4(3):26:1– 26:??, August 2011. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Agarwal:1994:EPA

[AGZ94a] R. C. Agarwal, F. G. Gustavson, and M. Zubair. An efficient parallel algorithm for the 3-D FFT NAS parallel benchmark. In IEEE [IEE94c], pages 129–133. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Agarwal:1994:HMA

[AGZ94b] R. C. Agarwal, F. G. Gustavson, and M. Zubair.

A high-performance matrixmultiplication algorithm on a distributed-memory parallel computer, using overlapped communication. IBM Journal of Research and Development, 38(6):673–681, November 1994. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL

                                        http://www.almaden.ibm.
                                        com/journal/rd38-6.html#
                                        three.

Ammarguellat:1990:ARI

[AH90] Zahira Ammarguellat and Luddy Harrison. Automatic recognition of induction variables and recurrence relations by abstract interpretation. Technical Report CSRD 980, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1990. 13 pp.

Andres:1993:UIF

[AH93] T. H. Andres and W. C. Hajas. Using iterated fractional factorial design to screen parameters in sensitivity analysis of a probabilistic risk assessment model. In Kusters et al. [KSW93], pages 328– 340. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Asai:1993:JMC

[AHAM93] K. Asai, K. Higuchi, M. Akimoto, and H. Matsumoto. The JAERI Monte Carlo machine [invited]. In Kusters et al. [KSW93], pages 341– 352. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Akimoto:1993:HHA

[AHFK93] M. Akimoto, K. Higuchi, M. Fujii, and S. Kambayashi. HASP: Human acts simulation program. In Kusters et al. [KSW93], pages 420– 431. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Arni:1994:ADE

[AHH94] V. V. Arni, T. L. Huntsberger, and B. A. Huntsberger. ADEAS: a distributed environment for acoustic simulation. In IEEE [IEE94c], pages 623–628. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Ahmed:1992:RHP

[Ahm92] Rana Ejaz Ahmed. Review of High-Performance Computer Architecture Ahmed, 2nd edn. (Stone, H. S; 1990). Computer, 25(4):111, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Asaoka:2002:EHJ

[AHOK02] Kae Asaoka, Akio Hirano, Yasuo Okabe, and Masanori Kanazawa. Evaluation of the HPF/JA extensions on Fujitsu VPP using the NAS parallel benchmarks. Lecture Notes in Computer Science, 2327:503–??, 2002.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2327/23270503.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270503.
pdf.

Aliaga:1997:PIG

[AHP97] J. I. Aliaga, V. Hernandez, and J. L. Perez. A parallel implementation of the general Lanczos method on the Cray T3D. Lecture Notes in Computer Science, 1215:168–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alef:1993:VTE

[AHSS93] M. Alef, C. P. Hugelmann, K. H. Schmidmeier, and D. Seldner. Visualization of technical electromagnetic devices developed at KfK. In Kusters et al. [KSW93], pages 797–?? ISBN 3-923704-11-9. LCCN ???? Two volumes.

Arno:1992:PSC

[AI92] Steven Arno and Ken Iobst. The PETASYS supercomputer and a class of pseudorandom number generators. Technical report SRC-TR-92-069, Supercomputing Research Center: IDA, Lanham, MD, USA, April 1992. 29 pp.

AIAA:1993:ACA

[AIA93] AIAA, editor. AIAA Computing in Aerospace 9 Conference: October 19–21, 1993, San Diego, CA, Papers 93-4469 — 93-4720 — American Institute of Aeronautics and Astronautics 1993. American Institute of Aeronautics and Astronautics, Washington, DC, USA, 1993. ISBN ???? LCCN ????

AIAA:1994:ASM

[AIA94] AIAA, editor. 32nd Aerospace Sciences Meeting and Exhibit: January 10–13, 1994, Reno, NV, Papers 94-0001 — 94-0872 — American Institute of Aeronautics and Astronautics. AIAA, Aerospace Center, 370 L’Enfant Promenade, SW, Washington, DC 20024-2518, USA, 1994. ISBN ???? LCCN ???? Twelve volumes.

Alexandrov:1997:SRI

[AISS97] Albert D. Alexandrov, Maximilian Ibel, Klaus E. Schauser, and Chris J. Scheiman. SuperWeb: research issues in Java-based global computing. Concurrency: practice and experience, 9(6): 535–553, June 1997. CO-DEN CPEXEI. ISSN 1040-3108. URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?ID=
13876;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13876&PLACEBO=IE.pdf.

Antoun:1993:ADA

[AJ93] N. Antoun and E. Jankovich. Automotive design applications of a non-linear implicit F.E.M. code. In Anonymous [Ano93-31], pages 709–718. ISBN 0-947719-62-8. LCCN ????

Andersson:1997:SCS

[AJ97] J. Andersson and B. Jacobson. A study of control strategies for hybrid vehicle propulsion systems using dynamic simulation. In Roller [Rol97], pages 289– 296. ISBN 0-947719-88-1 (paperback). LCCN ????

Ashcroft:1986:EEL

[AJFH86] Edward A. Ashcroft, R. Jagannathan, A. A. Faustini, and Ben M. Huey. Easyflow engines for Lucid: a family of supercomputer architectures based upon demanddriven and data-driven computation. Technical report CR-R 86063; TR 85-007, Dept. of Computer Science, College of Engineering and Applied Sciences, Arizona State University, Tempe, AZ, USA, 1986. 13 pp.

Allen:1987:ATF

[AK87] Randy Allen and Ken Kennedy. Automatic translation of Fortran programs to vector form. ACM Transactions on Programming Languages and Systems, 9(4):491–542, October 1987. CODEN ATPSDT.

ISSN 0164-0925 (print), 1558-4593 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0164-0925/
29875.html.

Axmann:1993:DTV

[AK93] J. K. Axmann and U. Kasemeyer. DIBU, a twodimensional vectorized diffusion code for the evaluation of the PROTEUS-PHASE-II highconverter experiments. In Kusters et al. [KSW93], pages 162–176. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Athas:1994:API

[AK94] W. C. Athas and J. G. Koller.

The architecture and programming of the ISI embeddable variant multicomputer. In IEEE [IEE94c], pages 134– 141. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Adeli:1995:CSO

[AK95] Hojjat Adeli and Sanjay Kumar. Concurrent structural optimization on massively parallel supercomputer. Journal of structural engineering, 121(11):1588– ??, November 1, 1995. CO-DEN JSENEI. ISSN 0733-9445.

Adams:1993:AES

[AKDM93] M. D. Adams, A. R.

Kerlavage, M. Dubnick, and R. F. Moreno. Analysis of expressed sequence tags from human brain cDNAs. In Lim et al. [L^{+}93], pages 113–120. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Anderson:1987:IOP

[AKG87] J. Wayne Anderson, Robert R. Kessler, and William F. Galway. Implementation and optimization of Portable Standard Lisp for the Cray. Proceedings of the Hawaii International Conference on System Science, 2:3–13, 1987. CODEN PHISD7. ISSN 0073-1129.

Armstrong:2006:CCM

[AKM^{+}06] Rob Armstrong, Gary Kumfert, Lois Curfman McInnes, Steven Parker, Ben Allan, Matt Sottile, Thomas Epperly, and Tamara Dahlgren. The CCA component model for high-performance scientific computing. Concurrency and Computation: Practice and Experience, 18(2):215– 229, February 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ando:1990:CSM

[AKT90] S. Ando, K. Kurimoto, and K. Taga. Crash simulation methods for vehicle development at Mazda. In Pitcher [Pit90], pages 475– 482. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Arrott:1992:SPV

[AL92a] M. Arrott and S. Latta. Supercomputers — perspectives on visualization. IEEE Spectrum, 29(9):61–65, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Arrott:1992:RCV

[AL92b] Matthew Arrott and Sara Latta. Researchers are counting on visualization to help them get the most out of supercomputers. IEEE Spectrum, 29(9):61–65, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Alabama-DEPA:19xx:ASA

[Alaxx] Alabama Supercomputer Authority, State of Alabama, Montgomery, 19xx. State of Alabama, Dept. of Examiners of Public Accounts, Montgomery, AL, USA.

Alestalo:1990:NWP

[Ale90]  M. Alestalo.  Numerical

weather prediction with the Finnish Cray. In Pitcher [Pit90], pages 171–179. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Aliabadi:1986:SPS

[Ali86] Adeeb Zarea Aliabadi. A supercomputer programming system. Thesis (Ph.D.), The Queen’s University of Belfast, Belfast, Northern Ireland, 1986. ?? pp.

Allen:1993:EPP

[All93] Sarah Allen. Evaluation of probation/parole scheduling via simulation. Research Report UMSI 93/115, University of Minnesota, Supercomputer Institute, Minneapolis, MN, USA, July 1993. 7 pp.

Abotel:1993:LET

[ALM93] K. N. Abotel, E. W. Larsen, and W. R. Martin. “Local” exponential transform methods for the Monte Carlo simulation of multigroup X,Ygeometry transport problems. In Kusters et al. [KSW93], pages 733–744. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Arbenz:1992:ADS

[ALMS92] P. Arbenz, H. P. Luthi, J. E. Mertz, and W. Scott. Applied distributed supercomputing in homogeneous networks. International Journal of High Speed Computing, 4 (2):87–108, June 1992. CO-DEN IHSCEZ. ISSN 0129-0533.

Attig:2001:GEL

[ALN^{+}01] N. Attig, Th. Lippert, H. Neff, J. Negele, and K. Schilling. Giant eigenproblems from lattice gauge theory on CRAY T3E systems. Computer Physics Communications, 142(1–3): 196–200, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465501003290.

Amestoy:2000:BRC

[ALPP00] Patrick Amestoy, Xiaoye Li, Fred Petry, and Marcin Paprzycki. Book reviews: Collection offers overview of research on structured matrices: High Performance Algorithms for Structured Matrix Problems, Advances in the Theory of Computation and Computational Mathematics, Volume 2; A comprehensive look at heterogeneous multidatabases: Management of Heterogeneous and Autonomous Database Systems; new version not sufficiently updated: Numerical Linear Algebra for High-Performance Computers. IEEE Concurrency, 8(4): 91–93, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL

http://dlib.computer.org/
pd/books/pd2000/pdf/p4091.
pdf.

Aluru:1996:PAL

[Alu96] S. Aluru. Parallel additive lagged Fibonacci random number generators. In ACM [ACM96], pages 102– 108. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Abe:1991:UHL

[AM91] M. Abe and T. Mimura. Ultrahigh-Speed HEMT LSI technology for supercomputer. IEEE Journal of Solid-State Circuits, 26(10): 1337–1344, October 1, 1991. CODEN IJSCBC. ISSN 0018-9200 (print), 1558-173X (electronic).

Adams:1993:UAS

[AM93a] B. T. Adams and J. E. Morel. An upscatter acceleration scheme for the multigroup Sn equations. In Kusters et al. [KSW93], pages 445–456. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Amphlett:1993:SLF

[AM93b] S. A. Amphlett and J. P. March. Studying low frequency vehicle phenomena using advanced modelling techniques. Part I — construction of a driveline model. In Anonymous [Ano93-31], pages 359–372. ISBN 0-947719-62-8. LCCN ????

Antoni:1993:HMN

[AM93c] John K. Antoni and Richard C. Metzger. Hypersphere mapper: a nonlinear programming approach to the hypercube embedding problem.

Journal of Parallel and Distributed Computing, 19(3): 262–??, November 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Armstrong:1994:ISI

[AM94] M. P. Armstrong and R. Marciano. Inverse-distance-weighted spatial interpolation using parallel supercomputers. Photogrammetric engineering and remote sensing, 60(9):1097–??, ???? 1994. ISSN 0099-1112.

Armstrong:1996:LIU

[AM96] M. P. Armstrong and R. J.

Marciano. Local interpolation using a distributed parallel supercomputer. International journal of geographical information systems, 10 (6):713–??, ???? 1996. ISSN 0269-3798.

Andersen:2015:MEL

[AM15] Timothy D. Andersen and Michael Mascagni. Memory efficient lagged-Fibonacci random number generators for GPU supercomputing. Monte Carlo Methods and Applications, 21(2):163–174, June 2015. CODEN MC-MAC6. ISSN 0929-9629 (print), 1569-3961 (electronic). URL http://

www.degruyter.com/view/
j/mcma.2015.21.issue-2/
mcma-2014-0017/mcma-2014-
0017.xml.

Ammarguellat:1989:NPC

[Amm89] Zahira Ammarguellat. Normalization of program control flow. Technical Report CSRD 885, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1989. 26 pp.

Ammarguellat:1990:CNA

[Amm90] Zahira Ammarguellat. A control-flow normalizations algorithm and its complexity. Technical Report CSRD 1042, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1990. 32 pp.

Ammarguellat:1992:CNA

[Amm92] Zahira Ammarguellat. A control-flow normalization algorithm and its complexity. Technical Report CSRD 1241, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1992. 38 pp.

Abraham:2015:GHP

[AMS^{+}15] Mark James Abraham, Teemu Murtola, Roland Schulz, Szilárd Páll, Jeremy C. Smith, Berk Hess, and Erik Lindahl. GROMACS: High performance molecular simulations through multi-level parallelism from laptops to supercomputers. SoftwareX, 1–2(??):13–18, September 2015. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711015000059.

Anastasio:1991:CSL

[Ana91] Thomas A. Anastasio. A critical survey of the literature on mapping static dataflow programs by level. Technical report SRC-TR-91-040, Supercomputing Research Center: IDA, Lanham, MD, USA, August 15, 1991. 14 pp.

Ananda:1994:DCC

[Ana94] A. L. Ananda. Distributed computing — client/server approach. In Balakrishnan [Bal94], pages 408–?? ISBN 0-07-462044-4. LCCN ????

Anderson:1988:PIP

[And88] Edward Charles Anderson. Parallel implementation of preconditioned conjugate gradient methods for solving sparse systems of linear equations. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 65 pp.

Andrews:1989:ITG

[And89] Phil Andrews. Integration of TEX and graphics at the Pittsburgh Supercomputing Center. TUGboat, 10(2):177– 178, July 1989. ISSN 0896-3207.

Anderson:1990:AIS

[And90a] Christine J. Anderson. Automated infrared scanning in Cray Y-MP production. Proceedings of the SPIE — The International Society for Optical Engineering, 1313: 207–216, 1990. CODEN PSISDG. ISBN 0-8194-0364-4. ISSN 0277-786X (print), 1996-756X (electronic).

Anderson:1990:RNG

[And90b] Stuart L. Anderson. Random number generators on vector supercomputers and other advanced architectures. SIAM Review, 32(2): 221–251, June 1990. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//www.jstor.org/stable/
2030521.

Andrews:1990:HTF

[And90c] John Barrett Andrews. A hardware tracing facility for a multiprocessing supercomputer. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1990. vi + 61 pp.

Anderson:2010:OLC

[And10] Mark Anderson. Optical lasers in a $100 cable. Really. IEEE Spectrum, 47(1): 24–25, January 2010. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2011:BBS

[And11]  Mark Anderson.  Better

benchmarking for supercomputers. IEEE Spectrum, 48 (1):12–14, January 2011. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:2020:WCA

[And20] M. Anderson. Will China attain exascale supercomputing in 2020? IEEE Spectrum, 57 (1):52–53, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Angouras:1991:SPP

[Ang91] George N. Angouras. Scheduling of parallel programs on multiprogrammed parallel processor systems. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1991. vii + 63 pp.

CSC:1992:AR

[Ann92] Annual report, 1992. Consortium Supercomputing Consortium, Pasadena, CA, USA.

Anonymous:1981:FR

[Ano81] Anonymous. For further reading. IEEE Spectrum, 18 (1):83–84, January 1981. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1985:NPC

[Ano85a] Anonymous. NBS parallel computer benchmark collection. Computers and Mathematics with Applications, 11 (12):1243, December 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122185901105.

Anonymous:1985:PSM

[Ano85b] Anonymous. Publications: Software Magazine (IEEE Computer Society); ZeroOne SUPERNET (supercomputer newsletter); technical reports from Argonne; Structured Fortran for Business (textbook); X/OPEN portability guide (common applications environment); Esprit ’84 status report (European technology research). ACM Fortran Forum, 4(3):14, October 1985. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1986:BII

[Ano86a] Anonymous. BT in India — IKBS funding — support for ADA — Anglo–French supercomputer. Data Processing, 28(2):58–63, March 1986. CODEN ???? ISSN 0011-684X (print), 1878-3058 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0011684X86900997.

Anonymous:1986:POS

[Ano86b] Anonymous. Prospects for an optical supercomputer. IEEE Spectrum, 23(8):44– 49, August 1986. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1986:Se

[Ano86c] Anonymous. Speakout. IEEE Spectrum, 23(6):18–19, June 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NEA

[Ano87a]  Anonymous.  NCSA ex-

pands academic affiliates program, SDSC installs molecular simulation library, Iowa simulation center goes from real time to the big time, Princeton and concurrent to codevelop Navier–Stokes supercomputer. Computers in Physics, 1(1):96– ??, November 1987. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4903444.

Anonymous:1987:PIc

[Ano87b] Anonymous. Papers are invited. IEEE Spectrum, 24 (3):55, March 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:PCC

[Ano87c]  Anonymous.  Princeton

and concurrent to codevelop Navier–Stokes supercomputer. Computers in Physics, 1(1):96–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

                                        //aip.scitation.org/doi/
                                        10.1063/1.4903447.

Anonymous:1987:TPA

[Ano87d] Anonymous. Twelve prestigious awards are announced: Winners of the IEEE’s specialist awards cover the fields of precision measurements, power transmission and distribution, electromagnetics, and supercomputer compilers. IEEE Spectrum, 24(8): 56–57, August 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1987:NMC

[Ano87e]  Anonymous.  The von

Neumann-Ulam Monte Carlo method for solving systems of linear algebraic equations on the ALLIANT FX/8. Technical Report CSRD 678, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 15 pp.

Anonymous:1988:AG

[Ano88a] Anonymous. The affordable gigaflop. Electronics and power, 34(3):123–??, March 10, 1988. CODEN ELPWAQ. ISSN 0013-5127.

Anonymous:1988:ASS

[Ano88b]  Anonymous.  Alliant to

ship second-generation minisupercomputers. Computers in Physics, 2(2):96– ??, March 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822712.

Anonymous:1988:ASEb

[Ano88c] Anonymous. Applications of supercomputers in engineering. Computers and Mathematics with Applications, 15(3):iv, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122188901782.

Anonymous:1988:ASEa

[Ano88d] Anonymous. Applications of supercomputers in engineering: Southampton University, England 5–7 September 1989. Computers and Mathematics with Applications, 15(2):iv, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122188900880.

Anonymous:1988:AAS

[Ano88e] Anonymous. ASE 89: Applications of supercomputers in engineering. Computers and Mathematics with Applications, 15(5):v, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122188901654.

Anonymous:1988:CPP

[Ano88f] Anonymous. Celerity packs power in departmental supercomputer. Computers in Physics, 2(1):84– ??, January 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822654.

Anonymous:1988:CCC

[Ano88g]  Anonymous.  Cray and

Convex compete for computing crown. Computers in Physics, 2(3):104– ??, May/June 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822716.

Anonymous:1988:FDS

[Ano88h] Anonymous. Fluid dynamics simulations up and running on ETA 10 and Ametek. Computers in Physics, 2(1):96–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.4822675.

Anonymous:1988:HSP

[Ano88i] Anonymous. House Subcommittee planning NSF supercomputing hearings. Computers in Physics, 2(2):96– ??, March 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822710.

Anonymous:1988:MPS

[Ano88j] Anonymous. Measuring the performance of supercomputers. Energy and technology review, 1988(4):1–??, May 1, 1988. ISSN 0884-5050.

Anonymous:1988:MES

[Ano88k] Anonymous. Memory enhancement for SCS supercomputers. Computers in Physics, 2(2):84– ??, March 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822686.

Anonymous:1988:OFI

[Ano88l] Anonymous. Ohio facility to install Cray Y-MP by mid-1989. Computers in Physics, 2(2):96–??, March 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4822709.

Anonymous:1988:SSSb

[Ano88m] Anonymous. Sandia SU-PERNET software yields supercomputer speeds. Computers in Physics, 2(5):104– ??, September 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4822770.

Anonymous:1988:SST

[Ano88n] Anonymous. Software and supercomputers tutor Carnegie-Mellon physics students. Computers in Physics, 2(1): 96–??, January 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4822674.

Anonymous:1988:SS

[Ano88o] Anonymous. Software for supercomputers. Computer, 21 (12):70–74, December 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1988:SRS

[Ano88p] Anonymous. Special report — software for supercomputers: Scientific supercomputer subcommittee of the IEEE committee on communications and information policy. Computer, 21(12):70– ??, December 1, 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1988:SAI

[Ano88q] Anonymous. The supercomputer and the automotive industry: Here’s a look at how the supercomputer is being used in crash simulation studies. Automotive Engineering, 96(11):56–62, November 1988. ISSN 0097-711X.

Anonymous:1988:SCF

[Ano88r] Anonymous. Supercomputer centers funding may increase by $20 million in FY1989. Computers in Physics, 2(4): 10–??, July 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822746.

Anonymous:1988:SVD

[Ano88s] Anonymous. Supercomputer vendor directory: a directory profiling 50 parallel processing hardware and 77 software vendors. Electronic Trend Publications, Saratoga, CA, USA, 1988. ISBN 0-914405-27-6 (paperback). various pp. LCCN ????

Anonymous:1988:SSJ

[Ano88t] Anonymous, editor. Supercomputing Symposium ’88, June 19–21, 1988, Edmonton, AB, Canada: proceedings. University of Alberta, Edmonton, AB, Canada, 1988. ISBN 0-88864-865-0. LCCN ????

Anonymous:1988:SSP

[Ano88u] Anonymous, editor. Supercomputing Symposium ’88, June 19–21, 1988, Edmonton, AB, Canada: proceedings. University of Calgary, Calgary, Alta., Canada, 1988. ISBN 0-88864-865-0.

Anonymous:1988:SAS

[Ano88v]  Anonymous.  SuperQuest

awarding supercomputer to high school scientists. Computers in Physics, 2(2):96– ??, March 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822711.

Anonymous:1988:TIC

[Ano88w] Anonymous, editor. Third International Conference on Supercomputing and Second World Supercomputer Exhibition, Hynes Convention Center, Boston, MA, May 15– 16, 1988: tutorial. International Supercomputing Institute, Inc., St. Petersburg, FL, USA, 1988.

Anonymous:1988:TPS

[Ano88x] Anonymous. Turn a PC into a supercomputer. Electronic Design, 36(26):89–??, December 28, 1988. CODEN ELODAW. ISSN 0013-4872.

Anonymous:1988:WIW

[Ano88y] Anonymous. What’s UP and down: Insidious as well as blatant changes have occurred in 25 years; our inventory reveals some surprises. IEEE Spectrum, 25(11):120– 123, November 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1989:TTM

[Ano89a] Anonymous. Are there too many supercomputer makers? Electronic business, 15 (4):23–??, February 20, 1989.

Anonymous:1989:APP

[Ano89b] Anonymous. Automatic parallel processing from Cray. Computers in Physics, 3(2): 96–??, March 1989. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822836.

Anonymous:1989:CG

[Ano89c] Anonymous. The color of geometry. Science News, 136(26 / 27):408–??, December 23, 1989. CO-DEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Anonymous:1989:CSP

[Ano89d] Anonymous. The computer spectrum: a perspective on the evolution of computing. Computer, 22(11):57– 63, November 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1989:CUF

[Ano89e] Anonymous. Cray under fire: Strategies for the supercomputer wars. Electronic business, 15(4):18–??, February 20, 1989.

Anonymous:1989:DSO

[Ano89f] Anonymous. DARPA studies optical, radar techniques for spotting submarines from air. Aviation week & space technology, 130(10):27–??, March 6, 1989. CODEN AWSTAV. ISSN 0005-2175.

Anonymous:1989:DD

[Ano89g] Anonymous. From dust to dust. Science News, 135(2): 24–??, January 14, 1989. CO-DEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Anonymous:1989:HCD

[Ano89h] Anonymous. High chip density boosts performance of interactive parallel supercomputer. Computer Design, 28(17):48–??, September 1, 1989. CODEN CMPDAM. ISSN 0010-4566.

Anonymous:1989:HHR

[Ano89i] Anonymous. A huge hoard of real estate will burden the government as it carries out the savings and loan bailout. Time, 133(18):54–??, May 1, 1989. CODEN TYMEA9. ISSN 0040-781X.

Anonymous:1989:ISS

[Ano89j] Anonymous. Introducing:

the super supercomputer. News in engineering, 61(3): 14–??, Summer 1989.

Anonymous:1989:IEE

[Ano89k] Anonymous. It’s end of an era for handmade Crays. Electronic Business, 15(4):20–??, February 20, 1989.

Anonymous:1989:LCC

[Ano89l] Anonymous. Look out, Cray — here comes a super-priced supercomputer. Electronics, 62(8):83–??, August 1, 1989. ISSN 0883-4989.

Anonymous:1989:NSP

[Ano89m] Anonymous. NASA supercomputer plan. Flight international, 135(4158):13–??, April 1, 1989.

Anonymous:1989:STE

[Ano89n] Anonymous. Sign of the times: Even Cray scales back R&D. Electronic business, 15 (16):61–??, August 7, 1989.

Anonymous:1989:SDS

[Ano89o] Anonymous. Supercomputer developments in the Soviet Union. Datamation, 35(14): 20–??, July 15, 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1989:SHU

[Ano89p] Anonymous. Supercomputer hardware: an update of the 1983 report’s summary and tables. Computer, 22(11): 63–68, November 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1989:S

[Ano89q] Anonymous. Supercomputing. Computer, 22(11):57– ??, November 1, 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1989:SCE

[Ano89r] Anonymous. Supercomputing in chemical engineering. Chemical engineering progress, 85(10):17–??, October 1, 1989. CODEN CEPRA8. ISSN 0360-7275.

Anonymous:1990:ATG

[Ano90a] Anonymous. Acousto/optical transducer gives holographic display: Experimental system uses a supercomputer and simplification of image data. Design news, 46(17): 64–??, September 3, 1990. CODEN DIGNAO. ISSN 0011-9407.

Anonymous:1990:ASS

[Ano90b] Anonymous. Alliant Sets Standards with Supercomputers. Computer Graphics World, 13(2):23–??, February 1990. CODEN CGWODH. ISSN 0271-4159.

Anonymous:1990:AI

[Ano90c] Anonymous. Authors’ index. In Pitcher [Pit90], pages 543–544. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Anonymous:1990:CS

[Ano90d] Anonymous. Colour section. In Pitcher [Pit90], pages 557–628. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Anonymous:1990:CCS

[Ano90e] Anonymous. Crystal clear storage: The holostore, a new mass storage device with supercomputer performance, could eliminate the I/O bottleneck. BYTE Magazine, 15(12):283–??, November 1, 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:FIC

[Ano90f] Anonymous, editor. The First International Conference on Supercomputing in Nuclear Applications (SNA ’90), March 12–16, 1990, Mito, Japan: proceedings. Nuclear Energy Data Center, Tokai-mura, Japan, 1990.

Anonymous:1990:FWC

[Ano90g] Anonymous, editor. First World Conference on Parallel Computing in Engineering and Engineering Education, Unesco, Paris, October 8–12, 1990. The Microcomputer Unit, London, UK, 1990. ISBN 1-873068-00-X. LCCN TA345.W67 1990.

Anonymous:1990:G

[Ano90h] Anonymous. The genius. Business week, 3157:80–??, April 30, 1990. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1990:GNT

[Ano90i] Anonymous. Gigabit network testbeds. Computer, 23(9): 77–80, September 1990. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:HUV

[Ano90j]  Anonymous.  HDTV up-

date: Video ‘supercomputer’ for advanced TV. Dealerscope merchandising, 32(1): 98–??, January 1, 1990.

Anonymous:1990:IVC

[Ano90k] Anonymous. Improved visualization capabilities for supercomputer users. Energy and technology review, pages 46–??, January 1, 1990. ISSN 0884-5050.

Anonymous:1990:IMP

[Ano90l] Anonymous. Internal motion of protein domains is probed by supercomputerbased method developed by Connecticut chemists. Chemical and engineering news, 68(50):20–??, December 10, 1990. CODEN CENEAR. ISSN 0009-2347.

Anonymous:1990:JSM

[Ano90m] Anonymous. Japan’s supercomputer market advances parallel processing. Signal, 44 (6):42–??, February 1, 1990.

Anonymous:1990:NRC

[Ano90n] Anonymous. New research consortium forms to use “world’s fastest computer,” an Intel supercomputer to be installed at Caltech. Chemical and engineering news, 68(50):20–??, December 10, 1990. CODEN CENEAR. ISSN 0009-2347.

Anonymous:1990:SPN

[Ano90o] Anonymous. Skystation puts new spin on a Sun: Accelerator gives supercomputer-like zip to SPARCstations without code alteration. Electronics, 63(10):87–??, October 1, 1990. ISSN 0883-4989.

Anonymous:1990:SED

[Ano90p] Anonymous. Stardent enters the desktop supercomputer market. Computer Graphics World, 13(10):25–??, October 1990. CODEN CG-WODH. ISSN 0271-4159.

Anonymous:1990:SSW

[Ano90q] Anonymous. Super sources: Who’s who in desktop supercomputing. BYTE Magazine, 15(5):258–??, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1990:MSG

[Ano90r] Anonymous. A US$67 million supercomputer grant from the Army. The Scientist (Philadelphia, PA), 4(7):5– ??, April 1990. ISSN 0890-3670.

Anonymous:1990:VOS

[Ano90s] Anonymous. Vendors object to supercomputer export plan. Computer Systems News, ??(455):6–7, February 1990. CODEN CSYND6. ISSN 0164-9981.

Anonymous:1990:WOD

[Ano90t] Anonymous, editor. Workshop on online documentation in the supercomputing environment: working paper, number NCAR/TN/353+PROC in NCAR technical note.

National Center for Atmospheric Research, Boulder, CO, USA, 1990.

Anonymous:1990:YSC

[Ano90u] Anonymous. The year in science — computers. Discover, 11(1):72–??, January 1, 1990.

Anonymous:1991:CP

[Ano91a] Anonymous. The CEDAR project. Technical Report CSRD 1122, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 51 pp.

Anonymous:1991:CS

[Ano91b] Anonymous. Computer and subsystems. Computer Design, 30(8):50–??, May 1, 1991. CODEN CMPDAM. ISSN 0010-4566.

Anonymous:1991:CAJ

[Ano91c] Anonymous. Computer: Applied Josephson junction device supercomputer. Science and technology in Japan, 10 (37):30–??, March 20, 1991. CODEN STJAE8. ISSN 0286-0406.

Anonymous:1991:CED

[Ano91d] Anonymous. Cray is eating dust. Business week, 3241: 88–??, November 25, 1991. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1991:DH

[Ano91e] Anonymous. Digital holography. Popular science, 238(1): 84–??, January 1, 1991.

Anonymous:1991:FRS

[Ano91f] Anonymous. FPS readies SuperCPU: Ships supercomputer based on Sun’s SPARC architecture. Information Week, 314:46–??, April 1, 1991. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1991:HRS

[Ano91g] Anonymous. Hospital records system gets boost from supercomputer. Modern healthcare, 21(26):29, July 1991. ISSN 0160-7480.

Anonymous:1991:IRD

[Ano91h]  Anonymous.  ISSCC re-

view: Digital technology: With more on-chip parallelism and closely coupled onchip memory, CPUs are hitting supercomputer-level performance numbers. Electronic Design, 39(3):53–??, February 14, 1991. CODEN ELODAW. ISSN 0013-4872. Anonymous:1991:JSK

[Ano91i] Anonymous. Japan’s supercomputer kick. Chilton’s automotive industries, 171(6): 185–??, June 1, 1991. CO-DEN CAUIEG. ISSN 0273-656X.

Anonymous:1991:NES

[Ano91j] Anonymous. National education supercomputer program. Energy and technology review, pages 19–??, April 1, 1991. ISSN 0884-5050.

Anonymous:1991:NSH

[Ano91k] Anonymous. Natural sciences: The human brain. German research: reports of the DFG, 1:27–??, 1991. ISSN 0172-1526.

Anonymous:1991:NCC

[Ano91l] Anonymous. Neural computer chip faster than supercomputer. Integrated circuits international, ????(????):4– ??, February 1991. ISSN 0263-6522.

Anonymous:1991:NTW

[Ano91m] Anonymous, editor. Nondestructive testing of wood: 8th International symposium — September 1991, Vancouver, WA. Washington State University, Pullman, WA, USA, 1991. ISBN ???? LCCN ????

Anonymous:1991:PRAa

[Ano91n] Anonymous. Perfect report 2: addendum 1. Technical Report CSRD 1052, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1991. 2 + [25] pp.

Anonymous:1991:PRAb

[Ano91o] Anonymous. Perfect report 2: addendum 2. Technical Report CSRD 1168, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1991. 20 pp.

Anonymous:1991:PSI

[Ano91p] Anonymous. Personal supercomputing with the Intel i860. BYTE Magazine, 16(1): 347–360, January 1991. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1991:PIS

[Ano91q] Anonymous, editor. Proceedings of the International Symposium on Supercomputing: Fukuoka, Japan, November 6–8, 1991. Kyushu University Press, Fukuoka, Japan, 1991. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Anonymous:1991:SLS

[Ano91r]  Anonymous.  A selected

list of supercomputer suppliers. Datamation, 37(4):76– ??, February 1991. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1991:SBG

[Ano91s] Anonymous. Supercomputer bill gets panel’s OK. Congressional quarterly weekly report, 49(13):794–??, March 30, 1991. ISSN 0010-5910.

Anonymous:1991:SIB

[Ano91t] Anonymous. Supercomputer imaging brings archaeologist face-to-face with mummy.

Research & Development, 33 (5):94, April 1991. CODEN REDEEA. ISSN 0746-9179.

Anonymous:1991:SR

[Ano91u] Anonymous. Supercomputer revolution. Chilton’s automotive industries, 171(5):55– ??, May 1991. CODEN CAUIEG. ISSN 0273-656X.

Anonymous:1991:SRS

[Ano91v] Anonymous. Supercomputer revolution: Supercomputing comes to the department level. Chilton’s automotive industries, 171(5):55–??, May 1, 1991. CODEN CAUIEG. ISSN 0273-656X.

Anonymous:1991:SVS

[Ano91w]  Anonymous.  Supercom-

puter voyages to the southern seas. Science, 254(5032):656– 662, November 1991. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1991:SHP

[Ano91x]  Anonymous.  Symposium

on high-performance computing for flight vehicles. Computers and Graphics, 15 (3):458, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1992:W

[Ano92a] Anonymous, editor. 11th Workshop: 1992, volume 6(1) of Speedup — Via Cantonale — 1992. Speedup, ????, 1992. ISBN ???? LCCN ????

Anonymous:1992:AS

[Ano92b] Anonymous. After the supercomputer. The Economist, 323(7757):102, May 1992. CODEN EONOEH. ISSN 0013-0613 (print), 1476-8860 (electronic).

Anonymous:1992:A

[Ano92c] Anonymous. Autorenverzeichnis. In Meuer [Meu92c], pages 245–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Anonymous:1992:DS

[Ano92d]  Anonymous.  Death of

the supercomputer. The Economist, 325(7787):79–??, November 1992. CODEN EONOEH. ISSN 0013-0613 (print), 1476-8860 (electronic).

Anonymous:1992:DNA

[Ano92e] Anonymous. Developing a Navier–Stokes algorithm for supercomputers. NASA tech briefs, 16(9):111–??, September 1992. CODEN NSTBAT. ISSN 0145-319X.

Anonymous:1992:EY

[Ano92f] Anonymous. Educator of the year. Electronic Learning, 12 (1):24–??, September 1, 1992. CODEN ELEADA. ISSN 0278-3258.

Anonymous:1992:ETI

[Ano92g] Anonymous, editor. Electromagnetic theory: International symposium — August 1992, Sydney, Australia, URSI International Symposium on Electromagnetic Theory 1992; 14th. URSI, ????, 1992. ISBN ???? LCCN ????

Anonymous:1992:EJE

[Ano92h]  Anonymous.  Electron-

ics: Japanese electric car to attempt 400 kph/Light-Emitting Silicon-Based material developed for supercomputer/transistor functionally comparable to human neuron cell. Science and technology in Japan, 11(42):44–??, June 30, 1992. CODEN STJAE8. ISSN 0286-0406.

Anonymous:1992:EN

[Ano92i] Anonymous. Enterprise networks. ComputerWorld, XXVI(42):77–??, October 1992. CODEN CMPWAB. ISSN 0010-4841.

Anonymous:1992:EDS

[Ano92j] Anonymous. European demand for supercomputers is growing. Business America: the magazine of international trade, 113(2):5–??, January 27, 1992. ISSN 0190-6275.

Anonymous:1992:FTS

[Ano92k] Anonymous. Fast track for supercomputer data. Datamation, 38(24):87–??, December 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:FMP

[Ano92l] Anonymous. Funding for a massively parallel supercomputer to advance the field of structural biology is being sought through a grand challenge grant proposal. Chemical and engineering news, 70 (9):25, March 2, 1992. CO-DEN CENEAR. ISSN 0009-2347.

Anonymous:1992:FT

[Ano92m] Anonymous. Future technologies. IEEE Spectrum, 29 (9):66–68, September 1, 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1992:HNM

[Ano92n] Anonymous. The hole of a new machine. Business week, 3296:124–??, December 7, 1992. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1992:JNJ

[Ano92o] Anonymous. Japan’s NEC joins bidding to supply NASA with new supercomputer. Aviation week & space technology, 136(17):21– ??, April 27, 1992. CODEN AWSTAV. ISSN 0005-2175.

Anonymous:1992:KSS

[Ano92p] Anonymous. KSR supercomputer scales up to 43 GFLOPS. Design news, 48(16):28–??, August 1992.

CODEN DIGNAO. ISSN 0011-9407.

Anonymous:1992:MFT

[Ano92q]  Anonymous.  Midrange/

mainframe: Fast track for supercomputer data. Datamation, 38(24):87–??, December 1, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MSO

[Ano92r] Anonymous. Molecular simulations open the friction frontier. Mechanical Engineering: the journal of the American Society of Mechanical Engineers, 114(9):60–??, September 1, 1992. CODEN MEENAH. ISSN 0025-6501.

Anonymous:1992:MM

[Ano92s] Anonymous. Moving to meta.

Information Week, 396:25– ??, October 19, 1992. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1992:NSR

[Ano92t] Anonymous. New supercomputer regulations. The Computer Lawyer, 9(7):39– ??, July 1992. CODEN CO-LAEB. ISSN 0742-1192.

Anonymous:1992:NTS

[Ano92u] Anonymous. The new technology of simulation. Automotive Engineering, 100(11): 49–??, November 1992. ISSN 0098-2571.

Anonymous:1992:OSC

[Ano92v] Anonymous. The Ohio supercomputer and die casting.

News in engineering, 64:8–??, Spring 1992.

Anonymous:1992:PPG

[Ano92w] Anonymous. Parallel processing gains in supercomputer market. Aviation week & space technology, 136(25):62– ??, June 22, 1992. CODEN AWSTAV. ISSN 0005-2175.

Anonymous:1992:PST

[Ano92x] Anonymous. Parallel supercomputers tackle tough research questions. Research & Development, 34(8):48–??, July 1992. CODEN RE-DEEA. ISSN 0746-9179.

Anonymous:1992:SPA

[Ano92y] Anonymous, editor. Signal processing applications and technology: International conference — November 1992, Boston, MA. DSP Associates, Newton MA, USA, 1992. ISBN ???? LCCN ????

Anonymous:1992:SCH

[Ano92z] Anonymous. Super clubs hit the sweet spot. Machine Design, 64(8):30–??, April 23, 1992. CODEN MADEAP. ISSN 0024-9114.

Anonymous:1992:SCa

[Ano92-27]  Anonymous.  Supercom-

puter cash. The Engineer, 275(7127):11–??, November 1992. CODEN ENGIAL. ISSN 0013-7758.

Anonymous:1992:SCb

[Ano92-28]  Anonymous.  Supercom-

puter choices. The Engineer, 275(7120):23–??, October 8, 1992. CODEN EN-GIAL. ISSN 0013-7758.

Anonymous:1992:SGC

[Ano92-29] Anonymous. Supercomputer giant Cray fights to keep up in MPP market. Electronics, 65(11):9–??, September 1992. ISSN 0883-4989.

Anonymous:1992:SGF

[Ano92-30]  Anonymous.  Supercom-

puter giants fight for piece of growing market in Korea. Electronics, 65(12):10– ??, September 28, 1992. ISSN 0883-4989.

Anonymous:1992:SR

[Ano92-31] Anonymous. Supercomputer row. The Engineer, 275 (7121):9–??, October 1992. CODEN ENGIAL. ISSN 0013-7758.

Anonymous:1992:Sa

[Ano92-32] Anonymous. Supercomputers. ComputerWorld, 26(47): 8–??, November 1992. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1992:Sb

[Ano92-33] Anonymous. Supercomputers. ComputerWorld, 26(47): 16–??, November 1992. CO-DEN CMPWAB. ISSN 0010-

Anonymous:1992:Sc

[Ano92-34] Anonymous. Supercomputers. ComputerWorld, 26(47): 33–??, November 1992. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1992:Sd

[Ano92-35] Anonymous. Supercomputers. ComputerWorld, 26(47): 35–??, November 1992. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1992:Se

[Ano92-36] Anonymous. Supercomputers. ComputerWorld, 26(47): 35–??, November 1992. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1992:Sf

[Ano92-37] Anonymous. Supercomputers. ComputerWorld, 26(47): 105–??, November 1992. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1992:SIH

[Ano92-38] Anonymous. Supercomputers image the human body in three dimensions. Science, 258(5083):747–??, October 1992. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1992:SKA

[Ano92-39] Anonymous. Supercomputers knock at IS doors. Datamation, 38(24):79–??, December 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SKD

[Ano92-40] Anonymous. Supercomputers Knock At IS Doors. Datamation, 38(24):79–??, December 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:SAH

[Ano92-41] Anonymous. Supercomputing at home. The Japan economic journal. Nihon keizai shimbun, 30(1528):8–??, August 1992. ISSN 0021-4388.

Anonymous:1992:SES

[Ano92-42] Anonymous. Superhuman effort: The story of one professor and his supercomputer demonstrates the effort needed to persevere with an invention while keeping it in the UK. The Engineer, 274(7094):28–??, April 2, 1992. CODEN ENGIAL. ISSN 0013-7758.

Anonymous:1992:TAN

[Ano92-43]  Anonymous.  Technology

in 1992 ascended to new heights. Electronic Design, 40(26):43–??, December 17, 1992. CODEN ELODAW. ISSN 0013-4872.

Anonymous:1992:TMT

[Ano92-44] Anonymous. Thinking Machines targets commercial users with a new supercomputer, but commerical software packages will not be available for at least a month. ComputerWorld, XXVI(42):

6–??, October 1992. CODEN CMPWAB. ISSN 0010-4841.

Anonymous:1992:PF

[Ano92-45] Anonymous. To probe further. IEEE Spectrum, 29(9): 76–??, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1992:WCB

[Ano92-46] Anonymous. Warren centre to boost supercomputer use. Process & control engineering, 45(5):6–??, May 1, 1992. ISSN 0816-8148.

Anonymous:1992:WAI

[Ano92-47]  Anonymous.  Wide Area

Information Servers: a supercomputer on every desk, 1992. 1 sound cassette.

Anonymous:1993:AMC

[Ano93a] Anonymous, editor. 7th Annual Midwest computer conference: March 1993, Whitewater, WI. University of Wisconsin, Whitewater, WI, USA, 1993. ISBN ???? LCCN ????

Anonymous:1993:APH

[Ano93b] Anonymous. Administration pushes high-tech initiative; NTTC and MIST cooperative for competitiveness; UPS chairman Kent Nelson receives technology award; US, Japan agree to optoelectronics project; supercomputing news; 1992 Manufacturing Intelligence Awards; IEEE sends technology list to Clinton administration; robot-in-residence and other smart machines; adapting to a changing world: An IEEE survey. IEEE expert: intelligent systems and their applications, 8(2):87–91, April 1993. CODEN IEEXE7. ISSN 0885-9000.

Anonymous:1993:AMP

[Ano93c] Anonymous, editor. Applications on massively parallel systems: 14th Workshop on vector and parallel computing — September 1993, Zurich, Switzerland, volume 7(2) of Speedup — Via Cantonale — 1993. Speedup, ????, 1993. ISBN ???? LCCN ????

Anonymous:1993:PSW

[Ano93d] Anonymous. Are pocket-size supercomputers on the way? on InfoWorld’s 15th anniversary, we look at the future of personal computing. InfoWorld, 15(50):48–??, December 1993. CODEN IN-WODU. ISSN 0199-6649.

Anonymous:1993:NR

[Ano93e]  Anonymous.  Nature re-

ports. Nature, 361(6410): 285–??, January 28, 1993. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:1993:C

[Ano93f] Anonymous. Conferences.

Computer, 26(1):112–??, January 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:CSC

[Ano93g] Anonymous, editor. Control systems and computer science: 9th International conference — May 1993, Bucharest. Polytechnical Institute of Bucharest, Bucharest, Roumania, 1993. ISBN ???? LCCN ????

Anonymous:1993:CUM

[Ano93h] Anonymous. Cray unveils mid-range supercomputers. Design news, 49(12):30–??, June 1993. CODEN DIG-NAO. ISSN 0011-9407.

Anonymous:1993:DCS

[Ano93i] Anonymous, editor. Dedicated conference on supercomputer applications in the automotive industries = 1993. 26th ISATA International Symposium on Automotive Technology and Automation, Aachen, Germany. Automotive Automation, Croydon, England, 1993. ISBN 0-947719-62-8 (paperback). LCCN ????

Anonymous:1993:DW

[Ano93j] Anonymous. Developments to watch. Business week, ?? (3321):83–??, May 1993. CO-DEN BUWEA3. ISSN 0007-7135.

Anonymous:1993:DS

[Ano93k] Anonymous. Double standard? The Japan economic journal. Nihon keizai shimbun, 31(1582):8–??, August 1993. ISSN 0021-4388.

Anonymous:1993:FFC

[Ano93l] Anonymous. Funding the fastest computer. Business week, 3307:40–??, March 1, 1993. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1993:HUG

[Ano93m] Anonymous. HP/UX gains super powers. Information Week, ??(423):14–??, May 1993. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1993:SPA

[Ano93n] Anonymous, editor. ICSPAT ’93: Proceedings of the fourth International Conference on Signal Processing Application and Technology: Santa Clara, California, USA, September 28–October 1, 1993. DSP Associates, Newton, MA, USA, 1993. ISBN ???? LCCN ???? Two volumes.

Anonymous:1993:ICM

[Ano93o] Anonymous. Individual chips muster to hit supercomputing aims. Signals, 47(10):45–??, June 1993. ISSN 0037-4938.

Anonymous:1993:TSY

[Ano93p]  Anonymous.  Is there a

supercomputer in your future? Production, 105(5):70– ??, May 1993. ISSN 0032-9819.

Anonymous:1993:MR

[Ano93q]  Anonymous.  Media re-

views. Computer, 26(4): 118–??, April 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MC

[Ano93r] Anonymous. Mix-and-match computing. Science News, 143(18):280–??, May 1993. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Anonymous:1993:MMA

[Ano93s] Anonymous. Moldflow mold analysis runs on Cray supercomputers. Modern plastics international, 23(6):8–??, June 1993. ISSN 0026-8283.

Anonymous:1993:NSH

[Ano93t]  Anonymous.  NCCS Sci-

ence Highlights, Supercomputing and Mass Storage Applications Supporting NASA Research for Earth and Space Sciences, FY 93. NASA, Washington, DC, USA, 1993. LCCN NAS 1.2: SCI 2/12 Govt Pub. 12/31/93 93-0706-P.

Anonymous:1993:PAM

[Ano93u] Anonymous. A pipeline approach to many-body problems. Physics world, 6(11): 32–??, November 1, 1993.

CODEN PHWOEW. ISSN 0953-8585.

Anonymous:1993:PVT

[Ano93v]  Anonymous.  PowerTools

from VXM technologies lets system managers create mixed clusters with the processing power of a supercomputer. Digital Review, 10(1): 7–??, January 1993. CODEN DIRVE5. ISSN 0739-4314.

Anonymous:1993:RS

[Ano93w] Anonymous. Redefining the supercomputer. Information Week, 420(420):24–??, April 12, 1993. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1993:R

[Ano93x] Anonymous. Review. New Scientist, 138(1869):38–??, April 1993. CODEN NWS-CAL. ISSN 0262-4079, 0028-6664.

Anonymous:1993:RC

[Ano93y] Anonymous. Rubbia’s concerto. IEE Review, 39(1):16– ??, January 1993. CODEN IEREEF. ISSN 0953-5683 (or 0013-5127??).

Anonymous:1993:RDP

[Ano93z] Anonymous. Russia debuts parallel supercomputer. Electronics, 66(21):3–??, November 8, 1993. ISSN 0883-4989.

Anonymous:1993:SUG

[Ano93-27] Anonymous. SGI unveils “graphics supercomputer”. Design news, 49(4):34–??, February 22, 1993. CODEN DIGNAO. ISSN 0011-9407.

Anonymous:1993:SUS

[Ano93-28] Anonymous. SGI unveils ‘graphics supercomputer’. Design news, 49(4):34–??, February 1993. CODEN DIGNAO. ISSN 0011-9407.

Anonymous:1993:SEP

[Ano93-29] Anonymous. Snapshot: Environmental protection agency. Government computer news, 12(1):12–??, January 1993. ISSN 0738-4300.

Anonymous:1993:SWb

[Ano93-30] Anonymous. SSI won’t die. Information Week, 411:15– ??, February 8, 1993. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1993:SAA

[Ano93-31]  Anonymous, editor.  Su-

percomputer applications in the automotive industries: ISATA International Symposium on Automotive Technology and Automation (26th: 1993: Aachen, Germany), ISATA -PROCEEDINGS-1993; 26th. Automotive Automation Ltd, Croydon, UK, 1993. ISBN 0-947719-62-8. LCCN ????

Anonymous:1993:SWa

[Ano93-32] Anonymous. A supercomputer in waiting. Information Week, 411(411):20–??, February 8, 1993. CODEN IN-FWE4. ISSN 8750-6874.

Anonymous:1993:SSa

[Ano93-33] Anonymous. Supercomputer sleeper. Business week, ?? (3331):72B, August 1993. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1993:SSF

[Ano93-34] Anonymous. Supercomputer speeds of five billion computation per second. Defense electronics, 25(3):20– ??, March 1993. ISSN 0278-3479.

Anonymous:1993:SSb

[Ano93-35] Anonymous. Supercomputer standoff. The Japan economic journal. Nihon keizai shimbun, 31(1575):8–??, June 1993. ISSN 0021-4388.

Anonymous:1993:SSc

[Ano93-36]  Anonymous.  Supercom-

puter system. Automotive Engineering, 101(11):14–??, November 1993. ISSN 0098-2571.

Anonymous:1993:SSM

[Ano93-37] Anonymous. Supercomputers simulate metal forming. Design news, 49(6):32–??, March 1993. CODEN DIG-NAO. ISSN 0011-9407.

Anonymous:1993:STC

[Ano93-38] Anonymous. Supercomputers talk common language; vendors target mainframe stronghold. Electronics, 66 (15):4–7, August 1993. ISSN 0883-4989.

Anonymous:1993:SEC

[Ano93-39] Anonymous, editor. Supercomputing Europe ’93 conference papers: fifth international exhibition and conference high-performance computing. Royal Dutch Fairs, Utrecht, The Netherlands, 1993.

Anonymous:1993:SFT

[Ano93-40] Anonymous. Supercomputing firm to tie in nationwide users via ATM. Network World, 10(41):8–??, October 1993. ISSN 0887-7661 (print), 1944-7655 (electronic).

Anonymous:1993:SST

[Ano93-41] Anonymous. Supercomputing specialists try out clusters and networks. Computers in Physics, 7(1): 6, January 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1993:TN

[Ano93-42]  Anonymous.  Technology

newsletter. Electronic Design, 41(4):23–??, February 1993. CODEN ELODAW. ISSN 0013-4872.

Anonymous:1993:TDB

[Ano93-43] Anonymous. Telemicrocopy — dialing for better research. Research & Development, 35 (12):66–??, October 25, 1993. CODEN REDEEA. ISSN 0746-9179.

Anonymous:1993:TT

[Ano93-44] Anonymous. Tire technology.

Popular mechanics, 170(2): 29–??, February 1993. ISSN 0032-4558.

Anonymous:1993:U

[Ano93-45] Anonymous. Update. Computer, 26(3):93–??, March 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:UAF

[Ano93-46] Anonymous. U.S. Air Force establishes supercomputing center. Defense electronics, 25(12):20–??, November 1993. ISSN 0278-3479.

Anonymous:1994:IAG

[Ano94a] Anonymous, editor. 27th ISATA, 1994, Aachen, Germany, 31st October–4th November: proceedings for the dedicated conferences on mechatronics and supercomputing applications in the transportation industries. Automotive Automation, Croydon, England, 1994. ISBN 0-947719-68-7. LCCN ????

Anonymous:1994:GUF

[Ano94b] Anonymous. A 65+ Gflop/s unstructured finite element simulation of chemically reacting flows on the Intel Paragon. In IEEE [IEE94e], pages 673–679. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ADP

[Ano94c] Anonymous. Adaptive data parallel methods for ecosystem monitoring. In IEEE [IEE94e], pages 281–290. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ALM

[Ano94d] Anonymous. Adaptive load migration systems for PVM. In IEEE [IEE94e], pages 390–399. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ASU

[Ano94e] Anonymous. Affinity scheduling of unbalanced workloads. In IEEE [IEE94e], pages 214–226. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:APUa

[Ano94f] Anonymous. Applicationspecific protocols for userlevel shared memory. In IEEE [IEE94e], pages 380– 389. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ABE

[Ano94g] Anonymous. Applications of boundary element methods on the Intel Paragon. In IEEE [IEE94e], pages 680–684. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:APUb

[Ano94h] Anonymous. Applications performance under OSF/1 AD and SUNMOS on Intel Paragon XP/S-15. In IEEE [IEE94e], pages 580–589. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:AVP

[Ano94i] Anonymous. Architecture of the VPP500 parallel supercomputer. In IEEE [IEE94e], pages 478–487. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:EPT

[Ano94j] Anonymous. Are expectations for parallelism too high? A survey of potential parallel users. In IEEE [IEE94e], pages 126–133. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ARC

[Ano94k] Anonymous. An atmospheric research center tried virtually every supercomputer under the sun before it made a decision. ComputerWorld, 28(31):55–??, August 1994. CODEN CMPWAB. ISSN 0010-4841.

Anonymous:1994:AHS

[Ano94l] Anonymous. Automated help system for a supercomputer. NASA tech briefs, 18(10): 102–??, October 1994. CO-DEN NSTBAT. ISSN 0145-319X.

Anonymous:1994:B

[Ano94m] Anonymous. Benchmarks. Datamation, 40(14):84–??, July 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:BBN

[Ano94n]  Anonymous.  Berkeley is

building a new supercomputer. The Chronicle of Higher Education, XLI(11): A19, November 1994. ISSN 0009-5982.

Anonymous:1994:BB

[Ano94o] Anonymous. Bits and bytes.

Business week, ??(3396): 116A, October 1994. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1994:BRP

[Ano94p]  Anonymous.  Book re-

view: Parallel supercomputing in MIMD architectures R. Michael Hord (ed.): Times Mirror International Publishers Ltd., c/o Excel Logistics, 3 Sheldon Way, Larkfield, Aylesford, Kent, ME20 6SF, England, ISBN 0-8493-4417-4, Price 61,00 Pounds Sterling. Journal of Computational and Applied Mathematics, 52(1–3): N3, July 20, 1994. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0377042794903697.

Anonymous:1994:BHC

[Ano94q] Anonymous. Building a.

high-performance collective communication library. In IEEE [IEE94e], pages 107–

116. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:CPV

[Ano94r] Anonymous. Cache performance in vector supercomputers. In IEEE [IEE94e], pages 255–264. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:CVS

[Ano94s] Anonymous. CERN vs Spain, NIST’s new boss, Thorp’s green light, EC budget, Princeton’s fusion record, Russia’s high energies, employment boom, supercomputing move. Physics world, 7(1):5–??, January 1994. CO-DEN PHWOEW. ISSN 0953-8585.

Anonymous:1994:CMR

[Ano94t] Anonymous. Communication and memory requirements as the basis for mapping and task and data parallel programs. In IEEE [IEE94e], pages 330–339. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:CCC

[Ano94u] Anonymous. A compilerdirected cache coherence scheme with improved and intertask locality. In IEEE [IEE94e], pages 773–782. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:CSM

[Ano94v] Anonymous. A computational steering model applied to problems in medicine. In IEEE [IEE94e], pages 540–549. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:CMS

[Ano94w] Anonymous. Computer modeling in science education: Toward a research and planning agenda. In IEEE [IEE94e], pages 5–6. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:C

[Ano94x] Anonymous. Conferences.

Computer, 27(1):97–??, January 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:CSP

[Ano94y] Anonymous. Control strategies for parallel mixed integer branch and bound. In IEEE [IEE94e], pages 41–48. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:CRS

[Ano94z] Anonymous. Cray Research shares programming model. ACM Fortran Forum, 13(1): 17, March 1994. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1994:CUF

[Ano94-27] Anonymous. Cray unveils the first in a new line of supercomputers designed for users with small budgets. Federal computer week, 8(29):33–??, September 1994. ISSN 0893-052X.

Anonymous:1994:DEN

[Ano94-28] Anonymous. The D editor: a new interactive parallel programming tool. In IEEE [IEE94e], pages 733–742. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:DSS

[Ano94-29] Anonymous. Data speeds from supercomputers to tape system. Design news, 49(6): 28–??, March 1994. CODEN DIGNAO. ISSN 0011-9407.

Anonymous:1994:DED

[Ano94-30] Anonymous. Design and evaluation of a DAMQ multiprocessor network with selfcompacting buffers. In IEEE [IEE94e], pages 713–722. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:DIM

[Ano94-31] Anonymous. Design and implementation of multicast operations for ATM-Based high performance computing. In IEEE [IEE94e], pages 164–173. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:DAN

[Ano94-32] Anonymous. Development and achievement of NAL numerical wind tunnel (NWT) and for CFD computations. In IEEE [IEE94e], pages 685–692. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:DOE

[Ano94-33] Anonymous. The development and operation of Edinburgh Parallel Computing and Centre’s Summer Scholarship Programme. In IEEE [IEE94e], pages 134–143. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:DEH

[Ano94-34] Anonymous. Distributed exploratorium for high performance computational and techniques. In IEEE [IEE94e], pages 117–125. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:DNC

[Ano94-35] Anonymous. Distributed network computing over local ATM networks. In IEEE [IEE94e], pages 154–163. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:DSA

[Ano94-36] Anonymous. DOE supercomputers to aid US E and P. The Oil and Gas Journal, 92 (21):29–??, May 1994. CO-DEN OIGJAV. ISSN 0030-1388.

Anonymous:1994:DFC

[Ano94-37] Anonymous. Dynamic fileaccess characteristics of a. production parallel and scientific workload. In IEEE [IEE94e], pages 640–649. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:DCI

[Ano94-38] Anonymous. Dynamic I/O characterization of I/O intensive scientific and applications. In IEEE [IEE94e], pages 660–669. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ECM

[Ano94-39] Anonymous. Editor-in chief’s message. Computer, 27(11): 5–??, November 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EM

[Ano94-40] Anonymous. Editor-in-chief’s message. Computer, 27(4): 6–??, April 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EAI

[Ano94-41] Anonymous. An efficient abstract interface for multidimensional array I/O. In IEEE [IEE94e], pages 650– 659. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:EAR

[Ano94-42]  Anonymous.  An effi-

cient algorithm for the runtime parallelization of and DOACROSS loops. In IEEE [IEE94e], pages 518–527. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ESC

[Ano94-43] Anonymous. Efficient and scalable cache coherence schemes for shared and memory hypercube multiprocessors. In IEEE [IEE94e], pages 498–507. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ECA

[Ano94-44] Anonymous. Efficient communication algorithms for pipeline multicomputers. In IEEE [IEE94e], pages 468– 477. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:EIM

[Ano94-45] Anonymous. Efficient implementation of the multigrid preconditioned and conjugate gradient method on distributed memory machines. In IEEE [IEE94e], pages 194–203. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:EPG

[Ano94-46] Anonymous. Efficient parallel global garbage collection on massively and parallel computers. In IEEE [IEE94e], pages 79–88. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:EUT

[Ano94-47] Anonymous. Enabling unimodular transformations. In IEEE [IEE94e], pages 753– 762. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:EMH

[Ano94-48] Anonymous. Environmental modeling helps clear the air. Mechanical Engineering: the journal of the American Society of Mechanical Engineers, 116(1):44–??, January 1994. CODEN MEENAH. ISSN 0025-6501.

Anonymous:1994:EAC

[Ano94-49]  Anonymous.  An experi-

mental APL compiler for a. distributed memory parallel and machine. In IEEE [IEE94e], pages 59–68. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ECD

[Ano94-50]  Anonymous.  Expressing

cross-loop dependencies through hyperplane data and dependence analysis. In IEEE [IEE94e], pages 508–517.

ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:EPP

[Ano94-51]  Anonymous.  EXTENT:

a portable programming environment for designing and implementing highperformance block-recursive algorithms. In IEEE [IEE94e], pages 49–58. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:FCG

[Ano94-52] Anonymous. Factors controlling generation and propagation of pacemaker and potentials in network models of mammalian SA node. In IEEE [IEE94e], pages 560–569. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:FRN

[Ano94-53] Anonymous. Fault-tolerant routing with non-adaptive wormhole algorithms and in mesh networks. In IEEE [IEE94e], pages 693–702. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:FTP

[Ano94-54] Anonymous. A fixed time performance evaluation of parallel CFD and applications. In IEEE [IEE94e], pages 18–23. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:GAM

[Ano94-55] Anonymous. Gallium arsenide makes news. Digital News and Review, 11(13):3– ??, July 1994. ISSN 1065-7452.

Anonymous:1994:GMS

[Ano94-56]  Anonymous.  Generalized

multiprocessor scheduling for directed acyclic and graphs. In IEEE [IEE94e], pages 237–246. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:GAG

[Ano94-57]  Anonymous.  Genetic al-

gorithms for graph partitioning and incremental and graph partitioning. In IEEE [IEE94e], pages 449–457. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:GAP

[Ano94-58] Anonymous. Global arrays:

a portable “Shared-Memory” programming model for distributed memory computers. In IEEE [IEE94e], pages 340–349. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:GOS

[Ano94-59]  Anonymous.  GRAPE-4:

a one-tflops special-purpose computer for astrophysical N-body problem. In IEEE [IEE94e], pages 429–438. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:HPC

[Ano94-60] Anonymous. The high performance computing and communications (HPCC) and program: Technologies for the national information and infrastructure. In IEEE [IEE94e], page 279. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:HPL

[Ano94-61] Anonymous. A high performance linear equation solver on the VPP500 and parallel supercomputer. In IEEE [IEE94e], pages 803–810. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:HPP

[Ano94-62] Anonymous. A high performance parallel algorithm for 1-D FFT. In IEEE [IEE94e], pages 34–40. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:IHS

[Ano94-63] Anonymous. Imagine holding a system with the power of a supercomputer in the palm of your hand. Digital News and Review, 11(14):15– ??, July 1994. ISSN 1065-7452.

Anonymous:1994:IPR

[Ano94-64] Anonymous. Implementation of a. portable and reproducible parallel and pseudorandom number generator. In IEEE [IEE94e], pages 311–319. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ILD

[Ano94-65]  Anonymous.  Improved

load distribution in parallel sparse Cholesky and factorization. In IEEE [IEE94e], pages 783–792. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:IPD

[Ano94-66] Anonymous. Improving the performance of DSM systems via compiler and involvement. In IEEE [IEE94e], pages 763–772. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:IT

[Ano94-67] Anonymous. Industry trends.

Computer, 27(11):6–??, November 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:INE

[Ano94-68]  Anonymous.  Integration

of numerical and experimental wind tunnels and (IofNEWT). In IEEE [IEE94e], page 7. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:IPH

[Ano94-69] Anonymous. Interpreting the performance of HPF/Fortran 90D. In IEEE [IEE94e], pages 743–752. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:IAH

[Ano94-70] Anonymous. Issues in applying high performance computing to real-time and systems. In IEEE [IEE94e], pages 1–2. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:IUE

[Ano94-71] Anonymous. Issues in undergraduate education in computational science and high performance computing. In IEEE [IEE94e], page 3. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:JOS

[Ano94-72] Anonymous. Japan opens supercomputer arena. Electronics, 67(4):2–??, February 28, 1994. ISSN 0883-4989.

Anonymous:1994:JNI

[Ano94-73] Anonymous. Joint NSF-NASA initiative in evaluation. In IEEE [IEE94e], pages 811–813. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:LAP

[Ano94-74] Anonymous. A library-based approach to portable, parallel, object-oriented programming: Interface, implementation, and application. In IEEE [IEE94e], pages 69–78. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:MSA

[Ano94-75] Anonymous, editor. Mechatronics and supercomputing applications in the transportation industries: ISATA International Symposium on Automotive Technology and Automation (27th: 1994: Aachen, Germany). Automotive Automation Limited, Croydon, UK, 1994. ISBN 0-947719-68-7. LCCN ????

Anonymous:1994:MNG

[Ano94-76] Anonymous. Meshes: The next generation. In IEEE [IEE94e], pages 275–276. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:MM

[Ano94-77] Anonymous. Miracle in Maui. Business week, ??(3390):84– ??, September 1994. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1994:NEP

[Ano94-78] Anonymous. NAS experiences with a. prototype cluster of workstations. In IEEE [IEE94e], pages 410–419. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:NCS

[Ano94-79] Anonymous. National Center for Supercomputing Applications. Industry Week, 243 (23):56–??, December 1994.

CODEN IWEEA4. ISSN 0039-0895.

Anonymous:1994:NR

[Ano94-80]  Anonymous.  Nature re-

ports. Nature, 371(6493):91– ??, September 8, 1994. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:1994:NRN

[Ano94-81] Anonymous. nCube releases new supercomputer. Information Week, 503(503):26– ??, November 28, 1994. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1994:NCW

[Ano94-82] Anonymous. A new computer workstation packs the power of a supercomputer into a desktop machine at a fraction of the cost. Chemical and engineering news, 72(43):34–??, October 24, 1994. CODEN CENEAR. ISSN 0009-2347.

Anonymous:1994:NCL

[Ano94-83] Anonymous. New Cray lowers ante for supercomputer entry. Design news, 49(21): 38–??, November 1994. CO-DEN DIGNAO. ISSN 0011-9407.

Anonymous:1994:NPA

[Ano94-84] Anonymous. Non-contiguous processor allocation algorithms for distributed and memory multicomputers. In IEEE [IEE94e], pages 227– 236. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:NSR

[Ano94-85]  Anonymous.  Nonnumeric

search results on the EM-4 distributed-memory and multiprocessor. In IEEE [IEE94e], pages 301–310. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:NAM

[Ano94-86] Anonymous. NSF has announced US$6 million in awards to make supercomputer centers more available to outside users. Federal computer week, 8(33):20–??, November 1994. ISSN 0893-052X.

Anonymous:1994:DCT

[Ano94-87] Anonymous. On the design of Chant: a talking threads package. In IEEE [IEE94e], pages 350–359. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:OSM

[Ano94-88] Anonymous. Optimal software multicast in wormholerouted multistage and networks. In IEEE [IEE94e], pages 703–712. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:ODR

[Ano94-89] Anonymous. Orientation determination in the 3D reconstruction of and icosahedral viruses using a. parallel computer. In IEEE [IEE94e], pages 550–559. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PTD

[Ano94-90] Anonymous. Paging tradeoffs in distributed sharedmemory multiprocessors. In IEEE [IEE94e], pages 590– 599. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PFI

[Ano94-91] Anonymous. A parallel formulation of interior point algorithms. In IEEE [IEE94e], pages 204–213. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PGA

[Ano94-92]  Anonymous.  A parallel

Gauss–Seidel algorithm for sparse power system and matrices. In IEEE [IEE94e], pages 184–193. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PIG

[Ano94-93] Anonymous. Parallel incremental graph partitioning using linear and programming. In IEEE [IEE94e], pages 458–467. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PIL

[Ano94-94] Anonymous. A parallel iterative linear solver for solving irregular grid and semiconductor device matrices. In IEEE [IEE94e], pages 24–33. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PLA

[Ano94-95]  Anonymous.  A parallel

Lauritzen-Spiegelhalter algorithm for probabilistic and inference. In IEEE [IEE94e], pages 320–329. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PPP

[Ano94-96] Anonymous. Parallel performance prediction using lost cycles analysis. In IEEE [IEE94e], pages 600–609. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PPSa

[Ano94-97] Anonymous. Parallel processing: Supercomputer software. DOE this month, 17(7):4–??, July 1, 1994.

Anonymous:1994:PPSb

[Ano94-98] Anonymous. Parallel protein structure determination from uncertain data. In IEEE [IEE94e], pages 570–579. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PMV

[Ano94-99] Anonymous. PARAMICS: Moving vehicles on the Connection Machine. In IEEE [IEE94e], pages 291–300. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PES

[Ano94-100] Anonymous, editor. Partial evaluation and semanticsbased program manipulation: Workshop — June 1994, Orlando, FL, number 9 in Technical Report- University of Melbourne Department of Computer Science 1994. University of Melbourne, Melbourne, Victoria, Australia, 1994. ISBN ???? LCCN ????

Anonymous:1994:PS

[Ano94-101] Anonymous. Pda to supercomputer. New electronics, 27(10):14–??, October 25, 1994. ISSN 0047-9624.

Anonymous:1994:PEI

[Ano94-102] Anonymous. Performance evaluation: Integrating techniques and tools into and environments and frameworks. In IEEE [IEE94e], pages 277–278. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PET

[Ano94-103] Anonymous. Performance evaluation of three distributed computing and environments for scientific applications. In IEEE [IEE94e], pages 400–409. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PHN

[Ano94-104] Anonymous. Performance of high-speed network I/O subsystems for: Case and study of a fibre channel network. In IEEE [IEE94e], pages 174–183. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PCM

[Ano94-105] Anonymous. The performance of the Cedar Multistage Switching Network. In IEEE [IEE94e], pages 265–274. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PDP

[Ano94-106] Anonymous. A portable debugger for parallel and distributed programs. In IEEE [IEE94e], pages 723–732. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:PND

[Ano94-107] Anonymous, editor. Proceedings of the 1994 NSF Design and Manufacturing Grantees Conference: Massachusetts Institute of Technology, the Laboratory for Manufacturing and Productivity, the Department of Ocean Engineering, The Materials Processing Center, Cambridge, Massachusetts, January 5–7, 1994. Society of Manufacturing Engineers, Dearborn, MI, USA, 1994. ISBN 0-87263-441-8. LCCN TS176.N72 1994.

Anonymous:1994:PTC

[Ano94-108] Anonymous, editor. Proceedings of the technical conference: International Electronics Packaging Conference, 14th, Atlanta, Georgia, September 25–28, 1994. International Electronics Packaging Society, Wheaton, IL, USA, 1994. ISBN 1-880433-16-8. LCCN ????

Anonymous:1994:RTD

[Ano94-109]  Anonymous.  The range

test: a dependence test for non-linear symbolic expressions. In IEEE [IEE94e], pages 528–537. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:RBH

[Ano94-110] Anonymous. Realizing the benefits of high performance computing for and solving environmental and other problems: What do we need and beyond faster hardware? In IEEE [IEE94e], page 280. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:RVP

[Ano94-111] Anonymous. Reducing the variance of point to point transfers in the IBM and 9076 parallel computer. In IEEE [IEE94e], pages 620–629. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:RCS

[Ano94-112] Anonymous. Run-time and compile-time support for adaptive irregular and problems. In IEEE [IEE94e], pages 97–106. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:SCSb

[Ano94-113] Anonymous. Scalability of the Cedar system. In IEEE [IEE94e], pages 247–254. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:SHE

[Ano94-114] Anonymous. A scalable highperformance environment for fluid flow and analysis on unstructured grids. In IEEE [IEE94e], pages 8–17. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:SPA

[Ano94-115] Anonymous. A scalable parallel algorithm for sparse Cholesky and factorization. In IEEE [IEE94e], pages 793–802. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:SPF

[Ano94-116] Anonymous. Scalable parallel formulations of the Barnes– Hut method for n-body simulations. In IEEE [IEE94e], pages 439–448. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:SUC

[Ano94-117]  Anonymous.  Scheduling

of unstructured communication on the Intel iPSC/ 860. In IEEE [IEE94e], pages 360–369. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:SMD

[Ano94-118] Anonymous. Selecting the “Best” MPP: Do benchmarks really tell the whole and story? In IEEE [IEE94e], page 4. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:SCO

[Ano94-119] Anonymous. Seymour Cray and other technology experts outline plans and visions for the future of supercomputing. ComputerWorld, 28(29): 121–??, July 1994. CODEN CMPWAB. ISSN 0010-4841. Anonymous:1994:SGA

[Ano94-120] Anonymous. Silicon graphics adds two entry-level models to its RISC-based supercomputers. ComputerWorld, 28(27):24–??, July 1994. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1994:SGP

[Ano94-121] Anonymous. Silicon Graphics is planning a one-two power punch to supercomputer competitors that already has analysts talking ringside. Digital Review, 11 (12):3–??, June 1994. CO-DEN DIRVE5. ISSN 0739-4314.

Anonymous:1994:SCS

[Ano94-122] Anonymous. Single chip supercomputer: Vector processing is key to high performance in floating point calculations. New electronics, 27 (4):7–??, April 1, 1994. ISSN 0047-9624.

Anonymous:1994:SCSa

[Ano94-123] Anonymous. Single chip supercomputer: Vector processing is key to high performance in floating point calculations. New electronics, 27 (4):7–??, April 1, 1994. ISSN 0047-9624.

Anonymous:1994:SLL

[Ano94-124] Anonymous. SLICC: a low latency interface for collective communications. In IEEE [IEE94e], pages 89–96. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:SS

[Ano94-125] Anonymous. Sorrowful supercomputers. The Economist, 332(7875):55–??, August 1994. CODEN EONOEH. ISSN 0013-0613 (print), 1476-8860 (electronic).

Anonymous:1994:SIP

[Ano94-126]  Anonymous.  Special is-

sue: panel sessions of the 1991 Workshop on Multithreaded Computers, November 22, 1991, Albuquerque, New Mexico, in conjunction with Supercomputing ’91. Computer architecture news, 22(1):2–33, 1994.

Anonymous:1994:SPH

[Ano94-127] Anonymous. Sunder: a programmable hardware prefetch architecture for numerical loops. In IEEE [IEE94e], pages 488–497. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:SSS

[Ano94-128] Anonymous. Super stars: Supercomputers throw their weight around wall street. Banking Technology, 11(7): 46–??, ???? 1994. CODEN BATEEM. ISSN 0266-0865.

Anonymous:1994:SCF

[Ano94-129] Anonymous. Supercomputer center’s fees supported. The Chronicle of Higher Education, XL(44):A23, July 1994. ISSN 0009-5982.

Anonymous:1994:SSM

[Ano94-130] Anonymous. Supercomputer sets mark at 1-Teraflop vector processing. Electronics, 67(21):2–??, November 14, 1994. ISSN 0883-4989.

Anonymous:1994:SEU

[Ano94-131] Anonymous. Supercomputers evolve as users multiply. Aviation week & space technology, 141(22):37–??, November 1994. CODEN AWSTAV. ISSN 0005-2175.

Anonymous:1994:SSI

[Ano94-132] Anonymous. Supercomputers to speed, improve reservoir simulation. The Oil and Gas Journal, 92(27):66– ??, July 1994. CODEN OIG-JAV. ISSN 0030-1388.

Anonymous:1994:SWH

[Ano94-133] Anonymous. Supercomputers: What happened? Information Week, ??(488):12– ??, August 1994. CODEN IN-FWE4. ISSN 8750-6874.

Anonymous:1994:SIC

[Ano94-134] Anonymous, editor. Supercomputing ’94: International conference — July 1994, Manchester. Univ of Manchester, Manchester, UK, 1994. ISBN ???? LCCN ????

Anonymous:1994:SRC

[Ano94-135] Anonymous. A supercomputing research center ATM network-interface board has a 1-Gbit/ s point-to-point link with 1.3-ms latency. Electronic engineering times, ?? (813):52–??, September 1994. ISSN 0192-1541.

Anonymous:1994:SSA

[Ano94-136] Anonymous. Supercomputing support for advanced biomedical imaging. In IEEE [IEE94e], pages 538–539. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:TBM

[Ano94-137] Anonymous. Tight binding molecular dynamics. In IEEE [IEE94e], pages 670–672. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:TSS

[Ano94-138] Anonymous. Time and/or space sharing in a. workstation cluster environment. In IEEE [IEE94e], pages 630–639. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:TNF

[Ano94-139] Anonymous. Tolerating node failures in cache only memory architectures. In IEEE [IEE94e], pages 370–379. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:TCO

[Ano94-140] Anonymous. Towards computations of ocean flows using Navier–Stokes and equations. In IEEE [IEE94e], pages 144–153. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:TDA

[Ano94-141] Anonymous. Truly distributionindependent algorithms for the N-body and problem. In IEEE [IEE94e], pages 420–428. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:U

[Ano94-142] Anonymous. Update. Computer, 27(2):84–??, February 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:UHS

[Ano94-143]  Anonymous.  Using high

speed networks to enable distributed parallel image and server systems. In IEEE [IEE94e], pages 610–619. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Anonymous:1994:VS

[Ano94-144] Anonymous. Virtual supercomputers. The Chronicle of Higher Education, XLI(13): A23, November 1994. ISSN 0009-5982.

Anonymous:1995:OSM

[Ano95a] Anonymous. As other supercomputer makers falter, Tera keeps trying. Puget Sound business journal, 15(49):4–??, April 1995. ISSN 8750-7757.

Anonymous:1995:CIS

[Ano95b] Anonymous. Call it the supernet. Business week, 3422: 120–??, May 1995. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1995:CCF

[Ano95c] Anonymous. Cray computer files for chapter 11, halting all work on its supercomputers. Electronic engineering times, ??(842):18–??, April 1995. ISSN 0192-1541.

Anonymous:1995:CEF

[Ano95d] Anonymous. Cray eliminates the fridge. Information Week, 519:126–??, March 20, 1995. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1995:CRR

[Ano95e] Anonymous. Cray Research rolls out ‘wireless’ supercomputer. Design news, 8(8):35– ??, ???? 1995. CODEN DIG-NAO. ISSN 0011-9407.

Anonymous:1995:GRR

[Ano95f] Anonymous. Cray Research rolls out ‘wireless’ supercomputer. Design news, 50(8): 35–??, April 1995. CODEN DIGNAO. ISSN 0011-9407.

Anonymous:1995:CRT

[Ano95g] Anonymous. Cray Research targets industrial users with its next generation of supercomputers. Computer-World, 29(10):61–??, March 1995. CODEN CMPWAB. ISSN 0010-4841.

Anonymous:1995:CSS

[Ano95h] Anonymous. Cray sets its sights on the federal market with the unveiling of its latest generation of vector supercomputers. Federal computer week, 9(5):33– ??, March 1995. ISSN 0893-052X.

Anonymous:1995:CSA

[Ano95i] Anonymous. Cray supercomputer aids in moldmaking. Modern plastics, 72(5):30–??, ???? 1995. CODEN MO-PLAY. ISSN 0026-8275.

Anonymous:1995:CCS

[Ano95j] Anonymous. Cray’s compact supercomputer for automotive. Chilton’s automotive industries, 175(1):26–??, January 1, 1995. CODEN CAUIEG. ISSN 0273-656X.

Anonymous:1995:CDU

[Ano95k]  Anonymous.  The cur-

tain drops on US supercomputer export restrictions. Computer fraud and security bulletin, ????(????):9– ??, November 1995. CO-DEN CFSBEK. ISSN 0142-0496 (print), 1878-3856 (electronic).

Anonymous:1995:DDS

[Ano95l] Anonymous. Dark days for science? federal budget cuts may threaten achievements ranging from supercomputers to atom smashers. will they imperil US science? two experts from the House of Representatives lock horns on the issue. Popular science, 247 (4):74–??, ???? 1995. ISSN 0161-7370.

Anonymous:1995:HS

[Ano95m] Anonymous. Der helvetische Supercomputer. Schweizerische technische Zeitschrift. Revue technique suisse. Rivista tecnica svizzera., 92(2): 25–??, ???? 1995. CODEN STZTA5. ISSN 0040-151X.

Anonymous:1995:DWa

[Ano95n] Anonymous. Developments to watch. Business week, ?? (3411):80–??, February 1995. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1995:DWb

[Ano95o] Anonymous. Developments to watch. Business week, ??(3443):140–??, September 1995. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1995:EPC

[Ano95p]  Anonymous.  Eight-way

processing Cray supercomputer designer, Steve Chen, is merging massively parallel processing and SMP. LAN times, 12(21):7–??, ???? 1995. ISSN 1040-5917.

Anonymous:1995:ESS

[Ano95q] Anonymous, editor. Electrochemical Society Spring Meeting: 187th Meeting — May 1995, Reno, NV, volume 95-1. The Electrochemical Society, Pennington, NJ, USA, 1995. ISBN 1-56677-107-2. LCCN ????

Anonymous:1995:EIS

[Ano95r] Anonymous. Embedded insights: The supercomputer is back. Electronic engineering times, ??(835):43–??, February 1995. ISSN 0192-1541.

Anonymous:1995:FC

[Ano95s] Anonymous. Feet of Cray.

Business week, 3419:42–??, April 10, 1995. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1995:FSN

[Ano95t]  Anonymous.  Fresh start

for NSF supercomputer centers. Science News, 148(26): 422–??, ???? 1995. CO-DEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Anonymous:1995:HST

[Ano95u] Anonymous, editor. Hardware and software trends in high-performance computing: 17th Workshop on vector and parallel computing — March 1995, Lugano, Switzerland, volume 9(1) of Speedup — Via Cantonale — 1995. Speedup, ????, 1995. ISBN ???? LCCN ????

Anonymous:1995:NNG

[Ano95v] Anonymous. In the news:

NSF grants $6 million; DoD’s five-year R&D plan; HPCC seeks new director; CERN offers supercomputing server; UK opens supercomputing center; NASA commercializes finite-element interface; Adam on the Internet; next year: Eve; when supercomputers aren’t enough. IEEE Computational Science & Engineering, 2(1):85–87, 89, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:NTF

[Ano95w] Anonymous. In the news: Thin-film lubricants may damage disk drives; protein structures calculated quickly; supercomputer looking for oil; modeling ceramics may improve yields; VLSI chip modeled after a leech; US Army studying imaging science; geomagnetic field reversals simulated; single-layer magnetism; National Medal of Science awarded to Herman A. Haus; distributed climate simulation; double bubble area is the smallest;l smart guitars. IEEE Computational Science & Engineering, 2(4):82–84, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:JBV

[Ano95x] Anonymous. Japan backs vector company; NSF under fire; supercomputer firms in trouble. Nature, 374(6521): 394–??, March 30, 1995. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:1995:LSD

[Ano95y] Anonymous. LANL supercomputer drives radiographic repository on TB. Radiology and imaging letter, 15(6):46– ??, April 1995. ISSN 0741-160X.

Anonymous:1995:PAP

[Ano95z] Anonymous. … Parallel and almost personal. Machine Design, 67(8):70–??, April 20, 1995. CODEN MADEAP. ISSN 0024-9114.

Anonymous:1995:LAN

[Ano95-27] Anonymous. Local area networks. Communications news (Geneva, IL), 32(3):67–??, March 1, 1995. CODEN CMUNA9. ISSN 0010-3632.

Anonymous:1995:LCS

[Ano95-28] Anonymous. Low-cost, compact supercomputers. Automotive Engineering, 103(1): 46–??, January 1995. ISSN 0098-2571.

Anonymous:1995:M

[Ano95-29]  Anonymous.  Manufacts.

Manufacturing systems, 13 (5):6–??, May 1995. CODEN MASYES. ISSN 0748-948X.

Anonymous:1995:MSH

[Ano95-30] Anonymous. Multiprocessor systems hit supercomputer speeds. Electronic Design, 43 (11):32–??, ???? 1995. CO-DEN ELODAW. ISSN 0013-4872.

Anonymous:1995:NNC

[Ano95-31]  Anonymous.  Nationwide

number cruncher: Using offthe-shelf kit, researchers are creating a virtual supercomputer reaching across the US. Communications international, 22(2):24–??, February 1995. ISSN 0305-2109.

Anonymous:1995:NIU

[Ano95-32]  Anonymous.  Networking:

IBM unleashes a suite of home-grown World Wide Web offerings encompassing notebooks to supercomputers; Ipswitch eyes Windows NT. Digital News and Review, 12(12):10–??, ???? 1995. ISSN 1065-7452.

Anonymous:1995:NSS

[Ano95-33] Anonymous. New supercomputer series sets industry pace. Electronics, 68(3):12– ??, February 13, 1995. ISSN 0883-4989.

Anonymous:1995:NPS

[Ano95-34] Anonymous. NEWS: Previews of Supercomputing ’95 and Plasma Physics Computer Show, review of Physics Computing ’95, corrections. Computers in Physics, 9(5): 475–??, ???? 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1995:NSH

[Ano95-35] Anonymous. NEWS: ‘reversed shear’ heats up research, supercomputing centers face uncertain future. Computers in Physics, 9(6): 575–??, ???? 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1995:NT

[Ano95-36] Anonymous. News trends.

Machine Design, 67(7):16– ??, April 1995. CODEN MADEAP. ISSN 0024-9114.

Anonymous:1995:NPC

[Ano95-37] Anonymous. NSF picks company to provide supercomputer network. The Chronicle of Higher Education, 41(33): A49, April 1995. ISSN 0009-5982.

Anonymous:1995:OCS

[Ano95-38] Anonymous, editor. Optical computing: summaries of the papers presented at the topical meeting, March 13–16, 1995, Salt Lake City, UT, USA, volume 10 of Technical Digest Series. Optical Society of America, Washington, DC, USA, 1995. ISBN 1-55752-389-4, 1-55752-390-8. LCCN ????

Anonymous:1995:PSA

[Ano95-39] Anonymous. Parallel supercomputers analyze suspensions. Tappi journal, 78(6):

30–??, ???? 1995. CODEN TAJODT. ISSN 0734-1415. Anonymous:1995:PSB

[Ano95-40] Anonymous. Parallel supercomputing is brought down to size. Signals, 50(2):61–??, ???? 1995. ISSN 0037-4938. Anonymous:1995:SSC

[Ano95-41] Anonymous. SCRI — Supercomputer Computations Research Institute. Florida State University research in review, 6(2):67–??, Spring 1995. ISSN 1043-4275.

Anonymous:1995:SB

[Ano95-42] Anonymous. Supercomputer for business. Information Week, 557(557):105–??, December 11, 1995. CODEN INFWE4. ISSN 8750-6874. Anonymous:1995:SCO

[Ano95-43] Anonymous. Supercomputing centre opened. Engineering, 236(3):6–??, ???? 1995. ISSN 0013-7782. Anonymous:1995:SMM

[Ano95-44] Anonymous. Supercomputing moves into the mainstream. Design news, ??(21): 25–??, ???? 1995. CODEN DIGNAO. ISSN 0011-9407. Anonymous:1995:SPS

[Ano95-45] Anonymous. Supercomputing: Panel suggests fewer NSF centers. Science, 269 (5228):1213–??, ???? 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1995:SAS

[Ano95-46]  Anonymous.  Swimming

against a supercomputing tide. Science, 269(5229): 1360–??, ???? 1995. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1995:TN

[Ano95-47]  Anonymous.  Technology

newsletter. Electronic Design, 43(11):27–??, May 30, 1995. CODEN ELODAW. ISSN 0013-4872.

Anonymous:1995:U

[Ano95-48] Anonymous. Update. Computer, 28(12):81–??, December 1, 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:USC

[Ano95-49] Anonymous. Update: Supercomputer centers. Computer, 28(12):81–??, December 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ISA

[Ano96a] Anonymous, editor. 29th International Symposium on Automotive Technology and Automation, Florence, Italy, 3rd–6th June 1996: fuzzy systems/soft computing in the automotive and transportation industries: supercomputer applications in the transportation industries. Automotive Automation, Croydon, England, 1996. ISBN 0-947719-81-4. LCCN ????

Anonymous:1996:AIS

[Ano96b]  Anonymous.  ACM/IEEE

Supercomputing ’96. Computer Graphics, 30(2):61– ??, May 1996. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Anonymous:1996:BRIi

[Ano96c]  Anonymous.  Book re-

view: An introduction to high-performance scientific computing: By Lloyd D. Fosdick, Elizabeth R. Jessup, Carolyn J. C. Schauble and Gitta Domik. MIT Press, Cambridge, MA. (1996). 760 pages. $55.00. Computers and Mathematics with Applications, 32(6): 135, September 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122196902576.

Anonymous:1996:CCU

[Ano96d] Anonymous. CAD/CAM Up-Date. Design engineering, 42 (6):12–??, June 1996. ISSN 0011-9342.

Anonymous:1996:CPS

[Ano96e] Anonymous. Call for papers: Special issue on irregular problems in supercomputing applications. Journal of Parallel and Distributed Computing, 39(2):192, December 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0167/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0167/production/pdf.

Anonymous:1996:CPSb

[Ano96f] Anonymous. Call for papers:

Special issue on irregular problems in supercomputing applications. Journal of Parallel and Distributed Computing, 39(2):192, December 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0167/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0167/production/pdf.

Anonymous:1996:C

[Ano96g] Anonymous. Commentary.

Business week, ??(3478):42– ??, June 1996. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1996:ETW

[Ano96h]  Anonymous.  Enabling

technologies for Web-based ubiquitous supercomputing. IEEE International Symposium on High Performance Distributed Computing, Proceedings, pages 112–119, 1996. CODEN PIDCFB. ISSN 1082-8907. IEEE catalog number 96TB100069.

Anonymous:1996:ESC

[Ano96i] Anonymous. Enhancing secondary chemistry instruction through supercomputing applications. In IEEE [IEE96a], pages 12–14.

Anonymous:1996:FSP

[Ano96j] Anonymous. French space plans under fire. supercomputing bids open in US. Nature, 384(6605):98, November 14, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:1996:GIC

[Ano96k] Anonymous, editor. Geo-Computations: International conference; 1st — September 1996, Leeds, volume 1 of GEOCOMPUTATION - PROCEEDINGS- 1996. [np], ????, 1996.

Anonymous:1996:GRS

[Ano96l] Anonymous, editor. Geologic remote sensing: practical solutions for real world problems: proceedings of the eleventh thematic conference, 27–29 February 1996, Las Vegas, Nevada, USA, volume 2 of Proceedings of the Thematic Conference on Geologic Remote Sensing 1996; conf 11. Environmental Research Institute of Michigan, Ann Arbor, MI, 1996. ISSN 1067-0106. LCCN QE 33.2 R4 G45 1996. Two volumes.

Anonymous:1996:GS

[Ano96m] Anonymous. Glueballs by supercomputer. The Sciences (New York), 36(3):10– ??, ???? 1996. CODEN SC-NCAD. ISSN 0036-861X.

Anonymous:1996:IQR

[Ano96n] Anonymous. In an insatiable quest for real-time information, analytics and more power, financial services firms are migrating towards supercomputers. Wall Street and Technology, 14(4):49–??, ???? 1996. CODEN WSTEE5. ISSN 1060-989X.

Anonymous:1996:INC

[Ano96o] Anonymous. Indian nuclear critic squeezed out. Japan denies supercomputer dumping charge. Nature, 381(6585): 723, June 27, 1996. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:1996:ITP

[Ano96p] Anonymous. Industry trends:

Palmtop wireless communications; TI superchip. Computer, 29(8):163–??, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LSA

[Ano96q] Anonymous, editor. Largescale structures in acoustics and electromagnetics: Symposium — September 26–27, 1994, Washington, DC. National Academy Press, Washington, DC, USA, 1996. ISBN 0-309-05337-4. LCCN TA646 .L35 1996.

Anonymous:1996:MSO

[Ano96r] Anonymous. Multiple servers offer supercomputer-class performance. Design news, 51(5):37–??, March 4, 1996. CODEN DIGNAO. ISSN 0011-9407.

Anonymous:1996:NAS

[Ano96s] Anonymous. A neighborly approach to supercomputing. Technology Review, 99(1):11– ??, ???? 1996. CODEN TEREAU. ISSN 0040-1692.

Anonymous:1996:NPR

[Ano96t] Anonymous. NEWS: Partnerships to replace supercomputer centers, simulating asteroid-impact probabilities Physics Computing ’96. Computers in Physics, 10(1): 9–??, January 1, 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1996:NTP

[Ano96u] Anonymous. News trends — ‘greatest plane ever’–or a pipe dream? Cryogenic cooler keeps picture clear for IR sensors. Solids and supercomputers to deliver cars in 24 months. Fluid power and motion shows come together in Chicago. New engines and brutal designs show up in Ford concept cars. Innovative design studio cuts product cycles in half. Machine Design, 68(8):24–??, ???? 1996. CODEN MADEAP. ISSN 0024-9114.

Anonymous:1996:Q

[Ano96v] Anonymous. Quanta. The Sciences (New York), 36(3): 10–??, May 1, 1996. CODEN SCNCAD. ISSN 0036-861X.

Anonymous:1996:RFR

[Ano96w] Anonymous, editor. Radio frequency radiation dosimetry workshop: present status and recommendations for future research: — 1996, AD REPORTS -NTIS- AD A 1996; AD-A309928. NTIS, [np]; [nd], 1996.

Anonymous:1996:RSG

[Ano96x]  Anonymous.  RANDOM

SAMPLES — getting a picture of sight. supercomputing program superseded. whistleblower woes. Science, 271 (5245):35–??, ???? 1996. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1996:RS

[Ano96y] Anonymous. Redefining the supercomputer. Science, 273 (5282):1655–??, ???? 1996. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1996:SIH

[Ano96z]  Anonymous.  Science in-

sights: Humans will remain the smartest entities on the planet, despite tough competition chess champion faced from a supercomputer. Chemical and engineering news: “news edition” of the American Chemical Society, 74(13):28–??, ???? 1996. ISSN 0009-2347.

Anonymous:1996:S

[Ano96-27] Anonymous. Science/technology. Chemical and engineering news, 74(13):28–??, March 25, 1996. CODEN CENEAR. ISSN 0009-2347.

Anonymous:1996:SBSa

[Ano96-28] Anonymous. SGI blends supercomputers — Silicon Graphics’ plan to blend its systems and Cray Research’s supercomputers will take a while. ComputerWorld, 30 (21):20–??, ???? 1996. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1996:SBSb

[Ano96-29] Anonymous. SGI buys supercomputer maker Cray. Info Canada, 21(5):24–??, ???? 1996. CODEN IFCAE3. ISSN 1187-7081.

Anonymous:1996:SDB

[Ano96-30] Anonymous. Sinking Deep Blue. Maclean’s, 109(9):60– ??, February 26, 1996. ISSN 0024-9262.

Anonymous:1996:SCS

[Ano96-31]  Anonymous.  Standards:

Cryptochip for supersafe communications. Computer, 29(2):78–??, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:SS

[Ano96-32] Anonymous. Super saviors.

Information Week, ??(568): 48–??, February 1996. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1996:SAN

[Ano96-33] Anonymous. Supercomputer aids nuclear testing. Electronic Design, 44(20):26–??, ???? 1996. CODEN ELO-DAW. ISSN 0013-4872.

Anonymous:1996:SDH

[Ano96-34] Anonymous. Supercomputer dispute hots up. France quells beef fears. Charity chief arrested. Nature, 382(6586): 5–??, July 4, 1996. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:1996:SIB

[Ano96-35]  Anonymous.  Supercom-

puter: IBM baut RISC-Grossrechner. Elektronik, 45 (18):8–??, ???? 1996. CO-DEN EKRKAR. ISSN 0013-5658.

Anonymous:1996:SI

[Ano96-36] Anonymous. Supercomputer installations. Research & Development, 38(5):9–??, ???? 1996. CODEN REDEEA. ISSN 0746-9179.

Anonymous:1996:SNS

[Ano96-37] Anonymous. Supercomputers — National Science Foundation overhauls its supercomputer center program. ComputerWorld, 30(3):41– ??, ???? 1996. CODEN CM-PWAB. ISSN 0010-4841.

Anonymous:1996:SAW

[Ano96-38] Anonymous. SuperComputers: Alive and well. Information Week, ??(591):26–??, August 1996. CODEN IN-FWE4. ISSN 8750-6874.

Anonymous:1996:SGN

[Ano96-39] Anonymous. Supercomputers grow into new roles and applications. Signals, 50(6):27–??, ???? 1996. ISSN 0037-4938.

Anonymous:1996:SAS

[Ano96-40] Anonymous. Supercomputing at Sandia. Geotimes, 41 (4):6–??, ???? 1996. ISSN 0016-8556.

Anonymous:1996:TFT

[Ano96-41]  Anonymous.  Time for

fair trading of supercomputers. Nature, 382(6586):

1, July 4, 1996. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:1996:USC

[Ano96-42] Anonymous. Update: Seymour Cray legacy. Computer, 29(11):88–??, November 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:VPC

[Ano96-43] Anonymous, editor. Vector and parallel computing: Workshop; 19th — March 1996, Basel, Switzerland, volume 10(1) of SPEEDUP - VIA CANTONALE- 1996; /2. SPEEDUP Society, ????, 1996. ISSN 1421-6337.

Anonymous:1996:YMD

[Ano96-44] Anonymous. Your move, Deep Blue. Business week, 3463:36–??, February 19, 1996. CODEN BUWEA3. ISSN 0007-7135.

Anonymous:1997:CSS

[Ano97a]  Anonymous.  Computers/

software — supercomputer grant. Research & Development, 39(5):9–??, ???? 1997. CODEN REDEEA. ISSN 0746-9179.

Anonymous:1997:CUM

[Ano97b] Anonymous. CS update:

Members elect officers and board members to begin terms in 1998; balloting and participation; Kenneth Laker wins in balloting for IEEE 1998 president-elect; 1987 Gorden Bell Prize winners announced at supercomputing. Computer, 30 (12):85–86, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://pdf.computer.org/
co/books/co1997/pdf/rz085.
pdf.

Anonymous:1997:EAC

[Ano97c] Anonymous, editor. European aerosol conference: — September 1997, Hamburg, Germany, volume 28 of Journal of Aerosol Science. Pergamon, New York, NY, USA, 1997. ISSN 0021-8502.

Anonymous:1997:HTS

[Ano97d] Anonymous. HARDWARE Top-500 supercomputer — Number-Cruncher berechnen den Aufprall eines eine Milliarden Tonnen schweren Kometen. Computerwoche, 24(28):35–??, ???? 1997. ISSN 0170-5121.

Anonymous:1997:FUS

[Ano97e]  Anonymous.  In focus:

University supercomputers compete to simulate nuclear weapons. Scientific American, 276(3):14– ??, March 1997. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.sciam.com/0397issue/
0397currentissue.html.

Anonymous:1997:IGS

[Ano97f] Anonymous. Internet goldmine: San Diego Supercomputer Center. Computers in Physics, 11(3):229– ??, ???? 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1997:INI

[Ano97g] Anonymous, editor. Issues for networked interpersonal communicators: Colloquium — May 1997, London, number 139 in COLLOQUIUM DIGEST- IEE 1997. IEE, London, UK, 1997. ISSN 0963-3308.

Anonymous:1997:NSC

[Ano97h] Anonymous. New supercomputer cracks ‘teraflops’ barrier. Design news, ??(1):36– ??, ???? 1997. CODEN DIG-NAO. ISSN 0011-9407.

Anonymous:1997:NIP

[Ano97i] Anonymous. NEWS — Internet 2, Physics Computing ’97, Supercomputing ’96. Computers in Physics, 11(1): 8–??, ???? 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1997:NAG

[Ano97j] Anonymous. News: American, German supercomputers get cozy; lung and weather simulations win Computerworld Smithsonian Award; State of the Field talks to feature HPC experts; UCAR and HP to collaborate on shared-memory systems; Intel Teraflops is up and running; Tera sets integersorting record. IEEE Concurrency, 5(3):78–80, July/ September 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1997/pdf/p3078.
pdf.

Anonymous:1997:NNC

[Ano97k] Anonymous. NEWS: NSF chooses supercomputing partnerships, conference looks at computing’s future fluid dynamics used in observatory siting. Computers in Physics, 11(3):216–??, ???? 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1997:NPP

[Ano97l] Anonymous. NEWS: Previews of PC ’97 and Tokyo supercomputing meeting; news update. Computers in Physics, 11(4):312– ??, ???? 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Anonymous:1997:NTW

[Ano97m] Anonymous. News trends — workstations harness new Pentium IIs. planetary rover. supercomputer pushes for speed. flat-panel display. crash dummy ‘organs’. improved displays. sophisticated design software. old bird, new missions. Machine Design, 69(12):22–??, ???? 1997. CODEN MADEAP. ISSN 0024-9114.

Anonymous:1997:NWS

[Ano97n] Anonymous. The next wave of supercomputing centers. Science, 275(5305):1412–??, ???? 1997. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1997:NDA

[Ano97o] Anonymous. NIST develops advanced user interface to supercomputing applications. Journal of research of the National Institute of Standards and Technology, 102(5):601–??, ???? 1997. CODEN JRITEF. ISSN 1044-677X (print), 2165-7254 (electronic).

Anonymous:1997:OOS

[Ano97p] Anonymous. OhioLINK and Ohio Supercomputer unite in access and delivery project. Library Hi Tech News, ?? (142):29–??, ???? 1997. ISSN 0741-9058.

Anonymous:1997:PJI

[Ano97q] Anonymous, editor. Proceedings of the Joint International Conference on Mathematical Methods and Supercomputing for Nuclear Applications, Saratoga Springs, New York, October 5-9, 1997. American Nuclear Society, La Grange Park, IL, USA, 1997. ISBN 0-89448-619-5. LCCN TK9006 .J67 1997 v.1-2. Two volumes.

Anonymous:1997:SDS

[Ano97r] Anonymous. San Diego Supercomputer Center. Computers in Physics, 11(3): 229–??, May 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822548.

Anonymous:1997:SRS

[Ano97s] Anonymous. Sandia running ‘fastest supercomputer in world’ — capable of one trillion floating point operations a second. Aviation week & space technology, 147(1): 65–??, ???? 1997. CODEN AWSTAV. ISSN 0005-2175.

Anonymous:1997:SCO

[Ano97t] Anonymous. Supercomputer cleans up one of the last vestiges of the Cold War. ComputerWorld, 31(49):53– ??, ???? 1997. CODEN CM-PWAB. ISSN 0010-4841.

Anonymous:1997:SFT

[Ano97u] Anonymous. Supercomputer faces test — the world’s fastest supercomputer will be put to the test to see if it can handle the job of ensuring the safety of the U.S. nuclear stockpile. Defense news, 12 (26):13–??, ???? 1997. ISSN 0884-139X.

Anonymous:1997:SSG

[Ano97v] Anonymous. Supercomputer software goes PC. Design news, ??(7):45–??, ???? 1997. CODEN DIGNAO. ISSN 0011-9407.

Anonymous:1997:SAT

[Ano97w] Anonymous. Supercomputing applications TC. Computer, 30(11):84–??, ???? 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:SC

[Ano97x] Anonymous. Supercomputing centers. Research & Development, 39(8):11–??, ???? 1997. CODEN REDEEA. ISSN 0746-9179.

Anonymous:1997:SCT

[Ano97y] Anonymous. Supercomputing centers throw big hardware at business processes. ComputerWorld, 31(27):69– ??, ???? 1997. CODEN CM-PWAB. ISSN 0010-4841.

Anonymous:1997:SIS

[Ano97z] Anonymous. Supercomputing: Illinois, San Diego centers win renewed NSF funding. Science, 276(5309):29– ??, ???? 1997. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1997:SRA

[Ano97-27] Anonymous. SUPERCOM-PUTING: Researchers at five universities will gain access to DOE’s biggest computers. Chemical and engineering news: “news edition” of the American Chemical Society, 75(32):11–??, ???? 1997. ISSN 0009-2347.

Anonymous:1997:TAF

[Ano97-28] Anonymous. Technical activities forum: Supercomputing Applications TC faces new challenges. Computer, 30(11):84, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://pdf.computer.org/
co/books/co1997/pdf/ry084.
pdf.

Anonymous:1997:TBS

[Ano97-29]  Anonymous.  Technology

breakthrough — specialized algorithms, combined with supercomputer technology promise improved cancer radiation therapies. verticallystacked multichip modules promise low-cost miniaturization of electronics in 3D. Electronic Design, 45(6):35– ??, ???? 1997. CODEN ELO-DAW. ISSN 0013-4872.

Anonymous:1997:VPC

[Ano97-30] Anonymous, editor. Vector and parallel computing: Joint workshop; 22nd — September 1997, Lausanne, Switzerland, volume 11(2) of SPEEDUP -VIA CANTONALE- 1997. SPEEDUP Society, ????, 1997. ISSN 1421-6337.

Anonymous:1998:ASM

[Ano98a] Anonymous, editor. Aerospace sciences meeting: 36th — January 1998, Reno, NV, number 540 in PAPERS-AMERICAN INSTITUTE of AERONAUTICS AND AS-TRONAUTICS 1998. AIAA, Aerospace Center, 370 L’Enfant Promenade, SW, Washington, DC 20024-2518, USA, 1998.

Anonymous:1998:CPA

[Ano98b] Anonymous. Call for papers: 1999 ACM International Conference on Supercomputing. ACM SIG-PLAN Notices, 33(12):15, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:CPAb

[Ano98c] Anonymous. Call for papers: 1999 ACM International Conference on Supercomputing. ACM SIG-PLAN Notices, 33(12):15, December 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1998:CUS

[Ano98d] Anonymous. CS update: Seymour Cray Award; Pioneer Award for Svoboda; Golden Core memberships. Computer, 31(1):95–96, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://pdf.computer.org/
                                        co/books/co1998/pdf/r1095.
                                        pdf.

Anonymous:1998:EBP

[Ano98e] Anonymous. From the editor: The Beowulf Project #46. Linux Journal, 45: ??, January 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1998:SPM

[Ano98f]  Anonymous.  Supercom-

puting partnerships move ahead; Pittsburgh, Cornell find new projects. Computers in Physics, 12(3): 218–??, May 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.168678.

Anonymous:1998:TSA

[Ano98g] Anonymous. The teraflop supercomputer APEmille: architecture, software and project status report. Computer Physics Communications, 110(1–3):216–219, May 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S001046559700180X.

Anonymous:1999:NFP

[Ano99] Anonymous. News: Supercomputing fashion: Mining scientific databases: Cluster computing: Los Alamos computing complex: a smarter DoD: Parallel processing: Grid community effort. IEEE Concurrency, 7(4):4–??, October/December 1999. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1999/pdf/p4004.
pdf.

Anonymous:2000:BRSb

[Ano00a]  Anonymous.  Book re-

view: Supercomputing, collision processes, and applications: Edited by Kenneth L. Bell, Keith A. Berrington, Derrick S. F. Crothers, Alan Hibbert and Kenneth T. Taylor. Kluwer Academic/ Plenum Publishers, New York. (1999). 284 pages. $125.00, GBP 81.25, NLG 255.00. Computers and Mathematics with Applications, 39(3–4):262, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122100900312.

Anonymous:2000:MNM

[Ano00b] Anonymous. Micro news: Motorola expands IP and SOC efforts; market benefits again; customized VLIW cores proposed; building the world’s fastest supercomputer; advancing wireless use. IEEE Micro, 20(1):4– 5, January/February 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2000/pdf/m1004.
                                        pdf.

Anonymous:2000:NAS

[Ano00c] Anonymous. News: Supercomputing, teraflop performance, embedded Java, middleware. IEEE Concurrency, 8(1):5–??, January/March 2000. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

                                        //dlib.computer.org/pd/
                                        books/pd2000/pdf/p1005.
                                        pdf.

Anonymous:2000:NST

[Ano00d] Anonymous. News: Supercomputing, teraflop performance, embedded Java, middleware. IEEE Concurrency, 8(1):5–??, January/March 2000. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/
books/pd2000/pdf/p1005.
pdf.

Anonymous:2001:CRW

[Ano01a] Anonymous. Conference report: 5th Workshop on Distributed Supercomputing Scalable Cluster Software. ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2001:ESL

[Ano01b] Anonymous. Express: Supercomputer auf Linux-Basis: 3300-Itanium-Prozessoren fur neuen Weltrekordrechner. Elektronik, 50(18):10–11, 2001. CODEN EKRKAR. ISSN 0013-5658.

Anonymous:2001:WSM

[Ano01c] Anonymous. A Windows for supercomputing: Microsoft is quietly launching an OS to handle Intel’s new, 64-bit Itanium CPU. PC World, 19(8):58–65, 2001. CO-DEN PCWDDV. ISSN 0737-8939 (print), 1944-9143 (electronic).

Anonymous:2002:MNI

[Ano02a] Anonymous. Micro news:

IBM’s Cell completes design phase; silver molecules render electroluminescent light source; next-generation disc storage; IBM electron microscope; Cornell clustersupercomputing expansion; silicon makes low-voltage gas sensor; NIST chemists explore plastics; micro bits. IEEE Micro, 22(5):9–11, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5009.
pdf.

Anonymous:2002:MNIa

[Ano02b] Anonymous. Micro news: IBM’s Cell completes design phase; silver molecules render electroluminescent light source; next-generation disc storage; IBM electron microscope; Cornell clustersupercomputing expansion; silicon makes low-voltage gas sensor; NIST chemists explore plastics; micro bits. IEEE Micro, 22(5):9–11, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5009.
pdf.

Anonymous:2003:MNIc

[Ano03a] Anonymous. Micro news: IBM ups the ante in silicon transistor speed; new benchmark suite based on highperformance computing applications, MPI and OpenMP [SPEC HPC2002]; EU OKs Hitachi, Mitsubishi Electric semiconductor joint venture; Intel launches Pentium 4 at 3.06 GHz; TSMC unveils viable 25nm transistors. IEEE Micro, 23(1):6–6, 87, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2003/pdf/m1006.
pdf.

Anonymous:2003:SLB

[Ano03b] Anonymous. Spectral lines boost your IQ, build a supercomputer. IEEE Spectrum, 40(7):7, July 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2005:SLNb

[Ano05] Anonymous. Spectral lines:

Nuclear testing goes virtual. IEEE Spectrum, 42 (12):10, December 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:EC

[Ano08a] Anonymous. Exascale computing by 2015? IEEE Spectrum, 45(12):12, December 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2008:UBC

[Ano08b] Anonymous. Update: A better camera pill. IEEE Spectrum, 45(8):12, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2009:CPSa

[Ano09] Anonymous. Call for papers for special issue on highperformance computing with accelerator. IEEE Transactions on Parallel and Distributed Systems, 20(6):912, June 2009. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Anonymous:2011:CSWb

[Ano11a] Anonymous. Chuck Seitz wins Cray Award. Computer, 44(11):70, November 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2011:SST

[Ano11b] Anonymous. The scientist as space tourist. IEEE Spectrum, 48(5):18, May 2011. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2014:EWS

[Ano14] Anonymous. Europe wants a smartphone supercomputer [news]. IEEE Spectrum, 51 (6):20, June 2014. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2016:NSS

[Ano16] Anonymous. Nominations are solicited for the Seymour Cray, Sidney Fernbach & Ken Kennedy awards. Computer, 49(12):89, December 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
co/2016/12/mco2016120089.
pdf.

Anonymous:2017:ISFi

[Ano17] Anonymous. IEEE Spectrum — front cover. IEEE Spectrum, 54(9):c1, September 2017. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2018:EGS

[Ano18]  Anonymous.  Erratum:

Glimpses of space-time beyond the singularities using supercomputers. Computing in Science and Engineering, 20(5):4, September/October 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/05/mcs2018050004.

html. See [Sin18].

Anonymous:2020:PSS

[Ano20] Anonymous. Preface: Summit and Sierra supercomputers. IBM Journal of Research and Development, 64(3/4):1– 4, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Anonymous:2022:VVSa

[Ano22] Anonymous. 2021 VGTC Visualization Service Award Loretta Auvil, National Center for Supercomputing Applications at the University of Illinois at Urbana Champaign. IEEE Transactions on Visualization and Computer Graphics, 28(1):xxvi, January 2022. CODEN ITVGEA. ISSN 1077-2626.

ANS:1992:TNG

[ANS92] ANS, editor. Proceedings of the Fifth International Topical Meeting on Reactor Thermal Hydraulics, NURETH-5: towards the next generation of nuclear power reactors, September 21–24, 1992, Little America Hotel, Salt Lake City, UT, USA, volume 146(1) of Nuclear Engineering and Design. The American Nuclear Society, LaGrange Park, IL, USA, 1992. ISBN 0-89448-178-9. ISSN 0029-5493. LCCN TK9202.I537 1992. Six volumes.

Abraham:1987:PGC

[AP87a] Santosh G. Abraham and Janak H. Patel. Parallel garbage collection on a virtual memory system. Technical Report CSRD 620; UILU-ENG-620, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 18 pp.

Allen:1987:DPF

[AP87b] Todd R. Allen and David A.

Padua. Debugging parallel Fortran on a shared memory machine. Technical Report CSRD-624, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 17 pp.

Abraham:1990:CBE

[AP90] Seth Abraham and Krishnan Padmanabhan. Constraint based evaluation of multicomputer networks. Technical Report CSRD 959, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1990. 33 pp.

Andrews:1991:AAP

[AP91] John B. Andrews and C. D.

(Constantine D.) Polychronopoulos. An analytical approach to performance/cost modeling of parallel computers. Technical Report CSRD 1110, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 27 pp.

Agrawal:1993:SIP

[AP93] Dharma P. Agrawal and Lalit M. Patnaik. Special issue on performance of supercomputers: Guest Editors’ introduction. Journal of Parallel and Distributed Computing, 19(3): 143–147, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

AbdelBaky:2012:EHP

[APK^{+}12] Moustafa AbdelBaky, Manish Parashar, Hyunjoo Kim, Kirk E. Jordan, Vipin Sachdeva, James Sexton, Hani Jamjoom, Zon-Yin Shae, Gergina Pencheva, Reza Tavakoli, and Mary F. Wheeler. Enabling highperformance computing as a service. Computer, 45(10): 72–80, October 2012. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Appleton:1995:CAS

[App95] Elaine Appleton. A crossmedia approach to saving the Chesapeake Bay. Environmental science and technology, 29(12):550A–??, December 1, 1995. CODEN ES-THAG. ISSN 0013-936X. BMS-CPSMA-NRC:1996:LSS

[App96] Board on Mathematical Sciences, Commission on Physical Sciences, Mathematics, and National Research Council Applications, editor. Large-scale structures in acoustics and electromagnetics: proceedings of a symposium held on September 26–27, 1994, in Washington, DC. National Academy Press, Washington, DC, USA, 1996. ISBN 0-309-05337-4. LCCN TA646 .L35 1996.

Araki:1991:LFC

[Ara91] K. Araki. Language features for concurrent programming. In Anonymous [Ano91q], pages 246–251. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Arabnia:1996:PDP

[Ara96] H. R. Arabnia, editor. Parallel and distributed processing techniques and applications: International conference — August 1996, Sunnyvale, CA, volume 2 of PDPTA -INTERNATIONAL CONFERENCE- 1996. CSREA, ????, 1996. ISBN 0-9648666-4-1, 0-9648666-2-5. LCCN ????

Arabnia:1997:HPC

[Ara97]  H. R. Arabnia.  High-

performance computing and applications in image processing and computer vision. Lecture Notes in Computer Science, 1336:72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aragon:2014:CIAb

[Ara14] Alejandro M. Aragón. A C++11 implementation of arbitrary-rank tensors for high-performance computing. Computer Physics Communications, 185(11): 3065–3066, November 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465514002586.

Arbeloa:1992:VFE

[Arb92] Francisco Jose Seron Arbeloa. Visualization and finite element techniques for seismic interpretation. Computers and Graphics, 16(4): 383–394, Winter 1992. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Ansaloni:1995:POQ

[ARE95] R. Ansaloni, E. Rossi, and S. Evangelisti. Porting and optimizing a quantumchemistry FCI algorithm on the Cray T3D. Lecture Notes in Computer Science, 919:488–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allsopp:2012:MDB

[ARF12] Nicholas Allsopp, Giancarlo Ruocco, and Andrea Fratalocchi. Molecular dynamics beyonds the limits: Massive scaling on 72 racks of a BlueGene/P and supercooled glass dynamics of a 1 billion particles system. Journal of Computational Physics, 231(8): 3432–3445, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S002199911200040X.

Arno:1988:IQF

[Arn88] Steven Arno. The imaginary quadratic fields of class number 4. Technical report SRC-TR-88-002, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. i + 17 pp.

Arno:1989:NPP

[Arn89] Steven Arno. A note on Perrin pseudoprimes. Technical report SRC-TR-89-014, Supercomputing Research Center: IDA, Lanham, MD, USA, September 26, 1989. 11 pp.

Arthur:1993:SPW

[Art93] Charles Arthur. Small is powerful: Will quantum electronics shrink the supercomputer? New Scientist, 140(1893):44–??, October 2, 1993. CODEN NWSCAL. ISSN 0262-4079, 0028-6664.

Arno:1992:DAN

[ARW92] Steven Arno, M. L. Robinson, and Ferrell S. Wheeler. On denominators of algebraic numbers and integer polynomials. Technical report SRC-TR-92-083, Supercomputing Research Center: IDA, Lanham, MD, USA, November 1992. 11 pp.

Altekar:1993:PTA

[ARW93a] S. D. Altekar, A. K. Ray, and B. R. Wienke. On the parallelization of a S_{n} transport algorithm on a CRAY Y MP. Parallel Computing, 19 (7):823–834, July 1993. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Arno:1993:IQF

[ARW93b] Steven Arno, M. L. Robinson, and Ferrell S. Wheeler. Imaginary quadratic fields with small odd class number. Technical report SRC-TR-93-102, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1993. 34 pp.

Anderson:1988:SST

[AS88] Edward Charles Anderson and Youcef Saad. Solving sparse triangular linear systems on parallel computers. Technical Report CSRD 794, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 6, 1988. i + 29 pp.

Amini:1993:SCA

[AS93] M. M. Amini and R. E.

Schooley. Supercomputing in corporate America: a sample survey. Information and management, 24(6):291–304, June 1993. ISSN 0378-7206.

Adeli:1998:HPC

[AS98] Hojjat Adeli and Roesdiman Soegiarso. High Performance Computing in Structural Engineering. CRC series on computer-aided engineering. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1998. ISBN 0-8493-2091-7. 249 pp. LCCN TA641 .A14 1999. US$89.95.

Almond:1999:UUA

[AS99] Jim Almond and Dave Snelling. UNICORE: uniform access to supercomputing as an element of electronic commerce. Future Generation Computer Systems, 15(5–6):539–548, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/21/17/abstract.
html.

Asanovic:1993:CAS

[Asa93a] Krste Asanovic. CNS-1 architecture specification: a connectionist network supercomputer. Technical report TR-93-94, International Computer Science Institute, Berkeley, CA, USA, April 1, 1993. iv + 57 pp.

Asanovic:1993:DCNa

[Asa93b]  Krste Asanovic.  Devel-

opment of a connectionist network supercomputer. Technical Report UCB/CSD 93/749, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, June 1993. 9 pp.

Asanovic:1998:VM

[Asa98] Krste Asanovic. Vector microprocessors. Ph.D. thesis, Computer Science Division, Univ. of California at Berkeley, Berkeley, CA, USA, May 1998. xxi + 278

                                        pp. URL http://sunsite.
                                        berkeley.edu/Dienst/UI/
                                        2.0/Describe/ncstrl.ucb/

CSD-98-1014. Also available as report no. UCB/CSD-98-1014.

Ashworth:1993:PPE

[Ash93]  M. Ashworth.  Parallel

processing in environmental modelling. In Hoffmann and Kauranne [HK93b], pages 1– 25. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Project:1998:GBC

[Asi98] Asian Technology Information Project. Global broadcast: CP-PACS: a look at Japanese supercomputing. IEEE Concurrency, 6(1): 89–91, January/March 1998. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1998/pdf/p1089.
pdf.

Abu-Sufah:1985:PPT

[ASK85] Walid Abu-Sufah and Alex Y.

Kwok. Performance prediction tools for Cedar: a multiprocessor supercomputer. ACM SIGARCH Computer Architecture News, 13(3): 406–413, June 1985. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Askew:1993:MCT

[Ask93] J. R. Askew. Monte Carlo techniques for the simulation of large reactors [invited]. In Kusters et al. [KSW93], pages 691–701. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Aslam:1991:ASD

[Asl91a] Sohail Aslam. The advanced software development and commercialization project: progress report PR-2. Technical Report CSRD 1129, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 77 pp.

Aslam:1991:ETH

[Asl91b]  Sohail Aslam.  Experi-

ments in thermal hydraulics simulation: multiprocessing COMMIX. Technical Report CSRD 1130, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 10 pp.

Abu-Sufah:1986:ERV

[ASM86] Walid Abdul-Karim Abu-Sufah and Allen Davis Malony. Experimental results for vector processing on the Alliant FX/8. Technical Report UILU-ENG-539, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 20 pp.

Apduhan:1991:EAT

[ASNT91] B. Apduhan, T. Sueyoshi, Y. Namiuchi, and T. Tezuka. Experiments and analysis toward distributed supercomputing on a distributed workstation environment. In Anonymous [Ano91q], pages 183–190. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Aspray:1993:TCC

[Asp93] William Aspray, editor. Technological competitiveness: contemporary and historical perspectives on the electrical, electronics, and computer industries: Proceedings of a conference held Oct. 10–11, 1991 at Rutgers University. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-0427-6. LCCN HD9696.C62 T39 1993.

Agrawal:1994:ERS

[ASS94] G. Agrawal, A. Sussman, and J. Saltz. Efficient runtime support for parallelizing block structured applications. In IEEE [IEE94c], pages 158–167. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Alef:1993:EPC

[ASSW93] M. Alef, W. Schmidt, D. Seldner, and T. Westermann. Electromagnetic particle-in-cell code development: Experiences with the implementation on massively-parallel systems. In Kusters et al. [KSW93], pages 124–134. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Auslander:1989:DCA

[AT89] Louis Auslander and Anna Tsao. A divide and conquer algorithm for the eigenproblem via complementary invariant subspace decomposition. Technical report SRC-TR-89-003, Supercomputing Research Center: IDA, Lanham, MD, USA, May 1989. 15 pp.

Auslander:1991:PE

[AT91] Louis Auslander and Anna Tsao. On parallelizable eigensolvers. Technical report SRC-TR-91-028, Supercomputing Research Center: IDA, Lanham, MD, USA, January 16, 1991. 12 pp.

Awaga:1993:BVC

[AT93a] Makoto Awaga and Hiromasa Takahashi. The μVP 64-bit vector coprocessor: a new implementation of high-performance numerical computation. IEEE Micro, 13(5):24–36, September/October 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Awaga:1993:MVB

[AT93b] Makoto Awaga and Hiromasa Takahashi. The μVP 64-bit vector coprocessor: a new implementation of high-performance numerical computation. IEEE Micro, 13(5):24–36, September/October 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Atapattu:1991:PPS

[Ata91] Dayanada W. Atapattu. Performance prediction of supercomputer programs. Thesis (Ph.D.), Indiana University, Bloomington, IN, USA, 1991. xii + 142 pp.

Almlof:1990:SCS

[ATL90] Jan Almlof, Donald G. Truhlar, and Terry P. Lybrand. Supercomputer chemistry: Structure, dynamics, and biochemical applications. Interdisciplinary science reviews: ISR, 15(3):252–??, September 1990. ISSN 0308-0188.

Araya:1993:DSI

[ATSA93] F. Araya, N. Terashita, T. Shimizu, and K. Asai. Development of software integration methodology for human-friendly and intelligent nuclear reactor design support system. In Kusters et al. [KSW93], pages 466– 478. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Attig:1996:QPC

[Att96] N. Attig. QCD on parallel computers at the HLRZ Supercomputing Centre. In Borcherds et al. [BBM96], pages 536–545. ISBN 83-902363-3-8. LCCN ????

ARC-FDD:1987:SAP

[AU87] Ames Research Center.Fluid Dynamics Division and United States.National Aeronautics and Space Administration and Scientific and Technical Information Branch, editors. Supercomputing in aerospace: proceedings of a conference held at NASA Ames Research Center, Moffett Field, California, March 10–12, 1987, number NASA CP-2454 in NASA Conference Publication. NASA Scientific and Technical Information Branch, Washington, DC, 1987.

Afuah:1990:ENS

[AU90] Alan N. Afuah and James M. Utterback. The emergence of a new supercomputer architecture. Sloan School of Management Working Paper WP 21-90; 90-3215BPS, The International Center for Research on the Management of Technology, Sloan School of Management, Massachusetts Institute of Technology, Cambridge, MA, USA, July 1990. 30 pp.

Afuah:1991:ENS

[AU91] Allan N. Afuah and James M. Utterback. The emergence of a new supercomputer architecture. Technological forecasting and social change, 40(4):315–328, December 1, 1991. CODEN TFSCB3. ISSN 0040-1625.

AustraliaParliament:1993:AEA

[Aus93] Australia.Parliament. The Auditor-General: efficiency audit: the purchase and use of a supercomputer ansto. Parliamentary paper; 222/1993 4004422857. Australian National Audit Office, Canberra, Australia, 1993.

ISBN 0-644-05253-8. xvii + 35 pp. LCCN ????

Appavoo:2008:PKB

[AUW08] Jonathan Appavoo, Volkmar Uhlig, and Amos Waterland. Project Kittyhawk: building a global-scale computer: Blue Gene/P as a generic computing platform. Operating Systems Review, 42(1):77–84, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Adams:2002:SCS

[AV02] Joel Adams and David Vos.

Small-college supercomputing: building a Beowulf cluster at a comprehensive college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):411–415, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Aparicio:1993:PSI

[AVS93] F. Aparicio, C. Vera, and J. L. San Roman. 93SF017 the program SINRAT III and its application for vehicle crashes. In Anonymous [Ano93-31], pages 251–256. ISBN 0-947719-62-8. LCCN ????

Arno:1991:SDR

[AW91] Steven Arno and Ferrell S.

Wheeler. Signed digit representations of minimal Hamming weight. Technical report SRC-TR-91-046, Supercomputing Research Center: IDA, Lanham, MD, USA, July 1991. 18 pp.

Andrews:1993:PSC

[AW93] Phil Andrews and Joel Welling. The Pittsburgh Supercomputing Center’s computer graphics environment. Computers and Graphics, 17 (1):5–8, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Abdelrahman:1994:DAD

[AW94] T. S. Abdelrahman and T. N. Wong. Distributed array data management on NUMA multiprocessors. In IEEE [IEE94c], pages 551–559. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Ao:2018:POH

[AYL^{+}18] Yulong Ao, Chao Yang, Fangfang Liu, Wanwang Yin, Lijuan Jiang, and Qiao Sun. Performance optimization of the HPCG benchmark on the Sunway TaihuLight Supercomputer. ACM Transactions on Architecture and Code Optimization, 15(1): 11:1–11:??, April 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Ashok:1994:ARS

[AZ94] I. Ashok and J. Zahorjan. Adhara: Runtime support for dynamic space-based applications on distributed memory MIMD multiprocessors. In IEEE [IEE94c], pages 168– 175. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Asenjo:1995:SLF

[AZ95] R. Asenjo and E. L. Zapata. Sparse LU factorization on the Cray T3D. Lecture Notes in Computer Science, 919:690–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asenjo:1999:PPL

[AZ99] R. Asenjo and E. L. Zapata. Parallel pivots LU algorithm on the Cray T3E. Lecture Notes in Computer Science, 1557:38–47, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amoretti:2013:EAC

[AZC13] Michele Amoretti, Francesco Zanichelli, and Gianni Conte. Efficient autonomic cloud computing using online discrete event simulation. Journal of Parallel and Distributed Computing, 73(6): 767–776, June 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731513000336.

Boley:1989:PIM

[B^{+}89] Daniel L. Boley et al., editors. Practical iterative methods for large scale computations: proceedings of the Minnesota Supercomputer Institute Workshop on Practical Iterative Methods for Large Scale Computations, Minneapolis, 23–25 October 1988, volume 53(1–3) of Computer physics communications ISSN: 0010-4655; v. 53, no. 1-3 (May 1989). North-Holland, Amsterdam, The Netherlands, 1989. ISBN 0-444-88023-2. LCCN QC 20 A1 C739 v.53 1989; QA297.8.M56 1988.

Bailey:1995:PPS

[B^{+}95] David H. Bailey et al., editors. Proceedings of the Seventh SIAM Conference on Parallel Processing for Scientific Computing: San Francisco, Calif., Feb 15–17, 1995. SIAM Press, Philadelphia, PA, USA, 1995. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Bernsten:1995:SNN

[BA95] K. N. Bernsten and P. Alstroem. Self-adaptive neural networks with random architecture. In Herrmann et al. [HWP95], pages 355– 360. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Bader:2008:HPC

[BA08] David A. Bader and Srinivas Aluru. High-performance computational biology. Parallel Computing, 34(11):613– 615, November 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Babaoglu:1992:PEP

[BAAD92] O. Babaoglu, L. Alvisi, A. Amoroso, and R. Davoli. Paralex: An environmental for parallel programming in distributed systems. In ACM [ACM92b], pages 178–187. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Becciani:1997:PTC

[BAAD^{+}97] U. Becciani, R. Ansaloni, V. Antonuccio-Delogu, G. Erbacci, M. Gambera, and A. Pagliaro. A parallel tree code for large N-body simulation: dynamic load balance and data distribution on a CRAY T3D system. Computer Physics Communications, 106(1–2): 105–113, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465597001021.

Baber:1990:HAD

[Bab90] Marc Baber. Hypertasking: automatic data parallel domain decomposition on the Intel parallel supercomputer. Technical report CS/E 90-006, Oregon Graduate Institute of Science and Technology, Dept. of Computer Science and Engineering, Beaverton, OR, May 1990. 16 pp.

Babcock:1994:CBS

[Bab94] Charles Babcock. Charles Babcock says that if the supercomputer makers can hang on long enough, demand for their products will go up. ComputerWorld, 28 (41):6–??, October 1994. CO-DEN CMPWAB. ISSN 0010-4841.

Bacon:1988:PSC

[Bac88] Ben Bacon. Pittsburgh Supercomputing Center to get gallium arsenide Cray 3 in 1990. Computers in Physics, 2(6):9–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

                                        //aip.scitation.org/doi/
                                        10.1063/1.4822801.

Bader:1999:ENA

[Bad99] David A. Bader. Editorial: a new, architectural paradigm for high-performance computing. Parallel and Distributed Computing Practices, 2(2):??, ???? 1999.

CODEN ???? ISSN 1097-

2803. URL http://www.cs.
okstate.edu/~pdcp/vols/
vol02/vol02no2editorial.
html.

Becciani:2001:YRF

[BAD01] U. Becciani and V. Antonuccio-Delogu. Are you ready to FLY in the universe? A multi-platform N-body tree code for parallel supercomputers. Computer Physics Communications, 136(1–2): 54–63, May 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465500002538.

Bader:2004:CBH

[Bad04] David A. Bader. Computational biology and highperformance computing. Communications of the ACM, 47 (11):34–41, November 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bader:2008:PCA

[Bad08] David A. Bader, editor.

Petascale computing: algorithms and applications. Chapman and Hall/CRC computational science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2008. ISBN 1-58488-909-

8. xlvi + 565 + 8 pp. LCCN QA76.88 .P475

2008. URL http://www.loc.

                                        gov/catdir/enhancements/
                                        fy0828/2007044024-d.html;
                                         http://www.loc.gov/catdir/
                                        toc/ecip083/2007044024.
                                        html.

Baer:2001:LEI

[Bae01] Troy Baer. Linux in education: Integrating a Linux cluster into a production high-performance computing environment. Linux Journal, 87:38, 40, 42, 44–48, July 2001. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bailey:1988:EHS

[Bai88] David H. Bailey. Extra high speed matrix multiplication on the Cray-2. SIAM Journal on Scientific and Statistical Computing, 9(3):603–607, May 1988. CODEN SIJCD4. ISSN 0196-5204.

Bailey:1992:MPS

[Bai92] David H. Bailey. Misleading performance in the supercomputing field. In Proceedings Supercomputing ’92, pages 155–158. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2630-5, 0-89791-537-2. LCCN QA76.5 S87a 1992.

Bailey:1997:TOO

[Bai97]  David H. Bailey.  Tech-

nical opinion: Onward to petaflops computing. Communications of the ACM, 40 (6):90–92, June 1997. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Baker:1993:IMC

[Bak93] R. S. Baker. Implementation of a Monte Carlo algorithm for neutron transport on a massively parallel SIMD machine. In Kusters et al. [KSW93], pages 366– 373. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Bakos:2010:HPH

[Bak10]  Jason D. Bakos.  High-

performance heterogeneous computing with the Convey HC-1. Computing in Science and Engineering, 12(6):80–87, November/ December 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Balvers:1993:FPS

[Bal93] R. P. A. Balvers. Floating point speed and accuracy on a standard automotive MCU. In Anonymous [Ano93-31], pages 907–914. ISBN 0-947719-62-8. LCCN ????

Balakrishnan:1994:CSE

[Bal94] N. Balakrishnan, editor.

Computer systems and education: International conference — June 1994, Bangalore, India. Tata McGraw-Hill, New Delhi, India, 1994. ISBN 0-07-462044-4. LCCN ????

Bic:1993:EUI

[BAM93] Lubomir Bic and Mayez Al-Mouhamed. The EM-4 under implicit parallelism. Journal of Parallel and Distributed Computing, 19(3): 255–??, November 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Bamforth:1997:JSS

[Bam97]  R. Bamforth.  Java —

from smartcard to supercomputer. In Anonymous [Ano97g], pages 1–?? ISSN 0963-3308.

Banerjee:1979:SOP

[Ban79] Utpal Banerjee. Speedup of ordinary programs. Ph.D. thesis, Dept. of Computer Science, Univ. of Illinois at Urbana-Champaign, Urbana-Champaign, October 1979.

Banerjee:1988:DAS

[Ban88]  Utpal Banerjee.  Depen-

dence analysis for supercomputing. The Kluwer international series in engineering and computer science. Parallel processing and fifth generation computing. Kluwer Academic, Boston, MA, USA, 1988. ISBN 0-89838-289-0. x + 155 pp. LCCN QA76.5 .B2641 1988.

Banerjee:1990:UTD

[Ban90] Utpal Banerjee. Unimodular transformations of double loops. Technical Report CSRD 1036, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1990. 28 pp.

Barnwell:1988:EID

[Bar88] T. P. (Thomas Pinkney)

Barnwell. Equipment for the implementation of a digital signal processing supercomputer for complex knowledgebased systems. Technical report, School of Electrical Engineering, Georgia Institute of Technology, Atlanta, GA, USA, 1988. 14 pp.

Barrett:1993:DAA

[Bar93a] P. A. Barrett. Dependable architectures for avionics applications. In Anonymous [Ano93-31], pages 503–510. ISBN 0-947719-62-8. LCCN ????

Barron:1993:LAU

[Bar93b] S. Barron. Linguistic approaches to understanding the meaning of DNA. In Lim et al. [L^{+}93], pages 33–34. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Baran:2000:NVI

[Bar00a] Nicholas Baran. News and views: Internet security in “Infantile state”; Hedy Lamarr: Spread spectrum pioneer; new weather supercomputer not so super; researchers like armies of cheap robots; Kurzweil predicts human brain uploads. Dr. Dobb’s Journal of Software Tools, 25(4):18, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVIa

[Bar00b] Nicholas Baran. News and views: Internet security in “Infantile state”; Hedy Lamarr: Spread spectrum pioneer; new weather supercomputer not so super; researchers like armies of cheap robots; Kurzweil predicts human brain uploads. Dr. Dobb’s Journal of Software Tools, 25(4):18, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVN

[Bar00c] Nicholas Baran. News and views: New modem standards should speed up Internet access; robocopter: AI lifts off; feet don’t fail me now; IBM claims world’s fastest supercomputer; new color displays based on lightemitting polymers; W3C moves forward with XLink. Dr. Dobb’s Journal of Software Tools, 25(9):18, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVNa

[Bar00d] Nicholas Baran. News and views: New modem standards should speed up Internet access; robocopter: AI lifts off; feet don’t fail me now; IBM claims world’s fastest supercomputer; new color displays based on lightemitting polymers; W3C moves forward with XLink. Dr. Dobb’s Journal of Software Tools, 25(9):18, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2001:NVW

[Bar01] Nicholas Baran. News and views: WSDL goes to W3C for standardization; shortchanging science; EUVL may keep Moore’s Law going; spy satellites to generate hightech jobs; Mexican government adopts Linux; supercomputer on a chip in the works; brain scan database goes public. Dr. Dobb’s Journal of Software Tools, 26(6): 18, June 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Bass:1995:GG

[Bas95a]  Thomas A. Bass.  Gene

genie. Wired, 3(8):114–??, August 1, 1995. CODEN WREDEM. ISSN 1059-1028 (print), 1078-3148 (electronic).

Bass:1995:GGI

[Bas95b] Thomas A. Bass. Gene genie — it’s a hundred times faster than the best serial supercomputer. it’s a trillion times denser than the best storge media. it’s a teaspoonful of DNA that’s a computer! and Leonard Adleman invented it. Wired, 3 (8):114–??, ???? 1995. CO-DEN WREDEM. ISSN 1059-1028 (print), 1078-3148 (electronic).

Bataineh:1999:PVI

[BAT99] Abdulla Bataineh, Mike Aamodt, and Kevin Thomas. A parallel and vector implementation of circuit simulation on Cray supercomputers. Parallel Algorithms and Applications, 14(2):109– 118, March 1999. CO-DEN PAAPEC. ISSN 1063-7192. URL http:

//www.informaworld.com/
smpp/content~content=a777957922.

Baucom:1988:RSP

[Bau88] Christianne Louise Baucom. Reduced systems and the preconditioned conjugate gradient method on a multiprocessor. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1988. ix + 63 pp.

Baum:1996:DBT

[Bau96] Eric Baum. DIALOG BOX — tomorrow’s supercomputer processors may be made of DNA. Windows Magazine, 7(6):57–??, ???? 1996. CODEN WINMEV. ISSN 1060-1066.

Barbieri:1987:EWI

[BB87] R. W. (Richard W.) Barbieri and N. L. Bonavito. Estimation of wind-wave interaction spectra by the Maximum Entropy method. Technical report SRC-TR-87-005, Supercomputing Research Center: IDA, Lanham, MD, USA, 1987. 24 pp.

Berrington:1990:SAM

[BB90] K. A. Berrington and P. G.

Burke. Supercomputational atomic and molecular physics. In Pitcher [Pit90], pages 159–169. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Bramley:1991:SOD

[BB91a] Randall Barry Bramley and James Bordner. Sequential optimization and data distribution for ARC2D on the Cedar hierarchical multiprocessor. Technical Report CSRD 1128, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1991. 48 pp.

Bremner:1991:TPG

[BB91b] Andrew Bremner and Duncan A. Buell. Three points of great height on elliptic curves and a computational method. Technical report SRC-TR-91-037, Supercomputing Research Center: IDA, Lanham, MD, USA, July 1991. 6 pp.

Bronson:1992:CSF

[BB92] Nathan D. Bronson and Duncan A. Buell. Congruential sieves on FPGA computers. Technical report SRC-TR-92-077, Supercomputing Research Center: IDA, Lanham, MD, USA, October 1992. 9 pp.

Bradley:1993:WTQ

[BB93] J. N. Bradley and C. M. Brislawn. Wavelet transformvector quantization compression of supercomputer ocean models. In Storer and Cohn [SC93], pages 224–233. ISBN 0-8186-3392-1, 0-8186-3391-3. ISSN 1068-0314. LCCN QA76.9.D33 D38 1993. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Borovski:1994:SRP

[BB94] B. Borovski and A. Badareen. Supercomputing with RISC processors. Comptes rendus de l’Academie bulgare des sciences: sciences mathematiques et naturelles, 47(7): 19–??, ???? 1994. ISSN 0861-1459.

Buehlmann:1998:SDI

[BB98] B. Buehlmann and H. Bieri.

Supercomputing at the desktop: An improved interface using Internet facilities. Lecture Notes in Computer Science, 1401:617–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1999:CCC

[BB99] Mark Baker and Rajkumar Buyya. Cluster computing: the commodity supercomputer. Software—Practice and Experience, 29(6): 551–576, May 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?ID=
61000492;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=61000492&
PLACEBO=IE.pdf.

Bokhari:2013:CCX

[BB13] Shahid H. Bokhari and Saniyah S. Bokhari. A comparison of the Cray XMT and XMT-2. Concurrency and Computation: Practice and Experience, 25(15):2123– 2139, October 2013. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barrett:1991:SAA

[BBB^{+}91] C. Barrett, F. Bobrowicz, R. G. Brickner, B. A.

Clark, R. Gupta, A. H. Hayes, H. Trease, and A. B. White. Supercomputing at Los Alamos National Laboratory. International Journal of Supercomputer Applications, 5(2):3–??, Summer 1991. CODEN IJSAE9. ISSN 0890-2720.

Bernholdt:2020:SMU

[BBB^{+}20a] David E. Bernholdt, Swen Boehm, George Bosilca, Manjunath Gorentla Venkata, Ryan E. Grant, Thomas Naughton, Howard P. Pritchard, Martin Schulz, and Geoffroy R. Vallee. A survey of MPI usage in the US exascale computing project. Concurrency and Computation: Practice and Experience, 32(3):e4851:1–e4851:??, February 10, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brown:2020:MEG

[BBB^{+}20b] Nick Brown, Brian Bainbridge, Ciarán Beggan, William Brown, Brian Hamilton, and Susan Macmillan. Modelling the Earth’s geomagnetic environment on Cray machines using PETSc and SLEPc. Concurrency and Computation: Practice and Experience, 32(20):e5660:1– e5660:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bambos:1996:SSS

[BBBC96] N. Bambos, J. Bannister, L. A. Bergman, and J. Cong. Supercomputer Supernet (SSN): a high-speed electro-optic campus and metropolitan network [2692-04]. In Cloonan [Clo96], pages 22–33. ISBN 0-8194-2066-2. ISSN 0361-0748. LCCN TA1632 .O672 1996.

Booth:1989:LSA

[BBC^{+}89] S. P. Booth, K. C. Bowler, D. J. Candlin, R. D. Kenway, B. J. Pendleton, A. M. Thornton, D. J. Wallace, J. Blair-Fish, and D. Roweth. Large scale applications of transputers in HEP: the Edinburgh Concurrent Supercomputer Project. Computer Physics Communications, 57(1–3):101–107, December 1, 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465589901975.

Bianchi:1992:ALS

[BBC92] Carlo Bianchi, Giuseppe Bruno, and Andrea Cividini. Analysis of large scale econometric models using supercomputer techniques. Computational Economics, 5(3): 271–??, August 1, 1992. CO-DEN CNOMEL. ISSN 0927-

Beccaria:1999:HPR

[BBC^{+}99] Matteo Beccaria, Guido Buresti, Alberto Ciampa, Giovanni Lombardi, Wolfgang Gentzsch, Hans-Georg Paap, and Andrea Viceré. Highperformance road-vehicle optimised aerodynamic design: Application of parallel computing to car design. Future Generation Computer Systems, 15(3):323–332, April 1, 1999. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/30/19/18/abstract.
html.

Brandt:2000:BGC

[BBC^{+}00] M. Brandt, J. Brooks, M. Cahir, T. Hewitt, E. Lopez-Pineda, and D. Sandness. The Benchmarker’s Guide for Cray SV1 Systems. Cray Inc., Seattle, WA, USA,

2000. URL http://www.
cray.com/craydoc/.

Brodlie:2005:SAR

[BBC^{+}05] Ken Brodlie, John Brooke, Min Chen, David Chisnall, Ade Fewings, Chris Hughes, Nigel W. John, Mark W. Jones, Mark Riding, and Nicolas Roard. State of the art reviews: Visual supercomputing: Technologies, applications and challenges. Computer Graphics Forum, 24(2):217–245, June 2005. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Billingsley:1992:SES

[BBD92] Keith R. Billingsley, Hilton U.

Brown, and Ed Derohanes. Scientific excellence in supercomputing: the 1990 IBM contest prize papers. Baldwin Press, Athens, GA, USA, 1992. ISBN ???? x + 890 pp. LCCN QA76.9.C65 C653 1992. Two volumes.

Bender:2008:CAP

[BBD^{+}08] Michael A. Bender, David P.

Bunde, Erik D. Demaine, Sándor P. Fekete, Vitus J. Leung, Henk Meijer, and Cynthia A. Phillips. Communicationaware processor allocation for supercomputers: Finding point sets of small average distance. Algorithmica, 50 (2):279–298, February 2008. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0178-4617&volume=50&
issue=2&spage=279.

Bailey:1994:NPB

[BBDS94] D. H. Bailey, E. Barszcz, L. Dagum, and H. D. Simon. NAS parallel benchmark results 3-94. In IEEE [IEE94c], pages 111–120. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Bercea:2020:OSS

[BBE^{+}20] G. T. Bercea, A. Bataev, A. E. Eichenberger, C. Bertolli, and J. K. O’Brien. An opensource solution to performance portability for Summit and Sierra supercomputers. IBM Journal of Research and Development, 64(3/4):12:1– 12:23, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Baylor:1995:PEP

[BBH95] S. J. Baylor, C. Benveniste, and Y. Hsu. Performance evaluation of a parallel I/O architecture. In Anonymous [Ano93-31], pages 404–413. ISBN 0-947719-62-8. LCCN ????

Bal:2000:DAS

[BBH^{+}00] Henri Bal, Raoul Bhoedjang, Rutger Hofman, Ceriel Jacobs, Thilo Kielmann, Jason Maassen, Rob van Nieuwpoort, John Romein, Luc Renambot, Tim Rühl, Ronald Veldema, Kees Verstoep, Aline Baggio, Gerco Ballintijn, Ihor Kuz, Guillaume Pierre, Maarten van Steen, Andy Tanenbaum, Gerben Doornbos, Desmond Germans, Hans Spoelder, Evert-Jan Baerends, Stan van Gisbergen, Hamideh Afsermanesh, Dick van Albada, Adam Belloum, David Dubbeldam, Zeger Hendrikse, Bob Hertzberger, Alfons Hoekstra, Kamil Iskra, Drona Kandhai, Dennis Koelma, Frank van der Linden, Benno Overeinder, Peter Sloot, Piero Spinnato, Dick Epema, Arjan van Gemund, Pieter Jonker, Andrei Radulescu, Cees van Reeuwijk, Henk Sips, Peter Knijnenburg, Michael Lew, Floris Sluiter, Lex Wolters, Hans Blom, Cees de Laat, and Aad van der Steen. The distributed ASCI Supercomputer project. Operating Systems Review, 34(4):76– 96, October 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bischof:2001:HTU

[BBHL01] Christian H. Bischof, H. Martin Bücker, Jörg Henrichs, and Bruno Lang. Handson training for undergraduates in high-performance computing using Java. Lecture Notes in Computer Science, 1947:306–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1947/19470306.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1947/19470306.
pdf.

Bohm:2008:FGP

[BBK^{+}08] E. Bohm, A. Bhatele, L. V.

Kalé, M. E. Tuckerman, S. Kumar, J. A. Gunnels, and G. J. Martyna. Fine-grained parallelization of the Car– Parrinello ab initio molecular dynamics method on the IBM Blue Gene/L supercomputer. IBM Journal of Research and Development, 52(1/2):159–??, January/ March 2008. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL

                                        http://www.research.ibm.
                                        com/journal/rd/521/bohm.
                                        html.

Belopolsky:1995:BPE

[BBL95] V. I. Belopolsky, B. N. Bezdenezhnykh, and O. V. Lovy. Brain potential and eye movement correlates of a fixational load under gazefree conditions. In Herrmann et al. [HWP95], pages 111–

116. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Borcherds:1996:PCJ

[BBM96] P. Borcherds, M. Bubak, and A. Maksymowicz, editors. Physics computing: Joint international conference; 8th — September 1996, Cracow, Poland, PHYSICS COMPUTING - INTERNATIONAL CONFERENCE-1996; 8th. Academic Computer Centre, CYFRONET, ????, 1996. ISBN 83-902363-3-8. LCCN ????

Boeres:2019:NAH

[BBM19] Cristina Boeres, Cristiana Bentes, and Edward D. Moreno. New advances in high-performance computing systems. Concurrency and Computation: Practice and Experience, 31(18): e5172:1–e5172:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Blue:1994:FBM

[BBS94] J. L. Blue, Isabel Beichl, and Francis Sullivan. Faster BKL Monte Carlo simulations. Technical report SRC-TR-94-125, Supercomputing Research Center: IDA, Lanham, MD, USA, September 8, 1994. 7 pp.

Brown:1990:RAO

[BBW90] A. Brown, A. Burns, and K. Wohlever. Research at the Ohio Supercomputer Center. International Journal of Supercomputer Applications, 4 (1):6–??, Spring 1990. CO-DEN IJSAE9. ISSN 0890-2720.

Brown:2019:LMR

[BBW19] Nick Brown, Michael Bareford, and Michèle Weiland. Leveraging MPI RMA to optimize halo-swapping communications in MONC on Cray machines. Concurrency and Computation: Practice and Experience, 31 (16):e5008:1–e5008:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bogoch:1990:SGP

[BBWR90] S. Bogoch, I. Bason, J. Williams, and M. Russell. Supercomputers get personal: The i860-based ComputeServer serves up power, not partitions. BYTE Magazine, 15(5):231–234, 236–237, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Bordawekar:1995:CSO

[BC95] R. Bordawekar and A. Choudhary. Communication strategies for out-of-core programs on distributed memory machines. In ACM [ACM95a], pages 395–403. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Boyle:2005:OQQ

[BCC^{+}05] P. A. Boyle, D. Chen, N. H. Christ, M. A. Clark, S. D. Cohen, C. Cristian, Z. Dong, A. Gara, B. Joó, C. Jung, C. Kim, L. A. Levkova, X. Liao, R. D. Mawhinney, S. Ohta, K. Petrov, T. Wettig, and A. Yamaguchi. Overview of the QCDSP and QCDOC computers. IBM Journal of Research and Development, 49 (2/3):351–365, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/boyle.pdf.

Basili:2008:UHP

[BCC^{+}08] Victor R. Basili, Jeffrey C.

Carver, Daniela Cruzes, Lorin M. Hochstein, Jeffrey K. Hollingsworth, Forrest Shull, and Marvin V. Zelkowitz. Understanding the high-performance-computing community: a software engineer’s perspective. IEEE Software, 25 (4):29–36, July/August 2008. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Belletti:2009:JFB

[BCC^{+}09] Francesco Belletti, Maria Cotallo, Andres Cruz, Luis Antonio Fernandez, Antonio Gordillo-Guerrero, Marco Guidetti, Andrea Maiorano, Filippo Mantovani, Enzo Marinari, Victor Martin-Mayor, Antonio Munoz-Sudupe, Denis Navarro, Giorgio Parisi, Sergio Perez-Gaviro, Mauro Rossi, Juan Jesus Ruiz-Lorenzo, Sebastiano Fabio Schifano, Daniele Sciretti, Alfonso Tarancon, Raffaele (lele) Tripiccione, Jose Luis Velasco, David Yllanes, and Gianpaolo Zanier. Janus: An FPGA-based system for high-performance scientific computing. Computing in Science and Engineering, 11(1):48–58, January/ February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Borchers:1997:SDA

[BCCG97] B. Borchers, S. H. Conrad, R. Cox, and R. J. Glass. A simulation and decision analysis approach to locating dense nonaqueous phase liquids in subsurface sediments. In Delic and Wheeler [DW97], pages 281– 286. ISBN 0-89871-378-1. LCCN ????

Blake:2005:TES

[BCCP05] R. J. Blake, P. V. Coveney, P. Clarke, and S. M. Pickles. The TeraGyroid experiment — Supercomputing 2003. Scientific Programming, 13(1):1–17, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bokhari:2014:MMM

[BÇG14] Shahid H. Bokhari, Ümit V. Çatalyürek, and Metin N. Gurcan. Massively multithreaded maxflow for image segmentation on the Cray XMT-2. Concurrency and Computation: Practice and Experience, 26(18):2836– 2855, December 25, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Blelloch:1993:IPN

[BCH^{+}93] Guy E. Blelloch, Siddhartha Chatterjee, Jonathan C. Hardwick, Jay Sipelstein, and Marco Zagha. Implementation of a portable nested data-parallel language. ACM SIGPLAN Notices, 28(7): 102–111, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2012:HPV

[BCH12] E. Wes Bethel, Hank Childs, and Charles Hansen. High Performance Visualization: Enabling Extreme-Scale Scientific Insight. Chapman and Hall/CRC Computational Science. Chapman and Hall/CRC, Boca Raton, FL, USA, 2012. ISBN 1-4398-7572-3. ???? pp. LCCN ????

Brase:2022:CHP

[BCH^{+}22] Jim Brase, Nancy Campbell, Barbara Helland, Thuc Hoang, Manish Parashar, Michael Rosenfield, James Sexton, and John Towns. The COVID-19 High-Performance Computing Consortium. Computing in Science and Engineering, 24(1):78–85, January/February 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Baillie:1994:CSM

[BCHH94] C. F. Baillie, G. Carr, L. Hart, and T. Henderson.

Comparison of shared memory and distributed memory parallelization strategies for grid-based weather forecast models. In IEEE [IEE94c], pages 560–567. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Brent:1994:IPS

[BCHJ94] R. Brent, A. Cleary, M. Hegland, and J. Jenkinson. Implementation and performance of Scalable Scientific Library subroutines on Fujitsu’s VPP500 parallelvector supercomputer. In IEEE [IEE94c], pages 526– 533. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Boyle:2013:CDI

[BCK13] P. A. Boyle, N. H. Christ, and C. Kim. Co-design of the IBM Blue Gene/Q Level 1 prefetch engine with QCD. IBM Journal of Research and Development, 57(1/2):13:1– 13:10, January–March 2013. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Berry:1991:SBC

[BCL91] Michael W. Berry, George Cybenko, and John Leonard Larson. Scientific benchmark characterizations. Technical Report CSRD 1183, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1991. 22 pp.

Barley:1990:HRS

[BCM90] J. J. Barley, M. A. Christie, and A. H. Muggeridge. High resolution simulation of multi-dimensional viscous fingering. In Pitcher [Pit90], pages 233–241. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Barton:1994:MPM

[BCM94] E. Barton, J. Cownie, and M. McLaren. Message passing on the Meiko CS-2. Parallel Computing, 20(4): 497–507, April 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bordawekar:1996:AOC

[BCR96] R. Bordawekar, A. Choudhary, and J. Ramanujam. Automatic optimization of communication in compiling outof-core stencil codes. In ACM [ACM96], pages 366– 373. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Behrendt:1993:RMI

[BCW93] F. Behrendt, C. Chevalier, and J. Warnatz. 93SC023 reaction mechanism of ignition and combustion of some compounds relevant to engine knock. In Anonymous [Ano93-31], pages 147–154. ISBN 0-947719-62-8. LCCN ????

Berghaus:2020:HTC

[BCW20] F. Berghaus, K. Casteels, and J. Weldon. High-throughput cloud computing with the cloudscheduler VM provisioning service. Computing and Software for Big Science, 4(1):??, December 2020. CO-DEN ???? ISSN 2510-2036 (print), 2510-2044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s41781-020-0036-1.

Blelloch:1995:SLR

[BCZ95] Guy E. Blelloch, Siddhartha Chatterjee, and Marco Zagha. Solving linear recurrences with loop raking. Journal of Parallel and Distributed Computing, 25(1): 91–97, February 15, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1995.1031/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1995.1031/production/pdf.

Becker:1993:PSH

[BD93a] J. C. Becker and L. Dagum. Particle simulation on heterogeneous distributed supercomputers. Concurrency: practice and experience, 5(4): 367–377, June 1993. CODEN CPEXEI. ISSN 1040-3108.

Borghi:1993:NST

[BD93b] R. Borghi and L. Delamare. 93SC025 numerical simulation of a turbulent flame propagating in a closed chamber. In Anonymous [Ano93-31], pages 163–182. ISBN 0-947719-62-8. LCCN ????

Boyd:1994:CKM

[BD94] E. L. Boyd and E. S. Davidson. Communication in the KSR1 MPP: Performance evaluation using synthetic workload experiments. In Anonymous [Ano94-134], pages 166–175. ISBN ???? LCCN ????

Bandyopadhyay:1994:SSS

[BDM94] S. Bandyopadhyay, B. Das, and A. E. Miller. Supercomputing with spin-polarized single electrons in a quantum coupled architecture. Nanotechnology (Bristol), 5(2): 113–??, ???? 1994. CODEN NNOTER. ISSN 0957-4484.

Boulet:1994:PT

[BDRR94] P. Boulet, A. Darte, T. Risset, and Y. Robert. (Pen)- ultimate tiling? In IEEE [IEE94c], pages 568–576. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Beetem:1985:GS

[BDW85] John Beetem, Monty Denneau, and Don Weingarten. The GF11 supercomputer. ACM SIGARCH Computer Architecture News, 13(3): 108–115, June 1985. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Bina:1988:FFB

[BE88] Eric Jon Bina and Perry A. Emrath. A faster fsck for BSD UNIX. Technical Report CSRD 823, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1988. 12 pp.

Blume:1992:PAP

[BE92] William Blume and R. Eigenmann. Performance analysis of parallelizing compilers on the Perfect BenchmarksTM programs. Technical Report CSRD 1218, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1992. 27 pp.

Bekakos:1993:PRR

[BE93a] M. P. Bekakos and O. B. Efremides. 93SC035 program representation of an R and F systolic algorithm. In Anonymous [Ano93-31], pages 199– 206. ISBN 0-947719-62-8. LCCN ????

Bekakos:1993:IMS

[BE93b] M. P. Bekakos and O. B.

Efremides. Implementation of a modified systolic solver on mesh architectures. In Anonymous [Ano93-31], pages 899–906. ISBN 0-947719-62-8. LCCN ????

Buell:1993:PFQ

[BE93c] Duncan A. Buell and Veikko Ennola. On a parameterized family of quadratic and cubic fields. Technical report SRC-TR-92-079, Supercomputing Research Center: IDA, Lanham, MD, USA, January 5, 1993. 17 pp.

Beasley:1990:LPC

[Bea90] J. E. Beasley. Linear programming on Cray supercomputers. OR: the journal of the Operational Research Society, 41(2):133– 139, February 1990. CO-DEN JORSDZ. ISSN 0160-5682 (print), 1476-9360 (electronic).

Becker:1989:DS

[Bec89a] Jordan Becker. Distributed supercomputing, 1989. 2 videocassettes (VHS).

Beckmann:1989:RSS

[Bec89b]  Carl J. Beckmann.  Re-

ducing synchronization and scheduling overhead in parallel loops. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1989. xii + 163 pp.

Beckmann:1990:CAS

[Bec90] Carl J. Beckmann. CARL: an architecture simulation language. Technical Report CSRD 1066, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1990. various pp.

Becker:2001:HBS

[Bec01] Don Becker. Home Beowulf systems, 2001. URL http:

//www.linuxshowcase.org/

tech.html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Bednar:1993:NSP

[Bed93] J. Bednar. New spark plugs. In Anonymous [Ano93-31], pages 303–310. ISBN 0-947719-62-8. LCCN ????

Buell:2007:GEI

[BEGGK07] Duncan Buell, Tarek El-Ghazawi, Kris Gaj, and Volodymyr Kindratenko. Guest Editors’ introduction: Highperformance reconfigurable computing. Computer, 40 (3):23–27, March 2007. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
co/2007/03/r3023.pdf.

Blume:1994:ADP

[BEH^{+}94] William Blume, Rudolf Eigenmann, Jay Hoeflinger, David Padua, Paul Petersen, Lawrence Rauchwerger, and Peng Tu. Automatic detection of parallelism: a grand challenge for high-performance computing. IEEE parallel and distributed technology: systems and applications, 2(3): 37–47, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Bernholdt:2002:CIH

[BEK02] David E. Bernholdt, Wael R.

Elwasif, and James A. Kohl. Communication infrastructure in high-performance component-based scientific computing. Lecture Notes in Computer Science, 2474: 260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.de/link/
service/series/0558/bibs/
2474/24740260.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740260.pdf.

Bell:1986:DPC

[Bel86] Alan G. Bell, editor. Digest of papers / Compcon 86, March 3–6, Spring; Thirtyfirst IEEE Computer Society International Conference, Cathedral Hill Hotel, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. CO-DEN PCICDQ. ISBN 0-8186-0692-4 (paperback), 0-8186-4692-6 (microfiche). LCCN QA75.5.C58 1986. IEEE Computer Society order number 692; IEEE catalog number 86CH2285-5.

Bell:1989:RSD

[Bel89] C. Gordon Bell. The 11 rules of supercomputer design, July 19, 1989. 1 videocassette (47 min.).

Bell:1992:PLW

[Bel92a] Trudy E. Bell. Physical limits will force 21st century supercomputer designers to make radical departures. IEEE Spectrum, 29(9): 72–75, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1992:SBT

[Bel92b] Trudy E. Bell. Supercomputers — beyond today’s supercomputers. IEEE Spectrum, 29(9):72–75, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bels:1993:CBS

[Bel93] F. Bels. Correlation between structural loading of vehicles on test tracks and during customer use by means of an online data-acquisition system and fatigue cycle matching. In Anonymous [Ano93-31], pages 249–256. ISBN 0-947719-62-8. LCCN ????

Bell:1996:VOS

[Bel96] Gordon Bell. Viewpoint:

1995 observations on supercomputing alternatives: Did the MPP bandwagon lead to a cul-de-sac? Communications of the ACM, 39(3):11–15, March 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
227235.html.

Bell:1998:SCP

[Bel98] Gordon Bell. A Seymour Cray perspective. World-Wide Web document, January 25, 1998. URL http://

research.microsoft.com/~
gbell/craytalk/.

Bell:1999:SDL

[Bel99] G. Bell. Supercomputing !D looking ahead. Lecture Notes in Computer Science, 1615:1–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bemmerl:1992:PXR

[Bem92]  Th. Bemmerl.  Paragon

XP/S. the road to teraFLOPS. In Meuer [Meu92c], pages 82–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Bennett:1990:SSb

[Ben90a] J. Bennett. Supercomputers and supercomputing. Search, 21(2):62–??, March 1990. ISSN 0004-9549.

Bennett:1990:SSa

[Ben90b] John M. (John Makepeace) Bennett. Supercomputers and supercomputing. Technical Report 735, Basser Dept. of Computer Science, University of Sydney, Sydney, NSW, Australia, 1990. ISBN 0-86758-410-6. 8 pp.

Bernhard:1982:CCS

[Ber82]  R. Bernhard.  Comput-

ers: Computing at the speed limit: Computers 1000 times faster than today’s supercomputers would benefit vital scientific applications. IEEE Spectrum, 19 (7):26–31, July 1982. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Berry:1986:PNA

[Ber86a] M. Berry. Parallel numerical algorithms on the CEDAR system. Technical Report CSRD-581, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. ii + 23 pp.

Berry:1986:PAC

[Ber86b] Michael Berry. Parallel algorithms on the CEDAR system. Technical Report CSRD 581, Center for Supercomputing Research and Development, University of Illinois, Urbana, IL, USA, 1986. ii + 23 pp.

Berry:1989:ADC

[Ber89a] Michael Berry. Algorithmic design on the CEDAR multiprocessor. Technical Report CSRD 851, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1989. 33 pp.

Berry:1989:PCB

[Ber89b] Michael Waitsel Berry. The PERFECT Club benchmarks: effective performance evaluation of supercomputers. Technical Report CSRD 827, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1989. v + 68 pp.

Berrington:1990:MHD

[Ber90a] K. A. Berrington. Multitasking the Householder diagonalization algorithm on the CRAY X-MP/48. In Evans and Wilson [EW90], chapter 13, pages 155–158. ISBN 0-306-43663-9, 1-4684-5820-5 (ebook). LCCN QA76.5 .S89437 1990.

Berry:1990:MSSb

[Ber90b] M. W. Berry. Multiprocessor sparse SVD methods for information retrieval applications. In Pitcher [Pit90], pages 133–144. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Berry:1990:MSSa

[Ber90c] Michael Waitsel Berry. Multiprocessor sparse SVD algorithms and applications. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1990. xviii + 174 pp.

Bergman:1995:GSD

[Ber95a] L. A. Bergman. Gigabit satellites in distributed supercomputing for global research. In IEEE [IEE95c], pages 71–76. ISBN 0-7803-2657-1 (hardcover), 0-8186-7029-0 (paperback), 0-7803-2658-X (microfiche). ISSN 1063-6390. LCCN QA 75.5 C58 1995. IEEE Computer Society Press order number PR07029. IEEE catalog number 95CH35737.

Bergmark:1995:OPC

[Ber95b] D. Bergmark. Optimization and parallelization of a commodity trade model for the IBM SP1/ 2 using parallel programming tools. In Anonymous [Ano93-31], pages 227–236. ISBN 0-947719-62-8. LCCN ????

Berlin:1996:MSC

[Ber96] F. Brett Berlin. In memorium: Seymour Cray, 1925– 1996. IEEE Computational Science & Engineering, 3 (4):90, 92, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Bernholdt:2007:SIC

[Ber07] David E. Bernholdt. Special issue: Component and framework technology in highperformance and scientific computing. Concurrency and Computation: Practice and Experience, 19(5):571– 572, April 10, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buzbee82a

[BEW82] B. Buzbee, R. Ewald, and J. Worlton. Japanese supercomputer technology. Science, 218(17):1189–93, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Burton:1991:HSU

[BF91] Vernon Burton and Terence Finnegan. Historians, supercomputing, and the U.S. manuscript census. Social science computer review, 9 (1):1–??, Spring 1991. CO-DEN SSCREH. ISSN 0894-4393.

Burtsev:1992:AMB

[BF92] V. S. Burtsev and V. B. Fedorov. Associative memory based on the principles of optical data processing for the new generation of supercomputers. Soviet journal of quantum electronics, 22(8): 735–??, August 1992. CO-DEN SJQEAF. ISSN 0049-1748.

Bollen:2011:HWT

[BFS11] Johan Bollen, Geoffrey Fox, and Prashant Raj Singhal. How and where the TeraGrid supercomputing infrastructure benefits science. Journal of Informetrics, 5 (1):114–121, January 2011. CODEN ???? ISSN 1751-1577 (print), 1875-5879 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1751157710000878.

Bhavsar:1991:SSJ

[BG91] Virendrakumar Chhabulal Bhavsar and Uday Govinddas Gujar, editors. Supercomputing Symposium ’91, June 3–5, 1991, Fredericton, NB, Canada: symposium proceedings [Fifth Annual Canadian Symposium on Supercomputing]. University of New Brunswick Press, Fredericton, NB, Canada, 1991. ISBN 0-920114-14-8. LCCN QA76.88.S873 1991. Fifth Annual Canadian Symposium on Supercomputing.

Bell:2002:WNH

[BG02] Gordon Bell and Jim Gray.

What’s next in high-performance computing? Communications of the ACM, 45(2):91– 95, February 2002. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burchert:2002:UAW

[BGH^{+}02] Frank Burchert, Stephan Gatzka, Christian Hochberger, Chang-Kun Lee, Ulrike Lucke, and Djamshid Tavangarian. Ubiquitous access to wide-area high-performance computing. Lecture Notes in Computer Science, 2299:209– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2299/22990209.
htm;   http://link.springer-

                                        ny.com/link/service/series/
                                        0558/papers/2299/22990209.
                                        pdf.

Bhanot:2005:OTL

[BGH^{+}05] G. Bhanot, A. Gara, P. Heidelberger, E. Lawless, J. C. Sexton, and R. Walkup. Optimizing task layout on the Blue Gene/L supercomputer. IBM Journal of Research and Development, 49 (2/3):489–500, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

                                        //www.research.ibm.com/
                                        journal/rd/492/bhanot.pdf.

Benodekar:1990:ITA

[BGIM90] R. Benodekar, A. D. Gosman, R. L. Issa, and T. Marinaccio. Industrial thermofluids analysis with STAR-CD. In Pitcher [Pit90], pages 181– 187. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Brune:1999:MCG

[BGKR99] Matthias Brune, Jörn Gehring, Axel Keller, and Alexander Reinefeld. Managing clusters of geographically distributed high-performance computers. Concurrency: practice and experience, 11(15):887–911, December 25, 1999. CO-DEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.

interscience.wiley.com/
cgi-bin/abstract/71005733/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71005733&PLACEBO=IE.
pdf.

Bertran:2011:LMD

[BGM^{+}11] Ramon Bertran, Marc Gonzàlez, Xavier Martorell, Nacho Navarro, and Eduard Ayguadé. Local memory design space exploration for high-performance computing. The Computer Journal, 54(5):786– 799, May 2011. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/content/54/5/786.full.
pdf+html.

Bachem:1996:MTS

[BGMR96] A. Bachem, C. Gawron, C. Moll, and M. Rickert. Mi-Croscopic traffic simulations of road networks using highperformance computers. Lecture Notes in Computer Science, 1067:306–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barnett:1994:ICC

[BGPS94] M. Barnett, S. Gupta, D. G.

Payne, and L. Shuler. Interprocessor collective communication library (InterCom). In IEEE [IEE94c], pages 357–

364. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Bjorklund:2019:FHF

[BGQ19] Andreas Björklund, Brajesh Gupt, and Nicolás Quesada. A faster Hafnian formula for complex matrices and its benchmarking on a supercomputer. ACM Journal of Experimental Algorithmics, 24(1):1.11:1–1.11:??, October 2019. CODEN ???? ISSN 1084-6654. URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3325111.

Brooks:1982:OCL

[BGS82] R. A. Brooks, R. P. Gabriel, and G. L. Steele, Jr. An optimizing compiler for lexically scoped LISP. ACM SIG-PLAN Notices, 17(6):261– 275, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bacon:1994:CTH

[BGS94] David F. Bacon, Susan L. Graham, and Oliver J. Sharp. Compiler transformations for high-performance computing. ACM Computing Surveys, 26 (4):345–420, December 1994. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0360-0300/
                                        197406.html.

Belcastro:2012:REA

[BGS^{+}12] Vincenzo Belcastro, Francesco Gregoretti, Velia Siciliano, Michele Santoro, Giovanni D’Angelo, Gennaro Oliva, and Diego di Bernardo. Reverse engineering and analysis of genome-wide gene regulatory networks from gene expression profiles using high-performance computing. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 9 (3):668–678, May 2012. CO-DEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Bigildeeva:1990:MSM

[BGT90] T. B. Bigil’deeva, T. V.

Ganzha, and A. A. Tret’yakov. A model and a solution method for the extremal location problem in supercomputer design. Soviet Physics—Doklady, 35(9): 779–??, September 1, 1990. CODEN SPHDA9. ISSN 0038-5689.

Boyle:1992:PFP

[BH92a] James M. Boyle and Terence J. Harmer. A practical functional program for the CRAY X-MP. Journal of Functional Programming, 2(1):81–126, January 1992. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:/

/www.cambridge.org/core/
product/64C00F14FB1E6B1A549A8F499BA0C576.

Burg:1992:ICS

[BH92b] H. C. Burg and J. Helin. 1991 International Conference on Supercomputing. Parallel Computing, 18(4): 467–472, April 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Baskett:1993:MDS

[BH93] F. Baskett and J. L. Hennessy. Microprocessors: From desktops to supercomputers. Science, 261(5123):864–??, August 1993. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Byrd:1995:MPA

[BH95] G. T. Byrd and M. A. Holliday. Multithreaded processor architectures. IEEE Spectrum, 32(8):38–46, August 1995. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Britt:2017:HPC

[BH17] Keith A. Britt and Travis S. Humble. High-performance computing with quantum processing units. ACM Journal on Emerging Technologies in Computing Systems (JETC), 13(3):39:1– 39:??, May 2017. CODEN ???? ISSN 1550-4832 (print), 1550-4840 (electronic).

Bhatkar:1994:CDA

[Bha94]  V. P. Bhatkar.  Centre

for development of advanced computing PARAM parallel supercomputer: Architecture, programming environment, and applications. In Siegel [Sie94], pages 388–389. ISBN 0-8186-5602-6, 0-8186-5601-8. ISSN 1063-7133. LCCN QA 76.58 I56 1994.

Bright:2005:BGC

[BHD^{+}05] A. A. Bright, R. A. Haring, M. B. Dombrowa, M. Ohmacht, D. Hoenicke, S. Singh, J. A. Marcella, R. F. Lembach, S. M. Douskey, M. R. Ellavsky, C. G. Zoellin, and A. Gara. Blue Gene/L compute chip: Synthesis, timing, and physical design. IBM Journal of Research and Development, 49 (2/3):277–287, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/bright.pdf.

Board:1994:SIM

[BHEG94] J. A. Board, Z. S. Hakura, W. D. Elliott, and D. C. Gray. Scalable implementations of multipole-accelerated algorithms for molecular dynamics. In IEEE [IEE94c], pages 87–94. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Bischof:1994:PPSb

[BHLST94] C. Bischof, S. Huss-Lederman, X. Sun, and A. Tsao. Parallel performance of a symmetric eigensolver based on the invariant subspace decomposition approach. In IEEE [IEE94c], pages 32–39. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Beecroft:1994:MCI

[BHM94a] J. Beecroft, M. Homewood, and M. McLaren. Meiko CS-2 interconnect Elan-Elite design. Parallel Computing, 20(10-11):1627–1638, November 1994. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Burns:1994:MIN

[BHM94b] Dan Burns, Frank Hady, and Ron Minnich. The memory integrated network interface. Technical report SRC-TR-94-135, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1994. 13 pp.

Birkeland:1998:USQ

[BHMH98] C. Birkeland, R. Holmestad, K. Marthinsen, and R. Høier. Use of supercomputers in quantitative electron diffraction. Journal of Scientific Computing, 13(1): 1–18, March 1998. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

                                        tronic). URL http://link.
                                        springer.com/content/pdf/

10.1023/A%3A1023232026276;
http://www.springerlink.
com/openurl.asp?genre=article&
issn=0885-7474&volume=13&
issue=1&spage=1-18.

Brooks:1992:NAD

[BHS92] Gary Brooks, Gilbert J.

Hansen, and Steve Simmons. A new approach to debugging optimized code. ACM SIGPLAN Notices, 27 (7):1–11, July 1992. CO-DEN SINODQ. ISBN 0-89791-475-9. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/pldi/143095/
p1-brooks/.

Bisseling:2002:FMF

[BHS^{+}02] Georg Bißeling, Hans-Christian Hoppe, Alexander Supalov, Pierre Lagier, and Jean Latour. Fujitsu MPI-2: Fast locally, reaching globally. Lecture Notes in Computer Science, 2474:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2474/24740401.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740401.pdf.

Bhuyan:1995:HPC

[Bhu95] Laxmi N. Bhuyan. Highperformance computer architecture. Future Generation Computer Systems, 11 (6):501–502, October 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bastian:1998:AMM

[BHW98] P. Bastian, W. Hackbusch, and G. Wittum. Additive and multiplicative multi-grid

— a comparison. Computing: Archiv fur informatik und numerik, 60(4):345–364, 1998. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL

http://www.springer.at/
springer.py?Page=10&Key=
362&cat=300607/tocs/springer.
py?Page=47&Key=340&cat=
3&id_abstract=3157&id_volume=
289&id_journal=8.

Boito:2018:CRP

[BIB^{+}18] Francieli Zanon Boito, Eduardo C. Inacio, Jean Luca Bez, Philippe O. A. Navaux, Mario A. R. Dantas, and Yves Denneulin. A checkpoint of research on parallel I/O for high-performance computing. ACM Computing Surveys, 51(2):23:1– 23:??, June 2018. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Bieterman:1988:PPC

[Bie88] Michael Bieterman. Parallel Programs on the Cray and on the Alliant: a Comparative Evaluation. IEEE, New York, NY, USA, 1988. ISBN 0-8186-0828-5. 82–83 pp. LCCN ???? IEEE Service Cent (cat n 88CH2539-5). Piscataway, NJ, USA.

Bina:1988:MUF

[Bin88] Eric Jon Bina. Modifications to the UNIX file system check program FSCK for quicker crash recovery. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA, August 1988. iv + 51 pp.

Bagheri:1994:CDS

[BIR94] B. Bagheri, A. Ilin, and L. Ridgway Scott. A comparison of distributed and shared memory scalable architectures. 1. KSR shared memory. In IEEE [IEE94c], pages 9–16. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Benodekar:1993:RSC

[BIRB93] R. W. Benodekar, R. I.

Issa, D. R. Robinson, and M. Bernaschi. 93SC029 running Star-CD on a cluster of IBM RS/ 6000 workstations. In Anonymous [Ano93-31], pages 283–288. ISBN 0-947719-62-8. LCCN ????

Bischof:1993:PPI

[Bis93]  Christian Bischof.  The

PRISM project: infrastructure and algorithms for parallel eigensolvers. Technical report SRC-TR-93-106, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1993. 9 pp.

Bischof:1994:CSM

[Bis94a] Christian Bischof. A case study of MPI: portable and efficient libraries. Technical report SRC-TR-94-130, Supercomputing Research Center: IDA, Lanham, MD, USA, 1994. 6 pp.

Bischof:1994:PPSa

[Bis94b] Christian Bischof. Parallel performance of a symmetric eigensolver based on the Invariant Subspace Decomposition approach. Technical report SRC-TR-94-117, Supercomputing Research Center: IDA, Lanham, MD, USA, March 1994. 8 pp.

Bischof:1994:SIS

[Bis94c] Christian Bischof. A study of the invariant subspace decomposition algorithm for banded symmetric matrices. Technical report SRC-TR-94-114, Supercomputing Research Center: IDA, Lanham, MD, USA, June 18, 1994. 6 pp.

Biswas:1994:FEE

[Bis94d] S. Biswas. A framework to explain expressibility. In Balakrishnan [Bal94], pages 290– 298. ISBN 0-07-462044-4. LCCN ????

Bideau:1996:GDM

[BISB96] D. Bideau, I. Ippolito, L. Samson, and G. G. Batrouni. Gravity driven motion of a single sphere on a rough inclined surface. In Wolf et al. [WSB96], pages 279–292. ISBN 981-02-2635-

7. LCCN ????

Bucher:1984:SVL

[BJ84] Ingrid Y. Bucher and Thomas L.

Jordan. Solving very large elliptic problems on a supercomputer with solid state disk. Journal of Computational Physics, 55(2):340– 345, August 1984. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0021999184900111.

Bennett:1993:AFS

[BJ93] S. Bennett and R. John.

The application of fuzzy set theory to vehicle brokerage. In Anonymous [Ano93-31], pages 205–210. ISBN 0-947719-62-8. LCCN ????

Beaty:1995:EAS

[BJ95] Steven J. Beaty and Gearold R.

Johnson. Effect of adding a scalar D-cache to the Cray-4 vector processor. IEEE International Conference on Algorithms and Architectures for Parallel Processing, 1:227– 230, 1995. IEEE catalog number 95TH0682-5.

Blount:1997:IAD

[BJH97] G. Blount, R. Jones, and D. Harmanto. An IDEF analysis of design for manufacture of automotive body panels. In Roller [Rol97], pages 185–192. ISBN 0-947719-88-1 (paperback). LCCN ????

Bartel:1995:SLP

[BJLW95] T. Bartel, J. E. Johannes, D. P. Lymberopoulos, and R. S. Wise. Simulation of low pressure plasma reactors on a massively parallel supercomputer using the direct simulation Monte Carlo (DSMC) method. In Anonymous [Ano95q], pages 184– 185. ISBN 1-56677-107-2. LCCN ????

Blaheta:2002:PHP

[BJS02] R. Blaheta, O. Jakl, and J. Starý. Parallel highperformance computing in geomechanics with inner/ outer iterative procedures. Lecture Notes in Computer Science, 2331:830–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2331/23310830.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2331/23310830.
pdf.

Bui:2016:ISD

[BJV^{+}16] Huy Bui, Eun-Sung Jung, Venkatram Vishwanath, Andrew Johnson, Jason Leigh, and Michael E. Papka. Improving sparse data movement performance using multiple paths on the Blue Gene/Q supercomputer. Parallel Computing, 51(??):3– 16, January 2016. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819115001167.

Bridgland:1996:STF

[BJZfDA96] Michael Bridgland, Robert Jamison, Jennifer. Zito, and Institute for Defense Analyses. The spanning trees forced by the path and the star. Technical report ????, Supercomputing Research Center : IDA, Lanham, MD, USA, May 31, 1996. 23 pp.

Baskett:1977:ECP

[BK77] F. Baskett and T. W. Keller.

An evaluation of the Cray-1 processor. In David J. Kuck, Duncan H. Lawrie, and Ahmed H. Sameh, editors, High speed computer and algorithm organization: [proceedings of the Symposium on High Speed Computer and Algorithm Organization, held at the University of Illinois April 13–15, 1977], pages 71–84. Academic Press, New York, NY, USA, 1977.

ISBN 0-12-427750-0. LCCN QA76.5 .S95 1977.

Bowler:1989:TQB

[BK89] Ken Bowler and Richard Kenway. The transputer, the quark and the black, black oil. Physics world, 2(4):28– ??, April 1, 1989. CODEN PHWOEW. ISSN 0953-8585.

Burd:1991:DSS

[BK91a] Stephen D. Burd and Suleiman K. Kassicieh. Decision support for supercomputer acquisition. Operations Research, 39(3):366–??, May 1, 1991. CODEN OPREAI. ISSN 0030-364X (print), 1526-5463 (electronic).

Burns:1991:SAO

[BK91b] Joseph A. Burns and Robert A. Klvoord. Supercomputer at observatory? A search for moons in planetary rings using the Cornell National Supercomputer Facility. Engineering: Cornell Quarterly, 25(2 / 3):7–??, Spring 1991. ISSN 0013-7871.

Barros:1993:PGS

[BK93] S. R. M. Barros and T. Kauranne. On the parallelization of global spectral Eulerian shallow-water models. In Hoffmann and Kauranne [HK93b], pages 36–43. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Baden:1995:PPP

[BK95a] Scott B. Baden and Scott R.

Kohn. Portable parallel programming of numerical problems under the LPAR system. Journal of Parallel and Distributed Computing, 27(1):38–55, May 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1995.1070/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1995.1070/production/pdf.

Borisyuk:1995:SNO

[BK95b] R. M. Borisyuk and Y. B.

Kazanovich. Synchronization of neural oscillators: Bifurcation analysis of a system with a central element. In Herrmann et al. [HWP95], pages 407–414. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Berglund:1997:MDE

[BK97] S. Berglund and J. Karlsson. A modular diesel engine toolbox for studies of charging and control system influence on emissions and performance. In Roller [Rol97], pages 281–288. ISBN 0-947719-88-1 (paperback). LCCN ????

Bader:2011:GEI

[BKK11] David A. Bader, David Kaeli, and Volodymyr Kindratenko. Guest Editor’s introduction: Special issue on high-performance computing with accelerators. IEEE Transactions on Parallel and Distributed Systems, 22(1): 3–6, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Bandman:1988:SPD

[BKM88] Olga Leonidovna Bandman, Vadim Evgenevich Kotov, and N. N. Mirenkov. Spetsializirovannye protsessory dlia vysokoproizvoditelnoi obrabotki dannykh. Izd-vo “Nauka,” Sibirskoe otd-nie, Novosibirsk, Russia, 1988. ISBN 5-02-028572-2. 204 pp. LCCN MLCS 94/16365 (Q).

Brandli:1993:EPD

[BKM93] N. Brandli, W. Kafer, and B. Malle. Enhancing product documentation with new information technology systems based on STEP. In Anonymous [Ano93-31], pages 429– 436. ISBN 0-947719-62-8. LCCN ????

Bhat:1994:RNN

[BKT94] C. L. Bhat, R. Koul, and A. K. Tickoo. Role of neural nets in sensitivity improvement of new generation gamma-ray telescopes. In Mahajan et al. [M^{+}94], pages 430–435. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Bradley:1991:FML

[BL91] David K. Bradley and John Leonard Larson. Finegrained measurements of loop performance on the Cray Y-MP. Technical Report CSRD 1120, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 11 pp.

Baker:1993:LET

[BL93] R. S. Baker and E. W.

Larsen. A “Local” exponential transform method for global variance reduction in Monte Carlo transport problems. In Kusters et al. [KSW93], pages 725– 732. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Blaskovich:1993:STS

[Bla93]  D. D. Blaskovich.  Su-

percomputing technologies for scientific productivity in earth sciences. In Zygielbaum [Zyg93], pages 27–28. ISBN 1-56396-094-X. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN QE48.8 .E27 1992. DOE-CONF-9202175.

Blackmon:1997:ACS

[Bla97] Kathryn L. Blackmon. Absorptive capacity and the survival of established firms in Schumpeterian environments the case of the supercomputer industry. Thesis (Ph.D.), University of North Carolina at Chapel Hill, Chapel Hill, NC, USA, 1997. viii, 188 pp.

Bobrowicz:1984:VMC

[BLFT84] F. W. Bobrowicz, J. E. Lynch, K. J. Fisher, and J. E. Tabor. Vectorized Monte Carlo photon transport. Parallel Computing, 1(3-4):295– 305, December 1984. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Bliss:1989:IFP

[Bli89] Brian Eugene Bliss. Instrumentation of Fortran programs for automatic roundoff error analysis and performance evaluation. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1989. xiii + 132 pp.

Bliss:1991:ISU

[Bli91] Brian Eugene Bliss. Interactive steering using the application executive. Technical Report CSRD 1149, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1991. iii + 30 pp.

Banavar:1994:TSC

[BLO94] G. Banavar, G. Lindstrom, and D. Orr. Type safe composition of object modules. In Balakrishnan [Bal94], pages 188–200. ISBN 0-07-462044-4. LCCN ????

Bauer:2021:SPL

[BLP^{+}21] Michael Bauer, Wonchan Lee, Manolis Papadakis, Marcin Zalewski, and Michael Garland. Supercomputing in Python with Legate. Computing in Science and Engineering, 23(4):73–79, July/ August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Blume:1992:SLA

[Blu92]  William Blume.  Success

and limitations in automatic parallelization of the Perfect BenchmarksTM programs. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1992. viii + 86 pp.

Bailey:2011:PTS

[BLW11] David H. Bailey, Robert F.

Lucas, and Samuel Watkins Williams, editors. Performance tuning of scientific applications, volume 11 of Chapman and Hall/CRC computational science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-1569-0 (hardback). ???? pp. LCCN Q183.9 .P47 2011.

Bossavit:1985:APT

[BM85] Alain Bossavit and Bertrand Meyer. An application of program transformation to supercomputer programming. Computer Physics Communications, 37(1–3): 27–38, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/001046558590133X.

Baudet:1987:SVS

[BM87] Gerard M. Baudet and Swaminathan Manohar. Supercomputing with VLSI: sorting. Technical report CS-87-12, Brown University, Dept. of Computer Science, Providence, RI, USA, 1987. 18 pp.

Bell:1990:SS

[BM90] C. G. Bell and G. Miranker. Shopping for supercomputers. Datamation, 36(20):76– 78, 80, October 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Baillie:1993:PWT

[BM93a] C. F. Baillie and A. E. MacDonald. Porting the well-posed topographical meteorological model to the KSR parallel supercomputer. In Hoffmann and Kauranne [HK93b], pages 26–35.

ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Borodovsky:1993:PGL

[BM93b] M. Borodovsky and J. McIninch. Prediction of gene locations using DNA Markov chain models. In Lim et al. [L^{+}93], pages 231–248. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Brackstone:1996:MMT

[BM96] M. Brackstone and M. Mc-Donald. The microscopic modelling of traffic flow: Weaknesses and potential developments. In Wolf et al. [WSB96], pages 151–166. ISBN 981-02-2635-7. LCCN ????

Barbero:1993:AFC

[BMCA93] J. I. Barbero, C. Muga, J. Calvo, and J. J. Anza. Automotive forging components design. In Anonymous [Ano93-31], pages 633–638. ISBN 0-947719-62-8. LCCN ????

Beckingsale:2020:UAF

[BMD^{+}20] D. A. Beckingsale, M. J.

McFadden, J. P. S. Dahm, R. Pankajakshan, and R. D. Hornung. Umpire: Applicationfocused management and coordination of complex hierarchical memory. IBM Journal of Research and Development, 64(3/4):00:1– 00:10, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Benoist:1993:CEI

[BMP93] P. Benoist, J. Mondot, and I. Petrovic. Calculational and experimental investigations of the void effect. A new model for leakage treatement of heterogeneous assemblies. In Kusters et al. [KSW93], pages 113–127. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Barkai:1985:STH

[BMR85] D. Barkai, K. J. M. Moriarty, and C. Rebbi. Supercomputers in theoretical high-energy physics. Computer Physics Communications, 38(1):1– 7, August/September 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0010465585900396.

Bell:1988:SO

[BMR88] C. G. Bell, G. S. Miranker, and J. J. Rubinstein. Supercomputing for one. IEEE Spectrum, 25 (4):46–50, April 1988. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baetke:1992:CAC

[BMS92] F. Baetke, B. Metzger, and P. Smith. CONVEX application compiler: a major step into the direction of automatic parallelization. In Meuer [Meu92c], pages 158–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Bhatt:1994:OIP

[BMSD94] H. S. Bhatt, Mritunjay, M. K.

Shah, and N. P. Darji. Optimisation of image processing software on VAX/VMS system configuration. In Mahajan et al. [M^{+}94], pages 293–300. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Bhatt:1994:SID

[BMSP94] H. S. Bhatt, I. C. Matieda, A. S. Shastry, and C. V. S. Prakash. Schedulers for IRS data product generation system. In Mahajan et al. [M^{+}94], pages 301–308. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Baxter:1996:EPF

[BMT96] Robert M. Baxter, Killian D. Murphy, and Shari M. Trewin. Experiences in parallelising FLITE3D on the Cray T3D. Concurrency: practice and experience, 8(10):741–755, December 1996. CO-DEN CPEXEI. ISSN 1040-3108. URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?ID=
23263.

Brennan:1991:SAS

[BMW91] Kevin F. Brennan, Nabil Mansour, and Yang Wang. Simulation of advanced semiconductor devices using supercomputers. Computer Physics Communications, 67 (1):73–92, August 1991. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0010465591902227.

Buttner:1999:APH

[BNSP99] Lars Büttner, Jörg Nolte, and Wolfgang Schröder-Preikschat. ARTS of PEACE

— a high-performance middleware layer for parallel distributed computing. Journal of Parallel and Distributed Computing, 59(2): 155–179, November 1999. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1999.1570/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1999.1570/production/pdf;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1999.1570/production/ref.

Borden:1989:GPG

[Bor89] Bruce S. Borden. Graphics processing on a graphics supercomputer. IEEE Computer Graphics and Applications, 9(4):56–62, July 1, 1989. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Borchers:1992:AMB

[Bor92] Robert R. Borchers. Albuquerque meeting brings news of supercomputer advances. Computers in Physics, 6(1): 8–??, January 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4823046.

Borchers:1993:SBH

[Bor93] Robert R. Borchers. SC’93 brings high-performance computing to Portland. Computers in Physics, 7(5):491– ??, September 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4823206.

Borchers:1994:NSF

[Bor94] R. R. Borchers. National science foundation (NSF) supercomputing metacenter. In Lundstrom [Lun94], pages 365–376. ISBN 0-8194-1681-9 (hardcover), 0-8194-1680-0 (softcover). LCCN TK5105.5.D45 1994.

Bordogna:2001:CE

[Bor01] J. Bordogna. The 21st century engineer. IEEE Spectrum, 38(1):17–19, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bataineh:1993:PVL

[BOS93] Abdulla Bataineh, Füsum Ozgüner, and Imre Szauter. Parallel and vector logic and fault simulation algorithms on the Cray Y-MP supercomputer. Simulation, 61(3):161– 169, September 1993. CO-DEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic).

Bose:1994:CCF

[Bos94a] A. K. Bose. Current challenges and future directions in electronic design automation. In Balakrishnan [Bal94], pages 221–223. ISBN 0-07-462044-4. LCCN ????

Bose:1994:PRA

[Bos94b] S. K. Bose. Parallel rendering approach for visualisation of molecules. In Mahajan et al. [M^{+}94], pages 40–?? ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Beletsky:1997:HML

[BOS97] D. Beletsky, W. P. O’Connor, and D. J. Schwab. Hydrodynamic modeling for the Lake Michigan Mass Balance Project. In Delic and Wheeler [DW97], pages 125– 128. ISBN 0-89871-378-1. LCCN ????

Botma:1996:CMI

[Bot96] H. Botma. Comparison of models for the instability of a traffic stream. In Wolf et al. [WSB96], pages 105– 118. ISBN 981-02-2635-7. LCCN ????

Bowen:1988:CMV

[Bow88] S. Bowen. Cray Y-MP — a VLSI supercomputer. IEEE, Piscataway, NJ, USA, 1988. ISBN 0-8186-0872-2. 21–23 pp. LCCN ???? Available from IEEE Service Cent (cat n 88CH2643-5). Piscataway, NJ, USA.

Boyd:2015:GSS

[Boy15]  J. Boyd.  Giving super-

computers a second wind [news]. IEEE Spectrum, 52 (6):20, June 2015. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1984:CCC

[BP84] W. K. Brown and P. R.

Pamidi. COSMIC/NASTRAN on the Cray computer systems. NASA conference publication, pages 47–53, 1984. CODEN NACPDX. ISSN 0191-7811. NTIS. Springfield, VA, USA.

Berry:1986:AES

[BP86] Michael W. Berry and Robert J. Plemmons. Algorithms and experiments for structural mechanics on high performance architectures. Technical Report CSRD-602, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 29 pp.

Beckmann:1989:EBS

[BP89a] Carl J. Beckmann and C. D. (Constantine D.) Polychronopoulos. The effect of barrier synchronization and scheduling overhead on parallel loops. Technical Report CSRD 842, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1989. 15 pp.

Brebbia:1989:ASE

[BP89b] C. A. Brebbia and A. Peters, editors. Applications of supercomputers in engineering: proceedings of the first international conference, Southampton, UK, September 1989. Elsevier, Amsterdam, The Netherlands, 1989. ISBN 0-444-88110-7, 1-85312-044-8, 0-945824-27-0. LCCN TA345 .A66 1989 [1-2] (1989). Two volumes.

Beckmann:1990:FBS

[BP90] Carl J. Beckmann and C. D. (Constantine D.) Polychronopoulos. Fast barrier synchronization hardware. Technical Report CSRD 986, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1990. 27 + [3] pp.

Beckmann:1991:BNF

[BP91a] Carl J. Beckmann and C. D. (Constantine D.) Polychronopoulos. Broadcast networks for fast synchronization. Technical Report CSRD 1070, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 21 + [15] pp.

Beckmann:1991:ESS

[BP91b] Carl J. Beckmann and C. D. (Constantine D.) Polychronopoulos. The effect of scheduling and synchronization overhead on parallel loop performance. Technical Report CSRD 1111, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 40 + [2] pp.

Beckmann:1992:MSD

[BP92] Carl J. Beckmann and C. D. (Constantine D.) Polychronopoulos. Microarchitecture support for dynamic scheduling of acyclic task graphs. Technical Report CSRD 1207, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1992. 30 + [6] pp.

Brebbia:1993:ASE

[BP93] C. A. Brebbia and H. Power, editors. Applications of supercomputers in engineering: Third International Conference on Applications of Supercomputers in Engineering, ASE/93, Conference held in Bath, UK, 27–29 September 1993. Elsevier Applied Science, London, UK, 1993. ISBN 1-85166-845-4, 1-85312-236-X, 1-56252-160-8. LCCN TA345.A66 1993.

Brilon:1996:ATF

[BP96] W. Brilon and M. Ponzlet. Application of traffic flow models. In Wolf et al. [WSB96], pages 23–40. ISBN 981-02-2635-7. LCCN ????

Buis:2006:PCF

[BPD06] Samuel Buis, Andrea Piacentini, and Damien Déclat. PALM: a computational framework for assembling high-performance computing applications. Concurrency and Computation: Practice and Experience, 18(2):231– 245, February 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bartel:1994:DMC

[BPJ94] T. J. Bartel, S. J. Plimpton, and C. R. Justiz. Direct Monte Carlo simulation of ionized rarefied flows on large MIMD parallel supercomputers. Progress in astronautics and aeronautics, 159 (??):155–??, ???? 1994. CO-DEN PAAEA9. ISSN 0079-6050.

Bricolo:1995:RIT

[BPL95] E. Bricolo, J. Pauls, and N. Logothetis. The role of inferior temporal cortex in visual object recognition. In Herrmann et al. [HWP95], pages 225–242. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Bancroft:1989:SVC

[BPM^{+}89] G. V. Bancroft, T. Plessel, F. Merritt, P. P. Walataka, and V. Watson. Scientific visualization in computational aerodynamics at NASA Ames Research Center. Computer, 22(8):89–95, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Borkowski:1994:EMV

[BPU94] W. Borkowski, E. Panczuk, and D. Uzycki. The effect of mechanical vibrations of a tracked vehicle on dynamical loadings of its power drive system. In Anonymous [Ano94-75], pages 495–502.

ISBN 0-947719-68-7. LCCN ????

Borkowski:1994:TCS

[BPUS94] W. Borkowski, L. Prochowski, D. Uzycki, and Z. Siejda. Truck carrying system and van body loadings modelling. In Anonymous [Ano94-75], pages 83– 90. ISBN 0-947719-68-7. LCCN ????

Bromnick:1997:MTD

[BPW97] P. Bromnick, R. J. Pearson, and D. E. Winterbone. Model of a turbocharged diesel engine with new intercooler model. In Roller [Rol97], pages 503– 510. ISBN 0-947719-88-1 (paperback). LCCN ????

Berger:1995:SAE

[BR95] S. B. Berger and D. J. Reis. Supercomputer algorithms for efficient linear octree encoding of three-dimensional brain images. Computer Methods and Programs in Biomedicine, 46(2):113–??, February 1, 1995. CO-DEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Braschi:1988:STS

[Bra88] Bertrand Braschi. Solving the traveling salesman problem with simulated annealing techniques on a concurrent supercomputer. Rapport de recherche RR 752-I, IMAG, Informatique et Mathématiques Appliquées de Grenoble, Université scientifique et medicale de Grenoble, Grenoble, France, November 1988. 21 pp.

Bramley:1989:BSA

[Bra89a] Randall Barry Bramley.

Block Stiefel acceleration. Technical Report CSRD 924, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1989. 18 pp.

Bramley:1989:SPP

[Bra89b] Randall Barry Bramley. On some parallel preconditioned CG schemes. Technical Report CSRD 939, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1989. 15 pp.

Bramley:1989:RPM

[Bra89c] Randall Barry Bramley. Row projection methods for linear systems. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1989. xv + 219 pp.

Bray:1989:NSM

[Bra89d] Olin H. Bray. New supercomputer market segments in manufacturing. Technical Report SAND–89-1452C; DE89-013070, Sandia National Laboratory, Albuquerque, NM, 1989. 9 pp.

Bradley:1991:SWD

[Bra91a] David K. Bradley. Supercomputer workload decomposition and analysis. Technical Report CSRD 1064, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 28 pp.

Brandt:1991:HPN

[Bra91b] L. E. Brandt. A history and prospectus for the NSF supercomputer centers. International Journal of Supercomputer Applications, 5(4): 4–9, Winter 1991. CODEN IJSAE9. ISSN 0890-2720.

Bramley:1992:OPA

[Bra92] Randall Bramley. An orthogonal projection algorithm for generalized Stokes problems. Technical Report CSRD 1190, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. 39 pp.

Brandel:1993:VCP

[Bra93] B. Brandel. VECFEM/S: a comfortable program for structural mechanics on vector processors and workstations. In Kusters et al. [KSW93], pages 786–?? ISBN 3-923704-11-9. LCCN ???? Two volumes.

Bramley:1994:BNR

[Bra94] Randall Bramley. Book news & reviews; High-Performance Computing, by Kevin Dowd. IEEE Computational Science & Engineering, 1(4):84– 85, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Brent:1987:UPS

[Bre87] Glen Alan Brent. Using program structure to achieve prefetching for cache memories. Technical Report CSRD-647, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. viii + 170 pp.

Briens:1990:CRS

[Bri90] Frederic Jean-Louis Briens.

Compositional reservoir simulation in parallel supercomputing environments. Thesis (Ph.D.), Texas A&M University, 1990. xiv + 193 pp.

Brown:2020:MLC

[BRL^{+}20] Nick Brown, Anna Roubícková, Ioanna Lampaki, Lucy Mac-Gregor, Michelle Ellis, and Paola Vera de Newton. Machine learning on Crays to optimize petrophysical workflows in oil and gas exploration. Concurrency and Computation: Practice and Experience, 32(20):e5655:1– e5655:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brobst:1986:IST

[Bro86]  Stephen A. Brobst.  In-

struction scheduling and token storage requirements in a dataflow supercomputer. Thesis (Elec. E.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1986. 5 + ii + 142 pp. Supervised by Arvind.

Brock:1991:SISa

[Bro91a] Jack L. Brock. Supercomputing in industry: statement for the record. Testimony / United States General Accounting Office GAO/T-IMTEC-91-3, United States General Accounting Office, Washington, DC, USA, March 5, 1991. 10 pp. Microfiche. Washington, DC: US GPO, 1991. 1 microfiche: negative; 11 x 15 cm. s 1991 dcu n.

Brock:1991:SISb

[Bro91b] Jack L. Brock. Supercomputing in industry: statement for the record by Jack L. Brock, Jr., Director, Government Information and Financial Management Issues, Information Management and Technology Division, before the Subcommittee on Technology and Competiveness, Committee on Science, Space, and Technology, United States House of Representatives. Technical Report GAO/T-IMTEC-91-5, U.S. General Accounting Office, Washington, DC, USA, March 7, 1991. 10 pp.

Brookshire:1991:SES

[Bro91c] Robert G. Brookshire. The Supercomputer Era / Sidney Karin and Norris Parker Smith. Social science computer review, 9(1):176–??, Spring 1991. CODEN SS-CREH. ISSN 0894-4393.

Brown:1991:EEI

[Bro91d] Larry D. Brown. Exploring the earth’s interior with the Cornell supercomputer. Engineering: Cornell Quarterly, 25(2 / 3):21–??, Spring 1991. ISSN 0013-7871.

Brown:1993:GSI

[Bro93] Timothy M. Brown, editor.

GONG 1992: seismic investigation of the sun and stars: proceedings of a conference held in Boulder, Colorado, August 11–14, 1993, volume 42 of Astronomical Society of the Pacific conference series. Astronomical Society of the Pacific, San Francisco, CA, USA, 1993.

ISBN 0-937707-61-9. LCCN QB539.O83 G66 1993.

Bromley:1996:QNG

[Bro96] B. C. Bromley. Quasirandom number generators for parallel Monte Carlo algorithms. Journal of Parallel and Distributed Computing, 38(1): 101–104, October 10, 1996. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0132/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0132/production/pdf;
                                         http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0133/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1996.0133/production/pdf.

Brooks:1997:SPE

[Bro97]  J. Brooks.  Single pro-

cessing element optimization techniques for the Cray T3D system. In Delic and Wheeler [DW97], pages 339– 352. ISBN 0-89871-378-1. LCCN ????

Brown:2000:MBP

[Bro00] Robert G. Brown. Maximizing Beowulf performance. In USENIX [USE00a], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        brownrobert.html.

Brown:2001:EHB

[Bro01] Robert G. Brown. eden:

a home Beowulf. ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-08/pdfs/brown.pdf.

Brock:2016:NFD

[Bro16] David C. Brock. The NSA’s frozen dream. IEEE Spectrum, 53(3):54–60, March 2016. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brock:2017:SCM

[Bro17a] David C. Brock. Seymour Cray: The man who brought style to supercomputers: The CDC 7600, released in 1969, featured blue-glass doors and walnut trim. IEEE Spectrum, ??(??):??, July 28, 2017. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic). URL http:/

/spectrum.ieee.org/geek-
life/history/seymour-cray-
the-man-who-brought-style-
to-supercomputers.

Brock:2017:SSC

[Bro17b] David C. Brock. Style in super-computing [past forward]. IEEE Spectrum, 54 (8):56, August 2017. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bruijs:1988:CTR

[Bru88] W. E. M. Bruijs. Computer time reduction by subcycling with an application to numerical crash analysis. In Marino [Mar88b], pages 347– 371. ISBN ???? LCCN TL240 .I528 1988.

Bruner:1990:CPS

[Bru90a] John D. Bruner. Chief: a parallel simulation environment for parallel systems. Technical Report CSRD 1050, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1990. 13 pp.

Bruner:1990:PUI

[Bru90b] John D. Bruner. PARSIM user interface reference manual. Technical Report CSRD 1002, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1990. 34 pp.

Brunet:1991:ATI

[Bru91] M.-C Brunet. An algorithm for transforming iterative numerical algorithms to adaptive equivalents (application to the Stokes problem). Technical Report CSRD 1063, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1991. 50 pp.

Berry:1987:MSS

[BS87a] Michael Berry and Ahmed Sameh. A multiprocessor scheme for the singular value decomposition. Technical Report CSRD 690, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 9 pp.

Brown:1987:HKM

[BS87b] Peter N. Brown and Youcef Saad. Hybrid Krylov methods for nonlinear system of equations. Technical Report CSRD 699, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1987. 42 + [4] pp.

Berry:1988:PAS

[BS88a] Michael Waitsel Berry and Ahmed Sameh. Parallel algorithms for the singular value and dense symmetric eigenvalue problem. Technical Report CSRD 761, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1988. 49 pp.

Bramley:1988:RPS

[BS88b] Randall Barry Bramley and Ahmed Sameh. A robust parallel solver for block tridiagonal systems. Technical Report CSRD 806, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 35 pp.

Bramley:1990:DDP

[BS90a] Randall Barry Bramley and Ahmed Sameh. Domain decomposition for parallel row projection algorithms. Technical Report CSRD 958, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1990. 13 pp.

Bramley:1990:RPM

[BS90b] Randall Barry Bramley and Ahmed Sameh. Row projection methods for large nonsymmetric linear systems. Technical Report CSRD 957, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1990. 31 pp.

Briscolini:1991:ACS

[BS91] M. Briscolini and P. Santangelo. Animation of computer simulations of twodimensional turbulence and three-dimensional flows. IBM Journal of Research and Development, 35(1/2):119–139, January/March 1991. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Bischof:1992:IUH

[BS92] C. Bischof and H. D. Simon. Implementation of the U.S. High Performance Computing and Communications Program. In Meuer [Meu92c], pages 198–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Berlin:1994:PESb

[BS94a] A. A. Berlin and R. J. Surati.

Partial evaluation for scientific computing: The supercomputer toolkit experience. In Anonymous [Ano94-100], pages 133–141. ISBN ???? LCCN ????

Berlin:1994:PESa

[BS94b] Andrew A. Berlin and Rajeev J. (Rajeev Jayantilal) Surati. Partial evaluation for scientific computing: the supercomputer toolkit experience. A.I. memo 1487, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, 1994. 10 pp.

Bhattacharya:1994:NAT

[BS94c] U. K. Bhattacharya and I. Sen Gupta. A new approach for test generation of CMOS combinational circuits. In Balakrishnan [Bal94], pages 383–384. ISBN 0-07-462044-4. LCCN ????

Biswas:1994:TML

[BS94d] C. Biswas and I. Sen Gupta. Technology mapping for lookup table based FPGAs using genetic algorithm. In Balakrishnan [Bal94], pages 236–246. ISBN 0-07-462044-4. LCCN ????

Bjoerstad:1994:UGS

[BS94e] P. E. Bjoerstad and R. Schreiber. Unstructured grids on SIMD torus machines. In IEEE [IEE94c], pages 658–665. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Blanch:1997:SIA

[BS97] J. O. Blanch and W. W. Symes. Simulation and inversion of acoustic response from metallic objects buried in attenuating sediments. In Delic and Wheeler [DW97], pages 317–328. ISBN 0-89871-378-1. LCCN ????

Bergmann:1998:HPC

[BS98a] N. W. Bergmann and P. R. Sutton. A high-performance computing module for a low Earth orbit satellite using reconfigurable logic. Lecture Notes in Computer Science, 1482:416–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bukhanovsky:1998:NSS

[BS98b] A. V. Bukhanovsky and A. M. Samsonov. Numerical simulation of strain solitons in elastic wave guides using vector and massive parallel supercomputers. Lecture Notes in Computer Science, 1401:869–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:2000:IQC

[BS00] Jon Baker and Matt Shirel.

Ab initio quantum chemistry on PC-based parallel supercomputers. Parallel Computing, 26(7–8): 1011–1024, July 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/35/21/42/29/34/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/42/29/
34/article.pdf.

Bogdanov:2001:UIP

[BS01] Alexander V. Bogdanov and Elena N. Stankova. The use of intrinsic properties of physical system for derivation of high-performance computational algorithms. Lecture Notes in Computer Science, 2110:204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2110/21100204.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2110/21100204.
                                        pdf.

Bokhari:2004:SAC

[BS04] Shahid H. Bokhari and Jon R. Sauer. Sequence alignment on the Cray MTA-2. Concurrency and Computation: Practice and Experience, 16(9):823–839, August 10, 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bernocchi:1993:NAI

[BSB93] A. Bernocchi, L. Sassi, and G. Bidini. Numerical analysis of the influence of the charge motion on the turbulence and velocity combustion in spark ignition engines. In Anonymous [Ano93-31], pages 823– 830. ISBN 0-947719-62-8. LCCN ????

Besaw:2020:CSM

[BSD^{+}20] N. Besaw, L. Scheidenbach, J. Dunham, S. Kaur, A. Ohmacht, F. Pizzano, and Y. Park. Cluster system management. IBM Journal of Research and Development, 64(3/4):7:1–7:9, May/ July 2020. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Bertran:2013:ALP

[BSJ^{+}13] R. Bertran, Y. Sugawara, H. M. Jacobson, A. Buyuktosunoglu, and P. Bose. Application-level power and performance characterization and optimization on IBM Blue Gene/Q systems. IBM Journal of Research and Development, 57(1/2):4:1–4:17, January–March 2013. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Baumann:1996:ART

[BSJW96] G. Baumann, T. Scheffler, I. M. Janosi, and D. E. Wolf. Angle of repose in a two-dimensional rotating drum model. In Wolf et al. [WSB96], pages 347– 352. ISBN 981-02-2635-7. LCCN ????

Bokma:1993:SSD

[BSKJ93] A. F. Bokma, A. Slade, S. R.

Kerridge, and K. Johnson. Strategies for the specification of distributed KBS using CONSENSUS. In Anonymous [Ano93-31], pages 543– 550. ISBN 0-947719-62-8. LCCN ????

Bischof:1994:PTT

[BSL94] C. Bischof, X. Sun, and B. Lang. Parallel tridiagonalization through two-step band reduction. In IEEE [IEE94c], pages 23–27. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Boender:1995:FIL

[BtR95] Henk Boender and H. J. J. te Riele. Factoring integers with large prime variations of the quadratic sieve. Report NM-R9513, CWI, Amsterdam, The Netherlands, July 1995. 27 pp.

Blom:1996:AVVa

[BTV96] J. G. Blom, R. A. Trompert, and J. G. Verwer. Algorithm 758: VLUGR2: a vectorizable adaptive-grid solver for PDEs in 2D. ACM Transactions on Mathematical Software, 22(3):302–328, September 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toms/1996-22-3/
p302-blom/.

Bucher:1983:CSS

[Buc83] I. Y. Bucher. The computational speed of supercomputers. In Proceedings of the 1983 ACM SIG-METRICS Conference on Measurement and Modeling of Computer Systems, August 29–31, 1983, Minneapolis, Minnesota, pages 151–165. ACM Press, New York, NY 10036, USA, 1983. ISBN 0-89791-112-1. LCCN QA76.9.E94 P48 v.11. ACM-SIGMETRICS Performance Evaluation Review, v.11, supplement.

Buell:1986:BRBa

[Bue86a] Duncan A. Buell. Book review: Supercomputers and parallel computation. Journal of the American Society for Information Science, 37 (2):91–92, March 1986. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buell:1986:ISC

[Bue86b] Duncan A. Buell. Integer squares with constant second difference. Technical report SRC-TR-86-001, Supercomputing Research Center: IDA, Lanham, MD, USA, 1986. 18 pp.

Buell:1991:BTE

[Bue91a] Duncan A. Buell. Broadcast and total exchange in supertoroidal networks. Technical report SRC-TR-91-048, Supercomputing Research Center: IDA, Lanham, MD, USA, November 6, 1991. 13 pp.

Buell:1991:SFB

[Bue91b]  Duncan A. Buell.  Su-

pertoroids, FFT butterflies, and cube-connected-cycles. Technical report SRC-TR-91-045, Supercomputing Research Center: IDA, Lanham, MD, USA, August 20, 1991. 27 pp.

Buell:1992:SS

[Bue92] Duncan A. Buell. Sorting on SPLASH 2. Technical report SRC-TR-92-078, Supercomputing Research Center: IDA, Bowie, MD, USA, September 29, 1992. 25 pp.

Bupuis:1987:SSC

[Bup87] M. Bupuis, editor. Supercomputer simulations in chemistry: proceedings of the Symposium on Supercomputer Simulations in Chemistry, held in Montreal, August 25-27, 1985, Lecture notes in chemistry; 44. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-17178-9 (paperback). LCCN QD462.A1 S98 1985.

Burkle:1988:BS

[Bür88] Martin Bürkle. Broadband services. Computer Networks and ISDN Systems, 16(1–2):135–136 (or 135– 137??), September 1988. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). Special Double Issue on RARE.

BMRC-DASETT:1991:MGW

[Bur91] Bureau of Meteorology Research Centre, Dept. of the Arts, Sport, the Environment Tourism and Territories, Melbourne, Victoria, Australia. Medium-range global weather prediction: a major scientific application of the supercomputer, 1991. 13 pp.

Burtsev:1993:PCU

[Bur93] V. S. Burtsev. Present capabilities in the use of optical information processing devices in supercomputer architecture (invited paper) [2051-02]. In Gulyaev and Pape [GP93c], pages 2–12. ISBN 0-8194-1310-0. ISSN 0361-0748. LCCN TA1630 .I568 1993.

Burkhart:1994:MPS

[Bur94a] H. Burkhart. Massively parallel systems: How to learn, how to program, and how to evaluate them. In Balakrishnan [Bal94], pages 7–15. ISBN 0-07-462044-4. LCCN ????

Burtsev:1994:AOM

[Bur94b] V. S. Burtsev. Application of optical methods of information processing in a supercomputer architecture. International journal of optoelectronics, 9(6):489–??, ???? 1994. CODEN IJOOEV. ISSN 0952-5432.

Burgess:2000:NCW

[Bur00] Mark Burgess. Needles in the Craystack: When machines get sick. ;login: the USENIX Association newsletter, 25(8): ??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/pdfs/burgess.pdf.

Burgess:2001:NCWf

[Bur01a] Mark Burgess. Needles in the Craystack: When machines get sick, epilogue. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-12/pdfs/burgess.pdf.

Burgess:2001:NCWa

[Bur01b] Mark Burgess. Needles in the Craystack: When machines get sick, part 2. ;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-02/pdfs/burgess.pdf.

Burgess:2001:NCWb

[Bur01c] Mark Burgess. Needles in the Craystack: When machines get sick, part 3. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-04/pdfs/burgess.pdf.

Burgess:2001:NCWc

[Bur01d] Mark Burgess. Needles in the Craystack: When machines get sick, part 4. ;login: the USENIX Association newsletter, 26(3):

??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/pdfs/burgess.pdf.

Burgess:2001:NCWd

[Bur01e] Mark Burgess. Needles in the Craystack: When machines get sick, part 5. ;login: the USENIX Association newsletter, 26(4): ??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/pdfs/burgess.pdf.

Burgess:2001:NCWe

[Bur01f] Mark Burgess. Needles in the Craystack: When machines get sick, part 7. ;login: the USENIX Association newsletter, 26(6):??, October 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-10/pdfs/burgess.pdf.

Butel:1992:CVC

[But92] R. Butel. Cray-2 versus CM-2 comparison using several polynomial benchmarks. Parallel Computing, 18(8): 931–945, August 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Buzbee:1984:GIS

[Buz84] B. Buzbee. Gaining insight from supercomputing. Proceedings of the IEEE, 72:19–21, 1984. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Buerger:1993:STP

[BV93] M. Buerger and H. Van Lengen. Solving techniques for the parallel finite-element method. In Anonymous [Ano93-31], pages 727–734. ISBN 0-947719-62-8. LCCN ????

Blom:1996:AVVb

[BV96] J. G. Blom and J. G. Verwer. Algorithm 759: VLUGR3: a vectorizable adaptive-grid solver for PDEs in 3D — part II. code description. ACM Transactions on Mathematical Software, 22(3):329– 347, September 1996. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toms/1996-22-3/
p329-blom/.

Bethel:2011:VSC

[BvRS^{+}11] E. Wes Bethel, John van Rosendale, Dale Southard, Kelly Gaither, Hank Childs, Eric Brugger, and Sean Ahern. Visualization at supercomputing centers: The tale of little big iron and the three skinny guys. IEEE Computer Graphics and Applications, 31(1):90–95, January/February 2011. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Buell:1988:MIA

[BW88] Duncan A. Buell and Robert L.

Ward. A multiprecise integer arithmetic package. Technical report SRC-TR-88-019, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 38 pp.

Bik:1994:NSA

[BW94] A. J. C. Bik and H. A. G.

Wijshoff. Nonzero structure analysis. In Anonymous [Ano94-134], pages 226–235. ISBN ???? LCCN ????

Burgee:1994:PMV

[BWGG94] S. L. Burgee, L. T. Watson, A. A. Giunta, and B. Grossman. Parallel multipoint variable-complexity approximations for multidisciplinary optimization. In IEEE [IEE94c], pages 734– 740. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Brown:2018:SDA

[BWHS18] Nick Brown, Michèle Weiland, Adrian Hill, and Ben Shipway. In situ data analytics for highly scalable cloud modelling on Cray machines. Concurrency and Computation: Practice and Experience, 30(1):??, January 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Barrett:1996:ELC

[BWO96] C. L. Barrett, M. Wolinsky, and M. W. Olesen. Emergent local control properties in particle hopping traffic simulations. In Wolf et al. [WSB96], pages 169–

174. ISBN 981-02-2635-7. LCCN ????

Basu:2017:CBC

[BWV^{+}17] Protonu Basu, Samuel Williams, Brian Van Straalen, Leonid Oliker, Phillip Colella, and Mary Hall. Compiler-based code generation and autotuning for geometric multigrid on GPU-accelerated supercomputers. Parallel Computing, 64(??):50–64, May 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819117300376.

Buell:1988:SLP

[BY88] Duncan A. Buell and Jeff Young. Some large primes and the Sierpinski problem. Technical report SRC-TR-88-004, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 6 pp.

Biham:1996:DPT

[BY96] O. Biham and N. Yoran. Dynamical phase transitions in two dimensional traffic models. In Wolf et al. [WSB96], pages 229–238. ISBN 981-02-2635-7. LCCN ????

Benchara:2021:NSD

[BY21] Fatéma Zahra Benchara and Mohamed Youssfi. A new scalable distributed kmeans algorithm based on cloud micro-services for highperformance computing. Parallel Computing, 101(??): Article 102736, April 2021. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819120301186.

Cisneros:1997:CCC

[C^{+}97] Gerardo Cisneros et al., editors. Computational chemistry and chemical engineering: proceedings of the Third UNAM-CRAY Supercomputing Conference: Universidad Nacional Autonoma de Mexico, 13-16 August 1996. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1997. ISBN 981-02-3220-9. LCCN QD39.3.E46U53 1996.

Carbonnell:1993:ULT

[CAB93] F. Carbonnell, N. Atassi, and J. Boree. Unsteady laminar and turbulent jets: a numerical and experimental study. In Anonymous [Ano93-31], pages 759–762. ISBN 0-947719-62-8. LCCN ????

Calahan:1981:PLA

[Cal81] Donald A. Calahan. Performance of linear algebra codes on the Cray-1. Society of Petroleum Engineers journal, 21(5):558–564, October 1981. CODEN SSPJDN. ISSN 0197-7520.

Calahan:1985:ASC

[Cal85a] D. A. Calahan. Analysis and Simulation of the Cray X-MP Memory System. IEEE, New York, NY, USA, 1985. ISBN 0-8186-0654-1. 568–574 pp. LCCN ???? IEEE Service Cent. Piscataway, NJ, USA.

Calahan:1985:TGS

[Cal85b] D. A. Calahan. Task granularity studies on a manyprocessor Cray X-MP. Parallel Computing, 2(2):109– 118, June 1985. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Calahan:1986:BLL

[Cal86a]  D. A. Calahan.  Block-

oriented, local-memory-based linear equation solution on the Cray-2: Uniprocessor algorithms. Proceedings of the International Conference on Parallel Processing, pages 375–378, 1986. CODEN PC-PADL. ISBN 0-8186-0724-6. ISSN 0190-3918. IEEE Service Cent. Piscataway, NJ, USA.

Calahan:1986:BOL

[Cal86b]  D. A. Calahan.  Block-

oriented, local-memory-based linear equation solution on the Cray-2: Uniprocessor algorithms. Proceedings of the International Conference on Parallel Processing, pages 375–378, 1986. CODEN PC-PADL. ISBN 0-8186-0724-6. ISSN 0190-3918. IEEE Service Cent. Piscataway, NJ, USA.

Calahan:1988:CMC

[Cal88] D. A. Calahan. Characterization of memory conflict loading on the Cray-2. Proceedings of the International Conference on Parallel Processing, 1:299–302, 1988. CODEN PCPADL. ISSN 0190-3918. Available from IEEE Service Cent (cat n 88CH2625-2). Piscataway, NJ, USA.

CADOCEOR:1991:SRS

[Cal91] California.Dept.of Commerce.Office of Economic Research. The State role in supercomputing. Sunnyvale, CA., June 24, 1991. 42 pp.

Calvin:1996:IPF

[Cal96] C. Calvin. Implementation of parallel FFT algorithms on distributed memory machines with a minimum overhead of communication. Parallel Computing, 22(9):1255– 1279, November 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Calamia:2011:CRI

[Cal11a] J. Calamia. China rising: international patent applications [the data]. IEEE Spectrum, 48(7):68, July 2011. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Calamia:2011:CGG

[Cal11b] J. Calamia. China’s Godson gamble. IEEE Spectrum, 48 (5):14–16, May 2011. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Calamia:2012:CHS

[Cal12] J. Calamia. China’s homegrown supercomputers. IEEE Spectrum, 49(1):60–62, January 2012. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cann:1992:RFD

[Can92] David Cann. Retire Fortran? A debate rekindled. Communications of the ACM, 35(8):81–89, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        135231.html.

Cap:1996:LNW

[Cap96] C. H. Cap. Large networks of workstations: the mythical poor man’s supercomputer? In Dekker et al. [DSZ96], pages 3–4. ISBN 0-444-82559-2. LCCN ????

Carlson:1988:CCP

[Car88] William W. Carlson. A C compiler and post-packer for horizon. Technical report SRC-TR-88-018, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 27 pp.

Carey:1989:PSM

[Car89a] Graham F. Carey. Parallel supercomputing: methods, algorithms, and applications. Wiley series in parallel computing. John Wiley and Sons, Inc., New York, NY, USA, 1989. ISBN 0-471-92436-9. x + 287 pp. LCCN QA76.5 .P314951 1989. US$58.55.

Carr:1989:SSG

[Car89b] J. A. Carr. SCRI’s semicomplete guide to UNIX on the ETA10 supercomputer. Technical Report FSU-SCRI-89-28, Florida State University, Tallahassee, FL, USA, February 13, 1989. 64 pp.

Carlson:1991:UHP

[Car91] David A. Carlson. Ultra high performance FFTs for the CRAY-2 and CRAY-YMP supercomputers. Technical report SRC-TR-91-053, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1991. 14 pp.

Carlson:1992:RSS

[Car92] William W. Carlson. RES: a simple system for distributed computing. Technical report SRC-TR-92-067, Supercomputing Research Center: IDA, Lanham, MD, USA, May 28, 1992. 43 pp.

Carlile:1993:ADC

[Car93]  Bradley R. Carlile.  Al-

gorithms and design: the Cray APP shared-memory system. In 1993 IEEE Compcon Spring (Feb 22–26 1993: San Francisco, CA, USA), pages 312–320. IEEE, Piscataway, NJ, USA, 1993. ISBN 0-7803-1294-5. LCCN ???? IEEE catalog number 93CH3251-6.

Carlson:1994:CRJ

[Car94a] D. Carlson. Cray Research J-90 computer system aircooled supercomputer packaging. In Anonymous [Ano94-108], pages 708–724. ISBN 1-880433-16-8. LCCN ????

Carlson:1994:CAF

[Car94b] David A. Carlson. Cross ambiguity functions on the MasPar MP-2. Technical report SRC-TR-94-131, Supercomputing Research Center: IDA, Lanham, MD, USA, October 1994. 10 pp.

Cass:2001:SNS

[Cas01] Stephen Cass. Super nets for supercomputers. IEEE Spectrum, 38(1):116–118, January 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2003:SS

[Cas03] Stephen Cass. Supercheap supercomputer. IEEE Spectrum, 40(7):17, July 2003. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Catlett:1992:SBR

[Cat92a] C. E. Catlett. Supercomputing — balancing resources. IEEE Spectrum, 29(9):48– 55, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Catlett:1992:GMF

[Cat92b]  Charles E. Catlett.  Gi-

ant memories and fast networks are cornerstones of the emerging US computing initiative. IEEE Spectrum, 29(9):48, 53–55, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cybenko:1989:HHA

[CB89] George Cybenko and Michael Berry. Hyperbolic Householder algorithms for factoring structured matrices. Technical Report CSRD 877, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1989. 29 pp.

Chalasani:1994:FWR

[CB94] S. Chalasani and R. V. Boppana. Fault-tolerant wormhole routing in tori. In Anonymous [Ano94-134], pages 146–155. ISBN ???? LCCN ????

Cisneros:1999:HPC

[CB99] Gerardo Cisneros and Jeff P. Brooks. High-performance computing at Silicon Graphics/Cray Research. Applied Numerical Mathematics: Transactions of IMACS, 30(1):125–135, May 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/apnum/sub/1999/
30/1/954.pdf.

Cirne:2000:ASP

[CB00] Walfredo Cirne and Francine Berman. Adaptive selection of partition size for supercomputer requests. Lecture Notes in Computer Science, 1911:187–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1911/19110187.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1911/19110187.
pdf.

Cirne:2002:UMI

[CB02] Walfredo Cirne and Francine Berman. Using moldability to improve the performance of supercomputer jobs. Journal of Parallel and Distributed Computing, 62(10): 1571–1601, October 1, 2002. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2002.1869;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2002.1869/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2002.1869/ref.

Chang:1990:SSI

[CBA90] Long C. (Long Chyr) Chang, Wolfgang W. Bein, and Edward Angel. Supercomputing surface intersections using subdivision. Technical report CS90-3, Dept. of Computer Science, College of Engineering, University of New Mexico, Albuquerque, NM, USA, July 20, 1990. 8 + 5 pp.

Chatterjee:2005:DEH

[CBB^{+}05] S. Chatterjee, L. R. Bachega, P. Bergner, K. A. Dockser, J. A. Gunnels, M. Gupta, F. G. Gustavson, C. A. Lapkowski, G. K. Liu, M. Mendell, R. Nair, C. D. Wait, T. J. C. Ward, and P. Wu. Design and exploitation of a high-performance SIMD floating-point unit for Blue Gene/L. IBM Journal of Research and Development, 49(2/3):377–391, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/chatterjee.
pdf.

Coteus:2005:PBG

[CBC^{+}05] P. Coteus, H. R. Bickford, T. M. Cipolla, P. G. Crumley, A. Gara, S. A. Hall, G. V. Kopcsay, A. P. Lanzetta, L. S. Mok, R. Rand, R. Swetz, T. Takken, P. La Rocca, C. Marroquin, P. R. Germann, and M. J. Jeanson. Packaging the Blue Gene/L supercomputer. IBM Journal of Research and Development, 49(2/3):213– 248, ???? 2005. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/coteus.pdf.

Chen:1993:PIC

[CBCH93] C. Y. Roger Chen, P. Bruce Berra, Alok N. (Alok Nidhi) Choudhary, and Salim Hariri, editors. Proceedings of the 1993 International Conference on Parallel Processing, August 16-20, 1993, Saint Charles, IL, USA. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1993. ISBN 0-8493-8983-6 (set), 0-8493-8984-4 (vol. 1), 0-8493-8985-2 (vol. 2), 0-8493-8986-0 (vol. 3). ISSN 0190-3918. LCCN QA76.58 .I55 1993 v.1-3 (c1993).

Carey:1992:VSA

[CBCJ92] G. F. Carey, E. Barragy, D. D. Cline, and W. Joubert. Vector-parallel supercomputing algorithms and applications. In Noye et al. [NBC92], pages 17–28. ISBN 0-86396-172-X. LCCN ????

Cybenko:1991:PCPb

[CBHS91] G. Cybenko, J. Bruner, S. Ho, and S. Sharma. Parallel computing and the Perfect Benchmarks. In Anonymous [Ano91q], pages 99– 110. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Cappos:2009:SPE

[CBKA09] Justin Cappos, Ivan Beschastnikh, Arvind Krishnamurthy, and Tom Anderson. Seattle: a platform for educational cloud computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1): 111–115, March 2009. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Costa:2013:AIE

[CBLS13] Rostand Costa, Francisco Brasileiro, Guido Lemos, and Dênio Sousa. Analyzing the impact of elasticity on the profit of cloud computing providers. Future Generation Computer Systems, 29(7): 1777–1785, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X13000058.

Costa:2005:AWT

[CBM^{+}05] A. Costa, U. Becciani, P. Miocchi, V. Antonuccio, R. Capuzzo Dolcetta, P. Di Matteo, and V. Rosato. Astrocomp: Web technologies for high performance computing on a network of supercomputers. Computer Physics Communications, 166(1):17–25, February 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465504005041.

Crouch:1991:FDT

[CBT91] P. E. Crouch, E. Brady, and D. J. Tylavsky. Frequency domain transient stability simulation of power systems: Implementation by supercomputer. IEEE transactions on power systems: a publication of the Power Engineering Society, 6(1):51–??, February 1, 1991. ISSN 0885-8950 (print), 1558-0679 (electronic).

Carlson:1988:FFT

[CC88a] David A. Carlson and John Michael Conroy. The fast Fourier transform and sparse matrix computations: a study of two applications on the Horizon supercomputer. Technical report SRC-TR-88-007, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 10 pp.

Clementi:1988:BAI

[CC88b] Enrico Clementi and S. (Steven)

Chin. Biological and artificial intelligence systems. ES-COM, Leiden, 1988. ISBN 90-72199-02-2. xxiv + 584 pp. LCCN ????

Chahande:1994:MAO

[CC94a] A. Chahande and P. Chahande. Module assignment to optimize the performance of parallel/distributed computing systems. In Mahajan et al. [M^{+}94], pages 416–426. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Chen:1994:CLE

[CC94b] D. Y. Chen and M. C. Chen.

CPAR — language extensions to C for irregular and adaptive parallel computations. In IEEE [IEE94c], pages 501–508. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Calvin:1996:PEM

[CC96] C. Calvin and L. Colombet. Performance evaluation and modeling of collective communications on Cray T3D. Parallel Computing, 22(10):1413–1427, December 15, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
parco/cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=10&aid=1097.

Clementi:1989:SSS

[CCC^{+}89] E. Clementi, S. Chin, G. Corongiu, J. H. Detrich, M. Dupuis, D. Folsom, G. C. Lie, D. Logan, and V. Sonnad. Supercomputing and supercomputers for science and engineering in general and for chemistry and biosciences in particular. International Journal of Quantum Chemistry, 35(1):3–89, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Carnes:2013:SLI

[CCD^{+}13] B. Carnes, B. Chan, E. W. Draeger, J.-L. Fattebert, L. Fried, J. Glosli, W. D. Krauss, S. H. Langer, R. Mc-Callen, A. A. Mirin, F. Najjar, A. L. Nichols, T. Oppelstrup, J. A. Rathkopf, D. Richards, F. Streitz, P. M. Vranas, J. J. Rice, J. A. Gunnels, V. Gurev, C. Kim, J. Magerlein, M. Reumann, and H.-F. Wen. Science at LLNL with IBM Blue Gene/ Q. IBM Journal of Research and Development, 57 (1/2):11:1–11:18, January– March 2013. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Cave:2017:TGH

[CCG^{+}17] Vincent Cavé, Romain Clédat, Paul Griffin, Ankit More, Bala Seshasayee, Shekhar Borkar, Sanjay Chatterjee, Dave Dunning, and Joshua Fryman. Traleika Glacier: a hardware–software co-designed approach to exascale computing. Parallel Computing, 64(??): 33–49, May 2017. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819117300169.

Ceruzzi:1990:RCB

[CCKSS90] Paul Ceruzzi, Martin Campbell-Kelly, Charles Susskind, and K. W. Smillie. Reviews: Ceruzzi: Beyond the Limits: Flight Enters the Computer Age; Staudenmaier: Technology’s Storytellers: Reweaving the Human Fabric; Bijker, Hughes, and Pinch: The Social Construction of Technological Systems: New Directions in the Sociology and History of Technology; Van Creveld, Martin. Technology and War: From 2000 B.C. to the Present; Rifkin and Harrar: The Ultimate Entrepreneur: The Story of Ken Olsen and Digital Equipment Corporation; Time-Life Books, eds. Understanding Computers Vol. 24: Illustrated Chronology and Index; Lee: Winning with People: The First 40 Years of Tektronix; Searle: The Bombsight War: Norden vs. Sperry; Kruh: Tales of Yardley: Some Sidelights to his Career; IEEE Scientific Supercomputer Subcommittee. The Computer Spectrum: A perspective on the Evolution of Computing; Forester. High-Tech Society; The Computer: A Hawkhill Learning Power Book; Molina: The Social Basis of the Microelectronics Revolution. Annals of the History of Computing, 12(2):147–152, April/June 1990. CODEN AHCOE5. ISSN 0164-1239. URL http:

//dlib.computer.org/an/
books/an1990/pdf/a2147b.
pdf;   http://www.computer.
org/annals/an1990/a2147babs.
htm.

Cappello:2011:PMV

[CCR11] Franck Cappello, Henri Casanova, and Yves Robert. Preventive migration vs. preventive checkpointing for extreme scale supercomputers. Parallel Processing Letters, 21(2):111–132, June 2011. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Clematis:1997:DNI

[CCSM97] A. Clematis, A. Coda, M. Spagnuolo, and M. Mineter. Developing non-local iterative parallel algorithms for GIS on Cray T3D using MPI. Lecture Notes in Computer Science, 1332:435–442, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chase:1992:CRS

[CCSR92] C. Chase, K. Crowley, J. Saltz, and A. Reeves. Compiler and runtime support for irregularly coupled regular meshes. In ACM [ACM92b], pages 438–446. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Clematis:1998:PFQ

[CCSS98] A. Clematis, A. Coda, M. Spagnuolo, and S. Spinello. Parallelising fuzzy queries for spatial data modelling on a Cray T3D. Lecture Notes in Computer Science, 1541:76–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhry:2005:HPT

[CCYT05] Shailender Chaudhry, Paul Caprioli, Sherman Yip, and Marc Tremblay. Highperformance throughput computing. IEEE Micro, 25(3): 32–45, May/June 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cagan:1993:CNA

[CCZ93] Leigh D. Cagan, Nicholas J. Carriero, and Stavros A. Zenios. A computer network approach to pricing Mortgage-Backed securities. Financial analysts journal, 49 (2):55–??, March 1, 1993. CODEN FIAJA4. ISSN 0015-198X.

Cheng:1992:SPS

[CD92] S. P. Cheng and S. Dandamudi. Scheduling in parallel systems with a hierarchical organization of tasks. In ACM [ACM92b], pages 377–386. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Carlson:1995:AT

[CD95a] William W. Carlson and Jesse M. Draper. AC for the T3D. Technical report SRC-TR-95-141, Supercomputing Research Center: IDA, Lanham, MD, USA, February 23, 1995. 25 pp.

Carlson:1995:DDA

[CD95b] William W. Carlson and Jesse M. Draper. Distributed data access in AC. ACM SIGPLAN Notices, 30(8):39– 47, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2008:TMS

[CD08] Yongzhi Chen and Yuefan Deng. Task mapping on supercomputers with cellular networks. Computer Physics Communications, 179(7): 479–485, October 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465508001513.

Chen:2009:DAC

[CD09] Yongzhi Chen and Yuefan Deng. A detailed analysis of communication load balance on BlueGene supercomputer. Computer Physics Communications, 180(8): 1251–1258, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465509000460.

Chandra:1994:EBS

[CDA94] Y. Chandra, U. C. Durgapal, and R. P. Arora. Extraction of background in SAR images. In Mahajan et al. [M^{+}94], pages 315–320. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Clementi:1987:LSC

[CDC^{+}87] E. Clementi, J. Detrich, S. Chin, G. Corongiu, D. Folsom, D. Logan, R. Caltabiano, A. Carnevali, J. Helin, M. Russo, A. Gnudi, and P. Palamidese. Large-scale computations on a scalar, vector and parallel “Supercomputer”. Parallel Computing, 5(1–2):13–44, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cohen:2006:SPG

[CDG^{+}06] Albert Cohen, Sébastien Donadio, Maria-Jesus Garzaran, Christoph Herrmann, Oleg Kiselyov, and David Padua. In search of a program generator to implement generic transformations for highperformance computing. Science of Computer Programming, 62(1):25–46, September 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chen:1984:MLA

[CDH84] Steven S. Chen, Jack J. Dongarra, and Christopher C. Hsiung. Multiprocessing linear algebra algorithms on the Cray X-MP-2: Experiences with small granularity. Journal of Parallel and Distributed Computing, 1(1): 22–31, August 1984. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chang:1994:APG

[CDMW94] C.-S. Chang, G. DeTitta, R. Miller, and C. Weeks. On the application of parallel genetic algorithms in X-Ray crystallography. In IEEE [IEE94c], pages 796– 802. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

CS-CSPUP:1990:CSI

[CDO90] College of Science, California State Polytechnic University, Pomona, Digital Equipment Corporation, and Oak Ridge National Laboratory, editors. Computational science in industry and the comprehensive university. California State Polytechnic University, Pomona, Pomona, CA, 1990.

Choi:1994:CNS

[CDPW94] J. Choi, J. J. Dongarra, R. Pozo, and D. W. Walker. Constructing numerical software libraries for highperformance computing environments. Lecture Notes in Computer Science, 879: 147–168, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colland:1996:NGH

[CDR96] P.-Y. Colland, A. Darte, and Y. Robert. A new guaranteed heuristic for the software pipelining problem. In ACM [ACM96], pages 261– 269. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Cliffe:1998:PIF

[CDS98] K. A. Cliffe, I. S. Duff, and J. A. Scott. Performance issues for frontal schemes on a cache-based high-performance computer. International Journal for Numerical Methods in Engineering, 42(1):127–143, 1998. CODEN IJNMBH. ISSN 0029-5981.

Choi:1994:PSP

[CDW94] J. Choi, J. J. Dongarra, and D. W. Walker. PB-BLAS: a set of parallel block basic linear algebra subprograms. In IEEE [IEE94c], pages 534– 541. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Cruz-Enriquez:2018:MSA

[CE18] Hector Cruz-Enriquez. Making supercomputing available to all Cuban researchers. Computing in Science and Engineering, 20(3):25–30, ???? 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8358046/.

Chen:2012:IBG

[CEH^{+}12] Dong Chen, Noel A. Eisley, Philip Heidelberger, Robert M. Senger, Yutaka Sugawara, Sameer Kumar, Valentina Salapura, David Satterfield, Burkhard Steinmacher-Burow, and Jeffrey Parker. The IBM Blue Gene/Q interconnection fabric. IEEE Micro, 32(1):32–43, January/February 2012. CO-DEN IAHCEX. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ceramalus:1995:HT

[Cer95] Nobilangelo Ceramalus. Horizons technology. Management, 42(4):30–??, May 1, 1995. ISSN 0025-1658.

Chamberlain:1992:UIN

[CF92] Robert F. Chamberlain and Charles M. (Charles Michael) Fiduccia. Universality of iterated networks. Technical report SRC-TR-92-084, Supercomputing Research Center: IDA, Lanham, MD, USA, November 1992. 64 pp.

Corbett:1994:DIV

[CF94] P. F. Corbett and D. G. Feitelson. Design and implementation of the Vesta parallel file system. In IEEE [IEE94c], pages 63–70. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Feng:2003:MCE

[cF03] Wu chun Feng. Making a case for efficient supercomputing. ACM Queue: Tomorrow’s Computing Today, 1(7): 54–64, October 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Cerin:2012:DGC

[CF12] Christophe Cérin and Gilles Fedak, editors. Desktop grid computing. Chapman and Hall/CRC numerical analysis and scientific computing. Taylor and Francis, London, UK and Boca Raton, FL, USA, 2012. ISBN 1-4398-6214-1 (hardback). ???? pp. LCCN QA76.9.C58 D48 2012.

Feng:2007:GLE

[cFC07] Wu chun Feng and Kirk Cameron. The Green500 list: Encouraging sustainable supercomputing. Computer, 40 (12):50–55, December 2007. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Csikor:2001:PPP

[CFH^{+}01] F. Csikor, Z. Fodor, P. Hegedüs, V. K. Horváth, S. D. Katz, and A. Piróth. The PMS project: Poor man’s supercomputer. Computer Physics Communications, 134(2): 139–149, February 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S001046550000182X.

Feng:2007:HPC

[cFM07] Wu chun Feng and Dinesh Manocha. High-performance computing using accelerators. Parallel Computing, 33(10– 11):645–647, November 2007. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Coletti:2020:TDL

[CFP20] M. Coletti, A. Fafard, and D. Page. Troubleshooting deep-learner training data problems using an evolutionary algorithm on Summit. IBM Journal of Research and Development, 64(3/4):1– 12, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Choudhary:1995:GEI

[CFS95] Alok Choudhary, Ian Foster, and Rick Stevens. Guest Editors’ introduction: Multimedia applications and high-performance computing. IEEE parallel and distributed technology: systems and applications, 3(2): 2–3, Summer 1995. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1995/pdf/h20002.
pdf.

Catlow:1990:MDS

[CFV^{+}90] C. R. A. Catlow, C. M. Freeman, B. Vessal, S. M. Tomlinson, and M. Leslie. Molecular dynamics studies of hydrocarbon diffusion in zeolites. In Pitcher [Pit90], pages 385–393. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Charlesworth:1986:IRV

[CG86] A. Charlesworth and J. Gustafson. Introducing replicated VLSI to supercomputing: the FPS-164/MAX scientific computer. Computer, 19(3): 10–23, 1986. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:1987:GEP

[CG87] Eleanor Chu and Alan George. Gaussian elimination with partial pivoting and load balancing on a multiprocessor. Parallel Computing, 5 (1-2):65–74, July 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing— issues in applied research and development (Loen, 1986).

Chen:1996:OIP

[CG96] R. T. Chen and P. S. Guilfoyle, editors. Optoelectronic interconnects and packaging: Conference — January 1996, San Jose, CA, volume 62 of CRITICAL RE-VIEWS of OPTICAL SCI-ENCE AND TECHNOL-OGY CR 1996. SPIE Optical Engineering Press, Bellingham, WA, USA, 1996. ISBN 0-8194-2017-4. LCCN ????

Cameron:2005:HPP

[CGFT05] Kirk W. Cameron, Virginia Tech Rong Ge, Virginia Tech Xizhou Feng, and Virginia Tech. Highperformance, power-aware distributed computing for scientific applications. Computer, 38(11):40–47, November 2005. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clements:1994:MDI

[CGHL94] D. Clements, M. Ganesh, S.- Y. Hwang, and E.-P. Lim. Myriad: Design and implementation of a federated database prototype. In Balakrishnan [Bal94], pages 103– 118. ISBN 0-07-462044-4. LCCN ????

Chen:1992:WPP

[CGL92] Hsin-Chu Chen, Hui Gao, and G. Lai. WHAMS3D project progress report PR-3: parallel implementations of WHAMS3D on two shared-memory multiprocessors. Technical Report CSRD 1248, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1992. 20 pp.

Cheng:19xx:IPL

[CGLxx] Franklin Y. Cheng, Jeng-Fuh Ger, and Dan. Li. INRESB-3D-SUPII program listing for supercomputer: general purpose program for inelastic analysis of RC and steel building systems for 3D static and dynamic loads and seismic excitation. Civil engineering study. Structural series 96-4, Dept. of Civil Engineering, University of Missouri-Rolla, Rolla, MO, USA, 19xx. iv + 114 pp.

Cheng:1996:ISP

[CGLY96] Franklin Y. Cheng, Jeng-Fuh Ger, Dan Li, and J. S. Yang. INRESB-3D-SUPII program listing for supercomputer: general purpose program for inelastic analysis of RC and steel building systems for 3D static and dynamic loads and seismic excitation. Technical Report NSF BCS 9001494, NSF MSS 9214664., Dept. of Civil Engineering, University of Missouri-Rolla, Rolla, MO, USA, 1996. iv + 114 pp. Final report prepared for the National Science Foundation.

Cerimele:1991:VNS

[CGM91] M. M. Cerimele, F. R. Guarguaglini, and L. Moltedo. Visualizations for a numerical simulation of a flame diffusion model. Computers and Graphics, 15(2): 231–235, 1991. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Chiu:2005:P

[CGR05] G. L.-T. Chiu, M. Gupta, and A. K. Royyurus. Preface. IBM Journal of Research and Development, 49 (2/3):191–193, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/preface.
pdf.

Chen:1991:WPP

[CGS91] Hsin-Chu Chen, Hui Gao, and Sanjay Sharma. WHAMS3D project progress report PR-2. Technical Report CSRD 1112, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 21 pp.

Chandra:1994:PEH

[CGSG94] R. Chandra, K. Gharachorloo, V. Soundararajan, and A. Gupta. Performance evaluation of hybrid hardware and software distributed shared memory protocols. In Anonymous [Ano94-134], pages 274–288. ISBN ???? LCCN ????

Chang:2005:SIS

[CGW05] Weng-Long Chang, Minyi Guo, and Jesse Wu. Solving the independent-set problem in a DNA-based supercomputer model. Parallel Processing Letters, 15 (4):469–479, December 2005. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Chen:1987:PQM

[CH87] M.-Q. Chen and S.-P. Han. A parallel quasi-Newton method for partially separable large scale minimization. Technical Report CSRD 689, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 25 pp.

Cabral:1989:VTA

[CH89a] B. Cabral and C. L. Hunter. Visualization tools at Lawrence Livermore National Laboratory. Computer, 22(8): 77–84, August 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:1989:PFS

[CH89b] Hsin-Chu Chen and Ai-Fang He. Performance of the finite strip method for structural analysis on a parallel computer. Technical Report CSRD 920, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1989. 14 pp.

Chow:1990:SSM

[CH90] Jyh-Herng Chow and Luddy Harrison. Switch-stacks: a scheme for microtasking nested parallel loops. Technical Report CSRD 1032, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1990. 10 pp.

Chow:1992:CAP

[CH92a] Jyh-Herng Chow and Williams Ludwell Harrison. Compile-time analysis of parallel programs that share memory. Technical Report CSRD 1166, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. 12 pp.

Chow:1992:GFA

[CH92b] Jyh-Herng Chow and Williams Ludwell Harrison. A general framework for analyzing shared-memory parallel programs. Technical Report CSRD 1239, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1992. 8 pp.

Chen:1994:EAC

[CH94] L.-L. Chen and W. L. Harrison. An efficient approach to computing fixpoints for complex program analysis. In Anonymous [Ano94-134], pages 98–106. ISBN ???? LCCN ????

Chow:1998:OFB

[CH98] Edmond Chow and Michael A. Heroux. An object-oriented framework for block preconditioning. ACM Transactions on Mathematical Software, 24 (2):159–183, June 1998. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toms/1998-24-2/
                                        p159-chow/.

Chiu:2010:MDS

[CH10] Matt Chiu and Martin C. Herbordt. Molecular dynamics simulations on highperformance reconfigurable computing systems. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4):23:1– 23:??, November 2010. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Chauvet:1984:MVM

[Cha84] Y. Chauvet. Multitasking a vectorized Monte Carlo algorithm on the CRAY X-MP/2.

CRAY Channels, 6(3):6–9, 1984. CODEN CRCHE8.

Chamberlain:1990:SPD

[Cha90] R. M. Chamberlain. Solving partial differential equations on a parallel supercomputer. In Anonymous [Ano90g], pages 115–120. ISBN 1-873068-00-X. LCCN TA345.W67 1990.

Chakraborty:1992:TCA

[Cha92a] Abhijeet Chakraborty. Transient circuit analysis on a vector supercomputer. Thesis (M.S. in engineering), University of Texas at Austin, Austin, TX, USA, 1992. ix + 45 pp.

Chan:1992:QQR

[Cha92b]  Tony F. Chan.  QMR-

CGSTAB: a quasi-minimal residual variant of the Bi-CGSTAB algorithm for nonsymmetric systems. Technical Report CSRD 1231, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1992. 15 pp.

Charlebois:1993:PMG

[Cha93] R. L. Charlebois. Physical mapping of genomes using the landmark strategy. In Lim et al. [L^{+}93], pages 219–230. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Chan:1994:PIR

[Cha94a] Tony K. Y. Chan, editor. Proceedings of 1994 IEEE Region 10’s ninth annual international conference: theme: Frontiers of computer technology, 22– 26 August, 1994, Singapore. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1863-3, 0-7803-1862-5, 0-7803-1864-1. LCCN QA75.5.P735 1994. IEEE catalog number 94CH3417-3.

Chao:1994:HPA

[Cha94b] S.-L. Chao. Heat pipe for aircooling of a 60 W GaAs gate array in a 737-LGA package for the Convex C4 supercomputer. In Guo [Guo94], pages 129–134. ISBN 7-80003-324-4. LCCN ????

Chen:1983:LSH

[Che83]  Steve Chen.  Large-scale

and high-speed multiprocessor system for scientific applications. In ?. Kwalik, editor, Proc. NATO Advanced Research Work on High Speed Computing (June, 1983), volume F7 of NATO ASI Series, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN ???? LCCN ???? Revised version reprinted in [Hwa84, pp. 46– 58].

Chen:1988:SDD

[Che88] Hsin-Chu Chen. The SAS domain decomposition method for structural analysis. Technical Report CSRD 754; UILU-ENG-88-8003, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. vi + 112 pp.

Chen:1989:MED

[Che89a] Ding-Kai Chen. MaxPar:

an execution driven simulator for studying parallel systems. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1989. vii + 67 pp.

Cheng:1989:VPC

[Che89b] H. Cheng. Vector pipelining, chaining, and speed on the IBM 3090 and Cray X-MP. Computer, 22(9):31–42, 44, 46, September 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cheong:1989:CCC

[Che89c] Hoichi Cheong. Compilerdirected cache coherence strategies for large-scale shared-memory multiprocessor systems. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1989. xi + 139 pp.

Chen:1990:TSC

[Che90a] Hsin-Chu Chen. Two special classes of matrices. Technical Report CSRD 979, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1990. 12 pp.

Chen:1990:CSD

[Che90b] M.-Q. Chen. Conjugate subspaces decomposition and its application in solving linear systems with many righthand sides. Technical Report CSRD 1024, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1990. 41 pp.

Chen:1990:PQM

[Che90c]  M.-Q. Chen.  A paral-

lel quasi-Newton method for partially separable large-scale minimization. Technical Report CSRD 1028, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1990. 22 pp.

Chen:1990:SBV

[Che90d] Philip C. Chen. Supercomputerbased visualization systems used for analyzing output data of a numerical weather prediction model. ACM SIGARCH Computer Architecture News, 18(3b):296– 309, September 1990. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Chen:1990:SVS

[Che90e] Philip C. Chen. Supercomputerbased visualization systems used for analyzing output data of a numerical weather prediction model. ACM SIGARCH Computer Architecture News, 18(3):296–??, September 1, 1990. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Chervin:1990:NEW

[Che90f] R. M. Chervin. Numerical exploration of the world ocean. In Pitcher [Pit90], pages 189– 192. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Chen:1991:CMD

[Che91] Hsin-Chu Chen. Circulative matrices of degree θ. Technical Report CSRD 1094, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1991. 20 pp.

Chen:1992:PSM

[Che92a] H.-C. Chen. Parallel SAS multicluster algorithms for solving linear systems with reflexive coefficient matrices. In ACM [ACM92b], pages 447–455. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Cheong:1992:LSSb

[Che92b] H. Cheong. Life span strategy — a compiler-based approach to cache coherence. In ACM [ACM92b], pages 139–148. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Cheong:1992:LSSa

[Che92c] Hoichi Cheong. Life span strategy: a compiler-based approach to cache coherence. Technical Report CSRD 1189, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1992. 10 pp.

Chen:1993:CLP

[Che93a] S.-S. Chen. Characterizing and learning of protein conformations. In Lim et al. [L^{+}93], pages 391–402.

ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Chen:1993:CDP

[Che93b] Yung-Chin Chen. Cache design and performance in a large-scale shared-memory multiprocessor system. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1993. 187 pp.

Chen:1994:FDB

[Che94a] Chih-Hsuan Chen. A finite difference: boundary element scheme for solving parabolic boundary value problems and supercomputing. Thesis (Ph.D.), Texas A&M University, College Station, TX, USA, 1994. viii + 89 pp.

Chen:1994:SVS

[Che94b] P. C. Chen. Supercomputing visualization systems for scientific data analysis and their applications to meteorology. In Grave et al. [GLH94], pages 99–110. ISBN 3-540-56147-1, 0-387-56147-1. LCCN T385 .V59 1994.

Chen:1996:CWS

[Che96] P. C. Chen. Climate and weather simulations and data visualization using a supercomputer, workstations, and microcomputers [2656-26]. In Grinstein and Erbacher [GE96], pages 254–264. ISBN 0-8194-2030-1. ISSN 0361-0748. LCCN TS510.S63 v.2656.

Chergui:1999:UPP

[Che99] J. Chergui. Using PMD to parallel solve largescale Navier–Stokes equations. performance analysis on SGI/CRAY-T3E machine. In Dongarra et al. [DLM99], pages 341–348. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Cheng:19xx:IUM

[Chexx] Franklin Y. Cheng. INRESB-3D-SUPII user’s manual: general purpose program for inelastic analysis of RC and steel building systems for 3D static and dynamic loads and seismic excitations (based on supercomputer and PC). Civil engineering study. Structural series 96-3, Dept. of Civil Engineering, University of Missouri-Rolla, Rolla, MO, USA, 19xx. xiii + 237 pp.

Chin:1981:PCD

[Chi81] R. C. Y. Chin. Parallel computation of a domain decomposition method. Technical Report CSRD 657, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1981. 7 pp.

Chiang:1986:NSG

[Chi86] Chen Yu Chiang. Numerical simulation of ground water contaminant transport on a supercomputer with injectionpumping networks using the modified MOC and MFE method. Thesis (Ph.D.), Rice University, Houston, TX, USA, 1986. viii + 143 pp.

Childs:1988:TC

[Chi88] Bart Childs. TEX on the Cray. TUGboat, 9(2):181, August 1988. ISSN 0896-3207.

Chinni:1990:BSE

[Chi90] M. Chinni, editor. Ballistics simulation: proceedings of the SCS Eastern Multiconference, 23–26 April, 1990, Nashville, Tennessee. Society for Computer Simulation, San Diego, CA, USA, 1990. ISBN 0-911801-69-3. LCCN ????

Chiueh:1995:POP

[Chi95] T.-C. Chiueh. Performance optimization for parallel tape arrays. In ACM [ACM95a], pages 375–384. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Chien:2000:SWC

[Chi00] Andrew A. Chien. Supercomputing on Windows clusters: Experience and future directions. In USENIX [USE00b], page ?? ISBN 1-880446-20-0. LCCN ???? URL http:

//db.usenix.org/events/
usenix-win2000/invitedtalks/
chien.ppt.

Chiu:2016:TGS

[Chi16]  Y. Chiu.  Has Taiwan

given up on supercomputing? IEEE Spectrum, 53(1): 16–17, January 2016. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chivers:2020:DCC

[Chi20] Ian D. Chivers. The diagnostic capability of the Cray, gfortran, Intel, Nag and Oracle Fortran compilers: Revision 1. ACM Fortran Forum, 39(1):3–7, March 2020. ISSN 1061-7264 (print), 1931-1311 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3432987.3432990.

Carrico:1993:CSA

[CHL93] C. B. Carrico, U. R. Hanebutte, and E. E. Lewis. Comparison of space-angle approximations in response matrix algorithms. In Kusters et al. [KSW93], pages 58–70. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Clark:1994:PMD

[CHMS94] T. W. Clark, R. V. Hanxleden, J. A. McCammon, and L. R. Scott. Parallelizing molecular dynamics using spatial decomposition. In IEEE [IEE94c], pages 95– 102. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Chou:1990:DNN

[Cho90a] Cheng-Taou Chou. Development of neural network signal processing algorithms on the Cray Y-MP8/864 supercomputer. Thesis (M.S.), Dept. of Electrical Engineering, Ohio State University, Columbus, OH, USA, 1990. viii + 60 pp. Advisor: Stanley C. Ahalt.

Chow:1990:PEL

[Cho90b] Jyh-Herng Chow. Parallel execution of LISP programs in the parallel run time environment. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1990. viii + 83 pp.

Christoph:1990:SCG

[Chr90]  G. Christoph.  Security

considerations of going to a UNIX based supercomputer operating system. In USENIX Association [USE90], pages 129–130. ISBN ???? LCCN QA 76.9 A25 U55 1990.

Christiansen:1992:IE

[Chr92] Donald Christiansen. But is it ethical? IEEE Spectrum, 29(9):25–??, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christidis:1993:PCA

[Chr93] Z. Christidis. Parallel computing at IBM research on geophysical applications. In Hoffmann and Kauranne [HK93b], pages 44–59. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Coteus:2013:PIB

[CHT^{+}13] P. W. Coteus, S. A. Hall, T. Takken, R. A. Rand, S. Tian, G. V. Kopcsay, R. Bickford, F. P. Giordano, C. M. Marroquin, and M. J. Jeanson. Packaging the IBM Blue Gene/Q supercomputer. IBM Journal of Research and Development, 57(1/2):2:1–2:13, January–March 2013. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Chu:1987:MIO

[Chu87] C. Henry Chu. A model for an intelligent operating system for executing tasks on a reconfigurable parallel architecture. Technical report SRC-TR-87-007, Supercomputing Research Center: IDA, Lanham, MD, USA, 1987. 31 + [6] pp.

Churbuck:1989:SYP

[Chu89] David Churbuck. Is that a supercomputer in your pocket? Forbes, 143(10):121–??, May 15, 1989. CODEN FORBA5. ISSN 0015-6914.

Chung:1991:SOS

[Chu91] Siu-Leung Chung. Supercomputer optimization for stochastic dynamic programming. Thesis (Ph.D. in mathematics), University of Illinois at Chicago, Chicago, IL, USA, 1991. x + 66 pp.

Cala:2013:CCF

[CHWW13] Jacek Cala, Hugo Hiden, Simon Woodman, and Paul Watson. Cloud computing for fast prediction of chemical activity. Future Generation Computer Systems, 29(7): 1860–1869, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X13000253.

Ciarcia:1988:CCC

[Cia88a] S. Ciarcia. Ciarcia’s circuit cellar. 3 A supercomputer. BYTE Magazine, 13(13): 327–339, December 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1988:Sa

[Cia88b] S. Ciarcia. A supercomputer.

1. BYTE Magazine, 13(10): 283–291, October 1988. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1988:Sb

[Cia88c] S. Ciarcia. A supercomputer. 2. BYTE Magazine, 13(12):399–406, November 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1988:CCCa

[Cia88d] Steve Ciarcia. Ciarcia’s circuit cellar: a supercomputer, part 1. BYTE Magazine, 13 (10):283–??, October 1, 1988. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1988:CCCb

[Cia88e] Steve Ciarcia. Ciarcia’s circuit cellar: a supercomputer, part 2. BYTE Magazine, 13(12):399–??, November 1, 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ciarcia:1988:CCCc

[Cia88f] Steve Ciarcia. Ciarcia’s circuit cellar: a supercomputer, part 3. BYTE Magazine, 13(13):327–??, December 0, 1988. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Cigarini:1997:CDD

[Cig97] M. Cigarini. CFD-applications in the design and development of automotive automatic transmissions. In Roller [Rol97], pages 265– 272 (or 511–518??). ISBN 0-947719-88-1 (paperback). LCCN ????

Cadenas:1993:GAM

[CJ93] J. M. Cadenas and F. Jimenez.

A genetic algorithm for the multiobjective solid transportation problem: a fuzzy approach. In Anonymous [Ano93-31], pages 327–334. ISBN 0-947719-62-8. LCCN ????

Chandru:1994:FDS

[CJ94] V. Chandru and V. S. Jayachandran. The fault diameter of star graph interconnection networks. In Balakrishnan [Bal94], pages 375–377. ISBN 0-07-462044-4. LCCN ????

Cheng:1994:HAI

[CJHH94] Y. Cheng, J. R. Jensen, T. L. Huntsberger, and B. A. Huntsberger. Hypercube algorithm for image component labeling. In IEEE [IEE94c], pages 259–262. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Christiansen:1990:CMC

[CK90] E. Christiansen and K. O.

Kortanek. Computing material collapse displacement fields on a Cray X-MP/ 48 by the LP primal affine scaling algorithm. Annals of Operations Research, 22: 355–376, 1990. CODEN AOREEV. ISSN 0254-5330 (print), 1572-9338 (electronic).

Carino:1992:EDP

[CK92a] F. Carino and P. Kostamaa. Exegesis of DBC/1012 and P-90 industrial supercomputer database machines. Lecture Notes in Computer Science, 605(??):877–892, ???? 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cybenko:1992:SRE

[CK92b] G. Cybenko and D. J. Kuck. Supercomputers — revolution or evolution? IEEE Spectrum, 29(9):39– 41, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cybenko:1992:CPN

[CK92c] George Cybenko and David J. Kuck. A computer paradigm is needed if massively parallel machines are to live up to their billing. IEEE Spectrum, 29(9):39–41, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cook:2019:EPC

[CKD^{+}19] Brandon Cook, Thorsten Kurth, Jack Deslippe, Pierre Carrier, Nick Hill, and Nathan Wichmann. Eigensolver performance comparison on Cray XC systems. Concurrency and Computation: Practice and Experience, 31(16):e4997:1– e4997:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coghlan:2013:AAI

[CKL^{+}13] S. Coghlan, K. Kumaran, R. M. Loy, P. Messina, V. Morozov, J. C. Osborn, S. Parker, K. M. Riley, N. A. Romero, and T. J. Williams. Argonne applications for the IBM Blue Gene/Q, Mira. IBM Journal of Research and Development, 57(1/2):12:1– 12:11, January–March 2013. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Cytron:1988:AMP

[CKM88] Ronald Gary Cytron, Steve Karlovsky, and Kevin P. McAuliffe. Automatic management of programmable caches: (extended abstract). Technical Report CSRD 780, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 17 pp.

Cybenko:1990:SPE

[CKPK90a] George Cybenko, Lyle Kipp, Lynn Pointer, and David Kuck. Supercomputer performance evaluation and the Perfect Benchmarks. ACM SIGARCH Computer Architecture News, 18(3b):254– 266, September 1990. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Cybenko:1990:SPEb

[CKPK90b] George Cybenko, Lyle Kipp, Lynn Pointer, and David Kuck. Supercomputer performance evaluation and the Perfect Benchmarks. ACM SIGARCH Computer Architecture News, 18(3b):254– 266, September 1990. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Clark:1999:NSF

[CKS99] David Clark, Vipin Kumar, and Gil Shif. News: Supercomputing fashion: scalable vectors, the new fall line; mining scientific datasets: Cluster computing: Add processors, save money; new computing complex construction contract [Los Alamos National Laboratory]; IEEE Technical Committee on Parallel Processing; A smarter department of defense; Grid community effort continues. IEEE Concurrency, 7(4):4– 8, October/December 1999. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1999/pdf/p4004.
pdf.

Chernykh:2021:HHC

[CKT21] Igor Chernykh, Igor Kulikov, and Alexander Tutukov. Hydrogen-helium chemical and nuclear galaxy collision: Hydrodynamic simulations on AVX-512 supercomputers. Journal of Computational and Applied Mathematics, 391(??):Article 113395, August 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042721000145.

Cantor:1991:FIC

[CL91] Charles R. Cantor and Hwa A. Lim, editors. The First International Conference on Electrophoresis, Supercomputing, and the Human Genome: proceedings of the April 10–13 conference at Florida State University, Tallahassee, Florida. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1991. ISBN 981-02-0273-3. LCCN QH445.2 .I58 1990.

Clauss:1996:CSL

[Cla96] P. Clauss. Counting solutions to linear and nonlinear constraints through Ehrhart polynomials: Applications to analyze and transform scientific programs. In ACM [ACM96], pages 278–285. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Clark:1997:ISN

[Cla97] David Clark. Industry spotlight: No nukes is good news for supercomputers [nuclear weapons simulation]. IEEE Concurrency, 5 (2):11–12, April/June 1997. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

                                        //dlib.computer.org/pd/
                                        books/pd1997/pdf/p2011.
                                        pdf.

Clark:1998:FSN

[Cla98] David Clark. Focus: Supercomputing: The next generation. IEEE Computational Science & Engineering, 5(4): 79–81, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1998/pdf/c4079.
                                        pdf.

Clarke:2018:TSC

[Cla18] Sam Clarke. Trust separation on the Cray XC40 using PBS Pro. Concurrency and Computation: Practice and Experience, 30(1): ??, January 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Calegari:2019:WPH

[CLB19] Patrice Calegari, Marc Levrier, and Paweł Balczyński.

Web portals for high-performance computing: a survey. ACM Transactions on the Web (TWEB), 13(1):5:1–5:??, February 2019. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Chen:2019:IOD

[CLF^{+}19] Yuedan Chen, Kenli Li, Xiongwei Fei, Zhe Quan, and Keqin Li. Implementation and optimization of a data protecting model on the Sunway TaihuLight supercomputer with heterogeneous many-core processors. Concurrency and Computation: Practice and Experience, 31(21):e4758:1– e4758:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Chang:1991:IPC

[CLmWH91] Pohua P. Chang, Daniel M.

Lavery, and Wen mei W. Hwu. The importance of prepass code scheduling for superscalar and superpipelined processors. Technical Report CSRD 1144, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 47 pp.

Cloonan:1996:OIB

[Clo96] T. J. Cloonan, editor. Optical interconnects in broadband switching architectures: 31 January–1 February, 1996, San Jose, California, number 2692 in Proceedings — SPIE The International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1996. ISBN 0-8194-2066-2. ISSN 0361-0748. LCCN TA1632 .O672 1996.

Courtois:1993:CMU

[CLP93] G. Courtois, E. Letang, and G. Poullot. Calculation methods used by IPSN in the criticality safety studies. In Kusters et al. [KSW93], pages 570–576. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Cavarec:1993:BCP

[CLPV93] C. Cavarec, J. C. Lefebvre, J. F. Perron, and D. Verwaerde. Benchmark calculations of power distributions within assemblies. In Kusters et al. [KSW93], pages 259– 271. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Casanova:2009:PA

[CLR09] Henri Casanova, Arnaud Legrand, and Yves Robert. Parallel algorithms, volume 3 of Chapman and Hall/CRC numerical analysis and scientific computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-58488-945-4. xv + 337 pp. LCCN QA76.642 .C39 2009. URL http:

//catdir.loc.gov/catdir/
toc/fy0805/2008019142.html.

Chen:2019:PAM

[CLY^{+}19] Yuedan Chen, Kenli Li, Wangdong Yang, Guoqing Xiao, Xianghui Xie, and Tao Li. Performance-aware model for sparse matrix–matrix multiplication on the Sunway TaihuLight supercomputer. IEEE Transactions on Parallel and Distributed Systems, 30(4):923–938, April 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic)ITDSEO. URL

https://ieeexplore.ieee.
org/document/8468040/.

Chauvet:1984:MCX

[CM84] Yves Chauvet and Gerard Meurant. Multitasking on the Cray X-MP. The Journal of Systems and Software, 6 (1-2):17–20, May 1984. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chauvet:1986:MCX

[CM86] Yves Chauvet and Gerard Meurant. Multitasking on the Cray X-MP. The Journal of Systems and Software, 6 (1–2):17–20, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cuccu:1993:TMS

[CM93] Fabrizio Cuccu and Laura Moltedo. Texture mapping for scientific visualization environments. Computers and Graphics, 17(2):131– 136, March–April 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Cosshall:1995:PPA

[CM95] W. L. Cosshall and I. Morrison. The porting of parallel applications from transputer based machines to the Intel Paragon supercomputer. In Prasanna [Pra95], pages 711– 716. ISBN 0-07-462332-X. LCCN ????

Coskun:2011:ASC

[CMAS11] Ayse K. Coskun, Jie Meng, David Atienza, and Mohamed M. Sabry. Attaining single-chip, high-performance computing through 3D systems with active cooling. IEEE Micro, 31(4):63–75, July/August 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chrisochoides:1994:PEL

[CMF94] N. Chrisochoides, N. Mansour, and G. Fox. Performance evaluation of load balancing algorithms for parallel single-phase iterative PDE solvers. In IEEE [IEE94c], pages 764–772. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Chen:1992:TDA

[CMHK92] W. Y. Chen, S. A. Mahlke, W.-M. W. Hwu, and T. Kiyohara. Tolerating data access latency with register preloading. In ACM [ACM92b], pages 104–113. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Cuccaro:1994:TTQ

[CMP94] Steven A. Cuccaro, Michael Mascagni, and Daniel V. Pryor. Techniques for testing the quality of parallel pseudorandom number generators. Technical report SRC-TR-94-128, Supercomputing Research Center: IDA, Lanham, MD, USA, October 4, 1994. 6 pp.

Cagetti:1993:SAM

[CMPR93] P. Cagetti, R. C. Michelini, F. Pampagnin, and R. Razzoli. SIRIAT: An animation module for virtual reality simulation of robotic manipulators. In Anonymous [Ano93-31], pages 609–616. ISBN 0-947719-62-8. LCCN ????

Commer:2008:MPE

[CNC^{+}08] M. Commer, G. A. Newman, J. J. Carazzone, T. A. Dickens, K. E. Green, L. A. Wahrmund, D. E. Willen, and J. Shiu. Massively parallel electrical-conductivity imaging of hydrocarbons using the IBM Blue Gene/L supercomputer. IBM Journal of Research and Development, 52(1/2):93–??, January/March 2008. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/521/commer.html.

Cullis:1990:RCP

[CNGR90] L. G. Cullis, M. A. Nash, B. Grayson, and M. Rapier. Realistic calculations in penetration mechanics and detonics. In Pitcher [Pit90], pages 243–252. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Crockett:1994:PPR

[CO94] Thomas W. Crockett and Tobias Orloff. Parallel polygon rendering for Message-Passing architectures. IEEE parallel and distributed technology: systems and applications, 2(2):17–28, Summer 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Chou:1993:EPD

[COC93] J. Chou, W. A. O’Neill, and H. D. Cheng. Evaluation of pavement distress using artificial neural networks. In Anonymous [Ano93-31], pages 319–326. ISBN 0-947719-62-8. LCCN ????

Cochran:2001:NVS

[Coc01] Shannon Cochran. News and views: Scientists seek immersive reality; USENIX names lifetime achievement recipients [the GNU Project and the Kerberos network authentication system]; robots need programmers; evangelizing the Semantic Web; get your supercomputer software free; Usenet creator Jim Ellis dies; DARPA funds FreeBSD security initiative. Dr. Dobb’s Journal of Software Tools, 26 (9):18, September 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.
ddj.com/.

Cochran:2002:NVC

[Coc02a] Shannon Cochran. News and views: China takes ACM programming contest; supercomputing Science Grid nearing birth; hacking the glow of the screen; PARC searching for sponsors; milestone in DNA computing; Holzmann receives ACM Software Award; games aren’t just for playing around. Dr. Dobb’s Journal of Software Tools, 27 (6):14, June 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Cochran:2002:NVCb

[Coc02b] Shannon Cochran. News and views: China takes ACM programming contest; supercomputing Science Grid nearing birth; hacking the glow of the screen; PARC searching for sponsors; milestone in DNA computing; Holzmann receives ACM Software Award; games aren’t just for playing around. Dr. Dobb’s Journal of Software Tools, 27 (6):14, June 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Cochran:2002:NVS

[Coc02c] Shannon Cochran. News and views: Stanford student winds collegiate tournament; Web services get real; new supercomputer on the block; 2001 Lovelace Award announced. Dr. Dobb’s Journal of Software Tools, 27(7): 14, July 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Cochran:2002:NVSc

[Coc02d] Shannon Cochran. News and views: Stanford student winds collegiate tournament; Web services get real; new supercomputer on the block; 2001 Lovelace Award announced. Dr. Dobb’s Journal of Software Tools, 27(7): 14, July 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Cochran:2003:NVG

[Coc03a] Shannon Cochran. News and views: Government ponders open-source strategy; POSIX, Single UNIX specification merged; IBM plans new supercomputers; robotic surgeons have a heart. Dr. Dobb’s Journal of Software Tools, 28(2):14, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/documents/
s=7790/ddj0302o/.

Cochran:2003:NVGb

[Coc03b] Shannon Cochran. News and views: Government ponders open-source strategy; POSIX, Single UNIX specification merged; IBM plans new supercomputers; robotic surgeons have a heart. Dr. Dobb’s Journal of Software Tools, 28(2):14, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/documents/
s=7790/ddj0302o/.

Cohen:1991:SAA

[Coh91]  Jarret S. Cohen.  Su-

percomputing in American academia. National forum, LXXI(3):42–??, Summer 1991. ISSN 0162-1831.

Collard:1994:STW

[Col94] J.-F. Collard. Space-time transformation of while-loops using speculative execution. In IEEE [IEE94c], pages 429– 436. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

CTRC:1989:ST

[Com89] Computer Technology Research Corp. Supercomputer technology. Computer Technology Research Corp., 6 N. Atlantic Wharf, Charleston, SC 29401-2150, USA, 1989. 152 pp.

Comerford:1992:HEG

[Com92a] Richard Comerford. Huge efforts are going into software for the new generation of supercomputers. IEEE Spectrum, 29(9):34– 38, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1992:SSB

[Com92b] Richard Comerford. Supercomputers — software on the brink. IEEE Spectrum, 29(9): 34–38, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Comerford:1998:CTA

[Com98] Richard Comerford. Computers [technology 1998 analysis and forecast]. IEEE Spectrum, 35(1):43–47, January 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Conroy:1986:NPC

[Con86] John Michael Conroy. A note on the parallel Cholesky factorization of wide banded matrices. Technical report SRC-TR-87-002, Supercomputing Research Center: IDA, Lanham, MD, USA, 1986. 21 pp.

Conroy:1987:PAS

[Con87a] John Michael Conroy. Parallel algorithms for the solution of narrow banded systems. Technical report SRC-TR-87-001, Supercomputing Research Center: IDA, Lanham, MD, USA, 1987. 24 pp.

CSR:1987:SRR

[Con87b] The Spang Robinson report on supercomputing and parallel processing, 1987. ISSN 0897-4047; 1053-1661. Spang Robinson, Manchester, MA, USA.

Conte:1988:STG

[Con88] Thomas Martin Conte. The simulation and tuning of the global memory subsystem of a multiprocessor. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1988. xi + 80 pp.

Connolly:1990:SSG

[Con90] J. W. D. Connolly. Scientific supercomputing in the 90’s: The grand challenges, the global village and the great debate. In Anonymous [Ano90g], pages 151– 158. ISBN 1-873068-00-X. LCCN TA345.W67 1990.

Connolly:1991:SSG

[Con91] J. W. D. Connolly. Scientific supercomputing in the ’90’s: The grand challenges, the global village and the great debate. In Haji-Sheikh et al. [HS^{+}91], pages 493– 500. ISBN 1-56032-066-4. LCCN TA330 .I58 1991.

Conroy:1994:DSL

[Con94] John Michael Conroy. Dataparallel sparse LU factorization. Technical report SRC-TR-94-125, Supercomputing Research Center: IDA, Lanham, MD, USA, October 6, 1994. 31 pp.

Constantinescu:2000:TSA

[Con00]  C. Constantinescu.  Ter-

aflops supercomputer: architecture and validation of the fault tolerance mechanisms. IEEE Transactions on Computers, 49(9): 886–894, September 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=869320.

Conway:2011:IAR

[Con11] Lynn Conway. IBM-ACS: Reminiscences and lessons learned from a 1960’s supercomputer project. Lecture Notes in Computer Science, 6875:185–224, 2011. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-24541-1_
15/.

Cook:1995:SCB

[Coo95] Nick Cook. Supercomputer centre a boost for UK stealth. expanding the scenario system synergy sought in NATO C3I. Jane’s defence weekly, 23(9):25–??, March 4, 1995. ISSN 0265-3818.

Coppola:1993:AOT

[Cop93] Peter Coppola. Apple’s OCE takes the labor out of collaborations. Data communications, 22(10):105–??, July 1993. CODEN DACODM. ISSN 0363-6399.

CNSF:1987:GSA

[Cor87] Cornell National Supercomputer Facility, Cornell University, Ithaca, NY, USA. Getting started at the CNSF, 1987. ?? pp.

Corcoran:1989:SBSa

[Cor89a] Elizabeth Corcoran. Science and business: Supercomputing. Scientific American, 260(2):70–72, February 1989. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.
nature.com/scientificamerican/
journal/v260/n2/pdf/scientificamerican0289-
70.pdf.

CNSF:1989:ARC

[Cor89b] Cornell National Supercomputer Facility, Cornell University, Ithaca, NY, USA. Abstracts; Research on the Cornell National Supercomputer Facility / Cornell University, 1989. ?? pp.

Chao:1989:RAC

[COS89] C. C. Chao, S. A. Orszag, and W. Shyy, editors. Recent advances in computational fluid dynamics: US/ ROC joint workshop — May 1988, Princeton, NJ, number 43 in Lecture Notes in Engineering 1989. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-50872-4, 0-387-50872-

4. ISSN 0176-5035. LCCN ????

Chetverushkin:2021:NSH

[COT21] Boris N. Chetverushkin, Olga G. Olkhovskaya, and Il’ya P. Tsigvintsev. Numerical solution of hightemperature gas dynamics problems on high-performance computing systems. Journal of Computational and Applied Mathematics, 390 (??):??, July 2021. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042720306658.

Cousins:1990:MMS

[Cou90] L. B. Cousins. Management of a major supercomputing centre. In Anonymous [Ano90g], pages 45–48. ISBN 1-873068-00-X. LCCN TA345.W67 1990.

Courtland:2011:SLG

[Cou11] Rachel Courtland. Superconductor logic goes lowpower. IEEE Spectrum, 48 (7):18–20, July 2011. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Courtland:2013:ISB

[Cou13]  Rachel Courtland.  Intel

strikes back [news]. IEEE Spectrum, 50(8):14, August 2013. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cox:1988:USV

[Cox88] Donna J. Cox. Using the supercomputer to visualize higher dimensions: An artist’s contribution to scientific visualization. Leonardo (Oxford, England), 21(3): 233–??, July 1, 1988. CO-DEN LEONDP. ISSN 0024-094X (print), 1530-9282 (electronic).

Conn:1992:PRSb

[CP92a] Harold E. Conn and Louis J.

Podrazik. Parallel recurrence solvers and related applications on PETASYS.

Technical report SRC-TR-92-080, Supercomputing Research Center: IDA, Bowie, MD, USA, November 1992. 11 pp.

Conn:1992:PRSa

[CP92b] Harold E. Conn and Louis J. Podrazik. Parallel recurrence solvers for vector and SIMD supercomputers. Technical report SRC-TR-92-058R, Supercomputing Research Center: IDA, Lanham, MD, USA, March 1992. 21 pp.

Conroy:1992:PBF

[CP92c] John Michael Conroy and Louis J. Podrazik. Parallel bisection for finding eigenvalues on vector and SIMD architectures (U). Technical report SRC-TR-92-074, Supercomputing Research Center: IDA, Lanham, MD, USA, September 1992. 18 pp.

Conn:1993:MMP

[CP93a] Harold E. Conn and Louis J. Podrazik. Matrix multiplication on PETASYS/ TERASYS. Technical report SRC-TR-93-089, Supercomputing Research Center: IDA, Lanham, MD, USA, March 1993. 18 pp.

Cwik:1993:CES

[CP93b] T. Cwik and J. Patterson. Computational electromagnetics and supercomputer architecture, volume 7 of Progress in electromagnetics research (PIER). EMW Publishing, Cambridge, MA, USA, 1993. ISBN ???? ISSN 1070-4698. xv + 353 pp. LCCN QC 759.6 P76 v.7 1993.

Chattopadhyay:1994:ESB

[CP94a] S. Chattopadhyay and P. Pol Chaudhuri. Efficient signatures of Boolean functions for rapid matching in antifuse based FPGA technology mapping. In Balakrishnan [Bal94], pages 224–235. ISBN 0-07-462044-4. LCCN ????

Chen:1994:NAI

[CP94b] Peter M. Chen and David A.

Patterson. A new approach to I/O performance evaluation: Self-scaling I/O benchmarks, predicted I/O performance. ACM Transactions on Computer Systems, 12(4):308–339, November 1994. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tocs/1994-12-4/
p308-chen/.

Conn:1994:PRS

[CP94c] H. E. Conn and L. J.

Podrazik. Parallel recurrence solvers for vector and SIMD supercomputers. Journal of Parallel and Distributed Computing, 23(3): 435–441, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1994.1153/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1994.1153/production/pdf.

Cornu:1996:CCT

[CP96] T. Cornu and M. Pahud. Contention in the Cray T3D communication network. Lecture Notes in Computer Science, 1124:689–??, ???? 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Checconi:2013:MDA

[CP13] F. Checconi and F. Petrini. Massive data analytics: The Graph 500 on IBM Blue Gene/Q. IBM Journal of Research and Development, 57(1/2):10:1–10:11, January– March 2013. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Cammarasana:2025:ACH

[CP25] Simone Cammarasana and Giuseppe Patané. Analysis and comparison of high-performance computing solvers for minimisation problems in signal processing. Mathematics and Computers in Simulation, 229:525–538, March 2025. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0378475424003926.

Courtois:1993:VIC

[CPR93] G. Courtois, G. Poullot, and J. Roussignol. Validation of IPSN computation methods used in criticality safety calculations. In Kusters et al. [KSW93], pages 559–

569. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Chan:1996:PNAb

[CPS96a] M. C. Chan, G. Pacifici, and R. Stadler. Prototyping network architectures on a supercomputer. In IEEE [IEE96b], pages 374–

382. ISBN 0-8186-7582-9. ISSN 1082-8907. LCCN ????

Chan:1996:PNAa

[CPS96b] Mun Choon Chan, Giovanni Pacifici, and Rolf Stadler. Prototyping network architectures on a supercomputer. Research report RC. International Business Machines Corporation. Research Division ; 20438 International Business Machines Corporation. Research Division. Research report ; RC 20438. RC 20438 (90380), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, April 23, 1996. 13 pp.

Cole:1989:SSC

[CR89] Henderson Cole and Ann Redfels. Science and supercomputing on the Cornell National Supercomputer Facility. Cornell Theory Center, Cornell University, Ithaca, NY, USA, 1989. 40 pp.

Charalambous:1994:PSR

[CR94] Hara Charalambous and Alyson Reeves. Poincaré series and resolutions of the residue field over monomial rings. Technical report SRC-TR-94-140, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1994. 12 pp.

Crawford:1991:WSP

[Cra91] Diane Crawford. From Washington: Supercomputing policy links science and the ‘C’ word. Communications of the ACM, 34(11):25–28, November 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        125512.html.

CRI:1992:ICY

[Cra92] Cray Research, Inc., Eagan, MN, USA. Introducing the Cray Y-MP EL supercomputer system, 1992. 8 pp.

Cray:1996:ITB

[Cra96] Seymour Cray. An imaginary tour of a biological computer (why computer professionals and molecular biologists should start collaborating): Remarks of Seymour Cray to the Shannon Center for Advanced Studies, University of Virginia May 30, 1996. World-Wide Web document., May 30, 1996. URL http:/

/americanhistory.si.edu/
csr/comphist/montic/cray.
htm.

Clausen:2010:PPL

[CRA10] Jonathan R. Clausen, Daniel A.

Reasor, Jr., and Cyrus K. Aidun. Parallel performance of a lattice-Boltzmann/ finite element cellular blood flow solver on the IBM Blue Gene/P architecture. Computer Physics Communications, 181(6):1013– 1020, June 2010. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465510000391.

Creecy:1991:TMM

[Cre91] Robert H. Creecy. Trading MIPS and memory for knowledge engineering: automatic classification of census returns on a massively parallel supercomputer. Technical report TMC-192, Thinking Machines Corp, Cambridge, MA, USA, April 1991. 22 pp.

Chuang:1994:EPU

[CRM94] Ling-Yu Chuang, Vernon Rego, and Aditya Mathur. Experiments with program unification on the Cray Y-MP. Concurrency: practice and experience, 6(1):33– 53, February 1994. CODEN CPEXEI. ISSN 1040-3108.

Chalasani:1994:FTR

[CRV94] S. Chalasani, C. S. Raghavendra, and A. Varma. Faulttolerant routing in MINbased supercomputers. Journal of Parallel and Distributed Computing, 22(2): 154–167, August 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1994.1078/production;
                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        1994.1078/production/pdf.

Chakrabarti:1994:RLT

[CRY94] S. Chakrabarti, A. Ranade, and K. Yelick. Randomized load-balancing for treestructured computation. In IEEE [IEE94c], pages 666– 673. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

CRI:1982:DSA

[CS82] Directory of supercomputer applications software: a resource provided to customers, 1982. Cray Research, Inc., Minneapolis, MN, USA.

Cheung:1984:ACX

[CS84] Tony Cheung and James E. Smith. Analysis of the Cray X-MP memory system. Proceedings of the International Conference on Parallel Processing, pages 499–505, 1984.

CODEN PCPADL. ISBN 0-8186-0560-X. ISSN 0190-3918. IEEE Service Cent. Piscataway, NJ, USA.

Cheung:1986:SSC

[CS86a] Tony Cheung and James E.

Smith. Simulation study of the Cray X-MP memory system. IEEE Transactions on Computers, C-35 (7):613–622, July 1986. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

CRI:1986:DSA

[CS86b] Cray Research, Inc and Trudy Sprague. Directory of supercomputer application software: a resource provided to customers. Cray Research, Inc., Minneapolis, MN, USA, 1986. 448 pp.

Chen:1988:MDM

[CS88] Hsin-Chu Chen and Ahmed Sameh. A matrix decomposition method for orthotropic elasticity problems. Technical Report CSRD 798, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 48 pp.

Chen:1989:DDM

[CS89] Hsin-Chu Chen and Ahmed Sameh. A domain decomposition method for 3D elasticity problems. Technical Report CSRD 890, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1989. 18 pp.

Chuan:1990:SCS

[CS90] C. H. Chuan and W. C. Schreiber. The simulation of convective solidification using a 3-D computer code vectorized for optimum performance on supercomputers. In Pitcher [Pit90], pages 431–442. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Carlson:1991:CUM

[CS91] William W. Carlson and Thomas Lawrence Sterling. Coincidizing: a unifying model for computing systems. Technical report SRC-TR-91-033, Supercomputing Research Center: IDA, Lanham, MD, USA, 1991. 18 pp.

Carlson:1993:ACL

[CS93a] William W. Carlson and Judith D. Schlesinger. AC: a C language and compiler for the CM-5 node architecture. Technical report SRC-TR-93-094, Supercomputing Research Center: IDA, Lanham, MD, USA, April 1993. 32 pp.

Chung:1993:ENN

[CS93b] S. L. Chung and R. Setiono. Efficient neural network training algorithm for the Cray Y-MP supercomputer. In IEEE [IEE93a], pages 1943–1946. ISBN 0-7803-1421-2, 0-7803-1423-9, 0-7803-1422-0. LCCN QA76.87 .I57 1993 v.1-3 (c1993). Three volumes. IEEE catalog number: 93CH3353-0.

Cross:1994:MMP

[CS94a] E. A. Cross and J. J. Spaeth.

Microlithics’ massively parallel supercomputer. In IEEE [IEE94b], pages 267– 273. ISBN 0-7803-1893-5, 0-7803-1894-3, 0-7803-1895-1. ISSN 0547-3578. LCCN TL 693 N37 1994. Two volumes.

Cundari:1994:QML

[CS94b] T. R. Cundari and L. A. Stohecker. Quantum modeling of lanthanide complexes on parallel supercomputers. In IEEE [IEE94c], pages 710– 717. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Chung:1995:ENN

[CS95] S. L. Chung and R. Setiono. Efficient neural network training on a Cray Y-MP. International Journal of High Speed Computing, 7(1):

109–??, 1995. CODEN IH-SCEZ. ISSN 0129-0533.

Carlson:1989:ELS

[CSB89] D. M. Carlson, D. C. Sullivan, and R. E. Bach. The ETA 10 Liquid-Nitrogen-Cooled supercomputer system. IEEE Transactions on Electron Devices, 36(8):1404– ??, August 1, 1989. CODEN IETDAI. ISSN 0018-9383.

Clark:2000:NBG

[CSFS00] David Clark, Keri Schreiner, Jennifer Ferrero, and Dale Strok. News: Blue Gene and the race toward petaflops capacity; embedded Java development moves ahead; putting teraflops to the test; Corba 3.0 on the way. IEEE Concurrency, 8(1):5– 10, January/March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/
books/pd2000/pdf/p1005.
pdf.

Culler:1999:PCA

[CSG99] David Culler, Jaswinder Pal Singh, and Anoop Gupta. Parallel Computer Architecture: a Hardware/Software Approach. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1999. ISBN 1-55860-343-3. xxix + 1025 pp. LCCN QA76.58 .C85 1999. US$89.95.

Clifton:1997:IBM

[CSPJ97] J. Clifton, P. Smith, R. Peck, and B. L. Jones. Intelligent battery manager. In Roller [Rol97], pages 229– 236. ISBN 0-947719-88-1 (paperback). LCCN ????

UIUC-CSRD:1989:CN

[CSR89] CSRD notes, 1989. University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA.

UIUC-CSRD:19xx:CB

[CSRxx] CSRD bulletin, 19xx. University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA.

Cubasch:1990:SGC

[CSRB90] U. Cubasch, B. D. Santer, E. Maier Reimer, and M. Böttinger. Sensitivity of a global coupled ocean atmosphere circulation model to a doubling of carbon dioxide. In Pitcher [Pit90], pages 347–351. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Carey:1992:SOF

[CSSY92] G. Carey, J. Schmidt, V. Singh, and D. Yelton. A scalable, object-oriented finite element solver for partial differential equations on multicomputers. In ACM [ACM92b], pages 387–396. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Chen:1989:ISG

[CSY89] D. K. Chen, H. M. Su, and Pen-Chung Yew. The impact of synchronization and granularity on parallel systems. Technical Report CSRD 942, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1989. 21 pp.

Chan:1993:STI

[CT93a] F. T. S. Chan and N. K. H. Tang. Simulation as a tool to improve the production in automotive industries. In Anonymous [Ano93-31], pages 839–848. ISBN 0-947719-62-8. LCCN ????

Chao:1993:ACM

[CT93b] Y. A. Chao and N. Tsoulfanidis. Application of conformal mapping to nodal methods for hexagonal nodes. In Kusters et al. [KSW93], pages 605–619. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Cote:1994:PSA

[CT94] J. Cote and S. J. Thomas. Parallel semi-Lagrangian advection on the sphere using PVM. In IEEE [IEE94c], pages 470–477. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Chen:2016:RSE

[CTD^{+}16] Juan Chen, Yuhua Tang, Yong Dong, Jingling Xue, Zhiyuan Wang, and Wenhao Zhou. Reducing static energy in supercomputer interconnection networks using topology-aware partitioning. IEEE Transactions on Computers, 65(8):2588– 2602, ???? 2016. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chalmers:1994:PPP

[CTM94] A. Chalmers, J. Tidmus, and R. Miles. Parallel processing for photo-realistic visualisation. In Mahajan et al. [M^{+}94], pages 27–33. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Coste:1993:NIS

[CTRR93] M. Coste, H. Tellier, P. Ribon, and C. Raepsaet. New improvements in the selfshielding formalism of the APOLLO2 code. In Kusters et al. [KSW93], pages 93–101. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Case:1990:USM

[CU90] C. R. Case and J. J. Ullo.

Use of supercomputing to model nuclear geophysical measurements. In Pitcher [Pit90], pages 443–453. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Cullati:1995:NMA

[Cul95a] A. G. Cullati, editor. 210th National meeting: August 1995, Chicago, IL, volume 35(2) of National Meeting — American Chemical Society Division of Environmental Chemistry. American Chemical Society, Chicago, IL, USA, 1995. ISBN ???? ISSN 0093-3066. LCCN ????

Cullati:1995:UEN

[Cul95b] A. G. Cullati. US EPA’s National Environmental Supercomputing Center. In 210th National meeting: August 1995, Chicago, IL [Cul95a], pages 166–168. ISBN ???? ISSN 0093-3066. LCCN ????

Cheong:1988:CCS

[CV88a] Hoichi Cheong and Alexander Veidenbaum. A cache coherence scheme with fast selective invalidation. Technical Report CSRD 714, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 8 pp.

Cheong:1988:PSM

[CV88b] Hoichi Cheong and Alexander Veidenbaum. The performance of software-managed multiprocessor caches on parallel numerical programs. Technical Report CSRD 621, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 317– 337 pp.

Cheong:1988:SDD

[CV88c] Hoichi Cheong and Alexander Veidenbaum. Stale data detection and coherence enforcement using flow analysis. Technical Report CSRD 736, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 138– 145 pp.

Cheong:1989:CCM

[CV89a] Hoichi Cheong and Alexander Veidenbaum. Compilerdirected cache management for multiprocessors. Technical Report CSRD 937, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1989. 19 pp.

Cheong:1989:VCA

[CV89b] Hoichi Cheong and Alexander Veidenbaum. A version control approach to cache coherence. Technical Report CSRD 832, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1989. 9 pp.

Chen:1990:CAS

[CV90] Yung-Chin Chen and Alexander Veidenbaum. Comparison and analysis of software and directory. Technical Report CSRD 1055, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1990. 29 pp.

Chen:1991:IWB

[CV91a] Yung-Chin Chen and Alexander Veidenbaum. An improved write buffer design for a write-through cache. Technical Report CSRD 1105, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 18 pp.

Chen:1991:SCS

[CV91b] Yung-Chin Chen and Alexander Veidenbaum. A software coherence scheme with the assistance of directories. Technical Report CSRD 1106, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 22 pp.

Chen:1992:DIT

[CV92a] Yung-Chin Chen and Alexander Veidenbaum. Design issues and their performance impact in systems with directory-based caches. Technical Report CSRD 1175, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1992. 33 pp.

Chen:1992:EWP

[CV92b] Yung-Chin Chen and Alexander Veidenbaum. An effective write policy for software coherence schemes. Technical Report CSRD 1199, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1992. 17 pp.

Collado-Vides:1993:LIB

[CV93] J. Collado-Vides. A linguistic integration of a biological database. In Lim et al. [L^{+}93], pages 487–500. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Chiueh:1995:CDS

[CV95] T.-C. Chiueh and M. Verma.

A compiler-directed distributed shared memory system. In ACM [ACM95a], pages 77–86. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Colley:1989:SMD

[CW89] Stephen R. Colley and Doran K. Wilde. Supercomputer memory design. High performance systems, 10(4):65–??, April 1, 1989. CODEN HP-SYEA. ISSN 1043-7282.

Castain:2008:ORT

[CWD^{+}08] R. H. Castain, T. S. Woodall, D. J. Daniel, J. M. Squyres, B. Barrett, and G. E. Fagg. The Open Run-Time Environment (Open-RTE): a transparent multicluster environment for high-performance computing. Future Generation Computer Systems, 24(2):153– 157, February 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Chen:1997:SCP

[CWLT97] R. T. Chen, L. Wu, F. Li, and S. Tang. Si COMS process compatible guided-wave multi-gbit/sec optical clock signal distribution system for Cray T-90 supercomputer. In Goodman [Goo97], pages 10– 24. ISBN 0-8186-7975-1, 0-8186-7974-3, 0-8186-7976-X. LCCN ????

Chen:1994:CTM

[CWW94] Y.-S. Chen, S.-D. Wang, and C.-M. Wang. Compiler techniques for maximizing finegrain and coarse-grain parallelism in loops with uniform dependences. In Anonymous [Ano94-134], pages 204–213. ISBN ???? LCCN ????

Chen:1991:ESD

[CY91] Ding-Kai Chen and Pen-Chung Yew. An empirical study on DOACROSS loops. Technical Report CSRD 1140, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1991. 21 pp.

Cybenko:1989:ASS

[Cyb89a] George Cybenko. Approximations by superpositions of a sigmoidal function. Technical Report CSRD 856, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1989. 15 pp.

Cybenko:1989:DNN

[Cyb89b] George Cybenko. Designing neural networks. Technical Report CSRD 934, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1989. 7 pp.

Cybenko:1990:SPEa

[Cyb90] George Cybenko. Supercomputer performance evaluation and the Perfect Benchmarks. Technical Report CSRD 965, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1990. 20 pp.

Cybenko:1991:PCPa

[Cyb91a] George Cybenko. Parallel computing and the Perfect Benchmarks. Technical Report CSRD 1191, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1991. 12 pp.

Cybenko:1991:SPT

[Cyb91b] George Cybenko. Supercomputer performance trends and the Perfect BenchmarksTM. Technical Report CSRD 1093, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 8 + 2 pp.

Cyr:1986:SMA

[Cyr86]  Joseph Cyr.  Structured

memory access architecture: an implementation and performance-evaluation. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. vi + 70 pp.

Che:2018:PSC

[CYXL18] Yonggang Che, Meifang Yang, Chuanfu Xu, and Yutong Lu. Petascale scramjet combustion simulation on the Tianhe-2 heterogeneous supercomputer. Parallel Computing, 77(??):101– 117, September 2018. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S016781911830190X.

Czernik:1993:IVH

[Cze93] D. E. Czernik. Intelligent vehicle highways systems — “To be or not to be?”. In Anonymous [Ano93-31], pages 13–22. ISBN 0-947719-62-8. LCCN ????

Czech:2016:IPC

[Cze16] Zbigniew J. Czech. Introduction to Parallel Computing. Cambridge University Press, Cambridge, UK, 2016. ISBN 1-107-17439-2 (hardcover), 1-316-79583-7 (e-book). xvii + 354 pp. LCCN QA76.58 .C975 2016.

Chang:1993:NSD

[CZRB93] C. S. Chang, Y. Zhang, B. Rogg, and K. N. C. Bray.

93SC024 numerical simulation of diesel spray antoignition. In Anonymous [Ano93-31], pages 155–162. ISBN 0-947719-62-8. LCCN ????

Dongarra:1995:HPC

[D^{+}95] J. J. Dongarra et al., editors. High performance computing: technology, methods, and applications (Advanced workshop, June 1994, Cetraro, Italy), volume 10 of Advances in Parallel Computing. Elsevier, Amsterdam, The Netherlands, 1995. ISBN 0-444-82163-5. ISSN 0927-5452. LCCN QA76.88.H55 1995.

Darbha:1994:STD

[DA94] S. Darbha and D. P. Agrawal. SDBS: a task duplication based optimal scheduling algorithm. In IEEE [IEE94c], pages 756–763. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Dettmann:1997:MSB

[DA97] E. H. Dettmann and M. A. Abdelrhman. Modeling short-term behavior of dredged material disposed in very shallow and very deep coastal waters. In Delic and Wheeler [DW97], pages 109– 116. ISBN 0-89871-378-1. LCCN ????

Doerfler:2018:ENC

[DAC^{+}18] Douglas Doerfler, Brian Austin, Brandon Cook, Jack Deslippe, Krishna Kandalla, and Peter Mendygral. Evaluating the networking characteristics of the Cray XC-40 Intel Knights Landingbased Cori supercomputer at NERSC. Concurrency and Computation: Practice and Experience, 30(1): ??, January 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dixon:1990:LDF

[DAF^{+}90] D. A. Dixon, J. Andzelm, G. Fitzgerald, E. Wimmer, and B. Delley. Local density functional methods for molecular systems: Benchmarks and testing on molecules of interest to the chemical industry. In Pitcher [Pit90], pages 285–295. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Dakshinamoorthy:1990:TPI

[Dak90] Ravichandran Dakshinamoorthy. Thirst for parallelism to increased supercomputer performance. Technical report, ????, ????, 1990. vii + 62 pp.

Dzwinel:1998:LSM

[DAKM98] W. Dzwinel, W. Alda, J. Kitowski, and J. Moscinski. Large-scale molecular dynamics experiments on Cray T3E system. Lecture Notes in Computer Science, 1401:881–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dallaire:1984:AUN

[Dal84] Gene Dallaire. American universities need greater access to supercomputers. Communications of the ACM, 27 (4):293–298, 1984. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Daly:1995:HSB

[Dal95] Jim Daly. How to staff IS in the boondocks. Forbes, pages 26–??, April 10, 1995. CO-DEN FORBA5. ISSN 0015-6914.

Dalton:1996:PFS

[Dal96] Rex Dalton. Partners fall out over supercomputer bid. Nature, 381(6578):104–??, May 9, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Damevski:2011:OEC

[Dam11] Kostadin Damevski. Offline enforcement of contracts for high-performance computing. Concurrency and Computation: Practice and Experience, 23(13):1465–1473, September 10, 2011. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Danait:1991:RTE

[Dan91] Sachin W. Danait. A real time expert system for performance monitoring and scheduling of a Cray Y-MP supercomputer. Thesis (M.S.), Texas A&M University, College Station, TX, 1991. viii + 76 pp.

Daoud:1988:HFS

[Dao88] Raja B. (Raja Bishara)

Daoud. High-speed fault simulation on the Cray X/MP supercomputer. Thesis (M.S.), Dept. of Electrical Engineering, Ohio State University, Columbus, OH, USA, 1988. x + 88 pp. Advisor: Fusun Ozguner.

Das:1994:PKW

[Das94]  Amit Das.  Productiv-

ity in knowledge work: a study of technical support in supercomputing. Thesis (Ph.D.), University of Minnesota, Minneapolis, MN, USA, 1994. vii + 118 pp.

Daukantas:1996:NSP

[Dau96] Patricia Daukantas. NSF seeks partnerships to replace supercomputer centers. Computers in Physics, 10(1):9–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

                                        //aip.scitation.org/doi/
                                        10.1063/1.4822361.

Daukantas:1997:SDI

[Dau97] Patricia Daukantas. San Diego, Illinois sites to lead NSF’s new supercomputing partnerships. Computers in Physics, 11(3): 216–??, May 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4822542.

Davidson:1986:DCM

[Dav86a] Edward Steinberg Davidson. Development of CEDAR multiprocessor supercomputer, 1986. 1 videocassette (50 min.).

Davis:1986:PCA

[Dav86b] Timothy Alden Davis. Psolve: a concurrent algorithm for solving sparse systems of linear equations. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. vii + 56 pp.

Davis:1987:FNS

[Dav87]  Stephen G. Davis.  Fo-

cus on NSF supercomputing. Computers in Physics, 1(1):18–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.4903429.

Davis:1989:PAS

[Dav89] Timothy Alden Davis. A parallel algorithm for sparse unsymmetric LU factorization. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1989. ix + 140 pp.

Davis:1992:BC

[Dav92]  Dwight Davis.  Big-time

changes. Computer Graphics World, 15(7):42–??, July 1992. CODEN CGWODH. ISSN 0271-4159.

Davis:2000:TVC

[Dav00] Robert A. Davis. THOR:

a versatile commodity component of supercomputer development. Linux Journal, 75:??, July 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Day:2012:SAW

[Day12] Charles Day. “Supercomputers are awesome and why I love what I DO!!!”. Computing in Science and Engineering, 14(1):88, January/ February 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dharne:1994:VMF

[DB94] S. P. Dharne and R. Bhargava. Visualisation of the moderator flow and temperature patterns in the Calandria of Indian PHWRs. In Mahajan et al. [M^{+}94], pages 81–?? ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Dzwinel:1995:PRM

[DB95] W. Dzwinel and J. Blasiak. Pattern recognition via molecular dynamics on vector supercomputers and networked workstations. Lecture Notes in Computer Science, 919 (919):508–??, ???? 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:2009:PPM

[DBK09] Kei Davis, Kevin J. Barker, and Darren J. Kerbyson. Performance prediction via modeling: a case study of the ORNL Cray XT4 upgrade. Parallel Processing Letters, 19(4):619–639, December 2009. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

DaLiao:1993:MMH

[DC93] M. Da Liao and V. Cossalter. A mathematical model for harmonic motors. In Anonymous [Ano93-31], pages 91– 98. ISBN 0-947719-62-8. LCCN ????

delRosario:1994:HPO

[dC94] Juan Miguel del Rosario and Alok N. Choudhary. High-performance I/O for massively parallel computers: problems and prospects. Computer, 27(3):59–68, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Filho:2001:UMI

[dCCF01] Walfredo da Costa Cirne Filho.

Using moldability to improve the performance of supercomputer jobs. Vita thesis (Ph.D.), University of California, San Diego, San Diego, CA, USA, 2001.

DantasDeMelo:1990:VMD

[DCG90] J. Dantas De Melo, J. L.

Calvet, and J. M. Garcia. Vectorization and multitasking of dynamic programming in control experiments on a Cray-2. Parallel Computing, 13(3):261–269, March 1990. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dietz:1993:WYR

[DCG93] H. G. Dietz, W. E. Cohen, and B. K. Grant. Would you run it here… or there? (AHS: Automatic heterogeneous supercomputing). In Chen et al. [CBCH93], pages II–217–II–221. ISBN 0-8493-8983-6 (set), 0-8493-8984-4 (vol. 1), 0-8493-8985-2 (vol.

2), 0-8493-8986-0 (vol. 3). ISSN 0190-3918. LCCN QA76.58 .I55 1993 v.1-3 (c1993).

Dietz:19xx:WYR

[DCGxx] H. G. Dietz, W. E. Cohen, and B. K. Grant. Would you run it here… or there? (AHS: Automatic heterogeneous supercomputing). Proceedings of the International Conference on Parallel Processing, ????(????):II–217, ???? 19xx. CODEN PC-PADL. ISSN 0190-3918.

Difilippo:1993:SPN

[DCW93] F. C. Difilippo, M. Caro, and T. Williams. Simulation of pulsed neutron source reactivity measurements. In Kusters et al. [KSW93], pages 597–604. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Davis:2007:HPC

[DCWH07] Larry P. Davis, Roy L. Campbell, Jr., William A. Ward, Jr., and Cray J. Henry. Highperformance computing acquisitions based on the factors that matter. Computing in Science and Engineering, 9(6):35–44, November/ December 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dave:1987:SMC

[DD87] Ameet K. Dave and Iain S. Duff. Sparse matrix calculations on the Cray-2.

Parallel Computing, 5(1-2): 55–64, July 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Davis:1988:PRD

[DD88] Timothy Alden Davis and Edward Steinberg Davidson. Pairwise reduction for the direct, parallel solution of sparse, unsymmetric sets of linear equations. Technical Report CSRD 786, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 24 pp.

Dayde:1990:UPL

[DD90] M. J. Daydé and I. S.

Duff. Use of parallel level 3 BLAS in LU factorization on three vector multiprocessors the ALLIANT FX/80, the CRAY-2, and the IBM 3090 VF. ACM SIGARCH Computer Architecture News, 18 (3b):82–95, September 1990. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Decker:1993:BIN

[DD93] W. J. Decker and J. J. Dorning. A block iterative nodal integral method for fluid dynamic problems. In Kusters et al. [KSW93], pages 208– 223. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Dayde:1999:RBB

[DD99] Michel J. Daydé and Iain S. Duff. The RISC BLAS: a blocked implementation of Level 3 BLAS for RISC processors. ACM Transactions on Mathematical Software, 25 (3):316–340, September 1999. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
articles/journals/toms/
1999-25-3/p316-dayde/p316-
dayde.pdf;   http://www.
acm.org/pubs/citations/
journals/toms/1999-25-3/
p316-dayde/;   http://www.
acm.org/pubs/citations/
journals/toms/1999-25-3/
p316-dayde/#abstract;
http://www.acm.org/pubs/
citations/journals/toms/
1999-25-3/p316-dayde/#indterms.

Decyk:2002:SMP

[DD02] Viktor K. Decyk and Dean E. Dauger. Supercomputing for the masses: a parallel Macintosh cluster. Lecture Notes in Computer Science, 2328:10– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2328/23280010.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2328/23280010.
pdf.

Dauger:2005:PPC

[DD05] Dean E. Dauger and Viktor K. Decyk. Plug-and-play cluster computing: Highperformance computing for the mainstream. Computing in Science and Engineering, 7(2):27–33, March/ April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/dl/mags/cs/
2005/02/c2027.htm;   http:
//csdl.computer.org/dl/
mags/cs/2005/02/c2027.pdf.

Dubois:2010:SMV

[DDB^{+}10] David Dubois, Andrew Dubois, Thomas Boorman, Carolyn Connor, and Steve Poole. Sparse matrix-vector multiplication on a reconfigurable supercomputer with application. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(1):2:1–2:??, January 2010. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

DHollander:1996:PCS

[DDC96] E. D’Hollander, K. De Bosschere, and J. Van Campenhout, editors. Parallel computing: the state-of-the-art and perspective: proceedings of the international conference ParCo95, Gent, Belgium, 19–22 September 1995, volume 11 of Advances in Parallel Computing — Amsterdam — 1996. North-Holland, Amsterdam, The Netherlands, 1996. ISBN 0-444-82490-1. LCCN QA76.58.P764 1995.

Donnarumma:1993:CES

[DDF93] A. Donnarumma, S. A. Donnarumma, and E. Freda. Computational and experimental study on kinematic and dynamic behaviour of a rack and pinion coupling subject to elastic constraints. In Anonymous [Ano93-31], pages 719–726. ISBN 0-947719-62-8. LCCN ????

Dally:1994:RRR

[DDHK94] W. J. Dally, L. R. Dennison, D. Harris, and K. Kan. The reliable router: a reliable and high-performance communication substrate for parallel computers. Lecture Notes in Computer Science, 853:241–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DDJStaff:1998:NVS

[DDJ98a] DDJ Staff. News and views: a Standard Linux? cryptography contest; drives get smaller and Smaller; Perl conference; really embedded systems; programmer shortage?; Beowulf: Linux clustering; Java SPEC released. Dr. Dobb’s Journal of Software Tools, 23(11):16, November 1998. CODEN DDJOEB.

ISSN 1044-789X. URL http:

//www.ddj.com/.

Staff:1998:NVK

[DDJ98b] DDJ Staff. News and views:

Kudos for free software pioneers; PSCs: Personal supercomputers; smart dialing; let it snow…; math for the Web; the taxman changes; advances in nanoelectromechanical technology; Tcl goes it alone. Dr. Dobb’s Journal of Software Tools, 23(5): 18, May 1998. CODEN DDJOEB. ISSN 1044-789X.

Drmanac:1993:SIC

[DDLV93] R. Drmanac, S. Drmanac, I. Labat, and A. Vicentic. SBH and the integration of complementary approaches in the mapping, sequencing, and understanding of complex genomes. In Lim et al. [L^{+}93], pages 121–134. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Desprez:1995:PSL

[DDT95] F. Desprez, J. J. Dongarra, and B. Tourancheau. Performance study of LU factorization with low communication overhead on multiprocessors. Parallel Processing Letters, 5(2):157–169, June 1995. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

DeRose:1991:POCa

[De 91a] Luiz A. De Rose. Parallel ocean circulation modeling on Cedar. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1991. ix + 77 pp.

DeRose:1991:POCb

[De 91b] Luiz A. De Rose. Parallel ocean circulation modeling on Cedar. Technical Report CSRD 1124, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 9 pp.

DeSario:1996:MIA

[De 96] M. De Sario, editor. Melecon ’96: industrial applications in power systems, computer science and telecommunications: Mediterranean electrotechnical conference; 8th — May 1996, Bari, Italy, PRO-CEEDINGS of MELECON MEDITERRANEAN ELEC-TROTECHNICAL CON-FERENCE 1996//V1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3110-9, 0-7803-3109-5, 0-7803-3111-7. LCCN ????

Dongarra:1984:SMA

[DE84] Jack J. Dongarra and Stanley C. Eisenstat. Squeezing the most out of an algorithm in CRAY FORTRAN.

ACM Transactions on Mathematical Software, 10(3):219– 230, September 1984. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Decker:1990:ILM

[Dec90] K. M. Decker. Igtlib — a library for Monte Carlo simulations of lattice gauge theory on CRAY computer systems. In Pitcher [Pit90], pages 395– 405. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Defend:1987:PRT

[Def87] Daniel Evan Defend. Parallel ray tracing in a vectormultiprocessing environment. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. viii + 59 pp.

Degen:1990:OPT

[Deg90] Gregg Francis Degen. Optimization of a power transient stability program on a vector supercomputer, theory and applications. Thesis (M.S.), San Diego State University, San Diego, CA, USA, 1990. x + 121 pp.

Dehn:1990:SMT

[Deh90] J. Dehn. Supercomputer modeling in terminal ballistics. In Chinni [Chi90], pages 57–62. ISBN 0-911801-69-3. LCCN ????

delGuercio:1989:SS

[del89] Gino del Guercio. Supercomputer showdown. World monitor (Boston, MA), 2(7):38– ??, July 1, 1989. ISSN 0897-9472.

Delacour:1997:SSL

[Del97] J. Delacour. SPEOS: Simulation of light systems in the automotive industry. In Roller [Rol97], pages 53–60. ISBN 0-947719-88-1 (paperback). LCCN ????

Demmel:1991:LPL

[Dem91] J. Demmel. LAPACK: a portable linear algebra library for high-performance computers. Concurrency: practice and experience, 3 (6):655–666, December 1991. CODEN CPEXEI. ISSN 1040-3108.

Dennis:1980:DFS

[Den80] Jack B. Dennis. Data flow supercomputers. Computer, 13 (11):48–56, November 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dent:1993:PIM

[Den93] D. W. Dent. Parallelization of the IFS model. In Hoffmann and Kauranne [HK93b], pages 73–87. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Deukmejian:1986:TGG

[Deu86] Governor George Deukmejian. Text of Governor George Deukmejian’s remarks at the dedication of the San Diego Supercomputer Center, September 8, 1986.

Deyhimy:1995:GAJ

[Dey95]  Ira Deyhimy.  Gallium

arsenide joins the giants. IEEE Spectrum, 32(2):33– 40, February 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Draper:1990:DMD

[DF90a] Richard N. Draper and Vance Faber. The diameter and mean diameter of supertoroidal networks. Technical report SRC-TR-90-004, Supercomputing Research Center: IDA, Lanham, MD, USA, January 30, 1990. 53 pp.

Dubey:1990:OP

[DF90b] Pradeep K. Dubey and Michael J. Flynn. Optimal pipelining. Journal of Parallel and Distributed Computing, 8(1):10– 19, January 1990. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Doallo:2012:SIE

[DF12] Ramón Doallo and Basilio B.

Fraguela. Special issue editorial: Accelerators for high-performance computing. Journal of Parallel and Distributed Computing, 72(9): 1055–1056, September 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731512001219.

Dongarra:2002:SPC

[DFF^{+}02] Jack Dongarra, Ian Foster, Geoffrey Fox, William Gropp, Ken Kennedy, Linda Torczon, and Andy White, editors. The Sourcebook of Parallel Computing. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2002. ISBN 1-55860-871-0. xvi + 842 + 8 pp. LCCN QA76.58 S638 2003. US$59.95.

Dickinson:1993:EPU

[DFS93] A. Dickinson, R. Ford, and D. F. Snelling. Experiences porting the unified model to the KSR-1. In Hoffmann and Kauranne [HK93b], pages 436–444. ISBN 981-02-1429-

4. LCCN QA76.58 E354 1992.

Dekker:1988:OLD

[DFSZ88] L. Dekker, E. E. E. Frietman, W. Smit, and J. C. Zuidervaart. Optical link in the Delft Parallel Processor — an example of MOMI-connection in MIMDsupercomputers. Future Generation Computer Systems, 4 (3):189–203, October 1988.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Drake:1993:MPV

[DFWW93] J. B. Drake, R. E. Flanery, D. W. Walker, and P. H. Worley. The message passing version of the parallel community climate model. In Hoffmann and Kauranne [HK93b], pages 500–513. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Desai:1995:CSS

[DG95] K. R. Desai and K. Ghose.

A comparative study of single hop WDM interconnections for multiprocessors. In ACM [ACM95a], pages 154– 163. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

DeDoncker:1996:PSP

[DGBE96] E. DeDoncker, A. Gupta, J. Ball, and P. Ealy. ParInt: a software package for parallel integration. In ACM [ACM96], pages 149–156. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

DeRose:1992:EOCa

[DGG92a] Luiz A. De Rose, Kyle A.

Gallivan, and E. J. (Efstratios J.) Gallopoulos. Experiments with an ocean circulation model on Cedar. Technical Report CSRD 1200, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1992. 12 pp.

DeRose:1992:EOCb

[DGG92b] L. DeRose, K. Gallivan, and E. Gallopoulos. Experiments with an ocean circulation model on CEDAR. In ACM [ACM92b], pages 397–408. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

DeRose:1993:SRP

[DGG93] L. DeRose, K. Gallivan, and E. Gallopoulos. Status report: Parallel ocean circulation on cedar. In Hoffmann and Kauranne [HK93b], pages 157–172. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Dhawan:2018:SSA

[DGG18] S. M. Dhawan, B. M. Gupta, and Ritu Gupta. Supercomputing: a scientometric assessment of global publications output during 2007–

16. Collnet Journal of Scientometrics and Information Management, 12(2): 197–213, 2018. CODEN ???? ISSN 0973-7766 (print), 2168-930X (elec-

tronic). URL http://www.
tandfonline.com/doi/abs/
10.1080/09737766.2018.1545395.

Dongarra:2020:NAH

[DGH20] Jack Dongarra, Laura Grigori, and Nicholas J. Higham. Numerical algorithms for high-performance computational science. Philosophical transactions of the Royal Society of London Series A, 378(2166):20190066:1– 20190066:18, 2020. ISSN 1364-503X (print), 1471-2962 (electronic).

Dousse:1993:FBD

[DGJG93] C. Dousse, J. Goette, M. Jacomet, and U. Grob. Fuzzylogic based decision-machine to solve the handover problem in a digital cellular radio system. In Anonymous [Ano93-31], pages 219–224. ISBN 0-947719-62-8. LCCN ????

Deegener:1993:RPM

[DGJK93] M. Deegener, G. Grosse, W. John, and B. Kuehnapfel. Rapid prototyping with MuSE. In Anonymous [Ano93-31], pages 625–632. ISBN 0-947719-62-8. LCCN ????

Dean:1989:SSP

[DGL89] E. Dean, R. Glowinski, and C. H. Li. Supercomputer solutions of partial differential equation problems in computational fluid dynamics and in control. Computer Physics Communications, 53 (1–3):401–439, May 1, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465589901768.

Davies:1990:DAS

[DGO90] A. M. Davies, R. B. Grzonka, and M. O’Neill. Development and application of 3D shallow sea models on the CRAY computer. In Pitcher [Pit90], pages 323–

334. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Dennis:1982:DFS

[DGT82] Jack B. (Jack Bonnell) Dennis, Guang Rong Gao, and Kenneth Wayne Todd. A data flow supercomputer. Computation Structures Group memo 213, Computation Structures Group, MIT Laboratory for Computer Science, Cambridge, MA, USA, March 1982. 35 pp.

Dennis:1984:MWD

[DGT84] Jack B. Dennis, Guang-Rong Gao, and Kenneth W. Todd. Modeling the weather with a data flow supercomputer. IEEE Transactions on Computers, C-33(7):592–603, July 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&
arnumber=5009332.

Domain:1994:UWP

[DGT94] C. Domain, J.-P. Gregoire, and B. Thomas. Ultrasonic wave propagation on parallel machines. In Anonymous [Ano94-134], pages 93– 97. ISBN ???? LCCN ????

Dongarra:1986:IDL

[DH86a] Jack J. Dongarra and Tom Hewitt. Implementing dense linear algebra algorithms using multitasking on the Cray X-MP-4 (or approaching the gigaflop). SIAM Journal on Scientific and Statistical Computing, 7(1):347–350, January 1986. CODEN SI-JCD4. ISSN 0196-5204.

Dongarra:1986:CCX

[DH86b] Jack J. Dongarra and Alan Hinds. Comparison of the Cray X-MP-4, Fujitsu VP-200, and Hitachi S-810/ 20. Simulation, 47(3):93– 107, September 1986. CO-DEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic).

Detert:1991:CXM

[DH91a] U. Detert and G. Hofemann. CRAY X-MP and Y-MP memory performance. Parallel Computing, 17(4–5): 579–590, July 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Detert:1991:CXY

[DH91b] Ulrich Detert and Gerd Hoffmann. Cray X-MP and Y-MP memory performance. Parallel Computing, 17(4-5): 579–590, July 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Diegele:1993:VMM

[DH93] E. Diegele and K. Hornberger. Viscoplastic material models in structural analysis of the first wall in a fusion reactor. In Kusters et al. [KSW93], pages 128– 136. ISBN 3-923704-11-9. LCCN ???? Two volumes.

DeVito:2013:TMS

[DHA^{+}13] Zachary DeVito, James Hegarty, Alex Aiken, Pat Hanrahan, and Jan Vitek. Terra: a multi-stage language for high-performance computing. ACM SIG-PLAN Notices, 48(6):105– 116, June 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dekker:1989:ASN

[DHD89] T. J. Dekker, W. Hoffmann, and P. P. M. De Rijk. Algorithms for solving numerical linear algebra problems on supercomputers. Future Generation Computer Systems, 4(4):255–263, March 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dongarra:1993:DSM

[DHHW93] J. J. Dongarra, R. Hempel, A. J. G. Hey, and D. W. Walker. A draft standard for message passing in a distributed memory environment. In Hoffmann and Kauranne [HK93b], pages 465– 481. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Dangelmaier:1997:SPP

[DHLS97] W. Dangelmaier, R. Holtkamp, T. Langemann, and H. J. Sobiech. Simulation of production planning and control system in the automotive industry. In Roller [Rol97], pages 161–168. ISBN 0-947719-88-1 (paperback). LCCN ????

Diede:1988:TGS

[DHM^{+}88] Tom Diede, Carl F. Hagenmaier, Glen S. Miranker, Jonathan J. Rubinstein, and William S. Worley, Jr. The Titan Graphics Supercomputer architecture. Computer, 21(9):13– 28, 30, September 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Douglas:1989:FRS

[DHT89] Ellen Douglas, Alan R.

Houser, and C. Roy Taylor. Final report on supercomputer research: 15 November 1983 to 31 May 1988. Technical Report CMU-CS-89-157, Carnegie Mellon University, Computer Science Dept, Pittsburgh, PA, USA, June 1989. various pp.

Dimpsey:1988:PAS

[DI88] R. T. Dimpsey and Ravishankar K. Iyer. Performance analysis of a shared memory multiprocessor: case study. Technical Report CSRD 779, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 8 pp.

Dickinson:1981:ONW

[Dic81] A. Dickinson. Optimizing numerical weather forecasting models for the Cray-1 and Cyber 205 computers. Computer Physics Communications, 26(3 & 4): 459–468, June 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Dickinson:1982:ONW

[Dic82] A. Dickinson. Optimizing numerical weather forecasting models for the Cray-1 and Cyber 205 computers. Computer Physics Communications, 26(3–4): 459–468, June 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465582901436.

Dickinson:1990:AFC

[Dic90] A. Dickinson. Autotasking forecast and climate models on a CRAY Y-MP. In Pitcher [Pit90], pages 35–

46. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Dick:1994:CUP

[Dic94] E. Dick. Chapter 4: a unified parallelizable multigrid approach to solve steady Euler

— and Navier–Stokes equations. In Murthy and Brebbia [MB94b], pages 75–96. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Dieckmann:1995:CAL

[Die95] U. Dieckmann. Coevolution as an autonomous learning strategy for neuromodules. In Herrmann et al. [HWP95], pages 427–432. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Dillmann:1993:UVR

[Dil93]  R. Dillmann.  The use

of virtual reality to support intelligent telerobot systems [invited]. In Kusters et al. [KSW93], pages 379–

382. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Ding:1991:CQP

[Din91] H. Q. Ding. Computing quark potential on a parallel supercomputer. Computer Physics Communications, 65 (1–3):92–99, April 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/001046559190159I.

Ding:1992:RAV

[Din92] Ailian Ding. Reduction and analysis of VLA maps for 281 radio-loud quasars using the Cray Y-MP supercomputer. Thesis (M.S.), University of Nevada, Las Vegas, Las Vegas, NV, USA, 1992. vii + 171 pp.

Ding:1993:MCS

[Din93] Hong Q. Ding. Monte Carlo simulations of quantum systems on massively parallel supercomputers. Computers in Physics, 7(6):687– ??, November 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.168476.

Diplock:1996:BNS

[Dip96] G. Diplock. Building new spatial interaction models using genetic programming and a supercomputer. In Anonymous [Ano96k], pages 213– 226.

Divins:1997:MAC

[Div97] D. Divins. Modelling an automotive cranking and charging system in a mixed domain simulator. In Roller [Rol97], pages 61–68. ISBN 0-947719-88-1 (paperback). LCCN ????

DeGloria:1994:TAS

[DJM94] A. De Gloria, M. R. (Mike R.)

Jane, and Daniele Marini, editors. Transputer applications and systems ’94: proceedings of the world transputer congress, 5–7 September 1994, Villa Erba, Cernobbio, Como, Italy, volume 41 of Transputer and Occam Engineering Series. IOS Press and Ohmsha, Amsterdam; Tokyo, 1994. ISBN 90-5199-177-0, 4-274-90004-5. ISSN 0925-4986. LCCN ????

Dash:1993:ITG

[DJSP93] S. K. Dash, B. Jha, S. Selvakumar, and M. Periasamy. Implementation of a T21 global spectral model on 16 nodes of parallel computers. In Hoffmann and Kauranne [HK93b], pages 60–72. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Deng:2001:PSB

[DK01] Yuefan Deng and Alex Korobka. The performance of a supercomputer built with commodity components. Parallel Computing, 27(1– 2):91–108, January 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-ng/
10/35/21/47/25/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/47/25/
26/article.pdf.

Damodaran-Kamal:1994:MSR

[DKF94] S. K. Damodaran-Kamal and J. M. Francioni. mdb: a semantic race detection tool for PVM. In IEEE [IEE94c], pages 702–709. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Dongarra:1986:SME

[DKH86] Jack J. Dongarra, Linda Kaufman, and Sven Hammarling. Squeezing the most out of eigenvalue solvers on high-performance computers. Linear Algebra and its Applications, 77(??): 113–136, May 1986. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0024379586901643.

Special volume on parallel computing.

Davidson:1986:STC

[DKLS86] E. Davidson, David J. Kuck, D. Lawrie, and Ahmed Sameh. Supercomputing tradeoffs and the Cedar system. Technical Report CSRD-577, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 14 pp.

Darema:1993:MCS

[DKS93] F. Darema, M. H. Kalos, and M. L. Simmons. Monte Carlo simulation on massively parallel machines [invited]. In Kusters et al. [KSW93], pages 353–355. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Doi:1990:SPV

[DL90] S. Doi and A. Lichnewsky.

Some parallel and vector implementations of preconditioned iterative methods on Cray-2. International Journal of High Speed Computing, 2(2):143–180, June 1990. CO-DEN IHSCEZ. ISSN 0129-0533.

Daoudi:1992:IBE

[DL92] E. M. Daoudi and J. Lobry.

Implementation of a boundary element method on distributed memory computers. Parallel Computing, 18(12): 1317–1324, December 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

DelCorral:1996:RIC

[DL96] A. Del Corral and J. Llaberia.

Reducing inter-vector-conflicts in complex memory systems. In ACM [ACM96], pages 382–389. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

DeLeege:1993:LWS

[DLG93] P. F. A. De Leege, J. M. Li, and I. C. Gauld. Linking WIMS with the SCALE depletion sequence. In Kusters et al. [KSW93], pages 577– 589. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Djurfeldt:2008:BSS

[DLJ^{+}08] M. Djurfeldt, M. Lundqvist, C. Johansson, M. Rehn, Ö. Ekeberg, and A. Lansner. Brain-scale simulation of the neocortex on the IBM Blue Gene/L supercomputer. IBM Journal of Research and Development, 52(1/2): 31–??, January/March 2008. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/521/djurfeldt.
html.

Ding:1998:ADA

[DLLG98] C. H. Q. Ding, P. M. Lyster, J. W. Larson, and J. Guo. Atmospheric data assimilation on distributed-memory parallel supercomputers. Lecture Notes in Computer Science, 1401:115–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1999:RAP

[DLM99] J. J. Dongarra, E. Luque, and Tomas Margalef, editors. Recent advances in parallel virtual machine and message passing interface: 6th European PVM/ MPI Users’ Group Meeting, Barcelona, Spain, September 26–29, 1999: proceedings, volume 1697 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Dikaiakos:1995:PPI

[DLMW95] M. D. Dikaiakos, C. Lin, D. Manoussaki, and D. E. Woodward. The portable parallel implementation of two novel mathematical biology algorithms in ZPL. In ACM [ACM95a], pages 365– 374. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Deng:1994:CTF

[DLPQ94] G. B. Deng, Y. Lecointe, J. Piquet, and P. Queutey. Chapter 3: Three-dimensional full Navier–Stokes solvers for incompressible flows past arbitrary geometries using supercomputers. In Murthy and Brebbia [MB94b], pages 41–74. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Davis:1993:SSR

[DLS93] A. B. Davis, S. Lovejoy, and D. Schertzer. Supercomputer simulation of radiative transfer inside multifractal cloud models. In Keevallik and Karner [KK93], pages 112– 115. ISBN 0-937194-28-X. LCCN QC912.3.I57 1992.

Dyer:1988:AFD

[DM88a] Stephen A. Dyer and L. Robert Morris. Afterword: Floating-Point digital signal processing chips, the end of the supercomputer era? IEEE Micro, 8(6):86–??, November/December 1, 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dyer:1988:AFP

[DM88b] Stephen A. Dyer and L. Robert Morris. Afterword: Floating-Point digital signal processing chips, the end of the supercomputer era? IEEE Micro, 8(6):86–??, November/December 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dawson:1993:HAP

[DM93] D. Dawson and F. W. Margrave. Hazards and accident prevention in the loading and unloading of commercial vehicles on roll on/ roll off ships. In Anonymous [Ano93-31], pages 527–534. ISBN 0-947719-62-8. LCCN ????

Daminelli:1996:PPSa

[DM96a] G. Daminelli and F. Mancosu. P-vision — Pirelli supercomputing in tyre technology. In Roller [Rol96], pages 87–96. ISBN 0-947719-81-4. LCCN ????

Daminelli:1996:PPSb

[DM96b] G. Daminelli and F. Mancosu. P-vision — Pirelli supercomputing in tyre technology. In Roller [Rol96], pages 87–96. ISBN 0-947719-81-4. LCCN ????

Destri:1996:BLA

[DM96c] G. Destri and P. Marenzoni. Benchmarking latticebased applications on parallel architectures. Parallel Processing Letters, 6 (3):309–320, September 1996. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Darnell:1992:ASC

[DMCK92] E. Darnell, J. M. Mellor-Crummey, and K. Kennedy. Automatic software cache coherence through vectorization. In ACM [ACM92b], pages 129–138. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Doriath:1993:EAE

[DMKW93] J. Y. Doriath, C. W. Mc-Callien, E. Kiefhaber, and U. Wehmann. ERANOS 1: The advanced European system of codes for reactor physics calculations. In Kusters et al. [KSW93], pages 177–186. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Doriath:1993:VNM

[DMPR93] J. Y. Doriath, F. Malvagi, G. Palmiotti, and J. M. Ruggieri. Variational nodal method (VNM) to solve 3-D transport equation. application to EFR design. In Kusters et al. [KSW93], pages 571–580. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Dubey:2021:PPE

[DMT^{+}21] Anshu Dubey, Lois Curfman McInnes, Rajeev Thakur, Erik W. Draeger, Thomas Evans, Timothy C. Germann, and William E. Hart. Performance portability in the exascale computing project: Exploration through a panel series. Computing in Science and Engineering, 23(5):46– 54, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dongarra:1987:CBP

[DMW87] J. Dongarra, J. L. Martin, and J. Worlton. Computer benchmarking: Paths and pitfalls: The most popular way of rating computer performance can confuse as well as inform; avoid misunderstanding by asking just what the benchmark is measuring. IEEE Spectrum, 24 (7):38–43, July 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DaLio:1993:EHM

[DNV93] M. Da Lio, A. Nista, and F. Viola. An electromagnetic harmonic motor with continuous control of position and torque. In Anonymous [Ano93-31], pages 487–494. ISBN 0-947719-62-8. LCCN ????

Daoud:1989:HVF

[DO89] R. Daoud and F. Ozguner.

Highly vectorizable fault simulation on the Cray X-MP supercomputer. IEEE transactions on computer-aided design of in, 8(12):1362–??, December 1, 1989.

Dongarra:1985:PVCc

[Don85]  J. J. Dongarra.  Perfor-

mance of various computers using standard linear equations software in a FOR-TRAN environment. ACM SIGARCH Computer Architecture News, 13(1):3– 11, March 1985. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Dongarra:1986:SHP

[Don86] Jack J. Dongarra. A survey of high performance computers. In Bell [Bel86], pages 8–11. CODEN PCICDQ. ISBN 0-8186-0692-4 (paperback), 0-8186-4692-6 (microfiche). LCCN QA75.5.C58 1986. IEEE catalog number 86CH2285-.

Dongarra:1987:EPC

[Don87] J. J. Dongarra. Experimental parallel computing architectures. Special topics in supercomputing. Elsevier Science Publishers, Amsterdam, The Netherlands, 1987. ISBN 0-444-70234-2 (U.S.). xiii + 303 pp. LCCN QA76.5.E985 1987.

Dongarra:1991:LPH

[Don91] J. Dongarra. LAPACK: a portable high performance numerical library for linear algebra. In Anonymous [Ano91q], pages 73–76. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Donnini:1992:CPP

[Don92a] A. Donnini, editor. Commercial parallel processing: Seminar — February 1992, London. Unicom, Uxbridge, UK, 1992. ISBN ???? LCCN ????

Donnini:1992:SRB

[Don92b] A. Donnini. Supercomputing and real-time business. In Commercial parallel processing: Seminar — February 1992, London [Don92a], pages 1–37. ISBN ???? LCCN ????

Dongarra:1993:LAL

[Don93a] J. Dongarra. Linear algebra libraries for highperformance computers: a personal perspective. IEEE parallel and distributed technology: systems and applications, 1(1):17–24, February 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL

http://www.netlib.org/utk/
people/JackDongarra/PAPERS/
Linear-Algebra-Libraries-
for-High-Performance-Computers.
pdf.

Donlin:1993:GSR

[Don93b] Mike Donlin. Graphics system rivals supercomputer performance. Computer Design, 32(6):90–??, June 1993. CODEN CMPDAM. ISSN 0010-4566.

Donlin:1993:TDT

[Don93c] Mike Donlin. Team’s divideand-conquer tactics yield super results. Computer Design, 32(12):91–??, December 1, 1993. CODEN CMPDAM. ISSN 0010-4566.

Donndelinger:1994:SS

[Don94]  Peter Donndelinger.  Su-

percomputer science. The Science teacher (Washington, DC), 61(3):16–??, March 1994. ISSN 0036-8555.

Downey:1998:LJS

[Dow98] Allen B. Downey. Lachesis: a job scheduler for the Cray T3E. Lecture Notes in Computer Science, 1459:47– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1459/14590047.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1459/14590047.
                                        pdf.

Dozier:1992:SSS

[Doz92] H. Dozier. Supercomputers — superchips for supercomputing. IEEE Spectrum, 29 (9):66–68, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davies:1990:DOF

[DP90] A. M. Davies and R. Proctor. Developing and optimizing a 3-D-spectral/finite difference hydrodynamic model for the Cray X-MP. Computers and Fluids, 18(3):259– 270, 1990. CODEN CPFLBI. ISSN 0045-7930.

Duke:1991:PCM

[DP91] D. W. (Dennis W.) Duke and Walter S. Pritchard, editors. Proceedings of the Conference on Measuring Chaos in the Human Brain, April 3–5, 1991, at the Supercomputer Computations Research Institute, Florida State University, Tallahassee, FL. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1991. ISBN 981-02-0701-8. LCCN QP376.5 .C66 1991.

DeRose:1996:MFT

[DP96] L. De Rose and D. Padua.

A MATLAB to Fortran 90 translator and its effectiveness. In ACM [ACM96], pages 309–316. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Disz:1997:UEI

[DPS97] T. Disz, M. E. Papka, and R. Stevens. UbiWorld: An environment integrating virtual reality, supercomputing and design. In Hensgen [Hen97], pages 46–59. ISBN 0-8186-7879-8, 0-8186-7881-X. LCCN ????

Duff:1981:ESM

[DR81] I. S. Duff and J. K. Reid.

Experience of sparse matrix codes on the Cray-1. Computer Physics Communications, 26(3 & 4):

293–302, June 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Duff:1982:ESM

[DR82] I. S. Duff and J. K. Reid. Experience of sparse matrix codes on the Cray-

1. Computer Physics Communications, 26(3–4):293– 302, June 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465582901199.

DuBois:1991:DED

[DR91] Andrew J. DuBois and John Rasure. Design and evaluation of a distributed asynchronous VLSI crossbar switch controller for a packet switched supercomputer network. ACM SIGARCH Computer Architecture News, 19 (4):69–79, June 1991. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Dowd:1993:CTF

[DR93] Kevin Dowd and David R. Radin. The Cray T3D: From fringe to forefront. The DEC Professional, 12(12): 26–??, December 1993. CO-DEN DECPDJ. ISSN 0744-9216.

Dutta-Roy:1998:VMD

[DR98] A. Dutta-Roy. Virtual meetings with desktop conferencing. IEEE Spectrum, 35 (7):47–56, July 1998. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Draper:1988:CHI

[Dra88] Jesse M. Draper. Compiling for the Horizon instruction set. Technical report SRC-TR-88-006, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 8 pp.

Draper:1989:EDA

[Dra89] Jesse M. Draper. Extending dependence analysis to accommodate synchronous machines. Technical report SRC-TR-89-004, Supercomputing Research Center: IDA, Lanham, MD, USA, August 1989. 10 pp.

Draper:1990:FDR

[Dra90a] Richard N. Draper. A fast distributed routing algorithm for supertoroidal networks. Technical report SRC-TR-91-032, Supercomputing Research Center: IDA, Lanham, MD, USA, July 24, 1990. 26 pp.

Draper:1990:SN

[Dra90b] Richard N. Draper. Supertoroidal networks. Technical report SRC-TR-90-005, Supercomputing Research Center: IDA, Lanham, MD, USA, January 30, 1990. ii + 26 pp.

Draper:1991:OSN

[Dra91a] Richard N. Draper. An overview of supertoroidal networks. Technical report SRC-TR-91-035, Supercomputing Research Center: IDA, Lanham, MD, USA, January 17, 1991. 14 pp.

Draper:1991:SAG

[Dra91b] Richard N. Draper. Synchronous algorithms on gcs(4,1,1). Technical report SRC-TR-92-057, Supercomputing Research Center: IDA, Lanham, MD, USA, October 23, 1991. 28 pp.

Draper:1994:CWC

[Dra94a] Richard N. Draper. Conferences & workshops: Computational Mechanics; SIAM; Supercomputing ’94; DAGS’94. IEEE Computational Science & Engineering, 1(4):78– 79, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Draper:1994:CWS

[Dra94b] Richard N. Draper. Conferences & workshops: Supercomputing ’93. IEEE Computational Science & Engineering, 1(1):85–86, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Drach:1995:HII

[Dra95] N. Drach. Hardware implementation issues of data prefetching. In ACM [ACM95a], pages 245–254. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Draper:1996:CWSb

[Dra96a] Richard N. Draper. Conferences & workshops: Supercomputing ’95. IEEE Computational Science & Engineering, 3(1):88–90, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Draper:1996:CWSa

[Dra96b] Richard N. Draper. Conferences & workshops: Supercomputing ’96. IEEE Computational Science & Engineering, 3(4):84–86, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Diaz:2008:SHP

[DRAB08] Javier Díaz, Eduardo Ros, Rodrigo Agís, and Jose Luis Bernier. Superpipelined high-performance opticalflow computation architecture. Computer Vision and Image Understanding: CVIU, 112(3):262–273, December 2008. CODEN CVIUF4. ISSN 1077-3142 (print), 1090-235X (electronic).

Dahm:2020:SCE

[DRB^{+}20] J. P. Dahm, D. F. Richards, A. Black, A. D. Bertsch, L. Grinberg, I. Karlin, S. Kokkila-Schumacher, E. A. León, J. R. Neely, R. Pankajakshan, and O. Pearce. Sierra Center of Excellence: Lessons learned. IBM Journal of Research and Development, 64(3/4):2:1–2:14, May/ July 2020. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

delRosario:1994:HIM

[dRC94] J. M. del Rosario and A. N. Choudhary. Highperformance I/O for massively parallel computers: problems and prospects. Computer, 27(3):59–68, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drouffe:1995:SNS

[Dro95] J.-M. Drouffe. Simulating the nervous system of the nematode C. Elegans. In Herrmann et al. [HWP95], pages 291–300. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Dhekne:1994:APC

[DRRM94] P. S. Dhekne, K. Ramesh, K. Rajesh, and S. M. Mahajan. Anupam parallel computer. In Mahajan et al. [M^{+}94], pages 3–12. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Diaz-del-Rio:2016:EAL

[dRSGS16] Fernando Diaz del Rio, Javier Salmeron-Garcia, and Jose Luis Sevillano. Extending Amdahl’s Law for the cloud computing era. Computer, 49(2):14–22, February 2016. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
co/2016/02/mco2016020014-
abs.html.

Dowdy:1999:SIH

[DRSS99] L. W. Dowdy, E. Rosti, G. Serazzi, and E. Smirni. Scheduling issues in highperformance computing. ACM SIGMETRICS Performance Evaluation Review, 26(4):60– 69, March 1999. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dongarra:1986:LAH

[DS86a] J. J. Dongarra and D. C.

Sorensen. Linear algebra on high-performance computers. Applied Mathematics and Computation, 20:57–88, 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dongarra:1986:FPA

[DS86b] J. J. Dongarra and D. C.

(Danny C.) Sorensen. A fully parallel algorithm for the symmetric eigenvalue problem. Technical Report CSRD-542, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 22 pp.

Decyk:1989:SC

[DS89] Viktor K. Decyk and Joan E. Slottow. Supercomputers in the classroom. Computers in Physics, 3(2):50– ??, March 1989. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.168316.

DasGupta:1994:SCB

[DS94a] S. Das Gupta and V. Sridhar. Spelling correction based on hidden Markov models. In Balakrishnan [Bal94], pages 163–173. ISBN 0-07-462044-4. LCCN ????

Demmel:1994:PGA

[DS94b] J. Demmel and S. Smith. Parallelizing a global atmospheric chemical tracer model. In IEEE [IEE94c], pages 718–725. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Dwivedi:1994:OIM

[DS94c] H. P. Dwivedi and G. Singh. 2-D object inspection model using computer vision. In Balakrishnan [Bal94], pages 370–372. ISBN 0-07-462044-4. LCCN ????

Dikaiakos:1996:PSC

[DS96a] M. D. Dikaiakos and J. Stadel. A performance study of cosmological simulations on message-passing and sharedmemory multiprocessors. In ACM [ACM96], pages 94– 101. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Duff:1996:DNF

[DS96b] I. S. Duff and J. A. Scott.

The design of a new frontal code for solving sparse, unsymmetric systems. ACM Transactions on Mathematical Software, 22(1):30–45, March 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toms/1996-22-1/
p30-duff/.

Dippel:1996:MBR

[DSB96] S. Dippel, L. Samson, and G. G. Batrouni. Motion of a ball on a rough inclined plane. In Wolf et al. [WSB96], pages 353–358. ISBN 981-02-2635-7. LCCN ????

Dongarra:2005:HPC

[DSSS05] Jack Dongarra, Thomas Sterling, Horst Simon, and Erich Strohmaier. Highperformance computing: Clusters, constellations, MPPs, and future directions. Computing in Science and Engineering, 7(2):51–59, March/ April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.

                                        computer.org/dl/mags/cs/
                                        2005/02/c2051.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/cs/2005/02/c2051.pdf.

Dekker:1996:HCT

[DSZ96] L. Dekker, W. Smit, and J. C. Zuidervaart, editors. HPCN challenges in telecomp and telecom: parallel simulation of complex systems and large-scale applications: International conference — June 1996, Delft, the Netherlands. Elsevier Science, Amsterdam; New York, 1996. ISBN 0-444-82559-2. LCCN ????

Dutt:1996:TAH

[DT96] S. Dutt and N. Trinh. Are there advantages to highdimension architectures?: Analysis of k-ary n-cubes for the class of parallel divideand-conquer algorithms (out of sequence due to late submission). In ACM [ACM96], pages 398–252. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Dongarra:1997:PTW

[DT97] J. J. Dongarra and Bernard Tourancheau, editors. Proceedings of the Third Workshop on Environments and Tools for Parallel Scientific Computing, volume 11(2) of International Journal of Supercomputer Applications and High Performance Computing. Sage Science Press, Thousand Oaks, CA, USA, 1997. LCCN QA 76.5 I55 v.11 no.2 1997.

Dennis:2008:SCS

[DT08] J. M. Dennis and H. M.

Tufo. Scaling climate simulation applications on the IBM Blue Gene/L system. IBM Journal of Research and Development, 52(1/2):117– ??, January/March 2008. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/521/dennis.html.

Desgagne:2000:PME

[DTV00] Michel Desgagné, Stephen Thomas, and Michel Valin. Performance of MC2 and the ECMWF IFS forecast model on the Fujitsu VPP700 and NEC SX-4M. Scientific Programming, 8(1):23–30, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C3%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Dubash:1987:SEB

[Dub87] Rumi Minoo Dubash. A study of extended Basic Linear Algebra Subprograms on the NEC SX-2 supercomputer. Thesis (M.S.), Dept. of Computer Science, University of Houston, Houston, TX, USA, 1987. xiv + 172 pp.

DuBois:1990:DSD

[DuB90] Andrew J. DuBois. Design and simulation of a distributed asynchronous VLSI crossbar switch controller for a packet switched supercomputer network. Thesis (M.S.), University of New Mexico, Albuquerque, NM, USA, 1990. viii + 95 pp.

Dueltgen:1989:TSS

[Due89] Ronald R. Dueltgen. The technophile — supercomputers and supercomputing. Database, 13(4):106–??, August 1989. CODEN DTB-SDQ. ISSN 0162-4105.

Duff:1982:SSLb

[Duf82] Iain S. Duff. The solution of sparse linear equations on the CRAY-1. CRAY Channels, 4(3):4–9, 1982. CODEN CRCHE8.

Duff:1984:SE

[Duf84] Iain S. Duff. Supercomputers in Europe. Parallel Computing, 1(3–4):321– 324, December 1984. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Duff:1985:USE

[Duf85]  Iain S. Duff.  The use

of supercomputers in Europe. Computer Physics Communications, 37(1–3): 15–25, July 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465585901328.

Duff:1990:SLS

[Duf90] Iain S. Duff. The solution of large-scale least-squares problems on supercomputers. Annals of Operations Research, 22:241–252, 1990. CODEN AOREEV. ISSN 0254-5330 (print), 1572-9338 (electronic).

Duff:1991:SSL

[Duf91]  I. S. Duff.  Solution of

sparse linear equations on supercomputers. In Anonymous [Ano91q], pages 252–?? ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Duff:2000:IHP

[Duf00] Iain S. Duff. The impact of high-performance computing in the solution of linear systems: trends and problems. Journal of Computational and Applied Mathematics, 123(1–2):515–530, 2000. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Numerical analysis 2000, Vol. III. Linear algebra.

Duke:1991:CSA

[Duk91] D. W. Duke. Computational science at the Supercomputer Computations Research Institute. International Journal of Supercomputer Applications, 5(3):4–??, Fall 1991. CODEN IJSAE9. ISSN 0890-2720.

Dumler:1997:ADL

[Dum97] Christopher M. Dumler. Anti-dumping laws trash supercomputer competition. Cato Institute briefing papers 32, Cato Institute, Washington, D.C., 1997. 14 pp.

Dunham:1992:SFW

[Dun92] Charles B. Dunham. Surveyor’s Forum: “What every computer scientist should know about floating-point arithmetic”. ACM Computing Surveys, 24(3):319, September 1992. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [Gol91b, Gol91a, Wic92].

Duncan:1999:BRU

[Dun99] Ralph Duncan. Book review: Understanding Parallel Supercomputing. IEEE Concurrency, 7(3):93–94, July/ September 1999. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1999/pdf/p3091.
pdf.

Dupuis:1986:SSC

[Dup86] M. (Michel) Dupuis, editor.

Supercomputer simulations in chemistry: proceedings of the Symposium on Supercomputer Simulations in Chemistry, held in Montreal August 25–27, 1985, volume 44 of Lecture notes in chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 3-540-17178-9. LCCN ????

Dupuis:1987:SSC

[Dup87] M. (Michel) Dupuis, editor. Supercomputer simulations in chemistry: proceedings of the Symposium on Supercomputer Simulations in Chemistry, held in Montreal, August 25–27, 1985, number 44 in Lecture notes in chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-17178-9 (paperback). LCCN QD462.A1S98 1985.

Dongarra:2012:HPC

[DvdS12] J. J. Dongarra and A. J.

van der Steen. Highperformance computing systems: Status and outlook. Acta Numerica, 21: 379–474, 2012. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Dunigan:2005:PEC

[DVWW05] Thomas H. Dunigan, Jr., Jeffrey S. Vetter, James B. White III, and Patrick H. Worley. Performance evaluation of the Cray X1 distributed shared-memory architecture. IEEE Micro, 25(1):30–40, January/ February 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/mi/
2005/01/m1030.htm;   http:
//csdl.computer.org/dl/
mags/mi/2005/01/m1030.pdf.

Delic:1997:ABC

[DW97] G. Delic and M. F. Wheeler, editors. – 1995 Aug: Bay City; MI. SIAM Press, Philadelphia, PA, USA, 1997. ISBN 0-89871-378-1. LCCN ????

Dmitruk:2001:SPF

[DWM^{+}01] P. Dmitruk, L.-P. Wang, W. H. Matthaeus, R. Zhang, and D. Seckel. Scalable parallel FFT for spectral simulations on a Beowulf cluster. Parallel Computing, 27(14):1921–1936, December 31, 2001. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
35/21/47/44/31/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/47/44/
31/article.pdf.

Dozier:1992:TQE

[DWV92] Harold Dozier, Jan Wikstrom, and Tony Vacca. In their quest for ever-higher processing rates, supercomputer designers are developing remarkable semiconductor technologies. IEEE Spectrum, 29(9):66–68, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Droegemeier:1993:ESA

[DXJM93] K. K. Droegemeier, M. Xue, K. Johnson, and K. Mills. Experiences with the scalableparallel ARPS cloud/mesoscale prediction model on massively parallel and workstation cluster architectures. In Hoffmann and Kauranne [HK93b], pages 99–129. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Davis:1990:SNP

[DY90] Timothy Alden Davis and Pen-Chung Yew. A stable non-deterministic parallel algorithm for general unsymmetric sparse LU factorization. Technical Report CSRD 908, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1990. 31 pp.

Deng:2013:ALP

[DZM^{+}13] Yuefan Deng, Peng Zhang, Carlos Marques, Reid Powell, and Li Zhang. Analysis of Linpack and power efficiencies of the world’s TOP500 supercomputers. Parallel Computing, 39(6– 7):271–279, June/July 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819113000513.

El-Araby:2009:EPR

[EAGEG09] Esam El-Araby, Ivan Gonzalez, and Tarek El-Ghazawi. Exploiting partial runtime reconfiguration for highperformance reconfigurable computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 1(4):21:1–21:??, January 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

El-Araby:2011:FEH

[EAMEG11] Esam El-Araby, Saumil G. Merchant, and Tarek El-Ghazawi. A framework for evaluating high-level design methodologies for highperformance reconfigurable computers. IEEE Transactions on Parallel and Distributed Systems, 22(1):33– 45, January 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Erbacci:1995:PCM

[EAMS95a] G. Erbacci, R. Ansaloni, B. Montagnini, and R. Scardovelli. Porting a coarsemesh neutron diffusion code on a Cray T3D massively parallel computer. Lecture Notes in Computer Science, 919:318–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erbacci:1995:PCN

[EAMS95b] G. Erbacci, R. Ansaloni, B. Montagnini, and R. Scardovelli. Porting a coarsemesh neutron diffusion code on a Cray T3D massively parallel computer. Lecture Notes in Computer Science, 919:318–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eigenmann:1991:ESP

[EB91] R. Eigenmann and William Blume. An effectiveness study of parallelizing compiler techniques. Technical Report CSRD 1090, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 9 pp.

Ebadifard:2018:PBT

[EB18] Fatemeh Ebadifard and Seyed Morteza Babamir. A PSO-based task scheduling algorithm improved using a load-balancing technique for the cloud computing environment. Concurrency and Computation: Practice and Experience, 30(12): ??, June 25, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

                                        onlinelibrary.wiley.com/
                                        doi/abs/10.1002/cpe.4368.

Eichenberger:2020:HCG

[EBB^{+}20] A. E. Eichenberger, G.-T. Bercea, A. Bataev, L. Grinberg, and J. K. O’Brien. Hybrid CPU/GPU tasks optimized for concurrency in OpenMP. IBM Journal of Research and Development, 64(3/4):13:1–13:14, May/ July 2020. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Ernenwein:1988:VSC

[EBS88] Rene Ernenwein, Marc Benard, and Isaiah Shavitt. Vectorizing a sequence of conditional branches: The calculation of the class index of two-electron repulsion integrals on Cray computers. Computer Physics Communications, 48(2):175– 180, February 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Eisenhauer:2002:NDR

[EBS02] Greg Eisenhauer, Fabián E. Bustamante, and Karsten Schwan. Native data representation: An efficient wire format for high-performance distributed computing. IEEE Transactions on Parallel and Distributed Systems, 13(12):1234–1246, December 2002. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
td/2002/12/l1234abs.htm;
 http://csdl.computer.org/
dl/trans/td/2002/12/l1234.
htm;   http://csdl.computer.
org/dl/trans/td/2002/12/
l1234.pdf.

Ecer:1996:PCP

[Ece96] A. Ecer, editor. Parallel CFD ’95: Parallel computational fluid dynamics, implementations and results using parallel computers: Conference — June 1995, Pasadena; CA, PARALLEL COMPU-TATIONAL FLUID DY-NAMICS 1995. Elsevier, Amsterdam, The Netherlands, 1996. ISBN 0-444-82322-0. LCCN ????

Eckland:1992:SGC

[Eck92a] Denise J. Eckland. Supercomputing and the grand challenges of science, 1992. 1 videodisc (50 min.) sd. 1/2 in.

Ecklund:1992:SGC

[Eck92b] Denise J. Ecklund. Supercomputing and the Grand Challenges of Science, 1992. 1 videocassette (59 min.).

Ecklund:1993:MPP

[Eck93] D. J. Ecklund. Massively parallel processing on the INTEL Paragon system: One tool in achieving the goals of the Human Genome Project. In Lim et al. [L^{+}93], pages 249–262. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Eichenberger:1995:OMS

[EDA95] A. E. Eichenberger, E. S. Davidson, and S. G. Abraham. Optimum modulo schedules for minimum register requirements. In ACM [ACM95a], pages 31–40. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Edelson:1992:SNS

[Ede92] B. I. Edelson. Supercomputer networking for space science applications: final report. George Washington University, Institute for Applied Space Research, Washington, DC, USA, February 21, 1992. v + 37 pp.

Edelman:1994:LNL

[Ede94a] A. Edelman. Large numerical linear algebra in 1994: The continuing influence of parallel computing. In IEEE [IEE94c], pages 781–787. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Edelsohn:1994:GEO

[Ede94b] D. J. Edelsohn. A generalized expression optimization hook for C++ on highperformance architectures. In IEEE [IEE94c], pages 381– 387. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Ellison:2010:SAW

[EDJ^{+}10] Steven Ellison, John Dean, Michael Johnson, Cindy Prebola, Charles Fabozzi, and Alexis Cenko. Supplying air warfare capability through high-performance computing. Computing in Science and Engineering, 12(5):18– 26, September/October 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Edwards:1997:CWH

[Edw97] Lynne K. Edwards. Conferences & workshops: High-Performance Computer Applications in the Behavioral Sciences. IEEE Computational Science & Engineering, 4(1):100–101, January/March 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1997/pdf/c1098.
pdf.

Edirisooriya:1993:EVA

[EE93] Samantha Edirisooriya and Geetani Edirisooriya. Enchancing vector access performance in Cray X-MP memory system. In 1993 IEEE Compcon Spring (Feb 22–26 1993: San Francisco, CA, USA), pages 569–576. IEEE, Piscataway, NJ, USA, 1993. ISBN 0-7803-1294-5. LCCN ???? IEEE catalog number 93CH3251-6.

Enenkel:2005:CMF

[EFG^{+}05] R. F. Enenkel, B. G. Fitch, R. S. Germain, F. G. Gustavson, A. Martin, M. Mendell, J. W. Pitera, M. C. Pitman, A. Rayshubskiy, F. Suits, W. C. Swope, and T. J. C. Ward. Custom math functions for molecular dynamics. IBM Journal of Research and Development, 49(2/3):465–474, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/enenkel.
pdf.

Eickermann:2000:GBS

[EFH^{+}00] Thomas Eickermann, Wolfgang Frings, Friedel Hossfeld, Stefan Posse, and Gernot Goebbels. Global broadcast: Supercomputerenhanced functional MRI of the human brain. IEEE Concurrency, 8(1):11–13, January/March 2000. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/
books/pd2000/pdf/p1011.
pdf.

Ebisuzaki:1991:GSP

[EFIM91] T. Ebisuzaki, T. Fukushige, T. Ito, and J. Makino. GRAPE; special purpose computer for simulations of many-body systems. In Anonymous [Ano91q], pages 9–18. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Ehrhardt:1993:RRT

[EFPSS93] J. Ehrhardt, F. Fischer, J. Paesler-Sauer, and O. Schuele. RODOS and RESY: Two integrated real-time on-line decision support systems for nuclear emergencies. In Kusters et al. [KSW93], pages 319– 330 (or 792–??). ISBN 3-923704-11-9. LCCN ???? Two volumes.

Eleftheriou:2005:SFF

[EFR^{+}05] M. Eleftheriou, B. G. Fitch, A. Rayshubskiy, T. J. C. Ward, and R. S. Germain. Scalable framework for 3D FFTs on the Blue Gene/L supercomputer: Implementation and early performance measurements. IBM Journal of Research and Development, 49(2/3):457– 464, ???? 2005. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/eleftheriou.
pdf.

El-Ghazawi:2008:PHP

[EGEAH^{+}08] Tarek El-Ghazawi, Esam El-Araby, Miaoqing Huang, Kris Gaj, Volodymyr Kindratenko, and Duncan Buell. The promise of high-performance reconfigurable computing. Computer, 41(2):69–76, February 2008. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eggers:1994:SCG

[Egg94]  Ron Eggers.  Supercom-

puter centers: a glimpse at tomorrow’s imaging technology. Photo Electronic Imaging, 37(3):16–??, March 1, 1994. CODEN PELIE6. ISSN 0146-0153.

Ellsworth:2006:CVP

[EGH^{+}06] David Ellsworth, Bryan Green, Chris Henze, Patrick Moran, and Timothy Sandstrom. Concurrent visualization in a production supercomputing environment. IEEE Transactions on Visualization and Computer Graphics, 12(5):997–1004, September/October 2006. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Eigenmann:2002:SHN

[EGJ^{+}02] Rudolf Eigenmann, Greg Gaertner, Wesley Jones, Hideki Saito, and Brian Whitney. SPEC HPC2002: The next high-performance computer benchmark. Lecture Notes in Computer Science, 2327:7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2327/23270007.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2327/23270007.
pdf.

Egecioglu:1987:FPP

[EGK87a] Omer Nuri Egecioglu, E. J.

(Efstratios J.) Gallopoulos, and Cetin K. Koc. Fast and practical parallel polynomial interpolation. Technical Report CSRD 646, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1987. 23 + [7] pp.

Egecioglu:1987:PHI

[EGK87b] Omer Nuri Egecioglu, E. J.

(Efstratios J.) Gallopoulos, and Cetin K. Koc. Parallel Hermite interpolation: an algebraic approach. Technical Report CSRD 671, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 27 pp.

Egecioglu:1989:FCD

[EGK89a] Omer Nuri Egecioglu, E. J.

(Efstratios J.) Gallopoulos, and Cetin K. Koc. Fast computation of divided differences and parallel Hermite interpolation. Technical Report CSRD 800, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1989. 19 pp.

Egecioglu:1989:PMF

[EGK89b] Omer Nuri Egecioglu, E. J. Efstratios J. Gallopoulos, and Cetin K. Koc. A parallel method for fast and practical high-order Newton interpolation. Technical Report CSRD 921, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1989. 22 pp.

Emrath:1988:ESA

[EGP88] Perry Alan Emrath, Sanjoy Ghosh, and David A. Padua. Event synchronization analysis for debugging parallel programs. Technical Report CSRD 839, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 16, 1988. 18 + [5] pp.

Emrath:1992:DNP

[EGP92] Perry A. Emrath, Sanjoy Ghosh, and David A. Padua. Detecting nondeterminacy in parallel programs. Technical Report CSRD 1118, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. 69–77 pp.

Eshaghian:1997:ASD

[EH97a] M. M. Eshaghian and L. Hai.

Application specific design of the optical communication topology in ORM. In Roller [Rol97], pages 399– 406. ISBN 0-947719-88-1 (paperback). LCCN ????

Eshaghian:1997:EEP

[EH97b] M. M. Eshaghian and L. Hai.

An efficient electro-optical parallel architecture. In Roller [Rol97], pages 391– 398. ISBN 0-947719-88-1 (paperback). LCCN ????

Eshaghian:1997:FPI

[EH97c] M. M. Eshaghian and L. Hai.

A fast parallel image convexity algorithm. In Roller [Rol97], pages 381–390. ISBN 0-947719-88-1 (paperback). LCCN ????

Eldredge:1997:HPP

[EHF^{+}97] Michael Eldredge, Thomas J. R. Hughes, Robert M. Ferencz, Steven M. Rifai, Arthur Raefsky, and Bruce Herndon. High-performance parallel computing in industry. Parallel Computing, 23(9):1217– 1233, November 3, 1997.

CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
parco/cas_sub/browse/browse.
cgi?year=1997&volume=23&
issue=9&aid=1207.

Elmasri:1995:TCL

[EHG95] N. Elmasri, H. H. J. Hum, and G. R. Gao. The threaded communication library: Preliminary experiences on a multiprocessor with dual-processor nodes. In ACM [ACM95a], pages 195– 199. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Etter:2001:ECH

[EHG01] Delores M. Etter, Charles J. Holland, and John Grosh. Export control of highperformance computing: Analysis and alternatives. Computing in Science and Engineering, 3(3):24–31, May/ June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://dlib.computer.org/
cs/books/cs2001/pdf/c3024.
pdf;   http://www.computer.
org/cse/cs1999c3024abs.
htm; pdf/c3024.pdf.

Ewinger:1989:MMM

[EHHS89] W. Ewinger, O. Haan, E. Haupenthal, and C. Siemers. Modelling and measurement of memory access in SIEMENS VP supercomputers. Parallel Computing, 11 (3):361–365, ???? 28, 1989. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Eswar:1994:MPS

[EHS94] K. Eswar, C.-H. Huang, and P. Sadayappan. Memoryadaptive parallel sparse Cholesky factorization. In IEEE [IEE94c], pages 317– 323. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

ETHZ:1991:SP

[Eid91] Eidgenossische Technische Hochschule Zurich. Supercomputing projects, 1989/90. Technical report, Informatik-Kommission der ETH Zurich, Zurich, 1991. x + 92 pp.

Eigenmann:1990:CFC

[Eig90a] R. Eigenmann. Cedar Fortran and its compiler. Technical Report CSRD 966, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1990. 12 pp.

Eigenmann:1990:CFR

[Eig90b] R. Eigenmann. Cedar Fortran and its restructuring compiler. Technical Report CSRD 1041, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1990. 22 pp.

Eigenmann:1991:EAP

[Eig91] R. Eigenmann. Experience in the automatic parallelization of four Perfect-Benchmark programs. Technical Report CSRD 1193, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1991. 19 pp.

Eigenmann:1992:TMO

[Eig92]  R. Eigenmann.  Toward

a methodology of optimizing programs for highperformance computers. Technical Report CSRD 1178, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1992. 17 pp.

Eigenmann:2001:PEB

[Eig01] Rudolf Eigenmann, editor. Performance evaluation and benchmarking with realistic applications. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-05066-8. 293 pp. LCCN QA76.9.E94 .P43 2001. US$40.00, UK£27.50.

Eijkhout:1990:API

[Eij90a] Victor Eijkhout. Analysis of parallel incomplete point factorizations. Technical Report CSRD 1045, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1990. 13 pp.

Eijkhout:1990:BMI

[Eij90b] Victor Eijkhout. Beware of modified incomplete point factorizations. Technical Report CSRD 1048, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1990. 8 pp.

Eijkhout:1991:BUM

[Eij91] Victor Eijkhout. Beware of unperturbed modified incomplete factorizations. Technical Report CSRD 1109, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 9 pp.

Eisele:1995:TLP

[Eis95] M. Eisele. Tentative learning principles of pyramidal neurons. In Herrmann et al. [HWP95], pages 415–420. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Eriksson:1997:MCP

[EJ97] A. Eriksson and B. Jacobson. Method for comparison of powertrain concepts using dynamic simulation. In Roller [Rol97], pages 273–

280. ISBN 0-947719-88-1 (paperback). LCCN ????

Eisenbeis:1990:CTO

[EJL90] C. Eisenbeis, W. Jalby, and A. Lichnewsky. Compiler techniques for optimizing memory and register usage on the Cray 2. International Journal of High Speed Computing, 2(2):193–??, June 1990. CODEN IHSCEZ. ISSN 0129-0533.

Evangelinos:1996:PCB

[EK96] Constantinos Evangelinos and George Em Karniadakis. Parallel CFD benchmarks on Cray computers. Parallel Algorithms and Applications, 9(3–4):273–298, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http:

//www.informaworld.com/
smpp/content~content=a778707985.

Eberl:1999:PCP

[EKTB99] M. Eberl, W. Karl, C. Trinitis, and A. Blaszczyk. Parallel computing on PC clusters — an alternative to supercomputers for industrial applications. In Dongarra et al. [DLM99], pages 493–

498. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Encarnacao:1990:DSA

[EKZ90] J. Encarnacao, G. Koberle, and Ning Zhang. Distributed supercomputing to achieve real-time representation and manipulation of complex scenes. Computers in industry, 14(1):23– 34, ???? 1990. CO-DEN CINUD4. ISSN 0166-3615 (print), 1872-6194 (electronic).

Eles:1993:LDS

[Ele93] P. Eles. Language and development system for supercomputer programming. In Anonymous [Ano93g], pages 8–15. ISBN ???? LCCN ????

Elmer:1993:SAT

[Elm93] K.-H. Elmer. Supercomputer applications in 2-D and 3-D transient dynamics. In Brebbia and Power [BP93], pages 313–322. ISBN 1-85166-845-4, 1-85312-236-X, 1-56252-160-8. LCCN TA345.A66 1993.

Elmer:1995:MDP

[Elm95a] K.-H. Elmer, editor. Machine dynamics and production automation: International workshop — April 1995, Szczecin, Poland. Technical University of Szczecin, Szczecin, Poland, 1995. ISBN 83-86359-12-9 (??bad ISBN??). LCCN ????

Elmer:1995:SFA

[Elm95b] K.-H. Elmer. Simulation with fast algorithms on supercomputer. In Machine dynamics and production automation: International workshop — April 1995, Szczecin, Poland [Elm95a], pages 111– 128. ISBN 83-86359-12-9 (??bad ISBN??). LCCN ????

Els:1989:CCS

[Els89] Sharon A. (Sharon Ann) Els. Configuration and competitiveness in the supercomputer industry. Thesis (M.S.), Massachusetts Institute of Technology, Sloan School of Management, Cambridge, MA, USA, 1989. 72 pp. Supervised by Donald Lessard.

Elster:2002:HPC

[Els02]  Anne C. Elster.  High-

performance computing: Past, present, and future. Lecture Notes in Computer Science, 2367:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2367/23670433.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2367/23670433.
pdf.

Elster:2021:EFA

[Els21] A. C. Elster. The European factor: From ARM to Atos. Computing in Science and Engineering, 23 (1):102–105, 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ewald:1978:HPG

[EM78] Robert H. Ewald and Lynn D.

Maas. A high performance graphics system for the CRAY-1. Computer Graphics, 12(3):76–81, August 1978. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Emrath:1991:MXP

[EM91] Perry A. Emrath and Bret A.

Marsolf. Mdb — Xylem parallel debugger user’s guide. Technical Report CSRD 1180, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 12, 1991. 27 pp.

Elzen:1994:SLS

[EM94a] Boelie Elzen and Donald MacKenzie. The social limits of speed: The development and use of supercomputers. IEEE annals of the history of computing, 16(1):46–61, Spring 1994. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

//dlib.computer.org/an/
books/an1994/pdf/a1046.

                                        pdf;   http://www.computer.
                                        org/annals/an1994/a1046abs.
                                        htm.

Engl:1994:IPO

[EM94b] Heinz W. Engl and Joyce McLaughlin, editors. Inverse problems and optimal design in industry: proceedings of the conference, July 8– 10, 1993, Philadelphia, PA, USA, volume 10 of European Consortium for Mathematics in Industry. Teubner, Stuttgart, Germany; Leipzig, Germany, 1994. ISBN 3-519-02179-X. LCCN ????

Emmen:1984:ISA

[Emm84] Ad Emmen. International supercomputer applications symposium. Parallel Computing, 1(3–4):317–319, December 1984. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167819184902618.

Emmen:1985:SAP

[Emm85] A. H. L. (Ad H. L.) Emmen, editor. Supercomputer applications: proceedings of the International Supercomputer Applications Symposium, Amsterdam, The Netherlands, November 7– 9, 1984. Elsevier Science Publishers, Amsterdam, The Netherlands, 1985. ISBN 0-444-87752-5. LCCN QA76.5 .I6 1984.

Emrath:1989:PL

[Emr89] Perry A. Emrath. Program laborious. Technical Report CSRD 873, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1989. 10 pp.

El-Moursy:2011:IPA

[EMS11] Ali A. El-Moursy and Fadi N.

Sibai. Image processing applications performance study on Cell BE and Blue Gene/L. Concurrency and Computation: Practice and Experience, 23(4):351–371, March 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Entacher:1999:CSR

[Ent99]  Karl Entacher.  On the

CRAY-system random number generator. Simulation, 72(3):163–169, March 1999. CODEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic). URL http://

sim.sagepub.com/content/
72/3/163.abstract.

Escaig:1991:ATM

[EO91] Yves Escaig and Wilfried Oed. Analysis tools for microand autotasking programs on Cray multiprocessor systems. Parallel Computing, 17(12): 1425–1433, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Eichenberger:2013:ELO

[EO13] A. E. Eichenberger and K. O’Brien. Experimenting with low-overhead OpenMP runtime on IBM Blue Gene/ Q. IBM Journal of Research and Development, 57(1/2):8:1–8:8, January– March 2013. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

EP:1997:FEA

[EP 97] EP Innovations, editor. Frontiers in Education 1997: 27th Annual Conference: proceedings, November 5–8, 1997, Pittsburgh, PA: Teaching and learning in an era of change. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-4087-6 (casebound), 0-7803-4086-8 (softbound), 0-7803-4088-4 (microfiche), 0-7803-4089-2 (CD-ROM). ISSN 0190-5848. LCCN T62 .F76 1997. Three volumes. IEEE catalog number: 97CH36099.

Egan:1994:PSD

[ER94] G. K. Egan and G. D. Riley. Parallelisation of the SDEM distinct element stress analysis code on the KSR-1. In Anonymous [Ano94-134], pages 85–92. ISBN ???? LCCN ????

Emmerich:1996:ATF

[ER96] H. Emmerich and E. Rank.

Analyzing traffic flow by a cellular automaton. In Wolf et al. [WSB96], pages 205– 210. ISBN 981-02-2635-7. LCCN ????

Ercegovac:1988:HSA

[Erc88] M. D. Ercegovac. Heterogeneity in supercomputer architectures. Parallel Computing, 7(3):367–372, September 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Erwin:1984:MYC

[Erw84] Dietmar W. Erwin. Making your Cray talk to your IBM and your users. In SEAS [SEA84], pages 342– 351. LCCN ????

El-Sayed:1988:FLC

[ES88] Mohamed E. M. El-Sayed.

Fatigue life consideration in automotive structural optimization. In Marino [Mar88b], pages 59–67. ISBN ???? LCCN TL240 .I528 1988.

Eisenbeis:1992:GAD

[ES92] C. Eisenbeis and J.-C. Sogno.

A general algorithm for data dependence analysis. In ACM [ACM92b], pages 292–302. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Eisenhauer:1996:DAP

[ES96] Greg Eisenhauer and Karsten Schwan. Design and analysis of a parallel molecular dynamics application. Journal of Parallel and Distributed Computing, 35(1): 76–90, May 25, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0070/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0070/production/pdf.

Endou:1993:CDA

[ESMH93] A. Endou, A. Saiki, T. Miki, and Y. Himeno. Conceptual design of autonomous operation system for nuclear power plants [invited]. In Kusters et al. [KSW93], pages 842– 851. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Ess:1990:FRC

[Ess90] Michael Ess. Folding RNA on the Cray-2. IEEE, Piscataway, NJ, USA, 1990. ISBN 0-8186-2056-0. 103–111 pp. LCCN ???? IEEE catalog number 90CH2916-5.

El-Sharkawy:1994:SDP

[ESTA94] M. El-Sharkawy, Wenlong Tsang, and M. Aburdene. Short data parallel vector Slant transform. Journal of Parallel and Distributed Computing, 23(1):

27–36, October 1994. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1116/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1116/production/pdf.

Esposito:1996:PHB

[ET96] A. Esposito and L. Tarricone. Parallel heuristics for bandwidth reduction of sparse matrices with IBM SP2 and Cray T3D. Lecture Notes in Computer Science, 1184:239–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1997:DBG

[Eva97] A. Evans. Deep Blue: Garry Kasparov was recently defeated by IBM’s Deep Blue supercomputer. Personal computer world, 20(8):108– ??, ???? 1997. CODEN PC-WODU. ISSN 0142-0232.

Efthivoulidis:1998:FTC

[EVM^{+}98] Giorgos Efthivoulidis, Evangelos A. Verentziotis, Apostolos N. Meliones, Theodora A. Varvarigou, Antonios Kontizas, Geert Deconinck, and Vincenzo De Florio. Faulttolerant communication in embedded supercomputing. IEEE Micro, 18(5):42–52, September/October 1998.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi1998/pdf/m5042.
                                        pdf;   http://www.computer.
                                        org/micro/mi1998/m5042abs.
                                        htm.

Evans:1990:SS

[EW90] R. G. Evans and S. Wilson, editors. Supercomputational Science. Springer US, Boston, MA, USA, 1990. ISBN 0-306-43663-9, 1-4684-5820-5 (ebook). 349 pp. LCCN QA76.5 .S89437 1990.

Ewald:1989:PFC

[Ewa89]  Robert H. Ewald.  Per-

spectives from the field: Cray Research. Computers in Physics, 3(1):33– ??, January 1989. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4822813.

Ewald:1996:LCR

[Ewa96] Bob Ewald. Letters: Cray Research stands by its claim. IEEE Computational Science & Engineering, 3(4): 95, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). The President of Cray Research comments on the dumping charges leveled against NEC in a proposed NCAR supercomputer acquisition. See [Smi96c] for an overview.

Ewing:1997:NMI

[Ewi97] R. E. Ewing. The need for multidisciplinary involvement in groundwater contaminant simulations. In Delic and Wheeler [DW97], pages 227–246. ISBN 0-89871-378-1. LCCN ????

Evangelinos:2013:DPC

[EWS^{+}13] C. Evangelinos, R. E. Walkup, V. Sachdeva, K. E. Jordan, H. Gahvari, I.-H. Chung, M. P. Perrone, L. Lu, L.-K. Liu, and K. Magerlein. Determination of performance characteristics of scientific applications on IBM Blue Gene/Q. IBM Journal of Research and Development, 57(1/2):9:1–9:12, January– March 2013. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Excell:1991:ONE

[EY91] P. S. Excell and K. W. Yip.

Optimisation of the numerical electromagnetics code on a four-processor Cray X-MP computer. IEE conference publication, 350:55–58, 1991. CODEN IECPB4. ISSN 0537-9987 (??invalid ISSN checksum??).

Fiedler:1993:NSA

[FA93] F. Fiedler and G. Adrian. Numerical simulation of atmospheric currents and turbulent transport of air pollution in the mesoscale gamma. In Kusters et al. [KSW93], pages 798–?? ISBN 3-923704-11-9. LCCN ???? Two volumes.

Fahringer:1994:UPG

[Fah94] T. Fahringer. Using the P03T to guide the parallelization and optimization effort under the Vienna Fortran compilation system. In IEEE [IEE94c], pages 437– 444. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Fields:1993:IGG

[FAKD93] C. Fields, M. D. Adams, A. R. Kerlavage, and M. Dubnick. Identification of genes in genomic and EST sequences. In Lim et al. [L^{+}93], pages 429–434. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Fang:1987:DPS

[Fan87] Zhixi Fang. Dynamic processor self-scheduling for general parallel nested loops. Technical Report CSRD-637, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 28 pp.

Farhat:1990:RSS

[Far90] C. Farhat. Redesigning the skyline solver for parallel/ vector supercomputers. International Journal of High Speed Computing, 2(3):223– 238, September 1990. CO-DEN IHSCEZ. ISSN 0129-0533.

Fatoohi:2010:ANA

[Fat10] Rod Fatoohi. Assessment of NSF application benchmarks on SGI Altix machines and Cray Opteron cluster. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 25(4): 315–329, 2010. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Fazio:1987:RMM

[Faz87] D. Fazio. It’s really much more fun building a supercomputer than it is simply inventing one. In IEEE, editor, Digest of papers: thirtysecond IEEE Computer Society international conference, Cathedral Hill Hotel, San Francisco, California, February 23–27, Compcon, spring 87: Intellectual leverage. Washington, DC, pages 102– 105. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-8186-0764-5 (paper), 0-8186-4764-7 (microfiche), 0-8186-8764-9 (case). LCCN QA76 .I49 1987. IEEE catalog number 87CH2409-1. Computer Society Order number 764.

Ferguson:1991:PTN

[FB91a] Helaman R. P. Ferguson and David H. Bailey. A polynomial time, numerically stable integer relation algorithm. Technical report SRC-TR-92-066, Supercomputing Research Center: IDA, Lanham, MD, USA, December 30, 1991. 14 pp.

Fouts:1991:ADG

[FB91b] D. J. Fouts and S. E. Butner. Architecture and design of a 500-MHz galliumarsenide processing element for a parallel supercomputer. IEEE Journal of Solid-State Circuits, 26(9):1199– 1211, September 1, 1991. CODEN IJSCBC. ISSN 0018-9200 (print), 1558-173X (electronic).

Filippone:1993:MCS

[FBA93] W. L. Filippone, R. S. Baker, and R. E. Alcouffe. The Monte Carlo/Sn hybrid method with orthogonal interface basis vectors. In Kusters et al. [KSW93], pages 714–724. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Frikel:1997:DNC

[FBB97] M. Frikel, S. Bourennane, and A. Bendjamaa. Detection of the noise correlation length. In Roller [Rol97], pages 491–494. ISBN 0-947719-88-1 (paperback). LCCN ????

Fraternali:2018:QIV

[FBCB18] Francesco Fraternali, Andrea Bartolini, Carlo Cavazzoni, and Luca Benini. Quantifying the impact of variability and heterogeneity on the energy efficiency for a nextgeneration ultra-green supercomputer. IEEE Transactions on Parallel and Distributed Systems, 29(7):1575– 1588, July 2018. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.
computer.org/csdl/trans/
td/2018/07/08081827-abs.
html.

Finnemann:1993:RLC

[FBGM93] H. Finnemann, H. Bauer, A. Galati, and R. Martinelli. Results of LWR core transient benchmarks [invited]. In Kusters et al. [KSW93], pages 243–258. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Finnemann:1993:FDS

[FBH93] H. Finnemann, R. Boer, and J. Huesken. Finite difference solution of the flux reconstruction problem in nodal reactor analysis. In Kusters et al. [KSW93], pages 533– 545. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Fink:1994:CID

[FBJ94a] S. J. Fink, S. B. Baden, and K. Jansen. Cluster identification on a distributed memory multiprocessor. In IEEE [IEE94c], pages 239– 246. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Fox:1994:ISO

[FBJ^{+}94b] G. C. Fox, E. Bogucz, D. A. Jones, K. Mills, M. Podgorny, and K. A. Hawick. InfoMall: a scalable organisation for the development of high-performance computing and communications — software and systems. Lecture Notes in Computer Science, 797:137–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fahringer:1992:APP

[FBZ92] T. Fahringer, R. Blasko, and H. P. Zima. Automatic performance prediction to support parallelization of Fortran programs for massively parallel systems. In ACM [ACM92b], pages 347–356. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Fickett:1993:GAA

[FC93] J. W. Fickett and M. J. Cinkosky. A genetic algorithm for assembling chromosome physical maps. In Lim et al. [L^{+}93], pages 273–286. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Feitelson:1995:PSM

[FCBH95a] Dror G. Feitelson, Peter F. Corbett, Sandra Johnson Baylor, and Yarsun Hsu. Parallel I/O subsystems in massively parallel supercomputers. IEEE parallel and distributed technology: systems and applications, 3(3):33–47, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1995/pdf/h30033.
pdf;   http://www.computer.
org/concurrency/pd1995/
p3033abs.htm.

Feitelson:1995:PIS

[FCBH95b] Dror G. Feitelson, Peter F.

Corbett, Sandra Johnson Baylor, and Yarsun Hsu. Parallel I/O subsystems in massively parallel supercomputers. IEEE parallel and distributed technology: systems and applications, 3 (3):33, ???? 1995. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Fernandes:1997:PDS

[FCD97] D. L. G. Fernandes, A. C.

Canale, and B. J. Da Fonseca. Performance and directional stability of combination vehicles during the braking process under any operation condition. In Roller [Rol97], pages 153– 160. ISBN 0-947719-88-1 (paperback). LCCN ????

Fabiani:1990:ICA

[FCGG90] G. Fabiani, L. Calori, A. Guidazzoli, and C. Giorgi. Interactive computer animations of biological and biomechanical structures in a supercomputing environment. In Pitcher [Pit90], pages 309– 314. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Feng:1993:SCE

[FD93] Z. Q. Feng and M. Domaszewski. Some computational examples of nonlinear contact problems in transportation industries. In Anonymous [Ano93-31], pages 701–708. ISBN 0-947719-62-8. LCCN ????

Fiorini:1997:AVS

[FD97] R. A. Fiorini and G. F. Dacquino. Animation validation by spacetime constraints paradigm for realistic simulation representation. In Roller [Rol97], pages 95–104. ISBN 0-947719-88-1 (paperback). LCCN ????

Freitas:2002:SCF

[FDD02] Conceição Freitas, Luís Dias, and Miguel Dias. Simulating cloth free-form deformation with a Beowulf cluster. Lecture Notes in Computer Science, 2474:96–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/

2474/24740096.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2474/24740096.pdf.

Fang:2007:PFN

[FDM07] Bin Fang, Yuefan Deng, and Glenn Martyna. Performance of the 3D FFT on the 6D network torus QCDOC parallel supercomputer. Computer Physics Communications, 176(8): 531–538, April 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465507000276.

Feautrier:1994:TAD

[Fea94] Paul Feautrier. Toward automatic distribution. Parallel Processing Letters, 4 (3):233–244, September 1994. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

FCCSET-SCEC:1987:USI

[Fed87a] Federal Coordinating Council for Science and Engineering Technology. Subcommittee on Science and Engineering Computing. The U.S. supercomputer industry. Technical Report DOE/ER-0362, Office of Science and Technology Policy, Executive Office of the President, Washington, DC, USA, December 1987. various pp.

FCCSETSSEC:1987:USI

[Fed87b] Federal Coordinating Council for Science, Engineering, and Technology. Subcommittee on Science and Engineering Computing. The US supercomputer industry. Executive Office of the President, 1987. various pp. Distributed to depository libraries in microfiche. “December 1987.” “DOE/ER-0362.”.

Feder:1996:DUS

[Fed96]  Toni Feder.  DOD up-

grades its supercomputing facilities, and forges new partnerships with academia. Physics Today, 49(9):77– ??, September 1996. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

                                        //www.physicstoday.org/
                                        resource/1/phtoad/v49/i9/
                                        p77_s1.

Feitelson:1994:TIM

[Fei94] D. G. Feitelson. Terminal I/O for massively parallel systems. In IEEE [IEE94c], pages 263–270. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Feitelson:2005:SIL

[Fei05] Dror G. Feitelson. The supercomputer industry in light of the Top500 data. Computing in Science and Engineering, 7(1):42–47, January/ February 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/dl/mags/cs/
2005/01/c1042.htm;   http:
//csdl.computer.org/dl/
mags/cs/2005/01/c1042.pdf.

Fakih:2020:GDE

[FEK20] Bilal Fakih, Didier El Baz, and Igor Kotenko. GRIDHPC: a decentralized environment for high performance computing. Concurrency and Computation: Practice and Experience, 32(10):e5320:1– e5320:??, May 25, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Feo:1992:CSP

[Feo92] John T. Feo. A comparative study of parallel programming languages: the Salishan problems, volume 6 of Special topics in supercomputing. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-88135-2. ix + 386 pp. LCCN QA76.642.C66 1992.

FIA:1983:FIA

[Fer83] Fermilab Industrial Affiliates. Fermilab Industrial Affiliates Roundtable on Supercomputer Developments in the Universities: [proceedings], Fermilab, May 1983. Technical report, Fermi National Accelerator Laboratory, Batavia, IL, USA, May 19, 1983. vii + 122 pp.

Fernbach:1984:SPP

[Fer84] Sidney Fernbach. Supercomputers — past, present, prospects. Future Generation Computer Systems, 1(1): 23–30, July 1984. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fernbach:1986:SCV

[Fer86] Sidney Fernbach. Supercomputers—Class VI Systems, Hardware and Software. North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-444-87981-1. vii + 251 pp. LCCN QA76.5.S896 1986.

Fernbach:1989:SS

[Fer89] Sidney Fernbach. Software for supercomputers. Technical Report CSRD 869, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1989. 70–74 pp.

Fet:1995:VPS

[Fet95] Yakov I. Fet. Vertical processing systems: a survey. IEEE Micro, 15(1):65–75, January/February 1, 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Foerster:1995:INM

[FFM95] H. Foerster, A. Frommer, and G. Mayer. Inexact Newton methods on a vector supercomputer. Journal of Computational and Applied Mathematics, 58(3): 237–??, April 20, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Fatoohi:1987:IFC

[FG87] Raad A. Fatoohi and Chester E.

Grosch. Implementation of a four color cell relaxation scheme on the MPP, FLEX/ 32, and Cray/2. Proceedings of the International Conference on Parallel Processing, pages 424–426, 1987. CO-DEN PCPADL. ISBN 0-271-00608-0. ISSN 0190-3918. IEEE Service Cent. Piscataway, NJ, USA.

Finnie:1992:BCU

[FG92] Finnie and Geidl. The benefits and costs of using the NSF supercomputer centers. CoED, II(4):69–??, October 1992. CODEN CWLJDP. ISSN 0736-8607.

Fitch:1993:VOE

[FG93] B. G. Fitch and M. E.

Giampapa. The Vulcan operating environment: a brief overview and status report. In Hoffmann and Kauranne [HK93b], pages 130– 143. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Feng:2006:APA

[FGC06] Xizhou Feng, Rong Ge, and Kirk W. Cameron. The Argus prototype: aggregate use of load modules as a high-density supercomputer. Concurrency and Computation: Practice and Experience, 18(15):1975–1987, December 25, 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fraysse:2009:ASF

[FGG09] Valérie Frayssé, Luc Giraud, and Serge Gratton. Algorithm 881: a set of flexible GMRES routines for real and complex arithmetics on high-performance computers. ACM Transactions on Mathematical Software, 35(2):13:1–13:12, July 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Foster:1997:MMC

[FGKT97] Ian Foster, Jonathan Geisler, Carl Kesselman, and Steven Tuecke. Managing multiple communication methods in high-performance networked computing systems. Journal of Parallel and Distributed Computing, 40(1): 35–48, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.

1996.1266/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1266/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1266/production/ref.

Frank:1990:EEP

[FGM90] G. Frank, E. J. (Efstratios J.) Gallopoulos, and Ulrike Meier. Experiments with elliptic problem solvers on the Cedar multicluster. Technical Report CSRD 902, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1990. 6 pp.

Fitch:2003:BMA

[FGM^{+}03] B. G. Fitch, R. S. Germain, M. Mendell, J. Pitera, M. Pitman, A. Rayshubskiy, Y. Sham, F. Suits, W. Swope, T. J. C. Ward, et al. Blue Matter, an application framework for molecular simulation on Blue Gene. Journal of Parallel and Distributed Computing, 63(7–8): 759–773, July/August 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Fairman:1995:CDP

[FH95] William Fairman and Randal Hoff. Cross-platform database programming. Dr. Dobb’s Journal of Software Tools, 20(3):36–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Fujino:1997:IMS

[FHKT97] S. Fujino, R. Himeno, A. Kojima, and K. Terada. Implementation of the multicolored SOR method on a vector supercomputer. IEICE transactions on information and systems, 80(4):518–??, ???? 1997. ISSN 0916-8532.

Fahringer:1995:UTD

[FHM95] T. Fahringer, M. Haines, and P. Mehrotra. On the utility of threads for data parallel programming. In ACM [ACM95a], pages 51– 59. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Fukushige:1999:HPS

[FHM99] Toshiyuki Fukushige, Piet Hut, and Junichiro Makino. High-performance specialpurpose computers in science. Computing in Science and Engineering, 1(2): 12–13, 16, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c2012.
                                        pdf.

Furukawa:1991:SNC

[FI91] M. Furukawa and M. Inoue. Supercomputing for numerical cascade tunnel. In Anonymous [Ano91q], pages 155– 164. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Fujii:1993:TNM

[FI93] M. Fujii and N. Ito. Two new methods for stereo-view problems: Two-stage dynamic programming model (TDM) and physical-space method (PSM). In Kusters et al. [KSW93], pages 677– 690. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Fiduccia:1990:BHO

[Fid90] Charles M. (Charles Michael)

Fiduccia. Bussed hypercubes and other pin-optimal networks. Technical report SRC-TR-90-009, Supercomputing Research Center: IDA, Lanham, MD, USA, March 21, 1990. 32 pp.

Fiduccia:1991:BIN

[Fid91] Charles M. (Charles Michael)

Fiduccia. Bussed interconnection networks from trees. Technical report SRC-TR-91-041, Supercomputing Research Center: IDA, Lanham, MD, USA, August 30, 1991. 34 pp.

Fiebrich:1986:SWV

[Fie86] Rolf-Dieter Fiebrich. A supercomputer workstation for VLSI CAD. Technical report TMC-27, Thinking Machines Corp, Cambridge, MA, USA, June 1986. 31–37 pp.

Fiedler:1993:CMA

[Fie93] F. Fiedler. Computer models applied to transport of air pollutants over complex terrain [invited]. In Kusters et al. [KSW93], pages 341–

352. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Fincham:1982:PDS

[Fin82] D. Fincham. Programs for the dynamic simulation of liquids and solids II. MDIONS: Rigid ions using the Ewald sum (vectorised version on the Cray-1). Computer Physics Communications, 25 (2):177–179, February 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465582900339.

AFD-OLA-SM:1994:MSC

[Fin94] Financial Audit Division, Office of the Legislative Auditor, State of Minnesota, Saint Paul, MN, USA. Minnesota Supercomputer Center, Inc. financial audit for the two years ended June 30, 1993, June 1994. 3 + 18 pp.

Fischetti:1983:US

[Fis83] M. A. Fischetti. The United States. IEEE Spectrum, 20 (11):51–69, November 1983. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fiduccia:1991:UMN

[FJ91] Charles M. (Charles Michael)

Fiduccia and Elaine M. Jacobson. Universal multistage networks via linear permutations. Technical report SRC-TR-91-050, Supercomputing Research Center: IDA, Lanham, MD, USA, December 23, 1991. 21 pp.

Freitag:1994:NTP

[FJP94] L. Freitag, M. Jones, and P. Plassmann. New techniques for parallel simulation of high-temperature superconductors. In IEEE [IEE94c], pages 726–733. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Fosdick:1996:IHP

[FJSD96] Lloyd D. Fosdick, Elizabeth R. Jessup, Carolyn J. C. Schauble, and Gitta Domik. An Introduction to High-Performance Scientific Computing. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-06181-3. xxiii + 760 pp. LCCN QA76.A594 1996. US$65.00. URL http://www.

mitpress.com/book-home.
tcl?isbn=0262061813.

Ferris:1995:CWF

[FJSP95] Michael Ferris and Jong-Shi-Pang. Conferences & workshops: First International Conference on Complementarity Problems; 11th ACM Symposium on Computational Geometry; High-Performance Computational Chemistry Tutorial and Workshop; 1995 IEEE Antennas and Propagation Society International Symposium; 1995 SIAM Annual Meeting. IEEE Computational Science & Engineering, 2(4):90–92, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Fruehauf:1993:IVC

[FK93] T. Fruehauf and K. Karlsson. Interactive visualization of CFD data with ISVAS. In Kusters et al. [KSW93], pages 779–780. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Foster:1998:CGD

[FK98] Ian Foster and Carl Kesselman. Computational grids: On-demand computing in science and engineering. Computers in Physics, 12(2): 109–??, March 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.168654.

Foster:1999:GBN

[FK99] Ian Foster and Carl Kesselman, editors. The Grid: blueprint for a new computing infrastructure. Morgan Kaufmann Publishers, San Francisco, CA, USA, 1999. ISBN 1-55860-475-8. xxiv + 677 pp. LCCN QA76.9.C58 G75 1999. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els033/98030570.
html;   http://www.loc.
gov/catdir/toc/els032/98030570.
html.

Foster:2004:GBN

[FK04] Ian Foster and Carl Kesselman. The Grid 2: Blueprint for a new computing infrastructure. Morgan Kaufmann Publishers, San Francisco, CA, USA, second edition, 2004. ISBN 1-55860-933-4. xxvii + 748 pp. LCCN QA76.9.C58 G75 2004. US$59.95.

Fisher:2008:TTC

[FKL^{+}08] R. T. Fisher, L. P. Kadanoff, D. Q. Lamb, A. Dubey, T. Plewa, A. Calder, F. Cattaneo, P. Constantin, I. Foster, M. E. Papka, S. I. Abarzhi, S. M. Asida, P. M. Rich, C. C. Glendenin, K. Antypas, D. J. Sheeler, L. B. Reid, B. Gallagher, and S. G. Needham. Terascale turbulence computation using the FLASH3 application framework on the IBM Blue Gene/L system. IBM Journal of Research and Development, 52(1/2):127– ??, January/March 2008. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

                                        //www.research.ibm.com/
                                        journal/rd/521/fisher.html.

Fritz:1992:CVP

[FL92] L. W. Fritz and J. R. Lucas, editors. 17th Congress Vol 29 pt B2, Systems for data processing and analysis: August 1992, Washington; DC, volume 29(2) of International Archives of Photogrammetry and Remote Sensing. The Committee of the Congress, Washington, DC, USA, 1992. ISBN ???? ISSN 0256-1840. LCCN ????

Freeh:2007:AET

[FLP^{+}07] Vincent W. Freeh, David K. Lowenthal, Feng Pan, Nandini Kappiah, Rob Springer, Barry L. Rountree, and Mark E. Femal. Analyzing the energy-time tradeoff in high-performance computing applications. IEEE Transactions on Parallel and Distributed Systems, 18(6): 835–848, June 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Flynn:1966:VHS

[Fly66] M. J. Flynn. Very high-speed computing systems. Proceedings of the IEEE, 54(12): 1901–1909, December 1966. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Foster:1993:MMP

[FM93] I. Foster and J. Michalakes.

MPMM: a massively parallel mesoscale model. In Hoffmann and Kauranne [HK93b], pages 354–363. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Fang:2007:FGP

[FMD07] Bin Fang, Glenn Martyna, and Yuefan Deng. A fine grained parallel smooth particle mesh Ewald algorithm for biophysical simulation studies: Application to the 6-D torus QCDOC supercomputer. Computer Physics Communications, 177(4): 362–377, August 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465507002445.

Fukuda:1991:TAP

[FMT91] Akira Fukuda, Kazuaki Murakami, and Shinji Tomita. Toward advanced parallel processing: exploiting parallelism at task and instruction levels. IEEE Micro, 11(4):16– 19, 50–61, July/August 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Furuta:1993:QIA

[FNK93] K. Furuta, K. Nakata, and S. Kondo. QCE: An intelligent aid towards selfexplanatory computer simulation [invited]. In Kusters et al. [KSW93], pages 367– 378. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Fichtner:1984:ISI

[FNP^{+}84] W. Fichtner, L. Nagel, R. Penumalli, W. Peterson, and J. D’Arcy. The impact of supercomputers on IC technology development and design. Proceedings of the IEEE, 72:76–112, 1984. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Fukuzaki:1993:DSS

[FNT93] T. Fukuzaki, N. Naito, and M. Tani. Development of safety support system for nuclear power plants [invited]. In Kusters et al. [KSW93], pages 383–387. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Fong:1985:NCT

[Fon85]  Kirby W. Fong.  NM-

FECC Cray time-sharing system. Software—Practice and Experience, 15(1):87– 103, January 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ford:1993:LED

[For93] W. K. Ford. Low-energy electron diffraction calculations using a parallel supercomputer. Surface science, 292(3):342–??, August 1993. CODEN SUSCAS. ISSN 0039-6028.

Forsell:2002:SHP

[For02] Martti Forsell. A scalable high-performance computing solution for networks on chips. IEEE Micro, 22(5):46– 55, September/October 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2002/pdf/m5046.
pdf;   http://www.computer.
org/micro/mi2002/m5046abs.
htm.

Foster:1993:FML

[Fos93] I. Foster. Fortran M as a language for building earth system models. In Hoffmann and Kauranne [HK93b], pages 144–151. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Foster:1996:HPD

[Fos96] I. Foster. High-performance distributed computing: The I-WAY experiment and beyond. Lecture Notes in Computer Science, 1123:3– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foster:2003:ITG

[Fos03] Ian Foster. Information technology: The grid: Computing without bounds. Scientific American, 288(4):

78–85, April 2003. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

                                        tronic). URL http://www.
                                        nature.com/scientificamerican/
                                        journal/v288/n4/pdf/scientificamerican0403-
                                        78.pdf.

Fox:1989:PCC

[Fox89] G. C. Fox. Parallel computing comes of age: Supercomputer level parallel computations at Caltech. Concurrency: practice and experience, 1(1):63–103, September 1989. CODEN CPEXEI. ISSN 1040-3108.

Fox:1990:GAQ

[Fox90a] D. J. Fox. Gaussian 90: Applying quantum chemistry to biology. In Pitcher [Pit90], pages 315–322. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Fox:1990:TS

[Fox90b]  Geoffrey Fox.  The ter-

aflop supercomputer. Computers in Physics, 4(1):112– ??, ???? 1990. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Fox:1997:JCS

[Fox97] G. C. Fox, editor. Java for computational science and engineering — simulation and modeling: Workshop — December 1996, Syracuse, KS, volume 9(6) of CON-CURRENCY 1997. John Wiley and Sons, Inc., New York, NY, USA, 1997. ISSN 1040-3108.

Fox:1998:EJH

[Fox98]  Geoffrey Fox.  Editorial:

Java for high-performance network computing. Concurrency: practice and experience, 10(11–13):821–824, September 1998. CO-DEN CPEXEI. ISSN 1040-3108. URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?ID=
10050427;   http://www3.
interscience.wiley.com/
cgi-bin/fulltext?ID=10050427&
PLACEBO=IE.pdf.  Spe-

cial Issue: Java for Highperformance Network Computing.

Furnari:1991:RTM

[FP91] Mario Furnari and C. D.

(Constantine D.) Polychronopoulos. Run time management of LISP parallelism and the hierarchical task graph program representation. Technical Report CSRD 1133, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 16 pp.

Francioni:2000:DSH

[FP00] Joan M. Francioni and Cherri M. Pancake. A debugging standard for highperformance computing. Scientific Programming, 8(2): 95–108, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.asp%
                                        3Fwasp=253x52trrm4r87tkuw1h%
                                        26referrer=parent%26backto=
                                        issue%2C3%2C3%3Bjournal%
                                        2C4%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Fincham:1981:MDS

[FR81] David Fincham and B. J. Ralston. Molecular dynamics simulation using the Cray-1 vector processing computer. Computer Physics Communications, 23(2):127– 134, July 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Fiduccia:1991:PS

[FR91] Charles M. (Charles Michael) Fiduccia and Kevin J. Rappoport. Perfect shifters. Technical report SRC-TR-91-031, Supercomputing Research Center: IDA, Lanham, MD, USA, February 6, 1991. 27 pp.

Floros:1995:ESE

[FR95] N. Floros and J. S. Reeve. Evaluation of a spectral element CFD code on parallel architectures. Parallel Computing, 21(7):1137– 1150, July 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Feitelson:1996:TCJ

[FR96a] D. G. Feitelson and L. Rudolph.

Towards convergence in job schedulers for parallel supercomputers. Lecture Notes in Computer Science, 1162:1– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fornasari:1996:CAC

[FR96b] N. Fornasari and S. Rovida.

Conjugate-gradients algorithms on a CRAY-T3D. Lecture Notes in Computer Science, 1067:668–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fornasari:1996:CGA

[FR96c] N. Fornasari and S. Rovida.

Conjugate-gradients algorithms on a CRAY-T3D. Lecture Notes in Computer Science, 1067:668–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1998:SII

[FR98] Afonso Ferreira and José Rolim. Special issue on irregular problems in supercomputing applications. Journal of Parallel and Distributed Computing, 50(1–2):

1–2, April/May 1998. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Frank:1990:ECM

[Fra90] George N. Frank. Experiments on the Cedar multicluster with parallel block cyclic reduction and an application to domain decomposition methods. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1990. vii + 69 pp.

Franco:1994:NSA

[Fra94] Jose Franco, editor. Numerical simulations in astrophysics: proceedings of the first UNAM-CRAY Supercomputing Workshop, “Numerical Simulations in Astrophysics: Modelling the Dynamics of the Universe” held in Mexico City, Mexico, July 26–30, 1993. Cambridge University Press, New York, NY, USA, 1994. ISBN 0-521-46238-X. LCCN QB460 .U53 1993.

Fried:1991:PSI

[Fri91] S. S. Fried. Personal supercomputing with the Intel i860: Crunching numbers with the i860. BYTE Magazine, 16(1):347–348, 350–352, 356, 356, 358, January 1991. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Fried:1994:SMP

[Fri94] Stephen S. Fried. Shared memory and PC supercomputing. Dr. Dobb’s Journal of Software Tools, 19(1):18– 20, 22, 24, 26, 28, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Friston:1995:FIC

[Fri95] K. J. Friston. Functional integration and connectivity in neuroimaging. In Herrmann et al. [HWP95], pages 119– 136. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Forrest:1988:NNM

[FRS^{+}88] B. M. Forrest, D. Roweth, N. Stroud, D. J. Wallace, and G. V. Wilson. Neural network models. Parallel Computing, 8(1-3):71–83, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fruehauf:1993:TTT

[Fru93] T. Fruehauf. Today’s and tomorrow’s techniques for the visualization of numerically generated flow data. In Kusters et al. [KSW93], pages 667–676. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Filaseta:1992:MEN

[FRW92] Michael Filaseta, M. L.

Robinson, and Ferrell S.

Wheeler. The minimal Euclidian norm of an algebraic number is effectively computable. Technical report SRC-TR-92-060, Supercomputing Research Center: IDA, Lanham, MD, USA, January 1992. 19 pp.

Frye:1997:MNC

[Fry97] A. Frye. Meeting the needs of a changing population — the challenges for the automotive industry. In Roller [Rol97], pages 45–52. ISBN 0-947719-88-1 (paperback). LCCN ????

Filbert:1993:NMT

[FS93a] D. Filbert and C. Schneider. New methods for the testing of electric low power motors. In Anonymous [Ano93-31], pages 99–106. ISBN 0-947719-62-8. LCCN ????

Fritscher:1993:PDC

[FS93b] J. F. Fritscher and F. Sukup. 93SC038 parallel distributed computing using PVM. In Anonymous [Ano93-31], pages 221–228. ISBN 0-947719-62-8. LCCN ????

Field:2018:CIV

[FSC18] L. Field, D. Spiga, and L. Cristella. CMS@home: Integrating the volunteer cloud and high-throughput computing. Computing and Software for Big Science, 2(1): ??, November 2018. CO-DEN ???? ISSN 2510-2036 (print), 2510-2044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s41781-018-0006-z.

Fen:1993:EMC

[FSGS93] V. G. Fen, I. R. Suslov, H. Gerwin, and W. Scherer. Effective method for calculations of the space-energy distribution of neutrons in complex geometry regions containing absorbers and holes. In Kusters et al. [KSW93], pages 590–596. ISBN 3-923704-11-9. LCCN ???? Two volumes.

FroeseFischer:1988:MCA

[FSY88] C. Froese Fischer, N. S. Scott, and J. Yoo. Multitasking the calculation of angular integrals on the Cray-2 and Cray X-MP. Parallel Computing, 8 (1-3):385–390, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fijalkowski:1993:CMC

[FT93a] B. T. Fijalkowski and K. I.

Trovato. A concept for a mechatronically controlled full-time 4WD × 4WB × 4WA × 4WS intelligent vehicle for drivers with special needs. In Anonymous [Ano93-31], pages 161–172. ISBN 0-947719-62-8. LCCN ????

Furtney:1993:WS

[FT93b] Mark Furtney and George Taylor. Of workstations and supercomputers. IEEE Spectrum, 30(5):64–68, May 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foster:1994:LAC

[FT94] I. T. Foster and B. R. Toonen. Load-balancing algorithms for climate models. In IEEE [IEE94c], pages 674– 681. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Farcy:1996:ISP

[FT96a] A. Farcy and O. Temam. Improving single-process performance with multithreaded processors. In ACM [ACM96], pages 350–357. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Foster:1996:ETW

[FT96b] I. Foster and S. Tuecke. Enabling technologies for Webbased ubiquitous supercomputing. In IEEE [IEE96b], pages 112–120. ISBN 0-8186-7582-9. ISSN 1082-8907. LCCN ????

Foster:1997:TUS

[FTT97] Ian Foster, George K. Thiruvathukal, and Steven Tuecke. Technologies for ubiquitous supercomputing: a Java interface to the Nexus communication system. Concurrency: practice and experience, 9(6):

465–475, June 1997. CO-DEN CPEXEI. ISSN 1040-3108. URL http://

www3.interscience.wiley.
com/cgi-bin/abstract?ID=
13871;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13871&PLACEBO=IE.pdf.

Fujino:1999:ECF

[Fuj99] Seiji Fujino. Estimation of conflict-free matrix product preconditioner on vector supercomputers. Applied Numerical Mathematics: Transactions of IMACS, 30(2– 3):257–266, June 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
apnum/cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=979.

Fujimoto:2011:NEP

[Fuj11] Keizo Fujimoto. A new electromagnetic particle-in-cell model with adaptive mesh refinement for highperformance parallel computation. Journal of Computational Physics, 230(23):8508– 8526, September 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0021999111004608.

Furber:2012:BB

[Fur12]  S. Furber.  To build a

brain. IEEE Spectrum, 49 (8):44–49, August 2012. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Foley:1994:NAS

[FV94] C. M. Foley and S. Vinnakota. Nonlinear analysis of structural frameworks using supercomputing techniques. In Khozeimeh [Kho94], pages 2058–?? ISBN 0-7844-0026-1. LCCN TA345.C657 1994. Two volumes.

Frye:1990:VCS

[FW90] Patrick J. Frye and Warren M. Washington. Visualization in computational science, 1990. 1 videocassette (ca. 75 min.).

Friebel:1996:QSA

[FWS96] W. Friebel, P. Wegner, and H. Simma. Quadrics supercomputing at Zeuthen. In Shellard and Nguyen [SN96], pages 902–906. ISBN 981-02-2783-3. LCCN ????

Fuchs:1995:ABB

[FWWD95] M. Fuchs, H.-A. Wischmann, M. Wagner, and R. Drenckhahn. Advanced biomagnetic and bioelectric reconstruction algorithms. In Herrmann et al. [HWP95], pages 161–174. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Foster:1994:CSI

[FXAC94] I. Foster, M. Xu, B. Avalani, and A. Choudhary. A compilation system that integrates High Performance Fortran and Fortran M. In IEEE [IEE94c], pages 293– 300. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Fujii:1992:NSB

[FY92] Kozo Fujii and Hideo Yoshihara. Navier/Stokes supercomputer benchmark — an update. Scientific information bulletin, 17(4):99–??, October 1, 1992. CODEN SINBEM. ISSN 1048-5678.

Fu:1996:RCP

[FY96] C. Fu and T. Yang. Runtime compilation for parallel sparse matrix computations. In ACM [ACM96], pages 237–244. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Fujino:1991:VCB

[FZM91] S. Fujino, S. Zhang, and M. Mori. Visualization of convergence behavior of Bi-CG STAB method. In Anonymous [Ano91q], pages 213–218. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Goyal:1984:PAF

[GA84] A. Goyal and T. Agerwala.

Performance analysis of future shared storage systems. IBM Journal of Research and Development, 28(1):95–108 (or 95–107??), January 1984.

CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Giladi:1995:SPE

[GA95] Ran Giladi and Niv Ahituv. SPEC as a performance evaluation measure. Computer, 28(8):33–42, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grunert:1997:WMV

[GA97] T. Grunert and R. Ali. The whirl modes of vibration of crankshafts in powertrain assemblies. In Roller [Rol97], pages 253–264. ISBN 0-947719-88-1 (paperback). LCCN ????

Guizzo:2012:RRW

[GA12] E. Guizzo and Evan Ackerman. The rise of the robot worker. IEEE Spectrum, 49(10):34–41, October 2012. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guest:1996:HPC

[GAB^{+}96] M. F. Guest, E. Apra, D. E. Bernholdt, H. A. Früchtl, R. J. Harrison, R. A. Kendall, R. A. Kutteh, X. Long, J. B. Nicholas, J. A. Nichols, H. L. Taylor, A. T. Wong, G. I. Fann, R. J. Littlefield, and J. Nieplocha. Highperformance computing in chemistry: NW Chem. Future Generation Computer Systems, 12(4):273–289, December 1996. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gallivan:1987:IHM

[Gal87] Kyle A. Gallivan. The impact of hierarchical memory systems on linear algebra algorithm design. Technical Report CSRD 625, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 63 pp.

Gallivan:1988:BCM

[Gal88a] Kyle A. Gallivan. Behavioral characterization of multiprocessor memory systems: a case study. Technical Report CSRD 808, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1988. 60 pp.

Gallivan:1988:PAC

[Gal88b]  Kyle A. Gallivan.  Per-

formance analysis on the CEDAR system. Technical Report CSRD 680, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 48 pp.

Gall:1989:SAM

[Gal89a] Vojtech Gall. Supercomputer assisted mine modeling. Thesis (M.S.E.M.), University of Alabama, Tuscaloosa, AL, USA, 1989. xiv + 119 pp.

Gallivan:1989:PPL

[Gal89b]  Kyle A. Gallivan.  Per-

formance prediction of loop constructs on multiprocessor hierarchical-memory systems. Technical Report CSRD 853, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1989. 19 pp.

Gallivan:1991:PBP

[Gal91] Kyle A. Gallivan. Preliminary basic performance analysis of the Cedar multiprocessor memory system. Technical Report CSRD 1116, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 26 pp.

Galli:1993:CPM

[Gal93]  P. Galli.  The contribu-

tion of polymeric materials to the future of the automotive industry. In Anonymous [Ano93-31], pages 5–12. ISBN 0-947719-62-8. LCCN ????

Galtier:1996:APT

[Gal96] J. Galtier. Automated partitioning techniques for solving partial differential equations on irregular adaptive meshes. In ACM [ACM96], pages 157–164. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Gannon:1986:RNL

[Gan86] Dennis B. Gannon. Restructuring nested loops on the Alliant Cedar cluster: a case study of Gaussian elimination of banded matrices. Technical Report CSRD-543, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 14 pp.

Gannon:1988:STB

[Gan88] Dennis B. Gannon. A software tool for building supercomputer applications. Technical Report CSRD 777, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1988. 12 pp.

Ganapathy:1994:VR

[Gan94a] S. K. Ganapathy. Virtual reality. In Balakrishnan [Bal94], pages 407–?? ISBN 0-07-462044-4. LCCN ????

Ganesan:1994:IPA

[Gan94b] S. Ganesan. Image processing architectures. In Balakrishnan [Bal94], pages 410–?? ISBN 0-07-462044-4. LCCN ????

Gao:1986:MPT

[Gao86] Guang R. Gao. A maximally pipelined tridiagonal linear equation solver. Journal of Parallel and Distributed Computing, 3(2): 215–235, June 1986. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Garrett:1992:VTS

[Gar92] P. Garrett. A vision of a teraflop supercomputing system. In Meuer [Meu92c], pages 193–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Garber:1999:NBA

[Gar99] Lee Garber. News briefs: Apple engineers a major turnaround; Java battles brew on several fronts; IBM and SGI unveil ultrafast supercomputers; Y2K briefs. Computer, 32(1): 23–25, January 1999. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co1999/pdf/r1023.
pdf.

Garber:2001:NBT

[Gar01] Lee Garber. News briefs:

Is tech downturn changing education and employment trends; HTMT promises high-performance computing; controversial software law [UCITA] hist resistance. Computer, 34(10): 19–21, October 2001. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/rx019.
pdf;   http://www.computer.
org/computer/co2001/rx019abs.
htm.

Gonzalez:1995:DCM

[GAV95] A. Gonzalez, C. Aliagas, and M. Valero. A data cache with multiple caching strategies tuned to different types of locality. In ACM [ACM95a], pages 338–347. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Gutbrod:1996:SGT

[GAW96a] F. Gutbrod, N. Attig, and M. Weber. The SU(2)- Lattice gauge theory simulation code on the Intel Paragon supercomputer. Parallel Computing, 22(3): 443–463, March 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gutbrod:1996:SLG

[GAW96b] F. Gutbrod, N. Attig, and M. Weber. The SU(2)- Lattice Gauge Theory simulation code on the Intel Paragon supercomputer. Parallel Computing, 22(3): 443–463, March 29, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        parco/cas_sub/browse/browse.
                                        cgi?year=1996&volume=22&
                                        issue=3&aid=1057.

Gao:1990:PSL

[GB90] Hui Gao and Michael Berry. Performance studies of LA-PACK on Alliant FX/80 and 1 CEDAR cluster. Technical Report CSRD 1001, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1990. 15 pp.

Gerla:1991:GDC

[GB91] M. Gerla and J. A. Bannister. A guide to data communications: high-speed local-area networks. IEEE Spectrum, 28 (8):26–31, August 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gupta:1992:MHS

[GB92] M. Gupta and P. Banerjee. A methodology for high-level synthesis of communication on multicomputers. In ACM [ACM92b], pages 357–367. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Ghosh:1996:ELM

[GB96] K. Ghosh and S. Breit. Evaluating the limits of message passing via the shared attraction memory on CC-COMA machines: Experiences with TCGMSG and PVM. In ACM [ACM96], pages 173–

180. ISBN 0-89791-803-

7. LCCN QA76.5 I61 1996. ACM order number 415961.

Gorobets:2022:HCP

[GB22] Andrey Gorobets and Pavel Bakhvalov. Heterogeneous CPU+GPU parallelization for high-accuracy scaleresolving simulations of compressible turbulent flows on hybrid supercomputers. Computer Physics Communications, 271(??):Article 108231, February 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S001046552100343X.

Giampapa:2005:BGA

[GBB^{+}05] M. E. Giampapa, R. Bellofatto, M. A. Blumrich, D. Chen, M. B. Dombrowa, A. Gara, R. A. Haring, P. Heidelberger, D. Hoenicke, G. V.

Kopcsay, B. J. Nathanson, B. D. Steinmacher-Burow, M. Ohmacht, V. Salapura, and P. Vranas. Blue Gene/L advanced diagnostics environment. IBM Journal of Research and Development, 49(2/):319–331, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/giampapa.
pdf.

Gara:2005:OBG

[GBC^{+}05] A. Gara, M. A. Blumrich, D. Chen, G. L.-T. Chiu, P. Coteus, M. E. Giampapa, R. A. Haring, P. Heidelberger, D. Hoenicke, G. V. Kopcsay, T. A. Liebsch, M. Ohmacht, B. D. Steinmacher-Burow, T. Takken, and P. Vranas. Overview of the Blue Gene/L system architecture. IBM Journal of Research and Development, 49(2/3):195–212, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/gara.pdf.

Gentzsch:1993:WCB

[GBF93] W. Gentzsch, U. Block, and F. Fersti. 93SC012 workstation clusters for batch and parallel computing. In Anonymous [Ano93-31], pages 265–274. ISBN 0-947719-62-8. LCCN ????

Ghafoor:1989:BFT

[GBG89] A. Ghafoor, T. R. Bashkow, and I. Ghafoor. Bisectional fault-tolerant communication architecture for supercomputer systems. IEEE Transactions on Computers, 38 (10):1425–1446, October 1, 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=35837.

Gunzinger:1996:PEH

[GBK^{+}96] A. Gunzinger, B. Baumle, M. Klebl, M. Kocheisen, P. Kohler, R. Morel, U. Muller, and M. Rosenthal. Programming environment for a High-Performance parallel supercomputer with intelligent communication. Scientific Programming, 5(1):25– 32, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gergel:2018:GNS

[GBS18] Victor Gergel, Konstantin Barkalov, and Alexander Sysoyev. Globalizer: A novel supercomputer software system for solving timeconsuming global optimization problems. Numerical Algebra, Control and Optimization, 8(1):47–62, ???? 2018. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http://

                                        aimsciences.org/article/
                                        doi/10.3934/naco.2018003.

Gokhale:1992:ICI

[GC92] Maya B. Gokhale and William W. Carlson. An introduction to compilation issues for parallel machines. Technical report SRC-TR-92-062, Supercomputing Research Center: IDA, Lanham, MD, USA, September 8, 1992. 38 pp.

GiaTong:1992:SPS

[GCB92] Vuong Gia Tong, Roger Chahine, and Stephen Behling. Supercomputing for power system analysis. IEEE computer applications in power: CAP, 5(3):45–49, July 1992. CODEN ICAPEH. ISSN 0895-0156.

Gao:1990:IBP

[GCP90] Hui Gao, Hsin-Chu Chen, and E. J. Plaskacz. Initial benchmarking and performance analysis for WHAMS3D. Technical Report CSRD 1054, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1990. 15 pp.

Gokulakrishnan:1994:GBA

[GCS94] S. Gokulakrishnan, C. Chellappon, and V. Sankaranarayanan. Genetic based adaptive machine learning classifier system for controlling robot performance. In Balakrishnan [Bal94], pages 154–162. ISBN 0-07-462044-4. LCCN ????

Gil-Costa:2020:HPC

[GCS20] Veronica Gil-Costa and Hermes Senger. High-performance computing for computational science. Concurrency and Computation: Practice and Experience, 32(20):e5904:1– e5904:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gokhale:2008:HTH

[GCY^{+}08] Maya Gokhale, Jonathan Cohen, Andy Yoo, W. Marcus Miller, Arpith Jacob, Craig Ulmer, and Roger Pearce. Hardware technologies for high-performance data-intensive computing. Computer, 41(4):60–68, April 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gowda:1994:ORU

[GD94a] K. C. Gowda and P. V. Desai. 2-D object recognition using self organising neural network. In Mahajan et al. [M^{+}94], pages 446–448. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Guillen:1994:CDM

[GD94b] P. Guillen and M. Dormieuz.

Chapter 2: Design of a 3D multidomain Euler code. In Murthy and Brebbia [MB94b], pages 21–40. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Green:1997:CMC

[GD97] M. L. Green and J. V. De-Pinto. Cross media coupling of mass balance models for toxic chemicals on massively parallel platforms. In Delic and Wheeler [DW97], pages 163–168. ISBN 0-89871-378-1. LCCN ????

Grinstein:1996:VDE

[GE96] Georges G. Grinstein and Robert F. Erbacher, editors. Visual data exploration and analysis III: 31 January–2 February, 1996, San Jose, California, number 2656 in Proceedings — SPIE The International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1996. ISBN 0-8194-2030-1. ISSN 0361-0748. LCCN TS510.S63 v.2656.

Gruner:2012:CBO

[GE12] Markus E. Gruner and Peter Entel. Competition between ordering, twinning, and segregation in binary magnetic 3d–5d nanoparticles: a supercomputing perspective. International Journal of Quantum Chemistry, 112(1):277– 288, January 2012. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Geist:2016:SMC

[Gei16]  A. Geist.  Supercomput-

ing’s monster in the closet. IEEE Spectrum, 53(3):30– 35, March 2016. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geller:2011:SET

[Gel11] Tom Geller. Supercomputing’s exaflop target. Communications of the ACM, 54 (8):16–18, August 2011. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gentzsch:1992:GCS

[Gen92] W. Gentzsch. Grand Challenges und Supercomputer. In Meuer [Meu92c], pages 1–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Gentzsch:1994:CRC

[Gen94] W. Gentzsch. Chapter 13:

Restructuring CFD algorithms for vector and parallel computers. In Murthy and Brebbia [MB94b], pages 309–330. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Gentzsch:1997:AHP

[Gen97] W. Gentzsch. Applications of high-performance computing. Lecture Notes in Computer Science, 1300:827–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

George:1994:VPR

[Geo94] S. M. George. VLSI placement and routing problems — the scope for its parallelisation. In Mahajan et al. [M^{+}94], pages 348–349. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Geppert:1999:CSB

[Gep99] L. Geppert. Champagne supercomputer on a beer budget. IEEE Spectrum, 36 (5):19–22, May 1999. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2000:MBG

[Gep00] L. Geppert. Microprocessors: the off-beat generation. IEEE Spectrum, 37 (7):44–49, July 2000. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geppert:2001:WMP

[Gep01] L. Geppert. World’s most powerful supercomputer hits full stride. IEEE Spectrum, 38(11):29–31, November 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gerz:1990:VVS

[Ger90] T. Gerz. Visualization of vortex structures in sheared and stratified flows by direct numerical simulation with very high resolution. In Pitcher [Pit90], pages 455– 466. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Getov:2015:SGC

[Get15] Vladimir Getov. Scientific grand challenges: Toward exascale supercomputing and beyond. Computer, 48(11): 12–14, November 2015. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
co/2015/11/mco2015110012.
html.

Geuss:1997:EMV

[Geu97] H. Geuss. Ergonoic modelling of vehicles with a computer based designing tool. In Roller [Rol97], pages 69–76. ISBN 0-947719-88-1 (paperback). LCCN ????

Gruber:1990:NSF

[GF90] K. Gruber and T. Frank.

Numerical simulation of the frontal impact with total and partial overlap of the front end. In Pitcher [Pit90], pages 219–231. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Ghuloum:1995:FPI

[GF95] Anwar M. Ghuloum and Allan L. Fisher. Flattening and parallelizing irregular, recurrent loop nests. ACM SIGPLAN Notices, 30(8):58– 67, August 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gabriel:2003:FTC

[GFB^{+}03] Edgar Gabriel, Graham E. Fagg, Antonin Bukovsky, Thara Angskun, and Jack J. Dongarra. A fault-tolerant communication library for Grid environments. In ????, editor, 17th Annual ACM International Conference on Supercomputing (ICS’03) International Workshop on Grid Computing and e-Science, June 21, 2003, San Francisco, page ?? ????, ????, 2003. ISBN ???? LCCN ???? URL http:

//www.netlib.org/netlib/
utk/people/JackDongarra/
PAPERS/FTMPI-SF-gabriel.
pdf.

Gabriel:2010:TPP

[GFBR10] Edgar Gabriel, Saber Feki, Katharina Benkert, and Michael M. Resch. Towards performance portability through runtime adaptation for high-performance computing applications. Concurrency and Computation: Practice and Experience, 22 (16):2230–2246, November 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Garg:1996:SST

[GFM96] R. Garg, J. Ferziger, and S. Monismith. Simulation of stratified turbulent channel flows on the Intel Paragon parallel supercomputer. In Ecer [Ece96], pages 481–488. ISBN 0-444-82322-0. LCCN ????

Guarna:1988:PUI

[GG88] Vincent A. Guarna and Yogesh Gaur. A portable user interface for a scientific programming environment. Technical Report CSRD 749, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1988. 10 pp.

Guccione:1995:SRA

[GG95] S. A. Guccione and M. J.

Gonzalez. Supercomputing with reconfigurable architectures. In Moore and Luk [ML95b], pages 389–398. CO-DEN LNCSD9. ISBN 3-540-60294-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.G36 I48 1995.

Garzon:1996:PIL

[GG96] E. M. Garzon and I. Garcia. Parallel implementation of the Lanczos method for sparse matrices: Analysis of data distributions. In ACM [ACM96], pages 294– 300. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Garcia-Garcia:1997:NSE

[GG^{+}97a] Fernando Garcia-Garcia et al., editors. Numerical simulations in the environmental and earth sciences: proceedings of the Second UNAM-CRAY Supercomputing Conference. Cambridge University Press, New York, NY, USA, 1997. ISBN 0-521-58047-1 (hardback). LCCN GE45.M37U53 1996.

Gruber:1997:SSP

[GG97b] R. Gruber and A. Gunzinger. The Swiss-Tx supercomputer project. In Anonymous [Ano97-30], pages 20– 22. ISSN 1421-6337.

Gordon:1995:LLC

[GGBR95] M. B. Gordon, D. R. Grempel, D. Berchier, and B. Raffin. Learning with a logistic cost function. In Herrmann et al. [HWP95], pages 349–354. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Gonzalez:2011:SWS

[GGC^{+}11] Juan Gonzalez, Judit Gimenez, Marc Casas, Miquel Moreto, Alex Ramirez, Jesus Labarta, and Mateo Valero. Simulating whole supercomputer applications. IEEE Micro, 31 (3):32–45, May/June 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Garg:1998:ALS

[GGG^{+}98] Sharad Garg, Robert Godley, Richard Griffiths, Andrew Pfiffer, Terry Prickett, David Robboy, Stan Smith, T. Mack Stallcup, and Stephan Zeisset. Achieving large scale parallelism through operating system resource management on the Intel TFLOPS supercomputer. Intel Technology Journal, (Q1):12, 1998. ISSN 1535-766X. URL

http://developer.intel.
com/technology/itj/q11998/
articles/art_3.htm;   http:
//developer.intel.com/technology/
itj/q11998/pdf/tos.pdf.

Gaur:1989:EPE

[GGJ89] Yogesh Gaur, Vincent A.

Guarna, and David Jablonowski. An environment for performance experimentation on multiprocessors. Technical Report CSRD 865, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1989. 8 pp.

Gitler:2020:LAR

[GGN20] Isidoro Gitler, Antônio Tadeu A. Gomes, and Sergio Nesmachnow. Latin America regional special section: Big trends: The Latin American supercomputing ecosystem for science. Communications of the ACM, 63 (11):66–71, October 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3419977.

Gornish:1990:CDP

[GGV90] Edward H. Gornish, Elana Denise Granston, and Alexander Veidenbaum. Compilerdirected data prefetching in multiprocessors with memory hierarchies. Technical Report CSRD 996, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1990. 15 pp.

Ganapol:1993:ANP

[GGW93a] B. D. Ganapol, J. C. Garth, and S. Woolf. Analytical neutral particle benchmarks in half-space geometry. In Kusters et al. [KSW93], pages 284–295. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Ganapol:1993:CEB

[GGW93b] B. D. Ganapol, J. C. Garth, and S. Woolf. A continuous energy benchmark for neutral particle motion in a semiinfinite medium. In Kusters et al. [KSW93], pages 272– 283. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Gao:2020:MES

[GGZ^{+}20] T. Gao, Y. Guo, B. Zhang, P. Cicotti, Y. Lu, P. Balaji, and M. Taufer. Memoryefficient and skew-tolerant MapReduce over MPI for supercomputing systems. IEEE Transactions on Parallel and Distributed Systems, 31(12): 2734–2748, 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Greenspan:1990:SDM

[GH90] Donald Greenspan and Larry F.

Heath. Supercomputer simulation of the modes of colliding microdrops of water. Technical report 273, University of Texas at Arlington, Dept. of Mathematics, Research Center for Advanced Study (RCAS), Arlington, TX, USA, 1990. 5 + 9 pp.

Greenspan:1991:SSM

[GH91] D. Greenspan and L. F.

Heath. Supercomputer simulation of the modes of colliding microdrops of water. Journal of physics D: applied physics, 24(11):2121–??, November 14, 1991. CODEN JPAPBE. ISSN 0022-3727.

Gehin:1993:TDF

[GH93] J. C. Gehin and A. F. Henry. Time-dependent discontinuity factors for transient nodal models. In Kusters et al. [KSW93], pages 496–506. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Gentzsch:1994:HCN

[GH94a] Wolfgang Gentzsch and Uwe Harms, editors. Highperformance computing and networking: International Conference and Exhibition, Munich, Germany, April 18– 20, 1994: proceedings, volume 797 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CO-DEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994). DM96.00. Two volumes.

Gentzsch:1994:HPCa

[GH94b] Wolfgang Gentzsch and Uwe Harms, editors. Highperformance computing and networking: International Conference and Exhibition, Munich, Germany, April 18– 20, 1994: proceedings, volume 796 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CO-DEN LNCSD9. ISBN 3-540-57980-X (Berlin), 0-387-57980-X (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994). DM96.00. Two volumes.

Gentzsch:1994:HPCb

[GH94c] Wolfgang Gentzsch and Uwe Harms, editors. Highperformance computing and networking: International Conference and Exhibition, Munich, Germany, April 18– 20, 1994: proceedings, volume 797 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CO-DEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994). DM96.00. Two volumes.

Gharachedaghi:1984:SNG

[Gha84] Afshin Gharachedaghi. A survey of next generation multiprocessor supercomputer projects. Thesis (M.S.), Santa Clara University, School of Engineering, Santa Clara, CA, USA, 1984. 144 pp.

Gharib:1996:PEP

[Gha96] Morteza Gharib. Perspective: The experimentalist and the problem of turbulence in the age of supercomputers. Journal of fluids engineering, 118 (2):233–??, ???? 1996. CO-DEN JFEGA4. ISSN 0098-2202.

Giorgino:2010:DCV

[GHdF10] Toni Giorgino, M. J. Harvey, and Gianni de Fabritiis. Distributed computing as a virtual supercomputer: Tools to run and manage large-scale BOINC simulations. Computer Physics Communications, 181(8): 1402–1409, August 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465510001189.

Gokhale:1995:PMT

[GHI95] Maya Gokhale, Bill Holmes, and Ken Iobst. Processing in memory: the Terasys massively parallel PIM array. Computer, 28(4):23– 31, April 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gokhale:1991:BUHb

[GHK^{+}91] Maya Gokhale, William Holmes, Andrew Kopser, Sara Lucas, Ronald Minnich, Douglas Sweely, and Daniel Lopresti. Building and using a highly parallel programmable logic array. Computer, 24(1):81– 89, January 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

George:1987:SCF

[GHNL87] Alan George, Michael T.

Heath, Esmond Ng, and Joseph Liu. Symbolic Cholesky factorization on a local-memory multiprocessor. Parallel Computing, 5 (1-2):85–95, July 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing— issues in applied research and development (Loen, 1986).

Gustafson:1986:AHV

[GHS86] John L. Gustafson, Stuart Hawkinson, and Ken Scott. The architecture of a homogeneous vector supercomputer. Journal of Parallel and Distributed Computing, 3 (3):297–304, September 1986. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Goel:1993:NSR

[GHW93] B. Goel, W. Hoebel, and H. Wuerz. Numerical simulation of radiation transport on supercomputers. In Kusters et al. [KSW93], pages 63–74.

ISBN 3-923704-11-9. LCCN ???? Two volumes.

Geuder:1994:SEC

[GHWZ94] U. Geuder, M. Haerdtner, B. Woerner, and R. Zink. Scalable execution control of grid-based scientific applications on parallel systems. In IEEE [IEE94c], pages 788– 795. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Gordon:1993:OMC

[GI93] R. S. Gordon and M. S. Imbadi. Optimisation of a modular closed circuit wind/ water tunnel using computational fluid dynamics. In Anonymous [Ano93-31], pages 787–794. ISBN 0-947719-62-8. LCCN ????

Gibert:1993:ODA

[Gib93]  P. Gibert.  Optimal de-

sign and automotive industry. In Anonymous [Ano93-31], pages 685–692. ISBN 0-947719-62-8. LCCN ????

Gibson:1995:NPC

[Gib95] Tamara L. Gibson. NAS parallel conjugate gradient benchmark on the Cray T3D. Technical report SRC-TR-94-129, Supercomputing Research Center: IDA, Lanham, MD, USA, January 20, 1995. 14 pp.

Gibbs:2001:CCS

[Gib01] W. Wayt Gibbs. Cybernetic cells: The simplest living cell is so complex that supercomputer models may never simulate its behavior perfectly. but even imperfect models could shake the foundations of biology. Scientific American, 285(2): 52–57, August 2001. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.
sciam.com/2001/0801issue/
2001/0801issue/0801hargrove.
html.

Guinea:1993:DAM

[GIBGA93] D. Guinea, A. Ibanez, P. Bustos, and M. C. Garcia-Alegre. Designing the autonomous mobile unit: The frontier of complexity and uncertainty. In Anonymous [Ano93-31], pages 189–196. ISBN 0-947719-62-8. LCCN ????

Giese:1996:SCR

[Gie96] G. Giese. Stochastic coefficient of restitution: a possible way out of inelastic collapse. In Wolf et al. [WSB96], pages 335–340. ISBN 981-02-2635-7. LCCN ????

Grinberg:2012:TCA

[GIF^{+}12] Leopold Grinberg, Joseph A.

Insley, Dmitry A. Fedosov, Vitali Morozov, Michael E. Papka, and George Em Karniadakis. Tightly coupled atomistic-continuum simulations of brain blood flow on petaflop supercomputers. Computing in Science and Engineering, 14(6):58–67, November/December 2012. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Giglio:1994:ASA

[Gig94] Dominick A. Giglio, editor. Algorithms for synthetic aperture radar imagery: 6–7 April 1994, Orlando, Florida, number 2230 in Proceedings — SPIE The International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1994. ISBN 0-8194-1534-0. ISSN 0361-0748. LCCN TK6592.S95 A44 1994.

Giloi:1988:STM

[Gil88] W. K. Giloi. SUPRENUM: a trendsetter in modern supercomputer development. Parallel Computing, 7(3):283– 296, September 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gilg:1992:NSM

[Gil92] A. Gilg. Numerische Simulation in der Mikroelektronik. In Meuer [Meu92c], pages 21–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Gillevet:1993:MGW

[Gil93] P. M. Gillevet. Multiplex genomic walking: Integration of the wet lab and computer lab into a single prototyping environment. In Lim et al. [L^{+}93], pages 197–206. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Giloi:1994:SSG

[Gil94a] W. K. Giloi. The SUPRENUM supercomputer: Goals, achievements, and lessons learned. Parallel Computing, 20(10– 11):1407–1425, November 1, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
parco/cas_sub/browse/browse.
cgi?year=1994&volume=20&
issue=10-11&aid=905.

Giloi:1994:PSA

[Gil94b] Wolfgang K. Giloi. Parallel supercomputer architectures and their programming models. Parallel Computing, 20(10–11):1443– 1470, November 1, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
parco/cas_sub/browse/browse.
cgi?year=1994&volume=20&
issue=10-11&aid=907.

Ginsberg:1982:SOE

[Gin82] Myron Ginsberg. Some observations on evaluation of a Cray-1 for an industrial research environment. Research Publication — General Motors Research Laboratories, May 1982. CODEN GMRLAW. ISSN 0099-9326.

Ginsberg:1993:CUS

[Gin93] Myron Ginsberg. Challenges to the use of supercomputers and scientific visualization for automotive applications. Computers and Graphics, 17(5):507–515, September– October 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Girkar:1991:FPT

[Gir91] Milind Girkar. Functional parallelism theoretical foundations and implementations. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1991. x + 95 pp.

Gisselquist:1986:ECC

[Gis86] Richard Gisselquist. An experimental C compiler for the Cray 2 computer. ACM SIG-PLAN Notices, 21(9):32–41, September 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gannon:1987:IMH

[GJ87] Dennis B. Gannon and William Jalby. The influence of memory hierarchy on algorithm organization: Programming FFTs on a vector multiprocessor. Technical Report CSRD 663, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 25 pp.

Gallivan:1988:POD

[GJG88] Kyle A. Gallivan, William Jalby, and Dennis B. Gannon. On the problem of optimizing data transfers for complex memory systems. Technical Report CSRD 802, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 238–253 pp.

Gallivan:1986:UBL

[GJM86] Kyle A. Gallivan, William Jalby, and Ulrike Meier. The use of BLAS3 in linear algebra on a parallel processor with a hierarchical memory. Technical Report CSRD-610, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 5 pp.

Ghodgaonkar:1994:MPP

[GJP94] M. D. Ghodgaonkar, A. K. Jethra, and S. S. Pande. Multitransputer parallel processing system (MTPPS). In Mahajan et al. [M^{+}94], pages 19–23. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Guidec:1996:OFS

[GJP96a] F. Guidec, J-M Jezequel, and J-L Pacherie. An objectoriented framework for supercomputing. The Journal of Systems and Software, 33(3): 239–??, ???? 1996. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Guidec:1996:OOF

[GJP96b] F. Guidec, J.-M. Jézéquel, and J.-L. Pacherie. An object-oriented framework for supercomputing. The Journal of Systems and Software, 33(3):239–251, June 1996. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Gaertel:1993:PIS

[GJS93] U. Gaertel, W. Joppich, and A. Schueller. Parallelization of the IFS: Status report and first results. In Hoffmann and Kauranne [HK93b], pages 152–156. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Gaertel:1994:MWF

[GJS94] U. Gaertel, W. Joppich, and A. Schueller. Medium-range weather forecast on parallel systems. In IEEE [IEE94c], pages 388–391. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Gallivan:1991:SBP

[GJW91] Kyle A. Gallivan, William Jalby, and Harry A. G. Wijshoff. Some basic performance measurements of the 16 × 16 Cedar configuration. Technical Report CSRD 1146, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1991. 83 pp.

Gelernier:1992:SRG

[GK92] D. Gelernier and D. Kaminsky. Supercomputing out of recycled garbage: Preliminary experience with Piranha. In ACM [ACM92b], pages 417–427. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Gupta:1993:PPL

[GK93] Anshul Gupta and Vipin Kumar. Performance properties of large scale parallel systems. Journal of Parallel and Distributed Computing, 19(3):

234–??, November 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gottlieb:2018:SEA

[GK18] Sigal Gottlieb and Gaurav Khanna. Supercomputingenabled advances in science and engineering. Computing in Science and Engineering, 20(4):8–9, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040008.
html.

Grimes:1987:SLD

[GKL^{+}87] Roger Grimes, Henry Krakauer, John Lewis, Horst Simon, and Su-Hai Wei. The solution of large dense generalized eigenvalue problems on the Cray X-MP/24 with SSD. Journal of Computational Physics, 69(2): 471–481, April 1987. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0021999187901781.

Galick:1991:ISE

[GKR91] Albert T. Galick, Thomas Kerkhoven, and Umberto Ravaioli. Iterative solution of the eigenvalue problem for a dielectric waveguide. Technical Report CSRD 1119, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1991. 20 pp.

Germann:2009:TMD

[GKS09] Timothy C. Germann, Kai Kadau, and Sriram Swaminarayan. 369 Tflop/s molecular dynamics simulations on the petaflop hybrid supercomputer ‘Roadrunner’. Concurrency and Computation: Practice and Experience, 21(17):2143–2159, December 10, 2009. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gmeiner:2014:PMH

[GKSR14] Björn Gmeiner, Harald Köstler, Markus Stürmer, and Ulrich Rüde. Parallel multigrid on hierarchical hybrid grids: a performance study on current high performance computing clusters. Concurrency and Computation: Practice and Experience, 26(1):217– 240, January 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gallopoulos:1988:BID

[GL88] E. J. (Efstratios J.) Gallopoulos and Daeshik Lee. Boundary integral domain decomposition on hierarchical memory multiprocessors. Technical Report CSRD 752, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 488–499 pp.

Grimsrud:1989:IDP

[GL89] A. Grimsrud and G. Lorig. Implementing a distributed process between workstation and supercomputer. In Brebbia and Peters [BP89b], pages 133–144. ISBN 0-444-88110-7, 1-85312-044-8, 0-945824-27-0. LCCN TA345 .A66 1989 [1-2] (1989). Two volumes.

Glowinski:1990:CMA

[GL90] R. Glowinski and A. Lichnewsky, editors. Computing methods in applied sciences and engineering: Proceedings of the Ninth International Conference on Computing Methods in Applied Sciences and Engineering, Paris, France, January 29– February 2, 1990, volume 9. SIAM Press, Philadelphia, PA, USA, 1990. ISBN 0-89871-264-5. LCCN QC39 .I49 1990.

Goto:1991:IJS

[GL91] Eiichi Goto and K. F. Loe. Issues in Josephson supercomuputer design. Studies in Josephson supercomputers. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1991. ISBN 981-02-0129-X. vii + 178 pp. LCCN TK7895.P3 I84 1991.

Gelberg:1992:PWV

[GL92] Larry Gelberg and Haim Levkowitz, editors. Volume visualization: Workshop — October 1992, Boston, MA, Workshop on Volume Visualization 1992. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-527-5 (soft cover), 0-89791-528-3 (hard cover). LCCN T385 .W75 1992.

Graf:1993:EIN

[GL93a] U. Graf and S. Langenbuch.

Experiences in implementing nuclear simulation codes on distributed and shared memory parallel computers. In Kusters et al. [KSW93], pages 17–28. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Grashoff:1993:ICI

[GL93b] H. Grashoff and J. A. Long.

Intelligent co-operative information systems conflict detection and resolution. In Anonymous [Ano93-31], pages 265–272. ISBN 0-947719-62-8. LCCN ????

Gropp:1994:SUT

[GL94] W. Gropp and E. Lusk. Scalable Unix tools on parallel processors. In IEEE [IEE94c], pages 56–62. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Gropp:1996:HMI

[GL96a] W. Gropp and E. Lusk. A high-performance MPI implementation on a sharedmemory vector supercomputer. Parallel Computing, 22 (11):1513–??, ???? 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gropp:1996:HPM

[GL96b] W. Gropp and E. Lusk. A high-performance MPI implementation on a sharedmemory vector supercomputer. Parallel Computing, 22 (11):1513–??, ???? 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gropp:1997:HPM

[GL97] William Gropp and Ewing Lusk. A high-performance MPI implementation on a shared-memory vector supercomputer. Parallel Computing, 22(11):1513–1526, January 26, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        parco/cas_sub/browse/browse.
                                        cgi?year=1997&volume=22&
                                        issue=11&aid=1113.

Glasser:1993:RMA

[Gla93] L. A. Glasser. A road map to ARPA involvement in electronic packaging. Computer, 26(4):82–86, April 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Glenn:1988:PPH

[Gle88] Ray R. Glenn. Performance prediction for the Horizon supercomputer. Technical report SRC-TR-88-012, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 4 pp.

Glenn:1991:CMH

[Gle91] Ray R. Glenn. Characterizing memory hot spots in a shared memory MIMD machine. Technical report SRC-TR-91-039, Supercomputing Research Center: IDA, Lanham, MD, USA, October 15, 1991. 24 pp.

Glendinning:1993:MMP

[Gle93] I. Glendinning. 93SC041 the MPI message passing interface. In Anonymous [Ano93-31], pages 229–236. ISBN 0-947719-62-8. LCCN ????

Grave:1994:VSC

[GLH94] M. Grave, Y. Le Lous, and W. T. Hewitt, editors. Visualization in scientific computing: 1st Workshop — 1991: Paris, Focus on Computer Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-56147-1, 0-387-56147-1. LCCN T385 .V59 1994.

Gloudeman:1984:AIS

[Glo84] I. Gloudeman. The anticipated impact of supercomputers on finite element analysis. Proceedings of the IEEE, 72:80–84, 1984. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Glowinski:1989:SFE

[Glo89] R. Glowinski. Supercomputing and the finite element approximation of the Navier– Stokes equations for incompressible viscous fluids. In Chao et al. [COS89], pages 277–315. ISBN 3-540-50872-4, 0-387-50872-4. ISSN 0176-5035. LCCN ????

George:2011:NGF

[GLS11] Alan George, Herman Lam, and Greg Stitt. Novo-G: At the forefront of scalable reconfigurable supercomputing. Computing in Science and Engineering, 13(1):82– 86, January/February 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Guarna:1987:IDP

[GM87] Vincent A. Guarna and Allen Malony. Instrumentation for the development of parallel programs. Technical Report CSRD 638, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1987. 3 pp.

Gokhale:1993:FCD

[GM93a] Maya B. Gokhale and Ron Minnich. FPGA computing in data parallel C. Technical report SRC-TR-93-097, Supercomputing Research Center: IDA, Lanham, MD, USA, April 5, 1993. 8 pp.

Groth:1993:PRO

[GM93b] A. Groth and Mertens. Process related optimization of the CAD/CAM-Application for simultaneous product finding processes. In Anonymous [Ano93-31], pages 43– 50. ISBN 0-947719-62-8. LCCN ????

Ghose:1994:UAT

[GM94a] K. Ghose and N. Mehdiratta. A universal approach for task scheduling for distributed memory multiprocessors. In IEEE [IEE94c], pages 577–584. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Gopinath:1994:VDC

[GM94b] K. Gopinath and R. Manohara Rao. Verification of the DASH cache consistency protocols. In Balakrishnan [Bal94], pages 16–31. ISBN 0-07-462044-4. LCCN ????

Goyal:1993:SIP

[GMBW93] G. Goyal, U. Mass, F. Behrendt, and J. Warnatz. 93SC019 simulation of ignition processes in one and two dimensional geometries. In Anonymous [Ano93-31], pages 123– 130. ISBN 0-947719-62-8. LCCN ????

George:2000:RTS

[GMF00] Alan D. George, Jeff Markwell, and Ryan Fogarty. Real-time sonar beamforming on high-performance distributed computers. Parallel Computing, 26(10):1231– 1252, August 15, 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/35/21/42/31/23/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/42/31/
23/article.pdf.

Gong:1994:CAF

[GMG94] C. Gong, R. Melhem, and R. Gupta. Compiler assisted fault detection for distributed-memory systems. In IEEE [IEE94c], pages 373– 380. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Greiss:1990:HDC

[GML90] H. Greiss, D. Mukhedkar, and P. J. Lagace. Heat dissipation computations of a HVDC ground electrode using a supercomputer. IEEE transactions on power delivery, 5(4):1802–??, October 1, 1990. CODEN IT-PDE5. ISSN 0885-8977 (print), 1937-4208 (electronic).

Gwun:1991:PRM

[GMMT91] O. Gwun, S. Murata, K. Murakami, and S. Tomita. A parallel rendering machine for high-speed ray tracing. In Anonymous [Ano91q], pages 173–182. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Gregoire:1997:PEA

[GMS97a] J. P. Gregoire, J. D. Mattei, and G. Simeoni. Parallelization of Estet-Astrid code on Cray C98. Lecture Notes in Computer Science, 1225:970–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gregoire:1997:PEC

[GMS97b] J. P. Gregoire, J. D. Mattei, and G. Simeoni. Parallelization of Estet-Astrid code on Cray C98. Lecture Notes in Computer Science, 1225:970–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunzinger:1993:ASP

[GMSB93] A. Gunzinger, U. Mueller, W. Scott, and B. Baeumle. Achieving supercomputer performance with a parallel array processor. In Anonymous [Ano93c], pages 55–58. ISBN ???? LCCN ????

Gillan:2011:SIJ

[GMSS^{+}11] Charles Gillan, Simon McIntosh-Smith, Nico Sanna, Stan Scott, and Thomas Steinke. Special issue of the Journal of Parallel and Distributed Computing: Novel architectures for high-performance computing. Journal of Parallel and Distributed Computing, 71(2):333, February 2011. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Gallivan:1991:MPS

[GMW91] Kyle A. Gallivan, Bret A. Marsolf, and Harry A. G. Wijshoff. MCSPARSE: a parallel sparse unsymmetric linear system solver. Technical Report CSRD 1142, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1991. 61 pp.

Gallivan:1994:PSU

[GMW94] K. A. Gallivan, B. A. Marsolf, and H. A. G. Wijshoff. The parallel solution of unsymmetric sparse linear systems using the H0* reordering and an associated factorization. In Anonymous [Ano94-134], pages 419–430. ISBN ???? LCCN ????

Ganapol:1993:SPN

[GNJW93] B. D. Ganapol, D. W. Nigg, S. N. Jahshan, and C. A.

Wemple. The searchlight problem for neutrons in a semi-infinite medium. In Kusters et al. [KSW93], pages 137–150. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Gokhale:1989:LDG

[Gok89]  Maya B. Gokhale.  The

logic description generator: a functional hardware description language. Technical report SRC-TR-89-008, Supercomputing Research Center: IDA, Lanham, MD, USA, November 13, 1989. 14 pp.

Gokhale:1990:LDG

[Gok90a]  Maya B. Gokhale.  The

Logic description generator. Technical report SRC-TR-90-011, Supercomputing Research Center: IDA, Lanham, MD, USA, June 19, 1990. 13 pp.

Gokhale:1990:SRL

[Gok90b] Maya B. Gokhale. SPLASH:

a reconfigurable linear logic array. Technical report SRC-TR-90-012, Supercomputing Research Center: IDA, Lanham, MD, USA, April 12, 1990. 16 pp.

Gokhale:1991:BUHa

[Gok91] Maya B. Gokhale. Building and using a highly parallel programmable logic array. Technical report SRC-TR-91-036, Supercomputing Research Center: IDA, Lanham, MD, USA, January 1991. 9 pp.

Gokhale:1992:MPP

[Gok92]  Maya B. Gokhale.  A

massively parallel processorin-memory array and its programming environment. Technical report SRC-TR-92-076, Supercomputing Research Center: IDA, Lanham, MD, USA, November 13, 1992. 19 pp.

Goldberg:1991:CWE

[Gol91a] David Goldberg. Corrigendum: “What every computer scientist should know about floating-point arithmetic”. ACM Computing Surveys, 23(3):413, September 1991. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [Gol91b, Wic92, Dun92].

Goldberg:1991:WEC

[Gol91b] David Goldberg. What every computer scientist should know about floating-point arithmetic. ACM Computing Surveys, 23(1):5–48, March 1991. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0360-0300/
                                        103163.html.  See also

[Gol91a, Dun92, Wic92].

Goldhirsch:1996:MKR

[Gol96] I. Goldhirsch. Microstructures and kinetics in rapid granular flows. In Wolf et al. [WSB96], pages 251– 266. ISBN 981-02-2635-7. LCCN ????

Goller:1999:PPS

[Gol99] A. Goller. Parallel processing strategies for large SAR image data sets in a distributed environment. Computing: Archiv fur informatik und numerik, 62(4):277–291, 1999. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer-ny.com/
link/service/journals/00607/
bibs/9062004/90620277.htm;
http://link.springer-ny.
com/link/service/journals/
00607/papers/9062004/90620277.
pdf.

Gonzalez:1993:ASA

[Gon93] O. V. Gonzalez. Application of severe accident code with graphic interface. In Kusters et al. [KSW93], pages 660–666. ISBN 3-923704-11-

9. LCCN ???? Two volumes.

Gooley:1988:PAU

[Goo88] Markian Myron Gooley. PrologX at the University of Illinois. Technical Report CSRD 718, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1988. 6 pp.

Goodman:1997:MPP

[Goo97] J. Goodman, editor. Massively parallel processing using optical interconnections:

International conference; 4th — June 1997, Montreal, Canada, number 4 in PRO-CEEDINGS of THE IN-TERNATIONAL CONFER-ENCE ON MASSIVELY PARALLEL PROCESSING USING OPTICAL INTER-CONNECTIONS 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7975-1, 0-8186-7974-3, 0-8186-7976-X. LCCN ????

Gornish:1989:CTA

[Gor89] Edward H. Gornish. Compile time analysis for data prefetching. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1989. x + 100 pp.

Goto:1991:CSE

[Got91a] E. Goto. Computational science and engineering. In Anonymous [Ano91q], pages 1–6. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Gottschewski:1991:SDG

[Got91b] Jurgen Gottschewski. Supercomputing during the German reunification. Technical report TR 91-5, Konrad-Zuse-Zentrum fur Informationstechnik Berlin, Berlin, Germany, June 1991. 6 pp.

Goudreau:1990:SNS

[Gou90] G. L. Goudreau. Supercomputing and nonlinear seismic structural response of freeway structures. Technical Report UCRL JC-104837, Lawrence Livermore National Laboratory, Livermore, CA, USA, 1990. 4 pp. Cover title. Prepared for presentation at the ASCE Structures Congress ’91, Indianapolis, Ind. “UCRL JC-104837.” Microfiche. Springfield, Va.: National Technical Information Service, 1992. 1 microfiche: negative; 11 x 15 cm.

Gajski:1985:EIM

[GP85] Daniel D. Gajski and Jih-Kwon Peir. Essential issues in multiprocessor systems. Computer, 18(6): 9–27, June 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Girkar:1988:PPP

[GP88] Milind Girkar and C. D.

(Constantine D.) Polychronopoulos. Partitioning programs for parallel execution. Technical Report CSRD 765, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 216–229 pp.

Glenn:1990:IMP

[GP90] Ray R. Glenn and Daniel V. Pryor. Instrumentation for a massively parallel MIMD application. Technical report SRC-TR-90-022, Supercomputing Research Center: IDA, Lanham, MD, USA, October 31, 1990. 22 pp.

Girkar:1991:FFP

[GP91a] Milind Girkar and C. D. (Constantine D.) Polychronopoulos. Formalizing functional parallelism. Technical Report CSRD 1141, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 31 pp.

Girkar:1991:HIR

[GP91b] Milind Girkar and C. D. (Constantine D.) Polychronopoulos. The HTG: an intermediate representation for programs based on control and data dependences. Technical Report CSRD 1046, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 30 pp.

Girkar:1991:ODC

[GP91c] Milind Girkar and C. D. (Constantine D.) Polychronopoulos. Optimization of data/ control conditions in task graphs. Technical Report CSRD 1234, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1991. 17 pp.

Gladwell:1993:PSA

[GP93a] Ian Gladwell and Marcin Paprzycki. Parallel solution of almost block diagonal systems on the Cray Y-MP using level 3 BLAS. Journal of Computational and Applied Mathematics, 45(1–2): 181–189, April 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Golini:1993:VPS

[GP93b] S. Golini and F. Papetti.

93SC007 validation of physical sub-models in a DI diesel engine. In Anonymous [Ano93-31], pages 139–146. ISBN 0-947719-62-8. LCCN ????

Gulyaev:1993:ICO

[GP93c] Yuri V. Gulyaev and Dennis R. Pape, editors. International Conference on Optical Information Processing: 2–7 August 1993, St. Petersburg, Russia, number 2051 in Proceedings — SPIE The International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1993. ISBN 0-8194-1310-0. ISSN 0361-0748. LCCN TA1630 .I568 1993.

Gall:2006:HSD

[GP06] R. Gall and D. Parsons. It’s hurricane season: Do you know where your storm is? IEEE Spectrum, 43(8): 27–32, August 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gajski:1982:SOD

[GPKK82] D. D. Gajski, D. A. Padua, D. J. Kuck, and R. H. Kuhn. A second opinion on data flow machines and languages. Computer, 15(2): 58–69, February 1982. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Golub:1986:PBS

[GPS86] Gene H. (Gene Howard) Golub, Robert J. Plemmons, and Ahmed Sameh. Parallel block schemes for large scale least squares computations. Technical Report CSRD-574, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 20 pp.

Gallivan:1990:PAD

[GPS90] Kyle A. Gallivan, Robert J. Plemmons, and Ahmed Sameh. Parallel algorithms for dense linear algebra computations. Technical Report CSRD 992, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1990. 82 pp.

Granston:1991:SPC

[GR91] Elana Denise Granston and Vincent Russo. Signaturebased polymorphism for C++. Technical Report CSRD 1104, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 15 pp.

Gupta:1994:DDM

[GR94] S. Gupta and E. Rothberg.

DME: A Distributed Matrix Environment. In IEEE [IEE94c], pages 629–636. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Grassl:1991:PPA

[Gra91] C. M. Grassl. Parallel performance of applications on supercomputers. Parallel Computing, 17(10–11):1257– 1273, December 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Granston:1992:RMA

[Gra92] Elana Denise Granston. Reducing memory access delays in large-scale shared-memory multiprocessors. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1992. v + 107 pp.

Graffunder:1993:BPI

[Gra93a] Sara K. Graffunder. Barrierbreaking performance for industrial problems on the Cray C916. Proceedings of the Supercomputing Conference, pages 516–519, 1993. CO-DEN 85ODA8. ISBN 0-8186-4340-4. ISSN 1063-9535.

Grave:1993:DVF

[Gra93b] Michel Grave. Distributed visualization in flow simulations. Computers and Graphics, 17(1):9–14, January– February 02, 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Grayson:1993:EER

[Gra93c] Lawrence P. Grayson, editor. Engineering education: renewing America’s technology: 23rd Annual conference on frontiers in education — November 6–9, 1993, Washington, DC, Frontiers in Education Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1482-4, 0-7803-1483-2, 0-7803-1484-0. ISSN 0190-5848. LCCN ???? IEEE Catalog number: 93CH3373-8.

Grayson:1994:FEP

[Gra94] Lawrence P. Grayson, editor. Frontiers in education: proceedings, twentyfourth annual conference, November 2–6, 1994, San Jose, California: educating engineers for world competition. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. CO-DEN PFECDR. ISBN 0-7803-2414-5, 0-7803-2413-7, 0-7803-2415-3. ISSN 0190-5848. LCCN T 62 F76 1994.

Greenspan:1988:MCS

[Gre88a] D. Greenspan. Mechanisms of capillarity via supercomputer simulation. Computers and Mathematics with Applications, 16(4):331–??, 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Greenspan:1988:QCV

[Gre88b] D. Greenspan. Quasimolecular channel and vortex street modeling on a supercomputer. Computers and Mathematics with Applications, 15(2):141–151, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0898122188900831.

Greenspan:1988:SSS

[Gre88c] D. Greenspan. Supercomputer simulation of sessile and pendent drops. Mathematical and computer modelling, 10(12):871–??, 1988. CODEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Greenspan:1989:PSB

[Gre89a] D. Greenspan. Particle simulation of biological sorting on a supercomputer. Computers and Mathematics with Applications, 18(9):823–??, 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Greenspan:1989:SSL

[Gre89b] D. Greenspan. Supercomputer simulation of liquid drop formation, fall, and collision. Applied mathematical modelling, 13(10):562– ??, October 1, 1989. CO-DEN AMMODL. ISSN 0307-904X (print), 1872-8480 (electronic).

Greenspan:1990:SSC

[Gre90a] D. Greenspan. Supercomputer simulation of colliding microdrops of water. Computers and Mathematics with Applications, 19(7): 91–97, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/089812219090197R.

Greenspan:1990:SSA

[Gre90b] D. Greenspan. Supercomputer studies in absolute minimization. Computers and Mathematics with Applications, 20(11):47–??, 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Gregoire:1990:EVC

[Gre90c] J. P. Gregoire. Efficient vectorization of the conjugate gradient method. In Pitcher [Pit90], pages 353– 359. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Green:1991:FFC

[Gre91a] Paul E. Green. The future of fibre-optic computer networks. Computer, 24(9):78– 87, September 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Greenspan:1991:SSL

[Gre91b] D. Greenspan. Supercomputer simulation of liquid drop formation on a solid surface. International Journal for Numerical Methods in Fluids, 13(7):895–??, October 1, 1991. CODEN IJNFDW. ISSN 0271-2091.

Greenstein:1994:SSE

[Gre94] S. Greenstein. From superminis to supercomputers: Estimating surplus in the computing market. Working paper series, ??(4899):ALL, ???? 1994. ISSN 0898-2937.

Gustafson:1991:DSF

[GREC91] John Gustafson, Diane Rover, Stephen Elbert, and Michael Carter. The design of a scalable, fixed-time computer benchmark. Journal of Parallel and Distributed Computing, 12(4): 388–401, August 1991. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Griffin:1986:EVA

[Gri86] Lisa Ann Willis Griffin. Explicit vectorization and application of a finite volume Euler equation solver on the NASA Langley VPS-32 supercomputer for transonic flow calculation. Thesis (M.S.), Mississippi State University. Department of Aerospace Engineering, Mississippi State, MS, USA, 1986. ix + 58 pp.

Grier:1988:SMC

[Gri88] David Alan Grier. Supercomputers and Monte Carlo experiments. Chance, 1(2): 19–28, Spring 1988. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Griffith:1990:SSS

[Gri90] Daniel A. Griffith. Supercomputing and spatial statistics: a reconnaissance. The Professional geographer: the journal of the Association of American Geographers, 42 (4):481–??, November 1990. ISSN 0033-0124.

Griebel:1992:CTS

[Gri92] M. Griebel. The combination technique for the sparse grid solution of PDEs on multiprocessor machines. Parallel Processing Letters, 2(1): 61–70, March 1992. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Grinstein:1993:SV

[Gri93] Georges G. Grinstein. Supercomputing and visualization, volume 17(1) of Computers and graphics. Pergamon Press, New York, NY, USA, January/February 1993. 117 pp.

Gross:1990:PSI

[Gro90] Phillip G. Gross, editor. Proceedings of the Seventeenth Internet Engineering Task Force, Pittsburgh Supercomputer Center, May 1–4, 1990. Corporation for National Research Initiatives, Reston, VA, USA, 1990.

Gross:1992:RSS

[Gro92a] G. Gross. Results of supercomputer simulations of meteorological mesoscale phenomena. Fluid dynamics research, 10(4/6):483–498, December 1992. ISSN 0169-5983.

Grossman:1992:SMR

[Gro92b] M. Grossman. Supercomputers — modeling reality. IEEE Spectrum, 29(9):56– 60, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grossman:1992:CAT

[Gro92c]  Morris Grossman.  The

challenges awaiting tomorrow’s supercomputers run the gamut from the grand to the mundane. IEEE Spectrum, 29(9):56–60, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Groscup:1993:IPX

[Gro93]  W. Groscup.  The Intel

Paragon[TM] XP/S supercomputer. In Hoffmann and Kauranne [HK93b], pages 173–187. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Gonzalez:1999:PPM

[GRRM99] J. A. Gonzalez, C. Rodriguez, J. L. Roda, and D. G. Morales. Performance and predictability of MPI and BSP programs on the CRAY T3E. In Dongarra et al. [DLM99], pages 27–34. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973 1999.

Gross:1993:FCV

[GRSS93] L. Gross, C. Roll, M. Schmauder, and W. Schoenauer. FIDISOL/ CADSOL and VECFEM-Powerful tools for the solution of partial differential equations. In Kusters et al. [KSW93], pages 785–?? ISBN 3-923704-11-9. LCCN ???? Two volumes.

Grund:1997:HLT

[Gru97] C. Grund. Hardware in the loop test bench for wheel slip systems. In Roller [Rol97], pages 113–120. ISBN 0-947719-88-1 (paperback). LCCN ????

Gallopoulos:1987:PBC

[GS87a] E. J. (Efstratios J.) Gallopoulos and Y. Saad. A parallel block cyclic reduction algorithm for the fast solution of elliptic equations. Technical Report CSRD 659, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 14 pp.

Gelberg:1987:SGE

[GS87b] Lawrence Gelberg and Thomas Stephenson. Supercomputing and graphics in the earth and planetary sciences. IEEE Computer Graphics and Applications, 7 (7):26–33, July 1987. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Girkar:1987:FVC

[GS87c] Milind Girkar and Milind Sohoni. On finding the vertex connectivity of graphs. Technical Report CSRD 669, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 5 pp.

Gallivan:1988:MCS

[GS88a] Kyle A. Gallivan and Ahmed Sameh. Matrix computation on shared-memory multiprocessors. Technical Report CSRD 760, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 70 pp.

Gallopoulos:1988:PBC

[GS88b] E. J. (Efstratios J.) Gallopoulos and Y. Saad. A parallel block cyclic reduction algorithm for the fast solution of elliptic equations. Technical Report CSRD 753, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 17 + [7] pp.

Gallopoulos:1989:PSP

[GS89a] E. J. (Efstratios J.) Gallopoulos and Y. Saad. On the parallel solution of parabolic equations. Technical Report CSRD 854, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1989. 22 pp.

Gallopoulos:1989:SFE

[GS89b] E. J. (Efstratios J.) Gallopoulos and Youcef Saad. Some fast elliptic solvers on parallel architectures and their complexities. Technical Report CSRD 862, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1989. 29 pp.

Gear:1989:SSS

[GS89c] C. W. Gear and H. D. Simon. Special section on sparse matrix algorithms on supercomputers. SIAM Journal on Scientific and Statistical Computing, 10(6):1135, ???? 1989. CODEN SIJCD4. ISSN 0196-5204.

Guerrini:1989:IRA

[GS89d] C. Guerrini and G. Spaletta.

Image reconstruction algorithm in tomography. A version for the Cray X-MP vector computer. Computers and Graphics, 13 (3):367–372, 1989. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Gallopoulos:1990:ESP

[GS90] E. J. (Efstratios J.) Gallopoulos and Y. Saad. Efficient solution of parabolic equations by polynomial approximation methods. Technical Report CSRD 969, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1990. 32 pp.

Gallopoulos:1992:ESP

[GS92a] E. J. (Efstratios J.) Gallopoulos and Y. Saad. Efficient solution of parabolic equations by Krylov approximation methods. Technical Report CSRD 1147, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. 30 pp.

Gurd:1992:MDP

[GS92b] J. R. Gurd and D. F. Snelling. Manchester Data-Flow: a progress report. In ACM [ACM92b], pages 216–225. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Gokhale:1993:DBC

[GS93] Maya B. Gokhale and Judith D. Schlesinger. A dataparallel bit-serial C (dbC). Technical report SRC-TR-93-096, Supercomputing Research Center: IDA, Lanham, MD, USA, May 1993. 14 pp.

Garg:1994:DON

[GS94a] S. Garg and H. A. Sholl. Determining the optimum number of processors to execute a scalable computation. In Balakrishnan [Bal94], pages 273–283. ISBN 0-07-462044-4. LCCN ????

Gee:1994:ECV

[GS94b] J. D. Gee and A. J. Smith.

The effectiveness of caches for vector processors. In Anonymous [Ano94-134], pages 333–343. ISBN ???? LCCN ????

Gokhale:1994:DPC

[GS94c] Maya B. Gokhale and Brian Schott. Data parallel C on a reconfigurable logic array. Technical report SRC-TR-94-121, Supercomputing Research Center: IDA, Lanham, MD, USA, October 1994. 23 pp.

Gross:1994:AIH

[GS94d] T. Gross and P. Steenkiste.

Architecture implications of high-speed I/O for distributedmemory computers. In Anonymous [Ano94-134], pages 176–185. ISBN ???? LCCN ????

Gupta:1994:TPS

[GS94e] V. Gupta and E. Schenfeld. Throughput performance of a synchronous, circuit-switched interconnection cached network. In Anonymous [Ano94-134], pages 246–255. ISBN ???? LCCN ????

Garg:2001:TOA

[GS01] Rajat P. Garg and Ilya Sharapov. Techniques for Optimizing Applications: High Performance Computing. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-093476-3. xliv + 616 pp. LCCN QA76.88 .G37

2002. URL books/apt.pdf;
 http://www.sun.com/books/
catalog/garg.html/index.

html. Part No. 806-6380-10 June 2001, Revision 01.

Garg:2006:OHR

[GS06] Rahul Garg and Yogish Sabharwal. Optimizing the HPCC randomaccess benchmark on Blue Gene/L supercomputer. ACM SIGMET-RICS Performance Evaluation Review, 34(1):369– 370, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Garza-Salazar:1995:RCH

[GSB95] D. A. Garza-Salazar and W. Boehm. Reducing communication by honoring multiple alignments. In ACM [ACM95a], pages 87–96. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Garcia:1994:ESU

[GSG^{+}94] G. D. Garcia, F. Suarez, M. Garcia, E. Lasso, R. Guzman, T. Carden, D. Watson, C. McGregor, F. Obeso, and M. Tarrio. An embedded solution using highperformance computing for cost effective on-line realtime monitoring of industrial processes. Lecture Notes in Computer Science, 796:40–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gallivan:1991:PDM

[GSZ91] Kyle A. Gallivan, Ahmed Sameh, and Zahari Zlatev. Parallel direct method codes for general sparse matrices. Technical Report CSRD 1143, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1991. 29 pp.

Golub:1991:IMC

[GT91] G. H. Golub and R. S. Tuminaro. Iterative methods for cyclically reduced nonself-adjoint linear systems. In Anonymous [Ano91q], pages 7–8. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Gruber:1994:PCI

[GT94] Ralf Gruber and Marco Tomassini, editors. Proceedings of the 6th Joint EPS-APS International Conference on Physics Computing: Physics Computing ’94, Palazzo dei Congressi, Lugano, Switzerland, 22–26 August 1994. European Physical Society, Geneva, Switzerland, 1994. ISBN 2-88270-011-3. LCCN ????

Gimenez:1997:SES

[GT97] C. Gimenez and G. N. Telles. Simultaneous engineering in subcontracting context. In Roller [Rol97], pages 201– 206. ISBN 0-947719-88-1 (paperback). LCCN ????

Granston:1991:DAS

[GTV91] Elana Denise Granston, Stephen W. Turner, and Alexander Veidenbaum. Design and analysis of a scalable, shared-memory system with support for burst traffic. Technical Report CSRD 1084, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 30 pp.

Guarna:1987:ACP

[Gua87a] Vincent A. Guarna. Analysis of C programs for parallelization in the presence of pointers. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1987. vii + 85 pp.

Guarna:1987:VPV

[Gua87b] Vincent A. Guarna. VPC, a proposal for a vector parallel C programming language. Technical Report CSRD 666, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. ii + 27 pp.

Guarna:1988:ECL

[Gua88a] Vincent A. Guarna. Extending the C language for parallel machines. Technical Report CSRD 722, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 22 pp.

Guarna:1988:FEP

[Gua88b] Vincent A. Guarna. Faust:

an environment for programming parallel scientific applications. Technical Report CSRD 757, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 8 pp.

Guarna:1988:FIE

[Gua88c] Vincent A. Guarna. Faust: an integrated environment for the development of parallel programs. Technical Report CSRD 825, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1988. 11 + [10] pp.

Guarna:1988:TAP

[Gua88d] Vincent A. Guarna. A technique for analyzing pointer and structure references in parallel restructuring compilers. Technical Report CSRD 721, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 9 pp.

Guelzow:1990:SEC

[Gue90] V. Guelzow. A supercomputing environment in climate research. In Pitcher [Pit90], pages 467–473. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Guidec:1996:OPS

[Gui96] F. Guidec. Object-oriented parallel software components for supercomputing. In D’Hollander et al. [DDC96], pages 247–254. ISBN 0-444-82490-1. LCCN QA76.58.P764 1995.

Guizzo:2005:IRS

[Gui05]  E. Guizzo.  IBM re-

claims supercomputer lead. IEEE Spectrum, 42(2):15– 16, February 2005. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2006:DS

[Gui06] Erico Guizzo. A desktop supercomputer. IEEE Spectrum, 43(1):54, January 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2008:WGS

[Gui08] Erico Guizzo. Winner: Geophysics solving the oil equation. IEEE Spectrum, 45(1): 32–36, January 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guizzo:2010:DQC

[Gui10] E. Guizzo. Does not quantum compute. IEEE Spectrum, 47 (1):42–43, January 2010. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gund:1988:SAM

[Gun88] Tamara Gund. Supercomputer applications in molecular modeling. IEEE engineering in medicine and biology magazine, 7(4):21–??, December 1, 1988. CODEN IEMBDE. ISSN 0739-5175.

Guo:1994:MPP

[Guo94]  Z.-Y. Guo, editor.  Mi-

croelectronic package and PCB technology: 1st International symposium — September 1994, Beijing, PRO-CEEDINGS OF THE IN-TERNATIONAL SYMPO-SIUM ON MICROELEC-TRONIC PACKAGE AND PCB TECHNOLOGY 1994; 1st. International Academic Publishers, Beijing, China, 1994. ISBN 7-80003-324-4. LCCN ????

Gupta:1988:CRM

[Gup88]  Sumnesh Gupta.  Com-

ments regarding Monte Carlo simulation of classical fluids on general purpose supercomputers. Computer Physics Communications, 50 (3):293–295, August 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465588901841.

Gupta:1994:SST

[Gup94] R. K. Gupta. Scientific supercomputing today: a perspective. In Balakrishnan [Bal94], pages 399–?? ISBN 0-07-462044-4. LCCN ????

Gurke:1988:ASE

[Gur88]  Renate Gurke.  Approxi-

mate solution of the Euclidean Traveling Salesman Problem on a Cray X-MP. Parallel Computing, 8(1-3): 177–183, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gurd:1994:SBB

[Gur94] J. R. Gurd. Supercomputing: big bang or steady state growth? ACM SIGARCH Computer Architecture News, 22(3):3–13, June 1994. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Gutbrod:1995:FRN

[Gut95] F. Gutbrod. A fast random number generator for the Intel Paragon supercomputer. Computer Physics Communications, 87(3):291– 306, June 1, 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/001046559500005Z.

Guzzi:1986:MRS

[Guz86] Mark David Guzzi. Multitasking runtime systems for the Cedar multiprocessor. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. v + 66 pp.

Guzzi:1987:CFP

[Guz87] Mark David Guzzi. Cedar Fortran programmer’s handbook. Technical Report CSRD 601, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. iv + 40 pp.

Guzzi:1988:CFO

[Guz88] Mark David Guzzi. Cedar Fortran and other vector and parallel Fortran dialects. Technical Report CSRD 731, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 8 pp.

Granston:1991:IHS

[GV91] Elana Denise Granston and Alexander Veidenbaum. An integrated hardware/software solution for effective management of local storage in high performance systems. Technical Report CSRD 1073, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1991. 32 pp.

Granston:1992:DRA

[GV92] Elana Denise Granston and Alexander Veidenbaum. Detecting redundant accesses to array data: extended version. Technical Report CSRD 1053, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1992. 37 pp.

Gonzalez-Velez:1996:DSP

[GV96a] H. Gonzalez-Velez. Designing a supercomputing policy for a developing country. In Roller [Rol96], pages 77–86. ISBN 0-947719-81-4. LCCN ????

Grayson:1996:HPP

[GV96b] Brian Grayson and Robert Van De Geijn. A high performance parallel Strassen implementation. Parallel Processing Letters, 6(1):3– 12, March 1996. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gregorio:1995:PNM

[GVBC95] J. A. Gregorio, F. Vallejo, R. Beivide, and C. Carrion. Petri net modeling of interconnection networks for massively parallel architectures. In ACM [ACM95a], pages 107–116. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Geers:1991:HEB

[GW91] N. Geers and W. Waelde.

Highly efficient basic numerical software for supercomputers. In Anonymous [Ano91q], pages 219–228. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Graf:1993:IEN

[GW93a] U. Graf and W. Werner. Improved efficiency in a numerical solution method for multidimensional hyperbolic fluid flow equations by a spatial discretisation operator of second order. In Kusters et al. [KSW93], pages 188– 197. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Groetzbach:1993:AFM

[GW93b] G. Groetzbach and M. Woerner. Analysis of flow mechanisms in Rayleigh-Benard convections at small Prandtl numbers. In Kusters et al. [KSW93], pages 236–247. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Gu:1993:NSA

[GW93c] C.-Y. Gu and E. Wilhelmsson. Numerical simulation of aerodynamic performance of high speed trains. In Anonymous [Ano93-31], pages 751– 758. ISBN 0-947719-62-8. LCCN ????

Geschiere:1995:ELG

[GW95] J. P. Geschiere and H. A. G. Wijshoff. Exploiting large grain parallelism in a sparse direct linear system solver. Parallel Computing, 21(8): 1339–1364, August 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Gentile:2004:PVS

[GW04] A. Gentile and D. S. Wills.

Portable video supercomputing. IEEE Transactions on Computers, 53 (8):960–973, August 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=1306990.

Gill:1993:FMT

[GWG93] A. Gill, J. Warnatz, and E. Guthell. 93SC022 flamelet modelling of turbulent diffusion flames in internal engine combustion. In Anonymous [Ano93-31], pages 115–122. ISBN 0-947719-62-8. LCCN ????

Groetzbach:1993:VTT

[GWH93] G. Groetzbach, M. Woerner, and E. Hesselschwerdt. Visualisation of three-dimensional time-dependent flow mechanisms in turbulent convection. In Kusters et al. [KSW93], pages 793–794. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Ghafoor:1992:DHS

[GY92] Arif Ghafoor and Jaehyung Yang. Distributed heterogeneous supercomputing management system. Technical report EE 92-45, Purdue University, School of Electrical Engineering, West Lafayette, IN, USA, October 1992. 32 pp.

Ghafoor:1993:DHS

[GY93a] Arif Ghafoor and Jaehyung Yang. A distributed heterogeneous supercomputing management system. Computer, 26(6):78–86, June 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gofuku:1993:MSO

[GY93b] A. Gofuku and H. Yoshikawa. A methodological study on organizing an intelligent CAD/CAE system for conceptual design of advanced nuclear reactor system. In Kusters et al. [KSW93], pages 454–465. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Gorb:2000:HPC

[GYL00] Leonid Gorb, Ilya Yanov, and Jerzy Leszczynski. High performance computing on the Cray T3E and IBM SP2 systems with the parallel version of GAUSSIAN 94. Parallel Computing, 26 (7–8):1043–1060, July 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-ng/
10/35/21/42/29/36/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/42/29/
36/article.pdf.

Gupta:1986:SLD

[GZA86] Viney K. Gupta, Scott D.

Zillmer, and Robert E. Allison. Solving large-scale dynamic systems using band Lanczos method in Rockwell NASTRAN on Cray X-MP. NASA conference publication, pages 236–246, 1986. CODEN NACPDX. ISSN 0191-7811.

Germain:2005:EPD

[GZE^{+}05] R. S. Germain, Y. Zhestkov, M. Eleftheriou, A. Rayshubskiy, F. Suits, T. J. C. Ward, and B. G. Fitch. Early performance data on the Blue Matter molecular simulation framework. IBM Journal of Research and Development, 49(2/3):447–455, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/germain.
pdf.

Gupta:1989:RMV

[GZR89] R. Gupta, A. Zorat, and I. V. Ramakrishnan. Reconfigurable multipipelines for vector supercomputers. IEEE Transactions on Computers, 38(9):1297–1307, September 1989. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=29468.

Gan:2022:TCG

[GZW^{+}22] Xinbiao Gan, Yiming Zhang, Ruibo Wang, Tiejun Li, Tiaojie Xiao, Ruigeng Zeng, Jie Liu, and Kai Lu. Tianhe-Graph: Customizing graph search for Graph500 on Tianhe supercomputer. IEEE Transactions on Parallel and Distributed Systems, 33(4): 941–951, April 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ha:1988:ENS

[Ha88]  Sung Nam Ha.  Experi-

mental numerical studies on a supercomputer of natural convection in an enclosure with localized heating. Thesis (Ph.D.), University of Texas at Arlington, Arlington, TX, USA, 1988. xv + 159 pp.

Ha:1990:ENS

[Ha90a] S. N. Ha. Experimental numerical studies on a supercomputer of natural convection in an enclosure with localized heating. Computers and Mathematics with Applications, 20(12):1–??, 1990. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Harrison:1990:CAV

[HA90b] Luddy Harrison and Zahira Ammarguellat. A comparison of automatic versus manual parallelization of the Boyer– Moore theorem prover. Technical Report CSRD 960, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1990. 17 pp.

Hejhal:1991:FCM

[HA91] Dennis A. Hejhal and Steven Arno. On Fourier coefficients of Maass waveforms for PSL(2,Z). Technical report SRC-TR-91-054, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1991. 57 pp.

Harrison:1992:PEV

[HA92] Williams Ludwell Harrison and Zahira Ammarguellat. A program’s eye view of Miprac. Technical Report CSRD 1227, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1992. 30 pp.

Hung:1993:PA

[HA93] S. L. Hung and H. Adeli.

Practical aspects. Neurocomputing, 5(6):287–??, December 1, 1993. CODEN NRCGEO. ISSN 0925-2312 (print), 1872-8286 (electronic).

Higuchi:1993:EPJ

[HAAS93] K. Higuchi, K. Asai, M. Akimoto, and S. Shingu. Effective performance of the JAERI Monte Carlo machine. In Kusters et al. [KSW93], pages 356–365. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Haberland:1986:SCS

[Hab86] J. Carl Haberland. Scientific computer systems corporation SCS-40, 1986. 1 videocassette (54 min.).

Haber:1989:SVR

[Hab89] R. B. Haber. Scientific visualization and the Rivers Project at the National Center for Supercomputing Applications. Computer, 22(8): 84–89, August 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Haber:1992:DER

[Hab92] Robert B. Haber. A distributed environment for runtime visualization and application steering in computational mechanics. Technical Report CSRD 1235, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1992. 39 pp.

Haefner:1991:FWS

[Hae91] James W. Haefner. Foodweb simulation on parallel computers: Inter-processor communication benchmarks.

Ecological Modelling, 54(1– 2):73–79, May 1991. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/030438009190099M.

Haghighat:1990:SDA

[Hag90] Mohammad Reza Haghighat.

Symbolic dependence analysis for high performance parallelizing compilers. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1990. xi + 115 pp.

Hagersten:2001:HPC

[Hag01]  Erik Hagersten.  High-

performance computers: Yesterday, today, and tomorrow. Lecture Notes in Computer Science, 1947:18–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1947/19470018.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1947/19470018.
pdf.

Hirzel:2013:ISP

[HAG^{+}13] M. Hirzel, H. Andrade, B. Gedik, G. Jacques-Silva, R. Khandekar, V. Kumar, M. Mendell, H. Nasgaard, S. Schneider, R. Soule, and K.-L. Wu. IBM Streams Processing Language: Analyzing Big Data in motion. IBM Journal of Research and Development, 57(3/4):7:1– 7:11, May–July 2013. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Haidar:1997:PFP

[Hai97] N. I. A. Haidar. Prediction of flow pressure losses in the Nubira air intake system. In Roller [Rol97], pages 431– 440. ISBN 0-947719-88-1 (paperback). LCCN ????

Hake:1989:LAS

[Hak89] J.-Fr. Hake. Linear algebra software on IBM and Cray computers. Journal of Computational and Applied Mathematics, 26(3): 311–325, July 1989. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Halford:1987:MSM

[Hal87] Robert J. Halford. Mass storage mechanization for Cray computer systems. Digest of Papers — IEEE Symposium on Mass Storage Systems, pages 52–57, 1987. CODEN DPISDX. ISBN 0-8186-0772-6. ISSN 1051-9173. IEEE Service Cent. Piscataway, NJ, USA.

Hall:1996:RSS

[Hal96] F. L. Hall. A review and some speculation about speed-flow relationships on freeways. In Wolf et al. [WSB96], pages 11–22. ISBN 981-02-2635-7. LCCN ????

Hammerslag:1990:FLB

[Ham90] David H. Hammerslag. Faust library browser: user’s manual. Technical Report CSRD 961, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1990. 15 + 4 pp.

Hamza:1994:PTI

[Ham94] M. H. Hamza, editor. Proceedings of the twelfth IASTED International Conference: Applied Informatics, Annecy, France, May 18–20, 1994, APPLIED INFORMATICS 1994. IASTED, Anaheim, CA, USA, 1994. ISBN 0-88986-190-0. ISSN 0013-5704. LCCN ????

Hanken:1989:DTM

[Han89] Jerry Hanken. Deep thought has miles to go. Chess life, 44 (3):22–??, March 1, 1989.

Han:1990:CLS

[Han90] Ningning Han. Computing local sequence similarities on a supercomputer. Technical report CS-90-52, Pennsylvania State University, Dept. of Computer Science, University Park, PA, USA, December 7, 1990. 16 pp.

Hansen:1994:SPL

[Han94] P. Brinch Hansen. SuperPascal — a publication language for parallel scientific computing. Concurrency: practice and experience, 6(5):461– 483, ???? 1994. CODEN CPEXEI. ISSN 1040-3108.

Hanson:2003:LST

[Han03] Floyd B. Hanson. Local supercomputing training in the computational sciences using remote national centers. Future Generation Computer Systems, 19(8): 1335–1347, November 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hanson:2020:CSS

[Han20] W. A. Hanson. The CORAL supercomputer systems. IBM Journal of Research and Development, 64(3/4):1:1– 1:10, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Harrison:1986:CLE

[Har86] W. Ludwell Harrison. Compiling LISP for evaluation on a tightly coupled multiprocessor. Technical Report CSRD-565, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. viii + 273 pp.

Harrison:1989:PPF

[Har89] B. K. Harrison. Performance of a process flowsheet system on a supercomputer. Computers & Chemical Engineering, 13(7):855–??, July 1, 1989. CODEN CCENDW. ISSN 0098-1354.

Harms:1990:SCI

[Har90] Uwe Harms, editor. Supercomputer and chemistry: IABG workshop, 1989. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 3-540-52915-2 (Berlin), 0-387-52915-2 (New York). LCCN QD39.3.E46 S926 1990.

Harms:1991:SCD

[Har91] Uwe Harms, editor. Supercomputer and chemistry 2: debis workshop 1990, Ottobrunn, November 19– 20, 1990. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-54411-9. LCCN QD39.3.E46 S925 1991. A debis workshop continues the serie [sic] of seminars organized by IABG.

Harrison:1994:CSP

[Har94a]  R. D. Harrison.  Com-

bat system prerequisites on supercomputer performance analysis. In Halang and Stoyenko [HS94b], pages 512– 513. ISBN 0-387-57558-8, 3-540-57558-8. ISSN 0258-1248. LCCN QA76.54.R4216 1994.

Hartwich:1994:RSF

[Har94b] P. Hartwich. Review of supercomputing in fluid flow. American Institute of Aeronautics and Astronautics Journal, 32(4):889–??, April 1994. CODEN AIAJAH. ISSN 0001-1452.

Harrison:1995:SEC

[Har95] Glennon J. Harrison. Supercomputer export controls. Washington, DC, USA, December 13, 1995. 6 pp. Major studies and issue briefs of the Congressional Research Service, supplement 95-1198 E.

Hastings:1984:UMC

[Has84] Chuck Hastings. Using a 16× 16 Cray multiplier as a 16bit microprocessor peripheral to perform 32-bit multiplication and division. Northcon — Conference Record, 1984. CODEN NCREDL.

Hasler:2017:SRC

[Has17] J. Hasler. Special report: Can we copy the brain? — A road map for the artificial brain. IEEE Spectrum, 54 (6):46–50, June 2017. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hawkinson:1986:HVA

[Haw86] Stuart Hawkinson. A homogeneous, vector architecture for scientific computing, 1986. 1 videocassette (50 min.).

Hawley:1988:BRS

[Haw88] John F. Hawley. Book review: Sidney Karin and Norris Parker Smith, The Supercomputer Era. Computers in Physics, 2(1):89– ??, January 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822667.

Hayes:1984:ASN

[Hay84]  Edward F. Hayes.  Ac-

cess to supercomputers: An NSF perspective (interview). Communications of the ACM, 27(4):299–303, 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hayes:1986:PED

[Hay86] Ann H. Hayes. Parallelprocessing experiences on the Denelcor HEP computer. The Journal of Systems and Software, 6(1–2):7–10, May 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Hayes:1989:IC

[Hay89] F. Hayes. Intel’s Cray-on-a-chip. BYTE Magazine, 14 (5):113–114, May 1989. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Haas:1989:MSD

[HB89] W. Haas and R. Brantner. Minisupercomputers and superworkstations in a “Distributed Supercomputing Environment”. In Brebbia and Peters [BP89b], pages 145– 156. ISBN 0-444-88110-7, 1-85312-044-8, 0-945824-27-0. LCCN TA345 .A66 1989 [1-2] (1989). Two volumes.

Habchi:1993:CCS

[HB93] C. Habchi and T. A. Baritaud. 93SC018 contribution of 3D CFD studies to the design cycles of new 2-Stroke DI engines. In Anonymous [Ano93-31], pages 103–114. ISBN 0-947719-62-8. LCCN ????

Hassen:1996:ITD

[HB96] S. B. Hassen and H. Bal. Integrating task and data parallelism using shared objects. In ACM [ACM96], pages 317–324. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Hendrickson:2008:GAH

[HB08] Bruce Hendrickson and Jonathan W. Berry. Graph analysis with high-performance computing. Computing in Science and Engineering, 10 (2):14–19, March/April 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Haring:2005:BGC

[HBB^{+}05] R. A. Haring, R. Bellofatto, A. A. Bright, P. G. Crumley, M. B. Dombrowa, S. M. Douskey, M. R. Ellavsky, B. Gopalsamy, D. Hoenicke, T. A. Liebsch, J. A. Marcella, and M. Ohmacht. Blue Gene/L compute chip: Control, test, and bring-up infrastructure. IBM Journal of Research and Development, 49(2/3):289–301, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/haring.pdf.

Hsu:1995:AEP

[HBCN95] Tai-Ran Hsu, Avram Bar-Cohen, and Wataru Nakayama, editors. Advances in electronic packaging, 1995: proceedings of the International Intersociety Electronic Packaging Conference, INTERpack ’95: presented at the International Intersociety Electronic Packaging Conference, March 26–30, 1995, Lahaina, Maui, Hawaii, volume 10 of ASME Publications EEP. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1995. ISBN 0-7918-1303-7. LCCN TK7870.15 .I578 1995 v.1-2. Two volumes.

Hawick:1993:PUM

[HBDS93] K. A. Hawick, R. S. Bell, A. Dickinson, and P. D. Surry. Parallelization of the unified model data assimilation scheme. In Hoffmann and Kauranne [HK93b], pages 188–203. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Hartenstein:1996:HPC

[HBKR96] Reiner W. Hartenstein, Jürgen Becker, Rainer Kress, and Helmut Reinig. Highperformance computing using a reconfigurable accelerator. Concurrency: practice and experience, 8(6): 429–443, July 1996. CO-DEN CPEXEI. ISSN 1040-3108. URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        23286.

Harrison:1991:DCP

[HC91] Williams Ludwell Harrison and Jyh-Herng Chow. Dynamic control of parallelism and granularity in executing nested parallel loops. Technical Report CSRD 1167, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1991. 8 pp.

Hide:1993:UVC

[HC93] W. Hide and L. Chan. Use of a visual comparative method to resolve conserved sequence motifs in proteins. In Lim et al. [L^{+}93], pages 505–512. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Haney:1999:SPH

[HC99] Scott Haney and James Crotinger. Scientific programming: How templates enable high-performance scientific computing in C++. Computing in Science and Engineering, 1(4):66–72, July/ August 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c4066.
pdf.

He:2018:PNU

[HCD^{+}18] Yun He, Brandon Cook, Jack Deslippe, Brian Friesen, Richard Gerber, Rebecca Hartman-Baker, Alice Koniges, Thorsten Kurth, Stephen Leak, Woo-Sun Yang, Zhengji Zhao, Eddie Baron, and Peter Hauschildt. Preparing NERSC users for Cori, a Cray XC40 system with Intel many integrated cores. Concurrency and Computation: Practice and Experience, 30 (1):??, January 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hsu:1995:DBS

[HCH95] F.-H. Hsu, M. S. Campbell, and A. J. Hoane. Deep Blue system overview. In ACM [ACM95a], pages 240– 244. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Han:1988:FTD

[HCL88] Tack-Don Han, David A. Carlson, and Steven P. Levitan. A fault tolerant design of the generalized cube network. Technical report SRC-TR-88-017, Supercomputing Research Center: IDA, Lanham, MD, USA, October 1988. 11 + [11] pp.

He:1994:SIP

[HCL94] Z. He, C. Cai, and T. Luo. Study and implement of parallel algorithms in image processing based on YH supercomputer. In IEEE [IEE94a], pages 757–759. ISBN 0-7803-1865-X, 0-7803-1866-8. LCCN TA1630.I58 1994. Two volumes. IEEE catalog number 94TH0638-7.

Ho:1995:MIS

[HCPS95] B. K. T. Ho, Z. Chen, R. K. Panwar, and R. M. Sadri. Medical image supercomputing in a PACS infrastructure [2435-71]. In Jost and Dwyer III [JD95], pages 536–546. ISBN 0-8194-1783-1. ISSN 0361-0748. LCCN R857.P52M426 1995.

Hipsz:1997:DAM

[HCV97] I. J. Hipsz, W. L. Cleghorn, and R. D. Venter. Determination of an alternative methodology for the initial design stage of automobile air vents. In Roller [Rol97], pages 359– 366. ISBN 0-947719-88-1 (paperback). LCCN ????

Houachi:1988:VHS

[HD88] M. F. Houachi and M. Degonde. Vehicle handling simulator. In Marino [Mar88b], pages 457–473. ISBN ???? LCCN TL240 .I528 1988.

Hwang:1989:PPS

[HD89] Kai Hwang and Doug DeGroot. Parallel processing for supercomputers and artificial intelligence. McGraw-Hill series in supercomputing and parallel processing. McGraw-Hill, New York, NY, USA, 1989. ISBN 0-07-031606-6. xix + 673 pp. LCCN QA76.5 .P314851 1989.

Horoi:1998:EIL

[HE98] M. Horoi and R. Enbody.

Efficient implementation of a Lanczos eigenvalue solver on a Cray T3E-900. Lecture Notes in Computer Science, 1401:907–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Healy:1991:PVA

[Hea91]  Brian E. Healy.  Paral-

lel and vector algorithms in nonlinear structural dynamics using the finite element method. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. v + 311 pp.

Hu:1996:CPC

[HEB96] Y. F. Hu, D. R. Emerson, and R. J. Blake. The communication performance of the Cray T3D and its effect on iterative solvers. Parallel Computing, 22(6):829– 844, September 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hooper:1993:ITS

[HED93] M. Hooper, I. Entwistle, and A. M. Dore. Interactive test scheduling returns control to the technician. In Anonymous [Ano93-31], pages 661– 668. ISBN 0-947719-62-8. LCCN ????

Hegland:1996:RCF

[Heg96] Markus Hegland. Real and complex fast Fourier transforms on the Fujitsu VPP 500. Parallel Computing, 22 (4):539–553, June 11, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
parco/cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=4&aid=1066.

Hehre:1986:MCR

[Heh86] Warren J. Hehre. Modeling chemical reactivity, 1986. 1 videocassette (50 min.).

Heinmets:1989:SAP

[Hei89] F. Heinmets. Supercomputer analysis of purine and pyrimidine metabolism leading to DNA synthesis. Cell biophysics, 14(3):283–??, June 1, 1989.

Heinzl:1990:DST

[Hei90] Carl G. (Carl George) Heinzl.

Diagnostics for the supercomputer toolkit MPCU module. Thesis (B.S.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1990. vi + 181 pp. Supervised by Harold Abelson.

HerrmannScheurer:1995:MCP

[HEJM95] A. K. Herrmann Scheurer, M. L. Egger, C. Joseph, and C. Morel. A Monte Carlo phantom simulator for positron emission tomography. In Herrmann et al. [HWP95], pages 205–209. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Helin:1992:PAC

[Hel92] J. Helin. Performance analysis of the CM-2, a massively parallel SIMD computer. In ACM [ACM92b], pages 45–52. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Helsel:1993:VBB

[Hel93] Sandra K. Helsel, editor. VR becomes a business: proceedings of Virtual Reality ’92: the 3rd annual conference and exhibition, San Jose, September 1992. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1993. ISBN 0-88736-854-9 (acid free paper). LCCN QA76.9. H85 V57 1992.

Helbing:1996:TMM

[Hel96] D. Helbing. Traffic modelling by means of physical concepts. In Wolf et al. [WSB96], pages 87–104. ISBN 981-02-2635-7. LCCN ????

Hemker:1984:MAR

[Hem84] P. W. Hemker. Multigrid algorithms run on supercomputers. CWI-Newsletter, 3: 25–30, 1984.

Henriquez:1991:SCE

[Hen91]  Adolfo Henriquez.  Sav-

ing computer and engineering time in reservoir simulation. Aalborg Universitetcenter og Hogskolesenteret i Rogaland, Rogaland, Norway, 1991. various pp.

Hensgen:1997:HCW

[Hen97] D. Hensgen, editor. Heterogeneous computing workshop: 6th — April 1997, Geneva, Switzerland, HET-EROGENEOUS COMPUT-ING WORKSHOP 1997; 6th. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7879-8, 0-8186-7881-X. LCCN ????

Herchuelz:1989:SSA

[Her89] P. Herchuelz. Supercomputers: Some aspects of their impact on communications. Computer Networks and ISDN Systems, 17(4–5): 328–331, October 10 1989. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Herbst:1990:MOM

[Her90a] K. Herbst. A more open market for supercomputers. Datamation, 36(18):123–125, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Herbst:1990:JMO

[Her90b] Kris Herbst. Japan: a more open market for supercomputers. Datamation, 36(18): 123–??, September 15, 1990. CODEN DTMNAT. ISSN 0011-6963.

Hernadi:1994:PNB

[Her94] Gyorgy Hernadi. Petri net based parallelization of the γ-CLF neural network on the KSR1 supercomputer. Thesis (M.S.), Dept. of Computer Science, University of Houston, Houston, TX, USA, 1994. x + 119 pp.

Herrmann:1995:FNN

[Her95] C. S. Herrmann. A fuzzy neural network for detecting graphoelements in EEGs. In Herrmann et al. [HWP95], pages 193–198. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Huber:1995:PHP

[HERC95] J. V. Huber, C. L. Elford, D. A. Reed, and A. A. Chien. PPFS: a high performance portable parallel file system. In ACM [ACM95a], pages 385–394. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Hess:1990:SIE

[Hes90] Karl Hess. Supercomputer images of electron devices physics. Physics Today, 43 (2):34–??, February 1, 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Hsiung:1993:PSO

[HES93] Ching-Kuo Hsiung and Mohamed E. M. El-Sayed. Parallel structural optimization on Cray X-MP. Advances in Engineering Software, 17 (3):135–139, 1993. CO-DEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic).

Hey:1990:STP

[Hey90] Anthony J. G. Hey. Supercomputing with transputers— past, present and future. ACM SIGARCH Computer Architecture News, 18(3b): 479–489, September 1990. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Hey:1994:GEP

[Hey94] T. Hey. The Genesis Esprit project — an overview. Parallel Computing, 20(10-11): 1605–1612, November 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Heydecker:1996:TRV

[Hey96] B. G. Heydecker. Treatment of random variability in traffic modelling. In Wolf et al. [WSB96], pages 119– 136. ISBN 981-02-2635-7. LCCN ????

Howell:1993:VBE

[HF93] L. J. Howell and L. L.

Frost. Vehicle body engineering — the past and future. In Anonymous [Ano93-31], pages 831–838. ISBN 0-947719-62-8. LCCN ????

Hester:1994:PPP

[HF94] P. D. Hester and W. J. Filip.

Preface: Power2 and PowerPC architecture and implementation. IBM Journal of Research and Development, 38(5):490–491, September 1994. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
almaden.ibm.com/journal/
rd38-5.html#one.

Henry:1998:PIT

[HFCM98] Greg Henry, Pat Fay, Ben Cole, and Timothy G. Mattson. The performance of the Intel TFLOPS supercomputer. Intel Technology Journal, (Q1):11, 1998. ISSN 1535-766X. URL

http://developer.intel.
com/technology/itj/q11998/
articles/art_2.htm;   http:
//developer.intel.com/technology/
itj/q11998/pdf/perf.pdf.

Hughes:1986:LVI

[HFH86] T. J. R. Hughes, R. M. Ferencz, and J. O. Hallquist. Large-scale vectorized implicit calculations in solid mechanics on a Cray X-MP/48 utilizing EBE preconditioned conjugate gradients. AMD (Symposia Series) (American Society of Mechanical Engineers, Applied Mechanics Division), 75:233– 277, 1986. CODEN AMD-VAS. ISSN 0160-8835.

Hughes:1987:LVI

[HFH87] Thomas J. R. Hughes, Robert M. Ferencz, and John O. Hallquist. Largescale vectorized implicit calculations in solid mechanics on a Cray X-MP/48 utilizing EBE preconditioned conjugate gradients. Computer Methods in Applied Mechanics and Engineering, 61(2): 215–248, March 1987. CO-DEN CMMECC. ISSN 0374-2830.

Herrmann:1996:DWD

[HFNP96] H. J. Herrmann, E. Flekkoey, K. Nagel, and G. Peng. Density waves in dry granular flow. In Wolf et al. [WSB96], pages 239–250. ISBN 981-02-2635-7. LCCN ????

Himeno:1994:CSU

[HFT94] R. Himeno, K. Fujitani, and M. Takagi. Chapter 6: Simulation of unsteady threedimensional flows around an automobile and its visualization. In Murthy and Brebbia [MB94b], pages 127–148. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Hazet:1988:SAC

[HG88] Bruno Hazet and Martine Guiraud. Some applications of crash simulation in car manufacturing industry. In International Conference on Supercomputing Applications in the Automotive Industry (2nd: 1988: Seville, Spain). Supercomputer applications in automotive research and engineering development, pages 21–33. Cray Research, Inc., Minneapolis, MN, USA, 1988.

Holland:2002:GEI

[HG02] Charles J. Holland and John Grosh. Guest Editors’ introduction: High-performance computing: Addressing defense needs of today and tomorrow. Computing in Science and Engineering, 4(2): 11–14, March/April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c2011abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2002/pdf/c2011.pdf.

Hemmings:1990:API

[HGB90] A. M. Hemmings, J. M. Goodfellow, and T. L. Blundell. The aspartic proteinases: Insights into inhibitor design from crystallography and computer simulation. In Pitcher [Pit90], pages 407–412. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Haussling:1994:CCI

[HGC94] H. J. Haussling, J. J. Gorski, and R. M. Coleman. Chapter 5: Computation of incompressible fluid flows about bodies with appendages. In Murthy and Brebbia [MB94b], pages 97–126. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Herbin:1988:PIM

[HGS88] Raphaèle Herbin, Stephane Gerbi, and Vijay Sonnad. Parallel implementation of a multigrid method on the experimental ICAP supercomputer. Applied Mathematics and Computation, 27(4 (part II)):281–312, September 1, 1988. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hung:1991:PCSa

[HGS91] G. G. Hung, Kyle A. Gallivan, and Resve A. Saleh. Parallel circuit simulation based on nonlinear relaxation methods. Technical Report CSRD 1113, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 5 pp.

Hutchinson:1993:SCP

[HH93] G. B. Hutchinson and M. R.

Hayden. SORFIND: a computer program that predicts exons in vertebrate genomic DNA. In Lim et al. [L^{+}93], pages 513–520. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Herpel:1993:FLA

[HHGS93] H. J. Herpel, S. K. Halgamuge, M. Glesner, and J. Stoecker. Fuzzy logic applied to control and data analysis problems in mechatronic systems. In Anonymous [Ano93-31], pages 225– 234. ISBN 0-947719-62-8. LCCN ????

Horiguchi:1994:ISP

[HHK94] S. Horiguchi, D. F. Hsu, and M. Kimura, editors. International Symposium on Parallel Architectures, Algorithms, and Networks (ISPAN): proceedings of the 1994, December 14–16, 1994, Kanazawa, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6507-6, 0-8186-6506-8. LCCN QA76.58.I5673 1994.

Huerta:2019:SHP

[HHK19] E. A. Huerta, Roland Haas, and Daniel S. Katz. Supporting high-performance and high-throughput computing for experimental science. Computing and Software for Big Science, 3(1): ??, December 2019. CO-DEN ???? ISSN 2510-2036 (print), 2510-2044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s41781-019-0022-7.

Hironaka:1991:SVP

[HHOM91] T. Hironaka, T. Hashimoto, K. Okazaki, and K. Murakami. A single-chip vectorprocessor prototype based on multithreaded streaming/ FIFO (MSFV) architecture. In Anonymous [Ano91q], pages 77–86. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Hironaka:1992:BVP

[HHOM92] T. Hironaka, T. Hashimoto, K. Okazaki, and K. Murakami. Benchmarking a vector-processor prototype based on multithreaded streaming/FIFO vector (MSFV) architecture. In ACM [ACM92b], pages 272–281. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Hargrove:2001:CDI

[HHS01a] William W. Hargrove, Forrest M. Hoffman, and Thomas Sterling. Computing: The do-it-yourself supercomputer. Scientific American, 285(2): 72–79, August 2001. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.
nature.com/scientificamerican/
journal/v285/n2/pdf/scientificamerican0801-
72.pdf.

Hargrove:2001:DIY

[HHS01b] William W. Hargrove, Forrest M. Hoffman, and Thomas Sterling. The do-it-yourself supercomputer: Scientists have found a cheaper way to solve tremendously difficult computational problems: connect ordinary PCs so that they can work together. Scientific American, 285(2): 72–79, August 2001. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.
sciam.com/2001/0801issue/
2001/0801issue/0801hargrove.
html.

Hagen:1993:VSD

[HHSW93] H. Hagen, S. Hahmann, P. Santerelli, and B. Wordenweber. Variational surface design and surface interrogation. In Anonymous [Ano93-31], pages 437–446. ISBN 0-947719-62-8. LCCN ????

Higuchi:1994:IPA

[HHT^{+}94] Tetsuya Higuchi, Kennichi Handa, Naoto Takahashi, Tatsumi Furuya, Hitoshi Iida, E. Sumita, O. Oi, and H. Kitano. The IXM2 parallel associative processor for AI. Computer, 27(11):53– 63, November 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hart:1990:CSC

[HHTD90] R. T. Hart, V. V. Hennebel, N. Thongpreda, and D. A. Dulitz. Computer simulation of cortical bone remodeling. In Pitcher [Pit90], pages 57–66. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Hibbard:2001:VFN

[Hib01]  Bill Hibbard.  Vis files:

NEC supercomputer visualization efforts. Computer Graphics, 35(2):10–13, May 2001. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Hicks:2018:IBC

[Hic18] Bryce Hicks. Improving I/O bandwidth with Cray DVS client-side caching. Concurrency and Computation: Practice and Experience, 30 (1):??, January 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anonymous:1992:HCR

[Hig92] High-performance computing review, 1992. ISSN 1068-0365. Publications and Communications, Inc, Austin, TX, USA.

Hillis:1991:MPS

[Hil91] W. Daniel Hillis. Massively parallel supercomputing the Connection Machine, 1991. 1 videocassette (53 min.).

Hillis:1992:MPS

[Hil92] W. Daniel Hillis. Massively parallel supercomputing: the Connection Machine, 1992. 1 videocassette (58 min.).

Hillman:1997:RSE

[Hil97] L. Hillman. Review: Scientist’s and engineer’s guide to workstations and supercomputers: coping with Unix, RISC, vectors, and programming, by Rubin H. Landau and Paul J. Fink, Jr. Scientific Programming, 6 (4):391–393, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Himeno:1993:CAA

[Him93] R. Himeno. 93SC045 CFD applications in automotive aerodynamic development at Nissan. In Anonymous [Ano93-31], pages 37–44. ISBN 0-947719-62-8. LCCN ????

Hinzmann:1993:FRP

[Hin93] Brock Hinzmann. The future of rapid prototyping: new technologies for design visualization and verification. Technical Report D93-1764, SRI International, Business Intelligence Program, 333 Ravenswood Avenue, Menlo Park, CA 94025-3493, USA, August 1993. 10 pp.

Hirschsohn:1992:PS

[Hir92a] Ian Hirschsohn. Personal supercomputing. Dr. Dobb’s Journal of Software Tools, 17 (6):16–18, 20, 22, 24, 26– 27, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Hirschsohn:1992:PSS

[Hir92b] Ian Hirschsohn. Personal supercomputing: Seamless portability. Dr. Dobb’s Journal of Software Tools, 17(7): 40, 42–44, 46, 48, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Hirschsohn:1992:PSV

[Hir92c] Ian Hirschsohn. Personal supercomputing: virtual memory, 64-bit. Dr. Dobb’s Journal of Software Tools, 17 (8):50, 52, 54, 56, 60, 62– 63, August 1992. CODEN DDJOEB. ISSN 1044-789X.

Hirsh:1994:ONS

[Hir94] R. S. Hirsh. An overview of the NSF supercomputer centers. In Thompson and Chimenti [TC94], pages 1–16. ISBN 0-306-44731-2. ISSN 0743-0760. LCCN ???? Two volumes.

Hanebutte:1994:SSP

[HJZ94] Ulf R. Hanebutte, Ronald D.

Joslin, and Mohammad Zubair. Scalability study of parallel spatial direct numerical simulation code on IBM SP1 parallel supercomputer. NASA contractor report NASA CR-194975; ICASE report no. 94-80, National Aeronautics and Space Administration, Langley Research Center, Hampton, VA, USA, 1994. ?? pp. Distributed to depository libraries in microfiche. Shipping list no.: 95-0419-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1995] 1 microfiche.

Herrmann:1993:WLS

[HK93a] H. J. Herrmann and F. Karsch, editors. Workshop on Large Scale Computational Physics on Massively Parallel Computers: HLRZ, KFA Julich, Germany, June 14–16, 1993, volume 4(6) of International Journal of Modern Physics C. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1993. ISBN 981-02-1643-2. ISSN 0129-1831 (print), 1793-6586 (electronic). LCCN QC20 .W67 1993.

Hoffmann:1993:PSA

[HK93b] Geerd-R. Hoffmann and Tuomo Kauranne, editors. Parallel supercomputing in atmospheric science: proceedings of the fifth ECMWF Workshop on the Use of Parallel Processors in Meteorology, Reading UK, November 23–27, 1992. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1993. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Hilliges:1996:DTF

[HK96] M. Hilliges and N. Koch. Dynamic traffic flow in freeway networks. In Wolf et al. [WSB96], pages 181– 186. ISBN 981-02-2635-7. LCCN ????

Huh:1997:SAI

[HK97] K. Huh and W. Kim. Supercomputing applications in internal combustion engine design and analysis. In IEEE [IEE97b], pages 178– 185. ISBN 0-8186-7901-8, 0-8186-7902-6, 0-8186-7903-4. LCCN ????

Hunding:1990:DSS

[HKG90] A. Hunding, S. A. Kauffman, and B. C. Goodwin. Drosophila segmentation: Supercomputer simulation of prepattern hierarchy. Journal of Theoretical Biology, 145(3):369–??, August 9, 1990. CODEN JTBIAP. ISSN 0022-5193.

Hiranandani:1994:CTB

[HKMCS94] S. Hiranandani, K. Kennedy, J. Mellor-Crummey, and A. Sethi. Compilation techniques for block-cyclic distributions. In Anonymous [Ano94-134], pages 392–403. ISBN ???? LCCN ????

Hossfeld:1989:MEA

[HKN89] F. Hossfeld, R. Knecht, and W. E. Nagel. Multitasking: experience with applications on a CRAY X-MP. Parallel Computing, 12(3):259– 283, December 1989. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Husmann:1988:ACF

[HKP88] Harlan Husmann, David J. Kuck, and David A. Padua. Automatic compound function definition for multiprocessors. Technical Report CSRD 784, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 34–41 pp.

Halin:1994:CFJ

[HKR94] Jurgen Halin, Walter J. Karplus, and R. Rimane, editors. CISS: First Joint Conference of International Simulation Societies proceedings: August 22–25, 1994, ETH Zurich, Zurich, Switzerland. Society for Computer Simulation, San Diego, CA, USA, 1994. ISBN 1-56555-031-5. LCCN ????

Hussaini:1993:ATC

[HKS93] M. Y. Hussaini, A. Kumar, and M. D. Salas, editors. Algorithmic trends in computational fluid dynamics: Workshop — September 15–17, 1991, Hampton, VA. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-94014-6, 3-540-94014-6. LCCN QA911 .A555 1993.

Hu:1994:OIC

[HKSY94] C. Hu, R. B. Kearfott, J. Sheldon, and Q. Yang. Optimizing INTBIS on the Cray Y-MP. Interval Computations = Interval’nye vychisleniia, 1994(4), 1994. ISSN 0135-4868.

Hiranandani:1992:ECO

[HKT92] S. Hiranandani, K. Kennedy, and C.-W. Tseng. Evaluation of compiler optimizations for Fortran D on MIMD distributed-memory machines. In ACM [ACM92b], pages 1–14. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Harms:1988:EBT

[HL88a] Uwe Harms and Hermann Luttermann. Experiences in benchmarking the three supercomputers CRAY-1M, CRAY-X/MP, FUJITSU VP-200 compared with the CYBER 76. Parallel Computing, 6(3):373–382, March 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hazet:1988:SAV

[HL88b] Bruno Hazet and Thierry Lambert. Simulations in acoustics and vibrations. In Marino [Mar88b], pages 475– 480. ISBN ???? LCCN TL240 .I528 1988.

Heinmets:1991:SDM

[HL91] F. Heinmets and R. H. Leary. A study of deoxyribonucleotide metabolism and its relation to DNA synthesis: Supercomputer simulation and Model-System analysis. Cell biophysics, 18(3): 263–??, June 1, 1991.

Hanebutte:1993:MPP

[HL93a] U. R. Hanebutte and E. E. Lewis. Massively parallel performance of neutron transport response matrix algorithms. In Kusters et al. [KSW93], pages 41–50. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Huss-Lederman:1993:MMI

[HL93b] Steven Huss-Lederman. Matrix multiplication on the Intel Touchstone Delta. Technical report SRC-TR-93-101, Supercomputing Research Center: IDA, Lanham, MD, USA, May 28, 1993. 18 pp.

Hamdi:1995:DLB

[HL95] M. Hamdi and C.-K. Lee. Dynamic load balancing of data parallel applications on a distributed network. In IEEE [IEE94c], pages 170– 179. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Halada:1996:PMS

[HL96] L. Halada and M. Lucka.

A partition method for solving block pentadiagonal linear system on Intel Hypercube iPSC/860. Lecture Notes in Computer Science, 1127:231–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heinmets:19xx:SDM

[HLxx] F. Heinmets and R. H.

Leary. A study of deoxyribonucleotide metabolism and its relation to DNA synthesis: Supercomputer simulation and model-system analysis. Cell biophysics, ???? (????):263–278, ???? 19xx. ISSN 0163-4992.

Harms:2018:TRI

[HLA^{+}18] Kevin Harms, Ti Leggett, Ben Allen, Susan Coghlan, Mark Fahey, Carissa Holohan, Gordon McPheeters, and Paul Rich. Theta: Rapid installation and acceptance of an XC40 KNL system. Concurrency and Computation: Practice and Experience, 30 (1):??, January 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Holm:1994:CSP

[HLB94] J. Holm, A. Lain, and P. Banerjee. Compilation of scientific programs into multithreaded and message driven computation. In IEEE [IEE94c], pages 518– 525. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Hammond:1995:IPI

[HLDS95] S. W. Hammond, R. D. Loft, J. M. Dennis, and R. K. Sato. Implementation and performance issues of a massively parallel atmospheric model. Parallel Computing, 21(10):1593– 1619, October 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Huss-Lederman:1993:CSP

[HLJT93] Steven Huss-Lederman, Elaine M. Jacobson, and Anna Tsao. Comparison of scalable parallel matrix multiplication libraries. Technical report SRC-TR-93-108, Supercomputing Research Center: IDA, Lanham, MD, USA, November 1993. 8 pp.

Hameetman:1997:EBR

[HLPP97] G. J. Hameetman, W. Loeve, G. Poppinga, and K. Potma. Early benchmark results on the NEC SX-4 supercomputer. In Schiano [Sch97a], pages 501–508. ISBN 0-444-82327-1. LCCN ???? Huss-Lederman:1993:PII

[HLTZ93] Steven Huss-Lederman, Anna Tsao, and Guodong Zhang. A parallel implementation of the invariant subspace decomposition algorithm for dense symmetric matrices. Technical report SRC-TR-93-091, Supercomputing Research Center: IDA, Lanham, MD, USA, February 1993. 9 pp.

Hady:1993:PCB

[HM93a] Frank Hady and B. L.

Menezes. The performance of crossbar-based binary hypercubes. Technical report SRC-TR-93-098, Supercomputing Research Center: IDA, Lanham, MD, USA, April 21, 1993. 21 pp.

Hebert:1993:DTG

[HM93b] A. Hebert and G. Mathonniere. Development of third generation SPH method for the homogenization of a PWR assembly. In Kusters et al. [KSW93], pages 558– 570. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Herbst:1993:CFL

[HM93c] A. Herbst and B. Malle.

Core functions of a long-term archiving system for product data. In Anonymous [Ano93-31], pages 381–388. ISBN 0-947719-62-8. LCCN ????

Huang:1997:CCA

[HM97] C. Huang and A. S. Mayer. Computational challenges associated with mathematical optimization of soil and groundwater remediation systems. In Delic and Wheeler [DW97], pages 287– 291. ISBN 0-89871-378-1. LCCN ????

Henshell:1993:CSA

[HMBS93] R. D. Henshell, P. C. Macey, I. E. Boston, and M. Surridge. Computational solutions for automotive analysis. In Anonymous [Ano93-31], pages 373–380. ISBN 0-947719-62-8. LCCN ????

Hollingsworth:1994:DPI

[HMC94] J. K. Hollingsworth, B. P. Miller, and J. Cargille. Dynamic program instrumentation for scalable performance tools. In IEEE [IEE94c], pages 841–850. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Hori:1997:PTP

[HMKI97] E. Hori, S. Mine, Y. Kuramitsu, and J. Inasaka. Packaging technology for the parallel supercomputer SX-4. In Suhir [Suh97], pages 73–80. ISBN 0-7918-1559-5. LCCN ????

Humble:2021:QCH

[HML^{+}21] Travis S. Humble, Alexander McCaskey, Dmitry I. Lyakh, Meenambika Gowrishankar, Albert Frisch, and Thomas Monz. Quantum computers for high-performance computing. IEEE Micro, 41(5):15–23, September/October 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hackstadt:1994:SPV

[HMM94] S. T. Hackstadt, A. D. Malony, and B. Mohr. Scalable performance visualization for data-parallel programs. In IEEE [IEE94c], pages 342– 349. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Hirata:1991:MPA

[HMNN91] H. Hirata, Y. Mochizuki, A. Nishimura, and Y. Nakase. A multithreaded processor architecture with simultaneous instruction issuing. In Anonymous [Ano91q], pages 87–96. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Hayes:1986:MBH

[HMS^{+}86a] John P. Hayes, Trevor Mudge, Quentin F. Stout, Stephen Colley, and John Palmer. A microprocessorbased hypercube supercomputer. IEEE Micro, 6(5):6– 17, September/October 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hayes:1986:MHS

[HMS^{+}86b] John P. Hayes, Trevor Mudge, Quentin F. Stout, Stephen Colley, and John Palmer. Microprocessorbased hypercube supercomputer. IEEE Micro, 6(5):6– 17, September/October 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hebeker:1993:NSK

[HMS93] F. K. Hebeker, R. Maly, and S. U. Schoeffel. 93SC028 numerical simulation of kock damage phenomena on an IBM RS/ 6000 workstation cluster. In Anonymous [Ano93-31], pages 275–282. ISBN 0-947719-62-8. LCCN ????

Homewood:1987:ITT

[HMSS87] Mark Homewood, David May, David Shepherd, and Roger Shepherd. The IMS T800 transputer. IEEE Micro, 7(5):10–26, September/October 1987. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hyatt:1990:CSD

[HN90] Robert M. Hyatt and Harry L. Nelson. Chess and supercomputers: Details about optimizing Cray Blitz. IEEE, Piscataway, NJ, USA, 1990. ISBN 0-8186-2056-0. 354–363 pp. LCCN ???? IEEE catalog number 90CH2916-5.

Henderson:1994:SHI

[HNS94] M. Henderson, B. Nickless, and R. Stevens. A scalable high-performance I/O system. In IEEE [IEE94c], pages 79–86. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Huang:2010:RCA

[HNS^{+}10] Miaoqing Huang, Vikram K.

Narayana, Harald Simmler, Olivier Serres, and Tarek El-Ghazawi. Reconfiguration and communicationaware task scheduling for high-performance reconfigurable computing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4):20:1– 20:??, November 2010. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Hiwa:1993:DAP

[HNST93] N. Hiwa, Y. Nobutoki, H. Sakamoto, and T. Terada. The development of the advanced protocol for automotive local area multiplexing network (advanced PALMNET). In Anonymous [Ano93-31], pages 889–898. ISBN 0-947719-62-8. LCCN ????

Ho:1988:ANA

[Ho88] Shou Sin Ho. acceSX network access system for Honeywell NEC SX-2 supercomputer. Thesis (M.S.), Dept. of Computer Science, University of Houston, Houston, TX, USA, 1988. viii + 68 pp.

Ho:1991:PII

[Ho91] Chung-Jang Ho. Parallel implementation of iterative methods on the Cray X-MP supercomputer. Where was this work produced??., 1991.

Ho:1992:MEI

[Ho92a] Samuel Ho. MaxPar extensions for isolating performance problems. Technical Report CSRD 1240, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1992. 19 pp.

Huang:1992:PPM

[HO92b] C.-M. Huang and D. P. O’Leary. Preconditioning parallel multisplittings for solving linear systems of equations. In ACM [ACM92b], pages 478–484. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Hockney:1985:MCC

[Hoc85] Roger W. Hockney. (r_{∞}, n_{1/2}, s_{1/2}) measurements on the 2-CPU CRAY X-MP. Parallel Computing, 2(1):1–14, March 1985. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hockney:1991:PPB

[Hoc91] R. Hockney. Performance parameters and benchmarking of supercomputers. Parallel Computing, 17(10–11):1111– 1130, December 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hockney:1994:CCM

[Hoc94] R. W. Hockney. The communication challenge for MPP: Intel Paragon and Meiko CS-2. Parallel Computing, 20(3): 389–398, March 1994. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hockney:1996:SCB

[Hoc96] Roger W. Hockney. The science of computer benchmarking. Software, environments, tools. SIAM Press, Philadelphia, PA, USA, 1996. ISBN 0-89871-363-3 (paperback). xiv + 129 pp. LCCN QA76.9.E94 H63 1996.

Hodek:1987:UIS

[Hod87] J. A. Hodek. The Unisys 1100/90 ISP system integrated supercomputing. In ICS 87. Second International Conference on Supercomputing. Proceedings. Supercomputing ’87, volume 1, page ?? ????, ????, ???? 1987. Three volumes.

Hoffmann:1993:PVR

[Hof93] G.-R. Hoffmann. PARMACS for viper: Report of an implementation. In Hoffmann and Kauranne [HK93b], pages 204–211. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Hoffmann:1994:HPC

[Hof94]  G.-R. Hoffmann.  High-

performance computing and networking for numerical weather prediction. Lecture Notes in Computer Science, 797:1–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haring:2012:IBG

[HOF^{+}12] Ruud A. Haring, Martin Ohmacht, Thomas W. Fox, Michael K. Gschwind, David L. Satterfield, Krishnan Sugavanam, Paul W. Coteus, Philip Heidelberger, Matthias A. Blumrich, Robert W. Wisniewski, Alan Gara, George Liang-Tai Chiu, Peter A. Boyle, Norman H. Chist, and Changhoan Kim. The IBM Blue Gene/Q compute chip. IEEE Micro, 32 (2):48–60, March/April 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hoganson:2002:HPC

[Hog02] Kenneth E. Hoganson. Highperformance computer architecture and algorithm simulator. ACM Journal on Educational Resources in Computing (JERIC), 2(1):131–148, March 2002. CODEN ???? ISSN 1531-4278.

Holcomb:1984:USI

[Hol84] Lee B. Holcomb. The U.S.

supercomputer industry: a strategy for the future. Thesis (M.S.), M.I.T., Sloan School of Management, Cambridge, MA, USA, 1984. 201 pp. Supervised by Mel Horwitch.

Holmes:1990:SC

[Hol90a] Bill Holmes. SPLASH cornerturn. Technical report SRC-TR-91-029, Supercomputing Research Center: IDA, Lanham, MD, USA, March 22, 1990. various pp.

Holowko:1990:PRA

[Hol90b] Paul Holowko. Parallel routing algorithm implementation on multiprocessing Cray Y-MP supercomputer. Thesis (M.S.), University of Cincinnati, Cincinnati, OH, USA, 1990. various pp.

Holmes:1993:SST

[Hol93]  Lewis M. Holmes.  Su-

percomputing specialists try out clusters and networks. Computers in Physics, 7 (1):6, January 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4823148.

Holmes:1994:SCE

[Hol94] Lewis M. Holmes. Supercomputing conference emphasizes scalable computing and networks. Computers in Physics, 8(1):8– ??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

                                        //aip.scitation.org/doi/
                                        10.1063/1.4823268.

Holmes:1995:SSH

[Hol95] Lewis M. Holmes. SC ’94 spotlights high-performance computing. Computers in Physics, 9(1):11–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

                                        https://aip.scitation.org/
                                        doi/10.1063/1.4823354.

Hoppe:1993:RPM

[Hop93] H.-C. Hoppe. The RAPS programming model O. In Hoffmann and Kauranne [HK93b], pages 212–239. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Hord:1982:IIFa

[Hor82a] R. Michael Hord. The Illiac IV, the first supercomputer. Computer Science Press, Inc., Rockville, MD, USA, 1982. ISBN 0-914894-71-4. xii + 350 pp. LCCN QA76.8.I5 .H67 1982.

Hord:1982:IIFb

[Hor82b] R. Michael Hord. The Illiac IV: the first supercomputer. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 3-540-11765-2. xii + 350 pp. LCCN ????

Hord:1990:PSS

[Hor90] R. Michael Hord. Parallel supercomputing in SIMD architectures. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1990. ISBN 0-8493-4271-6. 378 pp. LCCN QA76.5 .H675 1990.

Hord:1993:PSM

[Hor93] R. Michael Hord. Parallel supercomputing in MIMD architectures. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1993. ISBN 0-8493-4417-4. xviii + 401 pp. LCCN QA76.58.H67 1993.

Hori:1997:SSMb

[Hor97a] K. Hori. Supercomputer SX-4 multinode system. NEC Technical Journal = NEC giho, 50(11):82–??, ???? 1997. CODEN NECGEZ. ISSN 0285-4139.

Hori:1997:SSMa

[Hor97b] Kenichi Hori. Supercomputer SX-4 multinode system. Nippon Electric Company research and development, 38 (4):461–??, ???? 1997. CO-DEN NECRAU. ISSN 0048-0436.

Hord:1998:UPS

[Hor98] R. Michael Hord. Understanding parallel supercomputing. IEEE Press understanding science and technology series. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-7803-1120-5. ???? pp. LCCN QA76.58.H68 1998.

Hossfeld:1988:VS

[Hos88] F. Hossfeld. Vector-supercomputers. Parallel Computing, 7(3): 373–385, September 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hossfeld:1995:SBR

[Hos95] F. Hossfeld. On the supercomputational background of the Research Centre Juelich. In Herrmann et al. [HWP95], pages 435–448. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Hansen:1997:SSL

[HOSZ97] P. C. Hansen, Tz. Ostromsky, A. Sameh, and Z. Zlatev. Solving sparse linear least-squares problems on some supercomputers by using large dense blocks. BIT Numerical Mathematics, 37(3):535–558, September 1997. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.
mai.liu.se/BIT/contents/

bit37.html. Direct methods, linear algebra in optimization, iterative methods (Toulouse, 1995/1996).

Harrison:1988:PPA

[HP88a] Luddy Harrison and David A.

Padua. PARCEL: project for the automatic restructuring and concurrent evaluation of Lisp. Technical Report CSRD 653, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 527–538 pp.

Hutchings:1988:CTV

[HP88b] Barbara J. Hutchings and William Pien. Computation of three-dimensional vehicle aerodynamics using fluent/ BFC. In Marino [Mar88b], pages 233–255. ISBN ???? LCCN TL240 .I528 1988.

Haghighat:1991:SDA

[HP91] Mohammad Reza Haghighat and C. D. (Constantine D.) Polychronopoulos. Symbolic dependence analysis for high-performance parallelizing compilers. Technical Report CSRD 1197, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1991. 21 pp.

Haghighat:1992:SPA

[HP92] Mohammad Reza Haghighat and C. D. (Constantine D.) Polychronopoulos. Symbolic program analysis and optimization for parallelizing compilers. Technical Report CSRD 1237, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1992. 30 pp.

Hiller:1993:OAS

[HP93] M. Hiller and V. Pichler. An object-oriented approach for spatial vehicle dynamics simulation. In Anonymous [Ano93-31], pages 581–588. ISBN 0-947719-62-8. LCCN ????

Hu:1995:PMC

[HP95] Hong Hu and Jada M. Paysour. Panel method computational performance on CM-5 and Cray-YMP. Boundary Elements Communications, 6(2):51–54, March 1995. CODEN BECOFU. ISSN 1353-825X.

Hennessy:2003:CAQ

[HP03] John L. Hennessy and David A. Patterson. Computer Architecture—a Quantitative Approach. Morgan Kaufmann Publishers, San Francisco, CA, USA, third edition, 2003. ISBN 1-55860-596-7. xxi + 883 + A-87 + B-42 + C-1 + D-1 + E-1 + F-1 + G-1 + H-1 + I-1 + R-22 + I-44 pp. LCCN QA76.9.A73 P377 2003. US$89.95.

URL http://www.mkp.com/
books_catalog/catalog.asp?
ISBN=1-55860-596-7;   http:
//www.mkp.com/CA3. Ap-

pendix G, Vector Processors, is available electronically at the publisher’s Web site.

Holland:2004:GEI

[HP04] Charles J. Holland and Robert E. Peterkin, Jr. Guest Editors’ introduction: High-performance computing. Computing in Science and Engineering, 6(6):8–10, November/December 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

/csdl.computer.org/comp/
mags/cs/2004/06/c6008.pdf;
 http://csdl.computer.org/
dl/mags/cs/2004/06/c6008.
htm.

Hadri:2020:PSS

[HPH^{+}20] Bilel Hadri, Matteo Parsani, Maxwell Hutchinson, Alexander Heinecke, Lisandro Dalcin, and David Keyes. Performance study of sustained petascale direct numerical simulation on Cray XC40 systems. Concurrency and Computation: Practice and Experience, 32(20):e5725:1– e5725:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hasenfeld:1993:NAG

[HPLC93] A. Hasenfeld, E. Pepke, H. A. Lim, and C. R. Cantor. A new agarose gel model. In Lim et al. [L^{+}93], pages 501– 504. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Huedo:2001:IPM

[HPLT01] Eduardo Huedo, Manuel Prieto, Ignacio M. Llorente, and Francisco Tirado. Impact of PE mapping on Cray T3E message-passing performance. Lecture Notes in Computer Science, 1900:199– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1900/19000199.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1900/19000199.
                                        pdf.

Houstis:1988:SIC

[HPP88] E. N. Houstis, T. S. Papatheodorou, and C. D. Polychronopoulos, editors. Supercomputing: 1st International Conference, Athens, Greece, June 8–12, 1987: proceedings, volume 297 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. CODEN LNCSD9.

ISBN 0-387-18991-2, 3-540-18991-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.297. The conference was organized and sponsored by the Computer Technology Institute (C.T.I.) of Greece.

Hariri:1994:CSH

[HPPF94] S. Hariri, J.-B. Park, M. Parashar, and G. C. Fox. Communication system for highperformance distributed computing. Concurrency: practice and experience, 6(4):251– 270, June 1994. CODEN CPEXEI. ISSN 1040-3108.

Hock:1988:IMU

[HPS88] H. G. Hock, A. Poth, and W. Schrepfer. Influence of modeling undercarriage and power train components on the quality of numerical crash simulation. In Marino [Mar88b], pages 35–45. ISBN ???? LCCN TL240 .I528 1988.

Hatcher:1991:DPP

[HQ91] Philip J. Hatcher and Michael J.

Quinn. Data-Parallel Programming on MIMD Computers. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-08205-5. xiv + 231 pp. LCCN QA76.5.H42 1991. US$33.00. URL http://www.

mitpress.com/book-home.
tcl?isbn=0262082055.

Heath:1994:PFP

[HR94] M. T. Heath and P. Raghavan. Performance of a fully parallel sparse solver. In IEEE [IEE94c], pages 334– 341. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Huang:2004:HPP

[HR04] Chun-Hsi Huang and Sanguthevar Rajasekaran. Highperformance parallel biocomputing. Parallel Computing, 30(9–10):999–1000, September/October 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Hacker:2009:ACF

[HRC09] Thomas J. Hacker, Fabian Romero, and Christopher D. Carothers. An analysis of clustered failures on large supercomputing systems. Journal of Parallel and Distributed Computing, 69(7): 652–665, July 2009. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Halgamuge:1993:ATD

[HRG93] S. K. Halgamuge, T. A. Runkler, and M. Glesner. Autonomous truck driving based on on-line fuzzy control. In Anonymous [Ano93-31], pages 281–286. ISBN 0-947719-62-8. LCCN ????

Haji-Sheikh:1991:IMS

[HS^{+}91] A. Haji-Sheikh et al., editors. Integral methods in science and engineering-90: 2nd International conference on global methods — May 15– 18, 1990, University of Texas at Arlington, Arlington, TX, USA. Hemisphere Publishing Corporation, New York, NY, USA, 1991. ISBN 1-56032-066-4. LCCN TA330 .I58 1991.

Hady:1993:AVN

[HS93a] Frank Hady and David L.

Smitley. Adaptive vs. nonadaptive routing: an application driven case study. Technical report SRC-TR-93-099, Supercomputing Research Center: IDA, Lanham, MD, USA, March 30, 1993. 17 pp.

Horst:1993:STL

[HS93b] H. D. Horst and M. Saupe.

Simulation of a trailing load performed on a pulsation testing stand in comparison with FEM-Analysis. In Anonymous [Ano93-31], pages 573–580. ISBN 0-947719-62-8. LCCN ????

Hsu:1993:PCD

[HS93c] W.-C. Hsu and J. E.

Smith. Performance of cached DRAM organizations in vector supercomputers. ACM SIGARCH Computer Architecture News, 21(2): 327–336, May 1993. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Hafner:1994:EGB

[HS94a] H. Hafner and W. Schonauer. Expanding the gap between theoretical peak performance for supercomputer architectures. Scientific Programming, 3(2):157–??, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Halang:1994:RTC

[HS94b] Wolfgang A. Halang and Alexander D. Stoyenko, editors. Real time computing: Proceedings of the NATO Advanced Study Institute on Real Time Computing, held in Sint Maarten, Dutch Antilles, October 5–17, 1992, volume 127 of NATO ASI Series F Computer and Systems Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-57558-8, 3-540-57558-8. ISSN 0258-1248. LCCN QA76.54.R4216 1994.

Harrod:1994:NAC

[HS94c] W. J. Harrod and M. Sidani. Numerical algorithms for the Cray T3D. Lecture Notes in Computer Science, 879:304–311, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hixon:1994:UCF

[HS94d] D. Hixon and L. N. Sankar.

Unsteady compressible 2-D flow calculations on a MIMD parallel supercomputer. In AIAA [AIA94], pages 757– ?? ISBN ???? LCCN ???? Twelve volumes.

Hamdi:1995:EEH

[HS95a] M. Hamdi and S. W. Song.

Efficient embeddings into the hypercube using matrix transformations. In ACM [ACM95a], pages 280–288. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Hertzberger:1995:HCN

[HS95b] Bob Hertzberger and Giuseppe Serazzi, editors. Highperformance computing and networking: International Conference and Exhibition, Milan, Italy, May 3–5, 1995: proceedings, volume 919 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CO-DEN LNCSD9. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Hertzberger:1995:HPC

[HS95c] Bob Hertzberger and Giuseppe Serazzi, editors. Highperformance computing and networking: International Conference and Exhibition, Milan, Italy, May 3–5, 1995: proceedings, volume 919 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CO-DEN LNCSD9. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Helland:1996:ATL

[HS96] B. Helland and B. G. Summers. Assessment techniques for a learning-centered curriculum: Evaluation design for adventures in supercomputing. In Iskander [Isk96], pages 301–305. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN ????

Helland:19xx:ATL

[HSxx] B. Helland and B. G. Summers. Assessment techniques for a learning-centered curriculum: Evaluation design for adventures in supercomputing. Frontiers in Education Conference, ????(????): 301–??, ???? 19xx. CODEN PFECDR. ISSN 0190-5848.

Hsiao:1991:PSM

[Hsi91] David K. Hsiao. A parallel, scalable, microprocessorbased database computer for performance gains and capacity growth. IEEE Micro, 11(6):44–60, November/December 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hu:1995:OIC

[HSKY95] Chenyi Hu, Joe Sheldon, R. Baker Kearfott, and Qing Yang. Optimizing INTBIS on the Cray Y-MP. Reliable Computing = Nadezhnye vychisleniia, 1(3):265– 274, September 1995. CO-DEN RCOMF8. ISSN 1385-3139 (print), 1573-1340 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1385-3139&volume=1&
issue=3&spage=265.

Hsu:2015:WWW

[Hsu15] J. Hsu. When will we have an exascale supercomputer? [news]. IEEE Spectrum, 52 (1):13–16, January 2015. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hsu:2016:TPE

[Hsu16]  J. Hsu.  Three paths

to exascale supercomputing. IEEE Spectrum, 53(1):14– 15, January 2016. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hausheer:1990:SGR

[HSW^{+}90] F. H. Hausheer, U. Chandra Singh, A. L. Weis, J. Flory, J. Saxe, and K. Tufto. Supercomputer guided rational drug design in oncology and AIDS. In Pitcher [Pit90], pages 67–79. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Hintz:1972:CDS

[HT72] R. G. Hintz and D. P. Tate. Control Data STAR-100 processor design. In IEEE, editor, Digest of papers, 1972: innovative architecture. Continuing Compcon theme: innovation and change in computer design., pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1972. LCCN QA75.5 .C58 1972; TK7885.A1 C53 1972.

Hillis:1993:CCM

[HT93] W. Daniel Hillis and Lewis W. Tucker. The CM-5 Connection Machine: a scalable supercomputer. Communications of the ACM, 36 (11):31–40, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
163361.html.

Harris:1994:SDM

[HT94] T. J. Harris and N. P. Topham. The scalability of decoupled multiprocessors. In IEEE [IEE94c], pages 17–22. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Tang:1988:ECC

[hTD88] Ju ho Tang and Edward S.

Davidson. An evaluation of Cray-1 and Cray X-MP performance on vectorizable Livermore Fortran loops. Technical Report CSRD 785, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1988. 9 pp.

Hamano:1993:APM

[HTI93] K. Hamano, H. Tominaga, and H. Isizu. 93SC003 application of a plastic model test and CAE to weight reduction of vehicle body structure. In Anonymous [Ano93-31], pages 191–198. ISBN 0-947719-62-8. LCCN ????

Hart:1988:SUO

[HTV88] Richard T. Hart, Nisra Thongpreda, and William C. VanBuskirk. Supercomputer use in orthopaedic biomechanics research: Focus on functional adaptation of bone. IEEE engineering in medicine and biology magazine, 7(4):39–??, December 1, 1988. CODEN IEMBDE. ISSN 0739-5175.

Hall:1993:AVS

[HU93] B. B. Hall and T. R. Underhill. Active vehicle suspensions — some aspects of digital controller design. In Anonymous [Ano93-31], pages 147–154. ISBN 0-947719-62-8. LCCN ????

Huang:1992:TFG

[Hua92] Lihong Huang. Thin film growth: molecular dynamics simulation by supercomputer. Thesis (M.S. in mechanical engineering), University of Illinois at Chicago, Chicago, IL, USA, 1992. viii + 59 pp.

Hughes:1993:UAT

[Hug93] D. R. Hughes. The use of AI techniques and animated graphical simulation to control manufacturing cells. In Anonymous [Ano93-31], pages 567–572. ISBN 0-947719-62-8. LCCN ????

Hughes:1994:LJC

[Hug94] David Hughes. Labs join Cray team in software project. Aviation week & space technology, 140(25):55– 57, June 1994. CODEN AW-STAV. ISSN 0005-2175.

Hung:1990:PCS

[Hun90] G. G. Hung. A parallel circuit simulation using hierarchical relaxation. Technical Report CSRD 1014, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1990. 6 pp.

Hung:1991:PCSb

[Hun91] G. G. Hung. Parallel circuit simulation techniques using nonlinear relaxation. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1991. viii + 40 pp.

Hunding:1992:PFR

[Hun92] A. Hunding. Pattern formation of reaction-diffusion systems in 3 space coordinates. supercomputer simulation of Drosophila morphogenesis. Physica A, 188(1/3): 172–??, September 1992. CO-DEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic).

Hunding:1993:SST

[Hun93] A. Hunding. Supercomputer simulation of Turing structures in Drosophila morphogenesis. In Othmer et al. [OMM93], pages 149–160. ISBN 0-306-44661-8. ISSN 0258-1213. LCCN QH491 .N38 1993.

Huntsberger:1994:DAT

[Hun94] T. L. Huntsberger. Distributed algorithm for twodimensional global climate modeling. In IEEE [IEE94c], pages 392–396. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Huskamp:1986:MOS

[Hus86a] Jeffrey C. Huskamp. Modular operating system for the Cray-1. Software—Practice and Experience, 16(12): 1059–1076, December 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Husmann:1986:CMM

[Hus86b] Harlan Husmann. Compiler memory management and compound function definition for multiprocessors. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. vii + 192 pp.

Henry:1994:PUE

[HV94] G. Henry and R. Van de Geijn. A parallel unsymmetric eigensolver. In IEEE [IEE94c], pages 28–31. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Hill:1995:GSA

[HV95] S. L. Hill and A. E. P. Villa. Global spatio-temporal activity influenced by local kinetics in a simulated “Cortical” neural network. In Herrmann et al. [HWP95], pages 371– 378. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Hilka:1993:PDN

[HVSB93] M. Hilka, D. Veynante, A. Stoessel, and M. Baum. Parallel direct numerical simulation of turbulent premixed flames. In Anonymous [Ano93-31], pages 743–750. ISBN 0-947719-62-8. LCCN ????

Heroux:1991:PPC

[HVY91] Michael A. Heroux, Phuong Vu, and Chao Yang. A parallel preconditioned conjugate gradient package for solving sparse linear systems on a Cray Y-MP. Applied Numerical Mathematics: Transactions of IMACS, 8(2):93–115, 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Haeuser:1994:CSH

[HVZ94] J. Haeuser, A. Vinckier, and S. Zemsch. Chapter 12: Supercomputing in hypersonic flow with non-equilibrium effects. In Murthy and Brebbia [MB94b], pages 273–308. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88

Hyder:1996:SHN

[HW96] R. S. Hyder and D. A. Wood. Synchronization hardware for networks of workstations: Performance vs. cost. In ACM [ACM96], pages 245–397. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Hamrick:1997:CDO

[HW97] J. M. Hamrick and T. S. Wu. Computational design and optimization of the EFDC/HEM3D surface water hydrodynamic and eutrophication models. In Delic and Wheeler [DW97], pages 143–156. ISBN 0-89871-378-1. LCCN ????

Hager:2011:IHP

[HW11] Georg Hager and Gerhard Wellein. Introduction to high performance computing for scientists and engineers, volume 7 of Chapman and Hall/CRC computational science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-1192-X. xxv + 330 pp. LCCN QA76.88 .H34 2011.

Hwang:1984:TSD

[Hwa84] Kai Hwang, editor. Tutorial—Supercomputers: Design and Applications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. ISBN 0-8186-0581-2. viii + 640 pp. LCCN TK 7888.3 H82 1984.

Hwang:1985:MSS

[Hwa85]  K. Hwang.  Multiproces-

sor supercomputers for scientific… engineering applications. Computer, 18(6): 57–73, 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hill:1998:IGP

[HWG98] Jim Hill, Michael Warren, and Pat Goda. I’m not going to pay a lot for this supercomputer! Linux Journal, 45:??, January 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/2392.html.

Herrmann:1995:WSB

[HWP95] Hans J. Herrmann, D. E.

Wolf, and Ernst Poppel, editors. Workshop on Supercomputing in Brain Research: from tomography to neural networks, HLRZ, KFA Julich, Germany, November 21–23, 1994. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1995. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Hatton:1988:SKS

[HWS^{+}88] Les Hatton, Andy Wright, Stuart Smith, Gregg Parkes, Paddy Bennett, and Robert Laws. The Seismic Kernel System—a large-scale exercise in Fortran 77 portability. Software—Practice and Experience, 18(4):301– 329, April 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Describes portability issues in a 500,000+ line system.

Hsu:1989:SSN

[HY89] William Tsun-Yuk Hsu and Pen-Chung Yew. A simple synchonization network for large multiprocessor systems. Technical Report CSRD 841, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1989. 24 + [6] pp.

Hsu:1991:PHS

[HY91] William Tsun-Yuk Hsu and Pen-Chung Yew. The performance of hierarchical systems with wiring constraints. Technical Report CSRD 1061, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1991. 16 + [13] pp.

Hsu:1992:PEW

[HY92] William Tsun-Yuk Hsu and Pen-Chung Yew. Performance evaluation of wirelimited hierarchical networks. Technical Report CSRD 1179, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1992. 35 + [17] pp.

Hu:2020:MSS

[HYL^{+}20] Y. Hu, H. Yang, Z. Luan, L. Gan, G. Yang, and D. Qian. Massively scaling seismic processing on Sunway TaihuLight supercomputer. IEEE Transactions on Parallel and Distributed Systems, 31(5):1194–1208, May 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ido:1992:PSU

[IAIK92] S. Ido, K. Aoki, M. Ishine, and M. Kubota. Personal supercomputing by using transputer and Intel 80860 in plasma engineering. Computational mechanics, 9(5):305– 314, ???? 1992. CODEN CM-MEEE. ISSN 0178-7675.

Ishii:1992:OHS

[IAKH92] K. Ishii, H. Abe, S. Kawabe, and M. Hirai. Overview of the Hitachi S-3800 series supercomputer. In Meuer [Meu92c], pages 65–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Ijaz:2020:RHP

[IBBA20] Qaiser Ijaz, El-Bay Bourennane, Ali Kashif Bashir, and Hira Asghar. Revisiting the high-performance reconfigurable computing for future datacenters. Future Internet, 12(4):64, April 06, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/4/
                                        64.

Isaila:2011:DEM

[IBC^{+}11] Florin Isaila, Javier Garcia Blas, Jesus Carretero, Robert Latham, and Robert Ross. Design and evaluation of multiple-level data staging for Blue Gene systems. IEEE Transactions on Parallel and Distributed Systems, 22(6): 946–959, June 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

IBGT:2001:BGV

[IBM01a] IBM Blue Gene team. Blue Gene: a vision for protein science using a petaflop supercomputer. IBM Systems Journal, 40(2):310–327, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http:

                                        //www.research.ibm.com/
                                        journal/sj/402/allen.html;
                                         http://www.research.ibm.
                                        com/journal/sj/402/allen.
                                        pdf;   http://www.research.
                                        ibm.com/journal/sj/402/
                                        allen.txt.

team:2001:BGV

[IBM01b] IBM Blue Gene team. Blue Gene: a vision for protein science using a petaflop supercomputer. IBM Systems Journal, 40(2):310–327, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http:

//www.research.ibm.com/
journal/sj/402/allen.html;
http://www.research.ibm.
com/journal/sj/402/allen.
pdf;   http://www.research.
ibm.com/journal/sj/402/
allen.txt.

IBGT:2008:OIB

[IBM08] IBM Blue Gene team.

Overview of the IBM Blue Gene/P Project. IBM Journal of Research and Development, 52(1/2):199–??, January/March 2008. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL

http://www.research.ibm.
com/journal/rd/521/team.
html.

Team:2013:DIB

[IBM13a] IBM Blue Gene Team. Design of the IBM Blue Gene/ Q compute chip. IBM Journal of Research and Development, 57(1/2):1:1–1:13, January–March 2013. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Team:2013:IBG

[IBM13b] IBM Blue Gene Team. The IBM Blue Gene project. IBM Journal of Research and Development, 57(1/2):0:1–0:6, January–March 2013. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Team:2013:MVC

[IBM13c] IBM Blue Gene Team. Modeling, validation, and codesign of IBM Blue Gene/Q: Tools and examples. IBM Journal of Research and Development, 57(1/2):6:1–6:12, January–March 2013. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Iyer:2005:EDT

[IBP^{+}05] S. S. Iyer, J. E. Barth, Jr., P. C. Parries, J. P. Norum, J. P. Rice, L. R. Logan, and D. Hoyniak. Embedded DRAM: Technology platform for the Blue Gene/L chip. IBM Journal of Research and Development, 49 (2/3):333–350, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/iyer.pdf.

IEEE:1985:FIC

[IEE85] IEEE, editor. First International Conference on Supercomputing Systems: proceedings, St. Petersburg, Florida, USA, December 16–20, 1985. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-8654-5 (hard), 0-8186-0654-1 (paperback), 0-8186-4654-3 (Microfiche). LCCN QA76.5 .I5481 1985. IEEE catalog number 85CH2216-0. IEEE Computer Society order number 654.

IEEE:1989:SCR

[IEE89a] IEEE. Supercomputing: a collection of reports. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. various pp.

IEEE-SSS:1989:CSP

[IEE89b] IEEE Scientific Supercomputer Subcommittee. The computer spectrum: a perspective on the evolution of computing. Computer, 22 (11):57–63, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

IEEE:1990:PSN

[IEE90] IEEE, editor. Proceedings, Supercomputing ’90: November 12–16, 1990, New York Hilton at Rockefeller Center, New York, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2056-0 (IEEE paperback), 0-89791-412-0 (ACM paperback). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society order number 2056. IEEE 90CH2916-5.

IEEE-TCOS:1991:SAE

[IEE91] IEEE Computer Society.Technical Committee on Operating Systems. Supercomputing applications environment profile. Technical report, IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 1, 1991. 65 pp.

IEEE-TCOSA:1992:NSC

[IEE92] IEEE Computer Society.Technical Committee on Supercomputing and Applications. NSF supercomputer centers study. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. i + 29 pp.

IEEE:1993:INP

[IEE93a]  IEEE, editor.  IJCNN

’93 — Nagoya: proceedings of 1993 International Joint Conference on Neural Networks, Nagoya Congress Center, October 25–29, 1993, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1421-2, 0-7803-1423-9, 0-7803-1422-0. LCCN QA76.87 .I57 1993 v.1-3 (c1993). Three volumes. IEEE catalog number: 93CH3353-0.

IEEE:1993:PIS

[IEE93b] IEEE, editor. Proceedings / the 2nd International Symposium on High Performance Distributed Computing, July 20–23, 1993, Spokane, Washington, Cavanaugh’s Inn at the Park. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3900-8, 0-8186-3901-6. LCCN QA76.9.D5 I593 1993. IEEE catalog number 93TH0550-4.

IEEE:1993:PSPa

[IEE93c] IEEE, editor. Proceedings of the Scalable Parallel Libraries Conference: October 6–8, 1993, Mississippi State, Mississippi. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4980-1 (paper), 0-8186-4981-X (microfiche). LCCN QA76.58 .S34 1993.

IEEE:1993:PSPb

[IEE93d] IEEE Computer Society, editor. Proceedings, Supercomputing ’93: Portland, Oregon, November 15-19. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4342-0 (case). LCCN QA76.5 .S894 1993. ACM order number 415932. IEEE catalog number 93CH3342-3. IEEE Computer Society Press order number 4340-20.

IEEE:1994:IIS

[IEE94a] IEEE, editor. ISSIPNN ’94, 1994 International Symposium on Speech, Image Processing, and Neural Networks: proceedings, 13– 16 April 1994, Hong Kong Convention and Exhibition Centre, Hong Kong, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1865-X, 0-7803-1866-8. LCCN TA1630.I58 1994. Two volumes. IEEE catalog number 94TH0638-7.

IEEE:1994:PIN

[IEE94b] IEEE, editor. Proceedings of the IEEE 1994 National Aerospace and Electronics Conference, NAECON 1994, held at the Dayton Convention Center, May 23-27, 1994, Dayton, OH, USA, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1893-5, 0-7803-1894-3, 0-7803-1895-1. ISSN 0547-3578. LCCN TL 693 N37 1994. Two volumes.

IEEE:1994:PSH

[IEE94c]  IEEE, editor.  Proceed-

ings of the Scalable High-Performance Computing Conference, May 23–25, 1994, Knoxville, Tennessee. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Stevens:1994:HPD

[IEE94d] IEEE, editor. Proceedings of the Third IEEE International Symposium on High Performance Distributed Computing, August 2–5, 1994, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6395-2, 0-8186-6396-0. LCCN QA76.9.D5I328 1994.

IEEE:1994:PSW

[IEE94e]  IEEE, editor.  Proceed-

ings, Supercomputing ’94: Washington, DC, November 14–18, 1994, Supercomputing. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6605-6 (paper), 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case). ISSN 1063-9535. LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

IEEE:1995:ISI

[IEE95a] IEEE. 1003.10-1995 IEEE Standard for Information Technology — POSIX® — Based Supercomputing Application Environment Profile.

IEEE, New York, NY, USA, 1995. ISBN 1-55937-546-9 (print), 0-7381-0631-3 (electronic). 80 pp. US$71.00.

URL http://standards.
ieee.org/reading/ieee/std_
public/description/posix/
1003.10-1995_desc.html.

IEEE:1995:CPI

[IEE95b] IEEE, editor. Conference proceedings / the 1995 IEEE International Conference on Acoustics, Speech, and Signal Processing, May 9–12, 1995, Westin Hotel, Detroit, Michigan, USA, volume 4 of IEEE International Conference on Acoustics Speech and Signal Processing. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2432-3, 0-7803-2431-5, 0-7803-2433-1, 0-7803-2562-1. ISSN 0749-8411. LCCN TK 7882 S65 I16 1995. Five volumes.

IEEE:1995:DPC

[IEE95c] IEEE, editor. Digest of papers: Compcon ’95: technologies for the information superhighway: March 5–9, 1995, San Francisco, CA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2657-1 (hardcover), 0-8186-7029-0 (paperback), 0-7803-2658-X (microfiche). ISSN 1063-6390. LCCN QA 75.5 C58 1995. IEEE Computer Society Press order number PR07029. IEEE catalog number 95CH35737.

IEEE:1995:PFI

[IEE95d] IEEE, editor. Proceedings of the Fourteenth IEEE Symposium on Mass Storage Systems: storage — at the forefront of information infrastructures: second international symposium, Monterey, California, USA, September 11–14, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7064-9 (softcover), 0-7803-3098-6 (hardcover). ISSN 1051-9173. LCCN TK7895.M4 I5 1995. IEEE catalog number 95CB3586.

IEEE:1996:DTA

[IEE96a] IEEE, editor. Dual-use technologies and applications 6th Annual dual-use technologies and applications conference: Annual conference; 6th — June 1996, Syracuse, NY, ANNUAL DUAL USE TECHNOLOGIES AND AP-PLICATIONS CONFER-ENCE 1996; 6th. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996.

IEEE:1996:HPD

[IEE96b] IEEE, editor. High performance distributed computing: International symposium; 5th — August 1996, Syracuse, NY, PROCEED-INGS of THE IEEE INTER-NATIONAL SYMPOSIUM ON HIGH PERFORMANCE DISTRIBUTED COMPUT-ING 1996; 5th. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7582-9. ISSN 1082-8907. LCCN ????

IEEE:1996:IPP

[IEE96c] IEEE, editor. International parallel processing symposium: 10th — April 1996, Honolulu, HI, INTERNA-TIONAL PARALLEL PRO-CESSING SYMPOSIUM 1996; 10th. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7255-2, 0-8186-7257-9. LCCN ????

IEEE:1996:SPA

[IEE96d] IEEE, editor. Supercomputing ’96 proceedings of the 1996 ACM/IEEE Supercomputing Conference: November 17-22, 1996, Pittsburgh. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-89791-854-1. LCCN ???? Includes CD-ROM. ACM order no. 415962.IEEE Computer Society Press order no. RS00126.

IEEE:1997:HPC

[IEE97a] IEEE, editor. High performance computing: International conference; 4th — December 1997, Bangalore, India, INTERNA-TIONAL CONFERENCE ON HIGH PERFORMANCE COMPUTING 1094- 7256 1997; CONF 4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8068-7, 0-8186-8067-9, 0-8186-8069-5. LCCN ????

IEEE:1997:HAH

[IEE97b] IEEE, editor. HPC Asia ’97: High performance computing on the information superhighway: Conference — April 1997, Seoul, HPC ASIA 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7901-8, 0-8186-7902-6, 0-8186-7903-4. LCCN ????

Iobst:1995:PMT

[IGH95] Ken Iobst, Maya Gokhale, and Bill Holmes. Processing in memory: The Terasys massively parellel PIM array. Computer, 28(4):23– ??, April 1, 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Iffert:1994:OHP

[IH94] R. Iffert and U. Harms.

An open high-performance computing infrastructure in industry. Lecture Notes in Computer Science, 797: 109–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ide:2000:GMF

[IHE^{+}00] Nobuhiro Ide, Masashi Hirano, Yukio Endo, Shin ichi Yoshioka, Hiroaki Murakami, Atsushi Kunimatsu, Toshinori Sato, Takayuki Kamei, Toyoshi Okada, and Masakazu Suzuoki. 2.44 GFLOPS 300-MHz floatingpoint vector-processing unit for high-performance 3-D graphics computing. IEEE Journal of Solid-State Circuits, 35(7):1025–1033, July 2000. CODEN IJSCBC. ISSN 0018-9200 (print), 1558-173X (electronic). URL

http://www.hwswworld.com/
downloads/a7/1025ide.pdf.

Ishihata:1991:TGM

[IHIS91] H. Ishihata, T. Horie, S. Inano, and T. Shimizu. Third generation message passing computer AP1000. In Anonymous [Ano91q], pages 46–55. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Irmscher:1993:CDS

[IHK93] S. Irmscher, E. Hees, and T. Kutsche. A controlled damping system with continuous damping-force adjustment. In Anonymous [Ano93-31], pages 123–130. ISBN 0-947719-62-8. LCCN ????

Ishikawa:1993:IND

[IHSK93] M. Ishikawa, T. Hoshi, T. Sanda, and T. Kamel. Improvement of nuclear design methods for large LMFBR cores using cross-section adjustment [invited]. In Kusters et al. [KSW93], pages 593– 604. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Irtaza:2014:SIR

[IJM14] Aun Irtaza, M. Arfan Jaffar, and Muhammad Tariq Mahmood. Semantic image retrieval in a Grid computing environment using support vector machines. The Computer Journal, 57 (2):205–216, February 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/content/57/2/205.full.
pdf+html.

MIT:1994:IJS

[IJS94] The international journal of supercomputer applications and high performance computing, 1994. ISSN 1078-3482; 0890-2720. MIT Press, Cambridge, MA, USA.

Ibrahim:2014:TEY

[IJY^{+}14] Ayad Ibrahim, Hai Jin, Ali A.

Yassin, Deqing Zou, and Peng Xu. Towards efficient yet privacy-preserving approximate search in cloud computing. The Computer Journal, 57(2):241– 254, February 2014. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/content/57/2/241.full.
pdf+html.

Irani:1982:MDC

[IK82] K. B. Irani and N. G. Khabbaz. A methodology for the design of communication networks and the distribution of data in distributed supercomputer systems. IEEE Transactions on Computers, C-31(5):419–434, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=1676019.

Ito:1991:PEU

[IK91] N. Ito and Y. Kanada. Performance evaluation using random number generator and Ising Monte Carlo simulation — the application-based benchmark Test(AbBT). In Anonymous [Ano91q], pages 253–257. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Ishigami:1993:AES

[IK93] T. Ishigami and K. Kobayashi. Application of an expert system for real time prediction of reactor accident progression with uncertainty for radiological emergency response. In Kusters et al. [KSW93], pages 307–318. ISBN 3-923704-11-

9. LCCN ???? Two volumes.

Ikedo:1995:ASM

[Ike95] T. Ikedo. AIZU supercomputer: a massively parallel system for virtual reality problems. In Mirenkov et al. [M^{+}95], pages 54–62. ISBN 0-8186-7038-X. LCCN QA76.642.A43 1995.

Ina:1985:LSD

[IKM85] Hiroshi Ina, Sachio Kamiya, and Jiro Mikami. Languages and software development tools for supercomputers. Computer Physics Communications, 38(2):211–219, October/November 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465585900876.

Illert:1996:ASG

[Ill96] Chris Illert. The Australian Supercomputer Graphics Exhibition and First International Conchology Conference. Leonardo (Oxford, England), 29(2):158–??, ???? 1996. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282 (electronic).

Irvin:1996:MPD

[IM96] R. Bruce Irvin and B. P.

Miller. Mapping performance data for high-level and data views of parallel program performance. In ACM [ACM96], pages 69–77. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Ivanov:1993:MOI

[IMA93] K. N. Ivanov, M. A. Manolova, and T. G. Apostolov. Multilevel optimization of improved coarse-mesh steadystate diffusion calculations in hexagonal-Z geometry. In Kusters et al. [KSW93], pages 620–632. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Ikeda:1991:ASS

[IMF91] Minoru Ikeda, Masuhiro Mikami, and Kumiko Furuya. Atomic-Scale simulations for semiconductors by supercomputer. Fujitsu scientific and technical journal, 27(2):211– ??, Summer 1991. CODEN FUSTA4. ISSN 0016-2523.

Iori:1993:HFA

[IMP93] G. Iori, E. Marinari, and G. Parisi. Heteropolymer folding on a APE-100 supercomputer. In Herrmann and Karsch [HK93a], pages 275–284. ISBN 981-02-1643-2. ISSN 0129-1831 (print), 1793-6586 (electronic). LCCN QC20 .W67 1993.

Infante:1986:AIE

[Inf86] E. F. Infante. An agenda for improved evaluation of supercomputer performance: a report. Technical report, National Academy Press: [Available from the Energy Engineering Board], Washington, DC, USA, 1986. x + 58 pp.

Inadomi:2001:IEP

[INKN01] Yuichi Inadomi, Tatsuya Nakano, Kazuo Kitaura, and Umpei Nagashima. Increased efficiency of parallel calculations of fragment molecular orbitals by using fine-grained parallelization on a HITACHI SR8000 supercomputer. Lecture Notes in Computer Science, 2110:569–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2110/21100569.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2110/21100569.
pdf.

IEEE-TAB:1987:SIG

[Ins87a] Institute of Electrical and Electronics Engineers. Technical and Activities Board. Supercomputing: an informal glossary of terms. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. 20 pp.

IEEE:1987:SIG

[Ins87b] Institute of Electrical and Electronics Engineers. United States Activities Board. Supercomputing: an informal glossary of terms. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. 20 pp. LCCN QA76.5 .S896 1987 Bar.

IEEE:1990:SIG

[Ins90] Institute of Electrical and Electronics Engineers. United States Activities Board. Scientific Supercomputer Subcommittee. Supercomputing: an informal glossary of terms. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, second edition, 1990. 24 pp. LCCN QA76.5 .S896 1990. IEEE catalog no. UH0182-6.

IDC:1981:SMR

[Int81] International Data Corporation, Framingham, MA, USA. Supercomputer market: 1981 report, 1981. 47 pp.

Intel:1991:OIX

[Int91] Intel Corporation, Santa Clara, CA, USA. Overview of the i860 XP supercomputing microprocessor, 1991. various pp.

Intel:1992:MSP

[Int92] Intel Corporation. Multimedia and supercomputing processors. Intel, Santa Clara, CA, USA, 1992.

ISBN 1-55512-149-7. various pp. LCCN TK7895.M5 I584 1992.

Ipek:2019:MAD

[Ipe19] E. Ipek. Memristive accelerators for dense and sparse linear algebra: From machine learning to high-performance scientific computing. IEEE Micro, 39(1):58–61, January/February 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Illert:1995:NAS

[IS95] Chris Illert and Dave Sims.

In the news: Australian supercomputer graphics exhibition; VR World 95: The new element in VR is people. IEEE Computer Graphics and Applications, 15(4): 89–91, July 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Islam:2020:BHP

[IS20] R. Islam and G. Shah. Building a high-performance resilient scalable storage cluster for CORAL using IBM ESS. IBM Journal of Research and Development, 64(3/4): 4:1–4:9, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Isaksen:1993:PEO

[Isa93]  L. Isaksen.  Parallelizing

the ECMWF optimum interpolation analysis. In Hoffmann and Kauranne [HK93b], pages 240–249. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Iskander:1996:FET

[Isk96] M. F. Iskander, editor. Frontiers in education: Technologybased re-engineering engineering education: Annual conference; 26th — November 1996, Salt Lake City, UT, FRONTIERS IN ED-UCATION CONFERENCE 1996//V1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3349-7, 0-7803-3348-9, 0-7803-3350-0, 0-7803-3720-4. ISSN 0190-5848. LCCN ????

Ishigami:1993:DCP

[ITOK93] T. Ishigami, M. Tomizawa, K. Oyama, and K. Kobayashi. Development of a computer program streem for simulating transportation evacuation in emergency. In Anonymous [Ano93-31], pages 857–864. ISBN 0-947719-62-8. LCCN ????

Itoh:1987:TFL

[IU87] M. Itoh and K. Uchida. Trends in Fujitsu large scale computer technology. Parallel Computing, 5(1–2):105– 115, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Iwaya:1990:NSS

[Iwa90] A. Iwaya. NEC supercomputer SX-3 series architecture and software. In Glowinski and Lichnewsky [GL90], pages 353–359. ISBN 0-89871-264-5. LCCN QC39 .I49 1990.

Iwaya:1992:SES

[Iwa92] A. Iwaya. Supercomputing enhancements in support of large scale problem solving. In Loffler and Muller [LM92], page 29.1. ISBN ???? LCCN ????

Jacob:1992:DIY

[JA92a] Robert Jacob and John Anderson. Do-it-yourself massively parallel supercomputer does useful physics. Computers in Physics, 6(3): 244–??, May 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4823073.

Jacob:1992:DMP

[JA92b] Robert Jacob and John Anderson. Do-it-Yourself massively parallel supercomputer does useful physics. Computers in Physics, 6(3):244– ??, May 1, 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Jablonski:1988:SAR

[Jab88] Daniel G. Jablonski. Simulated annealing, reversible computation and neural nets. Technical report SRC-TR-88-015, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 58 pp.

Jablonowski:1990:GGM

[Jab90] David Jablonowski. GMB: Graph manager/browser. Technical Report CSRD 968, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1990. various pp.

Jezequel:1992:PEE

[JAB92] J.-M. Jezequel, F. Andre, and F. Bergheul. A parallel execution environment for sequential object oriented language. In ACM [ACM92b], pages 368–376. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Jabonsen:1993:CSI

[Jab93] K. P. Jabonsen. 93SC043 crash simulation implementation on a massively parallel computer. In Anonymous [Ano93-31], pages 257– 264. ISBN 0-947719-62-8. LCCN ????

Jackson:1985:DMD

[Jac85] Daniel Thomas Jackson.

Data movement in doall loops. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1985. iv + 70 pp.

Jalote:1994:SET

[Jal94] P. Jalote. Software engineering trend. In Balakrishnan [Bal94], pages 411–412. ISBN 0-07-462044-4. LCCN ????

James:1995:WGI

[Jam95]  C. James.  What goes

into an information warehouse? Computer, 28(8): 84–85, August 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Janosi:1996:HTP

[Jan96] I. M. Janosi. Highway traffic and price increase in the baking industry: Foundation of the systematic windshield dirtology;-). In Wolf et al. [WSB96], pages 187– 192. ISBN 981-02-2635-7. LCCN ????

Jarvis:2012:EPM

[Jar12] S. A. Jarvis. Editorial performance modelling, benchmarking and simulation of high-performance computing systems. The Computer Journal, 55(2):136– 137, February 2012. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/content/55/2/136.full.
pdf+html.

Jayasimha:1987:PAS

[Jay87] Doddaballapur N. Jayasimha. Parallel access to synchronization variables. Technical Report CSRD-630, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 9 pp.

Jayasimha:1988:CSP

[Jay88a] Doddaballapur N. Jayasimha. Communication and synchronization in parallel computation. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1988. xiv + 140 pp.

Jayasimha:1988:DS

[Jay88b] Doddaballapur N. Jayasimha. Distributed synchronizers. Technical Report CSRD 713, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 23–27 pp.

Johnson:1990:MRI

[JB90] R. F. Johnson, Jr. and D. G.

Brunder. Magnetic resonance image interpolation for 3-dimensional data analysis and display. In Pitcher [Pit90], pages 413–421. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Jungheim:1991:SHR

[JBI91] Louis N. Jungheim, Donald B. Boyd, and Joseph M. Indelicato. Synthesis, hydrolysis rates, supercomputer modeling, and antibacterial activity of bicyclic tetrahydropyridazinones. Journal of medicinal chemistry, 34(5): 1732–??, May 1, 1991. CO-DEN JMCMAR. ISSN 0022-2623.

Johnson:1997:AIS

[JBWB97] S. A. Johnson, D. T. Borup, J. W. Wiskin, and M. J. Berggren. Application of inverse scattering and other refraction corrected methods to environmental imaging with acoustic or electromagnetic energy. In Delic and Wheeler [DW97], pages 295–312. ISBN 0-89871-378-1. LCCN ????

Center:1987:SJN

[JC87a] Science at the John von Neumann National Supercomputer Center: annual research report, 1987. ISSN 1046-0632. Consortium for Scientific Computing, Princeton, NJ, USA.

JVNNSC:1987:SAJ

[JC87b] Science at the John von Neumann National Supercomputer Center: annual research report, 1987. ISSN 1046-0632. Consortium for Scientific Computing, Princeton, NJ, USA.

Jain:1994:PPN

[JC94a] A. Jain and N. S. Chaudhari. On parallel processing of natural languages using La-grammar formalism. In Mahajan et al. [M^{+}94], pages 407–412. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Jain:1994:UAP

[JC94b] A. Jain and N. S. Chaudhari. A unified approach for parallel parsing of ambiguous context-free grammars. In Mahajan et al. [M^{+}94], pages 401–406. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Joubert:1994:PSP

[JC94c] W. Joubert and G. F. Carey. PCG: a software package for the iterative solution of linear systems on scalar, vector and parallel computers. In IEEE [IEE94c], pages 811– 816. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Jung:1994:RMA

[JC94d] W. W. Jung and S. R. Cho.

Reduction method of aspiration wind noise influenced by doorframe stiffness and weather strip. In Anonymous [Ano94-75], pages 479–486. ISBN 0-947719-68-7. LCCN ????

Jung:1994:CAE

[JCJY94] E. W. Jung, B. L. Choi, J. W. Jeon, and Y. D. Yoo. A combination of analytical and experimental procedure for optimizing strength and fatigue performance of suspension system. In Anonymous [Ano94-75], pages 351– 358. ISBN 0-947719-68-7. LCCN ????

Jost:1995:MIP

[JD95] R. Gilbert Jost and Samuel J.

Dwyer III, editors. Medical imaging 1995. PACS design and evaluation: engineering and clinical issues: 28 February–2 March 1995, San Diego, California, number 2435 in Proceedings — SPIE The International Society for Optical Engineering. SPIE Optical Engineering Press, Bellingham, WA, USA, 1995. ISBN 0-8194-1783-1. ISSN 0361-0748. LCCN R857.P52M426 1995.

Jennions:1987:DC

[Jen87] Ian K. Jennions. Designing with the Cray. Leading Edge (Evendale OH), pages 24–27, Spring 1987. CODEN LEEDEE.

Jenkins:1988:NWJ

[Jen88] Richard A. Jenkins. The next wave of Japanese supercomputers. Computers in Physics, 2(2):34– ??, March 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4822679.

JPL:1990:SPJ

[Jet90] Jet Propulsion Laboratory (US). Supercomputing project: June 1989 through June 1990. Jet Propulsion Laboratory, Pasadena, CA, USA, 1990. 24 pp. Shipping list no. 91-0051-P. JPL 400-443 11/90.

JPL:1991:SPS

[Jet91] Jet Propulsion Laboratory (US). Supercomputing project: state-of-the-art supercomputing resources for JPL and Caltech. Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA, USA, September 1991. 10 pp.

JPL:1992:SPS

[Jet92] Jet Propulsion Laboratory (U.S.). Supercomputing project: state-of-the-art supercomputing resources for JPL and Caltech Campus. Technical Report JPL D-8875, Jet Propulsion Laboratory, Pasadena, CA, USA, 1992. 10 pp.

Jezequel:2000:OOF

[Jéz00] J.-M. Jézéquel. An objectoriented framework for data parallelism. ACM Computing Surveys, 32(1es), March 2000. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/surveys/2000-32-
1/p31-jezequel/. Article

No. 31.

Jablonowski:1988:DGT

[JG88] David Jablonowski and Vincent A. Guarna. A dynamic graph tool and its use in an integrated programming environment. Technical Report CSRD 746, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1988. ii + 26 pp.

Jackson:1999:ISA

[JG99] Larry S. Jackson and Ed Grossman. Integration of synchronous and asynchronous collaboration activities. ACM Computing Surveys, 31(2es): ??, June 1999. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL

http://www.acm.org/pubs/
articles/journals/surveys/
1999-31-2es/a12-jackson/
a12-jackson.pdf;   http:
//www.acm.org/pubs/citations/
journals/surveys/1999-31-
2es/a12-jackson/;   http:
//www.acm.org/pubs/citations/
journals/surveys/1999-31-
2es/a12-jackson/#abstract.

Jia-Hsu:1993:DMR

[JHGLG93] Y. Jia-Hsu, H. Gilson, K. Long, and R. A. Gibbs. Data management for resequencing DNA. In Lim et al. [L^{+}93], pages 207–218. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Joslin:1995:SPS

[JHZ95] Ronald D. Joslin, Ulf R. Hanebutte, and Mohammad Zubair. Scalability of parallel spatial direct numerical simulations on Intel Hypercube and IBM SP1 and SP2. Journal of Scientific Computing, 10(2): 233–269, June 1995. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02089951;   http:
//link.springer.com/content/
pdf/10.1007/BF02089951;
 http://www.springerlink.
com/openurl.asp?genre=article&
issn=0885-7474&volume=10&
issue=2&spage=233-269.

Ji:1991:BEM

[Ji91]  Link Ji.  Boundary ele-

ment methods and supercomputer computations for some distributed parameter control problems. Thesis (Ph.D.), Texas A&M University, 1991. viii + 107 pp.

Jian:1994:SAC

[Jia94] M. A. Jian. A systematic approach to the course design of data organization. In Balakrishnan [Bal94], pages 308–315. ISBN 0-07-462044-4. LCCN ????

Jih:1988:DMN

[jJ88] Tsae jiunn Jih. 3D dip moveout on the NEC SX-2 supercomputer. Thesis (M.S.), Dept. of Computer Science, University of Houston, Houston, TX, USA, 1988. viii + 78 pp.

Jain:1994:VML

[JJYL94] R. P. Jain, V. Jagannathan, R. D. S. Yadav, and S. V. Lawande. Visualisation and modelling of liquid poison injection shutdown system in PHWRs. In Mahajan et al. [M^{+}94], pages 75–80. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Jacobsen:2019:MCS

[JKL19] Douglas M. Jacobsen, Randy Kleinman, and Harold Longley. Managing a Cray supercomputer as a git branch. Concurrency and Computation: Practice and Experience, 31(16):e5092:1– e5092:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Joo:1993:NAC

[JKNK93] Hyung Kook Joo, Chang Hyo Kim, Jae Man Noh, and S.- H. Kim. A new approach to core-reflector boundary conditions for nodal reactor computations. In Kusters et al. [KSW93], pages 546– 557. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Johnson:1998:PVM

[JLC98] Andy Johnson, Jason Leigh, and Jim Costigan. Projects in VR: Multiway tele-immersion at Supercomputing 97. IEEE Computer Graphics and Applications, 18(4):6–9, July/ August 1998. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Johnsson:1989:ECGb

[JM89a] Lennart S. Johnsson and K. K. Mathur. Experience with the conjugate gradient method for stress analysis on a data parallel supercomputer. International journal for numerical methods in engineering, 27(3):523–546, December 1, 1989. See [JM89c].

Johnsson:1989:DSA

[JM89b] S. Lennart Johnsson and Kapil K. Mathur. Data structures and algorithms for the finite element method on a data parallel supercomputer. Technical report TMC-52, Thinking Machines Corp, Cambridge, MA, USA, 1989. 881–908 pp.

Johnsson:1989:ECGa

[JM89c] S. Lennart Johnsson and Kapil K. Mathur. Experience with the conjugate gradient method for stress analysis on a data parallel supercomputer. Technical report TMC-42, Thinking Machines Corp, Cambridge, MA, USA, 1989. 523–546 pp.

Johnsson:1990:DSA

[JM90] S. L. Johnsson and K. K.

Mathur. Data structures and algorithms for the finite element method on a data parallel supercomputer. International Journal for Numerical Methods in Engineering, 29 (4):881–??, March 1990. CO-DEN IJNMBH. ISSN 0029-5981.

Johnsson:1993:MPC

[JM93] S. L. Johnsson and K. K.

Mathur. Massively parallel computing: Mathematics and communications libraries. In Hoffmann and Kauranne [HK93b], pages 250–285. ISBN 981-02-1429-4. LCCN QA76.58 E354

Jorda:1995:SBS

[JML95] J. Jorda, A. Mzoughi, and D. Litaize. Semi-linear and bi-base storage schemes classes: General overview and case study. In ACM [ACM95a], pages 299–307. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Jorda:1996:PVP

[JML96] J. Jorda, A. Mzoughi, and D. Litaize. Performance of the vectorial process VEC-SM2 using serial multiport memory. In ACM [ACM96], pages 390–406. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Jamison:1998:VYF

[JNM^{+}98] J. Jamison, R. Nicklas, G. Miller, K. Thompson, R. Wilder, L. Cunningham, and C. Song. vBNS: not your father’s Internet. IEEE Spectrum, 35(7):38–46, July 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Joerg:1987:DPS

[Joe87] Christopher F. (Christopher Frank) Joerg. Design of a packet switched routing chip for the dataflow supercomputer. Thesis (B.S.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1987. 60 pp. Supervised by George A. Boughton.

JVNNSC:1986:JNN

[Joh86a] The John von Neumann National Supercomputer Center, Princeton, NJ, USA. The John von Neumann National Supercomputer Center, 1986. 4 pp.

XXX:1986:JNN

[Joh86b] The John von Neumann National Supercomputer Center, Princeton, NJ, USA. The John von Neumann National Supercomputer Center, 1986. 4 pp.

Johnson:1986:RAN

[Joh86c]  Gary M. Johnson.  Re-

mote access for NAS supercomputing in a university environment: final report covering the period April 1, 1985 through March 31, 1986. NASA contractor report NASA CR-177020, Colorado State University and Institute for Computational Studies; National Aeronautics and Space Administration; National Technical Information Service, distributor, Ft. Collins, CO, USA, 1986. ?? pp.

Johnson:1988:AES

[Joh88] Karen Lee Johnson. An advisory expert system for supercomputer job control language. Thesis (M.S.), Florida State University, Tallahassee, FL, USA, 1988. iv + 59 pp.

Johnsson:1990:DPS

[Joh90] S. Lennart Johnsson. Data parallel supercomputing. Technical report TMC-48, Thinking Machines Corp, Cambridge, MA, USA, 1990. 29 pp.

Johnston:1991:BMD

[Joh91] Anna M. Johnston. Basic methods for discrete logarithm computation. Technical report SRC-TR-91-047, Supercomputing Research Center: IDA, Lanham, MD, USA, October 15, 1991. 23 pp.

Johnston:1992:FLE

[Joh92] Anna M. Johnston. Fast logarithm and exponentiation approximations with applications. Technical report SRC-TR-92-070, Supercomputing Research Center: IDA, Lanham, MD, USA, June 2, 1992. 11 pp.

Johnson:1994:NES

[Joh94] D. LaMont Johnson. The National Education Supercomputer Program: a model for business and government involvement in education. Computers in the Schools, 11 (2):1–??, ???? 1994. ISSN 0738-0569.

Johnston:1997:RSC

[Joh97] W. W. Johnston. Rationale and strategy for a 21st Century scientific computing architecture: The case for using commercial symmetric multiprocessors as supercomputers. International Journal of High Speed Computing, 9(3): 191–222, 1997. CODEN IH-SCEZ. ISSN 0129-0533.

Jones:2018:EST

[JOK^{+}18] Terry Jones, George Ostrouchov, Gregory A. Koenig, Oscar H. Mondragon, and Patrick G. Bridges. An evaluation of the state of time synchronization on leadership class supercomputers. Concurrency and Computation: Practice and Experience, 30 (4):??, February 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/
doi/abs/10.1002/cpe.4341.

Jones:1989:EDC

[Jon89] T. Jones. Engineering design of the Convex C2. Computer, 22(1):36–44, January 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1996:TSL

[Jon96]  Anita K. Jones.  Tak-

ing stock, looking ahead: Part 2. modernizing highperformance computing for the military. IEEE Computational Science & Engineering, 3(3):71–74, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

                                        //www.computer.org/cse/
                                        cs1998/c3071abs.htm.

Jones:2003:MOC

[Jon03] Steve Jones. My other computer is a supercomputer. Linux Journal, 2003(115): 5, November 2003. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Jones:2019:SIP

[Jon19] K. E. Jones. Supercomputing improves predictions of fluid flow in rock. Computing in Science and Engineering, 21(6):74–76, November 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Jordan:1986:MPU

[Jor86] Harry Frederick Jordan. Multiprocessors and the principle of universal parallelism, 1986. 1 videocassette (50 min.).

Jordan:1987:PCL

[Jor87]  K. E. Jordan.  Perfor-

mance comparison of largescale scientific processors: Scalar mainframes, mainframes with vector facilities, and supercomputers. Computer, 20(3):10–23, March 1987. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Jalby:1990:SAI

[JP90] William Jalby and Bernard Philippe. Stability analysis and improvement of the Block Gram–Schmidt algorithm. Technical Report CSRD 913, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1990. 20 pp.

Jones:1994:PAA

[JP94] M. T. Jones and P. E. Plassmann. Parallel algorithms for the adaptive refinement and partitioning of unstructured meshes. In IEEE [IEE94c], pages 478–485. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Joiner:2008:EOT

[JPMG08] David Joiner, Charles Peck, Thomas Murphy, and Paul Gray. Education, outreach, and training for high-performance computing. Computing in Science and Engineering, 10(5):40– 45, September/October 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Joubert:1994:PCT

[JPTE94] G. R. Joubert, F. J. Peters, D. Trystram, and D. J.

Evans, editors. Parallel computing: trends and applications: proceedings of the international conference ParCo93, Grenoble, France, 7–10 September 1993, volume 9 of Advances in parallel computing. North-Holland, Amsterdam, The Netherlands, 1994. ISBN 0-444-81841-3. LCCN QA76.58 .P3794 1993.

Jensen:1991:FAA

[JR91] David W. Jensen and Daniel A. Reed. File archive activity in a supercomputer environment. Technical report UIUCDCS-R-91-1672; Tapestry technical report TTR91-29; UILU-ENG-91-1718, Dept. of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, April 1991. 29 pp.

Joseph:1994:VDI

[JR94] M. Joseph and P. B. Rao. Visual demonstration of the influence of equation of state on energy release in a core disruptive accident. In Mahajan et al. [M^{+}94], pages 54–59. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Jegou:1986:DSP

[JS86] Y. Jegou and A. Seznec. Data synchronized pipeline architecture: pipelining in multiprocessing environments. Journal of Parallel and Distributed Computing, 3(4): 508–526, December 1986.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Johnson:1995:SSI

[JS95] S. Johnson and S. Scott. A supercomputer system interconnect and scalable IOS. In IEEE [IEE95d], pages 357– 367. ISBN 0-8186-7064-9 (softcover), 0-7803-3098-6 (hardcover). ISSN 1051-9173. LCCN TK7895.M4 I5 1995. IEEE catalog number 95CB3586.

Jensen:1987:SRC

[JT87] Klavs F. Jensen and Donald G. Truhlar, editors. Supercomputer research in chemistry and chemical engineering: American Chemical Society Winter Symposium, March 16–17, 1987, Minneapolis, MN, number 353 in ACS symposium series ISSN: 0097-6156; 353. American Chemical Society, Washington, DC, USA, 1987. ISBN 0-8412-1430-1. ISSN 0097-6156. LCCN QD39.3.E46 S9271 1987. Developed from a symposium sponsored by the Division of Industrial and Engineering Chemistry of the American Chemical Society … [et. al.] at the Industrial and Engineering Chemistry Winter Symposium, Minneapolis, Minnesota, March 16-17, 1987.

Jablonowski:1991:VUM

[JT91] David Jablonowski and Allan Tuchman. Vista users manual. Technical Report CSRD 1068, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. various pp.

Jiang:2022:EAD

[JTX^{+}22] Bingting Jiang, Zhuo Tang, Xiong Xiao, Jing Yao, Ronghui Cao, and Kenli Li. Efficient and automated deployment architecture for OpenStack in TianHe SuperComputing environment. IEEE Transactions on Parallel and Distributed Systems, 33(8):1811–1824, 2022. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Jaffre:1993:AMT

[JV93] J. Jaffre and J.-L. Vaudescal. Arnoldi’s method for twogroup neutron diffusion. In Kusters et al. [KSW93], pages 512–520. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Juurlink:1998:QCP

[JW98] Ben H. H. Juurlink and Harry A. G. Wijshoff. A quantitative comparison of parallel computation models. ACM Transactions on Computer Systems, 16(3): 271–318, August 1998. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

                                        tronic). URL http://www.

acm.org:80/pubs/citations/
journals/tocs/1998-16-3/
p271-juurlink/.

Jahshan:1993:CAT

[JWG93] S. N. Jahshan, C. A. Wemple, and B. D. Ganapol. Comparison of analytical transport and stochastic solutions for neutron slowing down in an infinite medium. In Kusters et al. [KSW93], pages 742– 751. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Juang:1992:PCN

[JY92] F. Juang and Gung-Chung Yang. Performance characterization of N3S: a methodology and a case study. Technical Report CSRD 1198, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. 127 pp.

Komori:1991:HSA

[KA91] Yasuhiro Komori and Katsuhiko Akahori. High-Speed structural analysis program: POPLAS/FEM5 on supercomputer. Fujitsu scientific and technical journal, 27(1): 141–??, Spring 1991. CO-DEN FUSTA4. ISSN 0016-2523.

Kleijnen:1992:PNG

[KA92] J. P. C. Kleijnen and B. Annink. Pseudorandom number generators for supercomputers and classical computers: a practical introduction. European Journal of Operational Research, 63(1):76–85, November 25, 1992. CO-DEN EJORDT. ISSN 0377-2217 (print), 1872-6860 (electronic).

Kambayashi:1993:IUI

[KA93a] S. Kambayashi and Y. Abe. Instruction understanding for intelligent robots in nuclear facilities. In Kusters et al. [KSW93], pages 398–407. ISBN 3-923704-11-9. LCCN ???? Two volumes. Koschmieder:1993:VRS

[KA93b] D. Koschmieder and J. Altes. Visualization of results from structure mechanical studies. In Kusters et al. [KSW93], pages 781–?? ISBN 3-923704-11-9. LCCN ???? Two volumes.

Kortas:1996:PPM

[KA96] Samuel Kortas and Philippe Angot. A practical and portable model of programming for iterative solvers on distributed memory machines. Parallel Computing, 22(4):487–512, June 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kaegi:1995:TRO

[KABG95] A. Kaegi, N. Aboulenein, D. C. Burger, and J. R. Goodman. Techniques for reducing overheads of sharedmemory multiprocessing. In ACM [ACM95a], pages 11– 20. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Kacsuk:2002:HSG

[Kac02] P. Kacsuk. Hungarian supercomputing grid. Lecture Notes in Computer Science, 2330:671–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2330/23300671.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2330/23300671.
pdf.

Kading:1994:DDS

[Kad94] W. Kading. The Daimler-Benz driving simulator, technical setup and spectrum of application. In Anonymous [Ano94-75], pages 601–608. ISBN 0-947719-68-7. LCCN ????

Kahle:1991:WAI

[Kah91] Brewster Kahle. Wide area information servers a supercomputer on every desk, 1991. 1 videocassette (64 min.) sd. + col. 1/2 in.

Kahaner:1992:SJC

[Kah92] David K. Kahaner. Supercomputing Japan ’92 conference: Supercomputing Japan ’92 is described, with particular emphasis on Hitachi’s new 32-GFLOP supercomputer. Scientific information bulletin, 17(3):5–??, July 1, 1992. CODEN SINBEM. ISSN 1048-5678.

Kahaner:1993:ESS

[Kah93a] David K. Kahaner. Experiences of second summer student at NAL, STA Fellowship Program, 15 September 1993. Scientific information bulletin, 18(4):63–??, October 1, 1993. CODEN SINBEM. ISSN 1048-5678.

Kahaner:1993:SRS

[Kah93b] David K. Kahaner. Special report: Supercomputing — the view from Japan. IEEE Micro, 13(1):67–70, January/February 1993. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1993:SVJ

[Kah93c] David K. Kahaner. Supercomputing — the view from Japan. IEEE Micro, 13(1):67–70, January/ February 1993. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kahaner:1994:CJS

[Kah94] David K. Kahaner. CSE in Japan: Stiff competition: High-performance computing in Japan. IEEE Computational Science & Engineering, 1(2):84–86, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Kahaner:1997:GB

[Kah97]  David Kahaner.  Global

broadcast: Taiwan’s National Center For High-Performance Computing. IEEE Concurrency, 5(2):91– 93, April/June 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1997/pdf/p2091.
pdf.

Kamath:1986:SNS

[Kam86] Chandrika Kamath. Solution of nonsymmetric systems of equations on a multiprocessor. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. vii + 178 pp.

Kwack:2019:RAC

[KAMB19] JaeHyuk Kwack, Galen Arnold, Celso Mendes, and Gregory H. Bauer. Roofline analysis with Cray performance analysis tools (Cray-Pat) and roofline-based performance projections for a future architecture. Concurrency and Computation: Practice and Experience, 31 (16):e4963:1–e4963:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kang:2015:SIP

[Kan15] K. S. Kang. Scalable implementation of the parallel multigrid method on massively parallel computers. Computers and Mathematics with Applications, 70(11): 2701–2708, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122115003582.

Karlovsky:1989:AMP

[Kar89]  Steve Karlovsky.  Auto-

matic management of programmable caches: algorithms and experience. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1989. vi + 83 pp.

Karimi:1993:GDN

[Kar93] E. Karimi. GenEng: a dialogue-based natural language interface to the Gen-Bank. In Lim et al. [L^{+}93], pages 51–60. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Karia:1994:LBP

[Kar94] R. J. Karia. Load balancing of parallel volume rendering with scattered decomposition. In IEEE [IEE94c], pages 252–258. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Karlin:2010:AD

[Kar10]  S. Karlin.  Die another

day. IEEE Spectrum, 47(3): 24–25, March 2010. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karastoyanova:2013:SCS

[Kar13] Dimka Karastoyanova. Springer computing special issue: adaptation in service-oriented and Cloud Computing. Computing: Archiv fur informatik und numerik, 95(6): 449–451, June 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00607-013-0319-z.

Kaufman:1991:VV

[Kau91] Arie Kaufman. Volume visualization. IEEE Computer Society Press tutorial. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9020-8. ix + 479 pp. LCCN T385 .K375 1991.

Kauranne:1993:VAE

[Kau93a] T. Kauranne. 4D variational assimilation, ensemble forecasting and parallel computing. In Hoffmann and Kauranne [HK93b], pages 286– 311. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Kauranne:1993:EUP

[Kau93b] T. Kauranne. Experience on using parallel computers in meteorology: Summary of the discussion. In Hoffmann and Kauranne [HK93b], pages 514–522. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Kavenoky:1992:SAN

[Kav92] Alain Kavenoky. Supercomputing applied to nuclear reactors. Annals of nuclear energy, 19(10):679–??, October 1992. CODEN ANENDJ. ISSN 0306-4549.

Kazarinoff:1992:BRB

[Kaz92] Nicholas D. Kazarinoff. Book review: IBM’s 360 and Early 370 Systems (Emerson W. Pugh, Lyle R. Johnson, and John H. Palmer). SIAM Review, 34(1):158, March 1992. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kazic:1993:RAB

[Kaz93] T. Kazic. Reasoning about biochemical compounds and processes. In Lim et al. [L^{+}93], pages 35–50. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Karp:1988:CPF

[KB88] Alan H. Karp and Robert G. Babb. A comparison of 12 parallel Fortran dialects. IEEE Software, 5(5):52– 67, September 1988. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kauranne:1993:SEP

[KB93] T. Kauranne and S. R. M.

Barros. Scalability estimates of parallel spectral atmospheric models. In Hoffmann and Kauranne [HK93b], pages 312–328. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Kohn:1994:RPP

[KB94] S. R. Kohn and S. B. Baden.

A robust parallel programming model for dynamic nonuniform scientific computations. In IEEE [IEE94c], pages 509–517. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Kumar:1996:EVC

[KB96] A. Kumar and L. Bhuyan.

Evaluating virtual channels for cache-coherent sharedmemory multi-processors. In ACM [ACM96], pages 253– 260. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Krasowski:1997:UVD

[KB97] E. E. Krasowski and Z. Burski.

Using vibrations as diagnostic signals in monitoring and computer controlling of friction and wear processes of tractor engine units. In Roller [Rol97], pages 327–336. ISBN 0-947719-88-1 (paperback). LCCN ????

Kwack:2018:HHB

[KB18] JaeHyuk Kwack and Gregory H. Bauer. HPCG and HPGMG benchmark tests on multiple program, multiple data (MPMD) mode on Blue Waters — a Cray XE6/ XK7 hybrid system. Concurrency and Computation: Practice and Experience, 30 (1):??, January 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kehrer:2019:EPS

[KB19] Stefan Kehrer and Wolfgang Blochinger. Elastic parallel systems for high performance cloud computing: State-of-the-art and future directions. Parallel Processing Letters, 29(02):??, June 2019. ISSN 0129-6264 (print), 1793-642X (elec-

tronic). URL https://www.
worldscientific.com/doi/
10.1142/S0129626419500063.

Kuck:1974:MPO

[KBC^{+}74] D. Kuck, P. Budnik, S. C. Chen, E. Davis, Jr., J. Han, P. Kraska, D. Lawrie, Y. Muraoka, R. Strebendt, and R. Towle. Measurements of parallelism in ordinary FOR-TRAN programs. Computer, 7(1):37–46, January 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khier:1997:NSS

[KBD97] W. Khier, M. Breuer, and F. Durst. Numerical simulation of side wind effects on high speed trains. In Roller [Rol97], pages 441– 450. ISBN 0-947719-88-1 (paperback). LCCN ????

Kurzak:2010:SCM

[KBD10] Jakub Kurzak, David A.

Bader, and J. J. Dongarra, editors. Scientific computing with multicore and accelerators, volume 10 of Chapman and Hall/CRC computational science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-2536-X (hardback). xxxiii + 480 pp. LCCN Q183.9 .S325 2010.

Kerbyson:2013:PAT

[KBG^{+}13] Darren J. Kerbyson, Kevin J.

Barker, Diego S. Gallo, Dong Chen, Jose R. Brunheroto, Kyung Dong Ryu, George L. Chiu, and Adolfy Hoisie. A performance analysis of three generations of Blue Gene. Parallel Processing Letters, 23(4):1340007, December 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Kurzak:2008:PHP

[KBLD08] Jakub Kurzak, Alfredo Buttari, Piotr Luszczek, and Jack Dongarra. The PlayStation 3 for high-performance scientific computing. Computing in Science and Engineering, 10(3):84–87, May/June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kielmann:2002:PEH

[KBM^{+}02] Thilo Kielmann, Henri E. Bal, Jason Maassen, Rob van Nieuwpoort, Lionel Eyraud, Rutger Hofman, and Kees Verstoep. Programming environments for highperformance Grid computing: the Albatross Project. Future Generation Computer Systems, 18(8):1113–1125, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kerbyson:2014:PCC

[KBVH14] Darren J. Kerbyson, Kevin J. Barker, Abhinav Vishnu, and Adolfy Hoisie. A performance comparison of current HPC systems: Blue Gene/Q, Cray XE6 and InfiniBand systems. Future Generation Computer Systems, 30 (??):291–304, January 2014. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X13001337.

Katz:1989:SWC

[KC89] Alan Katz and Stephen H.

Casner. Supercomputer workstation communication. Special report ISI/SR-89-235, University of Southern California, Information Sciences Institute, Marina del Rey, CA, USA, June 1989. 24 pp.

Kratzer:1992:SMF

[KC92] Steven G. Kratzer and Andrew J. Cleary. Sparse matrix factorization on SIMD parallel computers. Technical report SRC-TR-92-063, Supercomputing Research Center: IDA, Lanham, MD, USA, April 10, 1992. 13 pp.

Kaftanoglu:1993:CMP

[KC93a] B. Kaftanoglu and N. Carkoglu.

93SC011 A computer-aided mechanical power transmission system design. In Anonymous [Ano93-31], pages 183– 190. ISBN 0-947719-62-8. LCCN ????

YongHeeKim:1993:PSA

[KC93b] Yong Hee Kim and Nam Zin Cho. Parallel Schwarz algorithm for the neutron diffusion equation on an MIMD architecture. In Kusters et al. [KSW93], pages 100– 111. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Kohl:1993:CCA

[KC93c] J. A. Kohl and T. L.

Casavant. 93SC005 creating custom, animated views for debugging and performance evaluation of massively parallel supercomputers for automotive applications. In Anonymous [Ano93-31], pages 207–212. ISBN 0-947719-62-8. LCCN ????

Karamcheti:1995:CAS

[KC95] Vijay Karamcheti and Andrew A. Chien. A comparison of architectural support for messaging in the TMC CM-5 and the Cray T3D. ACM SIGARCH Computer Architecture News, 23(2): 298–307, May 1995. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Koehler:2008:PAC

[KCG08] Seth Koehler, John Curreri, and Alan D. George. Performance analysis challenges and framework for high-performance reconfigurable computing. Parallel Computing, 34(4–5): 217–230, May 2008. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Krevat:2002:JSBa

[KCM02a] E. Krevat, J. G. Castaños, and J. E. Moreira. Job scheduling for the Blue-Gene/L system (research note). Lecture Notes in Computer Science, 2400:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2400/24000207.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2400/24000207.
pdf.

Krevat:2002:JSBb

[KCM02b] Elie Krevat, José G. Castaños, and José E. Moreira. Job scheduling for the Blue-Gene/L system. Lecture Notes in Computer Science, 2537:38–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2537/25370038.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2537/25370038.pdf.

Konuru:1994:UPP

[KCOP94] R. Konuru, J. Casas, S. Otto, and R. Prouty. A user-level process package for PVM. In IEEE [IEE94c], pages 48–55. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Koufaty:1995:DFS

[KCPT95] D. A. Koufaty, X. Chen, D. K. Poulsen, and J. Torrellas. Data forwarding in scalable shared-memory multiprocessors. In ACM [ACM95a], pages 255–264.

ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Kaufman:1993:VG

[KCY93] Arie Kaufman, Daniel Cohen, and Roni Yagel. Volume graphics. Computer, 26 (7):51–64, July 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:2014:AQS

[KCZJ14] Neeraj Kumar, Naveen Chilamkurti, Sherali Zeadally, and Young-Sik Jeong. Achieving quality of service (QoS) using resource allocation and adaptive scheduling in cloud computing with Grid support. The Computer Journal, 57(2):281–290, February 2014. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

                                        /comjnl.oxfordjournals.
                                        org/content/57/2/281.full.
                                        pdf+html.

Kaftanoglu:1993:CMS

[KD93] B. Kaftanoglu and H. Darendeliler. 93SC010 computeraided modelling of sheetmetal forming. In Anonymous [Ano93-31], pages 87– 94. ISBN 0-947719-62-8. LCCN ????

Klaassen:1995:PNM

[KDBG95] A. J. Klaassen, B. Delord, Y. Burnod, and E. Guigon.

Parallel neuron modelling using domain decomposition: Application toward learning sensori-motor sequences in prefrontal cortex. In Herrmann et al. [HWP95], pages 361–370. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Kogut:1989:SSS

[KDK89] J. B. Kogut, E. Dagotto, and A. Kocic. A supercomputer study of strongly coupled QED. Nuclear Physics B, 317 (2):271–??, May 1, 1989. CO-DEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic).

Kuck:1986:PST

[KDLS86] D. Kuck, E. Davidson, D. Lawrie, and A. Sameh. Parallel supercomputing today and the Cedar approach. Science, 231:967–974, 1986. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Kauzlaric:2014:SSP

[KDP^{+}14] David Kauzlarić, Marek Dynowski, Lars Pastewka, Andreas Greiner, and Jan G. Korvink. SYMPLER: SYMbolic ParticLE simulatoR with grid-computing interface. Computer Physics Communications, 185(3):1085– 1099, March 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465513004104.

Knott:1993:VCT

[KE93] D. Knott and M. Edenius. Validation of the CASMO-4 transport solution. In Kusters et al. [KSW93], pages 547–558. ISBN 3-923704-11-

9. LCCN ???? Two volumes.

Kedlaya:1992:PSC

[Ked92] Kiran Kedlaya. A procedure for solving certain simultaneous generalized Pell equations. Technical report SRC-TR-92-081, Supercomputing Research Center: IDA, Lanham, MD, USA, August 1992. 36 pp.

Kedlaya:1994:EIP

[Ked94] Kiran Kedlaya. Evaluating integer polynomials at the exponentials of algebraic numbers. Technical report SRC-TR-94-127, Supercomputing Research Center: IDA, Lanham, MD, USA, September 15, 1994. 9 pp.

Kelley:1985:CNL

[Kel85] John F. Kelley. CAL — a natural language program developed with the OZ paradigm: implications for supercomputing systems. Research Report RC 11324 (#50219), IBM Thomas J. Watson Research Center, Yorktown Heights, NY, 1985. 19 pp.

Kelly:1991:SSP

[Kel91] T. D. Kelly. Supercomputer study of a particle arch. In Haji-Sheikh et al. [HS^{+}91], pages 507–523. ISBN 1-56032-066-4. LCCN TA330 .I58 1991.

Kasyanov:1999:STS

[KEMB99] V. N. Kasyanov, V. A.

Evstigneev, J. V. Malinina, and J. V. Birjukova. Support tools for supercomputing and networking. Lecture Notes in Computer Science, 1593:1175–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennedy:1992:GEI

[Ken92] Ken Kennedy. Guest editor’s introduction to the special section on supercomputing. Communications of the ACM, 35(8):65, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
135229.html.

Kerschbaum:1994:PNT

[Ker94] W. K. Kerschbaum. ProSTEP-Introducing a new technology of product data management to industrial practice. In Anonymous [Ano94-75], pages 389–396. ISBN 0-947719-68-7. LCCN ????

Kumar:1994:RMA

[KESH94] B. Kumar, K. Eswar, P. Sadayappan, and C.-H. Huang. A reordering and mapping algorithm for parallel sparse Cholesky factorization. In IEEE [IEE94c], pages 803– 810. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Kumar:1995:CAP

[KESH95] B. Kumar, K. Eswar, P. Sadayappan, and C.-H. Huang. A clustering algorithm for parallel sparse Cholesky factorization. Parallel Processing Letters, 5(4):685–696, December 1995. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Kleinschmidt:1995:HNF

[KF95] A. Kleinschmidt and J. Frahm. Human neuroscience and functional MRI. current strategies and findings. In Herrmann et al. [HWP95], pages 45–60. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Kenway:1991:AVM

[KFB91] D. J. Kenway, C. F. Flatman, and W. N. Bauer. The architecture of the VisionSmart 3000 MIPS industrial supercomputer. In Anonymous [Ano91m], pages 265– ?? ISBN ???? LCCN ????

Kelley:1993:NCT

[KFF93a] E. F. Kelley, B. F. Field, and C. Fenimore. Nonlinear color transformations in real time using a video supercomputer. In Transforms and transportability of color: 1st Color imaging conference — November 1993, Scottsdale, AZ [KFF93b], pages 122–128. ISBN 0-89208-174-0. LCCN ????

Kelley:1993:TTC

[KFF93b] E. F. Kelley, B. F. Field, and C. Fenimore, editors. Transforms and transportability of color: 1st Color imaging conference — November 1993, Scottsdale, AZ. Society for Information Display and Society for Imaging Science and Technology, ????, 1993. ISBN 0-89208-174-0. LCCN ????

Komori:19xx:TSC

[KfGERJxx] Satoru Komori and Center for Global Environmental Research (Japan). Turbulence structure and CO2 transfer at the air-sea interface and turbulent diffusion in thermally-stratified flows. ???? ????, Center for Global Environmental Research, National Institute for Environmental Studies, Environment Agency of Japan, Tsukuba, Japan, 19xx. v + 51 pp.

Kelley:1994:DMA

[KFJB94] E. F. Kelley, B. F. Field, G. R. Jones, and P. A.

Boynton. Display modeling and an AMLCD model on a video supercomputer. Digest of technical papers, 25 (??):779–??, ???? 1994. ISSN 0097-966X (print), 2154-6738 (electronic), 2154-6746.

Khaleghzadeh:2020:NDP

[KFML20] Hamidreza Khaleghzadeh, Muhammad Fahad, Ravi Reddy Manumachu, and Alexey Lastovetsky. A novel data partitioning algorithm for dynamic energy optimization on heterogeneous highperformance computing platforms. Concurrency and Computation: Practice and Experience, 32(21):e5928:1– e5928:??, November 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kondo:2002:SCC

[KFN02] Masaaki Kondo, Motonobu Fujita, and Hiroshi Nakamura. Software-controlled on-chip memory for highperformance and low-power computing. ACM SIGARCH Computer Architecture News, 30(3):7–8, June 2002. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Koeninger:1994:SMM

[KFW94] R. Kent Koeninger, Mark Furtney, and Martin Walker. Shared memory MPP from Cray Research. Digital Technical Journal, 6(2):8– 21, Spring 1994. CODEN DTJOEL. ISSN 0898-901X.

Kannan:1994:TDI

[KG94] A. Kannan and T. V. Geetha.

Time dependent intelligent information management — systems. In Balakrishnan [Bal94], pages 85–92. ISBN 0-07-462044-4. LCCN ????

Kraske:1995:VAD

[KG95] Wolfgang F. Kraske and F. W. George III. VOXARall ATM distributed biomedical visualization: (1) local OC-3 linked workstation cluster (2) remote OC-3 linked 40 GFlops Cray T3D MPP. IEEE Symposium on Computer-Based Medical Systems, pages 249–257, 1995. CODEN PSCSFM. ISSN 1063-7125. IEEE catalog number 95CH35813.

Kaxiras:1996:GCC

[KG96] S. Kaxiras and J. R. Goodman. The GLOW cache coherence protocol extension for widely shared data. In ACM [ACM96], pages 35–43. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Karin:1998:HPC

[KG98] Sidney Karin and Susan Graham. The high-performance computing continuum. Communications of the ACM, 41 (11):32–35, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org:80/pubs/citations/
journals/cacm/1998-41-11/
p32-karin/.

Kohout:2003:HPC

[KG03] James Kohout and Alan D. George. A high-performance communication service for parallel computing on distributed DSP systems. Parallel Computing, 29(7):851– 878, July 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kleinrock:1996:SST

[KGB^{+}96] L. Kleinrock, M. Gerla, N. Bambos, J. Cong, E. Gafni, L. Bergman, J. Bannister, S. P. Monacos, T. Bujewski, P-C Hu, B. Kannan, B. Kwan, E. Leonardi, J. Peck, P. Palnati, and S. Walton. The supercomputer supernet testbed: A WDM-Based supercomputer interconnect. Journal of lightwave technology: a joint IEEE/OSA publication, 14 (6):1388–??, ???? 1996. ISSN 0733-8724 (print), 1558-2213 (electronic).

Khromov:1993:PMD

[KGKa93] V. V. Khromov, L. A. Goncharov, E. F. Kryuchkov, and G. V. and Tikhomirov. Probabilistic method of discrete ordinates (PMDO): Development and applications in global-local calculations. In Kusters et al. [KSW93], pages 605–616. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Kuo:1985:USN

[KGLA85] Franklin F. Kuo and Jose J.

Garcia-Luna-Aceves. USER-NET: a supercomputer network architecture. Future Generation Computer Systems, 1(3):161–168, February 1985. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Knudsen:1993:GCS

[KGS93] S. Knudsen, R. Guigo, and T. Smith. GeneID — a computer server for prediction of genes in DNA sequences. In Lim et al. [L^{+}93], pages 545– 554. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Kiefer:1987:CGI

[KH87] Dave Kiefer and John Heightley. Cray-3: A GaAs implemented supercomputer system. Technical Digest — GaAs IC Symposium (Gallium Arsenide Integrated Circuit), pages 3–6, 1987. CO-DEN TDGSEE. IEEE Service Cent. Piscataway, NJ, USA.

Kauranne:1993:PSA

[KH93] Tuomo Kauranne and G.-R.

Hoffmann, editors. Parallel supercomputing in atmospheric science: proceedings of the Fifth ECMWF Workshop on the Use of Parallel Processors in Meteorology, Reading, UK, November 23–27, 1992. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1993. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Kerry:1998:KIH

[KH98] K. E. Kerry and K. A. Hawick. Kriging interpolation on high-performance computers. Lecture Notes in Computer Science, 1401:429–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khan:1991:CSP

[Kha91] Fitratullah Khan. Chaos in spatial population dynamics: simulation on an IBM 3090 supercomputer and the development of an interactive graphics system. Thesis (M.S.), University of Kansas, Computer Science, Lawrence, KS, USA, 1991. ix + 95 pp.

Khan:1993:CSA

[Kha93] Shahin Khan. Cray S-MP architecture seamless vector and parallel processing in a SPARC environment. In 1993 IEEE Compcon Spring (Feb 22–26 1993: San Francisco, CA, USA), pages 605–607. IEEE, Piscataway, NJ, USA, 1993. ISBN 0-7803-1294-5. LCCN ???? IEEE catalog number 93CH3251-6.

Khan:1995:PDH

[Kha95] Zahira S. Khan. Performance of the distributed hash join algorithms in a distributed heterogeneous supercomputing environment. Thesis (Ph.D.), Temple University, Philadelphia, PA, USA, 1995. xxii + 220 pp.

Kielmann:2001:EJH

[KHBB01] Thilo Kielmann, Philip Hatcher, Luc Bougé, and Henri E. Bal. Enabling Java for high-performance computing. Communications of the ACM, 44(10): 110–117, October 2001. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kim:2014:ACT

[KHC14] Jongsung Kim, Bo Hong, and Naveen Chilamkurti. Advanced computer technologies and applications in Grid and cloud computing. The Computer Journal, 57 (2):181–182, February 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/content/57/2/181.full.
pdf+html.

Khermouch:1994:LC

[Khe94] G. Khermouch. Large computers. IEEE Spectrum, 31 (1):46–49, January 1994. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Khermouch:1995:LC

[Khe95] G. Khermouch. Large computers. IEEE Spectrum, 32 (1):48–51, January 1995. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karp:1995:SRG

[KHHS95] Alan H. Karp, Michael Heath, Don Heller, and Horst Simon. Special report: 1994 Gordon Bell Prize winners. Computer, 28(1): 68–74, January 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kulkarni:1994:CPP

[KHMD94] J. J. Kulkarni, G. Hari, R. S. Mundada, and P. S. Dhekne. Common platform for parallel and multiprocessing. In Mahajan et al. [M^{+}94], pages 382–388. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Kimura:1989:SDT

[KHN89] Fumitaka Kimura, Yasumasa Honjyo, and Akira Nakayama. Sequential-Type distance transformation by supercomputer. Systems and computers in Japan, 20(11): 90–??, November 1, 1989. CODEN SCJAEP. ISSN 0882-1666 (print), 1520-684X (electronic).

Khozeimeh:1994:CCE

[Kho94] Khalil Khozeimeh, editor.

Computing in civil engineering: proceedings of the First Congress held in conjunction with A/E/C Systems ’94 / sponsored by the Committee on Coordination Outside ASCE of the Technical Council on Computer Practices of the American Society of Civil Engineers, Washington, DC, June 20–22, 1994. American Society of Civil Engineers, New York, NY, USA, 1994. ISBN 0-7844-0026-1. LCCN TA345.C657 1994. Two volumes.

Kang:1988:FDA

[KHS88] Y. J. Kang, J. H. Herzog, and J. Spragins. FISHNET: a distributed architecture for high-performance local computer networks. IEEE Transactions on Computers, 37 (1):119–123, January 1988. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=75144.

Kaushik:1994:ACD

[KHSJ94] S. D. Kaushik, C.-H. Huang, P. Sadayappan, and R. W. Johnson. An approach to communication-efficient data redistribution. In Anonymous [Ano94-134], pages 364–373. ISBN ???? LCCN ????

Kristensen:2011:HPP

[KHV11] Mads Kristensen, Hans Happe, and Brian Vinter. Hybrid parallel programming for Blue Gene/P. Scalable Computing: Practice and Experience, 12(2):265–274, June 2011. CODEN ???? ISSN 1895-1767. URL http:/

                                        /www.scpe.org/index.php/
                                        scpe/article/view/719.

Kumar:2008:SMD

[KHZ^{+}08] S. Kumar, C. Huang, G. Zheng, E. Bohm, A. Bhatele, J. C. Phillips, H. Yu, and L. V. Kalé. Scalable molecular dynamics with NAMD on the IBM Blue Gene/L system. IBM Journal of Research and Development, 52(1/2):177– ??, January/March 2008. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

                                        //www.research.ibm.com/
                                        journal/rd/521/kumar.html.

Kim:1996:SPD

[Kim96] Hwa-Sung Kim. Scheduling of programs with diverse parallelism types in distributed heterogeneous supercomputing systems. Thesis (Ph.D.), Lehigh University, Bethlehem, PA, USA, 1996. xii, 174 pp.

Kindler:1996:DST

[Kin96] Thomas Paul Kindler. The development of supercomputing tools in a global atmospheric chemistry model and its application on selected problems in global atmospheric chemistry modeling. Thesis (Ph.D.), School of Earth and Atmospheric Sciences, Georgia Institute of Technology, Atlanta, GA, USA, 1996. xx + 207 pp.

Kirrmann:1989:MSR

[Kir89] H. Kirrmann. Multiprocessors and supercomputer research in Europe. IEEE Micro, 9(1):7–8, January/ February 1989. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kitai:1994:DSC

[KISY94] K. Kitai, T. Isobe, T. Sakakibara, and S. Yazawa. Distributed storage control unit for the Hitachi S-3800 multivector supercomputer. In Anonymous [Ano94-134], pages 1–10. ISBN ???? LCCN ????

Kightley:1985:CCG

[KJ85] J. R. Kightley and I. P. Jones.

Comparison of conjugate gradient preconditionings for three-dimensional problems on a Cray-1. Computer Physics Communications, 37 (1-3):205–214, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Kremer:1994:SPR

[KJ94] K. Kremer and K. F. A.

Juelich. Supercomputing in polymere research (invited paper). In Gentzsch and Harms [GH94a], pages 244– 253. CODEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994). DM96.00. Two volumes.

Kartashev:1982:DPM

[KK82] Svetlana P. Kartashev and Steven I. Kartashev. Designing and programming modern computers and systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1982. ISBN 0-13-201343-6 (vol. 1), 0-13-201435-1 (vol. 2). various pp. LCCN TK7885.D474 1982.

Kartashev:1985:SSP

[KK85] Svetlana P. Kartashev and Steven I. Kartashev, editors. Supercomputing systems: proceedings of the first international conference, St. Petersburg, Florida, U.S.A., December 16–20, 1985. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-444-87967-6. LCCN ????

Kartashev:1987:SPS

[KK87] Svetlana P. Kartashev and Steven I. Kartashev, editors. Supercomputing ’87: proceedings / Second International Conference on Supercomputing. International Supercomputing Institute, Inc., St. Petersburg, FL, USA, 1987. LCCN QA 76.5 I575 1987. Three volumes.

Kartashev:1988:SPN

[KK88] Svetlana P. Kartashev and Steven I. Kartashev, editors. Supercomputing ’88: proceedings / ICS 88, Third International Conference on Supercomputing. International Supercomputing Institute, Inc., St. Petersburg, FL, USA, 1988. LCCN QA 76.88 I58 1988. Three volumes. Spine title: ICS 88, Third International Conference on Supercomputing. Contents: v. 1. Supercomputing projects, applications, and artificial intelligence — v. 2. Technology assessment, industrial supercomputer outlooks, European supercomputing accomplishments, and performance and computations — v. 3. Supercomputer design, hardware and software.

Kartashev:1989:SSS

[KK89a] Svetlana P. Kartashev and Steven I. Kartashev, editors. Supercomputing ’89: supercomputing structures and computations: proceedings / Fourth International Conference on Supercomputing and Third World Supercomputer Exhibition, Santa Clara Convention Center, Santa Clara, CA, USA, April 30–May 5, 1989. International Supercomputing Institute, Inc., St. Petersburg, FL, USA, 1989. LCCN QA 76.5 I575 1989. Two volumes.

Kartashev:1989:SSR

[KK89b] Svetlana P. Kartashev and Steven I. Kartashev. Supercomputing systems: reconfigurable architectures, volume II of Designing and programming modern computer systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1989. ISBN 0-13-201435-1. xix + 428 pp. LCCN TK7885.D474 1982.

Klappholz:1989:CCF

[KK89c] David Klappholz and Xiangyun Kong. CFTP (Cray Fortran/ANSI FORTRAN ’77 prefine): a tool to aid in hand-parallelizing sequential code. Digest of Papers — IEEE Computer Society International Conference, pages 92–97, February 1989. CO-DEN DCSIDU. ISBN 0-8186-1909-0. Available from IEEE Service Cent. Piscataway.

Kartashev:1990:SSA

[KK90] Svetlana P. Kartashev and Steven I. Kartashev. Supercomputing systems: architectures, design, and performance. Van Nostrand Reinhold Co., New York, NY, USA, 1990. ISBN 0-442-25615-9. xxxiii + 622 pp. LCCN QA76.5 .S89841 1990.

Kuwahara:1992:PIW

[KK92] K. (Kunio) Kuwahara and Tsutomu Kambe, editors. Proceedings of the International Workshop on Supercomputing and Experiments in Fluid Dynamics: Nobeyama, Japan, 3– 5 September 1991, volume 10(4–6) of Fluid dynamics research. North-Holland, Amsterdam, The Netherlands, December 1992.

Keevallik:1993:ICP

[KK93] Sirje Keevallik and Olavi Karner, editors. IRS ’92: current problems in atmospheric radiation: proceedings of the International Radiation Symposium, Tallinn, Estonia, 3–8 August 1992. A. Deepak Pub., Hampton, VA, USA, 1993. ISBN 0-937194-28-X. LCCN QC912.3.I57 1992.

Khan:1995:PDP

[KK95a] Z. S. Khan and E. Kwatny.

Performance of a distributed pipelined multijoin algorithm in a distributed heterogeneous supercomputing environment. In Bailey et al. [B^{+}95], pages 289–290. ISBN 0-89871-344-7. LCCN QA76.58.S55 1995.

Knecht:1995:DLB

[KK95b] R. Knecht and G. A.

Kohring. Dynamic load balancing for the simulation of granular materials. In ACM [ACM95a], pages 164–169. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Karolyi:1996:LGS

[KK96a] A. Karolyi and J. Kertesz. Lattice gas simulations of granular media. In Wolf et al. [WSB96], pages 359– 366. ISBN 981-02-2635-7. LCCN ????

Krishnan:1996:APR

[KK96b] S. Krishnan and L. V. Kale. Automating parallel runtime optimizations using postmortem analysis. In ACM [ACM96], pages 221–228. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Kumar:1992:DGT

[KKB92] K. G. Kumar, D. Kulkarni, and A. Basu. Deriving good transformations for mapping nested loops on hierarchical parallel machines in polynomial time. In ACM [ACM92b], pages 82–92. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Kato:1997:AFC

[KKDO97] M. Kato, H. Kano, K. Date, and T. Oya. Analysis of flow with cavitation in diesel nozzle hole. In Roller [Rol97], pages 459–466. ISBN 0-947719-88-1 (paperback). LCCN ????

Koseki:1996:RAT

[KKF96] A. Koseki, H. Komatsu, and Y. Fukazawa. A register allocation technique using guarded PDG. In ACM [ACM96], pages 270–277. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Kel:1993:CAR

[KKKP93] A. E. Kel, N. A. Kolchanov, V. V. Kapitonov, and M. P. Ponomarenko. Computer analysis and recognition of functional sites via oligonucleotide pattern distributions. In Lim et al. [L^{+}93], pages 521–544. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Kolchanov:1993:POD

[KKPR93] N. A. Kolchanov, A. E. Kel, M. P. Ponomarenko, and A. G. Romachenko. Patterns of oligonucleotide distribution within DNA and RNA functional sites. In Lim et al. [L^{+}93], pages 445–464. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Kowalik:1999:EWC

[KL99] J. Kowalik and J. Lixvar. Editorial: a wakeup call for supercomputing. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-

2803. URL http://www.cs.

                                        okstate.edu/~pdcp/vols/
                                        vol02/vol02no4editorial.
                                        html.

Koniges:1995:PCR

[KLD95] Alice E. Koniges, Kevin R. Lind, and Paul Dubois. Parallelizing code for real applications on the T3D. Computers in Physics, 9(4): 399–??, July 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4823422.

Kocheisen:1994:HPN

[KLM94] M. Kocheisen, D. Lutziger, and U. A. Mueller. High performance neural net simulation on the MUSIC parallel supercomputer. In Halin et al. [HKR94], pages 367– 370. ISBN 1-56555-031-5. LCCN ????

Knecht:1990:PQDa

[KLN90a] S. Knecht, E. Laermann, and W. E. Nagel. Parallelizing QCD with dynamical fermions on a CRAY multiprocessor system. In Pitcher [Pit90], pages 193– 206. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Knecht:1990:PQDb

[KLN90b] Siegfried Knecht, Edwin Laermann, and Wolfgang E.

Nagel. Parallelizing QCD with dynamical fermions on a Cray multiprocessor system. Parallel Computing, 15(1–3): 3–20, September 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kothe:2019:ECU

[KLQ19] D. Kothe, S. Lee, and I. Qualters. Exascale computing in the United States. Computing in Science and Engineering, 21(1):17–29, January/February 2019. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Kim:1997:NST

[KLSC97] H. J. Kim, W. S. Lee, J. S.

Skim, and C. Choi. Numerical simulation of the thermal and fluid flow in engine room. In Roller [Rol97], pages 407– 412. ISBN 0-947719-88-1 (paperback). LCCN ????

Kao:1994:PIF

[KLY94] J. Kao, G. Li, and C. W.

Yang. Preconditioned iterative 3-D finite-difference migration or modeling on MPP systems. In IEEE [IEE94c], pages 601–606. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Kuba:1985:EML

[KM85] D. W. Kuba and K. J. M.

Moriarty. Efficient multitasking of the Su(3) lattice gauge theory algorithm on the Cray X-MP. Computer Physics Communications, 36 (4):351–362, June 1985. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465585900244.

Kohn:1989:III

[KM89] Les Kohn and Neal Margulis. Introducing the Intel i860 64-bit microprocessor. IEEE Micro, 9(4):15– 30, July/August 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kennedy:1992:OPD

[KM92] K. Kennedy and K. S. McKinley. Optimizing for parallelism and data locality. In ACM [ACM92b], pages 323–334. ISBN 0-89791-485-6 (paperback), 0-89791-486-

4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Kindratenko:2009:ITP

[KMB09] Volodymyr V. Kindratenko, Adam D. Myers, and Robert J. Brunner. Implementation of the two-point angular correlation function on a high-performance reconfigurable computer. Scientific Programming, 17(3): 247–259, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kieu:1996:LPI

[KMG96] T. D. (Tien D.) Kieu, B. H.

J. (Bruce H. J.) McKellar, and A. J. (Anthony J.) Guttmann, editors. Lattice 95: proceedings of the International Symposium on Lattice Field Theory, Melbourne, Australia, 11–15 July 1995, number 47 in Nuclear Physics B Proceedings Supplements 1996. North-Holland, Amsterdam, The Netherlands, 1996. ISBN ???? ISSN 0920-5632 (print), 1873-3832 (electronic). LCCN QC793.3.F5 I569 1995.

Katevenis:1997:TSH

[KMKD97] Manolis G. H. Katevenis, Evangelos P. Markatos, George Kalokerinos, and Apostolos Dollas. Telegraphos: a substrate for high-performance computing on workstation clusters. Journal of Parallel and Distributed Computing, 43 (2):94–108, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1997.1334/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1997.1334/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1997.1334/production/ref.

Kuksheva:2005:SSS

[KMN^{+}05] E. A. Kuksheva, V. E. Malyshkin, S. A. Nikitin, A. V. Snytnikov, V. N. Snytnikov, and V. A. Vshivkov. Supercomputer simulation of self-gravitating media. Future Generation Computer Systems, 21(5):749–757, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Keshk:1995:APS

[KMNT95] H. Keshk, S.-I. Mori, H. Nakashima, and S. Tomita. Amon: a parallel slice algorithm for wire routing. In ACM [ACM95a], pages 200–208. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Keshk:1996:APW

[KMNT96] H. Keshk, S.-I. Mori, H. Nakashima, and S. Tomita. Amon2: a parallel wire routing algorithm on a torus network parallel computer. In ACM [ACM96], pages 197–204. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Kommaraju:2020:SRL

[KMRR20] Ananda V. Kommaraju, Kristyn J. Maschhoff, Michael F. Ringenburg, and Benjamin Robbins. Scalable reinforcement learning on Cray XC. Concurrency and Computation: Practice and Experience, 32(20):e5636:1– e5636:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kobayashi:1994:CAC

[KMT94] T. Kobayashi, Y. Morinishi, and N. Tanaguchi. Chapter 7: Applications of LES to complicated flow fields. In Murthy and Brebbia [MB94b], pages 149–

168. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Karplus:1986:CDS

[KN86] Kevin Karplus and Alexandru Nicolau. A compilerdriven supercomputer. Applied Mathematics and Computation, 20(1–2):95–110, September 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0096300386901281.

Kampe:1988:PCC

[KN88] Frank C. Kampe and Tung M.

Nguyen. Performance comparison of the Cray-2 and Cray X-MP on a class of seismic data processing algorithms. Parallel Computing, 7(1):41–54, April 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Katouda:2016:MPA

[KNHN16] Michio Katouda, Akira Naruse, Yukihiko Hirano, and Takahito Nakajima. Massively parallel algorithm and implementation of RI-MP2 energy calculation for petascale many-core supercomputers. Journal of Computational Chemistry, 37 (30):2623–2633, November 15, 2016. CODEN JC-CHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Kennedy:1995:EAG

[KNS95] K. Kennedy, N. Nedeljkovic, and A. Sethi. Efficient address generation for blockcyclic distributions. In ACM [ACM95a], pages 180–184. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Kowalski:1997:TMC

[KNS97] Z. Kowalski, J. Nabrzyski, and M. Stroinski. Towards metacomputing — a case study of Poznan Supercomputing and Networking Center. Lecture Notes in Computer Science, ??(1225): 1016–??, ???? 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuo:1993:TMH

[KNWB93] H.-C. Kuo, A. Nilsson, D. Winklestein, and L. Bottomley. Traffic measurement on HIPPI links in a supercomputing environment. In Viniotis and Onvural [VO93], pages 199–224. ISBN 0-306-44486-0. LCCN TK5105 .A83 1993.

Kusakabe:1995:DLO

[KNYT95] S. Kusakabe, T. Nagai, Y. Yamashita, and R.-I. Taniguchi. A dataflow language with object-based extension and its implementation on a commercially available parallel machine. In ACM [ACM95a], pages 308– 317. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Karniadakis:1990:SSC

[KO90] G. Em Karniadakis and S. A.

Orszag. Spectral simulations of complex flows. In Pitcher [Pit90], pages 1–34. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Karniadakis:1993:NMF

[KO93a] George Em Karniadakis and Steven A. Orszag. Nodes, modes and flow codes. Physics Today, 46(3):34– ??, March 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Koshizuka:1993:CMT

[KO93b] S. Koshizuka and Y. Oka. A computational method for thermal-hydraulic analysis in three-dimensional complex geometries. In Kusters et al. [KSW93], pages 176– 187. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Koclas:1993:RKU

[Koc93] J. Koclas. Reactor kinetics using a three level time step hierarchy based on super nodal analysis. In Kusters et al. [KSW93], pages 581– 592. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Koeda:1996:OSV

[Koe96] Y. Koeda. Operating system of VX/VPP300/VPP700 series of vector-parallel supercomputer. Fujitsu, 47(6):442– ??, ???? 1996. CODEN FU-JTAR. ISSN 0016-2515.

Koeda:1997:OSV

[Koe97] Y. Koeda. Operating system of the VX/VPP300/VPP700 series of vector-parallel supercomputer systems. Fujitsu scientific and technical journal, 33(1):15–??, ???? 1997. CODEN FUSTA4. ISSN 0016-2523.

Kogge:1991:ASC

[Kog91] Peter M. Kogge. The architecture of symbolic computers. McGraw-Hill series in supercomputing and parallel processing. McGraw-Hill, New York, NY, USA, 1991. ISBN 0-07-035596-7. xxi + 739 pp. LCCN QA76.9.A73 K64 1991.

Kogge:2011:TF

[Kog11] P. Kogge. The tops in flops. IEEE Spectrum, 48(2): 48–54, February 2011. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kohring:1996:PDH

[Koh96] G. A. Kohring. Particle diffusion in a horizontal, rotating drum. In Wolf et al. [WSB96], pages 341–346. ISBN 981-02-2635-7. LCCN ????

Kokosinski:1994:MPG

[Kok94] Z. Kokosinski. Mask and pattern generation for associative supercomputing. In Hamza [Ham94], pages 324– 326. ISBN 0-88986-190-0. ISSN 0013-5704. LCCN ????

Kolodzey:1981:CCT

[Kol81] James S. Kolodzey. Cray-1 computer technology. IEEE transactions on components, hybrids, and manufacturing technology, CHMT-4(2):181– 186, June 1981. CODEN IT-TEDR. ISSN 0148-6411.

Konigsfeld:1987:ELA

[Kon87] Kris Gibson Konigsfeld. An Ethernet local area network controller with supporting operating system software. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 85 pp.

Konas:1991:PDEa

[Kon91a] Pavlos Konas. Parallel discrete event simulation on shared memory multiprocessors. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. ix + 79 pp.

Kondo:1991:SAA

[Kon91b] Jiro Kondo. Supercomputing: applications, algorithms, and architectures: for the future of supercomputing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. xvi + 218 + 4 pp. LCCN QA76.88 .S86 1991.

Konopka:1993:PCC

[Kon93] A. K. Konopka. Plausible classification codes and local compositional complexity of nucleotide sequences. In Lim et al. [L^{+}93], pages 69–88. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Konchady:1996:DSU

[Kon96] M. Konchady. Distributed supercomputing using ACTS. In IEEE [IEE96b], pages 172– 181. ISBN 0-8186-7582-9. ISSN 1082-8907. LCCN ????

Koornstra:1997:TFM

[Koo97] M. J. Koornstra. Trends and forecasts in motor vehicle kilometrage, road safety and environmental quality. In Roller [Rol97], pages 21–32. ISBN 0-947719-88-1 (paperback). LCCN ????

Kopetzky:1988:HSH

[Kop88] Daniel J. Kopetzky. Horse:

a simulation of the Horizon supercomputer. Technical report SRC-TR-88-013, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 2 pp.

Koppel:1991:PSS

[Kop91] Tom Koppel. Profile: Supercomputer solo. Scientific American, 264(3):34– 36, March 1991. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.
nature.com/scientificamerican/
journal/v264/n3/pdf/scientificamerican0391-
34.pdf.

Kopp:2000:MCC

[Kop00]  Carlo Kopp.  Managing

cluster computers. Dr. Dobb’s Journal of Software Tools, 25(7):21–26, July 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_07/cluster.txt.

Kortanek:1993:VSE

[Kor93] K. O. Kortanek. Vectorsupercomputer experiments with the primal affine linear programming scaling algorithm. SIAM Journal on Scientific Computing, 14(2): 279–294, March 1993. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Koss:1989:APS

[Kos89] Peter F. Koss. Application performance on supercomputers. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1989. iv + 50 pp.

Koski:1995:STL

[Kos95] Kimmo Koski. A step towards large scale parallelism: Building a parallel computing environment from heterogeneous resources. Future Generation Computer Systems, 11(4–5):491–498, August 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kowalik:1985:PMC

[Kow85] Janusz S. Kowalik. Parallel MIMD computation: the HEP supercomputer and its applications. The MIT Press series in scientific computation. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-11101-2. 411 pp. LCCN QA76.8.D436 P37 1985.

Kowalik:1986:BRB

[Kow86] Janusz Kowalik. Book review: Supercomputers and Parallel Computation (D. J. Paddon, ed.). SIAM Review, 28(1):109–111, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kowalik:1989:S

[Kow89a] Janusz S. Kowalik, editor.

Supercomputing, NATO ASI series. Series F, Computer and systems sciences; vol.

62. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-52691-9 (Berlin), 0-387-52691-9 (New York). LCCN QA76.5 .N344 1989. “Proceedings of the NATO Advanced Research Workshop on Supercomputing, held in Trondheim, Norway, June 19-23, 1989”–T.p. verso. “Published in cooperation with NATO Scientific Affairs Division.”.

Kowalik:1989:SPN

[Kow89b] Janusz S. Kowalik, editor.

Supercomputing: Proceedings of the NATO Advanced Research Workshop on Supercomputing, held in Trondheim, Norway, June 19–23, 1989, volume 62 of NATO ASI series. Series F, Computer and systems sciences.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-52691-9 (Berlin), 0-387-52691-9 (New York). LCCN QA76.5 .N344 1989.

King:1994:MDP

[KP94] P. D. King and M. E. Preston. Mechatronic design of piezoelectric-based small actuators. In Anonymous [Ano94-75], pages 107–114. ISBN 0-947719-68-7. LCCN ????

Keller:1995:CPS

[KP95] K. Keller and J. Poulton. Commercial packaging solutions for a research oriented graphics supercomputer. In Hsu et al. [HBCN95], pages 53–58. ISBN 0-7918-1303-7. LCCN TK7870.15 .I578 1995 v.1-2. Two volumes.

Kelly:1996:MCW

[KP96] W. Kelly and W. Pugh. Minimizing communication while preserving parallelism. In ACM [ACM96], pages 52–60. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Kale:1988:PEP

[KPS88] Laxmikant Vasudeo Kale, David A. Padua, and David C. Sehr. OR parallel execution of Prolog programs with side effects. Technical Report CSRD 740, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 12 pp.

Kennedy:1994:COS

[KR94a] K. Kennedy and G. Roth.

Context optimization for SIMD execution. In IEEE [IEE94c], pages 445–453. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Klingsporn:1994:NMD

[KR94b] M. Klingsporn and U. Renz.

Numerical modeling of diesel spray vaporization. In Anonymous [Ano94-75], pages 815–822. ISBN 0-947719-68-7. LCCN ????

Krafczyk:1994:SWC

[KR94c] M. Krafczyk and E. Rank.

Supercomputing on a workstation cluster: a parallelized lattice-gas solver for transient Navier–Stokes-flow. In Gentzsch and Harms [GH94a], pages 78–83. CO-DEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994). DM96.00. Two volumes.

Kwan:1994:PCS

[KR94d] T. T. Kwan and D. A.

Reed. Performance of the CM-5 scalable file system.

In Anonymous [Ano94-134], pages 156–165. ISBN ???? LCCN ????

Krause:1988:SSS

[Kra88] Karl Robert Krause. Simultaneous solution of systems of interlinked multistaged separators using the Cray X-MP supercomputer. Thesis (B.S.) in chemical engineering, University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA, 1988. 31 pp.

Kratzer:1990:MPS

[Kra90] Steven G. Kratzer. Massively parallel sparse-matrix computations. Technical report SRC-TR-90-008, Supercomputing Research Center: IDA, Lanham, MD, USA, February 1, 1990. 14 pp.

Kratzer:1992:SLF

[Kra92] Steven G. Kratzer. Sparse LU factorization on massively parallel SIMD computers. Technical report SRC-TR-92-072, Supercomputing Research Center: IDA, Lanham, MD, USA, April 16, 1992. 21 pp.

Kramer:1993:EBS

[Kra93] David Anthony Kramer. Efficient bit-parallel supercomputer architectures and algorithms. Thesis (Ph.D.), Princeton University, Princeton, NJ, USA, 1993. vi + 133 pp.

Kramer:1997:CSC

[Kra97] David Kramer. Computer sanctions ‘could affect US weapons research’. Nature, 386(6627):750, April 24, 1997. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Kranzlmuller:2001:NAS

[Kra01a] D. Kranzlmüller. Nondeterminism analysis on supercomputers and clusters. Parallel Processing Letters, 11 (2–3):251–??, 2001. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Kranzlmuller:2001:DMD

[Kra01b] Dieter Kranzlmüller. DeWiz — modular debugging for supercomputers and computational Grids. Lecture Notes in Computer Science, 2074:811– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2074/20740811.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2074/20740811.
pdf.

Kratzke:2020:VHC

[Kra20] Nane Kratzke. Volunteer down: How COVID-19 created the largest idling supercomputer on Earth. Future Internet, 12(6):98, June 06, 2020. CODEN ???? ISSN 1999-5903. URL https:/

                                        /www.mdpi.com/1999-5903/
                                        12/6/98.

Kremien:1995:SDS

[Kre95] O. Kremien. Scalability in distributed systems, parallel systems and supercomputers. Lecture Notes in Computer Science, 919(919): 532–??, ???? 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korzennik:1993:HMP

[KRJ93] S. G. Korzennik, E. J. Rhodes, and N. M. Johnson. Helioseismology on a massively parallel architecture: Reduction of 1024 by 1024 full-disk dopplergrams on Intel’s Touchstone Delta supercomputer. In Brown [Bro93], pages 461– 464. ISBN 0-937707-61-9. LCCN QB539.O83 G66 1993.

Kroj:1992:PSP

[Kro92] W. Kroj. Paralleles supercomputing Perspektiven für die 90er Jahre. In Meuer [Meu92c], pages 49–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Kazerouni:1994:DSP

[KRS94] L. Kazerouni, B. Rajan, and R. K. Shyamasundar. Derivation of systolic programs. In Mahajan et al. [M^{+}94], pages 350–358. ISBN 0-07-462240-

4. LCCN T385 .S37 1994. Rs387.00.

Krishnamoorthy:2013:SIJ

[KRS13] Sriram Krishnamoorthy, J. Ramanujam, and P. Sadayappan. A special issue of Journal of Parallel and Distributed Computing: Domainspecific languages and highlevel frameworks for highperformance computing. Journal of Parallel and Distributed Computing, 73(6): 895, June 2013. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S074373151300066X.

Kar:1994:VSA

[KRVJ94] S. Kar, Y. V. Rao, G. Valli, and D. G. Joshi. Vision system for automatic inspection of fuel pellets. In Mahajan et al. [M^{+}94], pages 284–292. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Kamath:1986:PMS

[KS86a] Chandrika Kamath and Ahmed Sameh. A projection method for solving nonsymmetric linear systems on multiprocessors. Technical Report CSRD-611, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 43 pp.

Kunkel:1986:OPS

[KS86b] S. R. Kunkel and J. E. Smith. Optimal pipelining in supercomputers. ACM SIGARCH Computer Architecture News, 14(2):404–411, June 1986. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Karin:1987:SE

[KS87a] Sidney Karin and Norris Parker Smith. The supercomputer era. Harcourt Brace Jovanovich, San Diego, CA, USA, 1987. ISBN 0-15-186787-9. x + 313 + 8 pp. LCCN QA76.5 .K356 1987.

Kuck:1987:SPE

[KS87b] David J. Kuck and Ahmed Sameh. A supercomputing performance evaluation plan. Technical Report CSRD 692, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 17 pp.

Kuehn:1988:HSS

[KS88] James T. Kuehn and Burton J. Smith. The Horizon supercomputing system: architecture and software. Technical report SRC-TR-88-009, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 7 pp.

Kumar:1990:SAT

[KS90] A. Kumar and K. R.

Swartzel. Supercomputing applications in thermal processing operations for foods. In Pitcher [Pit90], pages 81–90. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Kaufmann:1993:STS

[KS93a] William J. Kaufmann and Larry L. Smarr. Supercomputing and the transformation of science. Scientific American Library: Distributed by W.H. Freeman, New York, NY, USA, 1993. ISBN 0-7167-5038-4. xi + 238 pp. LCCN QA76.88 .K38 1993.

Kawamura:1993:LES

[KS93b] H. Kawamura and S. Satake. Large eddy simulation of turbulent flow in an annular tube. In Kusters et al. [KSW93], pages 248– 259. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Kessler:1993:CTN

[KS93c] R. E. Kessler and J. L.

Schwarzmeier. Cray T3D: a new dimension for Cray Research. In 1993 IEEE Compcon Spring (Feb 22–26 1993: San Francisco, CA, USA), pages 176–182. IEEE, Piscataway, NJ, USA, 1993.

ISBN 0-7803-1294-5. LCCN ???? IEEE catalog number 93CH3251-6.

Krishnamurthy:1994:OOT

[KS94a] M. V. Krishnamurthy and F. J. Smith. An object oriented tool for building knowledge based systems for quantitative problem solving in science and engineering. In Balakrishnan [Bal94], pages 119–129. ISBN 0-07-462044-4. LCCN ????

Kulkarni:1994:CCC

[KS94b] R. K. Kulkarni and Y. N. Srikant. C2G2: a complete code generator generator. In Balakrishnan [Bal94], pages 47–57. ISBN 0-07-462044-4. LCCN ????

Keifer:1995:IOC

[KS95] D. Keifer and V. W. Swanson. Implementation of optical clock distribution in a supercomputer. In Anonymous [Ano95-38], pages 260– 262. ISBN 1-55752-389-4, 1-55752-390-8. LCCN ????

Kurte:2019:PAO

[KSB^{+}19] Kuldeep Kurte, Jibonananda Sanyal, Anne Berres, Dalton Lunga, Mark Coletti, Hsiuhan Lexie Yang, Daniel Graves, Benjamin Liebersohn, and Amy Rose. Performance analysis and optimization for scalable deployment of deep learning models for country-scale settlement mapping on Titan supercomputer. Concurrency and Computation: Practice and Experience, 31(20):e5305:1– e5305:??, October 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kumar:1994:SMR

[KSH94] B. Kumar, P. Sadayappan, and C.-H. Huang. On sparse matrix reordering for parallel factorization. In Anonymous [Ano94-134], pages 431–438. ISBN ???? LCCN ????

Kharche:2008:SCE

[KSM^{+}08] Sanjay Kharche, Gunnar Seemann, Lee Margetts, Joanna Leng, Arun V. Holden, and Henggui Zhang. Simulation of clinical electrophysiology in 3D human atria: a high-performance computing and high-performance visualization application. Concurrency and Computation: Practice and Experience, 20 (11):1317–1328, August 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kurth:2019:TSP

[KSM^{+}19] Thorsten Kurth, Mikhail Smorkalov, Peter Mendygral, Srinivas Sridharan, and Amrita Mathuriya. Tensor-Flow at scale: Performance and productivity analysis of distributed training with Horovod, MLSL, and Cray PE ML. Concurrency and Computation: Practice and Experience, 31(16):e4989:1– e4989:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Khajeh-Saeed:2013:DNS

[KSP13] Ali Khajeh-Saeed and J. Blair Perot. Direct numerical simulation of turbulence using GPU accelerated supercomputers. Journal of Computational Physics, 235(??): 241–257, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0021999112006547.

Kerridge:1994:UML

[KSTB94] S. R. Kerridge, A. J. Slade, P. Thomas, and A. F. Bokma. Using machine learning for predictive maintenance in a condition monitoring system. In Anonymous [Ano94-75], pages 677–684. ISBN 0-947719-68-7. LCCN ????

Kerridge:1994:QMI

[KSTF94] S. R. Kerridge, A. J. Slade, P. Thomas, and R. Fuertes. Qualitative modelling for integrated reliability analysis in manufacturing. In Anonymous [Ano94-75], pages 67–

74. ISBN 0-947719-68-7.

LCCN ????

Kusters:1993:PJI

[KSW93] H. Kusters, E. Stein, and W. Werner, editors. Proceedings of the Joint International Conference on Mathematical Methods and Supercomputing in Nuclear Applications: M & C + SNA ’93, April 19–23, 1993, Congress and Exhibition Centre, Karlsruhe, Germany. Kernforschungzentrum Karlsruhe, Karlsruhe, Germany, 1993. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Kozdrowicki:1980:SGV

[KT80] E. W. Kozdrowicki and Douglas J. Theis. Second generation of vector supercomputers. Computer, 13(11): 71–83, November 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Korolev:1993:FDO

[KT93a] S. V. Korolev and V. G.

Tumanyan. Fractal dimensions of oligonucleotide compositions of DNA sequences. In Lim et al. [L^{+}93], pages 635–648. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Kume:1993:NSD

[KT93b] E. Kume and A. Takanishi. Numerical simulation for design of biped locomotion robots. In Kusters et al. [KSW93], pages 408– 419. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Kartsounis:1994:ACM

[KT94] D. Kartsounis and N. Tsabourakis. Advanced calibration and measurement techniques for high accuracy optical surface inspection. In Anonymous [Ano94-75], pages 257–264. ISBN 0-947719-68-7. LCCN ????

Kindratenko:2011:THP

[KT11] Volodymyr Kindratenko and Pedro Trancoso. Trends in high-performance computing. Computing in Science and Engineering, 13 (3):92–95, May/June 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kindratenko:2008:HPC

[KTG08] Volodymyr Kindratenko, George K. Thiruvathukal, and Steven Gottlieb. High-performance computing applications on novel architectures. Computing in Science and Engineering, 10(6):13–15, November/December 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kiker:1994:DSC

[KTK94] E. Kiker, F. Trzan, and Z. Klenovaek. Design solutions of the clamping device of CNC milling machine. In Anonymous [Ano94-75], pages 647–652. ISBN 0-947719-68-7. LCCN ????

Katoh:1993:AVR

[KTKK93] T. Katoh, K. Tanaka, Y. Kasai, and K. Kimura. Application of the virtual reality technology to the 3D-CAD system for nuclear plants. In Kusters et al. [KSW93], pages 654–659. ISBN 3-923704-11-

9. LCCN ???? Two volumes.

Kodama:2014:SRM

[KTN^{+}14] Chihiro Kodama, Masaaki Terai, Akira T. Noda, Yohei Yamada, Masaki Satoh, Tatsuya Seiki, Shin ichi Iga, Hisashi Yashiro, Hirofumi Tomita, and Kazuo Minami. Scalable rank-mapping algorithm for an icosahedral grid system on the massive parallel computer with a 3-d torus network. Parallel Computing, 40(8):362–373, August 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819114000659.

Kugo:1993:DIR

[KTNM93] T. Kugo, K. Tsuchihashi, M. Nakagawa, and T. Mori. Development of intelligent reactor design system IRDS. In Kusters et al. [KSW93], pages 199–209. ISBN 3-923704-11-

9. LCCN ???? Two volumes.

Kuck:1987:PST

[Kuc87] David J. Kuck. Parallel supercomputing today and the Cedar approach. Technical Report CSRD 652, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 22 pp.

Kuehn:1987:HAR

[Kue87] James T. Kuehn. Horizon assembler reference manual: with appendix, support software manual set. Technical report SRC-TR-87-004, Supercomputing Research Center: IDA, Lanham, MD, USA, 1987. 29 + [8] pp.

Kuebler:1992:P

[Kue92] F.-D. Kuebler. Positionspapiere. In Meuer [Meu92c], pages 240–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Kuehnapfel:1993:DPS

[Kue93] U. Kuehnapfel. Distributed processing in a STEP based multidisciplinary dynamics simulation environment, using KISMET as a realtime display processor. In Kusters et al. [KSW93], pages 795– 796. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Kulkarni:1994:MPL

[Kul94] S. Kulkarni. masterPlan — a plant layout design and walkthrough package (virtual reality). In Mahajan et al. [M^{+}94], pages 67–74. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Kumar:1991:DHP

[Kum91] Manish Kumar. Design of a high performance supercomputing node. Thesis (M.S.), Florida Institute of Technology, Melbourne, FL, USA, 1991. vii + 103 pp.

Kumar:1994:HIA

[Kum94] V. Kumar. A host interface architecture for HIPPI. In IEEE [IEE94c], pages 142– 149. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Kung:1984:SSW

[Kun84] S. Kung. On supercomputing with Systolic/Wavefront array processors. Proceedings of the IEEE, 72:867–884, 1984. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Kunert:1995:CSD

[Kun95] Del Kunert. Commercial supercomputer distributes data to cockpit. National defense (Washington), 80(509): 62–??, July 1, 1995. CODEN NTDFA2. ISSN 0092-1491.

Kuwahara:1992:FSS

[Kuw92] K. Kuwahara. Flow simulation on supercomputers and its visualization. International Journal of High Speed Computing, 4(1):49–??, March 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Kuwahara:1994:VCF

[Kuw94] K. Kuwahara. Visualization of computed flow structure: Flow simulation on supercomputers and its visualization. Denshi Joho Tsushin Gakkai shi = The journal of the Institute of Electronics, Information, and Communication Engineers, 77(7):752– ??, ???? 1994. ISSN 0913-5693.

Kim:1996:CAA

[KV96] J.-H. Kim and N. H. Vaidya. A cost-comparison approach for adaptive distributed shared memory. In ACM [ACM96], pages 44–51. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Knoesche:1995:RBA

[KVP95] T. Knoesche, M. Van Burik, and M. Peters. Reconstruction of brain activity from EEG. In Herrmann et al. [HWP95], pages 199– 204. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Kahaner:1992:SCU

[KW92] David K. Kahaner and Ulrich Wattenberh. The speed champion in uniprocessors, Japan is coming on strong in multiprocessing. IEEE Spectrum, 29(9):42– 47, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kohn:1993:A

[KW93] James Kohn and Winifred Williams. ATExpert. Journal of Parallel and Distributed Computing, 18(2): 205–222, June 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1057/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1057/production/pdf.

Koetter:1995:EES

[KW95] R. Koetter and J. Wickens. Effects of an excitatory synaptic conductance and a slow membrane potassium conductance in an inhibitory striatal network model. In Herrmann et al. [HWP95], pages 257–264. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Kubert:2011:USL

[KW11] Roland Kübert and Stefan Wesner. Using service level agreements in a highperformance computing environment. Scalable Computing: Practice and Experience, 12(2):164–177, June 2011. CODEN ???? ISSN 1895-1767. URL http:/

/www.scpe.org/index.php/
scpe/article/view/712.

Kindratenko:2010:HPC

[KWB^{+}10] Volodymyr Kindratenko, Robert Wilhelmson, Robert Brunner, Todd J. Martinez, and Wen mei Hwu. Highperformance computing with accelerators. Computing in Science and Engineering, 12 (4):12–16, July/August 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kanai:1994:ISS

[KWH94] H. Kanai, N. Wakabayashi, and N. Honda. An integrated supporting system for mechanical design and manufacturing process of NC machining using object oriented methodology. In Anonymous [Ano94-75], pages 335–342. ISBN 0-947719-68-7. LCCN ????

Kwok:1987:PAA

[Kwo87] Alex Yuen-Wai Kwok. A performance analysis of architectural scalability. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. xi + 220 pp.

Kompass:1992:TEO

[KWW92] E. J. Kompass, S. K. Whitlock, and T. J. Williams, editors. Technology for empowering the operator: making the new human organizations succeed in the CIM factory: proceedings of the eighteenth annual Advanced Control Conference, West Lafayette, Indiana, September 14–16, 1992. Purdue University, West Lafayette, IN, USA, 1992. ISBN 0-931682-34-7. LCCN ????

Kotoh:1990:AFS

[KY90] S. Kotoh and G. Yamanaka.

Air flow simulation in a residential room by a supercomputer. In Pitcher [Pit90], pages 113–124. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Konas:1991:PDEb

[KY91a] Pavlos Konas and Pen-Chung Yew. Parallel discrete event simulation on shared-memory multiprocessors. Technical Report CSRD 1079, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 15 pp.

Konas:1991:SPD

[KY91b] Pavlos Konas and Pen-Chung Yew. Synchronous parallel discrete event simulation on shared-memory multiprocessors. Technical Report CSRD 1159, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1991. 23 pp.

Kaburaki:1993:NSN

[KY93] H. Kaburaki and M. Yokokawa. Numerical simulation of near continuum gas flows by the direct simulation Monte Carlo method. In Kusters et al. [KSW93], pages 260– 270. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Kanada:1988:TS

[KYS88] Y. Kanada, C. K. Yuen, and Nihon Gakujutsu Shinkokai, editors. Trends in supercomputing. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1988. ISBN 9971-5-0831-1. LCCN QA76.5 .T691 1988. Proceedings of a seminar on supercomputing held at the Computer Centre of the University of Tokyo, 1988, sponsored by the Japan Society for the Promotion of Science.

Kuster:1994:IFF

[KZ94] U. Kuster and M. Zurn. Influence of Fortran 90 features on performance on Cray vector computer systems. Lecture Notes in Computer Science, 797:475–483, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1993:BSC

[L^{+}93] Hwa A. Lim et al., editors. The Second International Conference on Bioinformatics, Supercomputing, and Complex Genome Analysis: proceedings of the June 4–7, 1992 Conference at St. Petersburg Beach, Florida, USA. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1993. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Lyrintzis:1995:CAP

[L^{+}95] A. S. Lyrintzis et al., editors.

Computational aeroacoustics, 1995: presented at the 1995 ASME/JSME Fluids Engineering and Laser Anemometry Conference and Exhibition, August 13-18, 1995, Hilton Head, South Carolina, volume 219 of ASME -Publications- FED 1995. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1995. ISBN 0-7918-1474-2. LCCN TL574.N6C66 1995.

Luthi:1993:NSD

[LA93] H. P. Luthi and J. Almlof.

Network supercomputing: a distributed-concurrent direct SCF scheme. Theoretica Chimica Acta, 84(4):443–??, January 1993. CODEN TCHAAM. ISSN 0040-5744.

Ledeczi:1994:PAF

[LA94] A. Ledeczi and B. Abbott.

Parallel architectures with flexible topology. In IEEE [IEE94c], pages 271–276. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Laguna:2015:DHP

[LAdS^{+}15] Ignacio Laguna, Dong H. Ahn, Bronis R. de Supinski, Todd Gamblin, Gregory L. Lee, Martin Schulz, Saurabh Bagchi, Milind Kulkarni, Bowen Zhou, Zhezhe Chen, and Feng Qin. Debugging high-performance computing applications at massive scales. Communications of the ACM, 58(9):72– 81, September 2015. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://cacm.
acm.org/magazines/2015/
9/191185/fulltext.

Lagana:1989:SAR

[Lag89] Antonio Lagana, editor. Supercomputer algorithms for reactivity, dynamics, and kinetics of small molecules: Proceedings of the NATO Advanced Research Workshop on Supercomputer Algorithms for Reactivity, Dynamics, and Kinetics of Small Molecules, Colombella di Perugia, Italy, 30 August– 3 September 1988, volume 277 of NATO ASI series. Series C, Mathematical and physical sciences; vol. 277. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1989. ISBN 0-7923-0226-5 (U.S.). LCCN QD505.5 .N37 1988. “Proceedings of the NATO Advanced Research Workshop on Supercomputer Algorithms for Reactivity, Dynamics, and Kinetics of Small Molecules, Colombella di Perugia, Italy, 30 August– 3 September 1988”–T.p. verso. “Published in cooperation with NATO Scientific Affairs Division.”.

Lobosco:2002:JHP

[LAL02] M. Lobosco, C. Amorim, and O. Loques. Java for highperformance network-based computing: a survey. Concurrency and Computation: Practice and Experience, 14 (1):1–31, January 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/91014114/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=91014114&PLACEBO=IE.
pdf.

Lamonica:2014:SCP

[Lam14] M. Lamonica. Supercomputer cooling for photovoltaics [news]. IEEE Spectrum, 51(11):22, November 2014. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Langhammer:1992:PCA

[Lan92] F. Langhammer. Performance considerations of applications on second generation parallel computers. In Meuer [Meu92c], pages 173–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Lang:1993:ICS

[Lan93] U. Lang. Integration of CFD supercomputer simulations and visualization using high speed networks. In Kusters et al. [KSW93], pages 782–?? ISBN 3-923704-11-9. LCCN ???? Two volumes.

Landolt:1994:ANF

[Lan94] O. Landolt. Arbitrary nonlinear functions synthesis in analog CMOS technology. In Anonymous [Ano94-75], pages 211–218. ISBN 0-947719-68-7. LCCN ????

Lang:1994:SVS

[LAPR94] U. Lang, H. Aichele, H. Poehlmann, and R. Ruehle. Scientific visualization in a supercomputer network. In Grave et al. [GLH94], pages 3–9. ISBN 3-540-56147-1, 0-387-56147-1. LCCN T385 .V59 1994.

Larson:1984:MCX

[Lar84] John L. Larson. Multitasking on the Cray X-MP-2 multiprocessor. Computer, 17 (7):62–69, July 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lasaga:1992:IMM

[Las92] Antonio C. Lasaga. Ab initio methods in mineral surface reactions. Reviews of geophysics, 30(4):269–??, November 1, 1992.

Lathrop:2016:CAP

[Lat16]  Scott Lathrop.  A call

to action to prepare the high-performance computing workforce. Computing in Science and Engineering, 18(6):80–83, November/ December 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2016/06/mcs2016060080-
abs.html.

Lavery:1989:DHP

[Lav89] Daniel Michael Lavery. The design of a hardware performance monitor for the Cedar supercomputer. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA, May 1989. xii + 110 pp.

Lawrence:1989:BPC

[Law89] David D. Lawrence. Bootstrapping of prostate cancer survival data with Cox’s proportional hazards model on the Cray supercomputer. Thesis (M.S.), State University of New York at Buffalo, Buffalo, NY, USA, 1989. v + 81 pp.

Lawson:1990:ESD

[Law90] M. Lawson. Engine structural dynamics for aircraft safety on the CRAY. In Pitcher [Pit90], pages 265– 271. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Lawton:2000:TND

[Law00] George Lawton. Technology news: Distributed net applications create virtual supercomputers. Computer, 33 (6):16–20, June 2000. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //dlib.computer.org/co/
                                        books/co2000/pdf/r6016.
                                        pdf.

Layman:1991:CSA

[Lay91a] Mary Layman. Computational science at the San Diego Supercomputer Center: a bibliography. Technical Report ????, San Diego Supercomputer Center, San Diego, CA, USA, 1991. vi + 92 pp. This bibliography includes the approximately 3600 citations received at SDSC as of December 1990.

Layman:1991:LSS

[Lay91b] Mary Layman. Library services in a supercomputer center. RSR: Reference Services Review, 19(3):45–??, ???? 1991. ISSN 0090-7324.

Loveluck:1982:CSC

[LB82] J. M. Loveluck and E. Balcar. Computer simulation calculations for onedimensional magnetic systems, with some remarks on optimisation on the Cray computer. Computer Physics Communications, 27(4):335– 350, October 1982. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Letovsky:1993:CRP

[LB93] S. Letovsky and M. B.

Berlyn. CPROP: a rulebased program for constructing genetic maps. In Lim et al. [L^{+}93], pages 149–168. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Lain:1994:TOC

[LB94a] A. Lain and P. Banerjee.

Techniques to overlap computation and communication in irregular iterative applications. In Anonymous [Ano94-134], pages 236–245. ISBN ???? LCCN ????

Lalor:1994:ASR

[LB94b] N. Lalor and T. Bharj. The application of SEA to the reduction of passenger car interior noise. In Anonymous [Ano94-75], pages 343–350. ISBN 0-947719-68-7. LCCN ????

Lomdahl:1994:SPC

[LB94c] Peter S. Lomdahl and David M. Beazley. State-of-the-Art parallel computing — molecular dynamics on the Connection Machine. Los Alamos Science, 22:44–??, 1994. CO-DEN LASCDI. ISSN 0273-7116.

Lain:1996:CSH

[LB96] A. Lain and P. Banerjee. Compiler support for hybrid irregular accesses on multicomputers. In ACM [ACM96], pages 1–9. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Liang:1994:PEF

[LBT94] C. Liang, S. Bhattacharya, and J. Tan. Performance evaluation of fault-tolerant routing on star networks. In IEEE [IEE94c], pages 650– 657. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Lee:1990:NCS

[LC90] L. Lee and S. Christensen. The North Carolina Supercomputing Center: a study of economic development impact. International Journal of Supercomputer Applications, 4(4):3–??, Fall 1990. CODEN IJSAE9. ISSN 0890-2720.

Lim:1991:SAAa

[LC91] Swee Boon Lim and Michael W.

Condry. Supercomputing application access characteristics. Technical Report UIUCDCS-R-91-1708; UILU-ENG-91-1755, Dept. of Computer Science, University of Illinois at Urbana-Champaign, Urbana, IL, USA, October 1991. vi + 57 pp.

Lacroix:1993:PMH

[LC93] B. Lacroix and J.-J. Codani. Physical mapping of the Human Genome: Computational aspects. In Lim et al. [L^{+}93], pages 135–148. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Lee:1994:CTM

[LC94] S. C. Lee and D. Chen.

Chapter 9: Turbulence modelling for transition studies. In Murthy and Brebbia [MB94b], pages 185– 214. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Lehmann:1995:NAS

[LC95] C. Lehmann and M. Cottrell.

A new approach to stochastic integrate-and-fire neurone modelling. In Herrmann et al. [HWP95], pages 389– 396. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Lauria:1997:MFH

[LC97a] Mario Lauria and Andrew Chien. MPI-FM: High performance MPI on workstation clusters. Journal of Parallel and Distributed Computing, 40(1):4–18, January 10, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1264/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1264/production/pdf;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.1264/production/ref.

Li:1997:PFD

[LC97b] Y. Li and J. S. Chang. A practical finite difference scheme for advection on a sphere. In Delic and Wheeler [DW97], pages 31–46. ISBN 0-89871-378-1. LCCN ????

Lin:2012:REC

[LC12] Yi-Kuei Lin and Ping-Chen Chang. Reliability evaluation of a computer network in cloud computing environment subject to maintenance budget. Applied Mathematics and Computation, 219(8):3893–3902, December 15, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0096300312010259.

Lyu:1997:IPS

[LCD97] JrJung Lyu, Fu-Shine Peter Chou, and Jyh-Hong Ding. Improving the performance of a supercomputer system using quality engineering and simulation. Simulation, 69 (3):143–??, ???? 1997. CO-DEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic).

Lin:1987:ISA

[LCH87] W. T. Lin, C. Y. Chin, and C. Y. Ho. Integrating systolic arrays into a supersystem. Computer, 20(7): 100–101, July 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liddell:1996:HPC

[LCHS96] Heather Mary Liddell, A. Colbrook, B. Hertzberger, and P. Sloot, editors. Highperformance computing and networking: international conference and exhibition, HPCN EUROPE 1966, Brussels, Belgium, April 15–19, 1996: proceedings, volume 1067 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CO-DEN LNCSD9. ISBN 3-540-61142-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H52 1996.

Lindtjorn:2011:BTM

[LCP^{+}11] Olav Lindtjorn, Robert Clapp, Oliver Pell, Haohuan Fu, Michael Flynn, and Oskar Mencer. Beyond traditional microprocessors for geoscience highperformance computing applications. IEEE Micro, 31 (2):41–49, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:1990:CAP

[LCV90a] Ding Lee, A. S. Cakmak, and Robert Vichnevetsky, editors. Computational acoustics: proceedings of the 2nd IMACS Symposium on Computational Acoustics, Princeton, NJ, USA, 15–17 March, 1989. Elsevier Science Publishers, Amsterdam, The Netherlands, 1990. ISBN 0-444-88723-7 (set), 0-444-88720-2 (vol. 1). LCCN QC242.8.I54 1989. Three volumes.

Lee:1990:CAO

[LCV90b] Ding Lee, Ahmet S. Cakmak, and Robert Vichnevetsky, editors. Computational acoustics: ocean-acoustic models and supercomputing: proceedings of the 2nd IMACS Symposium on Computational Acoustics, Princeton, NJ, USA, 15-17 March, 1989.

North-Holland, Amsterdam, The Netherlands, 1990. ISBN 0-444-88723-7, 0-444-88720-2 (v.1), 0-444-88721-0 (v.2). LCCN QC242.8.I54 1989. Two volumes.

Lonsdale:1993:CSM

[LCVR93] G. Lonsdale, J. Clinckemaillie, S. Vlachoutsis, and J. F. De Ronde. 93SC006 crashworthiness simulation migration to distributed memory, MIMD machines. In Anonymous [Ano93-31], pages 237–244. ISBN 0-947719-62-8. LCCN ????

Lakshmivarahan:1990:ADP

[LD90] S. Lakshmivarahan and Sudarshan Kumar Dhall. Analysis and design of parallel algorithms: arithmetic and matrix problems. McGraw-Hill series in supercomputing and parallel processing. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-036139-8. xviii + 657 pp. LCCN QA76.642 .L35 1990.

Labat:1993:SOS

[LD93a] I. Labat and R. Drmanac.

Simulations of ordering and sequence reconstruction of random DNA clones hybridized with a small number of oligomeric probes. In Lim et al. [L^{+}93], pages 555–566. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Labeau:1993:SDP

[LD93b] P. E. Labeau and J. Devooght. Synthesis of distributions for probabilistic reactor dynamics. In Kusters et al. [KSW93], pages 303– 315. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Luding:1996:SGF

[LDMC96] S. Luding, J. Duran, T. Mazozi, and E. Clement. Simulations of granular flow: Cracks in a falling sandpile. In Wolf et al. [WSB96], pages 305– 310. ISBN 981-02-2635-7. LCCN ????

Leatherdale:2009:PPS

[Lea09]  Dik Leatherdale.  Pio-

neer profiles — Seymour Cray. Resurrection: The Bulletin of the Computer Conservation Society, ??(47): ??, Sumer 2009. ISSN 0958-7403. URL https://

                                        computerconservationsociety.
                                        org/resurrection/res47.
                                        htm#f.

Lee:1984:LDH

[Lee84] Stewart Lee. Low dielectric, high temperature substrate for the supercomputer. Technical paper IPC-TP-503, IPC, Evanston, IL, USA, 1984. 11 pp.

Lee:1986:SIG

[Lee86] Gyungho Lee. Some issues in general-purpose shared memory multiprocessing: Parallelism exploitation and memory access combining. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. v + 123 pp.

Lee:1987:LNN

[Lee87a] Daeshik Lee. The lambda network: a new permutation network. Technical Report CSRD 627, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 3 pp.

Lee:1987:ECD

[Lee87b] Roland L. Lee. The effectiveness of caches and data prefetch buffers in largescale shared memory multiprocessors. Technical Report CSRD 670; UILU-ENG-87-8005, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 140 pp.

Lee:1989:BMC

[Lee89] Lawrence A. Lee. Beyond megaflops: The Cornell National Supercomputer Facility. Educom review, 24(1): 38–??, Spring 1989. CODEN EDREEW. ISSN 1045-9146.

Lee:1990:RSP

[Lee90] Chia-Ling Lee. On runtime systems for parallel super-computers. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1990. viii + 82 pp.

Lee:1994:CRD

[Lee94] J. W. Lee. Concord: Rethinking the division of labor in a distributed shared memory system. In IEEE [IEE94c], pages 585–592. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Lee:1996:CSS

[Lee96] Chang-Ock Lee. A comparison of some standard elliptic solvers: CM-5 vs. Cray C-90. Journal of Scientific Computing, 11(2): 99–126, June 1996. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/BF02088819;   http:
//link.springer.com/content/
pdf/10.1007/BF02088819;
 http://www.springerlink.
com/openurl.asp?genre=article&
issn=0885-7474&volume=11&
issue=2&spage=99-126.

Leghart:1990:ST

[Leg90] Frank Leghart. Supercomputer technology. Computer Technology Research Corp., 6 N. Atlantic Wharf, Charleston, SC 29401-2150, USA, second edition, 1990. ISBN 0-927695-25-1. iii + 156 pp. LCCN MLCM 92/08627.

Legge:1994:IDI

[Leg94] D. I. Legge. Integration of design, inspection and quality management systems. In Anonymous [Ano94-75], pages 421–428. ISBN 0-947719-68-7. LCCN ????

Leiserson:1985:FTU

[Lei85]  C. E. Leiserson.  Fat-

trees: Universal networks for hardware-efficient supercomputing. IEEE Transactions on Computers, C-34(10):892– 901, October 1985. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Leiserson:1989:VTP

[Lei89] Charles Eric Leiserson. VLSI theory and parallel supercomputing. Technical Report MIT/LCS/TM; 402, Laboratory for Computer Science, Massachusetts Institute of Technology, Cambridge, MA, USA, May 25, 1989. 14 pp.

Leiserson:1991:VTP

[Lei91] C. E. Leiserson. VLSI theory and parallel supercomputing. In Rashid [Ras91], pages 29– 44. ISBN 0-201-52899-1. LCCN QA75.5 .C548 1990.

Luetkenhoener:1995:EMS

[LEMS95] B. Luetkenhoener, T. Elbert, E. Menninghaus, and O. Steinstraeter. Electro- and magnetoencephalographic source analysis: Current status and future requirements. In Herrmann et al. [HWP95], pages 175–192. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Leung:1990:IDP

[Leu90] Bruce P. Leung. Issues on the designs of parallelizing compilers. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1990. iv + 52 pp.

Leutzbach:1996:SRH

[Leu96] W. Leutzbach. Some remarks on the history of the science of traffic flow. In Wolf et al. [WSB96], pages 3–10. ISBN 981-02-2635-7. LCCN ????

Levine:1982:S

[Lev82] R. D. Levine. Supercomputers. Scientific American, 246 (1):112–125, January 1982. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Levin:1989:RGS

[Lev89]  E. Levin.  The role of

graphics super-workstations in a supercomputing environment. NASA contractor report NASA CR-188852, Research Institute for Advanced Computer Science, NASA Ames Research Center, Moffett Field, CA, USA, 1989. ?? pp. Also issued as RIACS technical report 89.37.

Levin:1998:RAS

[Lev98] Stewart A. Levin. Remark on Algorithm 622: a simple macroprocessor. ACM Transactions on Mathematical Software, 24(3):336–340, September 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/toms/1998-24-3/

p336-levin/. See [RRW84].

Lewin:1993:SSP

[Lew93] David I. Lewin. Supercomputing ’93 showcases parallel computing finalists. Computers in Physics, 7(6):614– ??, November 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4823228.

Lewin:1994:INF

[Lew94a] David I. Lewin. Interview:

NSF forsees bright future for its supercomputing centers. Computers in Physics, 8(2): 134–136, March 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Lewin:1994:NDC

[Lew94b] David I. Lewin. NCSA’s director champions broad access to supercomputers. Computers in Physics, 8 (5):496–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

                                        //aip.scitation.org/doi/
                                        10.1063/1.4823325.

Lewin:1994:NFB

[Lew94c] David I. Lewin. NSF foresees bright future for its supercomputing centers. Computers in Physics, 8(2):134– 136, March 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4823271.

Lewin:1996:ICM

[Lew96a] David I. Lewin. INTER-VIEW: Charles Mosher striking pay dirt in a supercomputer. Computers in Physics, 10(6):510–??, ???? 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Lewin:1996:SPD

[Lew96b] David I. Lewin. Striking pay dirt in a supercomputer. Computers in Physics, 10 (6):510–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.4822491.

Lewis:2017:PPR

[Lew17] Nicholas Lewis. Purchasing power: Rivalry, dissent, and computing strategy in supercomputer selection at Los Alamos. IEEE annals of the history of computing, 39(3):25–40, July/ September 2017. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL https://www.
computer.org/csdl/mags/
an/2017/03/man2017030025-
abs.html.

Levin:1986:SDC

[LEY86] E. Levin, C. K. Eaton, and Bruce Young. Scaling of data communications for an advanced supercomputer network. NASA contractor report NASA CR-180979, NASA, Washington, DC, USA, 1986. ?? pp.

Lublin:2003:WPS

[LF03] Uri Lublin and Dror G.

Feitelson. The workload on parallel supercomputers: modeling the characteristics of rigid jobs. Journal of Parallel and Distributed Computing, 63(11): 1105–1122, November 2003. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Li:2020:EAI

[LFJ^{+}20] Liandeng Li, Jiarui Fang, Jinlei Jiang, Lin Gan, Weijie Zheng, Haohuan Fu, and Guangwen Yang. Efficient AES implementation on Sunway TaihuLight supercomputer: a systematic approach. Journal of Parallel and Distributed Computing, 138(??):178–189, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0743731519301108.

Lee:1987:FPS

[LG87] Daeshik Lee and E. J. (Efstratios J.) Gallopoulos. Fast Poisson solver on irregular regions by boundary integralbased domain decomposition methods. Technical Report CSRD 687, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 4 pp.

Lagerstrom:1997:PPS

[LG97] R. N. Lagerstrom and S. K. Gipp. PScheD: Political scheduling on the CRAY T3E. Lecture Notes in Computer Science, 1291:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:2003:MLH

[LG03] Quanming Lu and Vladimir Getov. Mixed-language highperformance computing for plasma simulations. Scientific Programming, 11(1): 57–66, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Levinthal:1987:EHS

[LGG^{+}87] D. Levinthal, H. Goldman, C. Georgiopoulos, J. L. DeKeyser, S. Linn, S. Youssef, and M. F. Hodous. Experimental HEP supercomputing at F.S.U. Computer Physics Communications, 45(1–3):137–146, August 1, 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465587901482.

Lam:1986:FMEa

[LH86a] N. C. Lam and Chuck Hastings. Fast 16 × 16 ECL Cray multiplier. WESCON Conference Record, 1986. CO-DEN WCREDI. ISSN 1044-6036, 0083-8837.

Lam:1986:FMEb

[LH86b] N. C. Lam and Chuck Hastings. Fast 16 × 16 ECL Cray multiplier. Conference Record

— Midcon, 1986. CODEN MCORDY.

Lam:1986:FMEc

[LH86c] N. C. Lam and Chuck Hastings. Fast 16 × 16 ECL Cray multiplier. Northcon — Conference Record, pages 4.3.1–4.3.5, 1986. CODEN NCREDL.

Lam:1986:FMEd

[LH86d] N. C. Lam and Chuck Hastings. Fast 16 × 16 ECL Cray multiplier. Southcon Conference Record, 7, 1986. CO-DEN SCOREX.

Lam:1987:FME

[LH87] N. C. Lam and Chuck Hastings. Fast 16 multiplied by 16 ECL Cray MULTIPLIER. Conference Record — Electro, pages 11. 3. 1–11. 3. 8, 1987. CODEN ELCRDH.

Leland:1994:ESS

[LH94] R. Leland and B. Hendrickson. An empirical study of static load balancing algorithms. In IEEE [IEE94c], pages 682–685. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Lautrup:1995:MWS

[LHLM95] B. Lautrup, L. K. Hansen, I. Law, and N. Moerch. Massive weight sharing: a cure for extremely ill-posed problems. In Herrmann et al. [HWP95], pages 137–148. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Lirkov:2021:PAP

[LHPG21] Ivan Lirkov, Stanislav Harizanov, Marcin Paprzycki, and Maria Ganzha. Performance analysis of parallel high-resolution image restoration algorithms on Intel supercomputer. Concurrency and Computation: Practice and Experience, 33 (4):e5996:1–e5996:??, February 25, 2021. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:1989:IID

[Li89] Zhiyuan Li. Intraprocedural and interprocedural data dependence analysis for parallel computing. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1989. x + 161 pp.

Li:1991:CAE

[Li91] Zhiyuan Li. Compiler algorithms for event variable synchronization. Technical Report CSRD 1082, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 10 pp.

Li:1992:APP

[Li92] Z. Li. Array privatization for parallel execution of loops.

In ACM [ACM92b], pages 313–322. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Li:1995:CCO

[Li95] W. Li. Compiler cache optimizations for banded matrix problems. In ACM [ACM95a], pages 21–30. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Liddell:1996:HEH

[Lid96] H. Liddell, editor. HPCN Europe: High-performance computing and networking: International conference — April 1996, Brussels, number 1067 in LECTURE NOTES IN COMPUTER SCIENCE 1996. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61142-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

Liddell:1999:HPC

[Lid99] Heather M. Liddell. High-Performance Computing and Networking Europe 1998. Future Generation Computer Systems, 15(3):307– 308, April 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

19/19/30/19/31/abstract.
html.

Lien:1990:CCS

[Lie90] John Lien. CAD/CAM for supercomputer printed circuit board artwork, 1990. 1 videocassette (61 min.).

Liebman:1993:ANN

[Lie93] M. N. Liebman. Application of neural networks to the analysis of structure and function in biologically active macromolecules. In Lim et al. [L^{+}93], pages 331–348. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Liebsch:2020:CIA

[Lie20]  T. Liebsch.  Concurrent

installation and acceptance of Summit and Sierra supercomputers. IBM Journal of Research and Development, 64(3/4):6:1–6:8, May/ July 2020. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Lilja:1988:MRBa

[Lil88] David J. Lilja. Methods of reducing the branch penalty in pipelined processors. Technical Report CSRD 790, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 19, 1988. 30 pp.

Lilja:1989:EGP

[Lil89] David J. Lilja. Efficient generation of Poisson distributed random variables. Technical Report CSRD 900, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 31, 1989. 15 pp.

Lilja:1991:PPC

[Lil91] David J. Lilja. Processor parallelism consideration[s] and memory latency reduction in shared memory multiprocessors. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1991. x + 178 pp.

Lim:1991:CPP

[Lim91a] Hock-Beng Lim. Characterization of parallel program behavior on a sharedmemory multiprocessor. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1991. ix + 60 pp.

Lim:1991:SAAb

[Lim91b] Swee Boon Lim. Supercomputing application access characteristics. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA, 1991. x + 93 pp.

Lim:1993:SIC

[Lim93] Hwa A. Lim, editor. The Second International Conference on Bioinformatics, Supercomputing, and Complex Genome Analysis: proceedings of the June 4–7, 1992 Conference at St. Petersburg Beach, Florida, USA. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1993. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Lincoln:1982:TDT

[Lin82] N. R. Lincoln. Technology and design tradeoffs in the creation of a modern supercomputer. IEEE Transactions on Computers, C-31(5):349–362, May 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=1676013.

Lincoln:1983:SCC

[Lin83] N. Lincoln. Supercomputers = colossal computations + enormous expectations + renowned risk. Computer, 16(5):38–47, 1983. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:1989:SAP

[Lin89] Shun-Tien Lin. Supercomputer analysis of paperboard structures. Thesis (M.S.), Auburn University, Auburn, AL, USA, 1989. xii + 163 pp.

Liu:1995:WCD

[Liu95] Xiaomao Liu. Workstations cluster for distributed supercomputing. Mini-Micro Systems, 16(2):45–52, February 1995. CODEN XWJXEH. ISSN 1000-1220.

Liu:2012:HPC

[Liu12] Zhen Liu. High-performance computing in mobile services. ACM SIGMETRICS Performance Evaluation Review, 40 (1):3–4, June 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Littlehales:1997:MED

[LJ97] P. Littlehales and B. Jones. Motion engineering design instructor for manufacturing machine systems. In Roller [Rol97], pages 169– 176. ISBN 0-947719-88-1 (paperback). LCCN ????

Laifer:1993:DAT

[LK93] R. Laifer and A. Knocke. Distributed applications and their use by FORTRAN. In Kusters et al. [KSW93], pages 788–?? ISBN 3-923704-11-9. LCCN ???? Two volumes.

Lorenz:2005:VTB

[LKFU05] J. Lorenz, S. Kral, F. Franchetti, and C. W. Ueberhuber. Vectorization techniques for the Blue Gene/L double FPU. IBM Journal of Research and Development, 49 (2/3):437–446, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/lorenz.pdf.

Lewis:1994:DDA

[LKH94] C. P. Lewis, M. Kraft, and T. G. Hesketh. The development of a digital accelerometer. In Anonymous [Ano94-75], pages 873–880. ISBN 0-947719-68-7. LCCN ????

Luecke:2003:CPM

[LKJ03] Glenn R. Luecke, Marina Kraeva, and Lili Ju. Comparing the performance of MPICH with Cray’s MPI and with SGI’s MPI. Concurrency and Computation: Practice and Experience, 15 (9):779–802, August 10, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:1988:ADM

[LL88] Kyungsook Yoon Lee and Daeshik Lee. On the augmented data manipulator network in SIMD environments. Technical Report CSRD 773, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1988. 574–583 pp.

Lundberg:1994:OUB

[LL94] L. Lundberg and H. Lennerstad. An optimal upper bound on the minimal completion time in distributed supercomputing. In Anonymous [Ano94-134], pages 196–203. ISBN ???? LCCN ????

Lan:2008:AFM

[LL08] Zhiling Lan and Yawei Li. Adaptive fault management of parallel applications for high-performance computing. IEEE Transactions on Computers, 57(12): 1647–1660, December 2008. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=4531733.

Lyster:1995:ICT

[LLDF95] P. M. Lyster, P. C. Liewer, V. K. Decyk, and R. D. Ferraro. Implementation and characterization of threedimensional particle-in-cell codes on multiple-instruction-multiple-data massively parallel supercomputers. Computers in Physics, 9(4):420– ??, ???? 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Li:2009:FAR

[LLGS09] Yawei Li, Zhiling Lan, Prashasta Gujrati, and Xian-He Sun. Fault-aware runtime strategies for highperformance computing. IEEE Transactions on Parallel and Distributed Systems, 20(4): 460–473, April 2009. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Luo:2017:DNN

[LLL^{+}17] Tao Luo, Shaoli Liu, Ling Li, Yuqing Wang, Shijin Zhang, Tianshi Chen, Zhiwei Xu, Olivier Temam, and Yunji Chen. DaDianNao: A neural network supercomputer. IEEE Transactions on Computers, 66(1): 73–88, January 2017. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lloyd:1994:EQS

[Llo94] S. Lloyd. Envisioning a quantum supercomputer. Science, 263(5147):695–??, February 4, 1994. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Lang:1993:SVS

[LLR93a] Ulrich Lang, Ruth Lang, and Roland Rühle. Scientific visualization in a supercomputer network at RUS. Computers and Graphics, 17(1):15–22, January–February 02, 1993. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Lau:1993:CAP

[LLR93b] Richard L. Lau, Ding Lee, and Allan R. Robinson, editors. Computational acoustics: proceedings of the 3rd IMACS Symposium on Computational Acoustics, Cambridge, MA, USA, 26–28 June 1991. North-Holland, Amsterdam, The Netherlands, 1993. ISBN 0-444-89763-1 (set), 0-444-89761-5 (vol. 1), 0-444-89762-3 (vol. 2). LCCN QC242.8.I54 1991.

Ludwig:2002:TEC

[LLSR02] T. Ludwig, M. Lindermeier, A. Stamatakis, and G. Rackl. Tool environments in CORBA-based medical high-performance computing. Future Generation Computer Systems, 18(6): 841–847, May 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Liu:1992:SSP

[LLY92] Z. Liu, X. Li, and J.-H. You. On storage schemes for parallel array access. In ACM [ACM92b], pages 282–291. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Laminie:1990:SNE

[LM90a] Jacques Laminie and Ulrike Meier. Solving Navier–Stokes equations on the Cedar multicluster system. Technical Report CSRD 978, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1990. 6 pp.

Lue:1990:GSB

[LM90b] K. Lue and K. Miyai. Graphics supercomputer benchmark: Basic performance of two graphics supercomputers: Stellar GS1000 and Ardent Titan-2. Future Generation Computer Systems, 5 (4):387–394, January 1, 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Loffler:1992:BWC

[LM92] B. Loffler and A. Muller, editors. Basel world CFD user days 1992: 1st World applied computational fluid dynamics: 7th Annual conference — May 1992, Basel, Switzerland, Basel World User Days CFD 1992. IHFB, Basel, Switzerland, 1992. ISBN ???? LCCN ????

Lautard:1993:FPD

[LM93] J. J. Lautard and F. Moreau.

A fast 3D parallel diffusion solver based on a mixed dual finite element approximation.

In Kusters et al. [KSW93], pages 89–99. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Lue:19xx:GSB

[LMxx] K. Lue and K. Miyai. Graphics supercomputer benchmark. Future Generation Computer Systems, 5(4):387– ??, ???? 19xx. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lathrop:2008:HPC

[LM08] Scott Lathrop and Thomas Murphy. High-performance computing education. Computing in Science and Engineering, 10(5):9–11, September/October 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lu:2013:MLP

[LM13] Ligang Lu and Karen Magerlein. Multi-level parallel computing of reverse time migration for seismic imaging on Blue Gene/Q. ACM SIGPLAN Notices, 48 (8):291–292, August 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’13 Conference proceedings.

Leuschel:1998:CGP

[LMD98] Michael Leuschel, Bern Martens, and Danny De Schreye. Controlling generalization and polyvariance in partial deduction of normal logic programs. ACM Transactions on Programming Languages and Systems, 20(1): 208–258, January 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-
1/p208-leuschel/.

Lagace:1990:EEV

[LMH90] P. J. Lagace, D. Mukhedkar, and H. H. Hoang. Evaluation of the effect of vertical faults on the voltage distribution around HVDC electrodes using a supercomputer. IEEE transactions on power delivery, 5(3):1309– ??, July 1, 1990. CO-DEN ITPDE5. ISSN 0885-8977 (print), 1937-4208 (electronic).

Lubeck:1985:BCTb

[LMM85a] Olaf Lubeck, James Moore, and Raul Mendez. Benchmark Comparison of Three Supercomputers: Fujitsu VP-200, Hitachi S810/20, and Cray X-MP/2. IEEE, New York, NY, USA, 1985. ISBN 0-8186-0654-1. 320–329 pp. LCCN ???? IEEE Service Cent. Piscataway, NJ, USA.

Lubeck:1985:BCTa

[LMM85b] Olaf Lubeck, James Moore, and Raul H. Mendez. Benchmark comparison of three supercomputers: Fujitsu VP-200, Hitachi S810/20, and Cray X-MP/2. Computer, 18 (12):10–24, December 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lubeck:1986:PET

[LMM86] O. Lubeck, J. Moore, and R. Mendez. A performance evaluation of three supercomputers, Fujitsu XP-200, Hitachi S810/20, CRAY X-MP/24. Applied Mathematics and Computation, 20(??): 143–144, ???? 1986. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Larsen:1993:ADS

[LMM93] E. W. Larsen, J. E. Morel, and J. M. McGhee. Asymptotic derivation of the simplified PN equations. In Kusters et al. [KSW93], pages 718– 729. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Li:2021:SRS

[LMM^{+}21] Zhimin Li, Harshitha Menon, Dan Maljovec, Yarden Livnat, Shusen Liu, Kathryn Mohror, Peer-Timo Bremer, and Valerio Pascucci. SpotSDC: Revealing the silent data corruption propagation in high-performance computing systems. IEEE Transactions on Visualization and Computer Graphics, 27(10):3938–3952, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Leyland:1990:SIG

[LMP^{+}90] P. Leyland, S. Merazzi, P. Perrier, et al. Semiinteractive graphics via workstations from CRAY-2 for real-time industrial design. application: The European Space Shuttle HERMES. In Pitcher [Pit90], pages 361– 372. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Larson:1995:SMP

[LMT95] J. S. Larson, B. C. Massey, and E. Tick. Super Monaco: Its portable and efficient parallel runtime system. Lecture Notes in Computer Science, 966:527–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lilja:1988:MRBb

[LMY88] David J. Lilja, David Michael Marcovitz, and Pen-Chung Yew. Memory referencing behavior and cache performance in a shared memory multiprocessor. Technical Report CSRD 836, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 13, 1988. 31 pp.

Lagman:1994:DOA

[LN94] A. Lagman and W. A. Najjar. On the design of optimal adaptive routers for direct networks. In IEEE [IEE94c], pages 642–649. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Leyk:1996:ILM

[LO96] Z. Leyk and M. R. Osborne. Implementation of the Lanczos method for solving eigenproblems on the VPP500 supercomputer. In May and Easton [ME96], pages 487– 494. ISBN 981-02-2820-1. LCCN ????

Loo:1984:RPD

[Loo84] John Francis Loo. A registerlevel processor design for a massively parallel supercomputer. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA, 1984. iv + 97 pp.

Lopresti:1989:SCC

[Lop89] Daniel Philip Lopresti. Sequence comparison on commercial supercomputers. Technical report SRC-TR-89-010, Supercomputing Research Center: IDA, Lanham, MD, USA, October 1989. 35 pp.

Lou:1990:NIM

[Lou90a] Gang Lou. Nested iterative methods for a class of indefinite systems. Technical Report CSRD 1044, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1990. 23 pp.

Lou:1990:PMS

[Lou90b] Gang Lou. Parallel methods for solving linear systems via [overlapping] decomposition. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1990. v + 94 pp.

Lou:1992:SNR

[Lou92] Gang Lou. Some new results for solving linear systems arising from computational fluid dynamics problems. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. ix + 139 pp.

Lo:1986:SEP

[LP86] Sy-Shin Lo and Bernard Philippe. The symmetric eigenvalue problem on a multiprocessor. Technical Report CSRD-590, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 13 pp.

Laforenza:1990:STS

[LP90] Domenico Laforenza and Raffaele Perego, editors. Supercomputing tools for science and engineering [Sistemi informatici e calcolo parallelo]: Papers from the International Workshop held in Pisa, Italy, Dec. 4–7, 1989, number 53 in Collana scientifica. Franco Angeli, Milano, Italy, 1990. ISBN 88-204-6322-9. LCCN QA76.88.S88 1990.

Larriba-Pey:1994:APC

[LP94] Josep-L Larriba-Pey. An analysis of the parallel computation of arbitrarily branched cable neuron models. Technical report SRC-TR-94-134, Supercomputing Research Center: IDA, Lanham, MD, USA, November 1, 1994. 6 pp.

Lang:1995:PCS

[LPC^{+}95] Ulrich Lang, J. P. Peltier, Paul Christ, Stefan Rill, Dirk Rantzau, Harald Nebel, Andreas Wierse, Ruth Lang, Sylvain Causse, Frédéric Juaneda, Michel Grave, and Peter Haas. Perspectives of collaborative supercomputing and networking in European aerospace research and industry. Future Generation Computer Systems, 11 (4–5):419–430, August 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lange:2011:MOV

[LPD^{+}11] John R. Lange, Kevin Pedretti, Peter Dinda, Patrick G. Bridges, Chang Bae, Philip Soltero, and Alexander Merritt. Minimal-overhead virtualization of a large scale supercomputer. ACM SIG-PLAN Notices, 46(7):169– 180, July 2011. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1997:MOE

[LPLP97] K.-W. Lee, G.-J. Park, K. H.

Lee, and Y.-S. Park. The modelling and the optimization of an electric vehicle using joint analysis. In Roller [Rol97], pages 373– 380. ISBN 0-947719-88-1 (paperback). LCCN ????

Larriba-Pey:1994:GVS

[LPNRJ94] J.-L. Larriba-Pey, J. J.

Navarro, O. Roig, and A. Jorba. A generalized vision of some parallel bidiagonal systems solvers. In Anonymous [Ano94-134], pages 404–411. ISBN ???? LCCN ????

Lo:1986:MAS

[LPS86] Sy-Shin Lo, Bernard Philippe, and Ahmed Sameh. A multiprocessor algorithm for the symmetric tridiagonal eigenvalue problem. Technical Report CSRD-513, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 28 pp.

Lyttle:1990:MSP

[LPS90] R. W. Lyttle, R. H. Perrott, and P. Sritharan. Modeling SIMD-Type parallelism in Ada. Journal of Pascal, Ada and Modula-2, 9(2):10– ??, March 1, 1990. CODEN JOPAD5. ISSN 0735-1232.

Lewis:1994:MMC

[LPV94] J. G. Lewis, D. G. Payne, and R. A. Van de Geijn. Matrix-vector multiplication and conjugate gradient algorithms on distributed memory computers. In IEEE [IEE94c], pages 542–550. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Lu:2018:CRSb

[LQFC18] Yutong Lu, Depei Qian, Haohuan Fu, and Wenguang Chen. China region special section: Big trends: Will supercomputers be super-data and super-AI machines? Communications of the ACM, 61 (11):82–87, November 2018. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/cacm.acm.org/magazines/
2018/11/232217/fulltext.

Lewis:1988:DSM

[LR88a] Paul C. Lewis and Daniel A.

Reed. Debugging shared memory parallel programs. Technical Report CSRD 748, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 3 pp.

Li:1988:SDM

[LR88b] Zhi Yuan Li and Edward M.

Reingold. Solution of a divide-and-conquer maximin recurrence. Technical Report CSRD 835, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1988. 17 pp.

Ligon:1989:RSA

[LR89] W. B. Ligon and Umakishore Ramachandran. A reconfigurable supercomputer architecture. Technical report, School of Information and Computer Science, Georgia Institute of Technology, Atlanta, GA, USA, 1989. 18 + 3 pp.

Lanzatella:1990:SMI

[LR90] Thomas W. Lanzatella and Paul G. Rutherford. Storage management issues for Cray Research. Digest of Papers — IEEE Symposium on Mass Storage Systems, pages 176– 181, 1990. CODEN DPISDX. ISBN 0-8186-2034-x. ISSN 1051-9173. IEEE catalog number 90CH2844-9.

Lee:1992:CAP

[LR92] Y.-F. Lee and B. Ryder. A comprehensive approach to parallel data flow analysis. In ACM [ACM92b], pages 236–247. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Lichnewsky:1987:SS

[LS87] A. Lichnewsky and C. Saguez, editors. Supercomputing: state-of-the-art. Elsevier Science Publishers, Amsterdam, The Netherlands, 1987. ISBN 0-444-70320-9 (Elsevier). LCCN QA76.5 .S8981 1987. Proceedings of a seminar. “Institut national de recherche en informatique et en automatique.”.

Ladkin:1992:CAC

[LS92a] P. Ladkin and B. Simons. Compile-time analysis of communicating processes. In ACM [ACM92b], pages 248– 259. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Leiss:1992:ACY

[LS92b] Ernst L. Leiss and L.-T. Shih.

Autotasking on the Cray Y-MP: a performance evaluation on seismic codes. In Applied Computing: Technological Challenges of the 1990’s (Mar 1–3 1992: Kansas City, KS, USA), pages 825–830. ACM, New York, NY, USA, 1992. ISBN 0-89791-502-x. LCCN ????

Leung:1993:ENN

[LS93a] Chung Siu Leung and Rudy Setiono. Efficient neural network training algorithm for the Cray Y-MP supercomputer. Proceedings of the International Joint Conference on Neural Networks, 2:1943– 1946, 1993. ISBN 0-7803-1421-2.

Lie:1993:MAA

[LS93b] I. Lie and R. Skalin. A multidomain algorithm for advection problems and its application to atmospheric models. In Hoffmann and Kauranne [HK93b], pages 414– 435. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Lim:1993:HGF

[LS93c] H. A. Lim and I. N.

Shindyalov. Homologous gene family database compilation. In Lim et al. [L^{+}93], pages 403–412. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Li:1994:ECA

[LS94] H. Li and K. C. Sevcik. Exploiting cache affinity in software cache coherence. In Anonymous [Ano94-134], pages 264–273. ISBN ???? LCCN ????

Luecke:2004:PSS

[LSK04] Glenn R. Luecke, Silvia Spanoyannis, and Marina Kraeva. The performance and scalability of SHMEM and MPI-2 one-sided routines on a SGI Origin 2000 and a Cray T3E-600. Concurrency and Computation: Practice and Experience, 16(10):1037– 1060, August 25, 2004. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luo:2020:PEA

[LSS^{+}20] L. Luo, T. P. Straatsma, L. E. A. Suarez, R. Broer, D. Bykov, E. F. D’Azevedo, S. S. Faraji, K. C. Gottiparthi, C. De Graaf, J. A. Harris, R. W. A. Havenith, H. J. A. Jensen, W. Joubert, R. K. Kathir, J. Larkin, Y. W. Li, D. I. Lyakh, O. E. B. Messer, M. R. Norman, J. C. Oefelein, R. Sankaran, A. F. Tillack, A. L. Barnes, L. Visscher, J. C. Wells, and M. Wibowo. Pre-exascale accelerated application development: The ORNL Summit experience. IBM Journal of Research and Development, 64(3/4):11:1– 11:21, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Lin:1993:PCC

[LTD^{+}93] Mengjou Lin, Rose P. Tsang, David H. C. Du, Alan E. Klietz, and Stephen Saroff. Performance characteristics of the Connection Machine hypertree network. Journal of Parallel and Distributed Computing, 19(3): 245–??, November 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Lederman:1992:PER

[LTT92] Steven Lederman, Anna Tsao, and Thomas Turnbull. A parallelizable eigensolver for real diagonalizable matrices with real eigenvalues. Technical report SRC-TR-91-042, Supercomputing Research Center: IDA, Lanham, MD, USA, March 23, 1992. 22 pp.

Lu:1993:ESH

[Lu93] S.-Y. Lu. Exhaustive search of homologous regions between two large DNA sequences using an entropy measure. In Lim et al. [L^{+}93], pages 567–572. ISBN 981-02-1157-0. LCCN QH445.2 .I57

Lucas:1991:HMA

[Luc91] Robert F. Lucas. A hybrid multifrontal algorithm for factoring extremely sparse matrices. Technical report SRC-TR-91-052, Supercomputing Research Center: IDA, Lanham, MD, USA, October 24, 1991. 21 pp.

Lucas:2001:RSL

[Luc01] Mark Lucas. Remote sensing with Linux. Linux Journal, 82:168, 170, February 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Lucky:2009:WPR

[Luc09] R. W. Lucky. Wicked problems [reflections]. IEEE Spectrum, 46(7):27, July 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lumb:2001:LCH

[Lum01] Ian Lumb. Linux clustering for high-performance computing. ;login: the USENIX Association newsletter, 26 (5):??, August 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2001-08/pdfs/lumb.
                                        pdf.

Lundstrom:1990:SSP

[Lun90] S. F. Lundstrom. Supercomputing systems — a projection to 2000. Computing systems in engineering: an international journal, 1(2):145– 152, ???? 1990. CODEN COSEEO. ISSN 0956-0521.

Lundstrom:1994:DGI

[Lun94] Stephen F. Lundstrom, editor. Defining the global information infrastructure: infrastructure, systems, and services: proceedings of a conference held 1–3 November 1994, Boston, Massachusetts, volume 56 of Critical Reviews of Optical Science and Technology. SPIE Optical Engineering Press, Bellingham, WA, USA, 1994. ISBN 0-8194-1681-9 (hardcover), 0-8194-1680-0 (softcover). LCCN TK5105.5.D45 1994.

Luebeck:1996:SCP

[LUT96] S. Luebeck, K. D. Usadel, and B. Tadic. Self-organized criticality and phase transition in a stochastic sandpile automaton. In Wolf et al. [WSB96], pages 367– 371. ISBN 981-02-2635-7. LCCN ????

Langhammer:1992:ST

[LW92] F. Langhammer and F. Wray.

Supercomputing and transputers. In ACM [ACM92b], pages 114–128. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Liu:1994:BCC

[LW94] Xiao Liu and George L. Wilcox. Benchmarking of the CM-5 and the Cray machines with a very large backpropagation neural network. IEEE International Conference on Neural Networks — Conference Proceedings, 1: 22–27, 1994. IEEE catalog number 94CH3429-8.

Levesque:2011:HPC

[LW11] John M. Levesque and Gene Wagenbreth. High performance computing: programming and applications. Chapman and Hall/CRC computational science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2011. ISBN 1-4200-7705-8 (hardcover). xvi + 228 pp. LCCN QA76.88 .L48 2011.

Li:2020:QSC

[LWY^{+}20] R. Li, B. Wu, M. Ying, X. Sun, and G. Yang. Quantum supremacy circuit simulation on Sunway TaihuLight. IEEE Transactions on Parallel and Distributed Systems, 31(4):805–816, April 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2016:POL

[LXW^{+}16] Dali Li, Chuanfu Xu, Yongxian Wang, Zhifang Song, Min Xiong, Xiang Gao, and Xiaogang Deng. Parallelizing and optimizing large-scale 3D multi-phase flow simulations on the Tianhe-2 supercomputer. Concurrency and Computation: Practice and Experience, 28(5):1678– 1692, April 10, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Li:1988:EIA

[LY88a] Zhiyuan Li and Pen-Chung Yew. Efficient interprocedural analysis for program parallelization and restructuring. Technical Report CSRD 804, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 85–99 pp.

Li:1988:IAPa

[LY88b] Zhiyuan Li and Pen-Chung Yew. Interprocedural analysis and program restructuring for parallel programs. Technical Report CSRD 720, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 54 pp.

Li:1988:IAPb

[LY88c] Zhiyuan Li and Pen-Chung Yew. Interprocedural analysis for parallel computing. Technical Report CSRD 732, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 221– 228 pp.

Lilja:1990:CPE

[LY90a] David J. Lilja and Pen-Chung Yew. Comparing parallelism extraction techniques: superscalar processors, pipelined processors, and multiprocessors. Technical Report CSRD 954, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 8, 1990. 32 pp.

Lilja:1990:CDC

[LY90b] David J. Lilja and Pen-Chung Yew. A compiler-assisted directory-based cache coherence scheme. Technical Report CSRD 990, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1990. 49 pp.

Lilja:1990:ICB

[LY90c] David J. Lilja and Pen-Chung Yew. The interaction of cache block size and parallel loop scheduling strategy. Technical Report CSRD 989, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1990. 20 pp.

Lilja:1991:CHS

[LY91a] David J. Lilja and Pen-Chung Yew. Combining hardware and software cache coherence strategies. Technical Report CSRD 1057, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 10 pp.

Lim:1991:PPB

[LY91b] Hock-Beng Lim and Pen-Chung Yew. Parallel programs behavioral study on a shared multiprocessor. Technical Report CSRD 1062, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 1 + 29 pp.

Loo:1993:PEE

[LYC93] A. W. S. Loo, R. W. M. Yip, and C.-W. Chung. Performance evaluation of efficient parallel sorting algorithm for supercomputer. In Brebbia and Power [BP93], pages 111–120. ISBN 1-85166-845-4, 1-85312-236-X, 1-56252-160-8. LCCN TA345.A66 1993.

Lyu:1997:SAD

[LYKM97] M.-S. Lyu, S.-J. Yoon, Y.-G.

Ku, and J.-S. Maeng. Study on air-cleaner design for enhanced intake system. In Roller [Rol97], pages 413– 522. ISBN 0-947719-88-1 (paperback). LCCN ????

Lee:1987:MCD

[LYL87a] Roland L. Lee, Pen-Chung Yew, and Duncan Lawrie. Multiprocessor cache design considerations. Technical Report CSRD 640; UILU-ENG-640, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 10 pp.

Lee:1987:DPS

[LYL87b] Roland L. Lee, Pen-Chung Yew, and Duncan Hamish Lawrie. Data prefetching in shared memory multiprocessors. Technical Report CSRD-639, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 17 pp.

Loewe:1995:UTB

[LZ95] W. Loewe and W. Zimmermann. Upper time bounds for executing PRAM-Programs on the LogP-Machine. In ACM [ACM95a], pages 41– 50. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Li:2016:ROP

[LZF16] Gang Li, Qingpu Zhang, and Zhengqian Feng. Research on optimal path of data migration among multisupercomputer centers. Scientific Programming, 2016(??): 5018213:1–5018213:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/5018213/

.

Mahajan:1994:SSV

[M^{+}94] Sunita Mahajan et al., editors. Supercomputing for scientific visualisation: proceedings of the Seminar on Supercomputing for Scientific Visualisation, Bhabha Atomic Research Centre, Bombay, India, February 15–17, 1994. Tata McGraw-Hill Publishing Co. Ltd., New Delhi, India, 1994. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Mirenkov:1995:FAI

[M^{+}95] Nikolay Mirenkov et al., editors. The First Aizu International Symposium on Parallel Algorithms/Architecture Synthesis: proceedings, March 15–17, 1995, Aizu-Wakamatsu, Fukushima, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7038-X. LCCN QA76.642.A43 1995.

Magoules:2009:IGC

[M^{+}09] F. (Frédéric) Magoulès et al.

Introduction to grid computing. Chapman and Hall/CRC numerical analysis and scientific computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-7406-7 (hardcover). xxiii + 310 pp. LCCN QA76.9.C58 I5772 2009.

Murman:1985:PSC

[MA85] Earll M. Murman and Saul S. Abarbanel, editors. Progress and supercomputing in computational fluid dynamics: proceedings of U.S.-Israel Workshop, 1984, volume 6 of Progress in scientific computing; vol. 6. Birkhäuser Verlag, Basel, Switzerland, 1985. ISBN 0-8176-3321-9. LCCN QA911 .U21 1984.

Marcic:1997:CSC

[MA97] M. Marcic and J. Avsec. Computer simulation of the combustion of the injected fuel spray. In Roller [Rol97], pages 467–474. ISBN 0-947719-88-1 (paperback). LCCN ????

Ma:1999:CPP

[Ma99]  Sangback Ma.  Compar-

isons of the parallel preconditioners on the CRAY-T3E for large nonsymmetric linear systems. International Journal of High Speed Computing, 10(3):285–300, September 1999. CODEN IHSCEZ. ISSN 0129-0533.

Maas:1993:SCK

[Maa93a] U. Maas. 93SC021 simplification of chemical kinetics for the simulation of combustion processes. In Anonymous [Ano93-31], pages 131– 138. ISBN 0-947719-62-8. LCCN ????

Merino:1993:DVP

[MAA93b] F. Merino, C. Ahnert, and J. M. Aragones. Development and validation of the 3-D PWR core dynamics SIM-TRAN code. In Kusters et al. [KSW93], pages 646– 657. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Moreira:2005:BGP

[MAA^{+}05] J. E. Moreira, G. Almási, C. Archer, R. Bellofatto, P. Bergner, J. R. Brunheroto, M. Brutman, J. G. Castaños, P. G. Crumley, M. Gupta, T. Inglett, D. Lieber, D. Limpert, P. McCarthy, M. Megerian, M. Mendell, M. Mundy, D. Reed, R. K. Sahoo, A. Sanomiya, R. Shok, B. Smith, and G. G. Stewart. Blue Gene/L programming and operating environment. IBM Journal of Research and Development, 49 (2/3):367–376, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/moreira.
pdf.

MacDonald:1990:CCC

[Mac90] Tom MacDonald. Cray C: C in a 64-bit world. The Journal of C Language Translation, 1(4):310–315, March 1990. ISSN 1042-5721.

MacDonald:1991:CCF

[Mac91a] Tom MacDonald. Cray C and Fortran interlanguage communication. The Journal of C Language Translation, 2(4): 305–317, March 1991. ISSN 1042-5721.

MacKenzie:1991:IAL

[Mac91b] Donald MacKenzie. The influence of the Los Alamos and Livermore National Laboratories on the development of supercomputing. Annals of the History of Computing, 13(2):179–201, April/June 1991. CODEN AHCOE5. ISSN 0164-1239. URL http:

                                        //dlib.computer.org/an/
                                        books/an1991/pdf/a2179.
                                        pdf;   http://www.computer.
                                        org/annals/an1991/a2179abs.
                                        htm.

Mackerle:1992:FEB

[Mac92] J. Mackerle. Finite elements and boundary elements and supercomputing — a bibliography. Finite elements in analysis and design, 12(2): 151–160, October 1992. CO-DEN FEADEU. ISSN 0168-874X.

Macilwain:1996:CFF

[Mac96a]  Colin Macilwain.  Cray

files formal ‘dumping complaint’. Nature, 382(6590): 385, August 1, 1996. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Macilwain:1996:LGT

[Mac96b]  Colin Macilwain.  Liver-

more to get $100M supercomputer. Nature, 382(6590): 385, August 1, 1996. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Mackerle:1996:IFE

[Mac96c]  Jaroslav Mackerle.  Im-

plementing finite element methods on supercomputers, workstations and PCs: a bibliography (1985-1995). Engineering Computations, 13(1): 33–??, ???? 1996. CODEN ENCOEN. ISSN 0264-4401.

Machlis:1997:WSB

[Mac97a] Sharon L. Machlis. From workstation to supercomputer, block by block. Design news, ??(2):57–??, ???? 1997. CODEN DIGNAO. ISSN 0011-9407.

Macilwain:1997:ELU

[Mac97b] Colin Macilwain. Energy labs urged to boost supercomputing capability. Nature, 390(6661):651, December 18, 1997. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Macilwain:1997:USC

[Mac97c] Colin Macilwain. US supercomputer centres to concentrate on networking. Nature, 386(6625):528, April 10, 1997. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Mache:2008:GCU

[MAFW08] Jens Mache, Amy Apon, Thomas Feilhauer, and Barry Wilkinson. Grid computing at the undergraduate level: can we do it? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1): 381–382, March 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Magoules:2010:FGC

[Mag10] F. (Frédéric) Magoulès, editor. Fundamentals of grid computing: theory, algorithms and technologies. Chapman and Hall/CRC numerical analysis and scientific computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-0367-6 (hardcover). xxi + 298 pp. LCCN QA76.9.C58 F86 2010.

Mahabala:1994:CSE

[Mah94a] H. N. Mahabala. Computer Science education in India. In Balakrishnan [Bal94], pages 325–330. ISBN 0-07-462044-4. LCCN ????

Mahajan:1994:CPM

[Mah94b] S. M. Mahajan. A common programming model for distributed memory systems. In Mahajan et al. [M^{+}94], pages 359–366. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Maheshwari:1994:CAP

[Mah94c] P. Maheshwari. On the cost analysis of parallel functional programs for efficient dynamic execution. In Balakrishnan [Bal94], pages 201– 211. ISBN 0-07-462044-4. LCCN ????

Majumdar:1994:DPA

[Maj94]  J. Majumdar.  Develop-

ment of parallel algorithms for computer vision. In Mahajan et al. [M^{+}94], pages 267–274. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Malony:1986:CPE

[Mal86a] Allen Davis Malony. Cedar performance evaluation tools: a status report. Technical Report CSRD-582, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 32 pp.

Malony:1986:CPM

[Mal86b] Allen Davis Malony. Cedar performance measurements.

Technical Report CSRD-579, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 13 pp.

Malaney:1988:SNA

[Mal88a] Robert A. Malaney. Supercomputers and nuclear astrophysics. Computers in Physics, 2(6):40– ??, November 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822797.

Malony:1988:RPA

[Mal88b] Allen Davis Malony. Regular processor arrays. Technical Report CSRD 734, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 23 + [45] pp.

Malony:1989:JJE

[Mal89] Allen Davis Malony. JED: just an event display. Technical Report CSRD 887, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1989. 15 pp.

Malony:1990:PO

[Mal90] Allen Davis Malony. Performance observability. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1990. xiii + 288 pp.

Malony:1991:EPP

[Mal91] Allen D. Malony. Eventbased performance perturbation: a case story. Technical Report CSRD 1060, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 12 pp.

Manohar:1988:SV

[Man88] Swaminathan Manohar. Supercomputing with VLSI. Technical report CS-88-14, Brown University, Dept. of Computer Science, Providence, RI, USA, 1988. xvii + 275 pp.

MCC:1989:SAM

[Man89a] Manchester Computing Centre. Supercomputing at Manchester Computing Centre: a summary of peer reviewed research projects using the supercomputers at Manchester Computing Centre, 1987-1989. Manchester Computing Centre, Manchester, UK, 1989. x + 119 + [46] pp.

Manohar:1989:SV

[Man89b] Swaminathan Manohar. Supercomputing with VLSI. Thesis (Ph.D.), Brown University, Providence, RI, USA, 1989. xvii + 275 pp.

Mankofsky:1990:TDE

[Man90]  A. Mankofsky.  Three-

dimensional electromagnetic particle-in-cell simulations of physical devices. In Pitcher [Pit90], pages 373–384. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Mansfield:1991:DJP

[Man91] Lois Mansfield. Damped Jacobi preconditioning and coarse grid deflation for conjugate gradient iteration on parallel computers. SIAM Journal on Scientific and Statistical Computing, 12(6): 1314–1323, November 1991. CODEN SIJCD4. ISSN 0196-5204.

MCC:1992:SAM

[Man92] Manchester Computing Centre, Manchester, UK. Supercomputing at Manchester Computing Centre: a summary of peer reviewed research projects using the Amdahl VP1200 supercomputer at Manchester Computing Centre, 1989–1991, 1992. ix + 187 pp.

Martin:1985:DSR

[Mar85a] Joanne L. Martin. Developments and software requirements of the emerging National Supercomputer Research Centers. IEEE Software, 2(6):55–67, November 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Martin:1985:NSR

[Mar85b]  Joanne L. Martin.  Na-

tional Supercomputer Research Centers. IEEE Software, 2(6):55–67, November 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Marino:1986:SAA

[Mar86] C. (Carlos) Marino, editor. Supercomputer applications in automative research and engineering development: proceedings of the International Conference on Supercomputer Applications in the Automotive Industry, Zurich, Switzerland, October 1986, A Computational Mechanics publication. Computational Mechanics Publications, Southampton, UK, 1986. ISBN 0-905451-54-6 (Southampton), 0-931215-29-3 (Boston, Los Angeles). LCCN TL240.I5 1986.

Marcovitz:1988:MCP

[Mar88a] David Michael Marcovitz. A multiprocessor cache performance metric. Technical Report CSRD 813; UILU-ENG-88-8011, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. vi + 70 pp.

Marino:1988:SAA

[Mar88b] C. (Carlos) Marino, editor. Supercomputer applications in automotive research and engineering development: proceedings of the Second International Conference on Supercomputing Applications in the Automotive Industry, Seville, Spain, October 1988. Cray Research, Inc., Minneapolis, MN, USA, 1988. ISBN ???? LCCN TL240 .I528 1988.

Martin:1988:PES

[Mar88c] Joanne L. Martin. Performance evaluation of supercomputers, volume 4 of Special topics in supercomputing. Elsevier Science Publishers, Amsterdam, The Netherlands, 1988. ISBN 0-444-70448-5. x + 419 pp. LCCN QA76.9.E94 P441 1988.

Marshall:1990:CRC

[Mar90] T. Marshall. A calculating RISC: Coprocessors based on RISC engines will soon deliver supercomputer performance to your desktop. BYTE Magazine, 15(5):251– 254, 256, May 1990. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Marsolf:1991:LGP

[Mar91]  Bret A. Marsolf.  Large

grain parallel sparse system solver. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. viii + 80 pp.

Martin:1992:FPV

[Mar92] Jeanne Martin. Fortran 90 pointers vs. “Cray” pointers. ACM Fortran Forum, 11(2): 17–23, June 1992. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Marenzoni:1995:PAC

[Mar95] P. Marenzoni. Performance analysis of Cray T3D and Connection Machine CM-5: a comparison. Lecture Notes in Computer Science, 919:110–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marksteiner:1996:HPC

[Mar96] Peter Marksteiner. Highperformance computing — an overview. Computer Physics Communications, 97 (1–2):16–35, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0010465596000185.

Masdupuy:1991:UAI

[Mas91] F. Masdupuy. Using abstract interpretation to detect array data dependencies. In Anonymous [Ano91q], pages 19–27. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Masdupuy:1992:AOA

[Mas92] F. Masdupuy. Array operations abstraction using semantic analysis of trapezoid congruences. In ACM [ACM92b], pages 226–235. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Masi:1993:ISS

[Mas93] E. Masi. Intel Supercomputer Systems Division. In Hoffmann and Kauranne [HK93b], pages 329–344. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Mascagni:1994:FHQ

[Mas94a]  Michael Mascagni.  A

fast, high quality, and reproducible parallel lagged-Fibonacci pseudorandom number generator. Technical report SRC-TR-94-115, Supercomputing Research Center: IDA, Lanham, MD, USA, March 15, 1994. 18 pp.

Mascagni:1994:PPN

[Mas94b]  Michael Mascagni.  Par-

allel pseudorandom number generation using additive lagged-Fibonacci recursions. Technical report SRC-TR-94-133, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1, 1994. 15 pp.

Maslov:1995:EAD

[Mas95] V. Maslov. Enhancing array dataflow dependence analysis with on-demand global value propagation. In ACM [ACM95a], pages 265–269. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Morton:1985:ICT

[MAT85] Steven G. Morton, Enrique Abreu, and Fred Tse. ITT cap — toward a personal supercomputer. IEEE Micro, 5(6):37–49, November/December 1985. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Max:1981:VPM

[Max81] Nelson L. Max. Vectorized procedural models for natural terrain: Waves and islands in the sunset. Computer Graphics, 15(3):317– 324, August 1981. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

May:2001:PHP

[May01]  John M. May.  Paral-

lel I/O for high performance computing. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2001. ISBN 1-55860-664-5. xvii + 366 pp. LCCN QA76.88. M39 2001. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els031/00043508.
html;   http://www.loc.
gov/catdir/toc/els031/00043508.
html.

Melli:1989:SES

[MB89] Piero Melli and C. A. Brebbia, editors. Supercomputing in engineering structures. Springer-Verlag and Computational Mechanics Publications, Berlin, Germany / Heidelberg, Germany / London, UK / etc. and Southampton, UK, 1989. ISBN 1-85312-025-1 (??invalid checksum??), 0-945824-07-6 (U.S.), 3-540-50687-X (Germany), 0-387-50687-X (Germany). LCCN TA646 .S87 1989. Result of lectures presented at the Seminar on “Supercomputing in Engineering Structures”, held in Oberlech, Austria on July 11 to 15, 1988.

Murthy:1993:SFF

[MB93] T. K. S. Murthy and C. A. Brebbia, editors. Supercomputing in fluid flow: Papers presented at the International Seminar on Supercomputing in Fluid Flow held in Lovell, Massachusetts, USA, during 3–5 October 1989, International series on computational engineering. Elsevier Applied Science, London, UK, 1993. ISBN 1-85166-759-8 (Elsevier Applied Science), 1-85312-076-6 (Computational Mechanics, Southampton), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911 .S88 1993. Papers presented at the International Seminar on Supercomputing in Fluid Flow held in Lovell, Massachusetts, USA, during 3–5 October 1989.

May:1994:CVD

[MB94a] J. May and F. Berman.

Creating views for debugging parallel programs. In IEEE [IEE94c], pages 833– 840. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Murthy:1994:SFF

[MB94b] T. K. S. Murthy and C. A.

Brebbia, editors. Supercomputing in fluid flow: Papers presented at the International Seminar on Supercomputing in Fluid Flow held in Lovell, Massachusetts, USA, during 3–5 October 1989, International Series on Computational Engineering. Computational Mechanics Publications, Southampton, UK, 1994. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

McGary:1997:IPT

[MB97] J. E. McGary and A. L. Boyer. An interactive, parallel, three-dimensional fast Fourier transform convolution dose calculation using a supercomputer. Medical physics, 24(4):519–??, ???? 1997. ISSN 0094-2405.

Meira:2012:SIC

[MB12] Wagner Meira and Ricardo Bianchini. Special issue on computer architecture and high-performance computing. International Journal of Parallel Programming, 40(3):259–261, June 2012. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=40&
                                        issue=3&spage=259.

Marques:2020:TEE

[MB20] O. Marques and A. Barker. Training efforts in the Exascale Computing Project. Computing in Science and Engineering, 22(5):103–107, 2020. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mudge:1991:DM

[MBB^{+}91] Trevor N. Mudge, Richard B.

Brown, William P. Birmingham, Jeffrey A. Dykstra, Ayman I. Kayssi, Ronald J. Lomax, Oyekunle A. Olukotun, Karem A. Sakallah, and Raymond A. Milano. The design of a microsupercomputer. Computer, 24(1):57– 64, January 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mohr:1999:PHP

[MBD99] Bernd Mohr, Federico Bassetti, and Kei Davis. Parallel/high-performance objectoriented scientific computing. Lecture Notes in Computer Science, 1743:222–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1743/17430222.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1743/17430222.
pdf.

Muller:1992:ASP

[MBK^{+}92] Urs A. Muller, Bernhard Baumle, Peter Kohler, Anton Gunzinger, and Walter Guggenbuhl. Achieving supercomputer performance for neural net simulation with an array of digital signal processors. IEEE Micro, 12(5):55–65, September/ October 1, 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mahmoudi:2020:MPU

[MBM^{+}20] Sidi Ahmed Mahmoudi, Mohammed Amin Belarbi, Saïd Mahmoudi, Ghalem Belalem, and Pierre Manneback. Multimedia processing using deep learning technologies, high-performance computing cloud resources, and Big Data volumes. Concurrency and Computation: Practice and Experience, 32(17): e5699:1–e5699:??, September 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Morillon:1993:IFC

[MBN93] B. Morillon, J. P. Both, and J. C. Nimal. Importance function by collision probability for Monte-Carlo code TRIPOLI. In Kusters et al. [KSW93], pages 745– 755. ISBN 3-923704-11-9. LCCN ???? Two volumes.

McLoughlin:2005:FBC

[MBR05] Ian McLoughlin, Timo Bretschneider, and Bharath Ramesh. First Beowulf cluster in space. Linux Journal, 2005 (137):??, September 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Mueller:1992:ASP

[MBSK92] U. Mueller, B. Baemnfe, W. Scott, and P. Kohler. Achieving supercomputer performance with a DSP processor array. In Anonymous [Ano92y], pages 756– 763. ISBN ???? LCCN ????

Muraoka:2001:TAH

[MBSW01] Yoichi Muraoka, Randall Bramley, David F. Snelling, and Harry Wijshoff. Topic 07 applications on highperformance computers. Lecture Notes in Computer Science, 2150:358–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2150/21500358.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2150/21500358.
pdf.

Moret:2001:HPA

[MBW01] Bernard M. E. Moret, David A.

Bader, and Tandy Warnow. High-performance algorithm engineering for computational phylogenetics. Lecture Notes in Computer Science, 2074:1012–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2074/20741012.
htm;   http://link.springer-

                                        ny.com/link/service/series/
                                        0558/papers/2074/20741012.
                                        pdf.

Meng:2010:HPH

[MC10] Xiandong Meng and Vipin Chaudhary. A high-performance heterogeneous computing platform for biological sequence analysis. IEEE Transactions on Parallel and Distributed Systems, 21(9):1267– 1280, September 2010. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

McAulay:1992:OCA

[McA92]  A. D. McAulay.  Opti-

cal computer architectures for supervised learning systems. Computer, 25(5): 72–75, May 1992. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McBryan:1992:PSW

[McB92a] Oliver A. McBryan. Performance of the Shallow Water Equations on the Suprenum-1 parallel supercomputer. Technical report CU-CS-575-92, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, January 1992. 11 pp.

McBryan:1992:SPS

[McB92b] Oliver A. McBryan. Scaling performance of the Shallow Water Equations on the Suprenum-1 supercomputer.

Technical report CU-CS-637-92, University of Colorado, Boulder, Dept. of Computer Science, Boulder, CO, USA, December 1992. 13 pp.

McBryan:1993:PSW

[McB93] O. A. McBryan. Performance of the shallow water equations on the CM-200 and CM-5 parallel supercomputers. In Hoffmann and Kauranne [HK93b], pages 345– 353. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

McGarvey:2001:BCD

[MCB^{+}01] Brian McGarvey, Robert Cicconetti, Nathan Bushyager, Edan Dalton, and Manos Tentzeris. Beowulf cluster design for scientific PDE models. In USENIX [USE01], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
mcgarvey.html.

McCormick:1988:MMT

[McC88] S. F. (Stephen Fahrney) Mc-Cormick, editor. Multigrid methods: theory, applications, and supercomputing: Proceedings of the Third Copper Mountain Conference on Multigrid Methods, Copper Mountain, Colorado, April 5–10, 1987, volume 110 of Lecture notes in pure and applied mathematics. M. Dekker, New York, NY, USA, 1988. ISBN 0-8247-7979-7 (paperback). LCCN QA377 .M9431 1988. These papers stem from the Third Copper Mountain Conference on Multigrid Methods, which was held at Copper Mountain, Colorado, April 5–10, 1987.

McClaughry:1992:PPT

[McC92] Patrick E. McClaughry. PTOPP: a practical toolset for the optimization of parallel programs. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1992. vii + 22 pp.

McCann:1994:EDC

[McC94] Lester Ivan McCann. Enhanced database concurrency control in a supercomputer vector-processing environment. Thesis (Ph.D.), North Dakota State University, Fargo, ND, USA, 1994. xiv + 194 pp.

McDaniel:1985:NRE

[McD85] Timothy Alan McDaniel. Non-linear recurrences and EISPACK. Technical Report CSRD-511, UILU-ENG-85-8011, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1985. vi + 67 pp.

McDowell:1988:BRS

[McD88] Charles E. McDowell. Book review: Supercomputer Architectures by Paul B. Schneck (Kluwer Academic Publishers). ACM SIGARCH Computer Architecture News, 16(4):195–196, September 1988. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

McDonald:1990:NNA

[McD90]  Ryan O. McDonald.  A

neural network approach to phoneme recognition. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1990. xii + 99 pp.

McGuire:1987:MSC

[McG87]  Patrick J. McGuire.  A

measurement-based study of concurrency in a multiprocessor. Technical Report CSRD 674; UILU-ENG-87-2210, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 49 pp.

Mark:1991:QCR

[MCH91] Y.-S. Mark, S. A. Cuccaro, and P. G. Hipes. Quantum chemical reaction dynamics on a highly parallel supercomputer. Theoretica Chimica Acta, 79(3 / 4):225–??, 1991. CODEN TCHAAM. ISSN 0040-5744.

McKinley:1994:EAP

[McK94] K. S. McKinley. Evaluating automatic parallelization for efficient execution on shared-memory multiprocessors. In Anonymous [Ano94-134], pages 54–63. ISBN ???? LCCN ????

Moore:2007:VEF

[MCLK07] Nicholas Moore, Albert Conti, Miriam Leeser, and Laurie Smith King. Vforce: An extensible framework for reconfigurable supercomputing. Computer, 40(3):39– 49, March 2007. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McNamara:1987:SES

[McN87] Brendan McNamara. Supercomputer environments for science applications. Journal of Computational Physics, 73 (1):41–58, November 1987. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0021999187901045.

McTavish:1996:ECC

[McT96]  D. McTavish.  Equiva-

lence checking at Cray Research. IEEE Spectrum, 33 (6):71, June 1996. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

May:1998:HPE

[MCW98] Phil May, Sek M. Chai, and D. Scott Wills. HiPER-P: An efficient, high-performance router for multicomputer interconnection networks. Lecture Notes in Computer Science, 1417:103–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1417/14170103.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1417/14170103.
pdf.

Morris:1988:FPD

[MD88] L. R. Morris and S. A. Dyer.

Floating-point digital signalprocessing chips — the end of the supercomputer era. IEEE Micro, 8(6):86, November/December 1988. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Meleis:1994:OLR

[MD94] W. M. Meleis and E. S.

Davidson. Optimal local register allocation for a multipleissue machine. In Anonymous [Ano94-134], pages 107–116. ISBN ???? LCCN ????

Mahapatra:2004:AQE

[MD04] Nihar R. Mahapatra and Shantanu Dutt. Adaptive Quality Equalizing: Highperformance load balancing for parallel branch-and-bound across applications and computing systems. Parallel Computing, 30(7):867– 881, July 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Michalickova:2000:SAP

[MDH00] Katerina Michalickova, Moyez Dharsee, and Christopher W. V. Hogue. Sequence analysis on a 216-processor Beowulf cluster. In USENIX [USE00a], page ?? ISBN 1-880446-17-0. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/als2000/michalickova.
html.

Mehrotra:2016:PEA

[MDH^{+}16] Piyush Mehrotra, Jahed Djomehri, Steve Heistand, Robert Hood, Haoqiang Jin, Arthur Lazanoff, Subhash Saini, and Rupak Biswas. Performance evaluation of Amazon Elastic Compute Cloud for NASA high-performance computing applications. Concurrency and Computation: Practice and Experience, 28(4):1041– 1055, March 25, 2016. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Milojicic:2000:PM

[MDP^{+}00] Dejan S. Miloj́ičić, Fred Douglis, Yves Paindaveine, Richard Wheeler, and Songnian Zhou. Process migration. ACM Computing Surveys, 32(3):241–299, 2000. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL

http://www.acm.org/pubs/
articles/journals/surveys/
2000-32-3/p241-miloiic/
p241-miloiic.pdf;   http:
//www.acm.org/pubs/citations/
journals/surveys/2000-32-
3/p241-miloiic/.

Morel:1993:DSS

[MDW93] J. E. Morel, J. E. Dendy, and T. A. Wareing. Diffusionaccelerated solution of the 2-D Sn equations with bilineardiscontinuous differencing. In Kusters et al. [KSW93], pages 488–499. ISBN 3-923704-11-

9. LCCN ???? Two volumes.

McGrath:1987:UMC

[ME87] Robert Edward McGrath and Perry A. Emrath. Using memory in the Cedar system. Technical Report CSRD 655, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 27 pp.

Meier:1991:PPC

[ME91] Ulrike Meier and R. Eigenmann. Parallelization and performance of conjugate gradient algorithms on the Cedar hierarchical memory multiprocessor. Technical Report CSRD 1035, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1991. 11 pp.

May:1996:CTA

[ME96] R. L. May and A. K. Easton, editors. Computational techniques and applications: Biennial conference; 7th — July 1995, Melbourne, Australia, COMPUTATIONAL TECH-NIQUES AND APPLICA-TIONS 1995; 7th. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1996. ISBN 981-02-2820-1. LCCN ????

Mechoso:1995:HPC

[Mec95]  C. R. Mechoso.  High-

performance computing and networking for climate research. Lecture Notes in Computer Science, 919:142– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meher:1994:SAA

[Meh94] P. K. Meher. A systolic array architecture for parallel VLSI implementation of linear convolution of discrete sequences. In Mahajan et al. [M^{+}94], pages 413–415. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Mendez:1984:BJA

[Men84] R. H. Mendez. Benchmarks on Japanese and American supercomputers — preliminary results. IEEE Transactions on Computers, C-33(4):374, April 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=1676446. An ex-

panded version appeared in the SIAM News 17, No. 2, March, 1984, p. 3.

Mendez:1987:SJP

[Men87] R. (Raul) Mendez. Supercomputing in Japan: proceedings of the first APPI Workshop on Supercomputing. ISR Technical report ISDR TR 87-03, Institute for Supercomputing Research, Tokyo, Japan, 1987. 259 pp.

Mendez:1990:VS

[Men90] R. (Raul) Mendez, editor. Visualization in supercomputing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97149-1. LCCN QA76.5.V55 1990. Selected papers from a conference held August 1988, in Tokyo.

Merkey:1986:UDC

[Mer86] Phillip Merkey. Uniquely decodable codes on a circle. Technical report SRC-TR-86-003, Supercomputing Research Center: IDA, Lanham, MD, USA, 1986. 8 pp.

Merkle:2001:NWW

[Mer01] R. C. Merkle. Nanotechnology: what will it mean? IEEE Spectrum, 38(1):19– 21, January 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Messina:1993:CSC

[Mes93a] P. Messina. The concurrent supercomputing consortium: Year 1. IEEE parallel and distributed technology: systems and applications, 1(1): 9–16, February 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Messina:1993:KSD

[Mes93b] P. Messina. Keynote speech: Distributed supercomputing — the CASA Gigabit Testbed experience. In IEEE [IEE93b], pages 100–101. ISBN 0-8186-3900-8, 0-8186-3901-6. LCCN QA76.9.D5 I593 1993. IEEE catalog number 93TH0550-4.

Messina:1997:HPCa

[Mes97a] Paul Messina. High-performance computers: The next generation (Part I). Computers in Physics, 11(5):454– ??, September 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822591.

Messina:1997:HPCb

[Mes97b] Paul Messina. High-performance computers: The next generation (Part II). Computers in Physics, 11(6):598– ??, November 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822610.

Messina:2000:DSG

[Mes00] Paul Messina. Distributed supercomputing for the grid. Journal of Linux Technology, 1(1):30–??, 2000. ISSN 1527-2761.

Messina:2017:ECP

[Mes17]  Paul Messina.  The Ex-

ascale Computing Project. Computing in Science and Engineering, 19(3):63–67, May/June 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/03/mcs2017030063.
html.

Metropolis:1986:SCP

[Met86a] N. Metropolis, editor. Supercomputer conference proceedings: “Frontiers of supercomputing,” 15–19 August 1983, Los Alamos, New Mexico. University of California Press, Berkeley, CA, USA, 1986. ISBN 0-520-05190-4. LCCN QA76.5.S8947 1986. Papers presented at a conference co-sponsored by the Los Alamos National Laboratory and the National Security Agency.

Metropolis:1986:FS

[Met86b] N. (Nicholas) Metropolis, editor. Frontiers of supercomputing, volume 7 of Los Alamos series in basic and applied sciences 7. University of California Press, Berkeley, CA, USA, 1986. ISBN 0-520-05190-4. LCCN QA76.5 .F76 1983. Papers presented at a conference co-sponsored by the Los Alamos National Laboratory and the National Security Agency, held in Los Alamos on Aug. 15-19, 1983.

Meurant:1987:MCG

[Meu87] Gerard Meurant. Multitasking the conjugate gradient method on the Cray X-MP/ 48. Parallel Computing, 5 (3):267–280, November 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Meuer:1989:SAA

[Meu89a] H. W. (Hans W.) Meuer, editor. Supercomputer ’89: Anwendungen, Architekturen, Trends: Seminar, Mannheim, 8.–10. Juni 1989: Proceedings, volume 211 of Informatik-Fachberichte; 211. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-51310-8 (US), 3-540-51310-8 (Germany). LCCN QA76.5 .S8851 1989. English and German. On spine: Supercomputer ’89: proceedings.

Meurant:1989:PUC

[Meu89b] Gérard Meurant. Practical use of the conjugate gradient method on parallel supercomputers. Computer Physics Communications, 53 (1–3):467–477, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465589901793.

Meuer:1990:SAA

[Meu90] H. W. (Hans W.) Meuer, editor. Supercomputer ’90: Anwendungen, Architekturen, Trends: Mannheim, 21.–

23. Juni 1990: Proceedings, volume 250 of Informatik-Fachberichte; 250. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-52792-3. LCCN QA76.5 S89438 1990. German and English. Papers presented at the 5th Seminar “Supercomputer, Anwendung, Architekturen und Trends” sponsored by the Verein zur Forderung der Wissenschaftlichen Weiterbildung an der Universitat Mannheim.

Meuer:1991:SAA

[Meu91] H. W. (Hans W.) Meuer, editor. Supercomputer ’91: Anwendungen, Architekturen, Trends: Seminar, Mannheim, 20.–22. Juni 1991: Proceedings, number 278 in Informatik-Fachberichte. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-54231-0 (Berlin), 0-387-54231-0 (New York). LCCN QA76.5 .S87 1991. German and English.

Meuer:1992:HNS

[Meu92a] H. W. (Hans W.) Meuer. Heterogene Netze und Supercomputer. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55073-9 (Berlin). vi + 154 pp. LCCN TK5105.5.H48 1992.

Meuer:1992:SAAa

[Meu92b] H. W. (Hans W.) Meuer. Supercomputer, 1986-1990: Anwendungen, Architekturen, Trends, volume 3 of Fokus. K. G. Saur, München, Germany; New Providence, NJ, USA, 1992. ISBN 3-598-22401-X. 395 pp. LCCN QA76.88.S853 1992.

Meuer:1992:SAAb

[Meu92c] Hans-Werner W. Meuer, editor. Supercomputer ’92: Anwendungen, Architekturen, Trends: Seminar, Mannheim, 25.-27. Juni 1992, Informatik aktuell. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Meuer:1993:SAA

[Meu93] Hans-Werner Meuer, editor. Supercomputer ’93: Anwendungen, Architekturen, Trends: Seminar, Mannheim, 24.-26. Juni 1993, Informatik aktuell. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993.

Meuer:1995:SAA

[Meu95] H. W. (Hans W.) Meuer.

Supercomputer, 1995: Anwendungen, Architekturen, Trends, volume 13 of Fokus. K. G. Saur, München, Germany; New Providence, NJ, USA, 1995. ISBN 3-598-22412-5. 271 pp. LCCN QA76.88.S854 1995.

Miyata:1992:SSA

[MF92] K. Miyata and I. Fukai.

A supercomputer simulation of antenna radiation patterns and scattering fields by boundary-element method. In Anonymous [Ano92g], pages 176–178. ISBN ???? LCCN ????

Monahan:1993:EAS

[MF93] S. P. Monahan and W. L. Filippone. The exact adjoint for the SN/ diamond differenced Boltzmann/Spencer– Lewis equations in X-Y-Z geometry. In Kusters et al. [KSW93], pages 457–466. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Malaterre:1997:HAI

[MF97] G. Malaterre and J. Frechaux. How to assess and improve validity on a driving simulator? In Roller [Rol97], pages 137–144. ISBN 0-947719-88-1 (paperback). LCCN ????

Miyaji:1994:NCF

[MFK94] K. Miyaji, K. Fujii, and K. Karashima. Navier– Stokes computation of the flow over the delta wing with trailing-edge lateral blowing. In Anonymous [Ano94-75], pages 779–786. ISBN 0-947719-68-7. LCCN ????

Mueller-Gaertner:1995:CTN

[MG95] H.-W. Mueller-Gaertner. Current trends in neurosciences — implications for the Juelich program. In Herrmann et al. [HWP95], pages 449–455. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Martin:1994:CAA

[MGA94] A. Martin and M. C. Garcia-Alegre. Control architecture for autonomous mobile robots: Linguistic description of navigation and specific manoeuvring tasks. In Anonymous [Ano94-75], pages 273– 280. ISBN 0-947719-68-7. LCCN ????

Martinasso:2020:CPE

[MGS^{+}20] Maxime Martinasso, Miguel Gila, William Sawyer, Rafael Sarmiento, Guilherme Peretti-Pezzi, and Vasileios Karakasis. Cray programming environments within containers on Cray XC systems. Concurrency and Computation: Practice and Experience, 32 (20):e5543:1–e5543:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mukherjee:1994:EDP

[MH94] S. S. Mukherjee and M. D.

Hill. An evaluation of directory protocols for mediumscale shared-memory multiprocessors. In Anonymous [Ano94-134], pages 64–74. ISBN ???? LCCN ????

Mangun:1995:CEN

[MH95] G. R. Mangun and H.-J.

Heinze. Combining electrophysiology with neuroimaging in the study of human cognition. In Herrmann et al. [HWP95], pages 61–74. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Mehrotra:1996:EMA

[MH96] S. Mehrotra and L. Harrison. Examination of a memory access classification scheme for pointer-intensive and numeric programs. In ACM [ACM96], pages 133–

140. ISBN 0-89791-803-

7. LCCN QA76.5 I61 1996. ACM order number 415961.

Mattson:1998:OIT

[MH98] Timothy G. Mattson and Greg Henry. An overview of the Intel TFLOPS supercomputer. Intel Technology Journal, (Q1):12, 1998. ISSN 1535-766X. URL

                                        http://developer.intel.
                                        com/technology/itj/q11998/
                                        articles/art_1.htm;   http:
                                        //developer.intel.com/technology/
                                        itj/q11998/pdf/overview.
                                        pdf.

Michael:2018:EFS

[MH18] Scott Michael and Yun He. Editorial: Foreword to the special issue of the Cray User Group (CUG 2017). Concurrency and Computation: Practice and Experience, 30 (1):??, January 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Michaeli:1997:ITM

[MHE97] W. Michaeli, E. Haberstroh, and F. Ehrig. Injection transfer moulding of elastomers — simulation leads to better mould design. In Roller [Rol97], pages 475–

490. ISBN 0-947719-88-1 (paperback). LCCN ????

Malony:1991:TTV

[MHJ91] Allen D. Malony, David H.

Hammerslag, and David Jablonowski. TraceView: a trace visualization tool. Technical Report CSRD 1098, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1991. 22 pp.

Martonen:1997:CFT

[MHKY97] T. B. Martonen, D. Hwang, I. Katz, and Y. Yang. Cystic fibrosis: treatment with a supercomputer drug delivery model. In Power [Pow97], pages 359–364. CO-DEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic).

Moriarty:1984:EIL

[MHP84] K. J. M. Moriarty, M. Haraguchi, and C. Pangali. Efficient implementation of the SU(3) lattice gauge theory algorithm on the Fujitsu VP200 vector processor. Computer Physics Communications, 34(1–2):1–7, November/December 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/001046558490153X.

Masa:1994:HAN

[MHW94] Peter Masa, Klaas Hoen, and Hans Wallinga. A High-Speed analog neural processor. IEEE Micro, 14(3):40– 50, May/June 1, 1994. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

MITPress:1987:IJS

[M.I87] The International journal of supercomputer applications, 1987. ISSN 0890-2720. MIT Press, Cambridge, MA, USA.

Molvig:1993:DPN

[MI93] K. Molvig and R. A. Iannucci. 93SC044 digital physics. A new technology for fluid simulation. In Anonymous [Ano93-31], pages 77– 86. ISBN 0-947719-62-8. LCCN ????

Matsuoka:2001:TPE

[MI01] Satoshi Matsuoka and Shigeo Itou. Towards performance evaluation of highperformance computing on multiple Java platforms. Future Generation Computer Systems, 18(2):281–291, October 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/31/34/abstract.
                                        html.

MCNC-CC:1990:CCV

[Mic90] Microelectronics Center of North Carolina. Center for Communications, Research Triangle Park, NC, USA. Center for Communications videotape library: microelectronics, communications, supercomputing: Fall 1990, 1990. 20 pp.

Miklosko:1989:FAT

[Mik89] Jozef Miklosko. Fast algorithms and their implementation on specialized parallel computers, volume 5 of Special topics in supercomputing. North-Holland, Amsterdam, The Netherlands, 1989. ISBN 0-444-70141-9 (U.S.). xv + 261 pp. LCCN QA76.5 .F3541 1989.

Miki:1994:FVM

[Mik94] Y. Miki. A fast vectorized maze routing algorithm on a supercomputer. IE-ICE Transactions on Fundamentals of Electronics Communications and Computer Sciences, 77(12):2067–??, ???? 1994. CODEN IFE-SEX. ISSN 0916-8508 (print), 1745-1337 (electronic).

Miller:1987:NRS

[Mil87] Allan Ray Miller. Nonpreemptive run-time scheduling issues on a multitasked, multiprogrammed multiprocessor with dependencies, bidimensional tasks, folding and dynamic graphs. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development and Department of Computer Science, Urbana, IL, USA, 1987. v + 107 pp.

Miller:1988:SIC

[Mil88a] Warren F. Miller. Supercomputing: an indispensable component of applied research. Technical Report LA11392-MS; UC-705; CNSS papers 15, Center for National Security Studies, Los Alamos National Laboratory, Los Alamos, NM, USA, July 1988. viii + 4 pp. LA11392-MS. UC-705.

Milone:1988:MAL

[Mil88b] E. F. Milone. Modeling of asymmetric light curves of eclipsing binaries on the Cyber 205 supercomputer. Celestial mechanics, 45(1 / 3): 135–??, 1988. CODEN CLM-CAV. ISSN 0008-8714.

Miller:1990:IOB

[Mil90] Ethan L. Miller. Input/ output behavior of supercomputing applications: research project. Master of sciences, Plan II, University of California, Berkeley. Dept. of Electrical Engineering and Computer Sciences, Berkeley, CA, USA, 1990. 26 pp.

Miller:1991:IOB

[Mil91] Ethan L. Miller. Input/ output behavior of supercomputing applications. NASA contractor report NASA CR-187956; UCB/CSD 91/616, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, January 1991. 31 pp.

Miles:1993:BVP

[Mil93] Douglas Miles. Beyond vector processing: parallel programming on the Cray APP. In 1993 IEEE Compcon Spring (Feb 22–26 1993: San Francisco, CA, USA), pages 321– 328. IEEE, Piscataway, NJ, USA, 1993. ISBN 0-7803-1294-5. LCCN ???? IEEE catalog number 93CH3251-6.

Miller:1997:ALA

[Mil97a] N. L. Miller. An automated land analysis system (ALAS) for applications at a range of spatial scales: Watershed to global. In Delic and Wheeler [DW97], pages 169–176. ISBN 0-89871-378-1. LCCN ????

Milne:1997:RCC

[Mil97b] G. J. Milne. Reconfigurable custom computing as a supercomputer replacement. In IEEE [IEE97a], pages 260– 271. ISBN 0-8186-8068-7, 0-8186-8067-9, 0-8186-8069-5. LCCN ????

Miller:2017:PIT

[Mil17] Loren Miller. Product innovation through computational prototypes and supercomputing. Computing in Science and Engineering, 19(6):9–17, November/ December 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2017/06/mcs2017060009-
                                        abs.html.

MSI:1986:SAM

[Min86] Minnesota Supercomputer Institute. Supercomputing at Minnesota: carrying the tradition into tomorrow. Technical report, Minnesota Supercomputer Institute, Minneapolis, MN, USA, 1986. 31 pp.

MSI:1988:ARR

[Min88] Annual research report of the Minnesota Supercomputer Institute, 1988. Minnesota Supercomputer Institute, University of Minnesota, Minneapolis, MN, USA.

ML-OLA-PED:1992:UMS

[Min92] Minnesota.Legislature.Office of the Legislative Auditor and Program Evaluation Division. University of Minnesota supercomputing services. Minnesota document 92-0510, Program Evaluation Division, Office of the Legislative Auditor State of Minnesota, Saint Paul, MN, USA, October 1992. xiv + 48 + 5 pp.

Mirin:1988:PME

[Mir88] A. A. Mirin. Predicting multiprocessing efficiency on the Cray multiprocessors in a (CTSS) time-sharing environment/application to a 3-D magnetohydrodynamics code. Computers in Physics, 2(4):62–??, July 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.168304.

Mirin:1990:NER

[Mir90]  A. A. Mirin.  The Na-

tional Energy Research Supercomputer Center. International Journal of Supercomputer Applications, 4(3): 6–??, Fall 1990. CODEN IJSAE9. ISSN 0890-2720.

Miranker:1992:TIS

[Mir92] G. S. Miranker. Titan III supercomputer architectural overview. Computing: Archiv fur informatik und numerik, 48(1):39–60, March 1992. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Misztal:1990:RML

[Mis90] I. Misztal. Restricted maximum likelihood estimation of variance components in animal model using sparse matrix inversion and a supercomputer. Journal of dairy science, 73(1):163–??, January 1, 1990. CODEN JD-SCAE. ISSN 0022-0302.

Mitra:1988:MAM

[Mit88] Samir Girish Mitra. Measurementbased analysis of multiple errors and near-coincident fault discovery in a shared memory multiprocessor. Technical Report CSRD 743; UILU-ENG-88-8002, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. vii + 31 pp.

Mitropolski:1996:ISM

[Mit96] Y. I. Mitropolski. The interdependence of supercomputer and microprocessor development. In De Sario [De 96], pages 159–161. ISBN 0-7803-3110-9, 0-7803-3109-5, 0-7803-3111-7. LCCN ????

Mitchell:1998:SPS

[Mit98]  Bradley Mitchell.  Scal-

able platform services on the Intel TFLOPS supercomputer. Intel Technology Journal, (Q1):9, 1998. ISSN 1535-766X. URL

                                        http://developer.intel.
                                        com/technology/itj/q11998/
                                        articles/art_4.htm;   http:
                                        //developer.intel.com/technology/
                                        itj/q11998/pdf/sps.pdf.

Mahmassani:1990:MST

[MJH90] Hani S. Mahmassani, R. Jayakrishnan, and Robert Herman. Microscopic simulation of traffic in networks: Supercomputer experience. Journal of Computing in Civil Engineering, 4(1):1–??, January 1, 1990. CODEN JCCEE5. ISSN 0887-3801.

Marooney:1994:VPH

[MJRS94] C. J. Marooney, R. J. R.

Johns, M. Richardson, and M. Surridge. VECTIS: a parallel HPC combustion-flow simulation code. In Anonymous [Ano94-75], pages 735– 742. ISBN 0-947719-68-7. LCCN ????

Miller:1992:AFMa

[MK92a] Ethan L. Miller and Randy H.

Katz. An analysis of file migration in a Unix supercomputing environment. Technical Report UCB/CSD 92/712, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, November 1992. 12 pp. Supported in part by University Corporation for Atmospheric Research. S9128 Supported in part by NSF.

Miller:1992:AFMb

[MK92b] Ethan L. Miller and Randy H.

Katz. An analysis of file migration in a Unix supercomputing environment. NASA contractor report NASA CR-192908, Computer Science Division (EECS), University of California Berkeley, Berkeley, CA, USA, 1992. ?? pp. Distributed to depository libraries in microfiche. Shipping list no.:93-1028-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1993] 1 microfiche.

Machida:1993:CSV

[MK93] M. Machida and H. Kaburaki. Computer simulation of vortex motion in superconductors and superfluids. In Kusters et al. [KSW93], pages 271–279. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Moin:1997:TTS

[MK97] Parviz Moin and John Kim. Tackling turbulence with supercomputers. Scientific American, 276(1):62– 68, January 1997. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

                                        tronic). URL http://www.
                                        sciam.com/1997/0107issue/
                                        0197moin.html.

Murphy:2007:MAP

[MK07] R. C. Murphy and P. M. Kogge. On the memory access patterns of supercomputer applications: Benchmark selection and its implications. IEEE Transactions on Computers, 56(7): 937–945, July 2007. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&
                                        arnumber=4216292.

Misegades:1987:MFM

[MKB87] K. Misegades, L. Krause, and M. Booth. Microtasking of fluid mechanics codes on the Cray X-MP. American Society of Mechanical Engineers, Fluids Engineering Division (Publication) FED, 47:19–25, 1987. CODEN FEDSDL. ISSN 0888-8116.

Marshall:1990:VMS

[MKDY90] Robert Marshall, Jill Kempf, Scott Dyer, and Chieh-Cheng Yen. Visualization methods and simulation steering for a 3D turbulence model of Lake Erie. Computer Graphics, 24(2):89– 97, March 1990. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Morgan:1996:CCR

[MKfDA96] Dennis R. Morgan, Steven G.

Kratzer, and Institute for Defense Analyses. On a class of computationally-efficient, rapidly-converging, generalized NLMS algorithms. Technical report ????, Supercomputing Research Center : IDA, Lanham, MD, USA, May 1, 1996. 7 pp.

Madavan:1990:SAG

[MKG90] N. K. Madavan, P. Kelaita, and S. Gavali. Supercomputer applications in gas turbine flowfield simulation. International Journal of Supercomputer Applications, 4(2):

81–??, Summer 1990. CO-DEN IJSAE9. ISSN 0890-2720.

Martonen:1995:BAS

[MKHY95] T. B. Martonen, I. Katz, D. Hwang, and Y. Yang. Biomedical application of the supercomputer: targeted delivery of inhaled pharmaceuticals in diseased lungs. In Power and Hart [PH95], pages 241–248. ISBN 1-85312-321-8, 1-56252-245-0. LCCN R859.7.C65 I575 1995.

Maier:1997:PFT

[MKND97] R. S. Maier, Y. Kutsovsky, S. Nivarthi, and H. T. Davis. Pore-scale flow and transport calculations using the lattice Boltzmann method. In Delic and Wheeler [DW97], pages 259–264. ISBN 0-89871-378-1. LCCN ????

Milanesi:1993:GCT

[MKRI93] L. Milanesi, N. A. Kolchanov, I. B. Rogozin, and I. V. Ischenko. Genviewer: a computing tool for proteincoding regions prediction in nucleotide sequences. In Lim et al. [L^{+}93], pages 573–588. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Mummert:1996:FGP

[MKSF96] T. Mummert, C. Kosak, P. Steenkiste, and A. Fisher. Fine grain parallel communication on general purpose LANs. In ACM [ACM96], pages 341–349. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Martin:1989:SPN

[ML89] Joanne L. Martin and Stephen F.

Lundstrom, editors. Proceedings, supercomputing ’88 science and applications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-8923-4. LCCN QA76.5 .S894 1988. Two volumes. IEEE Computer Society order number 1923. IEEE order number 88TH0237-8.

Miften:1993:SQM

[ML93] M. M. Miften and E. W.

Larsen. A symmetrized quasidiffusion method for solving transport problems in multidimensional geometries. In Kusters et al. [KSW93], pages 707–717. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Martignon:1995:SIM

[ML95a] L. Martignon and K. B.

Laskey. Statistical inference methods for classifying higher order neural interactions. In Herrmann et al. [HWP95], pages 149–160. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Moore:1995:FLA

[ML95b] Will R. Moore and Wayne Luk, editors. Field-programmable logic and applications: 5th international workshop, FPL ’95, Oxford, United Kingdom, August 29–September 1, 1995: proceedings, volume 975 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. CO-DEN LNCSD9. ISBN 3-540-60294-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK7895.G36 I48 1995.

McCandless:1997:RAH

[ML97] B. C. McCandless and A. Lumsdaine. The role of abstraction in high-performance computing. Lecture Notes in Computer Science, 1343: 203–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malony:1990:TAPb

[MLR90a] Allen D. Malony, John L. Larson, and Daniel A. Reed. Tracing application program execution on the Cray X-MP and Cray 2. IEEE, Piscataway, NJ, USA, 1990. ISBN 0-8186-2056-0. 60–73 pp. LCCN ???? IEEE catalog number 90CH2916-5.

Malony:1990:TAPa

[MLR90b] Allen Davis Malony, John Leonard Larson, and Daniel A. Reed. Tracing application program execution on the Cray X-MP and Cray 2. Technical Report CSRD 985, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1990. 25 pp.

Maerean:2020:TAE

[MLWC20] S. Maerean, E. K. Lee, H.-F.

Wen, and I. Chung. Transformation of application enablement tools on CORAL systems. IBM Journal of Research and Development, 64(3/4):16:1–16:12, May/ July 2020. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Min:2010:SIA

[MLY10] Geyong Min, Keqiu Li, and Laurence T. Yang. Special issue: Advances in high-performance computing and communications. Concurrency and Computation: Practice and Experience, 22 (4):395–397, March 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

McNamara:1990:CSD

[MM90] B. McNamara and K. J. M.

Moriarty. Computer-aided software development tools for the supercomputer environment. International Journal of Supercomputer Applications, 4(4):56–??, Fall 1990. CODEN IJSAE9. ISSN 0890-

Marino:1991:PCI

[MM91a] D. Marino and G. Mastronardi, editors. Parallel computing: International workshop — September 1991, Trani, Italy. International Society for Mini and Microcomputers, Zurich, Switzerland, 1991. ISBN 0-88986-147-1. LCCN ????

Mitchell:1991:VOA

[MM91b] Michael J. Mitchell and J. Andrew McCammon. Vector optimization of AMBER 3.0 on the NEC SX-2/400 supercomputer. Computers and Chemistry, 15(1):79–??, ???? 1991. CODEN COCHDK. ISSN 0097-8485.

Makowitz:1993:CVP

[MM93a] H. Makowitz and G. L. Mesina. CERBERUS VER.02.6, a performance enhanced version of RELAP5/MOD2.5 suitable for parallel computing. In Kusters et al. [KSW93], pages 147–161. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Manning:1993:AAE

[MM93b] Serge M. Manning and David G. Meyer. Analysis of asynchronous execution streams with I-caching in massively parallel systems. Journal of Parallel and Distributed Computing, 19(3): 279–??, November 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mallinckrodt:1994:BRR

[MM94a] A. John Mallinckrodt and Susan McKay. Book review: Rubin H. Landau and Paul Fink, Jr. and Geert Wenes, A Scientist’s and Engineer’s Guide to Workstations and Supercomputers. Computers in Physics, 8(1):68– ??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.4823263.

Misra:1994:DSA

[MM94b] M. Misra and R. Moona. Design of systolic arrays for QR decomposition. In Balakrishnan [Bal94], pages 247–255. ISBN 0-07-462044-4. LCCN ????

Murthy:1994:TEA

[MM94c] K. N. Balasubramanya Murthy and C. Siva Ram Murthy. Two-sided elimination algorithm for parallel solution of linear equations using Householder reductions. In Balakrishnan [Bal94], pages 381– 382. ISBN 0-07-462044-4. LCCN ????

Murthy:1994:TSEa

[MM94d] K. N. Balasubramanya Murthy and C. Siva Ram Murthy. Two-sided elimination algorithm for parallel solution of linear equations using Householder reductions. In Balakrishnan [Bal94], pages 381– 382. ISBN 0-07-462044-4. LCCN ????

Moreira:2000:JPH

[MMG^{+}00] J. E. Moreira, S. P. Midkiff, M. Gupta, P. V. Artigas, M. Snir, and R. D. Lawrence. Java programming for high-performance numerical computing. IBM Systems Journal, 39(1):21–56, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http:

//www.research.ibm.com/
journal/sj/391/moreira.
html.

Montella:2018:MBP

[MMG^{+}18] Raffaele Montella, Livia Marcellino, Ardelio Galletti, Diana Di Luccio, Sokol Kosta, Giuliano Laccetti, and Giulio Giunta. Marine bathymetry processing through GPGPU virtualization in high performance cloud computing. Concurrency and Computation: Practice and Experience, 30(24):e4895:1– e4895:??, December 25, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Munz:1993:NSF

[MMHM93] C.-D. Munz, W. Maschek, S. Hirmer, and K. Mederacke. Numerical simulation of fluid flow around obstacles. In Kusters et al. [KSW93], pages 791–?? ISBN 3-923704-11-9. LCCN ???? Two volumes.

Moon:1997:ANN

[MMK97] S.-K. Moon, K.-S. Maeng, and Y.-J. Kim. Application of neural networks in development of vehicle durability testing method. In Roller [Rol97], pages 223– 228. ISBN 0-947719-88-1 (paperback). LCCN ????

Melnikov:1996:DES

[MMR96] V. A. Melnikov, Y. I.

Mitropolski, and G. V. Reznikov. Designing the Electronica SS BIS supercomputer. IEEE transactions on components, packaging, and manufacturing technology. Part A: a publication of the IEEE Components, Packaging, and Manufacturing Technology Society, 19(2): 151–??, June 1996. CODEN IMTAEZ. ISSN 1070-9886.

Martin:1993:EDP

[MMRL93] W. R. Martin, A. Majumdar, J. A. Rathkopf, and M. Litvin. Experiences with different parallel programming paradigms for Monte Carlo particle transport leads to a portable toolkit for parallel Monte Carlo. In Kusters et al. [KSW93], pages 418– 432. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Martin:1986:SPE

[MMW86] Joanne L. Martin and Dieter Mueller-Wichards. Supercomputer performance evaluation: status and directions. Research report RC 11893 (#53473), International Business Machines Inc., Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1986. 13 pp.

Miyama:1991:SSA

[MN91] Shoken M. Miyama and Mikio Nagasawa, editors. Symposium of Supercomputing Astronomy and Astrophysics in Japan: NAO, Mitaka, 1991, number 2 in NAO-TAP reports. National Astronomical Observatory, Mitaka, Japan, 1991. LCCN QB51.3.E43S95 1991. Cover title: Proceedings of Supercomputing Astronomy and Astrophysics Symposium in Japan.

Miller:1994:MDB

[MNB94] W. M. Miller, W. A. Najjar, and A. P. W. Boehm. A model for dataflow based vector execution. In Anonymous [Ano94-134], pages 11– 22. ISBN ???? LCCN ????

Martin:1986:MCP

[MNR86] William R. Martin, Paul F. Nowak, and James A. Rathkopf. Monte Carlo photon transport on a vector supercomputer. IBM Journal of Research and Development, 30(2):193–202, March 1986. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Maubert:1993:CBV

[MNV93] L. Maubert, A. Nouri, and T. Vergnaud. Criticality benchmarks validation of the Monte Carlo code TRIPOLI-2. In Kusters et al. [KSW93], pages 296–302. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Magoules:2009:GRM

[MNY09] F. (Frédéric) Magoulès, Thi-Mai-Huong Nguyen, and Lei Yu. Grid resource management: toward virtual and services compliant grid computing. Chapman and Hall/CRC numerical analysis and scientific computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-7404-0 (hardcover). xxi + 300 pp. LCCN QA76.9.C58 M34 2009. URL http:/

/www.loc.gov/catdir/toc/
fy0905/2009417808.html.

Meneses:2015:UMO

[MNZ^{+}15] Esteban Meneses, Xiang Ni, Gengbin Zheng, Celso L. Mendes, and Laxmikant V. Kale. Using migratable objects to enhance fault tolerance schemes in supercomputers. IEEE Transactions on Parallel and Distributed Systems, 26(7):2061– 2074, July 2015. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/trans/
                                        td/2015/07/06862914-abs.
                                        html.

Mendez:1988:JSA

[MO88] R. (Raul) Mendez and Steven A. Orszag. Japanese supercomputing: architecture, algorithms, and applications, volume 36 of Lecture notes in engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96765-6. iv + 161 pp. LCCN QA76.5 .J331 1988.

ePub:2012:SAI

[Mob12]  Mobi PDF ePub.  Sys-

tem administration of the IBM Watson supercomputer. Linux Journal, 2012(216): 1:1–1:??, April 2012. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Mohamed:1994:PCA

[Moh94] S. T. Mohamed. The proposed computer algorithm for computing the confidence limits of a repeatedly measured factor. In Anonymous [Ano94-75], pages 235–240. ISBN 0-947719-68-7. LCCN ????

Moitra:1993:AGM

[Moi93] A. Moitra. 93SC026 algebrain geometry modelling and grid generation for automobiles. In Anonymous [Ano93-31], pages 61–68. ISBN 0-947719-62-8. LCCN ????

Monagan:1988:AMC

[Mon88]  Michael Monagan.  An-

nouncement of Maple 4.0 for the Cray 2. Maple Newsletter, 0(2):??, January 1988. ISSN 1074-

3790. URL http://www.can.
nl/Systems_and_Packages/
Per_Purpose/General/Maple/
mtn/mtn2.html.

Monnier:1993:CCF

[Mon93] A. Monnier. Criticality calculations for fuel storage facilities with irradiated LMFBR fuel, using TRIMARAN-1. In Kusters et al. [KSW93], pages 766–778. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Moore:2006:MMS

[Moo06] Samuel K. Moore. Multimedia monster [supercomputer on a single chip]. IEEE Spectrum, 43(1):20– 23, January 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2008:MBN

[Moo08] Samuel K. Moore. Multicore is bad news for supercomputers. IEEE Spectrum, 45 (11):15, November 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:2011:CSP

[Moo11] Samuel K. Moore. China’s supercomputing prowess [the data]. IEEE Spectrum, 48 (4):64, April 2011. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Matsuda:1994:FPA

[MOOK94] H. Matsuda, G. J. Olsen, R. Overbeek, and Y. Kaneda. Fast phylogenetic analysis on a massively parallel machine. In Anonymous [Ano94-134], pages 297–302. ISBN ???? LCCN ????

Moreira:1992:MON

[Mor92a] Jose Eduardo Moreira. Multiple Omega networks for parallel processing. Technical Report CSRD 1214, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1992. 18 pp.

Morley:1992:EOS

[Mor92b] E. Morley. Empowering the operator with supercomputer technology. In Kompass et al. [KWW92], pages 85–90. ISBN 0-931682-34-7. LCCN ????

Mortensen:1992:JPC

[Mor92c] Paul Mortensen. Japanese physicists command powerful supercomputer resources. Computers in Physics, 6(4):

339–345, July 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Moreira:2001:BGM

[Mor01] Jose E. Moreira. Blue Gene:

a massively parallel system. Lecture Notes in Computer Science, 2073:10–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2073/20730010.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2073/20730010.
pdf.

Mount:1989:ETS

[Mou89] Ellis Mount. End-user training for sci-tech databases, volume 10(1) of Science and technology libraries. Haworth Press, New York, NY, USA, 1989. 128 pp.

Mount:1990:ETS

[Mou90] Ellis Mount. End-user training for sci-tech databases. Haworth Press, New York, NY, USA, 1990. ISBN 0-86656-963-4. 128 pp. LCCN Z 699.5 S3 E525 1990.

McKee:1996:DED

[MOWW96] S. McKee, C. W. Oliver, W. A. Wulf, and K. L. Wright. Design and evaluation of dynamic access ordering hardware. In ACM [ACM96], pages 125–132. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Malitz:1984:SSI

[MP84] Isaac Malitz and Robert Peters. The supercomputer snooper IBM-PC. DATA-MOST, Chatsworth, CA, USA, 1984. ISBN 0-8359-7148-1 (paperback). 184 pp. LCCN ????

Midkiff:1987:CASa

[MP87a] Samuel P. Midkiff and David A. Padua. Compiler algorithms for synchronization. Technical Report CSRD-595, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 51 pp.

Midkiff:1987:CASb

[MP87b] Samuel P. Midkiff and David A. Padua. Compiler algorithms for synchronization. Technical Report CSRD-643, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 52 pp.

Malony:1988:EAU

[MP88] Allen Davis Malony and Joseph R. Pickert. An environment architecture and its use in performance data analysis. Technical Report CSRD 829, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1988. 17 pp.

Midkiff:1990:ICO

[MP90] Samuel P. Midkiff and David A. Padua. Issues in the compile-time optimization of parallel programs. Technical Report CSRD 993, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1990. 26 pp.

Meyer:1991:FSG

[MP91a] Gerard G. L. Meyer and Louis J. Podrazik. A Frank-Wolfe / gradient projection method for large scale optimization. Technical report SRC-TR-91-034, Supercomputing Research Center: IDA, Lanham, MD, USA, January 28, 1991. iii + 31 pp.

Meyer:1991:PFG

[MP91b] Gerard G. L. Meyer and Louis J. Podrazik. A parallel Frank-Wolfe/gradient projection method for optimal control. Technical report SRC-TR-91-043, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1991. i + 15 pp.

Meyer:1991:SGP

[MP91c] Gerard G. L. Meyer and Louis J. Podrazik. A scaled gradient projection method for large scale optimization. Technical report SRC-TR-91-051, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1991. ii + 26 pp.

Midkiff:1991:CFS

[MP91d] Samuel P. Midkiff and David A. Padua. A comparison of four synchronization optimization techniques. Technical Report CSRD 1135, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 22 pp.

Moltedo:1991:STS

[MP91e] L. Moltedo and P. Palamidese. Software tools for supercomputing graphics. In Marino and Mastronardi [MM91a], pages 355–358. ISBN 0-88986-147-1. LCCN ????

Mulford:1992:SPA

[MP92] Carolyn Mulford and Douglas Price, editors. Summaries of proceedings: Access is the key: the Seventh Annual FLICC Forum on Federal Information Policies, March 20, 1990; building information superhighways, supercomputing networks and libraries: the Eight Annual FLICC Forum on Federal Information Policies, March 21, 1991. Federal Library and Information Center Committee, Library of Congress, Washington, DC, USA, 1992.

Malard:1994:EST

[MP94] J. Malard and C. C. Paige.

Efficiency and scalability of two parallel QR factorization algorithms. In IEEE [IEE94c], pages 615–622. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Midkiff:1989:CPU

[MPC89] Samuel P. Midkiff, David A.

Padua, and Ronald Gary Cytron. Compiling programs with user parallelism. Technical Report CSRD 911, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1989. 27 pp.

Meijer:1996:PMC

[MPG96] P. M. Meijer, S. Poedts, and J. P. Goedbloed. Parallel magnetohydrodynamics on the Cray T3D. Future Generation Computer Systems, 12 (4):307–323, December 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Martinez:1993:OCL

[MPH93] J. M. Martinez, C. Parey, and M. Houkari. Optimal control learning with artificial neural networks. In Kusters et al. [KSW93], pages 443– 453. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Moore:1987:CSV

[MPSB87] B. Moore, A. Padegs, R. Smith, and W. Bucholz. Concepts of the System/370 vector architecture. In The 14th Annual International Symposium on Computer Architecture: June 2–5, 1987, Pittsburgh, Pennsylvania: conference proceedings, pages 282– 293. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-89791-233-0. ISSN 0163-5964 (ACM), 0884-7495 (IEEE). LCCN QA76.9.A73 S97 1987. Also in Computer Architecture News, v. 15, no. 2.

Magoules:2012:CCD

[MPT12] F. (Frédéric) Magoulès, Jie Pan, and Fei Teng. Cloud Computing: Data-Intensive Computing and Scheduling. Chapman and Hall/CRC numerical analysis and scientific computing series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4665-0782-9 (hardback). ???? pp. LCCN QA76.585 .M34 2012.

Moriarty:1986:QFT

[MR86] K. J. M. Moriarty and C. Rebbi. Quantum field theory and supercomputers. Computer Physics Communications, 40(2–3): 181–188, June 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465586901074.

Moriarty:1987:LSQ

[MR87] K. J. M. Moriarty and C. Rebbi. Large-scale quantum field theory calculations on supercomputers. Computer Physics Communications, 47(1):75– 82, October 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465587900671.

McRae:1990:SSS

[MR90a] Gregory J. McRae and Armistead G. Russell. Smog, supercomputers and society. Computers in Physics, 4(3): 227–??, May 1990. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.4822910.

Moriarty:1990:SMS

[MR90b] K. J. M. Moriarty and C. Rebbi. Supercomputer methods for the solution of fundamental problems of particle physics. International Journal of Supercomputer Applications, 4(1):10– ??, Spring 1990. CODEN IJSAE9. ISSN 0890-2720.

Mueller:1995:EHP

[MR95] A. Mueller and R. Ruehl. Extending High Performance Fortran for the support of unstructured computations. In ACM [ACM95a], pages 127– 136. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Morales:1995:IKP

[MRAR95] D. Morales, J. Roda, F. Almeida, and C. Rodriguez. Integral knapsack problems: Parallel algorithms and their implementations on distributed systems. In ACM [ACM95a], pages 218–226. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Munir:2012:HPE

[MRGR12] Arslan Munir, Sanjay Ranka, and Ann Gordon-Ross. Highperformance energy-efficient multicore embedded computing. IEEE Transactions on Parallel and Distributed Systems, 23(4):684– 700, April 2012. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Mycek:2017:DPB

[MRL^{+}17] Paul Mycek, Francesco Rizzi, Olivier Le Maître, Khachik Sargsyan, Karla Morris, Cosmin Safta, Bert Debusschere, and Omar Knio. Discrete a priori bounds for the detection of corrupted PDE solutions in exascale computations. SIAM Journal on Scientific Computing, 39(1): C1–C28, ???? 2017. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Malony:1987:MPM

[MRM87] Allen Davis Malony, Daniel A.

Reed, and Patrick J. McGuire. MPF: a portable message passing facility for shared memory multiprocessors. Technical Report CSRD-651, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 17 + [8] pp.

Miranker:1988:SCS

[MRS88] Glen S. Miranker, Jon Rubinstein, and John Sanguinetti. Squeezing a Cray-class Supercomputer into a Single-user Package. IEEE, New York, NY, USA, 1988. ISBN 0-8186-0828-5. 452–456 pp. LCCN ???? IEEE Service Cent (cat n 88CH2539-5). Piscataway, NJ, USA.

Montagne:1994:MOG

[MRSB94] E. Montagne, M. Rukoz, R. Surós, and F. Breant. Modeling optimal granularity when adapting systolic algorithms to transputer based supercomputers. Parallel Computing, 20 (5):807–814, May 11, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
parco/cas_sub/browse/browse.
cgi?year=1994&volume=20&
issue=5&aid=840.

Morkoc:1984:HST

[MS84] H. Morkoc and P. M. Solomon. The HEMT: a superfast transistor: an experimental GaAs–AlGoAs device switches in picoseconds and generates little heat. This is just what supercomputers need. IEEE Spectrum, 21(2): 28–35, February 1984. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Meier:1988:BCG

[MS88] Ulrike Meier and Ahmed Sameh. The behavior of conjugate gradient algorithms on a multivector processor with a hierarchical memory. Technical Report CSRD 758, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 41 pp.

Munz:1993:HRH

[MS93] C.-D. Munz and L. Schmidt.

High resolution hydrodynamic schemes and their implementation on vector computers. In Kusters et al. [KSW93], pages 135–146. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Margrave:1994:ESA

[MS94a] F. W. Margrave and D. W.

Seward. Establishing the safety argument for the use of programmable electronics systems in transport. In Anonymous [Ano94-75], pages 519–526. ISBN 0-947719-68-7. LCCN ????

Meo:1994:AIM

[MS94b] S. Meo and M. Scarano.

An advanced implementation of a microprocessor anti-skid control for electrical traction drive. In Anonymous [Ano94-75], pages 155–160. ISBN 0-947719-68-7. LCCN ????

Moon:1994:ARS

[MS94c] B. Moon and J. Saltz. Adaptive runtime support for direct simulation Monte Carlo methods on distributed memory architectures. In IEEE [IEE94c], pages 176–183. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Moyer:1994:PSP

[MS94d] S. A. Moyer and V. S. Sunderam. PIOUS: a scalable parallel I/O system for distributed computing environments. In IEEE [IEE94c], pages 71–78. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Mahnke:1996:APF

[MS96] R. Mahnke and M. Seemann. Aggregation phenomena in a flow channel. In Wolf et al. [WSB96], pages 323– 328. ISBN 981-02-2635-7. LCCN ????

Maetani:1997:NAF

[MS97] N. Maetani and S. Shimojima. Numerical analysis of flow and heat transfer of finned heat exchanger used in large trucks. In Roller [Rol97], pages 451–458. ISBN 0-947719-88-1 (paperback). LCCN ????

Moreira:2007:BGS

[MSA^{+}07] José E. Moreira, Valentina Salapura, George Almasi, Charles Archer, Ralph Bellofatto, Peter Bergner, Randy Bickford, Mathias Blumrich, José R. Brunheroto, Arthur A. Bright, Michael Brutman, José G. Castaños, Dong Chen, Paul Coteus, Paul Crumley, Sam Ellis, Thomas Engelsiepen, Alan Gara, Mark Giampapa, Tom Gooding, Shawn Hall, Ruud A.

Haring, Roger Haskin, Philip Heidelberger, Dirk Hoenicke, Todd Inglett, Gerrard V. Kopcsay, Derek Lieber, David Limpert, Pat Mc-Carthy, Mark Megerian, Mike Mundy, Martin Ohmacht, Jeff Parker, Rick A. Rand, Don Reed, Ramendra Sahoo, Alda Sanomiya, Richard Shok, Brian Smith, Gordon G. Stewart, Todd Takken, Pavlos Vranas, Brian Wallenfelt, Michael Blocksome, and Joe Ratterman. The Blue Gene/L supercomputer: a hardware and software story. International Journal of Parallel Programming, 35(3):181–206, June 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=35&
issue=3&spage=181.

Mangione-Smith:1991:PCI

[MSAD91] William Mangione-Smith, Santosh G. Abraham, and Edward S. Davidson. A performance comparison of the IBM RS/6000 and the Astronautics ZS-1. Computer, 24 (1):39–46, January 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mangione-Smith:1992:RRP

[MSAD92] W. Mangione-Smith, S. G.

Abraham, and E. S. Davidson. Register requirements of pipelined processors. In ACM [ACM92b], pages 260–271. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

FAD-OLA-SM:19xx:MSC

[MSCxx] Minnesota Supercomputer Center, Inc. financial audit for the two years ended June 30, 19xx. Financial Audit Division, Office of the Legislative Auditor State of Minnesota, Saint Paul, MN, USA.

McNeil:1994:NNB

[MSGW94] A. R. McNeil, T. Sarkodie-Gyan, and A. Watson. A neural network based recognition scheme for the classification of planar shapes. In Anonymous [Ano94-75], pages 535–

542. ISBN 0-947719-68-7.

LCCN ????

McCurdy:2002:FDS

[MSK^{+}02] C. William McCurdy, Horst D. Simon, William T. C. Kramer, Robert F. Lucas, William E. Johnston, and David H. Bailey. Future directions in scientific supercomputing for computational physics. Computer Physics Communications, 147(1–2): 34–39, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S001046550200200X.

McIntosh-Smith:2020:BFG

[MSPPD20] Simon McIntosh-Smith, James Price, Andrei Poenaru, and Tom Deakin. Benchmarking the first generation of production quality Armbased supercomputers. Concurrency and Computation: Practice and Experience, 32 (20):e5569:1–e5569:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moriarty:1993:EMC

[MSTK93] K. J. M. Moriarty, S. Sanielevici, T. Trappenberg, and D. W. Kuba. Efficiently microtasked Cray Y-MP C90 version of the Kuba-Moriarty SU(3) gauge theory simulation program. Computer Physics Communications, 76 (1):87–97, June 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Moncrieff:1991:PPU

[MSW91] D. Moncrieff, V. R. Saunders, and S. Wilson. Parallel processing using macro-tasking in a multi-job environment on a Cray Y-MP computer. Parallel Computing, 17(6-7):733– 750, September 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Mattson:1996:TSA

[MSW96] T. G. Mattson, D. Scott, and S. Wheat. A TeraFLOP supercomputer in 1996: The ASCI TFLOP system. In IEEE [IEE96c], pages 84–93. ISBN 0-8186-7255-2, 0-8186-7257-9. LCCN ????

Martorell:2005:BGP

[MSW^{+}05] X. Martorell, N. Smeds, R. Walkup, J. R. Brunheroto, G. Almási, J. A. Gunnels, L. DeRose, J. Labarta, F. Escalé, J. Giménez, H. Servat, and J. E. Moreira. Blue Gene/L performance tools. IBM Journal of Research and Development, 49 (2/3):407–424, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

                                        //www.research.ibm.com/
                                        journal/rd/492/martorell.
                                        pdf.

Matsen:1986:SAA

[MT86] F. A. Matsen and T. Tajima, editors. Supercomputers— Algorithms, Architectures, and Scientific Computation. University of Texas Press, Austin, TX, USA, 1986. ISBN 0-292-70388-0. vi + 480 pp. LCCN QA76.5.S8945 1986.

Matsumoto:1991:DAB

[MT91] A. Matsumoto and T. Tsuda. Dependence analysis between pointer references in Pascal. In Anonymous [Ano91q], pages 28–37. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Morton:1996:MDO

[MT96] Donald J. Morton and John M. Tyler. Minimizing development overhead with partial parallelization. IEEE parallel and distributed technology: systems and applications, 4(3):15–24, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1996/pdf/p3015.
pdf;   http://www.computer.
org/concurrency/pd1996/
p3015abs.htm.

Morrow:1997:ICG

[MT97] D. Morrow and J. Thorp. Implementation of a conjugate gradient solver for MOD-FLOW on the Cray T3D. In Delic and Wheeler [DW97], pages 353–364. ISBN 0-89871-378-1. LCCN ????

Mikkonen:2013:CCI

[MT13] Tommi Mikkonen and Antero Taivalsaari. Cloud computing and its impact on mobile software development: Two roads diverged. The Journal of Systems and Software, 86(9): 2318–2320, September 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0164121213000241.

Matsen:1988:SAA

[MTH88] F. A. Matsen, T. Tajima, and Roger Haydock. Supercomputers: Algorithms, architectures, and scientific computation. Computers in Physics, 2(4):81–??, July 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4822755.

Michaels:1993:CAG

[MTHP93] G. S. Michaels, R. Taylor, R. Hagstrom, and M. Price. Comparative analysis of genomic data: a global look at structural and regulatory features. In Lim et al. [L^{+}93], pages 297–308. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Maldonado:1993:PIN

[MTK93] G. I. Maldonado, P. J. Turinsky, and D. J. Kropaczek. PWR in-core nuclear fuel management optimization utilizing nodal (non-linear NEM) generalized perturbation theory. In Kusters et al. [KSW93], pages 787– 798. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Mounes-Toussi:1994:ECO

[MTLL94] F. Mounes-Toussi, D. J. Lilja, and Z. Li. An evaluation of a compiler optimization for improving the performance of a coherence directory based cache coherence mechanism.

In Anonymous [Ano94-134], pages 75–84. ISBN ???? LCCN ????

Miura:1983:FVP

[MU83] K. Miura and K. Uchida.

FACOM vector processing system: VP100/200. In ?. Kwalik, editor, Proc. NATO Advanced Research Work on High Speed Computing (June), volume F7 of NATO ASI Series, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN ???? LCCN ???? Revised version reprinted in [Hwa84, pp. 59–73].

Mergen:2006:VHP

[MUKX06] Mark F. Mergen, Volkmar Uhlig, Orran Krieger, and Jimi Xenidis. Virtualization for high-performance computing. Operating Systems Review, 40(2):8–11, April 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Mulligan:1996:WSW

[Mul96] Jeffrey B. Mulligan. When are supercomputers worth the bother? Behavior research methods, instruments, and computers, 28(2):239– ??, ???? 1996. CODEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Mullins:2008:CM

[Mul08] J. Mullins. The Church of Microsoft. IEEE Spectrum, 45 (3):11–12, March 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mun:2004:MSS

[Mun04] Youngsong Mun. Modeling and simulation in supercomputing and telecommunications. Future Generation Computer Systems, 20(2): 179–180, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Murao:1991:VPS

[Mur91a]  H. Murao.  Vector pro-

cessing in symbolic determinant expansion on supercomputer. In Anonymous [Ano91q], pages 145–154. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Muraoka:1991:PDB

[Mur91b] Y. Muraoka. Panel discussion on benchmarking. In Anonymous [Ano91q], page 144. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Murray:1997:SSS

[Mur97] Charles J. Murray. The Supermen: the Story of Seymour Cray and the Technical Wizards Behind the Supercomputer. John Wiley and Sons, Inc., New York, NY, USA, 1997. ISBN 0-471-04885-2 (cloth). vii + 232 pp. LCCN TK7885.22.C73M87 1997.

Murphy:2006:EHP

[Mur06] Tom Murphy. Education:

High-performance computing in community colleges? IEEE Distributed Systems Online, 7(4):??, April 2006. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
ds/2006/04/o4003.pdf.

Murphy:2007:HPC

[Mur07] Tom Murphy. High-performance computing in high schools? IEEE Distributed Systems Online, 8(8):??, August 2007. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
ds/2007/08/mds2007080002.
pdf.

Murphy:2010:FOM

[Mur10]  K. Murphy.  Frank Op-

penheimer, the man who made science fun [review of “Something Incredible Wonderful Happens: Frank Oppenheimer and the World He Made Up” (Gell-Man, M.; 2009)]. IEEE Spectrum, 47 (3):25, March 2010. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Muthukrishnan:1994:OMV

[Mut94] C. R. Muthukrishnan. One man’s view of Computer Science. In Balakrishnan [Bal94], pages 256–?? ISBN 0-07-462044-4. LCCN ????

Murali:2016:QAF

[MV16] Prakash Murali and Sathish Vadhiyar. Qespera: an adaptive framework for prediction of queue waiting times in supercomputer systems. Concurrency and Computation: Practice and Experience, 28 (9):2685–2710, June 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

MouraSilva:1994:CSA

[MVS94] L. Moura Silva, B. Veer, and J. G. Silva. Checkpointing SPMD applications on transputer networks. In IEEE [IEE94c], pages 694– 701. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Mills:1981:CPD

[MW81] A. B. Mills and L. A. Wood. Cray-1: a powerful delivery system for engineering software. Advances in Engineering Software, 3(2):62– 66, April 1981. CODEN AESODT. ISSN 0141-1195, 0965-9978.

Madhavji:1982:CP

[MW82] N. H. Madhavji and I. R.

Wilson. Cray Pascal. ACM SIGPLAN Notices, 17(6):1– 14, June 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCammon:1988:SBP

[MW88] J. Andrew McCammon and George L. Wilcox, editors. Supercomputing in biology: proteins, nucleic acids, and water, selected papers from the Minnesota Conference, September 13–16, 1987, Minneapolis, Minnesota, USA, volume 1(4) of Journal of computer-aided molecular design. ESCOM, Leiden, The Netherlands, 1988. ISSN 0920-654x.

Magavi:1995:DIH

[MWB95] Sunil Magavi, Johnny Wong, and Prakash Bodla. Design and implementation of heterogeneous distributed multimedia system using Mosaic GSQL. Software—Practice and Experience, 25(11): 1223–1241, November 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Morton:1995:LLP

[MWO95] Don Morton, Kefei Wang, and David O. Ogbe. Lessons learned in porting Fortran/ PVM code to the Cray T3D. IEEE parallel and distributed technology: systems and applications, 3(1): 4–11, Spring 1995. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1995/pdf/h10004.
pdf;   http://www.computer.
org/concurrency/pd1995/
p1004abs.htm.

Martin:2018:HWP

[MWRK18] Steven Martin, Cary Whitney, David Rush, and Matthew Kappel. How to write a plugin to export job, power, energy, and system environmental data from your Cray XC system. Concurrency and Computation: Practice and Experience, 30 (1):??, January 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Myers:1986:GCS

[Mye86] W. Myers. Getting the cycles out of a supercomputer. Computer, 19(3):89–92, 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1992:S

[Mye92a] Ware Myers. Supercomputing 91. Computer, 25(1): 87–93, January 1992. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1992:SSR

[Mye92b]  Ware Myers.  Supercom-

puting 91: Special report: High-performance computing as a window into the future. Computer, 25(1):87– 90, January 1992. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1996:RIS

[Mye96] Ware Myers. On the road to the information superhighway. Computer, 29(4): 71–??, April 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marienhagen:1995:MCS

[MZ95] J. Marienhagen and A. Zippelius. Monte Carlo simulation of synaptic transmission at spinule- and nonspinule synapses. In Herrmann et al. [HWP95], pages 243–248. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Nodomi:1995:HPV

[N^{+}95] A. Nodomi et al. Hardware performance of the VPP500 parallel supercomputer. In Dongarra et al. [D^{+}95], pages 103–120. ISBN 0-444-82163-5. ISSN 0927-5452. LCCN QA76.88.H55 1995.

Noelting:1997:DPR

[NAAW97] S. Noelting, A. Alajbegovic, A. Anagost, and M. Wessels. A digital physics resoltion study of the flow over the ASMO-II body. In Roller [Rol97], pages 423– 430. ISBN 0-947719-88-1 (paperback). LCCN ????

Nagel:1988:UMC

[Nag88] Wolfgang E. Nagel. Using multiple CPUs for problem solving: Experiences in multitasking on the Cray X-MP/ 48. Parallel Computing, 8 (1-3):223–230, October 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nagel:1990:EAC

[Nag90] W. E. Nagel. Exploiting autotasking on a Cray Y-MP. an improved software interface to multitasking. Parallel Computing, 13(2):225– 233, February 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Nagy:1994:FEM

[Nag94] L. I. Nagy. Finite element method and computer applications in a new, faster design process. In Anonymous [Ano94-75], pages 915–937. ISBN 0-947719-68-7. LCCN ????

Nagasawa:1996:WBS

[Nag96a] M. Nagasawa. What to be supercomputers. Denshi Joho Tsushin Gakkai shi = The journal of the Institute of Electronics, Information, and Communication Engineers, 79(1):49–??, ???? 1996. ISSN 0913-5693.

Nagatani:1996:CBT

[Nag96b] T. Nagatani. Car bunching and traffic jams in cellular automation models. In Wolf et al. [WSB96], pages 57–72. ISBN 981-02-2635-7. LCCN ????

Nagel:1996:PHV

[Nag96c] K. Nagel. Particle hopping vs. fluid-dynamical models for traffic flow. In Wolf et al. [WSB96], pages 41–56. ISBN 981-02-2635-7. LCCN ????

Naik:1994:PNP

[Nai94] V. K. Naik. Performance of NAS Parallel Application Benchmark on IBM SP1. In IEEE [IEE94c], pages 121– 128. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Nalwa:1994:BCV

[Nal94] V. S. Nalwa. The basis of computer vision. In Balakrishnan [Bal94], pages 130– 144. ISBN 0-07-462044-4. LCCN ????

Nandakumar:1986:PPS

[Nan86] N. R. Nandakumar. Polynomial preconditioning of symmetric indefinite systems. Technical Report CSRD-580, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 34 pp.

Narasimhan:1995:NCS

[Nar95] V. L. Narasimhan. A new course on supercomputers and parallel architectures. IEEE transactions on education, 38(4):340–??, ???? 1995. CODEN IEEDAB. ISSN 0018-9359.

Nassersharif:1991:SEA

[Nas91]  B. Nassersharif.  Science

and engineering at the Texas A&M University Supercomputer Center. International Journal of Supercomputer Applications, 6(1):4– 12, Spring 1991. CODEN IJSAE9. ISSN 0890-2720.

NASA:1993:FNS

[NAS93] NASA, editor. Fifth NASA Symposium on VLSI design, the University of New Mexico, Albuquerque, New Mexico, November 4–5, 1993, NASA Publications N 1993; N94-21079-133; NASA contractor report NASA CR-194644. NASA, Washington, DC, USA, 1993. ISBN ???? LCCN ????

NSF:1984:NWS

[Nat84] National Science Foundation (U.S.), editor. NSF Workshop on Supercomputer Usage in Mechanics and Transport Phenomena, December 8–9, 1983, NCAR, Boulder, Colorado: [papers]. National Science Foundation, Washington, DC, USA, 1984.

NRAO:1985:SRA

[Nat85] National Radio Astronomy Observatory, Green Bank, WV, USA. A Supercomputer for radio astronomical imaging, 1985. 47 pp.

NCAR:1986:UNS

[Nat86a] National Center for Atmospheric Research (U.S.). UCAR and NCAR strategies in supercomputing. Technical report, National Center for Atmospheric Research, Boulder, CO, USA, September 16, 1986. xviii + 37 + [16] pp.

NCSA:1986:N

[Nat86b] News, 1986. National Center for Supercomputing Applications, Center for Supercomputing Research and Development, University of Illinois at Urbana-Champaign, Champaign, IL, USA.

NCSA:1986:UB

[Nat86c] User bulletin, 1986. National Center for Supercomputing Applications, Center for Supercomputing Research and Development, University of Illinois at Urbana-Champaign, Champaign, IL, USA.

NCSA:1986:ISP

[Nat86d] National Center for Supercomputing Applications, Center for Supercomputing Research and Development, University of Illinois at Urbana-Champaign, Champaign, IL, USA. Industrial supercomputing program, June 1986. 1 pp.

NCSA:1986:IN

[Nat86e] National Center for Supercomputing Applications, Center for Supercomputing Research and Development, University of Illinois at Urbana-Champaign, Champaign, IL, USA. Introducing the NCSA, November 1986. 1 pp.

NCSA:1986:NCS

[Nat86f] National Center for Supercomputing Applications, Center for Supercomputing Research and Development, University of Illinois at Urbana-Champaign, Champaign, IL, USA. National Center for Supercomputing Applications, December 1986. 5 pp.

NCSA:1986:SA

[Nat86g] National Center for Supercomputing Applications, Center for Supercomputing Research and Development, University of Illinois at Urbana-Champaign, Champaign, IL, USA. “Supercomputer Avenue”, November 1986. 1 pp.

Natori:1986:CGM

[Nat86h] Makoto Natori. Conjugate gradient method and supercomputer: PCG (Preconditioned Conjugate Gradient) [in Japanese]. Number 2 in Advances in numerical methods for large sparse sets of linear equations. Keio University, [Tokyo, Japan], 1986. 112 pp. LCCN QA218 .S95 1986. Text in Japanese. “PCG (Preconditioned Conjugate Gradient).”.

NCAR-SCD:1987:SVN

[Nat87a] Supercomputing: the view from NCAR, 1987. Scientific Computing Division, National Center for Atmospheric Research, Boulder, CO, USA.

NCSA:1987:A

[Nat87b] Access, 1987. ISSN 1064-9409. National Center for Supercomputing Applications, University of Illinois at Urbana-Champaign, Champaign, IL, USA.

NCSA:1987:AO

[Nat87c] National Center for Supercomputing Applications. Access online, 1987. ISSN 1064-9409.

NCSA:1987:NMS

[Nat87d] National Center for Supercomputing Applications. NCSA mass storage system. Technical Report 10.1, National Center for Supercomputing Applications, Center for Supercomputing Research and Development, University of Illinois at Urbana-Champaign, Champaign, IL, USA, February 1987. 1 pp.

Natarajan:1988:MNA

[Nat88a] Ramesh Natarajan. Matching numerical algorithms to supercomputer architectures. Research report RC 13625 (#60948), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, 1988. 24 pp.

NSF-CCSACMP:1988:ISC

[Nat88b] National Science Foundation (U.S.). Committee on Computer and Simulation and Analysis of Complex Material Phenomena. The Impact of supercomputing capabilities on U.S. materials science and technology: report of the Committee on Computer Simulation and Analysis of Complex Material Phenomena, volume 451 of Publication NMAB. National Academy Press, Washington, DC, USA, 1988. vii + 65 pp.

NASULGC-HETC:1989:SSR

[Nat89a] National Association of State Universities and Land-Grant and Colleges. Higher Education and Technology Committee. Supercomputing for the 1990s: a shared responsibility. Technical report, Office of Public Affairs/Office of Publications, University of Illinois at Urbana-Champaign, for the National Association of State Universities and Land-Grant Colleges, Champaign, IL, USA, January 1989. xiii + 52 pp.

NCSA:1989:ODC

[Nat89b] Online documentation on the Cray system, 1989. NCSA, Urbana, IL, USA.

NERSC:1990:ESS

[Nat90] National Energy Research Supercomputer Center, Livermore, CA, USA. Energy sciences supercomputing 1990: making an impact on scientific research, 1990. 59 pp.

NCAR-SCD:1991:SVN

[Nat91a] National Center for Atmospheric Research, Boulder, CO, USA. Supercomputing: the view from NCAR. FY90 review and FY91–92 development plan for the NCAR Scientific Computing Division, 1991. ix + 153 pp.

NCSA:1991:PDN

[Nat91b] National Center for Supercomputing Applications, Center for Supercomputing Research and Development, University of Illinois at Urbana-Champaign, Champaign, IL, USA. The process of discovery: NCSA science highlights, 1991. 32 pp.

NCSA:1992:UBN

[Nat92a] National Center for Supercomputing Applications. U.S. business and the National Science Foundation’s supercomputing centers. Technical report, University of Illinois, Office of Printing Sevices, Champaign, IL, USA, 1992. 27 pp.

NCSA:1992:PDN

[Nat92b] National Center for Supercomputing Applications, University of Illinois at Urbana-Champaign, Champaign, IL, USA. The process of discovery: NCSA science highlights, 1991, 1992. 32 pp.

NCSAEducation:1995:SM

[Nat95] National Center for Supercomputing Applications.Education and Outreach Group. Science for the millennium, 1995.

NCSA:19xx:ASU

[Natxxa] Applications software update, 19xx. National Center for Supercomputing Applications, University of Illinois at Urbana-Champaign, Champaign, IL, USA.

NCSA:19xx:N

[Natxxb] News, 19xx. ISSN 0891-0782. National Center for Supercomputing Applications, University of Illinois at Urbana-Champaign, Champaign, IL, USA.

NCSA:19xx:SH

[Natxxc] Science highlights, 19xx. National Center for Supercomputing Applications, University of Illinois at Urbana-Champaign, Champaign, IL, USA.

NCSA:19xx:SU

[Natxxd] Software update, 19xx. National Center for Supercomputing Applications, University of Illinois at Urbana-Champaign, Champaign, IL, USA.

NCSAUS:19xx:DL

[Natxxe]  Data link, 19xx.  ISSN

1064-9425. National Center for Supercomputing Applications, University of Illinois at Urbana-Champaign, Champaign, IL, USA.

NCSAUS:19xx:TRC

[Natxxf] Technical resources catalog, 19xx. ISSN 1064-9417. University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA.

Newton:1992:CGP

[NB92] P. Newton and J. C. Browne.

The CODE 2.0 graphical parallel programming language. In ACM [ACM92b], pages 167–177. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Nanda:1993:EMA

[NB93] Ashwini K. Nanda and Laxmi N. Bhuyan. Efficient mapping of applications on cache based multiprocessors. Journal of Parallel and Distributed Computing, 19(3): 179–??, November 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Niebur:1994:ESB

[NB94] E. Niebur and D. Brettle. Efficient simulation of biological neural networks on massively parallel supercomputers with hypercube architecture. Advances in neural information processing systems, ??(6):904–??, ???? 1994. ISSN 1049-5258.

Noye:1992:CTA

[NBC92] John Noye, Basil Benjamin, and Len Colgan, editors. Computational techniques and applications: proceedings of 5th International Computational Techniques and Applications Conference, held at The University of Adelaide, 14–17 July, 1991. Australian Mathematics Society, Adelaide, SA, Australia, 1992. ISBN 0-86396-172-X. LCCN ????

Noga:1996:DIE

[NBGS96] M. Noga, S. Bednarek, Z. Glowacz, and J. Skwarczynski. Determination of inductances of electrical machine with finite element method on Convex supercomputer. In Borcherds et al. [BBM96], pages 333– 338. ISBN 83-902363-3-8. LCCN ????

Nowatzyk:1995:SNW

[NBKP95a] A. G. Nowatzyk, M. C.

Browne, E. J. Kelly, and M. Parkin. S-connect: From networks of workstations to supercomputer performance. ACM SIGARCH Computer Architecture News, 23(2):71– 82, ???? 1995. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Nowatzyk:1995:CNW

[NBKP95b] Andreas G. Nowatzyk, Michael C.

Browne, Edmund J. Kelly, and Michael Parkin. Sconnect: from networks of workstations to supercomputer performance. ACM SIGARCH Computer Architecture News, 23(2):71– 82, May 1995. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Norton:2002:PUA

[NĆ02] Charles D. Norton and Thomas A. Ćwik. Parallel unstructured AMR and Gigabit networking for Beowulf-class clusters. Lecture Notes in Computer Science, 2328:552– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.

                                        com/link/service/series/
                                        0558/bibs/2328/23280552.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2328/23280552.
                                        pdf.

Navarro:1997:DDC

[NCDS97] H. A. Navarro, A. C. Canale, J. E. D’Elboux, and J. R. Saraiva. Driving dynamics of commercial vehicles. In Roller [Rol97], pages 145– 152. ISBN 0-947719-88-1 (paperback). LCCN ????

Neal:1988:RSH

[NCKMM88] Larry Neal, John Connolly, Doyle D. Knight, and David Matthews-Morgan. The role of supercomputers in higher education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 20(1):134, February 1988. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nicolas:1996:WIL

[NCVG96] M. Nicolas, J.-M. Chomaz, D. Vallet, and E. Guazzelli. Wave instability in liquidfluidized beds. In Wolf et al. [WSB96], pages 267– 278. ISBN 981-02-2635-7. LCCN ????

NYSERNet:1988:MEI

[NDLV88] New York State Education and Research Network (NYSERNet), William L. Dougan, Jim Lombardi, and Tom Vietorisz. Market and economic impact study. New York State Education and Research Network, New York, NY, USA, June 7, 1988. various pp.

Nedjah:2009:HPH

[NdMM09] Nadia Nedjah and Luiza de Macedo Mourelle. Highperformance hardware of the sliding-window method for parallel computation of modular exponentiations. International Journal of Parallel Programming, 37(6): 537–555, December 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=37&
issue=6&spage=537.

Neeman:1990:DAV

[Nee90a] Henry Neeman. A decomposition algorithm for visualization [i.e., visualizing] irregular grids. Technical Report CSRD 1071, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, November 1990. 8 pp.

Neeman:1990:VTT

[Nee90b] Henry Neeman. Visualization techniques for threedimensional flow fields. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1990. vii + 105 pp.

Neelakantan:1994:IES

[Nee94] K. Neelakantan. Indigenous efforts in supercomputing — Anurag’s efforts. In Mahajan et al. [M^{+}94], pages 13–18. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Anonymous:1991:NMS

[New91] New Mexico Supercomputing Challenge: finalist reports, 1991. Mew Mexico Technet, Albuquerque, NM, USA.

Newbury:1993:SAM

[New93] J. Newbury. The South African motor industry and development opportunities. In Anonymous [Ano93-31], pages 27–36. ISBN 0-947719-62-8. LCCN ????

NewMexicoTechnet:1995:NMH

[New95] New Mexico Technet, Inc. New Mexico high school supercomputing challenge: 1995 finalist reports. New Mexico Technet, Albuquerque, NM, USA, 1995. ca. 250 pp.

Newman:2014:SSS

[New14]  L. H. Newman.  Small

Swiss supercomputer shows the way ahead on efficiency. IEEE Spectrum, 51(2):13– 14, February 2014. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Netzer:1992:ERC

[NG92] Robert Netzer and Sanjoy Ghosh. Efficient race condition detection for sharedmemory programs with post/ wait conditions. Technical Report CSRD 1238, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1992. 7 pp.

Ng:1995:SAS

[Ng95] H. W. Ng. Supercomputer application to structural dynamic analysis of PWR piping system. Advances in Engineering Software, 22(2): 87–??, ???? 1995. CO-DEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic).

Navarro:1996:DPM

[NGDH96] J. J. Navarro, E. Gracia-Diego, and J. R. Herrero. Data prefetching and multilevel blocking for linear algebra operations. In ACM [ACM96], pages 109–116. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Navarro:1996:BAS

[NGLPJ96] J. J. Navarro, E. Garcia, J.- L. Larriba-Pey, and T. Juan. Block algorithms for sparse matrix computations on high performance workstations. In ACM [ACM96], pages 301– 308. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Niederberger:1999:HSS

[NGPH99] Ralph Niederberger, Helmut Grund, Eva Pless, and Ferdinand Hommes. High-speed supercomputer communications in broadband networks. Computer Networks (Amsterdam, Netherlands: 1999), 31(21):2309–2318, November 10, 1999. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/
10/15/22/32/40/38/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/15/22/32/40/
38/article.pdf.

Nagai:1991:PEM

[NH91] T. Nagai and Y. Hatano. Performance evaluation of mathematical functions. In Anonymous [Ano91q], pages 126– 133. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Nixon:1992:MST

[Nix92] J. B. Nixon. From minicomputer to supercomputer: a three-dimensional tidal model optimization. In Noye et al. [NBC92], pages 355– 364. ISBN 0-86396-172-X. LCCN ????

Navarro:1994:MFC

[NJL94] J. J. Navarro, T. Juan, and T. Lang. MOB forms: a class of multilevel block algorithms for dense linear algebra operations. In Anonymous [Ano94-134], pages 354–363. ISBN ???? LCCN ????

Nurkkala:1994:PHU

[NK94] T. Nurkkala and V. Kumar. The performance of a highly unstructured parallel algorithm on the KSR1. In IEEE [IEE94c], pages 215– 220. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Nieuwejaar:1996:GPF

[NK96] N. Nieuwejaar and D. Kotz.

The galley parallel file system. In ACM [ACM96], pages 374–381. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Nakashima:1995:SPV

[NKTT95] Y. Nakashima, T. Kitamura, H. Tamura, and M. Takiuchi. Scalar processor of the VPP500 parallel supercomputer. In ACM [ACM95a], pages 348–356. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Nakagawa:1993:WCC

[NM93] M. Nakagawa and T. Mori.

Whole core calculations of power reactors by Monte Carlo techniques. In Kusters et al. [KSW93], pages 702– 713. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Nation:1993:MEC

[NMS93] Wayne G. Nation, Anthony A. Maciejewski, and Howard Jay Siegel. A methodology for exploiting concurrency among independent tasks in partitionable parallel processing systems. Journal of Parallel and Distributed Computing, 19(3): 271–??, November 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

NCSA:1987:SAR

[NN87] Science: annual report to the National Science Foundation, 1987. National Center for Supercomputing Applications, University of Illinois at Urbana-Champaign, Champaign, IL, USA.

Natori:1988:SS

[NN88] Nakoto Natori and T. (Takeshi) Nodera, editors. Sparsity and supercomputer, volume 4 of Advances in numerical methods for large sparse sets of linear equations. Keio University, Tokyo, Japan, March 15, 1988. LCCN QA188 .S62 1988. English or Japanese; summaries in English. Proceedings of a one day symposium held in Keio University, March 15, 1988 and constituting the fourth Symposium of Large Sparse Sets of Linear Equations. PCG.

Nakamura:1990:SRS

[NN90] Takashi Nakamura and Mikio Nagasawa, editors. Status reports of supercomputing astrophysics in Japan: KEK, Tsukuba, August 31– September 2, 1989, volume 89-2 of KEK progress report. National Laboratory for High Energy Physics, Tsukuba-shi, Ibaraki-ken, Japan, 1990.

Nagashima:1990:IFA

[NNS^{+}90] Umpei Nagashima, Fumio Nishimoto, Takashi Shibata, Hiroshi Itoh, and Minoru Gotoh. An improvement of I/O function for auxiliary storage: parallel I/O for a large scale supercomputing. ACM SIGARCH Computer Architecture News, 18(3b):48–59, September 1990. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Nakano:1994:DAD

[NNSY94] M. Nakano, M. Nagamatsu, K. Suzuki, and T. Yoshimura. Development of acoustic double holography method. In Anonymous [Ano94-75], pages 463–470. ISBN 0-947719-68-7. LCCN ????

Nool:1995:EPB

[Noo95] Margreet Nool. Explicit parallel block Cholesky algorithms on the Cray APP. Applied Numerical Mathematics: Transactions of IMACS, 19(1-2):91–114, November 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Norrie:1984:SSA

[Nor84] C. Norrie. Supercomputers for superproblems: an architectural introduction. Computer, 17(3):62–74, 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

NCSC:1989:NCS

[Nor89] North Carolina Supercomputing Center, North Carolina. North Carolina Supercomputing Center, user guide, October 1989. 12 + [16] pp.

NCSC:1993:NCSb

[Nor93a] North Carolina Supercomputing Center, North Carolina. North Carolina Supercomputing Center, research abstracts catalog, January 1993. v + 161 pp.

NCSC:1993:NCSa

[Nor93b] North Carolina Supercomputing Center, Research Triangle Park, NC, USA. North Carolina Supercomputing Center, user guide, March 1993. various pp.

Norman:1996:PCM

[Nor96] Michael L. Norman. Probing cosmic mysteries by supercomputer. Physics Today, 49(10):42–??, October 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//www.physicstoday.org/
resource/1/phtoad/v49/i10/
p42_s1.

Norberg:1997:BRB

[Nor97a] Arthur L. Norberg. Book review: The Supermen: The Story of Seymour Cray and the Technical Wizards behind the Supercomputer by Charles J. Murray. Isis, 88(4):745– 746, December 1997. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.
jstor.org/stable/237893.

Noruk:1997:GMA

[Nor97b] J. S. Noruk. Gas metal arc penetration welding development utilizing neural nets. In Roller [Rol97], pages 215– 222. ISBN 0-947719-88-1 (paperback). LCCN ????

NCSC:19xx:NFN

[Norxx] NCSC flyer: newsletter of the North Carolina Supercomputing Center, 19xx. North Carolina Supercomputing Center, Research Triangle Park, NC, USA.

Norris:2003:WCN

[Nor03] William C. Norris. William C. Norris papers: 1946– 1998: CBI 164. World-Wide Web site, 2003. URL http:

//en.wikipedia.org/wiki/
William_Norris;   http:/
/news.zdnet.com/2100-9595_
22-6108755.html;   http:
//www.cbi.umn.edu/collections/
inv/cbi00164.html.

Nordrum:2017:LIU

[Nor17]  A. Nordrum.  A lan-

guage for the Internet of Underwater Things [news]. IEEE Spectrum, 54(9):9– 10, September 2017. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nowak:1993:GDA

[Now93] P. F. Nowak. A grey diffusion acceleration method for timedependent radiative transfer calculations: Analysis and application. In Kusters et al. [KSW93], pages 33–44. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Noor:1990:SLS

[NP90] Ahmed K. Noor and Jeanne M. Peters. Strategies for largescale structural problems on high-performance computers. ACM SIGARCH Computer Architecture News, 18(3b): 267–280, September 1990. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Nelson:1993:PEC

[NPS93] J. S. Nelson, S. J. Plimpton, and M. P. Sears. Planewave electronic-structure calculations on a parallel supercomputer. Physical review. B, Condensed matter, 47(4): 1765–??, January 15, 1993. ISSN 0163-1829.

Nishio:2020:ESI

[NPS^{+}20] Shin Nishio, Yulu Pan, Takahiko Satoh, Hideharu Amano, and Rodney Van Meter. Extracting success from IBM’s 20-qubit machines using error-aware compilation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 16(3):32:1– 32:25, July 2020. CODEN ???? ISSN 1550-4832. URL

https://dl.acm.org/doi/
abs/10.1145/3386162.

Nobile:1986:EIM

[NR86] A. Nobile and V. Roberto.

Efficient implementation of multidimensional Fast Fourier Transforms on a Cray X-MP. Computer Physics Communications, 40(2-3): 189–201, June 1986. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Noeth:2009:SSC

[NRM^{+}09] Michael Noeth, Prasun Ratn, Frank Mueller, Martin Schulz, and Bronis R. de Supinski. ScalaTrace: Scalable compression and replay of communication traces for high-performance computing. Journal of Parallel and Distributed Computing, 69(8): 696–710, August 2009. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Nobes:2000:CCF

[NRN00] Ross H. Nobes, Alistair P. Rendell, and Jarek Nieplocha. Computational chemistry on Fujitsu vector-parallel processors: Hardware and programming environment. Parallel Computing, 26(7–8): 869–886, July 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/35/21/42/29/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/42/29/
27/article.pdf.

Nakanishi:1995:ESC

[NRS95] H. Nakanishi, V. Rego, and V. Sunderam. On the effectiveness of superconcurrent computations on heterogeneous networks. Journal of Parallel and Distributed Computing, 24(2): 177–190, February 1, 1995. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1995.1017/production;
http://www.idealibrary.

com/links/doi/10.1006/jpdc.
1995.1017/production/pdf.

NASA:1986:EES

[NS86] National Aeronautics and Space Administration and Science Applications International Corporation. Earth and environmental science in the 1980’s. NASA contractor report NASA CR-4029; SAIC project no. 1-224-03-340-28, NASA, Washington, DC, USA, October 1986. various pp.

Nation:1988:PDP

[NS88] Wayne G. Nation and Howard Jay Siegel. Properties of disjoint paths in data manipulator networks. Technical report SRC-TR-88-001, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 25 pp.

Ninokata:1993:FRS

[NS93] H. Ninokata and T. Shimizu.

Fast reactor safety and computational thermo-fluid dynamics approaches [invited]. In Kusters et al. [KSW93], pages 151–163. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Nordlund:1996:SWS

[NSB96] A. Nordlund, R. F. Stein, and A. Brandenburg. Supercomputer windows into the solar convection zone. Bulletin of the Astronomical Society of India = Bharatiya Jyotir Vijyan Parishad, 24(2):261– ??, June 1, 1996. CODEN BANID3. ISSN 0304-9523.

Nilsson:1990:SSA

[NSF90] L. Nilsson, L. Sjöström, and B. Fredriksson. Supercomputing in structural analysis of aircrafts. In Pitcher [Pit90], pages 483–496. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Nabil:1995:CES

[NSH95] K. Nabil, A. Safwat, and Mohamed K. Hasan. Computational experience with a simultaneous transportation equilibrium model using the Cray Y-MP/2. Transportation Congress, Proceedings, 2: 1423–1431, 1995.

Numrich:1994:MCR

[NSP94] R. W. Numrich, P. L. Springer, and J. C. Peterson. Measurement of communication rates on the Cray T3D interprocessor network. Lecture Notes in Computer Science, 797:150–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielson:1990:VSC

[NSR90] Gregory M. Nielson, Bruce D. Shriver, and Lawrence J. Rosenblum. Visualization in scientific computing. IEEE Computer Society Press tutorial. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-8979-X (case), 0-8186-5979-3 (microfiche). xvi + 283 pp. LCCN T385 .V59 1990. IEEE catalog number EH0307-9.

Neeman:2008:SPE

[NSW08] Henry Neeman, Horst Severini, and Dee Wu. Supercomputing in plain English: teaching cyberinfrastructure to computing novices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):27–30, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:/

/ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2008.bib.

Neeman:1989:STA

[NT89] Henry Neeman and Allan Tuchman. Simulation time animation system. Technical Report CSRD 859, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1989.

[2] + 39 pp.

Nagoya:1991:TSA

[NU91] S. Nagoya and T. Ushijima. A theoretical stability analysis for a family of convection diffusion difference schemes using the supercomputing environment. In Anonymous [Ano91q], pages 236–245. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Numrich:1985:SA

[Num85] Robert W. Numrich, editor. Supercomputer applications. Plenum Press, New York, NY, USA; London, UK, 1985. ISBN 0-306-42013-9. LCCN QA76.5.S8944 1984a. Proceedings of the Supercomputer Applications Symposium cosponsored by the Purdue University Computing Center, the Purdue Center for Parallel and Vector Computing, and Control Data, held October 31– November 1, 1984, in West Lafayette, Indiana.

Nasser:1997:ICM

[NW97] S. H. Nasser and J. Weik. An integrated computer model to determine the ultimate design features of a very low fuel consumption car. In Roller [Rol97], pages 305– 318. ISBN 0-947719-88-1 (paperback). LCCN ????

Nakamura:2003:MEM

[NW03] Yoshiaki Nakamura and Chihiro Watanabe. Management and the effect of MITI’s R&D project: case study from a supercomputer project. Technovation, 23(3):221–238, March 2003. CODEN ???? ISSN 0166-4972 (print), 1879-2383 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0166497201000852.

Omran:1994:ESC

[OA94] R. A. Omran and M. A.

Aboelaze. An efficient single copy cache coherence protocol for multiprocessors with multistage interconnection network. In IEEE [IEE94c], pages 1–8. ISBN 0-8186-5680-8, 0-8186-5681-

6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Olszanskyj:1994:PWB

[OB94] S. J. Olszanskyj and A. W.

Bojanczyk. Parallel windowed block recursion least squares. In IEEE [IEE94c], pages 741–748. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

OBoyle:1995:CRS

[OB95] M. O’Boyle and F. Bodin.

Compiler reduction of synchronisation in shared virtual memory systems. In ACM [ACM95a], pages 318–

327. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Ohmacht:2005:BGC

[OBB^{+}05] M. Ohmacht, R. A. Bergamaschi, S. Bhattacharya, A. Gara, M. E. Giampapa, B. Gopalsamy, R. A. Haring, D. Hoenicke, D. J. Krolak, J. A. Marcella, B. J. Nathanson, V. Salapura, and M. E. Wazlowski. Blue Gene/L compute chip: Memory and Ethernet subsystem. IBM Journal of Research and Development, 49 (2/3):255–264, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/ohmacht.
pdf.

Olbrich:1994:BDM

[OBR94] T. Olbrich, D. A. Bradley, and A. M. D. Richardson. Built-in-self-test and diagnostics for microsystems. In Anonymous [Ano94-75], pages 511–518. ISBN 0-947719-68-7. LCCN ????

Orellana:2001:NMN

[OCVA01] Carlos J. García Orellana, Ramón Gallardo Caballero, Horacio M. González Velasco, and Francisco J. López Aligué. NeuSim: a modular neural networks simulator for Beowulf clusters. Lecture Notes in Computer Science, 2085:72–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2085/20850072.
htm;   http://link.springer-

ny.com/link/service/series/
0558/papers/2085/20850072.
pdf.

Ochs:1993:GEC

[OCVG93] R. S. Ochs, K. Conrow, S. Venkatasubramaniam, and S. Grover. A graphical environment for a computerized metabolic map. In Lim et al. [L^{+}93], pages 287–296. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Ozguner:1988:VFS

[OD88] F. Ozguner and R. Daoud.

Vectorized fault simulation on the Cray X-MP Supercomputer. IEEE, New York, NY, USA, 1988. ISBN 0-8186-0869-2. 198–201 pp. LCCN ???? Available from IEEE Service Cent (cat n 88CH2657-5). Piscataway, NJ, USA.

Okulicka-Dluzewska:2001:HPC

[OD01] Felicja Okulicka-Dluzewska.

High-performance computing in geomechanics by a parallel finite element approach. Lecture Notes in Computer Science, 1947:391–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1947/19470391.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1947/19470391.
pdf.

Ostromsky:2015:PID

[ODAZ15] Tzvetan Ostromsky, Ivan Dimov, Vassil Alexandrov, and Zahari Zlatev. Preparing input data for sensitivity analysis of an air pollution model by using highperformance supercomputers and algorithms. Computers and Mathematics with Applications, 70(11):2773– 2782, December 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122115003557.

Oed:1992:CMC

[Oed92a] W. Oed. Cray Y-MP C90: System features and early benchmark results (short communication). Parallel Computing, 18(8):947– 954, August 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Oed:1992:CYC

[Oed92b] Wilfried Oed. Cray Y-MP C90: System features and early benchmark results. Parallel Computing, 18(8):947– 954, August 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ocana:2020:BSG

[OGO^{+}20] Kary A. C. S. Ocaña, Marcelo Galheigo, Carla Osthoff, Luiz M. R. Gadelha, Fabio Porto, Antônio Tadeu A. Gomes, Daniel de Oliveira, and Ana Tereza Vasconcelos. BioinfoPortal: a scientific gateway for integrating bioinformatics applications on the Brazilian national high-performance computing network. Future Generation Computer Systems, 107(??):192–214, June 2020. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X19318230.

Ouenou-Gamo:1997:OTS

[OGOR97] S. Ouenou-Gamo, M. Ouladsine, and A. Rachid. Open thermodynamic system applied to intake and exhaust manifolds of a turbocharger diesel engine. In Roller [Rol97], pages 319–326. ISBN 0-947719-88-1 (paperback). LCCN ????

Ou:1995:ALT

[OGR95] C.-W. Ou, M. Gunwani, and S. Ranka. Architectureindependent locality-improving transformations of computational graphs embedded in k-dimensions. In ACM [ACM95a], pages 289–298. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Oyanagi:1990:SPR

[OGY90] Yoshio Oyanagi, Eiichi Goto, and N. Yoshida. Supercomputing pseudo random numbers: proposals on hardware and software. Technical report 90-012, University of Tokyo, Faculty of Science, Dept. of Information Science, Tokyo, Japan, April 1990. 6 pp.

Oyanagi:1991:SPR

[OGY91] Y. Oyanagi, E. Goto, and N. Yoshida. Supercomputing pseudo random numbers — proposals on hardware and software. In Anonymous [Ano91q], pages 97–98. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

OBoyle:1992:TAC

[OH92] M. O’Boyle and G. A. Hedayat. A transformational approach to compiling Sisal for distributed memory architectures. In ACM [ACM92b], pages 335–346. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Oliver:1994:ASK

[OHHIB94] C. E. Oliver, H. R. Hicks, M. Honey, and K. D. Iles-Brechack. Adventures in supercomputing: a K–12 program in computational science: An assessment. In Grayson [Gra94], pages 213– 217. CODEN PFECDR. ISBN 0-7803-2414-5, 0-7803-2413-7, 0-7803-2415-3. ISSN 0190-5848. LCCN T 62 F76 1994.

Oliver:1993:ASK

[OHIB93] C. E. Oliver, H. R. Hicks, and K. D. Iles-Brechak. Adventures in supercomputing: A K–12 program in computing and computational science. In Grayson [Gra93c], pages 694–697. ISBN 0-7803-1482-4, 0-7803-1483-2, 0-7803-1484-0. ISSN 0190-5848. LCCN ???? IEEE Catalog number: 93CH3373-8.

Ohr:1986:CSC

[Ohr86] Stephan Ohr. Computer scores with Cray compatibility. Electronic Design, 34(5): 61–62, March 1986. CODEN ELODAW. ISSN 0013-4872.

Ochi:1991:VAM

[OIY91] H. Ochi, N. Ishiura, and S. Yajima. A vector algorithm for manipulating Boolean functions based on shared binary decision diagrams. In Anonymous [Ano91q], pages 191–200. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Okumura:1993:TAS

[OK93] K. Okumura and T. Kuriyama.

93SC001 three-dimensional aerodynamic simulations of two-box automobile using turbulence model and QUICK scheme. In Anonymous [Ano93-31], pages 45–52. ISBN 0-947719-62-8. LCCN ????

Oed:1986:MMS

[OL86] W. Oed and O. Lange. Modelling, measurement, and simulation of memory interference in the Cray X-MP. Parallel Computing, 3 (4):343–358, October 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Olmsted:1996:TCA

[OLLG96] C. Olmsted, S. Li, T. Logan, and R. Guritz. Terrain correction algorithms for ERS-1 SAR images using a Cray parallel supercomputer to produce Alaskan land mosaics. In Anonymous [Ano96l], pages II–151– II–158. ISSN 1067-0106. LCCN QE 33.2 R4 G45 1996. Two volumes.

OHallaron:1994:CPP

[OLWW94] D. R. O’Hallaron, P. J. Lieu, L. P. Withers, and J. E. Whelchel. Computing the pipelined phase-rotation FFT. In IEEE [IEE94c], pages 462–469. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Ohta:1991:SAI

[OM91] M. Ohta and T. Maeno. Streamlined access for indirect addressing of an array. In Anonymous [Ano91q], pages 258–261. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Ohbuchi:1996:IMS

[OMA^{+}96] R. Ohbuchi, T. Miyazawa, M. Aono, A. Koide, M. Kimura, and R. Yoshida. Integrated medical-image system for cancer research and treatment. IBM Journal of Research and Development, 40 (2):185–210, March 1996. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL

http://www.almaden.ibm.
com/journal/rd40-2.html#
four.

Othmer:1993:ETA

[OMM93] Hans G. Othmer, Philip K.

Maini, and James D. Murray, editors. Experimental and theoretical advances in biological pattern formation: Proceedings of a NATO Advanced Research Workshop on Biological Pattern Formation, held August 27–31, 1992, in Oxford, England, United Kingdom, volume 259 of NATO ASI Series A Life Sciences. Plenum Press, New York, NY, USA; London, UK, 1993. ISBN 0-306-44661-8. ISSN 0258-1213. LCCN QH491 .N38 1993.

Ouisloumen:1993:ACP

[OMR93] M. Ouisloumen, G. Marleau, and R. Roy. Applying the collision probability method to hexagonal assemblies in two and three dimensions. In Kusters et al. [KSW93], pages 102–112. ISBN 3-923704-11-

9. LCCN ???? Two volumes.

Onstad:1988:PDT

[Ons88] David W. Onstad. Populationdynamics theory: the roles of analytical, simulation, and supercomputer models. Ecological Modelling, 43(1–2): 111–124, October 1988. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0304380088900750.

Orlando:1996:TNP

[OP96] S. Orlando and R. Perego. A template for non-uniform parallel loops based on dynamic scheduling and prefetching techniques. In ACM [ACM96], pages 117–

124. ISBN 0-89791-803-

7. LCCN QA76.5 I61 1996. ACM order number 415961.

Openshaw:1996:PSA

[Ope96] S. Openshaw. Parallel supercomputing applications in GIS. In Rumor et al. [RMO96], pages 661–670. ISBN 90-5199-268-8, 4-274-90098-3. LCCN ????

Opper:1995:LAN

[Opp95a]  M. Opper.  Learning in

artificial neural networks: The statistical mechanics approach. In Herrmann et al. [HWP95], pages 321–330. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Oppliger:1995:ISE

[Opp95b] R. Oppliger. Internet security enters the Middle Ages. Computer, 28(10):100– 101, October 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Olbrich:2001:USP

[OPR01] S. Olbrich, H. Pralle, and S. Raasch. Using streaming and parallelization techniques for 3D visualization in a high-performance computing and networking environment. Lecture Notes in Computer Science, 2110:231– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2110/21100231.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2110/21100231.
pdf.

EdwardOliver:1994:ASI

[ORS94] C. Edward Oliver, H. Richard, and B. G. Summers. Adventures in supercomputing: An innovative program for high school teachers. In Grayson [Gra94], pages 133–137. CO-DEN PFECDR. ISBN 0-7803-2414-5, 0-7803-2413-7, 0-7803-2415-3. ISSN 0190-5848. LCCN T 62 F76 1994.

Oliver:1994:ASI

[ORSS94] C. E. Oliver, H. Richard, B. G. Summers, and D. G. Staten. Adventures in supercomputing: An innovative program for high school teachers. Frontiers in Education Conference, ????(????): 133–??, ???? 1994. CODEN PFECDR. ISSN 0190-5848.

Ogino:1993:HFM

[OS93] T. Ogino and K. Sasaki. Human friendly man-machine system with advanced media technology [invited]. In Kusters et al. [KSW93], pages 617–629. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Oh:1994:PAL

[OS94] S. Oh and S. Y. Shin. A parallel algorithm for largescale linear programs with a special structure. In IEEE [IEE94c], pages 749– 755. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Ohta:1995:OTS

[OSKO95] H. Ohta, Y. Saito, M. Kainaga, and H. Ono. Optimal tile size adjustment in compiling general DOACROSS loop nests. In ACM [ACM95a], pages 270–279. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Ostanin:1994:ACV

[Ost94] A. Ostanin. Adaptive control of vibrations in multisupport vehicles. In Anonymous [Ano94-75], pages 181– 188. ISBN 0-947719-68-7. LCCN ????

Ortiz-Tapia:2007:DSA

[OT07]  Arturo Ortiz-Tapia.  A

door to state-of-the-art highperformance computing. IEEE Distributed Systems Online, 8(6):??, June 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
ds/2007/06/o6004.pdf.

Otero:2002:BSM

[Ote02] Glen Otero. The Beowulf state of mind. Linux Journal, 97:88–91, May 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//www.linuxjournal.com/
article.php?sid=5710.

Overill:1994:PPA

[OW94] Richard E. Overill and Stephen Wilson. Performance of parallel algorithms for the evaluation of power series. Parallel Computing, 20 (8):1205–1213, August 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Ohmacht:2013:IBG

[OWG^{+}13] M. Ohmacht, A. Wang, T. Gooding, B. Nathanson, I. Nair, G. Janssen, M. Schaal, and B. Steinmacher-Burow.

IBM Blue Gene/Q memory subsystem with speculative execution and transactional memory. IBM Journal of Research and Development, 57(1/2):7:1–7:12, January– March 2013. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Oyanagi:1999:DSJ

[Oya99] Yoshio Oyanagi. Development of supercomputers in Japan: Hardware and software. Parallel Computing, 25 (13–14):1545–1567, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/35/21/32/36/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/32/36/
25/article.pdf.

Oyanagi:2002:FS

[Oya02] Yoshio Oyanagi. Future of supercomputing. Journal of Computational and Applied Mathematics, 149(1): 147–153, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042702005265.

Ohno:2014:PMD

[OYK^{+}14] Yousuke Ohno, Rio Yokota, Hiroshi Koyama, Gentaro Morimoto, Aki Hasegawa, Gen Masumoto, Noriaki Okimoto, Yoshinori Hirano, Huda Ibeid, Tetsu Narumi, and Makoto Taiji. Petascale molecular dynamics simulation using the fast multipole method on K computer. Computer Physics Communications, 185(10): 2575–2585, October 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465514002082.

Oyake:1991:PIG

[OYWK91] I. Oyake, T. Yoshida, Y. Wauke, and A. Kawai. Parallel image generation for scientific visualization. In Anonymous [Ano91q], pages 165–

172. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Perry:1992:EMMa

[PA92] Tekla S. Perry and J. A.

Adam. Electronic mail — Email: pervasive and persuasive. IEEE Spectrum, 29(10): 22–23, October 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Palmer:1993:CGT

[PA93a] T. S. Palmer and M. L.

Adams. Curvilinear geometry transport discretizations in the “Thick” diffusion limit. In Kusters et al. [KSW93], pages 3–14. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Pena:1993:SPS

[PA93b] J. Pena and J. R. Alvarez. The safety parameters system SPS for the simulation of nuclear accidents. In Kusters et al. [KSW93], pages 331– 340. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Packard:1985:S

[Pac85] Edward Packard. Supeordenador, volume 23 of Elige tu propia venturia. Timun Mas, Barcelona, Spain, 1985. ISBN 84-7176-846-1. 118 pp. LCCN ????

Packard:1986:S

[Pac86] Edward Packard. La supercomputadora. Number 16 in Elige tu propia aventura. Editorial Atlantida, Buenos Aires, 1986. ISBN 950-08-0443-3. 118 pp. LCCN ????

Padua:1989:DPM

[Pad89]  David A. Padua.  The

Delta program manipulation system: preliminary design. Technical Report CSRD 880, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1989. 18 pp.

Palmer:2015:MBI

[Pal15]  Tim Palmer.  Modelling:

Build imprecise supercomputers. Nature, 526(7571): 32–33, September 29, 2015. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Pancake:1993:GEI

[Pan93] Cherri M. Pancake. Guest Editor’s introduction: The changing face of supercomputing. IEEE parallel and distributed technology: systems and applications, 1(4): 12–15, November 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Pandurangan:1996:EVH

[Pan96] Anand Pandurangan. Equalization of very high speed supercomputer wire data channels: a thesis presented to the faculty of the Graduate School, Tennessee Technological University. Thesis (M.S.), Tennessee Technological University, Cookeville, TN, USA, 1996. xiv + 145 pp.

Pandis:1997:FPS

[Pan97] S. N. Pandis. Formation and properties of secondary atmospheric aerosol: From the laboratory to the supercomputer. In Anonymous [Ano97c], pages S367–S370. ISSN 0021-8502.

Paprzycki:1992:CGE

[Pap92] Marcin Paprzycki. Comparisons of Gaussian elimination algorithms on a Cray Y-MP. Linear Algebra and its Applications, 172:57–??, July 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Paprzycki:1997:BRI

[Pap97] Marcin Paprzycki. Book review: An Introduction To High-performance Scientific Computing. IEEE Concurrency, 5(3):73–74, July/ September 1997. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

                                        //dlib.computer.org/pd/
                                        books/pd1997/pdf/p3073.
                                        pdf.

Papka:2016:ESH

[Pap16]  Michael E. Papka.  Ex-

panding the scope of highperformance computing facilities. Computing in Science and Engineering, 18 (3):84–87, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Parkinson:1986:PAP

[Par86] D. Parkinson. Performance analysis in a 4096 processor environment. The Journal of Systems and Software, 6 (1–2):11–15, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Parish:1990:CCS

[Par90a] Tom Parish. Crystal clear storage: The holostore, a new mass storage device with supercomputer performance, could eliminate the I/O bottleneck. BYTE Magazine, 15(12):283–288, November 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Parker:1990:PTE

[Par90b] Lloyd E. G. Parker. Preliminary testing and evaluation of new computer programs for traffic analysis: task, evaluation of supercomputer potential: Performance of traffic models TRANSYT-7F and INTEGRATION on supercomputers. Technical Report TDS-90-03, Research and Development Branch, Ontario Ministry of Transportation, Downsview, ON, Canada, December 1990. ii + 131 pp. Performance of traffic models TRANSYT-7F and INTE-GRATION on supercomputers.

Paruolo:1990:VEM

[Par90c] Giuseppe Paruolo. A vectorefficient and memory-saving interpolation algorithm for PIC codes on a Cray X-MP. Journal of Computational Physics, 89(2):462– 482, August 1990. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/002199919090153R.

Parisi:1994:RRC

[Par94] Franklin J. Parisi. A rose is a rose… but a Cray’s not always a Cray. Computers in Physics, 8(4): 380–??, July 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4823310.

Pasemann:1995:NDS

[Pas95] F. Pasemann. Neuromodules: a dynamical systems approach to brain modelling. In Herrmann et al. [HWP95], pages 331–348. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Pankratius:2012:FMS

[PATT12] Victor Pankratius, Ali-Reza Adl-Tabatabai, and Walter F. Tichy, editors. Fundamentals of multicore software development. Chapman and Hall/CRC computational science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-1273-X. xii + 317 pp. LCCN QA76.642 .F86 2012.

Paulson:2005:SSC

[Pau05] Linda Dailey Paulson. Squeezing supercomputers onto a chip. Computer, 38(1):21– ??, January 2005. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        co/2005/01/r1021.pdf;

http://csdl.computer.org/
dl/mags/co/2005/01/r1021.
htm.

Paulson:2008:NBG

[Pau08] Linda Dailey Paulson. News briefs: Group releases first 3-D chip standard; supercomputers get energy-efficient; company puts a new spin on music-editing software; attack hijacks domain names of Internet-addressing organization. Computer, 41(9):18– 20, September 2008. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulson:2009:NBS

[Pau09] Linda Dailey Paulson. News briefs: Start-up unveils flexible supercomputing approach. Computer, 42(1): 21–24, January 2009. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Payer:1997:NOE

[Pay97] E. Payer. NVH optimization of engines by means of numerical simulation techniques — state-of-the-art and future trends. In Roller [Rol97], pages 345–358. ISBN 0-947719-88-1 (paperback). LCCN ????

Packard:1984:S

[PB84] Edward Packard and Frank Bolle. Supercomputer, volume 39 of Choose your own adventure; 39. Bantam Doubleday Dell Publishing Group Inc., 666 Fifth Avenue, New York, NY 10130, USA, 1984. ISBN 0-553-24678-X. 118 pp. LCCN PS3566.A275 S975 1984 Spec Coll Eaton. US$1.95.

Polychronopoulos:1987:PAH

[PB87] C. D. (Constantine D.) Polychronopoulos and Utpal K. (Utpal Kumar) Banerjee. Processor allocation for horizontal and vertical parallelism and related speedup bounds. Technical Report CSRD-618, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 11 pp.

Pryor:1988:VMC

[PB88] Daniel V. Pryor and Patrick J. Burns. Vectorized Monte Carlo molecular aerodynamics simulation of the Rayleigh problem. Technical report SRC-TR-88-008, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 8 pp.

Pancake:1990:DPL

[PB90] Cherri M. Pancake and Donna Bergmark. Do parallel languages respond to the needs of scientific programmers? Computer, 23(12): 13–23, December 1990. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pantos:1994:SSS

[PB94a] E. Pantos and J. Bordas.

Supercomputer simulation of small angle X-ray scattering, electron micrographs and X-ray diffraction patterns of macromolecular structures. In Zaidi et al. [ZAS94], pages 77–82. ISBN 969-8117-05-9. ISSN 0033-4545. LCCN QP551 .I553 1993. Also in Pure and Applied Chemistry, vol. 66, no. 1 (1994).

Pascal:1994:NSC

[PB94b] H. Pascal and M. Buffat. Numerical simulation of a compressed turbulence for the flow modelling inside a reciprocating engine. In Anonymous [Ano94-75], pages 763– 770. ISBN 0-947719-68-7. LCCN ????

Pang:1998:SBD

[PB98] Yuan-Ping Pang and Stephen Brimijoin. Supercomputingbased dimeric analog approach for drug optimization. Parallel Computing, 24(9–10):1557–1566, September 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/parco/sub/1998/
24/9-10/1348.pdf.

Peters:1993:PIN

[PBDM93] A. Peters, H. Babovsky, H. Daniels, and T. Michl. 93SC031 the parallel implementation of NSFLEX on an IBM RS/ 60000 cluster. In Anonymous [Ano93-31], pages 289–?? ISBN 0-947719-62-8. LCCN ????

Pasquale:1991:SDW

[PBK91] Joseph Pasquale, Barbara Bittel, and Daniel Kraiman. A static and dynamic workload characterization study of the San Diego Supercomputer Center Cray X-MP. ACM SIGMETRICS Performance Evaluation Review, 19 (1):218–219, May 1991. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Poli:1996:ITA

[PBK96] D. J. Poli, M. S. Berry, and J. N. Kruchowski. IC technology and ASIC design for the Cray J90 supercomputer. IBM Journal of Research and Development, 40 (4):475–483, July 1996. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        almaden.ibm.com/journal/
                                        rd40-4.html#eight.

Packard:1987:S

[PBM87] Edward Packard, Frank Bolle, ill, and Inaki Mendiguren, tr. Superordenadorea, volume 3 of Aukeratu zeure abentura. ELKAR, Donostia, 1987. ISBN 84-7529-435-9. 118 pp. LCCN ????

Pfenning:1995:VSM

[PBM95] Jörg-Thomas Pfenning, Achim Bachem, and Ronald Minnich. Virtual shared memory programming on workstation clusters. Future Generation Computer Systems, 11 (4–5):387–399, August 1995. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Pryor:1993:UGA

[PC93] R. J. Pryor and D. D. Cline.

Use of a genetic algorithm to solve two-fluid flow problems on an NCUBE multiprocessor computer. In Kusters et al. [KSW93], pages 45–57. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Pancake:1994:WUN

[PC94a] C. M. Pancake and C. Cook.

What users need in parallel tool support: Survey results and analysis. In IEEE [IEE94c], pages 40–47. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Peltier:1994:NPC

[PC94b] J. P. Peltier and P. Christ.

Networking perspectives for collaborative supercomputing in European aerospace research and industry. In Gentzsch and Harms [GH94a], pages 53–59. CODEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1–2 (c1994). DM96.00. Two volumes.

Pilant:1997:PEI

[PC97] M. S. Pilant and H. Chen. Parameter estimation issues for flow in porous media with multiple length scales. In Delic and Wheeler [DW97], pages 329–338. ISBN 0-89871-378-1. LCCN ????

Papelis:1993:TAD

[PCK93] Y. E. Papelis, T. L. Casavant, and J. G. Kuhl. 93SC014 A tool-based approach to the design of hard real-time systems with varying temporal characteristics. In Anonymous [Ano93-31], pages 213– 220. ISBN 0-947719-62-8. LCCN ????

Perrott:1984:IPL

[PCM84] R. H. Perrott, D. Crookes, and P. Milligan. Implementing a parallel language on the Cray-1. Computer Physics Communications, 37 (1-3):119–124, July 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Peng:2019:CMC

[PCY^{+}19] Shaoliang Peng, Yingbo Cui, Shunyun Yang, Wenhe Su, Xiaoyu Zhang, Tenglilang Zhang, Weiguo Liu, and Xing-Ming Zhao. A CPU– MIC collaborated parallel framework for GROMACS on Tianhe-2 supercomputer. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 16(2):425– 433, March 2019. CO-DEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1109/
TCBB.2017.2713362.

Peters:1994:CAE

[PD94] M. Peters and A. M. Dore.

Can artificial experts work in the service bay? In Anonymous [Ano94-75], pages 669– 676. ISBN 0-947719-68-7. LCCN ????

Pacheco:1991:SPS

[PDR91] P. S. Pacheco, J. M. Del Rosario, and T. Rashid. SPICE03: a program for simulating integrated circuits on hypercubes. In Anonymous [Ano91q], pages 201– 205. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Panda:1994:MWM

[PDR94] D. K. Panda and V. A. Dixit-Radiya. Message-ordering for wormhole-routed multiport systems with link contention and routing adaptivity. In IEEE [IEE94c], pages 191–198. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Padua:1988:ADN

[PE88] David A. Padua and Perry A. Emrath. Automatic detection of nondeterminancy in parallel programs. Technical Report CSRD 717, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 5 pp.

Pottenger:1995:IRP

[PE95] B. Pottenger and R. Eigenmann. Idiom recognition in the Polaris parallelizing compiler. In ACM [ACM95a], pages 444–448. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Pelz:1993:PCF

[PEH93] R. B. Pelz, A. Ecer, and J. Hauser, editors. Parallel computational fluid dynamics ’92: proceedings of the Conference on Parallel CFD ’92: implementations and results using parallel computers, New Brunswick, NJ, USA, 18–20 May, 1992. North-Holland, Amsterdam, The Netherlands, 1993. ISBN 0-444-89986-3. LCCN QC150 .C66 1992.

Peisert:2017:SHP

[Pei17] Sean Peisert. Security in high-performance computing environments. Communications of the ACM, 60(9):72– 80, September 2017. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL http://cacm.
acm.org/magazines/2017/
9/220422/fulltext.

Peltier:1993:ACA

[Pel93a] J. P. Peltier, editor. AERO-DAYS ’93: Second Community Aeronautics RTD Conference: conference proceedings: Naples, 4–5 October 1993. Commission of the European Communities, Directorate-General Science, Research and Development, Luxembourg, 1993. ISBN 92-826-6301-9. LCCN ???? EUR 14977.

Peltier:1993:EAS

[Pel93b]  J. P. Peltier.  European

aeronautical supercomputing networking and the EC High Performance Computing and Networking Initiative. In AERODAYS ’93: Second Community Aeronautics RTD Conference: conference proceedings: Naples, 4–5 October 1993 [Pel93a], pages 371–372. ISBN 92-826-6301-9. LCCN ???? EUR 14977.

Pellegrini:1994:SMD

[Pel94] F. Pellegrini. Static mapping by dual recursive bipartitioning of process and architecture graphs. In IEEE [IEE94c], pages 486– 493. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Perry:1983:GKM

[Per83]  Tekla S. Perry.  Gov-

ermnent: Keyworth at midterm: an exclusive report from Presidential Science Advisor George A. Keyworth II on technology policy in the Reagan administration. IEEE Spectrum, 20(8):53– 59, August 1983. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1986:NSS

[Per86] Dennis G. Perry. Network support of supercomputers. Future Generation Computer Systems, 2(1):65–67, March 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Perry:1987:WPH

[Per87] Tekla S. Perry. At work or play, he’s the captain: Engineer-manager Howard Sachs races sailboats and builds machines. From microprocessors to supercomputers under often-daunting conditions. IEEE Spectrum, 24 (6):56–59, June 1987. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1989:IS

[Per89] Tekla S. Perry. Intel’s secret is out. IEEE Spectrum, 26 (4):22–28, April 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1993:EMW

[Per93a] Tekla S. Perry. Environment — modeling the world’s climate. IEEE Spectrum, 30 (7):33–37, July 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:1993:MC

[Per93b] Tekla S. Perry. Modeling the climate. IEEE Spectrum, 30 (7):33–37, July 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2002:VNV

[Per02] Tekla S. Perry. Vying for next video player crown. IEEE Spectrum, 39(6):19– 20, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2004:BFS

[Per04]  Tekla S. Perry.  Build-

ing a in a flash supercomputer. IEEE Spectrum, 41 (6):24–25, June 2004. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2006:BSF

[Per06]  Tekla S. Perry.  Build-

ing a supercomputer in a flash. IEEE Spectrum, 41 (6):24–25, June 2006. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Petersen:1983:VFN

[Pet83] W. P. Petersen. Vector Fortran for numerical problems on Cray-1. Communications of the ACM, 26(11):1008– 1021, November 1983. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Petersen:1989:PTS

[Pet89a] Paul M. Petersen. PDE tutor: a system for the automatic solution of partial differential equations. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1989. v + 50 pp.

Peterson:1989:SRS

[Pet89b] Victor L. Peterson. Supercomputer requirements for selected disciplines important to aerospace. Proceedings of the IEEE, 77(7):1038– ??, July 1, 1989. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Petrov:1997:PVS

[Pet97] V. Petrov. Physical volumic spaces, their artificial imitations and hypothetical spaces: Unified concept for volumic imaging. In Roller [Rol97], pages 105– 112. ISBN 0-947719-88-1 (paperback). LCCN ????

Pevzner:1993:DSO

[Pev93] P. A. Pevzner. DNA statistics, overlapping word paradox and Conway equation. In Lim et al. [L^{+}93], pages 61–68. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Porter:1990:SGS

[PF90] Stephen Porter and Richard Fichera. This survey of graphics supercomputer users provides a surprising market profile. Computer Graphics World, 13(8):90–??, August 1, 1990. CODEN CGWODH. ISSN 0271-4159.

Pfeiffer:1993:SS

[Pfe93] W. Pfeiffer. Scalable supercomputing. In Szeto and Rangayyan [SR93b], pages 581–?? ISBN 0-7803-1377-1, 0-7803-1378-X, 0-7803-1379-8, 0-7803-1411-5. LCCN R 856 A2 I344 1993. Three volumes: Part 1: Visualization, imaging, signal processing, modeling, neural networks; Part 2: Medical informatics, ethics, cardiology, instrumentation; Part 3: Biomechanics, rehabilitation, electrical phenomena, biomaterials.

Povinelli:1993:DIB

[PG93] C. M. Povinelli and R. A. Gibbs. Distribution of informative binary codes in mammalian DNA. In Lim et al. [L^{+}93], pages 589–596. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Pervez:2010:FMA

[PGK^{+}10] Salman Pervez, Ganesh Gopalakrishnan, Robert M. Kirby, Rajeev Thakur, and William Gropp. Formal methods applied to high-performance computing software design: a case study of MPI one-sided communication-based locking. Software—Practice and Experience, 40(1):23–43, January ??, 2010. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Padua:1987:SPE

[PGL87] David A. Padua, Vincent A. Guarna, and Duncan Lawrie. Supercomputer programming environments. Technical Report CSRD 673, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 25 pp.

Power:1995:CSB

[PH95] H. Power and R. T. Hart, editors. Computer simulations in biomedicine: International Conference on Computer Simulations in Biomedicine (3rd: June 1995: Milan, Italy). Computational Mechanics Publications, Southampton; Boston, 1995. ISBN 1-85312-321-8, 1-56252-245-0. LCCN R859.7.C65 I575 1995.

Peters:1997:PTT

[PH97] J. F. Peters and S. E.

Howington. Pre-asymptotic transport through porous media. In Delic and Wheeler [DW97], pages 271–280. ISBN 0-89871-378-1. LCCN ????

Park:2011:PHP

[PH11] Sang-Min Park and Marty A.

Humphrey. Predictable highperformance computing using feedback control and admission control. IEEE Transactions on Parallel and Distributed Systems, 22(3):396– 411, March 2011. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Philippe:1985:ASR

[Phi85] Bernard Philippe. Approximating the square root of the inverse of a matrix. Technical Report CSRD-508, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1985. iii + 17 pp.

Padua:1988:MLC

[PHK88] David A. Padua, W. Ludwell Harrison, and David J. Kuck. Machines, languages and compilers for parallel symbolic computing. Technical Report CSRD 729, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 453–469 pp.

Poole:1995:PIM

[PHVJ95] Eugene L. Poole, Michael A. Heroux, Pravin Vaidya, and Anil Joshi. Performance of iterative methods in ANSYS on Cray parallel/vector supercomputers. Computing systems in engineering: an international journal, 6(3): 251–259, June 1995. CODEN COSEEO. ISSN 0956-0521.

Pickover:1988:SRS

[Pic88] Clifford A. Pickover. A short recipe for seashell synthesis on a graphics supercomputer. Research report RC 14311 (#64103), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, December 20, 1988. 11 pp.

Pickover:1989:PRG

[Pic89] Clifford A. Pickover. Picturing randomness on a graphics supercomputer. Research report RC 14468 (#64759), IBM T.J. Watson Research Center, Yorktown Heights, NY, USA, 1989. iii + 8 pp.

Pickover:1991:PRG

[Pic91a] C. A. Pickover. Picturing randomness on a graphics supercomputer. IBM Journal of Research and Development, 35(1/2):227–230, January/March 1991. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Pickover:1991:ISS

[Pic91b] Clifford A. Pickover. An informal survey of the scientific and social impact of a soda can sized super supercomputer. Computers in Physics, 5(3):290–??, May 1, 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Pickover:1992:VMA

[Pic92] Clifford A. Pickover. A vacation on Mars — an artist’s journey in a computer graphics world. Computers and Graphics, 16(1):9–13, 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Pan:2004:PBC

[PIH04] Yi Pan, Constantinos S.

Ierotheou, and Majeed M. Hayat. Parallel bandwidth characteristics calculations for thin avalanche photodiodes on a SGI Origin 2000 supercomputer. Concurrency and Computation: Practice and Experience, 16(12):1207– 1225, October 2004. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pillai:1993:IS

[Pil93] M. G. G. Pillai. India’s supercomputers. Economic and political weekly, XXVIII(31): 1573–??, July 1993. ISSN 0012-9976.

Pini:1991:PAP

[Pin91] Giorgio Pini. Parallel algorithm for the partial eigensolution of sparse symmetric matrices on the Cray Y-MP. Parallel Computing, 17(4-5): 553–561, July 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Piner:1999:CSCl

[Pin99] Mary-Louise G. Piner. Computer society connection: Inaugural Cray award presented in a November ceremony; meritorious service and outstanding contribution awards show; appreciation of volunteer efforts; Computer recognizes reviewers’ expertise. Computer, 32 (12):86–90, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co1999/pdf/rz086.
pdf.

Piner:2001:CSCb

[Pin01] Mary-Louise G. Piner. Computer Society connection: Visionary Glen Culler named Cray Award winner; new IEEE membership option focuses on women in engineering; William R. Hewlett, 1913–2001; Cray Award recognizes creativity in highperformance computing; 2001 Fellows named for distinguished achievements; design competition teams selected for CSIDC 2001; Transactions on Software Engineering seeks Editor in Chief; new election process begins. Computer, 34 (2):84–90, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r2084.
pdf.

PSC:1986:PN

[Pit86] PSC news, 1986. Pittsburgh Supercomputing Center, Pittsburgh, PA, USA.

PSC:1987:PSC

[Pit87] Projects in scientific computing, 1987. ISSN 1048-2105. Pittsburgh Supercomputing Center, Pittsburgh, PA, USA.

PSC:1988:PSC

[Pit88] Pittsburgh Supercomputing Center, Pittsburgh, PA, USA. Pittsburgh Supercomputing Center, 1988. various pp.

Pittelli:1989:DBP

[Pit89] Frank M. Pittelli. Data base processing on the horizon. Technical report SRC-TR-89-001, Supercomputing Research Center: IDA, Lanham, MD, USA, January 1989. 14 pp.

Pitcher:1990:SES

[Pit90] E. J. (Eric J.) Pitcher, editor. Science and Engineering on Supercomputers: Proceedings of the Fifth International Conference, held in London, England, during October 22–24, 1990. Springer-Verlag and Computational Mechanics Publications, Berlin, Germany / Heidelberg, Germany / London, UK / etc. and Southampton, UK, 1990. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Pointer:1990:CSP

[PJ90] David Pointer and Greg Jaxon. Cedar synchronization processor instruction set reference. Technical Report CSRD 1017, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1990. 79 pp.

Perrott:1980:CSU

[PK80] R. H. Perrott and C. King.

Cray-1 simulation using Pascal-Plus. Proceedings of the International Conference on Parallel Processing, ?? (??):105–10, ???? 1980. CO-DEN PCPADL. ISSN 0190-3918.

Polychronopoulos:1987:GSS

[PK87] Constantine D. Polychronopoulos and David J. Kuck. Guided self-scheduling: a practical scheduling scheme for parallel supercomputers. IEEE Transactions on Computers, C-36(12): 1425–1439, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=5009495.

Pyle:1989:EPA

[PK89] L. Duane Pyle and Sang-Ha Kim. An exterior point algorithm for linear programming implemented on the SX-2 supercomputer. Technical report UH-CS-90-10, Dept. of Computer Science, University of Houston, Houston, TX, USA, November 1989. ii + 34 pp.

Pyle:1994:EPA

[PK94] L. Duane Pyle and S.-H. Kim. Exterior point algorithm for linear programming implemented on the SX-2 supercomputer. In Rice and De Millo [RD94], pages 111–138. ISBN 0-306-44697-9. LCCN QA76.S848 1994.

Perrott:1991:SDI

[PL91a] R. H. Perrott and T. F. Lunney. A syntax-directed integrated programming environment for developing SIMD supercomputer software. Software—Practice and Experience, 21(3):269– 286, March 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Perrott:1991:SIP

[PL91b] R. H. Perrott and T. F. Lunney. A syntax-directed integrated programming environment for developing SIMD supercomputer software. Software—Practice and Experience, 21(3):269– ??, March 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Phua:1991:SSC

[PL91c] Kang Hoh Phua and K. F. (Kia Fock) Loe, editors. Singapore Supercomputing Conference ’90: supercomputing for strategic advantage, 11– 12 December 1990. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1991. ISBN 981-02-0700-X. LCCN QA76.88.S56 1990.

Plank:1994:PRI

[PL94] J. S. Plank and K. Li. Performance results of ickp — a consistent checkpointer on the iPSC/860. In IEEE [IEE94c], pages 686–693. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Pascual:2019:EAS

[PLC^{+}19] Jose A. Pascual, Joshua Lant, Caroline Concatto, Andrew Attwood, Javier Navaridas, Mikel Luján, and John Goodacre. On the effects of allocation strategies for exascale computing systems with distributed storage and unified interconnects. Concurrency and Computation: Practice and Experience, 31(21):e4784:1– e4784:??, November 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Plis:1997:RSP

[Pli97] Y. M. Plis. Responses of structured plankton population to toxicant: Computational aspects of modeling. In Delic and Wheeler [DW97], pages 139–142. ISBN 0-89871-378-1. LCCN ????

Pankajakshan:2020:PSM

[PLS20] R. Pankajakshan, P.-H. Lin, and B. Sjögreen. Porting a 3D seismic modeling code (SW4) to CORAL machines. IBM Journal of Research and Development, 64(3/4):17:1– 17:11, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Poulos:2022:PSN

[PMC22] Alexandra Poulos, Sally A. McKee, and Jon C. Calhoun. Posits and the state of numerical representations in the age of exascale and edge computing. Software— Practice and Experience, 52 (2):619–635, February 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Plale:2021:RPH

[PMP21] Beth A. Plale, Tanu Malik, and Line C. Pouchard. Reproducibility practice in high-performance computing: Community survey results. Computing in Science and Engineering, 23(5):55– 60, September/October 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Punzo:1994:HRL

[PMS94] Giovanni Punzo, Federico Massaioli, and Sauro Succi. High-resolution lattice–Boltzmann computing on the IBM SP1 scalable parallel computer. Computers in Physics, 8 (6):705–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.168487.

Pang:2008:EIB

[PMS^{+}08] Y.-P. Pang, T. J. Mullins, B. A. Swartz, J. S. McAllister, B. E. Smith, C. J. Archer, R. G. Musselman, A. E. Peters, B. P. Wallenfelt, and K. W. Pinnow. EUDOC on the IBM Blue Gene/L system: Accelerating the transfer of drug discoveries from laboratory to patient. IBM Journal of Research and Development, 52(1/2): 69–??, January/March 2008. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL

http://www.research.ibm.
com/journal/rd/521/pang.
html.

Paczuski:1996:SCF

[PN96] M. Paczuski and K. Nagel.

Self-organized criticality and 1/f noise in traffic. In Wolf et al. [WSB96], pages 73–86. ISBN 981-02-2635-7. LCCN ????

Pandey:2013:CCS

[PN13] Suraj Pandey and Surya Nepal. Cloud computing and scientific applications — big data, scalable analytics, and beyond. Future Generation Computer Systems, 29(7): 1774–1776, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167739X13000861.

Puska:1993:RCC

[PNK93] E. K. Puska, J. Niemi, and H. Kontio. Reactor core calculations in plant analyser environment. In Kusters et al. [KSW93], pages 818–

830. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Poole:1988:SLS

[PO88] Eugene L. Poole and Andrea L. Overman. Solution of linear systems of equations with a structural analysis code on the NAS Cray-

2. NASA Contractor Reports, 4159, December 1988. CO-DEN NSCRAQ. ISSN 0565-7059.

Poeppel:1995:HSC

[Poe95] E. Poeppel. Homogeneity of space and continuity of time: Necessary prerequisites of perception? In Herrmann et al. [HWP95], pages 3–10. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Pointer:1989:PR

[Poi89] Lynn Pointer. Perfect report: 1. Technical Report CSRD 896, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1989. 20 pp.

Pointer:1990:PPE

[Poi90] Lynn Pointer. Perfect: performance evaluation for costeffective transformations, report 2. Technical Report CSRD 964, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1990. 57 pp.

Polychronopoulos:1986:PRS

[Pol86] C. D. (Constantine D.) Polychronopoulos. On program restructuring, scheduling, and communication for parallel process systems. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. viii + 256 pp.

Polychronopoulos:1987:ALO

[Pol87a] C. D. (Constantine D.) Polychronopoulos. Advanced loop optimizations for parallel computers. Technical Report CSRD 664, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 24 pp.

Polychronopoulos:1987:LCC

[Pol87b] C. D. (Constantine D.) Polychronopoulos. Loop coalescing: a compiler transformation for parallel machines. Technical Report CSRD-635, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 23 pp.

Polychronopoulos:1987:MAL

[Pol87c] C. D. (Constantine D.) Polychronopoulos. More on advanced loop optimizations. Technical Report CSRD 667, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 30 pp.

Polychronopoulos:1987:ARF

[Pol87d] Constantine D. Polychronopoulos. Automatic restructuring of Fortran programs for parallel execution. Technical Report CSRD 665, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 26 pp.

Polychronopoulos:1988:COE

[Pol88a] C. D. (Constantine D.) Polychronopoulos. Compiler optimizations for enhancing parallelism and their impact on architecture design. Technical Report CSRD 781, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. ?? pp.

Polychronopoulos:1988:IRO

[Pol88b] C. D. (Constantine D.) Polychronopoulos. The impact of run-time overhead on usable parallelism. Technical Report CSRD 778, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 108–112 pp.

Polychronopoulos:1988:MVM

[Pol88c] C. D. (Constantine D.) Polychronopoulos. Multiprocessing versus multiprogramming. Technical Report CSRD 861, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 8 pp.

Polychronopoulos:1988:PEP

[Pol88d] C. D. (Constantine D.) Polychronopoulos. Parafrase-2: an environment for parallelizing, partitioning, synchronizing, and scheduling programs on multiprocessors. Technical Report CSRD 837, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 10 pp.

Polychronopoulos:1988:TAC

[Pol88e] C. D. (Constantine D.) Polychronopoulos. Toward autoscheduling compilers. Technical Report CSRD 789, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 34 pp.

Polychronopoulos:1989:CRI

[Pol89] C. D. (Constantine D.) Polychronopoulos. Compile-time reduction of interprocessor communication for parallel computers. Technical Report CSRD 642, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1989. 34 pp.

Polychronopoulos:1990:ASC

[Pol90] C. D. (Constantine D.) Polychronopoulos. Auto scheduling: control flow and data flow come together. Technical Report CSRD 1058, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1990. 28 pp.

Pool:1996:FST

[Poo96a] R. Pool. Forget silicon, try DNA: a mere gram of the stuff of life could solve huge mathematical problems that would set a supercomputer in a spin — at least in theory. New Scientist, ??(2038): 26–??, ???? 1996. CODEN NWSCAL. ISSN 0262-4079, 0028-6664.

Pool:1996:CSF

[Poo96b] Robert Pool. Cover story:

Forget silicon, try DNA. New Scientist, 151(2038):26– ??, July 13, 1996. CODEN NWSCAL. ISSN 0262-4079, 0028-6664.

Pope:1991:WSC

[Pop91] Stephen B. Pope. What a supercomputer can reveal about turbulence. Engineering: Cornell Quarterly, 25(2 / 3):54–??, Spring 1991. ISSN 0013-7871.

Popova:1992:PSA

[Pop92] N. N. Popova. Problem-Oriented supercomputer architecture for a class of magnetohydrodynamic models. Computational mathematics and modeling, 3(1): 79–83, January 1, 1992. ISSN 1046-283X.

Pope:1997:MDN

[Pop97]  G. A. Pope.  Modeling

dense nonaqueous phase liquid field projects. In Delic and Wheeler [DW97], pages 187–200. ISBN 0-89871-378-1. LCCN ????

Porsching:1986:BRB

[Por86] Thomas A. Porsching. Book review: Progress and Supercomputing in Computational Fluid Dynamics: Proceedings of U.S.–Israel Workshop, 1984 (Earll M. Murman and Saul S. Abarbanel, eds.). SIAM Review, 28 (4):601, ???? 1986. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Porterfield:1989:SMI

[Por89] Allan Porterfield. Software methods for improvement of cache performance on supercomputer applications. Technical report COMP TR89-93, Rice University, Dept. of Computer Science, Houston, TX, USA, May 1989. ix + 149 pp.

Potter:1987:DSA

[Pot87] Jerry L. Potter. Data structures for associative supercomputers. Technical report SRC-TR-87-003, Supercomputing Research Center: IDA, Lanham, MD, USA, 1987. 23 pp.

Potter:1988:AT

[Pot88] Jerry L. Potter. The ASP tutorial. Technical report SRC-TR-88-005, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 30 pp.

Pountain:1986:PS

[Pou86] Dick Pountain. Personal supercomputers. BYTE Magazine, 11(7):363–368, July 1986. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pouzin:1988:INI

[Pou88] Louis Pouzin. Impact of national and international networks on the use of supercomputer centers. Computer Physics Communications, 49 (1):201–203, April 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465588902275.

Pountain:1994:LB

[Pou94a] Dick Pountain. The last bastion. BYTE Magazine, 19 (9):47–??, September 1, 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Pountain:1994:LBC

[Pou94b] Dick Pountain. The last bastion: “Computer on a chip” is old hat; are you ready for a “supercomputer on a chip”? BYTE Magazine, 19(9):47– ??, September 1994. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Power:1997:CSB

[Pow97] H. Power, editor. Computer simulations in biomedicine:

International conference; 4th — June 1997, Acquasparta, Italy, volume 28(6) of AD-VANCES IN ENGINEER-ING SOFTWARE 1997. Elsevier, Amsterdam, The Netherlands, 1997. CO-DEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic).

Pozrikidis:2013:XSC

[Poz13] C. Pozrikidis. XML in scientific computing. Chapman and Hall/CRC numerical analysis and scientific computing series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4665-1227-X (hardback). xv + 243 pages pp. LCCN Q183.9 .P69 2013.

Petersen:1991:EES

[PP91] Paul M. Petersen and David A. Padua. Experimental evaluation of some data dependence tests (extended abstracts). Technical Report CSRD 1080, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1991. 18 pp.

Petersen:1992:DDA

[PP92a] Paul M. Petersen and David A. Padua. Dynamic dependence analysis: a novel method for data dependence evaluation. Technical Report CSRD 1228, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1992. 21 pp.

Petersen:1992:MEP

[PP92b] Paul M. Petersen and David A.

Padua. Machine-independent evaluation of parallelizing compilers. Technical Report CSRD 1173, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. 21 pp.

Poon:1993:AGA

[PP93] P. W. Poon and G. T. Parks.

Application of genetic algorithms to in-core nuclear fuel management optimization. In Kusters et al. [KSW93], pages 777–786. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Pierra:1994:DEP

[PPG94] G. Pierra, J. C. Potier, and P. Girard. Design and exchange of parametric models for parts libraries. In Anonymous [Ano94-75], pages 397– 404. ISBN 0-947719-68-7. LCCN ????

Peiro:1990:CAF

[PPM90] J. Peiro, J. Peraire, and K. Morgan. The computation of 3D aerodynamical flows using unstructured meshes. In Pitcher [Pit90], pages 103– 111. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Pal:1994:CCA

[PPP94] Rajat Kumar Pal, S. Prasant Pal, and A. Pal. On the computational complexity of approximate area minimization in VLSI design. In Balakrishnan [Bal94], pages 378–380. ISBN 0-07-462044-4. LCCN ????

Patron:1995:ECT

[PPR95] M. Patron, T. Porcher, and F. Robin. An evaluation of the Cray T3D at CEA/CEL-V. Lecture Notes in Computer Science, 919:799–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierce:1994:PIN

[PR94a] P. Pierce and G. Regnier. The Paragon[TM] implementation of the NX message passing interface. In IEEE [IEE94c], pages 184–190. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Plata:1994:CSD

[PR94b] O. Plata and F. F. Rivera. Combining static and dynamic scheduling on distributedmemory multiprocessors. In Anonymous [Ano94-134], pages 186–195. ISBN ???? LCCN ????

Prasanna:1995:PPI

[Pra95] V. K. Prasanna, editor. Parallel processing: 1st International workshop — December 1994, Bangalore, India. Tata McGraw-Hill Pub. Co., New Delhi, India, 1995. ISBN 0-07-462332-X. LCCN ????

Press:1993:STS

[Pre93a] William H. Press. Supercomputing and the transformation of science by W. J. Kaufmann and L. L. Smarr. Nature, 362(6420): 507–??, April 8, 1993. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Prevost:1993:HSN

[Pre93b]  J. Prevost.  High speed

networking at the Commissariat à l’Energie Atomique [invited]. In Kusters et al. [KSW93], pages 10–16. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Primack:2012:US

[Pri12] J. R. Primack. The universe in a supercomputer. IEEE Spectrum, 49(10):42– 47, October 2012. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

PED-OLA-SM:1994:MSC

[Pro94] Program Evaluation Division, Office of the Legislative Auditor, State of Minnesota, Saint Paul, MN, USA. Minnesota Supercomputer Center, June 1994. xii + 24 + 2 pp.

Prokhorov:2001:CPR

[Pro01] V. V. Prokhorov. Computational portal: Remote access to high-performance computing. Lecture Notes in Computer Science, 2127:308–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2127/21270308.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2127/21270308.
pdf.

Phadke:1994:PPD

[PRS94] L. Phadke, R. Rastogi, and J. Sainis. Prospects and problems of developing automatic software for reading DNA sequence from auto-radiograms using Scanjet Digital Scanner. In Mahajan et al. [M^{+}94], pages 309–314. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Phadke:1994:DPI

[PRSS94] L. Phadke, R. Rastogi, S. Souche, and S. Shetiya. Development and parallelisation of image processing routines on BPPS. In Mahajan et al. [M^{+}94], pages 275–283. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Pryor:1994:IUP

[Pry94] Daniel V. Pryor. Implementation and usage of a portable and reproducible parallel pseudorandom number generator. Technical report SRC-TR-94-116, Supercomputing Research Center: IDA, Lanham, MD, USA, March 15, 1994. 16 pp.

Pittelli:1988:ATN

[PS88] Frank M. Pittelli and David Smitley. Analysis of a 3D toroidal network for a shared memory architecture. Technical report SRC-TR-88-011, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 14 pp.

Paprzycki:1994:SLR

[PS94a] M. Paprzycki and P. Stpiczynski. Solving linear recurrence systems on a Cray Y-MP. Lecture Notes in Computer Science, 879:416–424, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pozo:1994:LRS

[PS94b] R. Pozo and S. L. Smith.

Limited resource scheduling in multifrontal algorithms for sparse linear systems. In IEEE [IEE94c], pages 593– 600. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Papka:1996:UEI

[PS96] M. E. Papka and R. Stevens. UbiWorld: An environment integrating virtual reality, supercomputing and design. In IEEE [IEE96b], pages 306–

307. ISBN 0-8186-7582-9. ISSN 1082-8907. LCCN ???? Pickering:1998:MPM

[PS98] S. Pickering and I. K. Snook. A massively parallel molecular dynamics algorithm for the MasPar supercomputer. Computer Physics Communications, 108(2–3): 200–210, February 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465597001136.

Psarris:1992:EDD

[Psa92] K. Psarris. On exact data dependence analysis. In ACM [ACM92b], pages 303–312. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Piomelli:2001:LES

[PSB01] Ugo Piomelli, Alberto Scotti, and Elias Balaras. Largeeddy simulations of turbulent flows, from desktop to supercomputer (invited talk). Lecture Notes in Computer Science, 1981:551–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1981/19810551.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1981/19810551.
pdf.

Pan:2003:SHI

[PSG03] Yi Pan, Joseph J. S. Shang, and Minyi Guo. A scalable HPF implementation of a finite-volume computational electromagnetics application on a CRAY T3E parallel system. Concurrency and Computation: Practice and Experience, 15(6): 607–621, May 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pavlakos:1993:VMS

[PSM93] Constantine J. Pavlakos, Larry A. Schoof, and John F. Mareda. A visualization model for supercomputing environments. IEEE parallel and distributed technology: systems and applications, 1 (4):16–22, November 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Prodan:2012:EHP

[PSO12] Radu Prodan, Michael Sperk, and Simon Ostermann. Evaluating high-performance computing on Google App Engine. IEEE Software, 29 (2):52–58, March/April 2012. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

PSC:1990:SR

[PSOM90] Pittsburgh Supercomputing Center, San Diego Supercomputer Center, Ohio Supercomputer Graphics Project, and Media Magic. Supercomputing review, 1990. 1 videocassette (VHS) (30 min.).

Penha:2019:AAD

[PSS^{+}19] Jeronimo C. Penha, Lucas B. Silva, Jansen M. Silva, Kristtopher K. Coelho, Hector P. Baranda, José Augusto M. Nacif, and Ricardo S. Ferreira. ADD: Accelerator Design and deploy — a tool for FPGA high-performance dataflow computing. Concurrency and Computation: Practice and Experience, 31(18): e5096:1–e5096:??, September 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pao:1992:FNC

[PT92] Y.-H. Pao and Y. Takefuji. Functional-link net computing: theory, system architecture, and functionalities. Computer, 25(5): 76–79, May 1992. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Philippsen:1993:DMP

[PT93] M. Philippsen and W. Tichy.

Demonstration of massive parallel computer applications: Programming exercises for the massive parallel computer. In Kusters et al. [KSW93], pages 783– 784. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Ponnusamy:1993:EPE

[PTC^{+}93] Ravi Ponnusamy, Rajeev Thakur, Alok Choudhary, Kishore Velamakanni, and Zeki Bozkus. Experimental performance evaluation of the CM-5. Journal of Parallel and Distributed Computing, 19(3):192–??, November 1993. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Pryor:1993:TSS

[PTS93] Daniel V. Pryor, Mark R.

Thistle, and Nabeel Shirazi. Text searching on Splash 2. Technical report SRC-TR-93-095, Supercomputing Research Center: IDA, Lanham, MD, USA, May 1993. 9 pp.

Peinze:1989:SPS

[PTT89] Klaus Peinze, Ulrich Trottenberg, and Bernhard Thomas. Suprenum — the parallel supercomputer. Technische Mitteilungen Krupp, 1:15–??, June 1, 1989.

Puget:1994:IRS

[Pug94] D. Puget. Information retrieval systems with semantic and syntactic representation of texts. In Balakrishnan [Bal94], pages 73–84. ISBN 0-07-462044-4. LCCN ????

Peiron:1994:SAS

[PVA94] M. Peiron, M. Valero, and E. Ayguade. Synchronized access to streams in SIMD vector multiprocessors. In Anonymous [Ano94-134], pages 23–32. ISBN ???? LCCN ????

Padua:1986:ACOb

[PW86a] David A. Padua and Michael J. Wolfe. Advanced compiler optimization for supercomputers. Communications of the ACM, 29(12): 1184–1201, December 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        7904.html.

Padua:1986:ACO

[PW86b] David A. Padua and Michael J. Wolfe. Advanced compiler optimizations for supercomputers. Communications of the ACM, 29(12): 1184–1201, December 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/

toc/Abstracts/0001-0782/
7904.html.

Padua:1986:ACOa

[PW86c] David A. Padua and Michael Joseph Wolfe. Advanced compiler optimizations for supercomputers. Technical Report UILU-ENG-592, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 17 pp.

Prasanna:1994:SDP

[PW94] V. K. Prasanna and C.-L.

Wang. Scalable data parallel object recognition using geometric hashing on the CM-5. In IEEE [IEE94c], pages 817– 824. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Peszynska:2005:SIH

[PW05a] M. Peszynska and M. F.

Wheeler. Special issue: Highperformance computing in geosciences. Concurrency and Computation: Practice and Experience, 17(11):1363– 1364, September 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Plaszczak:2005:GCS

[PW05b] Pawel Plaszczak and Richard Wellner. Grid computing: the savvy manager’s guide. Elsevier/Morgan Kaufmann, San Francisco, CA, USA, 2005. ISBN 0-12-742503-9. 312 (est.) pp. LCCN QA76.9.C58 P65 2005. Poochinapan:2023:HPC

[PW23] Kanyuta Poochinapan and Ben Wongsaijai. Highperformance computing of structure-preserving algorithm for the coupled BBM system formulated by weighted compact difference operators. Mathematics and Computers in Simulation, 205: 439–467, March 2023. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0378475422003949.

Postma:1995:NMR

[PWVH95] E. O. Postma, E. H. Wolf, H. J. Van den Herik, and P. T. W. Hudson. The nature of memory representations. In Herrmann et al. [HWP95], pages 105–110. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Park:1997:RAV

[PYTL97] T.-W. Park, H.-J. Yim, T.- O. Tak, and K.-H. Lee. Ride analysis of a vehicle with a flexible body. In Roller [Rol97], pages 129–

136. ISBN 0-947719-88-1 (paperback). LCCN ????

Perry:1989:SEP

[PZ89] Tekla S. Perry and Glenn Zorpette. Supercomputer experts predict expansive growth. IEEE Spectrum, 26(2):26–33, February 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perrott:1986:SL

[PZA86] R. H. Perrott and A. Zarea-Aliabadi. Supercomputer languages. ACM Computing Surveys, 18(1):5–22, March 1986. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0360-0300/
6463.html.

Perrott:1987:SPD

[PZA87] R. H. Perrott and Adib Zarea-Aliabadi. A supercomputer program development system. Software— Practice and Experience, 17 (10):663–683, October 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Piccolo:1991:GWS

[PZGL91] F. Piccolo, V. Zecca, A. Grimaudo, and C. Loiodice. Graphic workstations and supercomputers: an integrated environment for simulation of fluid dynamics problems. IBM Journal of Research and Development, 35(1/2):167– 183, January/March 1991. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Peng:2020:HSC

[PZS^{+}20] Shaoliang Peng, Xiaoyu Zhang, Wenhe Su, Dong Dong, Yutong Lu, Xiangke Liao, Kai Lu, Canqun Yang, Jie Liu, Weiliang Zhu, and Dongqing Wei. Highscalable collaborated parallel framework for large-scale molecular dynamic simulation on Tianhe-2 supercomputer. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 17(3):804–816, May 2020. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1109/
                                        TCBB.2018.2805709.

Qatu:1992:SAS

[QB92] M. S. Qatu and A. M. Bataineh. Structural analysis of shallow shells on the Cray Y-MP supercomputer. Computers and Structures, 45(3): 453–459, October 1992. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Quisquater:1991:CLE

[QD91] Jean-Jacques Quisquater and Yvo G. Desmedt. Chinese Lotto as an exhaustive codebreaking machine. Computer, 24(11):14–22, November 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Quinn:1987:DEA

[Qui87] Michael J. (Michael Jay)

Quinn. Designing efficient algorithms for parallel computers. McGraw-Hill series in supercomputing and artificial intelligence. McGraw-Hill, New York, NY, USA, 1987. ISBN 0-07-051071-7. xvi + 288 pp. LCCN QA76.5 .Q561 1987.

Quinn:1995:CSV

[Qui95] Michael J. Quinn. Conferences: Supercomputing ’94: The view from the exhibition floor. IEEE parallel and distributed technology: systems and applications, 3(1):84–85, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1995/pdf/h10084.
pdf.

Rendell:2000:CCF

[R^{+}00] Alistair P. Rendell et al.

Computational chemistry on Fujitsu vector-parallel processors: Development and performance of applications software. Parallel Computing, 26(7–8): 887–911, July 2000. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/35/21/42/29/28/abstract.
html;   http://www.elsevier.

                                        nl/gej-ng/10/35/21/42/29/
                                        28/article.pdf.

Raath:1997:SLS

[Raa97]  A. D. Raath.  Service

load simulation in automotive structures and components for design, development and testing. In Roller [Rol97], pages 337–344. ISBN 0-947719-88-1 (paperback). LCCN ????

Ranganathan:1996:RCD

[RAES96] M. Ranganathan, A. Acharya, G. Edjlali, and A. Sussman. Runtime coupling of dataparallel programs. In ACM [ACM96], pages 229–236. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Raghavan:1994:DSG

[Rag94] P. Raghavan. Distributed sparse Gaussian elimination and orthogonal factorization. In IEEE [IEE94c], pages 607– 614. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Raghunathan:2006:MSD

[Rag06]  Sudarshan  Raghunathan.

Making a supercomputer do what you want: High-level tools for parallel programming. Computing in Science and Engineering, 8(5):70– 80, September/October 2006. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Rajamony:2011:PIP

[RAG11] R. Rajamony, L. B. Arimilli, and K. Gildea. PERCS: The IBM POWER7-IH highperformance computing system. IBM Journal of Research and Development, 55 (3):3:1–3:12, ???? 2011. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Ramakrishnan:1986:SIF

[Ram86] R. Ramakrishnan. Supercomputer implementation of finite element algorithms for high speed compressible flows. Thesis (Ph.D.), Old Dominion University, Norfolk, VA, USA, 1986. xii + 155 pp.

Ramaswamy:1988:SBS

[Ram88] Sitaram Ramaswamy. Supercomputer based second order design sensitivity analysis of constrained dynamic systems. Thesis (M.S.), Kansas State University, Manhattan, KS, USA, 1988. v + 123 pp.

Rambabu:1994:ANN

[Ram94] P. Rambabu. Artificial neural networks: Application to system identification and controls. In Mahajan et al. [M^{+}94], pages 429–?? ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Rauchwerger:1995:RMP

[RAP95] L. Rauchwerger, N. M. Amato, and D. A. Padua. Runtime methods for parallelizing partially parallel loops. In ACM [ACM95a], pages 137– 146. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Rashid:1991:CCS

[Ras91] Richard F. Rashid, editor. CMU computer science: a 25th anniversary commemorative: Technical symposium — 1991: Pittsburg, PA. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-52899-1. LCCN QA75.5 .C548 1990.

Rattner:1987:ATC

[Rat87] Justin Rattner. Architecture and technologies for concurrent supercomputing, October 13, 1987. 1 videocassette (VHS) (53 min.).

Rauchwerger:1991:PPP

[Rau91] Lawrence Rauchwerger. Π perfect: the portably instrumented Perfect Benchmarks. Technical Report CSRD 1150, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1991. 8 + 1 pp.

Ravikumar:1992:PDP

[Rav92] C. P. Ravikumar. Parallel r-dimensional placement on a vector minisupercomputer. Computer Systems Science and Engineering, 7 (3):147–151, July 1992. CO-DEN CSSEEI. ISSN 0267-6192.

Ravikumar:1995:PDP

[Rav95] C. P. Ravikumar. Parallel r-dimensional placement on a vector minisupercomputer. International Journal of Computer Systems Science and Engineering, 10(3): 138–143, July 1995. CODEN CSSEEI. ISSN 0267-6192.

Rawlings:1997:PMV

[Raw97] S. P. Rawlings. Predicting manufacturing variation in convertible roof systems. In Roller [Rol97], pages 177– 184. ISBN 0-947719-88-1 (paperback). LCCN ????

Radons:1995:SRS

[RBK95] G. Radons, V. Breuer, and J. Krueger. Simple reasons for spatio-temporal complexity of neuronal spike trains. In Herrmann et al. [HWP95], pages 249–256. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Renaud:1994:TPS

[RBL94] C. Renaud, F. Bricout, and E. Lepretre. Two parallel schemes for radiosity on the MP1. In Mahajan et al. [M^{+}94], pages 34–39. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Ramani:1994:KRS

[RC94] S. Ramani and R. Chandrasekar. A knowledge representation scheme for machine aided translation. In Balakrishnan [Bal94], pages 93–?? ISBN 0-07-462044-4. LCCN ????

Riley:2003:HPJ

[RCB03] Christopher J. Riley, Siddhartha Chatterjee, and Rupak Biswas. Highperformance Java codes for computational fluid dynamics. Concurrency and Computation: Practice and Experience, 15(3–5):395–415, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rahnejat:1997:NMD

[RCK97] H. Rahnejat, D. Centea, and P. Kelly. Non-linear multibody dynamic analysis for the study of in-cycle vibrations (whoop) of cable operated clutch systems. In Roller [Rol97], pages 245– 252. ISBN 0-947719-88-1 (paperback). LCCN ????

Rimpault:1993:RME

[RCR93] G. Rimpault, V. Colacioppo, and J. L. Rowlands. Recent methods in the European Cell Code ECCO with applications to the SE-FOR and SUPERPHENIX Doppler measurements. In Kusters et al. [KSW93], pages 187–198. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Russkov:2021:ARR

[RCS21] Alexander Russkov, Roman Chulkevich, and Lev N. Shchur. Algorithm for replica redistribution in an implementation of the population annealing method on a hybrid supercomputer architecture. Computer Physics Communications, 261(??): Article 107786, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465520303933.

Roska:1993:LCO

[RCZ93] Tamas Roska, Leon O. Chua, and Akos Zarandy. Language, compiler, and operating system for the CNN supercomputer. Memorandum UCB/ERL M93/34, Electronics Research Laboratory, College of Engineering, University of California, Berkeley, CA, April 30, 1993. 20 pp.

Rice:1994:SCS

[RD94] J. Rice and R. A. De Millo, editors. Studies in computer science: in honor of Samuel D. Conte: Proceedings of a conference held November 1– 3, 1989, West Lafayette, IN, USA, Software Science and Engineering. Plenum Press, New York, NY, USA; London, UK, 1994. ISBN 0-306-44697-9. LCCN QA76.S848

Rissland:2007:EFC

[RD07] Peter Rissland and Yuefan Deng. Electrostatic force computation for biomolecules on supercomputers with torus networks. Parallel Computing, 33(2):116– 123, March 2007. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rahali:1994:VES

[RDHC94] F. Rahali, J. D. Dessimoz, C. Hlavac, and J. D. Chatelain. Versatile 2D-Position estimating and seeking systems. In Anonymous [Ano94-75], pages 197–204. ISBN 0-947719-68-7. LCCN ????

Ruan:1993:DWC

[RDZ93] D. Ruan, C. De Raedt, and F. Zhang. Development work currently being carried out on the time-dependent finite-element diffusion code TRANSFUSION for nuclear oil well logging problems. In Kusters et al. [KSW93], pages 81–92. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Ramany:1994:IBV

[RE94] S. Ramany and D. Eager. The interaction between virtual channel flow control and adaptive routing in wormhole networks. In Anonymous [Ano94-134], pages 136–145. ISBN ???? LCCN ????

Redelfs:1991:NSF

[Red91]  Ann Redelfs.  The Na-

tional Science Foundation supercomputing centers, 1991. Technical report, Cornell Theory Center, Ithaca, NY, USA, 1991. 36 pp.

Reed:1988:DSA

[Ree88] Coke S. Reed. The dimension of a strange attractor. Technical report SRC-TR-88-014, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 6 pp.

Reinhardt:1985:DAM

[Rei85] Steve Reinhardt. Data-flow approach to multitasking on Cray X-MP computers. Operating Systems Review, 19(5): 107–114, 1985. CODEN OS-RED8. ISBN 0-89791-174-1. ISSN 0163-5980 (print), 1943-586X (electronic).

Reinhardt:1988:TPP

[Rei88] Steve Reinhardt. Two parallel processing aspects of the Cray Y-MP computer system. Proceedings of the International Conference on Parallel Processing, 1:311–314, 1988. CODEN PCPADL. ISSN 0190-3918. Available from IEEE Service Cent (cat n 88CH2625-2). Piscataway, NJ, USA.

Reins:1993:TQT

[Rei93] R. E. Reins. Total quality: a tool for corporate re-invention. In Anonymous [Ano93-31], pages 23– 26. ISBN 0-947719-62-8. LCCN ????

Renner:1997:PES

[Ren97] Rebecca Renner. Pump-and-treat enters the supercomputer age — Rebecca Renner reports on new “optimization techniques” being used to design more efficient cleanups. Environmental science and technology, 31(1): 30A, ???? 1997. CODEN ES-THAG. ISSN 0013-936X.

Report:1992:REW

[Rep92] Rubbia Report. Report of the EEC Working Group on High-Performance Computing. In Meuer [Meu92c], pages 213–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

UIUC-CSRD:19xx:RR

[Resxx] Research review, 19xx. University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA.

Resch:2001:TAH

[Res01]  Michael Resch.  Topic

07 applications on highperformance computers. Lecture Notes in Computer Science, 1900:479–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19000479.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19000479.
pdf.

Reuse:1992:P

[Reu92] B. Reuse. Positionspapiere.

In Meuer [Meu92c], pages 243–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Rau:1993:IPS

[RF93] Bantwal Ramakrishna Rau and Joseph A. Fisher. Instructionlevel parallelism: a special issue of The Journal of Supercomputing, volume SECS 235 of The Kluwer international series in engineering and computer science; SECS 235. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993. ISBN 0-7923-9367-8. 282 pp. LCCN QA76.58 .I49 1993. Reprinted from The Journal of Supercomputing, volume 7, number 1/2, 1993.

Redon:1994:SR

[RF94] X. Redon and P. Feautrier.

Scheduling reductions. In Anonymous [Ano94-134], pages 117–125. ISBN ???? LCCN ????

Rau:1987:DTR

[RFS87] Darwen Rau, Jose Antonio Baptista Fortes, and Howard Jay Siegel. Destination tag routing techniques based on a state model for the IADM network. Technical report SRC-TR-87-006, Supercomputing Research Center: IDA, Lanham, MD, USA, 1987. 53 pp.

Rothberg:1992:PIH

[RG92] E. Rothberg and A. Gupta. Parallel ICCG on a hierarchical memory multiprocessor — addressing the triangular solve bottleneck. Parallel Computing, 18(7): 719–741, July 1992. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Rill:1994:FAW

[RG94] S. Rill and R. Grosso. Future aerospace working scenarios using high speed networks and supercomputers applied to flow simulation for complete aircraft. Lecture Notes in Computer Science, 797:60–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rivas:2017:SFE

[RGH17] Juan M. Rivas, J. Javier Gutiérrez, and Michael González Harbour. A supercomputing framework for the evaluation of real-time analysis and optimization techniques. The Journal of Systems and Software, 124 (??):120–136, February 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0164121216302163.

Romero:2015:DPC

[RGL^{+}15] N. A. Romero, C. Glinsvad, A. H. Larsen, J. Enkovaara, S. Shende, V. A. Morozov, and J. J. Mortensen. Design and performance characterization of electronic structure calculations on massively parallel supercomputers: a case study of GPAW on the Blue Gene/P architecture. Concurrency and Computation: Practice and Experience, 27(1):69– 93, January 2015. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Rhea:1990:AS

[Rhe90]  John Rhea.  The air-

borne supercomputer. Air Force magazine, 73(5):162– ??, May 1, 1990. CO-DEN AFORCO. ISSN 0730-6784 (print), 1943-4782 (electronic).

Raafat:1996:DWP

[RHH96] T. Raafat, J. P. Hulin, and H. J. Herrmann. Density waves and pressure variations in dry granular flows in a vertical pipe. In Wolf et al. [WSB96], pages 317– 322. ISBN 981-02-2635-7. LCCN ????

Ryu:2013:IBG

[RIB^{+}13] K. D. Ryu, T. A. Inglett, R. Bellofatto, M. A. Blocksome, T. Gooding, S. Kumar, A. R. Mamidala, M. G. Megerian, S. Miller, M. T. Nelson, B. Rosenburg, B. Smith, J. Van Oosten, A. Wang, and R. W. Wisniewski. IBM Blue Gene/Q system software stack. IBM Journal of Research and Development, 57(1/2):5:1–5:12, January–March 2013. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Richards:1990:SDD

[Ric90a] W. G. Richards. Supercomputers in drug design. In Pitcher [Pit90], pages 423– 430. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Richardson:1990:CPC

[Ric90b] John L. Richardson. Computational physics on the CM-2 supercomputer. Technical report TMC-166, Thinking Machines Corp, Cambridge, MA, USA, September 1990. 26 pp.

Richardson:1991:CPC

[Ric91a] J. L. Richardson. Computational physics on the CM-2 supercomputer. Physics Reports, 207(3-5):305–320, September 1991. CODEN PRPLCM. ISSN 0370-1573.

Richardson:1991:VQS

[Ric91b] John L. Richardson. Visualizing quantum scattering on the CM-2 supercomputer. Computer Physics Communications, 63(1–3): 84–94, February 1, 1991. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/001046559190240L.

Rief:1993:MCP

[Rie93] H. Rief. Monte Carlo perturbation algorithms exploit new computer architectures. In Kusters et al. [KSW93], pages 395–405. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Rigault:1993:COS

[Rig93] P. Rigault. Clone ordering by simulated annealing: Application to the STS-Content map of chromosome 21. In Lim et al. [L^{+}93], pages 169–

176. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Ristic:1994:EMS

[Ris94] M. Ristic. Electronic management system for gas turbine powered hybrid vehicles.

In Anonymous [Ano94-75], pages 287–294. ISBN 0-947719-68-7. LCCN ????

Ritchie:1988:EUC

[Rit88a] Dennis M. Ritchie. Experience with Unicos on the Cray X-MP. Report, Bell Laboratories, Murray Hill, NJ 07974, September 1988. 4 pp. URL https:

                                        //www.bell-labs.com/usr/
                                        dmr/www/earlyunicos.pdf.

Ritchie:1988:ECX

[Rit88b] Dennis M. Ritchie. Experiences with the Cray X/MP. Web site, 1988. URL https:

                                        //www.bell-labs.com/usr/
                                        dmr/www/cray.html.

Ritchie:1988:GFU

[Rit88c] Dennis M. Ritchie. A guest facility for Unicos. Report, Bell Laboratories, Murray Hill, NJ 07974, September 1988. 4 pp. URL https:

                                        //www.bell-labs.com/usr/
                                        dmr/www/unicos.pdf.

Ritz:1997:ACD

[Rit97] J. Ritz. Automotive coatings — dedicated to car lifetime. In Roller [Rol97], pages 33– 44. ISBN 0-947719-88-1 (paperback). LCCN ????

Rivers:1990:SPM

[Riv90] Al Rivers. Supercomputer performance measurements, 1990. 1 videocassette (103 min.).

Rizzi:1994:CNC

[Riz94]  A. Rizzi.  Chapter 10:

Navier–Stokes computations of turbulent transonic flow around aircraft wings. In Murthy and Brebbia [MB94b], pages 215–240. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Reiter:2013:LCI

[RJ13] Edna E. (Edna Elizabeth)

Reiter and Clayton Matthew Johnson. Limits of computation: an introduction to the undecidable and the intractable. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4398-8206-1 (hardback). ???? pp. LCCN QA267.7 .R445 2013.

Resch:2022:BQC

[RK22] Salonik Resch and Ulya R.

Karpuzcu. Benchmarking quantum computers and the impact of quantum noise. ACM Computing Surveys, 54 (7):142:1–142:35, September 2022. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL

https://dl.acm.org/doi/
10.1145/3464420.

Rajesh:1994:PPS

[RKDM94] K. Rajesh, H. K. Kaura, P. S.

Dhekne, and S. M. Mahajan. PSIM — parallel simulator for BPPS. In Mahajan et al. [M^{+}94], pages 374–381. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Ramamoorthy:1977:PA

[RL77] C. V. Ramamoorthy and H. F. Li. Pipeline architecture. ACM Computing Surveys, 9(1):61–102, March 1977. CODEN CMSVAN. ISSN 0010-4892. See also [RL78].

Ramamoorthy:1978:CPA

[RL78] C. V. Ramamoorthy and H. F. Li. Corrigenda: “Pipeline architecture”. ACM Computing Surveys, 10(4): 508, December 1978. CO-DEN CMSVAN. ISSN 0010-4892. See [RL77].

Rihle:1990:SAS

[RL90a] R. Rihle and U. Lang. Scientific applications in a supercomputer environment. In Pitcher [Pit90], pages 91– 101. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Rizzoli:1990:ODY

[RL90b] V. Rizzoli and A. Lipparini. Optimization of the design yield of microwave integrated circuits on a CRAY supercomputer. In Pitcher [Pit90], pages 207–218. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Raghavendara:1996:PGS

[RL96] C. Raghavendara and T.-Y.

Lee. Parallel graphics on supercomputers. In Roller [Rol96], pages 67–76. ISBN 0-947719-81-4. LCCN ????

Rizzoli:1991:MPS

[RLC91] Vittorio Rizzoli, Alessandro Lipparini, and Alessandra Costanzo. Modern perspectives in Supercomputer-Aided microwave circuit design. International journal of microwave and millimete, 1 (2):201–??, April 1, 1991.

Rendell:1993:ECT

[RLKW93] A. P. Rendell, T. J. Lee, A. Komornicki, and S. Wilson. Evaluation of the contribution from triply excited intermediates to the fourth-order perturbation theory energy on Intel distributed memory supercomputers. Theoretica Chimica Acta, 84(4):271–??, January 1993. CODEN TCHAAM. ISSN 0040-5744.

Ruehle:1993:CVS

[RLW93] R. Ruehle, U. Lang, and A. Wierse. Cooperative visualization and simulation in a supercomputer environment.

In Kusters et al. [KSW93], pages 630–641. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Reed:1988:PDE

[RM88] Daniel A. Reed and Allen Davis Malony. Parallel discrete event simulation: the Chandy-Misra approach. Technical Report CSRD 767, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 6 pp.

Rudderman:1992:BFS

[RM92] Randal H. Rudderman and Robert L. Mullen. Biomechanics of the facial skeleton. Clinics in plastic surgery, 19: 11–??, January 1, 1992. CO-DEN CPSUDA. ISSN 0094-1298.

Reid-Miller:1996:LRL

[RM96] Margaret Reid-Miller. List ranking and list scan on the Cray C90. Journal of Computer and System Sciences, 53(3):344– 356, December 1996. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0022000096900744.

Ross:2012:TT

[RM12] P. E. Ross and Samuel K. Moore. Top tech 2012. IEEE Spectrum, 49(1):28– 29, January 2012. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roy:1993:CBT

[RMH93] R. Roy, G. Marleau, and A. Hebert. Consistent Bn theory for slab lattices. In Kusters et al. [KSW93], pages 476–487. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Reed:1987:PDE

[RMM87] Daniel A. Reed, Allen Davis Malony, and Bradley D. Mc-Credie. Parallel discrete event simulation: a shared memory approach. Technical Report CSRD-650, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 15 + [14] pp.

Roberts:2020:RIP

[RMM20] S. Roberts, C. Mann, and C. Marroquin. Redefining IBM power system design for CORAL. IBM Journal of Research and Development, 64(3/4):2:1–2:10, May/ July 2020. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Rumor:1996:GIJ

[RMO96] M. Rumor, R. McMillan, and H. F. L. Ottens, editors. Geographical information: Joint European conference; 2nd — March 1996, Barcelona, Spain. Ohmsha, IOS Press, 1996. ISBN 90-5199-268-8, 4-274-90098-3. LCCN ????

Rahnema:1993:IFM

[RMPW93] F. Rahnema, C. L. Martin, G. R. Parkos, and R. D. Williams. Influence function method for fast estimation of BWR core performance. In Kusters et al. [KSW93], pages 232–242. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Robinson:1985:SIM

[Rob85] Bruce Thomas Robinson. A survey of iterative methods for solving large sparse linear systems in the supercomputer environment. Thesis (M.S.), University of Colorado, Boulder, CO, USA, 1985. viii + 54 pp.

Robbins:1987:TLC

[Rob87] David F. Robbins. The TMAPOL language and compiler: an array processorbased approach to costeffective supercomputing. Thesis (M.S. in electrical engineering and computer science), University of Illinois at Chicago, Chicago, IL, USA, 1987. xiv + 227 pp.

Robinson:1989:CPS

[Rob89] M. L. Robinson. On certain properties stronger than uniform distribution modulo one. Technical report SRC-TR-89-002, Supercomputing Research Center: IDA, Lanham, MD, USA, March 3, 1989. 6 pp.

Robbins:1993:GIR

[Rob93] R. J. Robbins. Genome informatics: Requirements and challenges. In Lim et al. [L^{+}93], pages 17–32. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Rohl:1994:SLT

[Roh94] J. S. Rohl. Sigrid: a language for teaching formal program development. In Balakrishnan [Bal94], pages 299–307. ISBN 0-07-462044-4. LCCN ????

Rojek:2019:MLM

[Roj19] Krzysztof Rojek. Machine learning method for energy reduction by utilizing dynamic mixed precision on GPU-based supercomputers. Concurrency and Computation: Practice and Experience, 31(6):e4644:1–e4644:??, March 25, 2019. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Roller:1996:ATA

[Rol96] D. Roller, editor. Automotive technology and automation Fuzzy systems/soft computing in the automotive and transportation industries, Supercomputer applications in the transportation industries: International symposium; 29th — June 1996, Florence; Italy, number 3 in ISATA -PROCEEDINGS-1996. Automotive Association Limited, ????, 1996. ISBN 0-947719-81-4. LCCN ????

Roller:1997:SDV

[Rol97] D. (Dieter) Roller, editor. Simulation, diagnosis and virtual reality applications in the automotive industry including supercomputer applications. Automotive Automation, Croydon, 1997. ISBN 0-947719-88-1 (paperback). LCCN ????

Roskies:1989:SBS

[Ros89] Ralph Roskies. Supercomputing and biomedical science. Future Generation Computer Systems, 5(2–3): 197–205, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Rose:1993:HBG

[Ros93a] J. F. Rose. A hypertexi bibliography on genetic sequence analysis. In Lim et al. [L^{+}93], pages 597–608. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Roska:1993:ASC

[Ros93b] T. Roska. The analogic single-chip CNN visual supercomputer — a review. Lecture Notes in Computer Science, 719(719):813–821,

????  1993.  CODEN

LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosmond:1993:MNM

[Ros93c] T. E. Rosmond. Multitasking of the NOGAPS model. In Hoffmann and Kauranne [HK93b], pages 364–370. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Roska:1995:CUM

[Ros95] T. Roska. The CNN universal machine — a summary of an analogic supercomputer chip architecture for very high speed visual processing. In Vandoni and Verkerk [VV95], pages 295–298. ISBN 92-9083-069-7. ISSN 0304-2898. LCCN QC770 .E83 v.95, no.1.

Ross:2008:UCC

[Ros08] P. E. Ross. Update: A computer for the clouds. IEEE Spectrum, 45(8):11– 12, August 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ross:2009:CCK

[Ros09] P. E. Ross. Cloud computing’s killer app: Gaming. IEEE Spectrum, 46 (3):14, March 2009. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rothnie:1992:KSR

[Rot92] J. Rothnie. Kendall Square Research introduction to the KSR1. In Meuer [Meu92c], pages 104–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Rothberg:1994:PPB

[Rot94] E. Rothberg. Performance of panel and block approaches to sparse Cholesky factorization on the iPSC/860 and Paragon multiprocessors. In IEEE [IEE94c], pages 324– 333. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Roweth:1986:DPA

[Row86] D. Roweth. Design and performance analysis of Transputer arrays. The Journal of Systems and Software, 6 (1–2):21–22, May 1986. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rauchwerger:1994:PDT

[RP94] L. Rauchwerger and D. Padua. The PRIVATIZING DOALL test: a run-time technique for DOALL loop identification and array privatization. In Anonymous [Ano94-134], pages 33–43. ISBN ???? LCCN ????

Raghu:1994:TCU

[RPY94] P. P. Raghu, R. Poongodi, and B. Yagnanarayana. Texture classification using a combined self-organizing map and multilayer perceptron. In Balakrishnan [Bal94], pages 145–153. ISBN 0-07-462044-4. LCCN ????

Robbins:1989:CXM

[RR89] Kay A. Robbins and Steven Robbins. The Cray X-MP/ Model 24: a case study in pipelined architecture and vector processing, volume 374 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. CODEN LNCSD9. ISBN 0-387-97089-4. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 165 pp. LCCN QA76.8.C72 R631 1989. US$18.90.

Ram:1995:SDS

[RR95] P. Ram and D. K. Rand.

Satan: double-edged sword. Computer, 28(6):82–83, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raghavachari:1999:ALP

[RR99] Mukund Raghavachari and Anne Rogers. Ace: a language for parallel programming with customizable protocols. ACM Transactions on Computer Systems, 17(3):

202–248, August 1999. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tocs/1999-17-3/
p202-raghavachari/.

Ramesh:1994:PEB

[RRMD94] K. Ramesh, K. Rajesh, S. M. Mahajan, and P. S. Dhekne. Performance evaluation of BPPS-Linpack. In Mahajan et al. [M^{+}94], pages 367–373. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Riethmueller:1996:GFM

[RRSG96] T. Riethmueller, D. Rosenkranz, Schimansky, and L. Geier. Granular flow modelled by Brownian particles. In Wolf et al. [WSB96], pages 293– 298. ISBN 981-02-2635-7. LCCN ????

Rhodes:1993:XSB

[RRSS93] J. Rhodes, K. Rempe, K. Smith, and J. Stevens. X-IMAGE: A SIMULATE-3 based graphical user interface for interactive core design. In Kusters et al. [KSW93], pages 642–653. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Rice:1984:ASM

[RRW84] John R. Rice, Calvin Ribbens, and William A. Ward. Algorithm 622: a simple macroprocessor. ACM Transactions on Mathematical Software, 10 (4):410–416, December 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [Lev98].

Riganati:1984:S

[RS84] J. Riganati and P. Schneck.

Supercomputing. Computer, 17(10):97–113, 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rhoades:1985:EME

[RS85] Clifford E. Rhoades, Jr. and K. G. Stevens, Jr. Early MIMD experience on the Cray X-MP. Computer Physics Communications, 37 (1-3):215–221, July 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Robic:1993:HPC

[RS93] B. Robic and J. Silc. Highperformance computing on a honeycomb architecture. Lecture Notes in Computer Science, 734:1–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rawat:1994:PAQ

[RS94a] A. Rawat and A. Shukla.

Parallel algorithms for query optimisation for a computerised system developed using a data base management system. In Mahajan et al. [M^{+}94], pages 395–400. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Reeves:1994:SLP

[RS94b] Alyson Reeves and Mike Stillman. Smoothness of the lexicographic point. Technical report SRC-TR-94-139, Supercomputing Research Center: IDA, Lanham, MD, USA, November 16, 1994. 13 pp.

Reinefeld:1994:WBH

[RS94c] A. Reinefeld and V. Schnecke. Work-load balancing in highly parallel depth-first search. In IEEE [IEE94c], pages 773–780. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Roller:1994:PMM

[RSB94] D. Roller, M. Stolpmann, and M. Bihler. Product modelling as a major integration factor for future CAD systems. In Anonymous [Ano94-75], pages 51–58. ISBN 0-947719-68-7. LCCN ????

Riehle:1995:MRN

[RSRG95] A. Riehle, J. Seal, J. Requin, and S. Gruen. Multielectrode recording of neuronal activity in the motor cortex: Evidence for changes in the functional coupling between neurons. In Herrmann et al. [HWP95], pages 281– 288. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Raymond:1993:ARS

[RT93] P. Raymond and I. Toumi. An approximate Riemann solver for a two-fluid model. In Kusters et al. [KSW93], pages 198–207. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Rossi:1997:NEP

[RT97] T. Rossi and J. Toivanen.

Numerical experiments with a parallel fast direct elliptic solver on Cray T3E. Lecture Notes in Computer Science, 1300:722–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:2020:FSI

[RT20] Timothy W. Robinson and Abhinav Thota. Foreword to the special issue of the Cray User Group. Concurrency and Computation: Practice and Experience, 32 (20):e5755:1–e5755:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ruder:1990:SSC

[Rud90] Hanns Ruder. The supercomputer as spaceship — cosmic radiation on the screen. German research: reports of the DFG, 1:10–??, 1990. ISSN 0172-1526.

Ruehl:1992:ECG

[Rue92] R. Ruehl. Evaluation of compiler generated parallel programs on three multicomputers. In ACM [ACM92b], pages 15–24. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Ruhnau:1995:WLT

[Ruh95] E. Ruhnau. Which logical and temporal concepts should be used in neurocognitive modelling? In Herrmann et al. [HWP95], pages 83–88. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Ruighaver:1991:OMP

[Rui91] A. B. Ruighaver. The optoelectronic multicomputer project: a semi-dataflow approach to large-scale parallel processing. In Anonymous [Ano91q], pages 66–72. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Rulko:1993:VDM

[Rul93] R. P. Rulko. Variational derivation of the multigroup P2 equations and boundary conditions in planar geometry. In Kusters et al. [KSW93], pages 521–534. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Russell:1978:CCS

[Rus78] Richard M. Russell. The CRAY-1 computer system. Communications of the ACM, 21(1):63–72, January 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Robinson:1989:ENM

[RW89] M. L. Robinson and Ferrell S.

Wheeler. On the Euclidean norms of multiples of an integer polynomial. Technical report SRC-TR-89-015, Supercomputing Research Center: IDA, Lanham, MD, USA, 1989. 35 pp.

Rowlan:1994:PCL

[RW94a] J. S. Rowlan and B. T.

Wightman. PORTAL: a communication library for run-time visualization of distributed, asynchronous data. In IEEE [IEE94c], pages 350– 356. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Rutz:1994:MSD

[RW94b] R. Rutz and M. Winkler. Mechatronic suspension design using on-line optimization. In Anonymous [Ano94-75], pages 295–302. ISBN 0-947719-68-7. LCCN ????

Ragade:1994:NCM

[RWCA94] R. K. Ragade, M. Witten, M. A. Cassaro, and N. Y. Alemu. A nucleus of a cooperative model-based information management system for hazards and disaster emergency management. In Balakrishnan [Bal94], pages 351– 369. ISBN 0-07-462044-4. LCCN ????

Rasch:1998:DIS

[RWL^{+}98] J. Rasch, Colm T. Whelan, S. P. Lucey, C. Dal Cappello, and H. R. J. Walters. 6-dimensional integrals and supercomputers. Computer Physics Communications, 114(1–3): 378–384, November 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465598000691.

Reese:1994:PDR

[RWNJ94] E. A. Reese, H. Wilson, D. Nedwek, and J. Jex. A phase-tolerant 3.8GB/ s data-communication router for a multiprocessor supercomputer backplane. In Wuorinen [Wuo94], pages 296–297. ISBN 0-7803-1845-5, 0-7803-1844-7, 0-7803-1846-3. ISSN 0193-6530. LCCN ???? IEEE catalog number 94CH3410-8.

Ryan:1990:SBA

[Rya90] B. Ryan. Separated at birth: Although they’re the same age, PCs and supercomputers are now sharing more than just birthdays. BYTE Magazine, 15(5):207–208, 210, May 1990. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Ryabov:1992:SU

[Rya92] G. G. Ryabov. Supercomputing in the USSR. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 17(5):258, July 1992. CODEN PC-SODA. ISSN 0361-7688 (print), 1608-3261 (electronic).

Ryan:2013:CCS

[Rya13] Mark D. Ryan. Cloud computing security: the scientific challenge, and a survey of solutions. The Journal of Systems and Software, 86(9): 2263–2268, September 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0164121212003378.

Rau:1989:CDS

[RYYT89] B. Ramakrishna Rau, D. W. L.

Yen, Wei Yen, and R. A. Towle. The Cydra 5 departmental supercomputer: design philosophies, decisions, and trade-offs. Computer, 22 (1):12–35, January 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sincovec:1993:PPS

[S^{+}93] Richard F. Sincovec et al., editors. Proceedings of the Sixth SIAM Conference on Parallel Processing for Scientific Computing (March 1993, Norfolk, VA, USA). SIAM Press, Philadelphia, PA, USA, 1993. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Srini:1982:PAC

[SA82] Vason P. Srini and Jorge F. Asenjo. Performance analysis of Cray-1S using dataflow graphs. Modeling and Simulation, Proceedings of the Annual Pittsburgh Conference, pages 645–654, 1982. CO-DEN MSPCD4. ISBN 0-87664-713-1.

Srini:1983:ACA

[SA83] Vason P. Srini and Jorge F. Asenjo. Analysis of Cray-1S architecture. ACM SIGARCH Computer Architecture News, 11(3):194– 206, June 1983. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Sunwoo:1990:FCM

[SA90] Myung Hoon Sunwoo and J. K. Aggarwal. Flexibly coupled multiprocessors for image processing. Journal of Parallel and Distributed Computing, 10(2): 115–129, October 1990. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sundar:1994:PAC

[SA94] R. Sundar and G. Athithan. A parallel algorithm for constructing isovalue surfaces. In Mahajan et al. [M^{+}94], pages 41–46. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Sen:2010:CDL

[SA10a] S. K. Sen and Ravi P. Agarwal. Central difference limit for derivatives of ill-posed functions: Best approach in supercomputing era. Applied Mathematics and Computation, 217(7):3338–3348, December 1, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0096300310009306.

Sen:2010:ZCP

[SA10b] S. K. Sen and Ravi P. Agarwal. Zero-clusters of polynomials: Best approach in supercomputing era. Applied Mathematics and Computation, 215(12):4080–4093, February 15, 2010. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0096300309010066.

Saad:1987:DPN

[Saa87] Y. Saad. On the design of parallel numerical methods in message passing and shared memory environments. Technical Report CSRD-614, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 20 pp.

Saad:1988:PTN

[Saa88] Y. Saad. Preconditioning techniques for nonsymmetric and indefinite linear systems. Technical Report CSRD 792, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 18 + [3] pp.

Saad:1989:KSM

[Saa89] Youcef Saad. Krylov subspace methods on supercomputers. SIAM Journal on Scientific and Statistical Computing, 10(6):1200–1232, November 1989. CODEN SIJCD4. ISSN 0196-5204. Sparse matrix algorithms on supercomputers.

Saad:1993:SIP

[Saa93a] Y. Saad. Supercomputer implementations of preconditioned Krylov subspace methods. In Hussaini et al. [HKS93], pages 107–136. ISBN 0-387-94014-6, 3-540-94014-6. LCCN QA911 .A555 1993.

Saarinen:1993:CHT

[Saa93b]  M. Saarinen.  Computa-

tion of horizontal two-phase flow by PLIM. In Kusters et al. [KSW93], pages 280– 291. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Schulz:2005:SDB

[SAB^{+}05] Martin Schulz, Dong Ahn, Andrew Bernat, Bronis R. de Supinski, Steven Y. Ko, Gregory Lee, and Barry Rountree. Scalable dynamic binary instrumentation for Blue Gene/L. ACM SIGARCH Computer Architecture News, 33(5):9–14, December 2005. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Sistare:1994:SDM

[SABJ94] S. Sistare, D. Allen, R. Bowker, and K. Jourdenais. A scalable debugger for massively parallel message-passing programs. In IEEE [IEE94c], pages 825–832. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Stouch:1994:SSB

[SABK94] T. R. Stouch, H. E. Alper, and D. Bassolino-Klimas. Supercomputing studies of biomembranes. International Journal of Supercomputer Applications, 8(1):6–??, ???? 1994. CODEN IJSAE9. ISSN 0890-2720.

Shvedov:1993:COF

[SAGS93] M. O. Shvedov, V. A. Apse, L. A. Goncharov, and V. S. Shkolnik. Calculation of optimal fuel arrangement in fast reactor using the purposeful variant analysis method. In Kusters et al. [KSW93], pages 799–805. ISBN 3-923704-11-9. LCCN ???? Two volumes. Sahasrabuddhe:1994:EUC

[Sah94a] H. Sahasrabuddhe. Emerging uses of computers. In Balakrishnan [Bal94], pages 58–?? ISBN 0-07-462044-4. LCCN ????

Sahni:1994:CRB

[Sah94b] S. Sahni. Computing on reconfigurable bus architectures. In Balakrishnan [Bal94], pages 386–398. ISBN 0-07-462044-4. LCCN ????

Sahni:1995:DAF

[Sah95] S. Sahni. The DMBC: Architecture and fundamental operations. In ACM [ACM95a], pages 60–66. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Sakamura:2002:EMNb

[Sak02] Ken Sakamura. EIC’s message: a new definition for high-performance computing. IEEE Micro, 22 (2):2, March/April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2002/pdf/m2002.
                                        pdf;   http://www.computer.
                                        org/micro/mi2002/m2002abs.
                                        htm.

Saleh:1989:PCS

[Sal89] Resve A. Saleh. Parallel circuit simulation on supercomputers. Technical Report CSRD 898, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1989. 39 pp.

Salmelin:1995:MSH

[Sal95]  R. Salmelin.  MED in

the study in human cortical function. In Herrmann et al. [HWP95], pages 27–44. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Samba:1985:DIC

[Sam85] Augustine S. Samba. The design and implementation of cost-effective algorithms for direct solution of banded linear systems on the vector processor system 32 supercomputer. NASA contractor report NASA CR-176811, NASA, Washington, DC, USA, 1985. ?? pp.

Sameh:1991:AAG

[Sam91]  Ahmed Sameh.  Algo-

rithms and Applications Group: semiannual report july–december 1990. Technical Report CSRD 1127, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1991. 40 pp.

SDSC:1986:SAS

[San86] San Diego Supercomputer Center. Science at the San Diego Supercomputer Center: annual report to the National Science Foundation 1986. Technical Report GA-A 18847, San Diego Supercomputer Center, San Diego, CA, USA, 1986. various pp.

Sancken:1990:NSF

[San90] Paulette Sancken. The National Science Foundation supercomputing centers, 1990. National Center for Supercomputing Applications, University of Illinois at Urbana-Champaign, Champaign, IL, USA, 1990. 25 pp.

SDSC:1991:CSA

[San91] San Diego Supercomputer Center, San Diego, CA, USA. Computational science advances at the San Diego Supercomputer Center, 1991. 40 pp.

Sankoff:1993:MAG

[San93] D. Sankoff. Models and analyses of genomic evolution. In Lim et al. [L^{+}93], pages 177– 184. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Sanders:1995:HV

[San95] Lisa Sanders. From H-bombs to video. Forbes, 155(7):120– ??, March 27, 1995. CODEN FORBA5. ISSN 0015-6914.

Sarukkai:1990:PVP

[Sar90] Sekhar R. Sarukkai. Performance visualization and prediction of parallel supercomputer programs: an interim report. Technical report 318, Computer Science Dept., Indiana University, Bloomington, IN, USA, November 1990. 67 pp.

Sarma:1991:IGP

[Sar91] Joyashri Sarma. Implementation of graph partioning algorithm on multiprocessing CRAY Y/MP supercomputer. Thesis (M.S.), University of Cincinnati, Cincinnati, OH, USA, 1991. 96 + [63] pp.

Sato:1993:MPU

[Sat93] K. Sato. Motion planning in uncertain environment using the Laplacian potential field. In Kusters et al. [KSW93], pages 388–397. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Scott:1981:CRA

[SB81] N. S. Scott and P. G. Burke.

Calculation of R-matrix angular integrals on the Cray-1. Computer Physics Communications, 26(3 & 4): 419–421, June 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Scott:1982:CMA

[SB82a] N. S. Scott and P. G. Burke.

Calculation of R-matrix angular integrals on the Cray-1. Computer Physics Communications, 26(3–4):419– 421, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Scott:1982:CRM

[SB82b] N. S. Scott and P. G. Burke. The calculation of R-matrix angular integrals on the CRAY-1. Computer Physics Communications, 26 (3–4):419–421, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0010465582901370.

Sharma:1994:CEI

[SB94a] A. Sharma and R. K. Bagga. Computer education — Indian scenario. In Balakrishnan [Bal94], pages 331–340. ISBN 0-07-462044-4. LCCN ????

Solano:1994:ISC

[SB94b] I. Solano and P. Brunet. A 2D/3D integrated scheme for constructive constraint-based modelling. In Anonymous [Ano94-75], pages 455–462. ISBN 0-947719-68-7. LCCN ????

Steele:1994:ARR

[SB94c] Robert D. Steele and Paul G. Backes. Ada and real-time robotics: Lessons learned. Computer, 27(4):49–54, April 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sur:1994:ANF

[SB94d] S. Sur and A. P. W. Boehm.

Analysis of non-strict functional implementations of the dongarra-sorensen eigensolver. In Anonymous [Ano94-134], pages 412–418. ISBN ???? LCCN ????

Sohn:1996:STS

[SB96] A. Sohn and R. Biswas.

Satisfiability test with synchronous simulated annealing on the Fujitsu AP1000 massively-parallel multiprocessor. In ACM [ACM96], pages 213–220. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Schultheiss:2001:USP

[SB01] Bert C. Schultheiss and Erik H. Baalbergen. Utilizing supercomputer power from your desktop. Lecture Notes in Computer Science, 2110:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2110/21100052.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2110/21100052.
pdf.

Sirbu:2018:DDA

[SB18] Alina Sîrbu and Ozalp Babaoglu. A data-driven approach to modeling power consumption for a hybrid supercomputer. Concurrency and Computation: Practice and Experience, 30(9): ??, May 10, 2018. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL https://

onlinelibrary.wiley.com/
doi/abs/10.1002/cpe.4410.

Saarinen:1991:NSN

[SBC91] S. Saarinen, Randall Barry Bramley, and George Cybenko. The numerical solution of neural network training problems. Technical Report CSRD 1089, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1991. 30 pp.

Seyfarth:1994:GEE

[SBF94] B. R. Seyfarth, J. L. Bickham, and M. R. Fernandez. Glenda: An environment for easy parallel programming. In IEEE [IEE94c], pages 637– 641. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Shang:1980:PVT

[SBHW80] J. S. Shang, P. G. Buning, W. L. Hankey, and M. C. Wirth. Performance of a vectorized three-dimensional Navier–Stokes code on the Cray-1 computer. American Institute of Aeronautics and Astronautics Journal, 18(9): 1073–1079, September 1980. CODEN AIAJAH. ISSN 0001-1452.

Schmidt:1990:KBC

[SBJ90] F. Schmidt, B. Burger, and D. Jankov. Knowledge based control of large technical simulations. In Pitcher [Pit90], pages 335–346. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Siewiorek:1982:CSP

[SBN82] Daniel P. Siewiorek, C. Gordon Bell, and Allen Newell. Computer Structures: Principles and Examples. McGraw-Hill computer science series. McGraw-Hill, New York, NY, USA, 1982. ISBN 0-07-057302-6. xvi + 926 pp. LCCN QA76.9.A73

C65.  URL https://
archive.computerhistory.
org/resources/text/bell_
gordon/bell.computer_structures_
principles_and_examples.
1982.102630397.pdf.

Strassburger:1996:PFH

[SBSR96] G. Strassburger, A. Betat, M. A. Scherer, and I. Rehberg. Pattern formation by horizontal vibration of granular material. In Wolf et al. [WSB96], pages 329–

334. ISBN 981-02-2635-7. LCCN ????

Sprangers:1994:SOD

[SBW94] W. Sprangers, T. Bemmerl, and M. Weck. Shape optimal design on supercomputers. In Anonymous [Ano94-75], pages 849–856. ISBN 0-947719-68-7. LCCN ????

Schulthess:2019:RGB

[SBW^{+}19] T. C. Schulthess, P. Bauer, N. Wedi, O. Fuhrer, T. Hoefler, and C. Schär. Reflecting on the goal and baseline for exascale computing: A roadmap based on weather and climate simulations. Computing in Science and Engineering, 21(1):30– 41, January/February 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Smith:1993:PFG

[SBY93] M. A. Smith and Y. Bar-Yam. Pulsed field gel electrophoresis simulations in the diffusive regime. In Lim et al. [L^{+}93], pages 185–196. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Sayeed:2008:MHP

[SBZ^{+}08] Mohamed Sayeed, Hansang Bae, Yili Zheng, Brian Armstrong, Rudolf Eigenmann, and Faisal Saied. Measuring high-performance computing with real applications. Computing in Science and Engineering, 10 (4):60–70, July/August 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Stewart:1991:USE

[SC91a] Kris Stewart and Bob Clover.

Using supercomputing to enhance undergraduate education, 1991. 1 sound cassette (ca. 60 min.).

Stone:1991:CA

[SC91b] Harold S. Stone and John Cocke. Computer architecture in the 1990s. Computer, 24(9):30–38, September 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Swanson:1992:OSM

[SC92] C. D. Swanson and A. Chronopoulos. Orthogonal s-step methods for nonsymmetric linear systems of equations. In ACM [ACM92b], pages 456–465. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Storer:1993:DDC

[SC93] James A. Storer and Martin Cohn, editors. DCC ’93: Data Compression Conference (3rd: March 1993: Snowbird, Utah). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3392-1, 0-8186-3391-3. ISSN 1068-0314. LCCN QA76.9.D33 D38 1993. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Sorrentino:1997:ANN

[SC97] A. Sorrentino and A. Concilio. Artificial neural networks for structural systems identification. In Roller [Rol97], pages 237–244. ISBN 0-947719-88-1 (paperback). LCCN ????

Shriver:1999:SCC

[SC99] Bruce Shriver and Peter Capek. The Seymour Cray computer science and engineering award. Computer, 32(11):35, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co1999/pdf/ry034.
pdf;   http://www.computer.
org/computer/co1999/ry034abs.
htm.

Shen:2004:HPD

[SC04] Xiaohui Shen and Alok Choudhary. A high-performance distributed parallel file system for data-intensive computations. Journal of Parallel and Distributed Computing, 64(10):1157–1167, October 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Sharkawi:2020:CPO

[SC20] S. S. Sharkawi and G. A.

Chochia. Communication protocol optimization for enhanced GPU performance. IBM Journal of Research and Development, 64(3/4): 9:1–9:9, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Scanlon:1992:CTA

[Sca92] J. Scanlon. CAMPUS twotier architecture for massively parallel processing. In Meuer [Meu92c], pages 147–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Saini:2008:PES

[SCG^{+}08] Subhash Saini, Robert Ciotti, Brian T. N. Gunney, Thomas E. Spelce, Alice Koniges, Don Dossa, Panagiotis Adamidis, Rolf Rabenseifner, Sunil R. Tiyyagura, and Matthias Mueller. Performance evaluation of supercomputers using HPCC and IMB benchmarks. Journal of Computer and System Sciences, 74(6): 965–982, September 2008. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0022000007001006.

Sampson:2013:GEC

[SCG13a] A. Sampson, L. Ceze, and D. Grossman. Good-enough computing. IEEE Spectrum, 50(10):54–59, October 2013. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugavanam:2013:DLP

[SCG^{+}13b] K. Sugavanam, C.-Y. Cher, J. A. Gunnels, R. A. Haring, P. Heidelberger, H. M. Jacobson, M. K. McManus, D. P. Paulsen, D. L. Satterfield, Y. Sugawara, and R. Walkup. Design for low power and power management in IBM Blue Gene/Q. IBM Journal of Research and Development, 57(1/2):3:1–3:11, January–March 2013. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Schaefer:1987:PBI

[Sch87a] Mark Johannes Schaefer. A polynomial based iterative method for linear parabolic equations. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. v + 82 pp.

Schneck:1987:SA

[Sch87b] Paul B. Schneck. Supercomputer architecture. Number SECS 31 in The Kluwer international series in engineering and computer science; Parallel processing and fifth generation computing SECS 31. Kluwer Academic, Boston, MA, USA, 1987. ISBN 0-89838-234-4 (??invalid checksum??). 199 pp. LCCN QA76.5 .S27251 1987. US$45.00 (est.).

Schonauer:1987:SCV

[Sch87c] Willi Schonauer. Scientific computing on vector computers. Number 2 in Special topics in supercomputing. Elsevier Science Publishers, Amsterdam, The Netherlands, 1987. ISBN 0-444-70288-1. xii + 488 pp. LCCN QA76.5 .S27346 1987.

Schwandt:1987:IAB

[Sch87d] Hartmut Schwandt. Interval arithmetic block cyclic reduction on vector computers. Journal of Parallel and Distributed Computing, 4(5): 459–487, October 1987. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schachter:1988:BRH

[Sch88a] Lorne H. Schachter. Book review of High-Performance Computer Architecture by Harold S. Stone. Addison-Wesley 1987. ACM SIGARCH Computer Architecture News, 16(3):81–84, June 1988. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Schow:1988:AIC

[Sch88b] Peter H. Schow. Adaptation of the ISODATA clustering algorithm for vector supercomputer execution. Thesis (M.S.), Florida State University, Tallahassee, FL, USA, 1988. ix + 54 pp.

Schatz:1989:WWS

[Sch89a] Willie Schatz. Who’s winning the supercomputer race? Datamation, 35(14):18–21, July 15, 1989. CODEN DTMNAT. ISSN 0011-6963.

Schoen:1989:SSM

[Sch89b]  Martin Schoen.  Struc-

ture of a simple molecular dynamics FORTRAN program optimized for Cray vector processing computers. Computer Physics Communications, 52(2):175–185, January–February 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Schouten:1990:OIA

[Sch90a] Dale Allan Schouten. An overview of interprocedural analysis techniques for high performance parallelizing compilers. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1990. 2 + 62 pp.

Schrader:1990:ATD

[Sch90b] Jennifer A. Schrader. Associative template dataflow machine graph coordinator gate level description. Technical report SRC-TR-91-030, Supercomputing Research Center: IDA, Lanham, MD, USA, May 20, 1990. i + 39 + 25 pp.

Schuette:1990:BBD

[Sch90c] W. Schuette. Blade behaviour during birdstrike. In Pitcher [Pit90], pages 145– 157. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Schlenz:1992:PKA

[Sch92a] H. O. Schlenz. Parallelrechner für kommerzielle Anwendungen. In Meuer [Meu92c], pages 137–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Schonfeld:1992:TCL

[Sch92b] Jonathan F. Schonfeld. The theory of compensated laser propagation through strong thermal blooming. The Lincoln Laboratory journal, 5(1): 131–??, Spring 1992.

Schirm:1993:ETS

[Sch93a] L. Schirm. Emerging trends in supercomputing, with emphasis on image and signal processing. In Anonymous [Ano93n], pages 1092–1096. ISBN ???? LCCN ???? Two volumes.

Schneider:1993:TPS

[Sch93b]  R. Schneider.  Two-fluid

plasma simulations with explicit numerical methods. In Kusters et al. [KSW93], pages 71–80. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Schlesinger:1994:LCH

[Sch94a] Judith D. Schlesinger. A language for characterizing heterogeneous systems. Technical report SRC-TR-94-132, Supercomputing Research Center: IDA, Lanham, MD, USA, October 1994. 20 pp.

Schmidt:1994:HPC

[Sch94b] F. Schmidt. High-performance computing and networking as base element of the Environmental Information System UIS of Baden-Wurttemberg. Lecture Notes in Computer Science, 797:5–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneenman:1994:DSS

[Sch94c] Richard D. Schneenman. Distributed supercomputing software: experiences with the parallel virtual machine — PVM. Technical Report NISTIR 5381, U.S. Dept. of Commerce, National Institute of Standards and Technology, Gaithersburg, MD, USA, 1994. vi + 18 pp.

Souleyrette:1994:USI

[SCH94d] Reginald R. Souleyrette, Daniel R. Croce, and Zachary N. Hans. Use of supercomputer for interactive travel demand modeling through GIS. Technical report, Iowa Transportation Center, Iowa State University, Ames, IA, USA, 1994. 23 pp.

Schenfeld:1995:NTC

[Sch95a] E. Schenfeld. New trends in computer hardware technology for supercomputing. In Anonymous [Ano95u], pages 12–17. ISBN ???? LCCN ????

Schill:1995:IIG

[Sch95b] K. Schill. Inference by information gain: Towards a unification of the “Hybrid” approach. In Herrmann et al. [HWP95], pages 89–98. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Schroder:1995:AOD

[Sch95c] Florian Schröder. apE — the original dataflow visualization environment. Computer Graphics, 29(2):5–9, May 1995. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Schuele:1996:PLA

[Sch96] J. Schuele. Parallel Lanczos algorithm on a CRAY-T3D combining PVM and SHMEM routines. Lecture Notes in Computer Science, 1156:158–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiano:1997:PCC

[Sch97a] P. Schiano, editor. Parallel CFD: Conference — May 1996, Capri, Italy, PAR-ALLEL COMPUTATIONAL FLUID DYNAMICS 1996. North-Holland, Amsterdam, The Netherlands, 1997. ISBN 0-444-82327-1. LCCN ????

Schmeisser:1997:PSP

[Sch97b] K. Schmeisser. Process simulation for the production of CV-Joints. In Roller [Rol97], pages 367–372. ISBN 0-947719-88-1 (paperback). LCCN ????

Schmidt:1997:AGM

[Sch97c] J. H. Schmidt. 3-D adaptive grid modeling for groundwater mechanics. In Delic and Wheeler [DW97], pages 265– 270. ISBN 0-89871-378-1. LCCN ????

Schneider:2012:CST

[Sch12] D. Schneider. Could supercomputing turn to signal processors (again)? IEEE Spectrum, 49(10):13–14, October 2012. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:2018:USS

[Sch18]  D. Schneider.  U.S. su-

percomputing strikes back. IEEE Spectrum, 55(1):52– 53, January 2018. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schoening:2019:SSG

[Sch19] Timm Schoening. SHiPCC — a sea-going high-performance compute cluster for image analysis. Frontiers in Marine Science, 6, November 2019. ISSN 2296-7745.

Schneider:2022:EEU

[Sch22] David Schneider. The Exascale Era is upon us: The Frontier supercomputer may be the first to reach 1,000,000,000,000,000,000 operations per second. IEEE Spectrum, 59(1):34–35, January 2022. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

SAIC:1986:EES

[Sci86] Science Applications International Corporation. Earth and environmental science in the 1980’s: part 1. environmental data systems, supercomputer facilities, and networks. NASA contractor report NASA CR-4029, NASA, Washington, DC, USA, October 1986. various pp.

Shimojo:2000:SMD

[SCK^{+}00] Fuyuki Shimojo, Timothy J. Campbell, Rajiv K. Kalia, Aiichiro Nakano, Priya Vashishta, Shuji Ogata, and Kenji Tsuruta. A scalable molecular-dynamics algorithm suite for materials simulations: designspace diagram on 1024 Cray T3E processors. Future Generation Computer Systems, 17(3):279–291, November 1, 2000. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/28/29/abstract.
html.

Scott:1996:GC

[Sco96] Steve Scott. The GigaRing channel. IEEE Micro, 16(1):27–34, January/ February 1, 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Scroggs:1988:SPP

[Scr88] Jeffrey Scott Scroggs. The solution of a parabolic partial differential equation via domain decomposition: the synthesis of asymptotic and numerical analysis. Technical Report CSRD 791; UILU-ENG-88-8005, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. ix + 146 pp.

Shi:2012:VGA

[SCSL12] Lin Shi, Hao Chen, Jianhua Sun, and Kenli Li. vCUDA: GPU-accelerated high-performance computing in virtual machines. IEEE Transactions on Computers, 61(6):804–816, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Samaras:2001:SFI

[SCV01] William A. Samaras, Naveen Cherukuri, and Srinivas Venkataraman. Special feature: The IA-64 Itanium processor cartridge: For highperformance computing in a multiprocessing system environment, consider this innovative packaging solution. IEEE Micro, 21(1):82–89, January/February 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stricker:1988:NSC

[SD88] Rudolf Stricker and Alfred Dick. Numerical simulation of car aerodynamics and interior climate. In Marino [Mar88b], pages 99– 123. ISBN ???? LCCN TL240 .I528 1988.

Sguazzero:1992:PDC

[SD92] P. Sguazzero and R. Di Antonio. Parallel/distributed computing on clusters of high performance RISC workstations. In Meuer [Meu92c], pages 115–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Searls:1993:SPR

[SD93] D. Searls and S. Dong. A syntactic pattern recognition system for DNA sequences. In Lim et al. [L^{+}93], pages 89–102. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Srinath:1994:GDS

[SDB94] B. G. Srinath, S. P. Dharne, and R. Bhargava. A graphics display system for ISE. In Mahajan et al. [M^{+}94], pages 61–66. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Siegel:1993:VHS

[SDFP93] H. J. Siegel, H. G. Dietz, R. F. Freund, and C. Pangali. The virtual heterogeneous supercomputer: Can it be built? In IEEE [IEE93b], pages 30–33. ISBN 0-8186-3900-8, 0-8186-3901-6. LCCN QA76.9.D5 I593 1993. IEEE catalog number 93TH0550-4.

Steele:1998:SNS

[SDK98] C. Steele, J. Draper, and J. Koller. Safety net: Secure communications for embedded high-performance computing. Lecture Notes in Computer Science, 1388: 908–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schow:2010:GOB

[SDK10] C. Schow, F. Doany, and J. Kash. Get on the optical bus. IEEE Spectrum, 47(9): 32–56, September 2010. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strohmaier:1999:MHP

[SDMS99] Erich Strohmaier, Jack J.

Dongarra, Hans W. Meuer, and Horst D. Simon. The marketplace of high-performance computing. Parallel Computing, 25(13–14):1517–1544, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/35/21/32/36/24/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/32/36/
24/article.pdf;   http:/
/www.netlib.org/utk/people/
JackDongarra/PAPERS/pc-
benchmarking.pdf.

Silver:1990:EWS

[SE90] D. M. Silver and L. W.

Ehrlich. Electromagnetic wave scattering from periodic surfaces. In Pitcher [Pit90], pages 509–520. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Silberman:1992:AFM

[SE92] G. M. Silberman and K. Ebcioglu. An architectural framework for migration from CISC to higher performance platforms. In ACM [ACM92b], pages 198–215. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Sumiyoshi:1998:PPS

[SE98] K. Sumiyoshi and T. Ebisuzaki. Performance of parallel solution of a block-tridiagonal linear system on Fujitsu VPP500. Parallel Computing, 24(2):287–304, February 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/parco/sub/1998/
                                        24/2/1284.pdf.

SEAS:1984:PSA

[SEA84]  SEAS, editor.  Proceed-

ings — SEAS Anniversary Meeting, September 24-28, 1984, Distributed Intelligence, Garmisch-Partenkirchen, West Germany. SHARE European Association, Amsterdam (??), The Netherlands, 1984. LCCN ????

Seager:1986:PCG

[Sea86] Mark K. Seager. Parallelizing conjugate gradient for the Cray X-MP. Parallel Computing, 3(1):35– 47, March 1986. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Sehr:1988:OEP

[Seh88] David C. Sehr. OR-parallel execution of Prolog programs with side effects. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1988. vii + 110 pp.

Sehr:1992:APP

[Seh92] David C. Sehr. Automatic parallelization of Prolog programs. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1992. vi + 294 pp.

Sander:1998:HPC

[SEH98] V. Sander, D. Erwin, and V. Huber. High-performance computer management based on Java. Lecture Notes in Computer Science, 1401: 526–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sander:1999:HCM

[SEH99a] Volker Sander, Dietmar Erwin, and Valentina Huber. High-performance computer management based on Java. Future Generation Computer Systems, 15(3):425–??, ???? 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Sander:1999:HPC

[SEH99b] Volker Sander, Dietmar Erwin, and Valentina Huber. High-performance computer management based on Java. Future Generation Computer Systems, 15(3):425– 432, April 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/30/19/27/abstract.
                                        html.

Seibold:1994:PMD

[Sei94] W. Seibold. Product modelling for design, failure mode effect analysis and diagnosis with the model based functional analysis tool roon. In Anonymous [Ano94-75], pages 75–82. ISBN 0-947719-68-7. LCCN ????

Sela:1995:WFP

[Sel95] J. G. Sela. Weather forecasting on parallel architectures. Parallel Computing, 21(10): 1639–1654, October 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Serbedzija:1998:WSE

[Ser98] Nikola B. Serbedzija. The Web supercomputing environment. Computer Networks and ISDN Systems, 30 (1–7):742–744, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/comnet/sub/1998/
30/1-7/1838.pdf.

Schaar:1994:QPA

[SES94] M. A. Schaar, K. Efe, and W. Shang. Queueing performance analysis of co-scheduling in a pool of processors environment. In Anonymous [Ano94-134], pages 313–322. ISBN ???? LCCN ????

Scott:2009:THR

[SEV^{+}09] Stephen L. Scott, Christian Engelmann, Geoffroy R. Vallée, Thomas Naughton, Anand Tikotekar, George Ostrouchov, Chokchai Leangsuksun, Nichamon Naksinehaboon, Raja Nassar, Mihaela Paun, Frank Mueller, Chao Wang, Arun B. Nagarajan, and Jyothish Varma. A tunable holistic resiliency approach for high-performance computing systems. ACM SIGPLAN Notices, 44(4): 305–306, April 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schaefer:1982:CBS

[SF82] D. H. Schaefer and J. R. Fischer. Computers: Beyond the supercomputer: NASA is developing ultrafast machines to process satellite data that would choke a mere supercomputer. IEEE Spectrum, 19(3):32–37, March 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shur:1991:SSSb

[SF91] Michael Shur and T. A. Fjeldly, editors. Supercomputer simulation of semiconductor devices: proceedings of the Symposium on Supercomputer Simulation of Semiconductor Devices, Minneapolis, November 1990, volume 67(1) of Computer physics communications. North-Holland, Amsterdam, The Netherlands, 1991. LCCN QC1 .C742 v.67, no.1.

Schmitz:1993:MPM

[SF93a] F. Schmitz and U. Fischer. MCNP4, a parallel Monte Carlo implementation on a workstation network. In Kusters et al. [KSW93], pages 384–394. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Suslov:1993:SSP

[SF93b] I. R. Suslov and V. I. Folomeev. Solving the spacetime problem on the basis of quasi-stationary derivatives. In Kusters et al. [KSW93], pages 658–669. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Santana:2021:AMU

[SFC^{+}21] Alexandre Santana, Vinicius Freitas, Márcio Castro, Laércio L. Pilla, and Jean-François Méhaut. ARTful: a model for user-defined schedulers targeting multiple high-performance computing runtime systems. Software —Practice and Experience, 51(7):1622–1638, July 2021. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Smith:1994:PSM

[SFF94] W. Smith, T. R. Forester, and D. Fincham. Parallel supercomputing and molecular dynamics. In Gruber and Tomassini [GT94], pages 95–98. ISBN 2-88270-011-3. LCCN ????

Schoinas:1994:FAC

[SFL^{+}94] Ioannis Schoinas, Babak Falsafi, Alvin R. Lebeck, Steven K. Reinhardt, James R. Larus, and David A. Wood. Fine-grain access control for distributed shared memory. ACM SIGPLAN Notices, 29(11):297–306, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://

www.acm.org:80/pubs/citations/

                                        proceedings/asplos/195473/
                                        p297-schoinas/.

Saunders:1981:ACQ

[SG81] V. R. Saunders and M. F. Guest. Applications of the Cray-1 for quantum chemistry calculations. Computer Physics Communications, 26 (3 & 4):389–395, June 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Saunders:1982:ACQ

[SG82] V. R. Saunders and M. F. Guest. Applications of the Cray-1 for quantum chemistry calculations. Computer Physics Communications, 26 (3–4):389–395, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/0010465582901333.

Sarukkai:1992:PPV

[SG92a] S. R. Sarukkai and D. Gannon. Parallel program visualization using SIEVE.1. In ACM [ACM92b], pages 157–166. ISBN 0-89791-485-6 (paperback), 0-89791-486-

4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Schlesinger:1992:DRM

[SG92b] Judith D. Schlesinger and Maya B. Gokhale. DBC reference manual. Technical report SRC-TR-92-068, Supercomputing Research Center:

IDA, Lanham, MD, USA, June 1992. 102 pp.

Simoncini:1992:IMN

[SG92c] V. Simoncini and E. J. (Efstratios J.) Gallopoulos. An iterative method for nonsymmetric systems with multiple right-hand sides. Technical Report CSRD 1242, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1992. 22 pp.

Simoncini:1992:MHM

[SG92d] V. Simoncini and E. J.

(Efstratios J.) Gallopoulos. A memory-conserving hybrid method for solving linear systems with multiple right hand sides: extended abstract. Technical Report CSRD 1203, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1992. 9 pp.

Srinivasan:1994:DIR

[SG94a] B. Srinivasan and D. Ghazfan. Distributed information retrieval system. In Balakrishnan [Bal94], pages 409–?? ISBN 0-07-462044-4. LCCN ????

Srinivasan:1994:PAM

[SG94b] R. Srinivasan and C. D.

Geethalathai. Parallel architectures — a monograph. In Mahajan et al. [M^{+}94], pages 329–346. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Smith:1991:IDE

[SGB91] Stuart Smith, Georges G. Grinstein, and R. Daniel Bergeron. Interactive data exploration with a supercomputer. Technical report 91-11, Dept. of Computer Science, University of New Hampshire, Durham, NH, USA, 1991. 12 pp.

Summers:1997:ASS

[SGH97] B. G. Summers, E. Gentry, and B. Helland. Adventures in supercomputing: Scientific exploration in an era of change. In EP Innovations [EP 97], pages 1395– 1398. ISBN 0-7803-4087-6 (casebound), 0-7803-4086-8 (softbound), 0-7803-4088-4 (microfiche), 0-7803-4089-2 (CD-ROM). ISSN 0190-5848. LCCN T62 .F76 1997. Three volumes. IEEE catalog number: 97CH36099.

Smith:1993:ARG

[SGIS93] C. L. Smith, D. Grothues, T. Ito, and T. Sano. Accelerating the rate of genome mapping and sequencing: Will computational biology keep up? In Lim et al. [L^{+}93], pages 3–16. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Stunkel:2020:HSN

[SGS^{+}20] C. B. Stunkel, R. L. Graham, G. Shainer, M. Kagan, S. S. Sharkawi, B. Rosenburg, and G. A. Chochia. The high-speed networks of the Summit and Sierra supercomputers. IBM Journal of Research and Development, 64(3/4):3:1–3:10, May/ July 2020. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Sammur:1990:MSP

[SH90] N. M. Sammur and M. T.

Hagan. Mapping signal processing algorithms on parallel architectures. Journal of Parallel and Distributed Computing, 8(2): 180–185, February 1990. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Schonauer:1991:PES

[SH91] W. Schonauer and H. Hafner.

Performance estimates for supercomputers: the responsibilities of the manufacturer and of the user. Parallel Computing, 17(10–11):1131– 1149, December 1991. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Schoenauer:1993:EGB

[SH93] W. Schoenauer and H. Haefner.

Explaining the gap between theoretical peak performance and real performance for supercomputer architectures [invited]. In Kusters et al. [KSW93], pages 75–88. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Schoenauer:1994:SAT

[SH94a] W. Schoenauer and H. Haefner. Supercomputer architectures and their bottlenecks. In Joubert et al. [JPTE94], pages 411–420. ISBN 0-444-81841-3. LCCN QA76.58 .P3794 1993.

Schonauer:1994:EGB

[SH94b] W. Schonauer and H. Hafner. Explaining the gap between theoretical peak performance and real performance for supercomputer architectures. Scientific Programming, 3 (2):157–168, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Shah:1987:USS

[Sha87] Devendra Mangulal Shah. Using SX2 supercomputer system management facilities for management information purposes. Thesis (M.S.), Dept. of Computer Science, University of Houston, Houston, TX, USA, 1987. ix + 225 pp.

Shah:1989:NHT

[Sha89] R. K. Shah, editor. Numerical heat transfer with personal computers and supercomputing: presented at the 1989 National Heat Transfer Conference, Philadelphia, Pennsylvania, August 6–9, 1989, volume 110 of HTD; vol. 110. American Society of Mechanical Engineers, United Engineering Center, 345 E. 47th St., New York, NY 10017, USA, 1989. ISBN 0-7918-0352-X. LCCN TJ260 .N85 1989.

Sharma:1990:RVC

[Sha90] Sanjay Sharma. Real-time visualization of concurrent processes. Technical Report CSRD 999, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1990. 11 pp.

Shimotsuji:1992:RRS

[SHA^{+}92] Shigeyoshi Shimotsuji, Osamu Hori, Mieko Asano, Kaoru Suzuki, Fumihiko Hoshino, and Toshiaki Ishii. A robust recognition system for a drawing superimposed on a map. Computer, 25 (7):56–59, July 1992. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shah:1994:PNS

[Sha94a] D. N. Shah. Psoul — a novel semantic model for a sequential to parallel conversion automation. In Mahajan et al. [M^{+}94], pages 389–394. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Shapiro:1994:CFE

[Sha94b]  R. A. Shapiro.  Chap-

ter 1: Finite element algorithms for the Euler equations on the Connection Machine. In Murthy and Brebbia [MB94b], pages 1–20. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Sharp:1995:GC

[Sha95a] Oliver Sharp. The grand challenges. BYTE Magazine, 20(2):65–??, February 1995. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Sharp:1995:GCR

[Sha95b] Oliver Sharp. The grand challenges: Researchers are beginning to tackle problems in geography, weather, and other areas that require more computing capability than today’s most powerful computers can muster. here’s a look at the biggest of these challenges and the ways in which scientists are attacking them with supercomputers. BYTE Magazine, 20(2): 65–??, February 1995. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shapiro:1996:MS

[Sha96] E. Shapiro. The metamorphosis of the supercomputer. In Dekker et al. [DSZ96], pages 13–20. ISBN 0-444-82559-2. LCCN ????

Smitley:1991:HHN

[SHB91] David L. Smitley, Frank Hady, and Dan Burns. Hnet: a high-performance network evaluation testbed. Technical report SRC-TR-91-049, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1991. 14 pp.

Szczepanski:2013:DAV

[SHB^{+}13] A. F. Szczepanski, Jian Huang, T. Baer, Y. C. Mack, and S. Ahern. Data analysis and visualization in high-performance computing. Computer, 46(5): 84–92, May 2013. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shear:1990:EDD

[She90] David Shear. EDN’s DSPchip directory. EDN, 35(21): 171–??, October 11, 1990. CODEN EDNSBH. ISSN 0012-7515, 0364-6637.

Shenoy:1993:AHT

[She93] Meera Shenoy. ANURAG:

high-tech talk. Business India, 397:21–??, May 24, 1993. ISSN 0254-5268.

Stanley:1995:UPS

[SHG95] B. J. Stanley, C. Halloy, and G. Guiochon. Using parallel supercomputers to calculate surface energy distributions. Journal of Chemical Information and Computer Sciences, 35(1):110–??, ???? 1995. CODEN JCISD8. ISSN 0095-2338.

Shiba:1995:HSS

[Shi95] H. Shiba. High-speed supercomputer and studies strongly correlated quantummechanical systems. Gakujutsu geppo. Japanese scientific monthly, 48(5):12–??, May 1, 1995. ISSN 0387-2440.

Solt:2020:SFT

[SHL^{+}20] D. Solt, J. Hursey, A. Lauria, D. Guo, and X. Guo. Scalable, fault-tolerant job step management for highperformance systems. IBM Journal of Research and Development, 64(3/4):8:1– 8:9, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Sigurdsson:1997:IEC

[SHMH97] Steinn Sigurdsson, Bohr He, Rami Melhem, and Lars Hernquist. Implementing an efficient collisionless Nbody code on the Cray T3D. Computers in Physics, 11 (4):378–??, July 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

/aip.scitation.org/doi/
10.1063/1.168608.

Shankar:1993:SSA

[SHMR93] V. Shankar, W. F. Hall, A. Mohammadian, and C. Rowell. Science and supercomputing in the aerospace industry. In AIAA [AIA93], pages AIAA–93–4674. ISBN ???? LCCN ????

Shankar:1994:AAS

[SHMR94] V. Shankar, W. F. Hall, A. Mohammadian, and C. Rowell. Algorithmic aspects and supercomputing trends in computational electromagnetics. In Engl and McLaughlin [EM94b], pages 239–269. ISBN 3-519-02179-X. LCCN ????

Shankar:1996:AAS

[SHMR96] V. Shankar, W. F. Hall, A. Mohammadian, and C. Rowell. Algorithmic aspects and supercomputing trends in computational electromagnetics. In Anonymous [Ano96q], pages 103–121. ISBN 0-309-05337-4. LCCN TA646 .L35 1996.

Shores:1991:SDA

[Sho91] George Austin Shores. SUperNetwork and Distributed Application System (SUN-DAS): a theoretical supercomputer architecture. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA, 1991. vii + 87 pp.

Stefanovici:2015:BBB

[SHS15] Ioan Stefanovici, Andy Hwang, and Bianca Schroeder. Battling borked bits. IEEE Spectrum, 52(12):34–53, December 2015. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shu:1988:RPM

[Shu88] Wennie Shu. The reduceor process model for logic programs on parallel machines. Technical Report CSRD 737, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 17 pp.

Schwamborn:1994:CNS

[SHZK94] D. Schwamborn, A. Hilgenstock, H. Zimmermann, and W. Kordulla. Chapter 11: On numerical simulation of 3-D viscous transonic flows with separation. In Murthy and Brebbia [MB94b], pages 241–272. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Smitley:1990:BSC

[SI90] David L. Smitley and Ken Iobst. Bit-serial SIMD on the CM-2 and the Cray 2. Technical report SRC-TR-90-003, Supercomputing Research Center: IDA, Lanham, MD, USA, January 1990. 23 pp.

Smitley:1991:BSC

[SI91a] David Smitley and Kent Iobst. Bit-serial SIMD on the CM-2 and the Cray-2. Journal of Parallel and Distributed Computing, 11(2): 135–145, February 1991. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Smitley:1991:BSS

[SI91b] David Smitley and Kent Iobst. Bit-serial SIMD on the CM-2 and the Cray-2. Journal of Parallel and Distributed Computing, 11(2): 135–145, February 1991. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shindo:1995:HCA

[SIDH95] T. Shindo, H. Iwashita, T. Doi, and J. Hagiwara. HPF compiler for the AP1000. In ACM [ACM95a], pages 190–194. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Siegel:1990:INL

[Sie90]  Howard Jay Siegel.  In-

terconnection networks for large-scale parallel processing: theory and case studies. McGraw-Hill series in computer organization and architecture; McGraw-Hill series in supercomputing and parallel processing. McGraw-Hill, New York, NY, USA, second edition, 1990. ISBN 0-07-057561-4. xx + 390 pp. LCCN TK5105.5 .S54 1990.

Siegel:1994:PEI

[Sie94] Howard Jay Siegel, editor. Proceedings / Eighth International Parallel Processing Symposium, April 26–29, 1994, Cancun, Mexico, International Parallel Processing Symposium. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5602-6, 0-8186-5601-8. ISSN 1063-7133. LCCN QA 76.58 I56 1994.

Sigarch:1989:CP

[Sig89] Sigarch, editor. Conference proceedings. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-309-4. LCCN QA 76.88 I57 1989. Sponsored by ACM SIGARCH.

Sigarch:1990:CPI

[Sig90a]  Sigarch, editor.  Confer-

ence proceedings, 1990 International Conference on Supercomputing: June 11-15, 1990, Amsterdam, the Netherlands, Computer architecture news / ACM SIGARCH; v. 18, no. 3 (Sept. 1990). ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-369-8. LCCN QA1 .A85457. ACM order number: 415902.

SIGGRAPH:1990:SVR

[SIG90b] SIGGRAPH. SIGGRAPH video review: Supercomputing ’90 visualization theater, 1990. 1 videocassette (ca. 100 min.).

Sigarch:1995:CPI

[Sig95]  Sigarch, editor.  Confer-

ence proceedings of the 1995 International Conference on Supercomputing, Barcelona, Spain, July 3-7, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995.

Shindo:1994:TDL

[SIKD94] T. Shindo, H. Iwashita, S. Kaneshiro, and T. Doi. Twisted data layout. In Anonymous [Ano94-134], pages 374–381. ISBN ???? LCCN ????

Silcox:1991:MMS

[Sil91] John Silcox. The microstructure of materials: Studies using a new kind of electron microscope and a supercomputer. Engineering: Cornell Quarterly, 25(2 / 3):48–??, Spring 1991. ISSN 0013-7871.

Simon:1992:EMP

[Sim92a]  H. D. Simon.  Experi-

ence with massive parallelism for CFD applications at NASA Ames Research Center. In Meuer [Meu92c], pages 122–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Simon:1992:PCF

[Sim92b] Horst D. Simon, editor. Parallel Computational Fluid Dynamics: Implementations and Results. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-19326-4. x + 345 pp. LCCN QA911.P36 1992. US$50.00. URL http://www.

mitpress.com/book-home.
tcl?isbn=0262193264.

Simon:1997:SRR

[Sim97] Horst D. Simon. Site report: Reinventing the supercomputer center at NERSC. IEEE Computational Science & Engineering, 4(3): 14–17, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1997/pdf/c3014.
pdf.

Simoncini:2000:BRN

[Sim00] Valeria Simoncini. Book review: Numerical linear algebra for high-performance computers. Mathematics of Computation, 69(231):1307– 1309, July 2000. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/mcom/
2000-69-231/S0025-5718-
00-01249-7/bookrev-S0025-
5718-00-01249-7.html;
http://www.ams.org/mcom/
2000-69-231/S0025-5718-
00-01249-7/S0025-5718-00-
01249-7.dvi;   http://www.
ams.org/mcom/2000-69-231/
S0025-5718-00-01249-7/S0025-
5718-00-01249-7.pdf;   http:
//www.ams.org/mcom/2000-
69-231/S0025-5718-00-01249-
7/S0025-5718-00-01249-7.
ps;   http://www.ams.org/
mcom/2000-69-231/S0025-
5718-00-01249-7/S0025-5718-
00-01249-7.tex.

Singh:1994:CRT

[Sin94a] Avtar Singh, editor. Conference record of the Twentyeighth Asilomar Conference on Signals, Systems and Computers, October 30– November 2, 1994, Pacific Grove, California, Asilomar Conference on; Signals Systems and Computers. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6405-3, 0-8186-6406-1, 0-7803-2614-8. ISSN 1058-6393. LCCN TK 5102.5 A78 1994. Two volumes.

Singh:1994:PA

[Sin94b] G. S. Singh. Parallel algorithms. In Mahajan et al. [M^{+}94], pages 347–?? ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Sinha:1994:NF

[Sin94c] U. N. Sinha. Nal — Flosolver. In Mahajan et al. [M^{+}94], pages 24–?? ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Singh:2008:BDC

[Sin08a] Gary Singh. Back to digital crayons. IEEE Computer Graphics and Applications, 28(5):4–5, September/October 2008. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Singh:2008:UTH

[Sin08b] S. Singh. Update: Tata hopes its supercomputer is a money machine. IEEE Spectrum, 45(6):18, June 2008. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Singh:2018:GST

[Sin18] Parampreet Singh. Glimpses of space-time beyond the singularities using supercomputers. Computing in Science and Engineering, 20(4):26–38, July/ August 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2018/04/mcs2018040026.

html. See erratum [Ano18].

Sites:1978:PTS

[Sit78]  Richard L. Sites.  Pro-

gramming tools: statement counts and procedure timings. ACM SIGPLAN Notices, 13(12):98–101, December 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srinivas:1994:TGN

[SJA94] M. K. Srinivas, J. Jacob, and V. D. Agrawal. Test generation for non-scan sequential circuits. In Balakrishnan [Bal94], pages 373–?? ISBN 0-07-462044-4. LCCN ????

Smartt:1996:TWR

[SJD96] H. B. Smartt, J. A. Johnson, and S. A. David, editors. Trends in welding research: International conference; 4th — June 1995, Gatlingburg, TN. ASM, ????, 1996. ISBN 0-87170-567-2. LCCN ????

Stantchev:2009:UGP

[SJDV09] George Stantchev, Derek Juba, William Dorland, and Amitabh Varshney. Using graphics processors for highperformance computation and visualization of plasma turbulence. Computing in Science and Engineering, 11 (2):52–59, March/April 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sundar:1994:CEM

[SJPS94] N. S. Sundar, D. N. Jayasimha, D. K. Panda, and P. Sadayappan. Complete exchange in 2D meshes. In IEEE [IEE94c], pages 406– 413. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Sundar:1996:HAC

[SJPS96] N. S. Sundar, D. N. Jayasimha, D. K. Panda, and P. Sadayappan. Hybrid algorithms for complete exchange in 2D meshes. In ACM [ACM96], pages 181–188. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Stroschein:2005:BSC

[SJR05] Josh Stroschein, Doug Jennewein, and Joe Reynoldson. Bioinformatics supercomputing cluster: Applications of parallel computing. Linux Journal, 2005 (133):8, May 2005. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Sehr:1992:EIP

[SK92] David C. Sehr and Laxmikant Vasudeo Kale. Estimating the inherent parallelism in Prolog programs. Technical Report CSRD 1221, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1992. 8 pp.

Schlesinger:1993:AAD

[SK93a] Judith D. Schlesinger and James T. Kuehn. adl: an architecture description language: version 1.0. Technical report SRC-TR-93-104, Supercomputing Research Center: IDA, Lanham, MD, USA, September 1993. 50 pp.

Su:1993:OKB

[SK93b] S. Y. W. Su and N. Kamel.

An object-oriented knowledge base management technology for supporting scientific research and applications. In Lim et al. [L^{+}93], pages 309–330. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Schlesinger:1994:TMT

[SK94] Judith D. Schlesinger and Daniel J. Kopetzky. Terasys, microcode, and TWIST. Technical report SRC-TR-94-119, Supercomputing Research Center: IDA, Lanham, MD, USA, April 1994. 75 pp.

Sonoda:1993:PSM

[SKAT93] Y. Sonoda, S. Kanemoto, Y. Ando, and Y. Takayama. The plant status monitoring system for a boiling water reactor. In Kusters et al. [KSW93], pages 806– 817. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Schellingerhout:1989:CFC

[SKB89] N. W. Schellingerhout, L. P. Kok, and G. D. Bosveld. Configuration-space Faddeev calculations: Supercomputer accuracy on a personal computer. Physical review A: General physics, 40(10): 5568–??, November 15, 1989.

Samuel:2020:ARR

[SKB^{+}20] David Samuel, Syeduzzaman Khan, Cody J. Balos, Zachariah Abuelhaj, Anthony D. Dutoi, Chadi Kari, David Mueller, and Vivek K. Pallipuram. A2Cloud-RF: a random forest based statistical framework to guide resource selection for highperformance scientific computing on the cloud. Concurrency and Computation: Practice and Experience, 32 (24):e5942:1–e5942:??, December 25, 2020. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schraml:2002:HPC

[SKC02] Stephen J. Schraml, Kent D. Kimsey, and Jerry A. Clarke.

High-performance computing applications for survivabilitylethality technologies. Computing in Science and Engineering, 4(2):16–21, March/ April 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c2016abs.htm;   http://
dlib.computer.org/cs/books/
cs2002/pdf/c2016.pdf.

Skeel:1987:WIS

[Ske87] Robert D. Skeel. Waveform iteration and the shifted Picard splitting. Technical Report CSRD 700, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 40 pp.

Skeel:1989:MDS

[Ske89] Robert D. Skeel. Macromolecular dynamics on a shared-memory multiprocessor. Technical Report CSRD 929; Numerical Computing Group 89-5, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1989. 15 pp.

Sakakibara:1994:IMA

[SKIY94] T. Sakakibara, K. Kitai, T. Isobe, and S. Yazawa. An interprocessor memory access arbitrating scheme for the S-3800 vector supercomputer. In Horiguchi et al. [HHK94], pages 262–269. ISBN 0-8186-6507-6, 0-8186-6506-8. LCCN QA76.58.I5673 1994.

Sakakibara:1997:IMA

[SKIY97] T. Sakakibara, K. Kitai, T. Isobe, and S. Yazawa. Interprocessor memory access arbitrating scheme for TCMP type vector supercomputer. IEICE transactions on information and systems, 80(9): 925–??, ???? 1997. ISSN 0916-8532.

Shimada:1990:LSM

[SKK^{+}90] Y. Shimada, Y. Kobayashi, K. Kata, M. Kurano, and H. Takamizawa. Large scale multilayer Glass-Ceramic substrate for supercomputer. IEEE transactions on components, hybrids, and manufacturing technology, 13(4):751– ??, December 1, 1990. CO-DEN ITTEDR. ISSN 0148-6411.

Shen:2003:HPA

[SkLC^{+}03] Xiaohui Shen, Wei keng Liao, Alok Choudhary, Gokhan Memik, and Mahmut Kandemir. A high-performance application data environment for large-scale scientific computations. IEEE Transactions on Parallel and Distributed Systems, 14(12):1262–1274, December 2003. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
td/2003/12/l1262abs.htm;
 http://csdl.computer.org/
dl/trans/td/2003/12/l1262.
pdf.

Suganuma:1996:DGO

[SKN96] T. Suganuma, H. Komatsu, and T. Nakatani. Detection and global optimization of reduction operations for distributed parallel machines. In ACM [ACM96], pages 18–25. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Sehr:1991:FTF

[SKP91] David C. Sehr, Laxmikant Vasudeo Kale, and David A. Padua. Fortran-style transformations for functional programs: extended abstract. Technical Report CSRD 1171, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1991. 2 pp.

Sangireddy:2004:LPH

[SKS04] Rama Sangireddy, H. Kim, and A. K. Somani. Lowpower high-performance reconfigurable computing cache architectures. IEEE Transactions on Computers, 53(10): 1274–1290, October 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?tp=&
                                        arnumber=1327578.

ShantiKumar:1994:IPA

[SKSD94] C. Shanti Kumar, S. N. Kak, K. K. Sinha, and K. K. Dwivedy. Image processing of airborne geophysical data — a better visualisation, processing and analysis technique. In Mahajan et al. [M^{+}94], pages 321–326. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Schukin:1993:MMC

[SKVZ93] N. V. Schukin, A. S. Korsun, S. G. Vitruk, and V. G. Zimin. Mathematical models and computer codes for the analysis of advanced fast reactor dynamics. In Kusters et al. [KSW93], pages 670– 682. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Skytt:1994:SE

[Sky94] V. Skytt. Surface editing. In Anonymous [Ano94-75], pages 447–454. ISBN 0-947719-68-7. LCCN ????

Smitley:1988:CAH

[SL88] David L. Smitley and Insup Lee. Comparative analysis of hill climbing mapping algorithms. Technical report SRC-TR-88-016, Supercomputing Research Center: IDA, Lanham, MD, USA, October 1988. 24 pp.

Shyu:1990:SSC

[SL90] S. J. Shyu and R. C. T.

Lee. Solving the set cover problem on a supercomputer. Parallel Computing, 13(3): 295–300, March 1990. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Seznec:1992:IPS

[SL92] André Seznec and Jacques Lenfant. Interleaved parallel schemes: improving memory throughput on supercomputers. ACM SIGARCH Computer Architecture News, 20(2):246–255, May 1992. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Sato:1993:INC

[SL93] R. K. Sato and R. D. Loft.

Implementation of the NCAR CCM2 on the Connection Machine. In Hoffmann and Kauranne [HK93b], pages 371–393. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Siek:1999:SPM

[SL99] Jeremy G. Siek and Andrew Lumsdaine. Scientific programming: The matrix template library: Generic components for high-performance scientific computing. Computing in Science and Engineering, 1(6):70–78, November/December 1999. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1999/pdf/c6070.
                                        pdf;   http://www.computer.
                                        org/cse/cs1999/c6070abs.
                                        htm.

Shavit:1993:PPR

[SLB93] T. Shavit, L. Lee, and S. Breit. A practical parallel runtime environment on a multiprocessor with global address space. In Hoffmann and Kauranne [HK93b], pages 394–413. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Solovyev:1993:AFR

[SLML93] V. V. Solovyev, H. A. Lim, L. Milanesi, and C. Lawrence. Application of fractal representation of genetic texts for recognition of genome functional and coding regions. In Lim et al. [L^{+}93], pages 609– 622. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Su:1995:ACT

[SLRP95] E. Su, A. Lain, S. Ramaswamy, and D. J. Palermo. Advanced compilation techniques in the PARADIGM compiler for distributedmemory multicomputers. In ACM [ACM95a], pages 424– 433. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Sydow:1996:HPP

[SLS96] A. Sydow, T. Lux, and R.-P.

Schaefer. High-performance parallel computing for analyzing urban air pollution. Lecture Notes in Computer Science, 1030:417–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shen:1989:ESA

[SLY89] Zhiyu Shen, Zhiyuan Li, and Pen-Chung Yew. An empirical study on array subscripts and data dependences. Technical Report CSRD 840, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 16, 1989. 22 pp.

Shen:1990:ESF

[SLY90] Zhiyu Shen, Zhiyuan Li, and Pen-Chung Yew. An empirical study of Fortran programs for parallelizing compilers. Technical Report CSRD 983, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1990. 37 pp.

Sharma:1989:XTI

[SM89] Sanjay Sharma and Allen Davis Malony. Xcount: a tool to interpret counting related events. Technical Report CSRD 897, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1989. 7 pp.

Sherman:1992:GRW

[SM92] Arthur Sherman and Michael Mascagni. A gradient random walk method for two-dimensional reactiondiffusion equations. Technical report SRC-TR-92-059, Supercomputing Research Center: IDA, Lanham, MD, USA, March 2, 1992. 14 + [5] pp.

Srinivas:1994:CCC

[SM94] S. Srinivas and E. Malek. Compiler-directed cache coherence in multiprocessors: Classification, techniques and comparisons. In Balakrishnan [Bal94], pages 32–46. ISBN 0-07-462044-4. LCCN ????

Smarr:1993:IST

[Sma93] L. Smarr. Impact of supercomputing technology on research and development. In Thompson [Tho93c], pages 2572–2579. ISBN 0-86587-496-4. ISSN 0736-5721. LCCN ???? Two volumes.

Small:1995:MSN

[Sma95] C. H. Small. Multiprocessor “servers”: the new low-cost supercomputers. EDN, 40(6): 59–??, ???? 1995. CODEN EDNSBH. ISSN 0012-7515, 0364-6637.

Smarr:1990:GCC

[SMDL90] Larry Smarr, Gregory McRay, David Dixon, and Eric Lander. Grand challenges of computational science. Computer Graphics, 24(4):415– 417, August 1990. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Strohmaier:2015:TLP

[SMDS15] Erich Strohmaier, Hans W.

Meuer, Jack Dongarra, and Horst D. Simon. The TOP500 list and progress in high-performance computing. Computer, 48(11):42– 49, November 2015. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.
computer.org/csdl/mags/
co/2015/11/mco2015110042-
abs.html.

Schmidt:1985:ESN

[SMFG85] F. Schmidt, P. R. Mayer, G. Frey, and W. Giesser. Experiences in Solving the Neutron Diffusion Equation by the Finite Element Method on a Cray-1. ANS, La Grange Park, IL, USA, 1985. ISBN 0-89448-117-7. 12–21 pp. LCCN ????

Strout:1991:ECS

[SMH91] Robert E. Strout, II, James R.

McGraw, and Alan C. Hindmarsh. An examination of the conversion of software to multiprocessors. Journal of Parallel and Distributed Computing, 13(1):1– 16, September 1, 1991. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Smith:1981:AAH

[Smi81] B. J. Smith. Architecture and applications of the HEP multiprocessor system. In Tien F. Tao, editor, Real-time signal processing IV: August 25–28, 1981, San Diego, California, volume 298 of Proceedings of SPIE—the International Society for Optical Engineering, pages 241– 248. SPIE Optical Engineering Press, Bellingham, WA, USA, 1981. ISBN 0-89252-332-8. LCCN TK5102.5 .R43 1981.

Smith:1988:SR

[Smi88] Norris Parker Smith, editor. Supercomputing review. Supercomputing Review, San Diego, CA, USA, 1988. ISBN 1-871165-00-8. 252 pp. LCCN ????

Smitley:1989:DTH

[Smi89] David L. Smitley. Design tradeoffs for a high speed network node. Technical report SRC-TR-89-007, Supercomputing Research Center: IDA, Lanham, MD, USA, July 6, 1989. 21 pp.

Smith:1991:UMP

[Smi91] Mark Smith. Using massivelyparallel supercomputers to model stochastic spatial predator–prey systems. Ecological Modelling, 58(1–4): 347–367, November 1991. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0304380091900453.

Smitley:1992:PNU

[Smi92] David L. Smitley. Performance of networks using deflection routing. Technical report SRC-TR-92-082, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1992. 12 pp.

Smith:1993:MRC

[Smi93] K. Smith. Modern reactor core design codes and comparison to measured data [invited]. In Kusters et al. [KSW93], pages 479–495. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Smith:1995:ICS

[Smi95] Norris Parker Smith. Interfaces: Cray/SGI matchup adds drama as Kiwis win America’s Cup. IEEE Computational Science & Engineering, 2(2):72–73, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Smith:1996:DSC

[Smi96a] Norris Parker Smith. The death of Seymour Cray: a personal essay. Silicon Graphics World, 6(12):18, December 1996. ISSN 1057-7041.

Smith:1996:IHH

[Smi96b] Norris Parker Smith. Interfaces: Hardware for highperformance computing: Abstract progress, painful consolidation. IEEE Computational Science & Engineering, 3(2):104, 103, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Smith:1996:IRH

[Smi96c] Norris Parker Smith. Interfaces: Reciprocity in highperformance trade. IEEE Computational Science & Engineering, 3(3):88, 87, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). See letter from Cray Research President [Ewa96].

Smith:1996:SCP

[Smi96d] Norris Parker Smith. Seymour Cray: a personal essay. HP Chronicle, 14(1):14, December 1996. ISSN 0892-2829.

Smith:2001:CMM

[Smi01] Burton Smith. Cray MTA: Multithreading for latency response. Computer, 34 (4):69, April 2001. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r4059.
pdf;   http://www.computer.
org/computer/co2001/r4059abs.
htm.

Sporer:1988:IAS

[SMM88] M. Sporer, F. H. Moss, and C. J. Mathais. An introduction to the architecture of the Stellar Graphics supercomputer. In Digest of papers: intellectual leverage: Compcon spring 88, February 29–March 4, 1988, Thirty-Third IEEE Computer Society International Conference, Cathedral Hill Hotel, San Francisco, California., pages 464–

467. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0828-5 (paperback), 0-8186-4828-7 (microfiche), 0-8186-8828-9 (case). LCCN QA75.5 .C58 1988. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=
190.

Shchapov:2017:TPI

[SMM17] Vladislav A. Shchapov, Aleksei G. Masich, and Grigorii F. Masich. The technology of processing intensive structured dataflow on a supercomputer. The Journal of Systems and Software, 127(??):258–265, May 2017. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0164121216301054.

Sottile:2010:ICP

[SMR10] Matthew J. Sottile, Timothy G. Mattson, and Craig E. Rasmussen. Introduction to concurrency in programming languages. Chapman and Hall/CRC computational science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2010. ISBN 1-4200-7213-7 (hardcover). xii + 330 pp. LCCN QA76.7 .S62 2010.

Sterling:1995:ETP

[SMS95] Thomas Sterling, Paul Messina, and Paul H. Smith. Enabling Technologies for Petaflops Computing. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, 1995. ISBN 0-262-69176-0. x + 178 pp. LCCN QA76.885.S74 1995. US$30.00. URL http://www.

                                        mitpress.com/book-home.
                                        tcl?isbn=0262691760.

Shirley:1989:VVA

[SN89] Peter Shirley and Henry Neeman. Volume visualization at the center for supercomputing research and development. Technical Report CSRD 849, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1989. 14 pp.

Saletore:1995:MDP

[SN95a] V. A. Saletore and T. F.

Neff. Message-driven parallel computations on the MEIKO CS-2 parallel supercomputer. Lecture Notes in Computer Science, 919:241–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saletore:1995:MPC

[SN95b] V. A. Saletore and T. F.

Neff. Message-driven parallel computations on the MEIKO CS-2 parallel supercomputer. In Hertzberger and Serazzi [HS95b], pages 241–247. CO-DEN LNCSD9. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Shellard:1996:CHE

[SN96] R. Shellard and T. D.

Nguyen, editors. Computing in high energy physics: International conference — September 1995, Rio de Janeiro, Brazil, PROCEED-INGS of THE INTERNA-TIONAL CONFERENCE ON COMPUTING IN HIGH ENERGY PHYSICS 1995. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1996. ISBN 981-02-2783-3. LCCN ????

Snell:1994:ITS

[Sne94a] M. Snell. Industry trends: Supercomputing: same name, different game. Computer, 27(11):6–8, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Snell:1994:SSN

[Sne94b]  M. Snell.  Supercomput-

ing: same name, different game. Computer, 27(11): 6–8, November 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sanyal:2014:CBE

[SNEP14] Jibonananda Sanyal, Joshua New, Richard E. Edwards, and Lynne Parker. Calibrating building energy models using supercomputer trained machine learning agents. Concurrency and Computation: Practice and Experience, 26(13):2122–2133, September 10, 2014. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sasaki:1993:JVF

[SNK^{+}93] K. Sasaki, R. Nagai, N. Kikuzawa, M. Sawamura, and M. Takao. JAERIPULSE — a versatile FEL simulation code on a supercomputer. Nuclear instruments and methods in physics research. Section A, Accelerators, spectrometers, detectors and associated equipment, 331(1/3):450–??, July 1, 1993. CODEN NIMAER. ISSN 0168-9002, 0167-5087.

Sato:1995:UAB

[SNS95] H. Sato, T. Nanri, and M. Shimasaki. Using asynchronous and bulk communications to construct an optimizing compiler for distributed memory machines with consideration given to communication costs. In ACM [ACM95a], pages 185– 189. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Sakata:1997:PEW

[SNS^{+}97] Satoko Sakata, Umpei Nagashima, Mitsuhisa Sato, Satoshi Sekiguchi, and Haruo Hosoya. Performance evaluation of a workstation cluster, TMC CM-5, and Intel Paragon/XP using a parallel homology analysis program. Parallel Computing, 22(12): 1677–1693, February 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Snyder:1999:PGZ

[Sny99] Lawrence Snyder. A Programmer’s Guide to ZPL. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, 1999. ISBN 0-262-69217-1. viii + 154 pp. LCCN QA76.73.Z27S69 1999. US$30.00. URL http:

                                        //www.mitpress.com/book-
                                        home.tcl?isbn=0262692171.

Szauter:1991:MIH

[SO91] Imre F. Szauter and Fusun Ozguner. Multiprocessor implementation of a highspeed fault simulator on the CRAY Y-MP8/864 supercomputer. Thesis (M.S.), Dept. of Electrical Engineering, Ohio State University, Columbus, OH, USA, 1991. x + 88 pp. Advisor: F. Ozguner.

Suda:1995:ISH

[SO95] R. Suda and Y. Oyanagi. Implementation of sparta, a highly parallel circuit simulator by the preconditioned Jacobi method, on a distributed memory machine. In ACM [ACM95a], pages 209– 217. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Sobh:1992:IML

[Sob92] Nahil A. Sobh. Iterative methods for large scale static analysis of structures on a scalable multiprocessor supercomputer. NASA contractor report NASA CR-190369, Dept. of Mechanical Engineering and Mechanics, College of Engineering and Technology, Old Dominion University, Norfolk, VA, USA, 1992. ?? pp. Distributed to depository libraries in microfiche.

Shipping list no.: 92-2277-M. Microfiche. [Washington, DC?: National Aeronautics and Space Administration], 1992. 1 microfiche.

Sobh:1993:IML

[Sob93a] Nahil Atef Sobh. Iterative methods for large scale static analysis of structures on a scalable multiprocessor supercomputer. In Sincovec et al. [S^{+}93], pages 113–116. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Sobol:1993:EAM

[Sob93b] I. M. Sobol’. An efficient approach to multicriterial optimization. In Kusters et al. [KSW93], pages 27–32. ISBN 3-923704-11-9. LCCN ???? Two volumes.

SIAM:1994:PSW

[Soc94] Society for Industrial and Applied Mathematics, editor. Proceedings, Supercomputing ’94: Washington, DC, November 14-18, 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6606-4 (microfiche), 0-8186-6607-2 (case), 0-8186-6605-6 (paper). LCCN QA76.5 .S894 1994. IEEE catalog number 94CH34819.

Soerli:1994:CMM

[Soe94]  K. Soerli.  Chapter 8:

Mathematical modelling, numerical solution and visualization of steady threedimensional swirling fluid flow with turbulence and solid particles. In Murthy and Brebbia [MB94b], pages 169–184. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Solem:1984:MSM

[Sol84] Johndale C. Solem. MECA: a supercomputer for Monte Carlo. Technical Report LA-10005, Los Alamos National Laboratory, Los Alamos, NM, USA, April 1984. 5 pp.

Solbrig:1993:SPR

[Sol93] C. W. Solbrig. Simplified probabilistic risk assessment in fuel reprocessing. In Kusters et al. [KSW93], pages 316–327. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Soley:1994:TSO

[Sol94] C. Soley. From toaster to supercomputers: The object request broker. .EXE: the software developers’ magazine, 8 (10):44–??, ???? 1994. CO-DEN EXEEE5. ISSN 0268-6872.

Sommerville:2013:TCC

[Som13] Ian Sommerville. Teaching cloud computing: a software engineering perspective. The Journal of Systems and Software, 86(9): 2330–2332, September 2013. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0164121213000198.

Sun:1994:MID

[SP94] Y. B. Sun and G. A.

Parker. Mechatronic interface of a diaphragm-disc suspension valve. In Anonymous [Ano94-75], pages 131–138. ISBN 0-947719-68-7. LCCN ????

Smith:2012:OSH

[SP12] Melissa C. Smith and Gregory D. Peterson. Optimization of shared highperformance reconfigurable computing resources. ACM Transactions on Embedded Computing Systems, 11(2): 36:1–36:??, July 2012. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Anonymous:1987:STS

[Spe87] Special topics in supercomputing, 1987. Elsevier Science Publishers, Amsterdam, The Netherlands.

Sperling:1997:CIP

[Spe97] D. Sperling. Can industrygovernment partnerships produce the next generation vehicles? In Roller [Rol97], pages 11–20. ISBN 0-947719-88-1 (paperback). LCCN ????

Spector:2000:MBC

[Spe00] David Spector. Managing Beowulf clusters. Journal of Linux Technology, 1(1):18– ??, 2000. ISSN 1527-2761.

Stramaglia:1998:ISP

[SPGD98] S. Stramaglia, G. Pasquariello, L. Guerriero, and A. Distante. Interferometric SAR phase unwrapping by parallel tempering on a APE100/ quadrics supercomputer. Lecture Notes in Computer Science, 1401:898–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivas:1994:CAR

[SPK94] P. L. Srinivas, T. V. Prabhakar, and S. Kumar. COB-SQL — an aid for reengineering COBOL programs. In Balakrishnan [Bal94], pages 212–220. ISBN 0-07-462044-4. LCCN ????

Saldana:2010:MPM

[SPM^{+}10] Manuel Saldaña, Arun Patel, Christopher Madill, Daniel Nunes, Danyao Wang, Paul Chow, Ralph Wittig, Henry Styles, and Andrew Putnam. MPI as a programming model for high-performance reconfigurable computers. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 3(4):22:1– 22:??, November 2010. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Suits:2005:OMD

[SPP^{+}05] F. Suits, M. C. Pitman, J. W.

Pitera, W. C. Swope, and R. S. Germain. Overview of molecular dynamics techniques and early scientific results from the Blue Gene project. IBM Journal of Research and Development, 49 (2/3):475–487, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/492/suits.pdf.

Shiles:1990:PRS

[SPS90] G. Shiles, G. A. Pope, and K. Sepehrnoori. Petroleum reservoir simulation on the CRAY Y-MP. In Pitcher [Pit90], pages 497–507. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Shiles:1991:BUR

[SPS91] G. Shiles, G. A. Pope, and Kamy Sephernoori. Benchmark of UTCHEM reservoir simulator on various Cray computers. Communications in Applied Numerical Methods, 7(1):11–17, January 1991. CODEN CANMER. ISSN 0748-8025.

Sorli:1994:SHS

[SQM94] M. Sorli, G. Quaglia, and S. Mauro. Semi-active hydropneumatic suspension. In Anonymous [Ano94-75], pages 173–180. ISBN 0-947719-68-7. LCCN ????

Siddiqa:2019:CNE

[SQS^{+}19] Ayesha Siddiqa, Faisal Fayyaz Qureshi, Munam Ali Shah, Rahat Iqbal, Abdul Wahid, and Victor Chang. CCN: a novel energy efficient greedy routing protocol for green computing. Concurrency and Computation: Practice and Experience, 31(23):e4461:1– e4461:??, December 10, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sitaraman:1993:CBH

[SR93a] S. Sitaraman and F. Rahnema. Control blade history reactivity model for criticality calculations. In Kusters et al. [KSW93], pages 222– 231. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Szeto:1993:PAI

[SR93b] Andrew Y. J. Szeto and Rangaraj M. Rangayyan, editors. Proceedings of the Annual International Conference of the IEEE Engineering in Medicine and Biology Society: San Diego, California, USA, October 28–31, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1377-1, 0-7803-1378-X, 0-7803-1379-8, 0-7803-1411-5. LCCN R 856 A2 I344 1993. Three volumes: Part 1: Visualization, imaging, signal processing, modeling, neural networks; Part 2: Medical informatics, ethics, cardiology, instrumentation; Part 3: Biomechanics, rehabilitation, electrical phenomena, biomaterials.

Srinivasan:1994:CAL

[SR94] M. P. Srinivasan and T. Radhakrishnan. Computer assisted learning by the blind in rural India. In Balakrishnan [Bal94], pages 316–324. ISBN 0-07-462044-4. LCCN ????

Shoshani:2010:SDM

[SR10] Arie Shoshani and Doron Rotem, editors. Scientific data management: challenges, technology, and deployment. Chapman and Hall/CRC computational science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4200-6980-2. xxxii + 534 + 24 pp. LCCN Q183.9 .S33 2010.

Suzuki:1994:PCW

[SRBL94] L. R. C. Suzuki, R. Reid, T. J. Burns, and G. B. Lamont. Parallel computation of 3D wavelets. In IEEE [IEE94c], pages 454– 461. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Srivastava:1994:EIA

[Sri94] A. Srivastava. Evaluating intelligence of AI-based intelligent systems. In Balakrishnan [Bal94], pages 94–102. ISBN 0-07-462044-4. LCCN ????

Schwister:1990:EMS

[SS90a] B. Schwister and K. Solchenbach. A European made supercomputer. Future Generation Computer Systems, 5 (4):381–??, January 1, 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Schwister:1990:SEM

[SS90b] Bernd Schwister and Karl Solchenbach. SUPRENUM — a European made supercomputer. Future Generation Computer Systems, 5(4): 381–385, January 1, 1990. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Stevens:1990:CYU

[SS90c] K. G. Stevens, Jr. and Ron Sykora. Cray Y-MP: a user’s viewpoint. IEEE, Piscataway, NJ, USA, 1990. ISBN 0-8186-2028-5. 12–15 pp. LCCN ???? IEEE catalog number 90CH2843-1.

Sendyka:1994:AEI

[SS94] B. Sendyka and P. Sendyka.

Analysis of the effects of the introduction of fuel gap feeder into a constant depression carburettor. In Anonymous [Ano94-75], pages 311– 318. ISBN 0-947719-68-7. LCCN ????

Summers:1995:ASI

[SS95] Barbara G. Summers and David G. Staten. Adventures in supercomputing: An innovative program. Delta Kappa Gamma bulletin, 61 (3):43–??, Spring 1995. CO-DEN DKGBF7. ISSN 0011-8044.

Schadschneider:1996:CAT

[SS96a] A. Schadschneider and M. Schreckenberg. Cellular automata for traffic flow: Analytical results. In Wolf et al. [WSB96], pages 193–198. ISBN 981-02-2635-7. LCCN ????

Schneider:1996:GEE

[SS96b] R. Schneider and M. Schlenkrich.

Genecrunch and Europort, examples for hierarchical supercomputing at Silicon Graphics. In Anonymous [Ano96-43], pages 44–50. ISSN 1421-6337.

Skiles:1996:RMM

[SS96c] J. W. Skiles and C. H. Schulbach. Rehosting a minicomputer model on a supercomputer. Simulation, 66(1):

43–??, January 1, 1996. CO-DEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic).

Schwister:19xx:SEM

[SSxx] B. Schwister and K. Solchenbach. SUPRENUM — a European made supercomputer. Future Generation Computer Systems, 5(4):381–??, ???? 19xx. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Strohmaier:2007:AMP

[SS07] Erich Strohmaier and Hongzhang Shan. APEX-Map: a parameterized scalable memory access probe for highperformance computing systems. Concurrency and Computation: Practice and Experience, 19(17):2185–2205, December 10, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Sterling:2009:HPC

[SS09] Thomas Sterling and Dylan Stark. A high-performance computing forecast: Partly cloudy. Computing in Science and Engineering, 11 (4):42–49, July/August 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Smotherman:2010:ISP

[SS10] Mark Smotherman and Dag Spicer. IBM’s singleprocessor supercomputer efforts. Communications of the ACM, 53(12):28–30, December 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sterling:1999:HBB

[SSBS99] Thomas L. Sterling, John Salmon, Donald J. Becker, and Daniel F. Savarese. How to Build a Beowulf: a Guide to the Implementation and Application of PC Clusters. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, May 1999. ISBN 0-262-69218-X. xxi + 239 pp. LCCN QA76.58.H69 1999. US$30.00. URL http://www.

mitpress.com/book-home.
tcl?isbn=026269218X.

Saghi:1993:PPS

[SSG93] Gene Saghi, Howard Jay Siegel, and Jeffrey L. Gray. Predicting performance and selecting modes of parallelism: a case study using cyclic reduction on three parallel machines. Journal of Parallel and Distributed Computing, 19(3): 219–233, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1106/production;
http://www.idealibrary.

                                        com/links/doi/10.1006/jpdc.
                                        1993.1106/production/pdf.

Stunkel:1994:SHS

[SSGH94] C. B. Stunkel, D. G. Shea, D. G. Grice, and P. H. Hochschild. The SP1 highperformance switch. In IEEE [IEE94c], pages 150– 157. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Skipitaris:1996:EDF

[SSH96] D. Skipitaris, H. H. Simonsen, and M. Hermanrud. Experiences with DFS as a filesystem in the Norwegian MetaCenter for Supercomputing. In Liddell [Lid96], pages 992–993. ISBN 3-540-61142-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

Shelton:1994:FPS

[SSJL94] W. A. Shelton, G. M. Stocks, R. G. Jordan, and Y. Liu. First principles simulation of materials properties. In IEEE [IEE94c], pages 103– 110. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Shindyalov:1993:MJC

[SSKa93] I. N. Shindyalov, V. B. Streletc, N. A. Kolchanov, and L. and Milanesi. A method for joining contigs in automatic DNA sequencing.

In Lim et al. [L^{+}93], pages 413–418. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Segall:1997:SPD

[SSKR97] E. Segall, P. Steenkiste, N. Kumar, and A. Russell. Scalability of a portable distributed multiscale air quality model. In Delic and Wheeler [DW97], pages 81– 86. ISBN 0-89871-378-1. LCCN ????

Sikiotis:1990:FEB

[SSLR90] E. Sikiotis, V. Saouma, M. Long, and W. Rogger. Finite element based optimization of complex structures on a Cray X-MP supercomputer. Computers and Structures, 36(5):901–911, 1990. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Solovyev:1993:MSA

[SSM93] V. Solovyev, V. Streletc, and L. Milanesi. Multiple sequence alignment based on new approaches of tree construction and sequence comparison. In Lim et al. [L^{+}93], pages 419–428. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Stricker:1995:DSD

[SSOH95] T. Stricker, J. Stichnoth, D. O’Hallaron, and S. Hinrichs. Decoupling synchronization and data transfer in message passing systems of parallel computers. In ACM [ACM95a], pages 1–10. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Sarker:1993:SCB

[SSP93] Aroon Sarker, Meera Shenoy, and Devaprasad Purokayastha. Supercomputers: crashing big brothers’ party. Business India, 394:90–??, April 12, 1993. ISSN 0254-5268.

Sinvhal-Sharma:1991:PBS

[SSRL91] Priyamvada Sinvhal-Sharma, Lawrence Rauchwerger, and John Leonard Larson. Perfect BenchmarksTM: instrumented version. Technical Report CSRD 1152, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1991. 11 pp.

Sinvhal-Sharma:1990:CTB

[SSS90] Priyamvada Sinvhal-Sharma and Sanjay Sharma. CPROF: a trace based profiler for shared memory multiprocessor systems. Technical Report CSRD 1016, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1990. 15 pp.

Sander:1992:HGH

[SSS92] C. Sander, R. Schneider, and P. Stouten. Human genome and high performance computing in molecular biology. In Meuer [Meu92c], pages 32–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Schnupp:1994:XWW

[SSS94] P. Schnupp, H. Schawacht, and C. Schwieger. xQard — why we need an object oriented hypertext database for quality management. In Anonymous [Ano94-75], pages 59–66. ISBN 0-947719-68-7. LCCN ????

Salem:2020:SDS

[SSSR20] Farouk Salem, Florian Schintke, Thorsten Schütt, and Alexander Reinefeld. Scheduling data streams for low latency and high throughput on a Cray XC40 using Libfabric. Concurrency and Computation: Practice and Experience, 32(20):e5563:1– e5563:??, October 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Svede-Shvets:1996:OMO

[SSSSE96] V. N. Svede-Shvets, V. V.

Svede-Shvets, and L. C. Eisymont. Optoelectronic massparallel OPTOCOM supercomputer [2969-56]. In Alferov et al. [AGP96], pages 108–111. ISBN 0-8194-2375-0. ISSN 0361-0748. LCCN ????

Shirley:1990:PAD

[ST90] Peter Shirley and Allan Tuchman. A polygonal approximation to direct scalar volume rendering. Technical Report CSRD 1006, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1990. 8 pp.

Smith:1992:CMP

[ST92] J. E. Smith and W. R. Taylor. Characterizing memory performance in vector multiprocessors. In ACM [ACM92b], pages 35–44. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Scott:1994:ORC

[ST94] S. Scott and G. Thorson. Optimized routing in the Cray T3D. Lecture Notes in Computer Science, 853:281–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stanger:1988:NSP

[Sta88] John Stanger. A numerical simulation of the proposed European side impact test procedure. In International Conference on Supercomputing Applications in the Automotive Industry (2nd: 1988: Seville, Spain). Supercomputer applications in automotive research and engineering development, pages 9–20. Cray Research, Inc., Minneapolis, MN, USA, 1988.

Stadtherr:1994:SSS

[Sta94]  M. A. Stadtherr.  Su-

percomputing strategies for simulation and control of chemical process operations. In Anonymous [Ano94-107], pages 215–216. ISBN 0-87263-441-8. LCCN TS176.N72 1994.

Stalzer:1995:PFM

[Sta95a] Mark A. Stalzer. A parallel fast multipole method for the Helmholtz equation. Parallel Processing Letters, 5(2):263–274, June 1995. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Stalzer:1995:PFMa

[Sta95b] Mark A. Stalzer. A parallel fast multipole method for the Helmholtz equation. Parallel Processing Letters, 5(2):263–274, June 1995. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

SterlingHobe:1985:STS

[Ste85] Sterling Hobe Corporation.

Strategies and tactics of the supercomputer industry: Japan, EEC and selected other nations. Sterling Hobe, Washington, DC, USA, 1985. various pp.

Stephenson:1990:SCR

[Ste90] Frank Stephenson. Supercomputer Computations Research Institute: a national resource in supercomputing research. Technical report, Florida State University, Tallahassee, FL, USA, 1990. 24 pp.

Steele:1992:OCM

[Ste92] G. L. Steele, Jr. Overview of the connection machine model CM-5. In Meuer [Meu92c], pages 87–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Steele:1994:ACP

[Ste94a] C. S. Steele. Affinity: a concurrent programming environment. In IEEE [IEE94c], pages 365–372. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Steinmetz:1994:FGC

[Ste94b] M. Steinmetz. The formation of galaxies: a challenge for supercomputers — a simple task for GRAPE? Lecture Notes in Computer Science, 796:358–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stephens:1994:PBT

[Ste94c] R. Stephens. Parallel benchmarks on the transtech paramid supercomputer. In De Gloria et al. [DJM94], pages 136–146. ISBN 90-5199-177-0, 4-274-90004-5. ISSN 0925-4986. LCCN ????

Stephens:1994:MCC

[Ste94d] Sally Stephens. Modeling a comet’s collision. Mercury, 23(2):6–??, March 1, 1994. ISSN 0047-6773.

Stevens:1994:MSU

[Ste94e] R. Stevens. Multimedia supercomputing: The use of supercomputers to drive highperformance multimedia systems and virtual environments. In IEEE [IEE94d], pages 3–?? ISBN 0-8186-6395-2, 0-8186-6396-0. LCCN QA76.9.D5I328 1994.

Stevens:1994:HPC

[Ste94f] Rick Stevens. High-performance computing and communications. Future Generation Computer Systems, 10 (2–3):159–167, June 1994. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Steinbuechel:1995:TSS

[Ste95] N. V. Steinbuechel. Temporal system states in speech processing. In Herrmann et al. [HWP95], pages 75–82. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Steenkiste:1996:NBM

[Ste96] Peter Steenkiste. Networkbased multicomputers: a practical supercomputer architecture. IEEE Transactions on Parallel and Distributed Systems, 7(8):861– 875, August 1996. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//www.computer.org/tpds/
td1996/l0861abs.htm.

Sterling:2000:SCB

[Ste00] Thomas Sterling. Symbolic computing with Beowulfclass PC clusters. Lecture Notes in Computer Science, 1908:7–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1908/19080007.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1908/19080007.
pdf.

Sterling:2001:BCCc

[Ste01a]  Thomas Sterling.  Be-

owulf cluster computing at the Third Plateau. ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-08/pdfs/sterling.pdf.

Sterling:2001:BCCa

[Ste01b] Thomas L. Sterling, editor.

Beowulf Cluster Computing with Linux. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-69274-0. xxxiii + 496 pp. LCCN QA76.58 .B46 2002. US$42.95, UK£28.50.

Sterling:2001:BCCb

[Ste01c] Thomas L. Sterling, editor.

Beowulf Cluster Computing with Windows. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-69275-9. xxxiii + 445 pp. LCCN QA76.58 .B463 2002. US$42.95, UK£28.50.

Sterling:2002:BCC

[Ste02] Thomas Lawrence Sterling.

Beowulf cluster computing with Linux. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-69274-0. xxxiii + 496 pp. LCCN QA76.58 .B46 2002.

Sato:1998:NPL

[STH^{+}98] Mitsuhisa Sato, Hiroshi Tezuka, Atsushi Hori, Yutaka Ishikawa, Satoshi Sekiguchi, Hidemoto Nakada, Satoshi Matsuoka, and Umpei Nagashima. Ninf and PM-Communication libraries for global computing and highperformance cluster computing. Future Generation Computer Systems, 13(4–5): 349–359, March 11, 1998. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

www.elsevier.com/gej-ng/
10/19/19/28/19/27/abstract.
html.

SARA:1984:S

[Sti84] Supercomputer, 1984. ISSN 0168-7875. Amsterdam Universities Computing Centre (SARA), Amsterdam, Netherlands.

Stiff:1998:APS

[Sti98a] M. J. Stiff. (Astro)Physical supercomputing: Ada95 as a safe, object oriented alternative. Lecture Notes in Computer Science, 1411: 128–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stift:1998:APS

[Sti98b]  Martin J. Stift.  (As-

tro)Physical supercomputing: Ada95 as a safe, object oriented alternative. Lecture Notes in Computer Science, 1411:128–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1411/14110128.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1411/14110128.
pdf.

Shavlik:1993:UKN

[STN93] J. W. Shavlik, G. G. Towell, and M. O. Noordewier. Using knowledge-based neural networks to refine existing biological theories. In Lim et al. [L^{+}93], pages 377–390. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Stollenwerk:1995:SCN

[Sto95]  N. Stollenwerk.  Self-

controlling chaos in neuromodules. In Herrmann et al. [HWP95], pages 421–426. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Strok:1994:NJI

[Str94]  Dale C. Strok.  In the

news: Jupiter impacts: Resolution makes a big difference. supercomputer farming down under. HPF Forum welcomes comments. Smithsonian Awards honor computational scientists. low-life computer viruses. PVM developers get R&D-100 award. the eyes have it. neural nets detect breast cancer. better cars through cooperation. parallel version of global climate model. Lockheed to run Idaho National Engineering Lab. public-private partners: new drugs, new software. IEEE Computational Science & Engineering, 1(3):88–90, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Strenski:1997:ADC

[Str97] D. G. Strenski. Algorithm design for consistency of numerical results and optimal performance for the Cray T3D. In Delic and Wheeler [DW97], pages 365– 375. ISBN 0-89871-378-1. LCCN ????

Strohmaier:2003:WWH

[Str03] Erich Strohmaier. “Who’s who” in high performance computing: TOP500 celebrates 10th anniversary. Scientific Computing & Instrumentation, 20(9):28–30, 32, 48, August 2003. ISSN 1524-2560.

Strawn:2010:HPC

[Str10]  Roger Strawn.  High-

performance computing for rotorcraft modeling and simulation. Computing in Science and Engineering, 12(5): 27–35, September/October 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Strickland:2011:SPS

[Str11] E. Strickland. Supercomputers predict a stormy hurricane season. IEEE Spectrum, 48(7):11–12, July 2011. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strickland:2012:GBW

[Str12] E. Strickland. Good-bye, wheelchair. IEEE Spectrum, 49(1):30–32, January 2012. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sekiguchi:1995:HTS

[STSK95] S. Sekiguchi, H. Tsuzuki, F. Sakamoto, and K. Kobayashi. Hardware technology for supercomputer SX-4 series. NEC Technical Journal = NEC giho, 48(11):23–??, ???? 1995. CODEN NECGEZ. ISSN 0285-4139.

Stueben:1995:CTP

[Stu95] H. Stueben. The Cray T3D as a production machine at Konrad-Zuse-Zentrum Berlin. Lecture Notes in Computer Science, 919:347–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stueben:1997:HS

[Stu97] K. Stueben. HPC-Enabled simulation. In Roller [Rol97], pages 207–214. ISBN 0-947719-88-1 (paperback). LCCN ????

Studt:2003:IFN

[Stu03]  Tim Studt.  Ian Foster

named 2003 Innovator of the Year. Research & Development, 45(10):32–33, October 2003. CODEN REDEEA. ISSN 0746-9179. URL http:

//www.globus.org/.

Su:1992:MSD

[Su92] Hong-Men Su. On multiprocessor synchronization and data transfer. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. x + 188 pp.

Subramanian:1994:ISE

[Sub94] K. Subramanian. Information security education — the need for a specialised curriculum. In Balakrishnan [Bal94], pages 385–?? ISBN 0-07-462044-4. LCCN ????

Sugarman:1980:CSC

[Sug80a] R. Sugarman. Computers: Superpower computers: The omnipresent microprocessor can hardly supplant the large high-speed computer in modeling complex systems and phenomena. IEEE Spectrum, 17(4):28–34, April 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1980:TOS

[Sug80b] R. Sugarman. The technologist’s own ’supercomputer’. IEEE Spectrum, 17(9):49– 52, September 1980. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugla:1994:PAP

[Sug94] B. Sugla. Parallel application programming. In Balakrishnan [Bal94], pages 174–187. ISBN 0-07-462044-4. LCCN ????

Sugiyama:1996:DMC

[Sug96] Y. Sugiyama. Dynamical model for congestion of freeway traffic and its structural stability. In Wolf et al. [WSB96], pages 137– 150. ISBN 981-02-2635-7. LCCN ????

Suhir:1997:EPP

[Suh97]  E. Suhir, editor.  Elec-

tronic and photonic packaging: Pacific Rim/ASME international and intersociety conference — 1997 Jun, volume 19(1) of ASME - PUBLICATIONS- EEP 1997. ASME, ????, 1997. ISBN 0-7918-1559-5. LCCN ????

Sullivan:1991:VPI

[Sul91] Steve Sullivan. Vector and parallel implementations of the wavelet transform. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1991. viii + 64 pp.

Sullivan:1997:AEC

[Sul97] Francis Sullivan. From the Associate Editor-in-Chief: Will the circle be unbroken? (A meditation on the passing of Seymour Cray); editorial board changes. IEEE Computational Science & Engineering, 4(1):1, 4, 6, January/March 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs1997/pdf/c1001.
                                        pdf.

Sumner:1982:SST

[Sum82] F. Sumner. Supercomputer systems technology, volume 10(6) of State of the art report; ser. 10, no. 6. Pergamon Infotech, Maidenhead, Berkshire, England, 1982. ISBN 0-08-028569-4. iii + 455 pp. LCCN TK 7888.3 S84 1982.

Sundaram:1994:HIS

[Sun94] V. M. Sundaram. Harnessing the informal sector for information technology education. In Balakrishnan [Bal94], pages 341–350. ISBN 0-07-462044-4. LCCN ????

Anonymous:1987:S

[Sup87a] Supercomputing, 1987. Super-Computing Magazine, Inc, Sunnyvale, CA, USA.

SRC:1987:JS

[Sup87b] The Journal of supercomputing, 1987. ISSN 0920-8542 (print), 1573-0484 (electronic). Kluwer Academic Publishers, Dordrecht, The Netherlands.

Anonymous:1988:SN

[Sup88a] Supercomputing news, 1988. ISSN 0898-1426. Publications and Communications, Austin, TX, USA.

Anonymous:1988:SR

[Sup88b] Supercomputing review, 1988.

ISSN 1048-6836. London Manhattan Group of Companies, San Diego, CA, USA.

YIS:19xx:SPP

[Supxxa] Supercomputing and parallel processing today, 19xx. Yellowstone Information Services, Elkview, WY, USA.

SSI:19xx:AR

[Supxxb] Annual report, 19xx. Supercomputing Solutions, San Diego, CA, USA.

Suslov:1993:STE

[Sus93] I. R. Suslov. Solution of the transport equation in 2and 3-dimensional irregular geometry by the method of characteristics. In Kusters et al. [KSW93], pages 752– 763. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Suzuki:1989:SCI

[Suz89] Hiroshi Suzuki. A serial communication interface for a parallel simulation system. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1989. ix + 67 pp.

Sips:1996:ALE

[SVD96] H. J. Sips, K. Van Reeuwijk, and W. Denissen. Analysis of local enumeration and storage schemes in HPF. In ACM [ACM96], pages 10–17. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Saoudi:1995:MHR

[SVML95] A. Saoudi, A. Valda Ochoa, R. Mastrippolito, and C. Loc’h. A miniaturized high resolution gamma-camera to image in vivo dopamine D2 receptors in rat brain. In Herrmann et al. [HWP95], pages 99–104. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Svolopoulos:1993:NSM

[Svo93] Peter A. Svolopoulos. A numerical search for multistep methods using the Cray X-MP supercomputer. Technical report, ????, ????, 1993. vi + 61 pp.

Saad:1988:BPS

[SW88] Youcef Saad and Harry A. G. Wijshoff. Benchmark package for sparse matrix computations. Technical Report CSRD 787, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 10 pp.

Saunders:1991:SPC

[SW91] V. R. Saunders and S. Wilson. ‘Scavenger’ programming for the Cray X-MP computer. Parallel Computing, 17(9):1025–1034, November 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Smith:1994:PAT

[SW94] James E. Smith and Shlomo Weiss. PowerPC 601 and Alpha 21064: a tale of two RISCs. Computer, 27(6): 46–58, June 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaefer:1996:THG

[SW96] J. Schaefer and D. E. Wolf.

Transients in homogeneous granular pipe flow. In Wolf et al. [WSB96], pages 311– 316. ISBN 981-02-2635-7. LCCN ????

Sorgatz:1999:THP

[SW99] Andreas Sorgatz and Stefan Wehmeier. Towards highperformance symbolic computing: using MuPAD as a problem solving environment. Mathematics and Computers in Simulation, 49(3):235– 246, August 1999. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). Special issue on high performance symbolic computing.

Sartor:2010:MRE

[SW10a] Dale Sartor and Mark Wilson. Money for research, not energy bills: Finding energy and cost savings in high-performance computer facility designs. Computing in Science and Engineering, 12(6):11–19, November/ December 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Spiers:2010:HPC

[SW10b] Brad Spiers and Denis Wallez. High-performance computing on Wall Street. Computer, 43(12):53–59, December 2010. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Swarztrauber:1986:MF

[Swa86] Paul N. Swarztrauber. Multiprocessor FFTs. Technical Report CSRD-608, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 28 pp.

Sweatman:1994:DPB

[Swe94] W. L. Sweatman. The development of a parallel Nbody code for the Edinburgh Concurrent Supercomputer. Journal of Computational Physics, 111(1): 110–119, March 1, 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0021999184710485.

Salapura:2006:EWP

[SWG06] Valentina Salapura, Robert Walkup, and Alan Gara. Exploiting workload parallelism for performance and power optimization in Blue Gene. IEEE Micro, 26(5):67– 81, September/October 2006. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sitsky:1995:IPM

[SWJ95] D. Sitsky, D. Walsh, and C. Johnson. Implementation and performance of the MPI message passing interface on the Fujitsu AP1000 multicomputer. Australian Computer Science Communications, 17(1):475–481, ???? 1995. CODEN ACSCDD. ISSN 0157-3055.

Simmons:1991:PCT

[SWL^{+}91] Margaret L. Simmons, Harvey J. Wasserman, Olaf M. Lubeck, Christopher Eoyang, Raul Mendez, Hiroo Harada, and Misako Ishiguro. A performance comparison of three supercomputers: Fujitsu VP-2600, NEC SX-3, and Cray Y-MP. IEEE, Piscataway, NJ, USA, 1991. ISBN 0-8186-2158-3. 150–157 pp. LCCN ???? IEEE catalog number 91CH3058-5.

Simmons:1992:PCF

[SWL^{+}92] Margaret L. Simmons, Harvey J. Wasserman, Olaf M. Lubeck, Christopher Eoyang, Raul Mendez, Hiroo Harada, and Misako Ishiguro. A performance comparison of four supercomputers. Communications of the ACM, 35 (8):116–124, August 1992. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
135234.html.

Shan:2012:PEH

[SWS^{+}12] Hongzhang Shan, Nicholas J. Wright, John Shalf, Katherine Yelick, Marcus Wagner, and Nathan Wichmann. A preliminary evaluation of the hardware acceleration of the Cray Gemini interconnect for PGAS languages and comparison with MPI. ACM SIGMETRICS Performance Evaluation Review, 40(2):92– 98, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Segall:1997:EMC

[SWSR97] E. Segall, K. Walker, P. Steenkiste, and A. Russell. Environmental modeling on the C90-T3D heterogeneous system. In Delic and Wheeler [DW97], pages 87–94. ISBN 0-89871-378-1. LCCN ????

Su:1991:EDE

[SY91] Hong-Men Su and Pen-Chung Yew. Efficient doacross execution on distributed shared-memory multiprocessors. Technical Report CSRD 1072, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1991. 12 pp.

Sarukkai:1994:NPI

[SYG94] S. R. Sarukkai, J. Yan, and J. K. Gotwals. Normalized performance indices for message passing parallel programs. In Anonymous [Ano94-134], pages 323–332. ISBN ???? LCCN ????

Stredney:1992:SAB

[SYMT92] D. Stredney, R. Yagel, S. F.

May, and M. Torello. Supercomputer assisted brain visualization with an extended ray tracer. In Gelberg and Levkowitz [GL92], pages 33– 38. ISBN 0-89791-527-5 (soft cover), 0-89791-528-3 (hard cover). LCCN T385 .W75 1992.

Sameh:1989:SGS

[SZ89] Ahmed Sameh and Zahari Zlatev. Solving general sparse linear systems using conjugate gradient-type methods. Technical Report CSRD 895, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, July 1989. 11 pp.

Simunovic:1996:SAW

[SZ96] S. Simunovic and T. Zacharia. Supercomputing applications in welding simulations. In Smartt et al. [SJD96], pages 19–24. ISBN 0-87170-567-2. LCCN ????

Stankova:1998:NSM

[SZ98] E. N. Stankova and E. V. Zudilova. Numerical simulation by means of supercomputers. Lecture Notes in Computer Science, 1401: 901–??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siegel:2011:FFE

[SZ11] Stephen F. Siegel and Timothy K. Zirkel. FEVS: a Functional Equivalence Verification Suite for highperformance scientific computing. Mathematics in Computer Science, 5(4): 427–435, December 2011. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=1661-8270&volume=5&
                                        issue=4&spage=427.

Schulz-Ziemer:1995:HIP

[SZG95] G. Schulz-Ziemer and A. Geiger. HPF on Intel Paragon and CRAFT on CRAY T3D: basic performance measurements and experiments with a block-sparse CG-algorithm.

Lecture Notes in Computer Science, 919:618–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomko:1994:DPR

[TA94] K. A. Tomko and S. G. Abraham. Data and program restructuring of irregular applications for cache-coherent multiprocessors. In Anonymous [Ano94-134], pages 214–225. ISBN ???? LCCN ????

Tseng:1995:UCT

[TAAL95] C.-W. Tseng, J. M. Anderson, S. P. Amarasinghe, and M. S. Lam. Unified compilation techniques for shared and distributed address space machines. In ACM [ACM95a], pages 67– 76. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Taflove:1996:RES

[Taf96] A. Taflove. Re-inventing electromagnetics: Supercomputing solution of Maxwell’s equations via direct time integration of space grids. In Anonymous [Ano96w], pages 55–70.

Takagi:1993:OOO

[Tak93] T. Takagi. ODS: Overlapping oligonucleotide database with deductive engine for signal sequence search. In Lim et al. [L^{+}93], pages 263–272.

ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Taklanti:1994:CNS

[Tak94] A. Taklanti. Chapter 14: Numerical simulation of internal flows in the automotive industry. In Murthy and Brebbia [MB94b], pages 331–351. ISBN 1-85312-076-6 (Computational Mechanics, Southampton), 1-85166-759-8 (Elsevier Applied Science), 0-945824-59-9 (Computational Mechanics, Boston). LCCN QA911.S88 1993.

Taufer:2006:PPS

[TAKB06] Michela Taufer, Chahm An, Andreas Kerstens, and Charles L. Brooks, III.

Predictor@Home: a “Protein

Structure Prediction Supercomputer” based on global computing. IEEE Transactions on Parallel and Distributed Systems, 17(8):786– 796, August 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Treleaven:1991:VC

[TAM^{+}91] P. Treleaven, H. Aiso, M. A. Meth, A. K. Jones, R. H. Sprague, Jr., T. Van Duzer, G. E. Moore, A. C. Weaver, V. Cerf, M. R. Mercer, K. Boykin, D. Cheriton, T. G. Lewis, Luqi, and L. P. Deutsch. Viewpoints (computing in the 1990s). Computer, 24(9):98– 113, September 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tang:1987:DPP

[Tan87] Peiyi Tang. Deadlock prevention in processor selfscheduling for parallel nested loops. Technical Report CSRD-626, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 31 pp.

Tang:1989:PEV

[Tan89a] Ju-Ho Tang. Performance evaluation of vector machine architectures. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1989. ix + 152 pp.

Tang:1989:SDS

[Tan89b] Peiyi Tang. Self-scheduling, data synchronization and program transformation for multiprocessor systems. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1989. vi + 221 pp.

Tan:1995:DBC

[Tan95] C. J. Tan. Deep Blue: Computer chess and massively parallel systems. In ACM [ACM95a], pages 237– 239. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Taylor:1994:SRS

[Tay94] Peter Taylor. Site report: The San Diego Supercomputer Center. IEEE Computational Science & Engineering, 1(3):10–14, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Taylor:1995:SWM

[Tay95a] D. Taylor. Focus: Stormy weather main suspect in dinosaur disappearance. IEEE parallel and distributed technology: systems and applications, 3(4):12–13, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL

                                        http://dlib.computer.org/
                                        pd/books/pd1995/pdf/h40012.
                                        pdf.

Taylor:1995:GMT

[Tay95b] J. G. Taylor. Global modelling and testing of models of brain and mind. In Herrmann et al. [HWP95], pages 11–26. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Tuchman:1989:MVD

[TB89] Allan Tuchman and Michael Waitsel Berry. Matrix visualization in the design of numerical algorithms. Technical Report CSRD 826, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1989. 14 pp.

Thakur:1994:CRS

[TBC94] R. Thakur, R. Bordawekar, and A. Choudhary. Compiler and runtime support for out-of-core HPF programs. In Anonymous [Ano94-134], pages 382–391. ISBN ???? LCCN ????

Turcotte:1993:DDI

[TC93] Louis H. Turcotte and Bradley M. Comes. Delivering data interpretation: From GFLOPS to insight. Computers and Graphics, 17(1):23–30, January– February 02, 1993. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Thompson:1994:RPQ

[TC94] Donald O. Thompson and Dale E. Chimenti, editors. Review of progress in quantitative nondestructive evaluation: proceedings of the Thirteenth Symposium on Quantitative Nondestructive Evaluation, held August 1– 6, 1993, at Bowdoin College, Brunswick, Maine, volume 13. Plenum Press, New York, NY, USA; London, UK, 1994. ISBN 0-306-44731-2. ISSN 0743-0760. LCCN ???? Two volumes.

Thomas:2021:ISJ

[TC21] R. Thomas and S. Cholia. Interactive supercomputing with Jupyter. Computing in Science and Engineering, 23 (2):93–98, March/April 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Thakur:1994:RAR

[TCF94] R. Thakur, A. Choudhary, and G. Fox. Runtime array redistribution in HPF programs. In IEEE [IEE94c], pages 309–316. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Taylor:1993:MEI

[TCJS93] H. H. Taylor, D. Chin, A. W. Jessup, and D. Sarnoff. An MPEG encoder implementation on the Princeton Engine Video Supercomputer. In Storer and Cohn [SC93], pages 420–430. ISBN 0-8186-3392-1, 0-8186-3391-3. ISSN 1068-0314. LCCN QA76.9.D33 D38 1993. IEEE Computer Society Press order number 3392-02. IEEE catalog number 93TH0536-3.

Taraglio:1995:EIB

[TCM95] S. Taraglio, C. R. Casaccia, and F. Massaioli. An efficient implementation of a backpropagation learning algorithm on a quadrics parallel supercomputer. In Hertzberger and Serazzi [HS95b], pages 664–671. CO-DEN LNCSD9. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Tilakasiri:1990:SAS

[TD90] A. Tilakasiri and P. Du Bois.

A simulation of an automobile side impact with finite element dummies included in the analysis. In Pitcher [Pit90], pages 125– 131. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Tomko:1996:PDW

[TD96] K. A. Tomko and E. S.

Davidson. Profile driven weighted decomposition. In ACM [ACM96], pages 165– 172. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Tang:2013:JSA

[TDBL13] Wei Tang, Narayan Desai, Daniel Buettner, and Zhiling Lan. Job scheduling with adjusted runtime estimates on production supercomputers. Journal of Parallel and Distributed Computing, 73(7):926–938, July 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731513000312.

teRiele:1987:AAV

[tDv87] H. J. J. te Riele, T. J. Dekker, and H. A. van der Vorst, editors. Algorithms and applications on vector and parallel computers: Colloquium on Numerical Aspects of Vector and Parallel Processors (1985–1986, Amsterdam, The Netherlands), number 3 in Special topics in supercomputing. Elsevier Science Publishers, Amsterdam, The Netherlands, 1987. ISBN 0-444-70322-5 (U.S.). LCCN QA76.6 .A45851 1987.

SSRTCSST-USHR:1989:USI

[Tec89] Subcommittee on Science, Research, and Space Technology of the Committee on Science and U. S. House of Representatives Technology. U.S. supercomputer industry hearing before the Subcommittee on Science, Research, and Technology of the Committee on Science, Space, and Technology, U.S. House of Representatives, One Hundred First Congress, first session, June 20, 1989. Technical Report 45, United States Government Printing Office, Washington, DC, USA, 1989. iii + 140 pp.

Temperton:1983:CVC

[Tem83] Clive Temperton. Cray 1 v. Cyber 205: some comparisons. ACM SIGNUM Newsletter, 18(2):20–21, April 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Temperton:1988:IPF

[Tem88] Clive Temperton. Implementation of a prime factor FFT algorithm on CRAY-1. Parallel Computing, 6(1): 99–108, January 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Temperton:1989:FMC

[Tem89a] C. Temperton. Further measurements of (r_{∞}, n_{1/2}) on the CRAY-1 and CRAY X-MP. Parallel Computing, 11 (1):107–111, July 1989. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Temperton:1989:FMR

[Tem89b]  Clive Temperton.  Fur-

ther measurements of (r$- $INF$/, n$-$HLF$/) on the Cray-1 and Cray X-MP. Parallel Computing, 11(1): 107–111, July 1989. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tawbi:1992:PAL

[TF92] N. Tawbi and P. Feautrier. Processor allocation and loop scheduling on multiprocessor computers. In ACM [ACM92b], pages 63–71. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Tikekar:1994:PMS

[TF94] R. V. Tikekar and F. Fotouhi. PROGS: a modeling scheme for efficient querying in a multimedia information system. In Balakrishnan [Bal94], pages 374–?? ISBN 0-07-462044-4. LCCN ????

Tolmie:1995:HIJ

[TF95] Don Tolmie and Don Flanagan. HIPPI: It’s not just for supercomputers anymore. Data communications, 24(6): 107–??, May 1995. CODEN DACODM. ISSN 0363-6399.

Tonshoff:1997:HCU

[TF97] H. K. Tonshoff and Y. Falkenberg. High-speed-grinding of crankshafts using CBN-comb tools. In Roller [Rol97], pages 297–304. ISBN 0-947719-88-1 (paperback). LCCN ????

Thackston:2015:PLC

[TF15] Russell Thackston and Ryan C. Fortenberry. The performance of low-cost commercial cloud computing as an alternative in computational chemistry. Journal of Computational Chemistry, 36 (12):926–933, May 5, 2015. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Toomarian:1994:TPS

[TFB94a] N. Toomarian, A. Fijany, and J. Barhen. Time-parallel solution of linear partial differential equations on the Intel Touchstone Delta supercomputer. Concurrency: practice and experience, 6(8):641– 652, December 1994. CO-DEN CPEXEI. ISSN 1040-3108.

Toomarian:1994:TSL

[TFB94b] N. Toomarian, A. Fijany, and J. Barhen. Time-parallel solution of linear partial differential equations on the Intel Touchstone Delta supercomputer. Concurrency: practice and experience, 6(8):641– 652, ???? 1994. CODEN CPEXEI. ISSN 1040-3108.

Tokioka:19xx:TCE

[TfGERJxx] Tatsushi Tokioka and Center for Global Environmental Research (Japan). A transient CO2 experiment with the MRI CGCM: annual mean response. ???? ????, Center for Global Environmental Research, National Institute for Environmental Studies, Environment Agency of Japan, Tsukuba, Japan, 19xx. v + 86 pp.

Tamura:1994:PCV

[TFO94] Y. Tamura, K. Fujii, and T. Ogawa. Parallel computations and visualizations of flows around high-speed trains. In Anonymous [Ano94-75], pages 771–778. ISBN 0-947719-68-7. LCCN ????

Traenkle:1994:SME

[TFVK94] F. Traenkle, M. I. Frank, M. K. Vernon, and S. Kim. Solving microstructure electrostatics with MIMD parallel supercomputers and split-C. Journal of non-Newtonian fluid mechanics, 53(1):197– ??, ???? 1994. CODEN JNFMDI. ISSN 0377-0257.

Tomilinson:1994:MGA

[TG94] A. I. Tomilinson and V. K. Garg. Maintaining global assertions on distributed systems. In Balakrishnan [Bal94], pages 257–272. ISBN 0-07-462044-4. LCCN ????

Tang:1995:VBD

[TG95] P. Tang and N. Gao. Vectorization beyond data dependences. In ACM [ACM95a], pages 434–443. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Thakur:1996:EEP

[TGL96] R. Thakur, W. Gropp, and E. Lusk. An experimental evaluation of the parallel I/O systems of the IBM SP and Intel Paragon using a production application. Lecture Notes in Computer Science, 1127:24–??, 1996. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tang:2008:EET

[TGV08] Qinghui Tang, Sandeep Kumar S. Gupta, and Georgios Varsamopoulos. Energyefficient thermal-aware task scheduling for homogeneous high-performance computing data centers: a cyberphysical approach. IEEE Transactions on Parallel and Distributed Systems, 19(11): 1458–1472, November 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Tiwari:1994:DMA

[TH94] R. Tiwari and T. L. Huntsberger. A distributed memory algorithm for volume rendering. In IEEE [IEE94c], pages 247–251. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Thota:2019:SIC

[TH19] Abhinav Thota and Yun He.

Special issue of the Cray User Group (CUG 2018). Concurrency and Computation: Practice and Experience, 31 (16):e5117:1–e5117:??, August 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hagen:1990:DGW

[tHd90] P. J. W. ten Hagen, I. Herman, and J. R. G. de Vries. A dataflow graphics workstation. Computers and Graphics, 14(1):83–93, 1990. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Anonymous:1990:SRR

[The90a] The Spang Robinson report on high performance computing, 1990. ISSN 1053-1661. John Wiley and Sons, Inc., New York, NY, USA.

Theobald:1990:AFS

[The90b] Kevin Bryan Theobald. Adding fault-tolerance to a static data flow supercomputer. Thesis (M.S.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1990. 158 pp.

Theobald:1991:AFS

[The91] Kevin Bryan Theobald. Adding fault-tolerance to a static data flow supercomputer. Thesis (M.S.), Massachusetts Institute of Technology, Department of Electrical Engineering, Cambridge, MA, USA, 1991. 158 pp.

Taylor:1981:CCS

[THH81] D. J. Taylor, J. F. L. Hopkinson, and C. C. T. Henfrey. Cray-1S and the Cray service provided by the SERC at the Daresbury Laboratory. Computer Physics Communications, 26(3 & 4): 259–265, June 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Taylor:1982:CCS

[THH82] D. J. Taylor, J. F. L. Hopkinson, and C. C. T. Henfrey. The Cray-1s and the Cray service provided by the SERC at the Daresbury Laboratory. Computer Physics Communications, 26 (3–4):259–265, June 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465582901151.

TJHSSC-FCPS:1989:S

[Tho89] Thomas Jefferson High School for Science and Fairfax County Public Schools VA Technology. Supercomputing, 1989. 1 videocassette (30 min.).

Thorson:1990:SPS

[Tho90]  Greg Thorson.  Software

performance simulations of supercomputer memory systems, 1990. 1 videocassette (33 min.).

Thomas:1993:AIA

[Tho93a] J. B. Thomas. Artificial intelligence applications in the nuclear field: Achievements and prospects: The new challenge [invited]. In Kusters et al. [KSW93], pages 353– 366. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Thomborson:1993:DYW

[Tho93b] Clark D. Thomborson. Does your workstation computation belong on a vector supercomputer? Communications of the ACM, 36 (11):41–49, November 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        163363.html.

Thompson:1993:PIG

[Tho93c] Hugh A. Thompson, editor. Proceedings of the 1992 International Gas Research Conference: Orlando, Florida, United States, 16–19 November, 1992. Government Institutes Inc, Rockville, MD, 1993. ISBN 0-86587-496-4. ISSN 0736-5721. LCCN ???? Two volumes.

Thorndyke:1993:SPT

[Tho93d] Lloyd M. Thorndyke. Supercomputer packaging technologies compared. Scientific information bulletin, 18 (2):27–??, April 1, 1993. CO-DEN SINBEM. ISSN 1048-5678.

Thompson:1996:WFCa

[Tho96a] Tom Thompson. The world’s fastest computers. Australian Personal Computer, 17(2): 94–??, February 1, 1996. ISSN 0725-4415.

Thompson:1996:WFCb

[Tho96b] Tom Thompson. The world’s fastest computers. BYTE Magazine, 21(1):44–??, January 1, 1996. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tomizawa:1994:CST

[TIOK94] M. Tomizawa, T. Ishigami, K. Oyama, and K. Kobayashi. Comparative study of transportation evacuation simulation in emergency with Streem and Netsim. In Anonymous [Ano94-75], pages 865–872. ISBN 0-947719-68-7. LCCN ????

Temam:1994:UVL

[TJ94] O. Temam and Y. Jegou.

Using virtual lines to enhance locality exploitation. In Anonymous [Ano94-134], pages 344–353. ISBN ???? LCCN ????

Tuchman:1991:RVP

[TJC91a] Allan Tuchman, David Jablonowski, and George Cybenko. Runtime visualization of program data. Technical Report CSRD 1131, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 7 pp.

Tuchman:1991:SRD

[TJC91b] Allan Tuchman, David Jablonowski, and George Cybenko. A system for remote data visualization. Technical Report CSRD 1067, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1991. 41 pp.

Takai:1989:NTP

[TK89] Yoshiaki Takai and Toshiyasu Kunii. A new trend of the pipelined supercomputer. Technical report 89-023, University of Tokyo, Faculty of Science, Dept. of Information Science, Tokyo, Japan, September 1989. 33 pp.

Takizawa:1993:DPC

[TK93] A. Takizawa and S. Kondo. Development of the physicalcomponent-boundary-fitted-coordinate (PCBFC) method for the analysis of free surface flow. In Kusters et al. [KSW93], pages 164–175. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Tamura:1985:FVS

[TKI85] Hiroshi Tamura, Sachio Kamiya, and Takahiro Ishigai. FACOM VP-100/200: supercomputers with ease of use. Parallel Computing, 2 (2):87–107, June 1985. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Takeda:1993:HTC

[TKI93] T. Takeda, T. Kitada, and H. Ikeda. High-precision transport calculational methods of fast reactor cores [invited]. In Kusters et al. [KSW93], pages 730–741. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Toeroek:1996:PTT

[TKM96] J. Toeroek, J. Kertesz, and S. S. Manna. Phase transitions in a two-dimensional green wave traffic model. In Wolf et al. [WSB96], pages 217–228. ISBN 981-02-2635-7. LCCN ????

Thomadakis:1996:ESS

[TL96] M. E. Thomadakis and J.-C.

Liu. An efficient steepestedge simplex algorithm for SIMD computers. In ACM [ACM96], pages 286–293. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Tehrani-Movabed:1988:ATM

[TM88] Rouzbeh Tehrani-Movabed.

Atmospheric transmittance modelling using CRAY supercomputer. Thesis (M.S.), University of Texas at El Paso, El Paso, TX, USA, 1988. ix + 50 pp.

Tsai:1994:DSM

[TM94a] Y. Tsai and P. K. McKinley. A dominating set model for broadcast in all-port wormhole-routed 2D mesh networks. In Anonymous [Ano94-134], pages 126–135. ISBN ???? LCCN ????

Tsai:1994:EDN

[TM94b] Y.-J. Tsai and P. K. McKinley. An extended dominating node approach to collective communication in all-port wormhole-routed 2D meshes. In IEEE [IEE94c], pages 199–206. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

TalatOdman:1997:NAG

[TMAS97] M. Talat Odman, R. Mathur, K. Alapaty, and R. K. Srivastava. Nested and adaptive grids for multiscale air quality modeling. In Delic and Wheeler [DW97], pages 59–68. ISBN 0-89871-378-1. LCCN ????

Tiemann:1995:AIS

[TMHH95] B. Tiemann, H. V. Muehll, I. Hasler, and E. Hiltebrand. Architecture and implementation of a single-board desktop supercomputer. In Hertzberger and Serazzi [HS95b], pages 481–487. CO-DEN LNCSD9. ISBN 3-540-59393-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1995.

Turner:1994:CUP

[TMP94] C. J. Turner, D. Mosberger, and L. L. Peterson. Cluster-C0*: Understanding the performance limits. In IEEE [IEE94c], pages 229–238. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Tiotto:2020:OCO

[TMT^{+}20] E. Tiotto, B. Mahjour, W. Tsang, X. Xue, T. Islam, and W. Chen. OpenMP 4.5 compiler optimization for GPU offloading. IBM Journal of Research and Development, 64(3/4):14:1– 14:11, May/July 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Tarui:1992:ELM

[TNIA92] T. Tarui, T. Nakagawa, N. Ido, and M. Asaie. Evaluation of the lock mechanism in a snooping cache. In ACM [ACM92b], pages 53–62. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Torng:1987:ISI

[Tor87] H. C. Torng. Introduction to the special issue on supercomputing. IEEE Transactions on Computers, C-36(12):1393–1394, December 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=5009491.

Toussaint:1987:SQ

[Tou87] Doug Toussaint. Supercomputations in QCD. Computer Physics Communications, 45(1–3):111–120, August 1, 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0010465587901457.

Tray:2015:SSU

[TOWC15] Fred T. Tray, Thomas C. Oppe, William A. Ward, and Maureen K. Corcoran. A scalability study using supercomputers for huge finite element variably saturated flow simulations. Scalable Computing: Practice and Experience, 16(2):153–168, ???? 2015. CODEN ???? ISSN 1895-1767. URL https:/

/www.scpe.org/index.php/
scpe/article/view/1087.

Tomonoh:1996:LTC

[TOY96] Yuzuru Tomonoh, Kunio Ohno, and Seiken Yano. LSI technologies for CMOS-Based supercomputer SX-4. Nippon Electric Company research and development, 37 (4):508–??, ???? 1996. CO-DEN NECRAU. ISSN 0048-0436.

Tichy:1993:PPS

[TP93] W. F. Tichy and M. Philippsen.

Programming parallel supercomputers [invited]. In Kusters et al. [KSW93], pages 3–9. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Tu:1995:GSD

[TP95] P. Tu and D. Padua. Gated SSA-Based demand-driven symbolic analysis for parallelizing compilers. In ACM [ACM95a], pages 414–423. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Thompson:1997:CAG

[TP97] S. L. Thompson and D. Pollard. Computational aspects of the GENESIS earth systems modeling project. In Delic and Wheeler [DW97], pages 13–20. ISBN 0-89871-378-1. LCCN ????

Tu:2019:HPC

[TPJ^{+}19] Wanqing Tu, Florin Pop, Weijia Jia, Jie Wu, and Mauro Iacono. Highperformance computing in edge computing networks. Journal of Parallel and Distributed Computing, 123(??): 230, January 2019. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0743731518308086.

Thornton:1986:SIF

[TR86] Earl A. (Earl Arthur) Thornton and R. Ramakrishnan. Supercomputer implementation of finite element algorithms for high speed compressible flows progress report for the period ended June 30, 1986. NASA contractor report NASA CR-177065, Old Dominion University Research Foundation, Norfolk, VA, USA, June 30, 1986. 155 pp.

Travis:1989:EDD

[Tra89] Bill Travis. EDN’s DPS-chip directory. EDN, 34(21):125– ??, October 12, 1989. CO-DEN EDNSBH. ISSN 0012-7515, 0364-6637.

Trefil:1997:BDS

[Tre97] James S. Trefil. Between Darwin and the supercomputer. John Wiley and Sons, Inc., New York, NY, USA, 1997. ISBN 0-471-15536-5. ???? pp. LCCN BF444.T74 1997.

Triolet:1985:IAP

[Tri85] Remi Triolet. Interprocedural analysis for program restructuring with parafrase. Technical Report CSRD-538, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1985. 77 pp.

Trifonov:1993:DL

[Tri93] E. N. Trifonov. DNA as a language. In Lim et al. [L^{+}93], pages 103–112. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Tristram:1995:LBS

[Tri95a] C. Tristram. Load-Balancing Software: Load-balancing software promises to give you the performance clout of a supercomputer from the workstations you already have. UnixWorld’s Open Computing, 12(8):60–??, 1995. CO-DEN OPCOEB. ISSN 1072-4044.

Tristram:1995:LSL

[Tri95b] C. Tristram. Load-balancing software: Load-balancing software promises to give you the performance clout of a supercomputer from the workstations you already have. UnixWorld’s Open Computing, 12(8):60–??, ???? 1995. CODEN OPCOEB. ISSN 1072-4044.

Tristram:1995:TTK

[Tri95c] Claire Tristram. Ten things to know about … loadbalancing software. Open Computing, 12(8):60–??, August 1, 1995. CODEN OP-COEB. ISSN 1078-2370.

Tang:2013:TBS

[TRLD13] Wei Tang, Dongxu Ren, Zhiling Lan, and Narayan Desai. Toward balanced and sustainable job scheduling for production supercomputers. Parallel Computing, 39(12): 753–768, December 2013. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819113000999.

Truhlar:1988:BRS

[Tru88] Donald G. Truhlar. Book review: Scientific computing on vector computers: W. Schönauer, special topics in supercomputing series (volume 2), North-Holland, Amsterdam, 1987. Dfl.180.00 (hard cover). Computer Physics Communications, 50 (3):413, August 1988. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465588901944.

Thistle:1988:PAH

[TS88] Mark R. Thistle and Burton J. Smith. A processor architecture for horizon. Technical report SRC-TR-88-010, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 7 pp.

Thistle:1990:PPM

[TS90] Mark R. Thistle and Thomas Lawrence Sterling. A procedural programming methodology based on a dataflow computing model for the Zycad SDE. Technical report SRC-TR-90-001, Supercomputing Research Center: IDA, Lanham, MD, USA, January 11, 1990. 19 pp. See [TS91].

Thistle:1991:FGM

[TS91] Mark R. Thistle and Thomas Lawrence Sterling. A fine grain MIMD system with hybrid eventdriven/dataflow synchronization for bit-oriented computation. Technical report SRC-TR-91-038, Supercomputing Research Center: IDA, Lanham, MD, USA, August 22, 1991. 19 pp. Supersedes Technical Report SRC-TR-90-001 [TS90].

Takkella:1994:CEB

[TS94] S. S. Takkella and S. Seidel. Complete exchange and broadcast algorithms for meshes. In IEEE [IEE94c], pages 422–428. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Taylor:1994:DCS

[TSCG94] P. M. Taylor, K. G. Swift, A. J. Creed, and M. T. Griesser. A dynamically configurable surface based on shape-memory alloy (SMA) wires for virtual reality applications. In Anonymous [Ano94-75], pages 617–624. ISBN 0-947719-68-7. LCCN ????

Taylor:1994:RTA

[TSSK94] P. M. Taylor, K. G. Swift, D. R. Smith, and R. S.

Kalawsky. The reactabot: a tool for aiding the assessment of manual handling processes in assembly/ disassembly. In Anonymous [Ano94-75], pages 639–646. ISBN 0-947719-68-7. LCCN ????

Tsuda:1991:DIV

[Tsu91]  T. Tsuda.  Design and

implementation of a vectorizing compiler for the block-structured language PASCAL. In Anonymous [Ano91q], pages 38–45. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Tsyrkov:1994:NBS

[Tsy94] G. Tsyrkov. From the nuclear bomb to supercomputers. International Affairs (Royal Institute of International Affairs 1944–), ??(9): 41–??, ???? 1994. ISSN 0130-9641.

Tsao:1993:CAB

[TT93] Anna Tsao and Thomas Turnbull. A comparison of algorithms for banded matrix multiplication. Technical report SRC-TR-93-092, Supercomputing Research Center: IDA, Lanham, MD, USA, April 9, 1993. 9 pp.

Taboada:2011:DLC

[TTD^{+}11] Guillermo L. Taboada, Juan Touriño, Ramón Doallo, Aamir Shafi, Mark Baker, and Bryan Carpenter. Device level communication libraries for high-performance computing in Java. Concurrency and Computation: Practice and Experience, 23(18):2382– 2403, December 25, 2011. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Tian:2020:SSS

[TTM^{+}20] S. Tian, T. Takken, V. Mahaney, C. Marroquin, M. Schultz, M. Hoffmeyer, Y. Yao, K. O’Connell, A. Yuksel, and P. Coteus. Summit and Sierra supercomputer cooling solutions. IBM Journal of Research and Development, 64(3/4):5:1–5:12, May/ July 2020. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Tuchman:1991:VSR

[Tuc91] Allan Tuchman. Vista: a system for remote data visualization. Technical Report CSRD 1107, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1991. 6 pp.

Turchin:1979:SSB

[Tur79] V. F. Turchin. A supercompiler system based on the language REFAL. ACM SIG-PLAN Notices, 14(2):46–54, February 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turchin:1986:CS

[Tur86] Valentin F. Turchin. The concept of a supercompiler. ACM Transactions on Programming Languages and Systems, 8(3):292–325, July 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0164-0925/
                                        5957.html.

Turner:1989:SMI

[Tur89] Stephen W. Turner. Shared memory and interconnection network performance for vector multiprocessors. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1989. ix + 89 pp.

Turnbull:1990:SCS

[Tur90] Thomas Turnbull. Sequence comparison in a SIMD environment. Technical report SRC-TR-91-023, Supercomputing Research Center: IDA, Lanham, MD, USA, September 28, 1990. 16 pp.

Turner:1988:PSM

[TV88] Stephen W. Turner and Alexander Veidenbaum. Performance of a shared memory system for vector multiprocessors. Technical Report CSRD 745, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 11 pp.

Turner:1989:BTM

[TV89] Stephen W. Turner and Alexander Veidenbaum. Burst traffic in MIN-based sharedmemory systems. Technical Report CSRD 855, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, February 1989. 11 pp.

Trobec:2016:INP

[TVT^{+}16] Roman Trobec, Radivoje Vasiljević, Milo Tomasević, Veljko Milutinović, Ramon Beivide, and Mateo Valero. Interconnection networks in petascale computer systems: a survey. ACM Computing Surveys, 49(3):44:1– 44:??, November 2016. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Troyer:1992:IQM

[TW92] M. Troyer and D. Wuertz.

Implementation of the quantum Monte Carlo world line algorithm on the NEC SX-3 supercomputer. In Anonymous [Ano92a], pages 48–51. ISBN ???? LCCN ????

Torrellas:1996:OPD

[TY96] J. Torrellas and L. Yang. Optimizing primary data caches for parallel scientific applications: The pool buffer approach. In ACM [ACM96], pages 141–148. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Taiwo:1993:DSM

[TYK93] T. A. Taiwo, W. S. Yang, and H. S. Khalil. Direct solution of the mathematical adjoint equations for interface current nodal formulation. In Kusters et al. [KSW93], pages 507–521. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Tumanyan:1993:ENS

[TYKE93] V. G. Tumanyan, S. V. Yakovleva, Y. V. Krovatsky, and N. G. Esipova. The effects of nucleotide substitution on amino acid substitution. In Lim et al. [L^{+}93], pages 623–634. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Tyson:1991:MTR

[Tys91] Laura D’Andrea Tyson. Managing trade by rules and outcomes. California management review, 34(1):115– ??, Fall 1991. ISSN 0008-1256.

Tzeng:1985:FMI

[TYZ85] Nian-Feng Tzeng, Pen-Chung Yew, and Chuan-Qi Zhu.

Fault-diagnosis in a multiplepath interconnection network. Technical Report CSRD-518, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1985. i + 22 pp.

Tzeng:1988:RFI

[TYZ88] Nian-Feng Tzeng, Pen-Chung Yew, and Chuan-Qi Zhu. Realizing fault-tolerant interconnection networks via chaining. Technical Report CSRD 764, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 5 pp.

Tang:1989:PLL

[TYZ89] Peiyi Tang, Pen-Chung Yew, and Chuan-Qi Zhu. A parallel linked list for shared-memory multiprocessors. Technical Report CSRD 879, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1989. 17 pp.

Tang:1990:CTD

[TYZ90] Peiyi Tang, Pen-Chung Yew, and Chuan-Qi Zhu. Compiler techniques for data synchronization in nested parallel loops. Technical Report CSRD 1092, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1990. 10 pp.

Tang:1994:RDC

[TZ94] P. Tang and J. N. Zigman. Reducing data communication overhead for doacross loop nests. In Anonymous [Ano94-134], pages 44–53. ISBN ???? LCCN ????

Tzeng:1986:FMI

[Tze86] Nian-Feng Tzeng. Faulttolerant multiprocessor interconnection networks and their fault-diagnoses. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. vi + 130 pp.

Tzeng:1988:IPS

[Tze88]  Ping-Cheng Tzeng.  Im-

plementation of phase shift plus correction prestack migration methods on the NEC SX-2 supercomputer. Thesis (M.S.), Dept. of Computer Science, University of Houston, Houston, TX, USA, 1988. viii + 78 pp.

Tang:1988:AGD

[TZY88] Peiyi Tang, Chuan-Qi Zhu, and Pen-Chung Yew. Algorithms for generating datalevel synchronization instructions. Technical Report CSRD 733, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 26 pp.

Uchida:1986:FMT

[Uch86]  Hiroshi Uchida.  Fujitsu

machine translation system: ATLAS. Future Generation Computer Systems, 2(2):95– 100, June 1986. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Uchida:1996:HVV

[Uch96] N. Uchida. Hardware of VX/VPP300/VPP700 series of vector-parallel supercomputer systems. Fujitsu, 47(6): 434–??, ???? 1996. CODEN FUJTAR. ISSN 0016-2515.

Uchida:1997:HVV

[Uch97] N. Uchida. Hardware of VX/VPP300/VPP700 series of vector-parallel supercomputer systems. Fujitsu scientific and technical journal, 33 (1):6–??, ???? 1997. CODEN FUSTA4. ISSN 0016-2523.

Uberbacher:1993:GRA

[UEGM93] E. C. Uberbacher, J. R. Einstein, X. Guan, and R. J. Mural. Gene recognition and assembly in the GRAIL system: Progress and challenges. In Lim et al. [L^{+}93], pages 465–478. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Uenohara:1993:PCE

[Uen93] Y. Uenohara. Parallelism in continuous energy Monte Carlo method for neutron transport. In Kusters et al. [KSW93], pages 406–417. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Underwood:2009:SSL

[UHU09] Keith D. Underwood, K. Scott Hemmert, and Craig D. Ulmer. From silicon to science: The long road to production reconfigurable supercomputing. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 2(4): 26:1–26:??, September 2009. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

UH-CLRICIS:1989:RSS

[UL89] University of Houston — Clear Lake.Research Institute for Computing and Information Systems and Lyndon B.Johnson Space Center, editors. RICIS symposium ’89: Supercomputing: parallel and numerically intensive computing: research and applications, RICIS conference series. University of Houston, Houston, TX, USA, 1989.

Ullman:1983:STA

[Ull83] Jeffrey D. Ullman. Some thoughts about supercomputer organization. Technical Report STAN-CS-83-987, Stanford University, Dept. of Computer Science, Stanford, CA, USA, October 1983. 17 pp.

Ullman:1984:FSS

[Ull84] Jeffrey D. Ullman. Flux, sorting, and supercomputer organization for AI applications. Journal of Parallel and Distributed Computing, 1 (2):133–151, November 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Ulrich:2012:SC

[Ulr12] L. Ulrich. State of charge.

IEEE Spectrum, 49(1):56– 59, January 2012. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

MSI:1984:U

[UMS84] UMSI, 1984. University of Minnesota Supercomputer Institute, Minneapolis, MN, USA.

US-C-HCST-SSRT:1986:FSP

[Uni86a] United States.Congress.House.Committee on Science and Research Technology. Subcommittee on Science and Technology. Federal supercomputer programs and policies: hearing before the Subcomittee on Energy Development and Applications and the Subcommittee on Science, Research, and Technology of the Committee on Science and Technology, House of Representatives, Ninety-ninth Congress, first session, June 10, 1985. Washington, DC, USA, 1986. iv + 801 pp. Distributed to some depository libraries in microfiche. Shipping list no.: 86-76-P. “No. 44.” Item 1025-A-1, 1025-A-2 (microfiche).

US-C-HCSST-SSRT:1986:FSP

[Uni86b] United States.Congress.House.Committee on Science and Research Technology. Subcommittee on Science and Technology. Federal supercomputer programs and policies: hearing before the Subcommittee on Energy Development and Applications and the Subcommittee on Science, Research, and Technology of the Committee on Science and Technology, House of Representatives, Ninety-ninth Congress, first session, June 10, 1985. Washington, DC, USA, 1986. iv + 801 pp. Distributed to some depository libraries in microfiche. Shipping list no. 86-76-P. No. 44.

US-CHCST-SEDA:1986:FSP

[Uni86c] United States.Congress.House.Committee on Science and Technology. Subcommittee on Energy Development and Applications. Federal supercomputer programs and policies hearing before the Subcommittee on Energy Development and Applications and the Subcommittee on Science, Research, and Technology of the Committee on Science and Technology, House of Representatives, Ninety-ninth Congress, first session, June 10, 1985. United States Government Printing Office, Washington, DC, USA, 1986. iv + 801 pp.

NASA:1987:SAP

[Uni87a] United States.National Aeronautics and Space Administration. Scientific and Technical Information Branch, editor. Supercomputing in aerospace: proceedings of a symposium held at the NASA Ames Research Center, Moffett Field, California, March 10–12, 1987, number 2454 in NASA conference publication; 2454. NASA, Washington, DC, USA, 1987. ISBN ???? LCCN NAS 1.55:2454. Distributed to depository libraries in microfiche. Microfiche. [Washington, DC?: National Aeronautics and Space Administration], 1987. 4 microfiches.

UA-CS:1987:CSI

[Uni87b] University of Alberta.Computing Services. Cyber 205 supercomputer introduction. Technical Report R515.0187, Computing Services, University of Alberta, Edmonton, AB, Canada, January 22, 1987. iv + 50 pp.

UIUC-CSRD:1987:PPA

[Uni87c] University of Illinois at Urbana-Champaign.Center for and Supercomputing Research and Development, editor. Papers presented at the International Conference on Parallel Processing, August 18–21, 1987. University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. LCCN ????

US-C-HCSST-SSRT:1989:USI

[Uni89a] United States.Congress.House.Committee on Science, Space, and Research Technology. Subcommittee on Science and Technology. US supercomputer industry: hearing before the Subcommittee on Science, Research, and Technology of the Committee on Science, Space, and Technology, US House of Representatives, One Hundred First Congress, first session, June 20, 1989. Congressional Sales Office, US GPO, 1989. iii + 140 pp. Distributed to some depository libraries in microfiche. Shipping list no.: 89-710-P. “No. 45” Item 1025-A-1, 1025-A-2 (MF).

UTS-CHPC-SPSG:1989:SUT

[Uni89b] University of Texas System.

Center for High Performance Computing. Strategic Planning Study Group, Austin, TX, USA. Supercomputing in the University of Texas System, 1989. 49 pp.

US-DOE-OER-SCS:1991:PFE

[Uni91a] United States.Dept.of Energy.Office of Energy Research.Scientific Computing Staff, editor. Proceedings of the first energy research power supercomputer users symposium: May 21–22, 1991, Gaithersburg, Maryland. National Technical Information Service, Washington, DC, USA, 1991. LCCN DOC E 1.10:9105202 mf11 Microcopy Rm. Distributed to depository libraries in microfiche. Shipping list no.: 92-2142-M. “CONF-9105202.” Microfiche. [Washington, DC?]: Supt. of Docs., US GPO,

[1992] 3 microfiches: negative.

NASA:1991:SNS

[Uni91b] United States.National Aeronautics and Space Administration. [Supercomputer networking for space science applications]: [final report, 19 Jan. — 31 Dec. 1991]. NASA, Washington, DC, USA, 1991. ?? pp. Distributed to depository libraries in microfiche. Shipping list no.: 94-0264-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1994] 1 microfiche.

US-CHGO-LNSS:1992:AGU

[Uni92a] United States.Congress.House.Committee on Government and Operations. Legislation and National Security Subcommittee. Is the administration giving away the U.S. supercomputer industry? Hearings before the Legislation and National Security Subcommittee of the Committee on Government Operations, House of Representatives, One Hundred Second Congress, second session, July 1 and 8, 1992. United States Government Printing Office, Washington, DC, USA, 1992. ISBN 0-16-039092-3. iv + 452 pp. LCCN J61 .E9 102nd no.58.

US-C-HCGO-LNSS:1992:AGU

[Uni92b] United States.Congress.House.Committee on Government Operations.Legislation and National Security Subcommittee. Is the administration giving away the US supercomputer industry?: hearings before the Legislation and National Security Subcommittee of the Committee on Government Operations, House of Representatives, One Hundred Second Congress, second session, July 1 and 8, 1992. Supt. of Docs., Congressional Sales Office, 1992. iv + 452 pp. Distributed to some depository libraries in microfiche. Shipping list no.: 92-0580-P. Item 1016-A, 1016-B (MF).

US-DOC-TA:1992:GMS

[Uni92c] United States.Dept.of Commerce.Technology Administration. Global markets for supercomputers: the impact of the U.S.-Japan supercomputer procurement agreement, a report to the Congress. U.S. Dept of Commerce, Technology Administration, Washington, DC, USA, 1992. ix + 47 pp.

US-EPA-NDPD-NESC:1992:NNE

[Uni92d] United States.Environmental Protection Agency.National Data and Processing Division and National Environmental Supercomputing Center. NESC, National Environmental Supercomputing Center. Technical Report EPA 208-F-92-001, U.S. Environmental Protection Agency, Office of Administration and Resources Management, National Data Processing Division, Research Triangle Park, NC, USA, October 1, 1992. 12 pp.

US-EPA-NDPD:1992:NNE

[Uni92e] United States.Environmental Protection Agency.National Data Processing Division. NESC, National Environmental Supercomputing Center. Office of Administration and Resources Management, National Data Processing Division, 1992. 12 pp.

US-DOE-OSC:1993:RSE

[Uni93] United States.Dept.of Energy.Office of Scientific Computing. Requirements for supercomputing in energy research: the transition to massively parallel computing. Technical Report DOE/ER-0587, U.S. Dept. of Energy, Office of Energy Research, Office of Scientific Computing, Washington, DC, February 1993. ii + 73 pp. Distributed to depository libraries in microfiche. Shipping list no.: 93-1171-M. “February 1993.” “DOE/ER-0587.” Microfiche. [Washington, DC?]: Supt. of Docs., US GPO, [1993] 1 microfiche: negative.

US-GAO-RCED:1995:RES

[Uni95] United States.General Accounting Office.RCED. Relocation of Energy’s supercomputer. Technical Report GAO/RCED-96-55R. B-270732, United States. General Accounting Office. RCED, Washington, DC, USA, December 19, 1995. 5 + 2 pp.

US-EPA-OIG:1996:RAF

[Uni96] United States.Environmental Protection Agency.Office of the Inspector General. Report of audit: Final report of audit on the establishment of the National Environmental Supercomputing Facility in Bay City, Michigan. Technical Report E1BMF4-22-0359-6100306, U.S. Environmental Protection Agency, Washington, DC, USA, September 30, 1996. various pp.

USC-NCNS:1998:USE

[Uni98] United States Congress.House Committee on National Security. U.S. supercomputer export control policy: Committee on National Security, House of Representatives, One Hundred Fifth Congress, first session, hearing held November 13, 1997. Washington, DC, USA, 1998. iii + 113 pp. Shipping list no. 98-0179-P. H.N.S.C. no. 105-25.

Uthayopas:2001:FSR

[UP01] Putchong Uthayopas and Sugree Phatanapherom. Fast and scalable real-time monitoring system for Beowulf clusters. Lecture Notes in Computer Science, 2131:201– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2131/21310201.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310201.
pdf.

Uht:1987:CHS

[UPK87] Augustus Kinzel Uht, C. D.

(Constantine D.) Polychronopoulos, and John F. Kolen. On the combination of hardware and software concurrency extraction methods. Technical Report CSRD 694, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1987. 9 pp.

Usbeck:1995:NSA

[UR95] R. Usbeck and P. Rujan. Numerical simulations of the ACh-Driven synapse at the neuromuscular junction. In Herrmann et al. [HWP95], pages 265–272. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Uthayopas:2001:PCK

[US01] Putchong Uthayopas and Somsak Sriprayoonsakul. The PIRUN cluster at Kasetsart University: The most powerful supercomputer in Thailand. ;login: the USENIX Association newsletter, 26 (5):??, August 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2001-08/pdfs/utha.
                                        pdf.

USENIX:1990:USI

[USE90] USENIX Association, editor. UNIX Security II: USENIX workshop proceedings, August 27–28, 1990, Portland, Oregon. USENIX Association, Berkeley, CA, USA, 1990. ISBN ???? LCCN QA 76.9 A25 U55 1990.

Usevitch:1993:PRF

[Use93] Brian E. Usevitch. Perfect reconstruction filter banks for adaptive filtering and coding. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1993. vii + 110 pp.

USENIX:2000:PAL

[USE00a]  USENIX, editor.  Pro-

ceedings of the 4th Annual Linux Showcase and Conference, Atlanta, October 10– 14, 2000, Atlanta, Georgia, USA. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/

.

USENIX:2000:PUW

[USE00b] USENIX, editor. Proceedings of the 4th USENIX Windows Systems Symposium: August 3–4, 2000, Seattle, Washington, USA. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-20-0. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/usenix-win2000.

USENIX:2001:PAL

[USE01] USENIX, editor. Proceedings of the 5th Annual Linux Showcase and Conference, November 5–10, 2001, Oakland, CA. USENIX Association, Berkeley, CA, USA, 2001. ISBN ???? LCCN

????  URL http://www.
linuxshowcase.org/tech.
html.

Ujaldon:1996:EEE

[USZS96] M. Ujaldon, S. D. Sharma, E. L. Zapata, and J. Saltz. Experimental evaluation of efficient sparse matrix distributions. In ACM [ACM96], pages 78–85. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Uehara:1991:BVI

[UT91] T. Uehara and T. Tsuda. Benchmarking vector indirect load/store instructions. In Anonymous [Ano91q], pages 134–143. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Uthup:1994:ASV

[Uth94] B. Uthup. Anicol — scientific visualisation tool for CFD. In Mahajan et al. [M^{+}94], pages 60–?? ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Uchida:1993:VS

[UU93] K. Uchida and T. Utsumi. VPP500 supercomputer. In Hoffmann and Kauranne [HK93b], pages 445– 452. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Umeda:1994:ENN

[UU94] A. Umeda and K. Ueda. Effectiveness of the novel NRLM method for characterizing accelerometers verified by the round robin experiment organized by National Research Laboratory of Metrology. In Anonymous [Ano94-75], pages 241–248. ISBN 0-947719-68-7. LCCN ????

Ujaldon:1995:ERS

[UZ95] M. Ujaldon and E. L. Zapata. Efficient resolution of sparse indirections in dataparallel compilers. In ACM [ACM95a], pages 117–126. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

VanaVaramban:1994:VMT

[VA94] S. Vana Varamban and S. Anthonysamy. Visual modeling of thermochemical behaviour of nuclear fuels. In Mahajan et al. [M^{+}94], pages 47–53. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Vafidis:1988:SFD

[Vaf88] Antonios Vafidis. Supercomputer finite difference methods for seismic wave propagation. Thesis (Ph.D.), University of Alberta, Ottawa, ON, Canada, 1988. 184 pp.

Vagnetti:1988:SAP

[Vag88] F. Vagnetti, editor. Supercomputing in astrophysics: proceedings of a workshop held at Osservatorio di Roma Monteporizo Catone, 10-11 March 1988, volume 88/1 of Astronet special publication. Dipartimento di Fisica - II Universita di Roma, Rome, Italy, 1988. LCCN QB51.3.E43S87 1988.

Vidal-Ascon:1990:PPB

[VAGRMVA90] Luis Vidal-Ascon, Thomas Grace, Perpetua Ruiz-Mostacero, and Alberto Vidal-Ascon. A parallel preconditioned block conjugate gradient method for solving large systems of linear equations on MIMD supercomputer. Mathematical and computer modelling, 14(??):112–120, ???? 1990. CODEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Vajapeyam:1991:ILC

[Vaj91] Sriram Vajapeyam. Instructionlevel characterization of the Cray Y-MP processor. Ph.D. thesis, Computer Sciences Department, University of Wisconsin-Madison, Madison, WI, USA, 1991.

Valauskas:1994:MM

[Val94] Edward J. Valauskas. Mac monitor. Database, 17(4):86– ??, August 1994. CODEN DTBSDQ. ISSN 0162-4105.

VanZandt:1986:ADC

[Van86] John Van Zandt. The architecture of a dataflow computer, 1986. 1 videocassette (57 min.).

VanderSteen:1991:AEB

[Van91a] A. J. Van der Steen. The activities of the EuroBen benchmark group. In Anonymous [Ano91q], pages 111– 119. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

VanderVorst:1991:UBU

[Van91b] H. A. Van der Vorst. The use of Bi-CGSTAB for unsymmetric linear systems. In Anonymous [Ano91q], pages 206–212. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

VanDerSluis:1993:CSV

[Van93] J. Van Der Sluis. 93SF055 computer simulation and vehicle front optimisation. In Anonymous [Ano93-31], pages 245–250. ISBN 0-947719-62-8. LCCN ????

VanGemund:1994:CPM

[Van94] A. J. C. Van Gemund.

Compiling performance models from parallel programs. In Anonymous [Ano94-134], pages 303–312. ISBN ???? LCCN ????

VanGijzen:1995:PIS

[Van95a] M. Van Gijzen. Parallel iterative solution methods for linear finite element computations on the Cray T3D. Lecture Notes in Computer Science, 919:723–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanGijzen:1995:LSF

[van95b] M. B. van Gijzen. Large scale finite element computations with GMRES-like methods on a Cray Y-MP. Applied Numerical Mathematics: Transactions of IMACS, 19(1-2):

51–62, November 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vanhelsuwe:1997:CSC

[Van97] Laurence Vanhelsuwé. Cover story: Create your own supercomputer with Java. JavaWorld: IDG’s magazine for the Java community, 2 (1):??, January 1997. CO-DEN ???? ISSN 1091-

8906. URL http://www.
javaworld.com/javaworld/
jw-01-1997/jw-01-dampp.
htm.

Vanderbauwhede:2013:HPC

[Van13] Wim Vanderbauwhede, editor. High-performance Computing Using FPGAs. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-1790-2, 1-4614-1791-0. xi + 803 pp. LCCN ????

Vichnevetsky:1982:IWC

[VAS82] R. Vichnevetsky, W. F. Ames, and S. Sankar, editors. 10th IMACS World Congress: proceeding, August 8–13, 1982, Montréal, Canada. International Association for Mathematics and Computers in Simulation, Brussels, Belgium, 1982. LCCN TA343 .I523 1982. Five volumes.

Vos:1990:DMB

[VB90] J. B. Vos and C. M. Bergman.

The development of a multiblock flow solver for parallel computers. In Pitcher [Pit90], pages 521–532. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

VanBuul:1994:UMD

[VD94] K. Van Buul and M. De Krom. UNI-ECU, a modular development system for automotive applications. In Anonymous [Ano94-75], pages 881–888. ISBN 0-947719-68-7. LCCN ????

Veje:1996:KDW

[VD96] C. T. Veje and P. Dimon.

Kinematic density waves in a two-dimensional granular flow. In Wolf et al. [WSB96], pages 299–304. ISBN 981-02-2635-7. LCCN ????

vandeGeijn:1997:UP

[vdG97] Robert A. van de Geijn. Using PLAPACK. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, 1997. ISBN 0-262-72026-4. xvii + 194 pp. LCCN QA185.D37V36 1997. US$30.00. URL http://www.

mitpress.com/book-home.
tcl?isbn=0262720264.

Vetter:1991:NSE

[VDK91] Ronald J. Vetter, David H. C. Du, and Alan E. Klietz. Network supercomputing: experiments with a CRAY-2 to CM-2 HiPPI connection. AH-PCRC preprint 91-83, Army High Performance Computing Research Center, Minneapolis, MN, USA, 1991. iii + 17 pp.

Vetter:1992:NS

[VDK92] Ronald J. Vetter, David H. C. Du, and Alan E. Klietz. Network supercomputing. IEEE network, 6(3):38–45, May 1992. CODEN IENEET. ISSN 0890-8044.

vanderSteen:1996:ORS

[vdSD96a] Aad J. van der Steen and Jack Dongarra. Overview of recent supercomputers. NHSE Review, 1(1):??, February 10, 1996.

vanderSteen:1996:ORSb

[vdSD96b] Aad J. van der Steen and Jack Dongarra. Overview of recent supercomputers. NHSE Review, 1(1):??, February 10, 1996.

Vetter:2005:EHP

[VdSK^{+}05] Jeffrey S. Vetter, Bronis R. de Supinski, Lynn Kissel, John May, and Sheila Vaidya. Evaluating high-performance computers. Concurrency and Computation: Practice and Experience, 17(10):1239– 1270, August 25, 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vanderVorst:1991:IMS

[vdV91] H. A. van der Vorst. Iterative methods for the solution of large systems of equations on supercomputers. Adv. Water Resources, 13:137–146, 1991. CODEN AWREDI. ISSN 0309-1708 (print), 1872-9657 (electronic).

Veidenbaum:1985:COA

[Vei85] Alexander Veidenbaum. Compiler optimizations and architecture design issues for multiprocessors. Thesis (Ph.D.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1985. vi + 129 pp.

Verlo:1995:TSD

[Ver95] Alan Verlo. Topological surface deformation: an application of virtual reality with real-time supercomputing. Thesis (M.S. in electrical engineering and computer science), University of Illinois at Chicago, Chicago, IL, USA, 1995. vi + 66 pp.

Versteeg:1997:RUG

[Ver97] R. Versteeg. Requirements for the use of ground penetrating radar as an organic contaminant detection tool. In Delic and Wheeler [DW97], pages 313– 316. ISBN 0-89871-378-1. LCCN ????

Vetter:2012:REC

[Vet12] Jeffery S. Vetter. On the road to exascale computing: contemporary architectures in high performance computing. Chapman and Hall/CRC, Boca Raton, FL, USA, 2012. ISBN 1-4665-6834-8. ???? pp. LCCN ????

Vezolle:1995:PCE

[Vez95] P. Vezolle. Parallel computational electromagnetics on the CRAY T3D using boundary element method. Lecture Notes in Computer Science, 919:658–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veshagh:1993:DNF

[VF93] A. Veshagh and M. Fenton. 93SC040 development of a non-linear fluid flow model of an automotive engine lubrication system. In Anonymous [Ano93-31], pages 53– 60. ISBN 0-947719-62-8. LCCN ???? Venkateswaran:2004:MPN

[VFK^{+}04] N. Venkateswaran, Waran Research Foundation, Aditya Krishnan, S. Niranjan Kumar, Arrvindh Shriraman, and Srinivas Sridharan. Memory in processor: a novel design paradigm for supercomputing architectures. ACM SIGARCH Computer Architecture News, 32(3):19– 26, June 2004. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Vajapeyam:1993:TES

[VH93a] Stiram Vajapeyam and Wei-Chung C. Hsu. Toward effective scalar hardware for highly vectorizable applications. Journal of Parallel and Distributed Computing, 19 (3):147–162, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1100/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1100/production/pdf;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1101/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1101/production/pdf.

Vollmers:1993:VSS

[VH93b] H. Vollmers and A. Hilgenstock. Visualization and simulation of a 3-D stator/rotor flow field in an axial turbine. In Kusters et al. [KSW93], pages 789–790. ISBN 3-923704-11-9. LCCN ???? Two volumes.

VanDessel:1994:DOE

[VHJB94] M. Van Dessel, S. Henneberger, T. B. Johansson, and R. Belmans. Design and optimisation of extra small and micromotors. In Anonymous [Ano94-75], pages 115–

122. ISBN 0-947719-68-7. LCCN ????

Vick:1980:AAS

[VKK80] C. R. Vick, S. P. Kartashev, and S. I. Kartashev. Adaptable architectures for super systems. Computer, 13(11): 17–35, November 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vonLaszewski:1999:LCM

[vL99] Gregor von Laszewski. A loosely coupled metacomputer: co-operating job submissions across multiple supercomputing sites. Concurrency: practice and experience, 11(15):933–948, December 25, 1999. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/71005731/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71005731&PLACEBO=IE.
pdf.

Valero:1992:CAV

[VLA92] M. Valero, T. Lang, and E. Ayguade. Conflict-free access of vectors with powerof-two-strides. In ACM [ACM92b], pages 149–156. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Vanka:1987:VMF

[VM87] S. P. Vanka and K. P. Misegades. Vectorized multigrid fluid flow calculations on a Cray X-MP/48. International Journal for Numerical Methods in Fluids, 7(6):635– 648, June 1987. CODEN IJNFDW. ISSN 0271-2091.

VandeVen:1994:SAM

[VM94] P. Van de Ven and J. P.

Maes. Some advantages of monoacid/diacid corrosion inhibitors in engine coolants. In Anonymous [Ano94-75], pages 653–660. ISBN 0-947719-68-7. LCCN ????

Vazhkudai:2007:RTD

[VM07] Sudharshan Vazhkudai and Xiaosong Ma. Recovering transient data: automated on-demand data reconstruction and offloading for supercomputers. Operating Systems Review, 41(1):14– 18, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Villarino:1993:HUT

[VMS93] E. A. Villarino, E. Maartensson, and R. J. J. Stamm’ler. HELIOS: Usage of transformation laws for angularydependent collision probabilities. In Kusters et al. [KSW93], pages 433–444.

ISBN 3-923704-11-9. LCCN ???? Two volumes.

Vaughan-Nichols:2004:NTR

[VN04] Steven J. Vaughan-Nichols. New trends revive supercomputing industry. Computer, 37(2):10–??, February 2004. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2004/02/r2010.htm;   http:
//csdl.computer.org/dl/
mags/co/2004/02/r2010.pdf.

Vergnaud:1993:TMC

[VNB93] T. Vergnaud, J. C. Nimal, and J. P. Both. TRIPOLI-3: A Monte Carlo code with a powerful and automatic biasing. In Kusters et al. [KSW93], pages 756– 765. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Viniotis:1993:ATM

[VO93] Yannis Viniotis and Raif O. Onvural, editors. Asynchronous transfer mode networks: Proceedings of TRI-COMM ’93, held April 26– 27, 1993, in Raleigh, North Carolina. Plenum Press, New York, NY, USA; London, UK, 1993. ISBN 0-306-44486-0. LCCN TK5105 .A83 1993.

Voelcker:1989:TPC

[Voe89a]  J. Voelcker.  Technology

’89: personal computers. IEEE Spectrum, 26(1):31– 34, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Voelcker:1989:TS

[Voe89b] J. Voelcker. Technology ’89:

software. IEEE Spectrum, 26 (1):37–39, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vogel:1993:NVI

[Vog93] D. L. Vogel. A new, versatile iteration algorithm for the analytic nodal method. In Kusters et al. [KSW93], pages 522–532. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Voigt:1994:FCH

[Voi94] Robert G. Voigt. Focus on CSE: HPCC/IITA /NII: The alphabet soup of US high-performance computing. IEEE Computational Science & Engineering, 1(2): 80, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Vollum:1989:XPD

[Vol89] Charles Vollum. X parallel desktop supercomputer: Transputers play host. Digest of Papers — IEEE Computer Society International Conference, pages 61–62, February 1989. CODEN DCSIDU. ISBN 0-8186-1909-0. Available from IEEE Service Center, Piscataway, NJ, USA.

Venkateswaran:1993:NVP

[VPDA93] N. Venkateswaran, S. Pattabiraman, R. Devanthan, and A. Ahmed. A novel VLSI processor architecture for supercomputing arrays. In NASA [NAS93], page 8.6. ISBN ???? LCCN ????

Villaverde:2001:PPI

[VPGG01] K. Villaverde, E. Pontelli, H. Guo, and G. Gupta. PALS: An or-parallel implementation of Prolog on Beowulf architectures. Lecture Notes in Computer Science, 2237:27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2237/22370027.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2237/22370027.
pdf.

VanDriessche:1994:DLB

[VR94] R. Van Driessche and D. Roose. Dynamic load balancing with an improved spectral bisection algorithm. In IEEE [IEE94c], pages 494–500. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Vroom:1994:WIS

[Vro94] R. W. Vroom. What is the information that should be generated by the engineering departments of automotive supplier companies? In Anonymous [Ano94-75], pages 413– 420. ISBN 0-947719-68-7. LCCN ????

VanCamp:1993:UFS

[VRSG93] M. Van Camp, D. Ruan, A. Sohier, and P. Govaerts. The use of the fuzzy set theory to reduce uncertainties for nuclear emergency decision aiding systems. In Kusters et al. [KSW93], pages 432–442. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Vetter:1999:THP

[VS99] Jeffrey Vetter and Karsten Schwan. Techniques for high-performance computational steering. IEEE Concurrency, 7(4):63–74, October/December 1999. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/
books/pd1999/pdf/p4063.
pdf;   http://www.computer.
org/concurrency/pd1999/
p4063abs.htm.

VanVoorst:1994:PCW

[VSB94] B. Van Voorst, S. Seidel, and E. Barszcz. Profiling the communication workload of an iPSC/860. In IEEE [IEE94c], pages 221–228. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Vermaas:2021:SPS

[VSB^{+}21] J. V. Vermaas, A. Sedova, M. B. Baker, S. Boehm, D. M. Rogers, J. Larkin, J. Glaser, M. D. Smith, O. Hernandez, and J. C. Smith. Supercomputing pipelines search for therapeutics against COVID-19. Computing in Science and Engineering, 23(1):7–16, 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vajapeyam:1990:EOP

[VSH90] Sriram Vajapeyam, Gurindar S. Sohi, and Wei-Chung Hsu. Exploitation of operationlevel parallelism in a processor of the Cray X-MP. Proceedings — IEEE International Conference on Computer Design: VLSI in Computers and Processors, pages 20–23, September 1990. CO-DEN PIIPE6. ISBN O-8186-2079-X. IEEE catalog number 90CH2909-0.

Vajapeyam:1991:ESC

[VSH91] Sriram Vajapeyam, Gurindar S. Sohi, and Wei-Chung Hsu. An empirical study of the CRAY Y-MP processor using the Perfect Club benchmarks. ACM SIGARCH Computer Architecture News, 19(3):170–179, May 1991. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE). IEEE catalog number 91CH2995-9.

VanderWijngaart:1996:EIS

[VSM96] R. F. Van der Wijngaart, S. R. Sarukkai, and P. Mehra. The effect of interrupts on software pipeline execution on message-passing architectures. In ACM [ACM96], pages 189–196. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Venkateswaran:2007:FGSa

[VSM^{+}07a] N. Venkateswaran, Deepak Srinivasan, Madhavan Manivannan, T. P. Ramnath Sai Sagar, Shyamsundar Gopalakrishnan, VinothKrishnan Elangovan, Karthik Chandrasekar, Prem Kumar Ramesh, Viswanath Venkatesan, Arvindakshan Babu, and Sudharshan. Future generation supercomputers I: a paradigm for node architecture. ACM SIGARCH Computer Architecture News, 35(5):49–60, December 2007. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Venkateswaran:2007:FGSb

[VSM^{+}07b] N. Venkateswaran, Deepak Srinivasan, Madhavan Manivannan, T. P. Ramnath Sai Sagar, Shyamsundar Gopalakrishnan, VinothKrishnan Elangovan, Arvind M., Prem Kumar Ramesh, Karthik Ganesan, Viswanath Krishnamurthy, and Sivaramakrishnan. Future generation supercomputers II: a paradigm for cluster architecture. ACM SIGARCH Computer Architecture News, 35(5):61–70, December 2007. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Vetter:1994:MWW

[VSW94] Ronald J. Vetter, Chris Spell, and Charles Ward. Mosaic and the World Wide Web. Computer, 27(10): 49–57, October 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Villa:1995:SPA

[VT95] A. E. P. Villa and I. V. Tetko. Spatio-temporal patterns of activity controlled by system parameters in a simulated “Thalamo-Cortical” neural network. In Herrmann et al. [HWP95], pages 379– 388. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Villa:2012:FAS

[VTSM12] Oreste Villa, Antonino Tumeo, Simone Secchi, and Joseph B. Manzano. Fast and accurate simulation of the Cray XMT multithreaded supercomputer. IEEE Transactions on Parallel and Distributed Systems, 23(12): 2266–2279, December 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Voinovich:1998:UAS

[VTTS98] P. Voinovich, E. Timofeev, K. Takayama, and T. Saito. 3-D unstructured adaptive supercomputing for transient problems of volcanic blast waves. In Anonymous [Ano98a], page ALL.

Vuik:1993:SDI

[Vui93]  C. Vuik.  93SC016 the

solution of the discretized incomprehensible Navier– Stokes equations with iterative methods. In Anonymous [Ano93-31], pages 69– 76. ISBN 0-947719-62-8. LCCN ????

Vujic:1993:GAT

[Vuj93]  J. L. Vujic.  GTRAN2:

An advanced transport theory code for advanced assembly calculations. In Kusters et al. [KSW93], pages 695– 706. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Verma:1994:EAT

[VV94] R. M. Verma and P. J. Varman. Efficient archivable time index: a dynamic indexing scheme for temporal data. In Balakrishnan [Bal94], pages 59–72. ISBN 0-07-462044-4. LCCN ????

Vandoni:1995:CSC

[VV95] C. E. Vandoni and C. Verkerk, editors. 1994 CERN School of Computing: Sopron, Hungary, 28 August–10 September 1994: proceedings, number 95-01 in Proceedings — CERN School of Computing 1995; 17th. CERN, Geneva, Switzerland, 1995. ISBN 92-9083-069-7. ISSN 0304-2898. LCCN QC770 .E83 v.95, no.1.

Vogels:1995:NSC

[VVH95] M. E. S. Vogels, H. Van der Ven, and G. J. Hameetman. The need for super computers in aerospace research and industry. Lecture Notes in Computer Science, 919:448–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDrunen:1996:APS

[VVKB96] R. Van Drunen, C. Van Teylingen, M. Kroontje, and H. J. C. Berendsen. The Amfisbaena: a parallel supercomputer system based on i860 as a generic platform for molecular dynamics simulations. In Arabnia [Ara96], pages 637–644. ISBN 0-9648666-4-1, 0-9648666-2-5. LCCN ????

VanEngelen:1995:CPP

[VW95] R. Van Engelen and L. Wolters. A comparison of parallel programming paradigms and data distributions for a limited area numerical weather forecast routine. In ACM [ACM95a], pages 357–364. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

VanEngelen:1996:CGM

[VWC96] R. VanEngelen, L. Wolters, and G. Cats. CTADEL: a generator of multi-platform high performance codes for PDE-Based scientific applications. In ACM [ACM96], pages 86–93. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Vu:1988:CTS

[VY88] D. Vu and C. Yang. Comparing tridiagonal solvers on the CRAY X-MP/416 system. CRAY Channels, 9 (4):22–25, 1988. CODEN CRCHE8.

Wang:2012:CCM

[W^{+}12] Lizhe Wang et al., editors.

Cloud computing: methodology, systems, and applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-5641-9 (hardcover). xxxvii + 805 pp. LCCN QA76.585 .C575 2012.

Wazlowski:2005:VSB

[WAB^{+}05] M. E. Wazlowski, N. R.

Adiga, D. K. Beece, R. Bellofatto, M. A. Blumrich, D. Chen, M. B. Dombrowa, A. Gara, M. E. Giampapa, R. A. Haring, P. Heidelberger, D. Hoenicke, B. J. Nathanson, M. Ohmacht, R. Sharrar, S. Singh, B. D. Steinmacher-Burow, R. B. Tremaine, M. Tsao, A. R. Umamaheshwaran, and P. Vranas. Verification strategy for the Blue Gene/L chip. IBM Journal of Research and Development, 49(2/3):303–318, ???? 2005. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

                                        //www.research.ibm.com/
                                        journal/rd/492/wazlowski.
                                        pdf.

Wheeler:1997:NCA

[WABD97] M. F. Wheeler, T. Arbogast, S. Bryant, and C. N. Dawson. New computational approaches for chemically reactive transport in porous media. In Delic and Wheeler [DW97], pages 217– 226. ISBN 0-89871-378-1. LCCN ????

Wacker:1992:EH

[Wac92]  H.-M. Wacker.  Die eu-

ropaeische HPC-initiative. In Meuer [Meu92c], pages 233–?? ISBN 0-387-55709-1 (paperback), 3-540-55709-1 (Germany). LCCN QA76.88.S858 1992. German and English.

Wade:1986:DSS

[Wad86] Mitch Wade. Digital scene simulation at digital productions, 1986. 1 videocassette (50 min.).

Goddard:1989:ISC

[WAD^{+}89a] William D. Wilson, Robert J. Asaro, Robert W. Dutton, Juan M. Sanchez, David J. Srolovitz, Richard H. Boyd III, William A. Goddard, John R. Smith, and Wilhelm G. Wolfer. The impact of supercomputing capabilities on U.S. materials science and technology. Future Generation Computer Systems, 5(2–3):283–293, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wilson:1989:ISC

[WAD^{+}89b] William D. Wilson, Robert J.

Asaro, Robert W. Dutton, Juan M. Sanchez, David J. Srolovitz, Richard H. Boyd III, William A. Goddard, John R. Smith, and Wilhelm G. Wolfer. The impact of supercomputing capabilities on U.S. materials science and technology. Future Generation Computer Systems, 5(2–3):283–293, September 1989. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Wagner:1996:TSU

[Wag96] P. Wagner. Traffic simulations using cellular automata: Comparison with reality. In Wolf et al. [WSB96], pages 199–204. ISBN 981-02-2635-7. LCCN ????

Wait:2005:IPF

[Wai05] C. D. Wait. IBM PowerPC 440 FPU with complexarithmetic extensions. IBM Journal of Research and Development, 49(2/3):249–254, ???? 2005. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL

                                        http://www.research.ibm.
                                        com/journal/rd/492/wait.
                                        pdf.

Walker:1981:JEA

[Wal81] Grady M. Walker. Is the Josephson effect the answer to the supercomputer? Thesis (B.S.), East Texas State University, Commerce, TX, USA, 1981. ii + 65 pp.

Wallich:1985:MMS

[Wal85]  P. Wallich.  Minis and

mainframes: Superminicomputers push mainframe performance, mainframes operate at supercomputer speeds, and supercomputers reach 400 million operations per second. IEEE Spectrum, 22 (1):42–44, January 1985. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallace:1990:ST

[Wal90] D. J. Wallace. Supercomputing with transputers. Computing systems in engineering: an international journal, 1(1):131–142, ???? 1990. CO-DEN COSEEO. ISSN 0956-0521.

Wallace:1992:LSU

[Wal92] D. R. Wallace. Low-level scheduling using the hierarchical task graph. In ACM [ACM92b], pages 72– 81. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Wall:1995:RS

[Wal95] Alison Wall. A review of supercomputing 1991–1994. Engineering and Physical Sciences Research Council, Swindon, England, March 1995. ISBN 1-899371-20-6 (??invalid ISBN checksum??). ?? pp. LCCN ????

Wallich:2009:SGP

[Wal09] P. Wallich. The supercomputer goes personal. IEEE Spectrum, 46(4):64, April 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Waltz:2017:BMB

[Wal17] E. Waltz. Biocomputer and memory built inside living bacteria [news]. IEEE Spectrum, 54(9):11–12, September 2017. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weeks:2001:MCM

[WAM^{+}01] Michael Weeks, Jim Austin, Anthony Moulds, Aaron Turner, Zygmunt Ulanowski, and Julian Young. Mapping correlation matrix memory applications onto a Beowulf cluster. Lecture Notes in Computer Science, 2130:156– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2130/21300156.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2130/21300156.
pdf.

Warter:1989:EMC

[War89] Nancy Jeanne Warter. An environment monitor for the Cedar supercomputer. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1989. v + 84 pp.

Wareing:1993:NDA

[War93a]  T. A. Wareing.  New

diffusion-synthetic acceleration methods for the SN equations with corner balance spatial differencing. In Kusters et al. [KSW93], pages 500–511. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Warnatz:1993:DCH

[War93b] J. Warnatz. Detailed chemistry of hydrocarbon combustion and its coupling with flow processes. In Anonymous [Ano93-31], pages 95– 102. ISBN 0-947719-62-8. LCCN ????

Warren:2000:BG

[War00] Henry S. Warren, Jr. Blue Gene. Lecture Notes in Computer Science, 1940:32–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1940/19400032.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1940/19400032.
pdf.

Ward:2003:CKA

[War03] Bob Ward. Cray, Kanai awards recognize achievement in high-performance computing and distributed systems: [Monty Denneau, IBM Blue Gene architect; Steven S. Yao; Alfred Z. Spector]. Computer, 36(6): 89–91, June 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2003/06/r6089.htm;   http:
//csdl.computer.org/dl/
mags/co/2003/06/r6089.pdf.

Ward:2009:MWC

[War09] Bob Ward. Miura wins Cray Award. Computer, 42(12): 71–73, December 2009. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:2010:BGA

[War10] Bob Ward. Blue Gene architect Gara receives 2010 Cray Award. Computer, 43(11): 78–80, November 2010. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Washington:1996:DEM

[Was96a] David W. P. E. Washington. Discrete element modeling of dry granular material using a massively parallel supercomputer. Thesis (Ph.D.), New Jersey Institute of Technology, Department of Civil and Environmental Science, 1996. xiii, 133 pp.

Wasserman:1996:BTD

[Was96b]  H. Wasserman.  Bench-

mark tests on the Digital Equipment Corporation Alpha AXP 21164-Based AlphaServer 8400. In ACM [ACM96], pages 333–340. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Watson:1972:TAH

[Wat72] W. J. Watson. The TI ASC — a highly modular and flexible super processor architecture. In Proc. AFIPS Fall Joint Computer Conf., pages 221–228. ????, ????, 1972. ISBN ???? LCCN ????

Watanabe:1987:APN

[Wat87]  T. Watanabe.  Architec-

ture and performance of NEC supercomputer SX system. Parallel Computing, 5(1–2): 247–255, July 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Watanabe:1991:HNN

[Wat91]  T. Watanabe.  HIDM,

a new numerical scheme to solve partial differential equations. In Anonymous [Ano91q], pages 229–235. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Watson:1992:SIP

[Wat92a] G. F. Watson. Supercomputers — interconnections and packaging. IEEE Spectrum, 29(9):69–71, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watson:1992:PNS

[Wat92b] George F. Watson. Packaging the new semiconductors will also demand cutting-edge technologies. IEEE Spectrum, 29(9):69–71, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watts:1993:FGC

[Wat93]  Howard R. Watts.  Fu-

ture of GaAs in the Cray-3 and Cray-4 supercomputers. Technical Digest — GaAs IC Symposium (Gallium Arsenide Integrated Circuit), pages 137–139, 1993.

CODEN TDGSEE. ISBN 0-7803-1393-3. IEEE catalog number 93CH3346-4.

Watermann:1995:BKN

[Wat95]  A. Watermann.  Berech-

nung von Konvektionsstroemungen nach der Methode der Finiten Elemente auf dem massiv parallelen Supercomputer Intel/Paragon. Zeitschrift für Angewandte Mathematik und Mechanik, 75(??):S511, ???? 1995. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Wayner:1996:INN

[Way96] Peter Wayner. Inside the NC: Are network computers just stripped-down terminals? No way. The official NC platform covers everything from a set-top box to a Cray. BYTE Magazine, 21(11):105, 106, 108, 110, November 1996. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic). Discusses the importance of Java in a Network Computer.

Waz:1989:PFS

[Waz89] Joseph W. Waz Jr. Proposed federal supercomputer network highlights policy issues in telecommunications, copyright. The Computer Lawyer, 6(10):27–??, October 1, 1989. CODEN COLAEB. ISSN 0742-1192.

Williams:1985:SPL

[WB85] Elizabeth Williams and Frank Bobrowicz. Speedup predictions for large scientific parallel programs on Cray X-MP-like architectures. Proceedings of the International Conference on Parallel Processing, pages 541–543, 1985. CODEN PCPADL. ISBN 0-8186-0637-1. ISSN 0190-3918. IEEE Service Cent. Piscataway, NJ, USA.

Wait:1988:OBM

[WB88] R. Wait and N. G. Brown.

Overlapping block methods for solving tridiagonal systems on transputer arrays. Parallel Computing, 8(1-3): 325–333, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the International Conference on Vector and Parallel Processors in Computational Science, III (Liverpool, 1987).

Wallqvist:1987:EPP

[WBP87] A. Wallqvist, B. Berne, and C. Pangali. Exploiting physical parallelism using supercomputers: Two examples from chemical physics. Computer, 20(5):9–21, 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolters:1993:PIH

[WC93] L. Wolters and G. Cats. A parallel implementation of the HIRLAM model. In Hoffmann and Kauranne [HK93b], pages 486–499. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Wolters:1994:LAN

[WCG94] L. Wolters, G. Cats, and N. Gustafsson. Limited area numerical weather forecasting on a massively parallel computer. In Anonymous [Ano94-134], pages 289–296. ISBN ???? LCCN ????

Wu:1991:QCR

[WCHK91] Y. S. Mark Wu, S. A. Cuccaro, P. G. Hipes, and A. Kuppermann. Quantum chemical reaction dynamics on a highly parallel supercomputer. Theoretica Chimica Acta, 79(3):225–240, ???? 1991. CODEN TCHAAM. ISSN 0040-5744.

Wang:2018:IMI

[WCZ^{+}18] Zihao Wang, Yu Chen, Jingrong Zhang, Lun Li, Xiaohua Wan, Zhiyong Liu, Fei Sun, and Fa Zhang. ICON-MIC: Implementing a CPU/MIC collaboration parallel framework for ICON on Tianhe-2 supercomputer. Journal of Computational Biology, 25(3):270–281, March 2018. CODEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (electronic). URL

https://www.liebertpub.
com/doi/abs/10.1089/cmb.
2017.0151;   https://www.

liebertpub.com/doi/pdf/
10.1089/cmb.2017.0151.

Waddell:1993:MTT

[WD93a] M. W. Waddell and H. L.

Dodds. A method for transient, three-dimensional neutron transport calculations. In Kusters et al. [KSW93], pages 633–645. ISBN 3-923704-11-9. LCCN ???? Two volumes.

White:1993:SHB

[WD93b] J. R. White and P. M. Delmolino. Simple heuristics: a bridge between manual core design and automated optimization methods. In Kusters et al. [KSW93], pages 210– 221. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Womack:1994:PAC

[WD94] Lucas A. Womack and Richard N. Draper. A parallel approach to CRC encoding. Technical report SRC-TR-94-137, Supercomputing Research Center: IDA, Lanham, MD, USA, December 30, 1994. 5 pp.

Wearring:1997:VPV

[Wea97] C. Wearring. Virtual prototyping to validate functionality of product and process designs in the presence of dimensional variation. In Roller [Rol97], pages 85–94. ISBN 0-947719-88-1 (paperback). LCCN ????

Weber:1991:NSF

[Web91] T. A. Weber. The National Science Foundation Supercomputer Centers Program. International Journal of Supercomputer Applications, 5 (4):3, Winter 1991. CODEN IJSAE9. ISSN 0890-2720.

Weber:1993:CFD

[Web93] P. Weber. Computational fluid dynamics using FIDAP on a SNI S600/ 20 vector processor and a workstation. In Kusters et al. [KSW93], pages 787–?? ISBN 3-923704-11-9. LCCN ???? Two volumes.

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4): 348–361, October/December 1988. CODEN AHCOE5. ISSN 0164-1239. URL http:

                                        //dlib.computer.org/an/
                                        books/an1988/pdf/a4348.
                                        pdf;   http://www.computer.
                                        org/annals/an1988/a4348abs.
                                        htm.

Weiss:1989:SSA

[Wei89] S. Weiss. Scalar supercomputer architecture. Proceedings of the IEEE, 77(12): 1970–??, December 1, 1989. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Weisz:1990:FME

[Wei90] Juraj Weisz. S. F. Mccormick (Ed.): Multigrid Methods, Theory Applications, and Supercomputing. Acta Applicandae Mathematicae, 18(2): 183–184, February 1990. CO-DEN AAMADV. ISSN 0167-8019.

Weiss:1991:FDP

[Wei91] Ray Weiss. 32-bit floatingpoint DSP processors. EDN, 36(23):126–??, November 7, 1991. CODEN EDNSBH. ISSN 0012-7515, 0364-6637.

Weiss:1992:TRP

[Wei92] Ray Weiss. Third-generation RISC processors. EDN, 37 (7):96–??, March 30, 1992. CODEN EDNSBH. ISSN 0012-7515, 0364-6637.

Wenes:1994:BRB

[Wen94] Geert Wenes. Book review:

Projects in Scientific Computation, by Richard E. Crandall and A Scientist’s and Engineer’s Guide to Workstations and Supercomputers, by Rubin H. Landau and Paul J. Fink, Jr. Computers in Physics, 8(1):68– ??, ???? 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Westman:1989:ACX

[Wes89] James Allen Westman. Applications of a Cray X-MP supercomputer to digital image classification. Thesis (M.S.), University of Wisconsin — Madison, Madison, WI, USA, 1989. x + 262 pp.

Westropp:1996:SRP

[Wes96] John C. Westropp. Site report: Pittsburgh Supercomputing Center. IEEE Computational Science & Engineering, 3(1):8–12, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Wolski:1993:PPN

[WF93] Richard M. Wolski and John T. Feo. Program partitioning for NUMA multiprocessor computer systems. Journal of Parallel and Distributed Computing, 19(3): 203–??, November 1993. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Worley:1994:PST

[WF94] P. H. Worley and I. T. Foster. Parallel spectral transform shallow water model: a runtime-tunable parallel benchmark code. In IEEE [IEE94c], pages 207–214. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Wilkinson:2008:TTA

[WF08] Barry Wilkinson and Clayton Ferner. Towards a topdown approach to teaching an undergraduate grid computing course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):126–130, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Wang:2017:ISV

[WFJ^{+}17] Jianfei Wang, Fengfeng Fan, Li Jiang, Xiaoyao Liang, and Naifeng Jing. Incorporating selective victim cache into GPGPU for high-performance computing. Concurrency and Computation: Practice and Experience, 29(24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weinberg:1993:TDT

[WFT93] D. Weinberg, H.-H. Frey, and H. Tschoeke. A three dimensional transient calculation for the reactor model RA-MONA using the COMMIX-2(V) code. In Kusters et al. [KSW93], pages 292– 306. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Wallis:1982:PRS

[WG82] John R. Wallis and J. Rodney Grisham. Petroleum reservoir simulation on the Cray-1 and on the FPS-164. In Vichnevetsky et al. [VAS82], pages 308–310. LCCN TA343 .I523 1982. Five volumes.

Wong:1991:FHA

[WG91] W. F. Wong and E. Goto.

Fast hardware-based algorithms for elementary function computations. In Anonymous [Ano91q], pages 56–65. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Woerner:1993:CTM

[WG93a] M. Woerner and G. Groetzbach. Contributions to turbulence modelling of natural convection in liquid metals by direct numerical simulation. In Kusters et al. [KSW93], pages 224–235. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Wolcott:1993:IPU

[WG93b] P. Wolcott and S. E. Goodman. International perspectives: Under the stress of reform: High-performance computing in the former Soviet Union. Communications of the ACM, 36 (10):25–29, October 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        164063.html.

Weeks:1994:CUT

[WG94] N. J. Weeks and P. A. Galwas. Computervision universal translator and transfer system. In Anonymous [Ano94-75], pages 405–412. ISBN 0-947719-68-7. LCCN ????

Wong:1991:SBP

[WGOY91] W. F. Wong, E. Goto, Y. Oyanagi, and N. Yoshida.

Six benchmark problems for number crunchers. In Anonymous [Ano91q], pages 120– 125. ISBN 4-87378-284-8. LCCN QA76.88.I1991.

Wolfer:1993:IIS

[WGR93] J. Wolfer, T. Grace, and J. Roberge. An investigation of image synthesis software migration to the BBN TC2000 Butterfly II supercomputer. In Anonymous [Ano93a], pages 23–34. ISBN ???? LCCN ????

Wen:1991:PEW

[WGS91] Yen-Cheng Wen, Kyle A.

Gallivan, and Resve A. Saleh. Parallel event-driven waveform relaxation. Technical Report CSRD 1145, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1991. 5 pp.

Woo:2004:AAJ

[WGW04] Jongwook Woo, Jean-Luc L.

Gaudiot, and Andrew L. Wendelborn. Alias analysis in Java with referenceset representation for highperformance computing. International Journal of Parallel Programming, 32(1):39– 76, February 2004. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&

                                        issn=0885-7458&volume=32&
                                        issue=1&spage=39.

Wait:1993:NSH

[WH93] R. Wait and T. J. Harding. Numerical software for 3D hydrodynamic modelling using transputer arrays. In Hoffmann and Kauranne [HK93b], pages 453– 464. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Wang:1994:USM

[WH94] Z. P. Wang and D. R. Hayhurst. The use of supercomputer modelling of hightemperature failure in pipe weldments to optimize weld and heat affected zone materials property selection. Proceedings of the Royal Society of London. Ser, 446(1926): 127–??, July 8, 1994. ISSN 0080-4630.

Wang:1993:VPP

[WHBH93] Z. P. Wang, D. R. Hayhurst, B. A. Bilby, and J. C. Howard. Vectorization and parallel processing studies using a Cray X-MP in non-linear computational solid mechanics. Engineering Computations, 10(5):387– 395, October 1993. CODEN ENCOEN. ISSN 0264-4401.

Wheat:1983:KBM

[Whe83] Stephen Randolph Wheat. A kosloff/baysal method, 3D migration program implemented on the CYBER 205 supercomputer. Thesis (M.S.), Dept. of Computer Science, University of Houston-University Park, Houston, TX, USA, 1983. vii + 78 pp.

Wheeler:1989:DER

[Whe89] Ferrell S. Wheeler. The distribution of epacts in a random mapping and the effects of parallelism. Technical report SRC-TR-92-056, Supercomputing Research Center: IDA, Lanham, MD, USA, February 5, 1989. 20 pp.

Wu:1993:QEC

[WHL93] C. E. Wu, Y. Hsu, and Y.- H. Liu. A quantitative evaluation of cache types for high-performance computer systems. IEEE Transactions on Computers, 42(10): 1154–1162, October 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=257701.

Wu:1997:BEH

[WHMA97] T. S. Wu, J. M. Hamrick, S. C. McCutcheon, and R. B. Ambrose. Benchmarking the EFDC/HEM3D surface water hydrodynamic and eutrophication models. In Delic and Wheeler [DW97], pages 157–162. ISBN 0-89871-378-1. LCCN ????

Wholey:1992:ADM

[Who92]  S. Wholey.  Automatic

data mapping for distributedmemory parallel computers. In ACM [ACM92b], pages 25–34. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Wichmann:1992:SFW

[Wic92]  B. A. Wichmann.  Sur-

veyor’s Forum: “What every computer scientist should know about floating-point arithmetic”. ACM Computing Surveys, 24(3):319, September 1992. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). See [Gol91b, Gol91a, Dun92].

Wickman:1996:IOI

[Wic96] R. W. Wickman. Implementation of optical interconnects in the GigaRing supercomputer channel. In Chen and Guilfoyle [CG96], pages 343–365. ISBN 0-8194-2017-4. LCCN ????

Wiederhold:1987:FOD

[Wie87] Gio Wiederhold. File organization for database design. McGraw-Hill computer science series; McGraw-Hill series in computer organization and architecture; McGraw-Hill series in supercomputing and artificial intelligence; McGraw-Hill series in artificial intelligence. McGraw-Hill, New York, NY, USA, 1987. ISBN 0-07-070133-4. xiv + 619 pp. LCCN QA76.9.F5 W53 1987.

Wienke:1994:BDP

[Wie94] B. R. Wienke. Basic diving physics and applications. Diversification series. Best Pub. Co, Flagstaff, AZ, USA, 1994. ISBN 0-941332-41-1. 320 pp. LCCN VM981 .W54 1994. Wiedemann:1996:ERM

[Wie96] R. Wiedemann. Empirical research for modelling driving behaviour on motorways. In Wolf et al. [WSB96], pages 167–168. ISBN 981-02-2635-7. LCCN ????

Wijshoff:1989:ISB

[Wij89a] Harry A. G. Wijshoff. Implementing sparse BLAS primitives on concurrent/ vector processors: a case study. Technical Report CSRD 843, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1989. 51 pp.

Wijshoff:1989:SOU

[Wij89b] Harry A. G. Wijshoff. Symmetric orderings for unsymmetric sparse matrices. Technical Report CSRD 901, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1989. 14 pp.

Williams:1988:MTS

[Wil88a] Elizabeth A. Williams. Measurement of two scientific workloads using the Cray X-MP performance monitor. Technical report SRC-TR-88-020, Supercomputing Research Center: IDA, Lanham, MD, USA, 1988. 22 pp.

Wilson:1988:ISC

[Wil88b] William D. Wilson. The impact of supercomputing capabilities on U.S. materials science and technology: report of the. Technical Report NMAB 451, [Available from NTIS], Washington, DC, USA, 1988. vii + 65 pp.

Williams:1990:CFD

[Wil90a] B. R. Williams. Computational fluid dynamics and the CRAY-2 at RAE. In Pitcher [Pit90], pages 533– 541. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Williams:1990:IID

[Wil90b]  Peter L. Williams.  Is-

sues in interactive direct projection volume rendering of nonrect[i]linear meshed data sets. Technical Report CSRD 1059, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1990. 13 pp.

Williams:1991:ACG

[Wil91] Peter L. Williams. Applications of computational geometry to volume visualization. Technical Report CSRD 1117, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1991. 5 pp.

Williams:1992:ISN

[Wil92a] Peter L. Williams. Interactive splatting of nonrectilinear volumes. Technical Report CSRD 1217, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1992. 8 pp.

Williams:1992:VOM

[Wil92b] Peter L. Williams. Visibility ordering meshed polyhedra. Technical Report CSRD 1097, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. 28 pp.

Williams:1992:VOMa

[Wil92c] Peter L. Williams. Visibility ordering meshed polyhedra. Technical Report CSRD 1097, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1992. 28 pp.

Wiley:1993:UPP

[Wil93] R. L. Wiley. Using parallel processors for production — the Met. Office’s view. In Hoffmann and Kauranne [HK93b], pages 482– 485. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Wilkinson:1994:ECN

[Wil94] E. Wilkinson. Estimation of component noise from vehicle passby tests. In Anonymous [Ano94-75], pages 471– 478. ISBN 0-947719-68-7. LCCN ????

Wilson:1995:PPP

[Wil95] Gregory V. Wilson. Practical Parallel Programming. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, 1995. ISBN 0-262-23186-7. viii + 564 pp. LCCN QA76.642.W553 1995. US$55.00. URL http:

                                        //www.mitpress.com/book-
                                        home.tcl?isbn=0262231867.

Wilson:1996:BOH

[Wil96] Gregory V. Wilson. Bookshelf: An overview of highperformance computing for the uninitiated. IEEE Software, 13(2):118, March 1996. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Wilkinson:2010:GCT

[Wil10] Barry Wilkinson. Grid computing: techniques and applications. Chapman and Hall/CRC computational science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4200-6953-5 (hardcover). xxi + 365 pp. LCCN QA76.9.C58 W55 2010.

Winslett:2002:DDS

[Win02] Marianne Winslett. David DeWitt speaks out: on rethinking the CS curriculum, why the database community should be proud, why query optimization doesn’t work, how supercomputing funding is sometimes very poorly spent, how he’s not a good coder and isn’t smart enough to do DB theory, and more. SIGMOD Record (ACM Special Interest Group on Management of Data), 31 (2):50–62, June 2002. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Witten:1989:PIL

[Wit89] Matthew Witten. Peering inside living systems: physiology in a supercomputer, volume 587 of IMA preprint series. Institute for Mathematics and Its Application, University of Minnesota, Minneapolis, MN, USA, October 1989. 10 pp.

Wren:1994:LST

[WJ94] C. S. Wren and O. Johnson. Latest simulation techniques applied to intake and exhaust system design. In Anonymous [Ano94-75], pages 589– 600. ISBN 0-947719-68-7. LCCN ????

Wang:2009:GCI

[WJC09] Lizhe Wang, Wei Jie, and Jinjun Chen, editors. Grid computing: infrastructure, service, and applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-4200-6766-4 (hardcover). xvi + 512 pp. LCCN QA76.9.C58 G694 2009.

Wang:1995:NBA

[WK95] C.-M. Wang and C.-Y. Ku. A near-optimal broadcasting algorithm in all-port wormholerouted hypercubes. In ACM [ACM95a], pages 147–153. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Wesche:1997:HIV

[WKFFK97] G. Wesche, O. Kraemer-Fuhrmann, B. Froehlich, and F. R. Klimetzek. Highly interactive visualization of fluid dynamics in a virtual environment. In Roller [Rol97], pages 495–502. ISBN 0-947719-88-1 (paperback). LCCN ????

Wu:1997:CRE

[WKHS97] Xudong T. Wu, Prakashan P.

Korambath, Edward F. Hayes, and Danny C. Sorensen. Computation of rovibrational eigenvalues of van der Waals molecules on a CRAY T3D. Journal of Computational Physics, 138 (2):286–301, December 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0021999197958198.

Wang:2016:EDT

[WKL^{+}16] Ke Wang, Abhishek Kulkarni, Michael Lang, Dorian Arnold, and Ioan Raicu. Exploring the design tradeoffs for extreme-scale highperformance computing system software. IEEE Transactions on Parallel and Distributed Systems, 27(4):1070– 1084, April 2016. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://www.
computer.org/csdl/trans/
td/2016/04/07103354-abs.
html.

Woo:1983:BSE

[WL83] P. T. Woo and John M.

Levesque. Benchmarking a sparse elimination routine on the Cyber 205 and the Cray 1-S. Society of Petroleum Engineers journal, 23(5):743– 745, October 1983. CODEN SSPJDN. ISSN 0197-7520.

Walstrom:1994:MSC

[WL94] Tom Walstrom and Elliot Long. Minnesota supercomputer center. Technical report, Program Evaluation Division, Office of the Legislative Auditor State of Minnesota, Saint Paul, MN, USA, June 1994. xii + 24 + 2 pp.

Wilson:1996:PPU

[WL96] Gregory V. Wilson and Paul Lu, editors. Parallel Programming Using C++. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-73118-5. xxxiv + 758 pp. LCCN QA76.73.C153P365 1996. US$49.50. URL http://www.

                                        mitpress.com/book-home.
                                        tcl?isbn=0262731185. Fore-

word by Bjarne Stroustrup. Describes fifteen parallel programming systems based on C++.

Wang:2002:MSP

[WLCG02] P. Wang, Karen Y. Liu, Tom Cwik, and Robert Green. MODTRAN on supercomputers and parallel computers. Parallel Computing, 28 (1):53–64, January 2002. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
35/21/60/27/30/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/60/27/
30/00001685.pdf.

Wang:2000:ICV

[WLH00] K.-Y. Wang, D. J. Lary, and S. M. Hall. Improvement of a 3D CTM and a 4D variational data assimilation on a vector machine CRAY J90 through a multitasking strategy. Computer Physics Communications, 125(1–3): 142–153, March 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465599004695.

Wang:1995:IMA

[WLKI95] J. Wang, H. Lung, Y. Katsumata, and T. Ishigai. Implementing a 3D multigrid algorithm on Fujitsu’s vector parallel supercomputer. In Mirenkov et al. [M^{+}95], pages 107–113. ISBN 0-8186-7038-X. LCCN QA76.642.A43 1995.

Wei:2020:IDA

[WLLZ20] Jinxia Wei, Chun Long, Jiawei Li, and Jing Zhao. An intrusion detection algorithm based on bag representation with ensemble support vector machine in cloud computing. Concurrency and Computation: Practice and Experience, 32(24):e5922:1– e5922:??, December 25, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wimberly:1996:PTA

[WLN^{+}96a] Frank C. Wimberly, Michael H. Lambert, Nicholas A. Nystrom, Alex Ropelewski, and William Young. Porting third-party applications packages to the Cray T3D: Programming issues and scalability results. Parallel Computing, 22(8):1073–1089, October 30, 1996. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wimberly:1996:PTP

[WLN^{+}96b] Frank C. Wimberly, Michael H. Lambert, Nicholas A. Nystrom, Alex Ropelewski, and William Young. Porting third-party applications packages to the Cray T3D: Programming issues and scalability results. Parallel Computing, 22(8):1073– 1089, October 28, 1996. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        parco/cas_sub/browse/browse.
                                        cgi?year=1996&volume=22&
                                        issue=8&aid=1088.

Wheeler:1991:SIC

[WM91] Daniel D. Wheeler and Robert A. J. Matthews. Supercomputer investigations of a chaotic encryption algorithm. Cryptologia, 15 (2):140–152, April 1, 1991. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). URL http:

//www.informaworld.com/
smpp/content~content=a741902749~
db=all~order=page. chaotic

encryption algorithm; nonlinear pseudo-random number generator; chaos theory; cycling keys; low-precision arithmetic; numerical investigation; Cray Y-MP machine; cycling problem.

Williams:1992:VDO

[WM92] Peter L. Williams and Nelson Max. A volume density optical model. Technical Report CSRD 1216, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1992. 8 pp.

Wehner:1997:CSM

[WMBC97] M. F. Wehner, A. A. Mirin, J. H. Bolstad, and U. E. Creach. Climate systems modeling on massively parallel processing computers at Lawrence Livermore National Laboratory. In Delic and Wheeler [DW97], pages 21–

30. ISBN 0-89871-378-1. LCCN ????

Williams:1990:CTS

[WMK90] Elizabeth Williams, C. Thomas Myers, and Rebecca Koskela. The characterization of two scientific workloads using the Cray X-MP performance monitor. IEEE, Piscataway, NJ, USA, 1990. ISBN 0-8186-2056-0. 142–152 pp. LCCN ???? IEEE catalog number 90CH2916-5.

Wan:1996:BSI

[WMKS96] M. Wan, R. Moore, G. Kremenek, and K. Steube. A batch scheduler for the Intel Paragon MPP system with a noncontiguous node allocation algorithm. Lecture Notes in Computer Science, 1162:48– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woh:2010:MSN

[WMMC10] Mark Woh, Scott Mahlke, Trevor Mudge, and Chaitali Chakrabarti. Mobile supercomputers for the nextgeneration cell phone. Computer, 43(1):81–85, January 2010. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wanschura:1996:EAS

[WMR96] T. Wanschura, S. Migowsky, and P. Rujan. Effect of adaptive strategies on a simple model of freeway traffic. In Wolf et al. [WSB96], pages 211–216. ISBN 981-02-2635-7. LCCN ????

Wang:1992:SBL

[WN92] H. Wang and A. Nicolau.

Speedup of banded linear recurrences in the presence of resource constraints. In ACM [ACM92b], pages 466–477. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Wendykier:2010:PCH

[WN10] Piotr Wendykier and James G.

Nagy. Parallel Colt: a high-performance Java library for scientific computing and image processing. ACM Transactions on Mathematical Software, 37(3):31:1– 31:22, September 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Wang:1996:CPC

[WNKS96] Haigeng Wang, Alexandru Nicolau, Stephen Keung, and Kai-Yeung (Sunny) Siu. Computing programs containing band linear recurrences on vector supercomputers. IEEE Transactions on Parallel and Distributed Systems, 7(8):769–782 (or 769–781??), August 1996. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//www.computer.org/tpds/
td1996/l0769abs.htm.

Wong:1994:SPE

[WOG94] W. F. Wong, Y. Oyanagi, and E. Goto. Supercomputer performance evaluation using six benchmarks. In Chan [Cha94a], pages 1107–1111. ISBN 0-7803-1863-3, 0-7803-1862-5, 0-7803-1864-1. LCCN QA75.5.P735 1994. IEEE catalog number 94CH3417-3.

Wong:2000:SUB

[WOK^{+}00] Adrian Wong, Leonid Oliker, William Kramer, Teresa Kaltz, and David Bailey. System utilization benchmark on the Cray T3E and IBM SP. Lecture Notes in Computer Science, 1911:56–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1911/19110056.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1911/19110056.
                                        pdf.

Womack:1990:PUG

[Wom90] Lucas A. Womack. PlotTool user’s guide v. 1.1. Technical report SRC-TR-90-024, Supercomputing Research Center: IDA, Lanham, MD, USA, August 14, 1990. 35 pp.

Woodruff:1992:SCC

[Woo92]  S. B. Woodruff.  Some

computational challenges of developing efficient parallel algorithms for datadependent computations in thermal-hydraulics supercomputer applications. In ANS [ANS92], pages 463– 472 (or 1615–1624??). ISBN 0-89448-178-9. ISSN 0029-5493. LCCN TK9202.I537 1992. Six volumes.

Woodward:1993:SVS

[Woo93] P. R. Woodward. Scientific visualization of supercomputer simulations of fluid flow. AHPCRC preprint 93-040, Army High Performance Computing Research Center, Minneapolis, MN, USA, 1993. iii + 13 pp.

Woodruff:1994:SCC

[Woo94] S. B. Woodruff. Some computational challenges of developing efficient parallel algorithms for data-dependent computations in thermalhydraulics supercomputer applications. Nuclear engineering and design: an international journal devoted to the thermal, mechanical and structural problems of nuclear energy, 146(1/3):463– ??, February 1994. ISSN 0029-5493.

Woods:1996:ESC

[Woo96a] A. W. Woods. The effect of speed controls in a continuum model of traffic flow. In Wolf et al. [WSB96], pages 175–180. ISBN 981-02-2635-7. LCCN ????

Woodward:1996:PST

[Woo96b] Paul R. Woodward. Perspective on supercomputing: Three decades of change. Computer, 29(10): 99–111, October 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woo:2005:SAJ

[Woo05] Jongwook Woo. Static analysis for Java with alias representation reference-set in high-performance computing. Scalable Computing: Practice and Experience, 6(1):125–139, March 2005. CODEN ???? ISSN 1895-1767. URL http://

                                        www.scpe.org/vols/vol06/
                                        SCPE_6_1_10.pdf;   http:
                                        //www.scpe.org/vols/vol06/
                                        SCPE_6_1_10.zip.

Worlton:1981:PS

[Wor81] Jack Worlton. A philosophy of supercomputing. Report LA-8849-MS, National Technical Information Service, Washington, DC, USA, 1981. 31 pp.

Worlton:1984:USB

[Wor84] J. Worlton. Understanding supercomputer benchmarks. Datamation, 30(14):121–130, 1984. CODEN DTMNAT. ISSN 0011-6963.

Wehner:2009:RCC

[WOS09] M. Wehner, L. Oliker, and J. Shalf. A real cloud computer. IEEE Spectrum, 46(10):24–29, October 2009. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Westphal:1994:AFD

[WP94] H. Westphal and D. Popovic.

Application fundamentals of distributed, high-performance computing and networking systems. Lecture Notes in Computer Science, 797:36– ??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Webb:1992:DSG

[WQS92] C. Webb, F. Que, and P. R.

Senior. Dynamic simulation of gas-liquid dispersion behaviour in a 2-D bubble column using a graphics minisupercomputer. Chemical engineering science, 47(13 / 14):3305–??, 1992. CODEN CESCAC. ISSN 0009-2509.

Wichern:1995:ADD

[WR95] B. Wichern and P. Rujan.

Analysis of dopamine diffusion in the extracellular space of the retina. In Herrmann et al. [HWP95], pages 273– 280. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Westerink:1997:ICS

[WR97] J. J. Westerink and P. J.

Roache. Issues in convergence studies in geophysical flow computations. In Delic and Wheeler [DW97], pages 95–108. ISBN 0-89871-378-1. LCCN ????

Wright:2019:PMB

[Wri19] Steven A. Wright. Performance modeling, benchmarking and simulation of high performance computing systems. Future Generation Computer Systems, 92(??):900–902, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0167739X18328590.

Ward:1993:NMC

[WRW93] C. M. Ward, R. A. Rydin, and M. L. Woosley. A new method of calculating resonance neutron absorption in non-uniform heterogeneous arrays. In Kusters et al. [KSW93], pages 535–

546. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Weiss:1984:IILa

[WS84a] S. Weiss and J. E. Smith. Instruction issue logic in pipelined supercomputers. IEEE Transactions on Computers, C-33(11):1013–1022, November 1984. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=1676375.

Weiss:1984:IIL

[WS84b] Shlomo Weiss and James E.

Smith. Instruction issue logic for pipelined supercomputers. ACM SIGARCH Computer Architecture News, 12(3):110–118, June 1984. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Weiss:1984:IILb

[WS84c] Shlomo Weiss and James E.

Smith. Instruction issue logic for pipelined supercomputers. ACM SIGARCH Computer Architecture News, 12(3):110–118, June 1984. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Williamson:1984:NWP

[WS84d] David L. Williamson and Paul N. Swarztrauber. Numerical weather prediction model — computational aspects on the Cray-1. Proceedings of the IEEE, 72(1): 56–67, January 1984. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Weiss:1987:SSCa

[WS87a] Shlomo Weiss and James E.

Smith. A study of scalar compilation techniques for pipelined supercomputers. ACM SIGARCH Computer Architecture News, 15(5): 105–109, October 1987. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Weiss:1987:SSCb

[WS87b] Shlomo Weiss and James E. Smith. A study of scalar compilation techniques for pipelined supercomputers. ACM SIGPLAN Notices, 22 (10):105–109, October 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weiss:1987:SSCc

[WS87c] Shlomo Weiss and James E. Smith. A study of scalar compilation techniques for pipelined supercomputers. Operating Systems Review, 21(4):105–109, October 1987. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Weiss:1990:SSC

[WS90] Shlomo Weiss and James E. Smith. A study of scalar compilation techniques for pipelined supercomputers. ACM Transactions on Mathematical Software, 16(3):223– 245, September 1990. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/toms/1990-16-3/
                                        p223-weiss/.

Weiss:1993:BSP

[WS93] R. Weiss and W. Schoenauer.

Black-box solvers for partial differential equations. In Kusters et al. [KSW93], pages 29–40. ISBN 3-923704-11-9. LCCN ???? Two volumes.

White:1999:FUS

[WS99] Simon D. M. White and Volker Springel. Fitting the universe on a supercomputer. Computing in Science and Engineering, 1(2): 36–45, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c2036.
pdf;   http://www.computer.
org/cse/cs1999/c2036abs.
htm.

Wolf:1996:WOJ

[WSB96] D. E. Wolf, M. Schreckenberg, and A. Bachem, editors. Workshop — 1995 Oct: Julich; Germany. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1996. ISBN 981-02-2635-7. LCCN ????

Wasserman:1988:PMA

[WSL88] H. J. Wasserman, M. L. Simmons, and O. M. Lubeck. The performance of minisupercomputers: Alliant FX/ 8, Convex C-1, and SCS-40. Parallel Computing, 8(1–3): 285–293, October 1988. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Wennekers:1995:IRA

[WSP95] T. Wennekers, F. T. Sommer, and G. Palm. Iterative retrieval in associative memories by threshold control of different neural models. In Herrmann et al. [HWP95], pages 301–320. ISBN 981-02-2250-5. LCCN QP356.W67 1994.

Wu:2011:PCH

[WT11] Xingfu Wu and Valerie Taylor. Performance characteristics of hybrid MPI/OpenMP implementations of NAS parallel benchmarks SP and BT on large-scale multicore supercomputers. ACM SIGMETRICS Performance Evaluation Review, 38(4):56– 62, March 2011. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Wu:2013:PMH

[WT13] Xingfu Wu and Valerie Taylor. Performance modeling of hybrid MPI/OpenMP scientific applications on largescale multicore supercomputers. Journal of Computer and System Sciences, 79(8): 1256–1268, December 2013. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0022000013000639.

Wang:2002:DPC

[WTC^{+}02] Cho-Li Wang, Anthony T. C.

Tam, Benny W. L. Cheung, Wenzhang Zhu, and David C. M. Lee. Directed Point: a communication subsystem for commodity supercomputing with Gigabit Ethernet. Future Generation Computer Systems, 18(3):401– 420, January 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/32/33/abstract.
html.

Wunderlich:1989:IWC

[Wun89] Marvin C. Wunderlich. An investigation of Waring’s conjecture on the Connection FFMachine: SRC 1988 summer workshop. Technical report SRC-TR-89-011, Supercomputing Research Center: IDA, Lanham, MD, USA, December 18, 1989. 13 pp.

Wuorinen:1994:DTP

[Wuo94] John H. Wuorinen, editor.

Digest of technical papers: 1994 IEEE International Solid-State Circuits Conference, February 1994, Castine, ME, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-1845-5, 0-7803-1844-7, 0-7803-1846-3. ISSN 0193-6530. LCCN ???? IEEE catalog number 94CH3410-8.

Wojcik:1988:LEW

[WVBM88a] G. L. Wojcik, D. K. Vaughan, M. Barenberg, and J. Mould. Large-scale, explicit wave simulations on the Cray-2. Applied Numerical Mathematics: Transactions of IMACS, 4(1):47–70, March 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wojcik:1988:LSE

[WVBM88b] G. L. Wojcik, D. K. Vaughan, M. Barenberg, and J. Mould. Large-scale, explicit wave simulations on the Cray-2. Applied Numerical Mathematics: Transactions of IMACS, 4(1):47–70, March 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wenisch:2007:CSD

[WvTB^{+}07] Petra Wenisch, Christoph van Treeck, André Borrmann, Ernst Rank, and Oliver Wenisch. Computational steering on distributed systems: Indoor comfort simulations as a case study of interactive CFD on supercomputers. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 22(4):275–291, ???? 2007. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic). URL http:

//www.informaworld.com/
smpp/content~content=a779508970.

Weerawarana:1992:PCG

[WW92] Sanjiva Weerawarana and Paul S. Wang. A portable code generator for CRAY FORTRAN. ACM Transactions on Mathematical Software, 18(3):241–255, September 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toms/1992-18-3/
p241-weerawarana/.

Wang:2009:TAR

[WWJ09] Shuo Wang, Lei Wang, and Faquir Jain. Towards achieving reliable and highperformance nanocomputing via dynamic redundancy allocation. ACM Journal on Emerging Technologies in Computing Systems (JETC), 5(1):2:1–2:??, January 2009. CODEN ???? ISSN 1550-4832.

Weaver:1997:FTR

[WWKR97] J. W. Weaver, J. T. Wilson, D. H. Kampbell, and M. E. Randolph. Field-derived transformation rates for modeling natural bioattenuation of trichloroethene and its degradation products. In Delic and Wheeler [DW97], pages 177–186. ISBN 0-89871-378-1. LCCN ????

Whitson:1992:CIN

[WWTE92] George M. Whitson, Cathy Wu, John Taylor, and Adisorn Ermongkonchai. CANS: an interactive neural network system for Cray supercomputers. In Applied Computing: Technological Challenges of the 1990’s (Mar 1– 3 1992: Kansas City, KS, USA), pages 665–668. ACM, New York, NY, USA, 1992. ISBN 0-89791-502-x. LCCN ????

Wu:1993:PCA

[WWY93] C. H. Wu, C.-C. Wang, and I. Yazdanpanah. Protein classification artificial neural system: a filter program for database search. In Lim et al. [L^{+}93], pages 349–358. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Wallich:1987:MMV

[WZ87] P. Wallich and Glenn Zorpette. Minis and mainframes: VAX family grows at a dizzying pace minicomputer makers merge to survive minisupers emulate VAXes and Cray’s Connection Machine is an AI supercomputer. IEEE Spectrum, 24(1):28–31, January 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Walker:1997:SOP

[WZ97] M. B. Walker and J. Ziebarth. Selected outreach programs at the National Center for Supercomputing Applications. In EP Innovations [EP 97], pages 391–397. ISBN 0-7803-4087-6 (casebound), 0-7803-4086-8 (softbound), 0-7803-4088-4 (microfiche), 0-7803-4089-2 (CD-ROM). ISSN 0190-5848. LCCN T62 .F76 1997. Three volumes. IEEE catalog number: 97CH36099.

Wallich:1986:MMM

[WZB86] P. Wallich, Glenn Zorpette, and C. G. Bell. Minis and mainframes: Microcomputers as we know them may disappear as sophisticated microprocessors reach toward mainframe capability; more powerful supercomputers are introduced. IEEE Spectrum, 23(1):36–39, January 1986. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Xiao:1996:MOM

[XB96] Y. Y. Xiao and J. K. Bennett. Memory organization in multi-channel optical networks: NUMA and COMA revisited. In ACM [ACM96], pages 26–34. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Xin:1993:UBN

[XCLW93] Y. Xin, T. T. Carmeli, M. N.

Liebman, and G. L. Wilcox. Use of the backpropagation neural network algorithm for prediction of protein folding patterns. In Lim et al. [L^{+}93], pages 359–376. ISBN 981-02-1157-0. LCCN QH445.2 .I57 1992.

Xia:1988:PWC

[Xia88] Eugene Zhu Xia. Parallel waveform-relaxation-newton for circuit simulation. Technical Report CSRD 772; UILU-ENG-88-8004, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 61 pp.

Xu:1994:DRD

[XL94] C.-Z. Xu and F. C. M. Lau. Decentralized remapping of data parallel computations with the generalized dimension exchange method. In IEEE [IEE94c], pages 414– 421. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Xue:1992:MLP

[XMR92] G.-L. Xue, R. S. Maier, and J. B. Rosen. Minimizing the Lennard-Jones potential function on a massively parallel computer. In ACM [ACM92b], pages 409–416. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Xie:2020:COB

[XOZ^{+}20] Bing Xie, Sarp Oral, Christopher Zimmer, Jong Youl Choi, David Dillow, Scott Klasky, Jay Lofstead, Norbert Podhorszki, and Jeffrey S. Chase. Characterizing output bottlenecks of a production supercomputer: Analysis and implications. ACM Transactions on Storage, 15(4):26:1–26:39, February 2020. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3335205.

Xu:1991:SAC

[Xu91] Xinmin Xu. Stress analysis of cylinder-to-cylinder shell intersections by use of supercomputer. Thesis (M.S. in mechanical engineering), University of Illinois at Chicago, Chicago, IL, USA, 1991. ix + 70 pp.

Xia:2020:DAB

[XZC^{+}20] Chunwei Xia, Jiacheng Zhao, Huimin Cui, Xiaobing Feng, and Jingling Xue. DNNTune: Automatic benchmarking DNN models for mobile-cloud computing. ACM Transactions on Architecture and Code Optimization, 16(4):1– 26, January 2020. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3368305.

Yang:1992:DSI

[Y^{+}92] Y.-K. Yang et al. Development of supercomputer image processing software with X-Window user-interface for the processing of the remotely sensed data. In Fritz and Lucas [FL92], pages 235–239. ISBN ???? ISSN 0256-1840. LCCN ????

Yavuz:1993:ADT

[YA93] M. Yavuz and C. Aykanat. Alternating direction transport sweeps for the linear discontinuous SN method. In Kusters et al. [KSW93], pages 764–776. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Yang:1993:EET

[YAG93] J. Yang, I. Ahmad, and A. Ghafoor. Estimation of execution times on heterogeneous supercomputer architectures. In Chen et al. [CBCH93], pages I–219–I– 226. ISBN 0-8493-8983-6 (set), 0-8493-8984-4 (vol. 1), 0-8493-8985-2 (vol. 2), 0-8493-8986-0 (vol. 3). ISSN 0190-3918. LCCN QA76.58 .I55 1993 v.1-3 (c1993).

Yang:19xx:EET

[YAGxx] J. Yang, I. Ahmad, and A. Ghafoor. Estimation of execution times on heterogeneous supercomputer architectures. Proceedings of the International Conference on Parallel Processing, ????

(????):I–219, ???? 19xx. CO-DEN PCPADL. ISSN 0190-3918.

Yang:1990:DPD

[Yan90a] Gung-Chung Yang. DSPACK:

a parallel direct sparse matrix package for shared-memory multiprocessors. Technical Report CSRD 1085, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1990. 4 pp.

Yang:1990:PPS

[Yan90b] Gung-Chung Yang. PARASPICE:

a parllel [sic] circuit simulator for shared-memory multiprocessors. Technical Report CSRD 1088, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, June 1990. 6 pp.

Yang:1990:PPP

[Yan90c] Gung-Chung Yang. PARASPICE:

a portable parallel circuit simulator. Technical Report CSRD 1086, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1990. 11 pp. Published in Advances in Electrical Engineering Software (Electrosoft ’90): Proceedings of the First International Conference on Electrical Engineering Analysis and Design, Lowell, MA, pp. 3–14, August 1990. Editor: P.P. Silvester, Computationsl [sic] Mechanics Publications, Springer-Verlag.

Yang:1991:PSS

[Yan91] Gung-Chung Yang. Parallelizing Spice2 on sharedmemory multiprocessors. Technical Report CSRD 1148, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1991. 8 pp.

Yang:1992:PCG

[Yan92] Ulrike Meier Yang. Preconditioned conjugate gradientlike methods for nonsymmetric linear systems. Technical Report CSRD 1210, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1992. 54 pp.

Yang:1993:PCM

[Yan93] Qing Yang. Performance of cache memories for vector computers. Journal of Parallel and Distributed Computing, 19(3):163–178, November 1993. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.

1993.1102/production;
http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1993.1102/production/pdf.

Yang:1994:AMN

[Yan94]  Y. Yang.  An analysis

model on nonblocking multirate broadcast networks. In Anonymous [Ano94-134], pages 256–263. ISBN ???? LCCN ????

Yau:1988:NOS

[Yau88] Benjamin Oyman Yau. A network oriented SX-2 supercomputer access facility. Thesis (M.S.), Dept. of Computer Science, University of Houston, Houston, TX, USA, 1988. vi + 92 pp.

Young:1986:TFN

[YB86] Jeff Young and Duncan A.

Buell. The twentieth Fermat number is composite. Technical report SRC-TR-86-002, Supercomputing Research Center: IDA, Lanham, MD, USA, 1986. 2 pp.

Yew:1990:SSE

[YB90] Pen-Chung Yew and John D.

Bruner. SEE: a system evaluation environment for studying parallel systems. Technical Report CSRD 1051, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, December 1990. 8 pp.

Yeh:1997:NMF

[YCC97] G.-T. Yeh, J.-R. Cheng, and H.-P. Cheng. Numerical modeling of flow and contaminant transport in the subsurface. In Delic and Wheeler [DW97], pages 247– 258. ISBN 0-89871-378-1. LCCN ????

Yeich:1992:TP

[Yei92] Christopher R. Yeich. Tower of power. Chilton’s automotive industries, 172(6):69– ??, June 1, 1992. CODEN CAUIEG. ISSN 0273-656X.

Yew:1988:ACP

[Yew88] Pen-Chung Yew. Architecture of the Cedar parallel computer. Technical Report CSRD 609, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. 137–148 pp.

Yokono:1995:ISS

[YF95] Y. Yokono and H. Fujita. Interactive steering of supercomputing simulation for aerodynamic noise radiated from square cylinder. In Lyrintzis et al. [L^{+}95], pages 59–64. ISBN 0-7918-1474-2. LCCN TL574.N6C66 1995.

Yang:1998:SSE

[YF98] Tao Yang and Cong Fu. Space/time-efficient scheduling and execution of parallel irregular computations.

ACM Transactions on Programming Languages and Systems, 20(6):1195–1222, November 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-
6/p1195-yang/.

Feng:1989:SIS

[yFH89] Tse yun Feng and A. R. Hurson, editors. Special issue on supercomputer technology, volume 77(12) of Proceedings of the IEEE. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989.

Yamazaki:1993:PSG

[YFOT93] T. Yamazaki, M. Fujisaki, M. Okuda, and M. Takano. A parallelization study of the general purpose Monte Carlo code MCNP4 on a distributed memory highly parallel computer. In Kusters et al. [KSW93], pages 374– 383. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Yang:2013:AHA

[YFY^{+}13] Donghua Yang, Yuqiang Feng, Ye Yuan, Xixian Han, Jinbao Wang, and Jianzhong Li. Ad-hoc aggregate query processing algorithms based on bit-store for query intensive applications in cloud computing. Future Generation Computer Systems, 29(7): 1725–1735, September 2013. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167739X12000623.

Yang:1992:PST

[YG92] T. Yang and A. Gerasoulis. PYRROS: Static task scheduling and code generation for message passing multiprocessors. In ACM [ACM92b], pages 428–437. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Yang:1994:HPF

[YGSB94] S. X. Yang, D. Gannon, S. Srinivas, and F. Bodin. High Performance Fortran interface to the Parallel C++. In IEEE [IEE94c], pages 301–

308. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Yi:1990:OFS

[YH90] Kwang Keun Yi and Luddy Harrison. On-the fly [sic] circuit to measure the average working set size. Technical Report CSRD 1095, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1990. 4 pp.

Yi:1992:IDF

[YH92] Kwang Keun Yi and Williams Ludwell Harrison. Interprocedural data flow analysis for compile-time memory management. Technical Report CSRD 1244, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1992. 47 pp.

Yi:1993:PEV

[YHA93] S. Yi, H. H. Hilton, and M. F. Ahmad. Performance evaluation of viscoelastic finite element supercomputer algorithms. In Brebbia and Power [BP93], pages 497– 512. ISBN 1-85166-845-4, 1-85312-236-X, 1-56252-160-8. LCCN TA345.A66 1993.

Hsu:1992:IWC

[yHY92] William Tsun yuk Hsu and Pen-Chung Yew. The impact of wiring constraints on hierarchical network performance. Technical Report CSRD 1156, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, March 1992. 11 pp.

Hsu:1987:ESH

[yHYZ87] William Tsun yuk Hsu, Pen-Chung Yew, and Chuan-Qi Zhu. An enhancement scheme for hypercube interconnection networks. Technical Report CSRD-628, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 21 pp.

Yi:1990:OMM

[Yi90] Kwang Keun Yi. On-the-fly [methods] to measure the locality of programs. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1990. iv + 60 pp.

Yi:2011:PEG

[Yi11] Hongsuk Yi. Performance evaluation of GAIA supercomputer using NPB multizone benchmarks. Computer Physics Communications, 182(1):263–265, January 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465510003437.

Yan:1993:PRL

[YJD93] H. Yan, Z. Jin, and X. Ding. Preliminary results of a limited area model (YH model) on parallel computing environment. In Hoffmann and Kauranne [HK93b], pages 88–98. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Yuba:1987:JNP

[YK87] T. Yuba and H. Kashiwagi. The Japanese national project for new generation supercomputing systems. Parallel Computing, 4 (1):1–16, February 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Yamakado:1994:JSA

[YK94] M. Yamakado and Y. Kadomukai. A jerk sensor and its application to vehicle motion control systems. In Anonymous [Ano94-75], pages 139– 146. ISBN 0-947719-68-7. LCCN ????

Yang:2000:RPS

[YKB^{+}00] Pamela Yang, Timothy M.

Kunau, Bonnie Holte Bennett, Emmett Davis, and Bill Wren. Reconfigurable parallel sorting and load balancing on a Beowulf cluster: HeteroSort. Lecture Notes in Computer Science, 1800:862– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18000862.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000862.
pdf.

Yoshida:1996:DFM

[YKK96] A. Yoshida, K. Koshizuka, and H. Kasahara. Datalocalization for Fortran macrodataflow computation using partial static task assignment. In ACM [ACM96], pages 61–68. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Yang:1990:ICS

[YKY90] A. T. Yang, S. M. Kang, and Gung-Chung Yang. An integrated CAD system for device model design, parameter extraction and circuit simulation. Technical Report CSRD 1087, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1990. 12 pp.

Yasar:1993:RFP

[YMT93] O. Yasar, G. A. Moses, and T. J. Tautges. Radiationmagnetohydrodynamics of fusion plasmas on parallel supercomputers. In Kusters et al. [KSW93], pages 51–62. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Yasugi:1992:AON

[YMY92] M. Yasugi, S. Matsuoka, and A. Yonezawa. ABCL/ onEM-4: a new software/ hardware architecture for object-oriented concurrent computing on an extended dataflow supercomputer. In ACM [ACM92b], pages 93– 103. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Yang:1990:FSK

[YMZ90] H.-H. Yang and L. Manas-Zloczower. Flow simulations in the kneading discs region of a co-rotating twin screw extruder. In Pitcher [Pit90], pages 273–284. ISBN 1-85312-115-0 (Southampton), 0-945824-99-8 (Boston), 3-540-53226-9 (Heidelberg), 0-387-53226-9 (New York). LCCN QA76.5 .S355 1990.

Yoshida:2009:SSS

[Yos09] Naoki Yoshida. Supercomputer simulations of structure formation in the Universe. Computer Physics Communications, 180(4): 625–627, April 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465508004487.

Yamaguchi:1997:EAC

[YOY97] H. Yamaguchi, M. Osawa, and M. Yamamoto. Estimation of ambiguous change in 4WS control system using generalized likelihood ratio. In Roller [Rol97], pages 121–128. ISBN 0-947719-88-1 (paperback). LCCN ????

Yu:2012:HHC

[YQTV12] Weikuan Yu, Xinyu Que, Vinod Tipparaju, and Jeffrey S. Vetter. HiCOO:

Hierarchical cooperation for scalable communication in Global Address Space programming models on Cray XT systems. Journal of Parallel and Distributed Computing, 72(11):1481–1492, November 2012. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0743731512000342.

Yasar:1993:PKI

[YR93] O. Yasar and C. J. Rutland. Parallelization of KIVA-II on the iPSC-860 supercomputer. In Pelz et al. [PEH93], pages 419–426. ISBN 0-444-89986-

3. LCCN QC150 .C66 1992.

Yan:1994:UUS

[YS94] X. T. Yan and J. E. E. Sharpe. The use of a unified simulation system in the design of mechatronic schemes in automobile. In Anonymous [Ano94-75], pages 551–

558. ISBN 0-947719-68-7. LCCN ????

Yamada:1996:HTS

[YSK^{+}96] Masahiro Yamada, Fumihiko Sakamoto, Tsutomu Katoh, Takashi Oguri, Hironobu Ikeda, and Masuo Yamazaki. Hardware technologies for supercomputer SX-4. Nippon Electric Company research and development, 37(4):493– ??, ???? 1996. CODEN NECRAU. ISSN 0048-0436.

Yamana:1995:MUS

[YSKS95] H. Yamana, M. Sato, Y. Kodama, and H. Sakane. A macrotask-level unlimited speculative execution on multiprocessors. In ACM [ACM95a], pages 328–337. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Yvars:1997:UCP

[YSL97] P. A. Yvars, C. Salvan, and C. Lenguin. Using constraint programming techniques for integrating design knowledge in CAD/CAM systems. In Roller [Rol97], pages 193– 200. ISBN 0-947719-88-1 (paperback). LCCN ????

Yan:1994:MPM

[YSS94] J. C. Yan, M. A. Schmidt, and S. Sarukkai. Monitoring the performance of multidisciplinary applications on the iPSC/860. In IEEE [IEE94c], pages 277–284. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Yew:1987:DHA

[YTL87] Pen-Chung Yew, Nian-Feng Tzeng, and Duncan Lawrie. Distributing hot-spot addressing in large-scale multiprocessors. Technical Report CSRD-619, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1987. 8 pp.

Yuval:1977:CH

[Yuv77] G. Yuval. Cray-1 for 6000… 7000 hackers. Software —Practice and Experience, 7(3):427–428, June 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yew:1989:CPS

[YVC89] Pen-Chung Yew, Alexander Veidenbaum, and Hoichi Cheong. Chief: a parallel simulation environment for parallel systems. Technical Report CSRD 915; UILU-ENG-89-8005, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, August 1989. ii + 13 pp.

Yerkes:1994:IWS

[YW94] C. R. Yerkes and E. Webster. Implementation of wk synthetic aperture radar imaging algorithm on a massively parallel supercomputer [2230-12]. In Giglio [Gig94], pages 171–179. ISBN 0-8194-1534-0. ISSN 0361-0748. LCCN TK6592.S95 A44 1994.

Yerkes:1994:ISA

[YWD94] C. Yerkes, E. Webster, and P. D’Arnaud. Interferometric synthetic aperture radar processing on a massively parallel supercomputer. In Singh [Sin94a], pages 132– 136. ISBN 0-8186-6405-3, 0-8186-6406-1, 0-7803-2614-8. ISSN 1058-6393. LCCN TK 5102.5 A78 1994. Two volumes.

Yerkes:19xx:ISA

[YWDxx] C. Yerkes, E. Webster, and P. D’Arnaud. Interferometric synthetic aperture radar processing on a massively parallel supercomputer. Conference record, ????(????):132– ??, ???? 19xx. ISSN 1058-6393.

Yang:2012:RWE

[YWXZ12] Xuejun Yang, Zhiyuan Wang, Jingling Xue, and Yun Zhou. The reliability wall for exascale supercomputing. IEEE Transactions on Computers, 61(6):767–779, June 2012. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yoshida:1993:NMV

[YYK93] K. Yoshida, S. Yamamoto, and T. Kuriyama. New meshgeneration and verification of multidimensional port and in-cylinder flow simulation. In Anonymous [Ano93-31], pages 807–814. ISBN 0-947719-62-8. LCCN ????

Yu:2020:SBS

[YYW^{+}20] J. Yu, W. Yang, F. Wang, D. Dong, J. Feng, and Y. Li.

Spatially bursty I/O on supercomputers: Causes, impacts and solutions. IEEE Transactions on Parallel and Distributed Systems, 31(12): 2908–2922, 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yagawa:1993:PFE

[YYYS93] G. Yagawa, A. Yoshioka, S. Yoshimura, and N. Soneda. A parallel finite element method with a supercomputer network. Computers and Structures, 47(3): 407–??, May 1993. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Yu:2020:LSA

[YZL^{+}20] T. Yu, W. Zhao, P. Liu, V. Janjic, X. Yan, S. Wang, H. Fu, G. Yang, and J. Thomson. Large-scale automatic K-means clustering for heterogeneous many-core supercomputer. IEEE Transactions on Parallel and Distributed Systems, 31(5):997– 1008, May 2020. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zage:1982:IPS

[Zag82] Wayne M. Zage. Information processing in a supercomputer. Thesis (M.A.), Ball State University, Muncie, IN, USA, 1982. v + 52 pp.

Zaslavsky:1993:ASM

[Zas93] L. Y. Zaslavsky. An adaptive self-stabilized multigrid method for the multigroup neutron diffusion. In Kusters et al. [KSW93], pages 15–26. ISBN 3-923704-11-9. LCCN ???? Two volumes.

Zaidi:1994:PSR

[ZAS94] Zafar H. Zaidi, Atiya Abbasi, and David L. Smith, editors. Protein structurefunction relationship: proceedings of the International Symposium on Protein Structure Function Relationship, held in Karachi, Pakistan, 8–12 January. 1993, volume 66(1). TWEL Publishers, Karachi, Pakistan, 1994. ISBN 969-8117-05-9. ISSN 0033-4545. LCCN QP551 .I553 1993. Also in Pure and Applied Chemistry, vol. 66, no. 1 (1994).

Zitney:1995:PDS

[ZBLZ95] S. E. Zitney, L. Brull, L. Lang, and R. Zeller. Plantwide dynamic simulation on supercomputers: Modeling a Bayer distillation process. American Institute of Chemical Engineers symposium series, 91(304):313– ??, ???? 1995. CODEN AC-SSCQ. ISSN 0065-8812.

Zolfaghari:2019:HOA

[ZBN^{+}19] Hadi Zolfaghari, Barna Becsek, Maria G. C. Nestola, William B. Sawyer, Rolf Krause, and Dominik Obrist. High-order accurate simulation of incompressible turbulent flows on many parallel GPUs of a hybridnode supercomputer. Computer Physics Communications, 244(??):132–142, November 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0010465519301997.

Zanghirati:2000:CTI

[ZCPT00] G. Zanghirati, F. Cocco, G. Paruolo, and F. Taddei. A Cray T3E implementation of a parallel stochastic dynamic assets and liabilities management model. Parallel Computing, 26(5):539–567, March 2000. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:/

/www.elsevier.nl/gej-ng/
10/35/21/42/27/25/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/42/27/
25/article.pdf.

Zecca:1993:HPI

[Zec93] V. Zecca. High performance I/O in supercomputing applications. In Brebbia and Power [BP93], pages 439–

446. ISBN 1-85166-845-4, 1-85312-236-X, 1-56252-160-

8. LCCN TA345.A66 1993.

Zhang:2017:ESA

[ZEC^{+}17] Boyu Zhang, Trilce Estrada, Pietro Cicotti, Pavan Balaji, and Michela Taufer. Enabling scalable and accurate clustering of distributed ligand geometries on supercomputers. Parallel Computing, 63(??):38–60, April 2017. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167819117300261.

Zenios:1994:PSP

[Zen94] Stavros A. Zenios. Parallel and supercomputing in the practice of management science. Interfaces, 24 (5):122–??, September 1994. CODEN INFAC4. ISSN 0092-2102 (print), 1526-551X (electronic).

Zenios:1999:HPC

[Zen99] Stavros A. Zenios. Highperformance computing in finance: The last 10 years and the next. Parallel Computing, 25(13–14):2149– 2175, December 1999. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/35/21/32/36/47/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/35/21/32/36/
47/article.pdf.

Zeyher:1991:CCP

[Zey91] Allen Zeyher. Cray constructs powerful chemistry. Computers in Physics, 5(4):

369–??, July 1991. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:/

                                        /aip.scitation.org/doi/
                                        10.1063/1.4822999.

Zhao:2018:OCN

[ZFF^{+}18] Wenlai Zhao, Haohuan Fu, Jiarui Fang, Weijie Zheng, Lin Gan, and Guangwen Yang. Optimizing convolutional neural networks on the Sunway TaihuLight Supercomputer. ACM Transactions on Architecture and Code Optimization, 15(1): 13:1–13:??, April 2018. CO-DEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Zhu:2014:MSS

[ZGL14] Sheng-Xin Zhu, Tong-Xiang Gu, and Xing-Ping Liu. Minimizing synchronizations in sparse iterative solvers for distributed supercomputers. Computers and Mathematics with Applications, 67 (1):199–209, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122113006640.

Zabolitzky:1988:MCS

[ZH88] John G. Zabolitzky and Hans J. Herrmann. Multitasking case study on the Cray-2: The Q2R cellular automaton. Journal of Computational Physics, 76(2):426–447, June 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0021999188901507.

Zheng:1997:VSW

[Zhe97] Weimin Zheng. Virtual supercomputing on the World Wide Web. Thesis (M.S.), University of South Carolina, Columbia, SC, USA, 1997. 32 pp.

Zhong:1988:CVE

[Zho88] Jialin Zhong. Computer visualization of electron movement in gallium arsenide simulation. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1988. v + 34 pp.

Zima:1996:TSL

[Zim96] Hans P. Zima. Taking stock, looking ahead: Part 2. highperformance languages for parallel computing. IEEE Computational Science & Engineering, 3(3):63–65, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:

//www.computer.org/cse/
cs1998/c3063abs.htm.

Ziegler:1997:NMT

[ZL97] C. Kirk Ziegler and W. Lick. Numerical modeling of the transport and fate of hydrophobic contaminants and fine-grained sediments in surface waters. In Delic and Wheeler [DW97], pages 129– 138. ISBN 0-89871-378-1. LCCN ????

Zlatev:2001:ETL

[Zla01] Zahari Zlatev. Efficient treatment of large-scale air pollution models on supercomputers. Lecture Notes in Computer Science, 2074:82– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2074/20740082.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2074/20740082.
pdf.

Zhang:2021:CRSa

[ZLC21] Yiming Zhang, Kai Lu, and Wenguang Chen. China region special section: Hot topics: Processing extremescale graphs on China’s supercomputers. Communications of the ACM, 64(11): 60–63, November 2021. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/
3481614.

Zhang:2020:ECR

[ZLRC20] Yongmin Zhang, Xiaolong Lan, Ju Ren, and Lin Cai. Efficient computing resource sharing for mobile edge-cloud computing networks. IEEE/ACM Transactions on Networking, 28 (3):1227–1240, June 2020. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TNET.2020.2979807.

Zenios:1986:NNP

[ZM86] Stavros A. Zenios and John M. Mulvey. Nonlinear network programming on vector supercomputers: a study on the Cray X-MP. Operations Research, 34(5): 667–682, September–October 1986. CODEN OPREAI. ISSN 0030-364X (print), 1526-5463 (electronic).

Ziavras:1994:AMS

[ZM94] S. G. Ziavras and P. Meer.

Adaptive multiresolution structures for image processing on parallel computers. Journal of Parallel and Distributed Computing, 23(3): 475–483, December 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1994.1159/production;
http://www.idealibrary.

                                        com/links/doi/10.1006/jpdc.
                                        1994.1159/production/pdf.

Zitney:1996:MVF

[ZMDS96] S. E. Zitney, J. Mallya, T. A. Davis, and M. A. Stadtherr. Multifrontal vs frontal techniques for chemical process simulation on supercomputers. Computers & Chemical Engineering, 20(6):641– ??, ???? 1996. CODEN CCENDW. ISSN 0098-1354.

Zorpette:1989:TMM

[Zor89a] Glenn Zorpette. Technology ’89: minis and mainframes. IEEE Spectrum, 26 (1):29–31, January 1989. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1989:TME

[Zor89b] Glenn Zorpette. Technology ’89: the main event. IEEE Spectrum, 26(1):28, January 1989. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1990:TMM

[Zor90] Glenn Zorpette. Technology ’90: minis and mainframes. IEEE Spectrum, 27 (1):30–34, January 1990. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1991:TMM

[Zor91] Glenn Zorpette. Technology 1991: minis and mainframes. IEEE Spectrum, 28 (1):40–43, January 1991. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1992:RDM

[Zor92a]  Glenn Zorpette.  A re-

markable diversity of massively parallel machines is ushering in a new era in high-performance computing. IEEE Spectrum, 29(9):28– 33, September 1992. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1992:SPP

[Zor92b] Glenn Zorpette. Supercomputers — the power of parallelism. IEEE Spectrum, 29 (9):28–33, September 1992. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1993:HBI

[Zor93a]  Glenn Zorpette.  Henry

Burkhardt III. IEEE Spectrum, 30(3):62–??, March 1, 1993. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1993:LC

[Zor93b] Glenn Zorpette. Large computers. IEEE Spectrum, 30 (1):34–37, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zorpette:1993:TLC

[Zor93c] Glenn Zorpette. Technology 1993 — large computers. IEEE Spectrum, 30(1): 34–37, January 1993. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zitney:1993:SSD

[ZS93] Stephen E. Zitney and Mark A. Stadtherr. Supercomputing strategies for the design and analysis of complex separation systems. Industrial and engineering chemistry research, 32(4): 604–??, April 1993. CODEN IECRED. ISSN 0888-5885.

Zadzaonkar:1994:HCN

[ZS94a] A. S. Zadzaonkar and A. Shukla. Hidden control neural network architecture for high speed speech recognition. In Mahajan et al. [M^{+}94], pages 436–445. ISBN 0-07-462240-4. LCCN T385 .S37 1994. Rs387.00.

Ziavras:1994:HEH

[ZS94b] S. G. Ziavras and D. P. Shah. High-performance emulation of hierarchical structures on hypercube supercomputers. Concurrency: practice and experience, 6(2): 85–100, April 1994. CODEN CPEXEI. ISSN 1040-3108.

Ziavras:1994:HPE

[ZS94c] S. G. Ziavras and D. P.

Shah. High-performance emulation of hierarchical structures on hypercube supercomputers. Concurrency: practice and experience, 6(2): 85–100, April 1994. CODEN CPEXEI. ISSN 1040-3108.

Zola:2002:EJH

[ZW02] Jarosław Zola and Roman Wyrzykowski. EPL-Julia the high-performance library for evolutionary computations. Lecture Notes in Computer Science, 2328:652–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2328/23280652.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2328/23280652.
pdf.

Ziavras:2003:VAH

[ZWP03] Sotirios G. Ziavras, Qian Wang, and Paraskevi Papathanasiou. Viable architectures for high-performance computing. The Computer Journal, 46(1):36–54, January 2003. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_
journal/hdb/Volume_46/Issue_
01/460036.sgm.abs.html;
 http://www3.oup.co.uk/

                                        computer_journal/hdb/Volume_
                                        46/Issue_01/pdf/460036.
                                        pdf.

Zhang:1995:MSP

[ZX95] X. Zhang and Z. Xu. Multiprocessor scalability predictions through detailed program execution analysis. In ACM [ACM95a], pages 97– 106. ISBN 0-89791-728-6. LCCN QA 76.88 I57 1995. ACM order number: 415951.

Zygielbaum:1993:ESS

[Zyg93] Arthur Zygielbaum, editor. Earth and space science information systems: Proceedings of the International Space Year Conference on Earth and Space Science Information Systems, 10–13 Feb., 1992, number 283 in AIP Conference Proceedings. American Institute of Physics, Woodbury, NY, USA, 1993. ISBN 1-56396-094-X. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN QE48.8 .E27 1992. DOE-CONF-9202175.

Zhou:2016:IBS

[ZYL^{+}16] Zhou Zhou, Xu Yang, Zhiling Lan, Paul Rich, Wei Tang, Vitali Morozov, and Narayan Desai. Improving batch scheduling on Blue Gene/Q by relaxing network allocation constraints. IEEE Transactions on Parallel and Distributed Systems, 27(11):3269–3282, November 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL https://www.
computer.org/csdl/trans/
td/2016/11/07404249-abs.
html.